A Complete Bibliography of Publications in the SIAM Journal on Control and Optimization

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

24 August 2024 Version 3.22

Title word cross-reference

#15374 [Kum88]. #18045 [Cou83]. #9528 [FR82a].

(A, B) [Emr80a]. (A, B) [Hir81]. (ad_{f,g}) [Kre85]. (Ad_{f,g}), (ad_{f,g}) [Kre85]. (C, A) [ISW00]. (f, g) [DCBT88, DM89]. (J, J′) [CH05]. (A, B) [EH80]. (n, m) [Enq04]. (s, S) [BB09, BLS05a]. (sE − A) [Arm86]. + [GBCG98]. 1 [BP97, BE13, Dág06, DK89b, Gag16, Gue14, HH19, Hay19, HVDK19, KK17b, KR15, LY23, MSH91, MT22, RW01, VZ09, XZ20]. 2 [BBG12, BG13, FRZ93, JL93, KMC21, PP13, Set93, Zam98, ZH05]. 2 × 2 [CVKB13, PR14]. 3 [BRS11, Bre86, CC16, CM12, CYL96, GGMZRB20, Ner21, Wan02, ZM11]. 4 [CYL97b]. 5 [Sac19]. ^{∞} [GR02]. _{∞} [McE04]. B(t) [MN77a]. H_{∞} [Bar94]. H_{∞} [FD87]. H_{∈}/ℓ_{∞} [SKD99]. C^{1} [CB15, Hay19]. C^{2} [Zei01]. C^{2}, α [Nov04]. C^{∞} [Sus87c, GB84b]. C_{0} [JZ04, JZ09a]. C_{1} [JL98]. E^{f} [DQS16]. D [KK79]. D_{s} [QN82]. ẋ = [A + B(t)]x [MN77a]. ℓ^{1} [Sta93]. ℓ^{2} [FW92]. ℓ_{1} [HBS06]. ℓ_{1}/H_{2} [SDV97]. ℓ_{2} [HSP16]. ℓ_{2}(N_{0}) [Jac04]. ℓ_{2}(Z) [Jac04]. ϵ [Cha14, NC13]. η [CYL97a]. G [LZ21, PW23, PZLM24a, PZLM24b, YCRZ21, GHZ15]. Γ [BR01]. H [BCG84, GR02, GM11, PLSB14]. H^{1} [IK02b]. H^{2} [CVKB13, MRP99a, MRP99b, MT09b]. H^{2}/H^{∞} [DA02]. H^{∞} [BJ02, DRT11, FTZ87, FGT95, HM93, HW03, HP98, IM98, LH87, LB89, MP14, MRP99a, MRP99b, ÖT90, RNK91, Smi90, XB99]. H_{2} [Fer97a, FO98, GBGdO00, RTF21, TS95, dOG05]. H_{2}/H_{∞} [ANR08]. H_{∞} [Akç01, Bar95a, Bar95b, CF06, CdSBT09, EÖSÖ98, Gah94, GL97a, GBGdO00, Hew93, JB96, KNP91, NBF16, Ngu18, Pee20, Sch92a, Sch92b, SR97, ST90, Sto91b, Sto92, Tad89, Tad93, THA02, TF08, Van94, WFZ20, XC03, dOG05, FA13, ZC06, Yan95]. hp [WW16]. J [CF06, GGLD90, Gre92]. k [GBS22, LW07]. L^{0} [CW20]. L^{0}(Ω) [Wac19]. L^{1} [RW78]. L^{1}(R^{d}) [Bar23b]. L^{1} [BBB24]. L^{2} [Dun96, HH17, KV07a, LM13b]. L^{∞} [ARS09a, BDR94, Kat19, KN13, Kha94, MR05, Wan08, Dun95]. L^{p} [GLS05, Mos82a, Mos82b, Wac07, YC06]. L^{p}(Ω) [IK14]. L_{1} [CC13, Lin07]. L_{2} [Vid77, ZLD12]. L_{2}(0, T; L_{2}(Γ)) [LT86]. L_{∞} [ZLD12]. l_{∞}/H_{∞} [Szn95]. L_{∞}/L_{2} [HIK18]. L_{p} [BTBT88, CHL15, Cou00, Mun07, Ols76a, Pha00, NVD18]. λ [AK02b, IM08]. LC^{1} [Che96]. ≤ 1 [Lom21]. log x [CL87a]. LQ [NYZZ17, van95]. LQR [YZ97]. M^{2} [ZFKZ21]. L^{2} [Dun98]. R^{3} [Bon85, Sch88b]. R^{d} [ABBK20, AB20, Cir14, GM91b, GM93]. R^{n} [BBV11]. R^{N} [BBC14, de 98, CHS87, CC99, Fer78]. H(curl) [You13]. R^{d} [FH94]. H^{∞} [LH88, LAKG92]. H_{∈} [Red21, FCG18, ZHB21, ZGS99, dVC22, DDG12]. H_{∈}/H_{∞} [GPS95]. H_{∈}/H_{∞} [CZ01, Lei01]. H_{∞} [FCG18, FCC20, GZSAB13, GGLD90, Gre92, KS00, LMPR09, NR97, Rot96, SS99a, Sor96, TO15, ZHB21, BBT14, DDG12]. KL [KT05]. L^{∈} [TD94]. L_{∈} [SK94]. S [EMH08]. L^{1} [CCC16]. Sl(2, C) [ARS14]. Sl(2, R) ⊗ R^{2} [dRSV09]. H_{∞} [Ste15]. μ [ALY13, KHP06]. N [BP14a, BBGBO14, CFR13, FCGIP06, FR20, GX19, LAM19, LAM21, MD15, Uch78, Var76, CMK15, EMP^{+}16, Ete24, GBB18, pJY20, Kru00a, Kru00b, MP21, RW01, Top79, Vie00, VJ77, Yin99]. n + m [AA17]. n − 2 [pJY20]. ∇u [FT12, GT22a]. ν [HG02]. O(N^{3}) [Gal84]. P [SW11, CDCVV21, CKL16, Day09, MNC14, ZWYJ14]. Φ [DK78]. p ∈ [0, 1) [IK14]. Q [ABB02, DMZ22, LHH23, LC05]. q(x) [AS95]. QP [PTH88]. QQ_{c} + RP_{c} = Φ [Emr80b]. R^{3} [AS03]. R^{d} [DF98c]. S [FR82b, PDB23, ZZS14, ZzJh10, Cha11]. S^{2} [MSBC08]. S^{3} [BR08]. SL(2) [BR08]. SO(3) [BC05b, BR08, ZVBR14]. SO(n) [SL13]. t [ABP14]. θ [Fon89]. × [Cha91]. u [FT12, GT22a]. u_{t} = Δu − qu [AS95]. ε [AX05, CL19a, NA02]. W^{1,1} [BBV10]. x [FT12, GT22a]. XR + QY = Φ [ES81b]. Z^{2} [Lou94].

-Approximate [LW07]. -Block [JL93].

-Body [MD15].

-Brownian-Motion-Driven

[PZLM24a, PZLM24b]. -Class [Fon89]. -Control [Sch92a]. -Controllability [BCG84]. -Convergence [BR01]. -Convexity [DK78]. -Cost [IK02b]. -D [BP97, BE13, Dág06, FRZ93, Gue14, Hay19, KK17b, LY23, MP21, VZ09, XZ20]. -Determinacy [MSH91]. -Dimensional [Gag16, HH19, HVDK19, Kru00a, Kru00b, Sac19, Wan02, Yin99, BBGBO14, FCGIP06]. -Disk [LB89]. -Disturbance [HH17]. -Elliptic [LWY21, You13]. -Energy [Cha11]. -Equations [AK02b]. -Equilibria [NA02]. -Error [ARS09a]. -Estimates [MR05]. -Exact [Kha94]. -Expectation [LZ21]. -expectations [DQS16]. -Finite [WW16]. -Firm [CFR13]. -Free [PTH88]. -Gains [HSP16]. -Games [FR82b]. -gap [HG02]. -Groups [JZ09a]. -Hamiltonians [Zei01]. -Hedging [Pha00]. -Induced [SK94]. -Infinity [GM11]. -Instability [Vid77]. -Invariance [DCBT88, ISW00]. -Invariant [DM89, EH80, Hir81, Emr80a, ZVBR14]. -Laplace [CKL16]. -Laplacian [CDCVV21]. -Learning [ABB02, DMZ22, LHH23, LC05]. -Lemma [PDB23, ZZS14, ZzJh10]. -Level [EMP^{+}16, LAM19, LAM21]. -Local [BBB24]. -Lossless [CH05, Gre92]. -Mean-Variance [GHZ15]. -Minimization [CCC16]. -Monotone [Day09]. -Multipliers [RW78]. -Nash [NC13]. -Norm [Set93, CB15, Lin07, Mun07]. -norms [Red21]. -Null [Wan08]. -Optimal [AX05, CL19a, Cha14, HBS06, XB99, Pee20, FTZ87, GLS05, LH87, LH88, dVC22]. -Optimality [QN82]. -Optimization [JL98, Sch92b]. -Person [Top79, Uch78, BP14a]. -Player [Ete24, GBB18, Vie00, Var76]. -Positive [GBS22]. -Problem [Bar94, Bar95b, van95]. -Procedure [EMH08]. -Regularization [CC13]. -Regularized [ZH05]. -SDEs [YCRZ21]. -Semigroups [JZ04]. -space [CMK15]. -Spectral [CF06, GGLD90]. -Splines [SW11]. -Stability [KK79, Ngu18, NBF16, KT05]. -Stabilization [CHL15, YC06]. -State [FR20]. -Step [BBG12]. -Stochastic [PZLM24a, PZLM24b]. -Stopping [PW23]. -Synthesis [ALY13]. -Systems [FW92]. -theory [Wac07]. -Tracking [IM08]. -Values [KHP06].

0-1 [HJM07].

1 [IM05a, Kou24, MDW23, MMR17a, SGM19a, VGBS13a, Zas95a]. 1-D [CGW10, Mic09, SCK10, TL06]. 1-parameter [JR06a]. 14 [Kum88]. 15 [Cou83]. 16 [FR82a]. 18 [SC81]. 19 [GRC82]. 1D [WF24].

2 [IM05b, MMR17b, SGM19b, VGBS13b, Zas95b]. 20 [Col85, Mos82a]. 22 [Pen88]. 23 [Ber88, Cur85a]. 26 [AW90, Shi90]. 2D [Mar20]. 2D-Channel [Mar20].

3-Dimensional [MN08]. 30 [Flå95, ZZ94]. 32 [GK95].

56 [Cou83]. 58 [FR82a, Kum88].

82b [SC81]. 83a [Mos82a]. 83d [GRC82]. 84h [Col85]. 85i [Pen88]. 86a [Cur85a]. 86g [Ber88]. 89d [Shi90]. 89e [AW90].

92m [ZZ94]. 93c [Flå95]. 95b [GK95].

A. [JLZ98]. Abelian [Wee07]. Ability [ZY24]. Abnormal [Ava19, BM18a, BP02, Mon94]. Abnormality [FM22]. Abscissa [BO01, HHLP16]. Absence [Bal94, CMR19]. Absolute [Alt09, AR77, BP20b, MTS20, UP99, ZP03a, ZP05, vRM98]. Absorbing [FP19, HW05, KK16, PZ24, Ros00a, SLG84, Wat24]. Absorption [CC94, DEO24, GHS22, MNORP21]. Abstract [ADS24, Bal91, Ber98b, Cur76, CP77, FR00, IK07b, LR21, LLR01, Mah03, MT78, MS97a, TWW16, VP87, WZ17b, XLL08, Zho83, Zho84, Zol78]. Abstractions [BD19b]. Abstractly [Zit14]. Acausal [Kre87, NLW91]. Accelerated [SE21]. Acceleration [KN13, PJ92]. Accessibility [AS93, Gra08, Son88]. Accommodation [FLJ13]. Accumulation [AFK10]. Accuracy [ACV22, BKK02, vH10]. Accurate [NSC19]. Achieve [vH10]. Achieving [CC01, MYP14]. Acoustic [AN15a, AN15b, AY23a, AHR23, BP97, CCFV20, FB98, GLX18, GMN22, LBMH^{+}16, MNORP21]. Acoustics [AHH17, ALR99, HRPT21]. Across [WL95]. Acting [BLP14, GWY14]. Action [ALNP24, AS91, BS24, BF15, BP98b, BLS12, CP77, DM17, LC80, MD15, Yu20b, Yus97, dWvS00]. Active [BGR04, BCG^{+}17b, GV05, GZ13, GZ14b, HT10, IK04, IK07a, RV93]. Active-Set-Newton [HT10]. Activity [BØ94]. Actor [KB99, KT03d]. Actuated [MÖ14]. Actuation [BP14b, VS14]. Actuator [EM19, PTZ17, Tuc96, WQS^{+}24].

Actuators [AGN^{+}16, HH05]. Acyclic [KCP16, MYAS09]. Adaptation [CL19a, DBC12, MP11a]. Adapted [LS91a, LLC11]. Adaptive [AA17, Bas91, BC23a, BKR00, Ber98a, BP08a, BC21, BCC^{+}19b, BDRR98, BV82, BBE99, CZ95a, CK98, Che84, CG87a, CG87b, CG90, DK05, DCI07, DMPD94, DPDS98, DMPD00, EGR^{+}22, Ega80, EP93, FIR06, FB12, FK21, GAC24, GGW91, GRC81, GHX86, GL97c, GK11, Guo90, Guo94a, GG13, GZ19, GZ23, HN00, Heu94, IRS02, ITT04, JALG18, JGC^{+}23, KMP13, KKA23, Kum82, Kum83, Kum85, KH81a, LY91a, LY91b, LLMT02, LG24, LI94b, LR00b, LT03, Ma09, MZ18, MS92, MZ15, MV07a, MSVA09, Mor79, NK94, PP00, PO00, PLK89, PC00, PMI06, RMB12, RM92, RY95, Ris86, Rus86b, Rya98, Sch79, Tow96, VW08, WS13a, WLZ22, WL21, WTJY21, WHB92, XG18, XZ15, YZSS23, YLYC15, YHJ^{+}18, YC24, ZC96, ZK20, dBMS10, GRC82]. Adaptivity [HH08, HH10]. Addendum [AW90, Col85, Mar77, Mos82a, OZ03, SC81, Tri80, Ber88]. Addition [Nic96, Nic97]. Additional [Per12, RR94]. Additive [AW83, BHS19, CMR85, DSI22, FS23, FC05, IS12, JL16, NPY20, ZLZ18]. ADI [ORW13]. Adiabatic [ABS18, CM17]. Adjoint [BIP88, FB98, GKLW14, Hau81b, Hau81a, LR10, Ulb02, GX04]. Adjoint-Based [GKLW14]. Adjustment [KH81a, LZ13]. Adjustments [BBD18]. Admissibility [Bon89, Che16, CyY21, CFZ22, CZZ24, HK07, JPP14, Jun00, Wei89]. Admissible [BC11, Gra07, HR83a, HR83b, JZ99, Nar82]. Admission [AK99b]. Advances [HLDM20]. Advection [Liu08]. Adverse [AEÉ^{+}20]. Advertising [BCC^{+}19a]. Aeroacoustic [CRV09]. Aerodynamic [WCCL04]. Affects [TWB23]. Affine [AZ07, ABK21, BB18, Baş84, BP07b, BCS20, BG14, BL05, Cal09, CJT08, CW24b, DEJ23, FP96, Flo97, GP94, GEF17, GM09, JP15, JS84, LS99b, Luk85, Mal23, MTS20, MSBC08, NR17, TC14, UU00, VW02, YD08, YXC21, dBMS10, dRSV09, JR06a]. Affine-Scaling [UU00]. Against [BD01b, SSY24, BFPB13, CWC^{+}22]. Age [BG20, BG21]. Age-Structured [BG20, BG21]. Agent [CWC^{+}22, HPT07, LC19, LMA07a, LMA07b, LRTY22, LL21b, RJME09, Rud78, SC16a, YGC18, ZLZ18, KM18]. Agents [BLP14, CBB19, EP19, Ete19, HPS03, HM09, Kim22, NC13, SNB09]. Aggregation [BCY15]. Aggregative [BGP23, ZHH17]. Agreement [LFM07]. AIMD [RS07b]. Air [TW03]. Algebra [AD08, GEF17, GK85, pJY20, MS02]. Algebraic [AL01, BR17a, BR01, Bro22, Brü13, BC02b, CDJS23, CR21, CLT06, CF06, DAR15, DGJ93, Fla87, GL90, GL92, Gho88, GLR86, Gri93, GM11, HH78, Hei16, Hel76, Her80, IY12, JS90, KH79a, KH79b, KS90a, KH90, Lev85, tLR23, LMP95, Mar84, MG20, MG21, MH78, Oht08, Qua03a, Qua03b, RR92, RR12, RV19, SS92, Sha83a, Sha83b, SGM19a, TR01, TO15, WS92, Wim92, Wim94, Wim06, WM10, ZQ18, vN04, JS91]. Algebras [BP08b, CYL96, CYL97b, CY94, CS11, Cur11, Cur13, DTWY91, Her78, SY17, TWY90, Tan00, WY06, Cla86]. algèbres [Cla86]. Algebroids [GJ11]. Algorithm [AH21a, Alm01, AB12, AB17b, BI91, Bar17, BGP20a, BET07, BF22, Bon86, BH07, BCS21c, BC20, BCKM24, BQ99, Cha77, CHL82, Che84, CR93, CL06b, CM18, DN23, Don24, DMZ22, Dye77, EK85, FFR14, FR82b, FY76, FM00, FP99, FO98, Gal84, GCB06, GK87a, Gül91, Ha90, HK78, HLG93, HL88, HU19, Heu94, HP14, Jaf17, JL93, JH00, KSY23, KSS20, Kim22, LCF18, Lei01, LY12, LL13, Liu97b, LLT18, LGZ13, LT91, LB98, Luq84, MD89, Mey77a, NF13, PY16, PTH88, PLSB14, PM85, PQ98, Pul93, PV98, Roc76, Set93, SWW22, SZZ97, Tse91, Tse93, VGBS13a, VGBS13b, WZY09, WZHZ19, WCLC05, Yan96, YHH16, ZC96, ZRS97, CL78, Qua80]. algorithme [CL78, Qua80]. Algorithmic [RN13]. Algorithms [ABB01, ABB02, ABB15, AMP00, BBC03, Ber98a, BK02, BCS24, BK00, BLZ94, CB09, CLRJ13, Che98, CZA08, CHJ15, CC90b, DJ95, DHV98, Dun79, EK85, FK91, Fre99, GM91b, GM93, GP79, Guo94b, Han77, HW05, HM09, Izm96, KM07a, KB99, KT03d, KS15, KK76, KH79c, KY87b, LY91a, LN82, LR79b, Mal23, MTS15, MP11b, Pel00, PMW83, Sac85, SE21, SKXL17, Tam80, TS24, TDNT18, Tod83, TBT90, UUH99, WD07, WP76, YMH94, ZYHX18, dJ78]. Allen [CS15, FSH16]. Allocation [Cal09, RN13]. Allocations [Pap86a, SBT19]. Allowing [LW15c]. Almost [AJ16, BTY11, BC05a, Bou95, CWZZ17, CW09, CMT03, DI87, DHX08, DPDS98, GJ11, GMY16, JLZ98, Mao90, Mar99, Obe09, Pel00, Tre86, ZKZB20, ZWYJ14, vH06]. Almost-Riemannian [AJ16]. Along [AB09, AG97b, BS90, BS93, PV15]. Alphabet [XUP08]. Also [PV14]. Alternating [CL19a, PT09b]. Alternative [SB78]. Ambiguity [BZ15a, CS12b, HJ16, PW23, TCC15b, TCC19]. Ambiguous [FPW16]. Ambush [AA86, AFGT13]. American [Ach05, ASR^{+}13, Bay09b, BL02, BL09, Ham06, HT10]. Among [Rus76, SNB09, Van92]. Ampère [Sto04]. Amplitude [CRS21]. Analogue [MS07a, ZCC07]. Analyse [Pen89]. Analysis [AHJM10, ABL13, AMP00, AGL08, AE17, AFV17, AJ12, AAA12, AY23a, AGJK12, ANS14, AOS16, AHS13, ANM22, AH21b, Aus99, AHK24, BM06, BS16, BC18a, BKS10, BCG86, BO16, BKS19, BKP18, ByS17, Bro22, BCS21c, BO01, BL03, BEL02, CG00b, CJK13, CFT12, CNS10, CD13b, Cas12, CT14, CRT15, CC16, CWW18, CK23a, CK23b, CCJ23, CZA08, CFNG10a, CFNG10b, CR21, CLS21, CY24, CR93, Chr19, CL87b, CE06, CSA^{+}13, DF96, DF98a, DL09, DZ88, Doa21, DWZ^{+}16, EMH08, Efi12, Ega80, Ete19, FZ05, FM12, FB08, FL81, Fer97a, FPS17, FK21, FS10, FO98, GCB06, Gib81, GM05, GHZ14, GZ14a, Gri93, Grü09, GPSW10, GSS19, GMS22, GW19, HZ07, Hag02, HKB04, HYZ24, HM02, IST02, Izm96, Jou00a, Jou00b, KRS14, KMR10, KH81a, KY92, KM96, LS91b]. Analysis [LI00, LS05b, LHH23, LZ16a, LW15b, LL18b, LWLZ20, LZC21b, LB00, LH87, Liu95a, LZ21, LZSC24, LGZ13, LY01, LY02, Luq84, MG20, MG21, MRVM00, MRV13, MMÖ18, MADF98, McE04, MV11, Mey77d, MVBC18, MM11, MS96, NVD18, NT17, NY16, NBF16, Ngu16, Ngu18, OW96, OM14, PV13, PRS08, PV98, Qu14, Qua03a, Qua03b, RS07a, RSE20, RX21, Roc96, Ros00a, Sak02, Sha88a, SB92, Sha94, SL22, SO06a, Sok87, Sok88, SS19, Spi80, SZS21, TLC99, Tan88, TYW11, VW95, VW02, Wac07, WK13, WP18, WZ98, XW12, XG18, XZ15, YÖ08, YK91, YXJ11, Yu13, YC24, YS17, ZW21, ZH18, Zor17, dBLR14, CL06a, Pen89]. Analysis-Based [CR93]. Analytic [AS93, All99, BR87b, BD77, CQ05, CSES21, GB86, GL05, HJ15, Hua88, JS81, KS13, LH21, Las95, Mos80, PZ11, RSS22, Sus87b, Sus87d, TG12]. Analytical [Bay09a, FLUU17, QS13, RSMG17, WM12]. Analyticity [LT83a, Sei82]. Angle [ARS09b, BK21]. Angular [Eno94b, LAM19]. Anisotropic [MP11a, SL18]. Annealing [Cat98, CC88, CC94, CK89, DM99, FK91, FGQ97, GM93, Maz94, Roy89, Tro96]. Anodes [HS96]. Antagonisms [MMH19]. Anthropomorphic [BDGR12]. Anti [XZ20]. Anti-dampings [XZ20]. Antiangiogenic [LS07]. Anticipating [MA19]. Anticommutation [EMP^{+}16].

Antidamping [SCK10]. Antiderivatives [BLW07]. Antidisturbance [ZY24]. Antiplane [KL16]. Any [BM00b, BFZ10]. Anytime [CFZ10]. Aperiodic [LSL^{+}18]. Appearing [CT20, Mau77]. Application [ABB02, Alm08, AHS13, ARS09a, ABK19, BT11a, Bar77, BBGBO14, BCC^{+}19a, Ber77, BR17b, BASG21, BM18b, Car97a, Car97b, CM06, Chr77, DKK^{+}19, Emr80b, FP04, Fon89, GT17, GS19, GAM93, GM05, GMR22, GGMZRB20, GHS92, GX04, GY12, GY13, Hag05, HZ10a, HY22, Ho86, HZ05b, Hui92, HBS06, Ish88, JZLW18, KT03a, KO18, Kre77b, LXY16, LS09, LM16, LY19b, MMN09, Nag95, NNTD14, NS20, Nic00, Ouv78a, ÖB19, Pic93, RH06, SS09c, SWW22, Ste23, Ste92, Tam80, TY23, Was79, WR14, Wee18, XG03, ZES12, ZZS14, aJM18, dNCP20, Iof83]. Applications [AØ23, AK02a, ASTT09, AB10a, AAFC19, Art89a, BK82, BR10, BK80, BZ96, BHT22, BHL10, BR93, CL02, CFG21, Cer01b, CKT24, CSSZ20, CSES21, CLSL20, CW23c, CV89a, Cor94, DL06, Dat80, DEO24, DFŚ24, DM99, DK86, EGR^{+}22, FM12, FCG06, FS16, FB12, Gai04, GSW89, GK85, GL11, HPW10, HO01, HSBTZ13, Hei16, HL16, HS87, HJF10, HWX09, ITT04, JPP14, JLR09, Kam76, KG22, KJ07, KT03c, KK06, Lam87, Lam92, LM15, LL07, Li95, Li07, LFL23, LH24, Lü13, MM00, MH78, MP76a, MNPR20, McE95, MS97b, Mor79, MV91, NRB20, PS10a, PW99, PZLM24a, PZLM24b, Rap22, RKL16, Rya98, Sav05, SW78, SL18, SZSC21, SK94, Sle78, Son23, SV79, SCCL18, Sub81, Tse91, Tsi91, TW00, WX13, Wan22, WP98, WZ77, WA84, XLL08]. Applications [Yam89, YMH94, YZ16, YZSS23, YZ22, Yük17, ZC24, dBLR14, vWCET23]. Applied [BL03, Gon94, HLM^{+}18, Ho93, HV00, Oss01, Qu14, Wac19]. Appreciation [Ano95]. Approach [Aey85, AMP00, AEÉ^{+}20, AFV17, ATW93, AS21, AGJ^{+}13, AA92, Arm86, AG97b, AS11, Aub90a, AVCCD22, BD01a, Bal84, BH96, BSB^{+}97, Bar88, BBC14, BdP21, BS24, BTBT88, Ben96, BB09, BLS05a, BB06, BET07, Ber84, BP17, BFP08, BS93, BCG^{+}17b, Bli04, BPZ16, BGP23, BDT20, BKM14, BGL98, BEL02, CK98, CFT12, CMA08a, CLAL13, CLS12, CGS00, CDR08, CCHH19, CFNG10a, CFNG10b, CDM21b, CLMSR00, CLT24, CP82, CdVF05, CdSBT09, CK17b, Del86, DZ19, DR87, DG20, DK89a, EC22, El 17, Enq04, EB24, Fab08, FP02, FZ05, FHH22, FZ00, FFP^{+}09, Fio20, FV89, Flo99, FGS93, FGT95, FV95, FC05, Fra80, FB12, FK21, FT22, GR06, GQ09, GKY21, GMN22, GNP18, GVL12, GPS95, Gha11, Gho88, GB93, GFF17, GL97b, GO04, Gon94]. Approach [Gor19, GHS22, GZ14a, GGLD90, GLNT23, Gue03, Guo01, Gup20, Had06, HFY14, HJ15, HH78, HSZ15, Hel93, Hel76, Hen80, HRMK17, Hoe12, Hoe14, HR98, HMR15, Ich79, IS12, IM05a, IM05b, IK08, JS90, JQ97, KKP85, KT00, Kas09, Kas13, Kha03, KN03, KS98a, KZ00b, KK02, Kot94, Kum08, KS84b, LH21, LN82, Laz88, LD19, LNW23, LL20, LAKG92, LLC18, LMR97, LL21b, MP14, MM04, MPZ03, MRR16, MRRR19, MGMBN^{+}19, Mat98, MO02, MMN17, ML12, MR97, Mil17, MR86, MI15, MN03, MM11, MPC91, Mor98, MA01, MS04b, MT22, MLYF02, Mur89, NVD18, NS00, NT95, NP22, Ngu18, ÖT90, PS99, PZ11, PZ22, Pon09, PA93, PC00, PMI06, Pue09, Qua03a, Qua03b, RS07a, RSMG17, RBS21, RR12, RNK91, Reb95, RK89, Rot96, SS09b]. Approach [Sch83, SP02, SW11, SI00, SD13, SDSÁR00, SD15, ST99, ST09b, Ste15, Ste23, Stu15, Szn93, Szn95, Tak97, Tie14, Tse90, Und76, Val01, VB23, VC03, VG95, VK81, WOH18, WZ12b, Wee07, WKS86, WM10, WORO00, Xia19, Yam88, Yam89, YGC18, Yu15b, Yu20b, Zam98, ZP03a, Zha95, ZLD12, ZFKZ21, Zwa88a, dFN23, Dai88, Gho86b, JS91]. Approachability [SQS09]. Approaches [CZL21, Mey77d, Pan13]. Approximability [BH23]. Approximants [CL94, KNW95]. Approximate [Alm08, AM23, Ave16, Ave19, BLPP24, BH96, CM17, CDZ94, Den97, DL19, DMR19, El 17, GP98, GN03, HSA17, HW99, HU84, JL98, JT11, Kar11, KLR09, Kob77, LL00a, LR00a, LW07, Lev85, LR16, LR23a, LR23b, LL15, LLM15, Mah03, Man81, MS04a, Mun14, Mun07, Ner21, NLPT09, PP02, PT09b, RKL18, WYY21, War19, Zho83, dOP03, de 98]. Approximated [MR05]. Approximately [KR89, NS91]. Approximating [CSB15, HU82, Kum08, Mey77b, Mil88]. Approximation [ABB02, AMP05, APA08, Bag84, BFR21, BK82, BW89, BRR90, BF90a, BBC18, BGR90, Bis78a, BM00c, BH17, Bra98, BMN19, BvdMR05, BK02, BK99, BK00, BLS05b, CFN06, CDDP11, CR06, CMR07, CS10a, CGM11, CCK12, CKL16, CWW18, Che98, Chr83, CL94, CMT11, CL06b, Com00, CR17b, DGJ93, DGH09, DDS15, DF20, DR97, Doa21, DF98b, DA17, DJ98, Fer05, FMSV16, FO98, GJP88, GMP00, GR88a, GA91, GRT92, GCP88, GHZ14, GAHL09, GH96, GW96b, HLG93, HK21, Her78, HR21, HLZ15, HLZ20, IT87, Ito90, Ito96, IM98, IR00, Kam10, KT00, KS87a, KS90a, KLM20c, Kno82, KS15, Kun82, KW87b, KS13, KK76, KH79c, KS85c, KY87b, KY93, KVA96, Las84, LLMT02, LZ17, LTT24, LRT96, LZ94, Liu97b, LG20, MXZ19, Man21, MRV13, MD13, MDW23, MML21]. Approximation [MS98, Mik12b, MU22, MPS99, OCI00, Pep16, PTZ21, PJ92, Pro90, QC95, Red18b, Ros84b, SI03, SD15, SYZ13, Ste22, TLC99, Tib11, WZY09, Wat24, Whi80b, WŽ07, XB99, Yak93, Yu14, ZM11, ZW21, vTY99, DK89b, GR85a, GR85b]. Approximations [AFV17, AM89, BM19, BB78, Bar80c, Ben96, BHS05, BS90, BT98, Bor98, Bor00b, BR15b, CC99, Cro84, Dai12, Don96, DPR13, DK85, GM91a, Gib83, Gom89, GWY14, GP79, GKLM92, GW19, HI84, Her86, Her96, HHLP16, HY97a, IT86, JX22, KR80, KS89a, Kus79, KH81b, KS84b, Kus89, Kus02, Kus04, Kus07, Kus14, Las95, LS98, Lep90, Lóp12, Lóp13, MGHT19, MP99, MS92, MT94, MG20, MG21, MR78, MR79, Men89, Mor95, MA03, NP96, PRG16, PSV18, PP02, Ros84a, RS13, Sal85, SS09c, SP96, Sha18, SZ04, SK23, SS18, TD21, TA96, TPA97, VGBS13a, VGBS13b, VZ97b, Wri95, dFN23]. ARAT [DPR22, JN24]. Arbitrarily [CC01, PS04]. Arbitrary [BHL10, DEJ23, GW76, KvdSD21, LNT02, MCS13, MVZ^{+}09, OC23]. Arc [BHT87]. Arches [IST02]. Arcs [HW19]. Area [ASR^{+}13, Fer78]. Arenas [ZSS^{+}20]. Arguments [MR78, RM92]. Arise [MS98]. Arising [Ach05, Bag84, BMS15, CDDP11, Fag08, Fla84, Fla87, FK76, Gah94, Gen13, HLT20, IK94, JMP21, KSY23, KS98a, LB00, MZ97, NM10, RZ20, SP02, Sha85, Tes92]. Arm [BCW21]. Arma [WDS79, CZ10, Enq04, FZ23]. ARMA/ARMAX [CZ10]. ARMAX [Ber95, CG87b, CZ10, SK86]. Armed [Agr95]. Arrow [GS93, Jah88, Pet90]. Artificial [KKLO21]. Artstein [Tsi95]. ARX [GL22a]. Ascent [Tse90]. Asian [CK17b]. Aspects [CM10a, Sch89, Sha10, SS92, Tow96, dJ78]. Aspiration [CAS13]. Asplund [Sag15]. Assessment [OVS08]. Asset [Cal09, DL13a, KR04]. Assets [KMZ98]. Assignability [ABCD22, BBC^{+}19, Ho86, Reb89]. Assignment [Ari98, CM91, CW81, FR78a, GB93, HRW06, LT83b, LDZ10, MW09, NS14, RRW96, Ros92b, SRS15, Wan94, WCCL04, XS96, FR82a]. Assimilation [BB97, CB11, Pue09]. Assisted [CMT03]. Associated [ARS09b, BD83, BT03a, DMS02, Du15, FP11b, FL80, FL82, Gor15, HH76, HYZ11, HT22, Ich11, KMF19, Kun82, LZM05, LM00, Rus86a, TZK11, Teb22, YT13, Zha09, ZDM20, ZI00, ZTC16]. Assumption [BFZ10, DM02a, Stu15]. Assumptions [CCJ23, DL06, Hin14, Sch92b, VL78b, War76b, War83, Ye00b, Zas00]. Asymmetric [Alp10b, BFR21, Car07, GNLB14, Jac22]. Asymptotic [ACS00a, ABHY23, AG97a, Aus99, BM11, BTY11, BCL04, BA15, Bon89, BS11, BFP11, CC02, CK02, CLS12, CNS10, CG90, CLSL20, CGC17, CC94, CSB88, CS89, Cor99, Cza04b, DMK90, DMS02, DHC19, Dun81, DK85, FL81, FCJ14, FP99, FH17a, GGM01, Gla05, GS18a, Gor15, Grü98, GI02, HZ07, HMV05, HHSS23, HS09, HM09, IK02a, IK09, KT03a, KPCW22, Kra93, KH81a, KH81b, KY87b, KY93, LR00a, LCF18, Liu13, Luq84, MRS04, MGMBN^{+}19, Maz94, MC91, Mey09, MM80a, MN77b, MS04b, MR18b, OW93, OP96, PSJ18, Pel00, Pen89, PA99, Pfl86, PS04, Pri05, Rod82, Ros00a, SAM03, Sîr15, SZ04, Son83, TLC99, Tar85, WC89, WC91, WW23, ZC91, ZY07]. Asymptotical [DJ95]. Asymptotically [Bag81, CG87a, CR11b, GL97c, Her91, IP06b, KMP13, LW87, LY91a, MSVA09]. Asymptotics [ACKM22, GO11, PS17b, RBA22, ST13, VJBK24]. asymptotique [Pen89]. Asynchronous [BE87, Bor98, Bor00b, CMA08b, LMA07b, TBT90, XW12]. Atom [MR09, SSL11]. Atomless [FP19]. Attached [AE18]. Attack [ZM14]. Attack-Resilient [ZM14]. Attacks [CWC^{+}22]. Attainability [DDK12, KP97, Reb77]. Attainable [AM98, AC05, ADM17, Her78, WXZ15, WR14]. Attenuation [LD06, QLD01]. Attitude [BP14b, Bil19, ZTC16]. Attraction [BM00b, RS10, XWZ^{+}21, ZZCaL19]. Attractivity [SL22]. Attractors [LW12]. Augmented [BK97b, CZ99, DG79, Iof79b, IK90, IK96b, KSW18, NNTD14, Vol00, VW02]. Automata [Cha14, CIP96, TW94a]. Automatic [VS80]. Automaton [BP02]. Autonomous [BB00, Cam76, CO03b, Mai03, PP13, Val01, BGST23]. Autoregressive [BC00, KS90b, PO00]. Auxiliary [CM90]. Average [ABB01, ABFG^{+}93, ABBK20, Ber98c, BK24b, BHT10, Bor89, Bor00a, BGS19, CLLP13, CC96, CCHH19, CW23b, CW24a, CP13a, CD10, FZ09, FGL08, GR06, GC05, GHS12, HLL94, HLL96, HLVAC99, HLL00, HLGHLM03, JLL20, Küe07, Kus78a, Kus89, NC16, Ols17, PS17a, PRG16, Rob81, RFF15, Ste17, Ste22, TCC19, VA03, WC19, Yu20a, Yu20b, ZGS99]. Average-Delay [JLL20]. Average-Preserving [BHT10]. Average-Reward [CP13a]. Averaged [AMV20, BP13b, GW96b, Pel00]. Averaging [BB78, Bar93, Bul02, Cha87b, DS99b, GN02, HSBTZ13, KH81a, KY93, Liu97c, LK10, MP11b, OT09, PA99, PJ92, TLC99]. Averse [AKP21, BB23, ÇR14, CS12b, CHP^{+}11]. Aversion [AD18, BZ15a, Xia11]. Aviation [LMJ09]. Avoidance [GC22, Gor19]. Avoiding [BGSS21]. Aware [HJ15, ZM13]. Axes [Bil19]. Axis [BGMR94, BFP11, CFNG10b, Far08, Far13, Lou15, OS12b, OS14, OS19, Sch92a]. Axisymmetric [BP13a].

Back [BASG21, VS80, ÇD21, MSZ18].

Backstepping

[BHM99, CDM21b, CVKB13, MB04, MNB11, ÖB19, PB99a, WCS12, Xia19, ZD24]. Backstepping-like [ZD24]. Backward [BC16, BHL12, CE11, CF13, FHH22, FT04, GCL15, GHQ15, GT05, GT14, HPW10, HQZ16, HP09b, HJX18, HJX22, HWX09, JL23, KZ00b, KT03c, LZ18a, LY18, LX23, Lia15, LZ01, MT03, ØS10a, PW99, RV87, ST14, Tan03, TZ09, Tse00, WWX13, WZ17c, Wan22, WY21, YT13, YZ16, Yon10, ZDM20, ZS23]. Bailout [NPY20]. Baire [Zag00]. Balance [BÇS07, CK89, LY19a, SU21, ZPQ20]. Balanced [FS10, GO14, HO97, HSBTZ13, Hel93, LF87, MNC16, OMS90, Obe91, OW93, Red18a, RV09]. Balancing [OC04, Opm07, ZHC06]. Ball [AB18]. Banach [Cur13, Tri80, Ahm83, Ahm03, Ahm19, BR17b, BP08b, BC96b, BC96c, BC96d, BKS19, Bor78, BZ96, BS13b, BS00, CFG21, CLMSR00, CS11, Cur11, DM83, DPT04, Den98, FHT09, GST20, Gfr06, HK07, HS11, HP94, KSW18, KS18, Lev99, Mas08, NZ03, PS86, PD92, Phe85, Poo96, Sag17, Tri76, Tri77, Tri78, vN04]. Band [EAK86, Kus79, Kus14]. Banded [HBS06]. Bandit [Agr95, BCW21, MR90]. Bang [ASZ02, Ben76, BPV19a, Bre85, Cas12, CWW17, CWW18, CM95, CW24b, EL20, Fel03, LT13, MO04, MS97a, Nar84, Ols02, PS11, PS13, Pog17, PS18, Sar97, Sch88b, Sch80a, Shr81, Sus79, WXZ15, WY16]. Bang-Bang [BPV19a, CWW17, CW24b, EL20, Pog17, WXZ15, WY16]. Bang-Bang-Singular [PS18]. Bang-Singular-Bang [PS13]. Bank [CH98, CM07, CF14]. Barabanov [Mor24a, Pro22]. Barankin [GS93, Jah88, Pet90]. Barrier [Aus99, BB06, BT03a, Bou08, KM07a, PYM93, RLXS23]. Barriers [BB88, DL13b, SLG84]. Base [UMD88]. Based [AGJK12, BB78, BSB^{+}97, BKSW24, Bas91, Bat07, BBC18, BBT14, BB03, BP17, BCG^{+}17b, BC96b, BC96c, BK14, BH23, CLZH24, CLYL13, CZL21, CHL15, CR93, CE06, Cob12, CR13, CM18, DYZ12, FCG16, FM00, FV95, FS10, FO98, GL76, GZ12, GVL12, GZSAB13, GKLW14, GLNT23, GN03, Guo94a, GNLB14, HS15, HLT08, HLZ15, HQR24, HHT19, Jaf17, KK19b, KZ23, KK18, KK14, KGBS08, KV07b, LL13, LXY14, LW17, LMMY20, LGZ13, Mac14, MDB05, Mao90, MYAS09, MMN17, MA03, NDHL24, PYM93, Qua04, RSMG17, RTF21, SP96, Sha17, SX19, Sla03, Son23, SS18, TLC99, Wac07, WZD^{+}23, WF24, Wim06, WM12, WHGY22, XG18, YYH^{+}16, YCRZ21, YLL^{+}15, YHJ^{+}18, ZVBR14, ZHC06, ZZ13, ZG19, ZG21, Zor17, Zd83, dFN23, dMCJ22, dWvS00, vWCET23, JV06, Lei01]. Bases [Akç01, Bis83, JKZ21, NHG99, XCZ20]. Basic [AGL08, BKR00, BH00, Li94a, Per81]. Basin [RS10]. Basins [Aub01]. Basis [BVW00, CW08, Guo01, Guo02, GX04, HdHVW03, KNW95, XS96, XG03]. Batch [Bd95, FEBL13, GRR09]. Bathymetry [CRS21]. Bayes [MM00]. Bayesian [BBD18, FP11a]. be [BK03, ŁS85b]. Beads [SAMB14]. Beam [ÁRL21, Bea08, CF24, CMR17, CM98, Guo01, Guo02, GH04, GZ19, GZ23, HCG20, HT12, KR87a, KGBS08, LYLZ99, LL98, LZ18b, MM04, ÖH14, RL20, Tuc96, VS14, WC91, YW07, YZ22]. Beams [AT00a, AZ08, Ber02, CZ98, CDKP87, Con90, LLS93, Leu99, Mor98, MÖ14, NP96, Reb95, ST99, WXY05]. Beat [BV13b, NMBM98]. Becomes [PS17c]. Beginning [McS89]. Behavior [AMP10, BF76a, Cér00, CC94, FL81, FP99, GS18a, HHSS23, HMR15, HM09, IRT07, LP10, MGMBN^{+}19, McC76, MC91, Rei83, RX15, SW78, TW94a, TW94b, WC89, ZL01]. Behavioral [BFP08, Brü13, FZ00, GL97b, Hug16, IM05a, IM05b, NZ99, PS99, RW97b, SS09b, TLL16, TFT11, Val01, WORO00, Zam98]. Behaviors [BV13b, BMO15, CF18, FV98, NvdPS10, Obe09, RR12, RT04, SBBV15, Sha01, Sha14, TG13, WSR04, ZL02, Zer05]. Behaviour [Kra93]. Behaviours [FRZ93]. Beliefs [AY23b]. Bellman [BL08a, Pla21, ACL09, Aub02, ABL18, BD83, BBC14, BS13a, BBF18, BN16, CC04, DL06, Da 85, EJ89, FM00, Fra93, FL80, FL82, FMT10, Gar03, Gom20, HS12a, HS12b, HJX19, Ich11, KS99a, Kre00, LW14, LM82a, LM82b, LS85a, Lor87, Mal02, McE95, McE09, MO99, Nag96, Pen92, PZ14, Qiu18, Wat24, WWZ14, WY08]. Belonging [NY18]. Ben [HN96]. Ben-Tal [HN96]. Benchmark [BLY21]. Bending [Pan17]. Benjamin [Mic01]. Bernoulli [ASTT09, AT00a, ÁRL21, Guo02, GH04, GZ19, GZ23, Kim92, LT89, LL98, WC91]. Best [ALNP24, AK99b, BVW00, BRW05, CL94, HK94, Kle18, SMK18]. Best-Choice [HK94]. Best-Response [SMK18]. Better [Jac04]. Between [CFRV98, CF05, CIP96, DGKL10, GN18, HN96, JHMG06, Koe90, KZ00b, Pop97, Zwa88a, ASR^{+}13, BL00b, CV87, DV04, FG15, Gug03, GT08, KS84a, KS85a, LR23b, LFC21, McN84, NSW17, QW18, Ste17, TG13, Tri78, ZBA13]. Beyond [FK99, GS18a]. BFGS [Che96]. Bias [PRHL06, TD21, ZVBR14]. Biased [CK98, FB12, PC00]. BIBO [SWZ24]. Bidding [AX05, BF13]. Bifurcation [CL19b, GS19, GCSB99, Kan98a, Kan98b, Kan00, Poo96, ZR23]. Bifurcations [HKB04, JR06a, KL02]. Bilaplacian [Ant14]. Bilateral [BL04, CCT05, CCT07, IK04, RU21]. Bilevel [BM13, Ye95, Ye97]. bilinéaires [GB82a]. Bilinear [ARS09b, BMS82, BD11b, BW79, Boo82, BI08, BKP18, BGSS21, CS18, CGS00, CWW18, CT19, CZL21, Duc20, FM12, FZ19, Fla90, Fra80, Grü96, GGMZRB20, HSBTZ13, HMMS13, HJST24, IK07b, OMS90, Ouz10, PD79, RT04, Red18a, Red21, Sac97, Sle78, Tie14, Tie20, Wan98, WL17, dRSV09, GB82a]. Bimodal [dBMS10]. Binary [BP12, DV12, VP22b, Xie12]. Binomial [AZZ20, PP18]. Biochemical [Cha05]. Biological [CH12]. Biology [CG22]. Bioremediation [RRR16]. Bipartite [GH20, SZSC21]. Birkhoff [LL07]. Birth [KLM20a]. Bisection [WFH13]. Bisimulation [Gra07, LZC21b]. Bisimulations [LCW18, LZC22, PT09b].

Bistable [ALNP24]. Bit [Col12]. Bitangential [Ste13]. Black [Bur20, PDT19, FP11b]. Blackwell [GS93, HBY21, Jah88, Pet90, Yus97]. Blending [GG24]. Blind [CS12a]. Block [ALNP24, Com00, GLD^{+}91, HO97, II97, JL93, Nym94, Sta93]. Block-Balanced [HO97]. Block-Iterative [Com00]. Blowup [BL96, LW11a]. Bodies [Eno94a]. Body [CD13a, Eno94b, GJ20, GY98, MD15]. Bogdanov [GS19]. Bogolyubov [DPT04]. Bogolyubov-Type [DPT04]. Boltyanskii [Bre07]. Bolza [Cla76a, Dav07, HW19, Iof76, LR96, LR97, PP14b, You87]. Bona [Mic01]. Bond [LS91a]. Bonds [SPS04, SS06]. Book [BF13]. Books [GN18]. Boolean [LM13a, LW15b, LW17, LCW18, LZC21b, LZC22, LZH^{+}16, LSL^{+}18, WeFC22, ZZS16, ZLLC18]. Border [Gor19]. Borel [HLL94, HLL00, HLGHLM03, HBY21, Yu20a]. Borelian [Yus97]. Borkar [Yük24]. Both [DF98c, GR17b, GW18, GH04, Now89, OS02, Pau08, Ric77, TZ07]. Bottom [FLRO17]. Bouligand [RU19]. Bound [BW88, EÖSÖ98, HK94, JHMG06, Jou00a, Jou00b, KKA23, LH88, RSS22, Was79, ZK88]. Boundaries [ASR^{+}13, De 15, DFM15, DS19, FP14, KB81, Sto97]. Boundary [ASK05, AN15a, AN15b, AFT91, AT96, AX96, AB93, AB02a, AB03, AK99a, AM89, Ama04, Amm02, AC18b, AM98, AC05, AN21, AA17, ANS14, AFCLS22, Aub02, Aug19, AL00, AS95, BD01a, Bad09, BM86, Bar80b, Bar82, BT91, BF91, Bar95a, Bar12, Bar21, Bar89, BC81, BLR92, BB19, BBGBO14, BF76a, BGBPG04, BRW05, BFM23, BC02a, Bro87, Buc92, BN16, BGH94, CFZ17, CT96, CM06, Cas93, CF95, Cas97, CR06, CS10a, CZ98, CCFV20, Caz14, CRZ13, CGY03, CG10, CSFCL^{+}23, Che81b, CGY21, CM03a, CS77, CH94, CF14, CDM21b, CMR17, CDZ94, CLM24, CLT24, CFSGS15, CS15, CGS18, CBdN08, CGW10, CB15, CRV09, Cur85b, DZ76, DLP86, DGH09, Del04, DRT11, DF98c, DM17, DMPD94, DMPD98, EH10, Fab08, FG09, Fab92, Fag08, Far08, FSGHH20]. Boundary [FM94, Fen16, FH94, Fla84, Fla87, FNR19, FK76, FR78b, FH88, Fuj86, FGH98, FI98, FGH05, GT17, GLX18, GZ21, Gev81, GW76, GK86, GT93, GY11, GRS00, GLM19, GV05, GL11, GW18, GM13, GMS22, Gue14, Gug03, GLS05, GS18b, GH19, GHS92, GM00, GH04, GZ09, GY13, GZ14b, GZ19, GTX22, Hab02, Hac79, HYZ24, HRW88, Han94, HK21, HO22, Hay19, HLP08, Hu15, HW18, HVDK19, Ito98, Kap94, KMM07, KK17b, KK19a, KK19b, KB01, KMM11, KR87a, Kim93b, KC10, Kim90, Kom91, Kom97, KY05, KGBS08, KGS11, KMR10, KW87b, KW87a, KP94b, KV07a, Lag77, Lag78, Lag83a, Lag88, Lag89, LL00a, LD05, LT83a, LT83b, Las84, LT86, LT89, Las98, LZM05, LI00, LR03, LHX12, LR16, LY19a, LY23, Liu03, Liu23, LPS12, Lü15, LtLR21, MXZ19]. Boundary [MNORP21, MY93, Mar99, MRV13, MD15, MM23, MZ97, Mic09, MKK19, MS97a, Mol19, MR04b, Mur92, NT95, NR15, Nic96, Nic97, Nic00, NP06, NLW91, NMI15, NFI15, Now08, ÖH14, Pan98, Pan17, PU15, PS19, PS04, PT95, Rao98, Rav20, Ray06, RR85, RS15a, RL20, Ros00c, Rus86a, SW78, Sch80a, Sch80b, Sch86, SU21, Sei78, Ser03, SCK10, Sok88, Ste13, Sun82, Świ20, TL06, Tib11, VM00, VC22, WW16, WC89, WL95, Wan00, WXY05, WW23, Was79, Wat24, Wec84, Wec00, WCG14, Whi80a, WT83, XX19, YDZ98, Zha99, ZH03, ZPQ20, ZZ20, ZL07, ZG18, ZW18, Zua90, Cur85a, Sei82]. Boundary-Control [Świ20]. Boundary-Noise [Świ20]. Boundary-Value [Aub02]. Boundary/Point [DMPD98]. Bounded [BM00a, Ban19, BGP09, Boe05, BT17, CFZ17, Cel02b, CRZ13, Che79, Che81a, CW03, DV04, Day09, DWZ20, Far08, Fer78, FCJ14, FT12, GRR09, GT22a, GP21, GM99, HW19, IP06a, Jay10, Kub85, LtLR21, MPN98, Mau77, MB04, MS03, PP14b, PK23, Ram98, RS21, RZ09, RK89, Rus76, Tre86, WC96a, Wee03, Wee05, XS96, ZHH17, ZSS^{+}20]. Bounded-Energy [Jay10]. Bounded-Input [BM00a, CW03, WC96a]. Bounded-Output [CW03, WC96a]. Bounded-State [BM00a]. Boundedness [BP08b, FHMS18, GTSN13, GP94, HK19, JP01, PAS00, Xie12]. Bounding [NTP18]. Bounds [Bar80a, BKL^{+}22, BC24, CK17b, Den98, GL90, GCP88, HW03, JK20, KS18, KLM20c, LCC17, Li95, Mun07, RY95, RSS22, SZ04, SDSÁR00, Sus79, UUH99, UU00, YF97, dFN23, DK89b]. Boussinesq [CC18a, CSFCL^{+}23, FCGIP06, FI98, LI00, VC22]. Box [BLPP24, Bur20, CY99]. Box-Constrained [CY99]. Brachistochrone [Lip97]. Bracket [SS18]. Brackets [Liu97c, MPS99, MR18b, Sus83]. Branching [CDDP11]. Bregman [BBC03, Kiw97]. Brezis [Ste08]. Bridge [HRCK21]. Bridges [CGP22]. Brightness [ZVBR14]. Brightness/Depth [ZVBR14]. Brinkman [AHR23]. Brockett [CY94, DF16, Rya94, Tan00]. Broken [CO22]. Brownian [AC02, BJ17, DHPD00, DMPD12, EMP24, GKL20, HP09b, Mor10, PZLM24a, PZLM24b, RS21, Shr81, Van92, ZH24]. Broyden [Fon89]. Brunn [Bou17]. Bruno [GEF17]. BSDEs [AJK14, Ban19, CH13, FHT06, FHT09, FMT10, GT20, HY22, HLT20, JZ22, TZK11]. BSPDEs [Dok17]. Bucy [AC19, BD17a, DKT17, JKSZ21, New06]. Budget [TCC^{+}15a]. Building [MA19]. Bundle [HU19, Kiw89]. Burgers [AK16, BT17, Cha09, DD99, GG07, KR15]. Burgers-Type [Cha09]. BV [BS04, CKK17, Dus21]. BV-Functions [CKK17]. Byrnes [IST16]. Byzantine [BL15].

C [HMW12, HT10]. C-Stationarity [HMW12, HT10]. Caching [Lid15]. Cahn [CFSGS15, CS15, CGS18, FSH16, FRS16, HW12, HW14, HKW17, RS15b, Sca19]. Calculus [Bou13, CZ94, CZ95b, CHPA20, DM09, DHPD00, Eke77, Fer05, GM05, GLNT23, HO01, HJM07, Hen80, HU84, IZ00, McS89, Mur86, OC12, PM16, RRW96, Thi97, Ulb02, Urs82, WB87]. Calibration [Ach05]. Calmness [Bur91a]. Can [IP06b, Mur86]. Canal [Mot00]. Canals [CGHS09, LS02]. Cancellations [LH87]. Cancer [LS07]. Cannot [BK03]. Canonical [BCG86, Bol80, CS05, CW81, HO97, IT17, JH14, Kal82, Pan83, PD79, Reb88, TR02, Teg84, WD76, WF76, WA84, BR87a]. Canonicity [Gri88]. Capacitary [AS13]. Capacity [BC96a, CKT20, CKT24, HYZ24, MS05, MZ07]. Capillary [Rei95]. Capital [BLY21, FS19, TST07]. Capture [Aub01]. Carathéodory [BR08, MPV18]. Carbon [CFH09, HLG21]. Carbonation [AT17]. Cardiac [CSB15]. Carleman [AAFC19, BHL10, CMV08, DFLZ23, FCG06, GT17, LY19b, WD14, YZ22, Zha00]. Carleson [HR83b, HR83a, JPP14]. Carlo [BKSW24, CCHP12, HLCH18, KK76, Kus78b, KY92, YK91]. Carnot [BR08, Hak20, MPV18]. Cars [GW96a]. Cartograms [LL09]. Cascade [BRS14, HS15, KCBP21, MSR01]. Cascaded [GW18, VPJ19, dNCP20]. Cascades [AAS02]. Case [ALY13, AZZ20, Bar95a, BÇS07, BÇMS^{+}08, BGP20b, BS09a, Bor94, BGS19, Bos02, CC02, CZ95b, CO03b, Cel15, CHL82, Che98, CLW17, CL87b, Cur03, Cza04b, FZ05, Fat90, FGH98, FGH05, GGM01, Gla05, GL99, GP02, HCG20, HLY24, HO97, HG02, HW12, HPT07, HZ19, KS07, Kru00b, KY92, Kus04, LS91a, LT89, LS19a, LMA07a, LMA07b, NR15, NSW17, PS17a, PS18, QN82, RT16, Rai99, Sag15, SX19, Sul94, Sul98, Sus87c, Sus87d, Tad93, Teb11, VP07, VP15, VSB18, WC19, YH21, ZZ15, ZZ17, BR87a, GR85b]. Cases [SL18]. Casino [HHOZ19]. Catching [Car90]. Catching-Up [Car90]. Categories [SO13]. Cathodic [HS96]. Cauchy [BSI19, HK13, IS12, IS14, ISdCN99, NR21, WHC03]. Causal [Fei80, FGT95, HH81, Jak86, NZ99, Por77, PY96, Ver97, WYZ23, WSR04, Zam98]. Causality [AT92, AT94, Fei80, Gro21, Kan82]. Caustic [Sac19]. Cavities [AHJM10]. Cavity [BL14]. Cech [AR99b]. Celebrating [Yin22]. Cell [BKSW24, CFRV98, CCS13]. Center [ATV13, AD17]. Centers [PYM93, RSS22]. Central [CH98, ISdCN99, WD07, Ber98a]. Centralized [KS97b]. Cepstral [Enq04, ZZ23]. Certain [Bal78, Car76, CHL82, Kra93, McE07, MDW23, MOF77, MN77b, PB83, VL78a, YMH94, Zhu96]. Certainty [Hua10, KV13]. Certificates [MS22]. Certifying [LPJ20]. CFD [SS09c]. chaîne [CL78]. Chain [BBC18, BA15, CGM14, CCHP12, CUB20, CSV24, HFY14, HLCH18, Kum08, Lei87, PR01, TZ07, Wat24, Zar92]. Chained [LAN04, MS04b, CL78]. Chains [BV82, Bor83, Bor84, Bor89, Bor94, Bor00a, Bor07, BLS12, Cal18, Can84, CC96, CHL15, CH13, CK89, Ete24, GL97c, GC05, HYZ11, Ich16, PRHL06, PRHL08, Rot77a, TLPA00, YD03]. champs [Bac79]. Chance [Chr78]. Chandrasekhar [IP87]. Chaney [HN96]. Change [BL15]. Changepoint [VC03]. Changing [CMA08a, CMA08b, MMH19]. Channel [GTQ21, JGC^{+}23, Mar20, dOP03]. Channels [BC96a, CFZ10, CM76, MS05, MS07a, TZQ21, Uch81, Yan93, YL12, Yük16]. Chaos [DKT17, FK21]. Chaotic [XWPG23]. Chaperoning [Aub13]. Characteristic [BSI19, CGW10, HW19, IS12, LY23, LMYZ23, Zer05]. Characteristics [WZ01].

Characterization [AD08, All88, AW90, ABBK20, BL01, Bel00, BN97, BB04, BK16, CIH18a, CWC^{+}22, CZL21, CW03, CvdSC05, CA10, DZ19, EÖ88, EH80, ES81b, FP14, Fer97b, Gal00, GLD^{+}91, GdZ90, HP21, HO90a, HJZ17, Jah85, KC15, KS98b, KS99b, LZ17, MARRH23, Mar02, MGMBN^{+}19, MN07, Mik12a, MS02, MCWY14, RU21, SK94, Son83, Ste04, Vin80, WC96a, ZLD12, Zol81, Zwa88b]. Characterizations [CF91, LSW^{+}17, LWXY22, SW00, SW99, Zol78]. Characterize [MP21]. Characterizing [HLT08, Ken80, MA19, Pan13]. Chattering [Art89b]. Checks [BÇMS^{+}08]. Chemical [Bd95, ITT04]. Chemo [GGMZRB20]. Chemo-Repulsion [GGMZRB20]. Chemotaxis [GGC23]. Chemotaxis-Consumption [GGC23]. Choice [BGP20b, BC18b, GL97c, HK94]. Choices [LW15c]. Choquet [HWZ23]. Chordal [Sas14]. Circle [Mos82b, San96, Mos82a]. Circuit [van86]. Circular [ZC91]. Claims [EQ95, KK18]. Clamped [CMR17]. Clarke [Bre07]. Class [AD15, AL12, Ahm83, Ahm03, ACS00b, ADRU22, Alv01, AMR14, AC05, ANP83, AP94, ACPZ19, AK04, AC79, ÁRL21, BP10, Bal91, Bar83, BCL83, BM90, BR10, BG09, BC13, BKR17, BGL11, BGH94, CMV08, Car90, CT09, CK17a, CK22, CDR08, CC96, CLW17, CK78, CNS21, CW81, CB11, CA09, CdSBT09, Cra99, DM89, DMK90, De 11, De 00, DKT17, DHV98, DG79, Dmi92, DM15, DMS12, DA17, Dun81, DS01, EMP18, FFRR21, FR78a, FG15, FS16, FBJM14, Fon89, FV11, FHT06, FO16, Fuh76, GLZ04, GT93, GP79, GTSN13, Gor15, GZSAB13, GSS19, GH96, GT09, GTX22, GYBL15, HD11, HZ07, HL88, HY22, HL81, HS87, IK09, IK10, KG80, KM07a, KJ07, KK19a, KBMS20, KN14a, KG17, Kob79, KY05, KW87a, KV07a, Lag77]. Class [Lag78, Lau99, Lev87, LY19a, LŻ11, LLC18, LG07, Liu13, LLS21, LH24, LI94b, Ma92, MP14, MADF98, MK10, Mey77c, Mey77b, MR86, MNS91, NR21, NMBM98, Nov04, Nur06, OT94, Pav13, Pul93, QZ13, QW19, RW87b, RBS21, Reb88, Sla00, SW12, SZZ97, Sut20, TKH10, Teg84, TC14, Tie14, TY00, TWW16, VP22a, WD14, WZ09, WCLC05, XWPG23, Yam89, YMH94, Yan95, YW12, Yu15a, Yük17, ZC06, ZZS14, ZZ20, Zho83, dWvS00, dBLR14, CCHH19, FR82a, GB84b]. Classes [Car76, CE98, CR13, FM08, Gra07, GY12, pJY20, Koe90, Obe91, SPBF10]. Classical [BPS21, BH22, GDL11, HT12, HZ21, Mor98, Ngu18, Pav13, SM13, Vrd16, WR14]. Classification [AZ07, BP98a, CYL96, CYL97b, CY94, Dic76, DTWY91, LL20, Pic96, TR02, Tal05, Tan00, WY06, FFR14]. Classification/Estimation [FFR14]. Clearance [ADL23]. Clearing [FT22]. Clebsch [Lic13]. CLF [LSDS13, SD13, SD15]. Clock [XC18]. Closed [Bru76, BH18, DS09, GSW89, HK21, Hin05, LWZ22, MXZ19, MZ20, MPS99, Ols02, SD15, SY14, SLY16, WZ17c, WZ21, Zwa88a]. Closed-Form [MXZ19]. Closed-Loop [Bru76, LWZ22, MZ20, Ols02, SY14, SLY16, WZ21, Zwa88a]. Closest [DHK20]. Closure [Bal82, Bal84, Las97, Rai99]. Closures [HLT08, HO90a]. Cloud [ABHY23]. Cluster [CBB17a, EOF07, Kim21]. Clustering [BHPM04, BM06, KP04]. CMG [BP14b]. Coalitional [SBT19]. Cobb [MZ09]. COCOLOG [CW95, WC96b]. Coders [Yan93]. Codes [FNP^{+}22, Hut09]. Codesign [CWQ18]. Codimensions [Sac97]. Coding [CKT20, CKT24, CWQ18]. Coding/Control [CWQ18]. Coefficient [Aca93, BHOP14, CKL16, GB93, JZ09b, JZ21, LZ16c, Luk85, McN84, Riv90, XZ15]. Coefficients [Ahm81, AF18, AG19, BM86, BB19, BGR07, BD19a, Bis76, Bou93, BKM97, CK95, CDCCMG04, CDCCMG08, CLHH05, Cer01a, CPY09, DDK12, DKR08, DQ21, EB20, EB24, FL22, Gor19, GM09, Guo02, Hop87, HQZ16, HZ05b, KH81b, Lam87, LC01, LBMH^{+}16, LYZ23, Nag85, Nic96, Pie79, Tan03, Tan15, Toa11, WC91, WY21, YT13, Yao99, ZP03b, ZDM20]. Coercive [HLZ20, Idc98, MS10]. Coherent [ZLHZ12]. Cohomology [Rav00]. Cohypomonotone [CP04]. Coils [HS03]. Coincidence [BT91]. Collection [FPD08]. Collective [AHKS22]. Collectively [BLP14]. Collision [GC22]. Collocated [WCG14]. Collocation [HLM^{+}18, Red79, TKXP12]. Colocated [GZ05, CW06]. Colored [BR13, LR00a]. Coloring [KBvdSD17]. Combination [Jar93]. Combinatorial [BO14, Man21]. Combinatorics [CZL21]. Combinatorics-Based [CZL21]. Combined [AK99b, DQS16, KB76, Mor03, Zer03]. Commensurate [BFP11, Ngu18, Tad89]. Comment [Bal01]. Comments [Boo82, Cou83, Kot94, Wei77]. Common [Ahu16, FIR06, GNLB14, Gup20, Imm07, KRW04, MBC06, SSY24]. Communicating [CC96, KY87b, LC19]. Communication [CIH18a, FZ09, GAP^{+}09, JVB^{+}23, KvdSD21, Lin14, LLX11, LFH15, MS05, MS07a, RC16, SSK02, TZQ21, WQS^{+}24, Won09, Yan93, Yük16, ZZ13]. Communications [AK99b, BHT10, LCF18]. Commutation [EMP^{+}16]. Commutative [BRR02, GK85, Kha82, MA01, Sul94, Sul98]. commutatives [Cla86]. Commutativity [VP87]. Commuting [Lou94]. Compact [BJ02, BR15b, Duc20, Gib80, Gow89, GY12, KK16, Kum82, Lei85, Mis19, SD13]. Compactification [AR99b, Bal81, Ros94]. Compactness [AG88]. Companion [Hug16]. Company [CTZ03, CM07]. Comparison [Cel07, EK85, GT90, HBY21, Liu11, Mey77d, RTZ20, SQS09, Xie12, CL06a, CdLdDM02].

Compartmental [van98]. Compass

[HT92]. Compasses [MSJ15].

Compensation [BH86, Ber89b, Gho88, GP02, LMZ20, LR06, Ros94, WL24, WF76]. Compensator [GB93, JN19a, Las95, Sch83]. Compensators [Cob12, CR13, Cur84, CS86, Ito90, XB99]. Competition [CNS21, ELO22, EMO23, Win78]. Competitive [BS19b, CH96, EP19]. Complementarity [BP17, Bro22, CD80, FK99, FT77, Fuj80b, Gow89, HK78, Jer78, JQ97, Kan78, LT91, LC23, MS87, SP05, Wat79, WZC20, ZI00, ZL01, van86]. Complete [AJPR14, FN84, HPW83, HO90a, Hui92, Lep90, LWXY22, Nar84, RW76, SO06a]. Completely [AS13, BCM12a, Sun81, YD03]. Complex [ABK19, Ete20, FL17, HS13, JHMG06, LH11, TLL16, YCLK13, YLYC15, ZZCaL19]. Complexes [Lom12b]. Complexity [CZA08, FGJ^{+}19, JP15, Kam10, KOS04, MADF98, NNTD14, Ovs14, Won09]. Compliance [CD15]. Component [ADS24]. Components [JR93]. Composite [HP09a, SKS90]. Composition [Rus21]. Compositional [FM09, KG22]. Compound [AM15, BB09, BLS05a]. Compressible [AAR18, CRR12, CDM21b, Mol19, PS06, Tao18, Wan21]. Compression [BHM99, Hag02, RKL16]. Computability [Col11]. Computable [GL90, JK20]. Computation [BBBM08, BL89, Bol80, CH05, CM18, DZ91, FPT15, GL97a, HTY01, Kal82, KHJ14, RS10, RU19, WHC03, ZKW99]. Computational [AKO17, EK85, HD11, IY12, MPZ03, MADF98, NNTD14]. Computations [Hag02]. Compute [RBS21]. Computed [Pro22]. Computing [DA17, GO04, ID87, Kor20, Lin07, Yan96]. Concave [KSY23]. Concentration [XJ24]. Concept [BKR00, EPR14, KK14].

Conception [BK97a]. Concepts [BM99, Cur86, FP96]. Conceptual [VGBS13a]. Concerning [FGB95, JZ04, OT16]. Concrete [AT17]. Condition [Ahu16, Ama04, AAA12, AJK14, BM00a, BO14, BM94, BN16, CN11, Cel02b, CWQ18, CFSGS15, CS15, Col82a, Fen16, FT12, FN84, FT22, GL22a, GB82b, GW15, GT22a, GZSAB13, GX04, HHS18, HQZ16, HLP08, JL23, Jön09, KC10, Kra09, Las97, Lic13, MP76b, MS02, OP96, Pet85, Pom09, PS04, RH06, RL78, RW94, Rya94, Sîr14a, Sus78, Sus83, TWY90, Tsi93, VL78b, VZ97a, XG18, YH21, ZL01]. Conditional [AØ23, CW95, CEK03, Dun79, Dun80a, Dun86, GP98, Ham82, MSZ18, MY17, XUP08]. Conditionally [KM86]. Conditions [AN15a, AN15b, AT89, AX97, ACT10, AJ77, AX02, AMP05, AA17, ANS14, AR99b, ABK21, AK04, AL00, Bah08, Bal82, BLJ15, Bar81, Bar21, BLR92, Bat95, BJ04, BGR07, BT96, Ber84, BBV13, Bet19, BGMR94, BGBPG04, BDP14, BT03b, BT10, Bro87, BC23b, BPTT21, CM06, CDCCMG04, CTU00, CM02, CT09, CWW17, CW20, CCFV20, CM10b, Cha82, Cha87a, CH99, CLM91, Cha91, CSFCL^{+}23, Che04, CGC17, CFR13, Chr77, CV83, Cob10, CGS18, CKR22, CBdN08, CB15, DAR15, DM09, Den22, DO14, Du15, DJB21, DT92, Dun95, Dun96, Dun98, EMH08, EZB11, Emr82, Eng10, Fei80, FBJM14, FNR19, Fra84, FT13, FO18, Fuj86, GT17, GST20, GAC24, GLX18, GKY21, Gfr06, Gil77, GK86, GT93, Gon87b, GTSN13, GHQ15, Gra88, GKLW14, GY16b]. Conditions [Han86, HM93, Her76, HU78a, Hoe12, Hoe14, Iof79a, Iof79c, Iof79b, Ish88, IK11a, IS02, JZ99, Khá93, KWY08, KN14a, KTY18, KQWY20, KBvdSD17, KK14, KR17, KR13, Kus78a, LLS90, LD05, Las84, LT89, LS98, LP79a, LP81, Lew80, LY91c, LXZ11, LHX12, tLR23, LF15, LLLT18, LR96, LRW96, LY09, LZZ21, Mal97, MT78, Man81, MO02, MO04, McE95, MMR17b, MM23, Mey77a, MPT06, MY09, Mic80, Mil02, Mir86, Mor95, MW02, Muñ00, NR15, Now08, Now17, Ped98, Pen98, PS00, RT03, RW08, Sar97, Sch88a, Sei88, Sha84, SV97, SZ03, SZ96, Stu86, SW99, Sun82, TL94, TY23, Tib85, Tib11, TSS^{+}22, TR01, Tri76, Uch81, Urs82, VP87, VT95, VC22, Vin83a, Vin83b, WS90, War76b, War84, War86, War88, Ye95, Yon10, Yu13]. Conditions [ZZ88, Zei01, ZZ15, ZZ17, ZZ20, Zho93, Zho98, ZH02, Zhu00, ZLZ18, dRdPR08, dWvS00, van95]. Conduction [BF89b, EB20, EB24, FJ83, GH07]. Conductive [GBR18]. Conductivity [GH07, KS86, Vrd16]. Cone [ACT10, BBV11, Bre07, CWC^{+}22, DG90, GR02, GR21a, Pha00, SB92, WK77]. Cone-Invariant [CWC^{+}22]. Cones [Bor77, Bor78, DT92, Fra89, Gow89, MPV18, SH12]. Conewise [CPS06, SH12]. Confidence [ZHH17]. Configuration [CM03b, LM97a, RM98]. Confined [BBV11]. Confinement [CP13b]. Confirmation [BSB^{+}97]. Conflicting [War91]. Congested [BJO09]. Congestion [CJS08, KDB15, RS07b]. Conic [CZ04b, KC15, PHL16, WZY09]. Conical [All88, AW90, War88]. Conjecture [AFK10, Bay09a, CYL97a, Fre99, OT16, Ros00b]. Conjectures [MB12, Wyn09]. Conjugate [HPT07, KZ00a, Kaw01, MOF77, ZZ88]. Connected [BB19, CZ98, CDKP87, Cor99, Fio20, GX04, JR93]. Connection [BL05, CF05, NSW17, NM10]. Connections [GT08, KS84a, KS85a, LFC21, Ohs13]. Connectivity [BD17b, FL20, MG13, SNB09, SC12]. Consensus [APGC16, AB09, BGP09, BD21, BO14, CFT12, CMA08a, CMA08b, CSBC14, CF18, CLYL13, CBB17a, CLSL20, CHJ15, DV12, DZX24, FZ09, FC19, FCC20, GZ12, GL13, GP21, GH20, HM09, HLZ15, HQR24, Jaf17, LC19, LLC11, LLX11, LH11, LLJC18, LWDB20, LZ21, LGZ13, LL21b, MCL^{+}22, MG13, MBS13, MTS15, Med12, MMN09, MS18, OT09, Ols14, Ols17, QS13, RT16, SS09a, SJ13, TLL16, VJBK24, WSZ^{+}21, WQS^{+}24, WMWL22, XW12, XG18, YHJ^{+}18, ZZ13, ZLZ18]. Consensus-Based [GZ12, HQR24]. Consequence [BBB24]. Consequences [Wan08, Zua96, Zua98]. Conservation [AM98, ACCG18, AN21, ADM17, BC02a, Dus21, JS99, KY05, LGP15, Léa12, LY23, Per12, PU15, Ulb02]. Conservative [Liu97a, TW03]. Conserving [DvdS98]. Consistency [FP11a, HSJ19, Wri95]. Consistent [CG87a, CG87b, CE11, DHK20, Gom89, GSZ20, NP96, PP02, PEB19, SP96, VGBS13a, VGBS13b, WJW19]. Consisting [CZ98]. Constancy [Vin19]. Constant [AN20, Bar85a, BRW05, Can84, DKR08, Far08, FP99, Fre99, GH07, JZ09b, JH00, KCP16, KN77, KH81b, Luk85, Nik99, OPPU17, Pfl86, QWY21b, RS13, Cou83]. Constant-Gain [JH00]. Constants [Li94a]. Constituting [APH24]. Constrained [AM89, ABK21, AK15, AF21, Ban19, BCW21, BBBM08, BJ96, BT96, BK97b, BIK99, Ber76, BM00b, BTW82, BFZ10, BPZ16, BT03a, BC95b, BH07, BS11, BG20, BG21, BKS19, BB04, BF07, BDP24, BK02, Bud03, BKM14, Bur91b, CFG21, CK02, Cas97, CT14, Cel02a, CM95, CF05, CQ05, Cha87a, CHL82, Che96, CY99, Che04, CK03, CNP98, CS03, CMN20, CDP16, Cvi00, DL13b, Del22, DZ24, Di 91, DG89, DJZ96, DH98b, DKS^{+}11, DPR13, DPR22, Dun80b, Dun86, Dun98, DRDG08, Fon89, GB84a, GBBT78, GV03, GSZ20, GMS22, Gug10, GH96, GK11, GHS12, Hag79, HLM^{+}18, Han77, HJ13, He87, HLGHLM03, HK06, HH08, HH10, HQZ16, HZ05b, HPT07, Hua07, ITT04, IS02, JPSB24, JN19b, Kat19, KLM20c, KV07a, KS13, LM15, LSDS13, LZ18a, Lor87]. Constrained [LB98, MY09, MP11a, Mif77, MU22, MS04a, MN08, Nar82, NNTD14, NPR15, Nie21, OL77, Ouz10, PT87, PTH88, PP18, PS86, PD92, PZ20, PMW83, PAH88, RZ98, Rot96, Rou86, Sch09, SB80, SW11, SR97, Smi96, Sok87, ST16, SZ96, SZ97, Ste04, Stu15, SD19, Szn93, VGBS13a, VGBS13b, Vog90, WW02, WS13a, WYY21, War82, Wri89, Wri93, XWZ^{+}21, YZ97, dRdPR08, vDHV15, CTU00]. Constraint [AAC86, ACT10, AB05, Ban05, BV18, BDR02, BRS14, BBV10, BP17, BO16, BC96b, BC96c, BH07, BLZ17, CM90, DPT04, DO22, Dun00, FH20, GP79, GT90, Gün95, IK96a, KMF19, KP04, LN10, Lei06, MR16, MR17, MR18a, NS20, Son86b, TGSM18, ZZ15, ZRS97, Son86a]. Constraints [Alt09, AM95, AJK14, AR99a, AR00, AA97, AKP05, Aub02, ABSP08, Aub11, BB98, Bar94, BdP21, BSVD84, BB06, Ber92a, Ber98b, BP13a, Ber82b, BV10, BBV13, BK24b, BISW17, BdPV16, BC89, BZ99b, BJ10, BGP23, Bor94, Bou08, BEI10, BN13, ByS17, BT03b, BT10, BK83b, BQT08, BR07, BS06, CQSP00, CMS16, CSBC14, Cas86, Cas93, CF95, CRZ00, CM02, CL87a, CZ04b, CKL23, CTZ03, Chr78, CL87b, CdP10, CKR22, CRST18, CMOZ02, DXG15, De 83, Den22, DR08, DG85, DO14, DH12, Fag08, FSGHH20, FGG10, FGG11, FHH22, FS16, Fle85, Fra87b, FR99, FT13, FO18, FZZ18, FM22, GKY21, GZ14a, GY16b, GTT16b, Hal89, HMT18, Heu15, HT10, Hoe12, HS87, HN22, HY95, HHN18, HK19, IK04, Jer78, KC15, KP23]. Constraints [Kas02, KZ00a, Kha16, KN14a, KQWY20, KFP22, KS98a, KRS14, KKV11, KMR10, KR13, KLL10, KM07b, KV06, LZL02, LW15b, LL20, LLM^{+}24, LG07, LR97, Lor10, LW11b, LY01, LY02, LW15d, LC23, Mal97, MG20, MG21, MV08a, MV08b, MRV11, Mic80, MR04b, Mot95, Out00, PZ94, Pen98, PTZ21, Pha00, PZ22, PYM93, PA93, Pre96, PRHL08, PV98, RR94, RV00, RW78, RT03, RT06, RT07, RV93, Rus76, SU21, SS99b, SB92, ST00, SYZ12, Son23, Spi80, Sto76, TCC^{+}15a, Tam80, TL03, TLPA00, TKXP12, Tre99, Tse90, VW08, Wac16, Wan00, Wan02, Whi82, WK77, Wol12, WM12, XUP08, Yu95, YHH16, Zar94, Zei94, Zha13, Zhu00, dlRT07]. Construct [BF22]. Constructing [Fis11, Phe91]. Construction [CZA04a, DRW10, Gri83, Jac22, KKLO21, LW17, MI15, Mor79, SO13, WP98]. Constructive [CDR08, GK87a, MT22, NP22, ZL02]. Consumption [AMS96, ARS17, CLHH05, CSZ13, DJLY09, GGP14, GGC23, Ham21, HS12a, HS12b, HNS18, HK19, KLS87, LS91a, MZ09, OS02, SLG84, Zar92, Zar94]. Consumption-Investment [CLHH05, Ham21, Zar94]. Contact [BHOP14, BK21, CDD20, DTV18, Sac19]. Contagion [BLY19]. Contagious [BC17, QJL22]. Contain [MMK86]. Container [KLM04, SZ10]. Containing [Dol76b, LY09, Tol16]. Context [MW23, Per12, Rya94]. Contingent [EQ95, Fra89, KK18]. Continuation [CZ04a, DKRV13]. Continuity [AK15, AR77, BHT19, CDP23, CK03, CL94, Dav07, De 15, DMR19, GV03, Hag79, HBY23, JP01, MS87, Rot77b, TLPA00, Ye00b]. Continuous [AC13, AC17, ANP83, AP94, AS11, Ave16, Bag81, Ban19, BB00, BC21, BJ19, BHT10, Bro10, Cal18, CLS12, Car15, CG90, CLW17, CKL23, Chr83, Cor95, CdV02, CD10, DHX10, DS19, Don24, Ega80, EK97, FPT15, Fio20, FB01, FR05, FC05, GZLC17, GP22, GW84, GC06, GGW91, GC05, GHS12, GHZ15, GL19, HD11, HZ10a, HM14, HFY14, HWZ23, HJ21, HSZ15, HSZ24, IP06a, Jaf17, Jam14, JL98, KM16, KSY23, Kar89, Kim22, KN14b, KV13, Küe07, KS85c, Kus90, KOS04, LS05a, LFM07, LLM^{+}24, Lin07, Liu95a, LS99b, LK10, LB98, Mai13, MG13, Med12, MR90, MY17, MVBC18, NC16, OW96, OMK84, Opm07, OS14, OOS15, Pav84a, Per81, Phi94, PZ11, Piu15, Piu18, PZ22, PRHL06, PRHL08, PK23, Pul93, Pul95, Pul96, RSZ23, Rod82]. Continuous [RW92, SP02, SJ13, Sîr14a, SZX16, SR18, SV79, STZ24, TC14, TF08, TF11, TG23, TV02, TFL14, Vin83b, WZY09, WTJY21, Wim94, Wyn09, Xia11, Yan96, Yük24, Yus97, ZRSZG22, ZYHX18, Zha17b, ZH02, ZY03, ZH05, Zho07, ZLZ18, dBMS10]. Continuous-in-Action [Yus97]. Continuous-Mode [SZX16]. Continuous-Time [Ave16, Bag81, BJ19, BHT10, Cal18, Car15, CG90, CKL23, CdV02, DHX10, Ega80, FPT15, Fio20, FB01, FR05, FC05, GZLC17, GP22, GC06, GGW91, GC05, GHS12, GHZ15, GL19, HWZ23, Jaf17, KM16, KN14b, KV13, KOS04, LFM07, LLM^{+}24, Lin07, LK10, MG13, OW96, OMK84, Opm07, Pav84a, Phi94, PZ11, Piu15, Piu18, PZ22, PRHL06, PRHL08, SJ13, Sîr14a, STZ24, TF08, TF11, TG23, TV02, WTJY21, Wim94, Xia11, ZYHX18, Zha17b, ZH02, ZY03, ZH05, Zho07, ZLZ18, EK97, NC16]. Continuously [DG85]. Continuum [Agr95, ABS18, BB03, BF13, HO16, HTY01, NH12, NFI15, PYM93]. Continuum-Armed [Agr95]. Continuum-Parametrized [NH12]. Contrôlabilité [GB82a, Lob70, Lob76]. Contrôle [Tah92, LM84, Pen89, Qua80]. Contrôllabilité [Bon84]. Contract [HSM19]. Contracting [EP19]. Contraction [Ben78, LL07, LZ16a, NT18, Qu14, RS07b, VB23, Zor17, dBLR14]. Contractions [Bou93, Lou94]. Contractive [HMA81, LR79a]. Contractivity [PDB23]. Contracts [NS20]. Contrast [BCM13]. Contrasting [NS23]. Control [ASK05, AT89, ABVC19, AFT91, AT96, AZ07, AB18, AS20, AØ23, Ahm77, Ahm81, AX96, Ahm07, Ahm19, AW21, AT17, AGL08, ABCL17, AD98, ADRU22, AHKS22, AS13, AFOQ21, AFO22, AW90, ALNP24, Alm01, AHR20, AM89, AK99b, AK02a, Alv01, AB02b, AL08, AM17, APGC16, Ama04, AQ05, AY23a, AGJ^{+}13, AM98, AT92, AT94, AP92, Ang76, ACV22, AJK14, AF18, AKTW24, ANS14, AO15, AOS16, AKP21, AHS13, AMV20, ARS09a, AMPR15, AT00b, ANR08, AR99a, AR99b, AR00, Ara12, Ara13, Ara14, AB10a, ABK14, ACPZ19, AB20, AFCLS22, ADL23, ABK19, AMR20, ABK21, Art78, Art89a, AL02, AB10b, AA97, AKP05, AK04, AGS13, AC79, AK02b, AVCCD22, ABS18, AH21b, AMI20, ARS09b, AK16, BD01a, BM19, BDM07, Bah08, BP94, BRR02]. Control [Bal78, Bal84, BCD02, BKP14, Ban19, Ban05, BV18, BB78, BK82, BW89, BBBM08, BBD18, BLPdD15, Bar78, Bar80b, Bar81, Bar82, BD83, Bar88, Bar91, Bar94, Bar95a, BP97, Bar97, BCG^{+}17a, BRZ20, Bar23b, BC81, BLR92, BB18, Bar85b, BP88b, BDR94, BBC14, Bar85c, Bar93, BJ95, BL96, BJ96, BM90, BB08, BL18, BR87b, Bay09a, BRS14, BS16, BBB24, BEM13, BCP16, BCP17, BC23a, BdP21, BKR00, BA15, BCS17, BB97, Bel00, BKS10, BBGBO14, BBT14, Bv85, BE95, BN97, BCLY17, BBF18, Ber21, BDIW22, BDLW24, Ber92a, BT96, BK97b, Ber98b, BIK99, BZ99a, BL04, BF76a, Ber88, BS87, BGR04, BET07, Ber18, BF22, BV10, BV17a, Bet19, BC21, BFPB13, Bie95, BCC^{+}19b, Bil19, Bin76, BT98, BH00, BF90b, BV13b, Bis76]. Control [Bis77a, Bis78a, Bis78b, Bis82, BISW17, Bla97, BP07b, BC95a, BT97, BCS20, BdPV16, BV17b, BV77, BK80, Boe05, BJ02, BDRR98, BPZ16, BP13b, BCM12a, BHT19, BB23, BK21, BPV19a, BPV19b, Bon86, BC89, BC95b, BZ99b, BH07, BJ10, BS11, BDP14, BG20, BG21, Bon85, Bon91, Bd95, BS09a, BM13, BV82, Bor84, Bor89, Bor94, BB04, Bor07, BG07, BK10, BGS19, BK16, BKK02, BK13, BC05b, BD17b, BEI10, BT17, BT13, BT21, BDP24, Bra09, BK14, BKP18, BP20c, BK23, ByS17, BC02a, BR01, Bro87, BC23b, BDDSV05, Bro10, BG14, Bru76, BH18, BS13b, Buc92, BP98b, BM07, BN16, BJ17, Bud03, BR07, BGL11, BLS12, Bul01, Bul02, BL05, BCP09, BCF12, BKM14, BCKM24, BPTT21, BFO20, BIP88, BR88, Bus84, BM18b]. Control [BL08b, CC02, CLL20, CK95, CD13a, CZ95a, Cal09, Cal18, CFG21, CGW08, CBM23, Cam76, CK98, CGM98, CD89, CF91, CT96, CN11, CFS15, CQ15, CFM19, CK02, CLAL13, CY18, CLZH24, Car97a, Car97b, CDJS23, CHJ87, Car90, CFH09, Car15, Car76, CDCCMG04, CDCCMG08, Cas86, Cas93, CF95, Cas97, CRZ00, CTU00, CM02, CR06, CMR07, CT09, CS10a, CGM11, CCK12, Cas12, CCK13, CK14, CT14, CRT15, CKL16, CC16, CKK17, CWW17, CMT18, CWW18, CMR19, CK19, CW20, CT20, CK21, CK22, CK23a, CK23b, CCJ23, CW23a, CHY23, CZ04a, ÇR14, CO03b, CM95, CF05, CMP15, Cer01a, Cer01b, CIH18a, CG10, CT19, Cha87b, CE98, CKT20, CKT24, CC18b, Che79, Che81a, CM81, Che84, CG87a, CDKP87, CG87b, CLZ98, Che99, CZ00, CZ01, Che04, CZ04b, CCT05]. Control [CCT07, CY12, CG13, CWQ18, CR21, CSBR21, CM96a, CM03a, CS77, CH98, CJT08, CJM12, CK03, CDL85, CTZ03, CMR85, CNS21, Chr19, CW23c, Chr78, Chr80, Chr83, Cir14, CW81, CL86, CL87b, CW91, CN05, CdP10, CLdRdP11, CTW18, CCS13, Coh21, CFSGS15, CS15, CGS18, CGHS09, CP13b, CH78, Col85, Col88, CS89, CW09, CK09, CCS18, CSV24, CF13, Con90, CM76, Cor94, CG17b, CGW20, CMOZ02, CM03b, CvdSC05, CM18, CD10, CD12, CDP16, CRV09, Cou83, Cro84, CLZ23, CP77, Cur84, Cur85a, Cza04b, DDK12, DG14, DGO19, DM02a, DI87, Da 87, DI90, DD92, DI93, DD99, Dág04, DK05, Dai12, DXG15, DDT00, DGSW14, Dat80, DH84, Dav76b, DF96, DF98a, DGW10, DFM15, DEO24, DFŚ24, De 11, De 00, DGH09, DES09, DL09, Del86]. Control [DMS03, Den22, DZ24, DBC12, DZ19, DR05, Di 91, DR87, DS99a, DS07, DFP14, Dic76, DBS05, Die87, DJZ96, DCI07, DGQ13, DHKS18, DO14, DO22, Dok17, DR77, DF16, DF21, DD03, DWZ^{+}16, DTM22, DH93, Don96, DH98b, DKK^{+}19, DL19, DL20, DM17, Dre77, DNS17, DKS^{+}11, DM13b, Du15, DMZ22, DZX24, DM02b, DM04, DM06, DQS16, DMPD94, DPDS98, DMPD98, DMPD00, DMPD12, DPD13, Dun86, Dun96, Dun98, Dun00, DF98d, DJ98, DS01, Dup03, DRDG08, EGR^{+}22, Ega08, Ega80, EI13, ENJP88, EMP18, Ell77, EL20, EB20, EB24, Fab08, FG09, FR17, Fag08, FM12, Far08, FNA02, FSH16, FSGHH20, FMS20, Fat94, FM94, Fat97, FMP^{+}22, FGG10, FGG11, FT14, FFS20, FFRR21, FFR23, FL81, FR82a, FCG18, FZ19, FS21, Fer18, FR20]. Control [FPT15, FGL08, Fio20, Fis11, Flå95, Fla84, Fla87, Fla90, FP82, FV89, FM95, FHH97, FHH99, FP04, Flo97, Flo99, FM09, FR14, FM16, FB01, FD87, Fra84, Fra89, FR99, FT13, FO18, FZZ18, FK76, FB12, FP11b, FA13, FL80, FL82, FJ83, Fri84, Fri86, Fri87, FH88, FHY88, FRS16, FK21, FP03, FHT06, FHT09, FMT10, FO16, FN84, FT02, FGH98, FGH05, FM22, FMV24, GKB00, Gah94, GL97a, Gai92, GN02, Gai04, GR06, GQ09, GR17a, Gal00, GAC24, GLZ04, GCC07, GMCL10, GXZ22, GLS21, GT07, GK95, GS19, GVL12, Gen13, GMP00, GPB91, GPS95, GGF09, GGW91, GAM93, GAM97, Gib79, Gib83, GA91, GRT92, GRZ24, Gil77, GW76, Gli12, GM05, GG22, Goe05, GT93, GY11, GWY14, GHZ14, GTHLHR02]. Control [GRC81, GHX86, Goo76, Gor15, Goz91, GRS00, GM17a, GM17b, GM21, GHS22, Gra82, Gra07, GH15, GL97c, GV05, GK06, Gri83, Gro21, GLNT23, Grü96, Grü98, GPSW10, GG18, GSS19, Gu90, GW18, GM09, GM13, GT22b, GMS22, GQ11, Gue07, GLS05, Gug07, Gug10, GH19, GAHL09, Gui13, GGMZRB20, GGC23, GM11, GHS92, GH96, GK98, GM00, GK11, GW19, Guo94a, GC05, GZ05, GT08, GT09, GW09, GZ09, GZ13, GZ14b, GY16b, Ha88, Hab02, HT11, Hac79, HHI^{+}13, HLL^{+}17, HD11, HM03, HH76, HM76, HI84, HLM^{+}18, Hal89, HMA81, HKB04, HPW10, HCG20, HLZF21, Han94, HK21, HSBTZ13, HO22, HAB78, Hau81b, Hau81a, Hau87, Hau94, He87, HYZ11, Hei16, HS91, HSZ24, Her80, Her86, HP22, HHSS23, HLL94, HLL96, HLGH98, HLVAC99]. Control [HLR01, HLGHLM03, HMW12, HSW12, HSW15, HG02, HS13, HH08, HH10, HW12, HMMS13, HW14, HKW17, HK01, Hin05, Hir77, Hir79, Ho86, Ho90b, Ho93, Hoe12, Hoe14, HR21, HMRR10, HJF10, HN22, HQZ16, HR98, HTLV17, HLP08, HZ05b, HZ05a, HO08, HZ10b, HJZ12, HSP16, HJ16, HJZ17, HW18, HLDM20, HW20, HJX22, HT22, HCM05, Hua07, HWX09, HYZ21, HWY21, HV00, HBS06, Ich82, Idc98, Igl96, IY12, IM10, IP06b, Imm07, Iof24, IT87, IK94, IK96b, IM98, IK02a, IK02b, IK04, IK06, IK07b, IK09, IK10, IK14, Izm96, JMP21, JL23, JLZ98, JZ99, JR06a, JB96, JP15, JPS20, JS24, JX22, JK06, JGC^{+}23, JK20, JS81, JNKV21, Jun00, JM17, KM16, KR80, Kan98a, Kan98b, Kan00, KMM07, KY19, KY20, KJ07, KK24, KB81, Kas02, Kas09, Kaw11]. Control [KN13, KLR09, KT04, KS87b, KS92, KS95, KLM20a, Ken80, KS97a, Kha03, KNP91, KB01, KBDK03, KWY08, KTWY12, KN14a, KTY18, KQWY20, KFP22, KR87a, Kim92, Kim93a, Kim93b, KY23, KLM20c, KCBP21, Kno76, Kno77, Kno79, Kno82, KS98a, KS00, KV21, KvdSD21, Kog87, KT03b, KRS14, KM86, KM97, KHJ14, KK02, KKS93, KLN21, KL16, KT16, KL02, KM18, KKV11, KPZ18, KGBS08, KLM^{+}21, Kru00a, Kru00b, KMR10, KR13, Kul93, Kum82, Kum83, Kum85, KG95, KS97b, Kum13, Kum88, Kun82, KV07a, KLL10, KW13, KPV14, KTV16, KS13, KM04, KM07b, Kus89, KR89, Kus90, KM91, Kus95, Kus13, KOS04, Kwo80, Kwo91, Lag77, Lag83b, LLS93, LL00a, LY91a, LLS90, LD05, LG09, LM13a, LT83a, Las84, LS91b, LHPT08, LW15a, LZM05, LS07, Lee86, LI00, LS05b]. Control [LMJ09, LW07, LS99a, Lei06, LHB16, Leu99, LR79a, Lew80, LM97a, LV16, LSDS13, LPT22, LY91c, LLMT02, Li07, LW14, LW15b, LZ16b, LCW18, LL18a, LL18b, LD19, LS19a, LMZ20, LMMY20, LWLZ20, LZC21b, LWfZ22, LK23, LTT24, LWR24, Lia15, LWZ22, LZ01, LH87, LH88, LAKG92, LW11a, Lin14, LLC18, LM82a, LM82b, LS85a, LP86, LRT96, LZ94, Liu97a, LS99b, LZ06, LG07, Liu14a, LW15c, LFH15, LYB18, LWZC19, LLSX20, Liu23, LT23, LR94, LZ92, LT97, LRT98, LR00b, LM02, LT03, LM00, Lor10, LMPR09, Lou02, Lou03, Lou07, Lü15, LZH^{+}16, LSL^{+}18, LZZ21, LSW14, LC23, Ma92, MZ18, MP14, MM04, Mac14, MT09a, MPN98, MS92, Mal97, Mal23, Mal02, MZ15, Mam14, MS23, MBL00, MY93, MTS20, MPZ03, MG20, MG21, MRR13, MRVM00, MK90, MCPA17, Mas08, MSBC08, MR78, MR79, Mat05, Mau77]. Control [MP94, MO02, MO04, MRV13, MCS13, MMN17, MSSZ88, McC76, MY22, McN84, McS89, Mee89, MML21, MWY21, MV08a, MV08b, MV11, MMR17b, MMR17a, Men80a, MR17, MR18a, MS98, MS11, MW23, MP76b, MR04a, MR05, MPT06, MY09, MST16, MS17a, MW21, MP11a, MZ97, MN22, MP84, MT08, Mik12b, MM80b, Mil96, MR97, Mil02, Mil06, MY17, Mil88, MR86, MVBC18, Mir86, Mis19, MS97a, MS04a, MR04b, MW04, Mor79, MO99, Mor03, Mor10, MPS99, MS04b, Mos82a, Mot95, MS07b, MS10, MPR22, MS96, MT22, MLYF02, Mur89, Mur86, Nag85, Nag96, NR97, NK94, Nai87, NOV03, NS23, NSC19, NPR15, NWW17, NWW19, NMBM98, NP22, NYZZ17, NR17, NRL22, NSW17, NS86, NDHL24, NFI15, Niu21, Now08, NS14, OW00, Ohs13, ØS10a, ØS12a].

Control

[Ols02, OS12b, OS14, OS19, Orl04, OL77, OMFMC^{+}21, Oss01, OvS00, ÖT90, ÖB19, PNSL20, PS10a, PS17a, PB96, PKL13, Pan92, Pan98, PPZ13, Pan17, PT86, Pap89, PLSB14, PP14a, PHL16, Ped91, PT09a, Pen90, PW99, Pen20, PZLM24a, PZLM24b, Pen88, Per12, Pet04, PU15, PTZ21, PGB14, PW17, PRG16, Phu01, PDT19, PV13, PS99, PP02, POU11, PPTZ19, PZ20, PZ22, PQ00, Pla80, PS19, PT09b, PH92, PYM93, PW18, Pop95, Pop97, PZ13, PO00, PC00, PRHL08, PT06, PW78, PS87, PY96, Pro90, PV98, QW18, QWY21a, QW19, QV04, QR11, QV13, RRJ16, RM92, RT16, RSS18, RW87b, RMZ02, RV00, Rao98, RM98, RNK91, Rav20, RZ98, Reb88, Reb89, Reb77, Reb78, Red79, RSE20, Red18a, RR85, Rei86, Rei95, RCS01]. Control [RS15a, RV19, RZ20, RSZ23, RLXS23, RS21, RR13, RCA12, Ric77, Rif00, Rif02, Rin09, RHCK24, Ris86, Ris94, RV87, Rob81, RS15b, RW01, Roc87, Roc89, RW90, RTF21, Rog07, RL06, RT03, RT06, RT07, Ros84a, RK89, RS07b, Rot96, Rou21, Rub95, Rus86a, Rus88, Rya98, Sak02, Sak84, SDV97, SS99a, SZ19, Sap19, Sav05, Sca19, Sch89, Sch92a, Sch15a, Sch09, SW78, Sch80a, Sch80b, Sch86, Sch15b, SU21, SP96, Sei78, Sei88, SL98, Ser03, SRM16, SK86, SL18, SS20, Sha17, SX19, ST23, Sha85, Sha88b, SK95, SSO13, She85, Shi88, SZSC21, SSK02, SB78, Shr81, SR97, SV97, Sim76, Sîr15, SS24, SI03, Sle78, Smi90, Sok87, Sok88, Son86b, SS89, SY24a, STW20, Son23, Sor96, ST09b, Sta98b, SC81, SV79]. Control [Sta15, SZ96, Ste07, Ste23, Ste19, Ste22, Sto91a, ST90, Sto91b, Sto92, Sub81, SLY16, SX23, SY24b, SW23b, SAMB14, Sus87b, SS18, Sut20, Świ20, Szn93, TP05, TCC^{+}15a, Tak97, TR02, Tal05, TL94, Tan98, Tan03, Tan15, TC13, Tar85, Teb11, Teg84, Tes92, TW94a, TL09, TD94, Tib85, TLPA00, TKXP12, Tim14, TF08, Tol16, Ton03, TGSM18, Tre90, TWZ16, TZZ18, TS95, Trö99, Tsi91, Tsi92, TWW16, Tud90, TO15, Uch83, UP99, UP02, Urq05, VS80, Van94, VGBS13a, VGBS13b, Vel97, VW08, VS18, VS20, Vig97, VJ77, VL78a, Vin80, VK81, Vin83a, Vin83b, VW90, Vin93, Vin05, Vin19, VP22b, VSB18, Wac07, Wac12, WW16, Wac16, Wac19, WOH18, WS20, WS92, Wan98, Wan00, WW02, WW03, Wan08, WZ12b, WX13, WWX13]. Control [WL17, WZ17c, WHS19, WFZ20, WZ21, Wan22, WLZ22, WeFC22, WL24, WF24, WZ94, Was79, Wec84, Wee92, Wee05, Wee07, WC96b, WYY17, WD07, Wei89, WCG14, WTJY21, Wen15, Whi79, Whi80a, Whi82, WKS86, WP76, WHB92, Wol12, Won09, WRGA14, Wri95, WT83, WY08, WCS12, WM12, WJE16, WMWL22, WHGY22, WJZ12, XA93, XB99, XZ20, XYZ21, YL10, Yan14, YZSS23, YZZ01, YYH^{+}16, YCRZ21, Yon93, Yon13, YZ97, YDZ98, YLL^{+}15, You13, You17, YADF09, YCLK13, YLYC15, Yu15b, YHJ^{+}18, YH21, YC24, YL12, YS17, Yük19, Yük20, Yük24, Zas00, Zas06, ZR23, Zei84, ZZ88, Zei94, ZCGT10, ZZ94, ZMK21, ZC96, Zha98, ZJ06, ZC06, ZLHZ12, Zha13, ZZ15, Zha15, ZZS16, Zha18, ZG19, ZDM20, ZL21, ZHB21, ZG21, ZW21, ZS23, ZH24, Zho84, ZW18, Zho23a, Zhu92, Zhu96, ZRS97, Zhu00, ZM14, ZTC16, ZLLC18, Zhu21]. Control [Zol81, Zua96, Zua98, dMCJ22, dlRT07, dlRHM16, dBMS10, dVC22, vRM98, vWCET23, van87, vDHV15, BG88, EJ89, GR85a, GR85b, GRC82, KS84a, KS85a, LM84, Pen89, Qua80, Shi90, Son86a, Tah92, SZ97]. Control-Affine [AZ07, ABK21, BCS20, CJT08, JP15, LS99b, MSBC08, NR17, JR06a]. Control-Coding [CKT20, CKT24]. Control-Constrained [BKM14, vDHV15]. Control-Constraints [Den22]. Control-Dependent [HT22, Tes92]. Control-Lyapunov [KT04, Rif00, Rif02]. Control-Semilinear [Gra82]. Control-State [AR00, CRZ00, MG20, RT03, RT06, RT07, dlRT07].

Control-Strategy [Yu15b]. Control-Theoretic [Fio20]. Controllability [Aey85, AL12, AB03, ABCO17, AK99a, All99, AM23, AC18b, AN21, AZ08, Arm86, AFO86, Ava19, AMI20, AL00, AE18, ARS14, AD17, BM11, BMS82, BLPP24, BLJ15, BT91, Bar21, Bar23a, BP20a, BK97a, BM99, Bea08, BLT15, BB19, BBGBO14, BP14b, Bia83, BS90, BS93, BHS19, BPR04, BV13b, BGMS78, BCG84, BFZ10, BDFCGB12, Bon84, Boo82, BPS21, BFM23, BMN19, BL05, BGM97, CFZ17, CL02, CFM19, CS18, CZ98, CCM14, Caz14, CIP18, Cer07, CRZ13, CMP15, CC18a, CM12, CT19, Cha09, CSB15, CSZZ17, CSFCL^{+}23, CMC81, Che23, CZL21, Che76, CMK15, CM17, CRR12, CMR17, CDZ94, CLM24, CABP17, CL86, CLT24, CSV24, CT04, CGW10, CGR10, CN19, CGW20, CM03b, CRV09, DAR15, Da 16, DM89, DL09, DG16]. Controllability [DKR08, DG23, DL19, DFCGBZ02, Duc20, DMR19, ET15, Eno94a, Fab92, FW92, Far08, Far13, FGB95, FCG06, FCGIP06, FCLZ16, Fra87a, FYZ07, FL17, FKW23, FI98, FM22, Gag16, GCL15, GB82a, GB82b, GZ21, Gev81, GG07, Gon87b, Gra08, GH15, Gro12, Grü98, GL11, Gue07, Gue14, Gug03, GR21b, HZ95, HPS07, HMT18, Her76, Her82, Her96, Hey76, Hir76, Hir86, Hir90, Ho90b, Ho93, Hu15, Ich07, JPP14, JL76, Jaf20, JS90, JS91, JL14, JNKV21, Jou00a, Jou00b, JS84, KC15, Kam76, Kap94, Kha01, Kha03, Kha07, KBDK03, KMM11, KC10, Kim90, KY07, Kob78, Koe20, KLM04, KY05, Kra09, KP94a, Lag78, Lag89, LT89, Le 19, Léa12, LM22, Leq13, LM97a, LR03, LW15b, LR16, LY19a, LCZ20, LWLZ20, LZT20, LY23, LR23a, LFL23, LL24, LG07]. Controllability [LZ12, Liu14a, LL15, LLM15, LBMH^{+}16, Lob70, Lob76, Lom12a, LV97, LPS12, LM08, Lü14, LZH^{+}16, LW22, Luk85, Mac94, Mah03, MRS04, MT78, Man81, MR15, Mar20, Mar77, MRR13, MRR16, MRRR19, MRV03, MY79, Maz17, MM23, MZ97, Mic01, MZ06, Mic09, Mol19, Mor98, MR18b, Mot00, Mun14, Mur87, Nai87, Nar82, Nar84, Ner21, Nic96, Nic97, Nic00, OT83, Oss01, ÖH14, PPZ13, Pan17, PS86, Per12, PB83, PD92, Pos17, Pri05, PZ03, RS07a, RJME09, RS20, Riv90, RW97b, RL78, Ros00c, RZ09, RZ93, Sac97, Sch92c, SB80, Sei77, SMDB97, ST16, Son23, Son83, Son88, SC80, SC81, SGM19a, SGM19b, Sus78, Sus83, Sus87a, SIC00, TZ09, Tao18, TL06, Teb11, TC14, Tie14, Tie20, Tri76, Tri77, Tri80, Tuc96, TW00, TW03, UY79, VZ00]. Controllability [VLP24, Wan08, WD14, WYY21, WW23, War76b, War78a, War78b, War83, War85, War88, War89, War19, Wec82, Wec00, WZ09, WCG14, Whi80a, WM10, Xia19, Yam12, Yao99, Zha99, Zha01b, Zha18, ZG19, Zho83, Zho84, Zho07, ZL07, de 98, dRSV09, vN04, van82, Kre85, Kum88]. Controllable [Bol80, Cor95, Gra81, HZ21, HO90a, Jam14, MPS99, O’B79, OT16, OOS15, SA96, Sha14, SB94, Sun81, TR02, Tie14, Tri78, WR87]. Controllable/Observable [Bol80]. Controlled [AMV07, ABFG^{+}93, AB10a, AB12, AB17b, ABBK20, BD15, BC05a, BK24a, Bor83, BK95, Bor00a, BG05, BET10, BQT08, BIRS20, BCT10, CK23a, CM04, Cha10, CH99, Cha05, CM17, CTT16, CP16, CO22, CDM23, DS89, DMS12, Fle82, GKY21, GKP18, GR88b, GL97c, Guo02, HMV05, HSV00, HY97b, HMA97, Ich16, IRS02, KP97, KMF19, KSS20, KHJ14, Kor20, KM96, Lac17, LY20b, MSS96, MY22, MZ20, MPR17, Mor11, NF13, PRHL06, PM16, Ram94, RS15a, SY09, WZ12b, WZ12c, XLL08, Yon10, Yük16, Zit14, Zwa88b]. Controller [ABL13, AN06, BHM99, BH13, BDP24, CW95, ELO22, FIR06, GS19, Gre92, Imm07, KK92, Kob88, LŻ11, LH88, MT23, PB99a, Pee20, PD02, PLK89, RRJ16, RSV04, SS99a, Ste07, Szn95, TZQ21, WCLC05, XZGT23, YCY13]. Controller-and-Stopper [BH13, ELO22]. Controllers [ANM22, BT11a, Bar12, Bar20, Bar21, Bar77, BCL83, BR17a, FCG16, FL20, GMCL10, GHJ00, GP82, Hai86, KGS11, Lei01, MVHN09, MA17, NS00, PCP22, Pau17, PTY07, Qua04, RY95, RSMG17, SS20, ZY24]. Controlling [BKSW24, CKRT22, CBB19, Eva13, Wee18]. Controls [Ahm96, Ahm03, AC13, AC17, ABCO17, AC05, ACCG18, AN21, ABR00, Art78, AT82, Art89b, ARS14, BMRR20, Bal94, BSB^{+}97, Bar83, BL01, BS20a, BY13, Ber89a, BCS20, BGBPG04, BPV19a, BGP23, BR21, Bre85, Bre86, BLQR20, CL19a, CT14, CK17a, CK19, CT20, CK21, CSES21, CCW14, CMT11, CV90, CG22, CMN20, CW24b, Cos13, CR11b, Dág06, DI94, Di 91, DF94, Dus21, Fat90, Fel03, FCGO05, FH17b, Fu23, FT04, GRR09, GV03, GSW89, GS94, GJZ10, Gon87b, GWY14, GM13, GM99, Hau82, HL90, HS95b, Her91, HY97a, HY97b, Joh05, Kha01, KZ00b, KTV16, KS98b, KS99b, KR87b, KR88, Kus14, Lag83b, LT89, Las95, LXY16, LR16, LR23a, Lis14, LNP00, LT13, LY09, LW11b, LtLR21, Mil06, MV07b, MR96, MS10, NN90, NPS15, Nar82]. Controls [Nar84, Ngu22, Nic97, Now17, QWY21b, Ram98, RZ98, Reb78, RZ21, Res76, Ros92a, Ros00b, RS93, Sap19, Sar97, SB80, SZ82, Smi96, ST16, Tad88, VL78b, WD79, Wan02, WXY05, WZ12c, WXZ15, WY16, WYY21, Wan21, WYZ23, War77, War85, War91, Wee03, WY21, WP76, Yu14, ZZ15, ZZ17, Zho92, Zho93, Zho98, Zuy16, FCGIP06, HS95a, Sei82]. Controls/Observations [Las95]. Convection [CGS18, Liu23, WD14]. Convective [HLP08, RS15b, SS20]. Convergence [ATV13, AGL08, AM89, And14, AB09, ABK14, ABL18, Bal82, Bar80a, BE13, BBCS22, BP04, BP06, BKS10, BR82, BV13a, BB03, Bin76, BTW82, BM00c, BKK02, BC24, BK23, ByS17, BR01, BK02, CFN07, CMA08b, CDJS23, CDFP19, Che96, Che98, CC88, CL94, Cob10, Com00, DK05, DM99, DHV00, DV12, DMZ22, Dun79, Dun80a, Dun81, Dun00, DJ98, DS01, EHS07, EB20, FK91, Fat90, FY76, FMSV16, FK21, GP22, Ger92, Gib81, GRZ24, GLO19, Gül91, GY13, GZ13, HT11, HH76, HZ07, HLM^{+}18, HHF21, Heu94, IK07a, JS99, Jay10, JALG18, KP97, Kam10, KSS20, Koe79, KS15, Kus78b, KH79c, KS84b, KY93, KM96, Lei85, LZ16a, LX23, LZ21, LL94, LT91, LT92, Luq84, Mal23, MT94, MG20, MG21, MSS96, MBS13, MCS13, McE09, MK10]. Convergence [Mey77a, Mey77d, Mik12a, MB12, OT09, PY16, PG18a, PG18b, PV98, Qu14, RSZ23, RRS14, Sac85, SR04, SY09, SYZ13, Sor98, SR18, Sva11, TZZ22, TYW11, TG12, UUH99, UU00, VS18, VW02, VJBK24, WW16, WS13a, WP76, WHB92, Wri89, Xie12, YD03, YZ97, Yu15a, YHH16, YL12, ZKZB20, ZHB21, Zor17, DK89b, CKL16]. Convergences [Cou00, LR92]. Convergent [Bas91, BR07, CLRJ13, CS77, GHX86, KK03, PT87, PTH88, PLSB14, PQ98, QC95, Yak93, ZRS97]. Converges [BC20]. Converging [GL76]. Converse [EUMP18, EP24, HS11, HBY21, KT03a, KT04, LSW96, LE17, MK21, Wir05]. Convertible [SPS04, SS06]. Convex [AM89, ABB04, AAA12, AFO86, Bar78, Bar80b, BT96, BE87, BHT87, BF22, BP17, BCG^{+}17b, Bli04, BM13, BGL98, CK95, CC13, CSBC14, Che96, CLSL20, CB11, Com00, Cou00, Den97, Den98, Den22, DG20, EC22, Enq04, FV89, GBCG98, GM91a, GPB91, GPS95, GHZ98, Goe05, Goe13, GT90, Gül91, HM76, Hal89, HL88, HJ15, HLT08, HHF21, HW19, Heu15, HU82, HU84, HKL89, HLZ20, JHMG06, JG99, Kiw89, Kiw96a, Kiw96b, KG17, KHJ14, Kor20, KB76, KLN21, Kry98, Lei85, LP79a, LP81, LHS06, Li95, LP87, LLT18, LT92, MNC16, MW78, MOF77, Mor24a, MNS91, MS96, Nie21, O’N76, OC23, Pee20, PP14b, PZ11, PA93, PR07, Ram05, RZ14, RW76, Roc89, Rou21, SY21, Sch15a, SE21, SD13, Tan88, Tha86, Thi97, Tse90, Tse91]. Convex [Vin93, WS13a, WZHZ19, Wri95, YHH16, YS17, ZZ15, Zhu92, CL78, PV04]. convexe [CL78]. Convexity [Bal94, Cel07, DK78, FBJM14, Goe02, Hin14, Kno77, LN10, LY12, RW00a, RW00b, SPJ15, Tar90, WR14, Zas00]. Convolution [GMM91, Mos80]. Convolutional [FNP^{+}22, Hut09, VZ00]. Cooling [SSL11]. Cooperation [CNS21]. Cooperative [BCC^{+}19a, BCP09, Eng10, GL22a, GK02, LO15, YGC18, YXJ11, ZSS^{+}20]. Coordinate [BA08, HG90, MA01]. Coordinate-Free [MA01]. Coordinated [GAP^{+}09, HPS03]. Coordinates [BAP18]. Coordination [CB09, CAS13, CB05, CC90b, DJ13, GL06, HPT07, HM09]. Copepod [BBR18]. Copositivity [Pre96]. Coprime [Gre92, Mik07, OS12b, OS14, Qua03a, Sta98a]. Corank [BBG12, BG13]. Corrected [GŠZ05, GŠZ10]. Correction [AB17b, Bor07, KH69, Wei77]. Corrections [Cur13, DF98a, JLZ98]. Corrector [SZZ97]. Correlated [Ave19, Kwo91, VJBK24, WWX13]. Correlation [AKTW24, DS87, FPW16]. Correspondence [Bar77]. Corrigendum [BT10, Flå95, FL82, GRC82, JMPW23, Sul98]. Corrupted [GG13]. Coset [CDM90]. Cost [ABB01, AM23, And88, AMV20, ABFG^{+}93, Bar80b, BP14a, BD02, BBCS22, BBGBO14, Ber98c, BK24b, BM84, Bor84, Bor89, Bor00a, BM18b, Cam76, CK98, CW20, CDR08, CC96, CM81, CG87b, CL19b, CGW20, HRCK21, HSZ15, HLL96, HLVAC99, HLR01, HLGHLM03, HG02, IK02b, IK14, KKS93, KPZ18, KV07b, Kus78a, Kus89, LT83a, Lis14, ŁS85b, MWY21, Mil06, Ngu22, PS17a, PB96, PB99a, Pen20, PQ00, PC00, Rob81, SBR18, Sap19, Sla00, Sta78, Ste22, Stu16, THA02, Tre89, TCC19, Wac19, Wee03, XYZ21, Yu15a, Yu20a, Yu20b, GR85a, GR85b]. Cost-Biased [CK98, PC00]. Cost-Effective [CL19b]. Costate [BV10, VW95]. Costs [AMS96, ARS17, ABK14, BP10, BZ15b, BMS15, BC12, BHT87, BS17, BS19a, Cal09, CLZ98, CZ00, CSZ13, DJLY09, DPZ93, HLL94, HCM05, Kam03, LMKW17, LZ13, NA02, OS02, Pau08, PZ22, Ric77, Roc89, ST13, SY09, Tse90, XYZ21, Yu20b, Bla97]. Coulomb [BHOP14, Lip97]. Countable [HW05, NY16, NY18, XZ17, Yu20b]. Counterdiscounted [dVC22]. Counterexample [Fre99]. Counterexamples [BBV10, JZ04, Wyn09]. Counterparty [HL16]. Counting [AH14]. Coupled [AB02a, AB03, AA17, ABAC10, BD19b, CC02, CG17b, Cza04b, DXG15, DC06, Fu12, Ho93, HR21, HJX18, HCM05, KB16, KL99, KT16, LW14, LR16, tLR23, LR23a, LFM07, LZ19, Liu88, Liu14a, LLSX20, LtLR21, MZ18, OPPU17, PW99, QWY21b, Reb95, Sta15, SGM19a, SGM19b, Teb11, WYY21, WZ09, XG03, XZ20, YCY13, Zei92, ZHSP14, Zha18]. Coupling [CLLP13, CGR10, Fuh76, KB16, Le 19, LH11, SL13, ZL24]. Couplings [BAKS18, HP09a, KHP06, Kim21, LM22]. Cournot [AVJ21]. Covariance [BGL98, Enq04, Gen13, GGA^{+}17, KLN21, RKL16, RKL18, ZRS97, ZZ23]. Cover [PPZ06]. Coverage [BCF12, FGGH15, LL09, ZM13]. Covered [BLZ17]. Covering [RN13]. Cox [BPS05]. Crack [BB07, KL16]. Crack-Defect [KL16]. Cracks [AD96]. Crank [MV11, vDHV15]. Credit [BC18b]. Credit-Risk [BC18b]. Criteria [AL01, Bla76, DHX10, Dai12, DM97, FP19, GST20, HLL00, JL76, JN19b, Kre78, LR16, LSW^{+}17, LW24, MK21, MS97b, NLPT09, Vid77, Zei92, ZP03b, ZP05, ZD14]. Criterion [ABFG^{+}93, Bar80b, Bis77a, Bor89, CC96, CCHH19, CP13a, DZ76, DG20, FK83, FD87, GGMZRB20, HSZ15, HR83a, Küe07, Kum08, Mos82b, PB99a, SL22, XYZ21, Zei84, ZH05, HR83b, Mos82a]. Critic [KB99, KT03d]. Critic-Type [KB99]. Critical [AH14, Cer07, Kre00, MGMBN^{+}19]. Criticality [Bar17]. Crocco [MRV03]. Crocco-Type [MRV03]. Cross [BL14, CFNG10b, OVS08, SO06a]. Cross-Talking [OVS08, SO06a]. Crouch [Kup85]. Crowd [AD18, BFF20]. Crowds [TWB23]. CRRA [TZ10]. Crystallization [BASG21]. Crystals [SW23b]. CSVIU [dVC22]. CTL* [JK06]. Cubic [CGR10, CG17b]. Cuboidal [HLY24, YH21]. Cucker [LWR24, DHK20, DHJK20, SS21]. Cucker-Smale [LWR24]. curl [LWY21]. Current [GK06]. Currents [CFRV98]. Curry [Phe86]. Curse [McE07, McE09, MK10, Qu14]. Curse-of-Dimensionality [Qu14]. Curse-of-Dimensionality-Free [McE07, MK10, McE09]. Curvature [ATW93, Cha91, GL06, Liu11, MR15, SC06]. Curvature-Driven [ATW93]. Curve [DJ86, MZ15]. Curved [AST05, CS10a, CGM11, DGH09]. Curves [DIL15, ER88, HT92, KN13, MN08]. Cusp [GS19]. Cutting [Jer78, Jer80, LR79b]. Cutting-Plane [Jer80]. Cutting-Planes [Jer78]. CVaR [GZLC17]. CVS [SD13, SD15]. Cycle [Jön09, Tro96]. Cycles [Joh05, JM05, MMK86]. Cyclic [Bia83]. Cyclically [BLW07]. Cyclicity [Sch15b]. Cylindrical [BBGBO14, Rus86a].

D

[BP97, BRS11, BE13, CC16, CM12, CGW10, Dág06, FRZ93, Gue14, GGMZRB20, Hay19, KMC21, KK17b, KR15, LY23, Mic09, MP21, MT22, Ner21, PP13, RW01, SCK10, TL06, VZ09, XZ20, ZM11, Zam98]. DAE [KM07b]. Damped [Buc92, Bur20, CC01, FM08, GR21b, KMC21, Kle18, LL21a, LR06, MRR13, SMDB97]. Damping [AHR23, CO03a, CSC14, CSCV18, CCFV20, DDK10, Fre99, Fu12, GGH16, HLY24, Hua88, KMC21, LYLZ99, Liu97a, LL98, LZ16c, LZ18b, Nai95, Nie21, Ovs14, Sak84, SW23a, Sto91a, WC91, WXY05, YH21, ZC91]. dampings [XZ20]. dans [Tah92]. Data [Aus84, BP88a, BB97, BT21, BD01b, CB11, Enq04, EH10, FSGHH20, FS16, FA13, GN03, GS18b, Hag02, Hag05, HIK18, Heu94, Izm96, KK16, KLR09, Kob83, Kob84, Kor20, LLX11, LM02, LRT05, Log13, LSL^{+}18, MRP99a, MRP99b, MP21, NE04, NOV03, NLPT09, PU15, Pue09, SK94, Sta79, TWZ16, TS95, VW90, WOH18, Wak21, WF24, Yam12, YYH^{+}16, YC24, ZZ13, vWCET23]. Data-Based [vWCET23]. Datum [Ngu22]. Davis [Les12]. Dead [BV13b, NMBM98]. Dead-Beat [BV13b]. Deadbeat [CS84, Kim78]. Deadline [DY13]. Dealing [CF95, Lin14]. Debt [Fer18, FR20]. Debt-to-GDP [FR20]. Decay [ACS00a, AB09, CC01, CSC14, DDK10, FCJ14, Fre99, Fu12, GLX18, HZ22, Kim91, Kle18, LL21a, Liu88, LL98, LR06, MCP21, Sti18, Teb22]. Decays [Bur20]. Decentralization [BD19b]. Decentralized [CKT24, FLJ13, FL20, KS97b, Ma09, MZ18, OvS00, SSK02, Wan94, YS17, Yük20]. Decidable [BO14]. Decide [BO14, Son88]. Decision [ABB01, Alm08, AS91, AG97a, BS24, BM18b, Can84, CC96, CKT24, Cha14, CW23b, CW24a, CHJ15, CL06b, CP13a, DR06, DG16, DPR13, DG20, FKZ22, GHS12, GHZ15, GL19, HSA17, HJ13, HJ15, HW05, HP14, JV06, JN19b, KA06, Koe79, Koe90, KB99, Kre78, Kur89, LWDB20, PZ11, Piu15, Piu18, PZ22, PZ24, Rod82, Rot77a, SZ95, Sza95, VZ16, VZ21, Yu15a, Yu20a, Yu20b, Yus97, Zha17b]. Decision-Making [Cha14]. Decisions [Alp88, BP12, CSA^{+}13, KLS87, Sag13, Zer03]. Declustering [PDT19]. Decomposable [Laz88, YD03]. Decomposed [Ha88]. Decomposing [BP08b]. Decomposition [BCG86, CM90, CM18, CC90b, DGQ13, Dun91, Dun93, FR17, GFF17, JN19a, JH14, Kre77a, KLM^{+}21, LL00a, LF87, Leu99, LB89, MSv86, MNS91, O’N76, Ste92, Ter94, Tse91, Wim94, Zab78, Zab80, ZL02]. Decomposition/Coordination [CC90b]. Decompositions [KR80, Li07, Tro96]. Deconvolution [CF06, FP02]. Découplage [Cla86]. Decoupling [AF18, BKP90, CM00, CP95, DH84, DM89, DC88, EÖ88, FA89, GNP18, Hir81, HNv92, Hui92, II97, KS15, PNSL20, Tre86, WC81, Cla86]. Deep [BK24a]. Default [BLY19]. Defect [KL16]. Deficiency [GGA^{+}17]. Deficient [CL24]. Definable [BR15a]. Defined [CM16, CP76, Fer78, Ols76c]. Definite [Fle85, TR01]. Definiteness [AI02]. Definition [JALG18]. Definitions [Lew80]. Deflating [Ste23]. Deformations [GJ20]. Degeneracy [AA97, CFM19, Teb22]. Degenerate [ABCL17, AAFC19, BC05a, BFM23, BT03b, BT10, CF24, CMV08, Che98, DDK12, El 17, GLM21, GLZ04, GT20, Gue14, HLY24, HLT08, HL81, Ho86, HQZ16, HCM05, LY23, LY19b, Lou03, MCP21, Men80a, Men80b, OW00, RV00, Ste19, WD14, YZ23, Qua80]. Degenerating [CSB15]. dégénérée [Qua80]. Degradation [CZA04a]. Degree [BDIW22, BD19b, Gha11, HHI^{+}13, IRT07, KB81, LH88, MH78, MZ22b, Tha86]. Degrees [NCL13, Zaf03]. Delay [Alt09, ANM22, ALR99, CFN07, Car90, CMT18, CFNG10b, CLS21, CW91, DFŚ24, DKR08, DFP14, DC88, DTM22, EPR14, FHMS18, Fla90, Fre00, FIM09, FNV10, FKW23, FMT10, FMV24, GVL12, GLK12, GL97b, GM17a, GM17b, GM21, GTQ21, GM11, GMY16, Hab96, HLZF21, HZB13, HLDM20, HHT19, JLL20, Kar06, KK16, KPCW22, Kun82, KW78, LMMY20, LLLT18, Lou15, MCL^{+}22, MP99, Mam14, Mat05, MNB11, MNB12, MSR01, MN07, MVZ^{+}09, MZ22b, MSv86, NNT22, NS00, Ngo09, NBF16, NP06, NDHL24, PS10a, PPL09, Pla21, Pos17, RG14, RSMG17, RBS21, RX19, RW97b, Ros84b, Ros92a, SS20, SWW22, SFFC08, SO06a, SO06b, Tad97, TT20, VZ00, WZD^{+}23, WL19, WL21, WZ77, YC06, YÖ08, YW12, ZP05, ZL24]. Delay-Adaptive [WL21]. Delay-Based [RSMG17]. Delay-Differential [GL97b, RW97b, VZ00, WZ77]. Delay-Independent [MN07]. Delay-induced [Fre00]. Delay-Integral-Quadratic [Alt09]. Delayed [Bag84, BCY15, CFN06, DHKS18, DZX24, JLL20, KB16, Kar11, KKA23, LH14, LZH^{+}16, MP14, NR97, NYZZ17, Ros00b, SZ19, WL19, XZ20]. Delays [AE17, AHS13, BBC81, BL83, Bat19, BCLY17, BFPB13, BM84, BV17b, BHT19, BFP11, BIP88, CGM14, CMA08b, DLP86, Dat88, Del86, EOF07, FGG10, FGG11, FT14, FCJ14, FTZ87, GAP^{+}09, Gli12, GFF17, Had06, HRW88, Ich82, KvdSD21, Kwo80, LZ16b, Lin14, LLX11, LH11, LFH15, LLJC18, LRW96, ML13, MZ15, MT94, MMN17, MMÖ18, MERD02, MMN09, MZ22b, Mil88, NTP18, NT16, Ngu18, OVS08, Ouv78a, Pan83, QS13, RT98, SO06a, VK81, Vin19, WOH18, WGK11, XZGT23, YCY13, ZLD10, ZK20]. Demand [FFR23, FSS87, HYZ24, LYB18, PYB20, SZ10]. Demands [BB09, BLS05a, SPBF10]. Dennis [CDG15]. Densities [Bis78b, CL24, HLT08, HKW17, KSY23, Nur06]. Density [BP08a, BS03, CMS16, DHP19, HO01, HW12, MK21, RHCK24, Zhu94, vW84]. Dependence [BKL^{+}22, BV17a, DE14, LS99b, Ngu22, OT94, PT09a, PS06, RW92, SV79, Yük24]. Dependencies [BC18b]. Dependency [MVZ^{+}09]. Dependent [AT94, BK22a, BHOP14, BGP20b, BHT10, BJ10, BHHE03, BHT21, CG17a, DFŚ24, Dun00, EMP18, Ete19, Ete20, FCG18, FNV10, GXZ22, GdOB02, GGH16, GK85, GHS12, Heu94, HK01, HT22, KY07, KK06, KS84b, Kus13, Kus14, Lam92, LLC18, LLLT18, NY16, NNN21, PZ14, RTZ20, RX21, RTTY23, Sap19, Sha18, SX19, SYZ13, Tes92, VT95, WYZ23, YD08, ZBB07, VZ21, Zha17a]. Depending [CP98, FT12, GT22a, Mai03, Riv90]. Deployment [CB09]. Depth [ZVBR14]. Derivation [Dav78, MT78]. Derivative [AGJK12, Bou17, CDM21a, CC90a, DDS15, Del22, Del23, EI13, GS22, HU82, HU84, HN96, MVHN09, PS10b, RU19, Sla00, Sla03, SZ99, TD21, Toa11]. Derivatives [BFL16, CNS10, CHN94, CDD20, Del23, LCC17, LS85a, LL15, Pan13, RU21, LS86]. Descent [ATV13, Bar17, BK24a, DIL15, LT92, War77]. Described [BP94, Dol79, Gri83, Kan82, LL18b, PEGR18, VZ00]. Describing [Tim14]. Description [CdLdDM02, Dew76]. Descriptions [MCWY14]. Descriptor [BGM97, BKM97, CM00, CABP17, FMP^{+}22, FA89, HZ21, KS90b, KM04, LM08, LMPR09, Mur87, MV91, NLW91, RV09, IM05b]. Desensitizing [Teb11]. Design [AMP10, AGN^{+}16, AT92, AT94, APA08, AN06, ANM22, AB05, BA08, BSB^{+}97, BL14, BF91, Bar77, Bat07, BT97, BF07, BF89b, BGH94, CZ01, CHL15, CH12, CdSBT09, EM19, EUMP18, Enq04, FCG18, GCB06, GPB91, GLK22, Gho88, GB93, GW17, GS14, Grü09, GP02, HM03, HMA81, HG02, HL11, HS03, HJF10, HHT19, Jam91, KK19b, KP04, KX02, KX04, KV13, LG09, Las95, Lau99, LR10, LHX12, LW15b, LL18a, LMZ20, LŻ11, Liu88, LR92, MMN17, MRP99a, MRP99b, MPS99, MP07, NT95, NP96, NT16, PB99a, PTT17, PTZ17, QN82, RRJ16, Sag13, Sch83, Ste23, Szn95, TBP14, UMD88, VG95, VL23, VP22a, Vrd16, WSZ^{+}21, XG89, ZD24, dOG05, Gho86b]. Design-Dependent [AT94]. Design-Independent [AT92]. Designing [LMJ09, Lei01]. Designs [GS19]. Destabilized [KGBS08]. d’état [Loi88]. Detect [ELO22]. Detectability [ART21, AM81, CdV02, CdVF05, TSS^{+}22, WOH18]. Detecting [ET12, FFP^{+}09]. Detection [BP07a, BL15, EMP24, FLRO17, FV98, MS07a, VC03, ZHSP14]. Detector [RTF21]. Detector-Based [RTF21]. Determinacy [MSH91]. Determination [AD96, BW79, CFRV98, FT77, Kob76, Kob77, Kob79, Kob83, SO06b, XC18]. Determined [MR18b]. Determining [BP97, ZC91]. Deterministic [AGL08, Alm01, AB02b, Bar85b, Bar80c, BM99, BB93, BL18, Bla97, Cha87b, Cob12, CD03, CD08, CD10, CDP16, DA17, DC99, DS01, FP19, FGL08, FEBL13, GR06, GQ09, GW17, GV16, Gor15, GHS22, GAHL09, GGA^{+}17, HM08, Kre87, LRT96, Mah03, Nih83, Ols76b, Red18a, RX21, RW90, SS21, Zhu92, CL06a, GR85a, GR85b]. Development [EPR14, Win78]. Deviation [GSZ20, RBA22, WYZ17]. Deviation-Risk [GSZ20]. Deviations [BHHE03, DK85, DK86, DK89a, HYZ11, PZ04, DK89b]. Devices [LR10]. Diagonal [AP92]. Diagonally [Dun86, IK07a]. Dichotomy [ABCD22, BH18, Mor24b]. Dielectric [BDR02]. Diet [HBSW07]. Diffeomorphic [TY00]. Diffeomorphism [BAP18]. Difference [BKK02, BO09, CE11, De 92, DH22, EGR^{+}22, GLK12, KT05, LY90, LG20, MR78, Maz17, MML21, Men89, Ros84a, Tol16, Yu13]. Differences [CMR17, DNS17]. Different [ASR^{+}13, KK17b, WX13, WZ94, Zha15]. Differentiability [BFCLS97, CW22, DZ88, FLUU17, KO18, MP94, Pen84, Pen88, Sha88a, Stu15, War76b, War83, Ye00b]. Differentiable [Aey81, DG85, Ish88, Kre77a, PMW83]. Differential [Ach08, Ahm77, AMV07, AVZ20, Aub90a, Aub01, Ave19, BD15, BFR21, BHM99, BK82, BL83, BTY11, BRZ20, Bar20, Bar80c, Bas77, BL89, BP05, BY13, BKR00, BHS05, BCS15, BR17a, Ber86, BBV13, BH00, BD19a, Bis12, BR13, BG05, BHT22, BO09, Bre16, BR93, BS04, BS09b, Bro22, BS13b, BCR04, BQT08, BL08a, BHL12, BLP14, BM18b, CN11, CQ15, Car96, CP00, CQSP00, Car07, CS12a, CGS00, CW23a, CO03b, CPY09, CMC81, CM81, CR21, CY24, CDKV18, CLS97, CTW18, CH78, Col82a, Col82b, CF13, Cos13, DFŚ24, Del07, DS09, Dmi92, DE84, DF21, DF98b, DS99b, DTM22, DM15, DMS12, DA17, DRDG08, Ell76, ED81, Eng10, FGG10, FGG11, FHH22, FR82b, Fit79, FK83, Flo97, For77, Fra87b, FKW23, FT04, Gar76]. Differential [GT77, GR17b, Gev81, GLK22, GD88, GR05, Gib83, GW76, GL97b, Goe02, GYZ22, GR88b, Gra05, GS11, GM11, GMY16, HG90, Ham06, HBSW07, HJM07, Hei16, HHS18, HSW12, HSW15, HH08, HH10, Hop87, HP09b, HLDM20, Ich76, IRS02, Iof24, IT86, IT87, JB96, JLR09, Kan82, KK19a, KH90, KFP22, Kim93a, KS00, KZ00b, Kra93, Kum08, Kus79, Kus02, Kus04, Kus07, LT86, LZM05, LY90, Lei85, LXY16, Li07, LW12, LM16, LZ18a, LY18, LL18b, Lia15, LB00, LZ01, LW11a, LS85a, LS86, LZ19, LNP00, Liu97c, Liu14a, LG24, LR94, Lom18, Lü15, Luk85, Man04, Mao90, MG20, MG21, MGMBN^{+}19, MNC14, MNPR20, MWY21, MSR01, Moo21, Mor95, MW04, MA03, MN77a, MN77b, NN90, Nag95, NS05, Ngo11, NT17, NT18, Ngu16, ØS10a, Ols02]. Differential [PS17b, PD97, PW99, PZLM24a, PZLM24b, Pen84, PA99, PQ00, Pos17, Qia94, QW19, Qui92, RMZ02, Ram98, RT04, RV19, RW97b, RRS14, RS13, Rya98, Sal85, Sap19, SC16b, Ser09, Sha10, Sîr14b, Sor96, Sor98, SQS09, ST09b, SW21, SV79, Sti18, SY14, Sun21, SWW23, Tal91, Tan98, ThH07, Uch78, Uch79, Und76, UY79, Var76, VC03, VZ00, WW16, WS92, WL95, WS95, Wan00, WW02, Wan22, WT98, WZ77, Wil77b, Wol90, Wu18, YT13, YXJ11, Yon88a, Yon88b, Yon90, Yon02, Yon10, Yon13, You13, Yu15b, ZVBR14, Zha05, Zha17a, ZD14, Zho92, Zho93, Zho23a, aJM18, vRM98, BS18, Col85, Iof83, Pen88]. Differential-Algebraic [BR17a, Bro22, CR21, Hei16, MG20, MG21, RV19]. Differential-Difference [BO09]. Differential-Geometric [VC03]. Differentials [BM84]. Diffusing [CGP22]. Diffusion [Ahm07, AC13, AC17, AS88, AF18, AKTW24, AKP21, BP10, Ban78a, Ban78b, BN81, BB89, Bar21, BP20a, BDR94, BA15, BGR07, BB09, BLS05a, BD17a, Bis12, BDGR12, BKR17, Bud03, CLZH24, Cer01a, CSB15, CHS87, CTZ03, CTT16, DG14, DJK^{+}23, DGR18, DL13a, ET12, Fra03, Gad08, GH14, GV05, GW09, HL92, Hau81b, Hau81a, HLG21, Jaf17, JZ21, KK19a, KEK21, KMF19, Kus78a, Kus79, Le 19, LMMY20, LYB18, Liu23, McN84, MML21, MN22, Moo21, Mor11, NS23, NP82, NY16, NY18, NNN21, NF13, OW00, OPPU17, Osh06, Pau08, Ric77, Roy89, Rub95, Sak83, SS99a, SSW15, SX14a, Sha15, Wee05, Wee18, XZ17, YCY13, Zha98, ZES12, ZSX18, ZG18, ZY07, ZWYJ14, dOP03, Qua80]. Diffusions [AØ23, Alv01, AL08, AMR14, AMV20, AB10a, AB12, ABK14, AB17b, ACPZ19, ABBK20, AB20, BDM07, BC05a, Bay09b, BX12, BEM13, BS20b, Bis82, BB04, BG07, Bou08, BIRS20, CDP23, CWZZ17, CH99, CG13, DS87, DS89, DGW10, De 15, Ega08, Fle82, FP82, FV89, FT02, GXZ22, GS94, GH86, GAM93, GAM97, Hau82, Hau87, HO08, KSS20, Ma92, Men80a, Men80b, Nag85, Sha18, SWW22, SLG84, SY09, YX10, ZY09, Zhu21, BG88, ML12]. Diffusive [CGM98, CW08, War19]. Diffusive-Wave [CW08]. Digital [CFZ10, DCI07, Kob88, ZCC07]. Digraphs [MNC16]. Dilation [Lou94]. Dilemma [GBB18]. Dimension [BF07, BHL10, Bre86, CM10a, CYL96, CYL97b, CRR12, CDM21b, Dok17, FR17, HMR15, pJY20, Lis14, LNT02, Lom21, MARRH23, Mor24b, PY16, Sch88a, SY17, TY00, Wec82, WY06]. Dimensional [AZ07, Ahm96, AS88, Aih92, AMV20, AAFC19, AL02, Aug19, AK19, BT11a, BMRR20, BC96a, BMS05, BW89, BI91, BD83, Bar95b, BL01, BM90, Bas91, BSDR97, Bea08, BM09, BLT15, BBGBO14, BE95, BH86, BF90b, BV13b, BT13, BASG21, BL00a, BQT08, CT96, CHJ87, CK19, CK21, CZ04a, CC18b, CYL96, CYL97a, CYL97b, CW08, CY94, Coh97a, Coh97b, Cor99, CT04, CBdN08, CB15, CN19, CP76, Cur84, CS86, Da 87, Dág18, DE18, DMK90, De 15, DGH09, DRT11, DD03, DTWY91, DWZ^{+}16, DKS^{+}11, EK97, Eva13, Fab08, Fat94, Fat97, FFRR21, Fei76, FR78a, FG15, FCGIP06, FRZ93, FV95, FNR19, FGGH15, Fuh76, FT02, FGH98, FGH05, Gag16, GJP88, GZ21, GBR18, Gib80, Gib83, GR88a, Gla05, GH12, GCP88, GWY14, GT90, Grü09]. Dimensional [GKLM92, GT14, GLO19, GH96, GG13, GR21b, HH19, HM03, HP10, HRW88, Han94, HPS07, Ho90b, HVDK19, IST16, IP06a, Imm07, IP87, Ito90, JZ99, JNPS18, JMPW20, JMPW23, JB96, pJY20, Kam76, KK19b, KZ23, KY87a, KLR09, KB01, Kim91, Kim21, Kno76, Kno77, KS98a, KRT11, Kru00a, Kru00b, LP15, LH21, LY90, LS05b, Lev87, LV16, LPT22, LG20, LZ92, LI94b, LRT98, LR00b, LM02, LRT05, Log13, MZ15, MCS13, MS98, Mir21, MN08, Nag85, NZ99, Nar84, Nih83, OMS90, OW93, OW96, OW00, OC04, Opm07, OS10b, PEGR18, Pan92, PZ04, Pau17, PBPB15, PGB14, PS87, RA16, Ray06, Sac19, Sal85, SS99a, SWZ24, SY17, SS89, Sta93, Tak97, Tan00, TZK11, Teb22, Tud90, UU00, Van94, VB23, VLP24, VS20, WS20, WC89, Wan02]. Dimensional [WF24, WR00, WM10, WORO00, XWPG23, XB99, XS96, Yam88, Yam89, Yan95, Yin99, Yon93, Zha18, ZZ20, ZW18, Zwa88a, dOP03, CCM14, Ega08, FR82a, Fen16, MRR16, Nto12, UUH99]. Dimensionality [McE07, MK10, Qu14, TWY90, McE09]. Dimensions [AB06, AGN^{+}16, Bar91, CD91, CC04, Câr93, Cur86, FRS16, GW76, Gro12, GT05, GT22b, KS00, KS89b, KOS04, MMR17b, MMR17a, MS97b, RS15b, WS95, WWS21, WB87]. Diophantine [WA84]. Dirac [LZM05, NMI15]. Direct [APH24, DA02, Del22, Ega08, FK21, HK78, LF87, MP76b, O’M76, Sch83, VC03]. Directed [GH20, KCP16, MCL^{+}22, RT16, YLYC15, YHJ^{+}18, ZZ13]. Directional [BC96b, BC96c, CW22, DBC12, HSW12, HSW15, HU82, HU84, HN96, KPV14, LS85a, War89, LS86]. Directions [BDP24, LM97b, Mey83, Pav80, PV98, RT16, YL10]. Dirichlet [Ama04, AMPR15, ABSP08, BS18, Buc92, BN16, CT96, CR06, CGM11, Cel02a, CRV09, DGH09, FG09, FT02, GY11, GMS22, Gug15, GI02, GZ05, GZ09, HK13, HSW19, HK21, HR98, Idc98, KV07a, LT83a, LT83b, Las84, LT86, LT89, LR16, MRV13, Now08, Ped91, PS04, Rav20, Rus86a, Yan14, Zha09]. Disagreement [CF18, MMH19]. Discontinuities [ADL23]. Discontinuous [AX05, BGR07, BV17a, Bla97, BL11, CZ99, CMN20, Dav07, FMSV16, Gut90, Jam14, KW85, Lam87, Man04, Mar02, Nag85, PS10a, PQ00, Sta78, Ulb02, VP88, XYZ21, Ye00a]. Discount [CQ15, CK23a, CK23b, CCHH19, Ham21, LYZ15, SCFV97]. Discounted [CS87, CS89, CD12, CDP16, DR08, DPR13, FP19, GHZ15, GL19, HSZ15, HHSS23, HLL96, HLR01, JN24, Kum08, Kus89, LY20a, MSVA09, PZ11, SBR18, Wee03, XYZ21, dVC22]. Discounted-Cost [HLL96, HLR01]. Discounting [GQ09, JS24]. Discovering [SX14b]. Discrete [AD15, AS93, AD98, ABB12, AMV07, AM81, ABCD22, ABKM96, ABFG^{+}93, BCM^{+}17, BBC^{+}19, BCS21a, BCS21b, BKP90, BS86, BLJ15, BE13, BBT14, BR82, BGS19, BMO15, BHL10, BFO20, CGR96, CLYL13, CLW17, CW23b, Chr83, CL06b, CdVF05, DHX08, DS93, DGSW14, DGKL10, DR87, DS99a, DS07, DKR08, DFP14, Don96, DMZ22, DG20, Dür98, EDM05, FCG16, FCC20, Fis11, FKN96, FV11, Fuh76, GW16, GPS95, GBGdO00, GdOB02, GGF09, GB93, GR88a, GTHLHR02, GRC81, GRC82, GZSAB13, Gri93, GN03, GG18, GAHL09, GW19, HH76, HLZF21, HSP16, HZ19, Ich07, Igl96, IO96, Jac04, JS90, JK06, JGC^{+}23, KC15, KKP85, KF03, KT05, Kim78, KG19, Kob77, Kob83, Kob84, KL02, Kub85, KS81, KGM93, KG95, LD06, LK08, LHH23, LM13b, LL13, LZ16b, LMMY20]. Discrete [LTT24, Lin14, LM97b, LH14, LL19, LWY23, LH24, LM02, LDZ10, Ma09, MGHT19, MS92, MBL00, MPZ03, MS07a, MLZ17, Mey89, MVBC18, MM11, MNMS24, Mor95, Mor24b, MP21, NLPT09, NS18, NT16, NY18, OW93, OBL11, OC04, OS10b, OS12b, OL77, ÖW92, PP13, Pha00, PD92, PSV18, RG14, RW87a, RW87b, RW78, RW90, RL06, Ros84b, Rus88, SL16, SO13, SS09c, Sch15b, SL98, SK86, Sha17, SX19, SH12, Shi02, SB78, Sor98, Sto92, Sza95, TA81, TZ07, THA02, TW94b, TL09, Tie14, Tie20, TS95, TO15, Ver97, Viv03, VP15, WZD^{+}23, WC19, WL21, Wim92, Wim06, Wir98, WHGY22, Wyn09, XZGT23, YM92, YCRZ21, YLL^{+}15, YG95, Yu20b, dWvS00, JS91]. Discrete-Event [RL06, TW94b, TL09]. Discrete-Time [AD15, AS93, AD98, ABB12, AMV07, AM81, ABKM96, ABFG^{+}93, BKP90, BS86, BLJ15, BBT14, CLYL13, CW23b, CdVF05, DHX08, DS93, DR87, DS99a, DMZ22, DG20, EDM05, FCG16, FCC20, Fis11, FKN96, FV11, GW16, GPS95, GBGdO00, GdOB02, GGF09, GB93, GR88a, GTHLHR02, GZSAB13, Gri93, GN03, HLZF21, HSP16, HZ19, Ich07, Igl96, IO96, Jac04, JS90, JGC^{+}23, KC15, KKP85, KT05, Kim78, KG19, Kob77, Kob83, Kob84, KS81, LD06, LK08, LHH23, LL13, LMMY20, LTT24, LM97b, LH14, LL19, LH24, LM02, LDZ10, Ma09, MS92, MPZ03, MLZ17, MVBC18, MM11, MNMS24, Mor24b, NLPT09, NS18, NT16, OW93, OC04, RG14, SL16, SH12, SB78, TA81, TZ07, Tie14, Tie20, TS95, TO15, VP15, WC19, WL21, Wim92, Wir98, WHGY22, XZGT23, YCRZ21, YLL^{+}15]. Discrete-time [LZ16b, MGHT19, Wim06, JS91]. Discretionary [KW00, Mor10, Zer03]. Discretization [AOS16, BPV19a, BPV19b, Bro22, DNS17, LW15d, MV08a, MV08b, NPR15, PV13, PH92, PEB19, RV05, Red79, Sar22, Vel97, Vuo21, Wac12, WW16, vDHV15]. Discretizations [MRV11, VW95]. Discretized [Che04, DHV00, DKK^{+}19, DPDS98, EA21, Zhu96]. Disjunctive [Jer80]. Disk [HCG20, LB89]. Disk-Beam [HCG20]. Disorder [BP07a]. Dispersion [RZ93]. Dispersive [Lis14, VLP24]. Displacement [MW21, Pan17]. Displacements [PBPB15]. Dissipation [Sun96]. Dissipative [Alv00, ÁRL21, BS09a, BS13b, CBdN08, CB15, DGSW14, DHC19, FO16, GMN22, HM03, IK94, Jay10, MLYF02, PW02, Rap22, Reb95, Sun82]. Dissipativity [BBTN03, GG18, PB09, Pan98, TFT11]. Distance [BG98, DV04, Del04, GL92, GLD^{+}91, GY12, HBS06, JL93, KY23, LHS06, LR23b, OMFMC^{+}21, Pen13, RCA12, RSS22, Sas14, SS18, TG13, TCC15b, TCC19]. Distinguishable [AG95]. Distortion [JXZ24, SCCL18]. Distress [BC17]. Distributed [AT89, AFOQ21, AB09, BMS82, BGRW88, BRR90, Bar81, Bar83, BL89, BL15, BA15, BE87, BDRR98, ByS17, BP98b, BGH94, CD89, CSBC14, CMR07, CSC14, CMP15, Cha14, CKM20, CB05, CM18, CP77, Cur84, DG14, DXG15, Dav78, Dav79, DH09, Di 91, DJZ96, DMPD94, FFR14, FMS20, FCC20, Fla84, FA13, FRS16, Fuj80a, GL22a, GZ12, GK85, Gue03, GM99, HP10, Hel76, HW12, Ho90b, HY97a, HY97b, JVB^{+}23, KMP13, Kar06, Kha01, KN77, Kno79, Kob76, Kob77, Kob78, Kob79, Kob83, Kob84, Kob88, Kom97, KS85b, KM18, KY87b, Lag83b, Lam92, LD05, LG09, LCF18, LY91c, LLMT02, LR23a, Liu97a, LL98, LLX11, LW15c, LLSX20, LWDB20, LZ21, MM04, MYP14, MCPA17, MNC16, MMN09, MS90, MS17b, MS22, NOV03, OT09, Ols17, ÖT90, Pan92, PP10]. Distributed [Ped91, PS99, PW02, Pop95, PTZ17, Reb88, Reb89, RT98, RCS01, RC16, RS15b, RW92, Sca19, Sch83, SC12, SLQ^{+}23, SB94, Sok87, Sta15, Tib85, TG23, VS80, WZHZ19, WLZ22, WQS^{+}24, WCJP04, Won09, XC18, YZ97, Yu95, YLYC15, YHJ^{+}18, YHH16, ZYHX18, ZZ13, ZM13, ZM14, ZK20, Cou83, You87]. Distribution [ALNP24, AP20, ACCG18, AX05, BG98, BK21, CTZ03, CS87, Dor86, DMZ22, EGN85, Gen13, Man95, Maz94, NNN21, Ste92, Sun81]. Distributional [CW03, KY23, WC96a]. Distributionally [MU22]. Distributions [GKL20, Her86, Hir81, OT94, Ram94, Sny77, VP15, Kre85]. Disturbance [BKP90, CM00, CP95, DM89, DC10, EÖ88, EÖSÖ98, FLJ13, FA89, GG13, GZ13, GZ14b, GZ19, Hir81, HH17, HNv92, Hui92, KKA23, KLR09, LD06, Liu23, PNSL20, QLD01, SBBV15, Tre86, WFZ20, WC81, ZW18]. Disturbance-to-State [KKA23]. Disturbances [BGP09, BFPB13, FC05, GP21, Goo76, Gra07, GW18, KK17b, KR87b, MKK19, NK94, QV04, Rus86b, WTJY21]. Disturbed [ZD24]. Divergence [CA10, JPSB24, MG13]. Dividend [AM15, BP10, BM21, CGY21, CTZ03, DEO24, FS19, NPY20, Pau08, QJL22]. Dividends [AAM20]. Divisible [Lid15]. DNSS [ZCGT10]. Do [BGMR94, vH10]. Doeblin [Kur89]. Does [SA96]. Domain [ACS00a, BD01a, BC81, BFCLS97, BM00b, Bla76, BT17, BL00a, CFZ17, Cer07, CRZ13, Che79, Che81a, Cob06, CP82, Cor99, DDK10, DGQ13, FLUU17, Fuj86, Gon94, HH19, HLY24, HdHVW03, HRPT21, KX02, KX04, KLM^{+}21, LP15, LL00a, Leu99, LCZ20, LtLR21, MRP99b, NPS15, ÖB19, PS06, Reb95, Rei86, RZ06, Rot96, RZ93, Sla00, Sla03, SCK10, Sun96, WWY06, YH21, ZC91, Zwa88a]. Domains [AMPR15, Ban19, BF91, BBGBO14, BP13a, CN24, CS10a, CGM11, CSC14, CSCV18, CDZ94, DH84, Day09, DGH09, HBD21, II97, ISW00, MS23, Mar99, NT95, OC23, RV90, VO22, Wol12, XWZ^{+}21]. Dominance [CRST18, DR08, HJ13]. Dominance-Constrained [HJ13]. Dominant [IK07a]. Dominated [Nai87]. Domination [TY23]. Domination-Monotonicity [TY23]. Doob [JN19a]. d’optimalité [HU78a]. Double [BK80, CFSGS15, Cos13, HW12, Pog17, TYW11, ZRSZG22]. Double-Indexed [TYW11]. Double-Obstacle [Cos13, HW12]. Doubly [HPW10, OS12b, OS14, OS19, Qua03a, RRS14, Ste08]. Douglas [BC20, BCKM24, MZ09, Sva11]. DPP [NSW17]. Drag [BFCLS97, GHS92, PS10b]. Drift [BFM23, CY18, CI19, CYL96, CYL97a, CYL97b, CY94, ET12, FO16, Liu14a, MR15, McN84, Shr81, TZ07]. Drift-Diffusion [McN84]. Driftless [MPS99, Tie20]. Drifts [EMP24]. Driven [AC02, AC13, AC17, AW21, AHKS22, ATW93, Ban19, Bis12, BC17, BJ17, CH13, HYZ11, Heu94, HZ05a, HP09b, HT22, KR88, Kus89, MML21, MR96, Nie21, PZLM24a, PZLM24b, Ris94, SK95, TZ07, ZH24]. Driver [CSA^{+}13]. Drop [FZ09, GTQ21]. Droplet [LW15a]. Droplets [BK21]. Dropout [YYH^{+}16]. Dropouts [JGC^{+}23]. Dropping [GP79]. Drops [XZGT23]. Dual [BLW07, BS24, BTBT88, BIK99, DM15, EB24, GV05, HI84, HP87, Han77, HHF21, IK04, IK07a, JK20, KG17, KB76, LCF18, New06, New07, NDHL24, PW23, PA93, Sag17, Set93, TDBL22, Tse90, Wri95, ZBA13, Now17]. Dual-Gain [NDHL24]. Duality [AC79, BS80, Bal77, Bis78b, BM03, BCG^{+}17b, BL91, BO09, BKM14, BCKM24, BL00b, CN05, DGKL10, Die87, Eke77, Ete20, FGL08, FV89, FHM84, GR02, Goe13, HM76, HLT08, Iof79b, KLSX91, KA06, Le 19, LL00b, Lom21, Mat98, PP14b, PTZ21, PW78, Pul96, Roc89, RW00a, ST99, Vin93, WZY09, WK77, Zhu92]. Duality-Based [BCG^{+}17b]. Duality-Stability [Bal77]. Dubins [SC06]. Dubovitskii [IK08]. Duration [Ber85, HG90, Yon90, Ber88]. Durations [Aub13]. Dwell [DEJ23, LMYZ23, LLM^{+}24]. Dwell-Time [DEJ23]. Dwell-Times [LMYZ23]. Dynamic [AVJ21, AFV17, AJ12, Alm01, Ari98, AY23b, Bal78, Ban05, Bar91, Baş84, BY13, BCP16, BJ04, BR82, BR17a, Ber92b, BH86, Ber89b, Ber77, Ber98c, Ber18, BJ19, Bor89, Bor00a, Bor07, BT11b, BN13, Cal09, CF05, CLM91, CS05, CV87, CFSGS15, CS15, CGS18, CK17b, DR08, DI86, DQS16, EQ95, Emr80b, EK09, FT14, FG15, FFP^{+}09, GMCL10, GZLC17, GGF09, Gho88, Gom20, GAHL09, GYBL15, HS95b, HSD97, Hen83, HNv92, Hui92, Ich79, IS12, Ito98, KK92, KN14b, KS81, LL00a, Leu99, LZL02, LR23b, LLLT18, Log13, Ma09, MP14, MP76a, MM23, Mik07, MNMS24, MR96, NS12, ÖW92, PP18, Pha00, PW17, Pom09, Pop97, QN82, RRW96, Ros94, RV93, SY21, SBY23, Sch15a, SD98, SP02, Shi02, SPBF10, SBT19, ST02, ST09a, Sto91b, SZS21]. Dynamic [TST07, TTZ23a, Tan15, TZZ88, TLD24, Tsi95, TCC15b, TCC19, VP87, VL78b, Wag91, WL24, WF76, WY08, WZC20, WMWL22, Ye95, Ye97, Yük20, Zit14, dFN23, dMCJ22, TA96, TPA97]. Dynamic-Compensation [WL24]. Dynamical [AS11, BT11a, Bag81, BCL83, BSDR97, Ben96, Ber82a, BT03a, BC16, BF76b, BCF12, Cér00, CLSL20, CB05, Cro81, CI86, DSI22, FW93, Fio20, Fio21, GD88, GR88b, GP82, HBD21, HHF21, HSJ19, KG22, KK17a, KGM93, KG95, Li07, LH11, ML13, Mai03, MY93, NC13, NJD09, PPTZ19, PR07, Rao98, Rod21, SW11, SS19, TC14, TK81, TO15, Vuo21, Wil77a, ZBA13, ZJ89, ZZCaL19, ZL24, dFS15]. Dynamically [BJ02, CMA08a, CMA08b]. Dynamics [AHKS22, Art78, AG97b, AL02, AD18, BTB16, BBB24, BP14b, BHT10, BCS24, BJO09, BLZ94, CDJS23, CF18, CWC^{+}22, CM96a, CGC17, CdLdDM02, DPV00, EP93, Ete20, FFS20, FMSV16, GN18, GKP18, GK02, HO16, HY97a, HY97b, HCM05, HLZ15, IK94, JFB17, JR06b, KLM20a, KR85, Lac17, LM15, LZ06, LYB18, MYAS09, MMH19, MB12, NK94, Nie21, PW17, PS17c, QS13, RW00a, Sap19, SJ13, SR18, SMK18, TLL16, Ter94, TA21, Vig97, VP22a, Wir98, XJ24, ZG17, ZHH17, ZTC16]. Dynkin [Boe05, FT02, Grü13, HZ10a, HM14, LS19b, TV02, YT13, Yan14].

E. [BF89a]. Early [ASR^{+}13]. Easily [Pro22]. Eckstein [Bau09]. Economic [BØ94, DFŚ24, GL22b, Res76]. Economics [FP04]. Economy [FR20, GW20]. Effect [BFF20, CIH18b, DLP86, DPV00, Fra03, MCL^{+}22, MP11a, WCCL04, Zha08]. Effective [CL19b, GO04, TL09]. Effectiveness [GGH16, HRW88]. Effects [DKS^{+}11, DK86, GLS21, HSZ24, MÖ14, Tim14, VPJ19]. Efficiencies [Zhu94]. Efficiency [FBJ09, HS09, Kiw96a, Kiw96b, Pap86a, Pel00, TLC99, Zaf03, CR19]. Efficient [Bag81, BBBM08, BGP20a, Bor77, BDDSV05, CK02, CHJ15, CM07, DG90, FMS20, GL97c, KMP13, KR04, LW87, LY91a, Mey83, NM10, Tam80, ZSS^{+}20]. Efficiently [GL76]. Effort [AK99b, IK11b]. Eigenfrequencies [LYLZ99, WC91]. Eigenmodes [ZC91]. Eigenstructure [CM91, NS14]. Eigenvalue [CD15, CDM21a, CFNG10a, CFNG10b, CW81, Cob10, Jar93]. Eigenvalues [Ant14, Lou15, Pie79, QS13, VO22, dG06]. Eigenvector [McE04]. Eikonal [CM06]. Ekeland [GJN08, Qiu13, Ste08]. Elastic [CDD20, CRST18, DDK10, FSH16, Hua88, MY93, Mor98, NFI15, Rao98, Toa11]. Elastica [Ito98]. Elasticity [CDM21a, CG10, GGM01, GZ09, Oss01, Rao95, Sle78, Zua96, Zua98]. Elasto [MAJ18]. Elasto-plastic [MAJ18]. Elastodynamic [AK99a, Lag83a, SL18]. Elastodynamics [GY98]. Elastoplastic [GMR22]. Elastoplasticity [AS21, dlRHM16]. Electric [CFRV98]. Electrical [ADK08]. Electricity [AX02, AX05]. Electrodes [CFRV98]. Electron [FZ19]. Electronic [LR10]. electrostatic [KMM07]. Element [AGL08, AM89, BKR00, BGR04, BPV19b, ByS17, DGH09, GY11, GHZ14, GMS22, GW19, HR21, Kno82, Las95, LLMT02, LW15d, MRV13, McC09, MV07a, MV08a, MV08b, MRV11, NPR15, RV05, Tib11, VS18, WW16, XZ15]. Elementary [EP93, Sîr14b]. Elements [Bra09, CM81, Dol76b, DNS17, HR83a, Jah85, VW08, ZKW99, HR83b]. Elimination [BP04, BP06, Cob06]. Ellipsoid [EK85, Kiw89]. Ellipsoidal [KV06]. Elliptic [Aca93, AFO22, AO15, ABV13, Ber92a, BK97b, Ber98b, BC95b, BS11, BK13, BHL10, ByS17, CGM98, Cas86, Cas93, CF95, CTU00, CM02, CR06, CT09, CCK12, CK14, CT14, CT20, CCJ23, CZ99, Che99, CCT05, CW22, DK05, DGH09, DMS03, DJZ96, DGQ13, EL20, EHS07, Fuj86, GY11, GWY14, HK13, HL11, IK90, KN14a, KMR10, KW87b, KLL10, KS13, LWY21, LLMT02, Lou03, LY09, MRV13, MY09, Mic80, MS85, NPR15, PV13, PW18, QW19, Rai99, RV05, RT06, SW78, Sto91a, Sto97, VW08, WW16, WWY06, Wol12, YZ97, Tah92, You13]. elliptiques [Tah92]. Embeddable [HT92]. Embedding [BD01a, GTT16a, GTT16b, NT95, RTT20, Sla03, TFL14]. Embeddings [Hak20, JPP14]. Emerging [TLL16]. Emission [HLG21]. Employee [LS09]. Empty [SZ10]. Encoder [RRJ16]. End [Dun98, Fra87b, GKP18, KWY08, KQWY20, Mik15, Col82a]. End-Constrained [Dun98]. End-Time [KWY08]. Endogenous [GN18]. Endowment [BC12, TZ10]. Endpoint [Eno94b, ÖB19, Tre89, WL17, Zhu00]. Endpoints [Den22]. Endpoints-Constraints [Den22]. Ends [GH04]. Energetic [Duc20]. Energy [ACS00a, AB03, Bar80c, BD11b, CDCVV21, Cha10, Cha11, CDM23, DPZ91, DvdS98, DDK10, FMP^{+}22, Gha11, GL99, Gug07, HW12, Ich07, Jay10, Kim91, Kle18, Liu88, LL98, LE17, LR06, MS03, NCL13, NFI15, PPZ13, PZ03, RY95, Sha92, Ste17, VM00, ZM13]. Energy-Aware [ZM13]. Energy-Conserving [DvdS98]. Energy-Minimizing [Sha92]. Engel [BM18a]. Enhancement [Liu08]. Enlarged [ZSM10]. Ensemble [ABS18, Che23, DKT17, DHK20, LZT20, WL17, ZL21]. Ensembles [AS20]. Ensuring [ZFKZ21]. Enter [Pet88]. Entropic [RTTY23]. Entropy [BL91, CL87a, CK09, Col12, CH21b, Da 14, HLT08, Igl96, Kaw11, LY19a, LY23, New06, New07, Per12, STZ24, VL23, VP15, WHS19, XUP08]. Entropy-Based [HLT08]. Entropy-Like [BL91]. Entry [Zer03]. Enumerative [HV00]. Envelope [Bar77, JN19a, RW00b]. Envelopes [NR21]. Environment [CMA08a, CMA08b, FH17a]. Environments [SSZ11]. Enzymes [Ste92]. epi [CC99]. epi-Lipschitzian [CC99]. Epsilon [MS96]. Epsilon-Maximum [MS96]. Epstein [AH21b]. Equal [WGK11]. Equality [BT03b, BT10, CL87a, KZ00a, Pre96, Tre99]. Equals [CMS16]. Equation [ACS00a, AN15a, AN15b, Ahm77, Ahm07, ACL09, AMP10, AFO22, All88, AW90, ASTT09, Amm02, AC18b, AAR18, ABK19, ÁRL21, AK16, BHM99, Bar82, BD83, Bar91, BP97, BD11a, Bar20, BC81, BE13, BC18a, Bea08, BLT15, BGR07, BR82, BGR90, Ber21, BV17a, BPS05, BH00, BGMS78, BGBPG04, BCM12b, BJ10, BG05, Bou95, BT17, Bou13, BFM23, BK23, BHT21, Brü13, BK83b, Buc92, BHL12, Buc75, Buc77, BFO20, BR15b, BM18b, CFZ17, CF24, CC04, CS18, CM06, CRT15, CKL16, CMT18, CT20, CK23a, CC01, CZ04a, CCM14, CO03a, CSC14, CSCV18, CCFV20, Caz14, CIP18, CP98, Cel02b, Cer07, CRZ13, CC18a, Che79, Che81a, Che81b, CW08, Che76, CS77, CM17, CLT06, CMR17, CDZ94, CV83, CF06, CS15, CD03, CLZ23, CP76, Da 85]. Equation [Dág06, DE18, De 92, DGJ93, DM09, Del04, Del07, DFLZ23, DM17, DWZ20, DMR19, Emr80b, ES81b, Fab08, FG09, Fab92, Far08, Far13, Fen16, FZ19, Fla84, Fla87, Fla90, FM00, FBP99, Fre99, FL80, FL82, FL22, FI98, Gag16, GLM21, GL90, GLK22, GW17, Gla05, GG07, GLR86, Goz91, GM17a, GM21, GLM19, Gri83, GKLW14, GW18, GSZ20, GT05, GM09, GMS22, GLS05, Gug10, Gug15, GZ05, GG13, GZ19, GR21b, HH19, Had06, HH76, HLY24, HO22, HS12a, HS12b, Hau94, Hei16, Hel76, HK88, Ho90b, HP09b, HJX19, HT22, IY12, Ito96, IR00, IK07b, JL14, KS87a, KS90a, Kha94, Kha95, Kha01, Kha03, KS99a, Kim92, Kim93a, Kim93b, KKLO21, Kle18, KY07, Koe20, Kom91, Kou24, KKV11, KGS11, KW13, KTV16, LT89, LMMR15, LY90, Lev85, LPT22]. Equation [LYLZ99, LCZ20, LX23, LL24, LLC18, Liu03, LZ18b, LG20, Lou15, Mac94, Mal02, MT03, MRR13, MRRR19, Mar99, MRV03, MCPA17, MM23, Mic01, MZ06, Mic09, MN22, Mil02, Mis19, MS97a, MSv86, MO99, NPR15, Ngu22, NP06, OCI00, OZ00, OZ03, ÖB19, PY95, PCP22, Pav80, PZ14, POU11, PS04, QW18, RMZ02, RS15a, RS20, RS15b, Ros84a, Ros00c, RZ06, RZ09, RZ93, Sak02, SAM03, SZ19, Sca19, Sch80a, Sch80b, Sch86, Sei77, SZ82, SP02, Sha83a, Sha83b, Sha85, Sha86, SMDB97, SCK10, ST09a, Sor98, SV17, Sun96, SCFV97, TST07, TTZ23a, Tes92, Tib85, TR01, Und76, VZ09, VM00, VA03, VK81, WC89, WWY06, Wan08, WGK11, WZ12b, WY16, WYY21, Wat24, Wec84, WCG14, Wim92, Wim94, Wim06, WY08, Xia19, Xie12, XG03, XX19, Ye00a, Yu15b, YH21]. Equation [ZA10, Zei94, Zha99, Zha00, Zha01b, ZY08, ZDM20, ZQ18, ZC91, ZG18, Zua90, aJM18, de 98, vN04, EJ89, You87]. Equations [Aca93, AØ23, Ahm81, Ahm83, Ahm19, AS88, ABCL17, ACS00b, AFOQ21, AM23, AQ05, ABSP08, Aub11, AK02b, AK19, AHK24, Bad09, BMRR20, BD15, BP94, Bal91, BK82, BL83, BTY11, BRS11, Bar12, BRZ20, Bar21, Bar23a, Bar23b, BS16, BS13a, BK22a, BKR00, BK83a, BD11b, BBF18, BH86, Bet19, BHS19, BD19a, Bis77a, BISW17, BCG86, BR13, BRT15, BCM12a, BC95b, Bor89, BK95, Bor00a, BB04, BK14, Bre16, BKR17, BH23, BS13b, BvdMR05, BM07, BQT08, BL08a, BHL12, BN16, BSZ08, Bur20, CH21a, CGM98, CD89, CD91, CT96, CLAL13, CP00, CR11a, Cas93, CF95, Cas97, CTU00, CR06, CMR07, CT09, CK14, CC16, CKK17, CK17a, CK19, CK21, CK22, CW23a, CDR08, Cau13, Cer01b, CT19, CPY09, Cha09, CI19, CSZZ17, CC18b, CMC81, CM81]. Equations [CLW17, CR21, CY24, CDG15, CDKV18, CLM24, CTW18, CB11, CE11, CLT24, CF13, CK89, CPW88, Cor99, CT04, CG17b, Cur76, CI77, Cur03, CS11, Cur11, Cur13, DDK12, DL06, DI93, Dág18, DLP86, Dav78, Day09, DFŚ24, DH22, DI94, DRT11, DC06, DQ21, Dmi92, DO14, Dok05, DE84, DTM22, DL20, DM15, DMS12, DA17, DM13b, Du15, DMPD12, EGR^{+}22, EL20, FR17, Fag08, FGG10, FGG11, FGB95, FCLZ16, FM08, Fra93, FP11b, FNV10, Fri84, FYZ07, FL17, FKW23, FT04, FMT10, GL92, Gah94, GT17, GCL15, GLX18, Gar03, Gar76, Gev81, GGH16, Gib79, Gib83, GMM91, GW76, Gom20, GRS00, Gra05, GLNT23, GT14, Gue07, Gue14, GS18b, GI04, GM11, GK98, GK11, Guo02, GMY16, Gut90, HM03, HMV05, HPS07, HT12, HSW12, HSW15]. Equations [Hew93, HP09a, HH08, HH10, HMMS13, Hop87, HR98, HY95, HZ03, HMR15, HLDM20, Ich11, Ich79, Ich82, IRS02, IT86, IP87, IT87, IM98, JMP21, JB96, JQ97, Jun00, KMC21, Kap94, KK19a, KMF19, Kel24, KH90, KB01, KFP22, KC10, Kim90, KCBP21, KS98a, Koe20, KZ00b, KT03c, Kra93, KR15, KS90b, Kun82, KW85, KV07a, Kus79, Lag77, Lag78, Lag88, Lag89, LL00a, Lam92, Lar01, LT83a, LT83b, LT86, LS91b, LL21a, LL07, LI00, Lef11, LW07, LXY16, LHX12, LW14, LR16, LM16, LZ18a, LY18, tLR23, LR23a, Lia15, LFL23, LB00, LZ01, LW11a, LY20b, LM82a, LM82b, LS85a, Lis14, LZ19, Liu97c, LG07, Liu11, LZ12, LY19b, Liu23, LT23, Lor87, Lou03, LY09, LW11b, Lü13, Lü14, Lü15, LtLR21, LZZ21, LW22, Luk85, Lun91, Mah03, Mao90, MG20, MG21, MRR16]. Equations [MGMBN^{+}19, Mas08, MNC14, Maz17, McE95, MML21, MWY21, MY09, MST16, MS17a, MSR01, MERD02, MVZ^{+}09, Mol19, MR04b, MA03, MN77a, MN77b, MT22, Nag96, NN90, Nag95, NS05, NNSM08, Ngo11, NT17, NT18, NR15, Nic00, Now08, ØS10a, ORW13, Pap89, PS17b, PD97, Pen92, PW99, PZLM24a, PZLM24b, PA99, Phu01, Pla21, PS06, Pos17, PY96, PT95, QC95, QZ13, QWY21b, Qiu18, QW19, Rai99, Ram05, RR92, RZ98, Ray06, Rei83, RX15, Roc96, Rod21, RT07, Rub95, Rus86a, Sak84, Sal85, SW78, Sei78, SC16b, Sha10, SS20, Sha94, Shi02, Sor93, SW21, Ste08, Sti18, Świ20, Tan98, Tan03, TZ09, TZZ22, Tao18, TL06, Teb11, Tod83, Ton03, TWZ16, Trö99, UY79, VZ00, Vin77, VW95, Wac07, WW16, WS92, WL95, WS95, Wan00, Wan02]. Equations [WW02, WZ12c, WX13, WD14, WXZ15, WZ17c, WZ17b, Wan21, Wan22, WYZ23, WY23, War19, Was79, WWZ14, WS13b, Wol12, WA84, WHGY22, WJZ12, XA93, XG03, XZ20, YMH94, YT13, YZ16, Yao99, Yon93, Yon10, Yon13, You13, You17, Yu14, YZ22, YZ23, ZVBR14, Zha15, Zha18, ZD14, Zho83, Zho92, Zho93, ZL07, ZW18, Zho23a, aJM18, dlRT07, LS86, Sei82, Tah92, Tah92]. Equilibria [AY23b, Ave19, BL89, BCC^{+}19a, BB03, CC02, CR19, CJS08, Cza04b, DF20, DPR22, GNLB14, HW21, JN24, KK18, KDB15, MNMS24, NP22, NA02, SBR18, SSY24, WZ24]. Equilibrium [AVJ21, Bas77, BCL24, BK13, BCR04, CL18, Ete24, FT22, GY16b, HJ21, HT10, HJZ17, HZ19, Jac22, Kin84, LK11, Lóp12, Lóp13, MSZ18, MM21, MZ20, NLZK19, Out00, Ped98, Res76, SBY23, SS99b, TWZ21, Top79, Uch78, Uch79, WZ21, Zho23b]. Equilibriums [YXJ11]. Equity [MS11]. Equivalence [AD08, ABH98, Bon91, CD80, FG15, GdZ90, GLT95, HPW83, HO90a, Hua10, KV13, MRv94, PK95, Pom09, QW18, Uch81, VL78a, WZ12c, WX13, WZ94, Zha15, Zwa88a, vRM98]. Equivalent [BP17, GW96a, JZ99, Ols76b, Roc96, SBY23, YF97, ZzJh10, van95]. Equivariants [RR13]. Erdos [JNKV21]. Ergodic [ADRU22, AP20, AMV20, ABK14, ACPZ19, BP14a, BCL04, BG88, Bor94, BB04, Bor07, BG07, Bud03, BGL11, BLS12, CDR08, CKT20, CI19, Cir14, CH13, DPDS98, DMPD98, FHT09, GAM97, GR05, GM09, GT20, GC05, HLT20, Ich11, JL23, JN06, Kou24, Kru00b, Kum13, KY92, Kus04, LP86, MWY21, MR18a, Mey89, MO99, PS17a, PRHL08, RS93, Wee07, Yük16]. Ergodicity [BCP17, BBK00, CD08, GKY21, SWW22, XZ17, Xie12]. Errata [Ban78a, BP06, Cur85a, GM21]. Erratum [AC17, AW90, BCS21a, Ber88, BG21, Bor00b, Bor10, Buc77, FR82a, GK95, GP21, GŠZ10, Hau81a, HSW15, HR83b, JS91, Jou00a, KH79b, KX04, KS99b, LP81, Lob76, Lóp13, LY02, Pen88, PG18a, Shi90, Zab80, ZZ94, Zua98, Lob76]. Error [AFV17, AFOQ21, AFO22, ARS09a, Bar80a, Bar77, BPV19a, BPV19b, BC24, CR06, CMR07, CMR19, DGO19, Den98, Ger06, GCP88, GHZ14, GW19, GZ19, GGA^{+}17, HR21, Jou00a, Jou00b, KS18, KRS14, KR85, KMR10, LV16, Li95, LW15d, MRV13, McE04, MV08a, MV08b, MV11, MRV11, PV13, RY95, RV05, RZ20, TWZ16, XG89, YF97]. Errors [MD13, SD98, XC18]. Errors-in-Variables [SD98]. Escape [DK89a, DM97, Mil02]. Escaping [ABB15]. Essential [KMM11]. Essentially [LT92]. Estimatability [WR00]. Estimate [Aih92, AB09, Dmi92, Don96, DFLZ23, FL22, Lü13, NNT22]. Estimates [AFOQ21, AFO22, ARS09a, AAFC19, Bag81, Bar80a, BA15, BBGBO14, BBV10, BBV11, BPV19a, BPV19b, BHL10, CMV08, CR06, CMR07, CMR19, Cel07, CG87a, CG87b, DGO19, Dun81, DK86, GCL15, HLL^{+}17, HR21, Hu15, JV06, KMC21, LV16, LY19b, LW15d, MV08a, MV08b, MRV11, Men89, MR05, RV05, RZ20, Sor98, Teb22, TWZ16, VM00, WD14, WDS79, YZ22]. Estimating [ZZCaL19]. Estimation [ASR^{+}13, ABB12, AR21, BBC81, BL83, BM86, BN81, BS86, BSDR97, BFP09, BTBT88, Ber95, BP08a, BGP20a, BH99, BM03, CRZ^{+}18, CYL96, CYL97b, CZ10, CKM20, CY94, Cur76, DH09, DTWY91, EK97, EDM05, FB98, FS23, GAC24, GL22a, GZ12, GL95, GY16a, IS12, IK90, JVB^{+}23, pJY20, JZ09b, JZ21, KG80, KMP13, KM86, KH69, KW85, KW87a, KP94b, KW78, Lam87, Lam92, LM13b, LZSC24, MN03, Mur92, NR97, Pic93, PP00, PO00, QN82, RMB12, RY95, RKL16, Rus86b, Rus88, SW11, SY17, SCCL18, TWY90, Tan00, TSS^{+}22, VL23, VW95, WS13a, Wei77, WY06, XUP08, ZVBR14, ZFKZ21, FFR14]. Estimator [AB93, BSI19, DSI22, IS14, JKSZ21, NC16]. Estimators [Ger92, Ger06, PP00]. Euclidean [BM94, HBD21, KN13, Kim22, PDB23, RSS22, YXC21]. Euler [ASTT09, AT00a, ÁRL21, BCM12b, Bou13, Bro22, CP98, Cel02b, CO22, Cor99, DM09, DF98b, DKRV13, Gla05, GS18b, Guo02, GH04, GZ19, GZ23, Kim92, LT89, LL98, Mor95, Roc96, Sha18, VZ97a, WC91]. Eulerian [CFT12]. European [DPZ93]. Evaluation [LR23b, SYZ13]. Evaluations [VZ21]. Evasion [AL98, BB88, BB90, BB94, Ber86, Kaś77, Sor93, Sor98, TS18, Wen15, Yon88a, Zar84]. Even [BCS20, Sch86, WGK11]. Event [DGKL10, JK06, KEK21, KCY22, KGM93, KG95, Lin14, LH14, LLSX20, LZSC24, MBL00, ÖW92, RW87a, RW87b, RC16, RL06, Sch15b, SL98, TW94b, TL09, WS20, WMWL22, YYH^{+}16, YG95]. Event-Triggered [KEK21, KCY22, LLSX20, RC16, WS20, WMWL22, YYH^{+}16]. Events [CMA08b, SK95]. Eventually [BM11]. Every [MP21]. Evolution [Ahm81, Ahm83, Ahm19, ACS00b, BMRR20, BP94, BKP14, BW89, BMN19, BC23b, Bul01, CSZZ17, CGP22, Chr19, CLMSR00, Col11, Cur76, CP76, CI77, Del04, Don10, DL20, DM13b, FR17, FNV10, GGH16, GK11, GGA^{+}17, Had06, HM03, HMV05, HP94, Ich79, Ich82, JFB17, KV07a, LR21, LFL23, LL15, LLM15, LT23, LZZ21, Mah03, Pap89, Rin09, Sak84, Sei78, SV17, TBP18, Tar85, VS80, Vin77, VK81, VW95, WZ17b, XA93, XG03, Yon93, You17]. Evolutionary [Ara14, CC16, CK21, FGH05, GY16a, TLL16]. Evolutions [Aub13, Car12, MA19]. Exact [AE18, Bar23a, BLT15, Bon89, Bur91a, Bur91b, CS18, Car15, Cer07, CMC81, CyY21, CMT11, CLM24, CGW10, CRV09, DAR15, DL09, DG85, DG89, DR79, DTWY91, DG23, Duc20, DMR19, ET15, EK97, Fab92, FYZ07, FI98, GFF17, GL11, Gue14, HH19, HZ95, HPS07, Hei88, Ho90b, Hu15, Kap94, KMM07, Kha94, Kha95, KMM11, Kim92, KC10, KY07, KY05, Lag78, Lag89, LL00a, LT89, Laz88, LR03, LY19a, LY23, LL24, LBMH^{+}16, LV97, Lü14, LtLR21, Mac94, Man85, Mar20, MRRR19, MM23, Mol19, MT22, Nic96, Nic97, Nic00, ÖH14, Per12, RS07a, Ris86, Riv90, RL20, Ros00c, RZ09, RW94, Sch92c, SS99b, Sei78, SMDB97, Tao18, TL06, Tri77, Tri80, Tuc96, TW00, VLP24, WW23, Wec00, WCG14, Yao99, Zer05, Zha99, Zha01b, ZL07]. Exactly [LR23a]. Example [DLP86, DH98a, EP93, HM88]. Examples [ŁS85b, dNCP20]. Exchangeable [SSY24]. Excitation [BC00, DV04, GAC24, GL22a, KNW95, LC01, MCP21, RY95]. Excitations [CW09, UMD88]. Excitatory [KB16]. Excitatory/Inhibitory [KB16]. Excited [CS10b, Cur76, MCS13]. Execution [GH17]. Exercise [ASR^{+}13]. Exhaustive [SO06b]. Exhibiting [BF76a]. Existence [Ahm03, AP06, Ang76, Bal94, Bal01, Bar23b, Bat95, Bel00, Ben81, Ber85, BM13, BG05, Bud03, CG00a, CHJ87, Car90, CW23a, CM10a, CP02, CZ95b, CO03b, CM96b, CM05, DS93, DZ24, DF20, ENJP88, Ell76, Fat94, Fat97, FT77, FBJM14, FT04, GS94, GB84b, GGP02, Goo76, Gu90, GM13, GK98, GYBL15, Gup20, Hau82, HL90, HT12, HTY01, Hew93, HJX19, Hut09, Iof76, IPM10, JZ22, Jak80, KA06, KS81, Kur89, KS98b, KS99b, LZ17, LNT02, LT18, Lou07, LW11b, Mar02, MMR17a, MM80b, MS10, Mur86, NvS09, NP22, Nov04, Now89, PT09a, QZ13, Qua80, QR11, Reb78, RZ14, Rif00, RT06, SZ82, Tsi91, Tsi92, Uch78, Uch79, VC22, Vin83b, Wac12, WD79, Wu18, Yük20, Zar84, ZMK21, Zha17a, ZL01, ZQ18, Zho92, ZH02, vTY99]. existence [Ber88, BG88, HS12a, HS95a]. Exit [Ama04, BP88b, BA15, Bla97, BN16, Car97a, Car97b, HSW19, HHOZ19, HL81, HZ10b, Mal02, She85, Ye00a, Zer03]. Exit-Time [Car97a, Car97b]. Exogenous [AY23b, FZ23]. Exosystem [GZ23, LP15]. Exosystems [HP10, Pau17]. Exothermic [ITT04]. Expanded [PAH88]. Expansion [BS11, CSB88, Dür98, GI02, MZ07, SCFV97]. Expansions [Bul01, CHN94, JHB19, KM16]. Expectation [LZ21, PTZ21]. Expectations [CE11, DQS16]. Expected [AM17, AMV20, Cvi00, DG20, HOPS87, HK94, JN19b, Kre78, Lei06, WC19]. Expenditure [KM12a]. Experimental [BSB^{+}97, CH12, Sag13]. Expert [FPS17]. Explicit [CE98, Dup03, Heu15, KHP06, OC12, Sla00, VP07, Zha00]. Exploitation [KLM20b]. Exploiting [LM13b, WL24]. Exploratory [GRZ24, TZZ22]. Explosive [AQ05]. Exponent [AP20, AKP21]. Exponential [BRS11, Bv85, CWZZ17, Che16, CK03, CS87, Cob10, Col12, CG92, CVKB13, CA10, CW06, DGSW14, Dus21, Gra05, Grü98, Gug15, GG24, Guo02, GMY16, Hag05, HL16, HVDK19, HCM05, JALG18, JLL20, KK19b, KSS20, Kim91, Kle18, KLV09, KR15, LS09, LAM19, LY20a, LZ94, LL98, LLR01, Liu23, LH24, Mao90, MCP21, NNT22, OT94, Ouz10, PRS08, RA16, Reb95, RT98, Rod21, TZZ18, Urq05, WW16, WXY05, Wol90, XZ17, YW07, Zer05, Zha17b, ZW18, Zuy16]. Exponential-of-Integral [Bv85]. Exponentially [BMO15, GHX86, IP06b, JZ09a, LL19, MNC16, MNB12]. Exponentiated [PB96]. Exponents [AZ97, DHX10, VL23]. Expressing [BAP18]. Extended [ABVC19, AM82, AGJ^{+}13, BM06, BR87b, DKT17, EOF07, Fu23, Gal00, GV16, JXZ24, KK92, LAN04, SWP91, VZ97a, Zei84, ZYHX18, Zhu96]. Extending [MTS15, Pan13]. Extension [AD02, APH24, Bal77, BM90, BR17b, BC95b, BGL98, CLdRdP11, Fio21, KRW04, PMW83, PMH92, RKL16, RKL18, ZZ23, dG06]. Extensions [Fer81, JLZ98, Kiw96b, LC80, Tri76]. Exterior [CN11, DDK10, WWY06, War19]. External [And82, CLMSR00, Yam89]. Extraction [WCJP04]. Extragradient [Mai08]. Extragradient-Viscosity [Mai08]. Extremal [AKO17, BL01, BK95, BR88, DK78, PS18]. Extremality [Bal94, Urs82]. Extremals [BHT19, Bon85, BP02, Dun79, KS99a, Kog87, Kre77b, PS13, WZ94]. Extreme [AP92, Per81, PZ24]. Extremum [BT03b, BT10, PKL13, RSS18, Sut20, TDNT18].

Faà [GEF17]. Facial [Jer80]. Factor [CK23a, CK23b, DL13a, HNS18, HSD97, Nag03]. Factorization [AMZ04, BP08b, CL24, CH05, CF06, DH84, GK87b, GGLD90, Gre92, HH81, Hel76, JP01, JS81, Mik07, Mik12b, MS90, MPC91, Nur06, OS12b, SC23, WCJP04, ZS19, BR87a]. Factorizations [BGKV80, OS14, Pet04, PK87, Qua03a, Sta98a]. Factors [Emr80a, Fer97b, LMP95, Pet04]. Fail [ŁS85b]. Failure [TWZ21]. Failures [CIH18a]. Faithful [Ram05]. Families [BLPdD15, HL04, Hop87, JR06a, Jun00, KS99a, Lis14, NR21, PD97, RG14, Zit14, Bac79]. familles [Bac79]. Family [AMI20, PL15, PK95, Zor17]. Fan [GDL11]. Fan-Shaped [GDL11]. Farkas [Las97]. Fast [AG97b, AL02, BET07, BP13b, BLZ94, CZA04a, CM18, GFF17, Gra08, Hac79, Lis14, LL19, Mil06, NSC19, Ngu22, PA99, PRS08, RTF21, Vig97, Obe07]. Fast-Oscillating [BP13b]. Faster [PG18a, PG18b]. Fastly [GJZ10]. Fatigue [GS14]. Fattening [Liu11]. Fault [BL15, BCG^{+}17b, Ste23]. Fault-Tolerant [Ste23]. Favorable [Kul93, Lei06]. FBSDEs [LW14, TY23]. Feasibility [RW76, ZL01]. Feasible [CKR22, HK06, Li94a, LLT18, MSH91, MW78, Mey83, PT87, Per81, PV98, Rai99]. Feature [Gad08]. Fee [JS12]. Feed [VS80]. Feed-Back [VS80]. Feedback [AZ07, Ahm19, Alt09, AT00a, AC18b, AB02c, APA08, AA17, Ari98, Aug19, Bad09, BT11a, BH96, BW89, BI91, BSB^{+}97, Bar83, Bar95a, Bar12, Bar20, Bat95, BR87b, BCP16, BCC^{+}19a, Ber89a, Ber02, BM03, Bla76, Bon91, BK14, BKR17, BKP18, BGSS21, Bro10, BG14, BA84, BGH94, CG00b, CJK13, CGLM02, CGY03, CM91, CLM91, CCF07, CLW17, CK78, CM00, CW81, CLS97, CLRS00, CS03, CE06, Cob06, Con90, Cor95, CR17b, Cur85b, DN23, DLP86, Dat88, DTV18, De 83, DRT11, DI86, DBS05, DC88, DHKS18, Dol79, DTM22, Dre77, DS01, EGR^{+}22, EF09, EÖ88, Emr80b, EG02, FZ05, FP21, FG15, Flo97, Flo99, FGS93, FNR19, FIM09, FNV10, Fuj80a, GMCL10, GVL12, GGF09, Gho88, Gib80, Gla05, GR88b, GdZ90, GEF17, GP17, Grü00, Gu90, GCSB99]. Feedback [GDL11, Gug15, GS18b, GH04, GG13, GZ19, Ha88, HH81, HRW88, HH78, HRW06, Her80, Her91, Hey76, HS13, HP96, HMQD99, HJST24, HLDM20, Ich11, IP06a, IPM10, IK94, Ito98, IK09, Izm96, JLR09, JL18, JM17, KK92, Kan00, KT03a, Kar06, KJ07, KK19a, KK19b, KKA23, Kas02, KCBP21, Kob84, KGBS08, KM07b, KR87b, LCC17, LT83a, LT83b, Las98, Lef11, Lei01, LXZ11, LL13, LL18a, LD19, LMMY20, LAM21, LNP00, Liu03, LZ06, Liu13, Liu14b, LRW96, Log13, Luk85, Man04, Mee89, Mey01, MVHN09, MMK86, MS90, MR86, MV07b, MA17, MNMS24, MA01, Mos80, Mos82b, NE04, NRL22, NM08, Nie14, NDHL24, Now17, OPPU17, Ped91, PK95, Pon09, Pop95, PY96, QBP23, QLD01, RW87a, RRR16, Ray06, Ray10, RZ21, RL20, Ros92b, Ros94, Rya94]. Feedback [Sak83, Sak84, Sch92a, SS20, SS92, Sha17, SX19, Sha88b, SFFC08, ST90, Sto91b, Sto92, Sul94, Sul98, SAMB14, SIC00, TR02, Tal05, TZZ88, TP95, TF08, Tsi91, Tsi95, Uch81, Urq05, Wag91, Wan91, Wan98, WXY05, WYZ23, WL24, WL21, WC81, WHGY22, XX19, YCRZ21, YLL^{+}15, Yu15b, ZP05, ZC06, Zha08, ZY08, ZLHZ12, ZPQ20, ZLD12, ZG18, ZW18, ZD24, Zua90, Cur85a, CW06, Mos82a]. Feedbacks [BP98a, Dat88, FGG11, GM17b, GM21, GX04, Jam14, Kry98, MB04, MR09, NP06, Pep16, Rif02, Sch02, VM00, XS96]. Feedforward [Bat07, IP06a, IP06b, KG22, KK12]. Feedthrough [CG10]. Fees [Zar92]. Feller [FKZ22, JN06, NR21, Sha15, XZ17, ZY09]. FEM [AO15]. Ferromagnetic [CHY23]. Few [CBB19]. Fictitious [SPO22, SKXL17]. Fictitious-Play-Type [SKXL17]. Field [ABVC19, ACCD12, Ahu16, ADRU22, AHKS22, AS21, AM22, AD18, BCG^{+}17a, BP14a, BSS93, BTB16, BBC18, BCY15, BCLY17, BCR24, BGHR16, BCL22, BGP23, BR21, BDT20, BAKS18, BJ17, BS06, BPTT21, CH21a, CM16, CLLP13, CMS16, CD13b, CDFP19, CKM20, CNS21, DHK20, DMR19, FH17b, FH20, Fu23, FT22, GLS21, GMN22, GV16, GHS22, GX19, GHZ24, HHN18, HH17, HWY21, KBDK03, LM16, LWZ22, MSZ18, NWW17, NWW19, NC13, Nut18, RTTY23, RHCK24, SBR18, SY21, SSY24, SS21, Sar22, SC16a, SC19, SK23, SY24b, TTZ23a, TS24, TY23, WW03, WZ12a, WZ17a, WZ24, Yon13, You13, ZSX18, Bac79]. Fields [AB02c, AF18, DWZ^{+}16, Her78, HL04, LY12, LY23]. Filippov [Tol18]. Filter [AC19, BCL04, BBK00, BEL02, CA10, CdSBT09, Dav78, DDS15, Dok05, ENJP88, FR05, GP98, KW78, MY22, OP96, TD21, Vin77]. Filtering [ACKM22, AC02, AC06, AZ97, BB89, Bas91, BK95, Bou93, BVJ07, BLZ94, BL00b, CZ95a, CFN07, CGR96, CGS00, CGLM02, Cér00, DS87, De 92, FCC20, FP89, FM00, FR00, GP98, GJP88, GBGdO00, GdOB02, GTQ21, HH76, Ham82, HO01, Heu94, Ito96, IR00, JKSZ21, KSY23, KLV09, LMMR15, LMR97, Mar84, Ouv78a, Ouv78b, PZ04, PRS21, Pic93, PS88, RBA22, SM13, SC16b, SZX16, TA81, Tan00, VC03, YY05, YY08, YD03, Zha98, dOP03, dOG05]. Filters [ASP22, BK98, BK99, BK00, CFN06, CYL96, CYL97a, CYL97b, CY94, Coh97a, Coh97b, DKT17, EK97, pJY20, KBMS20, LR00a, Lev87, LZ16a, New00, New06, New07, Nih83, SVBZ09, WZY06, XG18, Zor17, vH10]. Filtration [LG07]. Final [AQ05, BGP23, FZZ18, MO02, MP76b, Sei88]. Finance [FP11b, ZES12]. Financial [DFŚ24, FP04]. Finding [ATV13, BLW07, BK08, CL18, GBBT78, KN13]. Fine [KY87a, Rou21]. Finetti [EMO23]. Finite [AGL08, AM89, ACPZ19, AZ97, AD17, AK19, BT11a, BMRR20, BB98, Bar80c, BB90, BS15, Bay09b, BC18a, BKR00, BG09, BE95, BGR04, BH86, BB00, BS19a, Bis83, BPV19b, BKK02, BT13, BT21, Bra09, ByS17, BFF20, CW95, CSC14, CSCV18, CCHH19, CM81, CYL96, CYL97a, CYL97b, CY12, CG13, CGY21, CIP96, CY94, CMR85, CMR17, Coh97a, Coh97b, Cor95, Cro81, CC87, Cur84, CS86, DM02a, DJLY09, Dat80, De 15, DGH09, Del19, DHP09, Doa21, DF21, DTWY91, DNS17, DPR13, EK97, FS19, FHH97, FHH99, FRZ93, FV98, GMP00, GBR18, GL13, GRZ24, GY11, GHZ14, Grü09, GKLM92, GMS22, GMR22, Gug07, GAHL09, GH96, GW19, HM03, Hai86, HXY21, HYZ24, Hei88, HW99, HR21, HJF10, Ito90, Jam91, Jam14, pJY20, Kar06, KLS87].

Finite [Kha16, KL21, Kno82, KLN21, Las95, LW10, Lei87, Lev87, LLMT02, LK23, LR79b, LCS96, Liu14b, LG20, LW15d, MP14, MPZ03, MNPR20, MRV13, McC09, MML21, MV07a, MV08a, MV08b, MRV11, Men89, MLZ17, Mor13, MS07b, NE04, NPR15, Nih83, ORW13, Orl04, OOS15, ÖB19, PS10a, Pan92, PR14, Pla80, PEB19, RV05, RSZ23, Rod82, RZ06, Sal85, SS99a, Sch92b, SY17, SB78, Son79, Ste11, TWY90, Tan00, TW94a, Tib11, VW08, VS18, WW16, WP18, WF24, WB87, Wee92, Wil77a, WCLC05, WJE16, XB99, XUP08, XZ15, YC06, YW12, Yon90, ZRSZG22, ZBA13, ZYGpJ23, ZD24, Kup85]. Finite- [BS15]. Finite-Alphabet [XUP08]. Finite-Differences [CMR17]. Finite-Differential [DF21]. Finite-Dimensional [AK19, BE95, BH86, CYL96, CYL97a, CYL97b, CY94, Coh97a, Coh97b, CS86, DTWY91, EK97, FRZ93, GKLM92, GH96, HM03, Ito90, Pan92, SS99a, Tan00, WF24, XB99]. Finite-Gain [LCS96, YC06]. Finite-Horizon [De 15, GRZ24, KLN21, MPZ03]. Finite-Level [MLZ17]. Finite-Precision [WCLC05]. Finite-Rank [ORW13]. Finite-Time [BB00, BFF20, Doa21, GL13, GMR22, HJF10, Jam14, Kar06, LK23, Liu14b, MNPR20, PR14, PEB19, Son79, WP18, WJE16, YW12, ZRSZG22, ZBA13, ZYGpJ23, ZD24]. Finitely [CM02, Dol76b, GTT16b, Kha95, Mey77b]. Finiteness [FK83]. Finslerian [LM00]. FIR [LM13b, MT09b]. Firm [CFR13]. First [ABCO17, AY23a, BB23, CMS16, CT09, CW20, Che04, Dok17, El 17, Fra84, GHZ15, Iof79a, IK96a, KL21, LH87, MCL^{+}22, Mai13, Maz94, MDW23, Now17, Sar97, SO06a, Tri78]. First- [AY23a, CT09, Sar97]. First-Order [BB23, El 17, IK96a, KL21, MCL^{+}22, Mai13, MDW23]. Fish [BS09b, GT07]. Fisher [HSJ19]. Fit [BP08a, DF98c, GT09, GW09, Ma92].

FitzHugh [BKR17, CRT15, SS20, BK14]. Fixed [Alm08, ARS17, BBCS22, Ber85, BH86, Ber89b, BC24, CPW88, EP24, EÖ88, Eno94b, Gen13, GL13, GH20, GM11, HG90, HP09a, Mai08, Mur89, OS02, Ols17, Pau08, PS18, PEB19, RY95, Ric77, Spi80, Ste17, VA03, WL17, YXC21, Ber88]. Fixed-Endpoint [Eno94b, WL17]. Fixed-Free [PS18]. Fixed-Order [BH86, Ber89b, GM11]. Fixed-Point [BC24, VA03]. Fixed-Time [EP24, GH20, PEB19]. Flat [NRB20]. Flatness [CLT24, MRR16, MRRR19, NR17, PR01, RM98, dSF01, vRM98]. Fleming [BSZ93, EK93]. Flexible [Bal78, BGRW88, CZ98, CCW14, CM98, GAM93, GA91, Guo01, Nai95, ZFKZ21]. Flextensional [AGN^{+}16]. Flock [CLG12, DHJK20]. Flocking [DHJK20]. Flocks [CKRT22]. Flow [AHJM10, AB02c, ACCG18, BFCLS97, BE87, BHT87, BFPB13, DH98a, FLUU17, FGH23, Gug03, GDL11, GG24, GHS92, Gut90, Hei16, HK01, Ito05, Liu08, MT03, Mar20, MMN09, Now17, Sla00, Sla03, SZ10, TBT90]. Flows [ATW93, ABB04, Ara13, BKP14, Bel23, BDK00, CN24, Cla82, GZ21, GM99, GM00, HY97a, HW18, KV07b, LS02, MCP21, Qu14, SS24]. Fluctuating [FSS87]. Fluctuations [EAK86]. Fluid [BKP14, Bra09, CM12, FNR19, GH12, HKW17, HK01, IK94, Leq13, Ray10, SZ10, Teb22, ZM11, ZG17, Del19]. Fluid-Structure [FNR19, Leq13, Ray10, Teb22]. Fluids [Ara12, Ara14, BDFCGB12, CFRV98, Sta15, YDZ98]. Flux [JS99, VP22b]. Focal [Zei92]. Fock [Zd83]. Focusing [DKL09]. Fokker [AØ23, Bar21, Bar23a, Bar23b, Ber21, BK23, LB00]. Fold [NR17]. Follicular [CCS13]. Föllmer [DJK^{+}23]. Follow [KR04]. Follower [CH94, DF20, FH20, GX19, LZ17, Pap81, WMWL22, Yon02, KS84a, KS85a]. Following [ABP14, GAP^{+}09, HK06, LWR24, Tse93, WD07]. fonctions [HU78b]. Footprint [LW15a]. Foraging [HBSW07]. Force [AT00a, Cha10, CCW14, GY98, Gue07, KB01, KBDK03, KGBS08]. Forced [BRT15]. Forces [CCM14]. Forcing [BR13, BRT15, Mey77d]. Forest [Alv04]. Form [Bol80, BDP14, BFM23, CS05, HS15, Hug16, IST16, IT17, JM05, Kal82, KK92, Kan98a, Kan98b, KS90b, LC05, LAN04, MXZ19, WCS12, Yan14, ZBB07]. Formation [CFH09, CBB19, EK09, GN18, HPT07, JZLW18, SS18, YXC21, ZM14]. Formations [AH14, CBB17b, YADF09]. Forms [AI02, CW81, FT02, HO97, KL02, MRv94, McE09, PPL09, Pul95, RT04, Reb88, TR02, WD76, WT98, WF76, Zha09]. Formula [AP20, AB17a, AB20, Bar88, Bou17, Igl96, LL07, MM00, Ris86, Sla00, Sla03, Ste07, Sto97, Wol90]. Formulae [KHP06, PP13, Pop95]. Formulas [BMS05, BLQR20, TA81]. Formulation [BK13, CDD20, CS02, DZ91, GK06, GHZ24, IK96a, LR10, PZ14, dlRHM16]. Formulations [BGS19, CB11, VL78a]. Fornasini [PEGR18]. Forth [BASG21]. forthe [FCGIP06]. Forward [AZZ20, BC16, FT04, GCL15, GTSN13, HJX18, HLT20, HJX22, LZ18a, LY18, ØS10a, PW99, Rus06, ST14, TZ09, Tse00, WWX13, WZ17c, WY21, Yon10, ZG17, ZS23]. Forward-Backward [BC16, FT04, HJX22, ØS10a, PW99, Tse00, WWX13, WZ17c, WY21, Yon10, ZS23]. Forwarding [VB23]. Fountains [CL02]. Four [GLD^{+}91, Nym94, Sta93]. Four-Block [Nym94, Sta93]. Fourier [Obe07, BGBPG04, NHG99]. Fourth [LW22, WCG14]. Fraction [Kha82, MCWY14]. Fractional [ACKM22, AC02, AO15, AOS16, AKP21, BJ17, DGO19, DHPD00, DMPD12, DPD13, FF22, GLK22, Gom20, GH20, GW19, HZ05a, HP09b, KFP22, Koe20, Liu14a, LL15, MML21, MZ06, MN22, NBF16, Ngu18, PNSL20, POU11, Qua03a, Qua03b, RGG19, SV17, TKH10, TT20, War19, Yam88, ZW21, ZH24, ZG18, ZS19]. Fractional-Order [Gom20, GH20, TT20]. Fracture [NWW17, NWW19]. Fractured [GLK22]. Fragments [WC96b]. Frames [WP98]. Framework [AAS02, BKSW24, Bas91, Bis12, Ete19, FLUU17, GŠZ05, GŠZ10, Gri93, MRP99a, NSW17, RCS01, TG23, Ves13, XC03, ZYL97, Zho07, Zd83]. Frameworks [SB78]. Fréchet [FLUU17]. Fredholm [GLM21, LY90]. Free [ANS14, AFCLS22, AMI20, AL00, BF91, BRW05, Cal18, Cér00, CGY21, CH94, CF14, De 15, DF98c, EH10, FZ19, FH94, FR78b, FH88, GZ21, GKP18, GH04, GTX22, HYZ24, HW12, JXZ24, KB81, KWY08, KQWY20, KLM04, MXZ19, MO02, McE07, McE09, MK10, MP76b, MA01, NT95, PTH88, PS18, Qu14, RV93, Sei88, Sto97, Tre89, WZHZ19]. Free-Boundaries [De 15]. Free-Boundary [AFCLS22, BF91]. Free-Endpoint [Tre89]. Free-Time [RV93]. Freedom [NCL13]. French [Bon84, GB82a, Lob70, Pen89, Qua80, Bac79, CL78, Cla86, HU78a, HU78b, LM84, Lob76, Loi88, Tah92]. Frequencies [AS91, Zer05]. Frequency [Bla76, BL00a, DE14, Hag02, HdHVW03, KvdSD21, LP15, LB89, MR09, PIH14, Reb95, Rus86b, WZ01, ZH02, Zwa88a]. Frequency-Scale [LB89]. Frequency/Period [Rus86b]. Friction [BHOP14, CDD20, CSFCL^{+}23, Lip97, NS18]. Frictional [CO03a, CCFV20]. Frictionless [SSL11]. Frictions [CR17a]. Frobenius [AD02]. Fromovitz [Gün95]. Fromowitz [dRdPR08]. Front [BSS93, DZ19]. Fronts [CM06]. Fuel [Ban05, GX19, IK02b, MS07b, Wee92]. Fukushima [FR17]. Full [AR21, AY23a, Ben76, Las98, QW19, RTT20]. Fuller [BCS20]. Fully [BFO20, CR11a, GW19, HW19, HJX18, Las95, LW14, PW99, Sor98, TP05].

Function

[ACS00b, AX05, Bad09, BSI19, Bar89, BF90a, BB18, BP98b, CFS15, Câr93, Car12, CQ05, CZA04a, CDV16, CL86, CMW06, CN10, CP16, Col82a, CPW88, Dav07, Del04, DG85, Doy94, El 17, FY76, Flo97, FBJM14, Fuj80b, GL99, GP17, Gui13, HG02, HU84, IS12, IK11a, JL76, JV06, KK23, KKP85, KF03, KM07a, Kim78, KG17, KKS93, KNW95, Laz88, LD19, LH14, LY01, LY02, MT78, MR90, Mey77d, Mey89, Mil02, MZ09, MS04b, Ngu16, NDHL24, OT83, Pen13, PQ98, RM92, RRR16, RL78, RSS22, RK89, SS89, SCCL18, Ste04, Wei05, Whi79, WZ98, Yan96]. Functional [AGJK12, BL83, Cam76, CDCCMG08, CW20, CM10a, CY24, CH78, DVPL24, EA21, Fuj80a, Gra88, IRS02, IT86, IT87, LH21, Lar01, MR78, MW04, NS05, NT17, NT18, OMFMC^{+}21, PS10b, PO00, Sal85, SP02, SWW22, Und76, UY79, ZD14]. Functional-Differential [IT86, MW04]. Functionals [Bal81, Bal86, BD11b, BC13, CDM21a, Cou00, Cra99, DDH15, DM09, Fer78, Fer81, FT12, GT22a, Hau78a, Iof77a, Iof77b, JT11, KV07b, MW22, MWY21, PS10a, Pen20, RZ14, Sla00, SYZ13, Sto76, Vil04, WK77, Zag00]. Functions [AJPR14, AAA12, AR77, BGKV80, BR88, CGW08, CKK17, CRZ^{+}18, CEK03, CS87, CLRS00, Cob10, CHPA20, DRW10, Del22, DG89, DFP14, DT92, Dun95, EUMP18, EGN85, ENW95, Fer78, FPT15, Gal00, GdOB02, Goe13, GP79, Ham21, HO01, HRW06, HdHVW03, HU82, HBY23, HKL89, JMPW20, JMPW23, JS81, KG22, KT04, KT05, Kiw97, Kob79, KL99, LHB16, LW12, LW17, ŁS85b, Mao90, MH78, MBC06, Mat76, Mif77, Mik07, MI15, MS02, Mos80, MR18b, OMS90, PB09, PP14b, PL15, PP02, PYM93, Pre96, RS10, RLXS23, Rif00, Rif02, Rot77b, SX14b, SW11, Stu86, Stu15, Stu16, TKH10, Thi91, Tse93, Tsi91, Tsi92, Vin83b, WS13a, WZHZ19, YD08, ZD14, ZK88, CL06a, HU78b, PQ00]. Fund [HY22].

Fundamental [DM15, MD15]. Funnel [Ber21, BDIW22, BDLW24, HHI^{+}13, IM10, RS15a]. Further [Guo94a, MA01, Red21, ZL24]. Fusion [LZSC24, WLZ22].

G. [JLZ98]. Gabriel [SO13]. Gain [CM91, CCF07, CE06, DRW10, GW15, GW16, Gho88, JL16, JM05, JH00, KJ07, KK19b, KEK21, KZ23, KLR09, LCS96, LFC21, LT97, LM02, LT03, LL21b, Mir21, NDHL24, Sch15a, TSNvL08, WL24, YC06, ZLD12, dFS15, CMT03]. Gain-Scheduling [Sch15a]. Gains [BI91, GW15, HSP16, MS90, TST07]. Galerkin [GJP88, Ito90, JV17, Las84, MV11, XB99]. Galerkin-like [JV17]. Galerkin-type [Ito90, XB99]. Gamblers [HHOZ19]. Gambling [HHOZ19, VZ16]. Game [Alp88, AL98, AFLO08, AA92, AVCCD22, BFR21, BB88, BB90, BB94, BC18a, BBC18, Boe05, BS09b, BF13, Car96, CS12a, DA17, ELO22, ED81, FHH22, Fit79, FT22, Gah94, Gen13, GHS22, GX19, Ham06, HZ10a, HM14, HBSW07, Jac22, Jer80, KK18, LM15, Lid15, LAKG92, LWZC19, MP14, MP76a, NC13, Nut18, PZ14, SS21, SC16a, SS06, SBT19, Son23, Wen15, Win78, Wu18, YT13, Yan14, Yon90, Yon02, Yu15b, Zha17a, ZG19, ZG21, Zho23b]. Game-Based [Son23, ZG19, ZG21]. Game-Theoretic [SBT19]. Games [ACCD12, Ahu16, AA86, AFK10, Alp10b, AFGT13, AKTW24, AM22, AY23b, Aub90a, Ave16, Ave19, BP14a, Bar80c, Bas77, BL89, BB93, BS15, BS20a, BTB16, BP05, BH13, BY13, BR10, BCY15, BCS15, BCLY17, BCC^{+}19a, BCR24, Ber85, Ber86, Ber92b, Bis77b, Bis12, BGP23, BDT20, BDP24, BS04, BAKS18, BCR04, BL08a, BLP14, CH21a, CM16, CQ15, CP00, CQSP00, Car07, CLS12, CLLP13, CMS16, CD13b, CCHH19, CDFP19, CLRJ13, CAS13, CIP96, CLS97, Cos13, Del07, DS09, DF20, DPR22, EP08, Ell76, Eng10, Ete24, FR82b, FTV98, For77, FH17b, FH20, Fu23, FT02, Gal79, GP22, GW84, GR17b, GR05, Goe02, GYZ22, GV16, Grü13, GS11, GX19, GTX22, GHZ24, GNLB14, HG90, HM08, HHS18, HLL00, HBY21, HBY23]. Games [HHN18, Hua10, HH17, Ich76, JB96, Jas02, JN06, JN24, Kin84, KS00, KDB15, Küe07, Kul93, KS81, Kum08, Kus02, Kus04, Kus07, LN82, Lar01, LS05a, Lei06, LC05, LM16, LY18, LS19b, Lid13, LS85a, LK11, LG24, Man04, MYAS09, MM21, MSVA09, MNMS24, Moo21, NH12, Now89, NA02, Ols02, Pap81, Pap86b, PB99b, PQ00, Qia94, Ram98, RTTY23, RHCK24, Ros00a, RSV04, SBR18, SBY23, SSY24, Sap19, Sar22, SPO22, SC19, SP02, Ser09, Shi02, Sîr14a, Sîr14b, Sîr15, SBT19, Sor96, Sor98, SQS09, SY14, Sun21, SWW23, SMK18, SCFV97, ThH07, TS24, Top79, TV02, Uch78, Uch79, Und76, Var76, VA03, Vie00, WZ12a, WZ24, WC19, Whi80b, Wil77b, YXJ11, Yon88a, Yon88b, Zar84, Zha05, ZM13, Ber88, LS86, TA96, TPA97]. Gamma [BLZ17, ST00]. Gap [AC18a, BF05, BFP08, CG00b, FGS93, FM22, IM08, IM10, JHMG06, JSV05, PB83, HG02]. Gap-Metric [CG00b]. Gaps [MPR22]. Gas [GLK22, GDL11, GG24, PS19]. Gasket [Liu18]. Gauss [BCS24, HLM^{+}18, Red79, SCCL18, WZC20]. Gaussian [CK98, CGR96, CKT20, CKT24, DPD13, EK97, FCC20, GA91, GNLB14, HO01, HHN18, HH17, HWY21, KLV09, Kum83, Kwo80, Kwo91, LP85, Lou94, MM00, NH12, OC04, Rud78, Uch81, VPJ19, WZ17a]. GDP [FR20]. General [Ahm03, Akç01, Bal84, BCW21, Bau09, BCS17, BS24, BR82, BT96, BC96b, Bor94, BHL12, BC02b, ÇD21, CKT24, CLYL13, CW24a, CR93, CH05, CD12, Cur76, DC06, DR77, ES09, FS16, FLUU17, Gar76, Gil77, GLD^{+}91, GHS12, GZ19, HSA17, Ham21, HNS18, HM83, HBY23, HS87, Iof76, Kiw96a, KLV09, KM12b, KY93, Lag77, Lag89, Lev01, Li07, LL20, LK11, LR97, LtLR21, MT03, Man81, MY93, MRP99a, Mir21, Nag03, NSW17, Pen90, Qiu13, Qua04, RW94, SLG84, Sus87a, Sus87d, Tad93, Tan03, Tan15, TK81, VZ97b, WZ21, XG18, Yam89, Yu13, YCLK13, ZZ17, ZSX18, ZSS^{+}20, IM05a, LM84]. générale [LM84]. généralisés [HU78b]. Generalization [CGW01, CC13, FEBL13, Ha90, Jah88, MP11b, PY95]. Generalizations [GS93, Tsi92, Tsi95]. Generalized [AN15a, AN15b, AJ77, ABH98, Arm86, BHM14, Ber86, BP17, BN13, BLSZ19, BN16, CHN94, CFR13, CLT06, CDG15, CDKV18, Cla76a, CC90a, DM99, Dmi92, Dol76a, DS99b, DM02b, EMH08, FR17, GS19, Gui13, HSW19, Hau94, HPW83, HSP16, HN96, HMA97, IO96, ISdCN99, JS86, JS24, JV17, Kiw97, Koe79, KH69, LS98, LS00, LYLZ99, LZ17, LY18, LWR24, LT18, LR96, LRT05, MR15, MTS20, MM21, Mey09, Mil96, Mun14, NHG99, Out00, ÖL89, Pan13, RMZ02, RU19, RU21, Rei83, RW90, RZ06, Rou86, SWY12, Sas14, SBBV15, Sha94, Sha88b, Tre90, Tre99, Wei77, Zar84, ZA10, ZP05, Zhu92, ZZ23, HU78b]. Generalizing [Brü13, PIH14]. Generated [Dok05, Dun80a, LP15]. Generating [Ahm77, CEK03, GL05, Cla86]. Generation [BM84, FV98, Hab02, VPJ19]. Generator [BSI19]. Generators [Fra87a, HJX22, Las95, Mot00, Nag85, Zab78, Zab80]. génératrices [Cla86]. Generic [Aey81, Ari98, BCS20, Bre86, BP98a, BA84, CM12, Cur03, EG02, OZ00, OZ03, Pic96, Sch88a, Zas00, vdW99]. Generically [Pro22]. Genericity [AMV07, AMV18, CR13, CW24b, Hab96, Sha14, Viv03]. Genetic [Fre99]. Genetics [EK93]. Gennes [SW23b]. Geodesic [Pen13]. Geodesics [BM18a, GGP02, Hak20, Sac19]. Geometric [Aey85, Arm86, BBR18, Bd95, CR21, CdDNdA20, CdLdDM02, CB05, EC22, FFP^{+}09, FGS93, GST20, GSW89, GZ12, Gha11, Gho88, GD88, Hen80, KG19, KR04, LGZ13, Mar84, Mor10, MLYF02, NS14, PNSL20, RN13, SS92, SD13, VC03, VG95, WKS86, YH21, Gho86b]. Geometrical [CWC^{+}22]. Geometrically [RL20]. Geometry [AM14, AJ16, BJ11, BG13, DTV18, HV00, LMP95, MH78, PPZ06, Sak02, Sha83a, Sha83b]. Georgiou [LFC21]. geostrophic [FCGO05]. Get [TW03]. Gevrey [ÁRL21]. Ginzburg [ASK05, DFLZ23, FL17]. Girsanov [Dav76b]. Given [BG98, Del19, IY12, Jay10, RTT20, WR87]. Global [Aey85, AN21, BLPP24, BCS15, BKS19, BEL02, CL02, Car15, Cha09, CSFCL^{+}23, CLW17, CBB17b, CHS87, CT04, DYZ12, DJK^{+}23, DCBT88, Den98, DFLZ23, Duc20, Dun81, Dun91, Efi12, EP93, FCG06, FT04, GB86, GM91b, GM93, GH86, HG92, HW03, Hir76, HD88, Hir90, HJX18, HJX22, JL18, KT03a, Kar06, KK12, KK17a, KK19a, KK19b, KPCW22, KY05, Kum88, LCC17, LB08, LSDS13, Liu13, Liu14b, MXZ19, MRS04, MCPA17, Mat98, NZ03, PCP22, QLD01, RV90, Rei83, RZ06, Rub95, SKS90, SD13, SDSÁR00, SD15, Tsi93, Tsi95, UUH99, Vin83a, WQS^{+}24, WL24, War89, YF97, YL10, YW07, ZY08, ZKZB20, ZHB21, ZS23, ZH24, ZLD10, ZD24, vH06]. Globally [HKL89, PTH88, PLSB14, PQ98, QC95, Yak93]. Goal [CJM12, HH08, HH10]. Goal-Oriented [HH08, HH10, CJM12]. Goals [MBL00, Res76]. Godsil [OT16]. Goldberg [Bay09a]. Goldstein [Dun80b]. Goodness [BP08a]. Goodness-of-Fit [BP08a]. Gordon [LYZ23]. Gordon-Wave [LYZ23]. Gossip [BCF12, MB16, XJ24]. Governed [Ahm77, Ahm07, AG97a, AQ05, ABV13, Bar81, Ber98b, BC23b, CM81, GY11, GMS22, GW19, He87, HR21, Kha03, KFP22, KLM^{+}21, LWY21, Lou02, NWW17, NWW19, RZ98, Wac07, WW02]. GPC [KS13]. Graded [ARS09a, BS90]. Gradient [ATV13, ABB04, BK24a, BK22a, Bel23, BR17b, BT03a, CR11a, CDCCMG08, CP98, Cel07, CvdSC05, DF16, DF21, DFCGBZ02, DNS17, Dun79, Dun80a, Dun81, Dun86, Dun91, Dun93, Dun96, Dun00, FP21, FPD08, FO98, GL76, GCB06, GRZ24, HXY21, IST96, JH00, KMF19, KS92, LW15d, Man95, MCP21, NNTD14, Phe85, Phe86, RSZ23, RX15, SE21, SS24, TDBL22, TD94, TGSM18, WP18, Wol12, YHH16, Zag00, ZKZB20]. Gradient-Like [BT03a]. Gradients [HU78b, PP02, Tre90, Tre99]. Gradostat [BRS14]. Gradual [PZ20, PZ22]. Gradual-Impulsive [PZ20, PZ22]. Graph [AJPR14, BF05, CG00a, CLZ23, DES09, DJB21, FIR06, KCP16, KBvdSD17, Les12, RJME09, RHCK24, ZZS16]. Graph-Coloring [KBvdSD17]. Graph-Structured [RHCK24]. Graph-Theoretic [DJB21, RJME09]. Graphical [BB93, CMA08a]. Graphon [CH21a]. Graphs [AA86, Duc20, GZ12, GL13, GK02, Ich16, JNKV21, Lei85, LGZ13, MCL^{+}22, MBS13, MML21, MS22, OT16, Ols17, RT16, RFF15, XJ24, vdSM13]. Grassmannians [Wan91, Wan94]. Gratings [BDR02]. Graves [CDV16]. Gravity [Rei95]. Gravity-Capillary [Rei95]. Greedy [BCS21c, BCS24]. Greitzer [BH00]. Grid [HK21, KN03]. Gridding [WOH18]. Grids [LXY14]. Gross [HMMS13]. Groundwater [AVCCD22]. Group [CBB17a, Coh97b, GEF17, NS23, PBPB15, YXC21]. groupes [Bon84]. Groups [AJ16, AD17, BO09, CDM90, Da 14, Da 16, Hak20, Hir77, JZ09a, LtLR21, MDB05, MS18, ZL21, dRSV09, Bon84]. Growing [LL19]. Growth [BK22a, CR11a, CDCCMG08, DM02a, DL06, DM09, FHT06, GKY21, GL22b, MZ09, Mor13, RX15, Zhu21]. Growth-Restricted [Zhu21]. Guarantee [ZHB21]. Guaranteed [AK99b, Bar77, BD02, LMYZ23, LŻ11]. Guaranteeing [GW15]. Guarantees [FGJ^{+}19]. Guarding [Rzy09]. Guidance [TS18]. Gurtin [CFZ22]. Gyroscopic [Fre00].

H [BSZ93, LWY21, McE04]. Habit [EK09]. Hadamard [CW22]. Half [Far08, Far13, Ros00c, VJ77]. Half-Axis [Far08, Far13]. Half-Line [Ros00c]. Half-Space [VJ77]. Halfline [FG09]. Hall [CFRV98]. Hambo [HdHVW03]. Hamilton [ACL09, Aub02, ABSP08, Aub11, ABL18, BS13a, BK22a, BV17a, BL08a, BN16, CD89, CD91, CT96, CP00, CR11a, Cer01b, CI19, CV83, CB11, Day09, DZ19, EJ89, Fag08, FM00, Fra93, FL80, FL82, FMT10, Gal00, Gar03, Gom20, GRS00, HS12a, HS12b, Hau94, HJX19, JB96, Kel24, KS99a, KS98a, Kre00, LW14, LM82a, LM82b, Lor87, McE95, McE09, MDW23, Mis19, OBL11, Pen92, Pla21, Qiu18, Ram05, RW00a, RW00b, Sak02, Shi02, Tan03, VW90, Wat24, WWZ14, WY08, Ye00a]. Hamiltonian [vdSM13, Aug19, Bar78, CL02, CvdS23, CM04, Cla82, CDM23, CI86, DF98d, FMP^{+}22, GO04, Gon87a, Gro16, JKZ21, LNW23, MM04, Mac14, MD13, PV15, RT04, Roc89, Roc96, Vin14, Vin19, VS14, WTJY21, WCS12, Zhu00]. Hamiltonians [CP00, Goe05, HHS18, HLZ20, JL23, WJZ12, Zei01]. Hammerstein [DCI07, FC19, MC12, ZZWY10]. Hand [BR15a, DDH15, Kus79, SV79]. Hankel [BR87a, CL94, FV95, GBS22, GO14, HIK18, HLG93]. Hankel-Norm [HLG93]. Hard [BO14, Far08, RY95, Wac19, WM12]. Hard-Thresholding [Wac19]. Hardening [HMW12, HS03, Wac12]. Harder [Son88]. Hardness [BT97]. Hardy [Mos80, WP98]. Harmonic [GG13, GZ19, SSL11, VPJ19, Zho07]. Harvesting [Alv04, BS09b, BCS13, BS19b, CG17a, DMS03, Fra03, Mor13, SSZ11]. Hausdorff [BG13, GY12, GY13, OMFMC^{+}21]. Hautus [JZ09a, Sha14]. Having [BB03, MWY21]. Hazard [JXZ24]. Healthcare [AH21b]. Heat [AMP10, BGR07, BHS19, BGBPG04, BF89b, BHT21, CZ04a, Caz14, CS77, CLT24, CT04, CG17b, DE18, DWZ20, FG09, FCLZ16, FJ83, GBR18, GW17, GH07, HO22, Kha94, Koe20, Liu03, LW11b, MS97a, Ngu22, POU11, QW18, QWY21b, RS15a, Sch80a, Sch80b, Sch86, Sei77, TWZ16, Wan08, WZ12b, WZ12c, WX13, WXZ15, WY16, WYY21, Wec84, YZ16, Yu14, Zha01b, de 98]. Heating [DKS^{+}11]. Heavy [AK99b, AK02a, JVB^{+}23, KR89, Kus95, KM96, MK90, MSS96, PR01, PS19, VJBK24]. Heavy-Tail [JVB^{+}23]. Heavy-Tailed [VJBK24]. Hedging [BB06, BLZ17, Cvi00, Kam03, KK18, KR04, LS09, Lim05, MTT08, Pha00, Sto04, Win82]. Hegselmann [DZX24, PDT19]. Heisenberg [NS23]. Helmholtz [PS04, SAM03]. Hemivariational [LL18b, LLM15, ZMK21]. Hemolysis [CN24]. Hereditary [Ang76, BB78, Col82a, Col82b, GMP00, Gib83, MT94, Mil88, Ouv78a, Pro90, Ros84a, Tad88, War78b, Col85]. Hermite [Cob06, HO22, HPW83]. Hermitian [GLR86]. Héroult [CFRV98]. Hessian [ABB04, DZ91, Han86, KG17, Nie21]. Hessian-Driven [Nie21]. Hessians [SS09c]. Heterogeneous [APGC16, BBB24, Ete19, HHN18, KCY22, Les12, LSW^{+}17, LL21b, Mir21, MB12, Nic00, SK23, WQS^{+}24, YGC18]. Hidden [Lid13, Lid15, Ris94, XUP08, YD03, ZW21]. Hidden-Memory [ZW21]. Hider [AA92, For77, Gal79]. Hierarchical [CKT20, MMH19, SZ95]. High [BI91, Bre85, CE06, CMT03, DE14, HL04, Izm96, JS12, Kim21, KK14, Kre77b, LS98, LS00, LYLZ99, LXZ11, NS23, PSV18, WL24, YL10, ZZS14]. High-Dimensional [Kim21]. High-Frequency [DE14]. High-Gain [CE06, WL24, CMT03]. High-Order [HL04, LS98, LS00, LXZ11, YL10, ZZS14]. High-Speed [Izm96]. High-watermark [JS12]. Higher [AB06, AMR20, Ber84, BCM12b, Brü13, CKR22, Dok17, Fer05, Gro12, KP23, Kim21, LCZ20, MTS20, MG20, MG21, MPR22, PY16, War86, War88]. Higher-Order [AMR20, Ber84, Brü13, Fer05, KP23, Kim21, War88]. Highly [DTM22, Liu97c, Mar02]. Hilbert [HR83b, Ahm81, Ahm07, Alv00, Bar97, BRZ20, BSZ08, Cer01b, CFZ22, CD80, DZ76, Da 85, DGJ93, Dol76a, Dol76b, DMPD12, FP03, GS94, Gib79, GRS00, HMA81, HS11, HU19, HR83a, Ich76, Ich07, Kub85, KS85c, LR79a, LLM15, Nak83, Ouv78a, Ouv78b, SW23a, TZZ18, Vol00, WK77, XG03]. Hildreth [LC80]. Hilliard [HW14, HKW17, CFSGS15, CGS18, FRS16, HW12, RS15b, Sca19]. Hilliard [?]NavierFrigeri:2016:ODC. Hirschorn [Kum88]. History [GHS12, VZ21]. History-Dependent [GHS12, VZ21]. Hitting [Maz94, NSC19]. Hitting-Time [Maz94]. HJB [BPS05, BISW17, BB04, BM07, BHL12, CDR08, Fab08, FR17, FGG10, FGG11, GM17a, GM21, GSZ20, HMR15, KMF19, Kou24, McE07, MK10, Świ20, TZZ22]. Hoffman [Jou00a, Jou00b]. Hold [Pep14]. Hölder [CR11a, DHC19]. Hole [PDT19, PS04]. Holomorphic [MH78]. Holonomic [HS09, KM97]. Homeomorphic [Fer97b]. homogénéisé [Pen89]. Homogeneity [CGW20, EPR14]. Homogeneous [AD02, AL12, APA08, ARS09b, CHL15, CI86, DM89, EUMP18, EA21, FCJ14, Grü00, Her96, HW12, HM02, Jam14, MPS99, PA99, QS13, SA96, Sha88b, SH12, dNCP20]. Homogenization [KS97a, NPS15, NS23, PT08, ST13]. Homogenized [Toa11, Pen89].

Homological [Lom21]. Homomorphisms [AD08]. Homotopies [HV00]. Homotopy [BH96, CY99, Tod83, Wat79]. Homotopy-Smoothing [CY99]. Hop [CIH18a]. Hopf [AMZ04, BR87a, Dav79, GEF17, HKB04, ZS19]. Horizon [AT96, ABB12, AFV17, Alm08, AB10b, AK16, BB98, BS15, Bay09b, BG09, BS17, BS19a, BGP20b, BKP18, BP20c, CD89, Car90, CH96, CK22, CK23a, CK23b, CZ04b, CY12, CG13, CGY21, CMR85, DJLY09, De 15, Di 91, DS99a, DS07, DHP09, FS21, FS19, FS23, FM95, FHH97, FHH99, GQ09, GMP00, GRZ24, Gri86, GPSW10, GM09, GAHL09, HXY21, HZ22, HYZ24, HW99, HYZ21, IK02a, IK06, KLS87, Ken80, Kha16, KLN21, KPZ18, KVA96, LW10, Lei87, LZ96, LRTY22, MP14, Mam14, MPZ03, MNMS24, Mor13, OC12, PS10a, Pla80, RV19, RSZ23, SB78, SR97, Ste11, TCC19, Wee92, Wee03, WY21, WJE16, XUP08, Zas95a, Zas95b, Zas97, dFN23, DM02a]. Horizons [AK04, BT21]. Houses [VZ16]. Howard [KSS20]. Hull [MW78]. Hulls [All88, AW90]. HUM [DL09]. Human [CJM12]. Hurwitz [AJ77]. Hybrid [AW21, All99, AG97a, AHS13, AK19, BM19, BBB24, BDDSV05, CT13, CZ98, Cha77, Col11, De 00, DR05, DTM22, DNS17, FHMS18, FGJ^{+}19, GP05, GD88, GS18a, Gra99, GZ14a, HT11, HLDM20, Hua07, LB08, LNP00, LH14, LT18, LLS21, Mai08, MZ97, NNN21, PLSB14, PS17c, Rao95, RSE20, SWY12, SX14b, TC13, WZY06, YZSS23, YD03, YLL^{+}15, ZH03, ZJ06, ZY07]. Hybrid-Event-Time [LH14]. Hydrodynamics [Rav20]. Hydrogen [GG24]. Hyperbolic [Ahm77, AB02a, AB03, ABCO17, AA17, BM86, Bar95a, BS04, Bro87, CMK15, CW81, CBdN08, CGW10, CVKB13, CB15, CN19, Dat88, FBP99, FYZ07, Fu12, Gev81, GL11, Gue14, GH19, HZ22, Hay19, Ho86, Hu15, HVDK19, Kha95, KL21, KY05, KLM^{+}21, Lag77, Lag78, LT86, LR03, LY19a, LY23, MSv86, MR04b, Pav13, PR14, PU15, PGB14, SU21, SL18, Sun82, Ulb02, VJ77, WW23, Wec82, XWPG23, XLL08, You17, ZPQ20]. Hyperbox [LSDS13]. Hypercausal [Hop84]. Hyperelasticity [LSW14]. Hyperstability [Bas91]. Hypoelliptic [AM23, BDGR12, LL21a, Sto04]. Hypotheses [Cla76b, MH11]. Hypothesis [Kur89]. Hysteresis [BF89b, FH88, GQ11, KT16, LMYZ23, LM02, MMK86, Tim14, TA21]. Hysteretic [EI13, ILR10, JLR09].

I/O [LF87]. ICDD [DGQ13]. Ichikawa [JLZ98]. Ideal [CP82, II97, ISW00]. Idempotent [KM16]. Identifiability [BEL02, Cou83, GH07, KN77, KS86, KW87b, Nak83]. Identifiable [Wri93]. Identification [Aca93, AS88, Akç01, AI02, AS95, BBC81, BK82, BGRW88, BRR90, Ban78a, Ban78b, BP88a, BB07, BV82, Che84, DY13, Din15, DPV00, FP08, FZ23, GAC24, Gut90, HSD97, HG02, HSV00, HSJ19, JS99, KN03, KK14, KS85b, KNW95, KS85c, LY91b, Lam92, ML13, Mor79, MC12, MCWY14, NP82, Pie79, RV05, Ros84b, RRKJ24, SB94, SV17, WeFC22, XCZ20, Yu95, ZZWY10, Zd83]. Identifying [CZ99, CF14, GY98]. Identity [Wen15]. Ignition [IK02b]. II [AN15b, AT94, ABK21, BC96a, BS19a, BC96c, Bru76, BK00, Car97b, CK23b, Che81a, CZ00, CFNG10b, CY94, Coh97b, Cza04b, DL13a, FGG11, FHH99, Gho88, GR85b, GM17b, GM21, HS12b, HS95b, Her76, Iof77b, Iof83, IT87, Kan98b, KS85a, Kiw96b, KY05, Kru00b, LM82b, LS86, MV08b, MW21, Mey03, MRP99b, NWW19, Nic97, Ouv78a, Pap86a, Qua03b, Red18a, Red18b, RS20, RW00b, Rot77a, Rot77b, SX19, Sha83b, SC06, SD15, Son86b, Sub81, TPA97, TW03, YY08, You17, ZZ17]. III [AR77, BC81, BC96d, Kre78]. IIR [Bas91]. iISS [Bat07, MI15]. Ill [CT06, Vog90]. Ill-Posed [Vog90, CT06]. Illiquid [GGP14]. Illposed [IK95]. Image [BDGR12, HN00, RKL16]. Images [ZVBR14]. Imaginary [BFP11, CFNG10b, Lou15, MGMBN^{+}19, RR12, Sch92a, SO06b]. Imaging [AGJK12, BCM13, KK14]. Immersion [BAP18, Jou03, Lev87, Oht08]. Immobile [Gal79]. Immunotherapy [Kas09]. Impact [BLZ17, BLSZ19, GH17, KV21]. Impedance [ADK08, BP97]. Imperfect [BB08, HD11, LCF18]. Imperfections [Amm02]. Implementable [VGBS13b]. Implementation [DC88, EUMP18, ZCC07, ZH05, dOG05]. Implementations [Kiw96b]. Implementing [PTY07]. Implications [KMZ98]. Implicit [ABL18, BKP90, Ber82a, Bro22, Dun80a, FL81, FKN96, Gri88, Mac14, TBP18, ZHB21, dSF01]. Implied [EZB11]. Implies [MRS04]. Imply [SA96]. Importance [Bal94]. Important [BL08b]. Imposing [MT09b]. Improved [CC18a, YCRZ21]. Improvement [AMV20, KSS20]. Improving [BK23]. Impulse [AL08, Ama04, AKP05, Bar85b, BS20a, BEM13, BCS17, CY18, CG13, CNS21, Cob06, Cos13, DGW10, Ega08, GRR09, GW09, HZ21, HSZ24, JPS20, JS24, LF87, LP86, MSSZ88, Men80a, MR17, MR18a, MS11, Mil96, PS10a, PS17a, PPTZ19, QWY21a, RZ20, Ric77, Rob81, SS09c, Ste22, TWZ16, WYY21, XYZ21, Zhu21]. Impulse-Equity [MS11]. Impulses [BR93, JLL20, RX21]. impulsional [LM84]. impulsionnel [LM84]. Impulsive [Ahm03, AMR20, DM13a, FMV24, LH11, LH14, LW24, MR96, PZ20, PZ22, RX21, SV97, TC13, WWS21, WŽ07, YCY13]. In-Domain [SCK10]. Inapproximability [AP20]. Inattentive [SRM16]. Incentive [Baş84]. Include [AFGT13]. Includes [Ste13]. Including [HW12, LI00]. Inclusion [BC16, CS05, Fra87b, SS03a, SV79, Vin14, Wol90]. Inclusions [AGJ^{+}13, ABAC10, AS11, Aub01, Aub02, BHS05, BBV13, CN11, CO03b, DF98b, DS99b, Don10, GLK12, GR88b, Iof24, JLR09, KT05, Lei85, Li07, LW12, LR94, Lor08, Mai13, MNPR20, Mor95, MW04, NLPT09, Ngu16, Qui92, RRS14, RS13, Rya98, SH12, Tal91, TL03, Tol18, Iof83]. Income [BGP20b]. Incommensurate [GH20]. Incomplete [CLHH05, Cvi00, EQ95, EGN85, FKZ22, GR17b, Grü13, Ham21, IK06, KLSX91, LN82, Lar01, NOV03, QV04, Ros00a, RSV04, Wu18]. Incompressible [Cor99, FLUU17, GHS92, ZM11]. Inconsistency [TWZ21, Zho23b]. Inconsistent [BCL24, CL18, DL20, Ham21, HJ21, HJZ12, HJZ17, HZ19, HW21, MZ20, Mil17, NP22, NLZK19, WZ21, WZ24, WYY17]. Incorrect [KY19, KY20, MY22]. Increasingly [MA19]. Incremental [ART21, AS21, Sto76]. Increments [Pav84a]. Indefinite [CLZ98, CZ00, Du15, Hew93, HZ03, MWY21, QZ13, RMZ02, Tre89, VSB18]. Independence [BSZ08, DHV00, KS92, Lau99, Vol00]. Independent [AE17, AT92, BC23b, CK17b, Ete24, KBvdSD17, MN07, Rin09]. Index [Bv85, MG20, MG21, Tie14]. Indexed [TYW11]. Indices [AMZ04, LMYZ23, MH78, Sha88b]. Indifference [HL16]. Indirect [AB02a, AB03, AT82, Che99, DAR15, NK94, Sag15]. Indistinguishable [AE95]. Individual [LC05]. Indivisible [SYZ12]. Induced [FPT15, LM97b, MP99, MD13, RW76, SK94, Szn93, Fre00]. Induction [DKS^{+}11]. Inductor [HS03]. Inequalities [AHR20, Ama04, Aus99, ABL18, BS80, Bar81, BCP17, BB09, BLS05a, Ber98b, BZ99a, BC95b, CHY23, CSSZ20, CY99, Che99, CCT07, Chr19, CW22, CW23c, Cos13, De 11, DVPL24, DKRV13, DWZ20, El 17, FSH16, FCG06, FP96, FR78b, Fri86, Fri87, GT17, GGF09, He87, HL11, KT00, KK06, Lar01, LWY21, Lev99, LY12, LXY14, LL18b, LP86, Liu95a, Liu95b, LLM15, Lou02, MS87, MD89, MP84, Mor03, Pen20, Pop97, RZ20, RSS22, SWY12, Shi88, ST96, Sub81, TBP18, Tse91, Vuo21, WZ17b, Yao99, Yao00, ZMK21, Zha00, ZJ06, Zha09, vWCET23, Shi90]. Inequality [Ach05, Ach08, ABSP08, BC23b, BS00, BLS05b, Den98, DG85, EMH08, GP94, HTY01, IS02, KZ00a, LMP95, LY01, LY02, PPZ13, PT87, PTH88, RH06, RW78, SS99c, Ste15, ST90, Tre99, YF97, YT13, YZ16, Yu20a, YHH16, ZLD12, vTY99]. Inequality-Constrained [IS02]. Inequaltity [WK77]. Inertial [HHF21, LM15, ZVBR14]. Inertial-Sensor [ZVBR14]. Inexact [CDG15, HU19, LS99a, NNTD14, VW02, WD07]. Infimum [MPR22]. infini [Loi88]. Infinite [AT96, Ahm96, AS88, Aih92, AFV17, Alm08, AMZ04, AB10b, AK04, Aug19, BC96a, BMS05, BW89, Bar78, BD83, Bar91, Bar95b, BL01, BM90, BB93, BB94, BS15, BSDR97, BM09, BH86, BS17, BGP20b, BT21, BKP18, BASG21, BL00a, BQT08, BK98, BK99, BK00, CD89, CD91, CT96, CC04, Câr93, CHJ87, Car90, CH96, CK22, CK23a, CK23b, CZ04b, CLYL13, CyY21, CFZ22, CdVF05, CP76, CS86, Cur86, Da 87, DHP19, Di 91, DS99a, DS07, DD03, Eva13, Fab08, FR17, Fat94, Fat97, FFRR21, Fei76, FR78a, FG15, FS21, FM95, FHH97, FHH99, FB01, Fuh76, GQ09, GJP88, Gib80, Gib83, GR88a, GCP88, GT90, Gri86, Grü09, GKLM92, GT05, GM09, GT14, GT22b, GLO19, Hak20, HP10, HCG20, HZ22, HLGH98, HMR15, HYZ21, Ich16]. Infinite [IST16, IP06a, Imm07, IP87, Ito90, JZ99, JNPS18, JMPW20, JMPW23, JB96, Kam76, KZ23, KY87a, KLR09, Ken80, KA06, Kno76, Kno77, KS98a, KS00, KRT11, KPZ18, KVA96, LP15, LH21, LY90, Lei87, LZ96, LK10, LZ92, LI94b, LRT98, LR00b, LM02, LRT05, Log13, Mam14, MARRH23, MS98, Mir21, MNMS24, MS97b, OMS90, OW93, OW96, OC12, OC04, Opm07, OS10b, OS12b, OS14, OS19, PS17b, Pau17, PGB14, Pla80, PMH92, PH92, PS87, RV19, Sch92b, SWZ24, Sha15, SWW22, SR97, Sta93, SS03b, Tak97, TW94a, TW94b, TF08, TF11, Tud90, TCC19, UUH99, UU00, Und76, Van94, VB23, VS20, VK81, WS20, WWS21, Wee03, WY21, WR00, XB99, XS96, Yam88, Yam89, Yan95, Yon93, Zas95a, Zas95b, Zas97, Zwa88a, dFN23]. infinite [BC96d, FR82a]. Infinite-Dimensional [Ahm96, AS88, Aih92, Aug19, BC96a, BMS05, BW89, Bar95b, BM90, BSDR97, BH86, BASG21, CT96, CP76, CS86, DD03, Eva13, Fab08, Fat94, Fat97, FFRR21, Fei76, FG15, Fuh76, GJP88, GR88a, GCP88, GT90, GKLM92, GLO19, HP10, IST16, IP06a, Imm07, Ito90, JZ99, JNPS18, JMPW20, JMPW23, JB96, Kam76, KZ23, KLR09, Kno76, Kno77, KS98a, KRT11, LP15, LY90, LZ92, LI94b, LRT98, LR00b, LM02, LRT05, Log13, MS98, Mir21, OMS90, OW93, OW96, OC04, Opm07, OS10b, Pau17, PGB14, SWZ24, Sta93, Tak97, Tud90, UU00, Van94, VB23, VS20, WS20, WR00, XB99, XS96, Yam88, Yam89, Yan95, Yon93, Zwa88a, UUH99, FR82a]. Infinite-Horizon [AT96, BS15, BKP18, CH96, HZ22, Ken80, LZ96, MNMS24, Pla80, Zas97]. Infinite-Time [CyY21, CFZ22]. Infinitely [Lid15, Rot77b]. Infinitesimal [CR93, Fra87a, MERD02]. Infinity [BGM97, GM11, MV91, Loi88]. Inflation [CH98]. Inflow [ACCG18]. Influence [BFF20, JFB17, Sap19, TWB23]. Information [AR21, AT92, AT94, BC96a, BFR21, Baş84, Ber92b, BC17, BMM85, Car07, CKT20, CIP96, Chr80, CFZ10, EGN85, FLJ13, FKZ22, FPS17, FPD08, GKY21, GR17b, Grü13, GNLB14, Gup20, HD11, Hau87, HN00, HBY21, HBY23, HSJ19, HO08, HWX09, Jac22, KCP16, LN82, Lar01, LNW23, MT09a, MT03, MTT08, MS07a, Nie14, Now89, ØS12a, Qia94, QV04, Ros00a, RSV04, SBY23, SZ19, TZ07, Wil77b, Wu18, XG18, XZ07, Yan93, YXJ11]. Informationally [Bas77]. Informed [BCR24]. Ingersoll [BPS05]. Ingham [KL99]. Ingham-Type [KL99]. Ingredients [Kou24]. Inhibitory [KB16]. Inhomogeneous [CCFV20, CR11b, DES09, Hay19, HVDK19, MST16, Osh06, SL18]. Initial [AFT91, Amm02, BG98, BGR07, Del04, Eva13, FSGHH20, FZZ18, Kob76, Kob77, LF15, MP21, Ngu22, OP96, Whi79, Yon10]. Initial-Final [FZZ18]. Initial-Terminal [Yon10]. Initial-Value [Whi79]. Initialized [Bel23]. Injection [CC18b, HG92, HS15, Jou03, OS10b]. Injection-Suction [CC18b]. Injections [FS19]. Injective [BAP18]. Inner [CCM14, Pet04]. Inner-Inner [Pet04]. Innovation [Uch81]. Innovations [VJBK24]. Input [AD08, ART21, AISW04, BM00a, BP88a, BRS14, BFP09, BFT96, Bos02, CT13, CRZ^{+}18, Cha05, CWQ18, CWKQ24, CHL15, CABP17, CW03, CM76, CS86, DM13a, DL13b, De 83, DBC12, Dew76, DMS02, Dun86, EP24, FCG18, FK83, FM16, GB86, GB93, Gra88, GP17, GLO19, Had06, HP21, HLZF21, HZB13, HG02, HR83a, HJF10, HJST24, HSP16, HHN18, ITT04, JLZ98, JNPS18, JMPW20, JMPW23, JLR09, Jay10, KMC21, Kar11, KKA23, Kob83, Kob84, KSW01, KM12b, LW15b, LZC21a, LCS96, LS99b, LF15, LRT98, LR00b, LM02, MRL12, MRS04, MARRH23, MT94, MRv94, MNB11, Mee89, MKK19, NVD18, NZ99, NR17, Nie14, OS10b, Pan92, PD79, Pet88, Por77, PC00, PS87, QN82, SL16, SW00, Sus83, Sus87b, Sus87c, Sus87d, Tad89, Tad97, TR02, Tal05]. Input [Teg84, TG12, Tsi93, WS95, WC96a, WWS21, Was79, WL19, WTJY21, WSR04, YC06, Yam88, Zam98, Zha08, ZZ20, ZLD10, HR83b, Mai03, NRL22, ZK20, Obe06]. Input-Output [AD08, AISW04, BP88a, DMS02, Gra88, LF15, MRv94, NVD18, OS10b, Por77, PC00, TG12]. Input-Output-to-State [KSW01, LZC21a]. Input-To-State [GLO19, CT13, Cha05, DM13a, EP24, HP21, HJF10, HJST24, HSP16, JNPS18, JMPW20, JMPW23, JLR09, MRS04, MARRH23, MKK19, SL16, WWS21, ZZ20]. Input/Output [ART21, BFT96, De 83, GP17, KM12b, NZ99, PD79, WS95, WSR04, Yam88, Zam98, Obe06]. Input/Output-to-State [ART21]. Inputs [AZ07, BBT14, CM17, CW03, DI87, DZA19, Dun98, FZ23, Gra07, JL16, JNKV21, LSDS13, MM80b, NT16, NRB20, SDSÁR00, TD94, WC96a, dFS15]. Insensitivity [Dág18]. Insensitizing [BGBPG04, Dág06, FCGO05]. Insights [RN13]. Inspection [Öze85]. Inspired [CG22, SS21]. Instabilities [BHM99, Fre00]. Instability [Cau13, HM83, Mat05, NP06, Vid77]. Installation [KV21]. Instantaneous [GH17, Hin05]. Instationary [KV07b]. Instructions [BCY15]. Insufficient [CIP96]. Insurance [AAM20, CGY21, CM07, MS11, NS20]. Integer [KLM20c, LR79b, MS23, Mey77c, MW78]. Integrability [BM18a, BC98, BCM12b, DvdS98, NMI15]. Integrable [BS09a]. Integral [Alt09, AISW04, AB05, AAS02, Bal81, Bal86, Bv85, BHS19, BR15b, CF95, CM10a, CW08, CHPA20, Cou00, DGO19, DO14, Fer81, FS16, Gib79, Gri83, Gug15, HHT19, Iof77a, Iof77b, JMP21, JT11, Jun00, LPT22, LY20b, LRT98, LR00b, LT03, MARRH23, Man21, PY96, Rya98, Sub81, Vil04, WZ17c, WYZ23, WY23, ZPQ20].

Integral-Input-to-State [AISW04]. Integral-Type [AB05]. Integrals [Bag84, Bou13, CP02, CEK03, Fer05, Hau78a, LM82a, LM82b]. Integrands [Bal81, Bal86, Cou00, LZ96, Mar02]. Integrated [LR23b]. Integration [Art89a, BDGK21, SP96, vdSR11]. Integrator [CHL15, DF16, ZRSZG22]. Integrators [CUB20]. Integro [Ach08, BS18, BHT22, Bre16, Kim93a, Wan22, ZVBR14, Zho23a]. Integro-Differential [Ach08, BHT22, Bre16, Kim93a, Wan22, Zho23a, BS18]. Integro/Partial [ZVBR14]. Integrodifferential [DI93, GLZ04]. Intensities [AW21, KR04]. Intensity [CDDP11]. Interacting [BR10, EP19, RL06]. Interaction [CLG12, FNR19, KL16, MG13, ZHH17]. Interactions [DG16, LBMH^{+}16, SS21, SK23, Teb22]. Interactive [New06]. Intercepting [FFP^{+}09]. Interconnected [ADS24, FFS20, GL05, GM11, HM83, KHP06, LD05, MI15, RG14, RLXS23, TG12, Vid77]. Interconnection [LG09, Obe09, RW01, TFT11, ZL02]. Interconnections [CJK13]. Interest [BPS05, KK02, LS91a]. Interface [BKP14, BGHR16, CFRV98, LZ16c, MPT06, MY09, Nic96, Nic97, DGQ13]. Interfaces [KP94b]. Interference [CIH18b, MN07]. Interior [BS11, CFM19, DHV98, HZ95, IST96, Jar93, LS99a, Mal23, MOF77, Set93, SS03b, UUH99, UU00, Wei05, ZL01]. Interior-Point [DHV98, Jar93, UU00, UUH99, ZL01]. Interleaving [Sch15b]. Intermediate [HHS18, LE17]. Intermittent [LO15, LMZ20]. Intermittently [HMV05]. Internal [ABCO17, And82, BRS11, BD11a, CLMSR00, EÖSÖ98, Fab92, Fu12, GZ23, HPS07, Imm07, Kha95, KB01, Kim92, KR15, LM22, LZ19, Nic97, NP06, PP10, PP14a, Qua03b, RS20, ZY08, Zha18]. Internally [GBS22, WZ12b, WZ12c]. Interplay [EAK86]. Interpolating [KN13]. Interpolation [BL00b, CQ05, GK87b, Gho86a, LM97b, Pav84a, Pav84b, Rav00, Rus06, SKD99, Ste13, Ste15, Tad89, Tha86]. Interpolator [KG80]. Interpolators [KG80]. Interpretations [BV10]. Intersection [Bre07]. Intersections [BHPM04]. Interval [AFLO08, BBTN03, BK98, BK99, BK00, DV12, LK10, MNB12, RZ09, WSZ^{+}21, Wil77a]. Intervals [Bar78, CHJ87, FP03, PIH14, PK23]. Intervention [LS19b]. Intractable [PT86]. Intrinsic [BLW07, BFP08, BR21, MDB05, RL20]. Introduction [BVJ07]. Invariance [BD17b, Coh97b, CK09, CCS18, Cur86, Da 14, DCBT88, Fon89, GTSN13, HMA97, ISW00, Kaw11, MRL12, MAG11, MH11, NS91, PZLM24a, PZLM24b, Ram94, Rya98, Sei87, WHS19, Zwa88a]. Invariant [BM18a, BC05b, BR08, CP00, CWC^{+}22, CWKQ24, DM89, DGKL10, Din15, ER88, Emr80a, FP03, GLK12, KPCW22, KHJ14, Kor20, Kra93, KV07b, KS84b, LM97b, OS10b, Pav80, QS13, RG14, RCS01, Sha18, Sny77, TK81, TSNvL08, VPJ19, VW02, WC96a, Yük24, ZVBR14, ZK88, Zwa88b, vdW99, Kre85, EH80, Hir81]. Invariants [BFZ00, Bis83, DGM89, HPW83, MH78, Nto12, Sha88b, TR02, WF76]. Invasion [ALNP24]. Inventory [BG09, BÇS07, BÇMS^{+}08, CS12b, FFR23, HSZ15, LYB18, CY18]. Inventory/Production [CY18]. Inverse [Ach05, Ach08, ASTT09, Amm02, BD15, BE13, BGR07, Cau13, Caz14, CT06, Doy94, FK96, FN84, GMM91, HT10, IK95, JHB19, Kat19, KS86, Nic00, NM10, PHL16, PA93, PM16, SSW15, ZG17]. Inverse-Square [Caz14]. Inverses [WDS79]. Inversion [AY23a, FKN96]. Invertibility [Fei80, Hir77, Hir79]. Invertible [Bat95, GY16a]. Investigation [AA97]. Investment [AMS96, ARS17, BM21, BBCS22, BS17, BS19a, BC17, CLHH05, CH09, CFR13, CF14, CSZ13, DJLY09, DK13, Dok05, FP14, FPS17, GGP14, Ham21, HYZ24, HLG21, JS12, KM12a, LZ13, MXZ19, MZ07, NS20, TST07, Zar92, Zar94, dMCJ22]. Investment-Consumption [AMS96, Zar92]. Investment/Consumption [GGP14]. Investments [KMZ98]. Investor [KLS87, LS91a]. Involving [BLP14, CSCV18, Cos13, DFCGBZ02, Hua10, MM80a, Pen20, SWY12, Teb22, WT83, YD03]. Inward [CI19, CKR22]. Irreducible [Mor24a]. Irregular [HBD21, IS02, Rei86]. Irreversible [CH09, CFR13, CF14, KM12a]. Isaacs’ [LS86, BL08a, CP00, DL06, HHS18, JB96, LS85a, PZ14, Pla21, Shi02, Sîr14a, Sor93, Sor98]. Isidori [IST16]. Isolated [Stu86]. Isolation [BCG^{+}17b]. Isometric [Hak20]. Isomorphism [GSW89, SBY23]. Isospectral [BDK00]. Isothermal [GS18b]. Isotropic [CD15]. ISS [AAS02, DRW10, GW16, KK17b]. Issue [BCP09]. Issues [Che23]. Itô [Bag84, DMS12, GM05, Hau78a, JPSB24, ØS12a]. Iterated [Ber98a, JH00, Liu97c, LL94, MZ22a]. Iterates [KY93]. Iteration [Alm08, AMV20, AB12, ABK14, AB17b, ADG^{+}11, Ber98c, CC96, DKL09, DBS05, McC09, Mun07, NF13, ORW13, SR04, TS24, Yu15a]. Iterations [BC24, BSZ08]. Iterative [Com00, DM83, DC06, FGQ97, Fis11, ID87, LP87, Pap86b, SWY12, SD19, Wac19, War82]. IV [CYL96, Sei77].

J [AW90, Ber88, BF89a, Col85, Cou83, Cur85a, FR82a, Flå95, GK95, GRC82, Kum88, Mos82a, Pen88, Shi90, SC81, ZZ94]. J. [BHJ83b, Dye77]. Jack [Ano95]. Jacobi [ACL09, Aub02, ABL18, BS13a, BL08a, BN16, CP00, EJ89, FM00, Fra93, FL80, FL82, FMT10, Gar03, Gom20, HS12a, HS12b, HJX19, JB96, KS99a, Kre00, LW14, LM82a, LM82b, Lor87, McE95, McE09, Pen92, Pla21, Qiu18, Shi02, Wat24, WWZ14, WY08, ABSP08, Aub11, BK22a, BV17a, CD89, CD91, CT96, CR11a, Cer01b, CI19, CV83, CB11, Day09, DZ19, Fag08, Gal00, GRS00, Hau94, Kel24, KS98a, MDW23, Mis19, OBL11, Ram05, RW00a, RW00b, Sak02, VW90, Ye00a, Zei84]. Jacobian [JL98]. Jammers [Yan93]. Jamming [CIH18b]. Jeffreys [CZZ24]. Joint [AJPR14, GX04, XZ20]. Joints [Reb95]. Jump [AØ23, AC13, AC17, AMR14, ACPZ19, Bay09b, BX12, BEM13, Bis12, BV77, BIRS20, CDP23, CZ95a, CWZZ17, CG13, CdV02, CdVF05, DGR18, Dav76a, DGW10, DL13a, EDM05, FB01, FR05, Fra03, Gad08, GGF09, Goo76, HLZF21, Hop87, HO08, HLG21, KMF19, Kus79, LD06, LK08, LYB18, ML12, Moo21, Mor11, Mur86, NC16, OM14, PZ20, RTF21, She85, TF08, TF11, WD79, WZ12a, WZ17a, XZ17, YX10, ZES12, ZSX18, ZWYJ14]. Jump-Diffusion [Bis12, DL13a, HLG21, KMF19, Kus79, Moo21, ZES12, ZSX18]. Jump-Diffusions [BIRS20, CDP23, ML12]. Jumping [FC05]. Jumps [BTY11, Lim05, ØS10a, SK94, STW20, Son23, TL94, TBP14, ZDM20, ZS23]. Junction [ACCG18, Lew80]. Junctions [BM19]. Jurdjevic [Flo02].

Kadomtsev [RS20]. Kalman

[AC19, ABP14, BD17a, Bol80, Bou93, CA10, DKT17, Dok05, GTQ21, GGA^{+}17, JKSZ21, JH14, LR16, New06, Zor17]. Kantorovich [BS03]. Karhunen [Dür98]. Karman [Las98, PT95]. Karoui [CF14]. Karush [Gün95, IK11a]. Kazantzis [AP06]. KdV [BT17, Mun14, PCP22]. Keeping [ZCGT10]. Kelvin [Bur20, EB24, HLY24, LL98, LZ16c, LZ18b, YH21]. Kernel [AF21, BP08a, BHS19, BP20b, BH17, PP00, RRKJ24]. Kernels [AS13, ACPZ19, Aub01, BCM12a, Che16, Din15, LR23b, MRL12]. Killing [ALNP24]. Kind [De 11, HJX19, LH87, LH88, WY08]. Kinds [WX13, Zha15]. Kinematic [BL05, HMW12, Wac12]. Kinetic [BK23, CO22]. Kinodynamic [ADL23]. Kirchhoff [AZ08, GZ14b]. Kitamura [Cou83]. KKT [KTY18]. Klein [LYZ23]. Kleinman [BSZ08]. Knothe [Mik12a]. Knowledge [FPD08, WP18, WL19]. Knowledge-Gradient [FPD08, WP18]. Known [BBE99, HK94, IRT07, KCP16]. Koopman [BH23]. Koopman-Based [BH23]. Korteweg [AC18b, CFZ17, Cer07, CRZ13, Gag16, KCBP21, MRRR19, MCPA17, ÖB19, Ros00c, RZ06, Sun96, Xia19, Zha99]. Krasovskii [DVPL24, EA21, HP21, MH11, RX19]. Krasovskii-Type [MH11]. Krause [CFT12, DZX24, PDT19]. Kravaris [AP06]. Kravaris/Luenberger [AP06]. Kreiss [AN20]. Kronecker [EOF07, IT17, MH78]. Krotov [Vin83b]. Kuhn [Gün95, IK11a, CFR13, DJ86, DT92, LP79a, LP81]. Kullback [BM18b, CBM23, JPSB24]. Kuramoto [CIP18, CMP15, DB12, GCL15, JHB19, Kim21, aJM18]. Kushner [IR00]. Kutta [SP96]. KW [Lom18]. Kyle [ÇD21, CKL23, MSZ18]. KYP [BM06, PIH14].

Labor [BGP20b]. Lack [ALR99, Koe20, Now89, Tri77, Tri80]. Lag [GZ14a]. Lagrange [Flå95, Ang76, AK15, Bar93, BCM12b, Bus84, CP98, Cel02b, DM09, Flå92, HM76, KG17, Mor95, Roc96, RT06, RT07, Thi97, Tre99, Trö99, VZ97a, War86]. Lagrangian [ABL13, BK97b, BGP23, BR93, CM04, CZ99, DZ91, Gag16, Han86, IK90, IK96b, KSW18, KM97, NNTD14, PAH88, PM16, RM98, Sar22, Stu15, Vol00, VW02]. Lagrangians [Cha10, CO22, Dav07, DG79, KO18, Mal02]. Lags [Qia94, Tad89, YXJ11]. Lambert [RSS22]. Laminated [WXY05]. Landau [ASK05, DFLZ23, FL17, SW23b]. Langevin [GXZ22]. Language [KGM93, WR87]. Laplace [CKL16, Dat80, DSI22, JPP14]. Laplacian [CDCVV21, DGO19]. Large [ACKM22, BM84, BCL22, BHHE03, BLP14, CC01, Cha14, DRW10, DA17, DK85, DK86, HYZ11, Her96, HMR15, Hua10, LS99a, Les12, LNW23, LP87, PZ04, RBA22, RX15, WZ12a, DK89a, DK89b]. Large-Population [Hua10, LNW23, WZ12a]. Large-Time [Her96]. LaSalle [BHJ83b, MH11]. Laser [FZ19]. Last [GH14]. Last-Passage [GH14]. Lattice [CD80, Sha01]. Lattices [CFG21]. Lavrentiev [BC13, CM96a, Fer05, HM88, Zas06, Zol92]. Law [Ahm19, ACCG18, ADM17, Ber98a, BC98, Cor95, GP98, JS99, JH00, Mil88, Ris86, WL21]. Laws [AM98, AN21, AK02b, Bat95, BKP18, BC02a, CE98, CR17b, Dic76, Dus21, GL97c, KY05, LGP15, Léa12, LY19a, LY23, MPS99, Per12, PU15, Pfl86, Pop95, SU21, Ulb02, ZPQ20]. Layers [Hab02]. Leader [FH20, LWR24, Pap81, WMWL22, Yon02]. Leader-Follower [FH20, Pap81, WMWL22, Yon02]. Leader-Following [LWR24]. Leaders [LWDB20]. Leadership [BCC^{+}19a, DZX24]. Leading [LY09]. Learning [ABB01, ABB02, BM00c, BF15, CLRJ13, CAS13, Don24, DMZ22, Ete24, FMS20, FEBL13, HWZ23, KB99, KDB15, KH81a, KOS04, LN82, LHH23, LO15, LC05, MYP14, SBT19, SN82, STZ24, TS24]. Least [All88, AW90, Bar80a, Ber98a, Che84, Guo94a, MD15, MT22, Set93, Vog90, WLZ22, WDS79, XG18, Yu13].

Least-Squares [All88, AW90, MT22, WLZ22, WDS79]. Lebesgue [Fer78, MW22]. Lefschetz [Sav05]. Left [Ban19, BM18a, BC05b, CG00a]. Left-Invariant [BM18a, BC05b]. Legendre [Eke77, IT86, IT87, Lic13]. Legendre-Tau [IT87]. Leibler [BM18b, CBM23, JPSB24]. Leibler-Quadratic [CBM23]. Lemma [BM06, CZ94, PIH14, PDB23, Van94, ZZS14, ZzJh10, Las97]. Length [Dun80a, Dun80b, Gug07, HT92, MPV18]. Lens [BR08]. Level [AB03, BH99, BS09a, DZA19, EMP^{+}16, GZ14a, Ish88, JR93, Kiw96a, LAM19, LAM21, Liu11, MAJ18, MLZ17, dG06]. Level-Set [GZ14a, Liu11, dG06]. Levitin [Iof79c]. Lévy [ACPZ19, BL02, BL11, DTM22, HHPY16, JPSB24, LYZ15, MM16, ØS12a, PYB20, Red18b, CT06]. Liao [DHX10]. Liao-Type [DHX10]. Lie [AD08, Bon84, BO09, JS91, AL01, AJ16, AD17, Bon84, BR01, CDM90, Cla86, Da 14, Da 16, GJ11, Her78, Her80, Hir77, JS90, Liu97c, MDB05, MPS99, MS02, MS18, MR18b, Sus83, SS18, ZL21]. Lie-algebraic [JS91, AL01, JS90]. Life [GS14, NS20]. Lifetime [BZ15a, BZ15b]. Lifted [MRP99b]. Lifting [MRP99a, MRP99b]. Light [Sag13]. Like [AS11, BT03a, BL91, Flo95, GL11, RS10, Son83, Wri89, XZGT23, GTSN13, JV17, LM22, Rod21, ZD24, Mai13, MR18b]. Likelihood [Aih92, AB93, DPDS98, EK97, Rus88]. Likely [CGP22]. Limit [AD98, ADRU22, AZ08, BJ11, BP20a, BBCS22, BN97, Ber98a, BGHR16, BG05, BF13, BLQR20, BK99, BK00, CQ15, DM02a, EB20, Fab92, GN02, Gai04, GN18, GT22b, JM05, Jön09, Kul93, Lac17, Léa12, MSBC08, MMK86, MPR17, NWW19, PTT17, PS17c, QR11, SY21, SSY24, Shi02, SCFV97, Wee07]. Limitations [DHKS18]. Limited [CFR13, DCI07, FLJ13, Kim22, LLX11, MS05, SNB09]. Limited-Range [SNB09]. Limiting [Dor86, JT11, HLR01, ZL01]. Limits [Vig97]. Line [AG95, AE95, AH14, BBBM08, BG98, BSDR97, Kus95, LA96, Ros00c, Rzy09]. linéaires [Cla86, Lob70, Lob76]. Linear [AD15, AT96, ACKM22, Ahm81, AC02, AC18a, AK99a, ABB12, All99, ALP10a, Alv01, AL08, AM81, And82, ANP83, AP92, AA92, AP94, AA17, ABCD22, Arm86, ACW83, AW83, AT82, AG19, AF21, Aug19, AD17, BCM^{+}17, BBC^{+}19, BA08, BRR02, BJ11, BGM05, BKP90, BV18, BK84, BI91, BFZ00, BS86, Bar85a, BBTN03, Bar97, BP14a, BL89, BB92, BCLY17, BFT96, BR17a, Ber82a, BHS19, Bie95, BW88, BT98, BD19a, Bis76, Bis77a, BCG84, BM00b, Bli04, BT97, BJ02, BDFCGB12, Bos02, BLZ17, BF07, Bou95, BMO15, BP20c, Bro22, BK83b, BKM14, BKM97, CK95, CL19a, CPS06, Cam76, CK98, CG00a, CG00b, CJK13, CLAL13, CGR96, CCM14, CIP18, CL87a, CG10, CLZ98, CZ00, CZ04b, CR21, Che23, CZZ24, CWKQ24, CS10b, CS02, CDL85]. Linear [CMR85, CABP17, CW81, CW03, CSV24, CM76, CN19, CdV02, CdVF05, CdSBT09, CD80, Cur85b, Cur03, CO05, DI87, DI90, DD92, Da 14, Da 16, DK05, DHX08, Dai12, DXG15, Dat80, De 00, Del86, Del07, DS09, DGKL10, DG16, DKR08, DFP14, Dic76, Die87, Din15, DHKS18, Dmi92, Doa21, Dok05, Dol76b, DV83, DL19, DL20, DMS12, Du15, DMZ22, DPR13, DMPD94, DMPD12, DPD13, DSI22, DRDG08, ER88, EMH08, ES81a, EK97, Emr82, EG02, FW93, FM12, FP21, FMS20, Fei76, FR78a, FKT86, FHH22, FCLZ16, FZ00, FGL08, FR00, FM16, FB01, FR05, FC05, Fra77, FD87, FIM09, FP03, GR06, GQ09, GCL15, GJP88, GPB91, GBGdO00, GdOB02, GC06, GGF09, GD88, GB93, Gib80, Gib83, GR88a, GA91, GRZ24, GFF17, GCP88, GK86, GYZ22, Gom89]. Linear [GL99, GH15, GY98, GK85, GP17, Gri93, GG18, Gu90, GMS22, GZ09, GHS12, GNLB14, GGA^{+}17, GW96b, HK07, Haa82, Had06, HLL^{+}17, Hal89, HMA81, HXY21, HH81, HH83, HLZF21, HZ22, Han94, HL92, HH78, Hel93, HLL94, HLL96, HLGH98, HMW12, HP94, HRPT21, HMQD99, Hop84, Hop87, HZ05a, HO08, HSP16, HJZ17, HHN18, HVDK19, Hua88, HH17, HYZ21, HWY21, HV00, Ich76, IS12, IY12, IM05a, IM05b, IP06b, IT87, II97, JLZ98, JL76, Jar93, JGC^{+}23, pJY20, JH14, JS81, JH00, Jou03, JM17, KC15, KP97, KH79a, KH79b, KKP85, Kan78, Kan82, KMP13, KY87a, Kha82, KY23, Kim78, KKLO21, KCBP21, KA06, KLV09, KZ00b, KT03b, KM86, Kom97, KL99, KS93, KLN21, KS15, Kre87, KMR10, Kub85, Kum83, Kun82, KTV16, KS13, KM04]. Linear [KS85c, Kwo80, Kwo91, Lag83a, LH21, LY91a, LY91b, LGP15, LCC17, LL21a, LY90, LD06, LB08, LK08, LR79a, LC01, Lev87, Li94a, LR03, LY18, LS19a, LR21, LNW23, LSW^{+}17, LWZ22, LB00, LZ01, LM97b, Lin07, LZ19, LNP00, LCS96, LLR01, LZ12, LW15c, LMKW17, LWY23, LT97, LRT98, LR00b, LM02, LT03, LRT05, Lom12b, Lom18, Lom21, Lou02, Lou03, LL21b, LT91, LT92, LB98, LC23, LDZ10, MSZ18, MP14, Mai03, MT23, MM79, MS87, MT78, Man81, MD89, MRRR19, MBC06, MBS13, MS05, MCS13, Maz17, MNB12, McC76, Mee89, MWY21, MS98, MLZ17, Mey77c, MZ97, Mil88, MR86, MVBC18, MSv86, MM11, Moo21, MN77b, Mor24a, Mor24b, NVD18, NE04, Nai87, Nak83, NC16, Ner21, NS05, NNSM08, Ngo09, Ngo11, NH12, NLZK19, NJD09, OT83, O’M76, Obe91, OW96]. Linear [OM14, Ols17, OC04, OS10b, OS19, Ouv78a, Ouv78b, PPZ13, PSJ18, Pau17, Pau19, PHL16, Pav84b, PPL09, Per81, Pet88, Pet02, PD92, Phi94, PSV18, PZ22, PK87, PRS08, PA93, PS87, Pro90, PK23, Pul93, Pul95, Pul96, QS13, QV13, RMB12, RRJ16, Rai99, RMZ02, RW97a, RT04, Rap22, RNK91, Red18b, RR85, RCS01, RV19, RSZ23, RN13, RC16, Ris86, Roc87, Roc89, RW90, RL78, RTF21, Rod21, RSS22, RW92, Ros00c, Rud78, RZ93, SKS90, Sak83, SD98, SS23, SB80, Sch02, SU18, SWZ24, Set93, Sha88b, SP05, SY17, SLQ^{+}23, SK94, Sny77, SS19, ST16, Son23, Sta93, Sta98a, Sta98b, SW12, SS03a, Sun81, SY14, SLY16, Sun21, SX23, SWW23, SY24b, STZ24, Tad88, Tad93, Tak97, THA02, Tan03, Tan15, TZZ88, Ter94]. Linear [Tod83, TF08, TF11, TSS^{+}22, Tre89, Tsa78, Tse90, Tse93, TW03, Uch83, Ver97, VL23, VT95, VLP24, VS20, VK81, VW95, VSB18, Wac12, WD76, WC96a, WZ17a, WYZ23, WL19, WY21, WL21, WR00, WZ09, WS13b, Wir98, Wir05, WORO00, WM12, WH20, WJZ12, XS96, XX19, YC06, Yam89, YZZ01, Yin99, Yon02, Yu15b, Zer05, ZP03b, Zha05, ZY08, ZLHZ12, Zha13, ZDM20, ZHB21, Zho84, ZK88, ZH05, Zho07, ZLD10, ZS19, ZJ89, Zhu96, ZK20, Zhu21, Zol78, Zua96, Zua98, van86, van98, vdW99, ABCO17, CCT05, CB15, Cur85a, EL20, HJZ12, Yon13]. Linear-Exponential [KLV09]. Linear-Gaussian [GNLB14]. Linear-Quadratic [AT96, BP14a, BL89, BCLY17, Bie95, Bis77a, BP20c, BK83b, BKM14, CZ04b, DK05, Del86, DMZ22, DPD13, FMS20, FHH22, Gib83, GR88a, GRZ24, GYZ22, HJZ17, IY12, JS81, KZ00b, Kun82, KM04, LNW23, LZ01, Moo21, NLZK19, RV19, Roc87, RW90, SLQ^{+}23, Sun21, SX23, SWW23, SY24b, STZ24, WYZ23, WY21, YZZ01, Yu15b, Zhu96, Zol78, DMPD12]. Linear-Quadratic-Gaussian [GA91, HHN18, HH17, HWY21, Kum83, Kwo80, NH12, Rud78, WZ17a]. Linearity [PP00]. Linearizable [Hir90, KR85, NR17]. Linearization [ABKM96, BH96, BR87b, CLM91, DN23, DTV18, HG92, KK92, NRL22, NM08, Nie14, QLD01, TR02, TGSM18, XG89]. Linearized [ASK05, BB97, CRR12, CDM21b, Cor94, KK24, KGS11, MRV03, Mic01, Xia19, aJM18]. Linearly [AF21, LHB16, LY23, Mau77, Nie21, RS21]. Link [Red21]. Lions [ZSS^{+}20]. Liouville [Far13, KS86, LL15]. Lipschitz [Ahm19, BFCLS97, BBT14, BR21, Cer01a, CW22, CDV16, Dav07, DS19, DQ21, DF98b, DS99b, DKK^{+}19, GV03, GT22a, Hag79, HY22, HRPT21, IPM10, JG99, LP10, Li94a, LK10, MS87, NT16, Rif00, Vin83b, Wol90, Ye00b]. Lipschitz-Continuity [Dav07]. Lipschitz-Type [JG99]. Lipschitzian [CV83, CV86, CC99, DH93, DH98b, ML13]. Liquid [KK24, KLM04, SW23b]. Liquid-Tank [KK24]. Liquidation [GHQ15]. Liquidity [CGY21]. Literature [Iof78]. Lloyd [PY16]. LMI [CdSBT09, GZSAB13, LHPT08, Lei01, TF08]. LMI-Based [Lei01]. LMI-Relaxations [LHPT08]. LMIs [AT00b]. Loans [ZZ09]. Local [AS03, AL12, ASR^{+}13, ABCD22, Ava19, AMI20, BBC^{+}19, Bad09, BLJ15, BBB24, Bea08, BLT15, BGR04, BP14b, Bia83, BGBPG04, BCL24, BTW82, BKS19, Bre86, BAKS18, BS06, CR11a, CRZ00, CV83, Col88, CVKB13, DN23, DDK10, Del19, DO22, DR79, Dun93, EB20, Fra87a, FO18, FGJ^{+}19, FHM84, FI98, GW84, Gon87b, GCSB99, Her76, Her78, Her82, Her96, Iof79a, Iof79c, Iof79b, Jaf20, Jak86, Jou00a, Jou00b, Kha07, KC10, Kra09, Kre00, KP94a, LS00, LY19a, LR92, LZ12, LZ16c, LZ18b, Mee89, Mol19, MA03, NM08, NS86, PS11, PS18, RZ09, Sch88b, Sch88a, Stu86, Sus78, Sus83, Sus87a, Tao18, WS90, ZPQ20, ZL07]. Localization [FCMA09, MS17b, PBPB15]. Localizations [SO13]. Localized [CSCV18, CCFV20, Caz14, DMR19, HLY24, KMC21, MS22]. Locally [ÁRL21, BDGK21, CSC14, Cor95, Gow89, Hir90, Ho90b, IPM10, Kha01, KK03, LR23a, Liu97a, LL98, LK10, ML13, PTH88, PLSB14, ZKZB20, Kre85]. Location [AHJM10, HH05]. Locations [WC91, WJE16]. Lock [Kam10]. Lock-In [Kam10]. Locomotion [CJM12, GJ20]. Locus [BGH94, Tie14]. Loève [Dür98]. Log [KSY23]. Log-Concave [KSY23]. Logarithm [Ber98a, JH00]. Logarithmic [AHR23, LL21a, Tse93]. Logic [CW95, JK06, RW87a, SZS21, WC96b]. Logical [LD19, LWLZ20, LLM^{+}24]. Logistic [BGP20a, CGM98, Mor13]. Lojasiewicz [LXY14]. Long [AMP10, Bel00, Bor89, BGS19, BLQR20, CLLP13, Cér00, CRS21, FGL08, GR06, JPS20, JS24, LZ13, PZ13, Rob81, Ste19, dVC22]. Long-Run [Bor89, BGS19, JPS20, JS24, dVC22]. Long-Term [LZ13]. Looking [ALNP24]. Loop [Bru76, CH96, Dav79, DS09, GSW89, HK21, Hin05, LWZ22, LRW96, MZ20, MNMS24, MPS99, Ols02, Rei86, SY14, SLY16, WZ21, WY21, Zwa88a]. Lorentz [MST16]. Losing [Sac19]. Loss [BET10, Cvi00, Lei06, Mor11]. Losses [GAP^{+}09, Lin14]. Lossless [CH05, EMH08, Gre92, PW02, RR12]. Lossy [FCC20]. Lost [HT92]. Lotka [HMT18, Pav13]. Low [BBR18, BL05, FLUU17, FGJ^{+}19, HK06, KLR09, KS89b, LT97, LM02, LT03, LL21b, Sch09, SU18, ZLD12, dOP03]. Low-Complexity [FGJ^{+}19]. Low-Gain [KLR09, LT97, LM02, LT03]. Low-Order [BL05]. Low-Rank [SU18]. Lower [AKN00, Bal81, Bal82, Bal84, CZ04a, Fer81, Fra93, GWY14, HBS06, Iof77a, Iof77b, KTWY12, Kre78, Zha00]. LP [BGS19]. LPM [GZLC17]. LPV [BM03]. LQ [FG09, HZ05b, HT22, MPZ03, MNMS24, OCI00, SZ97, TY23, WOH18, WTJY21]. LQ-Control [SZ97]. LQ-Optimal [WOH18]. LQG [GMP00, HHSS23, Hua10, Opm07, PC00, UP02]. LQR [DJZ96, QBP23]. LS [LL13]. LTI [AE17, BD01b, EOF07, MCWY14, RSMG17, SO06b, WQS^{+}24]. LTV [TZQ21]. Luenberger [AP06, And14, BASG21, MPI07]. Lur’e [BGLT16, GLO19, SL16, TBP14, ZP03a, ZP05, JM05, Pan98]. Lyapunov [AJPR14, APH24, AP20, AAA12, AZ97, BBC^{+}19, Bad09, Bat07, BD11b, BB03, BCG86, BH23, CGW08, CL06a, CHL15, CLRS00, CB15, DRW10, DVPL24, Efi12, EUMP18, EA21, EP24, FPT15, Flo95, Flo97, FGH23, GdOB02, GLK12, Goe13, GTSN13, HP21, HS11, HK88, JMPW20, JMPW23, KT03a, KZ23, KT04, KT05, KK06, LHB16, LW12, LW17, LD19, LSW96, LH14, Mao90, MBC06, MK21, MADF98, Mey89, MI15, MR18b, ORW13, PB09, RM92, RS10, RLXS23, Rif00, Rif02, RK89, Sha10, SX14b, Son83, SW00, Sti18, Tar90, Tsi91, Tsi92, VL23, Wir05, YD08, ZD14, ZK88, ZLD12]. Lyapunov-Based [Bat07, CHL15, KZ23, LW17]. Lyapunov-like [MR18b, Flo95, RS10, Son83, GTSN13]. Lyapunov-Type [Sha10]. Lyusternik [CDV16].

M [Fio21, Kum88, Wac16]. M-PID [Fio21]. M-stationarity [Wac16]. M1 [Coh21]. Machine [GL95]. Machines [CW95, FHH97, FHH99]. Magnetic [BCM13, MÖ14, WK13]. Magneto [Rav20]. Magneto-Hydrodynamics [Rav20]. Magnetohydrodynamic [HPS07, Lef11, Tao18]. Magnetometer [Zha08]. Magnetostatic [You13]. Magnitude [SDSÁR00]. Mahony [Mic01]. Maintaining [SNB09]. Maintenance [BD17b, KM12a]. Major [BCR24, BLP14, Hua10, NC13, SC16a]. Majorization [CWQ18]. Makers [Sza95]. Making [Cha14, LWDB20, SZ95, VL78b]. Malicious [CIH18a]. Malliavin [GM05]. Malmquist [XCZ20]. Management [BPS05, CGY21, DL13a, EGR^{+}22, Fer18, HY22]. Maneuvering [ZR23]. Mangasarian [dRdPR08, Gün95]. Manifold [CZ04a, GWY14, SM13]. Manifolds [ABHY23, AM14, BC95a, Bus84, DS89, DZ24, Fio20, Fio21, GC22, Gro16, ISdCN99, LY12, LWR24, Sac19, SS09a, TC13, TDNT18]. Manufacturing [FZ98, GAM93, Son93, Zha95]. Many [BC18a, CM02, Dol76b, GTT16b, Kha95, Rot77b, Zas06]. Map [LF87, PY95, Was79]. Mapping [BP08b, BO01, CDV16, JP01, Mey77d, ZI00]. Mappings [Ber77, BD77, GB86, JS86, MS97b, RS07b, SWY12, Tse00, Iof83]. Maps [ABB02, BC24, ER88, Fra87a, JL98, LL94, Pan13, PD79, RW97a, Yam88, vdSR11]. Marchesini [PEGR18]. Margin [Heu15, RBS21, VT95]. Marginal [BKL^{+}22, GTT16b, Mik15, HU78b]. marginales [HU78b]. Marginals [CGP22, RTT20]. Marine [BS19b]. Marked [CF13]. Market [AX05, BLZ17, BLSZ19, EQ95, FT22, GG09, KLSX91, SYZ12, ZSM10, BC18b]. Marketing [SZ95]. Markets [AX02, CLHH05, CR17a, Cvi00, GGP14, Ham21, NS18, TZ10]. Markov [KS99b, ABB01, Alm01, Alm08, AS91, AG97a, ABFG^{+}93, BS20a, BL18, BBC18, BS24, BBK00, BV82, Bor83, Bor84, Bor89, Bor94, Bor00a, BMT01, Bor07, Bud03, BLS12, BM18b, Cal18, CFN06, Can84, CC96, ÇR14, CCHP12, CW23b, CW24a, CTT16, CP13a, CH13, CK89, CdV02, CD03, CdVF05, CD08, CD10, CD12, CDP16, DR06, DS99a, DS07, DC99, DPR13, DG20, DPR22, DPDS98, EP08, EDM05, FF22, FKZ22, FB01, GGF09, GG09, Goo76, Gor15, GL97c, GC05, GHS12, GHZ15, GL19, GNLB14, HSA17, HLZF21, HJ13, HJ15, HYZ11, HSZ24, HLL94, HLL96, HLGH98, HLVAC99, HLGHLM03, HW99, HW05, HP14, HLCH18, Ich16, JV06, Jas02, JN19b, JN24, JR06b, KA06, Koe79, KB99, Küe07, Kum08, Kur89, KS98b, Lei87, LR23b, MSVA09, NC16]. Markov [NPY20, PZ11, Piu15, Piu18, PZ20, PZ22, PZ24, PRHL06, PRHL08, Ris94, Rod82, RTF21, Rot77a, SBR18, SRM16, SSO13, Sîr15, SCCL18, Ste22, Sza95, TZ07, TLPA00, TF11, VA03, VZ16, VZ21, WZ12a, WZ17a, Wat24, XUP08, YD03, Yu15a, Yu20a, Yu20b, Yus97, Zar92, Zha09, ZES12, Zha17b, ZSX18, Zit14]. Markov-Chain [Zar92]. Markov-Modulated [JR06b]. Markovian [BP05, CL18, ENJP88, FP99, FMSV16, FR05, FC05, GHQ15, HLZ15, JZ22, Koe90, LD06, LK08, LM84, LW24, Lou94, MBS13, MC91, OM14, TF08, WJW19, WC96b, Xie12, YZSS23, Yük17, ZSM10]. markovien [LM84]. Markowitz [ZY03]. Martingale [BCS18, KLSX91, MT03, ML12, Ouv78b, Sîr14a, Zit14]. Martingales [BK80, Dav76a, Gen13, GM05]. Maruyama [Sha18]. Mass [ATV13, BB19, BJO09, CZ98, CCS13, CM98, Guo01, HZ95, MD15, WQS^{+}24]. Mass-Spring [MD15]. Masses [AE18]. Master [BC18a, SZ19]. Matched [ZW18]. Matching [AK02b, DI86, Hui92, Kot94, KM12b, MPC91, Pet85, RKL18, SWP91, Sta93, TY00]. Material [CDCVV21, Lid15]. Materials [CNS10, CD15, OMFMC^{+}21]. Mathematical [Aus84, BP17, BF93, DvdS98, HT10, Hua88, IK11a, Ols76a, Out00, SS99b]. Matrices [All88, AW90, AP20, AMZ04, BC02b, BSS12, Cam76, CNS10, CLYL13, CDL85, CH05, DJB21, Emr80a, FF22, FV95, GGA^{+}17, HPW83, Hug16, Jar93, JL98, KK23, LF87, Lom13, NZ99, QWY21b, WS87, vW84]. Matrix [All88, AW90, BHPM04, BM06, Bis83, BGMS78, BK83b, CFNG10a, CCHP12, Cob10, DH22, DC06, Fle85, FV95, FO98, GGF09, IT17, JS81, Kha82, KRW04, Kra93, Lou15, LT91, MN77a, MCWY14, Nur06, Pav80, PD97, Pet88, RR92, RSS22, SC23, Sha85, Sha86, SY17, SZSC21, ST90, Sul94, Sul98, TR01, YMH94, vWCET23]. Matrix-Valued [JS81, Nur06]. Matroid [Mur89]. Matroid-Theoretic [Mur89]. Matroids [Mur87]. Maturity [CCS13]. Max [AGL08, ABB12, BN97, CLSL20, DZ88, DGKL10, DM15, FM00, FEBL13, GP82, HS91, Mat76, McC09, McE04, Qu14, GBCG98]. Max-Convex [CLSL20]. Max-Plus [AGL08, DGKL10, McC09, McE04, Qu14, DM15]. Max-Plus-Based [FM00]. Maxima [Bel23, McE09]. Maximal [Bau09, CYL96, CYL97b, ES09, GBBT78, GW15, Kre77b, MCL^{+}22, OvS00, Tse00, ZGS99, vH10]. Maximality [DEO24]. Maximally [BHM14, GHJ00]. Maximization [BC12, CDCVV21, CS87, DY13, EK09, Heu15, Kan78, KLSX91, KW00, NS18, ST14, SYZ12]. Maximizations [Bor77]. Maximizing [GCB06, HOPS87, IK95, PS17a, TZ10]. Maximum [ABVC19, Aih92, AB93, AMR20, AA97, AKP05, AK15, AK04, AMI20, BDM07, BDR02, BM90, BBB24, BCS15, BZ99a, BV17b, BB23, Bon86, BT13, CK95, CP98, CF05, Cla76b, CV87, Col82a, DYZ12, DM13b, DM06, DPDS98, EK97, ET12, FM12, Fat94, Fra87b, FO16, GJ11, GHS22, Gra99, HPW10, Hau78b, Hau87, HY95, HJ16, HJX18, HW20, HJX22, HCM05, HWX09, Iof24, Jar93, KSY23, KG19, KHJ14, KRT11, KPZ18, LS00, LY20b, Liu18, LT23, LT13, Lou03, MS04a, MS96, ØS10a, OL77, Pen90, Pop97, Rus88, Sag13, STW20, Sor98, SW21, Tad93, Tan98, VP88, WWX13, War76a, Wee18, ZES12, ZSX18, ZC24, ZS23, ZH24]. Maxmin [Win82]. MaxWeight [Mey09]. Maxwell [Kap94, Kim90, Lag89, MST16, Nic00, Rus86a, Wec00, WS13b, You17]. Mayer [Ang76, CFS15, DV83, FT13, PSV18, PS11]. McKean [AØ23, Ahm07, Bar23a, BS20b, BIRS20, CDJS23, DQ21, Lac17, LWfZ22, MZ20, PRS21, PW17, SC16b, SY24a, WHGY22]. McKendrick [HMT18]. McMillan [MH78]. McShane [BF89a]. MDPs [FP19, CS87]. MDS [Hut09]. Mean [ABVC19, ACCD12, Ahu16, ADRU22, AHKS22, AM22, AD18, BP14a, BTB16, BBC18, BCY15, BCLY17, BCR24, BCL22, BGP23, BR21, BDT20, BAKS18, BJ17, BPTT21, CH21a, CM16, CLLP13, CMS16, CD13b, CDFP19, CNS21, CMT03, DM83, FC05, FH17b, FH20, Fu23, FT22, GZLC17, GZ12, GV16, GHS22, GHZ15, GX19, GHZ24, HHN18, HH17, HWY21, JKSZ21, Kel24, Kub85, LZL02, LM16, LK23, LWZ22, Lim05, Liu11, LW15d, MSZ18, MTT08, NC16, NC13, Nut18, RTTY23, RHCK24, SBR18, SY21, SSY24, SS21, Sar22, SC16a, SC19, SK23, SY24b, TTZ23a, TS24, TY23, TFL14, VJBK24, WZ12a, WZ17a, WJW19, WZ24, Wee03, XG18, XZ07, XZZ20, Yon13, ZSX18, ZQ18, ZY03]. Mean-CVaR [GZLC17]. Mean-Field [ADRU22, AHKS22, AD18, BP14a, BTB16, BCL22, BR21, BDT20, BJ17, BPTT21, CD13b, FH20, GV16, GHZ24, HHN18, LM16, MSZ18, SBR18, SY21, SSY24, SK23, SY24b, TY23, WZ24, Yon13, ZSX18]. Mean-LPM [GZLC17]. Mean-Nonovershooting [LK23]. Mean-Square [ZQ18]. Mean-Variance [LZL02, Lim05, MTT08, WJW19, XZ07, XZZ20, ZY03]. Means [AT00a, BM11, Cor95, DR97, DG23, HMQD99, Kar06, Kar11, Mur87, Zha00, Zuy16]. Measurable [CP00, CSSZ20, Iof78, Kou24, Now85, VW90, Wag77, WZ17b, YZ16, Yu20a]. Measure [AK15, Ban19, BG13, BD77, BS09b, BCS13, CCK12, CCK13, CK14, CMT18, CK19, CK21, CSC14, CSCV18, CG17a, CG22, GY13, HSZ15, HR83a, KPV14, KTV16, KS84b, MT03, PV13, Sag15, Sag17, WHS19, HR83b]. Measure-Theoretic [WHS19].

Measure-Valued [BS09b, BCS13, CG17a]. Measurement [AC06, Bat19, CGM14, CMA08b, DSI22, EÖ88, GVL12, IS14, MP14, Sto91b, Sto92, WC81, ZLZ18]. Measurement-Based [GVL12]. Measurements [AD96, Bat07, CFRV98, DJ13, FLRO17, GY98, HM09, KP94b, LO15, NR97, RV05, SS18]. Measures [AW21, AS13, ACPZ19, AT82, BD21, BG05, BPTT21, CDDP11, FP03, GN02, Gai04, HN00, LHPT08, MS17b, NS12, Ped98, PZ24, Sha18, YS17, Yük24, Zit14]. Measuring [DV04]. mécaniques [Bon84]. Mechanical [AST05, Bul01, Bul02, CCC16, CdDNdA20, CO22, CMOZ02, CM03b, GLS21, KM97, LM97a, LAN04, MDB05, MLYF02, NL08, RR13, SS09b, Bon84]. Mechanics [Bra09, BR93]. Mechanism [RM92]. Mechanisms [WL24]. Media [DKL09, DES09, FB98, GMN22, GLK22, KMM07, LI00, Nic00]. Medium [CCFV20, Kim91, KKLO21, Sta15]. Membranes [LV97]. Memoriam [BHJ83a, BHJ83b]. Memory [ACS00a, BCT10, CSZZ17, GAC24, HCG20, LT18, YY08, ZW21]. Memoryless [Kim22, WL21]. Merit [PQ98]. Mesh [BSZ08, DHV00, Dun00, KS92, Lau99, MP11a, Vol00]. Mesh-Dependent [Dun00]. Mesh-Independence [Vol00]. Meshes [ARS09a]. Messages [Ols14]. Method [APH24, AGL08, AB03, AE17, ACV22, ABAC10, BP88b, BKSW24, BS13a, BZ15b, BBC18, BK22a, BB97, BGR90, Ber92a, Ber98c, Bis78a, BR17b, BM00c, BFL16, BP20c, BHL12, BS00, BLS05b, BFO20, Bur20, BK08, Cal09, CGW01, CGW08, Cha10, Che96, CZ99, CZ10, CLT06, CMT11, Coh21, DZ91, Del22, DGQ13, DS99b, DHV00, DKRV13, Dor90, DNS17, Dun80b, Dun96, DS01, Ega08, FP21, FB98, Flo97, FGJ^{+}19, GLX18, GPB91, GMM91, GG22, GY11, GS11, HLM^{+}18, HY22, HW19, IY12, IK90, IK04, ID87, Jar93, JX22, JV17, KSW18, KB81, KS92, KS95, Kiw89, Kob77, KW87a, KY92, Le 19, LC80, LR10, Mai08, MAJ18, McC09, McE07, McE09, MK10, Mey83, Mil06, MOF77, MNS91, Nag95, NP82, Nym94, PT87, PT08, Phe85, Phe86]. Method [PZ22, PMH92, PYM93, PTZ17, QC95, Qu14, RSZ23, RRKJ24, Rus21, Sag15, SWY12, Sch79, Sîr14b, Sîr15, Sla03, SS99c, Ste11, SD19, Sva11, TDBL22, TD94, TGSM18, Trö99, Tse00, Vog90, VZ97b, Vol00, Wac07, Wat79, WŽ07, WZC20, XZ15, YK91, ZBA13, ZC91, dG06]. Methodology [CW08]. Methods [ACCD12, ADK08, AKO17, AB02b, AF21, Aus99, BD01a, BB78, BLW07, Bau09, BKR00, BKS10, BGR04, Ber76, Ber82b, BE87, BHT87, Bis77b, Bis78b, BTW82, Bon89, BKS19, BH17, ByS17, BAKS18, BM18b, CY99, CDG15, Com00, CP04, DA02, DK05, Dav79, DVPL24, Dor86, Dun86, Dun91, Dun93, ES09, FK99, FB08, GL76, GB84a, Gho86a, Gho88, GD88, GGP02, GRZ24, GLK12, GK11, HZ07, HXY21, HHF21, Her80, HK01, IK96b, IK10, IK11b, IST96, ISdCN99, Jam91, KS18, KLSX91, KS87b, Kiw96a, Kiw96b, Kiw97, KB76, KKV11, KP94a, Kus78b, KH81a, Kus90, KM91, KM96, KVA96, Kus13, LS99a, LP87, LT92, Mal23, Mar84, McE04, MV07a, MKK19, NNTD14, Pol88, Ros84b, ST96, SY09, SYZ13, TDNT18, TF08, Tse90, UU00, VC03, VS18, VW02]. Methods [Wei05, Wri89, WM12, Yu13, ZKZB20, Gho86b]. Metric [ADG^{+}11, Bar17, BFP08, BD21, BZ96, BQ99, CG00b, FGS93, GB84a, Han77, IM08, IM10, LM00, MML21, Phe85, QV13, Sha88a, Bac79]. Metrically [CDKV18]. Metrics [AC18a, BBG12, BF05, BR08, DG16, JSV05, Qiu13]. métrique [Bac79]. Metrizable [Rou21]. Metropolis [GM93]. Metropolis-Type [GM93]. Metzler [DJB21]. Meyer [JN19a]. MF [GHZ24]. MF-OMO [GHZ24]. MFG [CR19]. MHD [GK06]. Microgrids [CM18].

Microstructures [Toa11].

Microswimmers [WK13]. Mild

[HMR15, Tri77, Tri80]. Miljutin [Iof79c]. Mills [CKRT22]. Milyutin [IK08]. Mimetic [DNS17]. MIMO [GZ13, SKD99, Yin99, ZYGpJ23]. Min [ABB12, BN97, DZ88, GP82, Mat76, PLSB14, FEBL13]. Min- [PLSB14]. Min-Max [ABB12, BN97, GP82, Mat76]. Mindlin [AZ08]. Minima [BF93, Del23, DR97, DR79, GT22a, NZ03, Stu86, SW99]. Minimal [AM22, BS80, BS86, BP88a, BRS14, Bis83, BFZ10, Bon85, Bd95, BS09a, CLZH24, Câr93, Cla76b, CMW06, Col12, DC88, FMP^{+}22, Fer97b, Fra82, GRR09, GBCG98, HCG20, HT92, IK11b, Jah85, JPSB24, LG09, LF87, Mar20, Pen13, QW18, RRR16, Ste04, TKH10, WZ12c, WZ98, dBMS10, GB84b]. Minimal-Time [AM22]. Minimality [Fei76, Lom12a, MP21, Zaf03]. Minimally [YADF09]. Minimax [AG88, AR99a, ABR00, Bar93, BJ95, CHL82, GTHLHR02, HP87, HK94, KY23, LA96, Now85, PMH92, Pol88, Stu15, UP99, UP02, Vin05]. Minimax-Optimal [HK94]. minimisation [CL78]. Minimization [Ber76, BK24b, BP17, BT03a, BL91, CCC16, DC10, DH12, Dun80b, ENW95, FKT86, GL76, Gül91, HLVAC99, HU19, HK06, IST96, Kiw89, Kiw97, KT03b, LZ18a, LT92, Man85, MNS91, NTP18, Nie21, Pfl86, Pre96, Sag17, Tad97, UUH99, Ver97, YF97, Zag00, CL78]. Minimizers [CM10a, CP02, CM05, FO18, Mar02, Mon94, MPV18, PV14, RZ14]. Minimizing [Alv00, BZ15a, BM94, CNP98, Cvi00, DK89a, GG09, Gli12, HOPS87, Jar93, KN13, LY20a, PW84, PQ98, Sha92, Wee03]. Minimum [AC13, AC17, ACS00b, AKN00, BEK89, Bar89, BF90a, BP07a, BA15, Bor84, CD13a, CC04, Car12, CP98, CN10, CP16, Cra99, DPZ91, DS89, DO14, DO22, GL99, Guo94a, HBS06, Iof79a, Iof79c, Iof79b, JKSZ21, Kur89, Kwo91, LM13a, MS03, NC16, Ngu16, Ovs14, PS13, Pog17, PLK89, SK86, SSL11, Ste17, VG95, VP15, Yu20b, ZTC16, Bac79]. Minimum-Time [Bar89, LM13a, Ovs14, SSL11, Ste17]. Minisum [CHL82]. Minkowski [Bou17]. Minor [BLP14, NH12, NC13]. Minorization [DS07]. Mismatched [Yan93]. Miss [Ben76]. Missing [KLM20a]. Mitter [CYL97a]. Mixed [AG88, ANR08, AR00, BdP21, BCC^{+}19a, BdPV16, BCL24, BGP23, CRZ00, CSCV18, CK03, CdP10, DO14, DO22, DNS17, FHH22, FNR19, GPS95, GV00, GY11, Ham06, HHN18, IT17, KMM11, KN14a, KQWY20, KLM20c, KKS93, MG20, MS11, NH12, NR15, NLZK19, RT07, SDV97, Szn95, Tib11, TV02, Ye00b, Yon10, Zei94, dlRT07]. Mixed-Hybrid [DNS17]. Mixed-Integer [KLM20c]. Mixed-Strategy [AG88]. Mixing [HW18, Liu08, Sîr14a, TYW11]. Mixture [CD15]. Mobile [AK02a, AA92, DH09, FFP^{+}09, For77, Gal79, Kha01, ZM13]. MOC [Gui13]. Modal [Bal78, CS84, Gib81]. Mode [BCG^{+}17a, CRZ^{+}18, EDM05, FEBL13, GW18, Kan00, KP04, LWY23, MVBC18, PKL13, POU11, SZX16]. Model [ASK05, ACL09, AMS96, AD08, ACV22, AGS13, ALR99, AM15, BM21, BHM99, BGRW88, BN81, Bar17, BRS14, BC23a, Bel00, BD11b, BDIW22, BDLW24, BCC^{+}19b, BGHR16, BDRR98, BHHE03, BK14, Bre16, CK02, CLZH24, CDFP19, CSB15, CG87b, CK03, CTZ03, Chr78, CCS13, CM07, CP82, CGW20, CRV09, CK17b, DD99, DXG15, DGR18, DL13a, DK13, DWZ^{+}16, DH12, DZX24, Enq04, FLJ13, FCGO05, FS10, GL22a, GLS21, GL22b, GSS19, GO14, GZ23, HL16, Heu15, HTLV17, HW20, Hua88, Hui92, Imm07, IK02b, KLM20a, Kha07, Kot94, KL16, KM12b, Kus78a, LWR24, LLSX20, Ma09, MZ07, Mey77a, MPC91, MZ09, Mor13, NWW17, NWW19, NP82, PB96, PP10, PP14a, PZ20, Ray10, RS21, Ris94, RL20, SWP91, SU18, SZ10, SN82, Sta93, SW23b, Tam80, TGSM18, WJW19, Win78, WRGA14]. Model [Xia11, XC03, YYH^{+}16, YG95, ZR23, ZCGT10, ZHC06, ZES12, ZSX18, ZW21, ZY03, Zhu92, dMCJ22, BR87a]. Model-Based [YYH^{+}16]. Model-Independent [CK17b]. Modeled [DMS12, LLC11]. Modeling [AT17, AD18, CLL20, CDKP87, GLK22, Hei88, HSZ24, Kus14, LXY16, MM04, MPN98, MÖ14, NY16, RA16, SSW15, VS14, WP98, Zer05]. Modelling [LS02, Sch92c]. Models [ACCG18, BP88a, BB08, BG09, Ber95, BT98, BK80, BL09, BL11, CFT12, ÇR14, ÇD21, CKT24, CS12b, CJM12, CKL23, CG22, DFŚ24, DK13, DI86, DPDS98, FGH23, Gom89, GLT95, GN03, GGC23, HNS18, HSD97, HSZ15, HW05, ITT04, KY20, KBDK03, Kim21, LYZ15, Lom18, MSZ18, MMN09, Mor11, Nag03, NLPT09, NPY20, PEGR18, PDT19, PT09b, PM85, PO00, SD98, SK95, TA21, VP87, WZ17a, XUP08, Zar92, Zar94]. Modes [EÖ88, GHZ98, Mur89]. modification [Loi88]. Modifications [Beh79]. Modified [Che84, DKL09, Dun86, JX22, LMPR09, ST96, Tse00, ZQ18]. Modular [RW87a]. Modulated [AJ12, FM16, GG09, JR06b]. Module [Kam76, WS87]. Modules [ISW00]. Moduli [KY87a]. Modulo [Jou03]. Molecule [BPS21]. Mollifier [ENW95]. Moment [AHKS22, AHS13, BRT15, CWZZ17, CEK03, HCG20, HLT08, PTZ17, RS07a, SI00, SI03, SRS15, SS19, VP22a, ZWYJ14, MNC14]. Moment-Driven [AHKS22]. Moments [BL08b, Ham82]. Momentum [LAM19, Sto04]. Monge [BS03, Sto04]. Monodomain [BK14]. Monotone [AS13, ABK14, ABAC10, AS11, Bar85c, BLW07, Bau09, BHM14, Ber77, BCM12a, BC16, BQ99, CH94, DDK12, Day09, DF20, Dol79, DJB21, ES09, FP89, GP22, LZ17, LT91, Mai08, Mai13, MD89, Roc76, SW11, Shi88, ST96, Tse00, Shi90, Vuo21, BBC03, KS84a]. Monotone-Follower [DF20, LZ17]. Monotonicity [Ahu16, CvdS23, FK99, GTT16a, JL23, MKK19, MS18, TY23]. Monotropic [ZYHX18]. Monster [AFLO08, Fit79]. Monte [HLCH18, BKSW24, CCHP12, KK76, Kus78b, KY92, YK91]. Moore [BH00]. Moral [JXZ24]. Moral-Hazard-Free [JXZ24]. Mordukhovich [Sd10]. Moreau [CP16]. Morphological [Lor08, Lor10]. Morse [BLPdD15, Li07, LW12]. Mortality [AH21b]. Mortensen [BK83a]. Mosco [Cou00]. Most [CGP22, LL18a]. Motion [AC02, BJ17, DHPD00, GJZ10, GL95, GH12, HP09b, JP15, LLT18, PZLM24a, PZLM24b, PS19, Shr81, ZH24]. Motions [DMPD12, Eno94b, HPS03, MM80b, Mor10, Van92, Vig97]. Motivated [GSS19, SW12]. Moving [ABHY23, ABB12, BLPP24, BC81, CN24, CCM14, CP13b, FS23, GTX22, HP22, HPS03, KCP16, Kha95, KN03, MRR13, MN22, Mur92]. Moving-Horizon [ABB12]. MP [NSW17]. MPC [FGJ^{+}19, Grü09, GPSW10, NNTD14]. MR [AW90, Ber88, Col85, Cou83, Cur85a, FR82a, Flå95, GK95, GRC82, Kum88, Mos82a, Pen88, Shi90, SC81, ZZ94]. Multi [CIH18a, CWQ18, CWC^{+}22, CWKQ24, CS84, GB93, HPT07, KK79, KM18, LC19, LMA07a, LMA07b, LZSC24, LL21b, Mai03, NS05, NR17, NRL22, RJME09, Rud78, Sta79, SC80, Tal05, YGC18, ZK20, ZLZ18]. Multi-Agent [CWC^{+}22, HPT07, LC19, LMA07a, LMA07b, LL21b, RJME09, Rud78, YGC18, ZLZ18, KM18]. Multi-Hop [CIH18a]. Multi-Input [CWQ18, CWKQ24, GB93, NR17, Tal05, Mai03, NRL22, ZK20]. Multi-Modal [CS84]. Multi-Output [CWKQ24, GB93]. Multi-Pair [SC80]. Multi-Parameter [KK79]. Multi-Perturbations [NS05]. Multi-Rate [Sta79]. Multi-sensor [LZSC24].

Multiagent [APGC16, BD17b, BD19b, Cha11, CLYL13, Ete20, GH20, GC22, LO15, LLX11, LWZC19, LWDB20, LZ21, LH24, Ma09, MZ18, MSJ15, MLZ17, SJ13, WZ12a, YHJ^{+}18, ZHH17, WSZ^{+}21, WMWL22]. Multiagents [LLC11, LLJC18, KCY22]. Multiarmed [BCW21, MR90]. Multibang [CTW18, Man21]. Multichain [HLGH98]. Multichained [Jam14]. Multiclass [KM96, MSS96, Mey01, Mey03]. Multiconnected [Gla05]. Multicontrol [DF94]. Multidelay [Pee20]. Multidimensional [AN15a, AN15b, BGM05, BH13, BTBT88, Ber21, BK16, BCS13, CG13, Che23, DGW10, FV98, FYZ07, GLX18, GQ11, GW09, GZ14b, HL16, HW21, KT03c, LZ12, Lom21, MS23, Obe06, Obe09, RKL16, RKL18, SO13, SBBV15, SR18, TZK11, WORO00, WSR04, Yan14, ZL02, Zer05, ZZ23, BG88, WHC03]. Multidirectional [AN06]. MultiDomains [FR14]. Multifunctions [Art89a]. Multigrid [ABV13, BKK02]. Multilayer [Bat19, ÖH14]. Multilayered [KMM07]. Multilevel [Rus21, SZ95]. Multilinear [CSBR21]. Multimarginal [HRCK21]. Multimodularity [dWvS00]. Multiobject [HSJ19]. Multiobjective [BJ04, BK13, CZ01, DZ19, FCG18, GAHL09, Gui13, KWY08, KQWY20, YZ23, Zhu00, KTY18, dMCJ22]. multipair [SC81]. Multiphase [BKP14, BV14]. MultiPhoton [SZX16]. Multiplayer [HM14, MYAS09]. Multiple [AD96, Bou13, CGM14, DHP09, ES81a, EL20, EOF07, EMP24, FP19, HPS03, LYZ15, LZ16b, Lid13, LWDB20, ML13, MBL00, MS05, NMI15, OVS08, PVZ09, PIH14, SX14b, SGM19a, SGM19b, ZLD10, dG06, CL06a]. Multiples [WGK11]. Multiplicative [DG16, DMS02, FCG16, FCG18, GT20, Hop87, IST96, JL18, Kha03, LZ16b, RTF21, ZHC06, ZLZ18]. Multiplicity [GGP02]. Multiplier [AK15, Ber76, BKS19, Bus84, CLdRdP11, Fer97a, HK06, KS18, KK76, Oss01, War78a, Ye00b]. Multipliers [Flå92, PV04, RW78, RT06, RT07, Rus76, Sch15a, Thi97, Tre99, War86, Flå95]. Multipopulation [AM22]. Multiprocesses [CV89a, CV89b]. Multirate [CQ05, MS05, TZZ88]. Multiresolution [HP14]. Multiresponse [Lee86]. Multiscale [CLAL13, CSA^{+}13, GN02]. Multistable [Efi12]. Multistage [BP17, O’N76, Ols76b, Ols76a, Ols76c]. Multistate [BC89, CDCCMG04, DF94]. Multitarget [CDDP11]. Multivalued [TBP14, ZI00]. Multivariable [BS86, DH84, Fra77, GBo97, GHJ00, LT03, ÖT90, SK86, WD76, WHB92]. Multivariate [BC12, CM92, DSI22, FZ23, IS14, LP85, Lom18, OT94]. Mutual [Yan93].

N [FCGIP06]. N-1 [FCGIP06]. Nagumo [BK14, BKR17, CRT15, SS20]. Naive [HHOZ19]. Nakagiri [Cou83]. Nakra [ÖH14]. Nash [AVJ21, BL89, BCC^{+}19a, BK13, BCR04, DF20, DPR22, Ete24, HK13, Hua10, Jac22, KK18, KDB15, LK11, MM21, MNMS24, NPvS13, NC13, Pap81, Pap86b, SBR18, SSY24, Uch78, Uch79]. Navier [HW14, HKW17, AFOQ21, AAR18, Bad09, BT11a, BRS11, BD11a, Bar12, BFCLS97, BB97, CMR07, CC16, CK19, CK21, Cau13, CSFCL^{+}23, CC18b, CRR12, CDM21b, DI94, DRT11, FGB95, FCGIP06, FLUU17, FRS16, FGH98, FGH05, GK98, GM99, GM00, Hin05, HY97a, HY97b, HR98, HW18, KC10, Mar20, Mol19, Ner21, NR15, Now17, PS06, Ray06, Sla03, Ton03, Wac07, Wan02, Wan21, dlRT07]. Navier-Slip-with-Friction [CSFCL^{+}23]. Near [ABK14, CK78, CL94, Dun79, GR17a, LWfZ22, SLQ^{+}23, Tie14, Tie20, WZY06, Zho98, Bel23]. Near-Best [CL94]. Near-Controllability [Tie14]. Near-Monotone [ABK14]. Near-Optimal [SLQ^{+}23, Zho98]. Near-Optimality [Zho98]. Nearest [GL92]. Nearly [CE11, KR87b, KR88, O’M76, Pic93, RS93, Tie14, VO22, YD03]. nécessaires [HU78a]. Necessarily [KPCW22]. Necessary [AX97, AX02, AR99b, BM00a, Bah08, Bar81, BJ04, Ber84, BBV13, BP20b, BO14, BDP14, Bro87, CM10b, Cha87a, CH99, Chr77, Den22, DO14, Emr82, Eng10, Fra84, FT13, FO18, Fuj86, GP05, Gil77, Gor19, GY16b, Iof79a, Iof79c, Iof79b, KWY08, KN14a, LLS90, LS98, LY91c, LR96, LZZ21, Man81, Mic80, MW02, Muñ00, Pom09, RW94, SV97, Stu86, TWY90, TL94, Urs82, VL78b, WS90, War76b, War84, Ye95, ZZ88, ZZ15, ZZ17, Zho93, Zho98, Zhu00, dRdPR08, HU78a]. Necessity [Pet85, Rya94, Zei94]. Needle [BK03]. Negative [AMR14, CP13b, ET12, HHPY16, KN14b, LYZ15, LY23, NPY20, RR12]. Neglecting [DPV00]. Negotiation [ZBA13]. Nehari [BMS05, CO05]. Neighboring [CKR22]. Neighbors [DHK20]. Nematic [SW23b]. Nested [Bis83, MPZ03, O’N76, Tam80]. Network [AC18b, AA92, BE87, BHT87, BFPB13, BD21, DH09, Ete20, FFP^{+}09, GL11, JNKV21, KM96, MS92, NJD09, PCP22, PBPB15, SPJ15, TBT90, XLL08, ZH18, ZZCaL19, dBLR14]. Networked [AHS13, CLZH24, CIH18b, CIH18a, CWQ18, FC19, GCC07, GMCL10, HM09, JGC^{+}23, KBvdSD17, Lin14, LFH15, LL21b, MT09a, RRJ16, SLQ^{+}23, TZQ21, YYH^{+}16, Yük19]. Networks [BM19, BGP09, BTB16, Bor83, BGL11, BCP09, BCF12, CM16, Cha14, Cha05, CGHS09, CSA^{+}13, CR11b, Dág04, DG16, Dol76a, Dol76b, DB12, Ete19, FFR14, FCMA09, FCC20, GP21, GH20, GP17, GDL11, GG24, HFY14, HLZ15, Izm96, JFB17, KG22, KZ23, LLS93, LM13a, LCF18, Les12, Leu99, LS02, LW15b, LW17, LCW18, LD19, LC19, LWLZ20, LZC21b, LZC22, LSW^{+}17, LLM^{+}24, LLC11, LH11, LLJC18, LZH^{+}16, LSL^{+}18, MK90, MNPR20, Mey01, Mey03, Mir21, MB12, NL08, PR14, SS09b, Sch92c, SZSC21, SN82, TLL16, TWB23, VZ09, WZHZ19, WeFC22, WM10, XC18, YCY13, YCLK13, YLYC15, ZZS16, ZM13, ZM14, ZLLC18, ZL24]. Neumann [CFZ17, CM06, CRV09, EH10, FSGHH20, HK13, HR98, KMR10, LT89, LI00, MR04b, Rus86a, Tib11, WL95]. Neural [KG22, MNPR20, YCY13]. Neutral [BFP11, CLS21, CY24, JL76, MERD02, MVZ^{+}09, MW04, NT17, NBF16, Ngu18, OT83, OVS08, PTT17, Pos17, RS07a, RBS21, RL78, Sha10, SO06a, SRS15, WZD^{+}23, Yam89]. Neutral-Type [RS07a, RBS21, SO06a, SRS15]. Nevanlinna [BL00b, GK87b, Rot96]. Newton [BCS24, HT10, MST16, Trö99, ADG^{+}11, AS11, BGP20a, Ber82b, BC98, BTW82, BSZ08, CDG15, DHV00, DKRV13, Dun80b, GG22, IK10, IK11b, KS87b, KS95, KKV11, Lau99, Mai13, PMH92, Sac85, UU00]. Newton-like [Mai13, AS11]. Newtonian [Ara14, LG07]. Nicolson [MV11, vDHV15]. Nilpotent [BBG12, Her86]. No [LZL02, NOV03, VL78b, Wil77b, AW90, Ber88, Col85, Cou83, Cur85a, FR82a, Flå95, GK95, GRC82, Kum88, Mos82a, Pen88, Shi90, SC81, Sun82, ZZ94]. No-Regret [NOV03]. No-Shorting [LZL02]. Nodal [LY23, WW23]. Nodes [LH11]. Noise [ACKM22, Ahu16, ALP10a, ACW83, AW83, BC96a, BD11a, BN97, BR13, Bor93, Bor10, BHHE03, Cal18, Cér00, CMT03, Cur76, DS87, DG16, DTM22, DMS02, DHC19, DF98d, DK86, FG09, FCG16, FP89, GM13, GT20, GT22b, Hop87, HT22, JVB^{+}23, JL16, JL18, KS84b, KR87b, KR88, Kus89, Kus14, Kwo91, LZ16b, LK23, MM00, MNC14, MNC16, MZ97, RBA22, Red18b, Świ20, Tes92, Tud90, VP15, Zha98, ZHC06, Zha08, dOP03]. Noise-Driven [Kus89]. Noise-Free [Cal18]. Noise-to-State [MNC14, MNC16]. Noises [ACKM22, DSI22, FCG18, HZ05a, IS12, IS14, LR00a, LZ21, RTF21, VJBK24, WWX13, WQS^{+}24, ZLZ18]. Noisy [Bat07, CFZ10, HXY21, HM09, MS07a, SSK02, WFH13]. Nominal [PSJ18]. Non [Ban19, BP05, BGST23, CPS06, CGR96, Cer01a, DQ21, GHQ15, GY16a, HY22, HO01, HBD21, HRPT21, JZ22, PDB23, WJW19, You17, Cla86, Lob70, Lob76, Qua80, Yük17, GX04, WW02, ZL01]. Non-autonomous [BGST23]. Non-Euclidean [HBD21, PDB23]. Non-Gaussian [CGR96, HO01]. Non-Lipschitz [Cer01a, DQ21, HY22, HRPT21]. Non-Markovian [BP05, GHQ15, JZ22, WJW19, Yük17]. Non-Quasi-Left-Continuous [Ban19]. Non-Smooth [You17]. Non-Time [GY16a]. Non-Zenoness [CPS06]. Nonanticipative [SCCL18]. Nonautonomous [AFT91, AT96, AHK24, Had06, KL21, MN77a, MN77b, Mor79, Reb77, Rod21, Sch02, Tal91]. Nonclassical [MT09a]. Noncoercive [Bal01, CP02, CM96b, Cra99, JMPW20, Mar02, JMPW23]. Noncoincidence [JT11]. Noncommensurate [MZ22b]. Noncommutative [BGM05, Cur11, Cur13, Cla86]. Noncompact [Fat97, Hag02, HS87]. Nonconcave [NS20]. Nonconforming [DNS17]. Nonconservative [LT89, LW24, Sle78]. Nonconstant [And88]. Noncontrollability [GV00]. Noncontrollable [BK83b]. Nonconvergence [BIP88]. Nonconvex [AM95, ADM17, Bal77, Bal81, Bal01, BHT22, CDCCMG04, CP02, CN05, CHPA20, Cra99, DFM15, DPT04, Eke77, FBP99, FBJM14, GCB06, HU19, LZ96, Mar02, Mat98, MW22, Mor95, Rin09, Rus21, Sag15, Sag17, SD19, TDBL22, Tol16, Tre99, UUH99, Vil04, VZ97a, Vin14]. Noncooperative [BS04, LG24, Whi80b]. Noncylindrical [HH19]. Nondegenerate [AB12, ABK14, AB17b, AKP05, DS87, FP96, FM22, Qua80]. Nondeterministic [KS97b, SK95]. Nondifferentiable [Aus84, Cha82, GM91a, MS85, PMW83, PM85, War78a, War78b, dBLR14]. Nondispersal [DHP19]. Nondissipative [CCW14, Gue03, LLR01]. Nondivergence [BFM23]. Nondominated [DK13, ES81a]. Nonergodic [BCL04, BGS19]. Nonexistence [Lou07, TWZ21]. Nonexpansive [ABB02, BC24, CQ15, QR11]. Nonexponentially [BMS05]. Nonholonomic [BC95a, BC98, BDK00, CdDNdA20, CdLdDM02, KM97, Liu97b, Sut20, YXC21, Zuy16]. Nonhomogeneous [HYZ11, Kim91, WC91]. Nonhyperbolic [Kre00]. Nonidentical [LH11]. Noninteracting [Ha88, NS86]. Noninteraction [Bat95, Wag91]. Noninterior [HK06]. Nonisolated [CC02, Cza04b]. NonLevel [RZ14]. Nonlinear [Aey85, Ahm83, AX96, AX97, AC06, Ahm07, AW21, ART21, AG97a, AM98, And14, ABH98, AAS02, AZ97, Aug19, BM00a, BT11a, BCD02, BHM99, BK82, BL83, BRR90, BB89, Bar82, Bar83, Bar88, BCG^{+}17a, BRZ20, Bar23b, BD02, Bat19, BR87b, BCP17, BS18, BLT15, BDIW22, Ber02, BBK00, BF05, Bla76, BGBPG04, BJ02, BFZ10, BHT19, Bon91, BK95, BT13, BH17, BBE99, BCS24, BLP14, BK98, BK99, BK00, Bus84, BLZ94, CGM14, CZ95a, CL02, CFN06, CFN07, CGM98, CD89, CR11a, CGLM02, Car15, CSC14, Cér00, Cer07, CRZ^{+}18, CT19, Cha77, CE98, CMC81, CYL96, CYL97a, CYL97b, CCF07, CY24, Che76, CY94, CK78, CE11, CV86, CBdN08, CA09, CdSBT09, CLZ23, DDK10, DM13a, DKS23, DS87, DCBT88, DL13b, DHV98, DI86, DGM89]. Nonlinear [DG79, DG85, DBS05, DCI07, Dol76b, DTM22, DH93, Don96, DH98b, DFCGBZ02, DM02b, EK85, EF09, Efi12, ET15, Eva13, FK99, FM94, FHMS18, FT77, Fle82, FM00, FKN96, FK83, Flo99, FT89, FGT95, Fon89, FM08, FB12, FP11b, FJ83, Fri84, FO16, Fuj80b, FS10, FHM84, GLZ04, GLX18, GKY21, GB82b, GK94, GT77, GP98, GT93, GH20, Gra05, GZSAB13, Gra81, Gra82, Gra07, GH15, GL05, GV05, Gri93, GKLW14, Grü98, GN03, Grü09, GPSW10, GH96, GZ13, Ha88, HT11, HK78, Ham82, HS15, HZB13, HBD21, Hay19, HT12, Her91, HSV00, Hir76, Hir79, Hir81, Hir86, HD87, HD88, HMQD99, HNv92, Hui92, HMA97, IRT07, IPM10, IK95, IK96b, Ito96, IR00, IK04, Jak80, Jak86, JS90, JB96, JSV05, Jay10, JQ97]. Nonlinear [JL16, JR93, Jou03, KMC21, KK92, Kan98a, Kan98b, KK16, KK17a, KK19a, KKA23, KBMS20, Kaś77, Kaw01, KCY22, Kno77, KS00, Kot94, KL16, KR85, KL02, KX02, KK03, KX04, KP94a, KY92, Las98, LHPT08, LMMR15, Laz88, Le 19, LMJ09, Lev87, LXZ11, LL13, LL18a, LMZ20, LMYZ23, LFM07, Liu95a, LZ06, LK11, Liu13, Liu14b, LLM15, LYB18, LLSX20, LFC21, Liu23, LMR97, MRL12, MARRH23, MM79, MPI07, Mar84, MR15, MRv94, MK21, MB16, MP94, MNB11, MADF98, McE04, Mey77c, MSR01, MM80b, Mil96, Mil17, MKK19, MN03, MPC91, Mot95, MR96, MS07b, MS10, Nag95, NS91, Nai95, NR21, NLPT09, New00, New07, NT17, NRL22, Nih83, NS86, NDHL24, NC13, Oht08, PEGR18, PB99a, PZ04, PRS21, PT09a, Pep14, Pet04, PAS00, PU15, Phe91, PP00]. Nonlinear [PT09b, PK95, PW18, Pon09, Poo96, Pop95, PR07, PMI06, QLD01, Reb77, RX19, RHCK24, RRS14, RT06, Ros84a, RRKJ24, RZ09, RS07b, Rub95, Sak02, SZ19, Sch79, Sei87, SC16b, Sha88a, SW23a, SI00, SI03, SD13, SDSÁR00, SD15, Sor96, Ste08, Sto91a, Stu15, SZZ97, TP05, TA81, TR02, Tal05, Tan00, TG12, Tib85, Tod83, TGSM18, TSS^{+}22, Tsi91, Tsi92, TT20, UU00, UY79, VM00, VB23, VC03, Vin80, Vin93, Viv03, Vog90, VS14, VJBK24, Wag91, WC89, WWS21, WLZ22, Wat79, WR14, Wil77a, WRGA14, WZC20, XG89, YMH94, YL10, YK91, YW12, YW07, YY05, YY08, YC24, ZA10, ZMK21, Zha98, ZC06, ZZS14, ZL01, ZYGpJ23, ZZ20, ZL07, ZD24, Zua90, Zd83, dOP03, vH10, vRM98, van82, Cla86, GK95, JS91, Kum88, Lob70, Lob76]. NonLinear [Yük16]. Nonlinearities [CG17b, LR00b, TBP14, ZR23]. Nonlinearity [GQ11, MMK86]. Nonlinearly [CT20, MZ18]. Nonlocal [AD18, BHS19, CM06, CR11a, CLLP13, CMT18, CM10a, CLW17, DG14, DKS^{+}11, EB20, EB24, FCLZ16, FR78b, FGH23, FRS16, GKLW14, MPT06, MY09, Muñ00, RS15b, SS20, SV17, ZHH17]. Nonlocalized [RS20]. Nonmatched [HKW17]. Nonmaximal [pJY20]. Nonminimal [WHB92]. Nonmonotone [BL11, TBP14, ZI00]. Nonnegative [DT92, IST96, JL18, SC23, TD94, Val01, BGLT16]. Nonnegativity [Pre96]. Nonoccurrence [BC13, Zas06]. Nonovershooting [LK23]. Nonparametric [Ban78a, Ban78b, BC23a, HSV00, PO00]. Nonpolyconvex [CM05]. Nonpositive [GL06, SC06]. nonquadratic [You87]. Nonqualified [AT89]. Nonrational [OMS90]. Nonregular [BdP21, Cel15, DO22, Hui92, JV17, NRL22]. Nonrobustness [LRW96]. Nonseparation [KR17, ZD24]. Nonsequential [AT92, AT94]. Nonsingular [Dun79, Emr80a, Sus87c]. Nonsmooth [AN06, ANR08, Bet19, CDR08, CHN94, Che04, CDG15, CC90a, DF21, DF98c, Fra84, GCB06, GVL12, HP22, HU19, HU82, HR21, IST02, IK11a, JG99, JL98, JQ97, Kiw89, LL00b, LZ16c, MS17a, MNS91, PZ94, Pen20, QC95, Rus21, Rya94, Sac85, Sd10, Stu86, Stu16, VC22, WZHZ19, WB87, War91, Wol12, ZR23, ZYHX18]. Nonsquare [Pet04]. Nonstandard [Pue09, van95]. Nonstationary [BRS11, BKR17, KR15, McE95, NP82, Now89, PD92, Rav20, Rot77a, Wac07, GR85b]. Nonstrict [Ste15]. Nonswitching [Gho88]. Nonsymmetric [BS15, VJ77]. Nontangency [BB03]. Nontangency-Based [BB03]. Nonuniform [DB12, FS23, KT03a, LL09, TSNvL08]. Nonuniformly [PS17a]. Nonunique [Bas77]. Nonuniqueness [Hen80]. Nonviscous [Cha09]. Nonzero [Ave19, BCR04, DF20, DPR22, HZ10a, HM14, Jac22, Kum08, Kus07, Man04, MM21, Ols02, Top79, Uch78, WC19]. Nonzero-Sum [Ave19, BCR04, DF20, DPR22, HZ10a, HM14, Jac22, Kus07, Man04, MM21, Ols02, Top79, WC19]. Norm [BDR94, CB15, DMS02, EÖSÖ98, FB08, GW76, HIK18, HLG93, Hay19, Khá93, Lin07, Mor24a, Pro22, QW18, RK89, Set93, SK94, Szn93, WZ12c, BR87a, Mun07]. Normal [BF76b, CL87b, CC99, Cou00, KK92, Kan98a, Kan98b, KL02, LC05, Mee89, PY95, Pan13, PQ98, TR02, ZBB07]. Normality [MPR22]. Normalized [Rot77a, XG18]. Norman [Coh97a]. Normed [GK85]. Norms [KK17b, PDB23, Red21]. Note [Bag81, Bau09, Ben78, CW23a, Che81b, De 15, DH22, Dye77, FK91, Gib80, JL14, Lag88, Tri77, Tri80, Uch79, Yük19]. Notions [Pet02]. Novel [GFF17, LR10, RSS22]. NP [BT97, BO14]. NP-Hard [BO14]. NP-Hardness [BT97]. Nuclear [BCM13, GO14]. Nudging [BASG21]. Null [AM23, BBGBO14, BHS19, CFM19, CIP18, CMP15, CSB15, Cor99, CGR10, FCLZ16, GCL15, Gue07, GR21b, Ich07, Koe20, KLM04, Le 19, Leq13, LY19a, LW22, MRR13, MRR16, MRV03, Mor98, MN08, PPZ13, PZ03, Rot77b, SB80, SGM19b, TZ09, UY79, Wan08, WD14, WY16, WYY21, Xia19, Zha01b, vN04]. Null-Controllability [AM23, CIP18, Koe20, Le 19]. Nulling [Ter94]. Number [ABCO17, BB94, BBR18, BCS20, BM94, BLP14, Del19, Dmi92, HK94, Sus79, vdW99]. numbers [Sav05]. Numerical [ACCD12, AM82, AI02, BFR21, BB78, BI91, BBC18, BV13a, BBR18, BO16, BDGK21, BK13, BHT21, BR07, BFO20, CR06, CMR07, CWW18, CMC81, CLT06, DS01, EUMP18, Gib83, GR88a, Grü96, JZ09b, JZ21, KG17, Kno79, KW13, Kus90, KM91, Kus02, Kus04, Kus07, Kus13, MT94, MRVM00, McE07, RW08, SWW22, SY09, WM12, dFN23, dJ78, Qua80]. Numerically [GL97a]. Numerics [Stu16]. numérique [Qua80]. Nutrient [GT07]. Nutrient-Phytoplankton-Zooplankton-Fish [GT07]. Nyquist [ZH05].

O [LF87]. O.D.E. [BM00c]. Object [Gad08, KK14]. Objective [Del22, ES81a, SDV97]. Objects [HK94, Lid13]. Oblique [Bar12, Rod21, TZK11]. Observability [Aey81, AB03, AK99a, ASTT09, AMV07, AMV18, AVZ20, Arm86, AG19, AF89, BCG84, BGST23, BGSS21, CPS06, CSSZ20, CyY21, CZZ24, CDM90, CS84, CdV02, DWZ20, EZB11, FS23, FL17, FL22, GST20, GCL15, GK94, GK95, GLM19, HH19, Kas13, Kha94, Kha95, KL99, LZ19, LZSC24, LPS12, LM08, Lü13, Mar77, OMK84, PI99, Pet02, Phu01, RW94, Sei77, Son79, Tri76, Viv03, WZ17b, XWPG23, YZ16, Yao99, Yao00, Zha00, ZBB07, van82]. Observable [BL18, Ben81, Bv85, Bol80, CE98, DH98a, HH83, HLR01, Kal82, MT23, VZ16, VZ21]. Observables [GR17a]. Observation [AQ05, BLR92, BS15, Cal18, CG10, Che16, CC87, Cur84, Dág04, DR87, DR77, DFLZ23, FCGO05, FP89, GZ05, GZ09, GG13, JZ04, KMM07, KS86, KS97b, KW85, LTT24, LWY23, Lü15, MSSZ88, Mil02, Ouv78a, PP14a, PW78, RS93, SM13, Sei77, SC19, Ste19, SX23, Uch81, WWX13, WCG14, WJE16, WHGY22, WJZ12, YL12, Zha98].

Observation-Innovation [Uch81]. Observations [AS95, Bor00a, Bor07, CFN06, DJZ96, ENJP88, IK06, Kha95, KS93, Kwo80, Las95, LMMY20, McC76, MR97, New00, PS88, Rud78, Sha17, SX19, TA81, WZD^{+}23, YY05, YCRZ21, YZ97, YDZ98, YLL^{+}15, ZZWY10, dOP03]. Observed [AC13, AC17, Ahm19, AD98, BEK89, BÇS07, BÇMS^{+}08, Bis82, CH99, Fle82, FP82, GP98, Hau82, JB96, KY19, MY22, Pic93, SC16a, Tan98, TL09, UP02, XZZ20, ZS23, ZH24]. Observer [AP06, APA08, BBT14, BAP18, CGM14, CW95, Cob12, CR13, FCG16, Fuj80a, GLK22, GBR18, GZSAB13, GG24, HM03, HMA81, HS15, Jam91, KX02, KX04, KGBS08, LHX12, MDB05, NT16, Rod21, TBP14, VP22a, WSZ^{+}21, WF24, XG89, ZD24]. Observer-Based [BBT14, Cob12, CR13, FCG16, GZSAB13, KGBS08, MDB05, WF24]. Observers [And14, AAR18, BA08, BR17a, BASG21, CRZ^{+}18, GK94, HZB13, HHT19, Kim78, Kob83, KR85, KK03, KGS11, MPI07, MNB12, Phe91, SO13, TYW11, XWPG23, van98, GK95]. Obstacle [BT91, BL04, Che99, CCT05, CCT07, Chr19, CW22, CW23c, CFSGS15, Cos13, HW12, Pen20, RU19, RU21, Sto97, TTZ23b, Wac16]. Obstacle-Type [Chr19, CW22, CW23c]. Obstruction [CT19]. Obstructions [BR01]. Obtain [GLX18, RY95]. Obtained [Oht08]. Occupation [LHPT08, PZ24, RRKJ24]. Occupational [BG05, GN02, Gai04]. Ocean [FCGO05]. Oceanography [BB97]. Octopus [CLL20]. ODE [DKS23, GW18, HR21]. ODE-PDE [DKS23]. ODEs [IK10, KT16, Zha13]. off [BF76a]. Offers [AX02]. Olech [CZ94]. OMO [GHZ24]. On-Line [BBBM08, BSDR97]. One [AZ07, AP20, ABCD22, AMV20, AL02, Bea08, BM09, BLT15, BBGBO14, BBT14, BDIW22, BRT15, CFM19, Car96, CM10a, CCM14, CMP15, CW08, CRR12, CDM21b, CT04, CBdN08, CB15, CN19, CM03b, Dág18, DE18, De 15, Del23, DF98b, DS99b, DWZ^{+}16, EMH08, Ega08, Fen16, FGGH15, Fu12, FT02, GRR09, GZ21, Gha11, Gue07, GG13, HG90, HRW88, Han94, Ho90b, KK19b, KB01, KBDK03, Kim91, LPT22, LY19a, Lis14, Liu14a, LG20, MRR16, MM16, Nag85, Nar84, NT16, NR17, OW00, Öze85, Pan17, PZ04, RM98, RSE20, Ros00a, SB94, TZK11, Teb22, TY00, WC89, Wec82, WM10, WORO00, WY08, Zha18, ZZ20, ZW18, dOP03]. One- [AZ07]. One-Dimensional [AMV20, AL02, Bea08, BM09, BLT15, BBGBO14, CW08, CT04, CBdN08, CB15, CN19, Dág18, DE18, De 15, DWZ^{+}16, FGGH15, FT02, GZ21, GG13, HRW88, Han94, Ho90b, KK19b, KB01, Kim91, LPT22, LG20, Nag85, Nar84, OW00, TZK11, Teb22, WC89, WM10, WORO00, Zha18, ZZ20, ZW18, dOP03, CCM14, Ega08, Fen16, MRR16]. One-Fold [NR17]. One-Sided [ABCD22, BBT14, CFM19, Del23, DS99b, LY19a, MM16, NT16, DF98b]. One-Unit [Öze85]. One-Vector-Lossless [EMH08]. Ones [Kha01]. Online [KDB15]. Only [BC20, ML13, PZ20]. onto [CMK15]. Open [CH96, Dav79, GY12, HT92, LS02, LWZ22, MNMS24, Ols02, Rei86, SY14, SLY16, VPJ19, WY21, ZCGT10, Zwa88a]. Open- [Ols02]. Open-Loop [CH96, Dav79, LWZ22, MNMS24, SY14, SLY16, WY21, Zwa88a]. Openness [GB82b]. Operating [And88]. Operation [CM07]. Operator [Ach08, Ahm81, BH96, BH23, CG10, DL09, Die87, DM15, Fra80, GT14, GQ11, GX04, HH76, Hag05, HIK18, HK21, HBS06, IM98, JMP21, MP99, MT78, OCI00, ORW13, PPZ13, RU19, RU21, Sti18, ZM14]. Operator-Valued [Ahm81, Sti18]. Operator-Vehicle [ZM14]. Operator/Norm [HIK18]. Operators [AO15, BLW07, Bau09, BHM14, BS18, BFT96, BT03a, BHL10, BCS24, BQ99, BR15b, BK08, CMV08, CFM19, CFNG10a, Che16, CW23c, CP04, CP76, CS86, DGJ93, Dol79, DMS02, ES09, Far13, FR78b, GO14, Hag02, HP94, Hop84, JZ99, JZ04, Jak86, Jun00, Kan82, KZ23, KW85, LZM05, LB00, Liu14a, Lou15, Mai08, PS87, Roc76, Tri76, Urq05, Wei89, WZ77, WJZ12, ZKW99, ZH02]. Opinion [BTB16, BHT10, CF18, HO16, JFB17, JNKV21, MMH19, MB12, SR18, XJ24, ZHH17]. Opinion-Dependent [BHT10]. Opinions [FPS17]. Opposite [Wim06]. Optical [BL03, Ito05, Kat19, ZLHZ12]. Optim [AW90, Ber88, Col85, Cou83, Cur85a, FR82a, Flå95, GK95, GRC82, Mos82a, Pen88, Shi90, SC81, ZZ94]. Optima [HW03, HWY21, WZ17a]. Optimal [AVJ21, AFT91, AS03, AB18, AØ23, Ahm77, Ahm96, Ahm03, AC06, Ahm07, Ahm19, AW21, AGL08, AAM20, AMP10, AFOQ21, AFO22, ALNP24, Alm01, AHR20, ARS17, AP20, Alv01, Alv04, AL08, AMR14, AQ05, AY23a, AGJ^{+}13, AP92, AX02, AX05, Ang76, AF18, ANS14, AO15, AOS16, Ant14, ARS09a, AMPR15, AR00, Ara12, Ara13, Ara14, AB05, AB10a, ABK19, AMR20, AB10b, AA97, AK04, AC79, AMI20, AM15, BM19, BDM07, BP10, BM21, BP94, BRR02, Bal84, Bal94, BR87a, BCD02, Ban19, Ban05, BW89, BI91, BL14, BBBM08, BBD18, BB89, BLPdD15, Bar83, BD83, Bar88, Bar91, BF91, BP97, Bar97, BRZ20, Bar20, Bar23b, BB18, BP88b, BDR94, BBC14, Bar93, BL96, BKL^{+}22, BM90, BL18, Bay09a, BS16, BBB24, BX12, BY14, BCS18, BBCS22]. Optimal [BdP21, BKR00, BMS15, Bel00, BS20b, Ben76, Bv85, Ber92a, BK97b, Ber98b, BIK99, BZ99a, BL04, Ber89a, BS87, BH86, Ber18, BF22, BV10, BV17a, BBR18, Bet19, BS17, BS19a, BGP20b, Bil19, Bin76, BF90b, Bis76, Bis77a, Bis78a, BP07b, BM84, BC17, BLY21, BdPV16, BV17b, BV77, BK80, BPZ16, BDGK21, BCM12a, BHT19, BB23, BK21, BPV19a, BPV19b, Bon86, BC89, BC95b, BZ99b, BH07, BJ10, BDP14, BG20, BG21, Bon91, BM13, BR21, BMT01, BGS19, BK16, BKK02, BC05b, Bou08, BEI10, BF07, BT17, BT13, BT21, BDT20, BL11, Bra09, BKP18, BP20c, BK23, BØ94, Bré79, ByS17, Bre86, BP98a, BCS13, BR01, BF89b, BC23b, BDDSV05, Bru76, BH18, BS13b, BN16, BLQR20, Bud03, BKM14, BCKM24]. Optimal [BPTT21, Bus84, CLL20, CK95, CDP23, CL19a, Cal09, Cal18, CBM23, Cam76, CGM98, CD89, CF91, CN11, CFS15, CQ15, CLAL13, CY18, Car97a, Car97b, CDJS23, CJS08, Car90, CFH09, CD15, CT09, CK14, CT14, CRT15, CKL16, CKK17, CK19, CW20, CT20, CK21, CK22, CK23a, CK23b, CCJ23, CHY23, CLHH05, CO03b, Cer01a, Cer01b, ÇD21, CPY09, Cha87b, CE98, CKT24, Cha14, CR17a, CM81, CG87a, CG87b, Che04, CCT05, CS12b, CCHP12, CY12, CGY21, CM96a, CH98, CJM12, CS02, CSZ13, CTZ03, CK78, Chr19, CW23c, Chr80, CSB88, CL86, CL87b, CV89a, CV89b, CV90, CW91, CN05, CdP10, CLdRdP11, CCS13, CG17a, CG22, CFSGS15, CS15, CGS18, CGHS09, CdDNdA20, CH78, Col88, CS89, CF13, CN19, CDP16, CR11b, CLZ23, DDK12, DG14]. Optimal [DGO19, DI87, Da 87, DI93, DJLY09, DYZ12, DGSW14, DF96, DF98a, De 15, DFM15, DS19, DFŚ24, De 11, De 92, De 83, De 00, DES09, DDS15, Del86, DMS03, Den22, DZ24, DK13, DI94, DBC12, DZ19, Di 91, DBS05, Die87, DHP09, DC10, DO14, DO22, Dok05, Dok17, DE84, Don24, Don96, DH98b, DKK^{+}19, DL20, DM17, DNS17, DKS^{+}11, DM13b, Du15, DQS16, DPDS98, Dun86, Dun96, Dun98, Dun00, DF98d, DJ98, DS01, DRDG08, EM19, EP08, ENJP88, Ell77, ED81, EL20, Eng10, Eno94b, FR17, FSH16, FSGHH20, Fat94, FM94, Fat97, FMP^{+}22, FGG10, FGG11, FP14, FT14, FL81, FZ19, FS21, Fer18, FS19, FR20, FGL08, Fis11, FP82, FSS87, FV89, FTV98, FTZ87, FR00, FR14, FH17a, FB01, Fra03, Fra87b, FO18]. Optimal [FZZ18, FGGH15, FK76, FL80, FL82, FJ83, Fri84, Fri86, Fri87, FHY88, FRS16, FK21, FT04, FHT06, FHT09, FMT10, FO16, FN84, FGH98, FGH05, FM22, FMV24, GKB00, GL97a, GR06, GQ09, GR17a, Gal84, Gal00, GV03, GLS21, GT07, GS94, GJP88, GAM93, GL13, GKP18, GKL20, GL06, Gib79, Gib81, Gib83, GR88a, Gil77, GW17, GJ20, Gli12, GH14, GM05, Goe05, GWY14, GHZ14, Goo76, GS14, Goz91, GM17a, GM17b, GM21, GH17, GV05, GK06, Gri83, Gro16, GLNT23, GG18, GSS19, Gu90, GM09, GM13, GT22b, GQ11, GLS05, Gug07, Gug10, GH19, GAHL09, Gui13, GGMZRB20, GGC23, GH96, GK98, GK11, GW19, GC05, GT08, GTT16a, GY16b, GTT16b, GY16a, GYBL15, Gup20, HRCK21, Hab02, HT11, HH76, HI84, HLM^{+}18]. Optimal [Hal89, HLZF21, HS12a, HS12b, HNS18, HL92, HAB78, Hau81b, Hau81a, Hau82, Hau87, HL90, HS95b, HYZ11, HHOZ19, HH05, HS91, HP22, HMW12, HSW12, HSW15, HG02, HK94, HH08, HH10, HL11, HW12, HMMS13, HW14, HKW17, HRMK17, Hin14, HK01, Hoe14, HR21, HS03, HMRR10, Hop87, HN22, HS96, HR98, HLP08, HPS03, HPT07, HZ10b, HJ16, HW18, HW20, HLCH18, HK19, HZ19, HYZ21, HW21, HLG21, HBS06, Ich11, Idc98, Iof24, IT87, IK96b, IK02a, IK02b, IK04, Ito05, IK07b, IK10, IK14, JLZ98, JS12, JX22, JGC^{+}23, JXZ24, JS81, KG80, KLS87, KN13, KS87b, KS92, KS95, KLM20a, Ken80, KS97a, Kha16, KWY08, KTWY12, KN14a, KTY18, KQWY20, KFP22, KLM20c, KA06, Kno76, Kno82, KS98a, KV21, Kog87, KT03b, KRS14, KM97, KKS93]. Optimal [KKV11, KPZ18, KLM^{+}21, Kru00a, Kru00b, KR13, Kul93, Kum83, KG95, Kun82, KV07b, KLL10, KW13, KPV14, KTV16, KM04, KM07b, KS98b, KS99b, KR87b, KR88, Kus89, KR89, KY93, LS91a, LN10, LLS90, Las84, LS91b, LHPT08, Lau99, LW10, LS07, LI00, LS05b, LD06, LK08, LW07, LS99a, Lei06, LXY16, Leu99, LS09, LYZ15, LR10, Lew80, LV16, LY91c, Li94a, LLMT02, LW14, LL18b, LS19a, LTT24, LH87, LH88, LW11a, LM82a, LM82b, LS85a, LR92, Liu08, LYB18, LT23, LR94, LT13, ŁS85b, LZ13, LM00, LR06, LMPR09, Lou02, Lou03, Lou07, LW11b, LZZ21, LSW14, LC23, MXZ19, MNORP21, MT09a, MT23, Mal97, Mal23, Mal02, Mam14, MT03, MS23, MBL00, MG20, MG21, MRVM00, MM21, Mas08, MSBC08, MR79, Mat05, Mau77, MO02, MO04, McS89, Mee89]. Optimal [MML21, MWY21, MV08a, MV08b, MV11, MMR17b, MMR17a, Men80a, Men80b, MR90, MR16, MS11, MW23, MP76b, MR04a, MR05, MPT06, MY09, MST16, MS17a, MW21, ML12, MP11a, MP84, MT08, Mik15, Mil17, MY17, Mil88, MS90, MSVA09, Mis19, MS97a, MA17, MM16, MS04a, MW04, MZ09, Mor13, Mot95, MS10, MS96, MP07, Mur86, Nag03, NN90, NS23, NSC19, NT95, NPR15, NWW17, NWW19, NP96, NS20, NYZZ17, NSW17, Now89, Nut18, OP96, Ohs13, OS02, ØS10a, ØS12a, OS10b, OS12b, OS14, OS19, OMFMC^{+}21, Osh06, ÖT90, Öze85, PS10a, PV15, PB96, Pap89, PW23, PLSB14, Pau08, PHL16, Pav84b, PT09a, Pen90, PW99, Pen20, PYB20, PU15, PTZ21, PVZ09, PW17, PRG16, Pic96, PV13, PTT17, PP02, PPTZ19, Pla80, PH92]. Optimal [PYM93, PW18, PZ13, PT06, PY96, Pul95, PV98, QBP23, QWY21b, QJL22, QV04, QR11, QV13, QN82, RRJ16, RRR16, RV00, Red79, RSE20, RV19, RTTY23, RN13, RCA12, Ric77, Rin09, Ris86, RH06, RV87, RS15b, RW78, Roc87, Roc89, RW90, Rog07, RT03, RT06, RT07, Rou21, RS93, SDV97, SY21, Sca19, Sch88b, Sch88a, Sch09, Sch80a, Sch15b, SU21, SP96, SZ82, Sei88, SL98, Ser03, SS20, ST23, Shi88, Shi90, SLQ^{+}23, SB78, SLG84, SR97, SV97, Sim76, SI03, Smi90, Sok87, Sok88, Son86a, Son86b, SSZ11, SYZ12, STW20, Sta98b, Sta78, SV79, SCCL18, Ste07, Ste92, Sto91a, Sto04, SLY16, SX23, SY24b, SW23b, Sus87b, Sus87c, Sus87d, Sza95, STZ24, TD21, TST07, TCC^{+}15a, TZ07, TTZ23a, THA02, TL94]. Optimal [Tan98, Tan03, TZK11, Tan15, TS24, TC13, TLD24, Tes92, Tha86, Thi91, TD94, TLPA00, TKXP12, Tim14, Tol16, Ton03, Tre90, TWZ16, TZZ18, TS95, TFT11, Trö99, TFL14, TWW16, Tud90, UMD88, UP99, Van92, VGBS13a, VGBS13b, VP07, VJ77, VL78a, Vin83a, Vin83b, VP88, VW90, Vin93, Vin05, Vin19, VP22b, Vrd16, VSB18, Wac07, Wac12, WW16, Wac16, Wac19, WOH18, WD79, Wan00, Wan02, WW02, WW03, Wan08, WZ12b, WX13, WXZ15, WY16, WL17, WZ17c, WZ21, Wan21, Wan22, WYZ23, Was79, WYY17, WY21, WD07, WKS86, Wil77a, WP76, Wol12, Wri95, WT83, WCLC05, WY08, WJE16, WH20, XA93, Yan93, Yan95, Ye97, Yon93, Yon13, YDZ98, You13, You17, Yu14, Yu15b, YS17, Yük19, Zas95a, Zas95b, Zas97, Zas00, Zas06, Zei84, ZZ88, Zei94, ZCGT10]. Optimal [ZZ92, ZMK21, Zha01a, ZJ06, Zha13, ZZ15, Zha15, ZZ17, ZDM20, ZKZB20, ZW21, ZS23, ZH24, Zho92, Zho93, Zho98, Zho23a, Zhu92, Zhu96, Zhu00, ZLLC18, Zhu21, Zol81, dFN23, dlRT07, dlRHM16, van87, vDHV15, EJ89, GR85a, GR85b, HS95a, KS84a, KS85a, Pen89, Qua80, Sei82, Tah92, ZZ94, dVC22, Pee20, Tah92, XB99]. Optimal-Harvesting [BCS13]. Optimality [AT89, ASZ02, AX97, ACT10, AR99b, ABK21, ARS09b, Bah08, BLJ15, BJ04, BB09, BLS05a, BT96, Bet19, BKK02, Bre85, BT03b, BT10, Bro87, BPTT21, CK02, Can84, CDCCMG04, CTU00, CM02, CT09, CKT20, CW23b, Chr77, CV83, CL86, DDT00, DMS03, Dun95, Dür98, FBJM14, FD87, FT13, FK96, FN84, GW84, Gfr06, GT93, GP79, GKLW14, GHS12, GHZ15, GY16b, Han86, HM93, HHPY16, HLL94, HLL96, HLVAC99, Heu15, Hoe12, Hoe14, Ish88, IS02, Joh05, Khá93, KN14a, KTY18, KQWY20, KK14, KR17, KR13, Kus78a, LLS90, LR00a, LS98, LWfZ22, LY09, Mal97, MYP14, MML21, MMR17b, Mey09, Mic80, Mir86, MW02, Muñ00, NM10, Now08, Now17, Pap86a, PP18, PS00, PS11, PS18, PRHL06, QN82]. Optimality [RT03, RW08, Rot77a, Sac19, SBY23, Sar97, Sha84, SI00, SZ03, SZ96, Tib85, VM00, VL78b, Vin83a, WS90, WZY06, Yon10, Yu20a, Yus97, Zei01, Zho98, dWvS00, HU78a]. Optimality-Conditions-Based [KK14]. optimisation [CL78]. Optimistic [SL22]. Optimizability [WR00]. Optimization [ADK08, AAC86, AS21, AG97a, ACT10, Ams13, ACCG18, AHH17, ABV13, ANR08, AST05, Art89a, AB10b, Bal77, BBC03, BP12, BdP21, BP04, BP06, Beh79, BJ04, BG09, BHOP14, BP13a, Ber84, Ber82b, BGHR16, BR17b, BLY19, BCL22, BTW82, BRW05, BC96b, BC96c, BC96d, BKS19, Bor78, BHT21, BH22, BV14, BCP09, BS06, Bur91b, BJO09, BGL98, CN24, CL87a, CHN94, Cha87a, CH99, Che96, CHS87, CR93, CNP98, CRS21, CM90, CMN20, Com00, CC90a, CHP^{+}11, CRST18, CB05, CC90b, DJK^{+}23, DDH15, DGR18, DG90, DDG12, Den97, DR08, DG89, DBS05, DH93, Dor86, Dor90, Doy94, Eke77, Enq04, EHS07, EGN85, FZ00, FFP^{+}09, Fer81, Fle85, FBJ09, FT89, FGT95, Fon89, FPW16, FH17a, Fuj86, GB84a, GZLC17, GMN22, GM91a, GM91b, GM93, GH86]. Optimization [GK87a, Gfr06, GR02, Gon94, GN03, GSZ20, GI04, GY12, GY13, GHZ24, GJN08, Han77, HJM07, HBD21, HHF21, HP96, HRPT21, HS87, HKL89, HQR24, IST02, IZ00, IK08, Ish88, Izm96, JS86, JL98, JR93, KSW18, Kar89, Kat19, Khá93, KP04, Kiw96a, Kiw96b, Kor20, KS07, KLN21, Kry98, Kum88, KW13, KK76, Kus78b, KS85c, KY92, LM15, LWY21, LY90, LB08, LCF18, Lei87, LM13b, LL20, LNT02, LE17, LLT18, LLLT18, LY01, LY02, LW15d, Man95, MNC16, Mat98, MAJ18, MV07a, MRV11, Mif77, Mil96, MR97, MU22, MS85, Nag03, Nov04, Now85, Nym94, Ols17, PT87, PTH88, PT08, PHL16, Pen84, Pen98, PT13, PMW83, PM85, Pol88, PAH88, PG18b, Qiu13, RV90, RHCK24, Rou86, RV93, Rus21, Sac85, SS09a, Sch92b, Sch15a, Sch79]. Optimization [SS23, Sha84, SW11, SWW22, SZ99, SZ03, Sd10, SDSÁR00, ST99, SS03b, Sto76, Stu16, Szn95, Tan88, TDNT18, Tib11, TG23, TFL14, VW08, VC22, WK13, WWY06, WS13a, WZHZ19, War78a, War82, War83, War84, War85, Wri89, Wri93, YM92, YMH94, YK91, Ye95, YHH16, YL12, ZM11, ZA10, ZYHX18, Zha09, ZHB21, dG06, dMCJ22, CL78, Pen88, PG18a]. Optimization-Based [GN03]. Optimizing [AN20, FP21, Fre99, Jaf17, MCL^{+}22, SV17]. Optimum [AGN^{+}16, PZ94, Sag13]. Option [ASR^{+}13, Bay09b, BB06, BET07, CT06, DPZ93, GG09, JR06b, LS09]. Options [Ach05, Bou08, BLZ17, BL02, BL09, CK17b, Ham06, HZ10a, HT10, ZCGT10]. Orbit [HO90a, ZTC16]. orbites [Bac79]. orbits [Bac79]. Order [ADK08, Ahm77, ABCO17, Alv00, AY23a, ANS14, AMR20, ABK21, BS86, BP88a, BLJ15, BS16, BD11b, BV13a, Ber84, BH86, Ber89b, Bet19, BB23, BDP14, BC16, BLSZ19, BFL16, Bre85, BF13, Brü13, BL05, CD91, CFS15, CMS16, CTU00, CM02, CT09, Cas12, CT14, CRT15, CWW17, CWW18, CW20, CHN94, Cha82, Cha87a, Che04, CZ10, CLYL13, CZZ24, CM96a, CKR22, CC90a, CK89, DL06, Den22, Dok17, Dun95, El 17, FG15, Fer05, Fio20, Fra84, FT13, FO18, FNV10, FK21, FS10, GL22a, GN18, GV00, Gfr06, GBR18, GGH16, GT93, Gom20, GH20, GRS00, GR21a, GM11, HD11, Han86, HMV05, HK01, HU82, HU84, HL04, Hoe12, Hoe14, HM02, HN96, Iof79a, Iof79b, IK96a, KP23, Kel24, KMM11, KN14a, KTY18, KQWY20, Kim93a, Kim21].

Order [KL21, KBvdSD17, Kre77b, KR13, LS98, LS00, Lew80, LXZ11, LCZ20, LR21, LLC18, Liu14a, LG20, LM08, LZZ21, LW22, MCL^{+}22, Mai13, MO02, MO04, MDW23, MZ06, MPR22, Nie21, Now17, PKL13, Pen98, PSV18, RGG19, RT03, RZ93, Sak84, Sar97, SNB09, SL18, SFFC08, SW23a, SO06a, ST09a, SV17, TKH10, Tri78, TT20, Vuo21, WL95, War84, War85, War86, War88, WCG14, WM10, XLL08, YL10, Zei01, Zha00, ZZS14, ZZ17, ZW21, ZZ15]. Orderability [AR77]. Orderable [AR77, Rot77b]. Ordering [HBY21]. Orders [WS95]. Ordinary [BR13, BM18b, CM81, DE84, Kus79, LW11a, MA03, MR96, WL95, War85]. Organization [KM18]. Organized [Bar17]. Oriented [BCS24, Del04, HH08, HH10, CJM12]. Orlicz [JNPS18]. Ornstein [AM23, Ber21, FP03]. Orr [GW18]. Orthant [Boo82, IST96, JL18, Sac97]. Orthogonal [SC23]. Orthonormal [Akç01, BVW00, HdHVW03, NHG99]. Oscillating [BDR02, BP13b, CZ04a, GJZ10, NPS15, Zuy16]. Oscillations [Art89b, Gra08, Hab02]. Oscillator [CSB88, Ste17]. Oscillatority [EF09]. Oscillators [CC02, Cza04b, DB12, Gib80, JHB19, KB16, VPJ19]. Oscillatory [GT17, Liu97c, MPS99]. Oseen [ZM11]. Osmolovskii [Iof79c]. Other [BL83, TBT90]. Outer [BLS05b, Com00, Da 14, GP79, VZ97b]. Output [AD08, APA08, AA17, AISW04, ABH98, BP88a, Ben81, BFT96, BDLW24, BGSS21, BA84, CM91, CWKQ24, CW08, CW03, De 83, Dew76, DCI07, DMS02, EG02, FG15, FC19, FK83, FIM09, GMCL10, GNP18, GB86, GGF09, GB93, Gra88, GEF17, GP17, Gu90, GCSB99, GG13, GZ23, Had06, HG92, HS15, HRW06, HS13, HP96, HD87, HD88, Hir02, IP06a, IPM10, Jou03, KCY22, KCBP21, Kob83, Kob84, KSW01, KM12b, LK08, Lei01, LXZ11, LZC21a, LM97b, LZ06, Liu13, LF15, LL21b, MT94, MPI07, MRv94, NVD18, NZ99, NMBM98, NDHL24, OS10b, OPPU17, PD79, Por77, PC00, PS87, Ros92b, SW00, SIC00, TBP18, TZZ88, TP95, Ter94, TG12, Tie20, TF08, VB23, Wan91, WS95, WC96a, WL24, WSR04, Yam88, YGC18, Zam98, ZBB07, ZW18, ZRS97, ZD24, Obe06]. Output-Controllability [Tie20]. Output-Dependent [ZBB07]. Output-Feedback [GMCL10, LXZ11, LZ06, WL24]. Output-Induced [LM97b]. Output-Near-Controllability [Tie20]. Output-Nulling [Ter94]. Output-to-State [ART21]. Outputs [HHT19, JL16, dFS15]. Outputs-Based [HHT19]. Overdamped [CZZ24]. Overlapping [BRR02, HO97]. Overtaking [CHJ87, HL92, HYZ21]. Ovulation [CCS13].

P [BHJ83b, Kup85]. Packet

[FZ09, GTQ21, XZGT23]. Packings

[DHP19]. Pair

[GL92, Kur89, SC80, Yu15b, Yu20b, ZZS16]. Pairs [CC02, Cza04b, DZ24, FV95, McN84, Por77, Wim06, XCZ20]. Panels [KV21]. Paper [Bau09, Kum88]. Papers [Mar77]. Parabolic [Ach05, Ach08, AFT91, AT96, AX96, Aih92, ACS00b, AM89, AQ05, AR99a, ABR00, AAFC19, ABK21, AK19, AHK24, BT11a, BK84, BW89, Bar81, BS16, BBGBO14, BBF18, Bet19, BPV19a, BPV19b, BJ10, BK16, CT96, CMV08, CFM19, CLAL13, Cas97, CCK13, CKK17, CK17a, CMT18, CMR19, CK22, CSES21, CCT07, CW22, CGR10, Cur84, DI93, Dok05, DFCGBZ02, FM94, FCG06, Fri84, Fri87, GLM21, GMM91, GW76, GT93, GHZ14, GLNT23, GSS19, GMS22, Gue07, Gue14, GK11, Hac79, HK21, HN22, HY95, HLZ20, KK17b, KK19b, Kha01, Kha03, KN77, Kno82, Kob84, KR13, KW85, KPV14, KS13, Lam87, LT83a, LT83b, Las84, LM22, LW07, LV16, LX23, Lis14, LZ12, Liu14a, LY19b, LW22, LW15d, Lun91, Mar77, MRR16, Mas08, MV07a, MV08a, MV08b]. Parabolic [MV11, MRV11, MS17a, MZ06, MI15, MKK19, Pan98, Pie79, Pop97, PTZ17, RZ98, RV87, Rod21, RT03, Rou21, SW78, SSW15, SZ82, Sok88, SGM19a, SGM19b, Sto97, TZ09, Trö99, TWW16, Wan00, WW02, WD14, WF24, Was79, Whi79, Whi80a, WT83, Yu95, YZ23, ZZ20, Zho23a, vDHV15, Cou83, Sei82]. Parabolic-like [Rod21]. Paradox [CGM11]. Parallel [AGS13, ABAC10, Fuh76, HL88, LY91a, Man95, MNS91, TBT90]. Parameter [AS88, Aih92, AD98, AC19, BBC81, BGRW88, BRR90, Bar83, BSDR97, BGP20a, BH99, BDRR98, BGH94, CZ95a, CD89, CG87a, CG87b, CP77, DJ95, Di 91, DPV00, DMPD94, EK97, FB98, FZ23, Fuj80a, GPB91, GdOB02, GK85, HP10, IK90, KG80, KMP13, KKA23, KK79, KN03, Kob76, Kob77, Kob78, Kob79, Kob83, Kob84, Kob88, KBvdSD17, KS85b, KK06, Kum82, KW85, KW87a, KS85c, Lam92, LY91c, MS90, PP10, Ped91, PTZ17, QN82, RY95, RV05, Reb88, Reb89, RT98, Ros84b, RW92, Sch83, Sok87, VW95, WHB92, WCJP04, Wir05, YD08, Yu95, ZGS99, Zhu96, JR06a, You87]. Parameter-Dependent [GK85, KK06, Lam92, YD08]. Parameter-Independent [KBvdSD17]. Parameter-Varying [Wir05]. Parameterization [Bay09a, Che23, Pon09]. Parameterizations [BEL02, WHB92]. Parameterized [AT00b, LL19, PK95]. Parameters [Bag81, BL83, BM86, BBD18, Bli04, Bou95, CZ99, Del19, DMZ22, Fla84, FB01, FC05, Gut90, KN77, KS85b, LS19a, ML13, Mai03, Mao90, Mur92, OPPU17, Pap86b, Pet88, VS80, WZ12a, WZ17a, Yu95, Zar92, Cou83]. Parametric [BL00a, DXG15, DBC12, DKRV13, Kan78, KTWY12, KW13, KS13, KY92, LDZ10, MP94, Poo96, SDSÁR00, Ste17, VPJ19, YK91]. Parametrically [BRT15]. Parametrization [Bar85a, GD88, Obe91, PTY07, SO13, Wim94, Wim06]. Parametrizations [HO97]. Parametrized [Art89a, Del23, KS99a, MZ18, NH12, SB92, Sha94]. Parentheses [GX04]. Pareto [BDGK21, DZ19, Eng10, GL06, MYP14, NOV03]. Pareto-Optimal [BDGK21]. Parisian [LY20a]. Part [LMA07a, LMA07b, Nai87, BC81, BS17, BS19a, BK00, Car97a, Car97b, CK23a, CK23b, Che81a, CFNG10a, CFNG10b, Coh97a, Coh97b, Gho88, IM05a, IM05b, IT87, Kan98a, Kan98b, Kiw96a, Kiw96b, Kou24, Kre87, Kru00a, Kru00b, MDW23, MV08a, MV08b, MMR17b, MMR17a, MW21, Mey01, Mey03, MRP99a, MRP99b, NWW19, Qua03a, Qua03b, Sha83a, Sha83b, SD13, SD15, SGM19a, SGM19b, TW03, VGBS13a, VGBS13b, Wac12, Zas95a, Zas95b, ZZ15, ZZ17]. Partial [Ahm77, BHM99, BK82, BTY11, BS15, BKR00, BH00, Bor00a, Bor07, BS13b, CW23a, CA09, DTV18, DR87, DL19, ENJP88, EÖSÖ98, FPS17, Gev81, GLK22, Gho86a, GM17a, GM21, GZ05, HD11, Hau87, HSW12, HSW15, HH08, HH10, HBY21, HO08, HWX09, KCP16, Kam03, Kan82, KK19a, KS97b, LT86, LXY16, LNW23, LTT24, LZ19, LV97, Lü15, MTT08, MSSZ88, NN90, Nag95, Nie14, ØS12a, QW19, RS93, SL13, SC19, Son23, SW21, SX23, TZ07, TP95, WW16, XZ07, YT13, You13, ZVBR14, Zho92, Zho93, aJM18]. Partially [AC13, AC17, Ahm19, AD98, BC96a, BEK89, BL18, Bv85, BÇS07, BÇMS^{+}08, Bis82, BZ99b, BBE99, CE98, CH99, Fle82, FP82, GP98, Hau82, HLR01, JB96, Jam14, KY19, LR06, MT23, MY22, SKS90, SC16a, Tan98, TL09, TBT90, UP02, VZ16, VZ21, XZZ20, ZS23, ZH24]. Particle [ASP22, BLPP24, BKSW24, BM09, BS20b, CDDP11, DDS15, TD21]. Particle-in-Cell [BKSW24]. Particles [CGP22, GH12].

Partitions [BCF12]. Partly [Dre77, ZP03b]. Parts [KC10, vdSR11]. Passage [GH14, GHZ15]. Passenger [LMJ09]. Passification [GCC07]. Passive [DJ13, Flo99, Pau19, SW12, WS13b]. Passivity [DF21, GCC07, Mac14, RGG19]. Passivity-Based [Mac14]. Past [NY16]. Past-Dependent [NY16]. Pasting [TWZ21]. Patch [Sei77]. Patches [RRR16]. Patchy [AB02c]. Path [AJPR14, BK22a, BGP20b, DFŚ24, EMP18, GAP^{+}09, GC05, HLVAC99, HK06, KN14b, NNN21, PB99b, PZ14, Phi94, RTZ20, RTTY23, Sap19, SYZ13, Tse93, WYZ23, WD07, Zha17a]. Path-Complete [AJPR14]. Path-Dependent [BK22a, DFŚ24, EMP18, NNN21, RTZ20, RTTY23, Sap19, SYZ13, WYZ23, Zha17a]. Path-Following [GAP^{+}09, HK06, Tse93]. Pathological [Bra98]. Paths [GW96a, ISdCN99]. Pathwise [BM07, DDT00, DS87, Kus89, PRHL08, Rog07]. Patterns [Chr80, NLW91]. Payment [AM15]. Payments [Pau08]. Payoff [GR05, HLL00, Kum08, MYAS09, MD13]. Payoff-Based [MYAS09]. Payoffs [BL11, BCR04, LK11]. PD [QBP23]. PDE [BSB^{+}97, DKS23, GGM01, GW19, HR21, Kat19, Kre00, MDW23, MPT06, MP11a, Mil17, MU22, Sto04, Stu15, TKXP12, VLP24, XWPG23]. PDE-Based [BSB^{+}97]. PDE-Constrained [MP11a, MU22]. PDEs [AA17, ABV13, AAFC19, ABK19, Cas12, DGH09, EK09, GS22, GY11, GSS19, HN22, HT22, HLZ20, IK96a, KK17b, KK19b, KEK21, KS13, LS19a, McE07, McE09, MK10, MKK19, RTZ20, WF24, WM10, ZZ20]. PDMPs [Ban19]. Peaking [Tre86]. Pedestrian [AD18]. Penalization [Ber92a, Bur91a, Bur91b, CM81, PT13, Rav20, SS99b]. Penalized [HR98, Lia15]. Penalties [BS11, DR79]. Penalty [Aus99, Ber76, Bon89, CDD20, DG85, DG89, Gug10, KB76, KW87a, Laz88, Man85, MOF77, RZ20, Ste11, SD19, Tse93, WK77]. Pencil [Arm86, KS90b]. Pencils [IT17]. Pendulum [Ovs14]. Peng [SW21]. Perfect [CM12, GH12, GNLB14, JN24, MW21]. Perforated [CDZ94]. Performance [AK99b, AZZ20, Bv85, BDLW24, Ber89b, BD01b, CABP17, DDG12, FZ05, Fer97a, FB12, GL97a, GZ12, HW03, HLT20, ILR10, JS12, KP04, LL18a, LŻ11, LGZ13, MT09a, Mun07, PSJ18, WRGA14, ZCC07, ZGS99]. Perimeter [Ams13, BO16]. Period [Rus86b]. Periodic [ACL09, Alt09, AB10b, BM11, BMRR20, Bar97, BGMS78, BCG84, BCG86, CQ05, Col88, CW09, DI87, Da 87, FB08, FKT86, FNP^{+}22, FHY88, GK87a, GL13, Gil77, GLT95, Hab02, IP06b, JLZ98, LW07, LCZ20, Lun91, LDZ10, Mar20, MM80b, MS02, Nai95, NPS15, NvdPS10, Öze85, PYB20, QWY21a, Rus86b, RZ93, Sha85, Sun96, Toa11, TZZ18, Tud90, UMD88, WS90, ZP03b, ZH02, ZH05, Zho07]. Periodical [GJ20]. Periodically [CG00a, CG00b, MK21, THA02]. Permanent [BT21]. Perpetual [BL02, SPS04]. Perron [AD02, BS13a, BZ15b, Sîr14b, Sîr15]. Persistence [BC00]. Persistency [KNW95]. Persistent [LC01, MG13, MCP21, MNC14, RY95, YADF09]. Persistently [CS10b, MCS13]. Person [BP12, CLS12, Kin84, LN82, PZ14, SS06, Sun21, Top79, Uch78, Zha05, BP14a]. Person-by-Person [BP12]. Perspective [ABVC19, Bal82, GL95, OVS08, RJME09, ZH18]. Pertaining [BA15]. Perturbation [AMP00, BB98, BF76a, CFNG10a, CFNG10b, CR93, DV83, FP02, FB08, FHH22, FS23, FR14, FPS17, Gli12, GYZ22, KK79, LWLZ20, MM79, MS96, NS00, O’B79, Reb78, Red18b, Sto97, TLC99, Tol18, Zha95, ZGS99]. Perturbation-Analysis-Based [TLC99]. Perturbations [AB02b, AG97b, BK24a, BG07, BK10, DD03, EAK86, FK76, Gra99, Gra81, Gra82, KK12, Kum13, NS91, NS05, Ped91, RSS18, Sei87, ST16, Son93, TWW16, Tud90, VT95, Yu14, Zho23a, Zol78]. Perturbative [Bur20]. Perturbed [AL02, Bin76, BC96b, BC96c, BC96d, CGW01, CUB20, CK78, Cob10, CMN20, CG92, CV86, CD12, DDG12, Don10, DMS02, DMS12, DF98d, FF22, Gai92, GN02, Gai04, GFF17, GO11, Gra05, HM08, HL81, HP14, JV17, LWfZ22, MSR01, MNS91, NTP18, PB96, Pen84, Shi02, Tsa78, Vig97, XWZ^{+}21, Pen88]. Pest [ALNP24, EGR^{+}22]. Petrov [MV11]. Petviashvili [RS20]. Phase [AMP10, AS21, AFCLS22, BCG^{+}17a, BSS93, BGHR16, BS06, BL03, CDCVV21, CWKQ24, FHY88, GLS21, GMN22, GBo97, HG90, Ich16, KBDK03, NWW17, NWW19, Sha85, Sha86, Vrd16, WW03, ZG17, ZLHZ12]. Phase-Field [AS21, BCG^{+}17a, BS06, GMN22, KBDK03, NWW17, NWW19, WW03]. Phasor [AJ12]. Phenomena [FM22, MN07]. Phenomenon [AA97, BC13, CM96a, Fer05, GH19, HH05, HM88, Zas06, Zol92]. Photon [DZA19]. Physical [DvdS98, EMP^{+}16, Ovs14]. Physically [SW12]. Phytoplankton [GT07]. PI [Kob88, LZ92]. PI-Control [LZ92]. PI-Controller [Kob88]. Pick [TR01, BL00b, GK87b, Rot96]. PID [Fio21, ANM22, Fio21, MCL^{+}22, ZY24]. Piecewise [Alm01, BL18, BW88, BP07b, BG14, Can84, CD03, CD08, CD10, CDP16, DC99, ER88, FP89, Gor15, GH07, HD11, HL92, HM08, HY97a, JZ09b, Li95, Nik99, Pro90, TC14, Tod83, dBMS10]. Piecewise-Affine [BP07b]. Piecewise-Constant [Can84, Nik99]. Piecewise-Deterministic [CD03, DC99]. Piecewise-Linear [Tod83]. Pieri [HV00]. Piezo [VS14]. Piezoelectric [AGN^{+}16, CNS10, KMM07, MÖ14, Tuc96]. Pinning [YCY13, YCLK13, ZL24]. Pinning-Impulsive [YCY13]. Pipeline [GG24]. Pipes [CGHS09]. Pipkin [CFZ22].

Piston [PS19]. Pitaevskii [HMMS13].

Placement

[BCM^{+}17, CLZH24, CWC^{+}22, EG02,

Gho86a, HRMK17, HV00, KRW04, WC81]. Planar [AB10b, Bos02, BP98a, CS89, HT92, Joh05, LLS93, MRL12, NMBM98, OOS15, Pic96, Tao18]. Planck [AØ23, Bar21, Bar23a, Bar23b, Ber21, BK23, LB00]. Plane [CM06, HPS03, JR06a, Jer80, LR79b, Sus87b, Sus87c, Sus87d, YADF09]. Planes [Jer78]. Planning [ACCD12, BSVD84, FSS87, FZ98, GJZ10, JP15, LLT18, RTTY23, Tam80, Zha95]. Plant [DC10, FLJ13, TZQ21]. Plants [BGMR94, GS19, NK94, Sch92a, YC24]. Plasma [BKSW24]. plastic [MAJ18]. Plasticity [HMW12, MW21, Wac12]. Plate [ASTT09, GZ14b, MP07, OZ00, OZ03, Pan17]. Plates [AL00, BF07, ET15, MY93, Rao98, ST99]. Play [ED81, GQ11, SPO22, SS09b, SKXL17]. Player [BG98, BC18a, BCR24, BLP14, CS12a, Ete24, GBB18, HG90, Hua10, Jac22, Var76, Vie00]. Players [AG95, AE95, Car96, NH12]. Plus [AGL08, DGKL10, FM00, McC09, McE04, Qu14, Cha10, DM15]. POD [AFV17]. Poincaré [CO22]. Point [AH21a, ABHY23, AB93, Alm08, Beh79, BB19, Ber85, BC24, Bré79, BS00, BQ99, CDDP11, CZ98, CBB17a, CGC17, CF13, DHX08, Del07, DHV98, DJZ96, DM17, DMPD94, DMPD98, Fat90, FK76, Gül91, Ha90, HZ07, HZ95, HQR24, ISdCN99, Jar93, KT00, Kha01, KQWY20, KRW04, Kin84, KKLO21, KS86, Kre00, KW87a, LS91a, LS99a, LY12, Liu88, LAN04, Luq84, Mai08, Mal23, MD15, Mey77d, MS97b, Roc76, SVBZ09, SS03b, Stu15, Sun96, SZZ97, Uch78, Uch79, UU00, VA03, Wan00, Wei05, Whi82, XZZ20, YZ97, YDZ98, ZCGT10, Ber88, UUH99, ZL01]. Point-to-Set [Mey77d]. Pointing [CKR22]. Points [AS20, AP92, BB18, Bia83, BDGK21, Bor77, CLSL20, CGC17, DG90, DS09, Fra87b, Grü98, HD87, Ho93, HPT07, KZ00a, Kaw01, LZT20, MP76a, Mik15, Per81, Red79, Rot77b, Ste13, SY14, Top79, TFL14, WCLC05, Yan96, ZZ88, Zei92]. Pointwise [AFO22, AT00a, ByS17, Cas86, Cas93, Con90, FT13, GS18a, GHZ14, HH10, HY95, KS95, Kha94, KN14a, KQWY20, KR13, LV16, LW15d, MRV11, MR04b, PY16, RV05, RT03, RT06, RT07, UUH99, UU00, Wol12, ZZ15, ZZ17]. Poisson [AM15, BKSW24, BP07a, BB09, BLS05a, BPR04, CD03, KR04, LMMR15, LS19b]. Pol [Fen16, LHX12]. Polarization [CNS10]. Pole [Ari98, BCM^{+}17, EG02, FR78a, Gho86a, HRW06, HV00, KRW04, KP04, LH87, LDZ10, RRW96, Ros92b, Wan94, WCCL04, WC81, WORO00, FR82a]. Pole-Clustering [KP04]. Pole-Zero [LH87]. Poles [BFP11, CPW88]. Policies [Ahm77, BP10, BLS12, Can84, CS12b, DS01, Ete24, FMS20, FP19, FP11a, GHS12, GYBL15, Ken80, KA06, Kru00a, Kru00b, KR89, Mey09, PYB20, Rot77a, SY21, TS24, Yu20a, Yük19, ZKZB20]. Policy [BB09, BLS05a, BK16, CL19a, FPD08, GRZ24, HSA17, HXY21, KSS20, Kur89, McC09, NF13, RSZ23, SR04, SZ10, TS24, WP18, Yük24, ZKZB20, ZHB21]. Polling [AK02a]. Pollution [AVCCD22]. Polyconvex [CM05, LSW14]. Polygonal [AMPR15, KLL10, Wol12]. Polygons [GCB06, LN10]. Polyhedra [ID87]. Polyhedral [AAA12, GMR22]. Polyhedrality [MW78]. Polyhedric [BZ99b]. Polyhedron [Dun95, Dun98]. Polynomial [AN15b, AE17, BKP18, CGR96, CP82, DM89, ER88, Emr80a, EH80, Emr80b, FK21, HHLP16, KHJ14, Lom13, MGHT19, MBC06, NMBM98, RS10, Son79, XWZ^{+}21, Zer05]. Polynomial-Exponential [Zer05]. Polynomials [BO01, Lou15]. Polytopes [vTY99]. Polytopic [NNT22, YD08].

Pontryagin [AK04, BBB24, BZ99a, BHT19, BB23, Bon86, BC95b, BDP14, BT13, Cas97, CRZ00, GJ11, HY95, KRT11, LY20b, Liu18, MS96, RZ98, Sag13]. Poor [Rud78]. Population [EK93, Fra03, Hua10, KLM20a, LXY16, LNW23, SK23, TA21, WZ12a]. Porous [Sta15]. Port [Aug19, CvdS23, CDM23, FMP^{+}22, JKZ21, MM04, Mac14, SZ10, VS14, vdSM13]. Port-Hamiltonian [vdSM13, Aug19, CvdS23, CDM23, FMP^{+}22, JKZ21, Mac14, VS14]. Portfolio [BC11, BPS05, BGP20b, BC18b, BLY19, BLY21, BR07, DGR18, DH12, FPW16, FH17a, GZLC17, Heu15, HZ05b, KLS87, KK02, KS07, LZL02, Nag03, OS02, WJW19, WYZ17, Xia11, XZ07, XZZ20, ZSM10, ZY03, dMCJ22]. Portfolios [FPS17, ML12]. Portrait [Sha85, Sha86]. Ports [DZ76]. Posed [Cur03, CO05, GS18a, GK86, Kob77, LT03, LRT05, Log13, Lü15, OS19, RZ06, Sta98a, Sta98b, TW03, Vog90, ZZ23, Zol81, CT06, CW06, WW02]. Posedness [ABL13, Aug19, BH07, BEL02, CHY23, CM03a, CSZ13, EB20, GT14, GZ09, IZ00, Lü15, PCP22, Smi90, TBP18, VS80, WZ09, WCG14, Zas00, LT18, Vil04]. Positioning [HS96]. Positive [All88, AW90, AI02, BA08, BGMR94, Boo82, CP13b, DG90, EGR^{+}22, FM12, FCJ14, FV11, GR21a, GBS22, Hew93, LY23, NTP18, NNT22, NS05, NNSM08, OM14, OMK84, OT94, PPL09, PYB20, Sac97, ZCGT10, van98]. Positively [RG14]. Positivity [HMT18, Mic80, MS18, Ngo09, RV09, Sch89]. Possessing [MC91]. Possible [AJ77, CLG12, LL19]. Possibly [ADS24, CHPA20]. Post [PT08]. Post-Treatment [PT08]. Posterior [KSY23]. Posteriori [AFOQ21, KRS14, KSY23]. Potential [BM09, Caz14, FY76, GP22, SMK18, TS24, Yu14]. Potentials [CFRV98, CFSGS15, CS15, DE14, DWZ20]. pour [HU78a]. Power [Ahm77, AEÉ^{+}20, BM84, DV04, DB12, FGT95, JFB17, KvdSD21, LXY14, MPN98, POU11, SI03, ZH18]. Power-Fractional [POU11]. Powers [AO15]. Practical [BM09, CH21b, GH20, LZ06, MS04b, YL10]. Prato [JLZ98]. Precision [KK14, Sch79, WCLC05]. Precommitment [Zho23b]. Precommitted [HHOZ19]. Preconditioners [NM10]. Predictable [AZZ20]. Predicted [Ben76]. Prediction [GH14, KH69, LY91b, Sei77, Wei77]. Prediction-Correction [KH69, Wei77]. Predictive [AHKS22, ACV22, BDIW22, BDLW24, CK02, CK03, CGW20, DXG15, GSS19, HTLV17, LLSX20, PGB14, TGSM18, WRGA14, YYH^{+}16, dMCJ22]. Predictor [SZZ97, XZGT23]. Predictor-Corrector [SZZ97]. Predictor-Like [XZGT23]. Predictors [Bat19, Kar11, ZK20]. Preferences [AH21b]. Preferred [BAP18]. Premium [JXZ24]. Preprocessing [New00]. Prescribed [BFZ00, BDLW24, CUB20, IRT07, ILR10, LL18a, LK23, NDHL24]. Prescribed-Performance [LL18a]. Prescribed-Time [LK23, NDHL24]. Prescribing [GH12]. Presence [AK02a, Amm02, DC10, GAP^{+}09, Gra08, GW18, LFH15, LM02, LZ13]. Preservation [EMP^{+}16, Pep16, TFL14]. Preserving [BHT10, Bre16, PH92, RV09]. Pressure [CMS16, CCS18, KC10, MW23]. Pressure-Robustness [MW23]. Pressures [GLK22]. Preview [MT09b, WFZ20]. Price [CMS16, CFH09, GH17, KV21]. Prices [AC79, CSZ13, CT06, KR04, KMZ98, SZ04]. Pricing [AEÉ^{+}20, BET07, Bou08, ÇD21, CS12b, DPZ93, EQ95, FT22, GG09, HL16, JR06b, SS06]. Primal [BLW07, BS24, BIK99, GV05, HHF21, IK04, IK07a, JK20, KB76, LCF18, LR79b, TDBL22, Wri95, dlRHM16]. Primal-Dual [BLW07, BS24, BIK99, GV05, HHF21, IK04, IK07a, LCF18, TDBL22, Wri95]. Prime [ABH98, MRv94]. Princess [Fit79, AFLO08]. Principal [CP82, II97, ISW00, LL20, LRTY22, Ohs13]. Principal-Agent [LRTY22]. Principle [ABVC19, AC13, AC17, AMR20, AA97, AKP05, AK15, AK04, BDM07, BEK89, BM90, BBB24, BY13, BZ10, BCS15, BZ99a, BV17b, BB23, Bon86, BC95b, BT11b, CK95, Cas97, CRZ00, CP98, CF05, CS05, Cla76b, CV87, Cob12, Col82a, CI77, Dav76b, DS89, DEO24, DO22, DM17, DM13b, DM06, DQS16, Dun91, Dun93, FM12, Fat94, Fra87b, FO16, GP05, Gom20, GJ11, GHS22, Gra99, GT09, GW09, GTT16a, GJN08, HPW10, Hau78b, Hau87, HY95, HJ16, HJX18, HW20, HJX22, HCM05, HWX09, KG19, KRT11, Kre77b, Lau99, LS00, LY20b, Liu11, Liu18, LT23, LT13, Lou03, Ma92, MD15, MS97a, MS04a, MS96, Muñ00, OL77, PP14a, Pen90, Pop97, Qiu13, RZ98, Ris94, Rya98, Sag17, Sch80a, STW20, Sor98, SS03a, SW21, Ste08, Tad93, TWZ21, Tan98]. Principle [VP88, War76a, WY08, Yon10, ZES12, ZSX18, ZC24, ZS23, ZH24, ZD24, Zor17, CL06a, BT13, EB24, Hua10, Sag13]. Principles [AISW04, AMI20, IZ00, Iof24, JXZ24, Lev01, MAG11, MW02, ØS10a, PZLM24a, PZLM24b, WWX13]. Priori [BPV19b, DGO19, Don96, GHZ14, HR21, KMR10, LV16, LW15d, McE95, MV08a, MV08b, MV11, MRV11, PV13, RV05, VL78b]. Prioritized [SK95]. Priority [HHS18, SPBF10]. Priors [KY19, MY22]. Prisoner [GBB18]. Pritchard [van95]. Probabilistic [CD13b, CIH18a, Cha14, CZA08, DXG15, FSGHH20, GR17b, HY22, HMR15, Jam91, LZC22, LLM^{+}24, Pla80]. Probabilistic-Variational [Jam91]. Probabilities [BA15, DK89a, FKZ22, HL81, JN06, Küe07, She85]. Probability [BG98, BD02, BZ15a, BZ15b, BD21, Bis77b, BPTT21, HO01, Kam10, LY20a, Sny77, TLD24]. Problem [Ach05, Ach08, ACCD12, AB18, Agr95, AAC86, AT17, AFOQ21, AFO22, Alp95, AKN00, AM17, AM95, Amm02, AJK14, ANS14, AO15, AOS16, AB05, ABK14, AKP05, AMI20, BB98, BD15, Bal01, BMS05, BKP90, BK84, BD83, BT91, Bar91, Bar94, Bar95b, BP97, BP20a, Bar89, Bar85c, BE13, BX12, BY14, BZ15b, BS18, BB92, BMS15, BB07, BB97, Bel00, BGR07, BE95, BCR24, BF76a, BS87, Ber98c, Bie95, BO16, BRW05, Bon85, Bon91, BCM13, BK16, BCS13, BHT21, BH22, BK83b, Buc92, BP98b, BN16, BJ17, Bud03, BR07, BGL11, BS00, BLS05b, BL03, BS03, BJO09, BGL98, CD13a, CGM98, CC04, CFS15, Car97a, Car97b, CGS00, CDCCMG08, Cas86, CMR07, CC16, CK23b, Cau13, CSCV18, CDFP19, CZ94, CZ95b, CP98]. Problem [Cel02a, Cer01b, CPY09, Che99, CCT07, CY12, CGY21, CHJ15, CH98, CH09, CF14, CY94, CMK15, CSZ13, Chr77, Cla76a, CG17a, CL06b, CT06, CP95, Cra99, CO05, Da 87, DD92, DA02, DF96, DF98a, DFM15, DEO24, Del86, DMS03, Die87, DC88, DHP09, DO22, Don24, DV83, DF94, DF98c, DFLZ23, DMS12, Dup03, EMO23, Fab08, FG09, FP02, Far13, FP08, FFS20, FL81, FKT86, FS21, Fer18, FS19, FH94, FT77, Fla84, FSS87, FTZ87, Fra77, FT13, FL80, FL82, FHY88, FN84, Fuj86, GKB00, GGP14, GPS95, GR88a, GW17, Gli12, GLD^{+}91, GY11, GP94, Goz91, GHQ15, Gri83, Gro16, GMS22, GAHL09, GI02, GGMZRB20, GGC23, GK98, GM99, GM00, GW19, GY13, GTT16a, GZ19, GX19, HRCK21, HLL^{+}17, Hal89, HLG93, HZ10a]. Problem [HYZ24, HT92, HS12a, HS12b, HNS18, HAB78, Hel76, Her91, HK94, HP09a, HT10, HR21, HMRR10, HQZ16, HY97a, HR98, HLP08, HS09, HZ10b, HJ16, HW20, HK19, Hui92, Idc98, IY12, Iof76, IT87, JL93, JS81, Kan78, Kaś77, KM12a, Kaw01, KS97a, KTWY12, KLV09, KV21, Koe90, KKS93, Kot94, Kou24, KL16, KO18, KS86, Kru00a, Kun82, KW85, KW87b, KW13, Kus78a, Kwo80, Kwo91, LT83a, Las84, LW10, LS07, Lee86, LZ17, LNW23, LMA07a, LMA07b, LP79b, LR96, LMPR09, Lou02, Lou07, LSW14, LT91, LC23, Mam14, MPZ03, MRVM00, MBS13, MZ20, MMR17b, MMR17a, Men80a, Men80b, MR90, MS11, MT08, Mik15, Mor10, MS07b, MP07, NWW17, NWW19, NYZZ17, NLZK19, Nic00, NS86, NPY20, O’M76, Ols76b, Ols76c, Osh06, Pan98, Pav84a, Pee20, Pie79]. Problem [PZ20, PS11, PS13, Pog17, PS87, Pue09, PPZ06, PM16, QBP23, RS07a, RU19, RU21, RV19, RS21, RN13, RTT20, RH06, RW78, RT03, RT06, RW92, Rus86a, Rzy09, SM13, SKD99, SBBV15, Sch80a, Ser03, SS20, Sha85, SB94, SSK02, SC06, SS89, Sta93, Sta15, SZ96, Ste13, Ste15, Ste92, Sto91b, Sto92, Tad93, TST07, TTZ23a, Tam80, THA02, Tan00, Tar85, Tes92, Ton03, Tre89, Tsi93, Tud90, TO15, Uch83, VP07, VG95, VT95, VK81, VSB18, Wac16, Wan08, WZ12b, Wat79, Wec00, Wee92, Wee05, Wee07, WHC03, Whi79, WY08, Yan14, YY05, YY08, ZZ92, Zer03, ZP03a, ZP05, Zhu92, Zhu96, ZRS97, Zhu00, ZZ23, Zol78, dVC22, van86, van95, GR85a, Pen89, Qua80, You87, ZZ94]. problème [Pen89, Qua80]. Problems [AT89, ABVC19, AFT91, AT96, AKO17, AGL08, AD98, ADRU22, AFV17, AM89, AS91, Alv01, AMR14, ASTT09, Ama04, AQ05, AP92, Ang76, AMPR15, AR99b, AR00, AFCLS22, AST05, AMR20, ABK21, AA97, AK15, AK04, AC79, Aub02, ABSP08, BM19, Bah08, Bal77, Bal82, BV18, BB78, BI91, BLPdD15, Bar78, Bar80b, Bar81, Bar82, Bar88, BF91, BB18, Bar85b, BP88b, BBC14, Bar93, BJ96, BM90, BBB24, BBCS22, Beh79, BCS17, BKS10, BTBT88, BG09, BHOP14, Ber92a, BT96, BK97b, Ber98b, BIK99, BL04, Ber82b, BE87, BHT87, BF22, BV10, BV17a, BK24b, BP17, Bin76, Bla97, BT97, BdPV16, BV17b, BCL24, BFZ10, BPZ16, BT03a, BB23, BCM12b, BPV19a, BPV19b, BZ99b, BH07, BS11, BDP14, BG20, BG21, BM13, BGS19]. Problems [BL91, BK13, BC16, BET10, BLSZ19, BF07, BT13, BT21, BKP18, BP20c, ByS17, BT03b, BT10, BR01, Bro87, BF89b, BC23b, BDDSV05, BM07, BV14, BKM14, CFG21, CT96, CLAL13, CHJ87, Car90, Car76, CDCCMG04, CD15, CF95, Cas97, CRZ00, CTU00, CM02, CT09, CS10a, CGM11, CCK12, Cas12, CCK13, CT14, CKL16, CWW17, CWW18, CMR19, CW20, CT20, CK22, CK23a, CK23b, CCJ23, CW23a, CLHH05, Cat98, CO03b, CM10b, Cel15, CF05, Cér00, Cer01a, CQ05, Cha82, CE98, CI19, CHL82, CDD20, CM81, Che04, CCT05, CM96a, CS02, CNP98, CL18, CNS21, Cir14, CW91, CdP10, CB11, Coh21, CS89, CW24b, CV86, CM96b, CD80, CM05, DDK12, DG14, DPZ91, DK05, DGSW14, Dat80, Dav07, De 15, Den22, DZ24, DHV98, DBC12, DZ19, DR87]. Problems [DG85, DJZ96, DGQ13, DO14, Dok05, Dok17, DK78, DPV00, DHV00, DL20, DM17, DNS17, DKS^{+}11, Du15, DM02b, DM04, DM06, Dun80b, Dun86, Dun96, Dun98, Dun00, DS01, Ega08, EÖ88, EMP18, EL20, EH10, EGN85, FK99, Fag08, Far08, Far13, Fat94, FM94, Fat97, FP08, FP14, FT14, FFRR21, FGL08, Fis11, Fla87, FBP99, FBJM14, FR14, FB01, Fra84, FO18, FZZ18, FK76, FP11b, FJ83, Fri84, FH88, FHT06, Fuj86, Fuj80b, FGH98, Gah94, GR06, GQ09, Gal84, GV03, GW96a, Gen13, Gfr06, Gho86a, Gib79, GRZ24, GW76, GHZ98, GG22, Goe05, GT93, GWY14, GHZ14, Gow89, GL99, GG18, GT22b, Gug10, GH19, Gui13, GH96, GK11, GT09, GY16b, GTX22, HK13, HT11, HK78, Hac79, HH76, HM76, Hag05, Ham21]. Problems [HSW19, HZ22, HK21, He87, HYZ11, HHF21, HJ21, HTY01, Her76, HW19, HHSS23, HLL96, Hoe12, HS87, HCM05, HWX09, HZ19, HYZ21, HW21, HQR24, HV00, Iof24, IK96a, IK95, IK96b, IK02a, IK04, IK09, IK11a, IK11b, IS02, JMP21, JX22, JQ97, KR80, KSW18, KS18, KK17a, Kar89, KZ00a, KS87b, KS92, KS95, KWY08, KN14a, KTY18, KQWY20, KFP22, KRW04, KG17, Kno79, Kno82, KN14b, Kog87, KRS14, KK02, KRT11, Kre78, KK06, KPZ18, KLM^{+}21, Kru00a, Kru00b, KMR10, KR13, Kum13, KW87a, KPV14, KS13, KM04, KM07b, Kus89, Kus90, KM91, Kus13, LL00a, LLS90, LS91b, Lau99, LI00, LS05b, LS99a, LZ96, Leu99, Lew80, LV16, LLMT02, LW14, LL18b, LH87, LH88, LRTY22, LRT96, Liu88, LNT02, LE17, LR97, Lóp12, Lóp13, Lor10, LZZ21]. Problems [LY01, LY02, LW15d, LB89, Ma92, MXZ19, Mai08, Mal23, Mal02, MS87, Mar02, MBL00, MG20, MG21, MM21, Mas08, MR78, Mat05, Mau77, MP94, MO02, MRV13, MADF98, McE04, MD15, MDW23, MML21, MV07a, MV08a, MV08b, MV11, MRV11, MR16, MR17, MR18a, MS98, MP76b, Mey77b, MR04a, MR05, MP11a, MMN09, Mik12b, Mil96, Mir86, Mor11, Mot95, MS10, MNS91, MS96, Mur92, Mur86, MN08, MS85, Nag03, NR97, NT95, NPR15, NR21, NP22, Nic96, Nic97, NSW17, NM10, Nov04, Now85, Nym94, PZ94, PT87, PT86, Ped91, PT09a, Pen90, Pen20, PP14b, Pen84, Pen98, PL15, PTZ21, PT13, Phi94, PV13, PSV18, PMW83, PMH92, PYM93, Pol88, PW18, PAH88, Poo96, Pop95, PA93, Pro90, PV98, QW18, Qua03a, Qua03b, QW19]. Problems [Qui92, QR11, QV13, RR94, RV00, RV05, RZ98, RSE20, RBA22, RZ20, RSZ23, Rob81, RW90, RT07, Ros84a, Rou86, Rou21, RV93, RS93, Rus76, Sac85, Sag15, Sag17, Sch09, Sch79, SSW15, SP96, Sei88, SV97, Sim76, Sîr14a, SS24, SI03, SRS15, Sle78, Smi90, Sok87, Sok88, SS99c, ST02, ST09a, SSZ11, STW20, Son79, Sor93, SV17, SV79, SZ97, Ste07, SS03b, Ste11, Sto97, SZZ97, SLY16, SY24b, Świ20, TTZ23b, Tal91, Tan03, TDBL22, TD94, TY23, Tib85, Tib11, TKXP12, Tim14, Tol16, TZZ18, TY00, Tsa78, Tse90, TBT90, TWW16, UU00, VS80, VC03, VW08, VS20, Vig97, VL78a, VW90, VZ97a, Vin19, Vog90, Vrd16, Wac07, WW16, Wac19, WL95, WX13, WZ17c, WZ21, Wan22, War76b, War78a, War82, War91, WZ94]. Problems [WYY17, WY21, WD07, Whi80a, WK77, Wri95, WT83, WCLC05, WZC20, YF97, YMH94, Ye95, Ye97, Ye00a, YD03, Yon93, Yon13, YZ97, You13, YZ23, ZG17, Zas97, Zas00, Zas06, ZZ88, Zei94, Zei01, ZYHX18, ZL02, Zha15, ZI00, ZL01, Zho23a, dFN23, dRdPR08, dWvS00, vTY99, vDHV15, GR85a, GR85b, KS84a, KS85a, Pen88]. Procedure [Bol80, EMH08, Fis11, MP11b, vW84, GR85a, GR85b]. Procedures [RN13, War82]. Process [AB93, AG97a, BDR94, Ber21, BV13a, BR15a, BHT22, Bré79, BASG21, CDDP11, CP16, CMN20, DR06, DG16, DSI22, GP98, GH07, Hau81b, Hau81a, HP22, HLG21, IS14, JV17, KR04, Koe79, LS91a, LFL23, Ric77, Rod82, Sza95, VZ21, WS90, Wee18]. Process-Controllability [LFL23]. Processes [ABB01, AC13, AC17, Alm01, Alm08, AZZ20, ABFG^{+}93, AFO86, BP10, Ban78a, Ban78b, BCW21, BB89, BL18, BA15, BS24, BR10, BD17a, BV77, BL02, BMN19, Bud03, BM18b, Cam76, CH96, CZ10, CW23b, CW24a, CC88, CC94, CTT16, CP13a, CF13, CT06, CD03, CD08, CD10, CD12, CDP16, Cur76, DM83, Dav76a, DH09, DS99a, DS07, DC99, DPR13, DG20, Dun81, DRDG08, EC22, EP08, EK93, FKZ22, FS21, FP03, Goe13, Gor15, GMR22, GW09, GHS12, GHZ15, GL19, HSA17, Hag79, HJ13, HJ15, Hau78a, Hau81b, Hau81a, HSZ24, HHPY16, HLL94, HLL96, HLGH98, HLVAC99, HLGHLM03, HSV00, HP14, HLT20, Ich11, JPSB24, JV06, JN19b, KMF19, KA06, KLV09, KB99, Kur89, KY93, Lag83b, LHS06, LP85, LLC11, Lou94, MM00]. Processes [Maz94, Mik12a, MM16, NF13, ØS12a, OC23, OPPU17, Osh06, Pau08, Pav84a, PYB20, PZ11, Piu15, Piu18, PZ22, PZ24, RW87b, Rin09, Roy89, Rus88, SRM16, SX14a, Sha15, Sha17, SX19, ST23, SSO13, She85, SVBZ09, SK23, SCCL18, Ste22, TYW11, VZ16, VP88, WD79, War78b, Wee05, XZ17, XZZ20, Yu15a, Yu20a, Yu20b, Yük17, Yus97, Zha09, Zha17b, Zhu92]. Processing [GP17]. Product [Bar88, Pop95, Wan94]. Production [BSVD84, FSS87, FZ98, MZ09, New06, New07, SZ95, Tam80, YXJ11, Zha95, CY18]. Production-Sales [Tam80]. Products [CLYL13]. Profile [LY23, WW23]. Profits [GW20]. Program [ALY13, BP17, HT10, Mey77c, MW78, Out00, Yük20, HU78a]. programme [HU78a]. Programming [AFV17, Alm01, ABB04, AA92, Aus84, Bar91, BY13, BCP16, BR82, Ber77, Ber98c, Ber18, BJ19, BCG^{+}17b, Bor89, Bor00a, Bor07, BT11b, BN13, BF93, CSBC14, CF05, Cha77, Cha82, CS02, CV87, CV86, CK17b, DHV98, DG79, DG85, DPR13, DG20, DQS16, ES81a, EK85, EQ95, EK09, FNA02, FT14, FL81, FGL08, Flå92, GR06, GQ09, GT77, GHZ98, Gom20, GT90, GAHL09, GHS12, HS95b, Hen83, HLL94, HLL96, HLGH98, Ich79, IK11a, KK17a, Kaw01, KA06, KB76, LPJ20, LC80, LP79a, LP81, LLS21, MD89, MS98, MR96, Ols76b, Ols76a, Ols76c, PV04, Per81, PW17, Phi94, PZ22, PH92, Pop97, RC16, RW87c, RW76, Roc87, SP02, ST02, ST09a, Sta93, TST07, Tak97, TTZ23a, Tan15, TLD24, Tse91, Tse93, TCC15b, TCC19]. Programming [VP87, VL78b, WZY09, WY08, YZZ01, Zit14, dFN23, Flå95]. Programs [ANP83, AP94, BW88, CC13, CD80, FHM84, Gri86, HL88, HW99, Ish88, JG99, Jer80, KM07a, Laz88, Lep90, Li94a, Li95, LP87, LR79b, Liu95a, LB98, MM79, MS87, Mey77c, MOF77, O’N76, Poo96, PR07, Pul93, Pul95, Pul96, SS99b, Set93, Sha88a, SB92, Sha94, SD19, Zas95a, Zas95b]. Progressive [STW20, ZS23]. Projected [Ber82b, BR17b, Dun81, HBD21, KS95, Ter94, TGSM18, Wir98]. Projection [BH86, BR15a, BK08, CM18, Dun91, Dun93, Dun96, Dun00, GB84a, KS92, Kiw96a, Kiw96b, Ouv78b, Phe85, Phe86, Rod21, ST96, SS99c, TD94, WK77]. Projection-Based [CM18]. Projection-Type [ST96]. Projections [CS18, ID87, Phe85, Sha88a]. Projective [Bau09, ES09]. Prolongation [NR17]. Prolongations [MTS20]. Promoting [MW22]. Proof [Bay09a, Bay09b, BO14, GŠZ05, GŠZ10, Ner21, Tar90]. Propagation [BSS93, DES09, DKT17, NWW17, NWW19, WGK11]. Proper [Bor77, DG90, KK23, NZ99, PK87, Zhu94]. Properly [Jah85]. Properness [Col88, Ros00b]. Properties [Ahm83, AGL08, AT92, AT94, ANP83, AB10a, AK15, BSI19, BD19a, BC00, BKK02, BD77, BHL12, BC02b, CD15, CG90, CDL85, CV90, CMW06, CS89, DHP19, DMS02, DK85, FL20, FP11b, GBo97, GH15, GG18, GLO19, GP82, HH76, Ham82, HBY23, IK02a, IK09, JALG18, JS84, KS89a, KLM20c, KH81b, KY87b, LCF18, Lor87, MR15, Mat76, MR04a, MRP99b, Pan92, Pap89, PAS00, QWY21b, Red21, Rod82, SBY23, SR04, Sha92, SL18, Sha15, Sha88a, SY24b, SKXL17, Uch83, Whi80a, XZ17, XA93, Yük16, ZI00, Zho84, ZH02, ZY07]. Property [Alv00, AJ77, BBK00, BP20c, CTT16, DS07, Guo02, LT13, SIC00, TZZ18, WXZ15, WY16, XG03, Zas97]. Proportional [ARS17, ABCD22, BBC^{+}19, BC12, CW08, EI13, LPT22, OS02, Pau08, Ric77, ZPQ20]. Proportional-Integral [ZPQ20]. Protection [HS96]. Protocols [LFH15, MCL^{+}22, Med12]. Provably [Bas91]. Providing [RS10]. Proximal [BAKS18, BS00, BQ99, BK08, CL87b, CP04, Gül91, Ha90, HZ07, ISdCN99, KT00, Kiw97, LY12, Luq84, Roc76, WZ98]. Proximal-Projection [BK08]. Proximity [MS22]. Pseudo [CTT16, ÖB19, Vuo21, Whi79, Whi80a, Yam88]. Pseudo-Backstepping [ÖB19]. Pseudo-Markov [CTT16]. Pseudo-monotone [Vuo21]. Pseudo-Parabolic [Whi79, Whi80a]. Pseudo-Rational [Yam88]. Pseudocontractive [SWY12]. Pseudodifferential [Ped91]. Public [Ave19, Fer18]. Public-Signal [Ave19]. Puiseux [SCFV97]. Pulse [AJ12, FM16, KB16, VP22b, XCZ20]. Pulse-Coupled [KB16]. Pulse-Modulated [AJ12]. Pulse-Width [FM16]. Purcell [BBR18]. Pure [BdPV16, DKR08, Hoe12, LL18a, MG21]. Pure-Feedback [LL18a]. Purification [Sag17]. Pursuit [Ber86, CQSP00, CLS97, KCP16, Sor93, Sor98, TS18, Wen15, Yon88b, Zar84, ZSS^{+}20]. Pursuit-Evasion [Sor98, TS18, Wen15, Zar84]. Pushed [BCL24]. Put [ASR^{+}13, Bay09b]. Pyragas [SS20].

q [DHK20]. q-Closest [DHK20]. Quadratic [AT96, ALY13, Alt09, AF21, BRR02, BV18, BI91, BP14a, BL89, BK22a, BCLY17, Bie95, Bis76, Bis77a, BP20c, BR93, BK83b, BP98b, BKM14, CL19a, CBM23, Cam76, CK98, CDCCMG08, CM81, CG87b, CLZ98, CZ00, CZ04b, CR17b, DM02a, DL06, DI90, DD92, DK05, DDH15, Del86, Del07, DS09, DH12, DL20, DMS12, Du15, DMZ22, DPD13, FMS20, FP08, FHH22, FZ00, FS16, FB01, FHT06, GM91a, Gib83, GR88a, GA91, GRZ24, GYZ22, GG18, GM09, HLL^{+}17, HXY21, HZ22, HP09a, HO08, HJZ17, HHN18, HJX22, HH17, HYZ21, HWY21, Ich76, Ich82, IY12, IT87, JZ22, JS81, KK92, KP04, KY23, KZ00b, KT03b, KMR10, Kum83, Kun82, KM04, Kwo80, Kwo91, LT83a, Lee86, LY90, LC80, Li95, LZ18a, LY18, LNW23, LWZ22, LZ01, LP87, LMKW17, McE09]. Quadratic [MWY21, Mil88, Moo21, NS00, NH12, NLZK19, OC04, PB96, Pap81, Pap86b, Pic93, Pre96, PS87, RRJ16, RMZ02, RV19, Ris86, RX15, Roc87, RW90, RW92, Rud78, SLQ^{+}23, Sta98b, ST90, SY14, SLY16, SD19, Sun21, SX23, SWW23, SY24b, SIC00, STZ24, Tan03, Tan15, Tre89, Tse93, Uch83, VK81, VSB18, WZ17a, WYZ23, WY21, WT98, WM12, YZZ01, Yon02, Yu15b, Zha05, ZDM20, ZS19, Zhu96, Zol78, vWCET23, DMPD12, HJZ12, Yon13]. Quadratically [SD19]. Quadratization [Car15]. Qualification [BC96b, BC96c, Gün95]. Qualifications [GT90]. Qualitative [BO16, Kha03, SQS09, TT20]. quantitative [Dai88]. Quantization [BMT01, New00, Yük17, Yük19]. Quantized [CB09, CF18, DJ13, FZ05, FCG16, JLR09, MLZ17, ZZ13, ZZWY10]. Quantizer [MT23, VL23]. Quantum [ABS18, BLPP24, BM09, BS09a, BPS21, BVJ07, DAR15, DWZ^{+}16, Duc20, EMP^{+}16, GVL12, GP17, IK09, LR10, LAM19, LAM21, MP14, MV07b, NJD09, Rav00, Ste17, VPJ19, VP22a, VP22b, Zha08, ZLHZ12]. Quasi [ABCO17, AHR20, Ama04, Ara14, ABL18, Ban19, BCP17, BB09, BLS05a, BTW82, BD01b, CCT05, CW22, CB15, Cos13, El 17, EL20, FCGO05, HVDK19, Ish88, KMM07, KS87b, LR03, LP86, LZ12, Lou02, Lou03, PV04, PTT17, PT06, Qiu13, RZ20, Sac85, Wac12, ZMK21, ZJ06, ZY08, You87]. Quasi-convex [PV04]. Quasi-Differentiable [Ish88]. Quasi-electrostatic [KMM07]. Quasi-geostrophic [FCGO05]. Quasi-hemivariational [ZMK21]. Quasi-Linear [HVDK19, LR03, LZ12, Lou02, Lou03, ZY08, ABCO17, CCT05, CB15, EL20]. Quasi-LTI [BD01b]. Quasi-Metrics [Qiu13]. Quasi-Neutral [PTT17]. Quasi-Newton [BTW82, KS87b, Sac85]. Quasi-Newtonian [Ara14]. Quasi-Optimal [PT06]. quasi-Riccati [You87]. Quasi-static [Wac12]. Quasi-Variational [Ama04, ABL18, BCP17, BB09, LP86, RZ20, ZJ06, AHR20, BLS05a, CW22, Cos13, El 17]. Quasiconvex [JS86]. Quasiconvexity [Cha91]. Quasiderivative [AG19]. Quasidifferentiable [Sha84]. Quasilinear [CF95, CT09, CHY23, CGW10, CVKB13, GL11, HN22, HY95, Hu15, KY05, LY23, LYZ23, Pen20, SZ82, TL06, VP22a, WW23, You13, ZPQ20]. Quaternion [FKW23]. Quenched [PZ04]. Queueing [BBC18, BGL11, Dup03, KM96, MK90, MSS96, Mey01, Mey03, Wee18]. Queues [CR93, KR89]. Quickest [BP07a, BL15, EMP24, ZHSP14]. Quinn [Flo02]. Quotients [TP05].

R [Kum88]. Rachford

[BC20, BCKM24, Sva11]. Radar [BL14]. Radial [KNW95]. Radially [ABB15, GKL20, Kru00b]. Radiation [DKS^{+}11, MPT06, MY09]. Radiative [GBR18]. Radiative-Conductive [GBR18]. Radii [HP96, KHP06, NS05, ST16, TF11]. Radio [AK02a]. Radius [AJPR14, CLG12, CLMSR00, TG12, ZGS99]. Rain [BÇMS^{+}08]. Ramsey [HK19]. Random [Ahm81, Ban19, BK22b, BR10, BC12, BD19a, Bis76, BHHE03, Bou93, Bou95, BK00, CK95, CC13, CSBC14, CIH18a, CPY09, DDH15, Dor86, Dor90, DMZ22, FCG16, GTQ21, HSZ24, HHS18, HZ05b, Kus79, LS91a, LCF18, LYZ15, LS19b, LRTY22, LL94, MS92, MBS13, OT16, SZ04, SSZ11, STW20, TZ10, TL94, Tan03, Tan15, TYW11, WY21, Xie12, XJ24, XZGT23, YT13, ZDM20, ZS23, dFS15]. Randomization [MTS15]. Randomized [Don24, KS81, PL15, Piu15]. Randomly [CL11, DHJK20, FSS87, TYW11]. Randomness [SSY24]. Range [BH18, Dun95, GBBT78, HBS06, Qua04, SNB09].

Ranges [AM82]. Rank [AP20, BP17, BRT15, CL24, CYL96, CYL97b, DGM89, GGA^{+}17, pJY20, MS02, ORW13, SU18, SD19, Tri76, SY17]. Rank-Deficient [CL24]. Rank-One [AP20]. Ranking [AKTW24]. Rao [ÖH14]. Rapid [Art89b, GLM21, Kom91, Kom97, Urq05, Ves13, Pen89]. rapides [Pen89]. Rapidly [BDR02, OPS87]. Ratcheting [AAM20, BLY21]. Rate [BC23b, BK02, BGL11, CY18, CDJS23, Cha05, Che98, CC88, DY13, Dun81, FF22, Fre99, Ger92, HLM^{+}18, KLM20a, Kle18, KH81a, KY93, LS91a, LYZ15, LLX11, LZ21, LR06, MCP21, MCS13, McE09, MK10, PH92, Rin09, SDSÁR00, Sta79, SCCL18, VJBK24, WS13a, YHH16]. Rate-Controlled [Cha05]. Rate-Independent [BC23b, Rin09]. Rate-Preserving [PH92]. Rates [ACS00a, BA15, BPS05, BK23, CMA08b, CSC14, Col12, DK05, Dun79, Dun80a, Dun00, DJ98, FCJ14, Fu12, GLX18, GP22, GL19, HHF21, Heu94, KK02, KS15, Kus78b, KH79c, MT94, NE04, PZ11, RS21, Sac85, SY09, YZ97, ZC91]. Ratio [FR20]. Rational [AMZ04, BVW00, BGL98, CL24, CH05, FO98, HO01, JS81, NZ99, NvS09, NHG99, RKL16, RKL18, TG13, WS92, WP98, Yam88, ZFKZ21, ZZ23]. Rationally [SRM16]. Ray [FZ19]. Razumikhin [RX19]. RBFNN [YC24]. Reach [Bro10, BG14, HOPS87, ZCGT10]. Reachability [BFZ10, BPZ16, BD19b, Col82b, EC22, EH80, FV11, GZ14a, KV02, KV06, NC16, OMK84, ÖL89, PI99, SC12, Yam89, Col85]. Reachable [BR88, DE18, Fra89, GO11, HO22, Kal82, KS89b, ŁS85b, MGHT19, NS91, Sch86, Sei87, Vin80, Wec84, Wol90]. Reachable/Observable [Kal82]. Reaching [CMA08a, CMA08b, OPS87]. Reaction [BKR17, CLZH24, Cer01a, CSB15, FA13, GV05, ITT04, KK19a, KEK21, Kha03, Le 19, Liu23, OPPU17, YCY13, ZG18]. Reaction-Diffusion [Cer01a, CSB15, GV05, KK19a, KEK21, OPPU17, YCY13]. Reactors [Bd95, GRR09]. Real [BGMR94, CFZ10, DBS05, EMP24, Fio20, KMZ98, LLT18, Mik12b, MR09, RR92, Ros92b, Sus87b, Sus87d, YY08, ZA10]. Real-Time [DBS05, EMP24, MR09, ZA10]. Real-Valued [CFZ10]. Realisation [GCP88]. Realizability [EMP^{+}16, OMK84]. Realizable [Piu18]. Realization [Bor93, Bor10, FRZ93, FS10, GBCG98, Gon87a, Gri88, GP17, Kre87, KS90b, KW13, LF87, LP79b, LP85, ML13, NvS09, NPvS13, NLW91, Pav80, PD79, Por77, RR12, TKH10, Val01, WCLC05, dJ78, vW84]. Realizations [BA08, Bis83, Cro81, CI86, CC87, FNP^{+}22, Fra82, GB86, Hel93, Hug16, Jak80, Jak86, MC91, NvS09, OMS90, OW93, Pan83, San96, Yam88, GB84b]. Rearrangement [LE17]. Rearrangements [BR88, SC12]. Rebalancing [LMKW17]. Receding [AK16, BP20c, IK02a, IK06]. Receding-Horizon [BP20c]. Reciprocal [MG13, San96]. Recognition [DF96, DF98a, Gad08]. Recognizability [BP02]. Recognizing [BSS12]. Reconstructibility [ZZS16]. Reconstruction [AGJ^{+}13, BDGR12, BCS21c, BCS24, HN00]. recours [HU78a]. Recourse [Lep90, Ols76b, Ols76a, RW76, RW78, HU78a]. Recovering [PSJ18, XZ15]. Recovery [FZ23, ZCC07]. Rectangular [LW11b, Zha13, ZC91]. Recurrence [AB10a, CL02, Ich11, MTS20, SX14a, ZY09]. Recursive [APA08, BN81, Che84, FP02, GM91b, Ger92, Ger06, Guo94b, HJ16, HW20, Jac22, JVB^{+}23, JX22, LY91a, LY91b, LT23, MC12, NSW17, WYY17, WY08, YMH94, dJ78]. Redheffer [Van94]. Reduce [Ben76]. Reduced [ABCO17, GL76, GBR18, Sto04, WTJY21, Lom13]. Reducible [JFB17].

Reducing [PZ20]. Reduction [BL14, BD11b, BF07, Bre16, CM04, DTV18, FS10, GK86, GO14, GHS92, HLCH18, HLG21, Iof79a, KM97, KV07b, LZC21b, LG20, MN08, NFI15, Ohs13, PZ22, SM13, SU18, XC03, ZHC06, ZJ89]. Reductions [BL05, BR87a]. Reference [BDRR98, GLK22]. Refined [LL24, Mor95, Mur87, YZ22, Yük20]. Refinement [HSA17, Sas14]. Reflectance [BDR02]. Reflected [Ser09, Shr81, TZK11, KS85a]. Reflecting [Bar21, Ser03, SLG84]. Reflection [Bar20, BISW17, Bou08, TZK11]. Reflective [HTLV17]. Reflexive [DM83]. Reformulation [BH22]. Refraction [HHPY16, LYZ15]. Regime [AM15, BLY19, BL09, CWZZ17, DH12, FS21, FR20, GGP14, Heu15, HLT20, LW10, NY18, SX14a, Sha15, Sha17, Sha18, SX19, ST23, SWW22, SY09, TZ10, VP07, WJW19, XZ17, YX10, ZZ09, ZSM10, ZES12, ZSX18, ZY03, ZY09, Zhu21, ZWYJ14]. Regime-Switching [BLY19, BL09, CWZZ17, DH12, FS21, GGP14, NY18, SX14a, Sha15, Sha17, Sha18, SX19, ST23, SWW22, SY09, WJW19, XZ17, YX10, ZSM10, ZES12, ZY09, ZWYJ14]. Regimes [Car76, DGR18, PVZ09, XZ17]. Region [DHV98, Kim90, Kiw89, KT16, Lag77, Lag89, RS10, Rus86a, ZZCaL19, UUH99]. Regional [BBC14, MRV03]. Regions [BHPM04, WZ17c]. Regression [BKS10, BS20b, BS24, GW96b]. Regressions [BGP20a]. Regret [NOV03]. Regular [BI08, CC02, CDKV18, Col82b, FMT10, GZ12, GMM91, Hal89, Hel76, LT97, LR00b, Lou94, NS86, Obe09, Pau17, PS00, RT06, Sch02, SZ97, Sto97, Sus87b, Tre89, VSB18, Col85]. Regularity [ACS00b, AMPR15, ADG^{+}11, BX12, BRW05, BR21, BZ96, BHL12, BS03, CFS15, Car97a, Car97b, CM06, CR11a, CHY23, Cel15, CV90, CMW06, DJ86, FGG10, FLUU17, GGP02, GT22a, Gor19, GL99, GBo97, GKLW14, GGMZRB20, GZ05, HK06, Kim93b, KW87a, KS13, LZ18b, Ngu16, PV15, QV13, RZ21, RT07, Sch09, Sei82, SS89, Tuc96, VT95]. Regularization [BKM97, CC13, CT06, FK99, Gli12, GT22b, HWZ23, KN03, KS85b, LP10, Man21, MS23, NWW19, PT13, Pue09, Vog90, ZHB21]. Regularizations [HT11]. Regularized [Aih92, Ams13, BP17, NWW17, NWW19, SS24, ZH05]. Regularizer [STZ24]. Regularizing [Fer81]. Regularly [PTY07]. Regulation [CW08, Emr82, Gib81, GZ19, GZ23, HP10, HL92, IP06a, KLM20b, LP15, LK08, LPT22, Liu13, Mey01, MPR17, Pau19, PMI06, TBP18, VB23, WL19, WL21, Wil77a, YGC18]. Regulator [AT96, AF21, Bal91, BK84, BI91, BBE99, BP98b, DMS12, FP08, Fra77, GR88a, Haa82, HXY21, Her91, O’M76, RW92, SLQ^{+}23, Tsa78]. Regulators [CLZ98, CZ00, Fio21, LW87, ZS19]. Reinforced [BR10]. Reinforcement [BM00c, BF15, Don24, FEBL13, HWZ23, STZ24]. Reinsurance [JXZ24, LY20a, MS11, QJL22]. Reinvestments [Pau08]. Rejecting [GZ14b]. Rejection [EÖSÖ98, GZ13, GZ14b, KLR09, Liu23, Rus86b, SBBV15]. Related [ACL09, Ama04, BB97, BRW05, BS09b, BJ17, DG90, DFM15, EL20, Fab08, FP08, FP96, FGL08, FBP99, FHT06, Kan78, Koe20, LZC21a, LB00, MRVM00, MDW23, SZ19, SIC00, Yon93]. Relation [BK83a, Kwo91]. Relations [Bal77, CFS15, EMP^{+}16, GH15, HSA17, Heu15, HN96, LZC21a, Pen84, Pen13, Red21, Rus76, Pen88]. Relationship [CV87, KZ00b, Pop97, Tri78]. Relationships [BL91, Koe90]. Relative [AB12, ABK14, AB17b, BDIW22, FKW23, HHI^{+}13, IRT07, Maz17, MZ22b, Pos17, VP15, XUP08, dSF01]. Relatively [BP07b, RW76]. Relativistic [MST16].

Relaxation [AT17, AB05, Bal81, BK24a, BJ96, BE87, BHT87, BP17, BS06, CDCCMG08, CSB88, FEBL13, FR99, Gon94, Laz88, MH11, Sag17, TA21, Tol16, VC22]. Relaxation-Compactification [Bal81]. Relaxations [LHPT08, Mey03, Rou21]. Relaxed [Ahm83, Ahm96, AC13, AC17, Art78, Art89b, BDM07, Bah08, BJ95, BDT20, BS13b, Car76, Col82a, Col82b, Fat94, GAC24, GW16, Lou07, Man21, PV14, Pap89, Piu15, RR94, RZ21, Ros92a, Ros00b, War77, WP76, XA93, Yük24, Zho92, Col85]. Relaxing [Lic13]. Relay [SFFC08]. Reliability [GS14]. Reliable [CFZ10, GL97a]. Remark [Fat97, Roy89, Wec82, Zha01b]. Remarkable [GGH16]. Remarks [AFCLS22, Kno77, Kob78, KR15, Kum88, Nag95, Tes92, Yam12]. Rendezvous [AG95, Alp95, AL98, AB06, AE95, BG98, CB09, Kim22, LWR24, LA96, LMA07a, LMA07b, XW12]. Rendezvous-Evasion [AL98]. Renewable [KLM20b]. Renewal [AHS13, BM21, HW99, SS19]. Rényi [JNKV21]. Reorientation [ZTC16]. Repeated [Alp88, CLS12, Lar01, SQS09]. Repetitive [DRDG08, Imm07]. Replacement [And88, Gal84, Öze85]. Replenishment [PYB20]. Repositioning [SZ10]. Representation [AM17, BET07, BLQR20, CR17a, CF14, Dav76a, Gai04, GR17b, Ger06, KN03, Lou94, Mis19, PP13, Qua03a, Qua03b, Whi80b, Yam88, Zam98, ZZS16, Iof83]. Representations [Bor93, Bor10, CG00a, DS93, DvdS98, Dun00, FW93, JSV05, Kha82, Lia15, NZ99, Ram05, RW00b, TG13, WZ94]. Represented [McE09]. Reproducibility [Gra88]. Repulsion [CC02, Cza04b, GGMZRB20]. Required [MP21]. Requirements [Heu15]. Reservoir [Lam87]. Resets [OM14, TCC^{+}15a].

Residence [DF98d]. Resilience [GH17]. Resilient [CKM20, ZM14]. Resistance [LGZ13, RFF15]. Resistance-Based [LGZ13]. Resolution [AGJK12, Qua80]. Resonance [BCM13, WL95]. Resonant [Bar97]. Resource [KLM20b, RN13]. Resources [BS19b, CFR13, RRR16]. Resp [LY23]. Respect [BFCLS97, BHT19, Bor77, Dat88, DE14, FLUU17, FS23, HRW88, HG02, LS99b, LRW96, OP96, OC23, RT98, RW92, Sla00, Sla03]. Response [CS84, DZA19, Hag02, LF87, SS09c, SMK18, ZH02]. Responses [WFH13]. Rest [Bia83]. Restarting [Doa21]. Restrictability [ÖW92]. Restricted [ARS14, CD13a, HO90a, KG80, Zhu21]. Restrictions [BF15, HG90, MAG11]. Result [Bal01, BCS17, BGBPG04, CF24, CKL16, CZ95b, Day09, FGB95, GV00, Mik12a, Nov04, Zua96, Zua98]. Results [ACS00b, Bal94, BDFCGB12, BM13, BK99, BK00, Cel07, CP13b, CM96b, DL06, Da 85, DG90, FCGIP06, Goz91, Guo94a, HT11, HM83, HS87, JS99, Kum85, Lou07, Mal23, MT94, MK21, MBS13, MH11, MM80a, MO99, NP06, NHG99, ÖH14, Pav84a, PT09a, PA99, RRS14, Ros92b, Sha14, Teb11, WŽ07, YCRZ21, Yük20, Zha05, ZJ89, Zhu94, ZzJh10, BG88]. Retarded [HP21, KS87a, KS89a, MT78, Man81, MGMBN^{+}19, Pep14, Pep16, SO06b]. Retirement [PW23]. retour [Loi88]. Retrieval [BL03]. Retrieving [CT06]. Return [Gui13, PW23, Loi88]. Revealing [Wu18]. Revenge [BB92]. Reversal [DKL09]. Reversible [CP13a, FP14, HYZ24, LD05, MZ07]. Review [MP07]. Revisited [Bie95, CK98, HZ21, LMR97, MNB11, PDB23, Pue09]. Reward [AB17a, ABBK20, CLRJ13, CP13a, DY13, DG20, JN19b, Küe07, MR90]. Reward-Rate [DY13]. Rewards [GC05]. Reynolds [BBR18]. RHC [AK19, AHK24]. Riccati [Cur13, BGMS78, BCG86, Bou95, BT21, BK14, Brü13, BK83b, BvdMR05, Buc75, Buc77, BSZ08, BR15b, CLT06, CF06, CP76, Cur03, CS11, Cur11, Dav78, De 92, DGJ93, DH22, Del07, DM15, DA17, Du15, Fla84, Fla87, GL90, GL92, Gah94, Gib79, Gib83, GLR86, GT05, GT14, HH76, HK21, Hei16, Hel76, Hew93, HP09a, HZ03, HT22, IY12, IO96, IM98, JMP21, KS87a, KS90a, KH90, KT03c, Kra93, LT83a, LT86, LL07, Lev85, LNW23, LLC18, LMP95, MO02, Nih83, OCI00, Pav80, PD97, QZ13, Qu14, RMZ02, RR92, Sha83a, Sha83b, Sha85, Sha86, Sti18, Tan03, Tes92, TR01, WYZ23, Wim92, Wim94, Wim06, WJZ12, Xie12, You87, Yu15b, Zei94, ZDM20, ZQ18, ZLD12, vN04]. Riccati-Based [BK14]. Riemann [LL15]. Riemannian [ATV13, ABHY23, ABB04, AM14, AJ16, BBG12, BM18a, BC95a, BG13, DZ24, Fio20, GGP02, GC22, ISdCN99, LY12, LWR24, PS88, Sac19, TC13, TDNT18]. Riesz [CW08, Guo01, Guo02, GX04, JKZ21, Wak21, XS96, XG03]. Riesz-Spectral [Wak21]. Right [BR15a, CG00a, DDH15, FKN96, KQWY20, Kus79, ÖB19, SV79]. Right-Hand [BR15a, DDH15]. Right-Inversion [FKN96]. Rightmost [QS13]. Rigid [CBB19, Eno94a, Eno94b]. Rigidity [JZLW18]. Ring [CP95, SAMB14, WZ77]. Rings [Emr82, Kha82, MA01, Oht08, Sul94, Sul98]. Risk [AD98, AB17a, AKP21, ABBK20, AB20, AGS13, AM15, BM21, BE95, BN97, BBF18, BPS05, BLY19, BB23, BK10, CCHH19, ÇR14, CS12b, CW23b, CW24a, CTZ03, CHP^{+}11, DM02a, DGR18, DL13a, DR06, DS99a, DS07, DH12, DM97, FM95, FHH97, FZ98, FHH99, GG09, GSZ20, GL19, HJ15, HL16, HHSS23, JPS20, KT03b, LZ16a, LZ18a, LR23b, LZ06, MY17, Nag96, Nag03, NS20, NS12, PB99a, QJL22, SSO13, WYZ17, Xia11, Zha95, BC18b]. Risk-Averse [AKP21, BB23, ÇR14]. Risk-Aware [HJ15]. Risk-Sensitive [AD98, AB17a, ABBK20, AB20, AGS13, BE95, BN97, BK10, CCHH19, CW23b, CW24a, DM02a, DL13a, DR06, DS99a, DM97, FM95, FHH97, FZ98, FHH99, GL19, HHSS23, JPS20, LZ16a, LZ06, Nag96, Nag03, PB99a, SSO13, Zha95]. Risks [CZA08]. Rissanen [Dye77]. Ritz [Las84, MR79]. RKH [MR86]. RKHS [BP20b]. Robin [JZ09b, CSFCL^{+}23, LtLR21, XZ15]. Robot [GK02]. Robotic [BCF12]. Robots [GL06]. Robust [AD15, Akç01, AR21, AAA12, AB02c, ABP14, BHPM04, BM06, BCD02, BBTN03, BY14, BCP16, BCP17, BC23a, BDLW24, BET07, Ber89b, BC21, BF05, BJ19, BCC^{+}19b, Bli04, BD01b, BL00a, CT13, CJK13, CM92, CL19b, CdSBT09, DM02a, DKS23, Dat88, DHKS18, DM97, Dup03, FNA02, FHMS18, Fer97a, FGS93, FK96, FIM09, FB12, GAC24, GBGdO00, GdOB02, GBB18, GHJ00, GP02, Hag05, HP10, HZB13, HP94, HRMK17, HH17, HHT19, JSV05, KLN21, LP15, LP10, LC19, LAM21, LWZ22, LSW96, LF15, LFC21, LAN04, LZ92, LDZ10, MPN98, MA19, MU22, NK94, NS18, NNSM08, NS14, Orl04, PW23, Pau17, Pau19, PD02, PLK89, Pri05, PT06, RM92, RCS01, SC12, SJ13, SFFC08, SBT19, Szn93, Szn95, TZQ21, TF11, TFT11, VPJ19, WW83, XUP08, XWZ^{+}21]. Robust [Yan96, YD08, ZK20, Zor17, vTY99]. Robustness [ABL13, ALR99, BD17b, BK98, CG00b, CG90, CZA04a, CZA08, DF21, FIR06, HRW88, IM08, IM10, IK95, Jön09, KY19, KY20, KT05, KY23, LRW96, MCL^{+}22, MZ15, MA19, MMÖ18, MTS15, MY22, MW23, MVHN09, RT98, SO06b, SR18, SKXL17, ZHB21]. Rocket [ZTC16]. Rod [Han94]. Roesser [PEGR18]. Role [AKTW24, CF18, SSY24, WRGA14]. Rolling [CMK15, Gro12, Gro16]. Roomy [Dew76]. Root [BHPM04, BM06, BGH94, RTT20, Tie14]. Root-Clustering [BHPM04, BM06]. Root-Locus [BGH94]. Roots [MGMBN^{+}19]. Rosenblatt [Mik12a]. Ross [BPS05]. Rotated [Oss01]. Rotating [BPS21, CCW14, HCG20]. Rotation [Bil19]. Rotors [Bay09a]. Rough [ACPZ19, BD15]. Routh [AJ77]. Routing [ABB15, Cha14, MS92, MK90, Mey01, Mey03, SPBF10, SN82]. Row [LC80]. Row-Action [LC80]. Ruckle [AFK10]. Ruin [BZ15a, BZ15b, LY20a]. Rule [Bus84, CLdRdP11, Dun80b, War78a, Zha01a]. Rules [BM11, ÇD21, CHPA20, Dor90, Dun80a, HHS18, HU84, LGP15, TWZ21, Ye00b]. Run [Bor89, BGS19, BLQR20, FGL08, GR06, JPS20, JS24, Rob81, Ste19, dVC22]. Runge [SP96]. Running [HS91, KPZ18, Sap19, TDBL22, Wee18]. Russian [Iof78].

S [Cou83, LL20]. S-Shaped [LL20]. Saddle [Beh79, Ber85, CGC17, Del07, DS09, HQR24, Stu15, SZZ97, SY14, WCLC05, Ber88]. Saddle-Point [CGC17]. Saddlepoints [BMM85]. Safe [RLXS23, RN13, GK02]. Safety [HD11]. Sakawa [Mar77]. Salami [ELO22]. Salamon [van95]. Sales [Tam80]. Sample [BCL22, GC05, HLVAC99, Kam10, Pep14, PRG16]. Sample-and-Hold [Pep14]. Sample-Path [GC05, HLVAC99]. Sampled [AMV18, AVZ20, BT21, BD01b, FS16, FA13, GN03, Hag02, Hag05, HIK18, KK16, KLR09, LM02, LRT05, Log13, LSL^{+}18, MRP99a, MRP99b, NLPT09, SK94, Sta79, TZZ88, TWZ16, TS95, WOH18, Wak21, WF24, Yam12, YC24]. Sampled-Data [BT21, BD01b, FS16, FA13, GN03, Hag02, Hag05, HIK18, KK16, KLR09, LM02, LRT05, Log13, LSL^{+}18, MRP99a, MRP99b, NLPT09, SK94, Sta79, TWZ16, TS95, WOH18, Wak21, WF24, Yam12, YC24]. Sampled-Time [AVZ20]. Sampling [BK00, CCHP12, DCI07, FP11a, KK12, LLM^{+}24, Pep16, RMB12, RW92, Sag13, TYW11, WŽ07, WRGA14]. Sandpile [Bar17]. Sandwich [LL00b, ÖH14]. Sard [Rei83]. Satellite [AK02a]. Satisfying [BBV10]. Saturated [Dus21, GBB18, PCP22, Sag15, Sag17, XX19]. Saturating [MCPA17]. Saturation [CHL15, De 00, LCS96, LRT98, WQS^{+}24, YC06]. Saturations [ZLD10]. Scalable [Cha14, NFI15, RBS21]. Scalar [AM98, AN21, ADM17, Bli04, Dus21, ET12, FZ05, Gfr06, Ho86, IS12, IS14, JS99, Léa12, Mos80, Per12, Rai99, Reb89, Sus83, Teg84, FCGIP06]. Scalar-Input [Sus83]. Scalarization [Khá93, TFL14]. Scale [DRW10, Doa21, GT22b, HYZ11, KP97, LS99a, Les12, LB89, MP11b, WZY06]. Scaled [Dun91, Dun93, HP09a]. Scales [BT13]. Scaling [UU00]. Scanning [Kha94]. Scattering [SW12, WS13b]. Scenario [Col88]. Schauder [KR80]. Schedule [KK12]. Schedulers [MBL00]. Scheduling [BB89, BM84, Cat98, KEK21, LFH15, MT23, Sch15a, STZ24]. Scheme [BF90a, BBC18, BV13a, BR07, CS77, DBS05, Doa21, HLZ20, LG20, MV11, MU22]. Schemes [ABL18, Baş84, DJ98, FGQ97, FEBL13, GP79, Grü09, GPSW10, Kun82, RZ20, Sor98]. Schmidt [DGJ93]. Scholes [FP11b]. Schrödinger [ABK19, BLT15, CS18, CGP22, CM17, CLZ23, DJK^{+}23, FZ19, HRCK21, IK07b, KGS11, LL21a, LT13, Lü13, Mac94, MM23, Phu01, RZ09, WCG14, Zha15]. Schrödinger-Type [LT13]. Schubert [RRW96]. Schur [FO98]. Schwarzkopf [CLdRdP11]. Screening [LMJ09]. SDE [BJ17]. SDEs [Bar23a, FF22, JL23, MSZ18, PRS21, WY21, YCRZ21]. Search [AA86, AG95, Alp95, Alp10b, AFGT13, AA92, AE95, AN06, BB92, For77, Gal79, Lid13, WFH13, WCLC05]. Second [ADK08, Ahm77, Alv00, AY23a, ANS14, ABK21, BLJ15, BS16, Bet19, BDP14, BC16, BLSZ19, BFL16, CD91, CFS15, CTU00, CM02, CT09, Cas12, CT14, CRT15, CWW17, CWW18, CW20, CHN94, Cha82, Cha87a, CLYL13, CZZ24, CM96a, CC90a, CMT03, DL06, De 11, Den22, Dun95, FG15, Fio20, FT13, FO18, FNV10, Gfr06, GGH16, GT93, GRS00, GR21a, Han86, HMV05, HK01, HU82, HU84, Hoe12, Hoe14, HM02, HN96, Iof79b, Kel24, KN14a, KTY18, KQWY20, Kim93a, KBvdSD17, KR13, LR21, LH88, LLC18, LM08, LZZ21, MO02, MO04, MDW23, Nie21, PKL13, Pen98, RT03, Sak84, Sar97, SNB09, SL18, SFFC08, SW23a, ST09a, Tha86, Tri78, Vuo21, WL95, War84, War85, WM10, WT83, XLL08, Zei01, ZZ15, ZZ17]. Second-Order [AY23a, ANS14, Bet19, BDP14, BLSZ19, CD91, CFS15, CT09, CT14, CWW17, CWW18, CHN94, Cha82, Cha87a, CLYL13, CM96a, CC90a, DL06, Den22, Dun95, Fio20, FT13, FO18, Gfr06, GT93, GR21a, HMV05, HU82, HU84, Hoe12, Hoe14, HM02, HN96, Kel24, KN14a, KTY18, KQWY20, Kim93a, KBvdSD17, LLC18, MDW23, Nie21, PKL13, Pen98, RT03, Sar97, SNB09, SL18, WL95, Zei01, ZZ17, ZZ15]. Secondary [KvdSD21]. Section [BL14]. Sector [Jön09]. Security [LMJ09]. Seeking [LK11, PKL13, RSS18, Sut20, TDNT18]. Segment [Rzy09]. Segmentations [Sha92]. Seidel [WZC20]. Selecting [ES81a]. Selection [AEÉ^{+}20, BC11, BVW00, BR07, CABP17, FK21, HBSW07, HZ05b, Iof78, LZL02, Now85, Tam80, Wag77, WJW19, WYZ17, Xia11, XZ07, XZZ20, ZSM10, ZY03]. Selections [Liu95a]. Self [Bar17, BR10, DHK20, HTLV17, KM18, KP87, LW87, RM92, RT16, GX04]. Self-Adjoint [GX04]. Self-Consistent [DHK20]. Self-Interacting [BR10]. Self-Organization [KM18]. Self-Organized [Bar17]. Self-Reflective [HTLV17]. Self-Stabilization [RM92]. Self-Tuning [KP87, LW87, RT16]. Selling [DYZ12, RH06, Zha01a]. Semantics [Col11].

Semi [BC96d, Can84, Fle85, GG09, Jas02, Kim92, KA06, LW24, OM14, PMH92, PH92, SS03b, VA03, Zha09, ZLD10]. Semi-Definite [Fle85]. Semi-Dirichlet [Zha09]. Semi-Global [ZLD10]. Semi-Infinite [PMH92, PH92, SS03b, BC96d]. Semi-Internal [Kim92]. Semi-Markov [Can84, GG09, Jas02, KA06, VA03]. Semi-Markovian [LW24, OM14]. Semialgebraic [Gho86b]. Semiconcave [Rif00, Rif02]. Semiconductor [PTT17]. Semicontinuity [Bal81, Bal84, Fer81, Iof77a, Iof77b, KTWY12]. Semicontinuous [AKN00, ENW95, Fra93]. Semicontractive [Ber18]. Semiconvex [Mif77]. Semidefinite [ALY13, All88, AW90, FNA02, HHLP16, Hew93, KM07a, LLS21, MGHT19, Wim94, YZZ01]. Semidiscrete [HW14, HKW17, Mic09]. Semidiscretization [LX23]. Semidiscretizations [CMR17]. Semidiscretized [LG20]. Semiexplicit [Bro22]. Semiglobal [ET15, JL14, SFFC08, TP95]. Semigroup [CLMSR00, Fle82, Ves13]. Semigroups [Ben78, BR82, BIP88, CSES21, Fra87a, JZ04, NR21, Pen13, Teb22, VB23, dRSV09]. Semilinear [AX96, ACS00b, AFO22, Bar95b, Bet19, BHS19, BGBPG04, BC89, BC95b, BS11, Bro87, Car12, Cas93, Cas97, CTU00, CM02, CR06, CK14, CKK17, CK17a, CMT18, CMR19, CK22, CW23a, CO03a, CCFV20, Che99, CLM24, CT04, DL09, DMPD98, DMPD00, EM19, FYZ07, Goz91, Gra82, HR21, HY95, HLZ20, JL14, KK19b, Kha01, Kha03, KN14a, KLM^{+}21, LW07, LPT22, LX23, LFL23, LY09, Mah03, MPT06, MY09, MS17a, MT22, Nai87, NPR15, QW19, RTZ20, RX15, RT07, Teb11, Trö99, WY16, Zha18, Zho83, Zho84, de 98]. Semimartingale [BC11, BK80, MT03, ST14]. Semimartingales [Mao90, Pic93].

Seminormality [Bal86]. Semipassive [DZ76]. Semipermeable [Car97a, Car97b]. Semisimple [AD17]. Semismooth [Cha87a, IK10, IK11b, KKV11, Mif77]. Semismoothness [CW23c]. Semiuniform [FKZ22]. Semivectorial [BM13]. Sense [AW83, PSJ18, Pep14]. Sensitive [AD98, AB17a, ABBK20, AB20, AGS13, BE95, BN97, BPS05, BLY19, BK10, CCHH19, CW23b, CW24a, DM02a, DGR18, DL13a, DR06, DS99a, DS07, DM97, FM95, FHH97, FZ98, FHH99, GL19, HHSS23, JPS20, LZ16a, LZ06, Nag96, Nag03, PB99a, SSO13, Zha95]. Sensitivity [AHJM10, BV10, BKP18, CFS15, Chr19, CL86, CW91, DZ88, FKT86, FTZ87, GL90, GK87a, GM05, GSS19, GI04, HK88, Jou00a, Jou00b, KH90, KY92, LS91b, Lev99, Lev01, LL18b, Liu95a, LY01, LY02, MERD02, NS00, Sha88a, SB92, Sha94, Sok87, Sok88, Spi80, Tad97, Tan88, Ulb02, Ver97, VG95, YMH94, YK91, Yan95]. Sensor [BB89, CWC^{+}22, DH09, DPV00, FFR14, FCMA09, FCC20, FFP^{+}09, HRMK17, Kha94, MS05, PBPB15, ZVBR14, ZM13, LZSC24]. Sensors [FGGH15, LZSC24]. Separable [BW88, HHF21, Sto76]. Separated [LB98, Pul95, Pul96, RRJ16, WZY09]. Separating [LZT20]. Separation [AISW04, Cob12, CI77, Dav76b, Les12, RN13, Ris94, Yük19]. Sequence [Gli12, WH20]. Sequences [Bis83, CNP98, Dun80a, GBCG98, PW84, VP15, VP22b]. Sequencing [Mey01, Mey03, Wen15]. Sequential [BMT01, DF96, DF98a, DY13, FCMA09, FNA02, FPD08, FP11a, GRR09, GYBL15, Kre78, KK76, Kus78b, MD89, RN13, Thi97, Whi80b, Zer03]. Serially [CDKP87, GX04]. Series [Bul01, CFNG10a, Cro81, CC87, FGT95, FV95, Fuh76, GL05, Hei88, Nih83, SCFV97, BC05b, Cla86, Kup85, Cla86]. Serre [SO13]. Server [AGS13, MS92]. Servomechanism [LI94b]. Set [ABL13, AHR20, AM98, AC05, ADM17, AR77, BT91, BB18, BP13a, BG05, BD77, BF15, BMN19, BK08, CDV16, DE18, ES81a, FL20, FPT15, FBJ09, Fra87a, Gai04, GH12, GZ14a, GV05, Gui13, Gün95, GJN08, HT92, HP22, HW05, HT10, IK08, IK04, IK07a, Jah85, KK16, Kas09, Kas13, KTWY12, KHJ14, Kum82, LY12, Liu11, LLT18, MP76a, MAJ18, Mey77d, MW78, MVBC18, Mis19, MS97b, NS91, NY18, Pan13, Qiu13, Reb77, RN13, Rot77b, Sei87, Szn93, Tre99, Urs82, Vin80, Wol90, XW12, Iof83, dG06]. Set-Covering [RN13]. Set-Induced [FPT15]. Set-Valued [ABL13, BD77, BMN19, BK08, FBJ09, Fra87a, Gui13, GJN08, IK08, Kas09, Kas13, LY12, MP76a, MVBC18, MS97b, Pan13, Qiu13, CDV16, XW12, Iof83]. Sets [Aub01, BJ11, BH99, BJ02, BR88, Cha91, CSSZ20, CP13b, CCS18, Del04, Dun95, EMH08, Fra89, GN02, GO11, GY12, Her78, JR93, JV17, KP97, KHP06, Kaw11, Kor20, KR17, KS89b, ŁS85b, MGHT19, MP76a, MSH91, NLPT09, Rai99, RG14, TSNvL08, Vin14, WZ17b, WR14, YZ16]. Setting [AVJ21, ALP10a, ABK19, BP05, BR17b]. Settling [AB09]. Setup [KM12b, XYZ21]. Several [Che76, EK85, Lom13, Van92]. Severe [LL18a]. Severely [QC95]. Shadow [AC79, CSZ13]. Shallow [Bel00, CGY03, Yao00]. Shallow-Water [Bel00]. Shannon [MS07a]. Shape [ADK08, ACT10, AHH17, ABV13, BL14, BHOP14, BP13a, BFL16, Bou17, BHT21, BH22, BV14, CN24, CDM21a, CDD20, CHP^{+}11, CRST18, DDK12, DDH15, DZ88, DZ91, Del23, Doy94, EHS07, GMN22, GL95, GI04, GK98, GY12, GY13, HL11, HRPT21, HS03, KL16, KO18, Lam92, Lau99, LWY21, LR92, LNT02, MY93, MAJ18, Nov04, PT08, PT13, PS10b, SS09c, SS23, Sok88, SZ99, SZ03, Stu15, Stu16, SS18, Tib11, Ton03, WK13, WWY06, ZM11, dG06]. Shape-Topological [KL16]. Shaped [ACS00a, AC18b, CIP18, GDL11, LL20, PCP22, PR14, PS04, Dág04, Mar99]. Shapes [GO11, LN10]. Shaping [BEL02, Cha10, CO22, CDM23, Gha11, NCL13]. Shaping-Filter [BEL02]. Shapley [BS15, VA03]. Sharing [BLS12, GW20]. Sharp [BLR92, BBGBO14, BGHR16, BF93, Fu12, GST20, Hu15, Li94a, NZ03, SW99, Teb22, WZ24]. Sharpness [BKL^{+}22]. Shear [Ara12, Ara13, CN24, CCW14, KGBS08]. Shear-Thickening [Ara13]. Shear-Thinning [Ara12]. Shedding [ASK05]. Shell [AHH17, Les12]. Shells [CGY03, Yao00]. Shielding [CHY23]. Shift [Fra80, MP99, ZLHZ12]. Shinzo [Cou83]. Ship [GS19, ZR23]. Shocks [MA19]. Shooting [BH07]. Short [Sac19]. Shortest [GW96a, KN14b, PB99b, Phi94]. Shorting [LZL02]. SIAM [AW90, Ber88, Col85, Cou83, Cur85a, FR82a, Flå95, GK95, GRC82, Kum88, Mos82a, Pen88, Shi90, SC81, ZZ94]. SICON [Yin22]. Side [BR15a, Ros00a, SV79]. Sided [ABCD22, BBT14, CFM19, Del23, DS99b, FFR23, KM12b, LY19a, MM16, NT16, RSS18, DF98b]. Sides [DDH15, GR17b, Kus79, Now89]. Siegel [KX04, KX02]. Sierpinski [Liu18]. Sieves [NP82]. Sign [Hew93, KS93, TR01]. Sign-Definite [TR01]. Sign-Observations [KS93]. Signal [Ave19, FCG16, FV98, Jac04, Jay10, JM17, MPN98]. Signaling [CKT24]. Signalling [CKT20]. Signals [BCL04, FM16, IP06a, IP06b, LP15]. Signed [AW21, GH20, SZSC21]. Signorini [BS87]. Similar [Lou15, RL06]. Similarity [BFZ00, GK86, HSA17]. Simple [Ben81, Ber82b, CP02, CMOZ02, Ete19, FH94, LM97a, MDB05, Pro22, Sîr15, ZC96]. Simplest [AJ77]. Simplex [Yan96]. Simplices [Bro10, BG14]. Simplicial [Yan96]. Simplicity [OZ00, OZ03]. Simplification [PB96]. Simplified [CSB15, Ste07]. Simplifying [FTV98]. Simply [Cor99]. Simulated [Cat98, CC94, CK89, DM99, FK91, Roy89, Tro96]. Simulation [Gra07, GH15, Hey76, JV06, Lam87, LZC21a, van86]. Simulation-based [JV06]. Simulations [CMC81]. Simultaneous [BGMR94, Dág18, Gho86a, Gho88, GB93, HMQD99, ID87, Niu21, SZ03, TZQ21, TW00, Gho86b]. Single [Ben81, BH07, Bos02, BGL11, DG23, FM16, Iof83, Kan00, MRL12, MS92, Mey77c, MR09, Nie14, RSV04, SO06b, SZ10, Sus87b, Sus87c, Sus87d, Tad97, TR02, VP22b, ZBB07]. Single-Atom [MR09]. Single-Flux-Quantum [VP22b]. Single-Input [FM16, MRL12, Nie14, Sus87b, Sus87c, Sus87d, TR02]. Single-Output [Ben81]. Single-Server [MS92]. Single-valued [Iof83]. Singular [AS13, Alv01, AB02b, AM17, AJK14, AG97b, BB98, BDM07, BK24a, BFT96, Ber82a, Boe05, Bon85, BCM13, BG07, BK10, BK16, BDP24, BP98b, BR07, Cam76, Car76, CY12, CJT08, CE06, Cob06, Cob12, Coh21, CS15, Cza04b, DFM15, DEO24, DD03, DV83, DM04, DM06, Dun79, EAK86, EMP18, FP02, FP08, FFS20, FFRR21, FFR23, FS21, Fer18, FR14, FK76, FH17b, Fu23, FS10, FT02, GL97a, Gli12, GYZ22, GHQ15, Gra99, Grü98, GT22b, GT08, GT09, Haa82, HLL^{+}17, Hag02, HS95a, HS95b, Her76, HO90a, HD87, HQZ16, IT17, JHMG06, KK23, KK79, KV21, Kre77b, Kru00a, Kru00b, Kum13, KR88, KM91, Kus13, Kus14, LLS90, Lew80, LY20b, Ma92, MW09, MK90, Mor10, MS07b, NS00, NF13, O’M76, OC12, ØS12a, ÖL89, PS13]. Singular [PS18, QBP23, Ram94, Reb78, Red18b, ST23, SS89, Son93, Sti18, ST90, Sto91b, Sun96, Tak97, Ter94, VS18, VS20, Wee07, WKS86, Yan14, Zha95, Dai88, KS84a, KS85a]. Singular-Controller [BDP24]. Singular-Point [Sun96]. Singularities [BCS20, BGSS21, Hir02, KS99a, Nic96, Nic97, Pic96]. Singularity [GS19]. Singularly [AL02, Bin76, CK78, Cob10, CG92, CD12, DDG12, Don10, DMS02, DMS12, Gai92, GN02, Gai04, GFF17, GO11, Gra05, HM08, HP14, LWfZ22, PB96, Shi02, Vig97]. SISO [GEF17, HO97, RSMG17]. Sivashinsky [CIP18, CMP15, GCL15, aJM18]. Size [Cha91, Pfl86, RS13]. Skew [HS13, LZM05, MN77a, ÖT90]. Skew-Symmetric [HS13, LZM05]. Skorohod [CR17a]. Skorokhod [GTT16a, GTT16b, Kru00a, RTT20]. Slice [Cou00]. Slicer [ELO22]. Slider [dNCP20]. Sliding [BCG^{+}17a, BK21, CRZ^{+}18, GHZ98, GW18, KP04, MVBC18, PKL13, POU11]. Sliding-Mode [MVBC18, POU11]. Slip [CSFCL^{+}23]. Slippage [CC18b]. Slipping [Gro12]. Slope [Cel02b, FT12, GT22a]. Slow [AG97b, GR17a, GFF17, HZ22, MG13]. Slow-Fast [GFF17]. Slowly [FH17a]. Smale [LWR24, DHK20, DHJK20, SS21]. Small [AHJM10, ACKM22, AL12, AD98, Amm02, BN97, BP14b, BS04, CNS10, CCF07, CRS21, DRW10, Dat88, DK86, FP89, GW15, GW16, GFF17, Jaf20, JL16, Joh05, JM05, KJ07, KK19b, KZ23, Kra09, KS89b, KH81a, LFC21, LRW96, Mir21, RBA22, Sac97, ST13, Tre90, TSNvL08, Zha98, dNCP20, dFS15, KLS87]. Small-Gain [CCF07, GW15, GW16, JL16, JM05, KJ07, KZ23, LFC21, TSNvL08, dFS15]. Small-Gain-Based [KK19b]. Small-Time [AL12, BP14b, Jaf20, Kra09, KS89b, dNCP20]. Smallest [CLG12]. Smarter [ZSS^{+}20]. Smith [Hug16, LFC21]. Smooth [Aey85, BZ10, BZ96, CRZ^{+}18, CC99, Cor94, DMK90, DF98c, Fio21, Goe05, GT09, GW09, HT11, KT05, LW12, LSW96, LT92, Ma92, Pol88, PS17c, QLD01, Rya94, TWZ21, YW07, You17, ZY08, van82]. Smoothed [TS24]. Smoothers [ABP14]. Smoothing [CY99, CF06, GM17a, GM21, MT09b, Ouv78a]. Smoothly [Tsi95]. Smoothness [BB18, Bay09b, CD15]. Snell [JN19a]. Sobolev [BRW05, Pav13, Sim76, WP98, WWZ14]. Social [BTB16, HWY21, JFB17, LWZ22, TWB23, WZ17a]. Soft [OMFMC^{+}21]. Solar [GW96a, KV21]. Solid [IK02b, LI00]. Solitons [Mun14]. Solution [All88, AW90, AHR20, BD83, BRS11, BHOP14, Bis12, BISW17, BP07b, BKK02, BK13, BDT20, BDDSV05, BHT21, CW23c, DA02, Dok05, DKRV13, DM15, DA17, Dup03, Ega08, Fab08, Fla90, Fra87b, Gri83, GK98, HK13, HS12a, HW99, HJX19, Hui92, JL93, Jun00, KS95, KTWY12, Kno79, KR15, Lev01, LY12, LZ18b, MP94, McE07, NLZK19, NSW17, O’M76, PT87, Pap81, Pap86b, Pee20, QBP23, RR94, RU19, RU21, RTT20, RS20, Ros00b, SKD99, SP02, Tad97, VP07, VS20, YY05, YY08, ZL02, ZQ18, Qua80]. Solution-Dependent [BHOP14]. Solutions [ACS00a, AKN00, AB02b, AP94, AMPR15, AW83, Aub11, Ave16, Bar85c, Bas77, BHM14, BS16, BBB24, BCP16, BMS15, BCS15, BCM12b, BS11, BK95, BZ96, BT11b, BKR17, BS04, BS09b, BCS13, BvdMR05, BL08a, BR15b, BIRS20, CP00, CR11a, CHJ87, Car90, CRZ00, CCK12, CCK13, CW23a, CO03b, CM10b, Cel15, CM95, CV83, CG17a, CS89, CV86, CdLdDM02, DD92, Dav76b, Day09, Den97, DR05, DC06, Dmi92, DS99b, DM02b, EÖ88, EL20, ES81b, Eng10, Fat97, FGG10, FP96, FT77, FM08, Fra93, Fre99, FP11b, FT04, FMT10, Gar76, GLD^{+}91, GLR86, GP94, GŠZ05, GŠZ10, Gro16, GDL11, Gui13, GGC23, Gup20, HSW19, Hau94, HT12, Hen80, HTY01, Heu15, Hew93, HZ10b, HMR15, HCM05, Hua07, IM98, IK09, KS99a, KM07b, LW07, Lei85, Lev99, Li94a]. Solutions [LW14, LM16, LY19a, LS19a, LY23, LS85a, LT18, Lor87, Mal02, MS87, McE95, MD15, MMR17a, Mik12b, Mil96, MRP99a, MT09b, MM16, NR97, NR21, OC12, OCI00, OvS00, Pav13, PD97, Per81, Per12, Pla21, PS06, PPZ06, Pul95, QZ13, Qiu18, RR92, RTZ20, RX15, RT07, Rou86, SBY23, SW78, Sha15, ST02, SY24a, Sor93, Sor96, SV79, Ste13, Ste15, Świ20, TTZ23b, TLPA00, TR01, Tri77, Tri80, Ulb02, VA03, WC89, WJW19, WWZ14, Wim92, Wim94, Wim06, YW07, Ye00a, YS17, Zas97, ZY08, ZYL97, Zhu92, ZFKZ21, dVC22, vTY99, LS86]. Solvabilities [SLY16]. Solvability [BS18, Cir14, Du15, FKN96, Kre00, KK06, SGM19a, van95]. Solvable [Alv01, AMR14, CNS21, Cro84, Da 16, Dai12, KV21, LHB16, Mey77c, Mey77b, Vie00]. Solver [HT10]. Solvers [KK17a]. Solving [AGL08, ASTT09, AS11, BKM14, Cat98, DM17, FR82b, GHZ98, GM17a, GM21, Hac79, HR98, IY12, Laz88, Mal23, MD89, PYM93, SWY12, SS03b, Und76, Wat79, YMH94, van86]. Some [AT89, ANP83, AAFC19, AK15, BM19, Bat95, BBC18, BT97, BDFCGB12, Boo82, BMM85, BD77, Bra98, BCT10, CF91, CTU00, CM10a, CDM21a, Cér00, CMC81, Cir14, CMW06, Da 85, FCGIP06, FM08, FL17, Goz91, Gue07, GY12, Ham82, HN96, Kno77, Kno79, Kob78, Kum85, Lis14, Liu14a, LY19b, ŁS85b, Lor87, Lü15, MR15, Mat76, McE95, Men89, MR16, MR17, MR18a, MP84, MO99, Muñ00, Pan92, Pap81, PT09a, Poo96, QR11, RR94, RBA22, RSZ23, Rob81, Sac85, Sei78, Sha10, Teb11, Teb22, Tes92, Tim14, TW00, VS80, WY16, WZ17b, Yam12, Zha05, ZI00, ZzJh10, Zol78, ABSP08]. Sommerfeld [GW18]. Sontag [Tsi93]. Sound [BGR04]. Source [ASTT09, AN21, GKL20, Nic00, Ulb02]. Sources [BMT01, KKLO21]. Space [Ahm81, Ahm83, Ahm07, ALP10a, AOS16, ABK19, BB98, BMS05, BP88a, BD21, BR17b, BPV19a, BPV19b, Bor78, BCF12, BPTT21, BSZ08, BR15b, CYL96, CYL97b, CLSL20, Che76, Col82a, CN19, CD12, CC87, DGJ93, Del86, DM15, DWZ20, DMPD12, Eno94a, FRZ93, FNP^{+}22, Fuj80b, Gad08, GS94, GPB91, GLK22, GL22b, GK11, HLG93, HO01, HO22, HZ21, HU19, HR83a, Ich76, Ich07, Jac04, JL76, JL93, KN13, Kim22, Kub85, KS85c, LR79a, LM13b, Lis14, MT78, MA19, MV07a, MV08a, MV08b, MR86, MSv86, NVD18, NR97, NY16, NA02, OT83, OMS90, Ols76a, Pan92, PW23, PV13, PS88, RA16, RNK91, RL78, SWZ24, Sha15, Son86b, Tad97, TTZ23b, TKH10, Ter94, VJ77, WS95, War19, Wec82, Wei05, WK77, Zd83, CMK15, HR83b, Son86a]. Space-Valued [GS94]. Spacecraft [BP14b]. Spaces [AS20, Ahm03, Ahm19, AM82, Alv00, ARS09b, Bar97, BRZ20, BM90, BS24, BC96b, BC96c, BC96d, BKS19, BZ96, BR08, BS13b, BQT08, BS00, CCK12, CCK13, CK14, Cer01b, CFZ22, Che23, CW24a, CDM90, CLMSR00, Cou00, CD80, Da 85, DM83, DPT04, Den98, DGKL10, FP03, FHT09, GST20, Gfr06, Gib79, GRS00, GT14, HK07, HMA81, HS11, HLL94, HLGH98, HLL00, HLGHLM03, HP94, HBY21, IK11a, JNPS18, KSW18, KS18, KY87a, Lam92, Lev99, LLM15, Mah03, MPN98, Mas08, MW22, MPV18, Mos80, Nak83, NZ03, Ouv78a, Ouv78b, Pav13, PS86, PD92, Phe85, Poo96, Sag15, Sag17, Sim76, SW23a, TLD24, TZZ18, Tri76, Tri77, Tri78, Tri80, Vol00, WP98, WC19, XG03, Yu20a, Yu20b, dWvS00, vN04]. Spanning [AB09, YHJ^{+}18]. Spanning-Trees [AB09]. Sparse [BT17, CCK12, CCK13, CT14, CRT15, CK17a, GLS21, HK21, JNKV21, LS19a, MS22, PDT19, PV13, Tod83, XCZ20]. Sparsity [CC13, HSW12, HSW15, KPV14, MW22, MS17b, PW18]. Sparsity-Promoting [MW22]. Spatial [ALNP24, AGN^{+}16, CDDP11, Cer07, DWZ^{+}16, FCLZ16, GY98, KM18, LX23, Liu14a, Mao90, MMR17b, MMR17a, MS17b, SVBZ09]. Spatially [BBB24, DH09, FB08, GK85, KN77, KS86, LD05, LWDB20, MS17b, MS22, RCS01, WM10, Cou83]. Spatially-Distributed [GK85]. Spatially-Varying [KN77, KS86, Cou83]. SPDEs [FO16, GHQ15, GM13, HQZ16]. Special [BCP09, Chr80, PT09a, YXC21]. Specialized [KM07a]. Species [GRR09, Le 19]. Specifications [JK06]. Specified [CM76]. Spectra [SO06b]. Spectral [AJPR14, BTBT88, BP08b, BO16, BC02b, CL24, CF06, Cur85b, EZB11, Fei76, Fer97b, GK87b, GGLD90, GLNT23, Hag02, Hel76, Ho86, JP01, JS81, KC15, KHP06, LB00, LMP95, LR06, LMR97, Mat98, Nur06, PI99, Pet04, Reb89, RKL16, SRS15, Wak21, WCJP04, Wir98, ZS19, ZFKZ21, Zwa88b, vW84, Cur85a]. Spectrally [AMR14, HHPY16, NPY20, PYB20]. Spectrum [ABCD22, BBC^{+}19, Hag05, KMM11, OZ00, OZ03, Sun81, XS96]. Speed [And14, AB09, CGW10, DF16, DF21, DV12, Izm96, KCP16, Kas02, MPR17, OT09]. Speed-Gradient [DF16, DF21]. Sphere [CN11, ZVBR14]. Spherical [BG13, Kim90, LV97, VO22]. Spike [WY23]. Spillover [HM03, HH05]. Spin [AD08, LAM21]. Spline [GMP00, KS87a, KS89a, Rus06]. Splines [SW11, Tha86, WDS79]. Splitting [ABAC10, Bau09, BGR90, ES09, Fra82, LMP95, LT91, Nag95, PRS08, Tse91, Tse00]. Splitting-Up [Nag95]. Spreading [Kas02]. Spring [MD15]. SQP [DHV98, DKK^{+}19, IK96b, LS99a, Trö99, Vol00, VW02, Wac07, Wri89]. SQP-Like [Wri89]. SQP-Method [Wac07]. Square [BM09, Caz14, FC05, GZ12, JKSZ21, Kub85, NC16, VJBK24, ZQ18]. Squares [All88, AW90, Bar80a, Ber98a, Che84, Guo94a, LPJ20, MT22, Vog90, WLZ22, WDS79, XG18, Yu13]. Squares-Based [XG18]. Squeeze [LL00b]. Squeezing [ZLHZ12]. Squire [GW18]. Stability [ACS00a, AN15a, AN15b, AD15, ABL13, AL01, AE17, ART21, AR21, AM82, AHR20, Alt09, AAA12, AGJK12, AB02c, And82, AMP05, AISW04, AAS02, ADS24, Aug19, Aus84, BM00a, Bal77, BTY11, Bat95, BCL04, BB07, BGR07, Ber89b, BB00, BB03, BF05, BFPB13, BGLT16, BD17a, BD19a, BP20b, Bla76, BR13, BRT15, Bli04, BFP11, Bos02, BBE99, BL00a, Bru76, Buc75, Buc77, CT13, CF24, CJK13, CT14, CRT15, CCJ23, CSCV18, CCFV20, CWZZ17, CL06a, Cha05, CCF07, CFNG10a, CFNG10b, CLS21, CY24, CGC17, CK03, CW22, CM92, CLMSR00, CW03, CSA^{+}13, CW24b, CG92, CB15, CD08, CA09, DDK12, DHX08, DHX10, DM13a, DKS23, DDS15, DKT17, DEJ23, DQ21, DC88, DHKS18, DF21, DF98b, DH93, DH98b, DKK^{+}19, DB12, DJB21, DHC19, DC99, Dus21]. Stability [EP24, Ega80, EÖSÖ98, Ete19, Ete20, FM12, FF22, FB08, FHMS18, Fel03, FG15, FCJ14, FS23, Flo95, FGS93, FM08, FC05, FB12, GW84, GT77, GC06, Gib81, GS18a, GP94, Gra05, GMR22, GLO19, Guo94b, Guo02, GMY16, HH76, Hag05, HP21, HMV05, Hay19, HP94, HP96, HP09a, HM02, HJF10, HJST24, JNPS18, JMPW20, JMPW23, JSV05, JG99, JLR09, JALG18, JL18, JLL20, Kan82, KS89a, KT03a, KPCW22, KHP06, KBMS20, KSS20, KK79, KBvdSD17, KSW01, Kub85, KGM93, KW78, Kwo80, LXY14, LW17, LZC21a, LSW96, LZ94, Liu95b, LLR01, LH14, LF15, LZ16c, LZ18b, LFC21, LLS21, LH24, LW24, LRW96, MARRH23, Mao90, MTS20, MNC14, MNPR20, MMN17, MMÖ18, MADF98, MY22, Med12, Mey09, MSR01, MN07, MVZ^{+}09, MVHN09, MM80a, MKK19, MS97b, MN77a]. Stability [MN77b, Mor24b, Mos80, Mos82b, MS22, Nai95, NLPT09, NS05, NNSM08, Ngo09, Ngo11, NT17, NBF16, NY18, NNN21, NP06, OW93, OW96, Obe06, OP96, OM14, OVS08, OC23, Orl04, Pau19, PPL09, Pen84, Pep16, PA99, PS13, Pog17, PDB23, PK23, QW19, QV13, Reb95, RT98, Reb78, RR85, RZ21, RX19, RX21, RW87c, RRS14, Sac19, Sal85, SL16, SO13, SWZ24, SX14a, SH12, SL22, SO06a, SO06b, SW00, SR18, Sta79, SV79, SZS21, Tan88, TBP14, TZQ21, Tim14, TF11, Tsi93, TW03, VT95, Vin77, Wag91, Wak21, WC96a, WWS21, WW23, WC81, WRGA14, YÖ08, YD08, YW12, YW07, YX10, Yük17, ZP03a, ZP03b, ZP05, ZZS14, ZD14, ZZ20, ZH18, ZWYJ14, aJM18, dBLR14, vH06, Dai88, KT05, Mos82a, Pen88, Ngu18]. Stabilizability [AM81, AK16, BMRR20, BC05a, Ben78, BP14b, Bou95, BA84, CRR12, CDM21b, Cur85b, Dai12, DS93, EC22, Gla05, GK85, Hab96, Ho93, JZ99, KB01, KGM93, Las98, LR79a, Lev85, LCZ20, LCS96, LWXY22, LWY23, LYZ23, Lun91, MVHN09, MÖ14, NE04, OC23, PK87, Pri05, RZ09, RZ93, Rya94, SA96, SC80, TC14, Tsi91, Tsi93, Tsi95, WOH18, Yin99, ZG21, Cur85a, SC81]. Stabilizable [IP06b, VSB18]. Stabilization [AD15, AB02a, ABCL17, AK99a, AT00a, AC18b, AHR23, AB02c, AA17, ACW83, ALR99, AK19, AHK24, BCS21a, BCS21b, Bad09, BT11a, BRS11, BD11a, Bar12, BC81, BLR92, BD02, BM09, Ber02, BM03, BDK00, BGMR94, Bra09, BKR17, CK17a, CIH18b, CGY03, Che79, Che81a, Che81b, CDKP87, CLW17, CBB17b, CWQ18, CS10b, CHL15, CUB20, CK78, CL19b, CLRS00, CS03, Col12, CH21b, Con90, CM98, CO22, Cor94, Cor95, Cor99, CVKB13, CMT03, DLP86, DMK90, Del19, DRT11, DTM22, DMZ22, EGR^{+}22, Fen16, Flo97, Flo99, FNR19, FK96, FIM09, FNV10, FGH23, Fuj80a, GLM21, GMCL10, GC06, Gha11, Gho86a, Gib80, GR88b, GZSAB13, Grü96, Grü98, Grü00, GN03, GCSB99, Gue03, GDL11, Gug15, GS18b, Guo90, Guo01, GH04, GG13, HM03, HCG20, HLZF21, HLY24, HRW88, HZ95]. Stabilization [HT12, HMQD99, Hop87, HVDK19, HLDM20, Ito98, JZLW18, JM17, KK23, Kap94, KT03a, Kar06, Kar11, KK12, KK16, KK19a, KK19b, KKA23, KL21, Kob84, Kom91, Kom97, KM12b, KR15, Lag83a, Lag88, LGP15, LCC17, LT83b, Lef11, LS02, LSDS13, LYLZ99, LXZ11, LL13, LZ16b, LD19, LAM19, LAM21, LLM^{+}24, LNP00, Liu03, Liu14b, LH14, LG24, LAN04, LRT05, Log13, LSL^{+}18, MZ18, MDB05, MRS04, MPI07, Mar99, MCPA17, MS05, MS07a, MCS13, MSR01, Mik07, Mil02, MA17, MA01, MS04b, Mot00, NR15, Nik99, Obe06, OS10b, OPPU17, OZ00, OZ03, Ouz10, OOS15, PEGR18, PCP22, Ped91, Pep14, PR14, Pet85, Pet88, PT06, PT95, PM16, QLD01, QWY21a, Qua03b, RM92, Rao95, Rao98, Ray06, Ray10, RLXS23, RL20, RZ06, RK89, SKS90, Sak83, SS92]. Stabilization [Sha17, SX19, SFFC08, Sle78, Smi96, SCK10, SD13, SDSÁR00, SD15, Sul94, Sul98, Sun82, TZZ88, TP95, TF11, Tow96, TFT11, Tsi92, UP99, Urq05, VZ09, Ves13, VLP24, WZD^{+}23, Wan98, WXY05, WGK11, WW83, WHGY22, XX19, XZ20, XZGT23, YC06, YW12, YCRZ21, YLL^{+}15, YH21, ZRSZG22, ZH03, ZPQ20, ZYGpJ23, ZK88, ZLD10, ZG18, ZW18, ZY09, ZD24, ZWYJ14, Zua90, Zuy16, aJM18, dNCP20, CW06]. Stabilize [LL19, RY95]. Stabilized [CMP15, Dat88, Rus06]. Stabilizer [Liu88, WC89]. Stabilizers [IPM10]. Stabilizing [ANM22, BCL83, Bre86, BP98a, CL11, CBB19, DA17, FCG16, FL20, Her80, Her91, Hew93, Jam14, MV07b, MPS99, OCI00, PTY07, Qua04, Rif02, Tad88, ZQ18]. Stable [BMS05, BK22b, Ber18, BMO15, BSS12, Col82b, Col85, Cur03, HO97, Hel76, HK88, JZ09a, KLR09, LZ06, LWDB20, MP76a, MNC16, MNB12, Mor79, NL08, PEB19, PK87, Qua04, RR92, WSR04]. Stable-Proper [PK87]. Stackelberg [BCY15, BCS15, BCLY17, BCC^{+}19a, BCR24, FHH22, LY18, Moo21, SWW23, Zho23b]. Stage [RW87c]. Stand [Alv04]. Standard [Ha88, HZ21, HBY21, Las97, Tad93]. Standards [MR09]. Star [ACS00a, AC18b, CIP18, LM22, Mar99, PCP22]. Star- [LM22]. Star-Shaped [ACS00a, AC18b, CIP18, PCP22, Mar99]. Stasis [Joh05]. State [ART21, ALP10a, AM89, AS91, AISW04, AR99a, AR00, ABK21, AA97, AKP05, AK15, AZ97, Aub13, BM00a, BB98, BSI19, BMS05, BP88a, Bar82, Bar95a, Bat19, BC18a, BFP09, BS24, BCLY17, Ber92a, BK97b, Ber98b, BV10, BBV10, BBV13, BK24b, Bis12, BISW17, BM03, BdPV16, BFZ10, BPZ16, BC89, BC95b, BH07, BJ10, BG20, BG21, BN13, BHHE03, ByS17, Bro10, BR07, CT13, CFG21, CQSP00, Cas86, Cas93, CF95, Cas97, CRZ00, CTU00, CM02, CMR07, CT14, CT20, CK23a, CK23b, CDFP19, CLM91, Cha05, CYL96, CYL97b, CM00, CL87b, CS03, CE06, Cob06, CKR22, CS84, CT04, CD12, Cur85b, DM13a, DL13b, Del19, Del86, Del22, DI86, Dic76, DO14, DH98b, DFLZ23, DFCGBZ02, DKS^{+}11, Dun00, EP24, EDM05, Ete19, Ete20, Eva13, Fag08]. State [FGG10, FGG11, FR20, FHH97, FHH99, FRZ93, FNP^{+}22, FR99, FO18, FH20, FM22, GV03, GXZ22, GNP18, Gli12, Goz91, Grü00, Gug10, GLO19, GY16a, Haa82, Had06, HP21, HLG93, HCG20, HO01, He87, HZ21, HW05, HH10, HMQD99, Hoe12, HJF10, HN22, HJST24, HY95, HSP16, HT22, HHT19, IS12, IK96a, JNPS18, JMPW20, JMPW23, JL93, JLR09, Jay10, pJY20, KMC21, KF03, KK92, Kan00, KKA23, KP23, KZ00a, Kha16, Kob76, Kob77, KS98a, KM86, KSW01, KMR10, KR13, KLL10, Kur89, KV06, KS84b, KR87b, Kus13, Kus14, Kwo80, Lam92, Lei87, LM13b, LW15b, LZC21a, LR23a, LFM07, LLC18, LG07, LLLT18, LZSC24, Lor10, LW15d, LB98, Mal97, MRS04, MARRH23, MG20, MG21, MA19, MNC14, MNC16, Mau77, MRV11, MY09, Mic80, MVHN09]. State [MR86, MKK19, MSv86, MA17, MR04b, Mot95, NVD18, NR97, NTP18, NPR15, NY16, NA02, OMS90, OM14, OL77, PT09a, PZ24, PYM93, PZ13, PV98, RMB12, RV00, RW97a, RA16, Rap22, RNK91, RZ98, RSE20, RV19, RX21, Rod82, RT03, RT06, RT07, Ros84a, RV93, Rus76, SL16, SL13, Sch92a, Sch09, Sch80b, SU21, SWZ24, Sha15, Sha18, SX19, SY17, Son86b, Ste04, ST90, Tad97, TL03, TBP14, TKH10, TP95, Tes92, TYW11, Tib85, TSS^{+}22, TZZ18, Tsi91, Tsi93, VK81, VP15, Wag91, WS95, Wan00, Wan02, WW02, WWX13, WWS21, WYZ23, Whi82, Wol12, WY06, WHGY22, XUP08, XWZ^{+}21, YLL^{+}15, Yu95, Zei94, ZC06, ZZ20, dlRT07, vdSR11, Cur85a, Loi88, Son86a]. State- [HT22, Tes92]. State-Action [AS91]. State-Constrained [BPZ16, BC95b, Cas97, CT14, Del22, DKS^{+}11, LB98, MY09, NPR15, RZ98, WW02, XWZ^{+}21, CTU00]. State-Constraints [HN22]. State-Control [Zei94]. State-Dependent [BHHE03, Ete19, Ete20, GXZ22, Kus13, Kus14, LLC18, LLLT18, RX21, Sha18, SX19]. State-Feedback [Sch92a]. State-Space [BB98, BMS05, CYL97b, FNP^{+}22, HLG93, HO01, JL93, LM13b, NVD18, NR97, OMS90, RA16, RNK91, SWZ24, Son86b, WS95, Son86a]. State-to-Output [GNP18]. States [BB94, Bia83, DDH15, DZ24, FZZ18, Mur86, NY18, Rai99, Sch09, Sch86, SP05, Ste17, Vie00, VPJ19, Wec84]. Static [BP07b, EF09, EG02, FP21, FG15, FL20, GP17, GYBL15, Gup20, HMW12, HS13, Lei01, dlRHM16, Loi88, Wac12]. Staticization [MZ22a]. Stationarity [BC23b, HMW12, HT10, Wac16]. Stationary [AFOQ21, AW83, BAKS18, CM16, CD89, CK19, Cau13, Cer01b, CNP98, Del19, DM17, DPR22, EH10, Ete24, FR05, Gla05, GK06, GK98, Heu94, LI00, LP85, Lou94, Pav84a, Sny77, Yan96, Yük16, dlRT07, GR85a]. statique [Loi88]. Statistical [Vig97]. Steady [CMR07, CT04, DG14, Haa82, PZ13, Sch80b, TZZ18]. Steady-State [CMR07, CT04, Sch80b, TZZ18]. Steepest [Bar17, War77]. Stefan [AFCLS22, BH22, FHY88]. Steklov [AKO17, VO22]. Stemming [VS20]. Step [BBG12, BG13, Dun80a, Dun80b, FP99, Hak20, Pfl86, RS13]. Step-Length [Dun80b]. Step-Size [Pfl86]. Steplength [Phe86]. Stepping [vDHV15]. Stepsize [Bon89]. Stiefel [AM14]. Stieltjes [Ngo11, Yon93]. Stochastic [AD15, ABB02, ABVC19, AØ23, AMP00, Ahm81, Ahm96, AC13, AC17, Ahm19, ACL09, AS88, Aih92, ADRU22, ALP10a, Alv01, AB02b, Alv04, AL08, AM17, AT92, AT94, AMP05, AHS13, AH21b, Ave16, Bag84, BDM07, Bal91, BFR21, BV18, BTY11, BS86, BP88a, BEK89, BD83, BRZ20, Bar20, BK24a, Bar80c, Baş84, BL89, BK97a, BM99, BB08, BD02, BP05, BS13a, BY13, BZ15b, BCS17, BKS10, Ben96, BHS05, BD11b, BSVD84, Bv85, BCS15, BGP20a, BS17, BS19a, BW88, BD19a, Bis76, Bis77a, Bis77b, Bis78a, Bis12, Bla76, BM84, BRT15, BCL22, BK80, Boe05, BPZ16, BCM12a, BB23, Bor83, Bor93, Bor98, BM00c, Bor00b, BG05, Bor10, BK10, BK16, BEI10, BET10, BLSZ19, Bra98, BC24, BBE99, BS13b, BK02, BCR04, BM07, BQT08]. Stochastic [BL08a, BHL12, BLP14, BN16, BJ17, BLQR20, BR07, CK95, CL19a, CFN07, CY18, CFH09, CGS00, CGLM02, CLHH05, Cer01a, Cer01b, CPY09, CE98, CEK03, CKT20, Cha14, CG90, Che98, CLZ98, CZ00, CZ04b, CY12, CLYL13, CLS21, CY24, CH09, CFR13, CF14, CR93, CMR85, Chr78, Chr80, Chr83, CE11, CF13, CRST18, CDM23, Cos13, CLZ23, CC90b, CI77, DD99, DXG15, DDT00, Dav76b, DY13, DFM15, DFŚ24, DR08, DR87, DQ21, DR97, Doa21, DHJK20, DTM22, DL19, DL20, DA17, DM13b, Du15, DHC19, DM02b, DM04, DM06, DQS16, DMPD94, DMPD98, DMPD00, DHPD00, DMPD12, DK85, DK86, DK89b, Ega08, ELO22, EAK86, EMP18, Ell76, Ell77, ED81, EK09, EGN85, Ete24, FR17, FGQ97, FFS20, FFRR21, FS21, Fer18, Flå92, Fla90, FSS87]. Stochastic [FZ98, FP04, FTV98, Flo95, Flo97, Flo99, Flo02, FP99, FMSV16, FPW16, FH17a, FC05, FZZ18, Fra82, FL17, FT04, FHT06, FMT10, FO16, GCL15, GKY21, Gar76, GM91b, GGW91, Gli12, GM05, GG22, GYZ22, Gom89, GTHLHR02, GRC81, GHX86, GRS00, GŠZ05, GŠZ10, GM17a, GM17b, GL22b, GH17, Gri86, GT05, GM09, GT14, GT22b, Guo90, Guo94b, GMY16, GX19, GTX22, GW96b, HLL^{+}17, Ham06, HPW10, HSBTZ13, HJ13, Hau78a, HAB78, Hau78b, Hau94, HS95b, HHS18, HLL00, HLR01, HP96, HP98, Hin14, HBY23, Hop87, HZ03, HZ05b, HZ05a, HP09b, HJZ12, HJ16, HJZ17, HJX18, HLDM20, HW20, HJX22, HT22, HCM05, HM09, HWX09, HLZ15, HK19, Ich79, IY12, JMP21, JN06, JN24, JX22, JL16, JGC^{+}23, JK20, JH00, KP97, KM16, Kam10, KY19]. Stochastic [KY20, KB81, KBMS20, Ken80, KV21, KZ00b, KT03c, KK02, KLN21, KS15, KPZ18, Kru00a, Kru00b, KV13, KS81, Kum85, Kum08, Kum13, KK76, KH79c, KH81b, KS84b, KS85c, KY87b, KR87b, Kus90, KM91, KY92, KY93, KVA96, Kus02, Kus04, Kus07, Kus13, Kus14, KW78, LH21, LY91a, LY91b, LN82, LR00a, LCF18, LXY16, LHB16, Lep90, LC01, LXZ11, LM16, LZ18a, LY18, LC19, LWfZ22, LX23, LTT24, Lia15, LL24, LB00, LZ01, LP79b, LM97b, LM82a, LM82b, LRT96, LZ06, LK10, LLC11, LK11, Liu14a, Liu18, LY19b, LLS21, LT23, LG24, Lü13, Lü14, Lü15, LZZ21, LW22, LC23, Ma92, Ma09, MZ18, Mah03, MS92, Man04, Mao90, Mas08, MNC14, MY22, Med12, MWY21, MS98, Mey89, MC91, MT08, Mik15, MR97, MR86, MSv86, Moo21, Mor11, Mor13, MS07b, MS10]. Stochastic [Nag85, NN90, Nag95, NE04, NP22, NYZZ17, NLZK19, NSW17, NC13, Now85, Now89, NA02, NJD09, ØS10a, ØS12a, Ols76b, Ols76a, Ols76c, PB96, PB99a, Pap81, PB99b, Pav80, Pav84b, Pel00, Pen90, Pen92, PW99, PZLM24a, PZLM24b, PTZ21, Pfl86, PW17, PJ92, PG18a, PG18b, QZ13, Qiu18, RM92, RSS18, RMZ02, Red21, RX19, RX21, RS21, RCA12, Ris86, Ris94, RW87c, RW76, RW90, Rog07, RSV04, RS13, RS93, Rus21, SY21, SBY23, SZ19, Sap19, SPO22, Sca19, SC16b, SZ95, Sha10, ST23, She85, SWW22, SSK02, SB78, Sîr14b, Sîr15, SS24, SPBF10, SS89, ST02, ST09a, SZ10, STW20, Son23, SK23, SW21, SY14, SLY16, Sun21, SX23, SWW23, SCFV97, Tak97, TL94, Tan98, TLC99, Tan03, ThH07, TZ09, Tan15, Tar85, TKXP12, Tsa78, Uch78, Uch79, Uch83, UP99]. Stochastic [UP02, Var76, Vie00, VS18, VS20, Vin77, VP15, VP22a, VZ97b, WOH18, WD79, WZD^{+}23, WWX13, WZ17c, WZHZ19, WFZ20, Wan21, WWS21, Wan22, WYZ23, WY23, WF24, War76a, Wee92, WYY17, WC19, WY21, Win78, WZ01, WY08, WCS12, WHGY22, Xie12, XC03, Yak93, YM92, YK91, YT13, Yan14, YZ16, YZZ01, YXJ11, Yon02, Yon10, Yon13, Yu15b, YZ22, YZ23, YHH16, YL12, Yük16, YS17, Yük17, Yük20, Zha95, ZC06, ZES12, ZZ15, Zha17a, ZZ17, ZSX18, ZG19, ZDM20, ZC24, ZD14, ZS23, ZH24, Zho92, Zho93, ZYL97, Zho98, vH06, vW84, CL06a, Flå95, GRC82, HS95a, HU78a, KS84a, KS85a, Qua80, SZSC21, GM21]. Stochastically [LLJC18]. stochastique [HU78a, Qua80]. Stock [DYZ12, KMZ98, LS09, RH06, Zha01a, ZZ09]. Stokes [AHJM10, AFOQ21, AAR18, Bad09, BT11a, BRS11, BD11a, Bar12, BP20a, BFCLS97, BB97, BP13a, CMR07, CC16, CK19, CK21, Cau13, CSSZ20, CC18b, CRR12, CDM21b, DI94, DRT11, FGB95, FCGIP06, FLUU17, FRS16, FGH98, FGH05, GI04, GK98, GM99, GM00, HW14, HKW17, Hin05, HY97a, HY97b, HR98, HW18, KC10, KO18, Mar20, Mol19, Ner21, NR15, NMI15, Now17, Oss01, PS06, Ray06, Sla00, Sla03, Ton03, Wac07, Wan02, Wan21, YDZ98, dlRT07].

Stokes-Type [BB97]. Stone [AR99b].

Stopped [Zhu92]. Stopper

[BH13, BDP24, ELO22]. Stopping

[Alv01, AMR14, BX12, BY14, BCS18, BS20b, BCL24, BDT20, BL11, CDP23, CPY09, CY12, CS02, CL18, De 15, DFM15, DS19, DR06, Don24, Dor90, DQS16, EP08, GKL20, HZ19, HW21, Jac22, KW00, LS05a, LW10, LS09, LYZ15, MXZ19, MM21, Men80b, MR16, Mil17, MM16, Mor03, Mor10, Nut18, ØS12a, Osh06, PS10a, PW23, PL15, RH06, ST23, SY09, Ste11, Sza95, TTZ23a, TWZ21, WZ24, Zer03, KS84a, KS85a]. Stopping/Stochastic [DQS16]. Storage [KF03, PB09]. Strategic [YS17]. Strategies [AA86, Ber92b, BC11, BCL24, ÇD21, CKT24, CR17a, CLS97, FTV98, GJ20, HK13, HHOZ19, HJ21, HHPY16, HK94, KMP13, KS81, LWZ22, LZ13, MSVA09, Nag03, Now89, Piu15, Piu18, PH92, QJL22, Sîr14b, Sîr15, SQS09, TV02, WZ21, YXJ11, Ye97, Yon90, Zho23b]. Strategy [AC06, AG88, BM21, BIK99, FH17a, FL80, FL82, GBB18, GV05, IK07a, LMZ20, YXC21, YXJ11, Yu15b]. Stratified [BBV13, Mir86]. Strengthened [Lic13]. Stress [BS06]. Stretching [Coh21]. Strict [Bah08, BP04, BP06, Cel07, Cha91, CLSL20, FBJ09, GG18, HPW83, OC23, ZL01]. Strictly [HH83, KK23, Mee89, Mor24a, SD13, Tse90, Yu20b]. String [AE18, Far08, FSGHH20, Gug07, GX04, HZ95, Liu88, LZ16c, LPS12, XG03]. Strings [BB19, Dág04, Ho93, Leu99, Sch92c, XLL08]. Strokes [BBR18]. Strong [AN15a, ASZ02, AMV18, AVZ20, BS16, BC96c, BC23b, Com00, Fei80, FO18, GGC23, Hir86, LR79a, LW12, Liu95b, MY79, MVZ^{+}09, MZ22b, PS11, PS18, Ris94, Sha15, SIC00, VZ16, VL78a, Wak21, WZY09, WR14, XZ17, Yin99, ZY09]. Strongly [ANM22, Buc92, Hut09, OCI00, Ros92a, Shi88, Vuo21, Shi90]. Structural [ALR99, BFZ00, BGHR16, BR17b, Buc75, Buc77, CS12b, Dai88, GNP18, Kot94, LT83b, LBMH^{+}16, MY79, MPC91, Mur87, Nto12, Pet85, PS13, Pog17, Uch83, WC91, WXY05]. Structurally [GR21b, MRR13, Mur89]. Structure [AS03, AMZ04, AP92, AHH17, Bac79, BGRW88, BK22b, Bre16, BCT10, CYL97a, CCW14, CN10, De 83, Del19, DTWY91, FV98, FNR19, HP09a, Imm07, Kam76, KG17, Kog87, KS89b, Leq13, LM13b, MVZ^{+}09, MT09b, MV91, OW96, Qua04, Ray10, Sal85, SBY23, Sch88b, Sch88a, SD98, Sha01, Sha88b, SY17, Sus87c, Sus87d, Teb22, Teg84, VS20, WORO00, Loi88, Del19]. Structure-Preserving [Bre16]. Structured [Aub11, BGM05, BFT96, BG20, BG21, CABP17, FHMS18, HRCK21, JHMG06, MA17, MV91, Pet02, RHCK24, SSW15, ST16, vdW99]. Structures [AT92, AT94, AST05, BM18a, DvdS98, GA91, HBY21, HBY23, LZM05, MT09a, NMI15, Oht08, TK81, UMD88, WSR04, WS87]. Student [ABP14]. Study [BBR18, CGM98, CL87b, DMS03, DJ95, EP93, Hag05, HPT07, HS09, MA01, Mur87, PA99, PW84, QS13, THA02]. Sturm [Far13, KS86]. Sub [AM14, BBG12, BM18a, BG13, Gug03, LM00, Sac19, SZSC21, GGP02]. Sub- [Gug03]. Sub-Finslerian [LM00]. Sub-Riemannian [AM14, BBG12, BM18a, BG13, Sac19, GGP02]. Sub-Super-stochastic [SZSC21]. Subadditive [BS80]. Subanalytic [ŁS85b]. Subderivatives [BZ96]. Subdifferential [CHPA20, Sd10, Thi97, Tol18]. Subdifferentials [JT11, Pen13, Thi91, Tol16]. Subdifferentiation [MW22]. Subdiffusion [JZ21]. Subdiffusions [ZC24]. Subgradient [CHL82, Kiw96a, Kiw96b, Roc96, Rus21, WZHZ19]. Subgradient-Free [WZHZ19]. Subgradients [ENW95]. Subject [AC13, AC17, ACT10, BK24b, CHL15, DO14, FCG18, GGC23, HL11, LCS96, LRT98, Mal97, MG20, MG21, MCP21, NT16, RCA12, SDSÁR00, TCC15b, TCC19, UUH99, YC06, dOG05]. Subjected [UMD88]. Subjective [AY23b]. Sublanguage [WR87]. Sublevel [WR14]. Submanifold [MA17]. Submersions [Gro16]. Submodular [DF20, Top79]. Suboptimal [CGS00, CO05, Fat90, Lei01]. Suboptimality [MD13, MW02]. Suboptimization [Gai92]. Subregion [Lag83b]. Subsets [ES81a, CC99]. Subsolutions [BF22, Fis11]. Subspace [Dun91, Dun93, Fon89, Ste23]. Subspaces [DM89, Emr80a, EH80, Fra82, Hey76, LSW^{+}17, LMP95, LM97b, ÖL89, RN13, Tie14, WXZ15, Zwa88b]. Successive [BT98, JX22, LCC17, QC95]. Suction [CC18b]. Suffice [BGMR94]. Suffices [AB10b]. Sufficiency [Cha82, FP19, Man85, PS00, Sta78, Zei84, Zei92, Zei94]. Sufficient [AX02, ANS14, ABK21, BM00a, Bah08, BLR92, Bet19, BP20b, BO14, CTU00, CWW17, CLM91, DT92, Dun96, Dun98, Emr82, Eng10, GST20, GT93, Gon87b, Gra88, GX04, Her76, Hoe14, Iof79a, Iof79c, Iof79b, Kra09, KR13, LP79a, LP81, Mal97, Man81, MO02, MO04, MP76b, Mir86, Now08, Now17, RH06, RL78, RT03, Sar97, Sei88, Stu86, SW99, Sus78, Sus83, TWY90, Uch81, VL78b, WS90, Zho98]. Sum [Ave19, BS20a, BHM14, BY13, Bis77b, Bis12, BDP24, BCR04, CLS12, CCHH19, CDV16, DF20, DPR22, GR17b, Goe02, GYZ22, Ham06, HZ10a, HM14, HHS18, HLL00, HBY21, HBY23, Jac22, Jas02, JN06, KS81, Kum08, Kus07, LN82, LS05a, LPJ20, LM16, Man04, MM21, MSVA09, Now89, Ols02, PZ14, Ros00a, SPO22, Shi02, Sîr14a, Sîr14b, Sun21, SWW23, Top79, Uch78, VA03, WC19, Yon90, Yu15b, Zha05, TA96, TPA97]. Summability [BP20b, NLPT09]. Summation [EOF07]. Summed [BK03].

Sums [Bau09, ES09, TYW11]. Super [BK22a, Nym94, SZSC21]. Super-Optimization [Nym94]. Super-Quadratic [BK22a]. Superconductivity [You17]. Superconductors [LWY21]. Superconvergence [DNS17, MR04a]. Supercritical [Gug03]. Superhedging [NS12]. Superior [CIP96]. Superlinear [UU00]. Superlinearly [PT87, PTH88, PQ98]. Superoptimal [HLG93, JL93]. Superquadratic [CR11a, JL23]. Superreplication [ST00]. Supervision [Cha14, TW94b]. Supervisory [FM09, JK06, KG95, KS97b, OvS00, RW87b, Sch15b, SK95]. Supplement [Iof78]. Supply [FMP^{+}22, HFY14]. Support [FV98, FZ23, Ngu22]. Supported [Lag83b]. Supports [AW21]. Suppression [CL19b]. Supremal [RZ14, WR87, ZKW99]. Supremum [AM17, GW76, ZKW99]. Supremum-Norm [GW76]. Sure [BTY11, BC05a, Bou95, CWZZ17, DHX08, GMY16, Mao90, Pel00, ZWYJ14]. Surely [CMT03]. Surface [BB97, FLRO17, HS03, KK24, KN03, KLM04]. Surfaces [CDP23, Car97a, Car97b, GS22, SC06, Sta78, Wri93]. Surjective [BAP18]. Surrogate [PV04]. Survey [ABFG^{+}93, Aub90b, Iof78, Kum85, LP87, Wag77]. Svaiter [Bau09]. Swarming [PG18b, PG18a]. Swarms [CKRT22, KM18]. Sweeping [BV13a, BHT22, BMN19, CP16, CMN20, GMR22, HP22, JV17]. Swimmer [HT92]. Swimmers [BBR18, CM12]. Swimming [Kha07]. Switch [AH21a]. Switched [AL01, BJ11, Bos02, CL11, CLS21, DHX08, DHX10, DDG12, DEJ23, FM12, FV11, GC06, HS11, HM02, HSP16, JM17, LD06, LK08, LPJ20, LW15b, LMZ20, LMYZ23, LH14, LLLT18, LH24, LW24, MAG11, MBC06, MMÖ18, MM11, NVD18, Orl04, PU15, PRS08, RX19, RX21, SX14b, SH12, SL22, SZS21, TS18, VGBS13a, VGBS13b, VL23, WH20, YÖ08, ZZS14, CL06a]. Switches [BCW21, RX21]. Switching [AB18, AM15, BM11, BCD02, BCP16, BCP17, BLY19, BL09, BØ94, CWZZ17, CSES21, Dai12, DEJ23, DHP09, DE84, DHJK20, DH12, FF22, FS21, FR20, GGP14, GAM93, GAM97, GT08, HP21, Heu15, Hin14, HLT20, Joh05, JM17, Kha16, LGP15, LW10, LHH23, LMZ20, LMMY20, LMYZ23, LLC11, LLJC18, LLLT18, McN84, MLZ17, Mor24a, Mor24b, NY16, NY18, PVZ09, Pog17, PRS08, PK23, QWY21b, RTF21, SU18, SX14a, Sha15, Sha17, Sha18, SX19, ST23, SWW22, SY09, SYZ12, SZS21, TZ10, ThH07, TZK11, Van92, VP07, WJW19, Wir05, WH20, XZ17, Xie12, YXC21, YZSS23, YX10, Yon90, ZC96, Zha98, ZZ09, ZSM10, ZES12, ZSX18, ZY03, ZY09, Zhu21, ZWYJ14]. Switching-Control [Dai12]. Switchings [Sus79]. Symbolic [FZ05, PT09b]. symétriques [Bac79]. Symmetric [AL98, BLW07, BPS21, Bra09, BF76b, Eno94a, FW93, GKL20, GLR86, HO16, HS13, Kru00b, LZM05, LT91, MN77a, PS88, SY21, Wec82, WCJP04, Bac79]. Symmetries [CBB17a, HRW06, van87]. Symmetrization [MTS15]. Symmetry [CM04, CO22, CMOZ02, DTV18, HZ10b, KM97, LE17, Ohs13]. Symplectic [Cla82, Sak02, WM12]. Synchronizable [LR23a]. Synchronization [CLG12, DB12, Fio20, GG24, JHB19, KB16, KCY22, LXY14, LR16, LR21, LSW^{+}17, LtLR21, Ma09, MLZ17, MR09, NL08, SL13, SK95, SPJ15, SAMB14, XC18, YCY13, YCLK13, YLYC15, ZH18, ZL24, dBLR14]. Synchronous [LMA07a]. Synchrony [Kim21]. Syntheses [LY90]. Synthesis [ALY13, BD83, BC05b, CLS97, Da 87, GKB00, Gre92, Hal89, Her80, IK94, LT83a, MT23, MBL00, MSBC08, MS90, NVD18, NJD09, Orl04, Pee20, Pic96, PS00, PD02, QS13, Qua03a, Qua03b, RRR16, SDV97, SE21, Sus87b, TL09, TFT11, dBMS10]. Synthesizing [XWZ^{+}21]. System [Ahm77, All99, AG97a, Alv00, AHR23, AZ08, ARS09b, BT11a, BHM99, BCG^{+}17a, BCP17, Ben96, BM00b, BP13b, BKK02, CY18, CZ98, CM95, CMP15, CG10, CSB15, CSSZ20, CSFCL^{+}23, Cha11, Che84, CRR12, CDM21b, CE06, CLT24, CFSGS15, CGS18, CGR10, CVKB13, DS93, DDK10, DKS23, Del19, DMS03, Din15, DZA19, DG23, EMH08, Ell77, Eno94a, FZ98, Flo99, FNR19, Fra80, FRS16, Fu12, FGH98, FGH05, Gai04, GBBT78, GT07, Gho88, GB93, GS18a, Gra99, GLT95, GV05, GK06, GP17, GGMZRB20, GX04, GGA^{+}17, Ha88, HCG20, HMT18, HSD97, HZ21, HPW83, HO90a, HW12, HW14, HKW17, Hin05, Ho86, HR21, IP06b, Jaf17, KMM07, KY20, KK24, KMM11, KCBP21, KT16, KLL10, KOS04, Lar01, Las98, Le 19, LS05b, LHH23, Leq13, Lev87]. System [LR16, LR21, tLR23, LR23a, LZ94, LLR01, LYZ23, LM00, LtLR21, MH78, MSS96, MSBC08, MML21, MZ97, MM80a, Mor24a, MP21, NL08, Ner21, Pet88, PAH88, QN82, Rao95, Reb78, RV87, RL78, RRKJ24, SFFC08, Sta79, Ter94, TW03, TS18, VC22, Vuo21, WW03, WCCL04, WP98, WDS79, WS13b, Wir98, WS87, XCZ20, XLL08, Yu14, ZKW99, ZH03, ZJ06, Zua96, Zua98, Zd83, vdW99, Gho86b]. Systematic [Ber84]. systemes [Bon84, Cla86, GB82a, Lob70, Lob76]. Systemic [QJL22]. Systems [AD15, AFT91, AT96, ABL13, Aey81, Aey85, AD02, AL01, AZ07, AL12, AX96, Ahm96, AX97, AC02, Ahm03, AC13, AC17, AW21, Aih92, AC18a, AB02a, AB03, ABCO17, AK99a, AS93, AD08, ABB12, AE17, AJ12, AK99b, AK02a, Alt09, APGC16, AMV07, AMV18, AVZ20, AC05, AM81, And82, ABCD22, AHS13, ANM22, AR99a, ABR00, ABKM96, ABH98, Arm86, ADL23, ACW83, AW83, Art78, AT82, Art89b, AL02, AG19, ADS24, AF89, Aub02, ABS18, Aug19, Ava19, AD17, BCM^{+}17, BBC^{+}19, BCS21a, BCS21b, BM00a, BA08, BT11a, Bag81, Bag84, BP94, BRR02, Bal78, BJ11, BMS82, BCD02, BGM05, BMS05, BKP90, BBC81, BK84, BM86, BW89, BRR90, BFZ00, BS86, Bar85a, BBTN03, BLJ15, Bar78, Bar83, Bar95b, Bar97, BL01, BCL83]. Systems [Baş84, BK97a, BM99, BD02, Bat07, Bat19, BR87b, BSDR97, BBC18, BA15, BFP09, BK22b, BS20b, BBGBO14, BBT14, Ben81, BD11b, Bv85, BÇS07, BÇMS^{+}08, BR17a, BDIW22, Ber82a, BH86, BB00, BB03, BF05, BGLT16, BPR04, BD19a, BF90b, BGMS78, BCG84, BC00, BM03, Bla76, BM84, Bli04, BC95a, BC98, BDK00, BCS20, BDRR98, BT03a, BGST23, BP13b, BC89, Bon91, BS09a, BFP11, BW79, Boo82, Bos02, BC05b, BD17b, BD19b, BC16, Bou95, BI08, BD01b, BO09, BH17, BR93, BC02a, BGSS21, BASG21, BF76b, BBE99, Bro22, BL00a, BDDSV05, Brü13, Bul01, Bul02, BL05, BCF12, BFF20, BIP88, Bus84, BCT10, BGM97, BKM97, BGH94, BL08b, CGM14, CT13, CZ95a, CL02, CFN06, CFN07, CGW01, CM16, CPS06, CvdS23, CD89, CG00a]. Systems [CG00b, CJK13, CLZH24, CGR96, CGS00, CGLM02, Car15, Cer01a, CQ05, CRZ^{+}18, CM04, CKT20, CL11, CSES21, CZ99, CCF07, CFNG10b, CLYL13, CCC16, CyY21, CSBR21, CLS21, CWC^{+}22, CFZ22, Che23, CZZ24, CWKQ24, CDM90, CM03a, CZL21, CJT08, CS10b, CDL85, CK78, CMR85, Chr83, CM00, CS05, CH12, CMT11, CABP17, CW81, CW03, Cob06, Col11, CdDNdA20, CH78, Col82a, Col82b, CW09, CK09, CSV24, CS84, CP82, CP95, CO22, CDM23, CM76, CG92, Cor94, Cor95, CBdN08, CGW10, CB15, CN19, CMOZ02, CM03b, CvdSC05, CB05, CdV02, CdVF05, CA09, CA10, CdSBT09, Cro84, CI86, CP76, CP77, Cur84, Cur85b, CS86, Cur03, CO05, DAR15, DI87, DI90, Da 14, Da 16, DHX08, DHX10, Dai12, DXG15, DS93, DvdS98, DRW10, DM13a, DV04, Dat88, DH84, Dav79]. Systems [DCBT88, DM89, DMK90, DL13b, DJ13, De 00, DDG12, DEJ23, Den98, Dew76, DR05, DGM89, Di 91, DGKL10, DKR08, DFP14, DJZ96, DCI07, Dol79, DD03, DV83, DL19, DFCGBZ02, DMS02, DMS12, Dre77, DJB21, DHC19, Duc20, DMPD94, DMPD98, DMPD00, DF98d, DSI22, DK86, Dus21, EM19, EF09, Efi12, EPR14, EA21, Ega80, EI13, EK97, EDM05, Emr82, EG02, EOF07, EP93, EMP^{+}16, Eva13, FW92, FW93, FZ05, FM12, FB08, FM94, FMP^{+}22, FP08, FHMS18, Fei76, FR78a, FKT86, FG15, FCG16, FC19, FCC20, FCG06, FZ00, FS16, FCJ14, FPT15, Fio20, Fio21, FA89, FKN96, FK83, Flo97, Flo99, FTZ87, FT89, FGS93, FV95, FV98, FV11, FM16, FR05, FC05, Fra89, FR99, Fre00, FA13, FZ23, FP03, Fuh76, Fuj80a]. Systems [FS10, FMV24, GKB00, Gai92, GN02, Gai04, GR17a, GLZ04, GCC07, GMCL10, GKY21, GGM01, GB82b, GK94, GW16, GK87a, GJP88, GMP00, GPB91, GPS95, GBGdO00, GdOB02, GC06, GGF09, GV00, Gha11, GBR18, GD88, GAM93, Gib83, GRT92, GFF17, GCP88, GL97b, GK86, Gom20, Gon87a, GH20, GTHLHR02, GR88b, GC22, GHX86, Goo76, GTSN13, GZ14a, GZSAB13, Gra81, Gra82, Gra88, GBo97, Gra07, GH15, GLT95, GL05, GK85, Gri88, Gri93, Gro21, Gro16, Grü96, Grü98, Grü00, GN03, Grü09, GR21a, GBS22, GKLM92, GL11, GMR22, Gue07, Gue03, GH19, GO14, GLO19, GM11, GHJ00, Guo90, GZ09, GZ13, GY16a, GP82, Ha88, HK07, Hab96, HHI^{+}13, Had06, HD11, Hag02, HIK18, HP21, HP10, HMA81, HH83, HPW10, HLZF21, HS11, HSBTZ13, HZB13]. Systems [HL92, HBD21, Hay19, HYZ11, HZ21, Hel93, Her86, HL81, HLR01, HM83, HO90a, HP96, HP09a, Hir76, Hir77, Hir79, Hir81, Hir86, HD87, HD88, Hir90, HMQD99, HR83a, Ho86, HM02, HJST24, HY97b, HSJ19, HZ05a, Hu15, HSP16, HJX18, HVDK19, HLT20, HJX22, Hua88, Hua07, HV00, HNv92, HMA97, HHT19, Ich07, IS12, Igl96, IRS02, IM05a, IM05b, IRT07, ILR10, IST16, IP06a, Imm07, IPM10, IP87, IK90, Ito90, IK94, II97, IK07a, JLZ98, JZ99, Jac04, JMPW20, JKZ21, JMPW23, JL76, Jak80, JS90, JR06a, Jam14, JLR09, Jay10, JALG18, JK06, JL16, JL18, JLL20, JGC^{+}23, JH14, Joh05, JM05, Jou03, JM17, JS84, KC15, KP97, Kam76, KH79a, KH79b, KKP85, KK92, Kan98a, Kan98b, Kan00, KG22, Kan82, KS87a, KS89a, KJ07]. Systems [Kar11, KK12, KK16, KPCW22, KHP06, KBMS20, KY87a, KLR09, Kha82, KB01, Kim78, KN77, KL21, Kno76, Kno79, Kob76, Kob77, Kob78, Kob79, Kob83, Kob84, Kob88, KBvdSD17, KvdSD21, KM86, Kom97, KL99, KY05, KM97, KS93, KHJ14, KLN21, KS85b, Kre77a, Kre87, KL02, KLM^{+}21, KP94a, Kub85, KV13, Kum83, KGM93, KG95, KS97b, KM04, KK76, KH81a, KS85c, KR87b, KR88, Kus89, KY92, Kus95, KVA96, Kus14, KW78, Kwo80, Lag83a, LH21, LY91a, LY91b, LGP15, Lam87, LD05, LCC17, Las84, LR00a, LZM05, LY90, LD06, LB08, LK08, LMJ09, LM22, LPJ20, LO15, LR79a, LC01, Lev87, LM97a, LSDS13, LY91c, LR03, Li07, LXZ11, LL13, LZ16b, LL18a, LY19a, LC19, LMZ20, LMMY20, LZC21a, LWfZ22, LY23, LMYZ23, LŻ11, LAM19, LAM21, LAKG92, LFM07, Lin14]. Systems [LLC18, LR23b, LM97b, Lin07, LZ19, LNP00, Liu88, LCS96, Liu97b, Liu97a, LS99b, LZ06, LK10, LLX11, Liu13, Liu14a, Liu14b, LH14, LL15, LW15c, LFH15, LF15, LLLT18, LT18, LWZC19, LL19, LLSX20, LWDB20, LZ21, LLS21, LWY23, LH24, LW24, LAN04, LZ92, LI94b, LRW96, LT97, LRT98, LR00b, LM02, LT03, LRT05, Log13, LT13, Lom12b, Lom18, Lom21, LR06, LM08, LMPR09, Lü15, LL21b, LDZ10, ML13, MZ18, MP14, Mac14, MRL12, MGHT19, MT09a, Mai03, Mai13, MDB05, MT23, MP99, Mal23, MAG11, MARRH23, MT78, Man81, MT94, MR15, MRv94, MTS20, Mar77, MH78, MP76a, MBC06, MK21, MS05, MCS13, MB04, MNB11, MNB12, MMN17, MMÖ18, McC76, MADF98, Mee89, MSJ15, MLZ17, Mey89, MC91, Mic80, MSR01, MZ22b, MM80a, MM80b, MMK86, Mil88, MS90]. Systems [MR86, MVBC18, MRP99a, MRP99b, MI15, Mir21, MV07b, MM11, MPC91, Moo21, Mor79, MPS99, MS04b, Mor24b, Mos80, Mos82b, MS17b, MS22, MR96, MC12, MCWY14, MLYF02, Mur87, Mur89, MV91, MS85, NVD18, NE04, Nai87, Nai95, Nak83, NOV03, NTP18, NNT22, NZ99, Nar82, Nar84, NC16, NvS09, NPvS13, NMBM98, NCL13, Ngo09, NY16, NT16, NBF16, NY18, Ngu18, NR17, NRL22, NM10, Nie14, NS86, NLW91, NDHL24, Niu21, NC13, Nto12, NJD09, O’B79, OT83, OMS90, Obe91, OW96, Obe06, OM14, Ohs13, OMK84, Oht08, OVS08, OC04, Opm07, OS10b, OS19, Orl04, Ouv78a, Ouz10, OOS15, ÖT90, ÖL89, Öze85, ÖW92, PNSL20, PEGR18, PB09, PP13, PB96, PB99a, Pan83, Pan92, Pan98, PPZ13, PSJ18, PP10, PP14a, PS17b, Pau17, Pau19, Pav13]. Systems [Pav84b, Ped91, PPL09, Pee20, Pep14, Pep16, PR14, PI99, Pet85, Pet02, Pet04, PR01, PAS00, PGB14, PD92, PRG16, PSV18, PS99, PW02, PD02, PPTZ19, Pla21, PQ00, Pla80, PT09b, PEB19, PK95, Pon09, PK87, PRS08, PR07, PC00, PT06, PW78, PS87, PTZ17, PK23, PM16, QLD01, QWY21a, QWY21b, QV04, QV13, RS07a, RJME09, RGG19, RG14, RW87a, RSMG17, RBS21, RW97a, Rap22, RM98, RS10, RNK91, Reb88, Reb89, RT98, Reb77, RSE20, Red18a, Red18b, Red21, RCS01, RV09, RV19, RX19, RX21, RLXS23, RR13, RN13, RCA12, Ris94, RW97b, RW01, Roc89, RTF21, RL06, Ros84b, RW92, Ros92a, Ros94, RK89, Rud78, SKS90, Sac97, Sak83, SS99a, ST14, SL16, Sch15b, SB80, Sch02, SU18, Sch83, SWZ24]. Systems [SL98, SA96, SZ95, SK86, SL18, Sha88b, SK95, SX14b, SP05, SH12, SLQ^{+}23, SB94, SL22, SW23a, SO06a, SO06b, SK94, SRS15, Sle78, Sny77, Sok87, Sok88, SD13, SDSÁR00, SD15, SS19, ST16, Son23, Son79, Sor96, ST09b, Sta98a, Sta98b, SW12, SC80, SS03a, SGM19a, SGM19b, Sto91a, Sun81, Sun82, SZS21, Sus83, Sus87b, Sus87c, Sus87d, Sut20, Szn93, TP05, Tad88, Tad89, Tad97, TR02, Tal05, THA02, TL94, Tan00, Tan03, ThH07, TBP14, TC13, TZZ88, Teg84, Ter94, TW94b, TL09, TG12, TC14, TZQ21, Tie14, Tie20, Tod83, TF08, TF11, TK81, TSS^{+}22, Tri78, Tsi91, TT20, TWW16, TO15, TSNvL08, UP99, UP02, Val01, VS80, Van94, VB23, VGBS13a, VGBS13b, Vel97, Ver97, VZ00, VL23, Vid77, VJ77, Vin80]. Systems [VK81, Viv03, VP15, VP22a, WS20, Wak21, WZD^{+}23, WD76, WS92, WC96a, Wan98, WZ12a, WWX13, WL17, WHS19, WFZ20, WSZ^{+}21, WWS21, WW23, WR14, Wec82, Wee18, WC96b, WL19, WL21, WR00, WZ09, WTJY21, WW83, Wil77a, WZ77, WCJP04, Wir98, Wir05, WŽ07, Won09, WORO00, WCS12, WJE16, WH20, WMWL22, WJZ12, XWPG23, XB99, XS96, XZGT23, XWZ^{+}21, YC06, Yam88, Yam89, Yam12, YÖ08, YL10, YGC18, YK91, Yan95, YD08, YW12, YXC21, YZSS23, YYH^{+}16, Yin99, YZ97, YLL^{+}15, YG95, Yu95, YHJ^{+}18, Yük16, Zam98, ZBA13, ZP03b, Zha95, ZHC06, ZJ06, ZC06, ZCC07, ZLHZ12, ZZS14, ZHSP14, Zha18, ZG19, ZPQ20, ZG21, ZGS99, ZZWY10, ZYGpJ23, ZS23, ZH24, Zho84, ZK88, ZH02, ZH05, Zho07, ZLD10, ZJ89, ZY07, ZH18, ZK20, ZD24, Zol81, ZWYJ14, ZLZ18, Zuy16, Zwa88b, Zwa88a, dNCP20, dSF01, dFS15]. Systems [dBMS10, dBLR14, dRSV09, dVC22, van95, vRM98, van98, van82, vdSM13, Bon84, CL06a, Che16, Cla86, Col85, Cou83, Cur85a, CW06, Dai88, FR82a, FCGIP06, GB82a, GB84b, GK95, HR83b, JS91, Kum88, Lob70, Lob76, Mos82a, SC81, You87].

Tail [JVB^{+}23]. Tailed [VJBK24]. Takagi [BMS05]. Takenaka [XCZ20]. Takens [GS19]. Tal [HN96]. Talking [OVS08, SO06a]. Tangent [Bor78, MPV18, Urs82]. Tangential [Oss01, Pan13]. Tank [KK24]. Tanks [BRS14]. Target [BEI10, BET10, BLSZ19, Car96, DF96, DF98a, Fat90, GKL20, KCP16, Mor11, Qui92, RS10, ST02, ST09a, Whi79]. Targets [ABB15, FFP^{+}09, Sch80b]. Tasks [Sch15b]. Tau [IT87, IT86]. Taxes [TST07]. Taylor [CHN94, JHB19, Lom12b]. Team [BP12, CL06b, Gup20, SSY24, dWvS00]. Team-Against-Team [SSY24]. Team-Optimal [Gup20]. Teams [GYBL15, Gup20, HBY23, SY21, SBY23]. Technique [Coh97a, PT13, Ste92, TFL14]. Techniques [BW89, BK97b, CM17, Coh97b, Flo95, Gug10, HP87, KV06, Pap86b, POU11, SS03b]. Technology [HLG21]. Telephone [SN82]. Temperate [Gor15]. Temperature [DHK20, GXZ22]. Temple [AC05]. Temporal [JK06, PR07, Yu13]. temps [Bac79]. Tension [KK24]. Tensor [RHCK24]. Tentacle [CLL20]. Term [AHR23, Bar21, BP20a, CMR19, CW20, CT19, CGR10, DFCGBZ02, Kha03, LZ13, LY09, NP06]. Terminal [AJK14, BV18, BMS15, BK83b, CM81, FSGHH20, FHH22, FH20, Gen13, GHQ15, HQZ16, LC23, PP18, PQ00, RV19, Son23, TZ07, WM12, Yon10]. Terminal-Perturbation [FHH22]. Terms [CM06, FCLZ16, GGH16, LT86, Ulb02, WD14, Zha00, Pet04]. Ternary [Ols14]. Test [BP08a, Bre85, JZ09a, Ner21, PI99, PP00, Sha14]. Tests [BB03, HW03]. th [MNC14, ZWYJ14]. th-Moment [ZWYJ14]. Their [Ahm19, Bal86, BKS10, Bis82, BR93, BCT10, CH21a, CK89, Dat88, GN03, HHSS23, HP96, IK11b, Kno79, NLPT09, PZLM24a, PZLM24b, PW78, SSY24, SO06b, TZZ22, Yam88, SY21]. Theorem [AD02, AG88, BCM^{+}17, BET07, BL00a, CR17a, CCF07, CF14, DGSW14, DPT04, DR06, EP24, FS16, Flo02, GS93, Gar76, Ger06, GŠZ05, GŠZ10, GM17b, GM21, HS12b, Iof76, Iof79a, Jah88, JL16, KS90a, KT03a, KJ07, tLR23, LSW96, Lor08, Mam14, Pet90, Rei83, Sha88b, Sta78, Sus79, Sus87a, Tar90, Tol18, Tsi92, Tsi95, Wir05, Zag00, dFS15, Kup85, Ber98a]. Theorems [Ang76, CMC81, CDG15, CDV16, DRW10, Die87, DTWY91, Doy94, Gar03, GW16, HS11, Hew93, Iof78, Kel24, KT04, Kin84, KL99, LL00b, Liu97c, LFC21, Mey89, Mir21, MB12, Mur86, Now85, RX19, Sha10, TSNvL08, Wag77, Zha15, ZYL97]. Theoretic [BMM85, BD77, DA17, DJB21, Fio20, LAKG92, LWZC19, Mur89, RJME09, SBT19, WHS19, WS87]. Theoretical [MRVM00, SB78, Zua96, Zua98]. Théorie [LM84, Bac79]. Theory [ALP10a, AT00b, Aub90b, AVCCD22, Bal91, Bal84, BK82, BRR90, BCW21, BSS93, Bas91, BP12, Bay09a, Ber84, Bis77b, BC96b, BC96c, Bd95, Bor93, Bor10, Bou17, BT21, CvdS23, CF91, Car97a, Car97b, CH96, CDM21a, CR21, CSBR21, CWKQ24, CE06, Cur76, CP77, DG90, DR77, DHPD00, Dun91, Dun93, DK86, DRDG08, EP19, Fat94, Fer97a, FP04, FT89, FD87, Fra80, Gen13, GL95, GA91, Gil77, Gon87a, GKLW14, HM76, HZ21, HSM19, HdHVW03, IM05a, IO96, IM98, JZLW18, JS81, JM05, KH79a, KH79b, Kar89, KZ23, Kaw01, KN13, KT03b, Kre77a, Kre87, Kwo80, Lac17, LP85, MH78, MP76a, MS07a, McS89, Mis19, NvS09, NPvS13, NRB20, NC13, NS14, OBL11, OCI00, Ouv78a, Ouv78b, PT86, PP10, Pul96, Rav00, RW00a, RW00b]. Theory [Sak02, Sav05, Sch89, SD98, SC16a, SC16b, SL98, Stu16, Sub81, TT20, VS80, Vin83a, Vin83b, VW90, WK77, Wir98, XW12, ZKW99, ZP03a, Bac79, LM84, Wac07]. Therapy [LS07]. There [Lim05]. Thermal [Ste17]. Thermistor [HMRR10, HLP08, LS05b, MMR17b, MMR17a]. Thermoelastic [AL00, GRT92, Han94, LZ94]. Thermostatic [BM19]. Thickening [Ara13]. Thickness [MP07]. Thin [BF07, Del04, TW03]. Thinning [Ara12, Bré79, CN24]. Third [RZ93]. Third-Order [RZ93]. Three [BGMR94, CD13a, CM17, CG17b, DGH09, DKS^{+}11, Eno94a, FGH05, GH12, LV16, MZ15, MMR17b, MMR17a, NCL13, PBPB15, RS15b, Sch88a, WX13]. Three-Body [CD13a]. Three-Dimensional [DGH09, DKS^{+}11, FGH05, GH12, MZ15, PBPB15]. Three-Inputs [CM17]. Threshold [NSC19]. Threshold-Hitting [NSC19]. Thresholding [Wac19]. Tight [KLM20c]. Tikhonov [CMR19]. Timber [Alv04]. Time [AD15, ACKM22, AB18, AL12, AC18a, ACS00b, AS93, AD98, ABB12, AMP10, Alt09, AKN00, Ama04, AMV07, AVZ20, AM81, And82, ABCD22, AOS16, ANM22, ABKM96, ABFG^{+}93, AM22, AK04, AG19, AMI20, ALR99, Ave16, AK19, BCM^{+}17, BBC^{+}19, BCS21a, BCS21b, BMRR20, Bag81, Bag84, BRR02, BKP90, BS86, BLJ15, Bar91, BL01, Bar89, BF90a, BP88b, BL96, Bat19, BRS14, Bay09b, BBT14, BG09, BR82, BV17a, BB00, BP14b, BC21, BFPB13, BJ19, BV13b, BC00, Bla97, BHT10, BV17b, BCL24, BPV19a, BPV19b, BJ10, Bon85, Bon91, Bd95, BS09a, Bor84, BGS19, BC05b, BI08, BT13, BT21, BMO15, Bre86, BP98a, Bro22, BHT21, Bru76, BN16, BK98, BK99, BK00, BLS12, BFF20, CGM14, CD13a, Cal18, CC04, CN11, CG00a]. Time [CG00b, CJK13, Car97a, Car97b, CP00, CLS12, CLLP13, Câr93, Car12, CHJ87, Car90, CGR96, Car15, CMT18, CO03b, Cér00, CIH18b, CG90, CZ04b, CFNG10b, CLYL13, CG13, CyY21, CLS21, CFZ22, CW23b, CWKQ24, CUB20, CKL23, CL18, Chr83, CL86, CW91, CW03, CG17a, CE11, Coh21, CMW06, CN10, CP16, Cor95, CR17b, CN19, CdV02, CdVF05, CA10, CR11b, DI90, DKL09, DHX08, DHX10, DS93, DGSW14, DLP86, Dat88, De 92, DL09, DEJ23, DR87, DS99a, DS07, DBS05, Din15, DC10, Doa21, Don24, DL20, DMZ22, DG20, DF98d, EPR14, EP24, Ega80, EDM05, EMP24, FKT86, FCG16, FCC20, FPT15, Fio20, Fis11, FM95, FKN96, FV11, FB01, FR05, FC05, FNV10, FP03, Fuh76, FMV24, GRR09, GZLC17, GP22, GVL12, GW16, GPS95, GBGdO00, GdOB02, GC06]. Time [GGF09, GGW91, GAP^{+}09, GGH16, GB93, GL13, GR88a, GLK12, GH14, GK86, GH20, GTHLHR02, GRC81, GHX86, GL99, GL22b, GZSAB13, Gra88, Gri93, GN03, GPSW10, GG18, GSZ20, GMR22, Gug07, GK11, GW19, Guo90, GC05, GHS12, GHZ15, GY16a, GL19, GGA^{+}17, Hab96, Hai86, Hal89, HZ10a, HM14, Ham21, HG92, HLZF21, HWZ23, HRW88, HZB13, HAB78, HYZ11, HHOZ19, HJ21, HOPS87, Hei88, Her96, HdHVW03, HSV00, HW99, HK01, HJF10, HZ10b, HJZ12, HMR15, Hu15, HSP16, HJZ17, HZ19, HW21, HHT19, Ich07, Igl96, IM05a, IM05b, IO96, IK10, Jac04, Jaf17, Jaf20, JS90, Jam91, Jam14, JGC^{+}23, JH14, KC15, KM16, KH79a, KH79b, KKP85, KF03, KT03a, Kar06, KPCW22, KT05, Kha95, KWY08, Kim78, KG19, KY07, KL21, Kno76, Kno82, Kob77, Kob83]. Time [Kob84, KN14b, Kra09, Kra93, KS89b, KL02, KLM^{+}21, KV13, Kul93, KS81, KW13, KM04, Kus78a, Kus89, Kus90, KOS04, LS05a, LM13a, Las84, LW10, LD06, LK08, LHH23, LL13, LMMY20, LK23, LTT24, LAKG92, LFM07, LW11a, LLM^{+}24, LM97b, Lin07, LK10, LH11, Liu13, Liu14b, LH14, LLT18, LLJC18, LLLT18, LL19, LH24, LR97, LR00b, LM02, LT13, LW11b, LW15d, Lun91, LDZ10, Ma09, ML13, MS92, Mal02, MZ15, Mam14, MARRH23, Mar20, MPZ03, MG13, MSBC08, MK21, MNPR20, Mat05, MO02, MO04, MB04, MMN17, MMÖ18, Maz94, Med12, Mee89, MML21, MZ20, MV07a, MV08a, MV08b, MRV11, Men80b, MR90, MLZ17, MP76b, Mey89, MZ22b, Mil17, MY17, MVBC18, MR09, MS97a, MP11b, MM11, MNMS24, MPS99, Mor24b, MS03, NK94, NNT22, NLPT09, NS18, NP22]. Time [NT16, Ngu16, NLZK19, NDHL24, OW93, OW96, OMK84, Ols17, OC04, Opm07, OS10b, OS12b, OS14, OS19, Orl04, Osh06, Ovs14, OOS15, PS10a, PS17a, PW23, Pav84a, PPL09, Pen13, PL15, Per81, PR14, PI99, PA99, PAS00, Pha00, Phi94, Pic96, PD02, PZ11, Piu15, Piu18, PZ22, Pla21, PS13, Pog17, PEB19, Pon09, PK87, PZ13, PRHL06, PRHL08, PK23, QS13, QW18, QWY21b, RG14, RRR16, RBS21, Rap22, RNK91, Rei95, RCS01, RV19, RSZ23, RX19, Riv90, RX15, RW78, RW90, Rod82, Ros92a, Rot96, RV93, SL16, Sch88b, Sch88a, Sch80a, Sei88, SK86, SS20, Sha17, SX19, SH12, Shi02, SJ13, SB78, SFFC08, SO06a, Sîr14a, SI00, SI03, Sny77, SS19, Son79, SR18, SS03a, SCCL18, SSL11, Ste17, Ste04]. Time [Ste22, Sto92, Sus87b, Sus87c, Sus87d, STZ24, TA81, TZ07, TWZ21, THA02, TDBL22, TBP14, TBP18, Ter94, Tie14, Tie20, TF08, TF11, TG23, TV02, TSS^{+}22, TS95, TFL14, TT20, TWW16, TO15, Vel97, VK81, VW90, Vin19, VP15, Wac12, WC96a, WW03, Wan08, WGK11, WZ12b, WZ12c, WXZ15, WY16, WP18, WZHZ19, WJW19, WZ21, WSZ^{+}21, WZ24, War19, Was79, Wee92, Wee03, WYY17, WC19, WL21, WTJY21, Wim92, Wim94, Wir98, WZ98, WJE16, WHGY22, Xia11, XZGT23, YC06, YÖ08, YW12, YCY13, YXJ11, YY08, Ye00a, YCRZ21, YLL^{+}15, Yu14, ZRSZG22, ZA10, ZCGT10, ZBA13, ZYHX18, ZZ13, Zha13, Zha17b, ZYGpJ23, ZW21, ZH02, ZY03, ZH05, Zho07, ZG18, Zho23a, Zho23b, ZJ89, ZTC16, ZD24, ZLZ18, dNCP20, vDHV15, Bac79, EK97, EJ89, GRC82, JS91, LZ16b, MGHT19, NC16, WZY06]. time [Wim06, DM02a, Ver97]. Time-Consistent [CE11, GSZ20, WJW19]. Time-Delay [ANM22, ALR99, CFNG10b, EPR14, Hab96, HZB13, HHT19, KPCW22, LLLT18, MZ22b, NNT22, NDHL24, PPL09, Pla21, RG14, RX19, SS20, YW12]. Time-Delayed [Bag84]. Time-Delays [LH11]. Time-Dependent [BJ10, BHT21, CG17a, FNV10, GGH16, HK01]. Time-Discontinuous [PS10a]. Time-Discretization [Bro22, Vel97]. Time-Domain [KLM^{+}21, Rot96]. Time-Fractional [GW19, MML21, ZW21]. Time-Inconsistency [TWZ21]. Time-Inconsistent [BCL24, CL18, DL20, Ham21, HJ21, HJZ12, HJZ17, HZ19, HW21, MZ20, Mil17, NP22, NLZK19, WZ21, WZ24, WYY17]. Time-Inhomogeneous [CR11b]. Time-Invariant [CWKQ24, GLK12, KPCW22, Kra93, OS10b, QS13, RCS01, Sny77, WC96a]. Time-Measurable [CP00]. Time-Minimal [Bd95, BS09a]. Time-Optimal [AB18, Bar91, BPV19a, BPV19b, BC05b, Bre86, BP98a, Hal89, IK10, MO04, Mee89, MS97a, Pic96, Sch88b, Sch88a, Sch80a, SI03, Sus87b, Sus87c, Sus87d, EJ89]. Time-Optimality [CL86]. Time-Periodic [Alt09, BMRR20, Lun91]. Time-Randomized [PL15]. Time-Space [GL22b]. Time-Stretching [Coh21]. Time-Varying [AD15, AC18a, AM81, And82, ABCD22, AG19, AK19, BCM^{+}17, BBC^{+}19, BCS21a, BCS21b, BRR02, BL01, Bat19, BC00, BI08, BMO15, CGM14, CG00a, CG00b, CJK13, CIH18b, CLS21, CW03, Cor95, CR17b, CA10, DI90, DS93, De 92, DC10, FKT86, GB93, GHX86, Gra88, Guo90, GGA^{+}17, HG92, Igl96, IM05a, IM05b, JH14, KH79a, KH79b, KKP85, Kar06, Kha95, KM04, LAKG92, Liu13, Liu14b, LLJC18, LR00b, MARRH23, MK21, MB04, MMN17, MMÖ18, MPS99, NK94, PI99, PA99, PAS00, PD02, Pon09, PK87, Rap22, RNK91, SS19, SS03a, SCCL18, THA02, TDBL22, TBP14, TBP18, Ter94, TSS^{+}22, WZHZ19, WSZ^{+}21, Wir98, WJE16, YCY13, ZZ13, Zha13, ZJ89]. Times [Alp10b, BP07a, BFZ10, CR93, CL18, GKP18, KN14b, LB08, LYZ15, LMYZ23, LS19b, SY09, TYW11]. Timing [HLG21]. Timoshenko [AZ08, KR87a, LLS93, Leu99, LZ18b, MM04, VS14]. Tip [CM98, Guo01, KGBS08]. Tip-Force [KGBS08]. Toeplitz [BC02b, HBS06, NHG99, ÖT90]. Tolerant [BL15, Ste23]. Tomography [ADK08, Kat19]. Tools [HD11, TP95]. Topological [AHJM10, AGJK12, CNS10, CT19, DHX08, Del22, Del23, GS22, GGM01, GI02, GI04, HJM07, KBvdSD17, KL16, KP94a, MH78, PS04, SAM03, SZ99, Toa11, Tow96]. Topologies [BF05, DHJK20, LLC11, LLJC18, MLZ17, YXC21, Yük24]. Topology [AS21, ACT10, Ams13, BR17b, BS06, Coh21, FIR06, GMN22, Gün95, Gup20, LG09, MM80a, SZ03, VC22, ZZ13]. Toroidal [RFF15]. Torque [CCW14]. Torus [DH98a, GR21b, SY24a, VLP24]. Total [Cha11, DG20, JN19b, MS23, PZ22, RCA12, Sto76, TCC15b, TCC19, Yu15a]. Trace [BRW05, FV95, Kim93b]. Tracing [BDGK21]. Track [IP06b]. Trackers [KP87]. Tracking [AMP00, ABHY23, AFO22, AG97b, Ber95, Ber98a, BP08a, BDLW24, BM00b, BLY21, BJ02, CDDP11, CC16, CG90, CdDNdA20, DJ95, DKRV13, Eno94b, EH10, GLK22, GM99, GM00, Guo94b, HL92, HD87, HD88, Hir02, HY97a, IRS02, ITT04, IRT07, IM08, ILR10, IK96b, KLR09, Liu23, MZ15, MW21, MT09b, ÖW92, PSJ18, POU11, PP00, SBBV15, SZSC21, YL10, YXC21, YZSS23, ZZ92, ZY24, ZZ94]. Traction [GY98]. Trade [GH17]. Trades [GN18]. Trading [CKL23, Kar89, Zha01a]. Traffic [AK99b, AK02a, CJS08, FGH23, KR89, Kus95, KM96, MK90, MSS96, MMN09, MPR17]. Trajectories [AS03, Ahm83, AG97b, BBV10, BBV11, BCM13, CF91, CJT08, CKR22, DKRV13, FL81, ISdCN99, KBDK03, LS99b, PV15, Pap89, Pog17, RA16, Roc89, Sch88b, Sch88a, Sus87c, Sus87d, VP88, XA93, ZCGT10, Zer05]. Trajectory [ASZ02, BS90, BS93, CdDNdA20, MMN17, MA03, MP21, PS11, RW01, SK95, Tad88, ZL01]. Trajectory-Based [MMN17, MA03]. Transaction [AMS96, ARS17, BP10, BZ15b, BMS15, BC12, BS17, BS19a, Cal09, CSZ13, DJLY09, DPZ93, Kam03, OS02, ST13, Zar92]. Transcendental [Gho86a]. Transfer [BGKV80, CPW88, GBR18, GP17, Hag05, HRW06, KK23, KKP85, MH78, Mik07, OMS90, TKH10]. Transfer-Function [KKP85]. Transfers [Izm96]. Transform [Dat80, HdHVW03, Obe07]. Transformation [Far13, GLM21, GEF17, Nih83, OMS90, WM12]. Transformations [ABH98, BA08, Cla82, SBY23]. Transforming [BAP18]. Transience [Ich11, Kre78]. Transient [ÇR14, DB12, GH14, IRT07, ILR10, LXY14]. Transients [KNP91]. Transit [KN14b]. Transition [CCHP12, FF22, FKZ22, JN06, Küe07, LB08]. Transitions [AHS13, Ich16, SS19, Ste17]. Transitivity [AS93, BW79]. Translation [KV07b]. Transmission [CFZ10, LL00a]. Transmutation [Mil06]. Transport [ABVC19, AVJ21, BP20a, BKL^{+}22, BCS18, BS03, CLW17, Fab08, FA13, GKP18, GKLW14, HRCK21, KY07, Lü14, TLD24]. Transportation [BJO09, CJS08, CSA^{+}13, LR23b, MT08, Mik15, RR94]. Transversality [AK04]. Transverse [DN23, MS02, MS04b, NM08, Nie14]. Traps [Bra98, SSL11]. Travel [Alp10b]. Traveling [ALNP24]. Treatment [EOF07, IK11b, LS07, MRVM00, PT08]. Tree [BB90, Dág04, GL11, HRCK21, LM22, PR14, SC12, YHJ^{+}18]. Tree-Like [GL11, LM22]. Tree-Shaped [PR14, Dág04]. Tree-Structured [HRCK21]. Trees [Alp10b, AB09, CIP18]. Trefftz [MR79]. Trend [ABP14, FFR23]. Trend-Following [ABP14]. Tresca [CDD20]. Triangular [HBS06, II97, JZ22, Pon09]. Triangulated [CBB17b]. Tribute [BSZ93, BF89a]. Trigger [YXJ11]. Triggered [KEK21, KCY22, LLSX20, RC16, WS20, WMWL22, YYH^{+}16]. Triggering [FM16]. Trotter [Pop95]. Trotter-Type [Pop95]. Truncation [Bar80a, GO14, Red18a, RV09]. Trunk [Kus95]. Trust [DHV98, Kiw89, UUH99]. Trust-Region [DHV98, UUH99]. Tucker [CFR13, DJ86, DT92, Gün95, IK11a, LP79a, LP81]. Tumor [GLS21]. Tuning [KP87, LW87, RT16, RSMG17]. Turbulence [DD99]. Turing [CLZH24]. Turnpike [BP20c, CH96, DGSW14, DR06, GG18, GH19, HZ22, HZ21, Mam14, SY24b, TZZ18, Zas97]. Twisting [Gro12]. Two [AZ07, AB93, AB03, AD96, AMP10, AGN^{+}16, AAFC19, AFCLS22, BSI19, Bas91, BHM14, BRS14, BP07a, BB19, Bil19, BF90b, BV13b, BS09a, Car96, CLS12, CD15, CDCVV21, CK19, CK21, CZ98, CDFP19, CC18b, Cor99, CG17b, DMK90, DGH09, DRT11, Doa21, DZA19, DM17, Eno94a, FFR23, FV95, FNR19, FK76, FHY88, FRS16, FGH98, GB84a, GS93, Gla05, GT22b, Gue07, GR21b, HHI^{+}13, HPS07, HYZ11, Imm07, Ish88, Jac22, JGC^{+}23, Joh05, KP97, Kin84, KM12b, KW87a, LN82, Le 19, LS05b, Mar77, MCS13, MD15, McN84, Mee89, Mik15, MP11b, Mor24b, NZ99, NCL13, Nto12, PEGR18, PZ14, RSS18, RRR16, RA16, Ray06, RW87c, SO06a, SS06, SS89, Sta15, Sun21, Sza95, Tol16, VP07, VLP24, Vrd16, Wan00, WZY06, XWPG23]. Two [YZ23, ZG17, Zha05, Zha15, dOP03]. Two- [DGH09]. Two-Channel [JGC^{+}23]. Two-Dimensional [AZ07, AAFC19, Bas91, BF90b, BV13b, CK19, CK21, CC18b, Cor99, DMK90, DRT11, FV95, FNR19, FGH98, Gla05, GR21b, HPS07, LS05b, MCS13, NZ99, PEGR18, RA16, Ray06, SS89, VLP24, XWPG23, dOP03, Nto12]. Two-Input [Mee89]. Two-Level [AB03, BS09a, DZA19, Ish88]. Two-Metric [GB84a]. Two-Person [CLS12, Kin84, LN82, SS06, Sun21, Zha05]. Two-Phase [AMP10, AFCLS22, CDCVV21, FHY88, Vrd16, ZG17]. Two-Photon [DZA19]. Two-Player [Jac22]. Two-Point [AB93, DM17, FK76, MD15]. Two-Regime [VP07]. Two-Scale [GT22b, KP97]. Two-Sided [FFR23, KM12b, RSS18]. Two-Stage [RW87c]. Two-State [BSI19]. Two-Time-Delay [SO06a]. Two-Time-Scale [Doa21, HYZ11, MP11b, WZY06]. Two-Valued [Joh05]. Twofold [GMP00]. Tychonov [Pue09]. Type [Ahm96, ACS00b, AHR23, AB05, AD18, BB97, Ben81, BKR17, CGM98, CT96, Cha09, CZZ24, Chr19, CW22, CW23c, CK89, DDK12, DI93, DHX10, DPT04, Fel03, Fen16, GM93, Ich11, Iof79c, IK96b, JG99, KN77, KL99, KB99, KK76, KH79c, LY90, LHX12, LR16, Liu11, LFC21, LT13, LW11b, MRV03, Mey77a, MH11, MS96, NBF16, Pop95, RS07a, RBS21, SC16b, Sha10, SO06a, SRS15, ST96, SV17, SKXL17, TY23, VL78b, WT83, WZC20, Zha13, ZSX18, ZD14, Cou83, DG16, Ito90, XB99, Red18a, Red18b, You17]. Type-II [You17]. Types [Imm07, ML13].

Uhlenbeck [AM23, Ber21, FP03]. UIO [Ste23]. Unbalanced [CGP22]. Unbounded [BY13, BMS15, CHJ87, CDR08, CSC14, CSCV18, CP77, Cur84, CS86, DD92, Die87, DF94, FCJ14, FNV10, FP03, Gon87b, GL19, HLL94, Kou24, KT16, Las95, LR94, MPR22, NA02, PPZ13, PP14a, PZ11, PW78, PS87, Ram98, RV90, RZ98, Reb89, Tol18, Tri76, Urq05, WZ94, Wei89, WJZ12, Yu20b]. Uncertain [AX97, ABB12, AA17, AKP21, ABR00, BBTN03, BCL83, Bli04, CDL85, CdSBT09, EMO23, FSGHH20, GPB91, GPS95, GR88b, GS18b, GP82, HP09a, KHP06, LL13, LC19, LS19a, LŻ11, Liu13, Liu14b, LL19, LT97, MVBC18, Orl04, PSJ18, Pet85, Pet88, Pet02, PRG16, PD02, POU11, QV04, RCA12, RK89, SL22, SFFC08, UP99, UP02, VP15, WW83, XUP08, YL10, ZP03b, ZHC06, ZGS99, ZYGpJ23, ZK88, ZD24]. Uncertainties [DXG15, FK21, LL18a, NNT22, TZQ21, VPJ19, WSZ^{+}21]. Uncertainty [AC19, AY23b, AF89, AHK24, BBD18, BKL^{+}22, Bat07, BC23a, BCC^{+}19b, BD01b, BØ94, BL00a, DK13, DC10, Fra03, HP96, HW20, HWY21, JKSZ21, KV02, NS12, Pet02, RCA12, RK89, UP99, YG95, dOG05]. Unconditional [DZX24]. Unconstrained [CDP16, Grü09, GPSW10, KS87b, Man95, TG23, WRGA14, YF97]. Uncontrollability [LHS06, Mai03]. Uncontrollable [GL92, Kan00, PB09, QLD01, Sha14]. Uncontrolled [Viv03]. Uncountable [HLGH98, NA02, WC19]. Undamped [GH04, KTV16]. Underactuated [CM03b, LAN04, RM98]. Underactuation [Gha11, NCL13]. Underactuations [SGM19a, SGM19b]. Underlying [AT82]. Undetectable [CWC^{+}22]. Undiscounted [ÇR14, PZ22, Pla80]. Unfolding [NPS15, NS23]. Unicycle [dNCP20]. Unification [ASP22]. Unified [BT21, FC05, Kas09, LN82, NVD18, Ngu18, SWY12, TG23, Tse90]. Uniform [AM23, AB10a, BP20a, BD01b, CSC14, CSB15, DL09, DDS15, DKT17, DHV00, FY76, GLX18, GG07, HRPT21, Ito98, JV06, KPCW22, KK06, Las98, Léa12, LR21, LZ94, LG20, Mic09, Mor98, MN77b, OOS15, PA99, Rao95, VZ16, WS13a, XZ20, ZD24, Zua90]. Uniformly [FP19, IP06a]. Unifying [AAS02]. Unilateral [SC23, War76b, War82]. Unions [BHPM04]. Unique [AD96, CZ04a, DH84, Mor24a, Pie79, Pro22]. Uniquely [MP21]. Uniqueness [AKN00, BMS15, BB07, CDFP19, DL06, Day09, Gar76, HW03, HJZ17, HJX19, Hua07, JZ22, Jak80, tLR23, McE95, SZ82, WHB92, Zua96, Zua98, GB84b]. Unit [Bon89, Bor84, Kus78a, Kus89, Öze85, PS17a, Ste22, Yan96, ZVBR14]. Unitary [Lou94]. Universal [CLS97, Tow96, Yük20]. Universally [DH98a, Yu20a]. Unknown [BC96a, BG98, BGP09, BFP09, BBT14, CRZ^{+}18, CLRJ13, DKL09, Dre77, DMZ22, ELO22, FFR23, GP21, GZ23, HZB13, HH17, KR04, KKLO21, Kob79, Mur92, NT16, Pap86b, RT16, YL10]. Unmatched [GZ19, WTJY21]. Unmixed [Wim92, Wim06]. Unmodeled [NK94]. Unobservable [BFP09]. Unreliable [FGGH15]. Unsafe [RN13]. Unsolvable [IY12]. Unstability [LPJ20]. Unstable [ADS24, Dav79, GTSN13, GKLM92, Liu03, TSNvL08, ZG18, ZW18]. Unsteady [AB09, FLUU17]. Unstructured [DM83]. Unweighted [ASP22]. Update [CR93]. Upper [BW88, DM09, JHMG06, Kre78, NR21, DK89b]. Use [Che84, MP11b, NP96]. Used [CDL85]. Using [AD08, ASTT09, BC18a, BH99, BDK00, BK21, Bon86, BB04, BDDSV05, CRZ^{+}18, CBB19, CHL15, CR93, CVKB13, DHX10, DH09, DZ19, Dus21, FB98, Fer97a, FFP^{+}09, FLUU17, FGS93, Fre99, GAC24, GM05, GO04, GP17, GM11, GZ23, HN00, JNKV21, KLR09, KA06, Kor20, KNW95, LF87, LMJ09, LPJ20, Lou07, MAJ18, Mey89, MRP99a, MRP99b, OMFMC^{+}21, Phe86, PRG16, PP00, PRS08, RM92, RY95, RSS22, SS99a, Sd10, SS18, Tse93, WTJY21, ZS19, Hug16, MMN17, Pan17]. Utilities [LT23]. Utility [BC12, CS87, EK09, GSZ20, HL16, Heu15, KLSX91, KW00, KK18, Kre78, LL20, NS18, PL15, ST14, SZ04, SYZ12, TZ10, WYZ17, Zha17b]. Utility-Based [KK18]. Utility-Deviation-Risk [WYZ17].

Vacations [AK02a]. Vakonomic [CdLdDM02]. Validity [BCM12b, CP98, Cel02b, GBBT78, YC24]. Valuation [KMZ98, LS09, ZZ09]. Value [AFT91, AB93, Alv04, AM17, Amm02, AB12, ABK14, AB17b, Aub02, BFR21, Bar89, BB18, BFT96, Ber85, Ber98c, Bro87, BP98b, CFS15, CLS12, CC96, CL86, DZ76, Dav07, DM17, El 17, Ell76, FBJM14, FK76, Fuj86, FS10, FGH98, Gal00, GR17b, Hag02, HBY23, JHMG06, JV06, KHP06, Koe79, KS81, KW87b, KW87a, Lag77, Lag78, LS05a, LY01, LY02, MW09, MD13, MD15, MY17, Mun07, NS20, NLW91, PQ00, QR11, SL22, SS89, Sti18, Thi91, VZ16, WL95, Whi79, WT83, Wu18, Yu15a, Zar84, Zha17a, Ber88]. Value-at-Risk [MY17, NS20]. Value-functions [PQ00]. Valued [ABL13, Ahm81, BD77, BS09b, BCS13, BMN19, BK08, CK19, CK21, CTW18, CG17a, CFZ10, EL20, FSH16, FBJ09, Fra87a, GS94, GT14, Gui13, GJN08, Hen83, IK08, Joh05, JS81, Kas09, Kas13, KL99, KPV14, KTV16, LY12, MP76a, MVBC18, MS97b, Nur06, Pan13, Qiu13, Sti18, CDV16, Iof83, XW12]. Values [BLQR20, BR15b, KHP06, LE17, LW15d, Sîr14a, TLPA00]. Vanishing [BP88b, BBCS22, BK22a, CQ15, CCHH19, CGP22, CGW10, GT22b, Ich07, Léa12, LK23, Mal02, Mic09, PPZ13, PZ03]. Variable [AW21, Aub13, BB19, BV10, BQ99, BKM97, Dic76, FT14, FL22, Guo02, Han77, HCG20, LZ16b, LBMH^{+}16, LYZ23, MP07, Mur92, RSE20, SFFC08, Spi80, Yao99, ZW21]. Variable-Order [ZW21]. Variables [Che76, Del86, Lom13, Mau77, SD98, VJ77]. Variance [Guo94a, GHZ15, HLVAC99, HLCH18, Kwo91, LZL02, Lim05, MT03, MTT08, OW00, PLK89, SK86, TFL14, WJW19, Wee03, XZ07, XZZ20, ZY03].

Variance-Minimization [HLVAC99]. Variance-Optimal [MT03]. Variants [DT92]. Variation [Boe05, Fer78, HW19, MS23, PP14b, Pic93, RCA12, TCC15b, TCC19, Wee03, Wee05]. Variational [Ach05, Ach08, ATW93, Ama04, AM95, AB17a, ABBK20, AB20, Aus99, ABL18, Bal01, Bar81, BCP17, BZ10, BB09, Ber98b, BZ99a, BS93, BCM12b, BC13, BC95b, BFL16, BC23b, BS00, BLS05b, BO01, BL03, CHY23, CM10b, CY99, Che99, CCT05, CCT07, Chr19, CW23c, Chr77, CM96b, De 11, DHV00, DKRV13, FSH16, FP96, FBP99, FBJM14, Fra84, FR78b, Fri86, Fri87, GW96a, GGP02, GO04, GP94, GJN08, HFY14, He87, Hel93, HTY01, HL11, Hoe12, Hoe14, IZ00, Jam91, KS18, KT00, KZ00a, KL16, Lar01, LWY21, LZ96, Lev99, LY12, LP86, Liu95a, Liu95b, Lou02, LY01, LY02, Mar02, MD89, MP84, MN03, MM11, MW02, Mor03, Muñ00, MN08, Ngu16, Pen20, Pop97, Qiu13, RZ20, RH06, RRS14, Roc96, Sag15]. Variational [Sag17, SWY12, Sar22, Shi88, ST96, SS99c, ST09b, TL03, TBP18, Tim14, Tse91, VZ97a, Vuo21, YF97, YT13, Yon10, Zas97, Zei01, ZJ06, Zha09, Zol78, vTY99, vDHV15, AHR20, BLS05a, CW22, Cos13, El 17, Shi90]. Variations [BFCLS97, BK03, Bou13, CZ94, CZ95b, DM09, Eke77, Fer05, FLUU17, Hen80, HL04, IZ00, McS89, Mur86, OC12, PM16, Sla00, Sla03, WY23]. Various [Koe90, KK06]. Varying [AD15, AC18a, AM81, And82, ABCD22, AG19, AK19, BCM^{+}17, BBC^{+}19, BCS21a, BCS21b, BRR02, BL01, Bat19, BC00, BJ02, BD19b, BI08, BMO15, CGM14, CG00a, CG00b, CJK13, CIH18b, CLS21, CW03, Cor95, CR17b, CA10, DI90, DS93, De 92, DC10, FKT86, FH17a, GB93, GK86, GHX86, Gra88, GPSW10, Guo90, GGA^{+}17, HG92, HSV00, Igl96, IM05a, IM05b, JH14, KH79a, KH79b, KKP85, Kar06, Kha95, KN77, KS86, KM04, LB08, LAKG92, Liu13, Liu14b, LLJC18, LR00b, MARRH23, MK21, MB04, MMN17, MMÖ18, MPS99, NK94, PI99, PA99, PAS00, PD02, Pon09, PK87, Rap22, RNK91, SS19, SS03a, SCCL18, THA02, TDBL22, TBP14, TBP18, Ter94, TSS^{+}22, WZHZ19, WSZ^{+}21, Wir98, Wir05, WJE16, YCY13, ZZ13, Zha13, ZJ89, Cou83, Gug07, Ver97]. vecteurs [Bac79]. Vector [AB02c, BZ10, BMT01, CTW18, CD80, DG90, Den97, DMR19, EMH08, FSH16, Gfr06, GJN08, Hen83, Her78, HL04, JS86, Khá93, KL99, LY12, Qiu13, RA16, RLXS23, Tan88, WCS12]. Vector-Exponential [RA16]. Vector-Valued [CTW18, FSH16, Hen83, KL99]. Vectorial [CZ95b, Kle18]. Vectors [ES81a, LP79a, LP81, Bac79]. Vehicle [ABB15, SPBF10, ZM14]. Velocity [CC16, CGS18, DZ91, Eno94b, GK06, GKLW14, GHS92, GM99, GM00, HY97a, RS15b, SS21, VM00, Vin14, YDZ98, ZY08, dG06]. Velocity-Current [GK06]. Verifiable [AMP05]. Verification [BCS17, DVPL24, FGG11, FM09, Gar03, GŠZ05, GŠZ10, GM17b, GM21, HSA17, LLS21, PR07, RM92, RL06, RW08, ZYL97, HS12b]. Version [CZ94]. Versions [Roc96, Tsi93]. Versus [BDP24, CLMSR00, Kha01, Ram98, CO03a, CNS21, GX19, PZ13]. Via [BP97, Bat95, CM17, Dok05, FR17, Gra99, Her96, Ito90, Kry98, LY90, MPS99, Sor98, Yan14, Zei92, AS13, AFV17, All88, AW90, AAA12, AY23a, AF18, AN06, ANR08, Ari98, AL00, BA08, BK24a, BS20b, Ber89b, BCG86, BCG^{+}17b, BPZ16, BDGR12, CZ95a, CSBC14, CL06a, Cha14, CWQ18, CR21, CF14, CB05, Cur85a, Cur85b, DJK^{+}23, Dav76b, DZ88, DI86, DM17, DK85, Eva13, FNA02, FMT10, FT02, GMCL10, GY98, GN03, GZ14b, HCG20, Her80, HLL96, HW18, JM17, Kan00, KZ23, KN13, Khá93, KP04, KN03, KM96, LHPT08, Lev85, LZ18a, LMYZ23, LY20a, LLC18, LNP00, Liu95a, Liu13, Liu14b, LH14, Liu23, LAN04, MPI07, MR15, MS05, MS07a, MPR17, MW02, NPS15, NS23, Ner21, NLPT09, NR17, NRL22, Nih83, POU11, PZ22, Pop95]. via [SO13, Sha94, SK95, SW11, SI03, SCCL18, Sto04, TDNT18, TLD24, TF08, Wan91, WS13a, XB99, XW12, YD08, YCY13, YZZ01, YCLK13, YHJ^{+}18, YH21, ZC96, ZLHZ12, ZD14, ZK88]. Viability [Aub90a, Aub90b, Aub01, Aub11, Kel24, Lor08, Tal91]. Viable [Lei85]. Vibrating [BB19, CMT11, FSGHH20, HZ95, Liu88, Nar82, Nar84, Sch92c, WCCL04]. Vibrational [Bul02]. Vibrations [Dág04, Pen89]. Vidyasagar [Tsi92]. View [DHX08, KRW04, SS23]. Viewpoint [AF21, Bur91b, Qua03a, Qua03b]. VII [CYL97a]. VIII [CYL97b]. Viot [EK93]. Viruses [Gor15]. Viscoelastic [BDFCGB12, CO03a, GLM19, LPS12, Pan17]. Viscoelasticity [HRW88, ZH03]. Viscosity [AKN00, AB02b, BP88b, Bar85c, BCP16, BK22a, BMS15, BB06, Bis12, BISW17, BZ96, BT11b, BL08a, BIRS20, CR11a, Day09, DR05, Fab08, FGG10, Gor19, GŠZ05, GŠZ10, HJX19, HCM05, Hua07, KS98a, Léa12, LW14, LS85a, Lor87, Mai08, Mal02, McE95, Mic09, NR21, NSW17, Pla21, Qiu18, RTZ20, ST02, SY24a, Sor93, Sor96, Świ20, TTZ23b, ZYL97, LS86]. Viscous [BH00, Cha09, CI19, CFSGS15, CGS18, GZ21, GHS92, KK24, Kou24, LYLZ99, PS19, WC91, ZM11]. Visibility [GCB06, Kim22]. Vision [GL95]. Vlasov [AØ23, Ahm07, Bar23a, BKSW24, BS20b, BIRS20, CDJS23, DQ21, Lac17, LWfZ22, MZ20, PRS21, PW17, SC16b, SY24a, WHGY22]. Voids [CNS10]. Voigt [Bur20, HLY24, LL98, LZ16c, LZ18b, YH21]. Volatilities [ASR^{+}13, FPW16]. Volatility [Ach05, BS17, BS19a, HJ16, HWY21, NS12, SZ04, TZ07]. Voltage [MÖ14]. Voltage-Actuated [MÖ14]. Volterra [BCM12a, Che16, CyY21, Cro81, CC87, Gri83, JMP21, Jun00, Kup85, LY20b, NNSM08, Ngo09, Ngo11, Nih83, Pav13, PY96, WZ17c, Wan22, WYZ23, WY23, Yon93]. Volume [AAC86, Amm02, GKY21]. Vortex [ASK05, DG23, KV07b]. Vries [AC18b, CFZ17, Cer07, CRZ13, Gag16, KCBP21, MRRR19, MCPA17, ÖB19, Ros00c, RZ06, Sun96, Xia19, Zha99]. VUCA [HSM19].

w.p [DK89b]. Walk [BÇS07, BHHE03]. Walks [BR10]. Walls [BLPP24]. Wardrop [CJS08]. Warga [Ano95]. Wasserstein [Bar17, BD21, KY23, TTZ23b]. Wastewater [MRVM00]. Water [Bel00, FLRO17, RRR16, RR85, Rei86, Rei95]. watermark [JS12]. Wave [ACS00a, AN15a, AN15b, ABCL17, Amm02, Bal91, BP97, BC81, BE13, BFM23, Buc92, BFO20, Bur20, CC01, CCM14, CO03a, CSC14, CSCV18, CCFV20, CT19, Che79, Che81a, Che81b, CW08, Che76, CDZ94, CLM24, Dág06, Dág18, DLP86, DM17, Fab92, Far13, Fen16, FCLZ16, FM08, Fre99, FL22, GT17, GLX18, GLM19, GLS05, Gug10, Gug15, GZ05, GG13, GR21b, HH19, HLY24, Ho90b, JL14, KMC21, KB01, Kim93b, KKLO21, Kle18, Kom91, KKV11, KW13, KTV16, Lag83b, Lag88, LL00a, LS91b, LL21a, LPT22, LHX12, LR16, LCZ20, tLR23, LR23a, LL24, LG20, LYZ23, LtLR21, MRR13, Mar99, Mic09, Mil02, Mot00, MT22, NP06, Niu21, Now08, POU11, SMDB97, SCK10, TL06, Teb11, VZ09, VM00, WC89, WGK11, XX19, XZ20, Yao99, YH21, Zha00, ZY08, Zha18, ZC91]. Wave [ZL07, ZW18, Zua90]. Waveform [AY23a]. Waveguides [DWZ^{+}16]. Wavelet [DK05, GK11]. Wavelets [WP98]. Waves [ALNP24, BLR92, Bel00, CRS21, DKL09, DL09, DE14, FLRO17, GMN22, MNORP21, Oss01, RR85, Rei86, Rei95]. Wazewski [Tol18]. Weak [Ahu16, Bal82, BY13, Ben78, BC96b, Bor78, BT11b, BN13, BF93, Coh21, CDM23, DO14, DQS16, JZLW18, KMC21, KT04, KM96, LM16, NZ03, Ouz10, PZ14, Rai99, SYZ13, SW99, Sva11, VL78a, WWZ14, Xie12, YD03, ZY09]. Weak-M1 [Coh21]. Weaker [LXZ11]. Weakest [Vin83b]. Weakly [AB02a, AB03, BC20, BSS12, Fu12, HCM05, Küe07, MYAS09, MS10, Qua03a, Ros92a, Ros00b]. Wealth [BMS15, PP18, TZ07]. Weapons [Win78]. Wei [Coh97a]. Weierstrass [MGMBN^{+}19]. Weight [CLZ98, CZ00, MNC16]. Weight-Balanced [MNC16]. Weighted [ASP22, Ber95, DR97, FT89, HK07, Tad97, TYW11, Ver97, Wyn09, ZZS16]. Weighting [NLW91]. Weights [MG13, MWY21, Sim76]. Weiss [Wyn09]. Well [ABL13, Aug19, BM09, BH07, BEL02, CHY23, CM03a, CSZ13, Cur03, CO05, EB20, GS18a, GK86, GT14, GZ09, IZ00, Kob77, LT18, LT03, LRT05, Log13, Lü15, Ols76c, OS19, PCP22, Smi90, Sta98a, Sta98b, TBP18, TW03, VS80, Vil04, WZ09, WCG14, Zas00, ZZ23, Zol81, CW06, WW02]. Well-Defined [Ols76c]. Well-Posed [Cur03, CO05, GS18a, GK86, Kob77, LT03, LRT05, Log13, Lü15, OS19, Sta98a, Sta98b, TW03, ZZ23, Zol81, CW06, WW02]. Well-Posedness [ABL13, Aug19, BH07, BEL02, CHY23, CM03a, CSZ13, EB20, GZ09, IZ00, Lü15, PCP22, Smi90, TBP18, VS80, WZ09, WCG14, Zas00, LT18, Vil04]. Wellposedness [Ahu16, HK07, Zol92]. Wendell [BSZ93]. Where [Mur86]. Which [MS98, MMK86, Pet88]. White [ALP10a, Bor93, Bor10, Cur76, DMS02, DF98d, HT22, JL16, Tud90, Uch81]. White-Noise [Bor93, Bor10]. Whole [DWZ20]. Whose [Mor24a, Nag85]. Wide [AW83, EAK86, KJ07, Kus79, Kus14]. Wide-Band [EAK86, Kus79, Kus14]. Wideband [KR87b, KR88, Kus89]. Width [FM16]. Wielandt [Les12]. Wiener [AMZ04, BR87a, Dav79, FC19, MC12, ZS19]. Will [CFNG10b]. Win [CIP96]. Window [CRZ^{+}18]. Wisdom [BFF20, TWB23]. Within [ZYL97, GŠZ05, GŠZ10]. Without [BGM97, HHS18, Las97, MV08a, SKD99, Sch92b, Stu15, ZL01, AG88, BFZ10, CDFP19, Cel07, Dav78, DM09, Gro12, HK19, KKA23, Kno77, Sîr14a, War76b, War83, War86, WL19, YY08, YH21, Zas00]. Wong [SY17]. Wonham [BCL04]. Workload [Mey03]. World [HSM19]. Worst [GP02, HG02, KS07, MCP21]. Worst-Case [HG02, KS07].

X [FZ19]. X-Ray [FZ19].

Y. [Mar77]. Yakubovich [PDB23, ZzJh10]. Years [Yin22]. Young [Ped98, Yük24].

Zakai [Ahm96, BK83a, Ito96, BGR90]. Zeno [GS18a, SP05]. Zenoness [CPS06]. Zero [BP20a, BS20a, BY13, BÇS07, BV13b, Bis77b, Bis12, BDP24, CLS12, CCHH19, CWC^{+}22, GR17b, Goe02, GYZ22, Ham06, HOPS87, HHS18, HLL00, HBY21, HBY23, Jas02, JN06, KS81, LN82, LS05a, LM16, LH87, LZ06, MSVA09, Now89, OPS87, PZ14, RV19, Ros00a, SKD99, SPO22, Shi02, Sîr14a, Sîr14b, Sun21, SWW23, VA03, WS87, Yon90, Yu15b, Zha05, TA96, TPA97]. Zero-Balance [BÇS07]. Zero-Diffusion [BP20a]. Zero-Dynamics [CWC^{+}22, LZ06]. Zero-Sum [BS20a, BY13, Bis77b, Bis12, BDP24, CLS12, CCHH19, GR17b, Goe02, GYZ22, Ham06, HHS18, HLL00, HBY21, HBY23, Jas02, JN06, KS81, LN82, LS05a, LM16, MSVA09, Now89, PZ14, SPO22, Shi02, Sîr14a, Sîr14b, Sun21, SWW23, VA03, Yon90, Yu15b, Zha05, TA96, TPA97]. Zero-Time-Controllability [BV13b]. Zeros [BK08, CFNG10b, Dmi92, LMP95, SWP91, Sch92a, Sch92b, vdW99]. Zin [AH21b]. Zooplankton [GT07]. Zowe [HN96]. Zubov [CGW01, CGW08, GS11].

References

Alpern:1986:ASS

[AA86] Steve Alpern and Miroslav Ašić. Ambush strategies in search games on graphs. SIAM Journal on Control and Optimization, 24(1):66–75, January 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Anderson:1992:LPA

[AA92] Edward J. Anderson and Miguel Aramendia. A linear programming approach to the search game on a network with mobile hider. SIAM Journal on Control and Optimization, 30(3):675–694, May 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Arutyunov:1997:IDP

[AA97] Aram V. Arutyunov and Sergei M. Aseev. Investigation of the degeneracy phenomenon of the maximum principle for optimal control problems with state constraints. SIAM Journal on Control and Optimization, 35(3):930–952, May 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                                     //epubs.siam.org/sam-bin/
                                                     dbq/article/26996.

Anfinsen:2017:AOF

[AA17] Henrik Anfinsen and Ole Morten Aamo. Adaptive output feedback stabilization of n+m coupled linear hyperbolic PDEs with uncertain boundary conditions. SIAM Journal on Control and Optimization, 55 (6):3928–3946, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ambrosino:2012:CCR

[AAA12] R. Ambrosino, M. Ariola, and F. Amato. A convex condition for robust stability analysis via polyhedral Lyapunov functions. SIAM Journal on Control and Optimization, 50 (1):490–506, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Aguilera:1986:OPV

[AAC86] N. Aguilera, H. W. Alt, and L. A. Caffarelli. An optimization problem with volume constraint. SIAM Journal on Control and Optimization, 24 (2):191–198, March 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Araruna:2019:CES

[AAFC19] F. D. Araruna, B. S. V.

Araújo, and E. Fernández-Cara. Carleman estimates for some two-dimensional degenerate parabolic PDEs and applications. SIAM Journal on Control and Optimization, 57(6):3985–4010, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Albrecher:2020:ORD

[AAM20] Hansjörg Albrecher, Pablo Azcue, and Nora Muler. Optimal ratcheting of dividends in insurance. SIAM Journal on Control and Optimization, 58(4):1822–1845, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Apte:2018:OCN

[AAR18] Amit Apte, Didier Auroux, and Mythily Ramaswamy. Observers for compressible Navier–Stokes equation. SIAM Journal on Control and Optimization, 56(2):1081–1104, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Arcak:2002:UII

[AAS02] Murat Arcak, David Angeli, and Eduardo Sontag. A unifying integral ISS framework for stability of nonlinear cascades. SIAM Journal on Control and Optimization, 40(6):1888–1904, November 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/38798.

Aihara:1993:MLE

[AB93] Shin Ichi Aihara and Arunabha Bagchi. Maximum likelihood estimator for two-point boundary value process. SIAM Journal on Control and Optimization, 31(1):147–158, January 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Alabau-Boussouira:2002:IBS

[AB02a] Fatiha Alabau-Boussouira. Indirect boundary stabilization of weakly coupled hyperbolic systems. SIAM Journal on Control and Optimization, 41(2):511–541, March 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38536.

Alvarez:2002:VSM

[AB02b] Olivier Alvarez and Martino Bardi. Viscosity solutions methods for singular perturbations in deterministic and stochastic control. SIAM Journal on Control and Optimization, 40(4):1159–1188, July 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36674.

Ancona:2002:FSP

[AB02c] Fabio Ancona and Alberto Bressan. Flow stability of patchy vector fields and robust feedback stabilization. SIAM Journal on Control and Optimization, 41(5): 1455–1476, September 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/39167.

Alabau-Boussouira:2003:TLE

[AB03] Fatiha Alabau-Boussouira. A two-level energy method for indirect boundary observability and controllability of weakly coupled hyperbolic systems. SIAM Journal on Control and Optimization, 42(3):871–906, May 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/40260.

Aranda:2005:ROD

[AB05] Ernesto Aranda and José C. Bellido. Relaxation of an optimal design problem with an integral-type constraint. SIAM Journal on Control and Optimization, 44 (1):195–209, January 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/41866.

Alpern:2006:RHD

[AB06] Steve Alpern and Vic Baston. Rendezvous in higher dimensions. SIAM Journal on Control and Optimization, 44(6): 2233–2252, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Angeli:2009:CSU

[AB09] David Angeli and Pierre-Alexandre Bliman. Convergence speed of unsteady distributed consensus: Decay estimate along the settling spanning-trees. SIAM Journal on Control and Optimization, 48(1):1–32, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Arapostathis:2010:URP

[AB10a] Ari Arapostathis and Vivek S.

Borkar. Uniform recurrence properties of controlled diffusions and applications to optimal control. SIAM Journal on Control and Optimization, 48(7):4181–4223, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Artstein:2010:POS

[AB10b] Zvi Artstein and Ido Bright.

Periodic optimization suffices for infinite horizon planar optimal control. SIAM Journal on Control and Optimization, 48(8):4963–4986, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Arapostathis:2012:RVI

[AB12] Ari Arapostathis and Vivek S.

Borkar. A relative value iteration algorithm for nondegenerate controlled diffusions. SIAM Journal on Control and Optimization, 50 (4):1886–1902, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See correction [AB17b].

Anantharam:2017:VFR

[AB17a] V. Anantharam and V. S. Borkar. A variational formula for risk-sensitive reward. SIAM Journal on Control and Optimization, 55(2):961–988, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Arapostathis:2017:CRV

[AB17b] Ari Arapostathis and Vivek S. Borkar. A correction to “A Relative Value Iteration Algorithm for Nondegenerate Controlled Diffusions”. SIAM Journal on Control and Optimization, 55(3):1711–1715, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [AB12].

Agrachev:2018:STO

[AB18] Andrei A. Agrachev and Carolina Biolo. Switching in timeoptimal problem with control in a ball. SIAM Journal on Control and Optimization, 56 (1):183–200, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Arapostathis:2020:VFR

[AB20] Ari Arapostathis and Anup Biswas. A variational formula for risk-sensitive control of diffusions in R^{d}. SIAM Journal on Control and Optimization, 58(1):85–103, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Attouch:2010:PSM

[ABAC10] Hédy Attouch, Luis M.

Briceño-Arias, and Patrick L. Combettes. A parallel splitting method for coupled monotone inclusions. SIAM Journal on Control and Optimization, 48(5):3246–3270, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Abounadi:2001:LAM

[ABB01] J. Abounadi, D. Bertsekas, and V. S. Borkar. Learning algorithms for Markov decision processes with average cost. SIAM Journal on Control and Optimization, 40(3):681–698, May 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36197.

Abounadi:2002:SAN

[ABB02] Jinane Abounadi, Dimitri P.

Bertsekas, and Vivek Borkar. Stochastic approximation for nonexpansive maps: Application to Q-learning algorithms. SIAM Journal on Control and Optimization, 41(1):1–22, January 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/34662.

Alvarez:2004:HRG

[ABB04] Felipe Alvarez, Jérôme Bolte, and Olivier Brahic. Hessian Riemannian gradient flows in convex programming. SIAM Journal on Control and Optimization, 43(2):477–501, March 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/41997.

Alessandri:2012:MMM

[ABB12] A. Alessandri, M. Baglietto, and G. Battistelli. Min-max moving-horizon estimation for uncertain discrete-time linear systems. SIAM Journal on Control and Optimization, 50 (3):1439–1465, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Agharkar:2015:VRA

[ABB15] Pushkarini Agharkar, Shaunak D. Bopardikar, and Francesco Bullo. Vehicle routing algorithms for radially escaping targets. SIAM Journal on Control and Optimization, 53(5):2934–2954, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Arapostathis:2020:VCR

[ABBK20] Ari Arapostathis, Anup Biswas, Vivek S. Borkar, and K. Suresh Kumar. A variational characterization of the risk-sensitive average reward for controlled diffusions on R^{d}. SIAM Journal on Control and Optimization, 58(6):3785–3813, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Anh:2022:PLA

[ABCD22] Pham The Anh, Artur Babiarz, Adam Czornik, and Thai Son Doan. Proportional local assignability of the dichotomy spectrum of onesided discrete time-varying linear systems. SIAM Journal on Control and Optimization, 60(3):1294–1319, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1410932.

Alabau-Boussouira:2017:CSD

[ABCL17] Fatiha Alabau-Boussouira, Piermarco Cannarsa, and Günter Leugering. Control and stabilization of degenerate wave equations. SIAM Journal on Control and Optimization, 55(3):2052–2087, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Alabau-Boussouira:2017:ICF

[ABCO17] Fatiha Alabau-Boussouira, Jean-Michel Coron, and Guillaume Olive. Internal controllability of first order quasilinear hyperbolic systems with a reduced number of controls. SIAM Journal on Control and Optimization, 55(1):300–323, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Arapostathis:1993:DTC

[ABFG^{+}93] Aristotle Arapostathis, Vivek S. Borkar, Emmanuel Fernández-Gaucherand, Mrinal K. Ghosh, and Steven I. Marcus. Discretetime controlled Markov processes with average cost criterion: a survey. SIAM Journal on Control and Optimization, 31(2):282–344, March 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Aranda-Bricaire:1998:ENS

[ABH98] E. Aranda-Bricaire and R. M. Hirschorn. Equivalence of nonlinear systems to prime systems under generalized output transformations. SIAM Journal on Control and Optimization, 37(1):118–130, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30706.

Ahn:2023:ATP

[ABHY23] Hyunjin Ahn, Junhyeok Byeon, Seung-Yeal Ha, and Jaeyoung Yoon. Asymptotic tracking of a point cloud moving on Riemannian manifolds. SIAM Journal on Control and Optimization, 61(4):2379–2406, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1523078.

Arapostathis:2014:CRV

[ABK14] Ari Arapostathis, Vivek S.

Borkar, and K. Suresh Kumar. Convergence of the relative value iteration for the ergodic control problem of nondegenerate diffusions under nearmonotone costs. SIAM Journal on Control and Optimization, 52(1):1–31, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Aronna:2019:OCP

[ABK19] M. Soledad Aronna, Joseph Frédéric Bonnans, and Axel Kröner. Optimal control of PDEs in a complex space setting: Application to the Schrödinger equation. SIAM Journal on Control and Optimization, 57 (2):1390–1412, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Aronna:2021:SCC

[ABK21] M. Soledad Aronna, J. Frédéric Bonnans, and Axel Kröner. State constrained controlaffine parabolic problems II: Second order sufficient optimality conditions. SIAM Journal on Control and Optimization, 59(2):1628–1655, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Aranda-Bricaire:1996:LDT

[ABKM96] E. Aranda-Bricaire, Ü. Kotta, and C. H. Moog. Linearization of discrete-time systems. SIAM Journal on Control and Optimization, 34(6): 1999–2023, November 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/26731.

Adly:2013:WPR

[ABL13] Samir Adly, Bernard Brogliato, and Ba Khiet Le. Wellposedness, robustness, and stability analysis of a setvalued controller for Lagrangian systems. SIAM Journal on Control and Optimization, 51(2):1592–1614, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Azimzadeh:2018:CIS

[ABL18] Parsiad Azimzadeh, Erhan Bayraktar, and George Labahn. Convergence of implicit schemes for Hamilton–Jacobi–Bellman quasi-variational inequalities. SIAM Journal on Control and Optimization, 56(6):3994– 4016, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Aravkin:2014:RTF

[ABP14] Aleksandr Y. Aravkin, James V.

Burke, and Gianluigi Pillonetto. Robust and trendfollowing Student’s t Kalman smoothers. SIAM Journal on Control and Optimization, 52(5):2891–2916, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Arada:2000:MCU

[ABR00] Nadir Arada, Maïtine Bergounioux, and Jean-Pierre Raymond. Minimax controls for uncertain parabolic systems. SIAM Journal on Control and Optimization, 38(5): 1481–1500, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34560.

Augier:2018:AEC

[ABS18] Nicolas Augier, Ugo Boscain, and Mario Sigalotti. Adiabatic ensemble control of a continuum of quantum systems. SIAM Journal on Control and Optimization, 56(6): 4045–4068, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Aubin:2008:DPS

[ABSP08] Jean-Pierre Aubin, Alexandre M. Bayen, and Patrick Saint-Pierre. Dirichlet problems for some Hamilton– Jacobi equations with inequality constraints. SIAM Journal on Control and Optimization, 47(5):2348–2380, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Antonietti:2013:MSO

[ABV13] Paola F. Antonietti, Alfio Borzì, and Marco Verani. Multigrid shape optimization governed by elliptic PDEs. SIAM Journal on Control and Optimization, 51(2):1417– 1440, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Acciaio:2019:EMF

[ABVC19] Beatrice Acciaio, Julio Backhoff-Veraguas, and René Carmona. Extended mean field control problems: Stochastic maximum principle and transport perspective. SIAM Journal on Control and Optimization, 57(6):3666–3693, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Aubin:1979:SPD

[AC79] J. P. Aubin and F. H. Clarke. Shadow prices and duality for a class of optimal control problems. SIAM Journal on Control and Optimization, 17(5):567–586, September 1979. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ahmed:2002:FLS

[AC02] N. U. Ahmed and C. D. Charalambous. Filtering for linear systems driven by fractional Brownian motion. SIAM Journal on Control and Optimization, 41(1):313–330, January 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36871.

Ancona:2005:AST

[AC05] Fabio Ancona and Giuseppe Maria Coclite. On the attainable set for Temple class systems with boundary controls. SIAM Journal on Control and Optimization, 43(6): 2166–2190, November 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40777.

Ahmed:2006:OMS

[AC06] N. U. Ahmed and C. D. Charalambous. Optimal measurement strategy for nonlinear filtering. SIAM Journal on Control and Optimization, 45(2): 519–531, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ahmed:2013:SMP

[AC13] N. U. Ahmed and Charalambos D. Charalambous. Stochastic minimum principle for partially observed systems subject to continuous and jump diffusion processes and driven by relaxed controls. SIAM Journal on Control and Optimization, 51(4): 3235–3257, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See erratum [AC17].

Ahmed:2017:ESM

[AC17] N. U. Ahmed and Charalambos D. Charalambous. Erratum: Stochastic minimum principle for partially observed systems subject to continuous and jump diffusion processes and driven by relaxed controls. SIAM Journal on Control and Optimization, 55(2): 1344–1345, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [AC13].

Akram:2018:GML

[AC18a] Muhammad Saeed Akram and Michael Cantoni. Gap metrics for linear time-varying systems. SIAM Journal on Control and Optimization, 56(2): 782–800, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ammari:2018:FSB

[AC18b] Kaïs Ammari and Emmanuelle Crépeau. Feedback stabilization and boundary controllability of the Korteweg–de Vries equation on a star-shaped network. SIAM Journal on Control and Optimization, 56(3): 1620–1639, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Allan:2019:PUK

[AC19] Andrew L. Allan and Samuel N.

Cohen. Parameter uncertainty in the Kalman–Bucy filter. SIAM Journal on Control and Optimization, 57(3): 1646–1671, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Acar:1993:ICE

[Aca93] Robert Acar. Identification of the coefficient in elliptic equations. SIAM Journal on Control and Optimization, 31(5): 1221–1244, September 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Achdou:2012:MFG

[ACCD12] Yves Achdou, Fabio Camilli, and Italo Capuzzo-Dolcetta. Mean field games: Numerical methods for the planning problem. SIAM Journal on Control and Optimization, 50 (1):77–109, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v50/i1/p77_s1.

Ancona:2018:OCL

[ACCG18] Fabio Ancona, Annalisa Cesaroni, Giuseppe M. Coclite, and Mauro Garavello. On the optimization of conservation law models at a junction with inflow and flow distribution controls. SIAM Journal on Control and Optimization, 56(5): 3370–3403, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Achdou:2005:IPP

[Ach05]  Yves Achdou.  An in-

verse problem for a parabolic variational inequality arising in volatility calibration with American options. SIAM Journal on Control and Optimization, 43(5): 1583–1615, September 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42442.

Achdou:2008:IPP

[Ach08] Yves Achdou. An inverse problem for a parabolic variational inequality with an integro-differential operator. SIAM Journal on Control and Optimization, 47(2):733–767, ???? 2008. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Afterman:2022:LFF

[ACKM22] Danielle Afterman, Pavel Chigansky, Marina Kleptsyna, and Dmytro Marushkevych. Linear filtering with fractional noises: Large time and small noise asymptotics. SIAM Journal on Control and Optimization, 60(3):1463–1487, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1360359.

Ahn:2009:HJB

[ACL09] Cheonghee Ahn, Hi Jun Choe, and Kijung Lee. On a Hamilton–Jacobi–Bellman equation related to a stochastic periodic model. SIAM Journal on Control and Optimization, 48(3):1309–1334, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Arapostathis:2019:ECC

[ACPZ19] Ari Arapostathis, Luis Caffarelli, Guodong Pang, and Yi Zheng. Ergodic control of a class of jump diffusions with finite Lévy measures and rough kernels. SIAM Journal on Control and Optimization, 57(2):1516–1540, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Aassila:2000:ASE

[ACS00a] M. Aassila, M. M. Cavalcanti, and J. A. Soriano. Asymptotic stability and energy decay rates for solutions of the wave equation with memory in a star-shaped domain. SIAM Journal on Control and Optimization, 38(5): 1581–1602, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34498.

Albano:2000:RRM

[ACS00b] Paolo Albano, Piermarco Cannarsa, and Carlo Sinestrari. Regularity results for the minimum time function of a class of semilinear evolution equations of parabolic type. SIAM Journal on Control and Optimization, 38(3):916–946, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33517.

Amstutz:2010:OCS

[ACT10] Samuel Amstutz and Marc Ciligot-Travain. Optimality conditions for shape and topology optimization subject to a cone constraint. SIAM Journal on Control and Optimization, 48(6):4056–4077, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Angelov:2022:AMP

[ACV22] Georgi Angelov, Alberto Domínguez Corella, and Vladimir M. Veliov. On the accuracy of the model predictive control method. SIAM Journal on Control and Optimization, 60 (4):2469–2487, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1460430.

Arnold:1983:SLS

[ACW83] L. Arnold, H. Crauel, and V. Wihstutz. Stabilization of linear systems by noise. SIAM Journal on Control and Optimization, 21(3):451–461, May 1983. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Alessandrini:1996:UDM

[AD96] Giovanni Alessandrini and Alvaro Diaz Valenzuela. Unique determination of multiple cracks by two measurements. SIAM Journal on Control and Optimization, 34(3):913–921, May 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26285.

Albertini:1998:SPL

[AD98] Francesca Albertini and Paolo Dai Pra. Small parameter limit for discrete-time partially observed risk-sensitive control problems. SIAM Journal on Control and Optimization, 37(1):1–32, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30780.

Aeyels:2002:EPF

[AD02] Dirk Aeyels and Patrick De Leenheer. Extension of the Perron–Frobenius theorem to homogeneous systems. SIAM Journal on Control and Optimization, 41(2):563–582, March 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36117.

Albertini:2008:IOM

[AD08] Francesca Albertini and Domenico D’Alessandro. Input-output model equivalence of spin systems: a characterization using lie algebra homomorphisms. SIAM Journal on Control and Optimization, 47(4):2016– 2043, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Aberkane:2015:RSR

[AD15] Samir Aberkane and Vasile Dragan. Robust stability and robust stabilization of a class of discrete-time timevarying linear stochastic systems. SIAM Journal on Control and Optimization, 53(1): 30–57, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ayala:2017:CLS

[AD17] Víctor Ayala and Adriano Da Silva. Controllability of linear systems on Lie groups with finite semisimple center. SIAM Journal on Control and Optimization, 55(2): 1332–1343, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Aurell:2018:MFT

[AD18] Alexander Aurell and Boualem Djehiche. Mean-field type modeling of nonlocal crowd aversion in pedestrian crowd dynamics. SIAM Journal on Control and Optimization, 56 (1):434–455, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Artacho:2011:MRN

[ADG^{+}11] F. J. Aragón Artacho, A. L.

Dontchev, M. Gaydu, M. H. Geoffroy, and V. M. Veliov. Metric regularity of Newton’s iteration. SIAM Journal on Control and Optimization, 49 (2):339–362, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i2/p339_s1.

Afraites:2008:SOS

[ADK08] L. Afraites, M. Dambrine, and D. Kateb. On second order shape optimization methods for electrical impedance tomography. SIAM Journal on Control and Optimization, 47(3):1556–1590, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Armstrong:2023:KCS

[ADL23] Niles Armstrong, Jory Denny, and Jeremy LeCrone. Kinodynamic control systems and discontinuities in clearance. SIAM Journal on Control and Optimization, 61 (4):2448–2467, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1516221.

Andreianov:2017:ASS

[ADM17] Boris P. Andreianov, Carlotta Donadello, and Andrea Marson. On the attainable set for a scalar nonconvex conservation law. SIAM Journal on Control and Optimization, 55(4): 2235–2270, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Albeverio:2022:MFL

[ADRU22] Sergio Albeverio, Francesco C. De Vecchi, Andrea Romano, and Stefania Ugolini. Meanfield limit for a class of stochastic ergodic control problems. SIAM Journal on Control and Optimization, 60(1):479–504, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1363479.

Atamas:2024:SAI

[ADS24] Ivan Atamas, Sergey Dashkovskiy, and Vitalii Slynko. Stability of abstract interconnected systems with a possibly unstable component. SIAM Journal on Control and Optimization, 62(1):581–599, February 2024. CODEN SJCODC. ISSN 1095-7138.

Anderson:1995:RSL

[AE95] Edward J. Anderson and Skander Essegaier. Rendezvous search on the line with indistinguishable players. SIAM Journal on Control and Optimization, 33(6): 1637–1642, November 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26070.

Alikoc:2017:PMS

[AE17] Baran Alikoc and Ali Fuat Ergenc. A polynomial method for stability analysis of LTI systems independent of delays. SIAM Journal on Control and Optimization, 55(4): 2661–2683, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Avdonin:2018:ECS

[AE18] Sergei Avdonin and Julian Edward. Exact controllability for string with attached masses. SIAM Journal on Control and Optimization, 56(2):945–980, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Alasseur:2020:ASA

[AEÉ^{+}20] Clémence Alasseur, Ivar Ekeland, Romuald Élie, Nicolás Hernández Santibáñez, and Dylan Possamaï. An adverse selection approach to power pricing. SIAM Journal on Control and Optimization, 58(2):686–713, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Aeyels:1981:GOD

[Aey81] Dirk Aeyels. Generic observability of differentiable systems. SIAM Journal on Control and Optimization, 19 (5):595–603, September 1981. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Aeyels:1985:GCS

[Aey85] Dirk Aeyels. Global controllability for smooth nonlinear systems: a geometric approach. SIAM Journal on Control and Optimization, 23 (3):452–465, May 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Aubin:1989:OSU

[AF89] Jean-Pierre Aubin and Halina Frankowska. Observability of systems under uncertainty. SIAM Journal on Control and Optimization, 27(5):949– 975, September 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ankirchner:2018:OCD

[AF18] Stefan Ankirchner and Alexander Fromm. Optimal control of diffusion coefficients via decoupling fields. SIAM Journal on Control and Optimization, 56(4):2959–2976, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Aubin-Frankowski:2021:LCL

[AF21] Pierre-Cyril Aubin-Frankowski.

Linearly constrained linear quadratic regulator from the viewpoint of kernel methods. SIAM Journal on Control and Optimization, 59(4): 2693–2716, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Araujo:2022:RCT

[AFCLS22] Raul K. C. Araújo, Enrique Fernández-Cara, Juan Límaco, and Diego A. Souza. Remarks on the control of twophase Stefan free-boundary problems. SIAM Journal on Control and Optimization, 60 (5):3078–3099, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1402261.

Alpern:2013:SGI

[AFGT13] Steve Alpern, Robbert Fokkink, Shmuel Gal, and Marco Timmer. On search games that include ambush. SIAM Journal on Control and Optimization, 51(6):4544–4556, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Alpern:2010:RCA

[AFK10] Steve Alpern, Robbert Fokkink, and Ken Kikuta. On Ruckle’s conjecture on accumulation games. SIAM Journal on Control and Optimization, 48(8): 5073–5083, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Alpern:2008:PMG

[AFLO08] Steve Alpern, Robbert Fokkink, Roy Lindelauf, and Geert-Jan Olsder. The “Princess and monster” game on an interval. SIAM Journal on Control and Optimization, 47(3): 1178–1190, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Aubin:1986:CCP

[AFO86] Jean-Pierre Aubin, Halina Frankowska, and Czesław Olech. Controllability of convex processes. SIAM Journal on Control and Optimization, 24(6):1192–1211, November 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Allendes:2022:EEP

[AFO22] Alejandro Allendes, Francisco Fuica, and Enrique Otárola. Error estimates for a pointwise tracking optimal control problem of a semilinear elliptic equation. SIAM Journal on Control and Optimization, 60(3):1763–1790, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1364151.

Allendes:2021:PEE

[AFOQ21] Alejandro Allendes, Francisco Fuica, Enrique Otarola, and Daniel Quero. A posteriori error estimates for a distributed optimal control problem of the stationary Navier–Stokes equations. SIAM Journal on Control and Optimization, 59 (4):2898–2923, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Acquistapace:1991:IBV

[AFT91] P. Acquistapace, F. Flandoli, and B. Terreni. Initial boundary value problems and optimal control for nonautonomous parabolic systems. SIAM Journal on Control and Optimization, 29(1): 89–118, January 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Alla:2017:EAP

[AFV17] A. Alla, M. Falcone, and S. Volkwein. Error analysis for POD approximations of infinite horizon problems via the dynamic programming approach. SIAM Journal on Control and Optimization, 55 (5):3091–3115, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Alpern:1988:MSM

[AG88] Steve Alpern and Shmuel Gal. A mixed-strategy minimax theorem without compactness. SIAM Journal on Control and Optimization, 26(6):1357– 1361, November 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Alpern:1995:RSL

[AG95] Steve Alpern and Shmuel Gal. Rendezvous search on the line with distinguishable players. SIAM Journal on Control and Optimization, 33(4):1270–1276, July 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26028.

Altman:1997:AON

[AG97a] Eitan Altman and Vladimir Gaitsgory. Asymptotic optimization of a nonlinear hybrid system governed by a Markov decision process. SIAM Journal on Control and Optimization, 35(6):2070–2085, November 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27998.

Artstein:1997:TFT

[AG97b] Zvi Artstein and Vladimir Gaitsgory. Tracking fast trajectories along a slow dynamics: a singular perturbations approach. SIAM Journal on Control and Optimization, 35(5):1487–1507, September 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28459.

Astrovskii:2019:OLT

[AG19] Anatoly I. Astrovskii and Ivan V. Gaishun. Observability of linear time-varying systems with quasiderivative coefficients. SIAM Journal on Control and Optimization, 57(3):1710–1729, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ammari:2013:NOC

[AGJ^{+}13] Habib Ammari, Pierre Garapon, François Jouve, Hyeonbae Kang, Mikyoung Lim, and Sanghyeon Yu. A new optimal control approach for the reconstruction of extended inclusions. SIAM Journal on Control and Optimization, 51(2): 1372–1394, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ammari:2012:SRA

[AGJK12] Habib Ammari, Josselin Garnier, Vincent Jugnon, and Hyeonbae Kang. Stability and resolution analysis for a topological derivative based imaging functional. SIAM Journal on Control and Optimization, 50(1):48–76, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicon/resource/
                                        1/sjcodc/v50/i1/p48_s1.

Akian:2008:MPF

[AGL08] Marianne Akian, Stéphane Gaubert, and Asma Lakhoua. The max-plus finite element method for solving deterministic optimal control problems: Basic properties and convergence analysis. SIAM Journal on Control and Optimization, 47(2):817–848, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Amigo:2016:ODF

[AGN^{+}16] R. Amigo, S. M. Giusti, A. A.

Novotny, E. C. N. Silva, and J. Sokolowski. Optimum design of flextensional piezoelectric actuators into two spatial dimensions. SIAM Journal on Control and Optimization, 54 (2):760–789, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Agrawal:1995:CAB

[Agr95] Rajeev Agrawal. The continuumarmed bandit problem. SIAM Journal on Control and Optimization, 33(6):1926–1951, November 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/23727.

Atar:2013:RSC

[AGS13] Rami Atar, Anindya Goswami, and Adam Shwartz. Risksensitive control for the parallel server model. SIAM Journal on Control and Optimization, 51(6):4363–4386, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Anderson:2014:CCF

[AH14] Brian D. O. Anderson and Uwe Helmke. Counting critical formations on a line. SIAM Journal on Control and Optimization, 52(1):219–242, ????

2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Aghaee:2021:SPA

[AH21a] Mahya Aghaee and William W. Hager. The switch point algorithm. SIAM Journal on Control and Optimization, 59(4): 2570–2593, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Aurand:2021:MHS

[AH21b] Joshua Aurand and Yu-Jui Huang. Mortality and healthcare: a stochastic control analysis under Epstein–Zin preferences. SIAM Journal on Control and Optimization, 59(5): 4051–4080, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Antil:2017:SOS

[AHH17] Harbir Antil, Sean Hardesty, and Matthias Heinkenschloss. Shape optimization of shell structure acoustics. SIAM Journal on Control and Optimization, 55(3):1347–1376, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Abda:2010:TSA

[AHJM10] A. Ben Abda, M. Hassine, M. Jaoua, and M. Masmoudi. Topological sensitivity analysis for the location of small cavities in Stokes flow. SIAM Journal on Control and Optimization, 48(5):2871–2900, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Azmi:2024:ARS

[AHK24] Behzad Azmi, Lukas Herrmann, and Karl Kunisch. Analysis of RHC for stabilization of nonautonomous parabolic equations under uncertainty. SIAM Journal on Control and Optimization, 62(1):220–242, January 2024. CODEN SJCODC. ISSN 1095-7138.

Albi:2022:MDP

[AHKS22] Giacomo Albi, Michael Herty, Dante Kalise, and Chiara Segala. Moment-driven predictive control of mean-field collective dynamics. SIAM Journal on Control and Optimization, 60(2):814–841, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1391559.

Ahmed:1977:OCG

[Ahm77]  N. U. Ahmed.  Optimal

control of generating policies in a power system governed by a second order hyperbolic partial differential equation. SIAM Journal on Control and Optimization, 15(6): 1016–1033, November 1977. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ahmed:1981:SCH

[Ahm81] N. U. Ahmed. Stochastic control on Hilbert space for linear evolution equations with random operator-valued coefficients. SIAM Journal on Control and Optimization, 19 (3):401–430, May 1981. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ahmed:1983:PRT

[Ahm83] N. U. Ahmed. Properties of relaxed trajectories for a class of nonlinear evolution equations on a Banach space. SIAM Journal on Control and Optimization, 21(6):953– 967, November 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ahmed:1996:ORC

[Ahm96] N. U. Ahmed. Optimal relaxed controls for infinitedimensional stochastic systems of Zakai type. SIAM Journal on Control and Optimization, 34(5):1592–1615, September 1996. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/26911.

Ahmed:2003:EOC

[Ahm03] N. U. Ahmed. Existence of optimal controls for a general class of impulsive systems on Banach spaces. SIAM Journal on Control and Optimization, 42(2):669–685, March 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39129.

Ahmed:2007:NDG

[Ahm07] N. U. Ahmed. Nonlinear diffusion governed by McKean– Vlasov equation on Hilbert space and optimal control. SIAM Journal on Control and Optimization, 46(1):356–378, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ahmed:2019:POS

[Ahm19] N. U. Ahmed. Partially observed stochastic evolution equations on Banach spaces and their optimal Lipschitz feedback control law. SIAM Journal on Control and Optimization, 57(5):3101–3117, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Alphonse:2020:SSS

[AHR20] Amal Alphonse, Michael Hintermüller, and Carlos N. Rautenberg. Stability of the solution set of quasi-variational inequalities and optimal control. SIAM Journal on Control and Optimization, 58(6): 3508–3532, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ammari:2023:LSA

[AHR23] Kaïs Ammari, Fathi Hassine, and Luc Robbiano. Logarithmic stabilization of an acoustic system with a damping term of Brinkman type. SIAM Journal on Control and Optimization, 61(1):170–197, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1430674.

Antunes:2013:SHS

[AHS13] Duarte Antunes, João P. Hespanha, and Carlos Silvestre. Stochastic hybrid systems with renewal transitions: Moment analysis with application to networked control systems with delays. SIAM Journal on Control and Optimization, 51(2):1481–1499, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ahuja:2016:WMF

[Ahu16] Saran Ahuja. Wellposedness of mean field games with common noise under a weak monotonicity condition. SIAM Journal on Control and Optimization, 54(1):30–48, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Arutyunov:2002:PDF

[AI02] Aram V. Arutyunov and Alexey F. Izmailov. Positive definiteness of forms: Numerical identification. SIAM Journal on Control and Optimization, 41(5):1567–1585, September 2002. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38992.

Aihara:1992:RML

[Aih92]  Shin Ichi Aihara.  Regu-

larized maximum likelihood estimate for an infinitedimensional parameter in stochastic parabolic systems. SIAM Journal on Control and Optimization, 30(4):745–764, July 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Angeli:2004:SPI

[AISW04] D. Angeli, B. Ingalls, E. D.

Sontag, and Y. Wang. Separation principles for inputoutput and integral-input-to-state stability. SIAM Journal on Control and Optimization, 43(1):256–276, January 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41904.

Anderson:1977:SPP

[AJ77] B. D. O. Anderson and E. I.

Jury. A “Simplest possible” property of the generalized Routh–Hurwitz conditions. SIAM Journal on Control and Optimization, 15(1): 177–184, January 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Almer:2012:DPA

[AJ12] Stefan Almér and Ulf T. Jönsson. Dynamic phasor analysis of pulse-modulated systems. SIAM Journal on Control and Optimization, 50 (3):1110–1138, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ayala:2016:ARG

[AJ16] Victor Ayala and Philippe Jouan. Almost-Riemannian geometry on Lie groups. SIAM Journal on Control and Optimization, 54(5):2919–2947, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ankirchner:2014:BST

[AJK14] Stefan Ankirchner, Monique Jeanblanc, and Thomas Kruse. BSDEs with singular terminal condition and a control problem with constraints. SIAM Journal on Control and Optimization, 52(2):893–913, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

alJamal:2018:LSP

[aJM18] Rasha al Jamal and Kirsten Morris. Linearized stability of partial differential equations with application to stabilization of the Kuramoto– Sivashinsky equation. SIAM Journal on Control and Optimization, 56(1):120–147, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ahmadi:2014:JSR

[AJPR14] Amir Ali Ahmadi, Raphaël M.

Jungers, Pablo A. Parrilo, and Mardavij Roozbehani. Joint spectral radius and pathcomplete graph Lyapunov functions. SIAM Journal on Control and Optimization, 52 (1):687–717, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Alabau:1999:BOC

[AK99a] Fatiha Alabau and Vilmos Komornik. Boundary observability, controllability, and stabilization of linear elastodynamic systems. SIAM Journal on Control and Optimization, 37(2):521–542, March 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31383.

Altman:1999:ACC

[AK99b] Eitan Altman and Harold J.

Kushner. Admission control for combined guaranteed performance and best effort communications systems under heavy traffic. SIAM Journal on Control and Optimization, 37(6):1780–1807, November 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33351.

Altman:2002:CPP

[AK02a] Eitan Altman and Harold J. Kushner. Control of polling in presence of vacations in heavy traffic with applications to satellite and mobile radio systems. SIAM Journal on Control and Optimization, 41(1):217–252, January 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35846.

Auckly:2002:EMC

[AK02b] David Auckly and Lev Kapitanski. On the λ-equations for matching control laws. SIAM Journal on Control and Optimization, 41(5): 1372–1388, September 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39330.

Aseev:2004:PMP

[AK04] Sergei M. Aseev and Arkady V. Kryazhimskiy. The Pontryagin maximum principle and transversality conditions for a class of optimal control problems with infinite time horizons. SIAM Journal on Control and Optimization, 43(3):1094–1119, May 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42751.

Arutyunov:2015:SCP

[AK15] A. V. Arutyunov and D. Yu.

Karamzin. On some continuity properties of the measure Lagrange multiplier from the maximum principle for state constrained problems. SIAM Journal on Control and Optimization, 53(4):2514–2540, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Azmi:2016:SBE

[AK16] Behzad Azmi and Karl Kunisch. On the stabilizability of the Burgers equation by receding horizon control. SIAM Journal on Control and Optimization, 54(3):1378–1405, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Azmi:2019:HFD

[AK19] Behzad Azmi and Karl Kunisch. A hybrid finitedimensional RHC for stabilization of time-varying parabolic equations. SIAM Journal on Control and Optimization, 57 (5):3496–3526, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Akcay:2001:GOB

[Akç01] Hüseyin Akçay. General orthonormal bases for robust identification in H_{∞}. SIAM Journal on Control and Optimization, 40(3):947–968, May 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/36039.

Alvarez:2000:ULS

[AKN00] Olivier Alvarez, Shigeaki Koike, and Isao Nakayama. Uniqueness of lower semicontinuous viscosity solutions for the minimum time problem. SIAM Journal on Control and Optimization, 38(2):470–481, March 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/31719.

Akhmetgaliyev:2017:CME

[AKO17] Eldar Akhmetgaliyev, Chiu-Yen Kao, and Braxton Osting. Computational methods for extremal Steklov problems. SIAM Journal on Control and Optimization, 55(2): 1226–1240, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Arutyunov:2005:NMP

[AKP05] A. Arutyunov, D. Karamzin, and F. Pereira. A nondegenerate maximum principle for the impulse control problem with state constraints. SIAM Journal on Control and Optimization, 43(5):1812–1843, September 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43006.

Antil:2021:RAC

[AKP21] Harbir Antil, Drew P. Kouri, and Johannes Pfefferer. Riskaverse control of fractional diffusion with uncertain exponent. SIAM Journal on Control and Optimization, 59(2): 1161–1187, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ankirchner:2024:RCD

[AKTW24] Stefan Ankirchner, Nabil Kazi-Tani, and Julian Wendt. The role of correlation in diffusion control ranking games. SIAM Journal on Control and Optimization, 62(3):1465– 1489, June 2024. CODEN SJ-CODC. ISSN 1095-7138.

Alpern:1998:SRE

[AL98] Steve Alpern and Wei Shi Lim.

The symmetric rendezvousevasion game. SIAM Journal on Control and Optimization, 36(3):948–959, May 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30977.

Avalos:2000:BCT

[AL00] George Avalos and Irena Lasiecka. Boundary controllability of thermoelastic plates via the free boundary conditions. SIAM Journal on Control and Optimization, 38(2):337–383, March 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33983.

Agrachev:2001:LAS

[AL01] Andrei A. Agrachev and Daniel Liberzon. Lie-algebraic stability criteria for switched systems. SIAM Journal on Control and Optimization, 40 (1):253–269, January 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36570.

Artstein:2002:SPC

[AL02] Zvi Artstein and Arie Leizarowitz. Singularly perturbed control systems with one-dimensional fast dynamics. SIAM Journal on Control and Optimization, 41(2):641–658, March 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39088.

Alvarez:2008:OSI

[AL08] Luis H. R. Alvarez and Jukka Lempa. On the optimal stochastic impulse control of linear diffusions. SIAM Journal on Control and Optimization, 47(2):703–732, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Aguilar:2012:STL

[AL12] Cesar O. Aguilar and Andrew D. Lewis. Small-time local controllability for a class of homogeneous systems. SIAM Journal on Control and Optimization, 50(3):1502–1517, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Allwright:1988:PSM

[All88] J. C. Allwright. Positive semidefinite matrices: Characterization via conical hulls and least-squares solution of a matrix equation. SIAM Journal on Control and Optimization, 26(3):537–556, May 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See erratum and addendum [AW90].

Allibert:1999:ACL

[All99] Brice Allibert. Analytic controllability of a linear hybrid system. SIAM Journal on Control and Optimization, 37(3):844–868, May 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32062.

Almudevar:2001:DPA

[Alm01]  Anthony Almudevar.  A

dynamic programming algorithm for the optimal control of piecewise deterministic Markov processes. SIAM Journal on Control and Optimization, 40(2):525– 539, March 2001. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36447.

Almudevar:2008:AFP

[Alm08] Anthony Almudevar. Approximate fixed point iteration with an application to infinite horizon Markov decision processes. SIAM Journal on Control and Optimization, 47(5): 2303–2347, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Almeida:2024:OCB

[ALNP24] Luis Almeida, Alexis Léculier, Grégoire Nadin, and Yannick Privat. Optimal control of bistable traveling waves: Looking for the best spatial distribution of a killing action to block a pest invasion. SIAM Journal on Control and Optimization, 62(2): 1291–1315, April 2024. CO-DEN SJCODC. ISSN 1095-7138.

Alpern:1988:GRD

[Alp88] Steve Alpern. Game with repeated decisions. SIAM Journal on Control and Optimization, 26(2):468–477, March 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Alpern:1995:RSP

[Alp95] Steve Alpern. The rendezvous search problem. SIAM Journal on Control and Optimization, 33(3):673–683, May 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24919.

Alpay:2010:LSS

[ALP10a] Daniel Alpay, David Levanony, and Ariel Pinhas. Linear stochastic state space theory in the white noise space setting. SIAM Journal on Control and Optimization, 48 (8):5009–5027, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Alpern:2010:SGT

[Alp10b] Steve Alpern. Search games on trees with asymmetric travel times. SIAM Journal on Control and Optimization, 48(8): 5547–5563, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicon/resource/
                                        1/sjcodc/v48/i8/p5547_s1.

Avalos:1999:LTD

[ALR99] George Avalos, Irena Lasiecka, and Richard Rebarber. Lack of time-delay robustness for stabilization of a structural acoustics model. SIAM Journal on Control and Optimization, 37(5):1394–1418, September 1999. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/33113.

Altshuller:2009:DIQ

[Alt09]  D. A. Altshuller.  Delay-

integral-quadratic constraints and absolute stability of timeperiodic feedback systems. SIAM Journal on Control and Optimization, 47(6):3185– 3202, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Alvarez:2000:MPS

[Alv00] Felipe Alvarez. On the minimizing property of a second order dissipative system in Hilbert spaces. SIAM Journal on Control and Optimization, 38(4):1102–1119, July 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33580.

Alvarez:2001:SSC

[Alv01] Luis H. R. Alvarez. Singular stochastic control, linear diffusions, and optimal stopping: a class of solvable problems. SIAM Journal on Control and Optimization, 39(6):1697–1710, November 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36782.

Alvarez:2004:SFS

[Alv04] Luis H. R. Alvarez. Stochastic forest stand value and optimal timber harvesting. SIAM Journal on Control and Optimization, 42(6): 1972–1993, November 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39345.

Agler:2013:CSQ

[ALY13] Jim Agler, Zinaida A. Lykova, and N. J. Young. A case of μ-synthesis as a quadratic semidefinite program. SIAM Journal on Control and Optimization, 51(3):2472–2508, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Anderson:1981:DST

[AM81] B. D. O. Anderson and J. B. Moore. Detectability and stabilizability of time-varying discrete-time linear systems. SIAM Journal on Control and Optimization, 19(1):20– 32, January 1981. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Allwright:1982:SES

[AM82] D. J. Allwright and A. I. Mees. Stability, extended spaces and numerical ranges. SIAM Journal on Control and Optimization, 20(3):328–337, May 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Alt:1989:CFE

[AM89] Walter Alt and Uwe Mackenroth. Convergence of finite element approximations to state constrained convex parabolic boundary control problems. SIAM Journal on Control and Optimization, 27(4):718–736, July 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Amar:1995:NVP

[AM95] Micòl Amar and Carlo Mariconda. A nonconvex variational problem with constraints. SIAM Journal on Control and Optimization, 33 (1):299–307, January 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/23504.

Ancona:1998:ASS

[AM98] Fabio Ancona and Andrea Marson. On the attainable set for scalar nonlinear conservation laws with boundary control. SIAM Journal on Control and Optimization, 36(1):290–312, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30440.

Autenried:2014:SRG

[AM14] Christian Autenried and Irina Markina. Sub-Riemannian geometry of Stiefel manifolds. SIAM Journal on Control and Optimization, 52(2):939–959, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Azcue:2015:ODP

[AM15] Pablo Azcue and Nora Muler.

Optimal dividend payment and regime switching in a compound Poisson risk model. SIAM Journal on Control and Optimization, 53(5):3270– 3298, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Alvarez:2017:ESR

[AM17] Luis H. R. Alvarez and Pekka Matomäki. Expected supremum representation of the value of a singular stochastic control problem. SIAM Journal on Control and Optimization, 55(6):3908–3927, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Arjmand:2022:MMT

[AM22] Saeed Sadeghi Arjmand and Guilherme Mazanti. Multipopulation minimal-time mean field games. SIAM Journal on Control and Optimization, 60(4):1942–1969, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1407306.

Alphonse:2023:ANC

[AM23] Paul Alphonse and Jérémy Martin. Approximate nullcontrollability with uniform cost for the hypoelliptic Ornstein–Uhlenbeck equations. SIAM Journal on Control and Optimization, 61 (3):1679–1711, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1487412.

Amadori:2004:QVI

[Ama04]  A. L. Amadori.  Quasi-

variational inequalities with Dirichlet boundary condition related to exit time problems for impulse control. SIAM Journal on Control and Optimization, 43(2):570–589, March 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40973.

Avakov:2020:LCF

[AMI20] Evgeny Avakov and Georgii Magaril-Il’yaev. Local controllability and a family of maximum principles for a free time optimal control problem. SIAM Journal on Control and Optimization, 58(6): 3212–3236, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ammari:2002:IIB

[Amm02] Habib Ammari. An inverse initial boundary value problem for the wave equation in the presence of imperfections of small volume. SIAM Journal on Control and Optimization, 41(4):1194–1211, July 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38424.

Aguech:2000:PAA

[AMP00] Rafik Aguech, Eric Moulines, and Pierre Priouret. On a perturbation approach for the analysis of stochastic tracking algorithms. SIAM Journal on Control and Optimization, 39(3):872–899, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/33385.

Andrieu:2005:SSA

[AMP05] Christophe Andrieu, Éric Moulines, and Pierre Priouret. Stability of stochastic approximation under verifiable conditions. SIAM Journal on Control and Optimization, 44 (1):283–312, January 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/41726.

Allaire:2010:LTB

[AMP10] Grégoire Allaire, Arnaud Münch, and Francisco Periago. Long time behavior of a twophase optimal design for the heat equation. SIAM Journal on Control and Optimization, 48(8):5333–5356, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Apel:2015:RSD

[AMPR15] Thomas Apel, Mariano Mateos, Johannes Pfefferer, and Arnd Rösch. On the regularity of the solutions of Dirichlet optimal control problems in polygonal domains. SIAM Journal on Control and Optimization, 53(6):3620–3641, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Alvarez:2014:CSO

[AMR14] Luis H. R. Alvarez, Pekka Matomäki, and Teppo A. Rakkolainen. A class of solvable optimal stopping problems of spectrally negative jump diffusions. SIAM Journal on Control and Optimization, 52(4):2224–2249, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Aronna:2020:HOM

[AMR20] M. Soledad Aronna, Monica Motta, and Franco Rampazzo. A higher-order maximum principle for impulsive optimal control problems. SIAM Journal on Control and Optimization, 58(2):814–844, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Akian:1996:ICM

[AMS96] Marianne Akian, José Luis Menaldi, and Agnès Sulem. On an investment-consumption model with transaction costs. SIAM Journal on Control and Optimization, 34(1):329– 364, January 1996. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/24715.

Amstutz:2013:RPT

[Ams13] Samuel Amstutz. Regularized perimeter for topology optimization. SIAM Journal on Control and Optimization, 51 (3):2176–2199, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ammar:2007:GDO

[AMV07] Sabeur Ammar, Mohamed Mabrouk, and Jean-Claude Vivalda. On the genericity of the differential observability of controlled discrete-time systems. SIAM Journal on Control and Optimization, 46(6): 2182–2198, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ammar:2018:GSO

[AMV18] Sabeur Ammar, Majid Massaoud, and Jean-Claude Vivalda. Genericity of the strong observability for sampled systems. SIAM Journal on Control and Optimization, 56(2): 1463–1490, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Anulova:2020:IIA

[AMV20] Svetlana V. Anulova, Hilmar Mai, and Alexander Yu. Veretennikov. On iteration improvement for averaged expected cost control for onedimensional ergodic diffusions. SIAM Journal on Control and Optimization, 58(4):2312– 2331, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Amparan:2004:WHF

[AMZ04] A. Amparan, S. Marcaida, and I. Zaballa. Wiener–Hopf factorization indices and infinite structure of rational matrices. SIAM Journal on Control and Optimization, 42(6): 2130–2144, November 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40259.

Apkarian:2006:CDN

[AN06] Pierre Apkarian and Dominikus Noll. Controller design via nonsmooth multidirectional search. SIAM Journal on Control and Optimization, 44(6):1923–1949, January 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Abbas:2015:MWEa

[AN15a] Z. Abbas and S. Nicaise. The multidimensional wave equation with generalized acoustic boundary conditions I: Strong stability. SIAM Journal on Control and Optimization, 53 (4):2558–2581, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Abbas:2015:MWEb

[AN15b] Z. Abbas and S. Nicaise. The multidimensional wave equation with generalized acoustic boundary conditions II: Polynomial stability. SIAM Journal on Control and Optimization, 53(4):2582–2607, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Apkarian:2020:OKC

[AN20] Pierre Apkarian and Dominikus Noll. Optimizing the Kreiss constant. SIAM Journal on Control and Optimization, 58(6):3342–3362, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ancona:2021:GCS

[AN21] Fabio Ancona and Khai T. Nguyen. On the global controllability of scalar conservation laws with boundary and source controls. SIAM Journal on Control and Optimization, 59(6):4314–4338, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Anderson:1982:IES

[And82] Brian D. O. Anderson. Internal and external stability of linear time-varying systems. SIAM Journal on Control and Optimization, 20(3):408–413, May 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Anderson:1988:RNO

[And88] R. F. Anderson. Replacement with nonconstant operating cost. SIAM Journal on Control and Optimization, 26(5): 1076–1098, September 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Andrieu:2014:CSN

[And14] Vincent Andrieu. Convergence speed of nonlinear Luenberger observers. SIAM Journal on Control and Optimization, 52 (5):2831–2856, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Angell:1976:ETH

[Ang76] Thomas S. Angell. Existence theorems for hereditary Lagrange and Mayer problems of optimal control. SIAM Journal on Control and Optimization, 14(1):1–18, January 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Appeltans:2022:ADS

[ANM22] Pieter Appeltans, Silviu-Iulian Niculescu, and Wim Michiels. Analysis and design of strongly stabilizing PID controllers for time-delay systems. SIAM Journal on Control and Optimization, 60(1):124–146, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M136726X.

Anonymous:1995:JWA

[Ano95] Anonymous. Jack warga: In appreciation. SIAM Journal on Control and Optimization, 33(2):347–350, March 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Anderson:1983:SPC

[ANP83] E. J. Anderson, P. Nash, and A. F. Perold. Some properties of a class of continuous linear programs. SIAM Journal on Control and Optimization, 21(5):758–765, September 1983. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Apkarian:2008:MCN

[ANR08] P. Apkarian, D. Noll, and A. Rondepierre. Mixed H_{2}/H_{∞} control via nonsmooth optimization. SIAM Journal on Control and Optimization, 47(3):1516–1546, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Antil:2014:OCF

[ANS14] Harbir Antil, Ricardo H. Nochetto, and Patrick Sodré. Optimal control of a free boundary problem: Analysis with second-order sufficient conditions. SIAM Journal on Control and Optimization, 52(5): 2771–2799, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Antunes:2014:OBE

[Ant14] Pedro R. S. Antunes. Optimal bilaplacian eigenvalues. SIAM Journal on Control and Optimization, 52(4):2250–2260, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Antil:2015:FOC

[AO15] Harbir Antil and Enrique Otárola. A FEM for an optimal control problem of fractional powers of elliptic operators. SIAM Journal on Control and Optimization, 53(6): 3432–3456, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Agram:2023:SFP

[AØ23] Nacira Agram and Bernt Øksendal. Stochastic Fokker– Planck equations for conditional McKean–Vlasov jump diffusions and applications to optimal control. SIAM Journal on Control and Optimization, 61(3):1472–1493, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1461034.

Antil:2016:STF

[AOS16] Harbir Antil, Enrique Otárola, and Abner J. Salgado. A space–time fractional optimal control problem: Analysis and discretization. SIAM Journal on Control and Optimization, 54(3):1295–1328, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Anderson:1992:EPL

[AP92] Edward J. Anderson and Andrew B. Philpott. Extreme points for linear optimal control problems with diagonal structure. SIAM Journal on Control and Optimization, 30(6):1385–1394, November 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Anderson:1994:SCC

[AP94] Edward J. Anderson and Andrew B. Philpott. On the solutions of a class of continuous linear programs. SIAM Journal on Control and Optimization, 32(5): 1289–1296, September 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/22721.

Andrieu:2006:EKK

[AP06] Vincent Andrieu and Laurent Praly. On the existence of a Kazantzis–Kravaris/ Luenberger observer. SIAM Journal on Control and Optimization, 45(2):432–456, January 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Altschuler:2020:LER

[AP20] Jason M. Altschuler and Pablo A. Parrilo. Lyapunov exponent of rank-one matrices: Ergodic formula and inapproximability of the optimal distribution. SIAM Journal on Control and Optimization, 58 (1):510–528, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Andrieu:2008:HAR

[APA08] Vincent Andrieu, Laurent Praly, and Alessandro Astolfi. Homogeneous approximation, recursive observer design, and output feedback. SIAM Journal on Control and Optimization, 47(4):1814–1850, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Alvergue:2016:CCH

[APGC16] Luis D. Alvergue, Abhishek Pandey, Guoxiang Gu, and Xiang Chen. Consensus control for heterogeneous multiagent systems. SIAM Journal on Control and Optimization, 54(3):1719–1738, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Akbarian:2024:CEL

[APH24] M. Akbarian, N. Pariz, and A. Heydari. Constituting an extension of Lyapunov’s direct method. SIAM Journal on Control and Optimization, 62(4):2346–2366, August 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1595242.

Amann:2005:OCP

[AQ05] H. Amann and P. Quittner. Optimal control problems with final observation governed by explosive parabolic equations. SIAM Journal on Control and Optimization, 44(4):1215–1238, July 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43345.

Aumann:1977:OSF

[AR77] Robert J. Aumann and Uriel G. Rothblum. Orderable set functions and continuity. III: Orderability and absolute continuity. SIAM Journal on Control and Optimization, 15(1):156–162, January 1977. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Arada:1999:MCP

[AR99a] Nadir Arada and Jean-Pierre Raymond. Minimax control of parabolic systems with state constraints. SIAM Journal on Control and Optimization, 38(1):254–271, January 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34141.

Arada:1999:NOC

[AR99b] Nadir Arada and Jean-Pierre Raymond. Necessary optimality conditions for control problems and the Stone–Cech compactification. SIAM Journal on Control and Optimization, 37(4):1011–1032, July 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33035.

Arada:2000:OCP

[AR00] N. Arada and J. P. Raymond. Optimal control problems with mixed control-state constraints. SIAM Journal on Control and Optimization, 39(5):1391–1407, September 2000. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/35792.

Allan:2021:RSF

[AR21] Douglas A. Allan and James B. Rawlings. Robust stability of full information estimation. SIAM Journal on Control and Optimization, 59(5): 3472–3497, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Arada:2012:OCS

[Ara12] Nadir Arada. Optimal control of shear-thinning fluids. SIAM Journal on Control and Optimization, 50(4):2515–2542, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Arada:2013:OCS

[Ara13] Nadir Arada. Optimal control of shear-thickening flows. SIAM Journal on Control and Optimization, 51(3):1940– 1961, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Arada:2014:OCE

[Ara14] Nadir Arada. Optimal control of evolutionary quasi-Newtonian fluids. SIAM Journal on Control and Optimization, 52(5):3401–3436, ???? 2014. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Ariki:1998:GPA

[Ari98] Susumu Ariki. Generic pole assignment via dynamic feedback. SIAM Journal on Control and Optimization, 36 (1):379–390, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27468.

Avalos:2021:GCL

[ÁRL21] G. Gómez Ávalos, J. Muñoz Rivera, and Z. Liu. Gevrey class of locally dissipative Euler–Bernoulli beam equation. SIAM Journal on Control and Optimization, 59(3): 2174–2194, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Armentano:1986:PCO

[Arm86] Vinícius A. Armentano. The pencil (sE−A) and controllability– observability for generalized linear systems: a geometric approach. SIAM Journal on Control and Optimization, 24 (4):616–638, July 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Apel:2009:EEG

[ARS09a] Thomas Apel, Arnd Rösch, and Dieter Sirch. L^{∞}-error estimates on graded meshes with application to optimal control. SIAM Journal on Control and Optimization, 48(3): 1771–1796, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ayala:2009:OHS

[ARS09b] V. Ayala, J. C. Rodríguez, and L. A. B. San Martín. Optimality on homogeneous spaces, and the angle system associated with a bilinear control system. SIAM Journal on Control and Optimization, 48(4):2636–2650, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ayala:2014:CRC

[ARS14] Victor Ayala, Rodolfo Ribeiro, Jr., and Luiz A. B. San Martin. Controllability on Sl(2, C) with restricted controls. SIAM Journal on Control and Optimization, 52(4): 2548–2567, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Altarovici:2017:OCI

[ARS17] Albert Altarovici, Max Reppen, and H. Mete Soner. Optimal consumption and investment with fixed and proportional transaction costs. SIAM Journal on Control and Optimization, 55(3):1673–1710, ???? 2017. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Artstein:1978:RCD

[Art78] Zvi Artstein. Relaxed controls and the dynamics of control systems. SIAM Journal on Control and Optimization, 16(5):689–701, September 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Artstein:1989:PIM

[Art89a] Zvi Artstein. Parametrized integration of multifunctions with applications to control and optimization. SIAM Journal on Control and Optimization, 27(6):1369–1380, November 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Artstein:1989:ROC

[Art89b] Zvi Artstein. Rapid oscillations, chattering systems, and relaxed controls. SIAM Journal on Control and Optimization, 27(5):940–948, September 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Allan:2021:NDI

[ART21] Douglas A. Allan, James Rawlings, and Andrew R. Teel. Nonlinear detectability and incremental input/ output-to-state stability. SIAM Journal on Control and Optimization, 59(4):3017–3039, ???? 2021. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Aihara:1988:IID

[AS88] S. I. Aihara and Y. Sunahara. Identification of an infinite-dimensional parameter for stochastic diffusion equations. SIAM Journal on Control and Optimization, 26(5):1062–1075, September 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Altman:1991:MDP

[AS91] Eitan Altman and Adam Shwartz. Markov decision problems and state-action frequencies. SIAM Journal on Control and Optimization, 29 (4):786–809, July 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Albertini:1993:DTT

[AS93] Francesca Albertini and Eduardo D. Sontag. Discretetime transitivity and accessibility: Analytic systems. SIAM Journal on Control and Optimization, 31(6):1599– 1622, November 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Avdonin:1995:IBO

[AS95] Sergei Avdonin and Thomas I. Seidman. Identification of q(x) in u_{t} = Δu − qu from boundary observations. SIAM Journal on Control and Optimization, 33(4):1247–1255, July 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24972.

Agrachev:2003:LSO

[AS03] Andrei A. Agrachev and Mario Sigalotti. On the local structure of optimal trajectories in R^{3}. SIAM Journal on Control and Optimization, 42(2):513–531, March 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40924.

Attouch:2011:CDN

[AS11] H. Attouch and B. F. Svaiter.

A continuous dynamical Newtonlike approach to solving monotone inclusions. SIAM Journal on Control and Optimization, 49(2):574–598, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sicon/
resource/1/sjcodc/v49/i2/
p574_s1.

Alfonsi:2013:CMC

[AS13] Aurélien Alfonsi and Alexander Schied. Capacitary measures for completely monotone kernels via singular control. SIAM Journal on Control and Optimization, 51(2): 1758–1780, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Agrachev:2020:CSE

[AS20] Andrei Agrachev and Andrey Sarychev. Control in the spaces of ensembles of points. SIAM Journal on Control and Optimization, 58(3): 1579–1596, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Almi:2021:TOI

[AS21] Stefano Almi and Ulisse Stefanelli. Topology optimization for incremental elastoplasticity: a phase-field approach. SIAM Journal on Control and Optimization, 59(1):339–364, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Aamo:2005:BCL

[ASK05] Ole Morten Aamo, Andrey Smyshlyaev, and Miroslav Krstic. Boundary control of the linearized Ginzburg– Landau model of vortex shedding. SIAM Journal on Control and Optimization, 43(6): 1953–1971, November 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/42601.

Abedi:2022:UWU

[ASP22] Ehsan Abedi, Simone Carlo Surace, and Jean-Pascal Pfister. A unification of weighted and unweighted particle filters. SIAM Journal on Control and Optimization, 60 (2):597–619, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1382404.

Ahmad:2013:AEB

[ASR^{+}13] Naveed Ahmad, Muhammad Shoaib Saleem, Nasir Rehman, Sultan Hussain, and Malkhaz Shashiashvili. Area estimation between the early exercise boundaries for the American put option with different local volatilities. SIAM Journal on Control and Optimization, 51(3):1988–2004, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Arnuautu:2005:OPC

[AST05] Viorel Arnuautu, Jürgen Sprekels, and Dan Tiba. Optimization problems for curved mechanical structures. SIAM Journal on Control and Optimization, 44(2):743–775, March 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42625.

Alves:2009:SIS

[ASTT09] Carlos Alves, Ana Leonor Silvestre, Takéo Takahashi, and Marius Tucsnak. Solving inverse source problems using observability. applications to the Euler–Bernoulli plate equation. SIAM Journal on Control and Optimization, 48 (3):1632–1659, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Agrachev:2002:SOB

[ASZ02] Andrei A. Agrachev, Gianna Stefani, and PierLuigi Zezza. Strong optimality for a bang– bang trajectory. SIAM Journal on Control and Optimization, 41(4):991–1014, July 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/38866.

Artstein:1982:LSI

[AT82] Zvi Artstein and Gilead Tadmor. Linear systems with indirect controls: The underlying measures. SIAM Journal on Control and Optimization, 20 (1):96–111, January 1982. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Abergel:1989:OCS

[AT89] F. Abergel and R. Temam. Optimality conditions for some nonqualified problems of distributed control. SIAM Journal on Control and Optimization, 27(1):1–12, January 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Andersland:1992:ISC

[AT92] Mark S. Andersland and Demosthenis Teneketzis. Information structures, causality, and nonsequential stochastic control I: Design-independent properties. SIAM Journal on Control and Optimization, 30(6):1447–1475, November 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Andersland:1994:ISC

[AT94] Mark S. Andersland and Demosthenis Teneketzis. Information structures, causality, and nonsequential stochastic control. II: Design-dependent properties. SIAM Journal on Control and Optimization, 32(6):1726–1751, November 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/21883.

Acquistapace:1996:IHL

[AT96] Paolo Acquistapace and Brunello Terreni. Infinite-horizon linear-quadratic regulator problems for nonautonomous parabolic systems with boundary control. SIAM Journal on Control and Optimization, 34(1):

1–30, January 1996. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ammari:2000:SBE

[AT00a] Kais Ammari and Marius Tucsnak. Stabilization of Bernoulli–Euler beams by means of a pointwise feedback force. SIAM Journal on Control and Optimization, 39(4):1160–1181, July 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/34931.

Apkarian:2000:PLC

[AT00b] Pierre Apkarian and Hoang Duong Tuan. Parameterized LMIs in control theory. SIAM Journal on Control and Optimization, 38(4):1241–1264, July 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/32612.

Aiki:2017:RCP

[AT17] Toyohiko Aiki and Sergey A. Timoshin. Relaxation for a control problem in concrete carbonation modeling. SIAM Journal on Control and Optimization, 55(6):3489–3502, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Afsari:2013:CGD

[ATV13] Bijan Afsari, Roberto Tron, and René Vidal. On the convergence of gradient descent for finding the Riemannian center of mass. SIAM Journal on Control and Optimization, 51(3):2230–2260, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Almgren:1993:CDF

[ATW93] Fred Almgren, Jean E. Taylor, and Lihe Wang. Curvaturedriven flows: a variational approach. SIAM Journal on Control and Optimization, 31 (2):387–438, March 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Aubin:1990:DGV

[Aub90a] Jean-Pierre Aubin. Differential games: a viability approach. SIAM Journal on Control and Optimization, 28(6):1294–1320, November 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Aubin:1990:SVT

[Aub90b] Jean-Pierre Aubin. A survey of viability theory. SIAM Journal on Control and Optimization, 28(4):749–788, July 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Aubin:2001:VKC

[Aub01]  Jean-Pierre Aubin.  Via-

bility kernels and capture basins of sets under differential inclusions. SIAM Journal on Control and Optimization, 40(3):853–881, May 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36968.

Aubin:2002:BVP

[Aub02] Jean-Pierre Aubin. Boundaryvalue problems for systems of Hamilton–Jacobi–Bellman inclusions with constraints. SIAM Journal on Control and Optimization, 41(2):425– 456, March 2002. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38151.

Aubin:2011:VSS

[Aub11] Jean-Pierre Aubin. Viability solutions to structured Hamilton–Jacobi equations under constraints. SIAM Journal on Control and Optimization, 49(5):1881–1915, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sicon/
resource/1/sjcodc/v49/i5/
p1881_s1.

Aubin:2013:CSE

[Aub13] Jean-Pierre Aubin. Chaperoning state evolutions by variable durations. SIAM Journal on Control and Optimization, 51(3):2081–2101, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Augner:2019:WPS

[Aug19] Björn Augner. Well-posedness and stability of infinitedimensional linear port-Hamiltonian systems with nonlinear boundary feedback. SIAM Journal on Control and Optimization, 57(3):1818–1844, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Auslender:1984:SMP

[Aus84]  Alfred Auslender.  Stabil-

ity in mathematical programming with nondifferentiable data. SIAM Journal on Control and Optimization, 22(2): 239–254, March 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Auslender:1999:AAP

[Aus99] A. Auslender. Asymptotic analysis for penalty and barrier methods in variational inequalities. SIAM Journal on Control and Optimization, 37(2):653–671, March 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/31090.

Avakov:2019:LCA

[Ava19] Evgeny Avakov. Local controllability for abnormal systems. SIAM Journal on Control and Optimization, 57(2): 947–962, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Augeraud-Veron:2022:GTA

[AVCCD22] Emmanuelle Augeraud-Véron, Catherine Choquet, Éloïse Comte, and Moussa M. Diédhiou. A game theory approach for the groundwater pollution control. SIAM Journal on Control and Optimization, 60(3):1667–1689, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/19M1278223.

Averboukh:2016:ASC

[Ave16] Yurii Averboukh. Approximate solutions of continuoustime stochastic games. SIAM Journal on Control and Optimization, 54(5):2629–2649, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Averboukh:2019:APS

[Ave19] Yurii Averboukh. Approximate public-signal correlated equilibria for nonzero-sum differential games. SIAM Journal on Control and Optimization, 57(1):743–772, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Acciaio:2021:CNE

[AVJ21] Beatrice Acciaio, Julio Backhoff Veraguas, and Junchao Jia. Cournot–Nash equilibrium and optimal transport in a dynamic setting. SIAM Journal on Control and Optimization, 59(3):2273–2300, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ammar:2020:SDO

[AVZ20] Sabeur Ammar, Jean-Claude Vivalda, and Basma Zitouni. Strong differential observability for sampled-time systems. SIAM Journal on Control and Optimization, 58(6):3814– 3841, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Arnold:1983:WSS

[AW83] L. Arnold and V. Wihstutz.

Wide sense stationary solutions of linear systems with additive noise. SIAM Journal on Control and Optimization, 21 (3):413–426, May 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Allwright:1990:EAP

[AW90] J. C. Allwright and K. C. Woodgate. Erratum and addendum: “Positive Semidefinite Matrices: Characterization via Conical Hulls and Least-Squares Solution of a Matrix Equation” [SIAM J. Control Optim. 26 (1988), no. 3, 537–556; MR 89e:65034]. SIAM Journal on Control and Optimization, 28(1):250– 251, January 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [All88].

Ahmed:2021:OCN

[AW21] N. U. Ahmed and Shian Wang. Optimal control of nonlinear hybrid systems driven by signed measures with variable intensities and supports. SIAM Journal on Control and Optimization, 59(6):4268– 4294, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ahmed:1996:NBC

[AX96] N. U. Ahmed and X. Xiang. Nonlinear boundary control of semilinear parabolic systems. SIAM Journal on Control and Optimization, 34(2):473–490, March 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25492.

Ahmed:1997:NUS

[AX97] N. U. Ahmed and X. Xiang.

Nonlinear uncertain systems and necessary conditions of optimality. SIAM Journal on Control and Optimization, 35(5):1755–1772, September 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28556.

Anderson:2002:NSC

[AX02] Edward J. Anderson and Huifu Xu. Necessary and sufficient conditions for optimal offers in electricity markets. SIAM Journal on Control and Optimization, 41(4):1212–1228, July 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36780.

Anderson:2005:OBE

[AX05] Edward J. Anderson and Huifu Xu. ε-optimal bidding in an electricity market with discontinuous market distribution function. SIAM Journal on Control and Optimization, 44(4):1391–1418, July 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42555.

Ammann:2023:AFW

[AY23a] Luis Ammann and Irwin Yousept. Acoustic full waveform inversion via optimal control: First- and secondorder analysis. SIAM Journal on Control and Optimization, 61(4):2468–2496, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1480045.

Arslan:2023:SEU

[AY23b] Gürdal Arslan and Serdar Yüksel. Subjective equilibria under beliefs of exogenous uncertainty for dynamic games. SIAM Journal on Control and Optimization, 61 (3):1038–1062, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1475946.

Atar:1997:LEF

[AZ97] Rami Atar and Ofer Zeitouni. Lyapunov exponents for finite state nonlinear filtering. SIAM Journal on Control and Optimization, 35(1):36–55, January 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/27204.

Agrachev:2007:FCC

[AZ07] Andrei Agrachev and Igor Zelenko. On feedback classification of control-affine systems with one- and two-dimensional inputs. SIAM Journal on Control and Optimization, 46(4): 1431–1460, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Araruna:2008:CKS

[AZ08] F. D. Araruna and E. Zuazua.

Controllability of the Kirchhoff system for beams as a limit of the Mindlin– Timoshenko system. SIAM Journal on Control and Optimization, 47(4):1909–1938, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Angoshtari:2020:PFP

[AZZ20] Bahman Angoshtari, Thaleia Zariphopoulou, and Xun Yu Zhou. Predictable forward performance processes: The binomial case. SIAM Journal on Control and Optimization, 58(1):327–347, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Byrnes:1984:OFG

[BA84] C. I. Byrnes and B. D. O. Anderson. Output feedback and generic stabilizability. SIAM Journal on Control and Optimization, 22(3):362–380, May 1984. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Back:2008:DPL

[BA08] Juhoon Back and Alessandro Astolfi. Design of positive linear observers for positive linear systems via coordinate transformations and positive realizations. SIAM Journal on Control and Optimization, 47 (1):345–373, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Befekadu:2015:AEE

[BA15] Getachew K. Befekadu and Panos J. Antsaklis. On the asymptotic estimates for exit probabilities and minimum exit rates of diffusion processes pertaining to a chain of distributed control systems. SIAM Journal on Control and Optimization, 53(4): 2297–2318, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bacciotti:1979:SMO

[Bac79]  Andrea Bacciotti.  Struc-

ture métrique des orbites de familles symétriques de champs de vecteurs et théorie du temps minimum. (French) [Metric structure of the orbits of symmetric families of field vectors and theory of minimum time]. SIAM Journal on Control and Optimization, 17 (2):311–319, March 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Badra:2009:LFL

[Bad09]  Mehdi Badra.  Lyapunov

function and local feedback boundary stabilization of the Navier–Stokes equations. SIAM Journal on Control and Optimization, 48(3): 1797–1830, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bagchi:1981:NAE

[Bag81] Arunabha Bagchi. A note on asymptotically efficient estimates of parameters in continuous-time dynamical systems. SIAM Journal on Control and Optimization, 19 (2):169–176, March 1981. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bagchi:1984:AII

[Bag84] Arunabha Bagchi. Approximation of Itô integrals arising in stochastic time-delayed systems. SIAM Journal on Control and Optimization, 22(6): 878–888, November 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bahlali:2008:NSO

[Bah08] Seid Bahlali. Necessary and sufficient optimality conditions for relaxed and strict control problems. SIAM Journal on Control and Optimization, 47(4):2078–2095, ????

2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Briceno-Arias:2018:PMS

[BAKS18] L. M. Briceño-Arias, D. Kalise, and F. J. Silva. Proximal methods for stationary mean field games with local couplings. SIAM Journal on Control and Optimization, 56(2): 801–836, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Balder:1977:EDS

[Bal77] E. J. Balder. An extension of duality-stability relations to nonconvex optimization problems. SIAM Journal on Control and Optimization, 15(2): 329–343, February 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Balas:1978:MCC

[Bal78] Mark J. Balas. Modal control of certain flexible dynamic systems. SIAM Journal on Control and Optimization, 16 (3):450–462, May 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Balder:1981:LSI

[Bal81] E. J. Balder. Lower semicontinuity of integral functionals with nonconvex integrands by relaxation-compactification. SIAM Journal on Control and Optimization, 19(4):533–542, July 1981. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Balder:1982:LCP

[Bal82] E. J. Balder. Lower closure problems with weak convergence conditions in a new perspective. SIAM Journal on Control and Optimization, 20 (2):198–210, March 1982. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Balder:1984:GAL

[Bal84] E. J. Balder. A general approach to lower semicontinuity and lower closure in optimal control theory. SIAM Journal on Control and Optimization, 22(4):570–598, July 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Balder:1986:SIF

[Bal86] E. J. Balder. On seminormality of integral functionals and their integrands. SIAM Journal on Control and Optimization, 24(1):95–121, January 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Balakrishnan:1991:SRT

[Bal91] A. V. Balakrishnan. Stochastic regulator theory for a class of abstract wave equations. SIAM Journal on Control and Optimization, 29(6):

1288–1299, November 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Balder:1994:NER

[Bal94] Erik J. Balder. New existence results for optimal controls in the absence of convexity: The importance of extremality. SIAM Journal on Control and Optimization, 32(3):890–916, May 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/19309.

Balder:2001:CER

[Bal01] Erik J. Balder. Comment on an existence result for a noncoercive nonconvex variational problem. SIAM Journal on Control and Optimization, 40(1):328–332, January 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37184.  See

[CM96b].

Banon:1978:ENI

[Ban78a] G. Banon. Errata: Nonparametric identification for diffusion processes. SIAM Journal on Control and Optimization, 16(6):860, November 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [Ban78b].

Banon:1978:NID

[Ban78b] G. Banon. Nonparametric identification for diffusion processes. SIAM Journal on Control and Optimization, 16(3): 380–395, May 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [Ban78a].

Bank:2005:OCU

[Ban05] Peter Bank. Optimal control under a dynamic fuel constraint. SIAM Journal on Control and Optimization, 44(4):1529–1541, July 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/61696.

Bandini:2019:CBD

[Ban19] Elena Bandini. Constrained BSDEs driven by a nonquasi-left-continuous random measure and optimal control of PDMPs on bounded domains. SIAM Journal on Control and Optimization, 57(6): 3767–3798, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bernard:2018:EOP

[BAP18] P. Bernard, V. Andrieu, and L. Praly. Expressing an observer in preferred coordinates by transforming an injective immersion into a surjective diffeomorphism. SIAM Journal on Control and Optimization, 56(3):2327–2352, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Barmish:1977:ECA

[Bar77]  B. Ross Barmish.  Enve-

lope correspondence and its application to the design of guaranteed error controllers. SIAM Journal on Control and Optimization, 15(6):1001– 1015, November 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Barbu:1978:CCP

[Bar78] Viorel Barbu. Convex control problems and Hamiltonian systems of infinite intervals. SIAM Journal on Control and Optimization, 16(6): 895–911, November 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Baram:1980:TEB

[Bar80a] Yoram Baram. Truncation error bounds and convergence of least squares estimates. SIAM Journal on Control and Optimization, 18(4):346–351, ???? 1980. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Barbu:1980:BCP

[Bar80b] Viorel Barbu. Boundary control problems with convex cost criterion. SIAM Journal on Control and Optimization, 18 (2):227–243, March 1980. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Barron:1980:FEA

[Bar80c] Emmanuel Nicholas Barron.

Finite energy approximations in stochastic and deterministic differential games. SIAM Journal on Control and Optimization, 18(3):245–263, May 1980. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Barbu:1981:NCD

[Bar81] Viorel Barbu. Necessary conditions for distributed control problems governed by parabolic variational inequalities. SIAM Journal on Control and Optimization, 19(1): 64–86, January 1981. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Barbu:1982:BCP

[Bar82] Viorel Barbu. Boundary control problems with nonlinear state equation. SIAM Journal on Control and Optimization, 20(1):125–143, January 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Barbu:1983:OFC

[Bar83] Viorel Barbu. Optimal feedback controls for a class of nonlinear distributed parameter systems. SIAM Journal on Control and Optimization, 21(6):871–894, November 1983. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Baratchart:1985:PLC

[Bar85a]  L. Baratchart.  On the

parametrization of linear constant systems. SIAM Journal on Control and Optimization, 23(5):752–773, September 1985. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Barles:1985:DIC

[Bar85b] G. Barles. Deterministic impulse control problems. SIAM Journal on Control and Optimization, 23(3):419–432, May 1985. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Barron:1985:VSM

[Bar85c] Emmanuel N. Barron. Viscosity solutions for the monotone control problem. SIAM Journal on Control and Optimization, 23(2):161–171, March 1985. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Barbu:1988:PFA

[Bar88] Viorel Barbu. A product formula approach to nonlinear optimal control problems. SIAM Journal on Control and Optimization, 26(3):497–520, May 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bardi:1989:BVP

[Bar89] Martino Bardi. A boundary value problem for the minimum-time function. SIAM Journal on Control and Optimization, 27(4):776–785, July 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Barbu:1991:DPE

[Bar91]  Viorel Barbu.  The dy-

namic programming equation for the time-optimal control problem in infinite dimensions. SIAM Journal on Control and Optimization, 29(2): 445–456, March 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Barron:1993:ALM

[Bar93] E. N. Barron. Averaging in Lagrange and minimax problems of optimal control. SIAM Journal on Control and Optimization, 31(6):1630– 1652, November 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Barbu:1994:PCC

[Bar94]  Viorel Barbu.  The H_{∞}-

problem with control constraints. SIAM Journal on Control and Optimization, 32(4):952–964, July 1994.

CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/22783.

Barbu:1995:BCS

[Bar95a] Viorel Barbu. H_{∞} boundary control with state feedback: The hyperbolic case. SIAM Journal on Control and Optimization, 33(3):684–701, May 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23859.

Barbu:1995:PID

[Bar95b]  Viorel Barbu.  The H_{∞}-

problem for infinite-dimensional semilinear systems. SIAM Journal on Control and Optimization, 33(4):1017–1027, July 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24405.

Barbu:1997:OCL

[Bar97] Viorel Barbu. Optimal control of linear periodic resonant systems in Hilbert spaces. SIAM Journal on Control and Optimization, 35(6): 2137–2156, November 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29478.

Barbu:2012:SNS

[Bar12] Viorel Barbu. Stabilization of Navier–Stokes equations by oblique boundary feedback controllers. SIAM Journal on Control and Optimization, 50(4):2288–2307, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Barbu:2017:SDA

[Bar17] Viorel Barbu. The steepest descent algorithm in Wasserstein metric for the sandpile model of self-organized criticality. SIAM Journal on Control and Optimization, 55(1): 413–428, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Barbu:2020:OFC

[Bar20] Viorel Barbu. Optimal feedback controllers for a stochastic differential equation with reflection. SIAM Journal on Control and Optimization, 58 (2):986–997, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Barbu:2021:CFP

[Bar21] Viorel Barbu. The controllability of Fokker–Planck equations with reflecting boundary conditions and controllers in diffusion term. SIAM Journal on Control and Optimization, 59(1):709–726, ???? 2021.

CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Barbu:2023:ECF

[Bar23a] Viorel Barbu. Exact controllability of Fokker–Planck equations and McKean–Vlasov SDEs. SIAM Journal on Control and Optimization, 61 (3):1805–1818, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M150784X.

Barbu:2023:EOC

[Bar23b] Viorel Barbu. Existence of optimal control for nonlinear Fokker–Planck equations in L^{1}(R^{d}). SIAM Journal on Control and Optimization, 61(3):1213–1230, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1485243.

Basar:1977:INE

[Bas77] Tamer Basar. Informationally nonunique equilibrium solutions in differential games. SIAM Journal on Control and Optimization, 15(4):636–660, July 1977. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Basar:1984:AIS

[Baş84] Tamer Başar. Affine incentive schemes for stochastic systems with dynamic information. SIAM Journal on Control and Optimization, 22(2): 199–210, March 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Basu:1991:FTD

[Bas91] Sankar Basu. A framework for two-dimensional hyperstability theory based provably convergent adaptive twodimensional IIR filtering. SIAM Journal on Control and Optimization, 29(6):1476– 1508, November 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Brivadis:2021:LOI

[BASG21] Lucas Brivadis, Vincent Andrieu, Ulysse Serres, and Jean-Paul Gauthier. Luenberger observers for infinite-dimensional systems, back and forth nudging, and application to a crystallization process. SIAM Journal on Control and Optimization, 59(2):857–886, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Battilotti:1995:NSI

[Bat95]  S. Battilotti.  Noninterac-

tion and stability via invertible feedback laws and some existence conditions. SIAM Journal on Control and Optimization, 33(1):107–125, January 1995. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/23988.

Battilotti:2007:LBD

[Bat07] Stefano Battilotti. Lyapunovbased design of iISS feedforward systems with uncertainty and noisy measurements. SIAM Journal on Control and Optimization, 46(1): 84–115, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Battilotti:2019:MSP

[Bat19] S. Battilotti. Multilayer state predictors for nonlinear systems with time-varying measurement delays. SIAM Journal on Control and Optimization, 57(3):1541–1566, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bauschke:2009:NPE

[Bau09] Heinz H. Bauschke. A note on the paper by Eckstein and Svaiter on “General Projective Splitting Methods for Sums of Maximal Monotone Operators”. SIAM Journal on Control and Optimization, 48(4): 2513–2515, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [ES09].

Bayen:2009:APR

[Bay09a] Térence Bayen. Analytical parameterization of rotors and proof of a Goldberg conjecture by optimal control theory. SIAM Journal on Control and Optimization, 47(6): 3007–3036, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bayraktar:2009:PSF

[Bay09b] Erhan Bayraktar. A proof of the smoothness of the finite time horizon American put option for jump diffusions. SIAM Journal on Control and Optimization, 48(2): 551–572, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Banks:1978:HCP

[BB78] H. T. Banks and J. A.

Burns. Hereditary control problems: Numerical methods based on averaging approximations. SIAM Journal on Control and Optimization, 16 (2):169–208, March 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Baston:1988:EGB

[BB88] V. J. Baston and F. A. Bostock. An evasion game with barriers. SIAM Journal on Control and Optimization, 26(5):1099–1105, September 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Baras:1989:OSS

[BB89] John S. Baras and Alain Bensoussan. Optimal sensor scheduling in nonlinear filtering of diffusion processes. SIAM Journal on Control and Optimization, 27(4):786–813, July 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Baston:1990:EGF

[BB90] V. J. Baston and F. A. Bostock. An evasion game on a finite tree. SIAM Journal on Control and Optimization, 28 (3):671–677, May 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Beck:1992:RLS

[BB92] Anatole Beck and Micah Beck. The revenge of the linear search problem. SIAM Journal on Control and Optimization, 30(1):112–122, January 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Baston:1993:IDG

[BB93] V. J. Baston and F. A. Bostock. Infinite deterministic graphical games. SIAM Journal on Control and Optimization, 31(6):1623–1629, November 1993. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Baston:1994:EGI

[BB94] V. J. Baston and F. A. Bostock. An evasion game with an infinite number of states. SIAM Journal on Control and Optimization, 32(2):471–479, March 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/18961.

Belmiloudi:1997:CMA

[BB97] Aziz Belmiloudi and Françoise Brossier. A control method for assimilation of surface data in a linearized Navier– Stokes-type problem related to oceanography. SIAM Journal on Control and Optimization, 35(6):2183–2197, November 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28613.

Bagagiolo:1998:SPF

[BB98] Fabio Bagagiolo and Martino Bardi. Singular perturbation of a finite horizon problem with state-space constraints. SIAM Journal on Control and Optimization, 36(6):2040–2060, November 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/31447.

Bhat:2000:FTS

[BB00] Sanjay P. Bhat and Dennis S. Bernstein. Finitetime stability of continuous autonomous systems. SIAM Journal on Control and Optimization, 38(3):751–766, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/32135.

Bhat:2003:NBL

[BB03] Sanjay P. Bhat and Dennis S. Bernstein. Nontangency-based Lyapunov tests for convergence and stability in systems having a continuum of equilibria. SIAM Journal on Control and Optimization, 42(5): 1745–1775, September 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/40711.

Borkar:2004:ECC

[BB04] Vivek Borkar and Amarjit Budhiraja. Ergodic control for constrained diffusions: Characterization using HJB equations. SIAM Journal on Control and Optimization, 43(4):1467–1492, July 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41761.

Bentahar:2006:BOH

[BB06] Imen Bentahar and Bruno Bouchard. Barrier option hedging under constraints: a viscosity approach. SIAM Journal on Control and Optimization, 45(5):1846–1874, ???? 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Belhachmi:2007:SUC

[BB07] Zakaria Belhachmi and Dorin Bucur. Stability and uniqueness for the crack identification problem. SIAM Journal on Control and Optimization, 46(1):253–273, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Basu:2008:SCI

[BB08] Arnab Basu and Vivek S.

Borkar. Stochastic control with imperfect models. SIAM Journal on Control and Optimization, 47(3):1274–1300, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Benkherouf:2009:OPC

[BB09] Lakdere Benkherouf and Alain Bensoussan. Optimality of an (s, S) policy with compound Poisson and diffusion demands: a quasi-variational inequalities approach. SIAM Journal on Control and Optimization, 48(2):756–762, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Barilari:2018:SPS

[BB18] Davide Barilari and Francesco Boarotto. On the set of points of smoothness for the value function of affine optimal control problems. SIAM Journal on Control and Optimization, 56(2):649–671, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

BenAmara:2019:BCT

[BB19] Jamel Ben Amara and Emna Beldi. Boundary controllability of two vibrating strings connected by a point mass with variable coefficients. SIAM Journal on Control and Optimization, 57(5):3360– 3387, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bonalli:2023:FOP

[BB23] Riccardo Bonalli and Benoît Bonnet. First-order Pontryagin maximum principle for risk-averse stochastic optimal control problems. SIAM Journal on Control and Optimization, 61(3):1881–1909, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1489137.

Bayen:2024:HMP

[BBB24] Térence Bayen, Anas Bouali, and Loïc Bourdin. The hybrid maximum principle for optimal control problems with spatially heterogeneous dynamics is a consequence of a Pontryagin maximum principle for L^{1}-local solutions. SIAM Journal on Control and Optimization, 62 (4):2412–2432, August 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M155311X.

Baotic:2008:ELC

[BBBM08] Mato Baotić, Francesco Borrelli, Alberto Bemporad, and Manfred Morari. Efficient on-line computation of constrained optimal control. SIAM Journal on Control and Optimization, 47(5): 2470–2489, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Banks:1981:PEI

[BBC81] H. T. Banks, J. A. Burns, and E. M. Cliff. Parameter estimation and identification for systems with delays. SIAM Journal on Control and Optimization, 19(6):791–828, November 1981. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bauschke:2003:BMO

[BBC03] Heinz H. Bauschke, Jonathan M. Borwein, and Patrick L. Combettes. Bregman Monotone optimization algorithms. SIAM Journal on Control and Optimization, 42(2):596– 636, March 2003. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/40712.

Barles:2014:BAR

[BBC14] G. Barles, A. Briani, and E. Chasseigne. A Bellman approach for regional optimal control problems in R^{N}. SIAM Journal on Control and Optimization, 52(3):1712– 1744, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bayraktar:2018:NSM

[BBC18] Erhan Bayraktar, Amarjit Budhiraja, and Asaf Cohen. A numerical scheme for a mean field game in some queueing systems based on Markov chain approximation method. SIAM Journal on Control and Optimization, 56(6):4017– 4044, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Babiarz:2019:PLA

[BBC^{+}19] Artur Babiarz, Irina Banshchikova, Adam Czornik, Evgenii Makarov, Michał Niezabitowski, and Svetlana Popova. Proportional local assignability of Lyapunov spectrum of linear discrete time-varying systems. SIAM Journal on Control and Optimization, 57(2):1355–1377, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bayraktar:2022:COI

[BBCS22] Erhan Bayraktar, Christoph Belak, Sören Christensen, and Frank T. Seifried. Convergence of optimal investment problems in the vanishing fixed cost limit. SIAM Journal on Control and Optimization, 60(5):2712–2736, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1462179.

Baradel:2018:OCU

[BBD18] Nicolas Baradel, Bruno Bouchard, and Ngoc Minh Dang. Optimal control under uncertainty and Bayesian parameters adjustments. SIAM Journal on Control and Optimization, 56 (2):1038–1057, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Brockwell:1999:SAR

[BBE99] A. Brockwell, K. Borovkov, and R. Evans. Stability of an adaptive regulator for partially known nonlinear stochastic systems. SIAM Journal on Control and Optimization, 37(5): 1553–1567, September 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/33151.

Bensoussan:2018:PBE

[BBF18] Alain Bensoussan, Dominic Breit, and Jens Frehse. Parabolic Bellman equations with risk control. SIAM Journal on Control and Optimization, 56(2):1535–1549, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Barilari:2012:SCN

[BBG12] Davide Barilari, Ugo Boscain, and Jean-Paul Gauthier. On 2-step, corank 2, nilpotent sub-Riemannian metrics. SIAM Journal on Control and Optimization, 50(1):559–582, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Benabdallah:2014:SEO

[BBGBO14] Assia Benabdallah, Franck Boyer, Manuel González-Burgos, and Guillaume Olive. Sharp estimates of the onedimensional boundary control cost for parabolic systems and application to the N-dimensional boundary null controllability in cylindrical domains. SIAM Journal on Control and Optimization, 52 (5):2970–3001, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bhatt:2000:MPE

[BBK00] A. G. Bhatt, A. Budhiraja, and R. L. Karandikar. Markov property and ergodicity of the nonlinear filter. SIAM Journal on Control and Optimization, 39(3):928–949, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35770.

Bettiol:2018:OSL

[BBR18] P. Bettiol, B. Bonnard, and J. Rouot. Optimal Strokes at low Reynolds number: a geometric and numerical study of copepod and Purcell swimmers. SIAM Journal on Control and Optimization, 56(3): 1794–1822, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Benallouch:2014:OBC

[BBT14] Mohamed Benallouch, M. Boutayeb, and H. Trinh. H_{∞} observerbased control for discrete-time one-sided Lipschitz systems with unknown inputs. SIAM Journal on Control and Optimization, 52(6):3751–3775, ???? 2014. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Barb:2003:RDI

[BBTN03] Florin Dan Barb, Aharon Ben-Tal, and Arkadi Nemirovski. Robust dissipativity of interval uncertain linear systems. SIAM Journal on Control and Optimization, 41(6): 1661–1695, November 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39817.

Bettiol:2010:TSS

[BBV10] Piernicola Bettiol, Alberto Bressan, and Richard Vinter. On trajectories satisfying a state constraint: W^{1,1} estimates and counterexamples. SIAM Journal on Control and Optimization, 48(7): 4664–4679, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bettiol:2011:ETC

[BBV11] Piernicola Bettiol, Alberto Bressan, and Richard B. Vinter. Estimates for trajectories confined to a cone in R^{n}. SIAM Journal on Control and Optimization, 49(1): 21–41, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i1/p21_s1.

Bettiol:2013:SNC

[BBV13] Piernicola Bettiol, Andrea Boccia, and Richard B. Vinter. Stratified necessary conditions for differential inclusions with state constraints. SIAM Journal on Control and Optimization, 51(5):3903–3917, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bardos:1981:CSW

[BC81] Claude Bardos and Goong Chen. Control and stabilization for the wave equation, Part III: Domain with moving boundary. SIAM Journal on Control and Optimization, 19(1):123–138, January 1981. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bonnans:1989:OCS

[BC89] Joseph Frédéric Bonnans and Eduardo Casas. Optimal control of semilinear multistate systems with state constraints. SIAM Journal on Control and Optimization, 27 (2):446–455, March 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bloch:1995:NCS

[BC95a] A. M. Bloch and P. E. Crouch.

Nonholonomic control systems on Riemannian manifolds. SIAM Journal on Control and Optimization, 33 (1):126–148, January 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/23533.

Bonnans:1995:EPP

[BC95b] Frédéric Bonnans and Eduardo Casas. An extension of Pontryagin’s principle for state-constrained optimal control of semilinear elliptic equations and variational inequalities. SIAM Journal on Control and Optimization, 33(1):274–298, January 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23777.

Baker:1996:ICC

[BC96a] C. R. Baker and I.-F. Chao. Information capacity of channels with partially unknown noise. II. infinite-dimensional channels. SIAM Journal on Control and Optimization, 34(4):1461–1472, July 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25296.

Bonnans:1996:POBa

[BC96b] J. Frédéric Bonnans and Roberto Cominetti. Perturbed optimization in Banach spaces I: a general theory based on a weak directional constraint qualification. SIAM Journal on Control and Optimization, 34(4):1151–1171, July 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26727.

Bonnans:1996:POBb

[BC96c] J. Frédéric Bonnans and Roberto Cominetti. Perturbed optimization in Banach spaces II: a theory based on a strong directional constraint qualification. SIAM Journal on Control and Optimization, 34(4):1172–1189, July 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26728.

Bonnans:1996:POBc

[BC96d] J. Frédéric Bonnans and Roberto Cominetti. Perturbed optimization in Banach spaces III: Semi-infinite optimization. SIAM Journal on Control and Optimization, 34(5):1555–1567, September 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27688.

Bloch:1998:NLI

[BC98] A. M. Bloch and P. E. Crouch.

Newton’s law and integrability of nonholonomic systems. SIAM Journal on Control and Optimization, 36(6): 2020–2039, November 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/29163.

Bittanti:2000:PEP

[BC00] Sergio Bittanti and M. C. Campi. Persistence of excitation properties for timevarying autoregressive systems. SIAM Journal on Control and Optimization, 39 (1):133–140, January 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/34348.

Bressan:2002:BCS

[BC02a] Alberto Bressan and Giuseppe Maria Coclite. On the boundary control of systems of conservation laws. SIAM Journal on Control and Optimization, 41(2):607–622, March 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/39252.

Bultheel:2002:ASP

[BC02b] A. Bultheel and P. Carrette. Algebraic and spectral properties of general Toeplitz matrices. SIAM Journal on Control and Optimization, 41(5): 1413–1439, September 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37718.

Bardi:2005:ASS

[BC05a] Martino Bardi and Annalisa Cesaroni. Almost sure stabilizability of controlled degenerate diffusions. SIAM Journal on Control and Optimization, 44(1):75–98, January 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43867.

Boscain:2005:TOS

[BC05b] Ugo Boscain and Yacine Chitour. Time-optimal synthesis for left-invariant control systems on series SO(3). SIAM Journal on Control and Optimization, 44(1):111–139, January 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/44153.

Biagini:2011:ASS

[BC11] Sara Biagini and Ales Cerný.

Admissible strategies in semimartingale portfolio selection. SIAM Journal on Control and Optimization, 49(1): 42–72, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i1/p42_s1.

Benedetti:2012:MUM

[BC12] Giuseppe Benedetti and Luciano Campi. Multivariate utility maximization with proportional transaction costs and random endowment. SIAM Journal on Control and Optimization, 50(3):1283– 1308, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bonfanti:2013:NLP

[BC13] Giovanni Bonfanti and Arrigo Cellina. The nonoccurrence of the Lavrentiev phenomenon for a class of variational functionals. SIAM Journal on Control and Optimization, 51 (2):1639–1650, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bot:2016:SOF

[BC16] Radu Ioan Bot and Ernö Robert Csetnek. Second order forward-backward dynamical systems for monotone inclusion problems. SIAM Journal on Control and Optimization, 54(3):1423–1443, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bo:2017:OIU

[BC17] Lijun Bo and Agostino Capponi. Optimal investment under information driven contagious distress. SIAM Journal on Control and Optimization, 55(2):1020–1068, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bayraktar:2018:AFS

[BC18a] Erhan Bayraktar and Asaf Cohen. Analysis of a finite state many player game using its master equation. SIAM Journal on Control and Optimization, 56(5):3538–3568, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bo:2018:PCM

[BC18b] Lijun Bo and Agostino Capponi. Portfolio choice with Market–Credit-Risk dependencies. SIAM Journal on Control and Optimization, 56 (4):3050–3091, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bui:2020:DRA

[BC20] Minh N. Bùi and Patrick L.

Combettes. The Douglas– Rachford algorithm converges only weakly. SIAM Journal on Control and Optimization, 58(2):1118–1120, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bhudisaksang:2021:ARC

[BC21] Theerawat Bhudisaksang and Álvaro Cartea. Adaptive robust control in continuous time. SIAM Journal on Control and Optimization, 59(5): 3912–3945, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bayraktar:2023:NAR

[BC23a] Erhan Bayraktar and Tao Chen. Nonparametric adaptive robust control under model uncertainty. SIAM Journal on Control and Optimization, 61(5):2737–2760, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1481609.

Brokate:2023:SSC

[BC23b] Martin Brokate and Constantin Christof. Strong stationarity conditions for optimal control problems governed by a rate-independent evolution variational inequality. SIAM Journal on Control and Optimization, 61 (4):2222–2250, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1494403.

Bravo:2024:SFP

[BC24] Mario Bravo and Roberto Cominetti. Stochastic fixedpoint iterations for nonexpansive maps: Convergence and error bounds. SIAM Journal on Control and Optimization, 62(1):191–219, January 2024. CODEN SJCODC. ISSN 1095-7138.

Bensoussan:2019:FSN

[BCC^{+}19a] Alain Bensoussan, Shaokuan Chen, Anshuman Chutani, Suresh P. Sethi, Chi Chung Siu, and Sheung Chi Phillip Yam. Feedback Stackelberg– Nash equilibria in mixed leadership games with an application to cooperative advertising. SIAM Journal on Control and Optimization, 57(5): 3413–3444, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bielecki:2019:ARC

[BCC^{+}19b] Tomasz R. Bielecki, Tao Chen, Igor Cialenco, Areski Cousin, and Monique Jeanblanc. Adaptive robust control under model uncertainty. SIAM Journal on Control and Optimization, 57(2):925–946, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ball:2002:ROS

[BCD02] Joseph A. Ball, Jerawan Chudoung, and Martin V. Day. Robust optimal switching control for nonlinear systems. SIAM Journal on Control and Optimization, 41(3):900–931, May 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37261.

Bullo:2012:GCC

[BCF12] Francesco Bullo, Ruggero Carli, and Paolo Frasca. Gossip coverage control for robotic networks: Dynamical systems on the space of partitions. SIAM Journal on Control and Optimization, 50(1):419–447, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bittanti:1984:COL

[BCG84] Sergio Bittanti, Patrizio Colaneri, and Guido Guardabassi. Hcontrollability and observability of linear periodic systems. SIAM Journal on Control and Optimization, 22(6): 889–893, November 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bittanti:1986:APL

[BCG86] Sergio Bittanti, Patrizio Colaneri, and Guido Guardabassi. Analysis of the periodic Lyapunov and Riccati equations via canonical decomposition. SIAM Journal on Control and Optimization, 24(6):1138– 1149, November 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Barbu:2017:SMC

[BCG^{+}17a] Viorel Barbu, Pierluigi Colli, Gianni Gilardi, Gabriela Marinoschi, and Elisabetta Rocca. Sliding mode control for a nonlinear phase-field system. SIAM Journal on Control and Optimization, 55(3): 2108–2133, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Blanchini:2017:AFI

[BCG^{+}17b] Franco Blanchini, Daniele Casagrande, Giulia Giordano, Stefano Miani, Sorin Olaru, and Vasso Reppa. Active fault isolation: a duality-based approach via convex programming. SIAM Journal on Control and Optimization, 55(3): 1619–1640, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Burachik:2024:OCD

[BCKM24] Regina S. Burachik, Bethany I.

Caldwell, C. Yalçin Kaya, and Walaa M. Moursi. Optimal control duality and the Douglas–Rachford algorithm. SIAM Journal on Control and Optimization, 62(1):680–698, February 2024. CODEN SJ-CODC. ISSN 1095-7138.

Barmish:1983:NCS

[BCL83] B. R. Barmish, M. Corless, and G. Leitmann. A new class of stabilizing controllers for uncertain dynamical systems. SIAM Journal on Control and Optimization, 21(2): 246–255, March 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Baxendale:2004:ASW

[BCL04] Peter Baxendale, Pavel Chigansky, and Robert Liptser. Asymptotic stability of the Wonham filter: Ergodic and nonergodic signals. SIAM Journal on Control and Optimization, 43(2):643–669, March 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/41692.

Bo:2022:LSM

[BCL22] Lijun Bo, Agostino Capponi, and Huafu Liao. Large sample mean-field stochastic optimization. SIAM Journal on Control and Optimization, 60(4):2538–2573, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1424937.

Bodnariu:2024:LTP

[BCL24] Andi Bodnariu, Sören Christensen, and Kristoffer Lindensjö. Local time pushed mixed equilibrium strategies for time-inconsistent stopping problems. SIAM Journal on Control and Optimization, 62(2):1261–1290, April 2024. CODEN SJCODC. ISSN 1095-7138.

Bensoussan:2017:LQM

[BCLY17] A. Bensoussan, M. H. M. Chau, Y. Lai, and S. C. P.

Yam. Linear-quadratic mean field Stackelberg games with state and control delays. SIAM Journal on Control and Optimization, 55(4):2748– 2781, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bonaccorsi:2012:OCS

[BCM12a] S. Bonaccorsi, F. Confortola, and E. Mastrogiacomo. Optimal control for stochastic Volterra equations with completely monotone kernels. SIAM Journal on Control and Optimization, 50(2):748–789, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bonfanti:2012:HIV

[BCM12b] Giovanni Bonfanti, Arrigo Cellina, and Marco Mazzola. The higher integrability and the validity of the Euler– Lagrange equation for solutions to variational problems. SIAM Journal on Control and Optimization, 50(2):888–899, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bonnard:2013:STC

[BCM13] Bernard Bonnard, Monique Chyba, and John Marriott. Singular trajectories and the contrast imaging problem in nuclear magnetic resonance. SIAM Journal on Control and Optimization, 51(2):1325– 1349, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Babiarz:2017:PPT

[BCM^{+}17] Artur Babiarz, Adam Czornik, Evgenii Makarov, Michal Niezabitowski, and Svetlana Popova. Pole placement theorem for discrete time-varying linear systems. SIAM Journal on Control and Optimization, 55(2):671–692, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bensoussan:2008:POI

[BÇMS^{+}08] Alain Bensoussan, Metin Çakanyildirim, J. Adolfo Minjárez-Sosa, Suresh P. Sethi, and Ruixia Shi. Partially observed inventory systems: The case of rain checks. SIAM Journal on Control and Optimization, 47(5):2490– 2519, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bullo:2009:SIC

[BCP09] Francesco Bullo, Jorge Cortés, and Benedetto Piccoli. Special issue on control and optimization in cooperative networks. SIAM Journal on Control and Optimization, 48 (1):vii, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bayraktar:2016:RFS

[BCP16] Erhan Bayraktar, Andrea Cosso, and Huyên Pham. Robust feedback switching control: Dynamic programming and viscosity solutions. SIAM Journal on Control and Optimization, 54(5):2594–2628, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bayraktar:2017:ERS

[BCP17] Erhan Bayraktar, Andrea Cossc, and Huyên Pham. Ergodicity of robust switching control and nonlinear system of quasi-variational inequalities. SIAM Journal on Control and Optimization, 55(3): 1915–1953, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Buckdahn:2004:NEP

[BCR04] Rainer Buckdahn, Pierre Cardaliaguet, and Catherine Rainer. Nash equilibrium payoffs for nonzerosum stochastic differential games. SIAM Journal on Control and Optimization, 43(2):624–642, March 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41155.

Bergault:2024:MFG

[BCR24] Philippe Bergault, Pierre Cardaliaguet, and Catherine Rainer. Mean field games in a Stackelberg problem with an informed major player. SIAM Journal on Control and Optimization, 62(3):1737–1765, June 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1615188.

Bensoussan:2007:POI

[BÇS07] Alain Bensoussan, Metin Çakanyildirim, and Suresh P. Sethi. Partially observed inventory systems: The case of zero-balance walk. SIAM Journal on Control and Optimization, 46(1):176–209, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bressan:2013:MOH

[BCS13] Alberto Bressan, Giuseppe Maria Coclite, and Wen Shen. A multidimensional optimalharvesting problem with measurevalued solutions. SIAM Journal on Control and Optimization, 51(2):1186–1202, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bensoussan:2015:MPG

[BCS15] Alain Bensoussan, Shaokuan Chen, and Suresh P. Sethi. The maximum principle for global solutions of stochastic Stackelberg differential games. SIAM Journal on Control and Optimization, 53(4):1956– 1981, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Belak:2017:GVR

[BCS17] Christoph Belak, Sören Christensen, and Frank Thomas Seifried. A general verification result for stochastic impulse control problems. SIAM Journal on Control and Optimization, 55(2):627–649, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bayraktar:2018:MOT

[BCS18] Erhan Bayraktar, Alexander M. G. Cox, and Yavor Stoev. Martingale optimal transport with stopping. SIAM Journal on Control and Optimization, 56(1):417–433, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Boarotto:2020:FSG

[BCS20] Francesco Boarotto, Yacine Chitour, and Mario Sigalotti. Fuller singularities for generic control-affine systems with an even number of controls. SIAM Journal on Control and Optimization, 58(2): 1207–1228, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Babiarz:2021:ESD

[BCS21a] Artur Babiarz, Adam Czornik, and Stefan Siegmund. Erratum: On stabilization of discrete time-varying systems. SIAM Journal on Control and Optimization, 59(3): 2410, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [BCS21b].

Babiarz:2021:SDT

[BCS21b] Artur Babiarz, Adam Czornik, and Stefan Siegmund. On stabilization of discrete timevarying systems. SIAM Journal on Control and Optimization, 59(1):242–266, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See erratum [BCS21a].

Buchwald:2021:AGR

[BCS21c] S. Buchwald, G. Ciaramella, and J. Salomon. Analysis of a greedy reconstruction algorithm. SIAM Journal on Control and Optimization, 59(6): 4511–4537, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Buchwald:2024:GNO

[BCS24] Simon Buchwald, Gabriele Ciaramella, and Julien Salomon. Gauss–newton oriented greedy algorithms for the reconstruction of operators in nonlinear dynamics. SIAM Journal on Control and Optimization, 62(3):1343–1368, June 2024. CODEN SJCODC. ISSN 1095-7138.

Buttazzo:2010:SSC

[BCT10] G. Buttazzo, G. Carlier, and R. Tahraoui. On some systems controlled by the structure of their memory. SIAM Journal on Control and Optimization, 48(8):5241–5253, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bao:2021:GTM

[BCW21] Wenqing Bao, Xiaoqiang Cai, and Xianyi Wu. A general theory of multiarmed bandit processes with constrained arm switches. SIAM Journal on Control and Optimization, 59 (6):4666–4688, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bensoussan:2015:MFS

[BCY15] A. Bensoussan, M. H. M.

Chau, and S. C. P. Yam. Mean field Stackelberg games: Aggregation of delayed instructions. SIAM Journal on Control and Optimization, 53(4): 2237–2266, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bradley:1977:SAM

[BD77] Mary Bradley and Richard Datko. Some analytic and measure theoretic properties of set-valued mappings. SIAM Journal on Control and Optimization, 15(4):625–635, July 1977. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Barbu:1983:SBE

[BD83] Viorel Barbu and Giuseppe Da Prato. Solution of the Bellman equation associated with an infinite dimensional stochastic control problem and synthesis of optimal control. SIAM Journal on Control and Optimization, 21(4):531–550, July 1983. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bonnard:1995:TGT

[Bd95] B. Bonnard and J. de Morant. Toward a geometric theory in the time-minimal control of chemical batch reactors. SIAM Journal on Control and Optimization, 33(5): 1279–1311, September 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/24133.

Badea:2001:BCA

[BD01a] L. Badea and P. Daripa. On a boundary control approach to domain embedding methods. SIAM Journal on Control and Optimization, 40(2):421–449, March 2001. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35738.

Bourdon:2001:URP

[BD01b] Sean E. Bourdon and Geir E.

Dullerud. Uniform robust performance against quasi-LTI uncertainty in sampleddata systems. SIAM Journal on Control and Optimization, 40(1):298–327, January 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36677.

Battilotti:2002:SPN

[BD02] S. Battilotti and A. De Santis. Stabilization in probability of nonlinear stochastic systems with guaranteed cost. SIAM Journal on Control and Optimization, 40(6): 1938–1964, November 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37502.

Barbu:2011:ISN

[BD11a] Viorel Barbu and Giuseppe Da Prato. Internal stabilization by noise of the Navier–Stokes equation. SIAM Journal on Control and Optimization, 49 (1):1–20, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.

                                        siam.org/sicon/resource/
                                        1/sjcodc/v49/i1/p1_s1.

Benner:2011:LEE

[BD11b] Peter Benner and Tobias Damm. Lyapunov equations, energy functionals, and model order reduction of bilinear and stochastic systems. SIAM Journal on Control and Optimization, 49(2): 686–711, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicon/resource/
                                        1/sjcodc/v49/i2/p686_s1.

Bailleul:2015:IPR

[BD15] I. Bailleul and J. Diehl. The inverse problem for rough controlled differential equations. SIAM Journal on Control and Optimization, 53(5): 2762–2780, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bishop:2017:SKB

[BD17a] Adrian N. Bishop and Pierre Del Moral. On the stability of Kalman–Bucy diffusion processes. SIAM Journal on Control and Optimization, 55(6): 4015–4047, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Boskos:2017:RIC

[BD17b] Dimitris Boskos and Dimos V. Dimarogonas. Robustness and invariance of connectivity maintenance control for multiagent systems. SIAM Journal on Control and Optimization, 55(3):1887–1914, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bishop:2019:SPS

[BD19a] Adrian N. Bishop and Pierre Del Moral. Stability properties of systems of linear stochastic differential equations with random coefficients. SIAM Journal on Control and Optimization, 57(2):1023–1042, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Boskos:2019:AVD

[BD19b] Dimitris Boskos and Dimos V.

Dimarogonas. Abstractions of varying decentralization degree for reachability of coupled multiagent systems. SIAM Journal on Control and Optimization, 57(5):3471–3495, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bishop:2021:NCW

[BD21] Adrian N. Bishop and Arnaud Doucet. Network consensus in the Wasserstein metric space of probability measures. SIAM Journal on Control and Optimization, 59(5): 3261–3277, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Broucke:2005:ESO

[BDDSV05] Mireille Broucke, Maria Domenica Di Benedetto, Stefano Di Gennaro, and Alberto Sangiovanni-Vincentelli. Efficient solution of optimal control problems using hybrid systems. SIAM Journal on Control and Optimization, 43(6): 1923–1952, November 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38309.

Boldrini:2012:SCR

[BDFCGB12] J. L. Boldrini, A. Doubova, E. Fernández-Cara, and M. González-Burgos. Some controllability results for linear viscoelastic fluids. SIAM Journal on Control and Optimization, 50(2): 900–924, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bolten:2021:TLP

[BDGK21] Matthias Bolten, Onur Tanil Doganay, Hanno Gottschalk, and Kathrin Klamroth. Tracing locally Pareto-optimal points by numerical integration. SIAM Journal on Control and Optimization, 59(5): 3302–3328, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Boscain:2012:AIR

[BDGR12] Ugo Boscain, Jean Duplaix, Jean-Paul Gauthier, and Francesco Rossi. Anthropomorphic image reconstruction via hypoelliptic diffusion. SIAM Journal on Control and Optimization, 50(3): 1309–1336, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Berger:2022:FMP

[BDIW22] Thomas Berger, Dario Dennstädt, Achim Ilchmann, and Karl Worthmann. Funnel model predictive control for nonlinear systems with relative degree one. SIAM Journal on Control and Optimization, 60 (6):3358–3383, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1431655.

Bloch:2000:SNS

[BDK00] Anthony M. Bloch, Sergey V.

Drakunov, and Michael K. Kinyon. Stabilization of nonholonomic systems using isospectral flows. SIAM Journal on Control and Optimization, 38(3):855–874, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33560.

Berger:2024:RFM

[BDLW24] Thomas Berger, Dario Dennstädt, Lukas Lanza, and Karl Worthmann. Robust funnel model predictive control for output tracking with prescribed performance. SIAM Journal on Control and Optimization, 62 (4):2071–2097, August 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1551195.

Bahlali:2007:RSM

[BDM07] Seid Bahlali, Boualem Djehiche, and Brahim Mezerdi. The relaxed stochastic maximum principle in singular optimal control of diffusions. SIAM Journal on Control and Optimization, 46(2):427–444, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bonnans:2014:SON

[BDP14] J. Frédéric Bonnans, Xavier Dupuis, and Laurent Pfeiffer. Second-order necessary conditions in Pontryagin form for optimal control problems. SIAM Journal on Control and Optimization, 52(6):3887– 3916, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Becerril:2021:OCN

[BdP21] J. A. Becerril and M. D. R. de Pinho. Optimal control with nonregular mixed constraints: an optimization approach. SIAM Journal on Control and Optimization, 59 (3):2093–2120, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bovo:2024:ZSS

[BDP24] Andrea Bovo, Tiziano De Angelis, and Jan Palczewski. Zero-sum stopper versus singularcontroller games with constrained control directions. SIAM Journal on Control and Optimization, 62(4):2203– 2228, August 2024. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1579558.

Boccia:2016:OCP

[BdPV16] A. Boccia, M. D. R. de Pinho, and R. B. Vinter. Optimal control problems with mixed and pure state constraints. SIAM Journal on Control and Optimization, 54 (6):3061–3083, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Barles:1994:OCN

[BDR94] Guy Barles, Christian Daher, and Marc Romano. Optimal control on the L^{∞} norm of a diffusion process. SIAM Journal on Control and Optimization, 32(3):612–634, May 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/22359.

Bao:2002:CMR

[BDR02] Gang Bao, David C. Dobson, and Karim Ramdani. A constraint on the maximum reflectance of rapidly oscillating dielectric gratings. SIAM Journal on Control and Optimization, 40(6): 1858–1866, November 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/37435.

Bohm:1998:MRA

[BDRR98] Michael Böhm, M. A. Demetriou, Simeon Reich, and I. G. Rosen. Model reference adaptive control of distributed parameter systems. SIAM Journal on Control and Optimization, 36(1):33–81, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/27971.

Bouveret:2020:MFG

[BDT20] Géraldine Bouveret, Roxana Dumitrescu, and Peter Tankov. Mean-field games of optimal stopping: a relaxed solution approach. SIAM Journal on Control and Optimization, 58(4):1795–1821, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bertsekas:1987:DAR

[BE87] Dimitri P. Bertsekas and Didier El Baz. Distributed asynchronous relaxation methods for convex network flow problems. SIAM Journal on Control and Optimization, 25(1): 74–85, January 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bensoussan:1995:FDR

[BE95] Alain Bensoussan and Robert J.

Elliott. A finite-dimensional risk-sensitive control problem. SIAM Journal on Control and Optimization, 33(6): 1834–1846, November 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25587.

Baudouin:2013:CIP

[BE13] Lucie Baudouin and Sylvain Ervedoza. Convergence of an inverse problem for a 1-D discrete wave equation. SIAM Journal on Control and Optimization, 51(1):556–598, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Beauchard:2008:LCO

[Bea08]  Karine Beauchard.  Lo-

cal controllability of a onedimensional beam equation. SIAM Journal on Control and Optimization, 47(3):1219– 1273, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Behar:1979:MOS

[Beh79] Isak Behar. Modifications in optimization and saddle point problems. SIAM Journal on Control and Optimization, 17 (6):673–688, November 1979. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bouchard:2010:OCU

[BEI10] Bruno Bouchard, Romuald Elie, and Cyril Imbert. Optimal control under stochastic target constraints. SIAM Journal on Control and Optimization, 48(5):3501–3531, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Baras:1989:POS

[BEK89] John S. Baras, Robert J. Elliott, and Michael Kohlmann. The partially observed stochastic minimum principle. SIAM Journal on Control and Optimization, 27(6):1279–1292, November 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Belmiloudi:2000:ECO

[Bel00] Aziz Belmiloudi. Existence and characterization of an optimal control for the problem of long waves in a shallowwater model. SIAM Journal on Control and Optimization, 39(5):1558–1584, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33596.

Byrnes:2002:IWP

[BEL02] Christopher I. Byrnes, Per Enqvist, and Anders Lindquist. Identifiability and well-posedness of shaping-filter parameterizations: a global analysis approach. SIAM Journal on Control and Optimization, 41(1):23–59, January 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38307.

Belabbas:2023:GFI

[Bel23] Mohamed-Ali Belabbas. On gradient flows initialized near maxima. SIAM Journal on Control and Optimization, 61 (5):2826–2848, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1450756.

Bayraktar:2013:ICJ

[BEM13] Erhan Bayraktar, Thomas Emmerling, and José-Luis Menaldi. On the impulse control of jump diffusions. SIAM Journal on Control and Optimization, 51(3):2612–2637, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Benes:1976:FBR

[Ben76] V. E. Benes. Full “Bang” to reduce predicted miss is optimal. SIAM Journal on Control and Optimization, 14(1): 62–84, January 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Benchimol:1978:NWS

[Ben78] Claude D. Benchimol. A note on weak stabilizability of contraction semigroups. SIAM Journal on Control and Optimization, 16(3):373–379, May 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bendsoe:1981:EOS

[Ben81] Martin Philip Bendsøe. On the existence of observable singleoutput systems of a simple type. SIAM Journal on Control and Optimization, 19(4): 555–559, July 1981. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Benaim:1996:DSA

[Ben96] Michel Benaim. A dynamical system approach to stochastic approximations. SIAM Journal on Control and Optimization, 34(2):437–472, March 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25353.

Bertsekas:1976:PMM

[Ber76] Dimitri P. Bertsekas. On penalty and multiplier methods for constrained minimization. SIAM Journal on Control and Optimization, 14(2): 216–235, February 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bertsekas:1977:MMA

[Ber77] Dimitri P. Bertsekas. Monotone mappings with application in dynamic programming. SIAM Journal on Control and Optimization, 15(3):438–464, May 1977. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bernhard:1982:SIL

[Ber82a] Pierre Bernhard. On singular implicit linear dynamical systems. SIAM Journal on Control and Optimization, 20(5):612–633, September 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bertsekas:1982:PNM

[Ber82b] Dimitri P. Bertsekas. Projected Newton methods for optimization problems with simple constraints. SIAM Journal on Control and Optimization, 20(2):221–246, March 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bernstein:1984:SAH

[Ber84] Dennis S. Bernstein. A systematic approach to higherorder necessary conditions in optimization theory. SIAM Journal on Control and Optimization, 22(2):211–238, March 1984. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Berkovitz:1985:EVS

[Ber85] Leonard D. Berkovitz. The existence of value and saddle point in games of fixed duration. SIAM Journal on Control and Optimization, 23(2): 172–196, March 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See erratum and addendum [Ber88].

Berkovitz:1986:DGG

[Ber86] Leonard D. Berkovitz. Differential games of generalized pursuit and evasion. SIAM Journal on Control and Optimization, 24(3):361–373, May 1986. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Berkovitz:1988:EAE

[Ber88] Leonard D. Berkovitz. Erratum and addendum: “The existence of value and saddle point in games of fixed duration” [SIAM J. Control Optim. 23 (1985), no. 2, 172–196; MR 86g:90141]. SIAM Journal on Control and Optimization, 26 (3):740–742, May 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [Ber85].

Berkovitz:1989:OFC

[Ber89a] Leonard D. Berkovitz. Optimal feedback controls. SIAM Journal on Control and Optimization, 27(5):991–1006, September 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bernstein:1989:RSP

[Ber89b] Dennis S. Bernstein. Robust stability and performance via fixed-order dynamic compensation. SIAM Journal on Control and Optimization, 27(2): 389–406, March 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bergounioux:1992:PMO

[Ber92a] Maïtine Bergounioux. A penalization method for optimal control of elliptic problems with state constraints.

SIAM Journal on Control and Optimization, 30(2):305– 323, March 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bernhard:1992:ISD

[Ber92b] P. Bernhard. Information and strategies in dynamic games. SIAM Journal on Control and Optimization, 30(1): 212–228, January 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bercu:1995:WET

[Ber95] B. Bercu. Weighted estimation and tracking for AR-MAX models. SIAM Journal on Control and Optimization, 33(1):89–106, January 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/22180.

Bercu:1998:CLT

[Ber98a] Bernard Bercu. Central Limit Theorem and law of iterated logarithm for least squares algorithms in adaptive tracking. SIAM Journal on Control and Optimization, 36(3):910–928, May 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29418.

Bergounioux:1998:OCP

[Ber98b] Maïtine Bergounioux. Optimal control of problems governed by abstract elliptic variational inequalities with state constraints. SIAM Journal on Control and Optimization, 36(1):273–289, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30261.

Bertsekas:1998:NVI

[Ber98c]  Dimitri P. Bertsekas.  A

new value iteration method for the average cost dynamic programming problem. SIAM Journal on Control and Optimization, 36(2):742–759, March 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29160.

Berrahmoune:2002:SBN

[Ber02] Larbi Berrahmoune. Stabilization of beams with nonlinear feedback. SIAM Journal on Control and Optimization, 41(4):1166–1193, July 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38852.

Bertsekas:2018:SOC

[Ber18] Dimitri P. Bertsekas. Stable optimal control and semicontractive dynamic programming. SIAM Journal on Control and Optimization, 56(1): 231–252, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Berger:2021:FCF

[Ber21] Thomas Berger. Funnel control of the Fokker–Planck equation for a multidimensional Ornstein–Uhlenbeck process. SIAM Journal on Control and Optimization, 59 (5):3203–3230, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bernhard:2007:RCA

[BET07] Pierre Bernhard, Naïma El Farouq, and Stéphane Thiery. Robust control approach to option pricing: a representation theorem and fast algorithm. SIAM Journal on Control and Optimization, 46(6): 2280–2302, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bouchard:2010:STP

[BET10] Bruno Bouchard, Romuald Elie, and Nizar Touzi. Stochastic target problems with controlled loss. SIAM Journal on Control and Optimization, 48(5):3123–3150, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Betz:2019:SOS

[Bet19]  Livia M. Betz.  Second-

order sufficient optimality conditions for optimal control of nonsmooth, semilinear parabolic equations. SIAM Journal on Control and Optimization, 57(6):4033–4062, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Berkey:1976:PPC

[BF76a] Dennis D. Berkey and Marvin I. Freedman. A perturbation problem from control exhibiting on and off the boundary behavior. SIAM Journal on Control and Optimization, 14(5):857–876, August 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Brockett:1976:NSD

[BF76b] Roger W. Brockett and Paul A. Fuhrmann. Normal symmetric dynamical systems. SIAM Journal on Control and Optimization, 14(1): 107–119, January 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Berkovitz:1989:TJM

[BF89a] L. D. Berkovitz and Wendell H. Fleming. A tribute to E. J. McShane. SIAM Journal on Control and Optimization, 27(5):909–915, September 1989. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Brokate:1989:ODH

[BF89b] Martin Brokate and Avner Friedman. Optimal design for heat conduction problems with hysteresis. SIAM Journal on Control and Optimization, 27(4):697–717, July 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bardi:1990:ASM

[BF90a] M. Bardi and M. Falcone. An approximation scheme for the minimum time function. SIAM Journal on Control and Optimization, 28(4):950–965, July 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bisiacco:1990:OCT

[BF90b] Mauro Bisiacco and Ettore Fornasini. Optimal control of two-dimensional systems. SIAM Journal on Control and Optimization, 28(3):582–601, May 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Barbu:1991:ODD

[BF91] Viorel Barbu and Avner Friedman. Optimal design of domains with free-boundary problems. SIAM Journal on Control and Optimization, 29 (3):623–637, May 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Burke:1993:WSM

[BF93] J. V. Burke and M. C. Ferris. Weak sharp minima in mathematical programming. SIAM Journal on Control and Optimization, 31(5):1340– 1359, September 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bian:2005:GTG

[BF05] W. Bian and M. French.

Graph topologies, gap metrics, and robust stability for nonlinear systems. SIAM Journal on Control and Optimization, 44(2):418–443, March 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42120.

Bouchitte:2007:ODT

[BF07] Guy Bouchitté and Ilaria Fragalà. Optimal design of thin plates by a dimension reduction for linear constrained problems. SIAM Journal on Control and Optimization, 46 (5):1664–1682, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bressan:2013:BGC

[BF13] Alberto Bressan and Giancarlo Facchi. A bidding game in a continuum limit order book. SIAM Journal on Control and Optimization, 51(5): 3459–3485, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bravo:2015:RLR

[BF15] Mario Bravo and Mathieu Faure. Reinforcement learning with restrictions on the action set. SIAM Journal on Control and Optimization, 53 (1):287–312, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bet:2022:ACS

[BF22] Gianmarco Bet and Markus Fischer. An algorithm to construct subsolutions of convex optimal control problems. SIAM Journal on Control and Optimization, 60 (5):2902–2926, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1402005.

Bello:1997:DDR

[BFCLS97] Juan Antonio Bello, Enrique Fernández-Cara, Jérôme Lemoine, and Jacques Simon. The differentiability of the drag with respect to the variations of a Lipschitz domain in a Navier– Stokes flow. SIAM Journal on Control and Optimization, 35(2):626–640, March 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27821.

Bullo:2020:FTI

[BFF20] Francesco Bullo, Fabio Fagnani, and Barbara Franci. Finite-time influence systems and the wisdom of crowd effect. SIAM Journal on Control and Optimization, 58(2): 636–659, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bouchitte:2016:VMS

[BFL16] Guy Bouchitté, Ilaria Fragalà, and Ilaria Lucardesi. A variational method for second order shape derivatives. SIAM Journal on Control and Optimization, 54(2):1056–1084, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Boutaayamou:2023:BCD

[BFM23] Idriss Boutaayamou, Genni Fragnelli, and Dimitri Mugnai. Boundary controllability for a degenerate wave equation in nondivergence form with drift. SIAM Journal on Control and Optimization, 61(4):1934–1954, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M151491X.

Burman:2020:FDN

[BFO20] Erik Burman, Ali Feizmohammadi, and Lauri Oksanen. A fully discrete numerical control method for the wave equation. SIAM Journal on Control and Optimization, 58(3): 1519–1546, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bian:2008:IBA

[BFP08] Wenming Bian, Mark French, and Harish K. Pillai. An intrinsic behavioral approach to the gap metric. SIAM Journal on Control and Optimization, 47(4):1939–1960, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bejarano:2009:UIS

[BFP09] Francisco J. Bejarano, Leonid Fridman, and Alexander Poznyak. Unknown input and state estimation for unobservable systems. SIAM Journal on Control and Optimization, 48(2):1155–1178, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bonnet:2011:SNS

[BFP11] Catherine Bonnet, André R. Fioravanti, and Jonathan R. Partington. Stability of neutral systems with commensurate delays and poles asymptotic to the imaginary axis.

SIAM Journal on Control and Optimization, 49(2):498– 516, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i2/p498_s1.

Bian:2013:SNF

[BFPB13] W. Bian, M. French, and A. Prügel-Bennett. Stability for network flow control against disturbances and time delays. SIAM Journal on Control and Optimization, 51(2): 1063–1098, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Banas:2021:NAV

[BFR21] Lubomír Banas, Giorgio Ferrari, and Tsiry A. Randrianasolo. Numerical approximation of the value of a stochastic differential game with asymmetric information. SIAM Journal on Control and Optimization, 59(2):1109–1135, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bercovici:1996:SSV

[BFT96] Hari Bercovici, Ciprian Foiaş, and Allen Tannenbaum. The structured singular value for linear input/output operators. SIAM Journal on Control and Optimization, 34(4):1392–1404, July 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/26882.

Baragana:2000:LSP

[BFZ00] I. Baragaña, V. Fernández, and I. Zaballa. Linear systems with prescribed similarity structural invariants. SIAM Journal on Control and Optimization, 38(4):1033–1049, July 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/33782.

Bokanowski:2010:RMT

[BFZ10] Olivier Bokanowski, Nicolas Forcadel, and Hasnaa Zidani. Reachability and minimal times for state constrained nonlinear problems without any controllability assumption. SIAM Journal on Control and Optimization, 48(7): 4292–4316, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Borkar:1988:ECM

[BG88] Vivek S. Borkar and Mrinal K. Ghosh. Ergodic control of multidimensional diffusions. I. The existence results. SIAM Journal on Control and Optimization, 26(1):112–126, January 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Baston:1998:RLW

[BG98] Vic Baston and Shmuel Gal.

Rendezvous on the line when the players’ initial distance is given by an unknown probability distribution. SIAM Journal on Control and Optimization, 36(6):1880–1889, November 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31413.

Borkar:2005:ELO

[BG05] Vivek Borkar and Vladimir Gaitsgory. On existence of limit occupational measures set of a controlled stochastic differential equation. SIAM Journal on Control and Optimization, 44(4):1436–1473, July 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/44347.

Borkar:2007:SPE

[BG07] Vivek S. Borkar and Vladimir Gaitsgory. Singular perturbations in ergodic control of diffusions. SIAM Journal on Control and Optimization, 46 (5):1562–1577, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Benkherouf:2009:COP

[BG09] Lakdere Benkherouf and Brian H. Gilding. On a class of optimization problems for finite time horizon inventory models. SIAM Journal on Control and Optimization, 48 (2):993–1030, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Boscain:2013:SHM

[BG13] Ugo Boscain and Jean-Paul Gauthier. On the spherical Hausdorff measure in step 2 corank 2 sub-Riemannian geometry. SIAM Journal on Control and Optimization, 51 (6):4450–4462, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Broucke:2014:RCS

[BG14] Mireille E. Broucke and Marcus Ganness. Reach control on simplices by piecewise affine feedback. SIAM Journal on Control and Optimization, 52 (5):3261–3286, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bonnans:2020:OCS

[BG20] J. Frédéric Bonnans and Justina Gianatti. Optimal control of state constrained age-structured problems. SIAM Journal on Control and Optimization, 58(4): 2206–2235, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bonnans:2021:EOC

[BG21] J. F. Bonnans and Justina Gianatti. Erratum: Optimal control of state constrained age-structured problems. SIAM Journal on Control and Optimization, 59(2): 1756, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bodart:2004:LRI

[BGBPG04] Olivier Bodart, Manuel González-Burgos, and Rosario Pérez-García. A local result on insensitizing controls for a semilinear heat equation with nonlinear boundary Fourier conditions. SIAM Journal on Control and Optimization, 43(3):955–969, May 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43161.

Byrnes:1994:RLB

[BGH94] Christopher I. Byrnes, David S.

Gilliam, and Jianqiu He. Root-locus and boundary feedback design for a class of distributed parameter systems. SIAM Journal on Control and Optimization, 32(5): 1364–1427, September 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/22262.

Blank:2016:SIL

[BGHR16] Luise Blank, Harald Garcke, Claudia Hecht, and Christoph Rupprecht. Sharp interface limit for a phase field model in structural optimization. SIAM Journal on Control and Optimization, 54(3):1558–1584, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bart:1980:FTF

[BGKV80] H. Bart, I. Gohberg, M. A. Kaashoek, and P. Van Dooren. Factorizations of transfer functions. SIAM Journal on Control and Optimization, 18(6): 675–696, November 1980. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Byrnes:1998:COA

[BGL98] Christopher I. Byrnes, Sergei V. Gusev, and Anders Lindquist. A convex optimization approach to the rational covariance extension problem. SIAM Journal on Control and Optimization, 37(1):211–229, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/32155.

Budhiraja:2011:ERC

[BGL11] Amarjit Budhiraja, Arka P. Ghosh, and Chihoon Lee. Ergodic rate control problem for single class queueing networks. SIAM Journal on Control and Optimization, 49(4): 1570–1606, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i4/p1570_s1.

Bill:2016:SNL

[BGLT16] Adam Bill, Chris Guiver, Hartmut Logemann, and Stuart Townley. Stability of Non-Negative Lur’e systems. SIAM Journal on Control and Optimization, 54(3):1176–1211, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Byers:1997:DSC

[BGM97] Ralph Byers, Ton Geerts, and Volker Mehrmann. Descriptor systems without controllability at infinity. SIAM Journal on Control and Optimization, 35(2):462–479, March 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26981.

Ball:2005:SNM

[BGM05] Joseph A. Ball, Gilbert Groenewald, and Tanit Malakorn. Structured noncommutative multidimensional linear systems. SIAM Journal on Control and Optimization, 44(4):1474–1528, July 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/44375.

Blondel:1994:SST

[BGMR94] V. Blondel, M. Gevers, R. Mortini, and R. Rupp. Simultaneous stabilization of three or more plants: Conditions on the positive real axis do not suffice. SIAM Journal on Control and Optimization, 32(2):572–590, March 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/21881.

Bittanti:1978:PSC

[BGMS78] S. Bittanti, G. Guardabassi, C. Maffezzoni, and L. Silverman. Periodic systems: Controllability and the matrix Riccati equation. SIAM Journal on Control and Optimization, 16(1):37–40, January 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bauso:2009:CNU

[BGP09] D. Bauso, L. Giarré, and R. Pesenti. Consensus for networks with unknown but bounded disturbances. SIAM Journal on Control and Optimization, 48(3):1756–1770, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bercu:2020:ESN

[BGP20a] Bernard Bercu, Antoine Godichon, and Bruno Portier. An efficient stochastic Newton algorithm for parameter estimation in logistic regressions. SIAM Journal on Control and Optimization, 58(1):348–367, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Biffis:2020:OPC

[BGP20b] Enrico Biffis, Fausto Gozzi, and Cecilia Prosdocimi. Optimal portfolio choice with path dependent labor income: the infinite horizon case. SIAM Journal on Control and Optimization, 58(4):1906–1938, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bonnans:2023:LAA

[BGP23] J. Frédéric Bonnans, Justina Gianatti, and Laurent Pfeiffer. A Lagrangian approach for aggregative mean field games of controls with mixed and final constraints. SIAM Journal on Control and Optimization, 61(1):105–134, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1407720.

Bensoussan:1990:AZE

[BGR90] A. Bensoussan, R. Glowinski, and A. Răşcanu. Approximation of the Zakaï equation by the splitting up method. SIAM Journal on Control and Optimization, 28(6):1420– 1431, November 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bermudez:2004:FEM

[BGR04] Alfredo Bermúdez, Pablo Gamallo, and Rodolfo Rodríguez. Finite element methods in local active control of sound. SIAM Journal on Control and Optimization, 43(2):437– 465, March 2004. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43178.

Benabdallah:2007:SDD

[BGR07] Assia Benabdallah, Patricia Gaitan, and Jérôme Le Rousseau. Stability of discontinuous diffusion coefficients and initial conditions in an inverse problem for the heat equation. SIAM Journal on Control and Optimization, 46 (5):1849–1881, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Banks:1988:IDP

[BGRW88] H. T. Banks, S. S. Gates, I. G. Rosen, and Y. Wang. The identification of a distributed parameter model for a flexible structure. SIAM Journal on Control and Optimization, 26 (4):743–762, July 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Borkar:2019:LFD

[BGS19] Vivek S. Borkar, Vladimir Gaitsgory, and Ilya Shvartsman. LP formulations of discrete time long-run average optimal control problems: The nonergodic case. SIAM Journal on Control and Optimization, 57(3):1783–1817, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Brivadis:2021:AOS

[BGSS21] Lucas Brivadis, Jean-Paul A.

Gauthier, Ludovic Sacchelli, and Ulysse Serres. Avoiding observability singularities in output feedback bilinear systems. SIAM Journal on Control and Optimization, 59(3): 1759–1780, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bombach:2023:ONA

[BGST23] Clemens Bombach, Fabian Gabel, Christian Seifert, and Martin Tautenhahn. Observability for non-autonomous systems. SIAM Journal on Control and Optimization, 61(1):315–341, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1485139.

Bernstein:1986:OPE

[BH86] Dennis S. Bernstein and David C. Hyland. The optimal projection equations for finite-dimensional fixedorder dynamic compensation of infinite-dimensional systems. SIAM Journal on Control and Optimization, 24(1): 122–151, January 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Banaszuk:1996:AFL

[BH96] Andrzej Banaszuk and John Hauser. Approximate feedback linearization: a homotopy operator approach. SIAM Journal on Control and Optimization, 34(5): 1533–1554, September 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26130.

Berg:1999:PEU

[BH99] Jordan M. Berg and Kenneth Holmström. On parameter estimation using level sets. SIAM Journal on Control and Optimization, 37(5): 1372–1393, September 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33634.

Birnir:2000:BCV

[BH00] Björn Birnir and Höskuldur Ari Hauksson. Basic control for the viscous Moore–Greitzer partial differential equation. SIAM Journal on Control and Optimization, 38(5): 1554–1580, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34518.

Bonnans:2007:WPS

[BH07] J. Frédéric Bonnans and Audrey Hermant. Well-posedness of the shooting algorithm for state constrained optimal control problems with a single constraint and control. SIAM Journal on Control and Optimization, 46(4):1398–1430, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bayraktar:2013:MCS

[BH13] Erhan Bayraktar and Yu-Jui Huang. On the multidimensional controller-and-stopper games. SIAM Journal on Control and Optimization, 51(2):1263–1297, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bouvrie:2017:KMA

[BH17] Jake Bouvrie and Boumediene Hamzi. Kernel methods for the approximation of nonlinear systems. SIAM Journal on Control and Optimization, 55(4):2460–2492, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Brunovsky:2018:OCD

[BH18] Pavol Brunovský and Mária Holecyová. Optimal control, dichotomy, and closed range. SIAM Journal on Control and Optimization, 56(2):633–648, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Brugger:2022:RCS

[BH22] Rahel Brügger and Helmut Harbrecht. On the reformulation of the classical Stefan problem as a shape optimization problem. SIAM Journal on Control and Optimization, 60(1):310–329, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1411007.

Breiten:2023:AKB

[BH23] Tobias Breiten and Bernhard Höveler. On the approximability of koopman-based operator Lyapunov equations. SIAM Journal on Control and Optimization, 61(5):3131–3155, October 2023. CODEN SJ-CODC. ISSN 1095-7138.

Boue:2003:LDR

[BHHE03] Michelle Boué, Daniel Hernández-Hernández, and Richard S.

Ellis. Large deviations for a random walk model with state-dependent noise. SIAM Journal on Control and Optimization, 42(3):810–838, May 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39661.

Banks:1983:M

[BHJ83a] H. T. Banks, H. G. Hermes, and M. Q. Jacobs. In memoriam. SIAM Journal on Control and Optimization, 21(6): vii–ix, ???? 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Banks:1983:MJP

[BHJ83b] H. T. Banks, H. G. Hermes, and M. Q. Jacobs. In memoriam: J. P. LaSalle. SIAM Journal on Control and Optimization, 21(6):vii–ix, November 1983. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Boyer:2010:DCE

[BHL10] Franck Boyer, Florence Hubert, and Jérôme Le Rousseau. Discrete Carleman estimates for elliptic operators in arbitrary dimension and applications. SIAM Journal on Control and Optimization, 48(8): 5357–5397, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Buckdahn:2012:RPG

[BHL12] Rainer Buckdahn, Jianhui Huang, and Juan Li. Regularity properties for general HJB equations: a backward stochastic differential equation method. SIAM Journal on Control and Optimization, 50(3):1466–1501, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Banaszuk:1999:BCN

[BHM99] Andrzej Banaszuk, Höskuldur Ari Hauksson, and Igor Mezic. A backstepping controller for a nonlinear partial differential equation model of compression system instabilities. SIAM Journal on Control and Optimization, 37(5): 1503–1537, September 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/31787.

Bauschke:2014:GSS

[BHM14] Heinz H. Bauschke, Warren L. Hare, and Walaa M. Moursi. Generalized solutions for the sum of two maximally monotone operators. SIAM Journal on Control and Optimization, 52(2):1034–1047, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Beremlijski:2014:SOC

[BHOP14] P. Beremlijski, J. Haslinger, J. V. Outrata, and R. Pathó. Shape optimization in contact problems with Coulomb friction and a solution-dependent friction coefficient. SIAM Journal on Control and Optimization, 52(5):3371–3400, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bachelier:2004:RRC

[BHPM04] Olivier Bachelier, Didier Henrion, Bernard Pradin, and Driss Mehdi. Robust rootclustering of a matrix in intersections or unions of regions. SIAM Journal on Control and Optimization, 43(3):1078–1093, May 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43236.

Benaim:2005:SAD

[BHS05] Michel Benaïm, Josef Hofbauer, and Sylvain Sorin. Stochastic approximations and differential inclusions. SIAM Journal on Control and Optimization, 44(1):328– 348, January 2005. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43930.

Biccari:2019:NCL

[BHS19] Umberto Biccari and Víctor Hernández-Santamaría. Null controllability of linear and semilinear nonlocal heat equations with an additive integral kernel. SIAM Journal on Control and Optimization, 57(4): 2924–2938, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bertsekas:1987:RMN

[BHT87] Dimitri P. Bertsekas, Patrick A. Hosein, and Paul Tseng. Relaxation methods for network flow problems with convex arc costs. SIAM Journal on Control and Optimization, 25(5): 1219–1243, September 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Blondel:2010:CTA

[BHT10] Vincent D. Blondel, Julien M. Hendrickx, and John N. Tsitsiklis. Continuous-time average-preserving opinion dynamics with opinion-dependent communications. SIAM Journal on Control and Optimization, 48(8):5214–5240, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bonalli:2019:CPE

[BHT19] Riccardo Bonalli, Bruno Hérissé, and Emmanuel Trélat. Continuity of Pontryagin extremals with respect to delays in nonlinear optimal control. SIAM Journal on Control and Optimization, 57(2): 1440–1466, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Brugger:2021:NST

[BHT21] Rahel Brügger, Helmut Harbrecht, and Johannes Tausch. On the numerical solution of a time-dependent shape optimization problem for the heat equation. SIAM Journal on Control and Optimization, 59 (2):931–953, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bouach:2022:NID

[BHT22] Abderrahim Bouach, Tahar Haddad, and Lionel Thibault. Nonconvex integro-differential sweeping process with applications. SIAM Journal on Control and Optimization, 60 (5):2971–2995, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1397635.

Banks:1991:NAO

[BI91] H. T. Banks and K. Ito. A numerical algorithm for optimal feedback gains in high dimensional linear quadratic regulator problems. SIAM Journal on Control and Optimization, 29(3):499–515, May 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bounit:2008:TVR

[BI08] Hamid Bounit and Abdelali Idrissi. Time-varying regular bilinear systems. SIAM Journal on Control and Optimization, 47(3):1097–1126, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bianchini:1983:LCR

[Bia83] Rosa Maria Bianchini. Local controllability, rest states and cyclic points. SIAM Journal on Control and Optimization, 21(5):714–720, September 1983. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bielecki:1995:LQC

[Bie95] Tomasz R. Bielecki. The linear-quadratic control problem revisited. SIAM Journal on Control and Optimization, 33(5):1425–1442, September 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/23389.

Bergounioux:1999:PDS

[BIK99] Maïtine Bergounioux, Kazufumi Ito, and Karl Kunisch. Primal-dual strategy for constrained optimal control problems. SIAM Journal on Control and Optimization, 37(4):1176–1194, July 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32860.

Billig:2019:OAC

[Bil19] Yuly Billig. Optimal attitude control with two rotation axes. SIAM Journal on Control and Optimization, 57(2): 1068–1093, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Binding:1976:SPO

[Bin76] Paul Binding. Singularly perturbed optimal control problems. I: Convergence. SIAM Journal on Control and Optimization, 14(4):591–612, July 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Burns:1988:NAS

[BIP88] John Burns, Kazufumi Ito, and Georg Propst. On nonconvergence of adjoint semigroups for control systems with delays. SIAM Journal on Control and Optimization, 26(6): 1442–1454, November 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Burzoni:2020:VSC

[BIRS20] Matteo Burzoni, Vincenzo Ignazio, A. Max Reppen, and H. M. Soner. Viscosity solutions for controlled McKean– Vlasov jump-diffusions. SIAM Journal on Control and Optimization, 58(3):1676–1699, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bismut:1976:LQO

[Bis76] Jean-Michel Bismut. Linear quadratic optimal stochastic control with random coefficients. SIAM Journal on Control and Optimization, 14 (3):419–444, May 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bismut:1977:OCL

[Bis77a] Jean-Michel Bismut. On optimal control of linear stochastic equations with a linearquadratic criterion. SIAM Journal on Control and Optimization, 15(1):1–4, January 1977. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bismut:1977:PTM

[Bis77b] Jean-Michel Bismut. Probability theory methods in zerosum stochastic games. SIAM Journal on Control and Optimization, 15(4):539–545, July 1977. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bismut:1978:AMO

[Bis78a] Jean-Michel Bismut. An approximation method in optimal stochastic control. SIAM Journal on Control and Optimization, 16(1):122–130, January 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bismut:1978:DMC

[Bis78b] Jean-Michel Bismut. Duality methods in the control of densities. SIAM Journal on Control and Optimization, 16(5):771–777, September 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bismut:1982:POD

[Bis82] Jean-Michel Bismut. Partially observed diffusions and their control. SIAM Journal on Control and Optimization, 20 (2):302–309, March 1982. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bistritz:1983:NBI

[Bis83] Yuval Bistritz. Nested bases of invariants for minimal realizations of finite matrix sequences. SIAM Journal on Control and Optimization, 21 (5):804–821, September 1983. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Biswas:2012:ZSS

[Bis12] Imran H. Biswas. On zero-sum stochastic differential games with jump-diffusion driven state: a viscosity solution framework. SIAM Journal on Control and Optimization, 50(4):1823–1858, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Biswas:2017:VSH

[BISW17] Anup Biswas, Hitoshi Ishii, Subhamay Saha, and Lin Wang. On viscosity solution of HJB equations with state constraints and reflection control. SIAM Journal on Control and Optimization, 55(1):365–396, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Barron:1995:RMC

[BJ95] E. N. Barron and R. Jensen. Relaxed minimax control. SIAM Journal on Control and Optimization, 33(4):1028– 1039, July 1995. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25053.

Barron:1996:RCC

[BJ96] E. N. Barron and R. Jensen. Relaxation of constrained control problems. SIAM Journal on Control and Optimization, 34(6):2077–2091, November 1996. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27432.

Bohacek:2002:NTC

[BJ02] Stephan Bohacek and Edmund Jonckheere. Nonlinear tracking over compact sets with linear dynamically varying H^{∞} control. SIAM Journal on Control and Optimization, 40(4):1042–1071, July 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35058.

Bellaassali:2004:NOC

[BJ04] Saïd Bellaassali and Abderrahim Jourani. Necessary optimality conditions in multiobjective dynamic optimization. SIAM Journal on Control and Optimization, 42(6): 2043–2061, November 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40657.

Bonnans:2010:OCP

[BJ10] J. Frédéric Bonnans and Pascal Jaisson. Optimal control of a parabolic equation with time-dependent state constraints. SIAM Journal on Control and Optimization, 48(7):4550–4571, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Balde:2011:GLS

[BJ11] Moussa Balde and Philippe Jouan. Geometry of the limit sets of linear switched systems. SIAM Journal on Control and Optimization, 49(3): 1048–1063, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicon/resource/
                                        1/sjcodc/v49/i3/p1048_s1.

Buckdahn:2017:MFS

[BJ17] Rainer Buckdahn and Shuai Jing. Mean-field SDE driven by a fractional Brownian motion and related stochastic control problem. SIAM Journal on Control and Optimization, 55(3):1500–1533, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bian:2019:CTR

[BJ19] Tao Bian and Zhong-Ping Jiang. Continuous-time robust dynamic programming. SIAM Journal on Control and Optimization, 57(6):4150–4174, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Buttazzo:2009:OPM

[BJO09] G. Buttazzo, C. Jimenez, and E. Oudet. An optimization problem for mass transportation with congested dynamics. SIAM Journal on Control and Optimization, 48(3): 1961–1976, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Boel:1980:SMS

[BK80] René Boel and Michael Kohlmann. Semimartingale models of stochastic optimal control, with applications to double martingales. SIAM Journal on Control and Optimization, 18(5):511– 533, September 1980. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Banks:1982:ATN

[BK82] H. T. Banks and K. Kunisch. An approximation theory for nonlinear partial differential equations with applications to identification and control. SIAM Journal on Control and Optimization, 20(6): 815–849, November 1982. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Benes:1983:RZM

[BK83a] Václav E. Beneš and Ioannis Karatzas. On the relation of Zakai’s and Mortensen’s equations. SIAM Journal on Control and Optimization, 21(3): 472–489, May 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Brunovsky:1983:MRE

[BK83b] Pavol Brunovský and Jozef Komorník. The matrix Riccati equation and the noncontrollable linear-quadratic problem with terminal constraints. SIAM Journal on Control and Optimization, 21 (2):280–288, March 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Banks:1984:LRP

[BK84] H. T. Banks and K. Kunisch. The linear regulator problem for parabolic systems. SIAM Journal on Control and Optimization, 22 (5):684–698, September 1984. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Borkar:1995:ESC

[BK95] Vivek S. Borkar and Sunil Kumar. On extremal solutions of controlled nonlinear filtering equations. SIAM Journal on Control and Optimization, 33(3):718–724, May 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/23583.

Bashirov:1997:CCS

[BK97a] Agamirza E. Bashirov and Kerim R. Kerimov. On controllability conception for stochastic systems. SIAM Journal on Control and Optimization, 35(2):384–398, March 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/26097.

Bergounioux:1997:ALT

[BK97b] Maïtine Bergounioux and Karl Kunisch. Augmented Lagrangian techniques for elliptic state constrained optimal control problems. SIAM Journal on Control and Optimization, 35(5):1524–1543, September 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29330.

Budhiraja:1998:RNF

[BK98] Amarjit Budhiraja and Harold J.

Kushner. Robustness of nonlinear filters over the infinite time interval. SIAM Journal on Control and Optimization, 36(5):1618–1637, September 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31848.

Budhiraja:1999:ALR

[BK99] Amarjit Budhiraja and Harold J.

Kushner. Approximation and limit results for nonlinear filters over an infinite time interval. SIAM Journal on Control and Optimization, 37(6):

1946–1979, November 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32817.

Budhiraja:2000:ALR

[BK00] Amarjit Budhiraja and Harold J. Kushner. Approximation and limit results for nonlinear filters over an infinite time interval: Part II, random sampling algorithms. SIAM Journal on Control and Optimization, 38(6):1874–1908, November 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/34993.

Buche:2002:RCC

[BK02] Robert Buche and Harold J. Kushner. Rate of convergence for constrained stochastic approximation algorithms. SIAM Journal on Control and Optimization, 40(4):1011– 1041, July 2002. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/36163.

Bianchini:2003:NVC

[BK03] Rosa-Maria Bianchini and Matthias Kawski. Needle variations that cannot be summed. SIAM Journal on Control and Optimization, 42 (1):218–238, January 2003.

CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40287.

Butnariu:2008:PPM

[BK08] Dan Butnariu and Gábor Kassay. A proximal-projection method for finding zeros of setvalued operators. SIAM Journal on Control and Optimization, 47(4):2096–2136, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Borkar:2010:SPR

[BK10] V. S. Borkar and K. Suresh Kumar. Singular perturbations in risk-sensitive stochastic control. SIAM Journal on Control and Optimization, 48(6):3675–3697, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Borzi:2013:FNS

[BK13] Alfio Borzì and Christian Kanzow. Formulation and numerical solution of Nash equilibrium multiobjective elliptic control problems. SIAM Journal on Control and Optimization, 51(1):718–744, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Breiten:2014:RBF

[BK14] Tobias Breiten and Karl Kunisch. Riccati-based feedback control of the monodomain equations with the Fitzhugh– Nagumo model. SIAM Journal on Control and Optimization, 52(6):4057–4081, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Boryc:2016:COP

[BK16] Marcin Boryc and Lukasz Kruk. Characterization of the optimal policy for a multidimensional parabolic singular stochastic control problem. SIAM Journal on Control and Optimization, 54(3): 1657–1677, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bonart:2021:OCS

[BK21] Henning Bonart and Christian Kahle. Optimal control of sliding droplets using the contact angle distribution. SIAM Journal on Control and Optimization, 59(2): 1057–1082, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bayraktar:2022:PDH

[BK22a] Erhan Bayraktar and Christian Keller. Path-dependent Hamilton–Jacobi equations with super-quadratic growth in the gradient and the vanishing viscosity method. SIAM Journal on Control and Optimization, 60(3):1690–1711, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1395557.

Belabbas:2022:SSR

[BK22b] Mohamed-Ali Belabbas and Artur Kirkoryan. On stable systems with random structure. SIAM Journal on Control and Optimization, 60(1):458–478, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1378363.

Breiten:2023:ICR

[BK23] Tobias Breiten and Karl K.

Kunisch. Improving the convergence rates for the kinetic Fokker–Planck equation by optimal control. SIAM Journal on Control and Optimization, 61(3):1557–1581, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1499698.

Bardi:2024:DRC

[BK24a] Martino Bardi and Hicham Kouhkouh. Deep relaxation of controlled stochastic gradient descent via singular perturbations. SIAM Journal on Control and Optimization, 62 (4):2229–2253, August 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1544878.

Bettiol:2024:ACM

[BK24b] Piernicola Bettiol and Nathalie Khalil. Average cost minimization problems subject to state constraints. SIAM Journal on Control and Optimization, 62(3):1884–1907, June 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1558124.

Borzi:2002:ACP

[BKK02] Alfio Borzì, Karl Kunisch, and Do Y. Kwak. Accuracy and convergence properties of the finite difference multigrid solution of an optimal control optimality system. SIAM Journal on Control and Optimization, 41(5):1477–1497, September 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/39343.

Bartl:2022:MDU

[BKL^{+}22] Daniel Bartl, Michael Kupper, Thibaut Lux, Antonis Papapantoleon, and Stephan Eckstein. Marginal and dependence uncertainty: Bounds, optimal transport, and sharpness. SIAM Journal on Control and Optimization, 60(1):410–434, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M144709X.

Byers:1997:RLD

[BKM97] Ralph Byers, Peter Kunkel, and Volker Mehrmann. Regularization of linear descriptor systems with variable coefficients. SIAM Journal on Control and Optimization, 35 (1):117–133, January 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27893.

Burachik:2014:DAS

[BKM14] R. S. Burachik, C. Y. Kaya, and S. N. Majeed. A duality approach for solving controlconstrained linear-quadratic optimal control problems. SIAM Journal on Control and Optimization, 52(3):1423– 1456, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Banaszuk:1990:DDP

[BKP90] A. Banaszuk, M. Kociecki, and K. M. Przyłuski. The disturbance decoupling problem for implicit linear discretetime systems. SIAM Journal on Control and Optimization, 28(6):1270–1293, November 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Banas:2014:CIE

[BKP14] L’ubomír Banas, Markus Klein, and Andreas Prohl. Control of interface evolution in multiphase fluid flows. SIAM Journal on Control and Optimization, 52(4):2284– 2318, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Breiten:2018:IHB

[BKP18] Tobias Breiten, Karl Kunisch, and Laurent Pfeiffer. Infinite-horizon bilinear optimal control problems: Sensitivity analysis and polynomial feedback laws. SIAM Journal on Control and Optimization, 56(5):3184–3214, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Becker:2000:AFE

[BKR00] Roland Becker, Hartmut Kapp, and Rolf Rannacher. Adaptive finite element methods for optimal control of partial differential equations: Basic concept. SIAM Journal on Control and Optimization, 39(1):113–132, January 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/35109.

Breiten:2017:FSN

[BKR17] Tobias Breiten, Karl Kunisch, and Sérgio S. Rodrigues. Feedback stabilization to nonstationary solutions of a class of reaction diffusion equations of FitzHugh–Nagumo type. SIAM Journal on Control and Optimization, 55(4): 2684–2713, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Belomestny:2010:RMS

[BKS10] Denis Belomestny, Anastasia Kolodko, and John Schoenmakers. Regression methods for stochastic control problems and their convergence analysis. SIAM Journal on Control and Optimization, 48(5): 3562–3588, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Borgens:2019:LGA

[BKS19] Eike Börgens, Christian Kanzow, and Daniel Steck. Local and global analysis of multiplier methods for constrained optimization in Banach spaces. SIAM Journal on Control and Optimization, 57(6):3694–3722, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bartsch:2024:CVP

[BKSW24] Jan Bartsch, Patrik Knopf, Stefania Scheurer, and Jörg Weber. Controlling a Vlasov– Poisson plasma by a particlein-cell method based on a Monte Carlo framework. SIAM Journal on Control and Optimization, 62(4):1977– 2011, August 2024. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1563852.

Banks:1983:EDO

[BL83] H. T. Banks and P. K. Daniel Lamm. Estimation of delays and other parameters in nonlinear functional differential equations. SIAM Journal on Control and Optimization, 21(6):895–915, November 1983. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Basar:1989:DCN

[BL89] Tamer Başar and Shu Li. Distributed computation of Nash equilibria in linear-quadratic stochastic differential games. SIAM Journal on Control and Optimization, 27(3):563–578, May 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Borwein:1991:DRE

[BL91] J. M. Borwein and A. S. Lewis. Duality relationships for entropy-like minimization problems. SIAM Journal on Control and Optimization, 29 (2):325–338, March 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Barron:1996:OCB

[BL96] Emmanuel N. Barron and Wenxiong Liu. Optimal control of the blowup time. SIAM Journal on Control and Optimization, 34(1):102–123, January 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24502.

Brosilow:2000:FDR

[BL00a] Coleman Brosilow and Marshall J. Leitman. A frequency domain robust stability theorem for infinite dimensional systems with parametric uncertainty. SIAM Journal on Control and Optimization, 39(5):1423–1440, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35755.

Byrnes:2000:DBF

[BL00b] Christopher I. Byrnes and Anders Lindquist. On the duality between filtering and Nevanlinna–Pick interpolation. SIAM Journal on Control and Optimization, 39(3):757–775, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/35111.

Barcenas:2001:CEC

[BL01] Diomedes Barcenas and Hugo Leiva. Characterization of extremal controls for infinite dimensional time-varying systems. SIAM Journal on Control and Optimization, 40(2):333–347, March 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32363.

Boyarchenko:2002:PAO

[BL02] S. I. Boyarchenko and S. Z. Levendorskii. Perpetual American options under Lévy processes. SIAM Journal on Control and Optimization, 40(6):1663–1696, November 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/37398.

Burke:2003:VAA

[BL03] James V. Burke and D. Russell Luke. Variational analysis applied to the problem of optical phase retrieval. SIAM Journal on Control and Optimization, 42(2):576–595, March 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/40643.

Bergounioux:2004:OCB

[BL04] Maïtine Bergounioux and Suzanne Lenhart. Optimal control of bilateral obstacle problems. SIAM Journal on Control and Optimization, 43 (1):240–255, January 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41691.

Bullo:2005:LOC

[BL05] Francesco Bullo and Andrew D. Lewis. Low-order controllability and kinematic reductions for affine connection control systems. SIAM Journal on Control and Optimization, 44(3):885–908, May 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42118.

Buckdahn:2008:SDG

[BL08a] Rainer Buckdahn and Juan Li. Stochastic differential games and viscosity solutions of Hamilton–Jacobi–Bellman– Isaacs equations. SIAM Journal on Control and Optimization, 47(1):444–475, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Byrnes:2008:IMS

[BL08b] Christopher I. Byrnes and Anders Lindquist. Important moments in systems and control. SIAM Journal on Control and Optimization, 47(5): 2458–2469, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Boyarchenko:2009:AOR

[BL09] Svetlana Boyarchenko and Sergei Levendorskii. American options in regime-switching models. SIAM Journal on Control and Optimization, 48 (3):1353–1376, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Boyarchenko:2011:OSL

[BL11] Svetlana Boyarchenko and Sergei Levendorskiĭ. Optimal stopping in Lévy models for nonmonotone discontinuous payoffs. SIAM Journal on Control and Optimization, 49(5):2062–2082, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sicon/
                                        resource/1/sjcodc/v49/i5/
                                        p2062_s1.

Bao:2014:OSD

[BL14] Gang Bao and Jun Lai. Optimal shape design of a cavity for radar cross section reduction. SIAM Journal on Control and Optimization, 52(4): 2122–2140, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bayraktar:2015:BFT

[BL15] Erhan Bayraktar and Lifeng Lai. Byzantine fault tolerant distributed quickest change detection. SIAM Journal on Control and Optimization, 53 (2):575–591, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bauerle:2018:OCP

[BL18] Nicole Bäuerle and Dirk Lange. Optimal control of partially observable piecewise deterministic Markov processes. SIAM Journal on Control and Optimization, 56(2):1441– 1462, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Blankenship:1976:FDS

[Bla76] Gilmer L. Blankenship. Frequency domain stability criteria for stochastic nonlinear feedback systems. SIAM Journal on Control and Optimization, 14(6):1107–1123, November 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Blanc:1997:DET

[Bla97] Alain-Philippe Blanc. Deterministic exit time control problems with discontinuous exit costs. SIAM Journal on Control and Optimization, 35(2):399–434, March 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26734.

Bliman:2004:CAR

[Bli04] Pierre-Alexandre Bliman. A convex approach to robust stability for linear systems with uncertain scalar parameters. SIAM Journal on Control and Optimization, 42(6): 2016–2042, November 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39869.

Barbero-Linnan:2015:SOC

[BLJ15] M. Barbero-Linñán and B. Jakubczyk. Second order conditions for optimality and local controllability of discrete-time systems. SIAM Journal on Control and Optimization, 53(1):352–377, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Buckdahn:2014:NSD

[BLP14] Rainer Buckdahn, Juan Li, and Shige Peng. Nonlinear stochastic differential games involving a major player and a large number of collectively acting minor agents. SIAM Journal on Control and Optimization, 52(1):451–492, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Barbero-Linnan:2015:MFO

[BLPdD15] María Barbero-Linñán, David Iglesias Ponte, and David Martín de Diego. Morse families in optimal control problems. SIAM Journal on Control and Optimization, 53(1):414–433, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Balmaseda:2024:GAC

[BLPP24] Aitor Balmaseda, Davide Lonigro, and Juan Manuel Pérez-Pardo. On global approximate controllability of a quantum particle in a box by moving walls. SIAM Journal on Control and Optimization, 62(2): 826–852, March 2024. CO-DEN SJCODC. ISSN 1095-7138.

Buckdahn:2020:RFL

[BLQR20] Rainer Buckdahn, Juan Li, Marc Quincampoix, and Jérôme Renault. Representation formulas for limit values of long run stochastic optimal controls. SIAM Journal on Control and Optimization, 58(4):1846–1873, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bardos:1992:SSC

[BLR92] Claude Bardos, Gilles Lebeau, and Jeffrey Rauch. Sharp sufficient conditions for the observation, control, and stabilization of waves from the boundary. SIAM Journal on Control and Optimization, 30(5):1024– 1065, September 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bensoussan:2005:OPC

[BLS05a] Alain Bensoussan, R. H. Liu, and Suresh P. Sethi. Optimality of an (s, S) policy with compound Poisson and diffusion demands: a quasi-variational inequalities approach. SIAM Journal on Control and Optimization, 44(5):1650–1676, September 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/44373.

Burachik:2005:OAM

[BLS05b] R. S. Burachik, J. O. Lopes, and B. F. Svaiter. An outer approximation method for the variational inequality problem. SIAM Journal on Control and Optimization, 43(6): 2071–2088, November 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41548.

Budhiraja:2012:ATS

[BLS12] Amarjit Budhiraja, Xin Liu, and Adam Shwartz. Action time sharing policies for ergodic control of Markov chains. SIAM Journal on Control and Optimization, 50(1):

171–195, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v50/i1/p171_s1.

Bouchard:2019:SOS

[BLSZ19] Bruno Bouchard, Grégoire Loeper, Halil Mete Soner, and Chao Zhou. Secondorder stochastic target problems with generalized market impact. SIAM Journal on Control and Optimization, 57(6):4125–4149, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Beauchard:2015:LEC

[BLT15] Karine Beauchard, Horst Lange, and Holger Teismann. Local exact controllability of a one-dimensional nonlinear Schrödinger equation. SIAM Journal on Control and Optimization, 53(5):2781–2818, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bauschke:2007:PDS

[BLW07] Heinz H. Bauschke, Yves Lucet, and Xianfu Wang. Primal-dual symmetric intrinsic methods for finding antiderivatives of cyclically monotone operators. SIAM Journal on Control and Optimization, 46(6):2031–2051, ???? 2007. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Bo:2019:RSP

[BLY19] Lijun Bo, Huafu Liao, and Xiang Yu. Risk sensitive portfolio optimization with default contagion and regimeswitching. SIAM Journal on Control and Optimization, 57 (1):366–401, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bo:2021:OTP

[BLY21] Lijun Bo, Huafu Liao, and Xiang Yu. Optimal tracking portfolio with a ratcheting capital benchmark. SIAM Journal on Control and Optimization, 59(3):2346–2380, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Byrnes:1994:NDF

[BLZ94] Christopher I. Byrnes, Anders Lindquist, and Yishao Zhou. On the nonlinear dynamics of fast filtering algorithms. SIAM Journal on Control and Optimization, 32(3):744–789, May 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/20242.

Bouchard:2017:HCO

[BLZ17] Bruno Bouchard, Grégoire Loeper, and Yiyi Zou. Hedging of covered options with linear market impact and gamma constraint. SIAM Journal on Control and Optimization, 55 (5):3319–3348, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Blankenship:1984:OSS

[BM84] G. L. Blankenship and J.-L.

Menaldi. Optimal stochastic scheduling of power generation systems with scheduling delays and large cost differentials. SIAM Journal on Control and Optimization, 22(1): 121–132, January 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Banks:1986:ECB

[BM86] H. T. Banks and K. A. Murphy. Estimation of coefficients and boundary parameters in hyperbolic systems. SIAM Journal on Control and Optimization, 24(5):926–950, July 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Basile:1990:EMP

[BM90] N. Basile and M. Mininni.

An extension of the maximum principle for a class of optimal control problems in infinite-dimensional spaces. SIAM Journal on Control and Optimization, 28(5):1113– 1135, September 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Braatz:1994:MEC

[BM94] Richard D. Braatz and Manfred Morari. Minimizing the Euclidean condition number. SIAM Journal on Control and Optimization, 32(6): 1763–1768, November 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/23868.

Bashirov:1999:CCD

[BM99] Agamirza E. Bashirov and Nazim I. Mahmudov. On concepts of controllability for deterministic and stochastic systems. SIAM Journal on Control and Optimization, 37(6): 1808–1821, November 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32184.

Bacciotti:2000:NSC

[BM00a] Andrea Bacciotti and Luisa Mazzi. A necessary and sufficient condition for boundedinput bounded-state stability of nonlinear systems. SIAM Journal on Control and Optimization, 39(2):478– 491, March 2000. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/35057.

Blanchini:2000:DAL

[BM00b] Franco Blanchini and Stefano Miani. Any domain of attraction for a linear constrained system is a tracking domain of attraction. SIAM Journal on Control and Optimization, 38(3):971–994, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34661.

Borkar:2000:DMC

[BM00c] V. S. Borkar and S. P. Meyn.

The O.D.E. method for convergence of stochastic approximation and reinforcement learning. SIAM Journal on Control and Optimization, 38(2):447–469, March 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33163.

Blanchini:2003:SLS

[BM03] Franco Blanchini and Stefano Miani. Stabilization of LPV systems: State feedback, state estimation, and duality. SIAM Journal on Control and Optimization, 42(1):76–97, January 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37228.

Bachelier:2006:RMR

[BM06] Olivier Bachelier and Driss Mehdi. Robust matrix rootclustering analysis through extended KYP lemma. SIAM Journal on Control and Optimization, 45(1):368–381, January 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Buckdahn:2007:PSC

[BM07] Rainer Buckdahn and Jin Ma. Pathwise stochastic control problems and stochastic HJB equations. SIAM Journal on Control and Optimization, 45 (6):2224–2256, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Beauchard:2009:PSQ

[BM09] Karine Beauchard and Mazyar Mirrahimi. Practical stabilization of a quantum particle in a one-dimensional infinite square potential well. SIAM Journal on Control and Optimization, 48(2):1179–1205, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bacciotti:2011:ACM

[BM11] Andrea Bacciotti and Luisa Mazzi. Asymptotic controllability by means of eventually periodic switching rules. SIAM Journal on Control and Optimization, 49(2):476–

497, ????  2011.  CO-

DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i2/p476_s1.

Bonnel:2013:SBC

[BM13] Henri Bonnel and Jacqueline Morgan. Semivectorial bilevel convex optimal control problems: Existence results. SIAM Journal on Control and Optimization, 50(6): 3224–3241, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Beschastnyi:2018:LIS

[BM18a] Ivan Beschastnyi and Alexandr Medvedev. Left-invariant sub-Riemannian Engel structures: Abnormal geodesics and integrability. SIAM Journal on Control and Optimization, 56 (5):3524–3537, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Busic:2018:ODE

[BM18b] Ana Busić and Sean Meyn.

Ordinary differential equation methods for Markov decision processes and application to Kullback–Leibler control cost. SIAM Journal on Control and Optimization, 56(1):343–366, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bagagiolo:2019:HTA

[BM19] Fabio Bagagiolo and Rosario Maggistro. Hybrid thermostatic approximations of junctions for some optimal control problems on networks. SIAM Journal on Control and Optimization, 57(4):2415–2442, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bai:2021:OID

[BM21] Lihua Bai and Jin Ma. Optimal investment and dividend strategy under renewal risk model. SIAM Journal on Control and Optimization, 59(6): 4590–4614, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Borden:1985:SIT

[BMM85] J. Martin Borden, David M. Mason, and Robert J. McEliece. Some information theoretic saddlepoints. SIAM Journal on Control and Optimization, 23(1):129–143, January 1985. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bressan:2019:ASP

[BMN19] Alberto Bressan, Marco Mazzola, and Khai T. Nguyen. Approximation of sweeping processes and controllability for a set-valued evolution. SIAM Journal on Control and Optimization, 57(4):2487–2514, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bourles:2015:ESL

[BMO15] H. Bourlès, B. Marinescu, and U. Oberst. Exponentially stable linear time-varying discrete behaviors. SIAM Journal on Control and Optimization, 53(5):2725–2761, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Badra:2020:STP

[BMRR20] Mehdi Badra, Debanjana Mitra, Mythily Ramaswamy, and Jean-Pierre Raymond. Stabilizability of time-periodic evolution equations by finite dimensional controls. SIAM Journal on Control and Optimization, 58(3):1735–1768, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ball:1982:CDB

[BMS82] J. M. Ball, J. E. Marsden, and M. Slemrod. Controllability for distributed bilinear systems. SIAM Journal on Control and Optimization, 20 (4):575–597, July 1982. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ball:2005:SSF

[BMS05] Joseph A. Ball, Kalle M.

Mikkola, and Amol J. Sasane. State-space formulas for the Nehari–Takagi problem for nonexponentially stable infinitedimensional systems. SIAM Journal on Control and Optimization, 44(2):531–563, March 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/43302.

Belak:2015:UUV

[BMS15] Christoph Belak, Olaf Menkens, and Jörn Sass. On the uniqueness of unbounded viscosity solutions arising in an optimal terminal wealth problem with transaction costs. SIAM Journal on Control and Optimization, 53(5):2878–2897, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Borkar:2001:OSV

[BMT01] Vivek S. Borkar, Sanjoy K. Mitter, and Sekhar Tatikonda. Optimal sequential vector quantization of Markov sources. SIAM Journal on Control and Optimization, 40(1):135– 148, January 2001. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36526.

Banon:1981:RED

[BN81] G. Banon and Hung T. Nguyen. Recursive estimation in diffusion model.

SIAM Journal on Control and Optimization, 19(5):676– 685, September 1981. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bensoussan:1997:MMC

[BN97] A. Bensoussan and H. Nagai. Min-max characterization of a small noise limit on risk-sensitive control. SIAM Journal on Control and Optimization, 35(4):1093–1115, July 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27942.

Bouchard:2013:WDP

[BN13] Bruno Bouchard and Marcel Nutz. Weak dynamic programming for generalized state constraints. SIAM Journal on Control and Optimization, 50(6):3344–3373, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Buckdahn:2016:GHJ

[BN16] Rainer Buckdahn and Tianyang Nie. Generalized Hamilton– Jacobi–Bellman equations with Dirichlet boundary condition and stochastic exit time optimal control problem. SIAM Journal on Control and Optimization, 54(2):602–631, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Brekke:1994:OSE

[BØ94] Kjell Arne Brekke and Bernt Øksendal. Optimal switching in an economic activity under uncertainty. SIAM Journal on Control and Optimization, 32(4):1021–1036, July 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/22983.

Burke:2001:VAA

[BO01] James V. Burke and Michael L. Overton. Variational analysis of the abscissa mapping for polynomials. SIAM Journal on Control and Optimization, 39(6):1651–1676, November 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/36765.

Bourles:2009:DDD

[BO09] Henri Bourlès and Ulrich Oberst. Duality for differentialdifference systems over lie groups. SIAM Journal on Control and Optimization, 48 (4):2051–2084, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Blondel:2014:HDC

[BO14] Vincent D. Blondel and Alex Olshevsky. How to decide consensus? A combinatorial necessary and sufficient condition and a proof that consensus is decidable but NPhard. SIAM Journal on Control and Optimization, 52(5): 2707–2726, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bogosel:2016:QNA

[BO16] Beniamin Bogosel and Édouard Oudet. Qualitative and numerical analysis of a spectral problem with perimeter constraint. SIAM Journal on Control and Optimization, 54 (1):317–340, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Boetius:2005:BVS

[Boe05] Frederik Boetius. Bounded variation singular stochastic control and Dynkin game. SIAM Journal on Control and Optimization, 44(4):1289– 1321, July 2005. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42904.

Boley:1980:KPC

[Bol80] Daniel L. Boley. On Kalman’s procedure for the computation of the controllable/observable canonical form. SIAM Journal on Control and Optimization, 18(6):624–626, November 1980. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Bonnard:1984:CSM

[Bon84] Bernard Bonnard. Contrôllabilité de systemes mécaniques sur les groupes de Lie. (French) [Controllability of mechanical systems on Lie groups]. SIAM Journal on Control and Optimization, 22(5):711– 722, September 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bonnard:1985:SET

[Bon85] Bernard Bonnard. On singular extremals in the time minimal control problem in R^{3}. SIAM Journal on Control and Optimization, 23(5):794– 802, September 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bonnans:1986:AOC

[Bon86] Joseph Frédéric Bonnans. On an algorithm for optimal control using Pontryagin’s maximum principle. SIAM Journal on Control and Optimization, 24(3):579–588, May 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bonnans:1989:AAU

[Bon89] Joseph Frédéric Bonnans. Asymptotic admissibility of the unit stepsize in exact penalty methods. SIAM Journal on Control and Optimization, 27(3):631–641, May 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bonnard:1991:FEN

[Bon91] B. Bonnard. Feedback equivalence for nonlinear systems and the time optimal control problem. SIAM Journal on Control and Optimization, 29(6):1300–1321, November 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Boothby:1982:SCP

[Boo82] William M. Boothby. Some comments on positive orthant controllability of bilinear systems. SIAM Journal on Control and Optimization, 20 (5):634–644, September 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Borwein:1977:PEP

[Bor77]  J. Borwein.  Proper effi-

cient points for maximizations with respect to cones. SIAM Journal on Control and Optimization, 15(1):57–63, January 1977. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Borwein:1978:WTC

[Bor78] J. Borwein. Weak tangent cones and optimization in a Banach space. SIAM Journal on Control and Optimization, 16(3):512–522, May 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Borkar:1983:CMC

[Bor83] Vivek S. Borkar. Controlled Markov chains and stochastic networks. SIAM Journal on Control and Optimization, 21 (4):652–666, July 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Borkar:1984:MCP

[Bor84] Vivek S. Borkar. On minimum cost per unit time control of Markov chains. SIAM Journal on Control and Optimization, 22(6):965–978, November 1984. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Borkar:1989:CMC

[Bor89] Vivek S. Borkar. Control of Markov chains with long-run average cost criterion: The dynamic programming equations. SIAM Journal on Control and Optimization, 27(3): 642–657, May 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Borkar:1993:WNR

[Bor93] Vivek S. Borkar. White-noise representations in stochastic realization theory. SIAM Journal on Control and Optimization, 31(5):1093–1102, September 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See erratum [Bor10].

Borkar:1994:ECM

[Bor94] Vivek S. Borkar. Ergodic control of Markov chains with constraints — the general case. SIAM Journal on Control and Optimization, 32 (1):176–186, January 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/21889.

Borkar:1998:ASA

[Bor98]  Vivek S. Borkar.  Asyn-

chronous stochastic approximations. SIAM Journal on Control and Optimization, 36(3):840–851, May 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28278. See er-

ratum [Bor00b].

Borkar:2000:ACD

[Bor00a] V. S. Borkar. Average cost dynamic programming equations for controlled Markov chains with partial observations. SIAM Journal on Control and Optimization, 39(3):673–681, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/34517. See cor-

rection [Bor07].

Borkar:2000:EAS

[Bor00b] Vivek S. Borkar. Erratum: “Asynchronous Stochastic Approximations”. SIAM Journal on Control and Optimization, 38(2):662–663, March 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/34591.  See

[Bor98].

Borkar:2007:DPE

[Bor07] Vivek S. Borkar. Dynamic programming for ergodic control of Markov chains under partial observations: a correction. SIAM Journal on Control and Optimization, 45(6): 2299–2304, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [Bor00a].

Borkar:2010:EWN

[Bor10] Vivek S. Borkar. Erratum: “White-Noise Representations in Stochastic Realization Theory”. SIAM Journal on Control and Optimization, 48(7): 4815–4819, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [Bor93].

Boscain:2002:SPS

[Bos02] Ugo Boscain. Stability of planar switched systems: The linear single input case. SIAM Journal on Control and Optimization, 41(1):89–112, January 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38283.

Bougerol:1993:KFR

[Bou93] Philippe Bougerol. Kalman filtering with random coefficients and contractions. SIAM Journal on Control and Optimization, 31(4):942–959, July 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bougerol:1995:ASS

[Bou95] Philippe Bougerol. Almost sure stabilizability and Riccati’s equation of linear systems with random parameters. SIAM Journal on Control and Optimization, 33(3):702–717, May 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/22906.

Bouchard:2008:ORD

[Bou08] Bruno Bouchard. Optimal reflection of diffusions and barrier options pricing under constraints. SIAM Journal on Control and Optimization, 47 (4):1785–1813, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bousquet:2013:EEM

[Bou13] Pierre Bousquet. The Euler equation in the multiple integrals calculus of variations. SIAM Journal on Control and Optimization, 51(2): 1047–1062, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Boulkhemair:2017:SDF

[Bou17]  A. Boulkhemair.  On a

shape derivative formula in the Brunn–Minkowski theory. SIAM Journal on Control and Optimization, 55(1):156–171, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Baram:1988:IMO

[BP88a] Y. Baram and B. Porat. Identification of minimal order state space models from stochastic input-output data. SIAM Journal on Control and Optimization, 26(1):56– 65, January 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Barles:1988:ETP

[BP88b] G. Barles and B. Perthame. Exit time problems in optimal control and vanishing viscosity method. SIAM Journal on Control and Optimization, 26(5):1133–1148, September 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Baker:1994:OCS

[BP94] T. E. Baker and E. Polak.

On the optimal control of systems described by evolution equations. SIAM Journal on Control and Optimization, 32(1):224–260, January 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/17411.

Barbu:1997:DAI

[BP97] V. Barbu and N. H. Pavel.

Determining the acoustic impedance in the 1-D wave equation via an optimal control problem. SIAM Journal on Control and Optimization, 35(5):1544–1556, September 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28369.

Bressan:1998:GCT

[BP98a] A. Bressan and B. Piccoli.

A generic classification of time-optimal planar stabilizing feedbacks. SIAM Journal on Control and Optimization, 36(1):12–32, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29111.

Bucci:1998:VFS

[BP98b] Francesca Bucci and Luciano Pandolfi. The value function of the singular quadratic regulator problem with distributed control action. SIAM Journal on Control and Optimization, 36(1):115–136, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29536.

Boscain:2002:ARA

[BP02] Ugo Boscain and Benedetto Piccoli. On automaton recognizability of abnormal extremals. SIAM Journal on Control and Optimization, 40(5):1333–1357, September 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38165.

Bednarik:2004:ESC

[BP04] Dusan Bednarík and Karel Pastor. Elimination of strict convergence in optimization. SIAM Journal on Control and Optimization, 43(3):1063–1077, May 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42417. See er-

rata [BP06].

Bayraktar:2005:SDG

[BP05] Erhan Bayraktar and H. Vincent Poor. Stochastic differential games in a non-Markovian setting. SIAM Journal on Control and Optimization, 43(5):1737–1756, September 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41763.

Bednarik:2006:EES

[BP06] Dusan Bednarík and Karel Pastor. Errata: Elimination of strict convergence in optimization. SIAM Journal on Control and Optimization, 45(1): 382–387, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [BP04].

Bayraktar:2007:QDM

[BP07a] Erhan Bayraktar and H. Vincent Poor. Quickest detection of a minimum of two Poisson disorder times. SIAM Journal on Control and Optimization, 46(1):308–331, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Blanchini:2007:ROC

[BP07b] Franco Blanchini and Felice Andrea Pellegrino. Relatively optimal control: a static piecewise-affine solution. SIAM Journal on Control and Optimization, 46(2): 585–603, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bercu:2008:KDE

[BP08a] Bernard Bercu and Bruno Portier. Kernel density estimation and goodness-of-fit test in adaptive tracking. SIAM Journal on Control and Optimization, 47(5): 2440–2457, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Boche:2008:BSF

[BP08b] Holger Boche and Volker Pohl. On the boundedness of the spectral factorization mapping on decomposing Banach algebras. SIAM Journal on Control and Optimization, 47(3): 1415–1427, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bai:2010:ODP

[BP10] Lihua Bai and Jostein Paulsen. Optimal dividend policies with transaction costs for a class of diffusion processes. SIAM Journal on Control and Optimization, 48(8):4987–5008, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bauso:2012:TTP

[BP12] D. Bauso and R. Pesenti. Team theory and person-by-person optimization with binary decisions. SIAM Journal on Control and Optimization, 50(5):3011–3028, ???? 2012.

CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bergounioux:2013:SOS

[BP13a] Maitine Bergounioux and Yannick Privat. Shape optimization with Stokes constraints over the set of axisymmetric domains. SIAM Journal on Control and Optimization, 51(1):599–628, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bombrun:2013:ACS

[BP13b] Alex Bombrun and Jean-Baptiste Pomet. The averaged control system of fastoscillating control systems. SIAM Journal on Control and Optimization, 51(3):2280– 2305, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bardi:2014:LQP

[BP14a] Martino Bardi and Fabio S.

Priuli. Linear-quadratic Nperson and mean-field games with ergodic cost. SIAM Journal on Control and Optimization, 52(5):3022–3052, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bhat:2014:STL

[BP14b] Sanjay P. Bhat and Aditya A.

Paranjape. Small-time local controllability and stabilizability of spacecraft attitude dynamics under CMG actuation. SIAM Journal on Control and Optimization, 52(2):797–820, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bi:2017:MCR

[BP17] Shujun Bi and Shaohua Pan. Multistage convex relaxation approach to rank regularized minimization problems based on equivalent mathematical program with a generalized complementarity constraint. SIAM Journal on Control and Optimization, 55 (4):2493–2518, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Barcena-Petisco:2020:UCS

[BP20a] Jon Asier Bárcena-Petisco. Uniform controllability of a Stokes problem with a transport term in the zero-diffusion limit. SIAM Journal on Control and Optimization, 58(3): 1597–1625, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bisiacco:2020:KAS

[BP20b] Mauro Bisiacco and Gianluigi Pillonetto. Kernel absolute summability is sufficient but not necessary for RKHS stability. SIAM Journal on Control and Optimization, 58(4):

2006–2022, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Breiten:2020:TPR

[BP20c] Tobias Breiten and Laurent Pfeiffer. On the turnpike property and the receding-horizon method for linear-quadratic optimal control problems. SIAM Journal on Control and Optimization, 58(2):1077– 1102, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Birtea:2004:CPS

[BPR04] Petre Birtea, Mircea Puta, and Tudor S. Ratiu. Controllability of Poisson systems. SIAM Journal on Control and Optimization, 43(3):937–954, May 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40125.

Bielecki:2005:RSP

[BPS05] Tomasz R. Bielecki, Stanley R.

Pliska, and Shuenn-Jyi Sheu. Risk sensitive portfolio management with Cox–Ingersoll– Ross interest rates: The HJB equation. SIAM Journal on Control and Optimization, 44(5):1811–1843, September 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/43795.

Boscain:2021:CQC

[BPS21] Ugo Boscain, Eugenio Pozzoli, and Mario Sigalotti. Classical and quantum controllability of a rotating symmetric molecule. SIAM Journal on Control and Optimization, 59 (1):156–184, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Burger:2021:MFO

[BPTT21] Martin Burger, René Pinnau, Claudia Totzeck, and Oliver Tse. Mean-field optimal control and optimality conditions in the space of probability measures. SIAM Journal on Control and Optimization, 59 (2):977–1006, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bonifacius:2019:EES

[BPV19a] Lucas Bonifacius, Konstantin Pieper, and Boris Vexler. Error estimates for space–time discretization of parabolic time-optimal control problems with bang-bang controls. SIAM Journal on Control and Optimization, 57(3): 1730–1756, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bonifacius:2019:PEE

[BPV19b] Lucas Bonifacius, Konstantin Pieper, and Boris Vexler. A priori error estimates for space–time finite element discretization of parabolic timeoptimal control problems. SIAM Journal on Control and Optimization, 57(1):129–162, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bokanowski:2016:SCS

[BPZ16] Olivier Bokanowski, Athena Picarelli, and Hasnaa Zidani. State-constrained stochastic optimal control problems via reachability approach. SIAM Journal on Control and Optimization, 54(5):2568–2593, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Burke:1999:VMP

[BQ99] J. V. Burke and Maijian Qian.

A variable metric proximal point algorithm for monotone operators. SIAM Journal on Control and Optimization, 37(2):353–375, March 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23554.

Buckdahn:2008:CSD

[BQT08] Rainer Buckdahn, Marc Quincampoix, and Gianmario Tessitore. Controlled stochastic differential equations under constraints in infinite dimensional spaces. SIAM Journal on Control and Optimization, 47(1):218–250, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bensoussan:1982:CDT

[BR82] A. Bensoussan and M. Robin. On the convergence of the discrete time dynamic programming equation for general semigroups. SIAM Journal on Control and Optimization, 20(5):722–746, September 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ball:1987:OHN

[BR87a] Joseph A. Ball and André C. M. Ran. Optimal Hankel norm model reductions and Wiener–Hopf factorization. I. The canonical case. SIAM Journal on Control and Optimization, 25(2):362– 382, March 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Baumann:1987:FCA

[BR87b] William T. Baumann and Wilson J. Rugh. Feedback control of analytic nonlinear systems by extended linearization. SIAM Journal on Control and Optimization, 25(5): 1341–1352, September 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Burton:1988:RSE

[BR88] G. R. Burton and E. P. Ryan.

On reachable sets and extremal rearrangements of control functions. SIAM Journal on Control and Optimization, 26(6):1481–1489, November 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bressan:1993:DSQ

[BR93] Alberto Bressan and Franco Rampazzo. On differential systems with quadratic impulses and their applications to Lagrangian mechanics. SIAM Journal on Control and Optimization, 31(5):1205– 1220, September 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Briani:2001:LAO

[BR01] Ariela Briani and Franco Rampazzo. Lie algebraic obstructions to Γ-convergence of optimal control problems. SIAM Journal on Control and Optimization, 40(2):370–392, March 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36356.

Budhiraja:2007:CNS

[BR07] Amarjit Budhiraja and Kevin Ross. Convergent numerical scheme for singular stochastic control with state constraints in a portfolio selection problem. SIAM Journal on Control and Optimization, 45(6): 2169–2206, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Boscain:2008:ICC

[BR08] Ugo Boscain and Francesco Rossi. Invariant Carnot– Carathéodory metrics on S^{3}, SO(3), SL(2), and lens spaces. SIAM Journal on Control and Optimization, 47(4): 1851–1878, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Benaim:2010:CSI

[BR10] Michel Benaïm and Olivier Raimond. A class of selfinteracting processes with applications to games and reinforced random walks. SIAM Journal on Control and Optimization, 48(7):4707–4730, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Blass:2013:SOD

[BR13] Timothy Blass and L. A. Romero. Stability of ordinary differential equations with colored noise forcing. SIAM Journal on Control and Optimization, 51(2):1099–1127, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bivas:2015:PPD

[BR15a] Mira Bivas and Nadezhda Ribarska. Projection process with definable right-hand side. SIAM Journal on Control and Optimization, 53(5): 2819–2834, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Burns:2015:SAR

[BR15b] John A. Burns and Carlos N.

Rautenberg. Solutions and approximations to the Riccati integral equation with values in a space of compact operators. SIAM Journal on Control and Optimization, 53(5): 2846–2877, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Berger:2017:ODC

[BR17a] Thomas Berger and Timo Reis. Observers and dynamic controllers for linear differential-algebraic systems. SIAM Journal on Control and Optimization, 55(6):3564– 3591, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Blank:2017:EPG

[BR17b] Luise Blank and Christoph Rupprecht. An extension of the projected gradient method to a Banach space setting with application in structural topology optimization. SIAM Journal on Control and Optimization, 55(3):1481–1499, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bonnet:2021:ILR

[BR21] Benoît Bonnet and Francesco Rossi. Intrinsic Lipschitz regularity of mean-field optimal controls. SIAM Journal on Control and Optimization, 59 (3):2011–2046, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Brandiere:1998:SPT

[Bra98] Odile Brandière. Some pathological traps for stochastic approximation. SIAM Journal on Control and Optimization, 36(4):1293–1314, July 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/30759.

Braack:2009:OCF

[Bra09] M. Braack. Optimal control in fluid mechanics by finite elements with symmetric stabilization. SIAM Journal on Control and Optimization, 48 (2):672–687, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bremaud:1979:OTP

[Bré79] P. Brémaud. Optimal thinning of a point process. SIAM Journal on Control and Optimization, 17(2):222–230, March 1979. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bressan:1985:HOT

[Bre85] Alberto Bressan. A high order test for optimality of bang–bang controls. SIAM Journal on Control and Optimization, 23(1):38–48, January 1985. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bressan:1986:GLT

[Bre86] Alberto Bressan. The generic local time-optimal stabilizing controls in dimension 3. SIAM Journal on Control and Optimization, 24(2): 177–190, March 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bressan:2007:ICC

[Bre07] Alberto Bressan. On the intersection of a Clarke cone with a Boltyanskii cone. SIAM Journal on Control and Optimization, 45(6):2054–2064, ???? 2007. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Breiten:2016:SPM

[Bre16] Tobias Breiten. Structurepreserving model reduction for integro-differential equations. SIAM Journal on Control and Optimization, 54(6): 2992–3015, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Brokate:1987:NOC

[Bro87] Martin Brokate. Necessary optimality conditions for the control of semilinear hyperbolic boundary value problems. SIAM Journal on Control and Optimization, 25(5): 1353–1369, September 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Broucke:2010:RCS

[Bro10] Mireille E. Broucke. Reach control on simplices by continuous state feedback. SIAM Journal on Control and Optimization, 48(5):3482–3500, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Brogliato:2022:AIE

[Bro22]  Bernard Brogliato.  Anal-

ysis of the implicit Euler time-discretization of semiexplicit differential-algebraic linear complementarity systems. SIAM Journal on Control and Optimization, 60 (4):2159–2183, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1396101.

Banks:1990:ATI

[BRR90] H. T. Banks, Simeon Reich, and I. G. Rosen. An approximation theory for the identification of nonlinear distributed parameter systems. SIAM Journal on Control and Optimization, 28(3):552–569, May 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bakule:2002:OQO

[BRR02] Lubomír Bakule, José Rodellar, and Josep M. Rossell. Overlapping quadratic optimal control of linear timevarying commutative systems. SIAM Journal on Control and Optimization, 40(5): 1611–1627, September 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36764.

Barbu:2011:IES

[BRS11] Viorel Barbu, Sérgio S. Rodrigues, and Armen Shirikyan. Internal exponential stabilization to a nonstationary solution for 3D Navier–Stokes equations. SIAM Journal on Control and Optimization, 49 (4):1454–1478, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sicon/
                                        resource/1/sjcodc/v49/i4/
                                        p1454_s1.

Bayen:2014:MTC

[BRS14] Terence Bayen, Alain Rapaport, and Matthieu Sebbah. Minimal time control of the two tanks gradostat model under a cascade input constraint. SIAM Journal on Control and Optimization, 52 (4):2568–2594, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Blass:2015:MSS

[BRT15] T. J. Blass, L. A. Romero, and J. R. Torczynski. On the moment stability of stochastic parametrically forced equations with rank one forcing. SIAM Journal on Control and Optimization, 53(4): 1842–1859, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Brunovsky:1976:CLT

[Bru76] Pavol Brunovský. The closedloop time optimal control. II: Stability. SIAM Journal on Control and Optimization, 14(1):156–162, January 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Brull:2013:GAR

[Brü13] Tobias Brüll. Generalizing the algebraic Riccati equation to higher-order behavioral systems. SIAM Journal on Control and Optimization, 51(3): 2544–2567, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bonder:2005:RFB

[BRW05] Julián Fernández Bonder, Julio D. Rossi, and Noemi Wolanski. Regularity of the free boundary in an optimization problem related to the best Sobolev trace constant. SIAM Journal on Control and Optimization, 44(5): 1614–1635, September 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/61361.

Barbu:2020:OCN

[BRZ20] Viorel Barbu, Michael Röckner, and Deng Zhang. Optimal control of nonlinear stochastic differential equations on Hilbert spaces. SIAM Journal on Control and Optimization, 58(4):2383–2410, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bachem:1980:MIS

[BS80] Achim Bachem and Rainer Schrader. Minimal inequalities and subadditive duality.

SIAM Journal on Control and Optimization, 18(4):437–443, ???? 1980. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Baram:1986:MOE

[BS86] Yoram Baram and Uri Shaked. Minimal order estimation of multivariable discrete-time stochastic linear systems. SIAM Journal on Control and Optimization, 24(4):817–820, July 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bermudez:1987:OCS

[BS87] A. Bermúdez and C. Saguez. Optimal control of a Signorini problem. SIAM Journal on Control and Optimization, 25 (3):576–582, May 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bianchini:1990:GAC

[BS90] Rosa Maria Bianchini and Gianna Stefani. Graded approximations and controllability along a trajectory. SIAM Journal on Control and Optimization, 28(4):903–924, July 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bianchini:1993:CAT

[BS93] Rosa Maria Bianchini and Gianna Stefani. Controllability along a trajectory: a variational approach. SIAM Journal on Control and Optimization, 31(4):900–927, July 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Burachik:2000:PPM

[BS00] Regina Sandra Burachik and Susana Scheimberg. A proximal point method for the variational inequality problem in Banach spaces. SIAM Journal on Control and Optimization, 39(5):1633–1649, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33974.

Buttazzo:2003:RTD

[BS03] Giuseppe Buttazzo and Eugene Stepanov. On regularity of transport density in the Monge–Kantorovich problem. SIAM Journal on Control and Optimization, 42(3):1044–1055, May 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38568.

Bressan:2004:SBS

[BS04] Alberto Bressan and Wen Shen. Small BV solutions of hyperbolic noncooperative differential games. SIAM Journal on Control and Optimization, 43(1):194–215, January 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/42558.

Burger:2006:PFR

[BS06] Martin Burger and Roman Stainko. Phase-field relaxation of topology optimization with local stress constraints. SIAM Journal on Control and Optimization, 45 (4):1447–1466, January 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bonnard:2009:TMC

[BS09a] Bernard Bonnard and Dominique Sugny. Time-minimal control of dissipative two-level quantum systems: The integrable case. SIAM Journal on Control and Optimization, 48(3):1289–1308, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bressan:2009:MVS

[BS09b] Alberto Bressan and Wen Shen. Measure-valued solutions for a differential game related to fish harvesting. SIAM Journal on Control and Optimization, 47(6):3118–3137, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bonnans:2011:AES

[BS11] J. Frédéric Bonnans and Francisco J. Silva. Asymptotic expansion for the solutions of control constrained semilinear elliptic problems with interior penalties. SIAM Journal on Control and Optimization, 49(6):2494–2517, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sicon/
resource/1/sjcodc/v49/i6/
p2494_s1.

Bayraktar:2013:SPM

[BS13a] Erhan Bayraktar and Mihai Sîrbu. Stochastic Perron’s method for Hamilton–Jacobi– Bellman equations. SIAM Journal on Control and Optimization, 51(6):4274–4294, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Brzezniak:2013:ORC

[BS13b] Zdzislaw Brzeźniak and Rafael Serrano. Optimal relaxed control of dissipative stochastic partial differential equations in Banach spaces. SIAM Journal on Control and Optimization, 51(3):2664–2703, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Basu:2015:FIH

[BS15] Arnab Basu and Lukasz Stettner. Finite- and infinitehorizon Shapley games with nonsymmetric partial observation. SIAM Journal on Control and Optimization, 53(6): 3584–3619, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bayen:2016:SOA

[BS16] Térence Bayen and Francisco J. Silva. Second order analysis for strong solutions in the optimal control of parabolic equations. SIAM Journal on Control and Optimization, 54(2):819–844, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bichuch:2017:OIT

[BS17] Maxim Bichuch and Ronnie Sircar. Optimal investment with transaction costs and stochastic volatility Part I: Infinite horizon. SIAM Journal on Control and Optimization, 55(6):3799–3832, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bayraktar:2018:SND

[BS18] Erhan Bayraktar and Qingshuo Song. Solvability of the nonlinear Dirichlet problem with integro-differential operators. SIAM Journal on Control and Optimization, 56 (1):292–315, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bichuch:2019:OIT

[BS19a] Maxim Bichuch and Ronnie Sircar. Optimal investment with transaction costs and stochastic volatility. Part II: Finite horizon. SIAM Journal on Control and Optimization, 57(1):437–467, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bressan:2019:CHM

[BS19b] Alberto Bressan and Vasile Staicu. On the competitive harvesting of marine resources. SIAM Journal on Control and Optimization, 57 (6):3961–3984, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Basu:2020:ZSM

[BS20a] Arnab Basu and Lukasz Stettner. Zero-sum Markov games with impulse controls. SIAM Journal on Control and Optimization, 58(1):580–604, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Belomestny:2020:OSM

[BS20b] Denis Belomestny and John Schoenmakers. Optimal stopping of McKean–Vlasov diffusions via regression on particle systems. SIAM Journal on Control and Optimization, 58 (1):529–550, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Belomestny:2024:PDR

[BS24] Denis Belomestny and John Schoenmakers. Primaldual regression approach for Markov decision processes with general state and action spaces. SIAM Journal on Control and Optimization, 62(1): 650–679, February 2024. CO-DEN SJCODC. ISSN 1095-7138.

Banks:1997:ECP

[BSB^{+}97] H. T. Banks, Ralph C. Smith, D. E. Brown, R. J. Silcox, and Vern L. Metcalf. Experimental confirmation of a PDE-based approach to design of feedback controls. SIAM Journal on Control and Optimization, 35(4):1263–1296, July 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28590.

Baumeister:1997:LPE

[BSDR97] J. Baumeister, W. Scondo, M. A. Demetriou, and I. G. Rosen. On-line parameter estimation for infinitedimensional dynamical systems. SIAM Journal on Control and Optimization, 35(2):678–713, March 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27092.

Bai:2019:PCF

[BSI19] Yu Bai, Jason L. Speyer, and Moshe Idan. Properties of the characteristic function generator of the two-state Cauchy estimator. SIAM Journal on Control and Optimization, 57 (6):3639–3665, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Barles:1993:FPP

[BSS93] G. Barles, H. M. Soner, and P. E. Souganidis. Front propagation and phase field theory. SIAM Journal on Control and Optimization, 31(2): 439–469, March 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Butkovic:2012:RWS

[BSS12] Peter Butkovic, Hans Schneider, and Sergei Sergeev. Recognizing weakly stable matrices. SIAM Journal on Control and Optimization, 50(5): 3029–3051, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bensoussan:1984:SPP

[BSVD84] A. Bensoussan, S. P. Sethi, R. Vickson, and N. Derzko. Stochastic production planning with production constraints. SIAM Journal on Control and Optimization, 22 (6):920–935, November 1984.

CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Berkovitz:1993:TWH

[BSZ93] L. D. Berkovitz, Steven E. Shreve, and William P. Ziemer. A tribute to Wendell H. Fleming. SIAM Journal on Control and Optimization, 31 (2):273–281, March 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Burns:2008:MIK

[BSZ08] J. A. Burns, E. W. Sachs, and L. Zietsman. Mesh independence of Kleinman–Newton iterations for Riccati equations in Hilbert space. SIAM Journal on Control and Optimization, 47(5):2663–2692, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Barbu:1991:BCC

[BT91] Viorel Barbu and Dan Tiba. Boundary controllability of the coincidence set in the obstacle problem. SIAM Journal on Control and Optimization, 29(5):1150–1159, September 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bergounioux:1996:GOC

[BT96] Maïtine Bergounioux and Dan Tiba. General optimality conditions for constrained convex control problems. SIAM Journal on Control and Optimization, 34(2):698–711, March 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26198.

Blondel:1997:NHS

[BT97] Vincent Blondel and John N.

Tsitsiklis. NP-Hardness of some linear control design problems. SIAM Journal on Control and Optimization, 35(6):2118–2127, November 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27263.

Birge:1998:SAL

[BT98] John R. Birge and Samer Takriti. Successive approximations of linear control models. SIAM Journal on Control and Optimization, 37 (1):165–176, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29386.

Bolte:2003:BOA

[BT03a] Jérôme Bolte and Marc Teboulle. Barrier operators and associated gradient-like dynamical systems for constrained minimization problems. SIAM Journal on Control and Optimization, 42(4):

1266–1292, January 2003. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Brezhneva:2003:OCD

[BT03b] Olga A. Brezhneva and Alexey A. Tret’yakov. Optimality conditions for degenerate extremum problems with equality constraints. SIAM Journal on Control and Optimization, 42(2):729–745, March 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/38848. See cor-

rigendum [BT10].

Brezhneva:2010:COC

[BT10] Olga Brezhneva and Alexey A. Tret’yakov. Corrigendum: Optimality conditions for degenerate extremum problems with equality constraints. SIAM Journal on Control and Optimization, 48(5):3670– 3673, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [BT03b].

Badra:2011:SPN

[BT11a] Mehdi Badra and Takéo Takahashi. Stabilization of parabolic nonlinear systems with finite dimensional feedback or dynamical controllers: Application to the Navier– Stokes system. SIAM Journal on Control and Optimization, 49(2):420–463, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sicon/
resource/1/sjcodc/v49/i2/
p420_s1.

Bouchard:2011:WDP

[BT11b] Bruno Bouchard and Nizar Touzi. Weak dynamic programming principle for viscosity solutions. SIAM Journal on Control and Optimization, 49(3):948–962, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sicon/
resource/1/sjcodc/v49/i3/
p948_s1.

Bourdin:2013:PMP

[BT13] Loïc Bourdin and Emmanuel Trélat. Pontryagin Maximum Principle for finite dimensional nonlinear optimal control problems on time scales. SIAM Journal on Control and Optimization, 51(5): 3781–3813, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Boulanger:2017:SOC

[BT17] Anne-Celine Boulanger and Philip Trautmann. Sparse optimal control of the KdV– Burgers equation on a bounded domain. SIAM Journal on Control and Optimization, 55 (6):3673–3706, ???? 2017.

CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bourdin:2021:URT

[BT21] Loïc Bourdin and Emmanuel Trélat. Unified Riccati theory for optimal permanent and sampled-data control problems in finite and infinite time horizons. SIAM Journal on Control and Optimization, 59 (1):489–508, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bauso:2016:ODS

[BTB16] D. Bauso, H. Tembine, and T. Basar. Opinion dynamics in social networks through meanfield games. SIAM Journal on Control and Optimization, 54(6):3225–3257, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ben-Tal:1988:DAM

[BTBT88] A. Ben-Tal, J. M. Borwein, and M. Teboulle. A dual approach to multidimensional L_{p} spectral estimation problems. SIAM Journal on Control and Optimization, 26(4): 985–996, July 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Boggs:1982:LCQ

[BTW82] Paul T. Boggs, Jon W. Tolle, and Pyng Wang. On the local convergence of quasi-Newton methods for constrained optimization. SIAM Journal on Control and Optimization, 20 (2):161–171, March 1982. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bao:2011:ASA

[BTY11] Jianhai Bao, Aubrey Truman, and Chenggui Yuan. Almost sure asymptotic stability of stochastic partial differential equations with jumps. SIAM Journal on Control and Optimization, 49(2):771– 787, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i2/p771_s1.

Bucy:1975:SSR

[Buc75] R. S. Bucy. Structural stability for the Riccati equation. SIAM Journal on Control, 13 (4):749–753, July 1975. CO-DEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic). See erratum [Buc77].

Bucy:1977:ESS

[Buc77] R. S. Bucy. Erratum: Structural stability for the Riccati equation. SIAM Journal on Control and Optimization, 15 (5):904, August 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [Buc75].

Bucci:1992:DBC

[Buc92] Francesca Bucci. A Dirichlet boundary control problem for the strongly damped wave equation. SIAM Journal on Control and Optimization, 30(5):1092–1100, September 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Budhiraja:2003:ECP

[Bud03] Amarjit Budhiraja. An ergodic control problem for constrained diffusion processes: Existence of optimal Markov control. SIAM Journal on Control and Optimization, 42(2):532–558, March 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/37907.

Bullo:2001:SEE

[Bul01] Francesco Bullo. Series expansions for the evolution of mechanical control systems. SIAM Journal on Control and Optimization, 40 (1):166–190, January 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/36479.

Bullo:2002:AVC

[Bul02] Francesco Bullo. Averaging and vibrational control of mechanical systems. SIAM Journal on Control and Optimization, 41(2):542–562, March 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36417.

Burke:1991:CEP

[Bur91a] J. V. Burke. Calmness and exact penalization. SIAM Journal on Control and Optimization, 29(2):493–497, March 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Burke:1991:EPV

[Bur91b] James V. Burke. An exact penalization viewpoint of constrained optimization. SIAM Journal on Control and Optimization, 29(4):968–998, July 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Burq:2020:DKV

[Bur20] Nicolas Burq. Decays for Kelvin–Voigt damped wave equations I: The black box perturbative method. SIAM Journal on Control and Optimization, 58(4):1893–1905, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bus:1984:LMR

[Bus84] J. C. P. Bus. The Lagrange multiplier rule on manifolds and optimal control of nonlinear systems. SIAM Journal on Control and Optimization, 22(5):740–757, September 1984. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Boel:1977:OCJ

[BV77] R. Boel and P. Varaiya. Optimal control of jump processes. SIAM Journal on Control and Optimization, 15(1): 92–119, January 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Borkar:1982:IAC

[BV82] Vivek Borkar and Pravin Varaiya. Identification and adaptive control of Markov chains. SIAM Journal on Control and Optimization, 20 (4):470–489, July 1982. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bensoussan:1985:OCP

[Bv85] A. Bensoussan and J. H. van Schuppen. Optimal control of partially observable stochastic systems with an exponentialof-integral performance index. SIAM Journal on Control and Optimization, 23(4):599–613, July 1985. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bettiol:2010:SIC

[BV10] Piernicola Bettiol and Richard B. Vinter. Sensitivity interpretations of the costate variable for optimal control problems with state constraints. SIAM Journal on Control and Optimization, 48(5):3297–3317, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bernicot:2013:CON

[BV13a] Frédéric Bernicot and Juliette Venel. Convergence order of a numerical scheme for sweeping process. SIAM Journal on Control and Optimization, 51(4):3075–3092, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bisiacco:2013:ZTC

[BV13b] Mauro Bisiacco and Maria Elena Valcher. Zero-time-controllability and dead-beat control of twodimensional behaviors. SIAM Journal on Control and Optimization, 51(1):195–220, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bucur:2014:MSO

[BV14] Dorin Bucur and Bozhidar Velichkov. Multiphase shape optimization problems. SIAM Journal on Control and Optimization, 52(6):3556–3591, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bettiol:2017:HJE

[BV17a] Piernicola Bettiol and Richard B.

Vinter. The Hamilton Jacobi equation for optimal control problems with discontinuous time dependence. SIAM Journal on Control and Optimization, 55(2):1199–1225, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Boccia:2017:MPO

[BV17b] A. Boccia and R. B. Vinter. The maximum principle for optimal control problems with time delays. SIAM Journal on Control and Optimization, 55(5):2905–2935, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bank:2018:LQS

[BV18] Peter Bank and Moritz Voß. Linear quadratic stochastic control problems with stochastic terminal constraint. SIAM Journal on Control and Optimization, 56(2):672–699, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bubak:2005:ASR

[BvdMR05] Pavel Bubák, Cornelis V. M. van der Mee, and André C. M. Ran. Approximation of solutions of Riccati equations. SIAM Journal on Control and Optimization, 44(4):1419–1435, July 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43684.

Bouten:2007:IQF

[BVJ07] Luc Bouten, Ramon Van Handel, and Matthew R. James. An introduction to quantum filtering. SIAM Journal on Control and Optimization, 46 (6):2199–2241, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bodin:2000:SBO

[BVW00] Per Bodin, Lars F. Villemoes, and Bo Wahlberg. Selection of best orthonormal rational basis. SIAM Journal on Control and Optimization, 38(4):995–1032, July 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32818.

Boothby:1979:DTB

[BW79] William M. Boothby and Edward N. Wilson. Determination of the transitivity of bilinear systems. SIAM Journal on Control and Optimization, 17 (2):212–221, March 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Birge:1988:SPL

[BW88] John R. Birge and Stein W.

Wallace. A separable piecewise linear upper bound for stochastic linear programs.

SIAM Journal on Control and Optimization, 26(3):725–739, May 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Banks:1989:OFC

[BW89] H. T. Banks and C. Wang. Optimal feedback control of infinite-dimensional parabolic evolution systems: Approximation techniques. SIAM Journal on Control and Optimization, 27(6):1182–1219, November 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bayraktar:2012:ROS

[BX12] Erhan Bayraktar and Hao Xing. Regularity of the optimal stopping problem for jump diffusions. SIAM Journal on Control and Optimization, 50(3):1337–1357, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bayraktar:2013:WDP

[BY13] Erhan Bayraktar and Song Yao. A weak dynamic programming principle for zerosum stochastic differential games with unbounded controls. SIAM Journal on Control and Optimization, 51(3): 2036–2080, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bayraktar:2014:ROS

[BY14] Erhan Bayraktar and Song Yao. On the robust optimal stopping problem. SIAM Journal on Control and Optimization, 52(5):3135–3175, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Brenner:2017:NCA

[ByS17] Susanne C. Brenner and Li yeng Sung. A new convergence analysis of finite element methods for elliptic distributed optimal control problems with pointwise state constraints. SIAM Journal on Control and Optimization, 55 (4):2289–2304, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Borwein:1996:VSV

[BZ96] Jonathan M. Borwein and Qiji J. Zhu. Viscosity solutions and viscosity subderivatives in smooth Banach spaces with applications to metric regularity. SIAM Journal on Control and Optimization, 34(5): 1568–1591, September 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26880.

Bergounioux:1999:PMP

[BZ99a] Maïtine Bergounioux and Housnaa Zidani. Pontryagin maximum principle for optimal control of variational inequalities. SIAM Journal on Control and Optimization, 37(4):1273–1290, July 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32808.

Bonnans:1999:OCP

[BZ99b] J. Frédéric Bonnans and Housnaa Zidani. Optimal control problems with partially polyhedric constraints. SIAM Journal on Control and Optimization, 37(6): 1726–1741, November 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33372.

Bednarczuk:2010:SVV

[BZ10] Ewa M. Bednarczuk and Dariusz Zagrodny. A smooth vector variational principle. SIAM Journal on Control and Optimization, 48(6):3735– 3745, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bayraktar:2015:MPL

[BZ15a] Erhan Bayraktar and Yuchong Zhang. Minimizing the probability of lifetime ruin under ambiguity aversion. SIAM Journal on Control and Optimization, 53(1):58–90, ????

2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Bayraktar:2015:SPM

[BZ15b] Erhan Bayraktar and Yuchong Zhang. Stochastic Perron’s method for the probability of lifetime ruin problem under transaction costs. SIAM Journal on Control and Optimization, 53(1):91–113, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Costa:2009:PSC

[CA09] Eduardo F. Costa and Alessandro Astolfi. Partial stability for a class of nonlinear systems. SIAM Journal on Control and Optimization, 47(6): 3203–3219, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Costa:2010:CED

[CA10] Eduardo F. Costa and Alessandro Astolfi. Characterization of exponential divergence of the Kalman filter for timevarying systems. SIAM Journal on Control and Optimization, 48(5):2917–2944, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Clark:2017:ISP

[CABP17] Andrew Clark, Basel Alomair, Linda Bushnell, and Radha Poovendran. Input selection for performance and controllability of structured linear descriptor systems. SIAM Journal on Control and Optimization, 55(1):457–485, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Calafiore:2009:ACM

[Cal09] Giuseppe Carlo Calafiore. An affine control method for optimal dynamic asset allocation with transaction costs. SIAM Journal on Control and Optimization, 48(4):2254–2274, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Calvia:2018:OCC

[Cal18] A. Calvia. Optimal control of continuous-time Markov chains with noise-free observation. SIAM Journal on Control and Optimization, 56(3): 2000–2035, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Campbell:1976:OCA

[Cam76] Stephen L. Campbell. Optimal control of autonomous linear processes with singular matrices in the quadratic cost functional. SIAM Journal on Control and Optimization, 14(6):1092–1106, November 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cantaluppi:1984:OPC

[Can84] Laurent Cantaluppi. Optimality of piecewise-constant policies in semi-Markov decision chains. SIAM Journal on Control and Optimization, 22(5):723–739, September 1984. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Carter:1976:SRC

[Car76] Thomas E. Carter. Singular regimes in certain classes of relaxed control problems. SIAM Journal on Control and Optimization, 14(2):359– 386, February 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Carlson:1990:ECO

[Car90] Dean A. Carlson. The existence of catching-up optimal solutions for a class of infinite horizon optimal control problems with time delay. SIAM Journal on Control and Optimization, 28(2): 402–422, March 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Carja:1993:MTF

[Câr93] Ovidiu Cârjă. The minimal time function in infinite dimensions. SIAM Journal on Control and Optimization, 31(5):1103–1114, September 1993. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Cardaliaguet:1996:DGT

[Car96] Pierre Cardaliaguet. A differential game with two players and one target. SIAM Journal on Control and Optimization, 34(4):1441–1460, July 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27223.

Cardaliaguet:1997:RSSa

[Car97a] Pierre Cardaliaguet. On the regularity of semipermeable surfaces in control theory with application to the optimal exit-time problem (Part I). SIAM Journal on Control and Optimization, 35(5): 1638–1652, September 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28729.

Cardaliaguet:1997:RSSb

[Car97b] Pierre Cardaliaguet. On the regularity of semipermeable surfaces in control theory with application to the optimal exit-time problem (Part II). SIAM Journal on Control and Optimization, 35(5): 1653–1671, September 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31215.

Cardaliaguet:2007:DGA

[Car07] P. Cardaliaguet. Differential games with asymmetric information. SIAM Journal on Control and Optimization, 46 (3):816–838, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Carja:2012:MTF

[Car12] Ovidiu Carja. The minimum time function for semilinear evolutions. SIAM Journal on Control and Optimization, 50 (3):1265–1282, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Carravetta:2015:GEQ

[Car15] Francesco Carravetta. Global exact quadratization of continuoustime nonlinear control systems. SIAM Journal on Control and Optimization, 53(1): 235–261, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Casas:1986:CEP

[Cas86] Eduardo Casas. Control of an elliptic problem with pointwise state constraints. SIAM Journal on Control and Optimization, 24(6):1309–1318, November 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Casas:1993:BCS

[Cas93] Eduardo Casas. Boundary control of semilinear elliptic equations with pointwise state constraints. SIAM Journal on Control and Optimization, 31 (4):993–1006, July 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Casas:1997:PPS

[Cas97] Eduardo Casas. Pontryagin’s principle for state-constrained boundary control problems of semilinear parabolic equations. SIAM Journal on Control and Optimization, 35(4):1297–1327, July 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/28363.

Casas:2012:SOA

[Cas12] Eduardo Casas. Second order analysis for bang–bang control problems of PDEs. SIAM Journal on Control and Optimization, 50(4):2355–2372, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chasparis:2013:ALC

[CAS13] Georgios C. Chasparis, Ari Arapostathis, and Jeff S. Shamma. Aspiration learning in coordination games. SIAM Journal on Control and Optimization, 51(1):465–490, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Catoni:1998:SSP

[Cat98]  Olivier Catoni.  Solving

scheduling problems by simulated annealing. SIAM Journal on Control and Optimization, 36(5):1539–1575, September 1998. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30781.

Caubet:2013:IIP

[Cau13] Fabien Caubet. Instability of an inverse problem for the stationary Navier–Stokes equations. SIAM Journal on Control and Optimization, 51(4): 2949–2975, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cazacu:2014:CHE

[Caz14] Cristian Cazacu. Controllability of the heat equation with an inverse-square potential localized on the boundary. SIAM Journal on Control and Optimization, 52(4): 2055–2089, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cortes:2005:CGO

[CB05] Jorge Cortés and Francesco Bullo. Coordination and geometric optimization via distributed dynamical systems. SIAM Journal on Control and Optimization, 44(5): 1543–1574, September 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42865.

Carli:2009:QCA

[CB09] Ruggero Carli and Francesco Bullo. Quantized coordination algorithms for rendezvous and deployment. SIAM Journal on Control and Optimization, 48(3):1251–1274, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Claudel:2011:CFD

[CB11] Christian G. Claudel and Alexandre M. Bayen. Convex formulations of data assimilation problems for a class of Hamilton–Jacobi equations. SIAM Journal on Control and Optimization, 49(2):383– 402, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i2/p383_s1.

Coron:2015:DBC

[CB15] Jean-Michel Coron and Georges Bastin. Dissipative boundary conditions for one-dimensional quasi-linear hyperbolic systems: Lyapunov stability for the C^{1}-norm. SIAM Journal on Control and Optimization, 53(3):1464–1483, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:2017:CCP

[CBB17a] Xudong Chen, Mohamed-Ali Belabbas, and Tamer Basar. Cluster consensus with point group symmetries. SIAM Journal on Control and Optimization, 55(6):3869–3889, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:2017:GST

[CBB17b] Xudong Chen, Mohamed-Ali Belabbas, and Tamer Basar. Global stabilization of triangulated formations. SIAM Journal on Control and Optimization, 55(1):172–199, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:2019:CSR

[CBB19] Xudong Chen, Mohamed-Ali Belabbas, and Tamer Ba. Controlling and stabilizing a rigid formation using a few agents. SIAM Journal on Control and Optimization, 57(1):104–128, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Coron:2008:DBC

[CBdN08] Jean-Michel Coron, Georges Bastin, and Brigitte d’Andréa Novel. Dissipative boundary conditions for one-dimensional nonlinear hyperbolic systems. SIAM Journal on Control and Optimization, 47(3):1460– 1498, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cammardella:2023:KLQ

[CBM23] Neil Cammardella, Ana Bušić, and Sean P. Meyn. Kullback– Leibler-quadratic optimal control. SIAM Journal on Control and Optimization, 61(5):3234– 3258, October 2023. CODEN SJCODC. ISSN 1095-7138.

Crouch:1987:OSR

[CC87] P. E. Crouch and P. C. Collingwood. The observation space and realizations of finite Volterra series. SIAM Journal on Control and Optimization, 25(2):316–333, March 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chiang:1988:CRA

[CC88] Tzuu-Shuh Chiang and Yun Shyong Chow. On the convergence rate of annealing processes. SIAM Journal on Control and Optimization, 26(6): 1455–1470, November 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cominetti:1990:GSO

[CC90a] R. Cominetti and R. Correa. A generalized secondorder derivative in nonsmooth optimization. SIAM Journal on Control and Optimization, 28(4):789–809, July 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Culioli:1990:DCA

[CC90b] J.-C. Culioli and G. Cohen.

Decomposition/coordination algorithms in stochastic optimization. SIAM Journal on Control and Optimization, 28(6):1372–1403, November 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chiang:1994:ABS

[CC94] Tzuu Shuh Chiang and Yunshyong Chow. The asymptotic behavior of simulated annealing processes with absorption. SIAM Journal on Control and Optimization, 32(5): 1247–1265, September 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/16653.

Cavazos-Cadena:1996:VIC

[CC96] Rolando Cavazos-Cadena. Value iteration in a class of communicating Markov decision chains with the average cost criterion. SIAM Journal on Control and Optimization, 34(6):1848–1873, November 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/19286.

Cornet:1999:SNA

[CC99] Bernard Cornet and Marc-Olivier Czarnecki. Smooth normal approximations of epi-Lipschitzian subsets of R^{n}. SIAM Journal on Control and Optimization, 37(3):710–730, May 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/32012.

Castro:2001:AAL

[CC01] Carlos Castro and Steven J. Cox. Achieving arbitrarily large decay in the damped wave equation. SIAM Journal on Control and Optimization, 39(6):1748–1755, ???? 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cabot:2002:ACP

[CC02] Alexandre Cabot and Marc-Olivier Czarnecki. Asymptotic control of pairs of oscillators coupled by a repulsion, with nonisolated equilibria I: The regular case. SIAM Journal on Control and Optimization, 41(4):1254–1280, July 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/38519.

Cannarsa:2004:BEM

[CC04] Piermarco Cannarsa and Ovidiu Cârjua. On the Bellman equation for the minimum time problem in infinite dimensions. SIAM Journal on Control and Optimization, 43(2):532–548, March 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41901.

Campi:2013:RCP

[CC13] M. C. Campi and A. Carè.

Random convex programs with L_{1}-regularization: Sparsity and generalization. SIAM Journal on Control and Optimization, 51(5):3532–3557, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Casas:2016:AVT

[CC16] Eduardo Casas and Konstantinos Chrysafinos. Analysis of the velocity tracking control problem for the 3D evolutionary Navier–Stokes equations. SIAM Journal on Control and Optimization, 54(1): 99–128, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cerpa:2018:CIB

[CC18a] Eduardo Cerpa and Emmanuelle Crépeau. On the controllability of the improved Boussinesq equation. SIAM Journal on Control and Optimization, 56(4):3035–3049, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chemetov:2018:ISC

[CC18b] Nikolai Chemetov and Fernanda Cipriano. Injectionsuction control for twodimensional Navier–Stokes equations with slippage. SIAM Journal on Control and Optimization, 56(2):1253–1281, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:2016:MMS

[CCC16] Z. Chen, J.-B. Caillau, and Y. Chitour. L^{1}-minimization for mechanical systems. SIAM Journal on Control and Optimization, 54(3):1245–1265, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:2007:SNF

[CCF07] Anna L. Chen, Gui-Qiang Chen, and Randy A. Freeman. Stability of nonlinear feedback systems: a new small-gain theorem. SIAM Journal on Control and Optimization, 46(6): 1995–2012, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cavalcanti:2020:SSW

[CCFV20] Marcelo Moreira Cavalcanti, Valéria Neves Domingos Cavalcanti, Cícero Lopes Frota, and André Vicente. Stability for semilinear wave equation in an inhomogeneous medium with frictional localized damping and acoustic boundary conditions. SIAM Journal on Control and Optimization, 58 (4):2411–2445, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cavazos-Cadena:2019:VDA

[CCHH19] Rolando Cavazos-Cadena and Daniel Hernández-Hernández. The vanishing discount approach in a class of zero-sum finite games with risk-sensitive average criterion. SIAM Journal on Control and Optimization, 57(1):219–240, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:2012:OTM

[CCHP12] Ting-Li Chen, Wei-Kuo Chen, Chii-Ruey Hwang, and Hui-Ming Pai. On the optimal transition matrix for Markov chain Monte Carlo sampling. SIAM Journal on Control and Optimization, 50 (5):2743–2762, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Casas:2023:NAS

[CCJ23] Eduardo Casas, Alberto Domínguez Corella, and Nicolai Jork. New assumptions for stability analysis in elliptic optimal control problems. SIAM Journal on Control and Optimization, 61(3):1394–1414, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M149199X.

Casas:2012:AEC

[CCK12] Eduardo Casas, Christian Clason, and Karl Kunisch. Approximation of elliptic control problems in measure spaces with sparse solutions. SIAM Journal on Control and Optimization, 50(4):1735–1752, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Casas:2013:PCP

[CCK13] Eduardo Casas, Christian Clason, and Karl Kunisch. Parabolic control problems in measure spaces with sparse solutions. SIAM Journal on Control and Optimization, 51 (1):28–63, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Castro:2014:CLO

[CCM14] Carlos Castro, Nicolae Cîndea, and Arnaud Münch. Controllability of the linear onedimensional wave equation with inner moving forces. SIAM Journal on Control and Optimization, 52(6):4027– 4056, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Clement:2013:OCC

[CCS13] Frédérique Clément, Jean-Michel Coron, and Peipei Shang. Optimal control of Cell mass and maturity in a model of follicular ovulation. SIAM Journal on Control and Optimization, 51(2):824–847, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Colonius:2018:IPC

[CCS18] Fritz Colonius, João A. N.

Cossich, and Alexandre J. Santana. Invariance pressure of control sets. SIAM Journal on Control and Optimization, 56(6):4130–4147, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:2005:OCO

[CCT05] Qihong Chen, Delin Chu, and Roger C. E. Tan. Optimal control of obstacle for quasilinear elliptic variational bilateral problems. SIAM Journal on Control and Optimization, 44(3):1067–1080, May 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/44307.

Chen:2007:BOC

[CCT07] Qihong Chen, Delin Chu, and Roger C. E. Tan. Bilateral obstacle control problem of parabolic variational inequalities. SIAM Journal on Control and Optimization, 46(4): 1518–1537, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:2014:NTS

[CCW14] Xin Chen, Boumediène Chentouf, and Jun-Min Wang. Nondissipative torque and shear force controls of a rotating flexible structure. SIAM Journal on Control and Optimization, 52(5):3287–3311, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cryer:1980:ELC

[CD80] C. W. Cryer and M. A. H. Dempster. Equivalence of linear complementarity problems and linear programs in vector lattice Hilbert spaces. SIAM Journal on Control and Optimization, 18(1):76–90, January 1980. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cannarsa:1989:NOC

[CD89] P. Cannarsa and G. Da Prato. Nonlinear optimal control with infinite horizon for distributed parameter systems and stationary Hamilton–Jacobi equations. SIAM Journal on Control and Optimization, 27(4):861–875, July 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cannarsa:1991:SOH

[CD91] Piermarco Cannarsa and Giuseppe Da Prato. Secondorder Hamilton–Jacobi equations in infinite dimensions. SIAM Journal on Control and Optimization, 29(2):474– 492, March 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Costa:2003:PEP

[CD03] Oswaldo L. V. Costa and François Dufour. On the Poisson equation for piecewisedeterministic Markov processes. SIAM Journal on Control and Optimization, 42(3):985–1001, May 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39352.

Costa:2008:SEP

[CD08] O. L. V. Costa and F. Dufour. Stability and ergodicity of piecewise deterministic Markov processes. SIAM Journal on Control and Optimization, 47(2):1053–1077, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Costa:2010:ACC

[CD10] O. L. V. Costa and F. Dufour. Average continuous control of piecewise deterministic Markov processes. SIAM Journal on Control and Optimization, 48(7):4262–4291, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Costa:2012:SPD

[CD12] O. L. V. Costa and F. Dufour. Singularly perturbed discounted Markov control processes in a general state space. SIAM Journal on Control and Optimization, 50(2):720–747, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Caillau:2013:MTC

[CD13a] J.-B. Caillau and B. Daoud. Minimum time control of the restricted three-body problem. SIAM Journal on Control and Optimization, 50(6): 3178–3202, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Carmona:2013:PAM

[CD13b] René Carmona and François Delarue. Probabilistic analysis of mean-field games. SIAM Journal on Control and Optimization, 51(4):2705–2734, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Casado-Diaz:2015:SPO

[CD15] Juan Casado-Díaz. Smoothness properties for the optimal mixture of two isotropic materials: The compliance and eigenvalue problems. SIAM Journal on Control and Optimization, 53(4):2319–2349, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cetin:2021:PRO

[ÇD21] Umut Çetin and Albina Danilova. On pricing rules and optimal strategies in general Kyle–Back models. SIAM Journal on Control and Optimization, 59(5):3973–3998, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Casado-Diaz:2004:OCN

[CDCCMG04] Juan Casado-Díaz, Julio Couce-Calvo, and José D. Martín-Gómez. Optimality conditions for nonconvex multistate control problems in the coefficients. SIAM Journal on Control and Optimization, 43(1):216–239, January 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41171.

Casado-Diaz:2008:RCP

[CDCCMG08] Juan Casado-Díaz, Julio Couce-Calvo, and José D. Martín-Gómez. Relaxation of a control problem in the coefficients with a functional of quadratic growth in the gradient. SIAM Journal on Control and Optimization, 47(3): 1428–1459, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Casado-Diaz:2021:MLE

[CDCVV21] Juan Casado-Diaz, Carlos Conca, and Donato Vásquez-Varas. The maximization of the p-laplacian energy for a two-phase material. SIAM Journal on Control and Optimization, 59(2):1497–1519, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chaudet-Dumas:2020:SDP

[CDD20] Bastien Chaudet-Dumas and Jean Deteix. Shape derivatives for the penalty formulation of elastic contact problems with Tresca friction. SIAM Journal on Control and Optimization, 58(6):3237–3261, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Colombo:2020:GOT

[CdDNdA20] Leonardo J. Colombo, David Martín de Diego, Aradhana Nayak, and Rodrigo T. Sato Martín de Almagro. Geometric optimal trajectory tracking of nonholonomic mechanical systems. SIAM Journal on Control and Optimization, 58(3): 1652–1675, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Caron:2011:PAI

[CDDP11] François Caron, Pierre Del Moral, Arnaud Doucet, and Michele Pace. Particle approximation of the intensity measures of a spatial branching point process arising in multitarget tracking. SIAM Journal on Control and Optimization, 49(4):1766– 1792, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i4/p1766_s1.

Cecchin:2019:CPM

[CDFP19] Alekos Cecchin, Paolo Dai Pra, Markus Fischer, and Guglielmo Pelino. On the convergence problem in mean field games: a two state model without uniqueness. SIAM Journal on Control and Optimization, 57(4):2443–2466, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cibulka:2015:INM

[CDG15] R. Cibulka, A. Dontchev, and M. H. Geoffroy. Inexact Newton methods and Dennis–Moré theorems for nonsmooth generalized equations. SIAM Journal on Control and Optimization, 53(2):1003–1019, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cardaliaguet:2023:ACR

[CDJS23] Pierre Cardaliaguet, Samuel Daudin, Joe Jackson, and Panagiotis E. Souganidis. An algebraic convergence rate for the optimal control of McKean–Vlasov dynamics. SIAM Journal on Control and Optimization, 61(6):3341– 3369, November 2023. CO-DEN SJCODC. ISSN 1095-7138.

Chen:1987:MSC

[CDKP87] G. Chen, M. C. Delfour, A. M. Krall, and G. Payre. Modeling, stabilization and control of serially connected beams. SIAM Journal on Control and Optimization, 25(3):526–546, May 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cibulka:2018:MRD

[CDKV18] R. Cibulka, A. Dontchev, M. I. Krastanov, and V. M. Veliov. Metrically regular differential generalized equations. SIAM Journal on Control and Optimization, 56(1):316–342, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chouinard:1985:PMU

[CDL85] L. G. Chouinard, J. P. Dauer, and G. Leitmann. Properties of matrices used in uncertain linear control systems. SIAM Journal on Control and Optimization, 23(3):381–389, May 1985. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cortes:2002:GDV

[CdLdDM02] J. Cortés, M. de León, D. Martín de Diego, and S. Martínez. Geometric description of Vakonomic and nonholonomic dynamics. Comparison of solutions. SIAM Journal on Control and Optimization, 41(5): 1389–1412, September 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36817.

Cheng:1990:OSL

[CDM90] D. Cheng, W. P. Dayawansa, and C. F. Martin. Observability of systems on Lie groups and coset spaces. SIAM Journal on Control and Optimization, 28(3):570–581, May 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Caubet:2021:SDS

[CDM21a] Fabien Caubet, Marc Dambrine, and Rajesh Mahadevan. Shape derivative for some eigenvalue functionals in elasticity theory. SIAM Journal on Control and Optimization, 59(2): 1218–1245, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chowdhury:2021:BSL

[CDM21b] Shirshendu Chowdhury, Rajib Dutta, and Subrata Majumdar. Boundary stabilizability of the linearized compressible Navier–Stokes system in one dimension by backstepping approach. SIAM Journal on Control and Optimization, 59(3):2147–2173, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cordoni:2023:WES

[CDM23] Francesco Cordoni, Luca Di Persio, and Riccardo Muradore. Weak energy shaping for stochastic controlled port-Hamiltonian systems. SIAM Journal on Control and Optimization, 61(5):2902–2926, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1482585.

Clarke:2010:OCP

[CdP10] Francis Clarke and M. R. de Pinho. Optimal control problems with mixed constraints. SIAM Journal on Control and Optimization, 48 (7):4500–4524, ???? 2010.

CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Costa:2016:CUO

[CDP16] O. L. V. Costa, F. Dufour, and A. B. Piunovskiy. Constrained and unconstrained optimal discounted control of piecewise deterministic Markov processes. SIAM Journal on Control and Optimization, 54(3): 1444–1474, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cai:2023:COS

[CDP23] Cheng Cai, Tiziano De Angelis, and Jan Palczewski. On the continuity of optimal stopping surfaces for jumpdiffusions. SIAM Journal on Control and Optimization, 61 (3):1513–1531, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1448094.

Cattiaux:2008:CAC

[CDR08] Patrick Cattiaux, Paolo Dai Pra, and Sylvie Rœlly. A constructive approach to a class of ergodic HJB equations with unbounded and nonsmooth cost. SIAM Journal on Control and Optimization, 47(5):2598–2615, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Coutinho:2009:RLF

[CdSBT09] Daniel F. Coutinho, Carlos E. de Souza, Karina A. Barbosa, and Alexandre Trofino. Robust linear H_{∞} filter design for a class of uncertain nonlinear systems: An LMI approach. SIAM Journal on Control and Optimization, 48 (3):1452–1472, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Costa:2002:ODC

[CdV02] Eduardo F. Costa and João B. R. do Val. On the observability and detectability of continuous-time Markov jump linear systems. SIAM Journal on Control and Optimization, 41(4):1295–1314, July 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/38546.

Cibulka:2016:LGT

[CDV16] R. Cibulka, A. Dontchev, and V. M. Veliov. Lyusternik– Graves theorems for the sum of a Lipschitz function and a set-valued mapping. SIAM Journal on Control and Optimization, 54(6):3273–3296, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Costa:2005:NAD

[CdVF05] Eduardo F. Costa, João B. R. do Val, and Marcelo D.

Fragoso. A new approach to detectability of discretetime infinite Markov jump linear systems. SIAM Journal on Control and Optimization, 43(6):2132–2156, November 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42992.

Cioranescu:1994:ABC

[CDZ94] Doina Cioranescu, Patrizia Donato, and Enrique Zuazua. Approximate boundary controllability for the wave equation in perforated domains. SIAM Journal on Control and Optimization, 32(1):35–50, January 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/22229.

Charalambous:1998:CNP

[CE98] Charalambos D. Charalambous and Robert J. Elliott. Classes of nonlinear partially observable stochastic optimal control problems with explicit optimal control laws. SIAM Journal on Control and Optimization, 36(2):542–578, March 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28732.

Cobb:2006:HGS

[CE06] Daniel Cobb and Jacob Eapen. High-gain state feedback analysis based on singular system theory. SIAM Journal on Control and Optimization, 44(6):2210–2232, January 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cohen:2011:BSD

[CE11] Samuel N. Cohen and Robert J. Elliott. Backward stochastic difference equations and nearly time-consistent nonlinear expectations. SIAM Journal on Control and Optimization, 49(1):125–139, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sicon/
                                        resource/1/sjcodc/v49/i1/
                                        p125_s1.

Charalambous:2003:CMG

[CEK03] C. D. Charalambous, R. J. Elliott, and V. Krishnamurthy. Conditional moment generating functions for integrals and stochastic integrals. SIAM Journal on Control and Optimization, 42(5): 1578–1603, September 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/33327.

Cellina:2002:CDP

[Cel02a]  Arrigo Cellina.  On a

constrained Dirichlet problem. SIAM Journal on Control and Optimization, 41(2):331–344, March 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38042.

Cellina:2002:BSC

[Cel02b]  Arrigo Cellina.  On the

bounded slope condition and the validity of the Euler– Lagrange equation. SIAM Journal on Control and Optimization, 40(4):1270–1279, July 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35466.

Cellina:2007:CRE

[Cel07] A. Cellina. Comparison results and estimates on the gradient without strict convexity. SIAM Journal on Control and Optimization, 46(2): 738–749, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cellina:2015:CRS

[Cel15] Arrigo Cellina. A case of regularity of solutions to nonregular problems. SIAM Journal on Control and Optimization, 53(5):2835–2845, ???? 2015.

CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cerou:2000:LTB

[Cér00] Frédéric Cérou. Long time behavior for some dynamical noise free nonlinear filtering problems. SIAM Journal on Control and Optimization, 38(4):1086–1101, July 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29012.

Cerrai:2001:OCP

[Cer01a] Sandra Cerrai. Optimal control problems for stochastic reaction-diffusion systems with non-Lipschitz coefficients. SIAM Journal on Control and Optimization, 39(6): 1779–1816, ???? 2001. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cerrai:2001:SHJ

[Cer01b] Sandra Cerrai. Stationary Hamilton–Jacobi equations in Hilbert spaces and applications to a stochastic optimal control problem. SIAM Journal on Control and Optimization, 40(3):824–852, May 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35994.

Cerpa:2007:ECN

[Cer07] Eduardo Cerpa. Exact controllability of a nonlinear Korteweg–de Vries equation on a critical spatial domain. SIAM Journal on Control and Optimization, 46(3):877–899, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cannarsa:1991:SCO

[CF91] Piermarco Cannarsa and Halina Frankowska. Some characterizations of optimal trajectories in control theory. SIAM Journal on Control and Optimization, 29(6):1322– 1347, November 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Casas:1995:DIS

[CF95] Eduardo Casas and Luis A.

Fernández. Dealing with integral state constraints in boundary control problems of quasilinear elliptic equations. SIAM Journal on Control and Optimization, 33(2):568–589, March 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/23463.

Cernea:2005:CBM

[CF05] Aurelian Cernea and Hélène Frankowska. A connection between the maximum principle and dynamic programming for constrained control problems. SIAM Journal on Control and Optimization, 44(2):673–703, March 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43058.

Colaneri:2006:ARE

[CF06] Patrizio Colaneri and Augusto Ferrante. Algebraic Riccati equation and J-spectral factorization for H_{∞} smoothing and deconvolution. SIAM Journal on Control and Optimization, 45(1):123–145, January 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Confortola:2013:BSD

[CF13] Fulvia Confortola and Marco Fuhrman. Backward stochastic differential equations and optimal control of marked point processes. SIAM Journal on Control and Optimization, 51(5):3592–3623, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chiarolla:2014:IFB

[CF14] Maria B. Chiarolla and Giorgio Ferrari. Identifying the free boundary of a stochastic, irreversible investment problem via the Bank–El karoui representation theorem. SIAM Journal on Control and Optimization, 52(2):1048–1070, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ceragioli:2018:CDR

[CF18] Francesca Ceragioli and Paolo Frasca. Consensus and disagreement: The role of quantized behaviors in opinion dynamics. SIAM Journal on Control and Optimization, 56 (2):1058–1080, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Camasta:2024:SRD

[CF24] Alessandro Camasta and Genni Fragnelli. A stability result for a degenerate beam equation. SIAM Journal on Control and Optimization, 62(1):630–649, February 2024. CODEN SJCODC. ISSN 1095-7138.

Calvia:2021:SCC

[CFG21] Alessandro Calvia, Salvatore Federico, and Fausto Gozzi. State constrained control problems in Banach lattices and applications. SIAM Journal on Control and Optimization, 59(6):4481–4510, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Carmona:2009:OSC

[CFH09] René Carmona, Max Fehr, and Juri Hinz. Optimal stochastic control and carbon price formation. SIAM Journal on Control and Optimization, 48 (4):2168–2190, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cannarsa:2019:NCP

[CFM19] Piermarco Cannarsa, Roberto Ferretti, and Patrick Martinez. Null controllability for parabolic operators with interior degeneracy and one-sided control. SIAM Journal on Control and Optimization, 57 (2):900–924, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Calzolari:2006:ANF

[CFN06] Antonella Calzolari, Patrick Florchinger, and Giovanna Nappo. Approximation of nonlinear filters for Markov systems with delayed observations. SIAM Journal on Control and Optimization, 45(2): 599–633, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Calzolari:2007:CNF

[CFN07] Antonella Calzolari, Patrick Florchinger, and Giovanna Nappo. Convergence in nonlinear filtering for stochastic delay systems. SIAM Journal on Control and Optimization, 46(5):1615–1636, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:2010:EPAa

[CFNG10a] Jie Chen, Peilin Fu, Silviu-Iulian Niculescu, and Zhihong Guan. An eigenvalue perturbation approach to stability analysis, Part I: Eigenvalue series of matrix operators. SIAM Journal on Control and Optimization, 48(8): 5564–5582, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v48/i8/p5564_s1.

Chen:2010:EPAb

[CFNG10b] Jie Chen, Peilin Fu, Silviu-Iulian Niculescu, and Zhihong Guan. An eigenvalue perturbation approach to stability analysis, Part II: When will zeros of time-delay systems cross imaginary axis? SIAM Journal on Control and Optimization, 48(8):5583– 5605, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v48/i8/p5583_s1.

Chiarolla:2013:GKT

[CFR13] Maria B. Chiarolla, Giorgio Ferrari, and Frank Riedel. Generalized Kuhn–Tucker conditions for N-firm stochastic irreversible investment under limited resources. SIAM Journal on Control and Optimization, 51(5):3863–3885, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cedio-Fengya:1998:DIB

[CFRV98] D. J. Cedio-Fengya, M. V. Romerio, and M. Vogelius. Determination of the interface between the fluids in a Hall–Héroult cell from measurements of electric potentials and currents on the electrodes. SIAM Journal on Control and Optimization, 37 (1):245–277, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/30632.

Cannarsa:2015:SOS

[CFS15] P. Cannarsa, H. Frankowska, and T. Scarinci. Second-order sensitivity relations and regularity of the value function for Mayer’s problem in optimal control. SIAM Journal on Control and Optimization, 53(6):3642–3672, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Colli:2015:OBC

[CFSGS15] Pierluigi Colli, M. Hassan Farshbaf-Shaker, Gianni Gilardi, and Jürgen Sprekels. Optimal boundary control of a viscous Cahn–Hilliard system with dynamic boundary condition and double obstacle potentials. SIAM Journal on Control and Optimization, 53(4):2696–2721, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Canuto:2012:EAA

[CFT12] C. Canuto, F. Fagnani, and P. Tilli. An Eulerian approach to the analysis of Krause’s consensus models. SIAM Journal on Control and Optimization, 50(1):243– 265, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v50/i1/p243_s1.

Como:2010:ART

[CFZ10] Giacomo Como, Fabio Fagnani, and Sandro Zampieri. Anytime reliable transmission of real-valued information through digital noisy channels. SIAM Journal on Control and Optimization, 48(6): 3903–3924, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Caicedo:2017:NBC

[CFZ17] Miguel Andres Caicedo, Roberto Capistrano Filho, and Bing-Yu Zhang. Neumann boundary controllability of the Korteweg–de Vries equation on a bounded domain. SIAM Journal on Control and Optimization, 55(6):3503–3532, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:2022:ITA

[CFZ22] Jian-Hua Chen, Lin Fu, and Hua-Cheng Zhou. Infinitetime admissibility of the Gurtin–Pipkin systems in Hilbert spaces. SIAM Journal on Control and Optimization, 60(1):505–529, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1403849.

Chen:1987:AOA

[CG87a] H. F. Chen and L. Guo. Asymptotically optimal adaptive control with consistent parameter estimates. SIAM Journal on Control and Optimization, 25(3):558–575, May 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:1987:OAC

[CG87b] Han Fu Chen and Lei Guo. Optimal adaptive control and consistent parameter estimates for ARMAX model with quadratic cost. SIAM Journal on Control and Optimization, 25(4):845–867, July 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:1990:CTS

[CG90] Han Fu Chen and Lei Guo.

Continuous-time stochastic adaptive tracking — robustness and asymptotic properties. SIAM Journal on Control and Optimization, 28(3): 513–527, March 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Corless:1992:ESS

[CG92] Martin Corless and Luigi Glielmo. On the exponential stability of singularly perturbed systems. SIAM Journal on Control and Optimization, 30(6):1338–1360, November 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cantoni:2000:ERL

[CG00a] Michael Cantoni and Keith Glover. Existence of right and left representations of the graph for linear periodically time-varying systems. SIAM Journal on Control and Optimization, 38(3):786–802, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34660.

Cantoni:2000:GMR

[CG00b] Michael Cantoni and Keith Glover. Gap-metric robustness analysis of linear periodically time-varying feedback systems. SIAM Journal on Control and Optimization, 38(3):803–822, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34659.

Chai:2010:FOL

[CG10] Shugen Chai and Bao-Zhu Guo. Feedthrough operator for linear elasticity system with boundary control and observation. SIAM Journal on Control and Optimization, 48(6): 3708–3734, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:2013:ICM

[CG13] Yann-Shin Aaron Chen and Xin Guo. Impulse control of multidimensional jump diffusions in finite time horizon. SIAM Journal on Control and Optimization, 51(3): 2638–2663, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Coclite:2017:TDO

[CG17a] G. M. Coclite and M. Garavello. A time-dependent optimal harvesting problem with measure-valued solutions. SIAM Journal on Control and Optimization, 55(2): 913–935, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Coron:2017:CTH

[CG17b] Jean-Michel Coron and Jean-Philippe Guilleron. Control of three heat equations coupled with two cubic nonlinearities. SIAM Journal on Control and Optimization, 55(2):989–1019, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Coclite:2022:MOC

[CG22] Giuseppe M. Coclite and Mauro Garavello. Measure optimal controls for models inspired by biology. SIAM Journal on Control and Optimization, 60(5):3051–3077, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M146332X.

Cherukuri:2017:SPD

[CGC17] Ashish Cherukuri, Bahman Gharesifard, and Jorge Cortés. Saddle-point dynamics: Conditions for asymptotic stability of saddle points. SIAM Journal on Control and Optimization, 55(1):486–511, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Colombo:2009:OCN

[CGHS09] R. M. Colombo, G. Guerra, M. Herty, and V. Schleper. Optimal control in networks of pipes and canals. SIAM Journal on Control and Optimization, 48(3):2032–2050, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Carravetta:2002:FNS

[CGLM02] F. Carravetta, A. Germani, R. Liptser, and C. Manes. Filtering of nonlinear stochastic feedback systems. SIAM Journal on Control and Optimization, 40(5):1576–1584, September 2002. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34714.

Canada:1998:SOC

[CGM98] A. Cañada, J. L. Gámez, and J. A. Montero. Study of an optimal control problem for diffusive nonlinear elliptic equations of logistic type. SIAM Journal on Control and Optimization, 36(4):1171–1189, July 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29332.

Casas:2011:PAD

[CGM11] Eduardo Casas, Andreas Günther, and Mariano Mateos. A paradox in the approximation of Dirichlet control problems in curved domains. SIAM Journal on Control and Optimization, 49(5):

1998–2007, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i5/p1998_s1.

Cacace:2014:CON

[CGM14] Filippo Cacace, Alfredo Germani, and Costanzo Manes. A chain observer for nonlinear systems with multiple time-varying measurement delays. SIAM Journal on Control and Optimization, 52(3): 1862–1885, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:2022:MLE

[CGP22] Yongxin Chen, Tryphon T.

Georgiou, and Michele Pavon. The most likely evolution of diffusing and vanishing particles: Schrödinger bridges with unbalanced marginals. SIAM Journal on Control and Optimization, 60(4):2016–2039, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1447672.

Carravetta:1996:PFL

[CGR96] Francesco Carravetta, Alfredo Germani, and Massimo Raimondi. Polynomial filtering for linear discrete time non-Gaussian systems. SIAM Journal on Control and Optimization, 34(5):

1666–1690, September 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/25289.

Coron:2010:NCP

[CGR10] Jean-Michel Coron, Sergio Guerrero, and Lionel Rosier. Null controllability of a parabolic system with a cubic coupling term. SIAM Journal on Control and Optimization, 48(8):5629– 5653, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicon/resource/
                                        1/sjcodc/v48/i8/p5629_s1.

Carravetta:2000:NSA

[CGS00] Francesco Carravetta, Alfredo Germani, and Marat K. Shuakayev. A new suboptimal approach to the filtering problem for bilinear stochastic differential systems. SIAM Journal on Control and Optimization, 38(4):1171–1203, July 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/32091.

Colli:2018:OVC

[CGS18] Pierluigi Colli, Gianni Gilardi, and Jürgen Sprekels. Optimal velocity control of a viscous Cahn–Hilliard system with convection and dynamic boundary conditions. SIAM Journal on Control and Optimization, 56(3):1665–1691, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Camilli:2001:GZM

[CGW01] Fabio Camilli, Lars Grüne, and Fabian Wirth. A generalization of Zubov’s method to perturbed systems. SIAM Journal on Control and Optimization, 40(2):496–515, March 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36316.

Camilli:2008:CLF

[CGW08] Fabio Camilli, Lars Grüne, and Fabian Wirth. Control Lyapunov functions and Zubov’s method. SIAM Journal on Control and Optimization, 47(1):301–326, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Coron:2010:EBC

[CGW10] Jean-Michel Coron, Oliver Glass, and Zhiqiang Wang. Exact boundary controllability for 1-D quasilinear hyperbolic systems with a vanishing characteristic speed. SIAM Journal on Control and Optimization, 48(5):3105–3122, ???? 2010. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Coron:2020:MPC

[CGW20] Jean-Michel Coron, Lars Grüne, and Karl Worthmann. Model predictive control, cost controllability, and homogeneity. SIAM Journal on Control and Optimization, 58(5): 2979–2996, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chai:2003:BFS

[CGY03] Shugen Chai, Yuxia Guo, and Peng-Fei Yao. Boundary feedback stabilization of shallow shells. SIAM Journal on Control and Optimization, 42(1):239–259, January 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/39715.

Chen:2021:FBP

[CGY21] Xiaoshan Chen, Chonghu Guan, and Fahuai Yi. A free boundary problem of liquidity management for optimal dividend and insurance in finite horizon. SIAM Journal on Control and Optimization, 59(4):2524–2545, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Colonius:1978:OCF

[CH78] F. Colonius and D. Hinrichsen. Optimal control of functional differential systems. SIAM Journal on Control and Optimization, 16(6): 861–879, November 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chiarolla:1994:FBM

[CH94] Maria B. Chiarolla and Ulrich G. Haussmann. The free boundary of the monotone follower. SIAM Journal on Control and Optimization, 32(3):690–727, May 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23237.

Carlson:1996:TTI

[CH96] D. Carlson and A. Haurie. A turnpike theory for infinitehorizon open-loop competitive processes. SIAM Journal on Control and Optimization, 34(4):1405–1419, July 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26515.

Chiarolla:1998:OCI

[CH98] Maria B. Chiarolla and Ulrich G. Haussmann. Optimal control of inflation: a central bank problem. SIAM Journal on Control and Optimization, 36(3):1099–1132, May 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30495.

Charalambous:1999:NCO

[CH99] Charalambos D. Charalambous and Joseph L. Hibey. Necessary conditions of optimization for partially observed controlled diffusions. SIAM Journal on Control and Optimization, 37(6): 1676–1700, November 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25937.

Chu:2005:CLF

[CH05] Delin Chu and Daniel W. C. Ho. Computation of the (J, J′)-lossless factorization for general rational matrices. SIAM Journal on Control and Optimization, 44 (1):240–267, January 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/60913.

Chiarolla:2009:SII

[CH09] Maria B. Chiarolla and Ulrich G. Haussmann. On a stochastic, irreversible investment problem. SIAM Journal on Control and Optimization, 48(2):438–462, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chung:2012:EDB

[CH12] Matthias Chung and Eldad Haber. Experimental design for biological systems. SIAM Journal on Control and Optimization, 50(1):471–489, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cohen:2013:EBD

[CH13] Samuel N. Cohen and Ying Hu. Ergodic BSDEs driven by Markov chains. SIAM Journal on Control and Optimization, 51(5):4138–4168, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Caines:2021:GMF

[CH21a] Peter E. Caines and Minyi Huang. Graphon mean field games and their equations. SIAM Journal on Control and Optimization, 59(6):4373– 4399, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Colonius:2021:EPS

[CH21b] Fritz Colonius and Boumediene Hamzi. Entropy for practical stabilization. SIAM Journal on Control and Optimization, 59(3):2195–2222, ???? 2021. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Chaney:1977:HAN

[Cha77] R. W. Chaney. A hybrid algorithm for nonlinear programming. SIAM Journal on Control and Optimization, 15 (3):384–399, May 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chaney:1982:SOS

[Cha82]  R. W. Chaney.  Second-

order sufficiency conditions for nondifferentiable programming problems. SIAM Journal on Control and Optimization, 20(1):20–33, January 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chaney:1987:SON

[Cha87a] Robin W. Chaney. Secondorder necessary conditions in constrained semismooth optimization. SIAM Journal on Control and Optimization, 25 (4):1072–1081, July 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chaplais:1987:ADO

[Cha87b] F. Chaplais. Averaging and deterministic optimal control. SIAM Journal on Control and Optimization, 25(3):767–780, May 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chavent:1991:NSC

[Cha91]  Guy Chavent.  New size

× curvature conditions for strict quasiconvexity of sets. SIAM Journal on Control and Optimization, 29(6):1348– 1372, November 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chaves:2005:ISS

[Cha05] Madalena Chaves. Inputto-state stability of ratecontrolled biochemical networks. SIAM Journal on Control and Optimization, 44(2):704–727, March 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43796.

Chapouly:2009:GCN

[Cha09] Marianne Chapouly. Global controllability of nonviscous and viscous Burgers-type equations. SIAM Journal on Control and Optimization, 48 (3):1567–1599, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chang:2010:MCL

[Cha10] Dong Eui Chang. The method of controlled Lagrangians: Energy plus force shaping. SIAM Journal on Control and Optimization, 48(8):4821–4845, ???? 2010. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Chazelle:2011:TEM

[Cha11] Bernard Chazelle. The total s-energy of a multiagent system. SIAM Journal on Control and Optimization, 49(4): 1680–1706, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicon/resource/
                                        1/sjcodc/v49/i4/p1680_s1.

Chattopadhyay:2014:SOD

[Cha14] Ishanu Chattopadhyay. Scalable ϵ-optimal decision-making and stochastic routing in large networks via distributed supervision of probabilistic automata. SIAM Journal on Control and Optimization, 52 (4):2512–2547, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chewning:1976:CNW

[Che76] William C. Chewning. Controllability of the nonlinear wave equation in several space variables. SIAM Journal on Control and Optimization, 14 (1):19–25, January 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:1979:CSW

[Che79] Goong Chen. Control and stabilization for the wave equation in a bounded domain. SIAM Journal on Control and Optimization, 17(1): 66–81, January 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:1981:CSW

[Che81a] Goong Chen. Control and stabilization for the wave equation in a bounded domain, Part II. SIAM Journal on Control and Optimization, 19 (1):114–122, January 1981. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:1981:NBS

[Che81b] Goong Chen. A note on the boundary stabilization of the wave equation. SIAM Journal on Control and Optimization, 19(1):106–113, January 1981. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:1984:RSI

[Che84] H. F. Chen. Recursive system identification and adaptive control by use of the modified least squares algorithm. SIAM Journal on Control and Optimization, 22 (5):758–776, September 1984. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:1996:CBM

[Che96] Xiaojun Chen. Convergence of the BFGS method for LC^{1} convex constrained optimization. SIAM Journal on Control and Optimization, 34(6): 2051–2063, November 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27482.

Chen:1998:CRS

[Che98] Han-Fu Chen. Convergence rate of stochastic approximation algorithms in the degenerate case. SIAM Journal on Control and Optimization, 36(1):100–114, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28173.

Chen:1999:IOC

[Che99] Qihong Chen. Indirect obstacle control problem for semilinear elliptic variational inequalities. SIAM Journal on Control and Optimization, 38 (1):138–158, January 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34337.

Chen:2004:FOC

[Che04] Xiaojun Chen. First order conditions for nonsmooth discretized constrained optimal control problems. SIAM Journal on Control and Optimization, 42(6):2004–2015, November 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41416.

Chen:2016:AOO

[Che16] Jian-Hua Chen. Admissibility of the observation operators for Volterra systems with exponential kernels. SIAM Journal on Control and Optimization, 54(2):587–601, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:2023:CIL

[Che23] Xudong Chen. Controllability issues of linear ensemble systems over multidimensional parameterization spaces. SIAM Journal on Control and Optimization, 61 (4):2425–2447, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1418691.

Carlson:1987:EOS

[CHJ87] D. A. Carlson, A. Haurie, and A. Jabrane. Existence of overtaking solutions to infinite dimensional control problems on unbounded time intervals. SIAM Journal on Control and Optimization, 25(6):1517– 1541, November 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chevalier:2015:EAC

[CHJ15] P.-Y. Chevalier, J. M. Hendrickx, and R. M. Jungers. Efficient algorithms for the consensus decision problem. SIAM Journal on Control and Optimization, 53(5):3104– 3119, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chatelon:1982:SAC

[CHL82] Jacques Chatelon, Donald Hearn, and Timothy J. Lowe. A subgradient algorithm for certain minimax and minisum problems — the constrained case. SIAM Journal on Control and Optimization, 20(4): 455–469, July 1982. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chitour:2015:SIC

[CHL15] Yacine Chitour, Mohamed Harmouche, and Salah Laghrouche. L_{p}-stabilization of integrator chains subject to input saturation using Lyapunov-based homogeneous design. SIAM Journal on Control and Optimization, 53(4):2406–2423, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chan:1994:GSO

[CHN94] W. L. Chan, L. R. Huang, and K. F. Ng. On generalized second-order derivatives and Taylor expansions in nonsmooth optimization. SIAM Journal on Control and Optimization, 32(3):591–611, May 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/22742.

Conti:2011:RAS

[CHP^{+}11] Sergio Conti, Harald Held, Martin Pach, Martin Rumpf, and Rüdiger Schultz. Risk averse shape optimization. SIAM Journal on Control and Optimization, 49(3):927– 947, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i3/p927_s1.

Correa:2020:SCR

[CHPA20] Rafael Correa, Abderrahim Hantoute, and Pedro Pérez-Aros. Subdifferential calculus rules for possibly nonconvex integral functions. SIAM Journal on Control and Optimization, 58(1):462–484, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Christopeit:1977:NOC

[Chr77] Norbert Christopeit. Necessary optimality conditions with application to a variational problem. SIAM Journal on Control and Optimization, 15(4):683–698, July 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Christopeit:1978:SCM

[Chr78]  Norbert Christopeit.  A

stochastic control model with chance constraints. SIAM Journal on Control and Optimization, 16(5):702–714, September 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Christopeit:1980:OSC

[Chr80] Norbert Christopeit. Optimal stochastic control with special information patterns. SIAM Journal on Control and Optimization, 18(5):559– 575, September 1980. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Christopeit:1983:DAC

[Chr83] Norbert Christopeit. Discrete approximation of continuous time stochastic control systems. SIAM Journal on Control and Optimization, 21(1): 17–40, January 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Christof:2019:SAO

[Chr19] Constantin Christof. Sensitivity analysis and optimal control of obstacle-type evolution variational inequalities. SIAM Journal on Control and Optimization, 57(1):192–218, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chiang:1987:DGO

[CHS87] Tzuu-Shuh Chiang, Chii-Ruey Hwang, and Shuenn Jyi Sheu. Diffusion for global optimization in R^{n}. SIAM Journal on Control and Optimization, 25 (3):737–753, May 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Caselli:2023:QVI

[CHY23] Gabriele Caselli, Maurice Hensel, and Irwin Yousept. Quasilinear variational inequalities in ferromagnetic shielding: Well-posedness, regularity, and optimal control. SIAM Journal on Control and Optimization, 61 (4):2043–2068, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1530483.

Curtain:1977:SPS

[CI77] Ruth F. Curtain and Akira Ichikawa. The separation principle for stochastic evolution equations. SIAM Journal on Control and Optimization, 15 (3):367–383, May 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Crouch:1986:DRH

[CI86] P. E. Crouch and M. Irving. Dynamical realizations of homogeneous Hamiltonian systems. SIAM Journal on Control and Optimization, 24 (3):374–395, May 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chasseigne:2019:EPV

[CI19] Emmanuel Chasseigne and Naoyuki Ichihara. Ergodic problems for viscous Hamilton–Jacobi equations with inward drift. SIAM Journal on Control and Optimization, 57(1):23–52, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cetinkaya:2018:PCR

[CIH18a] A. Cetinkaya, H. Ishii, and T. Hayakawa. A probabilistic characterization of random and malicious communication failures in multi-hop networked control. SIAM Journal on Control and Optimization, 56(5):3320–3350, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cetinkaya:2018:ETV

[CIH18b] Ahmet Cetinkaya, Hideaki Ishii, and Tomohisa Hayakawa. The effect of time-varying jamming interference on networked stabilization. SIAM Journal on Control and Optimization, 56(3):2398–2435, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chernorutskii:1996:SII

[CIP96] Vladimir Chernorutskii, Rauf Izmailov, and Alexei Pokrovskii. Superior information is insufficient to win in games between finite automata. SIAM Journal on Control and Optimization, 34(2):542–553, March 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23989.

Cazacu:2018:NCL

[CIP18] Cristian M. Cazacu, Liviu I.

Ignat, and Ademir F. Pazoto. Null-controllability of the linear Kuramoto–Sivashinsky equation on star-shaped trees. SIAM Journal on Control and Optimization, 56(4):2921– 2958, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cirant:2014:SSE

[Cir14] Marco Cirant. On the solvability of some ergodic control problems in R^{d}. SIAM Journal on Control and Optimization, 52(6):4001–4026, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cantoni:2013:RSA

[CJK13] Michael Cantoni, Ulf T. Jönsson, and Sei Zhen Khong. Robust stability analysis for feedback interconnections of time-varying linear systems. SIAM Journal on Control and Optimization, 51(1):353–379, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chitour:2012:OCM

[CJM12] Yacine Chitour, Frédéric Jean, and Paolo Mason. Optimal control models of goaloriented human locomotion. SIAM Journal on Control and Optimization, 50(1):147– 170, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v50/i1/p147_s1.

Carlier:2008:OTT

[CJS08] G. Carlier, C. Jimenez, and F. Santambrogio. Optimal transportation with traffic congestion and Wardrop equilibria. SIAM Journal on Control and Optimization, 47(3): 1330–1350, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chitour:2008:STC

[CJT08] Yacine Chitour, Frédéric Jean, and Emmanuel Trélat. Singular trajectories of controlaffine systems. SIAM Journal on Control and Optimization, 47(2):1078–1095, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chow:1978:NOF

[CK78] Joe H. Chow and Petar V.

Kokotovic. Near–optimal feedback stabilization of a class of nonlinear singularly perturbed systems. SIAM Journal on Control and Optimization, 16(5):756–770, September 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Connors:1989:SAT

[CK89] Daniel P. Connors and P. R.

Kumar. Simulated annealing type Markov chains and their order balance equations. SIAM Journal on Control and Optimization, 27(6):1440– 1461, November 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cadenillas:1995:SMP

[CK95] Abel Cadenillas and Ioannis Karatzas. The stochastic maximum principle for linear, convex optimal control with random coefficients. SIAM Journal on Control and Optimization, 33(2):590–624, March 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24072.

Campi:1998:ALQ

[CK98] M. C. Campi and P. R. Kumar. Adaptive linear quadratic Gaussian control: The cost-biased approach revisited. SIAM Journal on Control and Optimization, 36(6):1890–1907, November 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31749.

Cannon:2002:ECM

[CK02] Mark Cannon and Basil Kouvaritakis. Efficient constrained model predictive control with asymptotic optimality. SIAM Journal on Control and Optimization, 41(1):60–82, January 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35837.

Choi:2003:CES

[CK03] Jinhoon Choi and Wook Hyun Kwon. Continuity and exponential stability of mixed constrained model predictive control. SIAM Journal on Control and Optimization, 42(3):839–870, May 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38520.

Colonius:2009:IEC

[CK09] Fritz Colonius and Christoph Kawan. Invariance entropy for control systems. SIAM Journal on Control and Optimization, 48(3):1701–1721, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Casas:2014:OCS

[CK14] Eduardo Casas and Karl Kunisch. Optimal control of semilinear elliptic equations in measure spaces. SIAM Journal on Control and Optimization, 52(1):339–364, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Casas:2017:SSC

[CK17a] Eduardo Casas and Karl Kunisch. Stabilization by sparse controls for a class of semilinear parabolic equations. SIAM Journal on Control and Optimization, 55(1):512–532, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cox:2017:MIB

[CK17b] Alexander M. G. Cox and Sigrid Källblad. Modelindependent bounds for Asian options: a dynamic programming approach. SIAM Journal on Control and Optimization, 55(6):3409–3436, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Casas:2019:OCT

[CK19] Eduardo Casas and Karl Kunisch. Optimal control of the two-dimensional stationary Navier–Stokes equations with measure valued controls. SIAM Journal on Control and Optimization, 57(2): 1328–1354, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Casas:2021:OCT

[CK21] Eduardo Casas and Karl Kunisch. Optimal control of the two-dimensional evolutionary Navier–Stokes equations with measure valued controls. SIAM Journal on Control and Optimization, 59(3): 2223–2246, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Casas:2022:IHO

[CK22] Eduardo Casas and Karl Kunisch. Infinite horizon optimal control problems for a class of semilinear parabolic equations. SIAM Journal on Control and Optimization, 60 (4):2070–2094, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1464816.

Casas:2023:IHOa

[CK23a] Eduardo Casas and Karl Kunisch. Infinite horizon optimal control problems with discount factor on the state, Part I: Analysis of the controlled state equation. SIAM Journal on Control and Optimization, 61(3):1375–1393, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1490272.

Casas:2023:IHOb

[CK23b] Eduardo Casas and Karl Kunisch. Infinite horizon optimal control problems with discount factor on the state, Part II: Analysis of the control problem. SIAM Journal on Control and Optimization, 61(3):1438–1459, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1490296.

Casas:2017:OCS

[CKK17] Eduardo Casas, Florian Kruse, and Karl Kunisch. Optimal control of semilinear parabolic equations by BV-functions. SIAM Journal on Control and Optimization, 55(3):1752– 1788, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Casas:2016:AOC

[CKL16] Eduardo Casas, Peter I. Kogut, and Günter Leugering. Approximation of optimal control problems in the coefficient for the p-Laplace equation. I. Convergence result. SIAM Journal on Control and Optimization, 54(3): 1406–1422, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Choi:2023:TCC

[CKL23] Jin Hyuk Choi, Heeyoung Kwon, and Kasper Larsen. Trading constraints in continuoustime Kyle models. SIAM Journal on Control and Optimization, 61(3):1494–1512, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1446617.

Chen:2020:RDF

[CKM20] Yuan Chen, Soummya Kar, and José M. F. Moura. Resilient distributed field estimation. SIAM Journal on Control and Optimization, 58(3): 1429–1456, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Colombo:2022:SCH

[CKR22] Giovanni Colombo, Nathalie T.

Khalil, and Franco Rampazzo. State constraints, higher order inward pointing conditions, and neighboring feasible trajectories. SIAM Journal on Control and Optimization, 60(6):3326–3357, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1468796.

Carrillo:2022:CST

[CKRT22] José A. Carrillo, Dante Kalise, Francesco Rossi, and Emmanuel Trélat. Controlling swarms toward flocks and mills. SIAM Journal on Control and Optimization, 60 (3):1863–1891, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1404314.

Charalambous:2020:ECC

[CKT20] Charalambos D. Charalambous, Christos Kourtellaris, and Ioannis Tzortzis. Ergodic control-coding capacity of stochastic control systems: Information signalling and hierarchical optimality of Gaussian systems. SIAM Journal on Control and Optimization, 58(1):104–135, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Charalambous:2024:OCS

[CKT24] Charalambos D. Charalambous, Christos K. Kourtellaris, and Ioannis Tzortzis. Optimal control and signaling strategies of control-coding capacity of general decision models: Applications to Gaussian models and decentralized strategies. SIAM Journal on Control and Optimization, 62 (1):600–629, February 2024. CODEN SJCODC. ISSN 1095-7138.

Carasso:1978:AMC

[CL78] C. Carasso and P. J. Laurent. Un algorithme de minimisation en chaîne en optimisation convexe. (French) [A chained minimization algorithm for convex optimization]. SIAM Journal on Control and Optimization, 16(2): 209–235, March 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Clarke:1986:VFO

[CL86] Frank H. Clarke and Philip D. Loewen. The value function in optimal control: Sensitivity, controllability, and timeoptimality. SIAM Journal on Control and Optimization, 24 (2):243–263, March 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Censor:1987:OEL

[CL87a] Yair Censor and Arnold Lent. Optimization of “log x” entropy over linear equality constraints. SIAM Journal on Control and Optimization, 25 (4):921–933, July 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Clarke:1987:SCO

[CL87b] Frank H. Clarke and Philip D.

Loewen. State constraints in optimal control: a case study in proximal normal analysis. SIAM Journal on Control and Optimization, 25(6):1440– 1456, November 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chui:1994:CBH

[CL94] Charles K. Chui and Xin Li. Continuity of best Hankel approximation and convergence of near-best approximants. SIAM Journal on Control and Optimization, 32(6): 1769–1781, November 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/23224.

Caines:2002:GCN

[CL02] Peter E. Caines and Ekaterina S. Lemch. On the global controllability of nonlinear systems: Fountains, recurrence, and applications to Hamiltonian systems. SIAM Journal on Control and Optimization, 41(5): 1532–1553, September 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/36195.

Chatterjee:2006:SAD

[CL06a] Debasish Chatterjee and Daniel Liberzon. Stability analysis of deterministic and stochastic switched systems via a comparison principle and multiple Lyapunov functions. SIAM Journal on Control and Optimization, 45(1): 174–206, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cogill:2006:AAD

[CL06b] Randy Cogill and Sanjay Lall. An approximation algorithm for the discrete team decision problem. SIAM Journal on Control and Optimization, 45 (4):1359–1368, January 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chatterjee:2011:SRS

[CL11] Debasish Chatterjee and Daniel Liberzon. Stabilizing randomly switched systems. SIAM Journal on Control and Optimization, 49(5): 2008–2031, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicon/resource/
                                        1/sjcodc/v49/i5/p2008_s1.

Christensen:2018:FES

[CL18] Sören Christensen and Kristoffer Lindensjö. On finding equilibrium stopping times for time-inconsistent Markovian problems. SIAM Journal on Control and Optimization, 56 (6):4228–4255, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Caines:2019:SOL

[CL19a] Peter E. Caines and David Levanony. Stochastic εoptimal linear quadratic adaptation: an alternating controls policy. SIAM Journal on Control and Optimization, 57(2): 1094–1126, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Church:2019:CER

[CL19b] Kevin E. M. Church and Xinzhi Liu. Cost-effective robust stabilization and bifurcation suppression. SIAM Journal on Control and Optimization, 57(3):2240–2268, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cao:2024:SFR

[CL24] Wenqi Cao and Anders Lindquist. Spectral factorization of rank-deficient rational densities. SIAM Journal on Control and Optimization, 62(1):776–798, February 2024. CODEN SJCODC. ISSN 1095-7138.

Clarke:1976:GPB

[Cla76a] Frank H. Clarke. The generalized problem of Bolza. SIAM Journal on Control and Optimization, 14(4):682–699, July 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Clarke:1976:MPU

[Cla76b] Frank H. Clarke. The maximum principle under minimal hypotheses. SIAM Journal on Control and Optimization, 14(6):1078–1091, November 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Clarke:1982:HFS

[Cla82] Frank H. Clarke. On Hamiltonian flows and symplectic transformations. SIAM Journal on Control and Optimization, 20(3):355–359, May 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Claude:1986:DSN

[Cla86] Daniel Claude. Découplage des systèmes non linéaires, séries génératrices non commutatives et algèbres de Lie. (French) [Decoupling of nonlinear systems, noncommutative generating series, and Lie algebras]. SIAM Journal on Control and Optimization, 24 (3):562–578, May 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cao:2013:MAO

[CLAL13] Liqun Cao, Jianjun Liu, Walter Allegretto, and Yanping Lin. A multiscale approach for optimal control problems of linear parabolic equations. SIAM Journal on Control and Optimization, 50(6):3269– 3291, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Clarke:2011:ESM

[CLdRdP11] Francis Clarke, Yuri Ledyaev, and Maria do Rosário de Pinho. An extension of the Schwarzkopf multiplier rule in optimal control. SIAM Journal on Control and Optimization, 49(2):599–610, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sicon/
resource/1/sjcodc/v49/i2/
p599_s1.

Chen:2012:SPI

[CLG12] Ge Chen, Zhixin Liu, and Lei Guo. The smallest possible interaction radius for flock synchronization. SIAM Journal on Control and Optimization, 50(4):1950–1970, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Castaneda-Leyva:2005:OCI

[CLHH05] Netzahualcóyotl Castañeda-Leyva and Daniel Hernández-Hernández. Optimal consumptioninvestment problems in incomplete markets with stochastic coefficients. SIAM Journal on Control and Optimization, 44(4):1322–1344, July 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/44088.

Cacace:2020:MOC

[CLL20] Simone Cacace, Anna Chiara Lai, and Paola Loreti. Modeling and optimal control of an octopus tentacle. SIAM Journal on Control and Optimization, 58(1):59–84, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cardaliaguet:2013:LTA

[CLLP13] P. Cardaliaguet, J.-M. Lasry, P.-L. Lions, and A. Porretta. Long time average of mean field games with a nonlocal coupling. SIAM Journal on Control and Optimization, 51 (5):3558–3591, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Charlet:1991:SCD

[CLM91] B. Charlet, J. Lévine, and R. Marino. Sufficient conditions for dynamic state feedback linearization. SIAM Journal on Control and Optimization, 29(1):38–57, January 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Claret:2024:EBC

[CLM24] Sue Claret, Jérôme Lemoine, and Arnaud Münch. On the exact boundary controllability of semilinear wave equations. SIAM Journal on Control and Optimization, 62 (4):1953–1976, August 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1586598.

Clark:2000:SRI

[CLMSR00] Stephen Clark, Yuri Latushkin, Stephen Montgomery-Smith, and Timothy Randolph. Stability radius and internal versus external stability in Banach spaces: An evolution semigroup approach. SIAM Journal on Control and Optimization, 38(6):1757–1793, November 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34212.

Chapman:2013:CLA

[CLRJ13] Archie C. Chapman, David S.

Leslie, Alex Rogers, and Nicholas R. Jennings. Convergent learning algorithms for unknown reward games. SIAM Journal on Control and Optimization, 51(4):3154–3180, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Clarke:2000:FSL

[CLRS00] F. H. Clarke, Yu. S. Ledyaev, L. Rifford, and R. J. Stern. Feedback stabilization and Lyapunov functions. SIAM Journal on Control and Optimization, 39(1):25–48, January 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/35229.

Clarke:1997:SUF

[CLS97] F. H. Clarke, Yu. S. Ledyaev, and A. I. Subbotin. The synthesis of universal feedback pursuit strategies in differential games. SIAM Journal on Control and Optimization, 35(2):552–561, March 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/28397.

Cardaliaguet:2012:CTA

[CLS12] Pierre Cardaliaguet, Rida Laraki, and Sylvain Sorin. A continuous time approach for the asymptotic value in two-person zero-sum repeated games. SIAM Journal on Control and Optimization, 50(3): 1573–1596, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:2021:SAS

[CLS21] Huabin Chen, Cheng-Chew Lim, and Peng Shi. Stability analysis for stochastic neutral switched systems with timevarying delay. SIAM Journal on Control and Optimization, 59(1):24–49, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:2020:ACD

[CLSL20] Sheng Chen, Cheng-Chew Lim, Peng Shi, and Zhenyu Lu. Asymptotic consensus of dynamical points in a strict max-convex space and its applications. SIAM Journal on Control and Optimization, 58 (4):1984–2005, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chu:2006:NMG

[CLT06] Delin Chu, Wen-Wei Lin, and Roger C. E. Tan. A numerical method for a generalized algebraic Riccati equation. SIAM Journal on Control and Optimization, 45(4): 1222–1250, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Colle:2024:FAB

[CLT24] Blaise Colle, Jérôme Lohéac, and Takéo Takahashi. Flatness approach for the boundary controllability of a system of heat equations. SIAM Journal on Control and Optimization, 62(3):1766–1782, June 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1577833.

Chen:2017:GFS

[CLW17] Wenbin Chen, Chang Liu, and Zhiqiang Wang. Global feedback stabilization for a class of nonlocal transport equations: The continuous and discrete case. SIAM Journal on Control and Optimization, 55(2): 760–784, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:2013:CDT

[CLYL13] Yao Chen, Jinhu Lü, Xinghuo Yu, and Zongli Lin. Consensus of discrete-time second-order multiagent systems based on infinite products of general stochastic matrices. SIAM Journal on Control and Optimization, 51(4):3274–3301, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:1998:SLQ

[CLZ98] Shuping Chen, Xunjing Li, and Xun Yu Zhou. Stochastic linear quadratic regulators with indefinite control weight costs. SIAM Journal on Control and Optimization, 36(5):1685–1702, September 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31047.

Cui:2023:OCS

[CLZ23] Jianbo Cui, Shu Liu, and Haomin Zhou. Optimal control for stochastic nonlinear Schrödinger equation on graph. SIAM Journal on Control and Optimization, 61 (4):2021–2042, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1524175.

Cao:2024:MCP

[CLZH24] Yuexin Cao, Yibei Li, Lirong Zheng, and Xiaoming Hu. Minimal control placement of networked reaction–diffusion systems based on Turing model. SIAM Journal on Control and Optimization, 62 (3):1809–1831, June 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1616856.

Corfmat:1976:CLS

[CM76] J. P. Corfmat and A. S.

Morse. Control of linear systems through specified input channels. SIAM Journal on Control and Optimization, 14(1):163–175, January 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:1981:FET

[CM81] Goong Chen and Wendell H. Mills, Jr. Finite elements and terminal penalization for quadratic cost optimal control problems governed by ordinary differential equations. SIAM Journal on Control and Optimization, 19(6):744– 764, November 1981. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cohen:1990:OAC

[CM90] Guy Cohen and Bernadette Miara. Optimization with an auxiliary constraint and decomposition. SIAM Journal on Control and Optimization, 28(1):137–157, January 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Champetier:1991:EAG

[CM91] Calixte Champetier and Jean-François Magni. On eigenstructure assignment by gain output feedback. SIAM Journal on Control and Optimization, 29(4):848–865, July 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chui:1992:MRS

[CM92] Charles K. Chui and H. N. Mhaskar. On multivariate robust stability. SIAM Journal on Control and Optimization, 30(5):1190–1199, September 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cerf:1995:BBC

[CM95] Raphaël Cerf and Carlo Mariconda. On bang–bang constrained solutions of a control system. SIAM Journal on Control and Optimization, 33(2):554–567, March 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24203.

Cheng:1996:LPO

[CM96a] Chih-Wen Cheng and Victor J. Mizel. On the Lavrentiev phenomenon for optimal control problems with secondorder dynamics. SIAM Journal on Control and Optimization, 34(6):2172–2179, November 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26121.

Crasta:1996:ERN

[CM96b] Graziano Crasta and Annalisa Malusa. Existence results for noncoercive variational problems. SIAM Journal on Control and Optimization, 34(6):2064–2076, November 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/27820.  See

comment [Bal01].

Conrad:1998:SFB

[CM98] Francis Conrad and Ömer Morgül. On the stabilization of a flexible beam with a tip mass. SIAM Journal on Control and Optimization, 36(6):1962–1986, November 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/30236.

Chu:2000:DDD

[CM00] Delin Chu and Volker Mehrmann. Disturbance decoupling for descriptor systems by state feedback. SIAM Journal on Control and Optimization, 38(6): 1830–1858, November 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/33189.

Casas:2002:SOO

[CM02] Eduardo Casas and Mariano Mateos. Second order optimality conditions for semilinear elliptic control problems with finitely many state constraints. SIAM Journal on Control and Optimization, 40(5):1431–1454, September 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38201.

Cheng:2003:WPB

[CM03a] Ada Cheng and Kirsten Morris. Well-posedness of boundary control systems. SIAM Journal on Control and Optimization, 42(4):1244–1265, July 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38491.

Cortes:2003:CCM

[CM03b] Jorge Cortés and Sonia Martínez. Configuration controllability of mechanical systems underactuated by one control. SIAM Journal on Control and Optimization, 41(6):1901–1921, November 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37409.

Chang:2004:RCL

[CM04] Dong Eui Chang and Jerrold E. Marsden. Reduction of controlled Lagrangian and Hamiltonian systems with symmetry. SIAM Journal on Control and Optimization, 43(1):277–300, January 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41295.

Cupini:2005:EMP

[CM05] Giovanni Cupini and Elvira Mascolo. Existence of minimizers for polyconvex and nonpolyconvex problems. SIAM Journal on Control and Optimization, 44(4):1370–1390, July 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/61199.

Cardaliaguet:2006:REE

[CM06] Pierre Cardaliaguet and Claudio Marchi. Regularity of the eikonal equation with Neumann boundary conditions in the plane: Application to fronts with nonlocal terms. SIAM Journal on Control and Optimization, 45(3):1017– 1038, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Conlon:2007:MEO

[CM07] Joseph G. Conlon and Hyekyung Min. On a model for the efficient operation of a bank or insurance company. SIAM Journal on Control and Optimization, 46(6):1972–1994, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Castellanos:2010:SAE

[CM10a] Jesús Castellanos and Julio Muñoz. Some aspects of the existence of minimizers for a nonlocal integral functional in dimension one. SIAM Journal on Control and Optimization, 48(6):3838–3858, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cellina:2010:NCS

[CM10b] A. Cellina and M. Mazzola.

Necessary conditions for solutions to variational problems. SIAM Journal on Control and Optimization, 48(5): 2977–2983, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chambrion:2012:GCS

[CM12] Thomas Chambrion and Alexandre Munnier. Generic controllability of 3D swimmers in a perfect fluid. SIAM Journal on Control and Optimization, 50(5):2814–2835, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Camilli:2016:SMF

[CM16] Fabio Camilli and Claudio Marchi. Stationary mean field games systems defined on networks. SIAM Journal on Control and Optimization, 54(2): 1085–1103, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chittaro:2017:ACA

[CM17] Francesca Carlotta Chittaro and Paolo Mason. Approximate controllability via adiabatic techniques for the threeinputs controlled Schrödinger equation. SIAM Journal on Control and Optimization, 55 (6):4202–4226, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cortes:2018:PBD

[CM18] Andres Cortés and Sonia MartÍnez. A projection-based decomposition algorithm for distributed fast computation of control in microgrids. SIAM Journal on Control and Optimization, 56(2):583–609, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cao:2008:RCDa

[CMA08a] Ming Cao, A. Stephen Morse, and Brian D. O. Anderson. Reaching a consensus in a dynamically changing environment: a graphical approach. SIAM Journal on Control and Optimization, 47(2):575–600, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cao:2008:RCDb

[CMA08b] Ming Cao, A. Stephen Morse, and Brian D. O. Anderson. Reaching a consensus in a dynamically changing environment: Convergence rates, measurement delays, and asynchronous events. SIAM Journal on Control and Optimization, 47(2):601–623, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:1981:ECT

[CMC81] Goong Chen, Wendell H.

Mills, Jr., and Giovanni Crosta. Exact controllability theorems and numerical simulations for some nonlinear differential equations. SIAM Journal on Control and Optimization, 19(6):765– 790, November 1981. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chitour:2015:CRP

[CMK15] Y. Chitour, M. Godoy Molina, and P. Kokkonen. On the controllability of the rolling problem onto the hyperbolic nspace. SIAM Journal on Control and Optimization, 53(2): 948–968, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Colombo:2020:OPS

[CMN20] Giovanni Colombo, Boris S.

Mordukhovich, and Dao Nguyen. Optimization of a perturbed sweeping process by constrained discontinuous controls. SIAM Journal on Control and Optimization, 58(4):

2678–2709, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cortes:2002:SMC

[CMOZ02] J. Cortés, S. Martínez, J. P. Ostrowski, and H. Zhang. Simple mechanical control systems with constraints and symmetry. SIAM Journal on Control and Optimization, 41(3):851–874, May 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/38174.

Cerpa:2015:NCS

[CMP15] Eduardo Cerpa, Alberto Mercado, and Ademir F. Pazoto. Null controllability of the stabilized Kuramoto–Sivashinsky system with one distributed control. SIAM Journal on Control and Optimization, 53 (3):1543–1568, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chow:1985:ACS

[CMR85] Pao-Liu Chow, José-Luis Menaldi, and Maurice Robin. Additive control of stochastic linear systems with finite horizon. SIAM Journal on Control and Optimization, 23(6): 858–899, November 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Casas:2007:EEN

[CMR07] Eduardo Casas, Mariano Mateos, and Jean-Pierre Raymond. Error estimates for the numerical approximation of a distributed control problem for the steady-state Navier– Stokes equations. SIAM Journal on Control and Optimization, 46(3):952–982, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cindea:2017:BCF

[CMR17] Nicolae Cîndea, Sorin Micu, and Ionel Roventa. Boundary controllability for finitedifferences semidiscretizations of a clamped beam equation. SIAM Journal on Control and Optimization, 55(2):785–817, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Casas:2019:EES

[CMR19] Eduardo Casas, Mariano Mateos, and Arnd Rösch. Error estimates for semilinear parabolic control problems in the absence of Tikhonov term. SIAM Journal on Control and Optimization, 57(4): 2515–2540, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cardaliaguet:2016:FOM

[CMS16] Pierre Cardaliaguet, Alpár R.

Mészáros, and Filippo Santambrogio. First order mean field games with density constraints: Pressure equals price. SIAM Journal on Control and Optimization, 54(5):2672– 2709, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Crauel:2003:NAH

[CMT03] Hans Crauel, Iakovos Matsikis, and Stuart Townley. Noise assisted high-gain stabilization: Almost surely or in second mean. SIAM Journal on Control and Optimization, 42(5):1834–1853, September 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39390.

Cindea:2011:AME

[CMT11] Nicolae Cîndea, Sorin Micu, and Marius Tucsnak. An approximation method for exact controls of vibrating systems. SIAM Journal on Control and Optimization, 49(3): 1283–1305, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i3/p1283_s1.

Casas:2018:MCS

[CMT18] Eduardo Casas, Mariano Mateos, and Fredi Tröltzsch. Measure control of a semilinear parabolic equation with a nonlocal time delay. SIAM Journal on Control and Optimization, 56(6):4434–4460, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cannarsa:2008:CEC

[CMV08] P. Cannarsa, P. Martinez, and J. Vancostenoble. Carleman estimates for a class of degenerate parabolic operators. SIAM Journal on Control and Optimization, 47(1):1–19, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Colombo:2006:SNR

[CMW06] Giovanni Colombo, Antonio Marigonda, and Peter R. Wolenski. Some new regularity properties for the minimal time function. SIAM Journal on Control and Optimization, 44(6):2285–2299, January 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Clarke:2005:NDO

[CN05] F. H. Clarke and C. Nour.

Nonconvex duality in optimal control. SIAM Journal on Control and Optimization, 43(6):2036–2048, November 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42955.

Colombo:2010:SMT

[CN10] Giovanni Colombo and Khai T. Nguyen. On the structure of the minimum time function. SIAM Journal on Control and Optimization, 48(7): 4776–4814, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cannarsa:2011:ESC

[CN11] Piermarco Cannarsa and Khai T. Nguyen. Exterior sphere condition and time optimal control for differential inclusions. SIAM Journal on Control and Optimization, 49(6):2558–2576, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sicon/
                                        resource/1/sjcodc/v49/i6/
                                        p2558_s1.

Coron:2019:OTC

[CN19] Jean-Michel Coron and Hoai-Minh Nguyen. Optimal time for the controllability of linear hyperbolic systems in onedimensional space. SIAM Journal on Control and Optimization, 57(2):1127–1156, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Calisti:2024:SOH

[CN24] Valentin Calisti and Sárka Necasová. Shape optimization of hemolysis for shear thinning flows in moving domains. SIAM Journal on Control and Optimization, 62 (3):1546–1568, June 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1595485.

Chou:1998:MSS

[CNP98] Chin-Cheng Chou, Kung-Fu Ng, and Jong-Shi Pang. Minimizing and stationary sequences of constrained optimization problems. SIAM Journal on Control and Optimization, 36(6):1908–1936, November 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/29254.

Cardone:2010:AAP

[CNS10] G. Cardone, S. A. Nazarov, and J. Sokolowski. Asymptotic analysis, polarization matrices, and topological derivatives for piezoelectric materials with small voids. SIAM Journal on Control and Optimization, 48(6):3925–3961, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Christensen:2021:CVC

[CNS21] Sören Christensen, Berenice Anne Neumann, and Tobias Sohr. Competition versus cooperation: a class of solvable mean field impulse control problems. SIAM Journal on Control and Optimization, 59(5): 3946–3972, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cavalcanti:2003:FVV

[CO03a] Marcelo Moreira Cavalcanti and Higidio Portillo Oquendo. Frictional versus viscoelastic damping in a semilinear wave equation. SIAM Journal on Control and Optimization, 42 (4):1310–1324, January 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cellina:2003:ESD

[CO03b] Arrigo Cellina and António Ornelas. Existence of solutions to differential inclusions and to time optimal control problems in the autonomous case. SIAM Journal on Control and Optimization, 42(1):260–265, January 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/40804.

Curtain:2005:SNP

[CO05] Ruth F. Curtain and Mark R. Opmeer. The suboptimal Nehari problem for well-posed linear systems. SIAM Journal on Control and Optimization, 44(3):991–1018, May 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/44318.

Contreras:2022:CLS

[CO22] César Contreras and Tomoki Ohsawa. Controlled Lagrangians and stabilization of Euler–Poincaré mechanical systems with broken symmetry i: Kinetic shaping. SIAM Journal on Control and Optimization, 60(5):2684–2711, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1435458.

Cobb:2006:SFI

[Cob06] Daniel Cobb. State feedback impulse elimination for singular systems over a Hermite domain. SIAM Journal on Control and Optimization, 44(6): 2189–2209, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cobb:2010:ECC

[Cob10] Daniel Cobb. Eigenvalue conditions for convergence of singularly perturbed matrix exponential functions. SIAM Journal on Control and Optimization, 48(7):4327–4351, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cobb:2012:SOB

[Cob12]  Daniel Cobb.  Singular

observer-based compensators and the deterministic separation principle. SIAM Journal on Control and Optimization, 50(4):1921–1949, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

CohendeLara:1997:FDFa

[Coh97a] M. Cohen de Lara. Finitedimensional filters. Part I: The Wei–Norman technique. SIAM Journal on Control and Optimization, 35(3):980–1001, May 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27090.

CohendeLara:1997:FDFb

[Coh97b] M. Cohen de Lara. Finitedimensional filters. Part II: Invariance group techniques. SIAM Journal on Control and Optimization, 35(3):1002– 1029, May 1997. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27091.

Cohen:2021:SCP

[Coh21] Asaf Cohen. On singular control problems, the timestretching method, and the weak-M1 topology. SIAM Journal on Control and Optimization, 59(1):50–77, ???? 2021. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Colonius:1982:MPR

[Col82a] Fritz Colonius. The maximum principle for relaxed hereditary differential systems with function space end condition. SIAM Journal on Control and Optimization, 20 (5):695–712, September 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Colonius:1982:SRR

[Col82b] Fritz Colonius. Stable and regular reachability for relaxed hereditary differential systems. SIAM Journal on Control and Optimization, 20 (5):675–694, September 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Colonius:1985:ASR

[Col85] Fritz Colonius. Addendum:

“Stable and regular reachability of relaxed hereditary differential systems” [SIAM J. Control Optim. 20 (1982), no. 5, 675–694; MR 84h:49081]. SIAM Journal on Control and Optimization, 23(5):803– 807, September 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Colonius:1988:OPC

[Col88] Fritz Colonius. Optimal periodic control: a scenario for local properness. SIAM Journal on Control and Optimization, 26(2):452–464, March 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Collins:2011:SCE

[Col11]  Pieter Collins.  Semantics

and computability of the evolution of hybrid systems. SIAM Journal on Control and Optimization, 49(2):890– 925, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i2/p890_s1.

Colonius:2012:MBR

[Col12] Fritz Colonius. Minimal bit rates and entropy for exponential stabilization. SIAM Journal on Control and Optimization, 50(5):2988–3010, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Combettes:2000:SCB

[Com00] Patrick L. Combettes. Strong convergence of block-iterative outer approximation methods for convex optimization. SIAM Journal on Control and Optimization, 38(2):538–565, March 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32626.

Conrad:1990:SBP

[Con90] F. Conrad. Stabilization of beams by pointwise feedback control. SIAM Journal on Control and Optimization, 28 (2):423–437, March 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Coron:1994:LCS

[Cor94]  Jean-Michel Coron.  Lin-

earized control systems and applications to smooth stabilization. SIAM Journal on Control and Optimization, 32(2):358–386, March 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/22686.

Coron:1995:SFT

[Cor95] Jean-Michel Coron. On the stabilization in finite time of locally controllable systems by means of continuous timevarying feedback law. SIAM Journal on Control and Optimization, 33(3):804–833, May 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24049.

Coron:1999:NAS

[Cor99] Jean-Michel Coron. On the null asymptotic stabilization of the two-dimensional incompressible Euler equations in a simply connected domain. SIAM Journal on Control and Optimization, 37(6): 1874–1896, November 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34140.

Cosso:2013:SDG

[Cos13] Andrea Cosso. Stochastic differential games involving impulse controls and doubleobstacle quasi-variational inequalities. SIAM Journal on Control and Optimization, 51 (3):2102–2131, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Courdesses:1983:CIS

[Cou83] M. Courdesses. Comments on: “Identifiability of spatiallyvarying and constant parameters in distributed systems of parabolic type” [SIAM J. Control Optim. 15 (1977), no. 5, 785–802; MR 56 #18045] by Shinzo Kitamura and S. Nakagiri. SIAM Journal on Control and Optimization, 21 (3):410–412, May 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [KN77].

Couvreux:2000:MSC

[Cou00] Jérôme Couvreux. From the Mosco and Slice convergences for convex normal integrands to that of integral functionals on L_{p} spaces. SIAM Journal on Control and Optimization, 39(1):179–195, January 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32222.

Curtain:1976:IDR

[CP76] Ruth Curtain and A. J.

Pritchard. The infinitedimensional Riccati equation for systems defined by evolution operators. SIAM Journal on Control and Optimization, 14(5):951–983, August 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Curtain:1977:ATU

[CP77] Ruth F. Curtain and A. J.

Pritchard. An abstract theory for unbounded control action for distributed parameter systems. SIAM Journal on Control and Optimization, 15 (4):566–611, July 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Conte:1982:SPI

[CP82] G. Conte and A. M. Perdon. Systems over a principal ideal domain. A polynomial model approach. SIAM Journal on Control and Optimization, 20(1):112–124, January 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Conte:1995:DDP

[CP95] G. Conte and A. M. Perdon. The disturbance decoupling problem for systems over a ring. SIAM Journal on Control and Optimization, 33(3):750–764, May 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23563.

Cellina:1998:VMP

[CP98] Arrigo Cellina and Stefania Perrotta. On the validity of the maximum principle and of the Euler–Lagrange equation for a minimum problem depending on the gradient. SIAM Journal on Control and Optimization, 36(6): 1987–1998, November 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31131.

Cardaliaguet:2000:ISD

[CP00] Pierre Cardaliaguet and Slawomir Plaskacz. Invariant solutions of differential games and Hamilton–Jacobi–Isaacs equations for time-measurable Hamiltonians. SIAM Journal on Control and Optimization, 38(5):1501–1520, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29621.

Celada:2002:EMN

[CP02] P. Celada and S. Perrotta.

Existence of minimizers for nonconvex, noncoercive simple integrals. SIAM Journal on Control and Optimization, 41(4):1118–1140, July 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38799.

Combettes:2004:PMC

[CP04] Patrick L. Combettes and Teemu Pennanen. Proximal methods for cohypomonotone operators. SIAM Journal on Control and Optimization, 43(2):731–742, March 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42733.

Cogill:2013:RMD

[CP13a] Randy Cogill and Cheng Peng.

Reversible Markov decision processes with an averagereward criterion. SIAM Journal on Control and Optimization, 51(1):402–418, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Colombo:2013:CMS

[CP13b] R. M. Colombo and N. Pogodaev. On the control of moving sets: Positive and negative confinement results. SIAM Journal on Control and Optimization, 51(1):380–401, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Colombo:2016:MTF

[CP16] Giovanni Colombo and Michele Palladino. The minimum time function for the controlled Moreau’s sweeping process. SIAM Journal on Control and Optimization, 54(4): 2036–2062, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Camlibel:2006:CLS

[CPS06] M. Kanat Camlibel, Jong-Shi Pang, and Jinglai Shen. Conewise linear systems: Non-Zenoness and observability. SIAM Journal on Control and Optimization, 45(5):1769– 1800, ???? 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Conte:1988:FPT

[CPW88] G. Conte, A. M. Perdon, and B. F. Wyman. Fixed poles in transfer function equations. SIAM Journal on Control and Optimization, 26(2): 356–368, March 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chang:2009:OSP

[CPY09] Mou-Hsiung Chang, Tao Pang, and Jiongmin Yong. Optimal stopping problem for stochastic differential equations with random coefficients. SIAM Journal on Control and Optimization, 48(2):941–971, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chai:2005:MPS

[CQ05] Li Chai and Li Qiu. Multirate periodic systems and constrained analytic function interpolation problems. SIAM Journal on Control and Optimization, 43(6): 1972–1986, November 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43005.

Cannarsa:2015:VDL

[CQ15] P. Cannarsa and M. Quincampoix. Vanishing discount limit and nonexpansive optimal control and differential games. SIAM Journal on Control and Optimization, 53(4): 1789–1814, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cardaliaguet:2000:PDG

[CQSP00] Pierre Cardaliaguet, Marc Quincampoix, and Patrick Saint-Pierre. Pursuit differential games with state constraints. SIAM Journal on Control and Optimization, 39(5):1615–1632, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34932.

Chong:1993:OQU

[CR93] Edwin K. P. Chong and Peter J. Ramadge. Optimization of queues using an infinitesimal perturbation analysisbased stochastic algorithm with general update times. SIAM Journal on Control and Optimization, 31(3):698–732, May 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Casas:2006:EEN

[CR06] Eduardo Casas and Jean-Pierre Raymond. Error estimates for the numerical approximation of Dirichlet boundary control for semilinear elliptic equations. SIAM Journal on Control and Optimization, 45(5):1586–1611, ???? 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cardaliaguet:2011:HRV

[CR11a] Pierre Cardaliaguet and Catherine Rainer. Hölder regularity for viscosity solutions of fully nonlinear, local or nonlocal, Hamilton– Jacobi equations with superquadratic growth in the gradient. SIAM Journal on Control and Optimization, 49 (2):555–573, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i2/p555_s1.

Cudina:2011:AOC

[CR11b] Milica Cudina and Kavita Ramanan. Asymptotically optimal controls for timeinhomogeneous networks. SIAM Journal on Control and Optimization, 49(2):611–645, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sicon/
resource/1/sjcodc/v49/i2/
p611_s1.

Cobb:2013:GCO

[CR13] Daniel Cobb and Roengrut Rujanakraikarn. Genericity of classes of observer-based compensators. SIAM Journal on Control and Optimization, 51 (2):1035–1046, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cavus:2014:RAC

[ÇR14] Özlem Çavus and Andrzej Ruszczyński. Risk-averse control of undiscounted transient Markov models. SIAM Journal on Control and Optimization, 52(6):3935–3966, ???? 2014. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Chau:2017:SRT

[CR17a] Huy N. Chau and Miklós Rásonyi. Skorohod’s representation theorem and optimal strategies for markets with frictions. SIAM Journal on Control and Optimization, 55 (6):3592–3608, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Coron:2017:QAT

[CR17b] Jean-Michel Coron and Ivonne Rivas. Quadratic approximation and time-varying feedback laws. SIAM Journal on Control and Optimization, 55(6):3726–3749, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cardaliaguet:2019:EME

[CR19] Pierre Cardaliaguet and Catherine Rainer. On the (in)efficiency of MFG equilibria. SIAM Journal on Control and Optimization, 57(4):2292–2314, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:2021:GAD

[CR21] Yahao Chen and Witold Respondek. Geometric analysis of differential-algebraic equations via linear control theory. SIAM Journal on Control and Optimization, 59(1):

103–130, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Crasta:1999:MPC

[Cra99] Graziano Crasta. On the minimum problem for a class of noncoercive nonconvex functionals. SIAM Journal on Control and Optimization, 38 (1):237–253, January 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33070.

Crouch:1981:DRF

[Cro81] P. E. Crouch. Dynamical realizations of finite Volterra series. SIAM Journal on Control and Optimization, 19(2): 177–202, March 1981. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Crouch:1984:SAC

[Cro84] P. E. Crouch. Solvable approximations to control systems. SIAM Journal on Control and Optimization, 22(1): 40–54, January 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chowdhury:2012:CSL

[CRR12] S. Chowdhury, M. Ramaswamy, and J.-P. Raymond. Controllability and stabilizability of the linearized compressible Navier–Stokes system in one dimension. SIAM Journal on Control and Optimization, 50 (5):2959–2987, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cocquet:2021:OBL

[CRS21] Pierre-Henri Cocquet, Sebastian Riffo, and Julien Salomon. Optimization of bathymetry for long waves with small amplitude. SIAM Journal on Control and Optimization, 59(6):4429–4456, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Conti:2018:SDC

[CRST18] Sergio Conti, Martin Rumpf, Rüdiger Schultz, and Sascha Tölkes. Stochastic dominance constraints in elastic shape optimization. SIAM Journal on Control and Optimization, 56 (4):3021–3034, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Casas:2015:SOS

[CRT15] Eduardo Casas, Christopher Ryll, and Fredi Tröltzsch. Second order and stability analysis for optimal sparse control of the FitzHugh–Nagumo equation. SIAM Journal on Control and Optimization, 53 (4):2168–2202, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cot:2009:ECA

[CRV09] L. Cot, J.-P. Raymond, and J. Vancostenoble. Exact controllability of an aeroacoustic model with a Neumann and a Dirichlet boundary control. SIAM Journal on Control and Optimization, 48(3): 1489–1518, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Casas:2000:PPL

[CRZ00] E. Casas, J. P. Raymond, and H. Zidani. Pontryagin’s principle for local solutions of control problems with mixed control-state constraints. SIAM Journal on Control and Optimization, 39(4):1182–1203, July 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34562.

Cerpa:2013:BCK

[CRZ13] Eduardo Cerpa, Ivonne Rivas, and Bing-Yu Zhang. Boundary controllability of the Korteweg–de Vries equation on a bounded domain. SIAM Journal on Control and Optimization, 51(4):2976–3010, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chakrabarty:2018:UIE

[CRZ^{+}18] Ankush Chakrabarty, Ann E. Rundell, Stanisław H. Zak, Fanglai Zhu, and Gregery T. Buzzard. Unknown input estimation for nonlinear systems using sliding mode observers and smooth window functions. SIAM Journal on Control and Optimization, 56 (5):3619–3641, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chewning:1977:CSB

[CS77] William C. Chewning and Thomas I. Seidman. A convergent scheme for boundary control of the heat equation. SIAM Journal on Control and Optimization, 15(1): 64–72, January 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Conner:1984:SDR

[CS84] Luther T. Conner, Jr. and David P. Stanford. State deadbeat response and observability in multi-modal systems. SIAM Journal on Control and Optimization, 22(4):630–644, July 1984. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Curtain:1986:FDC

[CS86] R. F. Curtain and D. Salamon. Finite-dimensional compensators for infinite-dimensional systems with unbounded input operators. SIAM Journal on Control and Optimization, 24 (4):797–816, July 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chung:1987:DMD

[CS87] Kun Jen Chung and Matthew J.

Sobel. Discounted MDP’s: Distribution functions and exponential utility maximization. SIAM Journal on Control and Optimization, 25(1): 49–62, January 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Colonius:1989:APO

[CS89] Fritz Colonius and Malte Sieveking. Asymptotic properties of optimal solutions in planar discounted control problems. SIAM Journal on Control and Optimization, 27(3): 608–630, May 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cho:2002:LPF

[CS02] Moon Jung Cho and Richard H.

Stockbridge. Linear programming formulation for optimal stopping problems. SIAM Journal on Control and Optimization, 40(6): 1965–1982, November 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/37766.

Clarke:2003:SCF

[CS03] F. H. Clarke and R. J. Stern. State constrained feedback stabilization. SIAM Journal on Control and Optimization, 42(2):422–441, March 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/40453.

Chu:2005:CFI

[CS05] Delin Chu and Dragoslav D. Siljak. A canonical form for the inclusion principle of dynamic systems. SIAM Journal on Control and Optimization, 44(3):969–990, May 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/60961.

Casas:2010:ABC

[CS10a] Eduardo Casas and Jan Sokolowski. Approximation of boundary control problems on curved domains. SIAM Journal on Control and Optimization, 48(6):3746–3780, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chitour:2010:SPE

[CS10b] Yacine Chitour and Mario Sigalotti. On the stabilization of persistently excited linear systems. SIAM Journal on Control and Optimization, 48(6):4032–4055, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Curtain:2011:REB

[CS11] Ruth Curtain and Amol Sasane. On Riccati equations in Banach algebras. SIAM Journal on Control and Optimization, 49(2):464– 475, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i2/p464_s1.

Cardaliaguet:2012:DGB

[CS12a] Pierre Cardaliaguet and Anne Souquière. A differential game with a blind player. SIAM Journal on Control and Optimization, 50(4):2090–2116, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:2012:OSP

[CS12b] Xin Chen and Peng Sun. Optimal structural policies for ambiguity and risk averse inventory and pricing models. SIAM Journal on Control and Optimization, 50(1): 133–146, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v50/i1/p133_s1.

Colli:2015:OCA

[CS15] Pierluigi Colli and Jürgen Sprekels. Optimal control of an Allen–Cahn equation with singular potentials and dynamic boundary condition. SIAM Journal on Control and Optimization, 53(1):213–234, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Caponigro:2018:ECP

[CS18] Marco Caponigro and Mario Sigalotti. Exact controllability in projections of the bilinear Schrödinger equation. SIAM Journal on Control and Optimization, 56(4):2901–2920, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Como:2013:SAT

[CSA^{+}13] Giacomo Como, Ketan Savla, Daron Acemoglu, Munther A. Dahleh, and Emilio Frazzoli. Stability analysis of transportation networks with multiscale driver decisions. SIAM Journal on Control and Optimization, 51(1):230–252, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chuang:1988:AEO

[CSB88] Chien-Hsiung Chuang, Jason L. Speyer, and John V. Breakwell. An asymptotic expansion for an optimal relaxation oscillator. SIAM Journal on Control and Optimization, 26(3):678–696, May 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chaves-Silva:2015:UNC

[CSB15] Felipe Wallison Chaves-Silva and Mostafa Bendahmane. Uniform null controllability for a degenerating reactiondiffusion system approximating a simplified cardiac model. SIAM Journal on Control and Optimization, 53(6):3483– 3502, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Carlone:2014:DRC

[CSBC14] L. Carlone, V. Srivastava, F. Bullo, and G. C. Calafiore. Distributed random convex programming via constraints consensus. SIAM Journal on Control and Optimization, 52 (1):629–662, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:2021:MCS

[CSBR21] Can Chen, Amit Surana, Anthony M. Bloch, and Indika Rajapakse. Multilinear control systems theory. SIAM Journal on Control and Optimization, 59(1):749–776, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cavalcanti:2014:UDR

[CSC14] Marcelo M. Cavalcanti, Flávio R. Dias Silva, and Valéria N. Domingos Cavalcanti. Uniform decay rates for the wave equation with nonlinear damping locally distributed in unbounded domains with finite measure. SIAM Journal on Control and Optimization, 52 (1):545–580, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cavalcanti:2018:SMP

[CSCV18] Marcelo M. Cavalcanti, Flávio R. Dias Silva, Valéria N. Domingos Cavalcanti, and André Vicente. Stability for the mixed problem involving the wave equation, with localized damping, in unbounded domains with finite measure. SIAM Journal on Control and Optimization, 56(4):2802–2834, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chaves-Silva:2021:SCA

[CSES21] Felipe Wallison Chaves-Silva, Sylvain Ervedoza, and Diego A. Souza. Switching controls for analytic semigroups and applications to parabolic systems. SIAM Journal on Control and Optimization, 59(4): 2820–2849, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chaves-Silva:2023:GCB

[CSFCL^{+}23] F. W. Chaves-Silva, E. Fernández-Cara, K. Le Balc’h, J. L. F. Machado, and D. A. Souza. Global controllability of the Boussinesq system with Navierslip-with-friction and Robin boundary conditions. SIAM Journal on Control and Optimization, 61(2):484–510, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1425566.

Chaves-Silva:2020:OIM

[CSSZ20] Felipe W. Chaves-Silva, Diego A.

Souza, and Can Zhang. Observability inequalities on measurable sets for the Stokes system and applications. SIAM Journal on Control and Optimization, 58(4): 2188–2205, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Colonius:2024:CCL

[CSV24] Fritz Colonius, Alexandre J.

Santana, and Eduardo C. Viscovini. Chain controllability of linear control systems. SIAM Journal on Control and Optimization, 62 (4):2387–2411, August 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1626347.

Choi:2013:SPW

[CSZ13] Jin Hyuk Choi, Mihai Sîrbu, and Gordan Zitković. Shadow prices and well-posedness in the problem of optimal investment and consumption with transaction costs. SIAM Journal on Control and Optimization, 51(6):4414–4449, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chaves-Silva:2017:CEE

[CSZZ17] Felipe W. Chaves-Silva, Xu Zhang, and Enrique Zuazua. Controllability of evolution equations with memory. SIAM Journal on Control and Optimization, 55(4):2437–2459, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cannarsa:1996:IDH

[CT96] Piermarco Cannarsa and Maria Elisabetta Tessitore. Infinite-dimensional Hamilton– Jacobi equations and Dirichlet boundary control problems of parabolic type. SIAM Journal on Control and Optimization, 34(6):1831–1847, November 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/26335.

Coron:2004:GSS

[CT04] Jean-Michel Coron and Emmanuel Trélat. Global steadystate controllability of onedimensional semilinear heat equations. SIAM Journal on Control and Optimization, 43(2):549–569, March 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42471.

Cont:2006:RLP

[CT06] Rama Cont and Peter Tankov.

Retrieving Lévy processes from option prices: Regularization of an ill-posed inverse problem. SIAM Journal on Control and Optimization, 45 (1):1–25, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Casas:2009:FSO

[CT09] Eduardo Casas and Fredi Tröltzsch. First- and secondorder optimality conditions for a class of optimal control problems with quasilinear elliptic equations. SIAM Journal on Control and Optimization, 48 (2):688–718, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cai:2013:RIS

[CT13] Chaohong Cai and Andrew R.

Teel. Robust input-to-state stability for hybrid systems. SIAM Journal on Control and Optimization, 51(2):1651– 1678, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Casas:2014:SOS

[CT14] Eduardo Casas and Fredi Tröltzsch. Second-order and stability analysis for stateconstrained elliptic optimal control problems with sparse controls. SIAM Journal on Control and Optimization, 52 (2):1010–1033, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chambrion:2019:TOC

[CT19] Thomas Chambrion and Laurent Thomann. A topological obstruction to the controllability of nonlinear wave equations with bilinear control term. SIAM Journal on Control and Optimization, 57(4): 2315–2327, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Casas:2020:OCP

[CT20] Eduardo Casas and Fredi Tröltzsch. On optimal control problems with controls appearing nonlinearly in an elliptic state equation. SIAM Journal on Control and Optimization, 58(4):1961–1983, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Claisse:2016:PMP

[CTT16] Julien Claisse, Denis Talay, and Xiaolu Tan. A pseudo-Markov property for controlled diffusion processes. SIAM Journal on Control and Optimization, 54(2):1017– 1029, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Casas:2000:SOS

[CTU00] Eduardo Casas, Fredi Tröltzsch, and Andreas Unger. Second order sufficient optimality conditions for some stateconstrained control problems of semilinear elliptic equations. SIAM Journal on Control and Optimization, 38(5): 1369–1391, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32491.

Clason:2018:VVM

[CTW18] Christian Clason, Carla Tameling, and Benedikt Wirth. Vector-valued multibang control of differential equations. SIAM Journal on Control and Optimization, 56(3):2295– 2326, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Choulli:2003:DMO

[CTZ03] Tahir Choulli, Michael Taksar, and Xun Yu Zhou. A diffusion model for optimal dividend distribution for a company with constraints on risk control. SIAM Journal on Control and Optimization, 41(6):1946–1979, November 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38266.

Chitour:2020:SPC

[CUB20] Yacine Chitour, Rosane Ushirobira, and Hassan Bouhemou. Stabilization for a perturbed chain of integrators in prescribed time. SIAM Journal on Control and Optimization, 58(2):1022–1048, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Curtain:1976:ETA

[Cur76] Ruth F. Curtain. Estimation theory for abstract evolution equations excited by general white noise processes. SIAM Journal on Control and Optimization, 14(6):1124– 1150, November 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Curtain:1984:FDC

[Cur84]  Ruth F. Curtain.  Finite

dimensional compensators for parabolic distributed systems with unbounded control and observation. SIAM Journal on Control and Optimization, 22 (2):255–276, March 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Curtain:1985:ESL

[Cur85a]  Ruth F. Curtain.  Er-

rata: “On stabilizability of linear spectral systems via state boundary feedback” [SIAM J. Control Optim. 23 (1985), no. 1, 144–152; MR 86a:93088]. SIAM Journal on Control and Optimization, 23 (6):978, November 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [Cur85b].

Curtain:1985:SLS

[Cur85b] Ruth F. Curtain. On stabilizability of linear spectral systems via state boundary feedback. SIAM Journal on Control and Optimization, 23(1): 144–152, January 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See errata [Cur85a].

Curtain:1986:ICI

[Cur86] Ruth F. Curtain. Invariance concepts in infinite dimensions. SIAM Journal on Control and Optimization, 24 (5):1009–1030, July 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Curtain:2003:RES

[Cur03] Ruth F. Curtain. Riccati equations for stable well-posed linear systems: The generic case. SIAM Journal on Control and Optimization, 42(5): 1681–1702, September 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39936.

Curtain:2011:REN

[Cur11] Ruth Curtain. Riccati equations on noncommutative Banach algebras. SIAM Journal on Control and Optimization, 49(6):2542–2557, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sicon/
resource/1/sjcodc/v49/i6/

p2542_s1. See corrections [Cur13].

Curtain:2013:CRE

[Cur13] Ruth Curtain. Corrections to “Riccati Equations on Noncommutative Banach Algebras”. SIAM Journal on Control and Optimization, 51(4): 3351–3354, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [Cur11].

Clarke:1983:LOC

[CV83] Frank H. Clarke and Richard B. Vinter. Local optimality conditions and Lipschitzian solutions to the Hamilton–Jacobi equation. SIAM Journal on Control and Optimization, 21 (6):856–870, November 1983.

CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cornet:1986:LSP

[CV86] B. Cornet and Jean-Philippe Vial. Lipschitzian solutions of perturbed nonlinear programming problems. SIAM Journal on Control and Optimization, 24(6):1123–1137, November 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Clarke:1987:RBM

[CV87] Frank H. Clarke and Richard B.

Vinter. The relationship between the maximum principle and dynamic programming. SIAM Journal on Control and Optimization, 25(5): 1291–1311, September 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Clarke:1989:AOM

[CV89a] Frank H. Clarke and Richard B.

Vinter. Applications of optimal multiprocesses. SIAM Journal on Control and Optimization, 27(5):1048–1071, September 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Clarke:1989:OM

[CV89b] Frank H. Clarke and Richard B.

Vinter. Optimal multiprocesses. SIAM Journal on Control and Optimization, 27(5):

1072–1091, September 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Clarke:1990:RPO

[CV90] F. H. Clarke and R. B. Vinter. Regularity properties of optimal controls. SIAM Journal on Control and Optimization, 28(4):980–997, July 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Camlibel:2023:PHS

[CvdS23] M. K. Camlibel and A. J. van der Schaft. Port-Hamiltonian systems theory and monotonicity. SIAM Journal on Control and Optimization, 61(4):2193–2221, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1503749.

Cortes:2005:CGC

[CvdSC05] Jorge Cortés, Arjan van der Schaft, and Peter E. Crouch. Characterization of gradient control systems. SIAM Journal on Control and Optimization, 44(4):1192–1214, July 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/42556.

Cvitanic:2000:MEL

[Cvi00]  Jaksa Cvitanic.  Minimiz-

ing expected loss of hedging in incomplete and constrained markets. SIAM Journal on Control and Optimization, 38(4):1050–1066, July 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34185.

Coron:2013:LES

[CVKB13] Jean-Michel Coron, Rafael Vazquez, Miroslav Krstic, and Georges Bastin. Local exponential H^{2} stabilization of a 2 × 2 quasilinear hyperbolic system using backstepping. SIAM Journal on Control and Optimization, 51(3): 2005–2035, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Clarke:1981:CCF

[CW81] B. M. N. Clarke and D. Williamson.

Control canonical forms and eigenvalue assignment by feedback for a class of linear hyperbolic systems. SIAM Journal on Control and Optimization, 19(6):711–729, November 1981. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Clarke:1991:SOC

[CW91] F. H. Clarke and P. R. Wolenski. The sensitivity of optimal control problems to time delay. SIAM Journal on Control and Optimization, 29(5):1176– 1215, September 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Caines:1995:CCO

[CW95] Peter E. Caines and Suning Wang. COCOLOG: a conditional observer and controller logic for finite machines. SIAM Journal on Control and Optimization, 33(6): 1687–1715, November 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/22663.

Cobb:2003:CBI

[CW03] Daniel Cobb and Chi-Jo Wang. A characterization of bounded-input boundedoutput stability for linear time-varying systems with distributional inputs. SIAM Journal on Control and Optimization, 42(4):1222–1243, July 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38483.

Curtain:2006:ESW

[CW06] Ruth F. Curtain and George Weiss. Exponential stabilization of well-posed systems by colocated feedback. SIAM Journal on Control and Optimization, 45(1):273–297, January 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chentouf:2008:RBM

[CW08] Boumediène Chentouf and Jun-Min Wang. A Riesz basis methodology for proportional and integral output regulation of a one-dimensional diffusive-wave equation. SIAM Journal on Control and Optimization, 47(5):2275–2302, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Colonius:2009:CSA

[CW09] Fritz Colonius and Tobias Wichtrey. Control systems with almost periodic excitations. SIAM Journal on Control and Optimization, 48(2): 1055–1079, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Casas:2020:FSO

[CW20] Eduardo Casas and Daniel Wachsmuth. First and second order conditions for optimal control problems with an L^{0} term in the cost functional. SIAM Journal on Control and Optimization, 58(6): 3486–3507, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Christof:2022:LSH

[CW22] Constantin Christof and Gerd Wachsmuth. Lipschitz stability and Hadamard directional differentiability for elliptic and parabolic obstacle-type quasivariational inequalities. SIAM Journal on Control and Optimization, 60(6):3430–3456, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1419635.

Casas:2023:NES

[CW23a] Eduardo Casas and Daniel Wachsmuth. A note on existence of solutions to control problems of semilinear partial differential equations. SIAM Journal on Control and Optimization, 61(3):1095–1112, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1486418.

Chen:2023:RSA

[CW23b] Xian Chen and Qingda Wei. Risk-sensitive average optimality for discrete-time Markov decision processes. SIAM Journal on Control and Optimization, 61(1):72–104, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1476757.

Christof:2023:SSO

[CW23c] Constantin Christof and Gerd Wachsmuth. Semismoothness for solution operators of obstacle-type variational inequalities with applications in optimal control. SIAM Journal on Control and Optimization, 61(3):1162–1186, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1467365.

Chen:2024:RSA

[CW24a] Xian Chen and Qingda Wei.

Risk-sensitive average Markov decision processes in general spaces. SIAM Journal on Control and Optimization, 62 (4):2115–2147, August 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M156118X.

Corella:2024:SGB

[CW24b] Alberto Domínguez Corella and Gerd Wachsmuth. Stability and genericity of bangbang controls in affine problems. SIAM Journal on Control and Optimization, 62 (3):1669–1689, June 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1586446.

Chen:2022:GCS

[CWC^{+}22] Jianqi Chen, Jieqiang Wei, Wei Chen, Henrik Sandberg, Karl Henrik Johansson, and Jie Chen. Geometrical characterization of sensor placement for cone-invariant and multi-agent systems against undetectable zero-dynamics attacks. SIAM Journal on Control and Optimization, 60(2):890–916, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1403618.

Chen:2024:PTM

[CWKQ24] Wei Chen, Dan Wang, Sei Zhen Khong, and Li Qiu. A phase theory of multi-input multioutput linear time-invariant systems. SIAM Journal on Control and Optimization, 62 (2):1235–1260, April 2024. CODEN SJCODC. ISSN 1095-7138.

Chen:2018:MCM

[CWQ18] Wei Chen, Songbai Wang, and Li Qiu. A majorization condition for multi-input networked stabilization via coding/control codesign. SIAM Journal on Control and Optimization, 56(5):3129–3148, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Casas:2017:SSO

[CWW17] Eduardo Casas, Daniel Wachsmuth, and Gerd Wachsmuth. Sufficient second-order conditions for bang-bang control problems. SIAM Journal on Control and Optimization, 55(5): 3066–3090, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Casas:2018:SOA

[CWW18] Eduardo Casas, Daniel Wachsmuth, and Gerd Wachsmuth. Secondorder analysis and numerical approximation for bang– bang bilinear control problems. SIAM Journal on Control and Optimization, 56(6): 4203–4227, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chao:2017:ASM

[CWZZ17] Zhen Chao, Kai Wang, Chao Zhu, and Yanling Zhu. Almost sure and moment exponential stability of regime-switching jump diffusions. SIAM Journal on Control and Optimization, 55(6):3458–3488, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chiou:1994:FDF

[CY94] Wen Lin Chiou and Stephen S.-T. Yau. Finite-dimensional filters with nonlinear drift. II: Brockett’s problem on classification of finite-dimensional estimation algebras. SIAM Journal on Control and Optimization, 32(1):297–310, January 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/20166.

Chen:1999:HSM

[CY99] Xiaojun Chen and Yinyu Ye. On homotopy-smoothing methods for box-constrained variational inequalities. SIAM Journal on Control and Optimization, 37(2):589–616, March 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http://

                                        epubs.siam.org/sam-bin/
                                        dbq/article/31590.

Chen:2012:PSS

[CY12] Xiaoshan Chen and Fahuai Yi. A problem of singular stochastic control with optimal stopping in finite horizon. SIAM Journal on Control and Optimization, 50(4): 2151–2172, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cao:2018:ODR

[CY18] Ping Cao and Dacheng Yao. Optimal drift rate control and impulse control for a stochastic Inventory/Production system. SIAM Journal on Control and Optimization, 56(3): 1856–1883, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:2024:SAN

[CY24] Huabin Chen and Chenggui Yuan. Stability analysis for nonlinear neutral stochastic functional differential equations. SIAM Journal on Control and Optimization, 62(2): 924–952, March 2024. CO-DEN SJCODC. ISSN 1095-7138.

Chen:1996:FDF

[CYL96] Jie Chen, Stephen S.-T. Yau, and Chi-Wah Leung. Finitedimensional filters with nonlinear drift IV: Classification of finite-dimensional estimation algebras of maximal rank with state–space dimension 3. SIAM Journal on Control and Optimization, 34(1):179–198, January 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25131.

Chen:1997:FDFa

[CYL97a] Jie Chen, Stephen S.-T. Yau, and Chi-Wah Leung. Finitedimensional filters with nonlinear drift VII: Mitter conjecture and structure of η. SIAM Journal on Control and Optimization, 35(4):1116–1131, July 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27283.

Chen:1997:FDFb

[CYL97b] Jie Chen, Stephen S. T. Yau, and Chi-Wah Leung. Finitedimensional filters with nonlinear drift VIII: Classification of finite-dimensional estimation algebras of maximal rank with state-space dimension 4. SIAM Journal on Control and Optimization, 35(4):1132–1141, July 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/27332.

Chen:2021:ITA

[CyY21] Jian-Hua Chen and Nian yu Yi. Infinite-time admissibility and exact observability of Volterra systems. SIAM Journal on Control and Optimization, 59(2):1275–1292, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cellina:1994:VOL

[CZ94] Arrigo Cellina and Sandro Zagatti. A version of Olech’s lemma in a problem of the calculus of variations. SIAM Journal on Control and Optimization, 32(4):1114–1127, July 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/23466.

Caines:1995:ACJ

[CZ95a] Peter E. Caines and Ji Feng Zhang. On the adaptive control of jump parameter systems via nonlinear filtering. SIAM Journal on Control and Optimization, 33(6): 1758–1777, November 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/23867.

Cellina:1995:ERP

[CZ95b] Arrigo Cellina and Sandro Zagatti. An existence result in a problem of the vectorial case of the calculus of variations. SIAM Journal on Control and Optimization, 33(3):960–970, May 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24764.

Castro:1998:BCH

[CZ98] Carlos Castro and Enrique Zuazua. Boundary controllability of a hybrid system consisting in two flexible beams connected by a point mass. SIAM Journal on Control and Optimization, 36(5):1576–1595, September 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31637.

Chen:1999:ALM

[CZ99] Zhiming Chen and Jun Zou. An augmented Lagrangian method for identifying discontinuous parameters in elliptic systems. SIAM Journal on Control and Optimization, 37(3):892–910, May 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/31860.

Chen:2000:SLQ

[CZ00] Shuping Chen and Xun Yu Zhou. Stochastic linear quadratic regulators with indefinite control weight costs. II. SIAM Journal on Control and Optimization, 39 (4):1065–1081, July 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/34657.

Chen:2001:MBC

[CZ01] Xiang Chen and Kemin Zhou. Multiobjective H_{∈}/H_{∞} control design. SIAM Journal on Control and Optimization, 40(2):628–660, March 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/34637.

Castro:2004:UCC

[CZ04a] Carlos Castro and Enrique Zuazua. Unique continuation and control for the heat equation from an oscillating lower dimensional manifold. SIAM Journal on Control and Optimization, 43(4):1400–1434, July 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43031.

Chen:2004:SLQ

[CZ04b] Xi Chen and Xun Yu Zhou.

Stochastic linear-quadratic control with conic control constraints on an infinite time horizon. SIAM Journal on Control and Optimization, 43(3):1120–1150, May 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42952.

Chen:2010:NMO

[CZ10] Han-Fu Chen and Wen-Xiao Zhao. New method of order estimation for ARMA/ARMAX processes. SIAM Journal on Control and Optimization, 48 (6):4157–4176, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:2004:FCR

[CZA04a] Xinjia Chen, Kemin Zhou, and Jorge L. Aravena. Fast construction of robustness degradation function. SIAM Journal on Control and Optimization, 42(6):1960–1971, November 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/40923.

Czarnecki:2004:ACP

[Cza04b] Marc-Olivier Czarnecki. Asymptotic control of pairs of oscillators coupled by a repulsion, with nonisolated equilibria II: The singular case. SIAM Journal on Control and Optimization, 42(6):2145–2171, November 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/42320.

Chen:2008:PRA

[CZA08] Xinjia Chen, Kemin Zhou, and Jorge Aravena. Probabilistic robustness analysis— risks, complexity, and algorithms. SIAM Journal on Control and Optimization, 47 (5):2693–2723, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Cheng:2021:CBA

[CZL21] Gong Cheng, Wei Zhang, and Jr-Shin Li. Combinatoricsbased approaches to controllability characterization for bilinear systems. SIAM Journal on Control and Optimization, 59(5):3574–3599, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Chen:2024:AOJ

[CZZ24] Jian-Hua Chen, Xian-Feng Zhao, and Hua-Cheng Zhou. Admissibility and observability of Jeffreys type of overdamped second order linear systems. SIAM Journal on Control and Optimization, 62 (1):466–486, February 2024. CODEN SJCODC. ISSN 1095-7138.

DaPrato:1985:SRB

[Da 85] G. Da Prato. Some results on Bellman equation in Hilbert spaces. SIAM Journal on Control and Optimization, 23(1): 61–71, January 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

DaPrato:1987:SOC

[Da 87] G. Da Prato. Synthesis of optimal control for an infinite dimensional periodic problem. SIAM Journal on Control and Optimization, 25(3):706–714, May 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

DaSilva:2014:OIE

[Da 14] Adriano J. Da Silva. Outer invariance entropy for linear systems on Lie groups. SIAM Journal on Control and Optimization, 52(6):3917–3934, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

DaSilva:2016:CLS

[Da 16] Adriano Da Silva. Controllability of linear systems on solvable Lie groups. SIAM Journal on Control and Optimization, 54(1):372–390, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

DaSilveira:2002:SBP

[DA02] M. A. Da Silveira and R. Ades. The solution of the H^{2}/H^{∞} problem by direct methods. SIAM Journal on Control and Optimization, 41 (4):1089–1117, July 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36761.

Dragan:2017:CSS

[DA17] Vasile Dragan and Samir Aberkane. Computing the stabilizing solution of a large class of stochastic game theoretic Riccati differential equations: a deterministic approximation. SIAM Journal on Control and Optimization, 55 (2):650–670, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dager:2004:OCV

[Dág04]  René Dáger.  Observation

and control of vibrations in tree-shaped networks of strings. SIAM Journal on Control and Optimization, 43(2):590–623, March 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42184.

Dager:2006:ICD

[Dág06] René Dáger. Insensitizing controls for the 1-D wave equation. SIAM Journal on Control and Optimization, 45(5): 1758–1768, ???? 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dager:2018:SIO

[Dág18] René Dáger. Simultaneous insensitivity for one-dimensional wave equations. SIAM Journal on Control and Optimization, 56(2):890–912, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dai:1988:SSS

[Dai88] Li Yi Dai. Structural stability for singular systems — a quantitative approach. SIAM Journal on Control and Optimization, 26(3):557–568, May 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dai:2012:CSS

[Dai12] Xiongping Dai. Criteria of stabilizability for switchingcontrol systems with solvable linear approximations. SIAM Journal on Control and Optimization, 50(3):1358–1388, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

DAlessandro:2015:EAC

[DAR15] Domenico D’Alessandro, Francesca Albertini, and Raffaele Romano. Exact algebraic conditions for indirect controllability of quantum systems. SIAM Journal on Control and Optimization, 53(3):1509–1542, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Datko:1980:AFL

[Dat80] Richard Datko. Applications of the finite Laplace transform to linear control problems. SIAM Journal on Control and Optimization, 18(1): 1–20, January 1980. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Datko:1988:AFS

[Dat88] Richard Datko. Not all feedback stabilized hyperbolic systems are robust with respect to small time delays in their feedbacks. SIAM Journal on Control and Optimization, 26 (3):697–713, May 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Davis:1976:RMJ

[Dav76a] M. H. A. Davis. The representation of martingales of jump processes. SIAM Journal on Control and Optimization, 14 (4):623–638, July 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Davis:1976:SPS

[Dav76b] M. H. A. Davis. The separation principle in stochastic control via Girsanov solutions. SIAM Journal on Control and Optimization, 14(1): 176–188, January 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Davis:1978:DFD

[Dav78] Jon H. Davis. A distributed filter derivation without Riccati equations. SIAM Journal on Control and Optimization, 16 (4):584–592, July 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Davis:1979:WHM

[Dav79] Jon H. Davis. Wiener–Hopf methods for open-loop unstable distributed systems. SIAM Journal on Control and Optimization, 17(6):713– 728, November 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Davini:2007:BPD

[Dav07]  Andrea Davini.  Bolza

problems with discontinuous Lagrangians and Lipschitzcontinuity of the value function. SIAM Journal on Control and Optimization, 46(5): 1897–1921, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Day:2009:URM

[Day09] Martin V. Day. A uniqueness result for p-monotone viscosity solutions of Hamilton– Jacobi equations in bounded domains. SIAM Journal on Control and Optimization, 47 (6):3167–3184, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dorfler:2012:STS

[DB12] Florian Dörfler and Francesco Bullo. Synchronization and transient stability in power networks and nonuniform Kuramoto oscillators. SIAM Journal on Control and Optimization, 50(3):1616–1642, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Deshpande:2012:DIA

[DBC12] Saurabh A. Deshpande, Dominique Bonvin, and Benoît Chachuat. Directional input adaptation in parametric optimal control problems. SIAM Journal on Control and Optimization, 50(4):1995–2024, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

diBernardo:2014:CAC

[dBLR14] Mario di Bernardo, Davide Liuzza, and Giovanni Russo. Contraction analysis for a class of nondifferentiable systems with applications to stability and network synchronization. SIAM Journal on Control and Optimization, 52 (5):3203–3227, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

diBernardo:2010:MCS

[dBMS10] Mario di Bernardo, Umberto Montanaro, and Stefania Santini. Minimal control synthesis adaptive control of continuous bimodal piecewise affine systems. SIAM Journal on Control and Optimization, 48(7): 4242–4261, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Diehl:2005:RTI

[DBS05] Moritz Diehl, Hans Georg Bock, and Johannes P. Schlöder. A real-time iteration scheme for nonlinear optimization in optimal feedback control. SIAM Journal on Control and Optimization, 43(5):1714–1736, September 2005. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/40071.

Dion:1988:MDD

[DC88] J. M. Dion and C. Commault. The minimal delay decoupling problem: Feedback implementation with stability. SIAM Journal on Control and Optimization, 26(1):66–82, January 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dufour:1999:SPD

[DC99] François Dufour and Oswaldo L. V. Costa. Stability of piecewise-deterministic Markov processes. SIAM Journal on Control and Optimization, 37(5):1483–1502, September 1999. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/33089.

Ding:2006:ISG

[DC06] Feng Ding and Tongwen Chen. On iterative solutions of general coupled matrix equations. SIAM Journal on Control and Optimization, 44(6): 2269–2284, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Djouadi:2010:TVO

[DC10] Seddik M. Djouadi and Charalambos D. Charalambous. Time-varying optimal disturbance minimization in presence of plant uncertainty. SIAM Journal on Control and Optimization, 48(5):3354– 3367, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dayawansa:1988:GIN

[DCBT88] W. Dayawansa, D. Cheng, W. M. Boothby, and T. J. Tarn. Global (f, g)-invariance of nonlinear systems. SIAM Journal on Control and Optimization, 26(5):1119–1132, September 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ding:2007:ADC

[DCI07] Feng Ding, Tongwen Chen, and Zenta Iwai. Adaptive digital control of Hammerstein nonlinear systems with limited output sampling. SIAM Journal on Control and Optimization, 45(6):2257–2276, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

DaPrato:1992:USL

[DD92] G. Da Prato and M. C.

Delfour. Unbounded solutions to the linear quadratic control problem. SIAM Journal on Control and Optimization, 30 (1):31–48, January 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

DaPrato:1999:CSB

[DD99] Giuseppe Da Prato and Arnaud Debussche. Control of the stochastic Burgers model of turbulence. SIAM Journal on Control and Optimization, 37(4):1123–1149, July 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31130.

Donchev:2003:SPI

[DD03] T. D. Donchev and A. L. Dontchev. Singular perturbations in infinite-dimensional control systems. SIAM Journal on Control and Optimization, 42(5):1795–1812, September 2003. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41608.

Deaecto:2012:POS

[DDG12] Grace S. Deaecto, Jamal Daafouz, and José C. Geromel. H_{2} and H_{∞} performance optimization of singularly perturbed switched systems. SIAM Journal on Control and Optimization, 50(3):1597– 1615, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dambrine:2015:SOQ

[DDH15] Marc Dambrine, Charles Dapogny, and Helmut Harbrecht. Shape optimization for quadratic functionals and states with random right-hand sides. SIAM Journal on Control and Optimization, 53(5): 3081–3103, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Daoulatli:2010:LED

[DDK10] M. Daoulatli, B. Dehman, and M. Khenissi. Local energy decay for the elastic system with nonlinear damping in an exterior domain. SIAM Journal on Control and Optimization, 48(8):5254–5275, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

DApice:2012:OCP

[DDK12] Ciro D’Apice, Umberto De Maio, and Ol’ga P. Kogut. Optimal control problems in coefficients for degenerate equations of monotone type: Shape stability and attainability problems. SIAM Journal on Control and Optimization, 50(3):1174–1199, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

DelMoral:2015:USP

[DDS15] Pierre Del Moral, Arnaud Doucet, and Sumeetpal S. Singh. Uniform stability of a particle approximation of the optimal filter derivative. SIAM Journal on Control and Optimization, 53(3): 1278–1304, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

DaiPra:2000:POS

[DDT00] Paolo Dai Pra, Giovanni B. Di Masi, and Barbara Trivellato. Pathwise optimality in stochastic control. SIAM Journal on Control and Optimization, 39(5): 1540–1557, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/33477.

DeSantis:1983:OIO

[De 83] Romano M. De Santis. Optimal input/output feedback with structure constraints. SIAM Journal on Control and Optimization, 21(3):427–442, May 1983. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

DeNicolao:1992:TVR

[De 92] Giuseppe De Nicolao. On the time-varying Riccati difference equation of optimal filtering. SIAM Journal on Control and Optimization, 30(6):

1251–1269, November 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

deTeresa:1998:ACS

[de 98]  Luz de Teresa.  Approx-

imate controllability of a semilinear heat equation in R^{N}. SIAM Journal on Control and Optimization, 36(6): 2128–2147, November 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32204.

DeSchutter:2000:OCC

[De 00] Bart De Schutter. Optimal control of a class of linear hybrid systems with saturation. SIAM Journal on Control and Optimization, 39(3):835–851, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35464.

DeLosReyes:2011:OCC

[De 11] Juan Carlos De Los Reyes.

Optimal control of a class of variational inequalities of the second kind. SIAM Journal on Control and Optimization, 49(4):1629–1658, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sicon/
resource/1/sjcodc/v49/i4/
p1629_s1.

DeAngelis:2015:NCF

[De 15] Tiziano De Angelis. A note on the continuity of freeboundaries in finite-horizon optimal stopping problems for one-dimensional diffusions. SIAM Journal on Control and Optimization, 53(1):167–184, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dolcetta:1984:OSO

[DE84] I. Capuzzo Dolcetta and L. C. Evans. Optimal switching for ordinary differential equations. SIAM Journal on Control and Optimization, 22(1): 143–161, January 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dehman:2014:DHF

[DE14] Belhassen Dehman and Sylvain Ervedoza. Dependence of high-frequency waves with respect to potentials. SIAM Journal on Control and Optimization, 52(6):3722–3750, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Darde:2018:RSO

[DE18] Jérémi Dardé and Sylvain Ervedoza. On the reachable set for the one-dimensional heat equation. SIAM Journal on Control and Optimization, 56(3):1692–1715, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

DellaRossa:2023:SSA

[DEJ23] Matteo Della Rossa, Lucas N. Egidio, and Raphaël M. Jungers. Stability of switched affine systems: Arbitrary and dwell-time switching. SIAM Journal on Control and Optimization, 61(4):2165–2192, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1482226.

Delfour:1986:LQO

[Del86] M. C. Delfour. The linearquadratic optimal control problem with delays in state and control variables: a state space approach. SIAM Journal on Control and Optimization, 24(5):835–883, July 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Delfour:2004:ODF

[Del04]  Michel C. Delfour.  Ori-

ented distance function and its evolution equation for initial sets with thin boundary. SIAM Journal on Control and Optimization, 42(6): 2286–2304, November 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41194.

Delfour:2007:LQD

[Del07] Michel C. Delfour. Linear quadratic differential games: Saddle point and Riccati differential equation. SIAM Journal on Control and Optimization, 46(2):750–774, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Delay:2019:LSF

[Del19] Guillaume Delay. Local stabilization of a Fluid–Structure system around a stationary state with a structure given by a finite number of parameters. SIAM Journal on Control and Optimization, 57(6): 4063–4098, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Delfour:2022:TDS

[Del22] Michel C. Delfour. Topological derivative of state-constrained objective functions: a direct method. SIAM Journal on Control and Optimization, 60(1):22–47, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1368732.

Delfour:2023:OSD

[Del23] Michel C. Delfour. One-sided derivative of parametrized minima for shape and topological derivatives. SIAM Journal on Control and Optimization, 61(3):1322–1349, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1483001.

Deng:1997:ASC

[Den97] Sien Deng. On approximate solutions in convex vector optimization. SIAM Journal on Control and Optimization, 35(6):2128–2136, November 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29256.

Deng:1998:GEB

[Den98]  Sien Deng.  Global er-

ror bounds for convex inequality systems in Banach spaces. SIAM Journal on Control and Optimization, 36(4):1240–1249, July 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29364.

Deng:2022:SON

[Den22] Li Deng. Second-order necessary conditions for optimal control problems with endpoints-constraints and convex control-constraints. SIAM Journal on Control and Optimization, 60(5):2737–2758, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1378260.

DeAngelis:2024:MPS

[DEO24] Tiziano De Angelis, Erik Ekström, and Marcus Olofsson. The maximality principle in singular control with absorption and its applications to the dividend problem. SIAM Journal on Control and Optimization, 62(1):91–117, January 2024. CODEN SJCODC. ISSN 1095-7138.

Deckelnick:2009:OCP

[DES09] Klaus Deckelnick, Charles M. Elliott, and Vanessa Styles. Optimal control of the propagation of a graph in inhomogeneous media. SIAM Journal on Control and Optimization, 48(3):1335–1352, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dewilde:1976:IOD

[Dew76] P. Dewilde. Input–output description of roomy systems. SIAM Journal on Control and Optimization, 14(4):712–736, July 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dorroh:1994:MMP

[DF94] J. R. Dorroh and Guillermo Ferreyra. A multistate, multicontrol problem with unbounded controls. SIAM Journal on Control and Optimization, 32(5):1322–1331, September 1994. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/22982.

Davis:1996:TRP

[DF96] Mark H. A. Davis and Mohammad Farid. A target recognition problem: Sequential analysis and optimal control. SIAM Journal on Control and Optimization, 34(6): 2116–2132, November 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27369. See cor-

rection [DF98a].

Davis:1998:CTR

[DF98a] Mark H. A. Davis and Mohammad Farid. Corrections to “A Target Recognition Problem: Sequential Analysis and Optimal Control”. SIAM Journal on Control and Optimization, 36(4):1481–1484, July 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32030.  See

[DF96].

Donchev:1998:SEA

[DF98b] Tzanko Donchev and Elza Farkhi. Stability and Euler approximation of onesided Lipschitz differential inclusions. SIAM Journal on Control and Optimization, 36(2):780–796, March 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/29369.

Dorroh:1998:FBP

[DF98c] J. R. Dorroh and Guillermo Ferreyra. A free boundary problem in R^{d} with both smooth and nonsmooth fit. SIAM Journal on Control and Optimization, 36 (2):579–589, March 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/30144.

Dunyak:1998:ORT

[DF98d] James P. Dunyak and Mark I. Freidlin. Optimal residence time control of Hamiltonian systems perturbed by white noise. SIAM Journal on Control and Optimization, 36 (1):233–252, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/29165.

Dolgopolik:2016:SGC

[DF16] M. V. Dolgopolik and A. L. Fradkov. Speed-gradient control of the Brockett integrator. SIAM Journal on Control and Optimization, 54(4): 2116–2131, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dianetti:2020:NSS

[DF20] Jodi Dianetti and Giorgio Ferrari. Nonzero-sum submodular monotone-follower games: Existence and approximation of Nash equilibria. SIAM Journal on Control and Optimization, 58(3):1257–1288, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dolgopolik:2021:FDN

[DF21] M. V. Dolgopolik and A. L.

Fradkov. Finite-differential nonsmooth speed-gradient control: Stability, passivity, robustness. SIAM Journal on Control and Optimization, 59(2):1370–1392, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Doubova:2002:CPS

[DFCGBZ02] A. Doubova, E. Fernández-Cara, M. González-Burgos, and E. Zuazua. On the controllability of parabolic systems with a nonlinear term involving the state and the gradient. SIAM Journal on Control and Optimization, 41(3):798–819, May 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38646.

Dou:2023:GCE

[DFLZ23] Fangfang Dou, Xiaoyu Fu, Zhonghua Liao, and Xiaomin Zhu. Global Carleman estimate and state observation problem for Ginzburg–Landau equation. SIAM Journal on Control and Optimization, 61 (5):2981–2996, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1522516.

DeAngelis:2015:NSS

[DFM15] Tiziano De Angelis, Giorgio Ferrari, and John Moriarty. A nonconvex singular stochastic control problem and its related optimal stopping boundaries. SIAM Journal on Control and Optimization, 53(3): 1199–1223, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

deFrutos:2023:OBN

[dFN23] Javier de Frutos and Julia Novo. Optimal bounds for numerical approximations of infinite horizon problems based on dynamic programming approach. SIAM Journal on Control and Optimization, 61(2):415–433, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1459290.

Diblik:2014:NCF

[DFP14] J. Diblík, M. Feckan, and M. Pospísil. On the new control functions for linear discrete delay systems. SIAM Journal on Control and Optimization, 52(3):1745–1760, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

deFreitas:2015:SGT

[dFS15] Michael Marcondes de Freitas and Eduardo D. Sontag. A small-gain theorem for random dynamical systems with inputs and outputs. SIAM Journal on Control and Optimization, 53(4): 2657–2695, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

DeFeo:2024:OCS

[DFŚ24] Filippo De Feo, Salvatore Federico, and Andrzej Świech. Optimal control of stochastic delay differential equations and applications to pathdependent financial and economic models. SIAM Journal on Control and Optimization, 62(3):1490–1520, June 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1553960.

DiPillo:1979:NCA

[DG79] G. Di Pillo and L. Grippo. A new class of augmented Lagrangians in nonlinear programming. SIAM Journal on Control and Optimization, 17 (5):618–628, September 1979. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

DiPillo:1985:CDE

[DG85] G. Di Pillo and L. Grippo. A continuously differentiable exact penalty function for nonlinear programming problems with inequality constraints. SIAM Journal on Control and Optimization, 23 (1):72–84, January 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

DiPillo:1989:EPF

[DG89] G. Di Pillo and L. Grippo. Exact penalty functions in constrained optimization. SIAM Journal on Control and Optimization, 27(6):1333–1360, November 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dauer:1990:PPE

[DG90] Jerald P. Dauer and Richard J. Gallagher. Positive proper efficient points and related cone results in vector optimization theory. SIAM Journal on Control and Optimization, 28(1): 158–172, January 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

deGournay:2006:VEL

[dG06] Frédéric de Gournay. Velocity extension for the level-set method and multiple eigenvalues in shape optimization. SIAM Journal on Control and Optimization, 45(1): 343–367, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

DElia:2014:ODC

[DG14] Marta D’Elia and Max Gunzburger. Optimal distributed control of nonlocal steady diffusion problems. SIAM Journal on Control and Optimization, 52(1):243–273, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Diallo:2016:CMN

[DG16] Tidiane Diallo and Dan Goreac. Controllability metrics on networks with linear decision process–type interactions and multiplicative noise. SIAM Journal on Control and Optimization, 54(6): 3126–3151, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dufour:2020:CPA

[DG20] François Dufour and Alexandre Genadot. A convex programming approach for discrete-time Markov decision processes under the expected total reward criterion. SIAM Journal on Control and Optimization, 58(4):2535–2566, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dorsz:2023:ECV

[DG23] Justine Dorsz and Olivier Glass. Exact controllability of the vortex system by means of a single vortex. SIAM Journal on Control and Optimization, 61(6):3316–3340, November 2023. CODEN SJCODC. ISSN 1095-7138.

Deckelnick:2009:FEA

[DGH09] Klaus Deckelnick, Andreas Günther, and Michael Hinze. Finite element approximation of Dirichlet boundary control for elliptic PDEs on two- and three-dimensional curved domains. SIAM Journal on Control and Optimization, 48(4): 2798–2819, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

DeSantis:1993:AAR

[DGJ93] A. De Santis, A. Germani, and L. Jetto. Approximation of the algebraic Riccati equation in the Hilbert space of Hilbert–Schmidt operators. SIAM Journal on Control and Optimization, 31(4):847–874, July 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

DiLoreto:2010:DBI

[DGKL10] Michael Di Loreto, Stéphane Gaubert, Ricardo D. Katz, and Jean-Jacques Loiseau. Duality between invariant spaces for max-plus linear discrete event systems. SIAM Journal on Control and Optimization, 48(8):5606– 5628, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v48/i8/p5606_s1.

DiBenedetto:1989:RIN

[DGM89] M. D. Di Benedetto, J. W.

Grizzle, and C. H. Moog. Rank invariants of nonlinear systems. SIAM Journal on Control and Optimization, 27 (3):658–672, May 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

DElia:2019:PEE

[DGO19] Marta D’Elia, Christian Glusa, and Enrique Otárola. A priori error estimates for the optimal control of the integral fractional Laplacian. SIAM Journal on Control and Optimization, 57(4):2775–2798, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Discacciati:2013:ICD

[DGQ13] Marco Discacciati, Paola Gervasio, and Alfio Quarteroni.

The Interface Control Domain Decomposition (ICDD) method for elliptic problems. SIAM Journal on Control and Optimization, 51(5):3434– 3458, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Das:2018:RSP

[DGR18] Milan Kumar Das, Anindya Goswami, and Nimit Rana. Risk sensitive portfolio optimization in a jump diffusion model with regimes. SIAM Journal on Control and Optimization, 56(2):1550–1576, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Damm:2014:ETT

[DGSW14] Tobias Damm, Lars Grüne, Marleen Stieler, and Karl Worthmann. An exponential turnpike theorem for dissipative discrete time optimal control problems. SIAM Journal on Control and Optimization, 52(3):1935–1957, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Davis:2010:ICM

[DGW10] Mark H. A. Davis, Xin Guo, and Guoliang Wu. Impulse control of multidimensional jump diffusions. SIAM Journal on Control and Optimization, 48(8):5276–5293, ???? 2010. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Datta:1984:DMC

[DH84] K. B. Datta and M. L. J.

Hautus. Decoupling of multivariable control systems over unique factorization domains. SIAM Journal on Control and Optimization, 22(1):28– 39, January 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dontchev:1993:LSN

[DH93] Asen L. Dontchev and William W.

Hager. Lipschitzian stability in nonlinear control and optimization. SIAM Journal on Control and Optimization, 31 (3):569–603, May 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

DeStefano:1998:EUO

[DH98a] Alisa DeStefano and G. R.

Hall. An example of a universally observable flow on the torus. SIAM Journal on Control and Optimization, 36(4):1207–1224, July 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30841.

Dontchev:1998:LSS

[DH98b] A. L. Dontchev and W. W.

Hager. Lipschitzian stability for state constrained nonlinear optimal control. SIAM Journal on Control and Optimization, 36(2):698–718, March 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29931.

Demetriou:2009:ESD

[DH09] Michael A. Demetriou and Islam I. Hussein. Estimation of spatially distributed processes using mobile spatially distributed sensor network. SIAM Journal on Control and Optimization, 48(1): 266–291, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Donnelly:2012:QRM

[DH12] Catherine Donnelly and Andrew J. Heunis. Quadratic risk minimization in a regimeswitching model with portfolio constraints. SIAM Journal on Control and Optimization, 50(4):2431–2461, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

DelMoral:2022:NRM

[DH22] Pierre Del Moral and Emma Horton. A note on Riccati matrix difference equations. SIAM Journal on Control and Optimization, 60 (3):1393–1409, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1437226.

Duc:2019:ASS

[DHC19] Luu Hoang Duc, Phan Thanh Hong, and Nguyen Dinh Cong. Asymptotic stability for stochastic dissipative systems with a Hölder noise. SIAM Journal on Control and Optimization, 57(4):3046–3071, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dong:2020:SFC

[DHJK20] Jiu-Gang Dong, Seung-Yeal Ha, Jinwook Jung, and Doheon Kim. On the stochastic flocking of the Cucker–Smale flock with randomly switching topologies. SIAM Journal on Control and Optimization, 58(4):2332–2353, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dong:2020:CSE

[DHK20] Jiu-Gang Dong, Seung-Yeal Ha, and Doheon Kim. On the Cucker–Smale ensemble with the q-closest neighbors in a self-consistent temperature field. SIAM Journal on Control and Optimization, 58(1): 368–392, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dmitrishin:2018:LRS

[DHKS18] D. Dmitrishin, P. Hagelstein, A. Khamitova, and A. Stokolos. Limitations of robust stability of a linear delayed feedback control. SIAM Journal on Control and Optimization, 56(1):148–157, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Djehiche:2009:FHO

[DHP09] Boualem Djehiche, Said Hamadène, and Alexandre Popier. A finite horizon optimal multiple switching problem. SIAM Journal on Control and Optimization, 48(4):2751–2770, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Delyon:2019:NDP

[DHP19] Alexandre Delyon, Antoine Henrot, and Yannick Privat. Nondispersal and density properties of infinite packings. SIAM Journal on Control and Optimization, 57(2): 1467–1492, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Duncan:2000:SCF

[DHPD00] Tyrone E. Duncan, Yaozhong Hu, and Bozenna Pasik-Duncan. Stochastic calculus for fractional Brownian motion I. theory. SIAM Journal on Control and Optimization, 38(2):582–612, March 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34171.

Dennis:1998:TRI

[DHV98] J. E. Dennis, Matthias Heinkenschloss, and Luís N. Vicente. Trust-region interiorpoint SQP algorithms for a class of nonlinear programming problems. SIAM Journal on Control and Optimization, 36(5):1750–1794, September 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27903.

Dontchev:2000:UCM

[DHV00] A. L. Dontchev, W. W. Hager, and V. M. Veliov. Uniform convergence and mesh independence of Newton’s method for discretized variational problems. SIAM Journal on Control and Optimization, 39(3):961–980, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33857.

Dai:2008:ASS

[DHX08] Xiongping Dai, Yu Huang, and MingQing Xiao. Almost sure stability of discretetime switched linear systems: a topological point of view. SIAM Journal on Control and Optimization, 47(4):2137– 2156, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dai:2010:CSC

[DHX10] Xiongping Dai, Yu Huang, and MingQing Xiao. Criteria of stability for continuoustime switched systems by using liao-type exponents. SIAM Journal on Control and Optimization, 48(5):3271–3296, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

DiBlasio:1991:OCI

[Di 91] G. Di Blasio. Optimal control with infinite horizon for distributed parameter systems with constrained controls. SIAM Journal on Control and Optimization, 29(4): 909–925, July 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

DiBenedetto:1986:MNM

[DI86] Maria Domenica Di Benedetto and Alberto Isidori. The matching of nonlinear models via dynamic state feedback. SIAM Journal on Control and Optimization, 24(5): 1063–1075, July 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

DaPrato:1987:OCL

[DI87] G. Da Prato and A. Ichikawa.

Optimal control of linear systems with almost periodic inputs. SIAM Journal on Control and Optimization, 25(4):1007–1019, July 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See correction [JLZ98].

DaPrato:1990:QCL

[DI90] Giuseppe Da Prato and Akira Ichikawa. Quadratic control for linear time-varying systems. SIAM Journal on Control and Optimization, 28(2): 359–381, March 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

DaPrato:1993:OCI

[DI93] Giuseppe Da Prato and Akira Ichikawa. Optimal control for integrodifferential equations of parabolic type. SIAM Journal on Control and Optimization, 31(5):1167–1182, September 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Desai:1994:OCN

[DI94] Mihir Desai and Kazufumi Ito.

Optimal controls of Navier– Stokes equations. SIAM Journal on Control and Optimization, 32(5):1428–1446, September 1994. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/22497.

Dickinson:1976:CLS

[Dic76] Bradley W. Dickinson. The classification of linear state variable control laws. SIAM Journal on Control and Optimization, 14(3):467–477, May 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dietrich:1987:DTO

[Die87] H. Dietrich. Duality theorems for an optimal control problem with a linear unbounded operator. SIAM Journal on Control and Optimization, 25(5): 1192–1201, September 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Drusvyatskiy:2015:CD

[DIL15] D. Drusvyatskiy, A. D. Ioffe, and A. S. Lewis. Curves of descent. SIAM Journal on Control and Optimization, 53(1): 114–138, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dinuzzo:2015:KLT

[Din15] Francesco Dinuzzo. Kernels for linear time invariant system identification. SIAM Journal on Control and Optimization, 53(5):3299–3317, ???? 2015. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

deJong:1978:NAR

[dJ78] Lieuwe Sytse de Jong. Numerical aspects of recursive realization algorithms. SIAM Journal on Control and Optimization, 16(4):646–659, July 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). Reprinted and cited in [?].

Dontchev:1986:RKT

[DJ86] A. L. Dontchev and H. Th.

Jongen. On the regularity of the Kuhn–Tucker curve. SIAM Journal on Control and Optimization, 24(2):169– 176, March 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Delyon:1995:ASP

[DJ95] Bernard Delyon and Anatoli Juditsky. Asymptotical study of parameter tracking algorithms. SIAM Journal on Control and Optimization, 33(1):323–345, January 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23895.

Dupuis:1998:RCA

[DJ98] Paul Dupuis and Matthew R.

James. Rates of convergence for approximation schemes in optimal control. SIAM Journal on Control and Optimization, 36(2):719–741, March 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26778.

DePersis:2013:CPS

[DJ13] Claudio De Persis and Bayu Jayawardhana. Coordination of passive systems under quantized measurements. SIAM Journal on Control and Optimization, 50(6):3155–3177, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Duan:2021:GTS

[DJB21] Xiaoming Duan, Saber Jafarpour, and Francesco Bullo. Graph-theoretic stability conditions for Metzler matrices and monotone systems. SIAM Journal on Control and Optimization, 59(5):3447–3471, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dai:2023:GOS

[DJK^{+}23] Yin Dai, Yuling Jiao, Lican Kang, Xiliang Lu, and Jerry Zhijian Yang. Global optimization via Schrödinger– Föllmer diffusion. SIAM Journal on Control and Optimization, 61(5):2953–2980, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1477738.

Dai:2009:FHO

[DJLY09] Min Dai, Lishang Jiang, Peifan Li, and Fahuai Yi. Finite horizon optimal investment and consumption with transaction costs. SIAM Journal on Control and Optimization, 48(2):1134–1154, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ding:1996:CLP

[DJZ96] Zhonghai Ding, Link Ji, and Jianxin Zhou. Constrained LQR problems in elliptic distributed control systems with point observations. SIAM Journal on Control and Optimization, 34(1):264–294, January 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25379.

Dolecki:1978:CEP

[DK78] Szymon Dolecki and Stanisław Kurcyusz. On Φconvexity in extremal problems. SIAM Journal on Control and Optimization, 16(2): 277–300, March 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dupuis:1985:SAL

[DK85] Paul Dupuis and Harold J.

Kushner. Stochastic approximations via large deviations: Asymptotic properties. SIAM Journal on Control and Optimization, 23(5):675– 696, September 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dupuis:1986:LDE

[DK86] Paul Dupuis and Harold J. Kushner. Large deviations estimates for systems with small noise effects, and applications to stochastic systems theory. SIAM Journal on Control and Optimization, 24(5):979–1008, July 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dupuis:1989:MEP

[DK89a] Paul Dupuis and Harold Kushner. Minimizing escape probabilities: a large deviations approach. SIAM Journal on Control and Optimization, 27 (2):432–445, March 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dupuis:1989:SAL

[DK89b] Paul Dupuis and Harold J. Kushner. Stochastic approximation and large deviations: upper bounds and w.p.1 convergence. SIAM Journal on Control and Optimization, 27(5):1108–1135, September 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dahmen:2005:AWM

[DK05] Wolfgang Dahmen and Angela Kunoth. Adaptive wavelet methods for linear-quadratic elliptic control problems: Convergence rates. SIAM Journal on Control and Optimization, 43(5):1640–1675, September 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41919.

Denis:2013:OIU

[DK13] Laurent Denis and Magali Kervarec. Optimal investment under model uncertainty in nondominated models. SIAM Journal on Control and Optimization, 51(3):1803–1822, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dontchev:2019:LSD

[DKK^{+}19] Asen L. Dontchev, Ilya Kolmanovsky, Mikhail I. Krastanov, Marco Nicotra, and Vladimir M. Veliov. Lipschitz stability in discretized optimal control with application to SQP. SIAM Journal on Control and Optimization, 57 (1):468–489, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dahl:2009:FWU

[DKL09] Matias F. Dahl, Anna Kirpichnikova, and Matti Lassas. Focusing waves in unknown media by modified time reversal iteration. SIAM Journal on Control and Optimization, 48 (2):839–858, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Diblik:2008:CLD

[DKR08] Josef Diblík, Denys Ya. Khusainov, and M. Růžičková. Controllability of linear discrete systems with constant coefficients and pure delay. SIAM Journal on Control and Optimization, 47(3):1140– 1149, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dontchev:2013:ENC

[DKRV13] A. L. Dontchev, M. I. Krastanov, R. T. Rockafellar, and V. M. Veliov. An Euler– Newton continuation method for tracking solution trajectories of parametric variational inequalities. SIAM Journal on Control and Optimization, 51(3):1823–1840, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Druet:2011:OCT

[DKS^{+}11] Pierre-Étienne Druet, Olaf Klein, Jürgen Sprekels, Fredi Tröltzsch, and Irwin Yousept. Optimal control of threedimensional state-constrained induction heating problems with nonlocal radiation effects. SIAM Journal on Control and Optimization, 49(4): 1707–1736, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i4/p1707_s1.

Dashkovskiy:2023:RSN

[DKS23] Sergey Dashkovskiy, Oleksiy Kapustyan, and Vitalii Slynko. Robust stability of a nonlinear ODE-PDE system. SIAM Journal on Control and Optimization, 61 (3):1760–1777, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1467535.

DelMoral:2017:SUP

[DKT17] Pierre Del Moral, Aline Kurtzmann, and Julian Tugaut. On the stability and the uniform propagation of chaos of a class of extended ensemble Kalman–Bucy filters. SIAM Journal on Control and Optimization, 55(1):119–155, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

DaLio:2006:URS

[DL06] Francesca Da Lio and Olivier Ley. Uniqueness results for second-order Bellman–Isaacs equations under quadratic growth assumptions and applications. SIAM Journal on Control and Optimization, 45 (1):74–106, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dehman:2009:AHC

[DL09] B. Dehman and G. Lebeau. Analysis of the HUM control operator and exact controllability for semilinear waves in uniform time. SIAM Journal on Control and Optimization, 48(2):521–550, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Davis:2013:JDR

[DL13a] Mark Davis and Sébastien Lleo. Jump-diffusion risksensitive asset management II: Jump-diffusion factor model. SIAM Journal on Control and Optimization, 51(2):1441– 1480, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

DeDona:2013:BSI

[DL13b] José A. De Doná and Jean Lévine. On barriers in state and input constrained nonlinear systems. SIAM Journal on Control and Optimization, 51(4):3208–3234, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dou:2019:PAC

[DL19] Fangfang Dou and Qi Lu. Partial approximate controllability for linear stochastic control systems. SIAM Journal on Control and Optimization, 57(2):1209–1229, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dou:2020:TIL

[DL20] Fangfang Dou and Qi Lü.

Time-inconsistent linear quadratic optimal control problems for stochastic evolution equations. SIAM Journal on Control and Optimization, 58(1):485–509, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Datko:1986:EET

[DLP86] R. Datko, J. Lagnese, and M. P. Polis. An example on the effect of time delays in boundary feedback stabilization of wave equations. SIAM Journal on Control and Optimization, 24(1):152–156, January 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

delosReyes:2016:OCS

[dlRHM16] Juan Carlos de los Reyes, Roland Herzog, and Christian Meyer. Optimal control of static elastoplasticity in primal formulation. SIAM Journal on Control and Optimization, 54(6):3016–3039, ????

2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

delosReyes:2007:OCS

[dlRT07] J. C. de los Reyes and F. Tröltzsch. Optimal control of the stationary Navier– Stokes equations with mixed control-state constraints. SIAM Journal on Control and Optimization, 46(2):604–629, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

David:1983:UMI

[DM83] A. J. David and G. G. L. Meyer. Unstructured mean iterative processes in reflexive Banach spaces. SIAM Journal on Control and Optimization, 21(1):140–152, January 1983. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dayawansa:1989:DDI

[DM89] Wijesuriya P. Dayawansa and Clyde F. Martin. Disturbance decoupling, (f, g)-invariant and controllability subspaces of a class of homogeneous polynomial systems. SIAM Journal on Control and Optimization, 27(1):108–119, January 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dupuis:1997:RSR

[DM97] Paul Dupuis and William M. McEneaney. Risk-sensitive and robust escape criteria. SIAM Journal on Control and Optimization, 35(6): 2021–2049, November 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28162.

DelMoral:1999:CAG

[DM99] P. Del Moral and L. Miclo.

On the convergence and applications of generalized simulated annealing. SIAM Journal on Control and Optimization, 37(4):1222–1250, July 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31398.

DaLio:2002:FTH

[DM02a] Francesca Da Lio and William M.

McEneaney. Finite Time– Horizon risk-sensitive control and the robust limit under a quadratic growth assumption. SIAM Journal on Control and Optimization, 40(5): 1628–1661, September 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34515.

Dufour:2002:GSN

[DM02b] F. Dufour and Boris M. Miller.

Generalized solutions in nonlinear stochastic control problems. SIAM Journal on Control and Optimization, 40(6): 1724–1745, November 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/37422.

Dufour:2004:SSC

[DM04] F. Dufour and B. Miller. Singular stochastic control problems. SIAM Journal on Control and Optimization, 43(2):708–730, March 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/41271.

Dufour:2006:MPS

[DM06] F. Dufour and B. Miller. Maximum principle for singular stochastic control problems. SIAM Journal on Control and Optimization, 45(2): 668–698, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Degiovanni:2009:ELE

[DM09] Marco Degiovanni and Marco Marzocchi. On the Euler– Lagrange equation for functionals of the calculus of variations without upper growth conditions. SIAM Journal on Control and Optimization, 48 (4):2857–2870, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dashkovskiy:2013:ISS

[DM13a] Sergey Dashkovskiy and Andrii Mironchenko. Input-to-state stability of nonlinear impulsive systems. SIAM Journal on Control and Optimization, 51(3):1962–1987, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Du:2013:MPO

[DM13b] Kai Du and Qingxin Meng. A maximum principle for optimal control of stochastic evolution equations. SIAM Journal on Control and Optimization, 51(6):4343–4362, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dower:2015:MPD

[DM15] Peter M. Dower and William M.

McEneaney. A max-plus dual space fundamental solution for a class of operator differential Riccati equations. SIAM Journal on Control and Optimization, 53(2):969–1002, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dower:2017:STP

[DM17] Peter M. Dower and William M.

McEneaney. Solving two-point boundary value problems for a wave equation via the principle of stationary action and optimal control. SIAM Journal on Control and Optimization, 55(4):2151–2205, ????

2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

deMelo:2022:MDO

[dMCJ22] Maisa Kely de Melo, Rodrigo T. N. Cardoso, and Tales A. Jesus. MultiObjective dynamic optimization of investment portfolio based on model predictive control. SIAM Journal on Control and Optimization, 60(1):104–123, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1346420.

Dmitruk:1992:GEN

[Dmi92] A. V. Dmitruk. A generalized estimate of the number of zeros for solutions of a class of linear differential equations. SIAM Journal on Control and Optimization, 30(5): 1087–1091, September 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dayawansa:1990:ASC

[DMK90] W. P. Dayawansa, C. F. Martin, and G. Knowles. Asymptotic stabilization of a class of smooth two-dimensional systems. SIAM Journal on Control and Optimization, 28(6): 1321–1349, November 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Duncan:1994:ABP

[DMPD94] T. E. Duncan, B. Maslowski, and B. Pasik-Duncan. Adaptive boundary and point control of linear stochastic distributed parameter systems. SIAM Journal on Control and Optimization, 32(3):648–672, May 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/22872.

Duncan:1998:EBP

[DMPD98] T. E. Duncan, B. Maslowski, and B. Pasik-Duncan. Ergodic boundary/point control of stochastic semilinear systems. SIAM Journal on Control and Optimization, 36(3):1020–1047, May 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30319.

Duncan:2000:ACS

[DMPD00] T. E. Duncan, B. Maslowski, and B. Pasik-Duncan. Adaptive control for semilinear stochastic systems. SIAM Journal on Control and Optimization, 38(6):1683–1706, November 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/35182.

Duncan:2012:LQC

[DMPD12] T. E. Duncan, B. Maslowski, and B. Pasik-Duncan. Linearquadratic control for stochastic equations in a Hilbert space with fractional Brownian motions. SIAM Journal on Control and Optimization, 50(1): 507–531, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Duprez:2019:AEC

[DMR19] Michel Duprez, Morgan Morancey, and Francesco Rossi. Approximate and exact controllability of the continuity equation with a localized vector field. SIAM Journal on Control and Optimization, 57(2):1284–1311, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dragan:2002:API

[DMS02] Vasile Dragan, Toader Morozan, and Peng Shi. Asymptotic properties of inputoutput operators norm associated with singularly perturbed systems with multiplicative white noise. SIAM Journal on Control and Optimization, 41(1):141–163, January 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/36944.

Delgado:2003:SOH

[DMS03] M. Delgado, J. A. Montero, and A. Suárez. Study of the optimal harvesting control and the optimality system for an elliptic problem. SIAM Journal on Control and Optimization, 42(5): 1559–1577, September 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41090.

Dragan:2012:LQR

[DMS12] Vasile Dragan, Hiroaki Mukaidani, and Peng Shi. The linear quadratic regulator problem for a class of controlled systems modeled by singularly perturbed Itô differential equations. SIAM Journal on Control and Optimization, 50 (1):448–470, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Du:2022:LAD

[DMZ22] Kai Du, Qingxin Meng, and Fu Zhang. A Q-learning algorithm for discrete-time linearquadratic control with random parameters of unknown distribution: Convergence and stabilization. SIAM Journal on Control and Optimization, 60(4):1991–2015, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1379605.

DSouza:2023:ALT

[DN23] Rollen S. D’Souza and Christopher Nielsen. An algorithm for local transverse feedback linearization. SIAM Journal on Control and Optimization, 61(3):1248–1272, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1444588.

dAndrea-Novel:2020:STS

[dNCP20] Brigitte d’Andréa Novel, Jean-Michel Coron, and Wilfrid Perruquetti. Small-time stabilization of homogeneous cascaded systems with application to the unicycle and the slider examples. SIAM Journal on Control and Optimization, 58(5):2997–3018, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Droniou:2017:GDM

[DNS17] Jérome Droniou, Neela Nataraj, and Devika Shylaja. The gradient discretization method for optimal control problems, with superconvergence for nonconforming finite elements and mixed-hybrid mimetic finite differences. SIAM Journal on Control and Optimization, 55(6):3640–3672, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dmitruk:2014:NCW

[DO14] A. V. Dmitruk and N. P. Osmolovskii. Necessary conditions for a weak minimum in optimal control problems with integral equations subject to state and mixed constraints. SIAM Journal on Control and Optimization, 52 (6):3437–3462, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dmitruk:2022:LMP

[DO22] Andrei Dmitruk and Nikolai Osmolovskii. Local minimum principle for an optimal control problem with a nonregular mixed constraint. SIAM Journal on Control and Optimization, 60(4):1919–1941, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1411469.

Doan:2021:FTA

[Doa21]  Thinh T. Doan.  Finite-

time analysis and restarting scheme for linear two-time-scale stochastic approximation. SIAM Journal on Control and Optimization, 59(4): 2798–2819, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

deOliveira:2005:FDS

[dOG05] Maurício C. de Oliveira and José C. Geromel. H_{2} and H_{∞} filtering design subject to implementation uncertainty. SIAM Journal on Control and Optimization, 44(2):515–530, March 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42472.

Dokuchaev:2005:OSI

[Dok05] Nikolai Dokuchaev. Optimal solution of investment problems via linear parabolic equations generated by Kalman filter. SIAM Journal on Control and Optimization, 44(4):1239–1258, July 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42557.

Dokuchaev:2017:FOB

[Dok17] Nikolai Dokuchaev. First order BSPDEs in higher dimension for optimal control problems. SIAM Journal on Control and Optimization, 55(2): 818–834, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dolezal:1976:GHN

[Dol76a]  Vaclav Dolezal.  General-

ized Hilbert networks. SIAM Journal on Control and Optimization, 14(1):26–41, January 1976. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Dolezal:1976:LHN

[Dol76b] Vaclav Dolezal. Linear Hilbert networks containing finitely many nonlinear elements. SIAM Journal on Control and Optimization, 14(3):478–494, May 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dolezal:1979:FSD

[Dol79] Vaclav Dolezal. Feedback systems described by monotone operators. SIAM Journal on Control and Optimization, 17 (3):339–364, May 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dontchev:1996:PED

[Don96] Asen L. Dontchev. An a priori estimate for discrete approximations in nonlinear optimal control. SIAM Journal on Control and Optimization, 34(4):1315–1328, July 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26948.

Donchev:2010:SPE

[Don10] T. Donchev. Singularly perturbed evolution inclusions. SIAM Journal on Control and Optimization, 48(7):4572– 4590, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dong:2024:ROS

[Don24]  Yuchao Dong.  Random-

ized optimal stopping problem in continuous time and reinforcement learning algorithm. SIAM Journal on Control and Optimization, 62 (3):1590–1614, June 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1516725.

deOliveira:2003:ANF

[dOP03] Paula Milheiro de Oliveira and Jean Picard. Approximate nonlinear filtering for a two-dimensional diffusion with one-dimensional observations in a low noise channel. SIAM Journal on Control and Optimization, 41(6): 1801–1819, November 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36392.

Dorea:1986:LDR

[Dor86] Chang C. Y. Dorea. Limiting distribution for random optimization methods. SIAM Journal on Control and Optimization, 24(1):76–82, January 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dorea:1990:SRR

[Dor90] C. C. Y. Dorea. Stopping rules for a random optimization method. SIAM Journal on Control and Optimization, 28(4):841–850, July 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Doyen:1994:IFT

[Doy94] Luc Doyen. Inverse function theorems and shape optimization. SIAM Journal on Control and Optimization, 32(6): 1621–1642, November 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/22634.

Duncan:2013:LQF

[DPD13] Tyrone E. Duncan and Bozenna Pasik-Duncan. Linear-quadratic fractional Gaussian control. SIAM Journal on Control and Optimization, 51(6):4504– 4519, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Duncan:1998:DML

[DPDS98] T. E. Duncan, B. Pasik-Duncan, and Ł. Stettner. Discretized maximum likelihood and almost optimal adaptive control of ergodic Markov models. SIAM Journal on Control and Optimization, 36(2):422–446, March 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29836.

Dufour:2013:FLP

[DPR13] François Dufour and Tomás Prieto-Rumeau. Finite linear programming approximations of constrained discounted Markov decision processes. SIAM Journal on Control and Optimization, 51(2): 1298–1324, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dufour:2022:SMN

[DPR22] François Dufour and Tomás Prieto-Rumeau. Stationary Markov Nash equilibria for nonzero-sum constrained ARAT Markov games. SIAM Journal on Control and Optimization, 60(2):945–967, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M144565X.

DeBlasi:2004:BTT

[DPT04] F. S. De Blasi, G. Pianigiani, and A. A. Tolstonogov. A Bogolyubov-type theorem with a nonconvex constraint in Banach spaces. SIAM Journal on Control and Optimization, 43(2):466–476, March 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42315.

Dontchev:2000:ENS

[DPV00] Asen L. Dontchev, Michael P.

Polis, and Vladimir M. Veliov. On the effect of neglecting sensor dynamics in parameter identification problems. SIAM Journal on Control and Optimization, 38(4):1309–1321, July 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34808.

DaPrato:1991:MEP

[DPZ91] G. Da Prato, A. J. Pritchard, and J. Zabczyk. On minimum energy problems. SIAM Journal on Control and Optimization, 29(1):209–221, January 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Davis:1993:EOP

[DPZ93] Mark H. A. Davis, Vassilios G. Panas, and Thaleia Zariphopoulou. European option pricing with transaction costs. SIAM Journal on Control and Optimization, 31(2): 470–493, March 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ding:2021:SSM

[DQ21] Xiaojie Ding and Huijie Qiao. Stability for stochastic McKean–Vlasov equations with non-Lipschitz coefficients. SIAM Journal on Control and Optimization, 59 (2):887–905, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dumitrescu:2016:WDP

[DQS16] Roxana Dumitrescu, Marie-Claire Quenez, and Agnès Sulem. A weak dynamic programming principle for combined optimal stopping/ stochastic control with E^{f} - expectations. SIAM Journal on Control and Optimization, 54(4):2090–2115, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dolecki:1977:GTO

[DR77] Szymon Dolecki and David L. Russell. A general theory of observation and control. SIAM Journal on Control and Optimization, 15(2): 185–220, February 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dolecki:1979:EPL

[DR79] Szymon Dolecki and Stefan Rolewicz. Exact penalties for local minima. SIAM Journal on Control and Optimization, 17(5):596–606, September 1979. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

DiMasi:1987:ADT

[DR87] Giovanni B. Di Masi and Wolfgang J. Runggaldier. An approach to discretetime stochastic control problems under partial observation. SIAM Journal on Control and Optimization, 25(1): 38–48, January 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dippon:1997:WMS

[DR97] J. Dippon and J. Renz.

Weighted means in stochastic approximation of minima. SIAM Journal on Control and Optimization, 35(5): 1811–1827, September 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28378.

Dharmatti:2005:HCS

[DR05] Sheetal Dharmatti and Mythily Ramaswamy. Hybrid control systems and viscosity solutions. SIAM Journal on Control and Optimization, 44(4):1259–1288, July 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/61807.

Denardo:2006:TTR

[DR06] Eric V. Denardo and Uriel G.

Rothblum. A turnpike theorem for a risk-sensitive Markov decision process with stopping. SIAM Journal on Control and Optimization, 45(2): 414–431, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dentcheva:2008:SDO

[DR08] Darinka Dentcheva and Andrzej Ruszczyński. Stochastic dynamic optimization with discounted stochastic dominance constraints. SIAM Journal on Control and Optimization, 47(5):2540–2556, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dymkov:2008:COC

[DRDG08] M. Dymkov, E. Rogers, S. Dymkou, and K. Galkowski. Constrained optimal control theory for differential linear repetitive processes. SIAM Journal on Control and Optimization, 47(1):396–420, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

dePinho:2008:NCC

[dRdPR08] Maria do Rosário de Pinho and Javier F. Rosenblueth. Necessary conditions for constrained problems under Mangasarian– Fromowitz conditions. SIAM Journal on Control and Optimization, 47(1):535–552, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Drenick:1977:FCP

[Dre77] R. F. Drenick. Feedback control of partly unknown systems. SIAM Journal on Control and Optimization, 15(3): 506–519, May 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

doRocio:2009:SAG

[dRSV09] Osvaldo G. do Rocio, Alexandre J. Santana, and Marcos A. Verdi. Semigroups of affine groups, controllability of affine systems and affine bilinear systems in Sl(2, R) ⊗ R^{2}. SIAM Journal on Control and Optimization, 48(2): 1080–1088, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dharmatti:2011:FBS

[DRT11] Sheetal Dharmatti, Jean-Pierre Raymond, and Laetitia Thevenet. H^{∞} feedback boundary stabilization of the two-dimensional Navier– Stokes equations. SIAM Journal on Control and Optimization, 49(6):2318–2348, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sicon/
resource/1/sjcodc/v49/i6/
p2318_s1.

Dashkovskiy:2010:SGT

[DRW10] Sergey N. Dashkovskiy, Björn S.

Rüffer, and Fabian R. Wirth.

Small gain theorems for large scale systems and construction of ISS Lyapunov functions. SIAM Journal on Control and Optimization, 48(6): 4089–4118, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Davis:1987:PNF

[DS87] M. H. A. Davis and M. P. Spathopoulos. Pathwise nonlinear filtering for nondegenerate diffusions with noise correlation. SIAM Journal on Control and Optimization, 25(2): 260–278, March 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Davis:1989:MPC

[DS89] M. H. A. Davis and M. P. Spathopoulos. On the minimum principle for controlled diffusions on manifolds. SIAM Journal on Control and Optimization, 27(5):1092–1107, September 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dale:1993:SES

[DS93] Wilbur N. Dale and Malcolm C. Smith. Stabilizability and existence of system representations for discretetime time-varying systems. SIAM Journal on Control and Optimization, 31(6):1538– 1557, November 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

DiMasi:1999:RSC

[DS99a] G. B. Di Masi and Ł. Stettner. Risk-sensitive control of discrete-time Markov processes with infinite horizon. SIAM Journal on Control and Optimization, 38(1):61–78, January 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32061.

Donchev:1999:AMO

[DS99b] Tzanko Donchev and Iordan Slavov. Averaging method for one-sided Lipschitz differential inclusions with generalized solutions. SIAM Journal on Control and Optimization, 37(5):1600–1613, September 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32137.

Dupuis:2001:COF

[DS01] Paul Dupuis and Adam Szpiro. Convergence of the optimal feedback policies in a numerical method for a class of deterministic optimal control problems. SIAM Journal on Control and Optimization, 40(2):393–420, March 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/34496.

DiMasi:2007:IHR

[DS07] Giovanni B. Di Masi and Łukasz Stettner. Infinite horizon risk sensitive control of discrete time Markov processes under minorization property. SIAM Journal on Control and Optimization, 46 (1):231–252, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Delfour:2009:LQD

[DS09] Michel C. Delfour and Olivier Dello Sbarba. Linear quadratic differential games: Closed loop saddle points. SIAM Journal on Control and Optimization, 47(6):3138–3166, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

DeAngelis:2019:LCO

[DS19] Tiziano De Angelis and Gabriele Stabile. On Lipschitz continuous optimal stopping boundaries. SIAM Journal on Control and Optimization, 57 (1):402–436, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

daSilva:2001:RFF

[dSF01] Paulo Sérgio Pereira da Silva and Carlos Corrêa Filho. Relative flatness and flatness of implicit systems. SIAM Journal on Control and Optimization, 39(6):1929–1951, ???? 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Duong:2022:MEL

[DSI22] Nhattrieu C. Duong, Jason L.

Speyer, and Moshe Idan. Multivariate estimator for linear dynamical systems with additive Laplace measurement and process noises. SIAM Journal on Control and Optimization, 60(5):3127–3147, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1391237.

Dunn:1992:VKT

[DT92] J. C. Dunn and T. Tian.

Variants of the Kuhn–Tucker sufficient conditions in cones of nonnegative functions. SIAM Journal on Control and Optimization, 30(6):1361– 1384, November 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dong:2022:SHN

[DTM22] Hailing Dong, Juan Tang, and Xuerong Mao. Stabilization of highly nonlinear hybrid stochastic differential delay equations with Lévy noise by delay feedback control. SIAM Journal on Control and Optimization, 60(6):3302–3325, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1480392.

DeDona:2018:SRC

[DTV18] J. De Doná, N. Tehseen, and P. J. Vassiliou. Symmetry reduction, contact geometry, and partial feedback linearization. SIAM Journal on Control and Optimization, 56(1): 201–230, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dong:1991:SCT

[DTWY91] Rui-Tao Dong, Luen-Fai Tam, Wing Shing Wong, and Stephen S.-T. Yau. Structure and classification theorems of finite-dimensional exact estimation algebras. SIAM Journal on Control and Optimization, 29(4):866–877, July 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Du:2015:SCI

[Du15] Kai Du. Solvability conditions for indefinite linear quadratic optimal stochastic control problems and associated stochastic Riccati equations. SIAM Journal on Control and Optimization, 53(6): 3673–3689, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Duca:2020:GEC

[Duc20] Alessandro Duca. Global exact controllability of bilinear quantum systems on compact graphs and energetic controllability. SIAM Journal on Control and Optimization, 58 (6):3092–3129, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dunn:1979:RCC

[Dun79] J. C. Dunn. Rates of convergence for conditional gradient algorithms near singular and nonsingular extremals. SIAM Journal on Control and Optimization, 17 (2):187–211, March 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dunn:1980:CRC

[Dun80a] J. C. Dunn. Convergence rates for conditional gradient sequences generated by implicit step length rules. SIAM Journal on Control and Optimization, 18(5):473–487, September 1980. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dunn:1980:NMG

[Dun80b] J. C. Dunn. Newton’s method and the Goldstein step-length rule for constrained minimization problems. SIAM Journal on Control and Optimization, 18(6):659–674, November 1980. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Dunn:1981:GAC

[Dun81] J. C. Dunn. Global and asymptotic convergence rate estimates for a class of projected gradient processes. SIAM Journal on Control and Optimization, 19(3):368–400, May 1981. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dunn:1986:DMC

[Dun86]  J. C. Dunn.  Diagonally

modified conditional gradient methods for input constrained optimal control problems. SIAM Journal on Control and Optimization, 24(6): 1177–1191, November 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dunn:1991:SDP

[Dun91] J. C. Dunn. A subspace decomposition principle for scaled gradient projection methods: Global theory. SIAM Journal on Control and Optimization, 29(5):1160– 1175, September 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dunn:1993:SDP

[Dun93] J. C. Dunn. A subspace decomposition principle for scaled gradient projection methods: Local theory. SIAM Journal on Control and Optimization, 31(1):219–246, January 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dunn:1995:SOO

[Dun95] Joseph C. Dunn. Second-order optimality conditions in sets of L^{∞} functions with range in a polyhedron. SIAM Journal on Control and Optimization, 33(5):1603–1635, September 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24031.

Dunn:1996:SCG

[Dun96] J. C. Dunn. On L^{2} sufficient conditions and the gradient projection method for optimal control problems. SIAM Journal on Control and Optimization, 34(4):1270–1290, July 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26612.

Dunn:1998:SCE

[Dun98]  J. C. Dunn.  L^{2} suffi-

cient conditions for endconstrained optimal control problems with inputs in a polyhedron. SIAM Journal on Control and Optimization, 36(5):1833–1851, September 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/28851.

Dunn:2000:SCR

[Dun00] J. C. Dunn. On state constraint representations and mesh-dependent gradient projection convergence rates for optimal control problems. SIAM Journal on Control and Optimization, 39(4):1082–1111, July 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/35165.

Dupuis:2003:ESR

[Dup03] Paul Dupuis. Explicit solution to a robust queueing control problem. SIAM Journal on Control and Optimization, 42(5):1854–1875, September 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/39575.

Dur:1998:ODK

[Dür98] Arne Dür. On the optimality of the discrete Karhunen– Loève expansion. SIAM Journal on Control and Optimization, 36(6):1937–1939, November 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/31575.

Dus:2021:BES

[Dus21] Mathias Dus. BV exponential stability for systems of scalar conservation laws using saturated controls. SIAM Journal on Control and Optimization, 59(2):1656–1679, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dontchev:1983:SPM

[DV83] A. L. Dontchev and V. M.

Veliov. Singular perturbation in Mayer’s problem for linear systems. SIAM Journal on Control and Optimization, 21 (4):566–581, July 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Date:2004:MDB

[DV04] P. Date and G. Vinnicombe.

Measuring distance between systems under bounded power excitation. SIAM Journal on Control and Optimization, 43(3):922–936, May 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41011.

Draief:2012:CSB

[DV12] Moez Draief and Milan Vojnović. Convergence speed of binary interval consensus. SIAM Journal on Control and Optimization, 50(3): 1087–1109, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

doVal:2022:OCP

[dVC22] João B. R. do Val and Daniel S. Campos. The H_{∈}- optimal control problem of CSVIU systems: Discounted, counterdiscounted, and longrun solutions. SIAM Journal on Control and Optimization, 60(4):2320–2343, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1434593.

Dalsmo:1998:RIM

[DvdS98] Morten Dalsmo and Arjan van der Schaft. On representations and integrability of mathematical structures in energy-conserving physical systems. SIAM Journal on Control and Optimization, 37(1):54–91, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31203.

Diab:2024:VML

[DVPL24] Ali Diab, Giorgio Valmorbida, and William Pasillas-Lépine. Verification methods for the Lyapunov–Krasovskii functional inequalities. SIAM Journal on Control and Optimization, 62(2):877–902, March 2024. CODEN SJ-CODC. ISSN 1095-7138.

deWaal:2000:CTP

[dWvS00] Peter R. de Waal and Jan H.

van Schuppen. A class of team problems with discrete action spaces: Optimality conditions based on multimodularity. SIAM Journal on Control and Optimization, 38(3):875–892, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30145.

Dong:2016:SCM

[DWZ^{+}16] Wen-Bin Dong, Re-Bing Wu, Wei Zhang, Chun-Wen Li, and Tzyh-Jong Tarn. Spatial control model and analysis of quantum fields in one-dimensional waveguides. SIAM Journal on Control and Optimization, 54(3):1352– 1377, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Duan:2020:OIH

[DWZ20] Yueliang Duan, Lijuan Wang, and Can Zhang. Observability inequalities for the heat equation with bounded potentials on the whole space. SIAM Journal on Control and Optimization, 58(4):1939–1960, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dai:2015:DMP

[DXG15] Li Dai, Yuanqing Xia, and Yulong Gao. Distributed model predictive control of linear systems with stochastic parametric uncertainties and coupled probabilistic constraints. SIAM Journal on Control and Optimization, 53 (6):3411–3431, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dayanik:2013:RRM

[DY13] Savas Dayanik and Angela J. Yu. Reward-rate maximization in sequential identification under a stochastic deadline. SIAM Journal on Control and Optimization, 51(4): 2922–2948, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dye:1977:NAJ

[Dye77] B. R. Dye. A note on an algorithm by J. Rissanen. SIAM Journal on Control and Optimization, 15(1):73–78, January 1977. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Dai:2012:OSS

[DYZ12] Min Dai, Zhou Yang, and Yifei Zhong. Optimal stock selling based on the global maximum. SIAM Journal on Control and Optimization, 50(4): 1804–1822, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

DAmato:1976:BVC

[DZ76] L. P. D’Amato and A. H. Zemanian. A boundary value criterion for semipassive Hilbert ports. SIAM Journal on Control and Optimization, 14(4): 763–768, July 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Delfour:1988:SSA

[DZ88] M. C. Delfour and J.-P.

Zolésio. Shape sensitivity analysis via min max differentiability. SIAM Journal on Control and Optimization, 26 (4):834–862, July 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Delfour:1991:VML

[DZ91] Michel C. Delfour and Jean-Paul Zolésio. Velocity method and Lagrangian formulation for the computation of the shape Hessian. SIAM Journal on Control and Optimization, 29(6):1414–1442, November 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Desilles:2019:PFC

[DZ19] Anna Désilles and Hasnaa Zidani. Pareto front characterization for multiobjective optimal control problems using Hamilton–Jacobi approach. SIAM Journal on Control and Optimization, 57 (6):3884–3910, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Deng:2024:EOP

[DZ24] Li Deng and Xu Zhang. Existence of optimal pairs for optimal control problems with states constrained to Riemannian manifolds. SIAM Journal on Control and Optimization, 62(4):2098–2114, August 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1584095.

Dong:2019:RTL

[DZA19] Z. Y. Dong, G. F. Zhang, and N. H. Amini. On the response of a two-level system to twophoton inputs. SIAM Journal on Control and Optimization, 57(5):3445–3470, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Du:2024:UCC

[DZX24] Linglong Du, Jianwen Zhu, and Feng Xie. The unconditional consensus control through leadership for the delayed Hegselmann–Krause model. SIAM Journal on Control and Optimization, 62 (4):2297–2318, August 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1588858.

Efimov:2021:LKF

[EA21] Denis Efimov and Alexander Y. Aleksandrov. Lyapunov– Krasovskii functional for discretized homogeneous systems. SIAM Journal on Control and Optimization, 59(4): 2546–2569, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

El-Ansary:1986:ISP

[EAK86] Mohamed El-Ansary and Hassan Khalil. On the interplay of singular perturbations and wide-band stochastic fluctuations. SIAM Journal on Control and Optimization, 24(1): 83–94, January 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Evgrafov:2020:NCC

[EB20] Anton Evgrafov and José C.

Bellido. Nonlocal control in the conduction coefficients: Well-posedness and convergence to the local limit. SIAM Journal on Control and Optimization, 58(4):1769–1794, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Evgrafov:2024:NKP

[EB24] Anton Evgrafov and JoséC.

Bellido. The nonlocal Kelvin Principle and the dual approach to nonlocal control in the conduction coefficients. SIAM Journal on Control and Optimization, 62(1):487–508, February 2024. CODEN SJ-CODC. ISSN 1095-7138.

Eising:2022:GAC

[EC22] Jaap Eising and M. Kanat Camlibel. A geometric approach to convex processes: From reachability to stabilizability. SIAM Journal on Control and Optimization, 60 (4):2574–2600, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1426183.

Elliott:1981:OPS

[ED81] R. J. Elliott and M. H. A. Davis. Optimal play in a stochastic differential game. SIAM Journal on Control and Optimization, 19(4):543–554, July 1981. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Elliott:2005:SME

[EDM05] Robert J. Elliott, François Dufour, and W. P. Malcolm. State and mode estimation for discrete-time jump Markov systems. SIAM Journal on Control and Optimization, 44(3):1081–1104, May 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/44262.

Efimov:2009:ONS

[EF09] Denis V. Efimov and Alexander L. Fradkov. Oscillatority of nonlinear systems with static feedback. SIAM Journal on Control and Optimization, 48(2):618–640, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Efimov:2012:GLA

[Efi12] Denis Efimov. Global Lyapunov analysis of multistable nonlinear systems. SIAM Journal on Control and Optimization, 50(5):3132–3154, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Eremenko:2002:PPS

[EG02] A. Eremenko and A. Gabrielov.

Pole placement by static output feedback for generic linear systems. SIAM Journal on Control and Optimization, 41(1):303–312, January 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39191.

Egardt:1980:SAC

[Ega80] Bo Egardt. Stability analysis of continuous-time adaptive control systems. SIAM Journal on Control and Optimization, 18(5):540–558, September 1980. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Egami:2008:DSM

[Ega08] Masahiko Egami. A direct solution method for stochastic impulse control problems of one-dimensional diffusions. SIAM Journal on Control and Optimization, 47(3):1191– 1218, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ermoliev:1985:SOP

[EGN85] Yu. Ermoliev, A. Gaivoronski, and C. Nedeva. Stochastic optimization problems with incomplete information on distribution functions. SIAM Journal on Control and Optimization, 23(5):697– 716, September 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Edholm:2022:SAF

[EGR^{+}22] C. J. Edholm, C. Guiver, R. Rebarber, B. Tenhumberg, and S. Townley. Stabilization by adaptive feedback control for positive difference equations with applications in pest management. SIAM Journal on Control and Optimization, 60(4):2214–2245, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1398240.

Emre:1980:PCI

[EH80] E. Emre and M. L. J. Hautus. A polynomial characterization of (A, B)-Invariant and reachability subspaces. SIAM Journal on Control and Optimization, 18(4):420–436, ???? 1980. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Eppler:2010:TND

[EH10] Karsten Eppler and Helmut Harbrecht. Tracking Neumann data for stationary free boundary problems. SIAM Journal on Control and Optimization, 48(5):2901–2916, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Eppler:2007:CES

[EHS07] Karsten Eppler, Helmut Harbrecht, and Reinhold Schneider. On convergence in elliptic shape optimization. SIAM Journal on Control and Optimization, 46(1):61–83, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ekanayake:2013:PDC

[EI13] Dinesh B. Ekanayake and Ram V. Iyer. Proportional derivative control of hysteretic systems. SIAM Journal on Control and Optimization, 51 (5):3415–3433, ???? 2013.

CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Evans:1989:HJB

[EJ89] L. C. Evans and M. R. James. The Hamilton–Jacobi– Bellman equation for timeoptimal control. SIAM Journal on Control and Optimization, 27(6):1477–1489, November 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ecker:1985:CCE

[EK85] J. G. Ecker and M. Kupferschmid. A computational comparison of the ellipsoid algorithm with several nonlinear programming algorithms. SIAM Journal on Control and Optimization, 23(5):657– 674, September 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ethier:1993:FVP

[EK93] S. N. Ethier and Thomas G. Kurtz. Fleming–Viot processes in population genetics. SIAM Journal on Control and Optimization, 31(2): 345–386, March 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Elliott:1997:EFD

[EK97] Robert J. Elliott and Vikram Krishnamurthy. Exact finitedimensional filters for maximum likelihood parameter estimation of continuoustime linear Gaussian systems. SIAM Journal on Control and Optimization, 35(6): 1908–1923, November 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29255.

Englezos:2009:UMH

[EK09] Nikolaos Englezos and Ioannis Karatzas. Utility maximization with habit formation: Dynamic programming and stochastic PDEs. SIAM Journal on Control and Optimization, 48(2):481–520, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ekeland:1977:LDN

[Eke77] Ivar Ekeland. Legendre duality in nonconvex optimization and calculus of variations. SIAM Journal on Control and Optimization, 15(6): 905–934, November 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

ElFarouq:2017:DFO

[El 17] Naïma El Farouq. Degenerate first-order quasi-variational inequalities: an approach to approximate the value function. SIAM Journal on Control and Optimization, 55(4): 2714–2733, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Emamizadeh:2020:BBM

[EL20] Behrouz Emamizadeh and Yichen Liu. Bang-bang and multiple valued optimal solutions of control problems related to quasi-linear elliptic equations. SIAM Journal on Control and Optimization, 58(2):1103–1117, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Elliott:1976:EVS

[Ell76] Robert Elliott. The existence of value in stochastic differential games. SIAM Journal on Control and Optimization, 14 (1):85–94, January 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Elliott:1977:OCS

[Ell77] Robert J. Elliott. The optimal control of a stochastic system. SIAM Journal on Control and Optimization, 15(5): 756–778, August 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ekstrom:2022:HDS

[ELO22] Erik Ekström, Kristoffer Lindensjö, and Marcus Olofsson. How to detect a salami slicer: a stochastic controllerand-stopper game with unknown competition. SIAM Journal on Control and Optimization, 60(1):545–574, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M139044X.

Edalatzadeh:2019:OAD

[EM19] M. Sajjad Edalatzadeh and Kirsten A. Morris. Optimal actuator design for semilinear systems. SIAM Journal on Control and Optimization, 57(4):2992–3020, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ebihara:2008:GPI

[EMH08] Yoshio Ebihara, Katsutoshi Maeda, and Tomomichi Hagiwara. Generalized S-procedure for inequality conditions on one-vector-lossless sets and linear system analysis. SIAM Journal on Control and Optimization, 47(3):1547–1555, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ekstrom:2023:FPU

[EMO23] Erik Ekström, Alessandro Milazzo, and Marcus Olofsson. The de Finetti problem with uncertain competition. SIAM Journal on Control and Optimization, 61(5):2997–3017, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1524849.

Espinosa:2016:PRP

[EMP^{+}16] Luis A. Duffaut Espinosa, Z. Miao, I. R. Petersen, V. Ugrinovskii, and M. R. James. Physical realizability and preservation of commutation and anticommutation relations for n-level quantum systems. SIAM Journal on Control and Optimization, 54 (2):632–661, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Elie:2018:CPD

[EMP18] Romuald Elie, Ludovic Moreau, and Dylan Possamaï. On a class of path-dependent singular stochastic control problems. SIAM Journal on Control and Optimization, 56(5): 3260–3295, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ernst:2024:QRT

[EMP24] P. A. Ernst, H. Mei, and G. Peskir. Quickest real-time detection of multiple Brownian drifts. SIAM Journal on Control and Optimization, 62(3):1832–1856, June 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1587750.

Emre:1980:NFP

[Emr80a] E. Emre. Nonsingular factors of polynomial matrices and (A, B)-invariant subspaces. SIAM Journal on Control and Optimization, 18(3):288–296, May 1980. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Emre:1980:PEA

[Emr80b] E. Emre. The polynomial equation QQ_{c} + RP_{c} = Φ with application to dynamic feedback. SIAM Journal on Control and Optimization, 18(6):611–620, November 1980. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Emre:1982:NSC

[Emr82] E. Emre. On necessary and sufficient conditions for regulation of linear systems over rings. SIAM Journal on Control and Optimization, 20(2): 155–160, March 1982. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Engwerda:2010:NSC

[Eng10] Jacob Engwerda. Necessary and sufficient conditions for Pareto optimal solutions of cooperative differential games. SIAM Journal on Control and Optimization, 48(6):3859– 3881, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

ElKaroui:1988:EOM

[ENJP88] Nicole El Karoui, Du′ Hu̇ù Nguyen, and Monique Jeanblanc-Picqué. Existence of an optimal Markovian filter for the control under partial observations. SIAM Journal on Control and Optimization, 26(5): 1025–1061, September 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Enos:1994:CST

[Eno94a] Michael J. Enos. Controllability of a system of two symmetric rigid bodies in three space. SIAM Journal on Control and Optimization, 32(4):1170–1185, July 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/19557.

Enos:1994:OAV

[Eno94b]  Michael J. Enos.  Opti-

mal angular velocity tracking with fixed-endpoint rigid body motions. SIAM Journal on Control and Optimization, 32(4):1186–1193, July 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/19545.

Enqvist:2004:COA

[Enq04]  P. Enqvist.  A convex

optimization approach to ARMA(n, m) model design from covariance and cepstral data. SIAM Journal on Control and Optimization, 43(3):1011–1036, May 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39975.

Ermoliev:1995:MSF

[ENW95] Yuri M. Ermoliev, Vladimir I.

Norkin, and Roger J-B. Wets. The minimization of semicontinuous functions: Mollifier subgradients. SIAM Journal on Control and Optimization, 33(1):149–167, January 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23836.

Eldem:1988:DDP

[EÖ88] Vasfi Eldem and A. Bülent Özgüler. Disturbance decoupling problems by measurement feedback: a characterization of all solutions and fixed modes. SIAM Journal on Control and Optimization, 26(1): 168–185, January 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ergenc:2007:EKS

[EOF07] Ali Fuat Ergenc, Nejat Olgac, and Hassan Fazelinia. Extended Kronecker summation for cluster treatment of LTI systems with multiple delays. SIAM Journal on Control and Optimization, 46(1):143–155, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Eldem:1998:PDR

[EÖSÖ98] Vasfi Eldem, Hitay Özbay, Hasan Selbuz, and Kadri Özçaldiran. Partial disturbance rejection with internal stability and H_{∞} norm bound. SIAM Journal on Control and Optimization, 36 (1):180–192, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/28765.

Espana:1993:GDA

[EP93] Martín D. España and Laurent Praly. On the global dynamics of adaptive systems: a study of an elementary example. SIAM Journal on Control and Optimization, 31(5):1143– 1166, September 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ekstrom:2008:OSG

[EP08] Erik Ekström and Goran Peskir. Optimal stopping games for Markov processes. SIAM Journal on Control and Optimization, 47(2):684–702, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Elie:2019:CTC

[EP19] Romuald Elie and Dylan Possamaï. Contracting theory with competitive interacting agents. SIAM Journal on Control and Optimization, 57(2): 1157–1188, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Efimov:2024:CLT

[EP24] Denis Efimov and Andrey Polyakov. On converse Lyapunov theorem for fixed-time input-to-state stability. SIAM Journal on Control and Optimization, 62(1):118–134, January 2024. CODEN SJCODC. ISSN 1095-7138.

Efimov:2014:DHC

[EPR14] D. Efimov, W. Perruquetti, and J.-P. Richard. Development of homogeneity concept for time-delay systems. SIAM Journal on Control and Optimization, 52(3):1547–1566, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

ElKaroui:1995:DPP

[EQ95] Nicole El Karoui and Marie-Claire Quenez. Dynamic programming and pricing of contingent claims in an incomplete market. SIAM Journal on Control and Optimization, 33(1):29–66, January 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/23257.

Eaves:1988:IPC

[ER88] B. Curtis Eaves and Uriel G. Rothblum. Invariant polynomial curves of piecewise linear maps. SIAM Journal on Control and Optimization, 26 (2):401–416, March 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ecker:1981:SSS

[ES81a] J. G. Ecker and Nancy E. Shoemaker. Selecting subsets from the set of nondominated vectors in multiple objective linear programming. SIAM Journal on Control and Optimization, 19(4):505–515, July 1981. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Emre:1981:ECS

[ES81b] E. Emre and L. M. Silverman. The equation XR + QY = Φ: a characterization of solutions. SIAM Journal on Control and Optimization, 19(1): 33–38, January 1981. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Eckstein:2009:GPS

[ES09] Jonathan Eckstein and B. F. Svaiter. General projective splitting methods for sums of maximal monotone operators. SIAM Journal on Control and Optimization, 48(2):787–811, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See note [Bau09].

Espinosa:2012:DMS

[ET12] Gilles-Edouard Espinosa and Nizar Touzi. Detecting the maximum of a scalar diffusion with negative drift. SIAM Journal on Control and Optimization, 50(5):2543–2572, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Eller:2015:SEC

[ET15] Matthias Eller and Daniel Toundykov. Semiglobal exact controllability of nonlinear plates. SIAM Journal on Control and Optimization, 53(4): 2480–2513, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Etesami:2019:SFS

[Ete19] S. Rasoul Etesami. A simple framework for stability analysis of state-dependent networks of heterogeneous agents. SIAM Journal on Control and Optimization, 57(3):1757– 1782, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Etesami:2020:DSC

[Ete20] S. Rasoul Etesami. Duality and stability in complex multiagent state-dependent network dynamics. SIAM Journal on Control and Optimization, 58(6):3062–3091, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Etesami:2024:LSN

[Ete24] S. Rasoul Etesami. Learning stationary Nash equilibrium policies in n-player stochastic games with independent chains. SIAM Journal on Control and Optimization, 62(2): 799–825, March 2024. CO-DEN SJCODC. ISSN 1095-7138.

Efimov:2018:HLF

[EUMP18] D. Efimov, R. Ushirobira, J. A. Moreno, and W. Perruquetti. Homogeneous Lyapunov functions: From converse design to numerical implementation. SIAM Journal on Control and Optimization, 56(5):3454–3477, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Evans:2013:CNI

[Eva13] Neil D. Evans. Controlling nonlinear infinite-dimensional systems via the initial state. SIAM Journal on Control and Optimization, 51(5):3757– 3780, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

ElAlaoui:2011:SCI

[EZB11] Fatima-Zahrae El Alaoui, Hans Zwart, and Ali Boutoulout. Spectral conditions implied by observability. SIAM Journal on Control and Optimization, 49(2):672–685, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sicon/
resource/1/sjcodc/v49/i2/
p672_s1.

Fletcher:1989:DDD

[FA89] L. R. Fletcher and A. Aasaraai.

On disturbance decoupling in descriptor systems. SIAM Journal on Control and Optimization, 27(6):1319–1332, November 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fridman:2013:SDD

[FA13] Emilia Fridman and Netzer Bar Am. Sampled-data distributed H_{∞} control of transport reaction systems. SIAM Journal on Control and Optimization, 51(2):1500– 1527, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fabre:1992:EBC

[Fab92] Caroline Fabre. Exact boundary controllability of the wave equation as the limit of internal controllability.

SIAM Journal on Control and Optimization, 30(5):1066– 1086, September 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fabbri:2008:VSA

[Fab08] Giorgio Fabbri. A viscosity solution approach to the infinitedimensional HJB equation related to a boundary control problem in a transport equation. SIAM Journal on Control and Optimization, 47(2): 1022–1052, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Faggian:2008:HJE

[Fag08] Silvia Faggian. Hamilton– Jacobi equations arising from boundary control problems with state constraints. SIAM Journal on Control and Optimization, 47(4):2157–2178, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fardigola:2008:CPS

[Far08] L. V. Fardigola. Controllability problems for the string equation on a half-axis with a boundary control bounded by a hard constant. SIAM Journal on Control and Optimization, 47(4):2179–2199, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fardigola:2013:TOS

[Far13] Larissa V. Fardigola. Transformation operators of the Sturm–Liouville problem in controllability problems for the wave equation on a halfaxis. SIAM Journal on Control and Optimization, 51(2): 1781–1801, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fattorini:1990:CSC

[Fat90] H. O. Fattorini. Convergence of suboptimal controls: The point target case. SIAM Journal on Control and Optimization, 28(2):320–341, March 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fattorini:1994:ETM

[Fat94] H. O. Fattorini. Existence theory and the maximum principle for relaxed infinitedimensional optimal control problems. SIAM Journal on Control and Optimization, 32(2):311–331, March 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/22024.

Fattorini:1997:RES

[Fat97]  H. O. Fattorini.  A re-

mark on existence of solutions of infinite-dimensional noncompact optimal control problems. SIAM Journal on Control and Optimization, 35(4):1422–1433, July 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28788.

Fernandez-Berdaguer:1998:PEA

[FB98] Elena M. Fernández-Berdaguer. Parameter estimation in acoustic media using the adjoint method. SIAM Journal on Control and Optimization, 36(4):1315–1330, July 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29932.

Fragoso:2001:OCC

[FB01] Marcelo D. Fragoso and Jack Baczynski. Optimal control for continuous-time linear quadratic problems with infinite Markov jump parameters. SIAM Journal on Control and Optimization, 40 (1):270–297, January 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36748.

Fardad:2008:PMS

[FB08] Makan Fardad and Bassam Bamieh. Perturbation methods in stability and norm analysis of spatially periodic systems. SIAM Journal on Control and Optimization, 47(2): 997–1021, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

French:2012:BAN

[FB12] Mark French and Wenming Bian. A biased approach to nonlinear robust stability and performance with applications to adaptive control. SIAM Journal on Control and Optimization, 50(3): 1220–1243, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Flores-Bazan:2009:SES

[FBJ09] Fabián Flores-Bazán and Bienvenido Jiménez. Strict efficiency in set-valued optimization. SIAM Journal on Control and Optimization, 48(2): 881–908, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Flores-Bazan:2014:CVF

[FBJM14] F. Flores-Bazán, A. Jourani, and G. Mastroeni. On the convexity of the value function for a class of nonconvex variational problems: Existence and optimality conditions. SIAM Journal on Control and Optimization, 52(6): 3673–3693, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Flores-Bazan:1999:NVP

[FBP99] Fabián Flores-Bazán and Stefania Perrotta. Nonconvex variational problems related to a hyperbolic equation. SIAM Journal on Control and Optimization, 37(6): 1751–1766, November 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/33229.

Fragoso:2005:UAS

[FC05] Marcelo D. Fragoso and Oswaldo L. V. Costa. A unified approach for stochastic and mean square stability of continuous-time linear systems with Markovian jumping parameters and additive disturbances. SIAM Journal on Control and Optimization, 44(4):1165–1191, July 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/43475.

Feng:2019:OCN

[FC19] Wenhui Feng and Han-Fu Chen. Output consensus of networked Hammerstein and Wiener systems. SIAM Journal on Control and Optimization, 57(2):1230–1254, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Feng:2020:DGC

[FCC20] Yu Feng, Zhuoming Chen, and Xiang Chen. Distributed H_{∞} Gaussian consensus filtering for discrete-time systems over lossy sensor networks. SIAM Journal on Control and Optimization, 58(1):34–58, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fernandez-Cara:2006:GCI

[FCG06] Enrique Fernández-Cara and Sergio Guerrero. Global Carleman inequalities for parabolic systems and applications to controllability. SIAM Journal on Control and Optimization, 45(4):1395–1446, January 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Feng:2016:OBS

[FCG16] Yu Feng, Xiang Chen, and Guoxiang Gu. Observerbased stabilizing controllers for discrete-time systems with quantized signal and multiplicative random noise. SIAM Journal on Control and Optimization, 54(1):251–265, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Feng:2018:MSC

[FCG18] Yu Feng, Xiang Chen, and Guoxiang Gu. Multiobjective H_{∈}/H_{∞} control design subject to multiplicative input dependent noises. SIAM Journal on Control and Optimization, 56(1):253–271, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fernandez-Cara:2006:SCR

[FCGIP06] Enrique Fernández-Cara, Sergio Guerrero, Oleg Yu. Imanuvilov, and Jean-Pierre Puel. Some controllability results forthe N-dimensional Navier–Stokes and Boussinesq systems with N-1 scalar controls. SIAM Journal on Control and Optimization, 45(1): 146–173, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fernandez-Cara:2005:CIO

[FCGO05] Enrique Fernández-Cara, Galina C. Garcia, and Axel Osses. Controls insensitizing the observation of a quasi-geostrophic ocean model. SIAM Journal on Control and Optimization, 43(5):1616–1639, September 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43360.

Feyzmahdavian:2014:ASD

[FCJ14] Hamid Reza Feyzmahdavian, Themistoklis Charalambous, and Mikael Johansson. Asymptotic stability and decay rates of homogeneous positive systems with bounded and unbounded delays. SIAM Journal on Control and Optimization, 52(4):2623–2650, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fernandez-Cara:2016:NCL

[FCLZ16] Enrique Fernández-Cara, Qi Lü, and Enrique Zuazua. Null controllability of linear heat and wave equations with nonlocal spatial terms. SIAM Journal on Control and Optimization, 54(4):2009–2019, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fang:2009:SLS

[FCMA09] J. Fang, M. Cao, A. S. Morse, and B. D. O. Anderson. Sequential localization of sensor networks. SIAM Journal on Control and Optimization, 48 (1):321–350, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Francis:1987:LCT

[FD87] Bruce A. Francis and John C.

Doyle. Linear control theory with an H_{∞} optimality criterion. SIAM Journal on Control and Optimization, 25 (4):815–844, July 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fonteneau:2013:MMG

[FEBL13] R. Fonteneau, D. Ernst, B. Boigelot, and Q. Louveaux.

Min max generalization for deterministic batch mode reinforcement learning: Relaxation schemes. SIAM Journal on Control and Optimization, 51(5):3355–3385, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Feintuch:1976:SMI

[Fei76] Avraham Feintuch. Spectral minimality for infinitedimensional linear systems. SIAM Journal on Control and Optimization, 14(5):945– 950, August 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Feintuch:1980:SCC

[Fei80] Avraham Feintuch. Strong causality conditions and causal invertibility. SIAM Journal on Control and Optimization, 18 (3):317–324, May 1980. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Felgenhauer:2003:SBB

[Fel03] Ursula Felgenhauer. On stability of bang–bang type controls. SIAM Journal on Control and Optimization, 41(6): 1843–1867, November 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39927.

Feng:2016:SOD

[Fen16] Hongyinping Feng. Stabilization of one-dimensional wave equation with van der Pol type boundary condition. SIAM Journal on Control and Optimization, 54(5):2436–2449, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ferro:1978:FDF

[Fer78]  Francesco Ferro.  Func-

tionals defined on functions of bounded variation in R^{n} and the Lebesgue area. SIAM Journal on Control and Optimization, 16(5):778– 789, September 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ferro:1981:LSO

[Fer81] Francesco Ferro. Lower semicontinuity, optimization and regularizing extensions of integral functionals. SIAM Journal on Control and Optimization, 19(4):433–444, July 1981. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Feron:1997:ARP

[Fer97a] Eric Feron. Analysis of robust H_{2} performance using multiplier theory. SIAM Journal on Control and Optimization, 35(1):160–177, January 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/26650.

Ferrante:1997:HCM

[Fer97b] Augusto Ferrante. A homeomorphic characterization of minimal spectral factors. SIAM Journal on Control and Optimization, 35(5):1508– 1523, September 1997. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ferriero:2005:AHO

[Fer05] Alessandro Ferriero. The approximation of higher-order integrals of the calculus of variations and the Lavrentiev phenomenon. SIAM Journal on Control and Optimization, 44(1):99–110, January 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/43772.

Ferrari:2018:OMP

[Fer18] Giorgio Ferrari. On the optimal management of public debt: a singular stochastic control problem. SIAM Journal on Control and Optimization, 56(3):2036–2073, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fan:2022:SFS

[FF22] Xiliang Fan and Zhenzhen Feng. Stability of fractional SDEs with Markov switching perturbed by transition rate matrices. SIAM Journal on Control and Optimization, 60(5):2835–2858, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1418897.

Ferrari:2009:GOA

[FFP^{+}09] Silvia Ferrari, Rafael Fierro, Brent Perteet, Chenghui Cai, and Kelli Baumgartner. A geometric optimization approach to detecting and intercepting dynamic targets using a mobile sensor network. SIAM Journal on Control and Optimization, 48(1):292–320, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fagnani:2014:DCE

[FFR14] Fabio Fagnani, Sophie M. Fosson, and Chiara Ravazzi. A distributed Classification/Estimation algorithm for sensor networks. SIAM Journal on Control and Optimization, 52(1):189–218, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Federico:2023:TSS

[FFR23] Salvatore Federico, Giorgio Ferrari, and Neofytos Rodosthenous. Two-sided singular control of an inventory with unknown demand trend. SIAM Journal on Control and Optimization, 61 (5):3076–3101, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1442115.

Federico:2021:CID

[FFRR21] Salvatore Federico, Giorgio Ferrari, Frank Riedel, and Michael Röckner. On a class of infinite-dimensional singular stochastic control problems. SIAM Journal on Control and Optimization, 59(2): 1680–1704, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Federico:2020:SSC

[FFS20] Salvatore Federico, Giorgio Ferrari, and Patrick Schuhmann. A singular stochastic control problem with interconnected dynamics. SIAM Journal on Control and Optimization, 58(5):2821–2853, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fabbri:2009:LPH

[FG09] G. Fabbri and B. Goldys. An LQ problem for the heat equation on the halfline with Dirichlet boundary control and noise. SIAM Journal on Control and Optimization, 48(3):1473–1488, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Feng:2015:SEB

[FG15] Hongyinping Feng and Bao-Zhu Guo. On stability equivalence between dynamic output feedback and static output feedback for a class of second order infinite-dimensional systems. SIAM Journal on Control and Optimization, 53(4): 1934–1955, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fernandez-Cara:1995:RCC

[FGB95] E. Fernández-Cara and M. González-Burgos. A result concerning controllability for the Navier– Stokes equations. SIAM Journal on Control and Optimization, 33(4):1061–1070, July 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25381.

Federico:2010:HEO

[FGG10] Salvatore Federico, Ben Goldys, and Fausto Gozzi. HJB equations for the optimal control of differential equations with delays and state constraints, I: Regularity of viscosity solutions. SIAM Journal on Control and Optimization, 48(8): 4910–4937, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Federico:2011:HEO

[FGG11] Salvatore Federico, Ben Goldys, and Fausto Gozzi. HJB equations for the optimal control of differential equations with delays and state constraints, II: Verification and optimal feedbacks. SIAM Journal on Control and Optimization, 49(6): 2378–2414, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicon/resource/
                                        1/sjcodc/v49/i6/p2378_s1.

Frasca:2015:OOD

[FGGH15] Paolo Frasca, Federica Garin, Balázs Gerencsér, and Julien M. Hendrickx. Optimal onedimensional coverage by unreliable sensors. SIAM Journal on Control and Optimization, 53(5):3120–3140, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fursikov:1998:BVP

[FGH98] A. V. Fursikov, M. D. Gunzburger, and L. S. Hou. Boundary value problems and optimal boundary control for the Navier–Stokes system: The two-dimensional case. SIAM Journal on Control and Optimization, 36(3):852–894, May 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/27337.

Fursikov:2005:OBC

[FGH05] A. V. Fursikov, M. D. Gunzburger, and L. S. Hou. Optimal boundary control for the evolutionary Navier– Stokes system: The threedimensional case. SIAM Journal on Control and Optimization, 43(6):2191–2232, November 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40080.

Friedrich:2023:LSN

[FGH23] Jan Friedrich, Simone Göttlich, and Michael Herty. Lyapunov stabilization for nonlocal traffic flow models. SIAM Journal on Control and Optimization, 61(5):2849–2875, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M152181X.

Frick:2019:LCM

[FGJ^{+}19] Damian Frick, Angelos Georghiou, Juan L. Jerez, Alexander Domahidi, and Manfred Morari. Low-complexity method for hybrid MPC with local guarantees. SIAM Journal on Control and Optimization, 57(4):2328–2361, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Finlay:2008:DLP

[FGL08] Luke Finlay, Vladimir Gaitsgory, and Ivan Lebedev. Duality in linear programming problems related to deterministic long run average problems of optimal control. SIAM Journal on Control and Optimization, 47(4):1667–1700, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fang:1997:AIS

[FGQ97] Haitao Fang, Guanglu Gong, and Minping Qian. Annealing of iterative stochastic schemes. SIAM Journal on Control and Optimization, 35(6):1886–1907, November 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/29367.

Foias:1993:RSF

[FGS93] Ciprian Foiaş, Tryphon T. Georgiou, and Malcolm C. Smith. Robust stability of feedback systems: a geometric approach using the gap metric. SIAM Journal on Control and Optimization, 31(6):1518– 1537, November 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Foias:1995:NOC

[FGT95] Ciprian Foiaş, Caixing Xing Gu, and Allen Tannenbaum.

Nonlinear H^{∞} optimization: a causal power series approach. SIAM Journal on Control and Optimization, 33 (1):185–207, January 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/23616.

Friedman:1988:CFB

[FH88] Avner Friedman and Karl-Heinz Hoffmann. Control of free boundary problems with hysteresis. SIAM Journal on Control and Optimization, 26 (1):42–55, January 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ferreyra:1994:SFB

[FH94] Guillermo Ferreyra and Omar Hijab. A simple free boundary problem in R^{d}. SIAM Journal on Control and Optimization, 32(2):501–515, March 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23221.

Fouque:2017:AOS

[FH17a] Jean-Pierre Fouque and Ruimeng Hu. Asymptotic optimal strategy for portfolio optimization in a slowly varying stochastic environment. SIAM Journal on Control and Optimization, 55(3):1990–2023, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fu:2017:MFG

[FH17b] Guanxing Fu and Ulrich Horst. Mean field games with singular controls. SIAM Journal on Control and Optimization, 55(6):3833–3868, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fu:2020:MFL

[FH20] Guanxing Fu and Ulrich Horst. Mean-field leaderfollower games with terminal state constraint. SIAM Journal on Control and Optimization, 58(4):2078–2113, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fleming:1997:RSC

[FHH97] W. H. Fleming and D. Hernández-Hernández. Risk-sensitive control of finite state machines on an infinite horizon I. SIAM Journal on Control and Optimization, 35(5): 1790–1810, September 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/29162.

Fleming:1999:RSC

[FHH99] Wendell H. Fleming and Daniel Hernández-Hernández. Risk-sensitive control of finite state machines on an infinite horizon II. SIAM Journal on Control and Optimization, 37(4):1048–1069, July 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32149.

Feng:2022:BSD

[FHH22] Xinwei Feng, Ying Hu, and Jianhui Huang. Backward Stackelberg differential game with constraints: a mixed terminal-perturbation and linear-quadratic approach. SIAM Journal on Control and Optimization, 60(3):1488– 1518, ???? 2022. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1340769.

Fujiwara:1984:LDN

[FHM84] O. Fujiwara, S.-P. Han, and O. L. Mangasarian. Local duality of nonlinear programs. SIAM Journal on Control and Optimization, 22(1): 162–169, January 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fei:2018:SRS

[FHMS18] Weiyin Fei, Liangjian Hu, Xuerong Mao, and Mingxuan Shen. Structured robust stability and boundedness of nonlinear hybrid delay systems. SIAM Journal on Control and Optimization, 56(4): 2662–2689, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fuhrman:2006:CSO

[FHT06] Marco Fuhrman, Ying Hu, and Gianmario Tessitore. On a class of stochastic optimal control problems related to BS-DEs with quadratic growth. SIAM Journal on Control and Optimization, 45(4):1279– 1296, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fuhrman:2009:EBO

[FHT09] Marco Fuhrman, Ying Hu, and Gianmario Tessitore. Ergodic BSDEs and optimal ergodic control in Banach spaces. SIAM Journal on Control and Optimization, 48(3): 1542–1566, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Friedman:1988:OPC

[FHY88] Avner Friedman, Shao Yun Huang, and Jiong Min Yong. Optimal periodic control for the two-phase Stefan problem. SIAM Journal on Control and Optimization, 26(1): 23–41, January 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fursikov:1998:LEB

[FI98] A. V. Fursikov and O. Yu.

Imanuvilov. Local exact boundary controllability of the Boussinesq equation. SIAM Journal on Control and Optimization, 36(2):391–421, March 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29679.

French:2009:RSL

[FIM09] Mark French, Achim Ilchmann, and Markus Mueller. Robust stabilization by linear output delay feedback. SIAM Journal on Control and Optimization, 48(4):2533–2561, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fiori:2020:CTA

[Fio20] Simone Fiori. A controltheoretic approach to the synchronization of secondorder continuous-time dynamical systems on real connected Riemannian manifolds. SIAM Journal on Control and Optimization, 58(2):787–813, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fiori:2021:EPR

[Fio21] Simone Fiori. Extension of PID regulators to dynamical systems on smooth manifolds (M-PID). SIAM Journal on Control and Optimization, 59 (1):78–102, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

French:2006:RGT

[FIR06] Mark French, Achim Ilchmann, and Eugene P. Ryan. Robustness in the graph topology of a common adaptive controller. SIAM Journal on Control and Optimization, 45(5): 1736–1757, ???? 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fischer:2011:IPC

[Fis11] Markus Fischer. An iterative procedure for constructing subsolutions of discretetime optimal control problems. SIAM Journal on Control and Optimization, 49(6): 2608–2626, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i6/p2608_s1.

Fitzgerald:1979:PMD

[Fit79]  Carl H. Fitzgerald.  The

princess and monster differential game. SIAM Journal on Control and Optimization, 17(6):700–712, November 1979. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Friedman:1983:NOC

[FJ83] Avner Friedman and Li Shang Jiang. Nonlinear optimal control problems in heat conduction. SIAM Journal on Control and Optimization, 21(6): 940–952, November 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Freedman:1976:SPT

[FK76] Marvin I. Freedman and James L. Kaplan. Singular perturbations of twopoint boundary value problems arising in optimal control. SIAM Journal on Control and Optimization, 14(2): 189–215, February 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fliess:1983:FCN

[FK83] Michel Fliess and Ivan Kupka.

A finiteness criterion for nonlinear input–output differential systems. SIAM Journal on Control and Optimization, 21(5):721–728, September 1983. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Faigle:1991:NCS

[FK91] U. Faigle and W. Kern. Note on the convergence of simulated annealing algorithms. SIAM Journal on Control and Optimization, 29(1):153– 159, January 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Freeman:1996:IOR

[FK96] R. A. Freeman and P. V. Kokotovic. Inverse optimality in robust stabilization. SIAM Journal on Control and Optimization, 34(4):1365–1391, July 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25873.

Facchinei:1999:BMR

[FK99] Francisco Facchinei and Christian Kanzow. Beyond monotonicity in regularization methods for nonlinear complementarity problems. SIAM Journal on Control and Optimization, 37(4):1150–1161, July 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32293.

Frison:2021:CAA

[FK21] Lilli Frison and Christian Kirches. Convergence analysis and adaptive order selection for the polynomial chaos approach to direct optimal control under uncertainties. SIAM Journal on Control and Optimization, 59(1):509–533, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fliegner:1996:SRI

[FKN96] T. Fliegner, Ü. Kotta, and H. Nijmeijer. Solvability and right-inversion of implicit nonlinear discrete-time systems. SIAM Journal on Control and Optimization, 34(6): 2092–2115, November 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27950.

Feintuch:1986:SMP

[FKT86] Avraham Feintuch, Pramod Khargonekar, and Allen Tannenbaum. On the sensitivity minimization problem for linear time-varying periodic systems. SIAM Journal on Control and Optimization, 24(5): 1076–1085, July 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fu:2023:RCQ

[FKW23] Teng Fu, Kit Ian Kou, and Jinrong Wang. Relative controllability of quaternion differential equations with delay. SIAM Journal on Control and Optimization, 61(5):2927–2952, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1544684.

Feinberg:2022:MDP

[FKZ22] Eugene A. Feinberg, Pavlo O.

Kasyanov, and Michael Z.

Zgurovsky. Markov decision processes with incomplete information and semiuniform Feller transition probabilities. SIAM Journal on Control and Optimization, 60 (4):2488–2513, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1442152.

Friedman:1980:OSC

[FL80] Avner Friedman and Pierre Louis Lions. The optimal strategy in the control problem associated with the Hamilton–Jacobi– Bellman equation. SIAM Journal on Control and Optimization, 18(2):191–198, March 1980. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See corrigendum [FL82].

Feinstein:1981:AAB

[FL81] C. D. Feinstein and D. G. Luenberger. Analysis of the asymptotic behavior of optimal control trajectories: The implicit programming problem. SIAM Journal on Control and Optimization, 19(5):561– 585, September 1981. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Friedman:1982:COS

[FL82] Avner Friedman and Pierre-Louis Lions. Corrigendum:

The optimal strategy in the control problem associated with the Hamilton–Jacobi– Bellman equation. SIAM Journal on Control and Optimization, 20(1):153–154, January 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [FL80].

Fu:2017:COS

[FL17] Xiaoyu Fu and Xu Liu. Controllability and observability of some stochastic complex Ginzburg–Landau equations. SIAM Journal on Control and Optimization, 55(2): 1102–1127, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Feng:2020:CPS

[FL20] Han Feng and Javad Lavaei.

Connectivity properties of the set of stabilizing static decentralized controllers. SIAM Journal on Control and Optimization, 58(5):2790–2820, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fu:2022:OEW

[FL22] Xiaoyu Fu and Zhonghua Liao.

Observability estimate for the wave equation with variable coefficients. SIAM Journal on Control and Optimization, 60(4):2344–2372, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1468231.

Flandoli:1984:REA

[Fla84] Franco Flandoli. Riccati equation arising in a boundary control problem with distributed parameters. SIAM Journal on Control and Optimization, 22 (1):76–86, January 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Flandoli:1987:ARE

[Fla87] Franco Flandoli. Algebraic Riccati equation arising in boundary control problems. SIAM Journal on Control and Optimization, 25(3):612–636, May 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Flandoli:1990:SCB

[Fla90] Franco Flandoli. Solution and control of a bilinear stochastic delay equation. SIAM Journal on Control and Optimization, 28(4):936–949, July 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Flaam:1992:LMS

[Flå92] Sjur Didrik Flåm. Lagrange multipliers in stochastic programming. SIAM Journal on Control and Optimization, 30 (1):1–10, January 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See corrigendum [Flå95].

Flaam:1995:CLM

[Flå95] Sjur Didrik Flåm. Corrigendum: “Lagrange multipliers in stochastic programming” [SIAM J. Control. Optim. 30 (1992), no. 1, 1–10; MR 93c:90051]. SIAM Journal on Control and Optimization, 33(2):667–671, March 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/15567.  See

[Flå92].

Fleming:1982:NSC

[Fle82] Wendell H. Fleming. Nonlinear semigroup for controlled partially observed diffusions. SIAM Journal on Control and Optimization, 20 (2):286–301, March 1982. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fletcher:1985:SDM

[Fle85] R. Fletcher. Semi-definite matrix constraints in optimization. SIAM Journal on Control and Optimization, 23(4): 493–513, July 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Farokhi:2013:DDA

[FLJ13] Farhad Farokhi, Cédric Langbort, and Karl H. Johansson. Decentralized disturbance accommodation with limited plant model information. SIAM Journal on Control and Optimization, 51(2): 1543–1573, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Florchinger:1995:LLT

[Flo95] Patrick Florchinger. Lyapunovlike techniques for stochastic stability. SIAM Journal on Control and Optimization, 33(4):1151–1169, July 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25230.

Florchinger:1997:FSA

[Flo97] Patrick Florchinger. Feedback stabilization of affine in the control stochastic differential systems by the control Lyapunov function method. SIAM Journal on Control and Optimization, 35(2):500–511, March 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27996.

Florchinger:1999:PSA

[Flo99] Patrick Florchinger. A passive system approach to feedback stabilization of nonlinear control stochastic systems. SIAM Journal on Control and Optimization, 37(6): 1848–1864, November 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31748.

Florchinger:2002:SJQ

[Flo02]  Patrick Florchinger.  A

stochastic Jurdjevic–Quinn theorem. SIAM Journal on Control and Optimization, 41(1):83–88, January 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37078.

Fontelos:2017:BDT

[FLRO17] M. A. Fontelos, R. Lecaros, J. C. López Ríos, and J. H. Ortega. Bottom detection through surface measurements on water waves. SIAM Journal on Control and Optimization, 55(6):3890–3907, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fischer:2017:FDU

[FLUU17] Michael Fischer, Florian Lindemann, Michael Ulbrich, and Stefan Ulbrich. Fréchet differentiability of unsteady incompressible Navier–Stokes flow with respect to domain variations of low regularity by using a general analytical framework. SIAM Journal on Control and Optimization, 55(5):

3226–3257, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fattorini:1994:OPN

[FM94] H. O. Fattorini and T. Murphy. Optimal problems for nonlinear parabolic boundary control systems. SIAM Journal on Control and Optimization, 32(6):1577–1596, November 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/23795.

Fleming:1995:RSC

[FM95] Wendell H. Fleming and William M. McEneaney. Risksensitive control on an infinite time horizon. SIAM Journal on Control and Optimization, 33(6):1881–1915, November 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/25872.

Fleming:2000:MPB

[FM00] Wendell H. Fleming and William M. McEneaney. A max-plus-based algorithm for a Hamilton–Jacobi–Bellman equation of nonlinear filtering. SIAM Journal on Control and Optimization, 38(3):683–710, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33243.

Fragnelli:2008:SSS

[FM08] Genni Fragnelli and Dimitri Mugnai. Stability of solutions for some classes of nonlinear damped wave equations. SIAM Journal on Control and Optimization, 47(5): 2520–2539, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Flordal:2009:CVS

[FM09] Hugo Flordal and Robi Malik. Compositional verification in supervisory control. SIAM Journal on Control and Optimization, 48(3): 1914–1938, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fainshil:2012:MPS

[FM12] Lior Fainshil and Michael Margaliot. A maximum principle for the stability analysis of positive bilinear control systems with applications to positive linear switched systems. SIAM Journal on Control and Optimization, 50(4): 2193–2215, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Foroush:2016:TCS

[FM16] Hamed Shisheh Foroush and Sonia Martínez. On triggering control of single-input linear systems under pulsewidth modulated DoS signals. SIAM Journal on Control and Optimization, 54(6): 3084–3105, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fusco:2022:NAC

[FM22] Giovanni Fusco and Monica Motta. Nondegenerate abnormality, controllability, and gap phenomena in optimal control with state constraints. SIAM Journal on Control and Optimization, 60(1):280–309, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1382465.

Faulwasser:2022:OCP

[FMP^{+}22] Timm Faulwasser, Bernhard Maschke, Friedrich Philipp, Manuel Schaller, and Karl Worthmann. Optimal control of port-Hamiltonian descriptor systems with minimal energy supply. SIAM Journal on Control and Optimization, 60(4):2132–2158, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1427723.

Fattahi:2020:ELD

[FMS20] Salar Fattahi, Nikolai Matni, and Somayeh Sojoudi. Efficient learning of distributed linear-quadratic control policies. SIAM Journal on Control and Optimization, 58(5): 2927–2951, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fort:2016:CMS

[FMSV16] G. Fort, E. Moulines, A. Schreck, and M. Vihola. Convergence of Markovian stochastic approximation with discontinuous dynamics. SIAM Journal on Control and Optimization, 54 (2):866–893, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fuhrman:2010:SED

[FMT10] Marco Fuhrman, Federica Masiero, and Gianmario Tessitore. Stochastic equations with delay: Optimal control via BSDEs and regular solutions of Hamilton–Jacobi– Bellman equations. SIAM Journal on Control and Optimization, 48(7):4624–4651, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fusco:2024:OIC

[FMV24] Giovanni Fusco, Monica Motta, and Richard Vinter. Optimal impulsive control for time delay systems. SIAM Journal on Control and Optimization, 62(4):2012–2035, August 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/24M1632450.

Fujii:1984:COC

[FN84] Takao Fujii and Masaru Narazaki. A complete optimality condition in the inverse problem of optimal control. SIAM Journal on Control and Optimization, 22(2): 327–341, March 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fares:2002:RCS

[FNA02] B. Fares, D. Noll, and P. Apkarian. Robust control via sequential semidefinite programming. SIAM Journal on Control and Optimization, 40(6): 1791–1820, November 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/37348.

Fornasini:2022:SSR

[FNP^{+}22] E. Fornasini, D. Napp, R. Pereira, R. Pinto, and P. Rocha. State-space realizations of periodic convolutional codes. SIAM Journal on Control and Optimization, 60(5):2927–2946, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1452172.

Fournie:2019:FST

[FNR19] Michel Fournié, Moctar Ndiaye, and Jean-Pierre Raymond. Feedback stabilization of a two-dimensional fluidstructure interaction system with mixed boundary conditions. SIAM Journal on Control and Optimization, 57(5): 3322–3359, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fridman:2010:SSO

[FNV10] Emilia Fridman, Serge Nicaise, and Julie Valein. Stabilization of second order evolution equations with unbounded feedback with time-dependent delay. SIAM Journal on Control and Optimization, 48(8): 5028–5052, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fulcheri:1998:MRA

[FO98] Pascale Fulcheri and Martine Olivi. Matrix rational H_{2} approximation: a gradient algorithm based on Schur analysis. SIAM Journal on Control and Optimization, 36(6): 2103–2127, November 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28423.

Fuhrman:2016:SMP

[FO16] Marco Fuhrman and Carlo Orrieri. Stochastic maximum principle for optimal control of a class of nonlinear SPDEs with dissipative drift. SIAM Journal on Control and Optimization, 54(1):341–371, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Frankowska:2018:SLM

[FO18] Hélène Frankowska and Nikolai P. Osmolovskii. Strong local minimizers in optimal control problems with state constraints: Second-order necessary conditions. SIAM Journal on Control and Optimization, 56(3):2353–2376, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fontecilla:1989:SIB

[Fon89] Rodrigo Fontecilla. Subspace invariance of Broyden’s θ-class and its application to nonlinear constrained optimization. SIAM Journal on Control and Optimization, 27(1): 43–52, January 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Foreman:1977:DSG

[For77] Joe G. Foreman. Differential search games with mobile hider. SIAM Journal on Control and Optimization, 15(5):

841–856, August 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fleming:1982:OCP

[FP82] Wendell H. Fleming and Étienne Pardoux. Optimal control for partially observed diffusions. SIAM Journal on Control and Optimization, 20 (2):261–285, March 1982. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fleming:1989:PMF

[FP89] W. H. Fleming and É. Pardoux. Piecewise monotone filtering with small observation noise. SIAM Journal on Control and Optimization, 27(5): 1156–1181, September 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ferris:1996:NSR

[FP96] M. C. Ferris and J. S. Pang.

Nondegenerate solutions and related concepts in affine variational inequalities. SIAM Journal on Control and Optimization, 34(1):244–263, January 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26169.

Fort:1999:ABM

[FP99] Jean-Claude Fort and Gilles Pagès. Asymptotic behavior of a Markovian stochastic algorithm with constant step. SIAM Journal on Control and Optimization, 37(5): 1456–1482, September 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32861.

Fagnani:2002:SPA

[FP02] Fabio Fagnani and Luciano Pandolfi. A singular perturbation approach to a recursive deconvolution problem. SIAM Journal on Control and Optimization, 40(5): 1384–1405, September 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36825.

Fuhrman:2003:LCS

[FP03] Marco Fuhrman and Anna Maria Paganoni. Linear control systems on unbounded time intervals and invariant measures of Ornstein–Uhlenbeck processes in Hilbert spaces. SIAM Journal on Control and Optimization, 42(5): 1776–1794, September 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41465.

Fleming:2004:ASC

[FP04] Wendell H. Fleming and Tao Pang. An application of stochastic control theory to financial economics. SIAM Journal on Control and Optimization, 43(2):502–531, March 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41906.

Favini:2008:QRP

[FP08] A. Favini and L. Pandolfi.

A quadratic regulator problem related to identification problems and singular systems. SIAM Journal on Control and Optimization, 47(2): 768–790, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Frazier:2011:CSB

[FP11a] Peter I. Frazier and Warren B.

Powell. Consistency of sequential Bayesian sampling policies. SIAM Journal on Control and Optimization, 49(2): 712–731, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i2/p712_s1.

Frey:2011:NBS

[FP11b] Rüdiger Frey and Ulrike Polte. Nonlinear Black– Scholes equations in finance: Associated control problems and properties of solutions.

SIAM Journal on Control and Optimization, 49(1):185– 204, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicon/resource/
                                        1/sjcodc/v49/i1/p185_s1.

Federico:2014:COB

[FP14] Salvatore Federico and Huyên Pham. Characterization of the optimal boundaries in reversible investment problems. SIAM Journal on Control and Optimization, 52(4): 2180–2223, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Feinberg:2019:SDP

[FP19] Eugene A. Feinberg and Alexey Piunovskiy. Sufficiency of deterministic policies for atomless discounted and uniformly absorbing MDPs with multiple criteria. SIAM Journal on Control and Optimization, 57(1):163–191, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fatkhullin:2021:OSL

[FP21] Ilyas Fatkhullin and Boris Polyak. Optimizing static linear feedback: Gradient method. SIAM Journal on Control and Optimization, 59 (5):3887–3911, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Frazier:2008:KGP

[FPD08] Peter I. Frazier, Warren B.

Powell, and Savas Dayanik. A knowledge-gradient policy for sequential information collection. SIAM Journal on Control and Optimization, 47(5): 2410–2439, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fouque:2017:PAI

[FPS17] J.-P. Fouque, A. Papanicolaou, and R. Sircar. Perturbation analysis for investment portfolios under partial information with expert opinions. SIAM Journal on Control and Optimization, 55(3): 1534–1566, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fiacchini:2015:CSI

[FPT15] Mirko Fiacchini, Christophe Prieur, and Sophie Tarbouriech. On the computation of set-induced control Lyapunov functions for continuous-time systems. SIAM Journal on Control and Optimization, 53(3):1305–1327, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fouque:2016:POA

[FPW16] Jean-Pierre Fouque, Chi Seng Pun, and Hoi Ying Wong. Portfolio optimization with ambiguous correlation and stochastic volatilities. SIAM Journal on Control and Optimization, 54(5):2309–2338, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Feintuch:1978:PAC

[FR78a] Avraham Feintuch and Moshe Rosenfeld. On pole assignment for a class of infinite dimensional linear systems. SIAM Journal on Control and Optimization, 16(2): 270–276, March 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See erratum [FR82a].

Friedman:1978:FBV

[FR78b] Avner Friedman and Maurice Robin. The free boundary for variational inequalities with nonlocal operators. SIAM Journal on Control and Optimization, 16(2): 347–372, March 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Feintuch:1982:EPA

[FR82a] Avraham Feintuch and Moshe Rosenfeld. Erratum: “On pole assignment for a class of infinite-dimensional systems” [SIAM J. Control. Optim. 16 (1978), no. 2, 270–276; MR 58 #9528]. SIAM Journal on Control and Optimization, 20(4):602, July 1982. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [FR78a].

Filar:1982:ASG

[FR82b] Jerzy A. Filar and T. E. S.

Raghavan. An algorithm for solving S-games and differential S-games. SIAM Journal on Control and Optimization, 20(6):763–769, November 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Frankowska:1999:RCS

[FR99] Hélène Frankowska and Franco Rampazzo. Relaxation of control systems under state constraints. SIAM Journal on Control and Optimization, 37(4):1291–1309, July 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33101.

Fomin:2000:AOL

[FR00] Vladimir N. Fomin and Michael V. Ruzhansky. Abstract optimal linear filtering. SIAM Journal on Control and Optimization, 38(5): 1334–1352, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/34778.

Fragoso:2005:SFC

[FR05] Marcelo D. Fragoso and Nei C. S. Rocha. Stationary filter for continuous-time Markovian jump linear systems. SIAM Journal on Control and Optimization, 44(3):801–815, May 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/43625.

Forcadel:2014:SPO

[FR14] Nicolas Forcadel and Zhiping Rao. Singular perturbation of optimal control problems on MultiDomains. SIAM Journal on Control and Optimization, 52(5):2917–2943, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fabbri:2017:HEI

[FR17] Giorgio Fabbri and Francesco Russo. HJB equations in infinite dimension and optimal control of stochastic evolution equations via generalized Fukushima decomposition. SIAM Journal on Control and Optimization, 55(6): 4072–4091, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ferrari:2020:OCD

[FR20] Giorgio Ferrari and Neofytos Rodosthenous. Optimal control of debt-to-GDP ratio in an N-state regime switching economy. SIAM Journal on Control and Optimization, 58 (2):755–786, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Francis:1977:LMR

[Fra77] Bruce A. Francis. The linear multivariable regulator problem. SIAM Journal on Control and Optimization, 15(3): 486–505, May 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Frazho:1980:SOA

[Fra80] Arthur E. Frazho. A shift operator approach to bilinear system theory. SIAM Journal on Control and Optimization, 18(6):640–658, November 1980. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Frazho:1982:MSS

[Fra82] Arthur E. Frazho. On minimal splitting subspaces and stochastic realizations. SIAM Journal on Control and Optimization, 20(4):553–562, July 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Frankowska:1984:FON

[Fra84]  Halina Frankowska.  The

first order necessary conditions for nonsmooth variational and control problems. SIAM Journal on Control and Optimization, 22(1):1–12, January 1984. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Frankowska:1987:LCI

[Fra87a] Halina Frankowska. Local controllability and infinitesimal generators of semigroups of set-valued maps. SIAM Journal on Control and Optimization, 25(2):412–432, March 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Frankowska:1987:MPO

[Fra87b] Halina Frankowska. The maximum principle for an optimal solution to a differential inclusion with end points constraints. SIAM Journal on Control and Optimization, 25(1):145–157, January 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Frankowska:1989:CCR

[Fra89] Halina Frankowska. Contingent cones to reachable sets of control systems. SIAM Journal on Control and Optimization, 27(1):170–198, January 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Frankowska:1993:LSS

[Fra93] Hélène Frankowska. Lower semicontinuous solutions of Hamilton–Jacobi–Bellman equations. SIAM Journal on Control and Optimization, 31(1): 257–272, January 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Framstad:2003:OHJ

[Fra03] Nils Chr. Framstad. Optimal harvesting of a jump diffusion population and the effect of jump uncertainty. SIAM Journal on Control and Optimization, 42(4):1451–1465, January 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Freitas:1999:ORD

[Fre99] Pedro Freitas. Optimizing the rate of decay of solutions of the wave equation using genetic algorithms: a counterexample to the constant damping conjecture. SIAM Journal on Control and Optimization, 37(2):376–387, March 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32944.

Freitas:2000:DII

[Fre00] Pedro Freitas. Delay-induced instabilities in gyroscopic systems. SIAM Journal on Control and Optimization, 39 (1):196–207, January 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/35793.

Friedman:1984:NOC

[Fri84]  Avner Friedman.  Nonlin-

ear optimal control problems for parabolic equations. SIAM Journal on Control and Optimization, 22(5):805– 816, September 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Friedman:1986:OCV

[Fri86] Avner Friedman. Optimal control for variational inequalities. SIAM Journal on Control and Optimization, 24(3): 439–451, May 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Friedman:1987:OCP

[Fri87] Avner Friedman. Optimal control for parabolic variational inequalities. SIAM Journal on Control and Optimization, 25(2):482–497, March 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Frigeri:2016:ODC

[FRS16] Sergio Frigeri, Elisabetta Rocca, and Jürgen Sprekels. Optimal distributed control of a nonlocal Cahn–Hilliard/ Navier–Stokes system in two dimensions. SIAM Journal on Control and Optimization, 54 (1):221–250, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fornasini:1993:SSR

[FRZ93] E. Fornasini, P. Rocha, and S. Zampieri. State space realization of 2-D finite-dimensional behaviours. SIAM Journal on Control and Optimization, 31(6):1502– 1517, November 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fujimoto:2010:BRM

[FS10] Kenji Fujimoto and Jacquelien M. A. Scherpen. Balanced realization and model order reduction for nonlinear systems based on singular value analysis. SIAM Journal on Control and Optimization, 48(7): 4591–4623, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fetzer:2016:GIQ

[FS16] M. Fetzer and C. W. Scherer.

A general integral quadratic constraints theorem with applications to a class of sampled-data systems. SIAM Journal on Control and Optimization, 54(3):1105–1125, ???? 2016. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Ferrari:2019:ODP

[FS19] Giorgio Ferrari and Patrick Schuhmann. An optimal dividend problem with capital injections over a finite horizon. SIAM Journal on Control and Optimization, 57(4): 2686–2719, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Feng:2021:OSC

[FS21] Qian Feng and Jinghai Shao. Optimal singular control problem in infinite horizon for stochastic processes with regime-switching. SIAM Journal on Control and Optimization, 59(2):906–930, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Flayac:2023:NOM

[FS23] Emilien Flayac and Iman Shames. Nonuniform observability for moving horizon estimation and stability with respect to additive perturbation. SIAM Journal on Control and Optimization, 61 (5):3018–3050, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1475132.

Farshbaf-Shaker:2020:ONB

[FSGHH20] M. Hassan Farshbaf-Shaker, Martin Gugat, Holger Heitsch, and René Henrion. Optimal Neumann boundary control of a vibrating string with uncertain initial data and probabilistic terminal constraints. SIAM Journal on Control and Optimization, 58 (4):2288–2311, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Farshbaf-Shaker:2016:OCE

[FSH16] M. Hassan Farshbaf-Shaker and Claudia Hecht. Optimal control of elastic vectorvalued Allen–Cahn variational inequalities. SIAM Journal on Control and Optimization, 54 (1):129–152, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fleming:1987:OSP

[FSS87] W. H. Fleming, S. P. Sethi, and H. M. Soner. An optimal stochastic production planning problem with randomly fluctuating demand. SIAM Journal on Control and Optimization, 25(6):1494– 1502, November 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fisher:1977:NCP

[FT77] M. L. Fisher and J. W.

Tolle. The nonlinear complementarity problem: Existence and determination of solutions. SIAM Journal on Control and Optimization, 15 (4):612–624, July 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Foias:1989:WOT

[FT89] Ciprian Foiaş and Allen Tannenbaum. Weighted optimization theory for nonlinear systems. SIAM Journal on Control and Optimization, 27(4): 842–860, July 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fukushima:2002:DGD

[FT02] Masatoshi Fukushima and Michael Taksar. Dynkin games via Dirichlet forms and singular control of onedimensional diffusions. SIAM Journal on Control and Optimization, 41(3):682–699, May 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38713.

Fuhrman:2004:EOS

[FT04] Marco Fuhrman and Gianmario Tessitore. Existence of optimal stochastic controls and global solutions of forward-backward stochastic differential equations. SIAM Journal on Control and Optimization, 43(3):813–830, May 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42866.

Fiaschi:2012:BSC

[FT12] A. Fiaschi and G. Treu. The bounded slope condition for functionals depending on x, u, and ∇u. SIAM Journal on Control and Optimization, 50 (2):991–1011, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Frankowska:2013:PSO

[FT13] Hélène Frankowska and Daniela Tonon. Pointwise second-order necessary optimality conditions for the Mayer problem with control constraints. SIAM Journal on Control and Optimization, 51(5):3814– 3843, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Federico:2014:DPO

[FT14] Salvatore Federico and Elisa Tacconi. Dynamic programming for optimal control problems with delays in the control variable. SIAM Journal on Control and Optimization, 52(2):1203–1236, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fujii:2022:MFG

[FT22] Masaaki Fujii and Akihiko Takahashi. A mean field game approach to equilibrium pricing with market clearing condition. SIAM Journal on Control and Optimization, 60(1):259–279, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1326295.

Flesch:1998:SOS

[FTV98] J. Flesch, F. Thuijsman, and O. J. Vrieze. Simplifying optimal strategies in stochastic games. SIAM Journal on Control and Optimization, 36(4):1331–1347, July 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/31194.

Foias:1987:OSP

[FTZ87] Ciprian Foiaş, Allen Tannenbaum, and George Zames. On the H^{∞}-optimal sensitivity problem for systems with delays. SIAM Journal on Control and Optimization, 25(3): 686–705, May 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fu:2012:SDR

[Fu12] Xiaoyu Fu. Sharp decay rates for the weakly coupled hyperbolic system with one internal damping. SIAM Journal on Control and Optimization, 50(3):1643–1660, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fu:2023:EMF

[Fu23]  Guanxing Fu.  Extended

mean field games with singular controls. SIAM Journal on Control and Optimization, 61(1):285–314, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1384750.

Fuhrmann:1976:SPC

[Fuh76]  Paul A. Fuhrmann.  On

series and parallel coupling of a class of discrete time infinite-dimensional systems. SIAM Journal on Control and Optimization, 14(2):339– 358, February 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fujii:1980:FSD

[Fuj80a] Nobuo Fujii. Feedback stabilization of distributed parameter systems by a functional observer. SIAM Journal on Control and Optimization, 18 (2):108–120, March 1980. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fujimoto:1980:NCP

[Fuj80b] Takao Fujimoto. Nonlinear complementarity problems in a function space. SIAM Journal on Control and Optimization, 18(6):621–623, November 1980. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fujii:1986:NCD

[Fuj86] Nobuo Fujii. Necessary conditions for a domain optimization problem in elliptic boundary value problems. SIAM Journal on Control and Optimization, 24(3):346–360, May 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fleming:1989:CDA

[FV89] Wendell H. Fleming and Domokos Vermes. Convex duality approach to the optimal control of diffusions. SIAM Journal on Control and Optimization, 27(5):1136– 1155, September 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fornasini:1995:MPT

[FV95] Ettore Fornasini and Maria Elena Valcher. Matrix pairs in twodimensional systems: An approach based on trace series and Hankel matrices. SIAM Journal on Control and Optimization, 33(4):1127–1150, July 1995. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24632.

Fornasini:1998:MSF

[FV98] Ettore Fornasini and Maria Elena Valcher. Multidimensional systems with finite support behaviors: Signal structure, generation, and detection. SIAM Journal on Control and Optimization, 36(2):760–779, March 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29204.

Fornasini:2011:RCD

[FV11] Ettore Fornasini and Maria Elena Valcher. Reachability of a class of discrete-time positive switched systems. SIAM Journal on Control and Optimization, 49(1):162–184, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sicon/
resource/1/sjcodc/v49/i1/
p162_s1.

Fagnani:1992:CS

[FW92] Fabio Fagnani and Jan C.

Willems. Controllability of ℓ^{2}-systems. SIAM Journal on Control and Optimization, 30(5):1101–1125, September 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fagnani:1993:RSL

[FW93] Fabio Fagnani and Jan C. Willems. Representations of symmetric linear dynamical systems. SIAM Journal on Control and Optimization, 31(5):1267–1293, September 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fisher:1976:UCP

[FY76] Lloyd Fisher and S. J. Yakowitz. Uniform convergence of the potential function algorithm. SIAM Journal on Control and Optimization, 14 (1):95–103, January 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fu:2007:ECM

[FYZ07] Xiaoyu Fu, Jiongmin Yong, and Xu Zhang. Exact controllability for multidimensional semilinear hyperbolic equations. SIAM Journal on Control and Optimization, 46(5): 1578–1614, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fleming:1998:RSP

[FZ98] W. H. Fleming and Q. Zhang. Risk-sensitive production planning of a stochastic manufacturing system. SIAM Journal on Control and Optimization, 36(4):1147–1170, July 1998. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31034.

Ferrante:2000:LQO

[FZ00] Augusto Ferrante and Sandro Zampieri. Linear quadratic optimization for systems in the behavioral approach. SIAM Journal on Control and Optimization, 39(1):159–178, January 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35243.

Fagnani:2005:SAP

[FZ05] Fabio Fagnani and Sandro Zampieri. A symbolic approach to performance analysis of quantized feedback systems: The scalar case. SIAM Journal on Control and Optimization, 44(3):816–866, May 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43431.

Fagnani:2009:ACP

[FZ09] Fabio Fagnani and Sandro Zampieri. Average consensus with packet drop communication. SIAM Journal on Control and Optimization, 48(1): 102–133, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Feng:2019:OBC

[FZ19] Binhua Feng and Dun Zhao. Optimal bilinear control for the X-ray free electron laser Schrödinger equation. SIAM Journal on Control and Optimization, 57(5):3193–3222, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Fu:2023:SRP

[FZ23] Yanxin Fu and Wenxiao Zhao. Support recovery and parameter identification of multivariate ARMA systems with exogenous inputs. SIAM Journal on Control and Optimization, 61(3):1835–1860, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M146346X.

Frankowska:2018:SOC

[FZZ18] Hélène Frankowska, Haisen Zhang, and Xu Zhang. Stochastic optimal control problems with control and initial-final states constraints. SIAM Journal on Control and Optimization, 56(3):1823– 1855, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gibson:1991:ATL

[GA91] J. S. Gibson and A. Adamian. Approximation theory for linear-quadratic-Gaussian control of flexible structures.

SIAM Journal on Control and Optimization, 29(1):1–37, January 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gallegos:2024:REC

[GAC24] Javier Gallegos and Norelys Aguila-Camacho. Relaxed excitation conditions for robust identification and adaptive control using estimation with memory. SIAM Journal on Control and Optimization, 62(1):1–21, January 2024. CO-DEN SJCODC. ISSN 1095-7138.

Gadat:2008:JDF

[Gad08] Sébastien Gadat. Jump diffusion over feature space for object recognition. SIAM Journal on Control and Optimization, 47(2):904–935, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gagnon:2016:LCD

[Gag16]  Ludovick Gagnon.  La-

grangian controllability of the 1-dimensional Korteweg–de Vries equation. SIAM Journal on Control and Optimization, 54(6):3152–3173, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gahinet:1994:GRE

[Gah94] Pascal Gahinet. On the game Riccati equations arising in H_{∞} control problems. SIAM Journal on Control and Optimization, 32(3):635–647, May 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/19900.

Guigue:2009:DDP

[GAHL09] A. Guigue, M. Ahmadi, M. J. D. Hayes, and R. G. Langlois. A discrete dynamic programming approximation to the multiobjective deterministic finite horizon optimal control problem. SIAM Journal on Control and Optimization, 48(4):2581–2599, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gaitsgory:1992:SSP

[Gai92]  Vladimir Gaitsgory.  Sub-

optimization of singularly perturbed control systems. SIAM Journal on Control and Optimization, 30(5):1228– 1249, September 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gaitsgory:2004:RLO

[Gai04] Vladimir Gaitsgory. On a representation of the limit occupational measures set of a control system with applications to singularly perturbed control systems. SIAM Journal on Control and Optimization, 43(1):325–340, January 2004. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42418.

Gal:1979:SGM

[Gal79] Shmuel Gal. Search games with mobile and immobile hider. SIAM Journal on Control and Optimization, 17(1): 99–122, January 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gal:1984:AOR

[Gal84] Shmuel Gal. An O(N^{3}) algorithm for optimal replacement problems. SIAM Journal on Control and Optimization, 22(6):902–910, November 1984. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Galbraith:2000:EHJ

[Gal00] Grant N. Galbraith. Extended Hamilton–Jacobi characterization of value functions in optimal control. SIAM Journal on Control and Optimization, 39(1):281–305, January 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34788.

Ghosh:1993:OCS

[GAM93] Mrinal K. Ghosh, Aristotle Arapostathis, and Steven I. Marcus. Optimal control of switching diffusions with application to flexible manufacturing systems. SIAM Journal on Control and Optimization, 31(5):1183–1204, September 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ghosh:1997:ECS

[GAM97] Mrinal K. Ghosh, Aristotle Arapostathis, and Steven I. Marcus. Ergodic control of switching diffusions. SIAM Journal on Control and Optimization, 35(6): 1952–1988, November 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29930.

Ghabcheloo:2009:CPF

[GAP^{+}09] R. Ghabcheloo, A. P. Aguiar, A. Pascoal, C. Silvestre, I. Kaminer, and J. Hespanha. Coordinated path-following in the presence of communication losses and time delays. SIAM Journal on Control and Optimization, 48(1):234–265, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gard:1976:GUT

[Gar76] Thomas C. Gard. A general uniqueness theorem for solutions of stochastic differential equations. SIAM Journal on Control and Optimization, 14 (3):445–457, May 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Garavello:2003:VTH

[Gar03] Mauro Garavello. Verification theorems for Hamilton– Jacobi–Bellman equations. SIAM Journal on Control and Optimization, 42(5): 1623–1642, September 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39268.

Gauthier:1982:CSB

[GB82a] J. P. Gauthier and G. Bornard.

Contrôlabilité des systèmes bilinéaires. (French) [Controllability of bilinear systems]. SIAM Journal on Control and Optimization, 20(3):377–384, May 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gauthier:1982:OCC

[GB82b] J. P. Gauthier and G. Bornard.

An openness condition for the controllability of nonlinear systems. SIAM Journal on Control and Optimization, 20(6):808–814, November 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gafni:1984:TMP

[GB84a] Eli M. Gafni and Dimitri P.

Bertsekas. Two-metric projection methods for constrained optimization. SIAM Journal on Control and Optimization, 22(6):936–964, November 1984. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gauthier:1984:EUM

[GB84b] J. P. Gauthier and G. Bornard. Existence and uniqueness of minimal realizations for a class of C^{∞} systems. SIAM Journal on Control and Optimization, 22(4):666–670, July 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gauthier:1986:GRA

[GB86] J. P. Gauthier and G. Bornard. Global realizations of analytic input–output mappings. SIAM Journal on Control and Optimization, 24(3):509–521, May 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ghosh:1993:SCA

[GB93] Bijoy K. Ghosh and Paul R. Bouthellier. Simultaneous coefficient assignment of discrete-time multi-input multi-output linear timevarying system: a new approach for compensator design. SIAM Journal on Control and Optimization, 31(6): 1438–1461, November 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Giordano:2018:RSS

[GBB18] Giulia Giordano, Dario Bauso, and Franco Blanchini. A robust saturated strategy for n-player Prisoner’s Dilemma. SIAM Journal on Control and Optimization, 56(5):3478– 3498, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gal:1978:FMR

[GBBT78] Shmuel Gal, Boris Bachelis, and Aharon Ben-Tal. On finding the maximal range of validity of a constrained system. SIAM Journal on Control and Optimization, 16(3):473–503, May 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gaubert:1998:MMR

[GBCG98] Stéphane Gaubert, Peter Butkovic, and Raymond Cuninghame-Green. Minimal (max,+ ) realization of convex sequences. SIAM Journal on Control and Optimization, 36(1):137–147, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28534.

Geromel:2000:RFD

[GBGdO00] José C. Geromel, Jacques Bernussou, Germain Garcia, and Maurício C. de Oliveira. H_{2} and H_{∞} robust filtering for discrete-time linear systems. SIAM Journal on Control and Optimization, 38(5):1353–1368, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32737.

Grasse:1997:RPP

[GBo97] Kevin A. Grasse and Jonathan R. Bar-on. Regularity properties of the phase for multivariable systems. SIAM Journal on Control and Optimization, 35(4):1366–1386, July 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27924.

Ghattassi:2018:ROO

[GBR18] Mohamed Ghattassi, Mohamed Boutayeb, and Jean Rodolphe Roche. Reduced order observer of finite dimensional radiative-conductive heat transfer systems. SIAM Journal on Control and Optimization, 56(4):2485–2512, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Grussler:2022:IHP

[GBS22] Christian Grussler, Thiago Burghi, and Somayeh Sojoudi. Internally Hankel k-positive systems. SIAM Journal on Control and Optimization, 60 (4):2373–2392, ???? 2022.

CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1404685.

Guo:2005:OCE

[GC05] Xianping Guo and Xi-Ren Cao. Optimal control of ergodic continuous-time Markov chains with average samplepath rewards. SIAM Journal on Control and Optimization, 44(1):29–48, January 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42087.

Geromel:2006:SSC

[GC06] José C. Geromel and Patrizio Colaneri. Stability and stabilization of continuous-time switched linear systems. SIAM Journal on Control and Optimization, 45(5):1915–1930, ???? 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Goodman:2022:CAM

[GC22] Jacob R. Goodman and Leonardo J. Colombo. Collision avoidance of multiagent systems on Riemannian manifolds. SIAM Journal on Control and Optimization, 60(1):168–188, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1411056.

Ganguli:2006:MVN

[GCB06] Anurag Ganguli, Jorge Cortés, and Francesco Bullo. Maximizing visibility in nonconvex polygons: Nonsmooth analysis and gradient algorithm design. SIAM Journal on Control and Optimization, 45(5): 1657–1679, ???? 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gao:2007:PPN

[GCC07] Huijun Gao, Tongwen Chen, and Tianyou Chai. Passivity and passification for networked control systems. SIAM Journal on Control and Optimization, 46(4):1299–1322, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gao:2015:OEN

[GCL15] Peng Gao, Mo Chen, and Yong Li. Observability estimates and null controllability for forward and backward linear stochastic Kuramoto– Sivashinsky equations. SIAM Journal on Control and Optimization, 53(1):475–500, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Glover:1988:RAL

[GCP88] Keith Glover, Ruth F. Curtain, and Jonathan R. Partington. Realisation and approximation of linear infinitedimensional systems with error bounds. SIAM Journal on Control and Optimization, 26 (4):863–898, July 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gu:1999:BSL

[GCSB99] Guoxiang Gu, Xiang Chen, Andrew G. Sparks, and Siva S. Banda. Bifurcation stabilization with local output feedback. SIAM Journal on Control and Optimization, 37(3):934–956, May 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32092.

Ghosh:1988:DGM

[GD88] B. K. Ghosh and W. P.

Dayawansa. Differential geometric methods in hybrid parametrization of linear dynamical systems. SIAM Journal on Control and Optimization, 26(5):1149–1174, September 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gugat:2011:GFF

[GDL11] Martin Gugat, Markus Dick, and Günter Leugering. Gas flow in fan-shaped networks: Classical solutions and feedback stabilization. SIAM Journal on Control and Optimization, 49(5):2101–2117, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sicon/
                                        resource/1/sjcodc/v49/i5/
                                        p2101_s1.

Geromel:2002:RFD

[GdOB02] J. C. Geromel, M. C. de Oliveira, and J. Bernussou. Robust filtering of discrete-time linear systems with parameter dependent Lyapunov functions. SIAM Journal on Control and Optimization, 41(3):700–711, May 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/36630.

Gracia:1990:CFE

[GdZ90] J. M. Gracia, I. de Hoyos, and I. Zaballa. A characterization of feedback equivalence. SIAM Journal on Control and Optimization, 28(5): 1103–1112, September 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gray:2017:SOA

[GEF17] W. Steven Gray and Kurusch Ebrahimi-Fard. SISO output affine feedback transformation group and its Faà di Bruno Hopf algebra. SIAM Journal on Control and Optimization, 55(2):885–912, ????

2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gensbittel:2013:CCP

[Gen13] Fabien Gensbittel. Covariance control problems over martingales with fixed terminal distribution arising from game theory. SIAM Journal on Control and Optimization, 51(2): 1152–1185, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gerencser:1992:RCR

[Ger92] László Gerencsér. Rate of convergence of recursive estimators. SIAM Journal on Control and Optimization, 30(5): 1200–1227, September 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gerencser:2006:RTE

[Ger06] László Gerencsér. A representation theorem for the error of recursive estimators. SIAM Journal on Control and Optimization, 44(6):2123– 2188, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Geveci:1981:BCH

[Gev81]  Tunc Geveci.  Boundary

controllability of hyperbolic partial differential equations. SIAM Journal on Control and Optimization, 19(3):359–367, May 1981. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Glizer:2017:NAE

[GFF17] Valery Y. Glizer, Emilia Fridman, and Yuri Feigin. A novel approach to exact slow-fast decomposition of linear singularly perturbed systems with small delays. SIAM Journal on Control and Optimization, 55(1):236–274, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gfrerer:2006:SOO

[Gfr06]  Helmut Gfrerer.  Second-

order optimality conditions for scalar and vector optimization problems in Banach spaces. SIAM Journal on Control and Optimization, 45(3): 972–997, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Glass:2007:UCB

[GG07] O. Glass and S. Guerrero. On the uniform controllability of the Burgers equation. SIAM Journal on Control and Optimization, 46(4): 1211–1238, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ghosh:2009:RMO

[GG09] Mrinal K. Ghosh and Anindya Goswami. Risk minimizing option pricing in a semi-Markov modulated market. SIAM Journal on Control and Optimization, 48(3):1519–1541, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Guo:2013:AOF

[GG13] Wei Guo and Bao-Zhu Guo.

Adaptive output feedback stabilization for one-dimensional wave equation with corrupted observation by harmonic disturbance. SIAM Journal on Control and Optimization, 51 (2):1679–1706, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Grune:2018:TPS

[GG18] Lars Grüne and Roberto Guglielmi. Turnpike properties and strict dissipativity for discrete time linear quadratic optimal control problems. SIAM Journal on Control and Optimization, 56 (2):1282–1302, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gobet:2022:NMS

[GG22] Emmanuel Gobet and Maxime Grangereau. Newton method for stochastic control problems. SIAM Journal on Control and Optimization, 60 (5):2996–3025, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1408567.

Gugat:2024:OPF

[GG24] Martin Gugat and Jan Giesselmann. An observer for pipeline flow with hydrogen blending in gas networks: Exponential synchronization. SIAM Journal on Control and Optimization, 62(4):2273–2296, August 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1563840.

Gurumoorthy:2017:RDK

[GGA^{+}17] Karthik S. Gurumoorthy, Colin Grudzien, Amit Apte, Alberto Carrassi, and Christopher K. R. T. Jones. Rank deficiency of Kalman error covariance matrices in linear time-varying system with deterministic evolution. SIAM Journal on Control and Optimization, 55(2):741–759, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Guillen-Gonzalez:2023:OCP

[GGC23] Francisco Guillén-González and AndréLuiz Corrêa Vianna Filho. An optimal control problem subject to strong solutions of chemotaxis-consumption models. SIAM Journal on Control and Optimization, 61 (5):3156–3182, October 2023. CODEN SJCODC. ISSN 1095-7138.

Geromel:2009:DOF

[GGF09] José C. Geromel, Alim P. C.

Gonçalves, and André R. Fioravanti. Dynamic output feedback control of discrete-time Markov jump linear systems through linear matrix inequalities. SIAM Journal on Control and Optimization, 48(2): 573–593, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ghisi:2016:RET

[GGH16] Marina Ghisi, Massimo Gobbino, and Alain Haraux. The remarkable effectiveness of time-dependent damping terms for second order evolution equations. SIAM Journal on Control and Optimization, 54(3):1266–1294, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Green:1990:SFA

[GGLD90] Michael Green, Keith Glover, David Limebeer, and John Doyle. A J-spectral factorization approach to H_{∞}. SIAM Journal on Control and Optimization, 28(6):1350– 1371, November 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Garreau:2001:TAP

[GGM01] Stéphane Garreau, Philippe Guillaume, and Mohamed Masmoudi. The topological asymptotic for PDE systems: The elasticity case. SIAM Journal on Control and Optimization, 39(6):1756–1778, ???? 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Guillen-Gonzalez:2020:RCC

[GGMZRB20] Francisco Guillen-Gonzalez, Exequiel Mallea-Zepeda, and Maria Angeles Rodriguez-Bellido. A regularity criterion for a 3D chemo-repulsion system and its application to a bilinear optimal control problem. SIAM Journal on Control and Optimization, 58(3): 1457–1490, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Giambo:2002:EMR

[GGP02] Roberto Giambò, Fabio Giannoni, and Paolo Piccione. Existence, multiplicity, and regularity for sub-Riemannian geodesics by variational methods. SIAM Journal on Control and Optimization, 40(6): 1840–1857, November 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36724.

Gassiat:2014:ICP

[GGP14] Paul Gassiat, Fausto Gozzi, and Huyên Pham. Investment/consumption problem in illiquid markets with regimeswitching. SIAM Journal on Control and Optimization, 52 (3):1761–1786, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gevers:1991:CTS

[GGW91] M. Gevers, G. C. Goodwin, and V. Wertz. Continuoustime stochastic adaptive control. SIAM Journal on Control and Optimization, 29(2): 264–282, March 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Geman:1986:DGO

[GH86] Stuart Geman and Chii-Ruey Hwang. Diffusions for global optimization. SIAM Journal on Control and Optimization, 24(5):1031–1043, July 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gunzburger:1996:FDA

[GH96] Max D. Gunzburger and L. Steven Hou. Finitedimensional approximation of a class of constrained nonlinear optimal control problems. SIAM Journal on Control and Optimization, 34(3):1001–1043, May 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26236.

Guo:2004:BFS

[GH04] Faming Guo and Falun Huang. Boundary feedback stabilization of the undamped Euler– Bernoulli beam with both ends free. SIAM Journal on Control and Optimization, 43 (1):341–356, January 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/38096.

Gutman:2007:IPC

[GH07] Semion Gutman and Junhong Ha. Identifiability of piecewise constant conductivity in a heat conduction process. SIAM Journal on Control and Optimization, 46(2):694–713, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Glass:2012:PMS

[GH12] O. Glass and T. Horsin. Prescribing the motion of a set of particles in a threedimensional perfect fluid. SIAM Journal on Control and Optimization, 50(5):2726– 2742, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Glover:2014:OPL

[GH14] Kristoffer Glover and Hardy Hulley. Optimal prediction of the last-passage time of a transient diffusion. SIAM Journal on Control and Optimization, 52(6):3833–3853, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Grasse:2015:SRC

[GH15] Kevin A. Grasse and Nancy Ho. Simulation relations and controllability properties of linear and nonlinear control systems. SIAM Journal on Control and Optimization, 53(3):1346–1374, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Graewe:2017:OTE

[GH17] Paulwin Graewe and Ulrich Horst. Optimal trade execution with instantaneous price impact and stochastic resilience. SIAM Journal on Control and Optimization, 55 (6):3707–3725, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gugat:2019:TPO

[GH19] Martin Gugat and Falk M.

Hante. On the turnpike phenomenon for optimal boundary control problems with hyperbolic systems. SIAM Journal on Control and Optimization, 57(1):264–289, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gong:2020:PFT

[GH20] Ping Gong and Qing-Long Han. Practical fixed-time bipartite consensus of nonlinear incommensurate fractionalorder multiagent systems in directed signed networks. SIAM Journal on Control and Optimization, 58(6):3322– 3341, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gharesifard:2011:SSO

[Gha11] Bahman Gharesifard. Stabilization of systems with one degree of underactuation with energy shaping: a geometric approach. SIAM Journal on Control and Optimization, 49(4):1422–1434, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sicon/
resource/1/sjcodc/v49/i4/
p1422_s1.

Gungah:2000:MRC

[GHJ00] S. K. Gungah, G. D. Halikias, and I. M. Jaimoukha. Maximally robust controllers for multivariable systems. SIAM Journal on Control and Optimization, 38(6): 1805–1829, November 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/35055.

Ghosh:1986:TIM

[Gho86a] B. K. Ghosh. Transcendental and interpolation methods in simultaneous stabilization and simultaneous partial pole placement problems. SIAM Journal on Control and Optimization, 24(6):1091– 1109, November 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ghosh:1986:ASS

[Gho86b] Bijoy K. Ghosh. An approach to simultaneous system design. I. Semialgebraic geometric methods. SIAM Journal on Control and Optimization, 24 (3):480–496, May 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ghosh:1988:ASS

[Gho88] Bijoy K. Ghosh. An approach to simultaneous system design. Part II: Nonswitching gain and dynamic feedback compensation by algebraic geometric methods. SIAM Journal on Control and Optimization, 26 (4):919–963, July 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Graewe:2015:NML

[GHQ15] Paulwin Graewe, Ulrich Horst, and Jinniao Qiu. A non-Markovian liquidation problem and backward SPDEs with singular terminal conditions. SIAM Journal on Control and Optimization, 53(2): 690–711, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gunzburger:1992:BVC

[GHS92] Max D. Gunzburger, LiSheng Hou, and Thomas P. Svobodny. Boundary velocity control of incompressible flow with an application to viscous drag reduction. SIAM Journal on Control and Optimization, 30(1):167–181, January 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Guo:2012:LPC

[GHS12] Xianping Guo, Yonghui Huang, and Xinyuan Song. Linear programming and constrained average optimality for general continuous-time Markov decision processes in history-dependent policies. SIAM Journal on Control and Optimization, 50(1): 23–47, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v50/i1/p23_s1.

Graewe:2022:MPA

[GHS22] Paulwin Graewe, Ulrich Horst, and Ronnie Sircar. A maximum principle approach to a deterministic mean field game of control with absorption. SIAM Journal on Control and Optimization, 60 (5):3173–3190, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1412451.

Goodwin:1986:SAC

[GHX86] Graham C. Goodwin, David J.

Hill, and Xian Ya Xie. Stochastic adaptive control for exponentially convergent time-varying systems. SIAM Journal on Control and Optimization, 24(4):589–603, July 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Glazos:1998:SMS

[GHZ98] Michael P. Glazos, Stefen Hui, and Stanislaw H. Zak. Sliding modes in solving convex programming problems. SIAM Journal on Control and Optimization, 36(2):680–697, March 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25588.

Gong:2014:PEA

[GHZ14] Wei Gong, Michael Hinze, and Zhaojie Zhou. A priori error analysis for finite element approximation of parabolic optimal control problems with pointwise control. SIAM Journal on Control and Optimization, 52(1):97–119, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Guo:2015:FPM

[GHZ15] Xianping Guo, Xiangxiang Huang, and Yi Zhang. On the first passage g-mean-variance optimality for discounted continuous-time Markov decision processes. SIAM Journal on Control and Optimization, 53(3):1406–1424, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Guo:2024:MOO

[GHZ24] Xin Guo, Anran Hu, and Junzi Zhang. MF-OMO: An optimization formulation of meanfield games. SIAM Journal on Control and Optimization, 62(1):243–270, January 2024. CODEN SJCODC. ISSN 1095-7138.

Guillaume:2002:TAE

[GI02] Ph. Guillaume and K. Sid Idris. The topological asymptotic expansion for the Dirichlet problem. SIAM Journal on Control and Optimization, 41(4):1042–1072, July 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/38419.

Guillaume:2004:TSS

[GI04] Ph. Guillaume and K. Sid Idris. Topological sensitivity and shape optimization for the Stokes equations. SIAM Journal on Control and Optimization, 43(1):1–31, January 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41121.

Gibson:1979:RIE

[Gib79] J. S. Gibson. The Riccati integral equations for optimal control problems on Hilbert spaces. SIAM Journal on Control and Optimization, 17 (4):537–565, July 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gibson:1980:NSI

[Gib80] J. S. Gibson. A note on stabilization of infinite dimensional linear oscillators by compact linear feedback. SIAM Journal on Control and Optimization, 18(3):311–316, May 1980. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gibson:1981:AOM

[Gib81] J. S. Gibson. An analysis of optimal modal regulation: Convergence and stability. SIAM Journal on Control and Optimization, 19(5):686– 707, September 1981. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gibson:1983:LQO

[Gib83] J. S. Gibson. Linear-quadratic optimal control of hereditary differential systems: Infinite dimensional Riccati equations and numerical approximations. SIAM Journal on Control and Optimization, 21 (1):95–139, January 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gilbert:1977:OPC

[Gil77]  Elmer G. Gilbert.  Opti-

mal periodic control: a general theory of necessary conditions. SIAM Journal on Control and Optimization, 15(5): 717–746, August 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Grabowski:2011:PMP

[GJ11] Janusz Grabowski and Michał Jóźwikowski. Pontryagin maximum principle on almost Lie algebroids. SIAM Journal on Control and Optimization, 49(3):1306–1357, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sicon/
resource/1/sjcodc/v49/i3/
p1306_s1.

Giraldi:2020:PBD

[GJ20] Laetitia Giraldi and Frédéric Jean. Periodical body deformations are optimal strategies for locomotion. SIAM Journal on Control and Optimization, 58(3):1700–1714, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gutierrez:2008:SVE

[GJN08] C. Gutiérrez, B. Jiménez, and V. Novo. A set-valued Ekeland’s variational principle in vector optimization. SIAM Journal on Control and Optimization, 47(2):883–903, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Germani:1988:GAO

[GJP88] A. Germani, L. Jetto, and M. Piccioni. Galerkin approximation for optimal linear filtering of infinite-dimensional linear systems. SIAM Journal on Control and Optimization, 26(6):1287–1305, November 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gauthier:2010:MPF

[GJZ10] Jean-Paul Gauthier, Bronisław Jakubczyk, and Vladimir Zakalyukin. Motion planning and fastly oscillating controls. SIAM Journal on Control and Optimization, 48(5): 3433–3448, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Green:1985:SLS

[GK85] William L. Green and Edward W. Kamen. Stabilizability of linear systems over a commutative normed algebra with applications to spatiallydistributed and parameterdependent systems. SIAM Journal on Control and Optimization, 23(1):1–18, January 1985. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gohberg:1986:SRT

[GK86] I. Gohberg and M. A. Kaashoek. Similarity and reduction for time varying linear systems with well-posed boundary conditions. SIAM Journal on Control and Optimization, 24(5):961–978, July 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Georgiou:1987:CAS

[GK87a] Tryphon T. Georgiou and Pramod P. Khargonekar. A constructive algorithm for sensitivity optimization of periodic systems. SIAM Journal on Control and Optimization, 25(2):334–340, March 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Georgiou:1987:SFN

[GK87b] Tryphon T. Georgiou and Pramod P. Khargonekar. Spectral factorization and Nevanlinna–Pick interpolation. SIAM Journal on Control and Optimization, 25(3): 754–766, May 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gauthier:1994:OON

[GK94] J. P. Gauthier and I. A. K.

Kupka. Observability and observers for nonlinear systems. SIAM Journal on Control and Optimization, 32(4):975–994, July 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/22179. See er-

ratum [GK95].

Gauthier:1995:EOO

[GK95] J.-P. Gauthier and I. A. K.

Kupka. Erratum: “Observability and observers for nonlinear systems” [SIAM J. Control Optim. 32 (1994), no. 4, 975–994; MR 95b:93025]. SIAM Journal on Control and Optimization, 33(4):1277, July 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [GK94].

Gunzburger:1998:EOS

[GK98] Max D. Gunzburger and Hongchul Kim. Existence of an optimal solution of a shape control problem for the stationary Navier–Stokes equations. SIAM Journal on Control and Optimization, 36(3):895–909, May 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27612.

Ghrist:2002:SCR

[GK02] Robert W. Ghrist and Daniel E. Koditschek. Safe cooperative robot dynamics on graphs. SIAM Journal on Control and Optimization, 40(5):1556–1575, September 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36844.

Griesse:2006:OCS

[GK06] Roland Griesse and Karl Kunisch. Optimal control for a stationary MHD system in velocity-current formulation. SIAM Journal on Control and Optimization, 45(5): 1822–1845, ???? 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gunzburger:2011:STA

[GK11] Max D. Gunzburger and Angela Kunoth. Space–time adaptive wavelet methods for optimal control problems constrained by parabolic evolution equations. SIAM Journal on Control and Optimization, 49(3):1150–1170, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sicon/
resource/1/sjcodc/v49/i3/
p1150_s1.

Gabasov:2000:SPO

[GKB00] R. Gabasov, F. M. Kirillova, and N. V. Balashevich. On the synthesis problem for optimal control systems. SIAM Journal on Control and Optimization, 39(4):1008–1042, July 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36124.

Ghoussoub:2020:OBS

[GKL20] Nassif Ghoussoub, Young-Heon Kim, and Tongseok Lim. Optimal Brownian stopping when the source and target are radially symmetric distributions. SIAM Journal on Control and Optimization, 58(5): 2765–2789, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gu:1992:FDA

[GKLM92] G. Gu, P. P. Khargonekar, E. B. Lee, and P. Misra. Finite-dimensional approximations of unstable infinitedimensional systems. SIAM Journal on Control and Optimization, 30(3):704–716, May 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Groschel:2014:RTA

[GKLW14] Michael Gröschel, Alexander Keimer, Günter Leugering, and Zhiqiang Wang. Regularity theory and adjoint-based optimality conditions for a nonlinear transport equation with nonlocal velocity. SIAM Journal on Control and Optimization, 52(4):2141–2163, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ghoussoub:2018:OTC

[GKP18] Nassif Ghoussoub, Young-Heon Kim, and Aaron Zeff Palmer. Optimal transport with controlled dynamics and free end times. SIAM Journal on Control and Optimization, 56(5):3239–3259, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Garcia:2021:ECC

[GKY21] Nicolás Garcia, Christoph Kawan, and Serdar Yüksel. Ergodicity conditions for controlled stochastic nonlinear systems under information constraints: a volume growth approach. SIAM Journal on Control and Optimization, 59 (1):534–560, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gabay:1976:ECM

[GL76] Daniel Gabay and David G.

Luenberger. Efficiently converging minimization methods based on the reduced gradient. SIAM Journal on Control and Optimization, 14(1): 42–61, January 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gahinet:1990:CBS

[GL90] P. Gahinet and A. J. Laub.

Computable bounds for the sensitivity of the algebraic Riccati equation. SIAM Journal on Control and Optimization, 28(6):1461–1480, November 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gahinet:1992:ARE

[GL92] P. Gahinet and A. J. Laub. Algebraic Riccati equations and the distance to the nearest uncontrollable pair. SIAM Journal on Control and Optimization, 30(4):765–786, July 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ghosh:1995:PTM

[GL95] B. K. Ghosh and E. P. Loucks.

A perspective theory for motion and shape estimation in machine vision. SIAM Journal on Control and Optimization, 33(5):1530–1559, September 1995. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24685.

Gahinet:1997:NRC

[GL97a] Pascal Gahinet and Alan J. Laub. Numerically reliable computation of optimal performance in singular H_{∞} control. SIAM Journal on Control and Optimization, 35(5): 1690–1710, September 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26995.

Glusing-Luerssen:1997:BAD

[GL97b] Heide Glüsing-Lüerssen. A behavioral approach to delaydifferential systems. SIAM Journal on Control and Optimization, 35(2):480–499, March 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/28186.

Graves:1997:AEA

[GL97c] Todd L. Graves and Tze Leung Lai. Asymptotically efficient adaptive choice of control laws in controlled Markov chains. SIAM Journal on Control and Optimization, 35(3):715–743, May 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27544.

Gozzi:1999:RMT

[GL99] Fausto Gozzi and Paola Loreti.

Regularity of the minimum time function and minimum energy problems: The linear case. SIAM Journal on Control and Optimization, 37(4):1195–1221, July 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31276.

Gray:2005:GSI

[GL05] W. Steven Gray and Yaqin Li. Generating series for interconnected analytic nonlinear systems. SIAM Journal on Control and Optimization, 44(2):646–672, March 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43007.

Ghrist:2006:NCP

[GL06] Robert Ghrist and Steven M.

Lavalle. Nonpositive curvature and Pareto optimal coordination of robots. SIAM Journal on Control and Optimization, 45(5):1697–1713, ???? 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gu:2011:EBC

[GL11] Qilong Gu and Tatsien Li. Exact boundary controllability for quasilinear hyperbolic systems on a tree-like network and its applications. SIAM Journal on Control and Optimization, 49(4):1404– 1421, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicon/resource/
                                        1/sjcodc/v49/i4/p1404_s1.

Ghosh:2013:OFT

[GL13] Supratim Ghosh and Ji-Woong Lee. Optimal finitetime consensus on fixed and periodic graphs. SIAM Journal on Control and Optimization, 51(5):3946–3970, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Guo:2019:RSD

[GL19] Xianping Guo and Zhong-Wei Liao. Risk-sensitive discounted continuous-time Markov decision processes with unbounded rates. SIAM Journal on Control and Optimization, 57(6):3857–3883, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gan:2022:DOE

[GL22a] Die Gan and Zhixin Liu. Distributed order estimation of ARX model under cooperative excitation condition. SIAM Journal on Control and Optimization, 60(3):1519–1545, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1421362.

Gozzi:2022:SME

[GL22b] Fausto Gozzi and Marta Leocata. A stochastic model of economic growth in timespace. SIAM Journal on Control and Optimization, 60(2):620–651, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1414206.

Glass:2005:ASS

[Gla05] Olivier Glass. Asymptotic stabilizability by stationary feedback of the two-dimensional Euler equation: The multiconnected case. SIAM Journal on Control and Optimization, 44(3):1105–1147, May 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43115.

Glover:1991:CAS

[GLD^{+}91] K. Glover, D. J. N. Limebeer, J. C. Doyle, E. M. Kasenally, and M. G. Safonov. A characterization of all solutions to the four block general distance problem. SIAM Journal on Control and Optimization, 29 (2):283–324, March 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Glizer:2012:SSO

[Gli12] Valery Y. Glizer. Stochastic singular optimal control problem with state delays: Regularization, singular perturbation, and minimizing sequence. SIAM Journal on Control and Optimization, 50 (5):2862–2888, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gielen:2012:LMT

[GLK12] R. H. Gielen, M. Lazar, and I. V. Kolmanovsky. Lyapunov methods for timeinvariant delay difference inclusions. SIAM Journal on Control and Optimization, 50 (1):110–132, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicon/resource/
                                        1/sjcodc/v50/i1/p110_s1.

Ghaffour:2022:RTO

[GLK22] Lilia Ghaffour and Taous-Meriem Laleg-Kirati. Reference tracking and observer design for space fractional partial differential equation modeling gas pressures in fractured media. SIAM Journal on Control and Optimization, 60(3):1613–1641, ????

2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1424810.

Green:2019:BOV

[GLM19] Walton Green, Shitao Liu, and Mishko Mitkovski. Boundary observability for the viscoelastic wave equation. SIAM Journal on Control and Optimization, 57(3):1629–1645, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gagnon:2021:FTR

[GLM21] Ludovick Gagnon, Pierre Lissy, and Swann Marx. A Fredholm transformation for the rapid stabilization of a degenerate parabolic equation. SIAM Journal on Control and Optimization, 59(5): 3828–3859, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Grubisic:2023:OCP

[GLNT23] Luka Grubisić, Martin Lazar, Ivica Nakić, and Martin Tautenhahn. Optimal control of parabolic equations — a spectral calculus based approach. SIAM Journal on Control and Optimization, 61 (5):2802–2825, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1449762.

Guiver:2019:IDL

[GLO19] Chris Guiver, Hartmut Logemann, and Mark R. Opmeer. Infinite-dimensional Lur’e systems: Input-to-state stability and convergence properties. SIAM Journal on Control and Optimization, 57(1): 334–365, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gohberg:1986:HSS

[GLR86] I. Gohberg, P. Lancaster, and L. Rodman. On Hermitian solutions of the symmetric algebraic Riccati equation. SIAM Journal on Control and Optimization, 24(6): 1323–1334, November 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gugat:2005:OBC

[GLS05] M. Gugat, G. Leugering, and G. Sklyar. L^{p}-optimal boundary control for the wave equation. SIAM Journal on Control and Optimization, 44(1):49–74, January 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/41921.

Garcke:2021:SOC

[GLS21] Harald Garcke, Kei Fong Lam, and Andrea Signori. Sparse optimal control of a phase field tumor model with mechanical effects. SIAM Journal on Control and Optimization, 59(2): 1555–1580, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Grasselli:1995:SEP

[GLT95] Osvaldo M. Grasselli, Sauro Longhi, and Antonio Tornambè. System equivalence for periodic models and systems. SIAM Journal on Control and Optimization, 33(2):455–468, March 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/23457.

Gao:2018:NMO

[GLX18] Yuan Gao, Jin Liang, and Ti-Jun Xiao. A new method to obtain uniform decay rates for multidimensional wave equations with nonlinear acoustic boundary conditions. SIAM Journal on Control and Optimization, 56(2):1303–1320, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gao:2004:CND

[GLZ04] Hang Gao, Peidong Lei, and Bo Zhang. A class of nonlinear degenerate integrodifferential control systems. SIAM Journal on Control and Optimization, 43(3):986–1010, May 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/41767.

Gaudioso:1991:QAC

[GM91a] Manlio Gaudioso and Maria Flavia Monaco. Quadratic approximations in convex nondifferentiable optimization. SIAM Journal on Control and Optimization, 29(1):58–70, January 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gelfand:1991:RSA

[GM91b] Saul B. Gelfand and Sanjoy K. Mitter. Recursive stochastic algorithms for global optimization in R^{d}. SIAM Journal on Control and Optimization, 29(5):999–1018, September 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gelfand:1993:MTA

[GM93] Saul B. Gelfand and Sanjoy K. Mitter. Metropolis-type annealing algorithms for global optimization in R^{d}. SIAM Journal on Control and Optimization, 31(1):111–131, January 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gunzburger:1999:VTP

[GM99] M. D. Gunzburger and S. Manservisi. The velocity tracking problem for Navier–Stokes flows with bounded distributed controls. SIAM Journal on Control and Optimization, 37(6): 1913–1945, November 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33740.

Gunzburger:2000:VTP

[GM00] M. D. Gunzburger and S. Manservisi.

The velocity tracking problem for Navier–Stokes flows with boundary control. SIAM Journal on Control and Optimization, 39(2):594–634, March 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35377.

Gobet:2005:SAU

[GM05] Emmanuel Gobet and Remi Munos. Sensitivity analysis using Itô–Malliavin calculus and martingales, and application to stochastic optimal control. SIAM Journal on Control and Optimization, 43(5):1676–1713, September 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41905.

Guatteri:2009:IHE

[GM09] Giuseppina Guatteri and Federica Masiero. Infinite horizon and ergodic optimal quadratic control for an affine equation with stochastic coefficients. SIAM Journal on Control and Optimization, 48(3): 1600–1631, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gumussoy:2011:FOI

[GM11] Suat Gumussoy and Wim Michiels. Fixed-order Hinfinity control for interconnected systems using delay differential algebraic equations. SIAM Journal on Control and Optimization, 49(5): 2212–2238, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicon/resource/
                                        1/sjcodc/v49/i5/p2212_s1.

Guatteri:2013:EOC

[GM13] Giuseppina Guatteri and Federica Masiero. On the existence of optimal controls for SPDEs with boundary noise and boundary control. SIAM Journal on Control and Optimization, 51(3):1909–1939, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gozzi:2017:SOCa

[GM17a] Fausto Gozzi and Federica Masiero. Stochastic optimal control with delay in the control I: Solving the HJB equation through partial smoothing. SIAM Journal on Control and Optimization, 55(5): 2981–3012, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See errata [GM21].

Gozzi:2017:SOCb

[GM17b] Fausto Gozzi and Federica Masiero. Stochastic optimal control with delay in the control II: Verification theorem and optimal feedbacks. SIAM Journal on Control and Optimization, 55(5): 3013–3038, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See errata [GM21].

Gozzi:2021:ESO

[GM21] Fausto Gozzi and Federica Masiero. Errata: Stochastic optimal control with delay in the control I: Solving the HJB equation through partial smoothing, and Stochastic optimal control with delay in the control II: Verification theorem and optimal feedbacks. SIAM Journal on Control and Optimization, 59(4): 3096–3101, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [GM17a, GM17b].

Gao:2010:SNC

[GMCL10] Huijun Gao, Xiangyu Meng, Tongwen Chen, and James Lam. Stabilization of networked control systems via dynamic output-feedback controllers. SIAM Journal on Control and Optimization, 48 (5):3643–3658, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gilliam:1991:ICM

[GMM91] D. S. Gilliam, B. A. Mair, and C. F. Martin. An inverse convolution method for regular parabolic equations. SIAM Journal on Control and Optimization, 29(1):71–88, January 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Garcke:2022:PFA

[GMN22] Harald Garcke, Sourav Mitra, and Vanja Nikolić. A phase-field approach to shape and topology optimization of acoustic waves in dissipative media. SIAM Journal on Control and Optimization, 60 (4):2297–2319, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1449294.

Germani:2000:TSA

[GMP00] A. Germani, C. Manes, and P. Pepe. A twofold spline approximation for finite horizon LQG control of hereditary systems. SIAM Journal on Control and Optimization, 39(4):1233–1295, July 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33746.

Gudoshnikov:2022:FTS

[GMR22] Ivan Gudoshnikov, Oleg Makarenkov, and Dmitrii Rachinskii. Finitetime stability of polyhedral sweeping processes with application to elastoplastic systems. SIAM Journal on Control and Optimization, 60 (3):1320–1346, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1388796.

Gudi:2022:FEA

[GMS22] Thirupathi Gudi, Gouranga Mallik, and Ramesh Ch. Sau. Finite element analysis of a constrained Dirichlet boundary control problem governed by a linear parabolic equation. SIAM Journal on Control and Optimization, 60 (6):3262–3288, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M145361X.

Guo:2016:ASE

[GMY16] Qian Guo, Xuerong Mao, and Rongxian Yue. Almost sure exponential stability of stochastic differential delay equations. SIAM Journal on Control and Optimization, 54(4):1919–1933, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gaitsgory:2002:MSP

[GN02] Vladimir Gaitsgory and Minh-Tuan Nguyen. Multiscale singularly perturbed control systems: Limit occupational measures sets and averaging. SIAM Journal on Control and Optimization, 41(3):954–974, May 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/39305.

Grune:2003:OBS

[GN03] Lars Grüne and Dragan Nesić. Optimization-based stabilization of sampled-data nonlinear systems via their approximate discrete-time models. SIAM Journal on Control and Optimization, 42(1):98–122, January 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/40258.

Gayduk:2018:EFL

[GN18] Roman Gayduk and Sergey Nadtochiy. Endogenous formation of limit order books: Dynamics between trades. SIAM Journal on Control and Optimization, 56(3):1577– 1619, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gupta:2014:CIB

[GNLB14] Abhishek Gupta, Ashutosh Nayyar, Cédric Langbort, and Tamer Basar. Common information based Markov perfect equilibria for linear-Gaussian games with asymmetric information. SIAM Journal on Control and Optimization, 52 (5):3228–3260, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Garone:2018:SAS

[GNP18] Emanuele Garone, Lorenzo Ntogramatzidis, and Fabrizio Padula. A structural approach to state-to-output decoupling. SIAM Journal on Control and Optimization, 56 (5):3816–3847, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gomes:2004:CEH

[GO04] Diogo A. Gomes and Adam M.

Oberman. Computing the effective Hamiltonian using a variational approach. SIAM Journal on Control and Optimization, 43(3):792–812, May 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41762.

Goncharova:2011:ASS

[GO11] Elena Goncharova and Alexander Ovseevich. Asymptotics for shapes of singularly perturbed reachable sets. SIAM Journal on Control and Optimization, 49(2):403– 419, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i2/p403_s1.

Guiver:2014:MRB

[GO14] Chris Guiver and Mark R. Opmeer. Model reduction by balanced truncation for systems with nuclear Hankel operators. SIAM Journal on Control and Optimization, 52(2): 1366–1401, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Goebel:2002:CZS

[Goe02]  Rafal Goebel.  Convex-

ity in zero-sum differential games. SIAM Journal on Control and Optimization, 40(5): 1491–1504, September 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36073.

Goebel:2005:COC

[Goe05] Rafal Goebel. Convex optimal control problems with smooth Hamiltonians. SIAM Journal on Control and Optimization, 43(5):1787–1811, September 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41158.

Goebel:2013:LFD

[Goe13] Rafal Goebel. Lyapunov functions and duality for convex processes. SIAM Journal on Control and Optimization, 51 (4):3332–3350, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gombani:1989:CAL

[Gom89] Andrea Gombani. Consistent approximations of linear stochastic models. SIAM Journal on Control and Optimization, 27(1):83–107, January 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gomoyunov:2020:DPP

[Gom20] Mikhail I. Gomoyunov. Dynamic programming principle and Hamilton–Jacobi– Bellman equations for fractionalorder systems. SIAM Journal on Control and Optimization, 58(6):3185–3211, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Goncalves:1987:RTH

[Gon87a] J. Basto Gonçalves. Realization theory for Hamiltonian systems. SIAM Journal on Control and Optimization, 25 (1):63–73, January 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Goncalves:1987:SCL

[Gon87b] J. Basto Gonçalves. Sufficient conditions for local controllability with unbounded controls. SIAM Journal on Control and Optimization, 25(6): 1371–1378, November 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

GonzalezdePaz:1994:RAA

[Gon94] R. B. González de Paz. A relaxation approach applied to domain optimization. SIAM Journal on Control and Optimization, 32(1):154–169, January 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/18223.

Goor:1976:EOC

[Goo76] Robert M. Goor. Existence of an optimal control for systems with jump Markov disturbances. SIAM Journal on Control and Optimization, 14(5):899–918, August 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Goreac:2015:ACC

[Gor15] Dan Goreac. Asymptotic control for a class of piecewise deterministic Markov processes associated to temperate viruses. SIAM Journal on Control and Optimization, 53(4):1860–1891, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Goreac:2019:BAN

[Gor19] Dan Goreac. Border avoidance: Necessary regularity for coefficients and viscosity approach. SIAM Journal on Control and Optimization, 57 (6):4175–4204, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gowda:1989:CPL

[Gow89] M. Seetharama Gowda. Complementarity problems over locally compact cones. SIAM Journal on Control and Optimization, 27(4):836–841, July 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gozzi:1991:SRO

[Goz91] Fausto Gozzi. Some results for an optimal control problem with semilinear state equation. SIAM Journal on Control and Optimization, 29(4):751–768, July 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gonzaga:1979:CDS

[GP79] C. Gonzaga and E. Polak. On constraint dropping schemes and optimality functions for a class of outer approximations algorithms. SIAM Journal on Control and Optimization, 17 (4):477–493, July 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gutman:1982:PMM

[GP82] S. Gutman and Z. Palmor. Properties of min-max controllers in uncertain dynamical systems. SIAM Journal on Control and Optimization, 20(6):850–861, November 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gowda:1994:BSS

[GP94] M. Seetharama Gowda and Jong-Shi Pang. On the boundedness and stability of solutions to the affine variational inequality problem. SIAM Journal on Control and Optimization, 32(2):421–441, March 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/22888.

Gegout-Petit:1998:AFC

[GP98] A. Gegout-Petit. Approximate filter for the conditional law of a partially observed process in nonlinear filtering. SIAM Journal on Control and Optimization, 36(4):1423–1447, July 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28704.

Gutman:2002:WCD

[GP02] Shaul Gutman and Eytan Paldi. Worst case design for robust compensation. SIAM Journal on Control and Optimization, 41(1):278–302, January 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36357.

Garavello:2005:HNP

[GP05] Mauro Garavello and Benedetto Piccoli. Hybrid necessary principle. SIAM Journal on Control and Optimization, 43(5): 1867–1887, September 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41621.

Grivopoulos:2017:LQS

[GP17] Symeon Grivopoulos and Ian Petersen. Linear quantum system transfer function realization using static networks for input/output processing and feedback. SIAM Journal on Control and Optimization, 55 (5):3349–3369, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Giarre:2021:ECN

[GP21] Laura Giarré and Raffaele Pesenti. Erratum: Consensus for networks with unknown but bounded disturbances. SIAM Journal on Control and Optimization, 59(2):1757, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gao:2022:CTC

[GP22] Bolin Gao and Lacra Pavel. Continuous-time convergence rates in potential and monotone games. SIAM Journal on Control and Optimization, 60(3):1712–1731, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1381873.

Geromel:1991:CPS

[GPB91] J. C. Geromel, P. L. D. Peres, and J. Bernussou. On a convex parameter space method for linear control design of uncertain systems. SIAM Journal on Control and Optimization, 29(2):381–402, March 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Geromel:1995:CAM

[GPS95] J. C. Geromel, P. L. D. Peres, and S. R. Souza. A convex approach to the mixed H_{∈}/H_{∞} control problem for discrete-time uncertain systems. SIAM Journal on Control and Optimization, 33(6): 1816–1833, November 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/23823.

Grune:2010:AUN

[GPSW10] Lars Grüne, Jürgen Pannek, Martin Seehafer, and Karl Worthmann. Analysis of unconstrained nonlinear MPC schemes with time varying control horizon. SIAM Journal on Control and Optimization, 48(8):4938–4962, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gaitsgory:2009:LPA

[GQ09] Vladimir Gaitsgory and Marc Quincampoix. Linear programming approach to deterministic infinite horizon optimal control problems with discounting. SIAM Journal on Control and Optimization, 48(4):2480–2512, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gudovich:2011:OCH

[GQ11] Anastasia Gudovich and Marc Quincampoix. Optimal control with hysteresis nonlinearity and multidimensional play operator. SIAM Journal on Control and Optimization, 49 (2):788–807, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.

                                        siam.org/sicon/resource/
                                        1/sjcodc/v49/i2/p788_s1.

Gonzalez:1985:DCPa

[GR85a] R. Gonzalez and E. Rofman. On deterministic control problems: an approximation procedure for the optimal cost. I. The stationary problem. SIAM Journal on Control and Optimization, 23(2): 242–266, March 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gonzalez:1985:DCPb

[GR85b] R. Gonzalez and E. Rofman. On deterministic control problems: an approximation procedure for the optimal cost. II. The nonstationary case. SIAM Journal on Control and Optimization, 23 (2):267–285, March 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gibson:1988:NAI

[GR88a] J. S. Gibson and I. G. Rosen. Numerical approximation for the infinitedimensional discrete-time optimal linear-quadratic regulator problem. SIAM Journal on Control and Optimization, 26 (2):428–451, March 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Goodall:1988:FCD

[GR88b] D. P. Goodall and E. P.

Ryan. Feedback controlled differential inclusions and stabilization of uncertain dynamical systems. SIAM Journal on Control and Optimization, 26(6):1431–1441, November 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ghulchak:2002:DCO

[GR02] Andrey Ghulchak and Anders Rantzer. Duality in H ∞ cone optimization. SIAM Journal on Control and Optimization, 41(1):253–277, January 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36961.

Ghosh:2005:DGE

[GR05] Mrinal K. Ghosh and K. S. Mallikarjuna Rao. Differential games with ergodic payoff. SIAM Journal on Control and Optimization, 43(6): 2020–2035, November 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40451.

Gaitsgory:2006:LPA

[GR06] Vladimir Gaitsgory and Sergey Rossomakhine. Linear programming approach to deterministic long run average problems of optimal control. SIAM Journal on Control and Optimization, 44(6):2006– 2037, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gaitsgory:2017:NOC

[GR17a] Vladimir Gaitsgory and Sergey Rossomakhine. On near optimal control of systems with slow observables. SIAM Journal on Control and Optimization, 55(3):1398–1428, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gensbittel:2017:PRV

[GR17b] Fabien Gensbittel and Catherine Rainer. A probabilistic representation for the value of zero-sum differential games with incomplete information on both sides. SIAM Journal on Control and Optimization, 55(2):693–723, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Grussler:2021:SOC

[GR21a] Christian Grussler and Anders Rantzer. On second-order cone positive systems. SIAM Journal on Control and Optimization, 59(4):2717–2739, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Guzman:2021:NCS

[GR21b] Patricio Guzman and Lionel Rosier. Null controllability of the structurally damped wave equation on the twodimensional torus. SIAM Journal on Control and Optimization, 59(1):131–155, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Grasse:1981:PNC

[Gra81] Kevin A. Grasse. Perturbations of nonlinear controllable systems. SIAM Journal on Control and Optimization, 19 (2):203–220, March 1981. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Grasse:1982:NPC

[Gra82] Kevin A. Grasse. Nonlinear perturbations of controlsemilinear control systems. SIAM Journal on Control and Optimization, 20(3):311–327, May 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Grasse:1988:SCF

[Gra88]  Kevin A. Grasse.  Suffi-

cient conditions for the functional reproducibility of timevarying, input-output systems. SIAM Journal on Control and Optimization, 26(1): 230–249, January 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Grammel:1999:MPH

[Gra99] G. Grammel. Maximum principle for a hybrid system via singular perturbations. SIAM Journal on Control and Optimization, 37(4):1162–1175, July 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/33264.

Grammel:2005:ESN

[Gra05] G. Grammel. Exponential stability of nonlinear singularly perturbed differential equations. SIAM Journal on Control and Optimization, 44(5): 1712–1724, September 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/61459.

Grasse:2007:SBN

[Gra07] Kevin A. Grasse. Simulation and bisimulation of nonlinear control systems with admissible classes of inputs and disturbances. SIAM Journal on Control and Optimization, 46 (2):562–584, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Grammel:2008:ACP

[Gra08] G. Grammel. Accessibility and controllability in the presence of fast oscillations. SIAM Journal on Control and Optimization, 47(4):1879–1887, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Goodwin:1981:DTS

[GRC81] Graham C. Goodwin, Peter J. Ramadge, and Peter E. Caines. Discrete time stochastic adaptive control. SIAM Journal on Control and Optimization, 19(6): 829–853, November 1981. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See corrigendum [GRC82].

Goodwin:1982:CDT

[GRC82] Graham C. Goodwin, Peter J. Ramadge, and Peter E. Caines. Corrigendum: “Discrete time stochastic adaptive control” [SIAM J. Control Optim. 19 (1981), no. 6, 829–853; MR 83d:93070]. SIAM Journal on Control and Optimization, 20(6):893, November 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [GRC81].

Green:1992:CSL

[Gre92] Michael Green. H_{∞} controller synthesis by J-lossless coprime factorization. SIAM Journal on Control and Optimization, 30(3):522–547, May 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gripenberg:1983:CSO

[Gri83] Gustaf Gripenberg. The construction of the solution of an optimal control problem described by a Volterra integral equation. SIAM Journal on Control and Optimization, 21 (4):582–597, July 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Grinold:1986:IHS

[Gri86] Richard C. Grinold. Infinite horizon stochastic programs. SIAM Journal on Control and Optimization, 24(6): 1246–1260, November 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Grimm:1988:RCI

[Gri88] José Grimm. Realization and canonicity for implicit systems. SIAM Journal on Control and Optimization, 26(6): 1331–1347, November 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Grizzle:1993:LAF

[Gri93] J. W. Grizzle. A linear algebraic framework for the analysis of discrete-time nonlinear systems. SIAM Journal on Control and Optimization, 31 (4):1026–1044, July 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Grong:2012:CRT

[Gro12] Erlend Grong. Controllability of rolling without twisting or slipping in higher dimensions. SIAM Journal on Control and Optimization, 50(4): 2462–2485, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Grong:2016:SHS

[Gro16] Erlend Grong. Submersions, Hamiltonian systems, and optimal solutions to the rolling manifolds problem. SIAM Journal on Control and Optimization, 54(2):536–566, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Grochowski:2021:CCS

[Gro21] Marek Grochowski. Causality and control systems. SIAM Journal on Control and Optimization, 59(1):320–338, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gajardo:2009:MTS

[GRR09] P. Gajardo, H. Ramírez C., and A. Rapaport. Minimal time sequential batch reactors with bounded and impulse controls for one or more species. SIAM Journal on Control and Optimization, 47 (6):2827–2856, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gozzi:2000:SOH

[GRS00] Fausto Gozzi, Elisabeth Rouy, and Andrzej Swiech. Second order Hamilton–Jacobi equations in Hilbert spaces and stochastic boundary control. SIAM Journal on Control and Optimization, 38(2):400–430, March 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32490.

Gibson:1992:ACT

[GRT92] J. S. Gibson, I. G. Rosen, and G. Tao. Approximation in control of thermoelastic systems. SIAM Journal on Control and Optimization, 30(5): 1163–1189, September 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Grune:1996:NSB

[Grü96] Lars Grüne. Numerical stabilization of bilinear control systems. SIAM Journal on Control and Optimization, 34(6): 2024–2050, November 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/27229.

Grune:1998:ACE

[Grü98]  Lars Grüne.  Asymptotic

controllability and exponential stabilization of nonlinear control systems at singular points. SIAM Journal on Control and Optimization, 36(5):1485–1503, September 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31591.

Grune:2000:HSF

[Grü00] Lars Grüne. Homogeneous state feedback stabilization of homogeneous systems. SIAM Journal on Control and Optimization, 38(4):1288–1308, July 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34930.

Grune:2009:ADU

[Grü09] Lars Grüne. Analysis and design of unconstrained nonlinear MPC schemes for finite and infinite dimensional systems. SIAM Journal on Control and Optimization, 48(2): 1206–1228, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Grun:2013:DGI

[Grü13] Christine Grün. On Dynkin games with incomplete information. SIAM Journal on Control and Optimization, 51 (5):4039–4065, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Giegrich:2024:CPG

[GRZ24] Michael Giegrich, Christoph Reisinger, and Yufei Zhang. Convergence of policy gradient methods for finite-horizon exploratory linear-quadratic control problems. SIAM Journal on Control and Optimization, 62(2):1060–1092, March 2024. CODEN SJCODC. ISSN 1095-7138.

Gallagher:1993:TGT

[GS93] Richard J. Gallagher and Ossama A. Saleh. Two generalizations of a theorem of Arrow, Barankin, and Blackwell. SIAM Journal on Control and Optimization, 31(1): 247–256, January 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gatarek:1994:EOC

[GS94] Dariusz Gatarek and Jarosław Sobczyk. On the existence of optimal controls of Hilbert space-valued diffusions. SIAM Journal on Control and Optimization, 32(1):170–175, January 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/22626.

Grune:2011:DGZ

[GS11] Lars Grüne and Oana Silvia Serea. Differential games and Zubov’s method. SIAM Journal on Control and Optimization, 49(6):2349–2377, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sicon/
resource/1/sjcodc/v49/i6/
p2349_s1.

Gottschalk:2014:ORD

[GS14] Hanno Gottschalk and Sebastian Schmitz. Optimal reliability in design for fatigue life. SIAM Journal on Control and Optimization, 52(5): 2727–2752, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Goebel:2018:PAS

[GS18a] Rafal Goebel and Ricardo G.

Sanfelice. Pointwise asymptotic stability in a hybrid system and well-posed behavior beyond Zeno. SIAM Journal on Control and Optimization, 56(2):1358–1385, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gugat:2018:BFS

[GS18b] Martin Gugat and Rüdiger Schultz. Boundary feedback stabilization of the isothermal Euler equations with uncertain boundary data. SIAM Journal on Control and Optimization, 56(2):1491–1507, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gazor:2019:BCD

[GS19] Majid Gazor and Nasrin Sadri. Bifurcation controller designs for the generalized cusp plants of Bogdanov–Takens singularity with an application to ship control. SIAM Journal on Control and Optimization, 57 (3):2122–2151, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gangl:2022:TDP

[GS22] Peter Gangl and Kevin Sturm. Topological derivative for PDEs on surfaces. SIAM Journal on Control and Optimization, 60(1):81–103, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1339040.

Grune:2019:SAO

[GSS19] Lars Grüne, Manuel Schaller, and Anton Schiela. Sensitivity analysis of optimal control for a class of parabolic PDEs motivated by model predictive control. SIAM Journal on Control and Optimization, 57(4):2753–2774, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gallaun:2020:SCS

[GST20] Dennis Gallaun, Christian Seifert, and Martin Tautenhahn. Sufficient criteria and sharp geometric conditions for observability in Banach spaces. SIAM Journal on Control and Optimization, 58(4): 2639–2657, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gardner:1989:GIA

[GSW89] Robert B. Gardner, William F.

Shadwick, and George R. Wilkens. A geometric isomorphism with applications to closed loop controls. SIAM Journal on Control and Optimization, 27(6):1361–1368, November 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gozzi:2005:CPS

[GŠZ05] Fausto Gozzi, Andrzej Šwiech, and Xun Yu Zhou. A corrected proof of the stochastic verification theorem within the framework of viscosity solutions. SIAM Journal on Control and Optimization, 43(6): 2009–2019, November 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42818. See er-

ratum [GŠZ10].

Gozzi:2010:ECP

[GŠZ10] Fausto Gozzi, Andrzej Šwiech, and Xun Yu Zhou. Erratum: “A corrected proof of the stochastic verification theorem within the framework of viscosity solutions”. SIAM Journal on Control and Optimization, 48(6):4177–4179, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [GŠZ05].

Gu:2020:CUD

[GSZ20] Jia-Wen Gu, Shijing Si, and Harry Zheng. Constrained utility deviation-risk optimization and time-consistent HJB equation. SIAM Journal on Control and Optimization, 58 (2):866–894, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gauvin:1977:DSN

[GT77] Jacques Gauvin and Jon W. Tolle. Differential stability in nonlinear programming. SIAM Journal on Control and Optimization, 15(2):294– 311, February 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gowda:1990:CCQ

[GT90] M. Seetharama Gowda and Marc Teboulle. A comparison of constraint qualifications in infinite-dimensional convex programming. SIAM Journal on Control and Optimization, 28(4):925–935, July 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Goldberg:1993:SOS

[GT93] H. Goldberg and F. Tröltzsch.

Second-order sufficient optimality conditions for a class of nonlinear parabolic boundary control problems. SIAM Journal on Control and Optimization, 31(4):1007–1025, July 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Guatteri:2005:BSR

[GT05] Giuseppina Guatteri and Gianmario Tessitore. On the backward stochastic Riccati equation in infinite dimensions. SIAM Journal on Control and Optimization, 44 (1):159–194, January 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42550.

Garvie:2007:OCN

[GT07] Marcus R. Garvie and Catalin Trenchea. Optimal control of a nutrient-phytoplankton-zooplankton-fish system. SIAM Journal on Control and Optimization, 46(3):775–791, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Guo:2008:CBS

[GT08] Xin Guo and Pascal Tomecek.

Connections between singular control and optimal switching. SIAM Journal on Control and Optimization, 47(1):

421–443, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Guo:2009:CSC

[GT09] Xin Guo and Pascal Tomecek. A class of singular control problems and the smooth fit principle. SIAM Journal on Control and Optimization, 47(6):3076–3099, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Guatteri:2014:WPO

[GT14] Giuseppina Guatteri and Gianmario Tessitore. Well posedness of operator valued backward stochastic Riccati equations in infinite dimensional spaces. SIAM Journal on Control and Optimization, 52(6): 3776–3806, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gal:2017:CIW

[GT17] Ciprian G. Gal and Louis Tebou. Carleman inequalities for wave equations with oscillatory boundary conditions and application. SIAM Journal on Control and Optimization, 55(1):324–364, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Guatteri:2020:EBM

[GT20] Giuseppina Guatteri and Gianmario Tessitore. Ergodic BSDEs with multiplicative and degenerate noise. SIAM Journal on Control and Optimization, 58(4):2050–2077, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Giannetti:2022:LRM

[GT22a] F. Giannetti and G. Treu. On the Lipschitz regularity for minima of functionals depending on x, u, and ∇u under the bounded slope condition. SIAM Journal on Control and Optimization, 60 (3):1347–1364, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1396617.

Guatteri:2022:SLT

[GT22b] Giuseppina Guatteri and Gianmario Tessitore. Singular limit of two-scale stochastic optimal control problems in infinite dimensions by vanishing noise regularization. SIAM Journal on Control and Optimization, 60(1):575–596, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1408488.

Gonzalez-Trejo:2002:MCD

[GTHLHR02] J. I. González-Trejo, O. Hernández-Lerma, and L. F. Hoyos-Reyes. Minimax control of discrete-time stochastic systems. SIAM Journal on Control and Optimization, 41(5): 1626–1659, September 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38383.

Gu:2021:KFR

[GTQ21] Guoxiang Gu, Yang Tang, and Feng Qian. Kalman filtering over the random delay and packet drop channel. SIAM Journal on Control and Optimization, 59(4): 2454–2476, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gorban:2013:LLC

[GTSN13] A. N. Gorban, I. Tyukin, E. Steur, and H. Nijmeijer. Lyapunov-like conditions of forward invariance and boundedness for a class of unstable systems. SIAM Journal on Control and Optimization, 51(3):2306–2334, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Guo:2016:MPO

[GTT16a] Gaoyue Guo, Xiaolu Tan, and Nizar Touzi. On the monotonicity principle of optimal Skorokhod embedding problem. SIAM Journal on Control and Optimization, 54(5): 2478–2489, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Guo:2016:OSEa

[GTT16b] Gaoyue Guo, Xiaolu Tan, and Nizar Touzi. Optimal Skorokhod embedding under finitely many marginal constraints. SIAM Journal on Control and Optimization, 54 (4):2174–2201, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Guo:2022:CSG

[GTX22] Xin Guo, Wenpin Tang, and Renyuan Xu. A class of stochastic games and moving free boundary problems. SIAM Journal on Control and Optimization, 60(2):758–785, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1322558.

Gu:1990:ELO

[Gu90] Guoxiang Gu. On the existence of linear optimal control with output feedback. SIAM Journal on Control and Optimization, 28(3):711–719, May 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Guesmia:2003:NAS

[Gue03]  Aissa Guesmia.  A new

approach of stabilization of nondissipative distributed systems. SIAM Journal on Control and Optimization, 42(1):24–52, January 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39497.

Guerrero:2007:NCS

[Gue07] Sergio Guerrero. Null controllability of some systems of two parabolic equations with one control force. SIAM Journal on Control and Optimization, 46(2):379–394, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gueye:2014:EBC

[Gue14]  Mamadou Gueye.  Exact

boundary controllability of 1-D parabolic and hyperbolic degenerate equations. SIAM Journal on Control and Optimization, 52(4):2037–2054, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gugat:2003:BCB

[Gug03] Martin Gugat. Boundary controllability between sub- and supercritical flow. SIAM Journal on Control and Optimization, 42(3):1056–1070, May 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40966.

Gugat:2007:OEC

[Gug07] Martin Gugat. Optimal energy control in finite time by varying the length of the string. SIAM Journal on Control and Optimization, 46(5): 1705–1725, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gugat:2010:PTS

[Gug10] Martin Gugat. Penalty techniques for state constrained optimal control problems with the wave equation. SIAM Journal on Control and Optimization, 48(5):3026–3051, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gugat:2015:ESW

[Gug15] Martin Gugat. Exponential stabilization of the wave equation by Dirichlet integral feedback. SIAM Journal on Control and Optimization, 53(1): 526–546, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Guigue:2013:SVR

[Gui13] A. Guigue. Set-valued return function and generalized solutions for multiobjective optimal control problems (MOC). SIAM Journal on Control and Optimization, 51 (3):2379–2405, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Guler:1991:CPP

[Gül91] Osman Güler. On the convergence of the proximal point algorithm for convex minimization. SIAM Journal on Control and Optimization, 29(2): 403–419, March 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gunzel:1995:TKK

[Gün95] Harald Günzel. On the topology of the Karush–Kuhn– Tucker set under Mangasarian– Fromovitz constraint qualification. SIAM Journal on Control and Optimization, 33(6): 1847–1856, November 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/25779.

Guo:1990:AST

[Guo90]  Lei Guo.  On adaptive

stabilization of time-varying stochastic systems. SIAM Journal on Control and Optimization, 28(6):1432–1451, November 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Guo:1994:FRL

[Guo94a] Lei Guo. Further results on least squares based adaptive minimum variance control. SIAM Journal on Control and Optimization, 32 (1):187–212, January 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/22960.

Guo:1994:SRS

[Guo94b] Lei Guo. Stability of recursive stochastic tracking algorithms. SIAM Journal on Control and Optimization, 32(5):1195–1225, September 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/22560.

Guo:2001:RBA

[Guo01] Bao-Zhu Guo. Riesz basis approach to the stabilization of a flexible beam with a tip mass. SIAM Journal on Control and Optimization, 39(6):1736–1747, November 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35488.

Guo:2002:RBP

[Guo02] Bao-Zhu Guo. Riesz basis property and exponential stability of controlled Euler–Bernoulli beam equations with variable coefficients. SIAM Journal on Control and Optimization, 40(6): 1905–1923, November 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37251.

Gupta:2020:ETO

[Gup20] Abhishek Gupta. Existence of team-optimal solutions in static teams with common information: a topology of information approach. SIAM Journal on Control and Optimization, 58(2):998–1021, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gutman:1990:IDP

[Gut90] S. Gutman. Identification of discontinuous parameters in flow equations. SIAM Journal on Control and Optimization, 28(5):1049–1060, September 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Geymonat:2000:NRS

[GV00] Giuseppe Geymonat and Vanda Valente. A noncontrollability result for systems of mixed order. SIAM Journal on Control and Optimization, 39(3):661–672, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34832.

Galbraith:2003:LCO

[GV03] Grant N. Galbraith and Richard B. Vinter. Lipschitz continuity of optimal controls for state constrained problems. SIAM Journal on Control and Optimization, 42(5): 1727–1744, September 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40471.

Griesse:2005:PDA

[GV05] R. Griesse and S. Volkwein.

A primal-dual active set strategy for optimal boundary control of a nonlinear reactiondiffusion system. SIAM Journal on Control and Optimization, 44(2):467–494, March 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43869.

Gomes:2016:EDM

[GV16] Diogo A. Gomes and Vardan K. Voskanyan. Extended deterministic meanfield games. SIAM Journal on Control and Optimization, 54(2):1030–1055, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ge:2012:QMB

[GVL12] Shuzhi Sam Ge, Thanh Long Vu, and Tong Heng Lee. Quantum measurement-based feedback control: a nonsmooth time delay control approach. SIAM Journal on Control and Optimization, 50(2):845–863, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Glashoff:1976:BCP

[GW76] Klaus Glashoff and Norbert Weck. Boundary control of parabolic differential equations in arbitrary dimensions: Supremum-norm problems. SIAM Journal on Control and Optimization, 14(4):662–681, July 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gates:1984:LSO

[GW84] D. J. Gates and M. Westcott. Local stability and optimality in continuous games. SIAM Journal on Control and Optimization, 22(2):181– 198, March 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gates:1996:SCV

[GW96a] D. J. Gates and M. Westcott. Solar cars and variational problems equivalent to shortest paths. SIAM Journal on Control and Optimization, 34(2):428–436, March 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26027.

Gyorfi:1996:ASA

[GW96b] László Györfi and Harro Walk.

On the averaged stochastic approximation for linear regression. SIAM Journal on Control and Optimization, 34(1):31–61, January 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/22666.

Guo:2009:SFP

[GW09] Xin Guo and Guoliang Wu.

Smooth fit principle for impulse control of multidimensional diffusion processes. SIAM Journal on Control and Optimization, 48(2):594–617, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Geiselhart:2015:MGG

[GW15] Roman Geiselhart and Fabian Wirth. On maximal gains guaranteeing a small-gain condition. SIAM Journal on Control and Optimization, 53(1): 262–286, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Geiselhart:2016:RIS

[GW16] Roman Geiselhart and Fabian R. Wirth. Relaxed ISS small-gain theorems for discrete-time systems. SIAM Journal on Control and Optimization, 54(2): 423–449, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gimperlein:2017:DOD

[GW17] Heiko Gimperlein and Alden Waters. A deterministic optimal design problem for the heat equation. SIAM Journal on Control and Optimization, 55(1):51–69, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gu:2018:SMC

[GW18] Jian-Jun Gu and Jun-Min Wang. Sliding mode control of the Orr–Sommerfeld equation cascaded by both the Squire equation and ODE in the presence of boundary disturbances. SIAM Journal on Control and Optimization, 56 (2):837–867, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gunzburger:2019:EAF

[GW19] Max Gunzburger and Jilu Wang. Error analysis of fully discrete finite element approximations to an optimal control problem governed by a time-fractional PDE. SIAM Journal on Control and Optimization, 57(1):241–263, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Guasoni:2020:SPS

[GW20] Paolo Guasoni and Gu Wang.

Sharing profits in the sharing economy. SIAM Journal on Control and Optimization, 58(6):3559–3585, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gong:2014:AEO

[GWY14] Wei Gong, Gengsheng Wang, and Ningning Yan. Approximations of elliptic optimal control problems with controls acting on a lower dimensional manifold. SIAM Journal on Control and Optimization, 52 (3):2008–2035, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Guo:2004:SCR

[GX04] Bao-Zhu Guo and Yu Xie. A sufficient condition on Riesz basis with parentheses of Non– Self-Adjoint operator and application to a serially connected string system under joint feedbacks. SIAM Journal on Control and Optimization, 43(4):1234–1252, July 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/42035.

Guo:2019:SGF

[GX19] Xin Guo and Renyuan Xu. Stochastic games for fuel follower problem: N versus mean field game. SIAM Journal on Control and Optimization, 57(1):659–692, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gao:2022:SDT

[GXZ22] Xuefeng Gao, Zuo Quan Xu, and Xun Yu Zhou. State-dependent temperature control for Langevin diffusions. SIAM Journal on Control and Optimization, 60 (3):1250–1268, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1429424.

Grasselli:1998:ISB

[GY98] Maurizio Grasselli and Masahiro Yamamoto. Identifying a spatial body force in linear elastodynamics via traction measurements. SIAM Journal on Control and Optimization, 36(4):1190–1206, July 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/30028.

Gong:2011:MFE

[GY11] Wei Gong and Ningning Yan.

Mixed finite element method for Dirichlet boundary control problem governed by elliptic PDEs. SIAM Journal on Control and Optimization, 49(3): 984–1014, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i3/p984_s1.

Guo:2012:SCC

[GY12] Bao-Zhu Guo and Dong-Hui Yang. Some compact classes of open sets under Hausdorff distance and application to shape optimization. SIAM Journal on Control and Optimization, 50(1): 222–242, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v50/i1/p222_s1.

Guo:2013:CBH

[GY13] Bao-Zhu Guo and Dong-Hui Yang. On convergence of boundary Hausdorff measure and application to a boundary shape optimization problem. SIAM Journal on Control and Optimization, 51(1): 253–272, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Guo:2016:OSEb

[GY16a] Bao-Zhu Guo and Huaiqiang Yu. Optimal state estimation for non-time invertible evolutionary systems. SIAM Journal on Control and Optimization, 54(5):2754–2786, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Guo:2016:NOC

[GY16b] Lei Guo and Jane J. Ye. Necessary optimality conditions for optimal control problems with equilibrium constraints. SIAM Journal on Control and Optimization, 54 (5):2710–2733, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Gupta:2015:EOP

[GYBL15] Abhishek Gupta, Serdar Yüksel, Tamer Basar, and Cédric Langbort. On the existence of optimal policies for a class of static and sequential dynamic teams. SIAM Journal on Control and Optimization, 53(3):1681–1712, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Goldys:2022:SPZ

[GYZ22] Beniamin Goldys, James Yang, and Zhou Zhou. Singular perturbation of zerosum linear-quadratic stochastic differential games. SIAM Journal on Control and Optimization, 60(1):48–80, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1401802.

Guo:2005:RWE

[GZ05] Bao-Zhu Guo and Xu Zhang.

The regularity of the wave equation with partial Dirichlet control and colocated observation. SIAM Journal on Control and Optimization, 44(5): 1598–1613, September 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/61070.

Guo:2009:WPS

[GZ09] Bao-Zhu Guo and Zhi-Xiong Zhang. Well-posedness of systems of linear elasticity with Dirichlet boundary control and observation. SIAM Journal on Control and Optimization, 48(4):2139–2167, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Garin:2012:MSP

[GZ12] Federica Garin and Sandro Zampieri. Mean square performance of consensus-based distributed estimation over regular geometric graphs. SIAM Journal on Control and Optimization, 50(1):306–

333, ????  2012.  CO-

DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicon/resource/
                                        1/sjcodc/v50/i1/p306_s1.

Guo:2013:CNA

[GZ13] Bao-Zhu Guo and Zhi-Liang Zhao. On convergence of the nonlinear active disturbance rejection control for MIMO systems. SIAM Journal on Control and Optimization, 51 (2):1727–1757, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Granato:2014:LSA

[GZ14a] G. Granato and H. Zidani. Level-set approach for reachability analysis of hybrid systems under lag constraints. SIAM Journal on Control and Optimization, 52(1):606–628, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Guo:2014:ADR

[GZ14b] Bao-Zhu Guo and Hua-Cheng Zhou. Active disturbance rejection control for rejecting boundary disturbance from multidimensional Kirchhoff plate via boundary control. SIAM Journal on Control and Optimization, 52(5): 2800–2830, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Guo:2019:AEF

[GZ19] Wei Guo and Hua-Cheng Zhou. Adaptive error feedback regulation problem for an Euler–Bernoulli beam equation with general unmatched boundary harmonic disturbance. SIAM Journal on Control and Optimization, 57(3): 1890–1928, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Geshkovski:2021:COD

[GZ21] Borjan Geshkovski and Enrique Zuazua. Controllability of one-dimensional viscous free boundary flows. SIAM Journal on Control and Optimization, 59(3):1830–1850, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Guo:2023:ORE

[GZ23] Bao-Zhu Guo and Ren-Xi Zhao. Output regulation for Euler–Bernoulli beam with unknown exosystem using adaptive internal model. SIAM Journal on Control and Optimization, 61 (4):2088–2113, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1501805.

Gao:2017:DML

[GZLC17] Jianjun Gao, Ke Zhou, Duan Li, and Xiren Cao. Dynamic mean-LPM and mean-CVaR portfolio optimization in continuous-time. SIAM Journal on Control and Optimization, 55(3):1377–1397, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Grandvallet:2013:NLC

[GZSAB13] B. Grandvallet, A. Zemouche, H. Souley-Ali, and M. Boutayeb. New LMI condition for observer-based H_{∞} stabilization of a class of nonlinear discrete-time systems. SIAM Journal on Control and Optimization, 51(1):784–800, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ha:1988:SDS

[Ha88] In Joong Ha. The standard decomposed system and noninteracting feedback control of nonlinear systems. SIAM Journal on Control and Optimization, 26(5):1235– 1249, September 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ha:1990:GPP

[Ha90] Cu D. Ha. A generalization of the proximal point algorithm. SIAM Journal on Control and Optimization, 28 (3):503–512, May 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Haas:1982:SSS

[Haa82] Violet B. Haas. The singular steady state linear regulator. SIAM Journal on Control and Optimization, 20(2): 247–257, March 1982. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Haussmann:1978:NST

[HAB78] U. G. Haussmann, W. J.

Anderson, and A. Boyarsky. A new stochastic time optimal control problem. SIAM Journal on Control and Optimization, 16(1):1–15, January 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Habets:1996:GST

[Hab96] Luc C. G. J. M. Habets. On the genericity of stabilizability for time-delay systems. SIAM Journal on Control and Optimization, 34(3):833–854, May 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24799.

Habbal:2002:GOP

[Hab02] Abderrahmane Habbal. Generation of optimal periodic oscillations for the control of boundary layers. SIAM Journal on Control and Optimization, 41(3):712–722, May 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/35489.

Hackbusch:1979:FSP

[Hac79] Wolfgang Hackbusch. On the fast solving of parabolic boundary control problems. SIAM Journal on Control and Optimization, 17(2):231– 244, March 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hadd:2006:EEA

[Had06] Said Hadd. An evolution equation approach to nonautonomous linear systems with state, input, and output delays. SIAM Journal on Control and Optimization, 45(1): 246–272, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hager:1979:LCC

[Hag79] William W. Hager. Lipschitz continuity for constrained processes. SIAM Journal on Control and Optimization, 17(3): 321–338, May 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hagiwara:2002:SAS

[Hag02] Tomomichi Hagiwara. Spectral analysis and singular value computations of the noncompact frequency response and compression operators in sampled-data systems. SIAM Journal on Control and Optimization, 41(5): 1350–1371, September 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39480.

Hagiwara:2005:SSS

[Hag05]  Tomomichi Hagiwara.  A

study on the spectrum of the sampled-data transfer operator with application to robust exponential stability problems. SIAM Journal on Control and Optimization, 44 (1):313–327, January 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43682.

Haimo:1986:FTC

[Hai86] V. T. Haimo. Finite time controllers. SIAM Journal on Control and Optimization, 24 (4):760–770, July 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hakavuori:2020:IGI

[Hak20]  Eero Hakavuori.  Infinite

geodesics and isometric embeddings in Carnot groups of step 2. SIAM Journal on Control and Optimization, 58(1): 447–461, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Halicka:1989:RSL

[Hal89]  Margaréta Halická.  Reg-

ular synthesis for the linear time-optimal control problem with convex control constraints. SIAM Journal on Control and Optimization, 27 (2):308–329, March 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hammer:1982:SPC

[Ham82] Jacob Hammer. On some properties of conditional moments in nonlinear filtering. SIAM Journal on Control and Optimization, 20(4):497–505, July 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hamadene:2006:MZS

[Ham06] S. Hamadène. Mixed zero-sum stochastic differential game and American game options. SIAM Journal on Control and Optimization, 45(2):496– 518, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hamaguchi:2021:TIC

[Ham21]  Yushi Hamaguchi.  Time-

inconsistent consumptioninvestment problems in incomplete markets under general discount functions. SIAM Journal on Control and Optimization, 59(3):2121–2146, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Han:1977:DVM

[Han77] Shih-Ping P. Han. Dual variable metric algorithms for constrained optimization. SIAM Journal on Control and Optimization, 15(4):546–565, July 1977. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Han:1986:HLS

[Han86] S.-P. Han. On the Hessian of Lagrangian and second order optimality conditions. SIAM Journal on Control and Optimization, 24(3):339–345, May 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hansen:1994:BCO

[Han94] Scott W. Hansen. Boundary control of a one-dimensional linear thermoelastic rod. SIAM Journal on Control and Optimization, 32(4):1052– 1074, July 1994. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/22260.

Haussmann:1978:FIP

[Hau78a] U. G. Haussmann. Functionals of Itô processes as stochastic integrals. SIAM Journal on Control and Optimization, 16 (2):252–269, March 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Haussmann:1978:SMP

[Hau78b] U. G. Haussmann. On the stochastic maximum principle. SIAM Journal on Control and Optimization, 16(2): 236–251, March 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Haussmann:1981:EAP

[Hau81a] U. G. Haussmann. Erratum: On the adjoint process for optimal control of diffusion processes. SIAM Journal on Control and Optimization, 19(5): 710, September 1981. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [Hau81b].

Haussmann:1981:APO

[Hau81b]  U. G. Haussmann.  On

the adjoint process for optimal control of diffusion processes. SIAM Journal on Control and Optimization, 19(2): 221–243, March 1981. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [Hau81a].

Haussmann:1982:EOC

[Hau82] U. G. Haussmann. On the existence of optimal controls for partially observed diffusions. SIAM Journal on Control and Optimization, 20(3):385–407, May 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Haussmann:1987:MPO

[Hau87] U. G. Haussmann. The maximum principle for optimal control of diffusions with partial information. SIAM Journal on Control and Optimization, 25(2):341–361, March 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Haussmann:1994:GSH

[Hau94] Ulrich G. Haussmann. Generalized solutions of the Hamilton–Jacobi equation of stochastic control. SIAM Journal on Control and Optimization, 32(3):728–743, May 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/22785.

Hayat:2019:BSD

[Hay19] Amaury Hayat. Boundary stability of 1-D nonlinear inhomogeneous hyperbolic systems for the C^{1} norm. SIAM Journal on Control and Optimization, 57(6):3603–3638, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hauswirth:2021:PDS

[HBD21] Adrian Hauswirth, Saverio Bolognani, and Florian Dörfler. Projected dynamical systems on irregular, non-Euclidean domains for nonlinear optimization. SIAM Journal on Control and Optimization, 59(1):635–668, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hurak:2006:MDR

[HBS06] Zdenek Hurak, Albrecht Böttcher, and Michael Sebek. Minimum distance to the range of a banded lower triangular Toeplitz operator in ℓ_{1} and application in ℓ_{1}- optimal control. SIAM Journal on Control and Optimization, 45(1):107–122, January 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hamelin:2007:DSD

[HBSW07] Frédéric Hamelin, Pierre Bernhard, A. J. Shaiju, and Éric Wajnberg. Diet selection as a differential foraging game. SIAM Journal on Control and Optimization, 46(5):1539– 1561, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hogeboom-Burr:2021:CIS

[HBY21] Ian Hogeboom-Burr and Serdar Yüksel. Comparison of information structures for zerosum games and a partial converse to Blackwell ordering in standard Borel spaces. SIAM Journal on Control and Optimization, 59(3):1781–1803, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hogeboom-Burr:2023:CPV

[HBY23] Ian Hogeboom-Burr and Serdar Yüksel. Continuity properties of value functions in information structures for zerosum and general games and stochastic teams. SIAM Journal on Control and Optimization, 61(2):398–414, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1480707.

Han:2020:SRD

[HCG20] Zhong-Jie Han, Boumediene Chentouf, and Huan Geng. Stabilization of a rotating disk-beam system with infinite memory via minimal state variable: a moment control case. SIAM Journal on Control and Optimization, 58(2): 845–865, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Huang:2005:DSC

[HCM05] Minyi Huang, Peter E. Caines, and Roland P. Malhamé. Degenerate stochastic control problems with exponential costs and weakly coupled dynamics: Viscosity solutions and a maximum principle. SIAM Journal on Control and Optimization, 44 (1):367–387, January 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/41764.

Hirschorn:1987:OTN

[HD87] R. Hirschorn and J. Davis. Output tracking for nonlinear systems with singular points. SIAM Journal on Control and Optimization, 25(3):547–557, May 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hirschorn:1988:GOT

[HD88] R. M. Hirschorn and J. H. Davis. Global output tracking for nonlinear systems. SIAM Journal on Control and Optimization, 26(6):1321– 1330, November 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hafner:2011:CTS

[HD11] Michael R. Hafner and Domitilla Del Vecchio. Computational tools for the safety control of a class of piecewise continuous systems with imperfect information on a partial order. SIAM Journal on Control and Optimization, 49(6): 2463–2493, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

                                        tronic). URL http://epubs.

siam.org/sicon/resource/
1/sjcodc/v49/i6/p2463_s1.

Heuberger:2003:OBF

[HdHVW03] Peter S. C. Heuberger, Thomas J.

de Hoog, Paul M. J. Van den Hof, and Bo Wahlberg. Orthonormal basis functions in time and frequency domain: Hambo transform theory. SIAM Journal on Control and Optimization, 42(4): 1347–1373, January 2003. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

He:1987:SCC

[He87] Zheng Xu He. State constrained control problems governed by variational inequalities. SIAM Journal on Control and Optimization, 25(5):1119– 1144, September 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Heij:1988:EMF

[Hei88] Christiaan Heij. Exact modeling of a finite time series. SIAM Journal on Control and Optimization, 26(1): 83–111, January 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Heiland:2016:DAR

[Hei16] Jan Heiland. A differentialalgebraic Riccati equation for applications in flow control. SIAM Journal on Control and Optimization, 54(2):718–739, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Helton:1976:SFA

[Hel76] J. William Helton. A spectral factorization approach to the distributed stable regular problem; the algebraic Riccati equation. SIAM Journal on Control and Optimization, 14 (4):639–661, July 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Helmke:1993:BRL

[Hel93] U. Helmke. Balanced realizations for linear systems: a variational approach. SIAM Journal on Control and Optimization, 31(1):1–15, January 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Henri:1980:NSC

[Hen80] Dominique Henri. Nonuniqueness of solutions in the calculus of variations: a geometric approach. SIAM Journal on Control and Optimization, 18(6):627–639, November 1980. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Henig:1983:VVD

[Hen83] Mordechai I. Henig. Vectorvalued dynamic programming. SIAM Journal on Control and Optimization, 21(3):490–499, May 1983. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hermes:1976:LCS

[Her76] Henry Hermes. Local controllability and sufficient conditions in singular problems. II. SIAM Journal on Control and Optimization, 14(6):1049– 1062, November 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hermes:1978:LAV

[Her78] Henry Hermes. Lie algebras of vector fields and local approximation of attainable sets. SIAM Journal on Control and Optimization, 16(5):715– 727, September 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hermes:1980:SSF

[Her80] H. Hermes. On the synthesis of a stabilizing feedback control via Lie algebraic methods. SIAM Journal on Control and Optimization, 18(4): 352–361, ???? 1980. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hermes:1982:LC

[Her82] Henry Hermes. On local controllability. SIAM Journal on Control and Optimization, 20 (2):211–220, March 1982. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hermes:1986:NAC

[Her86] Henry Hermes. Nilpotent approximations of control systems and distributions. SIAM Journal on Control and Optimization, 24(4):731–736, July 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hermes:1991:ASF

[Her91] Henry Hermes. Asymptotically stabilizing feedback controls and the nonlinear regulator problem. SIAM Journal on Control and Optimization, 29(1):185–196, January 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hermes:1996:LTL

[Her96] Henry Hermes. Large-time local controllability via homogeneous approximations. SIAM Journal on Control and Optimization, 34(4):1291–1299, July 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/26805.

Heunis:1994:RCA

[Heu94] Andrew Heunis. Rates of convergence for an adaptive filtering algorithm driven by stationary dependent data. SIAM Journal on Control and Optimization, 32(1):116–139, January 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/16328.

Heunis:2015:UMR

[Heu15] Andrew J. Heunis. Utility maximization in a regime switching model with convex portfolio constraints and margin requirements: Optimality relations and explicit solutions. SIAM Journal on Control and Optimization, 53(4): 2608–2656, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hewer:1993:ETP

[Hew93] Gary Hewer. Existence theorems for positive semidefinite and sign indefinite stabilizing solutions of H_{∞} Riccati equations. SIAM Journal on Control and Optimization, 31(1): 16–29, January 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Heymann:1976:CSF

[Hey76] Michael Heymann. Controllability subspaces and feedback simulation. SIAM Journal on Control and Optimization, 14 (4):769–789, July 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Han:2014:VAC

[HFY14] Ke Han, Terry L. Friesz, and Tao Yao. A variational approach for continuous supply chain networks. SIAM Journal on Control and Optimization, 52(1):663–686, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Haji-Ghassemi:1990:DGF

[HG90] K. Haji-Ghassemi. On differential games of fixed duration with phase coordinate restrictions on one player. SIAM Journal on Control and Optimization, 28(3):624–652, May 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hammouri:1992:GTV

[HG92] H. Hammouri and J. P. Gauthier. Global time-varying linearization up to output injection. SIAM Journal on Control and Optimization, 30(6): 1295–1310, November 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hildebrand:2002:ICO

[HG02] Roland Hildebrand and Michel Gevers. Identification for control: Optimal input design with respect to a worstcase ν-gap cost function. SIAM Journal on Control and Optimization, 41(5): 1586–1608, September 2002.

CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39986.

Hager:1976:CSP

[HH76] William W. Hager and Larry L.

Horowitz. Convergence and stability properties of the discrete Riccati operator equation and the associated optimal control and filtering problems. SIAM Journal on Control and Optimization, 14(2): 295–312, February 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hautus:1978:LFA

[HH78] M. L. J. Hautus and Michael Heymann. Linear feedback — an algebraic approach. SIAM Journal on Control and Optimization, 16(1):83–105, January 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hammer:1981:CFL

[HH81] Jacob Hammer and Michael Heymann. Causal factorization and linear feedback. SIAM Journal on Control and Optimization, 19(4):445–468, July 1981. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hammer:1983:SOL

[HH83] Jacob Hammer and Michael Heymann. Strictly observable linear systems. SIAM Journal on Control and Optimization, 21(1):1–16, January 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hebrard:2005:SPO

[HH05] Pascal Hébrard and Antoine Henrot. A spillover phenomenon in the optimal location of actuators. SIAM Journal on Control and Optimization, 44(1):349–366, January 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43624.

Hintermuller:2008:GOA

[HH08] M. Hintermüller and R. H. W. Hoppe. Goal-oriented adaptivity in control constrained optimal control of partial differential equations. SIAM Journal on Control and Optimization, 47(4):1721–1743, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hintermuller:2010:GOA

[HH10] Michael Hintermüller and Ronald H. W. Hoppe. Goaloriented adaptivity in pointwise state constrained optimal control of partial differential equations. SIAM Journal on Control and Optimization, 48(8):5468–5487, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Huang:2017:RMF

[HH17] Jianhui Huang and Minyi Huang. Robust mean field linear-quadratic-Gaussian games with unknown L^{2}-disturbance. SIAM Journal on Control and Optimization, 55(5):2811– 2840, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Haak:2019:EOD

[HH19] Bernhard H. Haak and Duc-Trung Hoang. Exact observability of a 1-dimensional wave equation on a noncylindrical domain. SIAM Journal on Control and Optimization, 57 (1):570–589, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

He:2021:CRI

[HHF21] Xin He, Rong Hu, and Ya Ping Fang. Convergence rates of inertial primal-dual dynamical methods for separable convex optimization problems. SIAM Journal on Control and Optimization, 59(5): 3278–3301, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hackl:2013:FCS

[HHI^{+}13] Christoph M. Hackl, Norman Hopfe, Achim Ilchmann, Markus Mueller, and Stephan Trenn. Funnel control for systems with relative degree two. SIAM Journal on Control and Optimization, 51(2): 965–995, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hess:2016:SAP

[HHLP16] Roxana Heß, Didier Henrion, Jean-Bernard Lasserre, and Tien Son Pham. Semidefinite approximations of the polynomial abscissa. SIAM Journal on Control and Optimization, 54(3):1633–1656, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hu:2018:LQG

[HHN18] Ying Hu, Jianhui Huang, and Tianyang Nie. Linearquadratic-Gaussian mixed mean-field games with heterogeneous input constraints. SIAM Journal on Control and Optimization, 56(4):2835– 2877, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

He:2019:OET

[HHOZ19] Xue Dong He, Sang Hu, Jan Obłój, and Xun Yu Zhou. Optimal exit time from casino gambling: Strategies of precommitted and naive gamblers. SIAM Journal on Control and Optimization, 57(3):

1845–1868, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hernandez-Hernandez:2016:ORS

[HHPY16] Daniel Hernández-Hernández, José-Luis Pérez, and Kazutoshi Yamazaki. Optimality of refraction strategies for spectrally negative Lévy processes. SIAM Journal on Control and Optimization, 54(3): 1126–1156, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hernandez-Hernandez:2018:ZSS

[HHS18] Daniel Hernández-Hernández and Mihai Sîrbu. Zero-sum stochastic differential games without the Isaacs condition: Random rules of priority and intermediate Hamiltonians. SIAM Journal on Control and Optimization, 56(3): 2095–2119, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hernandez-Hernandez:2023:RSL

[HHSS23] Daniel Hernández-Hernández and Pedro Salazar-Sánchez. Risk-sensitive LQG discounted control problems and their asymptotic behavior. SIAM Journal on Control and Optimization, 61(3):1136–1161, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1459253.

Huong:2019:IOB

[HHT19] D. C. Huong, V. T. Huynh, and H. Trinh. Integral outputs-based robust state observers design for time-delay systems. SIAM Journal on Control and Optimization, 57 (3):2214–2239, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hager:1984:DAO

[HI84] William W. Hager and George D. Ianculescu. Dual approximations in optimal control. SIAM Journal on Control and Optimization, 22(3):423–465, May 1984. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hagiwara:2018:HON

[HIK18] Tomomichi Hagiwara, Akira Inai, and Jung Hoon Kim. The L_{∞}/L_{2} Hankel operator/norm of sampled-data systems. SIAM Journal on Control and Optimization, 56(5): 3685–3707, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hinze:2005:ICL

[Hin05]  M. Hinze.  Instantaneous

closed loop control of the Navier–Stokes system. SIAM Journal on Control and Optimization, 44(2):564–583, January 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hinz:2014:OSS

[Hin14] Juri Hinz. Optimal stochastic switching under convexity assumptions. SIAM Journal on Control and Optimization, 52 (1):164–188, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hirschorn:1976:GCN

[Hir76] Ronald M. Hirschorn. Global controllability of nonlinear systems. SIAM Journal on Control and Optimization, 14 (4):700–711, July 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [Kum88].

Hirschorn:1977:ICS

[Hir77] Ronald M. Hirschorn. Invertibility of control systems on Lie groups. SIAM Journal on Control and Optimization, 15(6):1034–1049, November 1977. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hirschorn:1979:INC

[Hir79] R. M. Hirschorn. Invertibility of nonlinear control systems. SIAM Journal on Control and Optimization, 17(2): 289–297, March 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hirschorn:1981:IDD

[Hir81] R. M. Hirschorn. (A, B)- Invariant distributions and disturbance decoupling of nonlinear systems. SIAM Journal on Control and Optimization, 19(1):1–19, January 1981. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hirschorn:1986:SCN

[Hir86] Ronald M. Hirschorn. Strong controllability of nonlinear systems. SIAM Journal on Control and Optimization, 24 (2):264–275, March 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hirschorn:1990:GCL

[Hir90] R. M. Hirschorn. Global controllability of locally linearizable systems. SIAM Journal on Control and Optimization, 28(3):540–551, May 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hirschorn:2002:OTT

[Hir02] R. M. Hirschorn. Output tracking through singularities. SIAM Journal on Control and Optimization, 40(4):993–1010, July 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/35487.

Haskell:2013:SDC

[HJ13] William B. Haskell and Rahul Jain. Stochastic dominanceconstrained Markov decision processes. SIAM Journal on Control and Optimization, 51 (1):273–303, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Haskell:2015:CAA

[HJ15] William B. Haskell and Rahul Jain. A convex analytic approach to risk-aware Markov decision processes. SIAM Journal on Control and Optimization, 53(3):1569–1598, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hu:2016:SMP

[HJ16] Mingshang Hu and Shaolin Ji. Stochastic maximum principle for stochastic recursive optimal control problem under volatility ambiguity. SIAM Journal on Control and Optimization, 54(2):918–945, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

He:2021:EST

[HJ21] Xue Dong He and Zhao Li Jiang. On the equilibrium strategies for time-inconsistent problems in continuous time. SIAM Journal on Control and Optimization, 59(5):3860– 3886, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hong:2010:FTI

[HJF10] Yiguang Hong, Zhong-Ping Jiang, and Gang Feng. Finitetime input-to-state stability and applications to finite-time control design. SIAM Journal on Control and Optimization, 48(7):4395–4418, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hassine:2007:DCT

[HJM07] Maatoug Hassine, Sophie Jan, and Mohamed Masmoudi. From differential calculus to 0-1 topological optimization. SIAM Journal on Control and Optimization, 45(6):1965– 1987, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hosfeld:2024:ISS

[HJST24] René Hosfeld, Birgit Jacob, Felix L. Schwenninger, and Marius Tucsnak. Input-to-state stability for bilinear feedback systems. SIAM Journal on Control and Optimization, 62(3):1369–1389, June 2024. CODEN SJCODC. ISSN 1095-7138.

Hu:2018:GSM

[HJX18] Mingshang Hu, Shaolin Ji, and Xiaole Xue. A global stochastic maximum principle for fully coupled forward– backward stochastic systems. SIAM Journal on Control and Optimization, 56(6):4309– 4335, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hu:2019:EUV

[HJX19] Mingshang Hu, Shaolin Ji, and Xiaole Xue. The existence and uniqueness of viscosity solution to a kind of Hamilton–Jacobi–Bellman equation. SIAM Journal on Control and Optimization, 57 (6):3911–3938, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hu:2022:GSM

[HJX22] Mingshang Hu, Shaolin Ji, and Rundong Xu. A global stochastic maximum principle for forward-backward stochastic control systems with quadratic generators. SIAM Journal on Control and Optimization, 60(3):1791–1818, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M137238X.

Hu:2012:TIS

[HJZ12] Ying Hu, Hanqing Jin, and Xun Yu Zhou. Timeinconsistent stochastic Linear– Quadratic control. SIAM Journal on Control and Optimization, 50(3):1548–1572, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hu:2017:TIS

[HJZ17] Ying Hu, Hanqing Jin, and Xun Yu Zhou. Timeinconsistent stochastic linearquadratic control: Characterization and uniqueness of equilibrium. SIAM Journal on Control and Optimization, 55(2):1261–1279, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Habetler:1978:DAN

[HK78] G. J. Habetler and M. M. Kostreva. On a direct algorithm for nonlinear complementarity problems. SIAM Journal on Control and Optimization, 16(3):504–511, May 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hewer:1988:SSL

[HK88] Gary Hewer and Charles Kenney. The sensitivity of the stable Lyapunov equation. SIAM Journal on Control and Optimization, 26(2): 321–344, March 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hill:1994:MOS

[HK94] T. P. Hill and D. P. Kennedy.

Minimax-optimal strategies for the best-choice problem when a bound is known for the expected number of objects. SIAM Journal on Control and Optimization, 32(4):937–951, July 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23472.

Hinze:2001:SOM

[HK01] Michael Hinze and Karl Kunisch. Second order methods for optimal control of timedependent fluid flow. SIAM Journal on Control and Optimization, 40(3):925–946, May 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36181.

Hintermuller:2006:FNP

[HK06] M. Hintermüller and K. Kunisch. Feasible and noninterior path-following in constrained minimization with low multiplier regularity. SIAM Journal on Control and Optimization, 45(4):1198–1221, January 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Haak:2007:WAW

[HK07] Bernhard H. Haak and Peer Chr.

Kunstmann. Weighted admissibility and wellposedness of linear systems in Banach spaces. SIAM Journal on Control and Optimization, 45(6): 2094–2118, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Habbal:2013:NDN

[HK13] A. Habbal and M. Kallel. Neumann–Dirichlet Nash strategies for the solution of elliptic Cauchy problems. SIAM Journal on Control and Optimization, 51(5):4066–4083, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Huang:2019:OCS

[HK19] Yu-Jui Huang and Saeed Khalili. Optimal consumption in the stochastic Ramsey problem without boundedness constraints. SIAM Journal on Control and Optimization, 57 (2):783–809, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Harbrecht:2021:SGA

[HK21] Helmut Harbrecht and Ilja Kalmykov. Sparse grid approximation of the Riccati operator for closed loop parabolic control problems with Dirichlet boundary control. SIAM Journal on Control and Optimization, 59(6): 4538–4562, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hamzi:2004:ACH

[HKB04] Boumediene Hamzi, Wei Kang, and Jean-Pierre Barbot. Analysis and control of Hopf bifurcations. SIAM Journal on Control and Optimization, 42(6): 2200–2220, November 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37271.

Hu:1989:OGC

[HKL89] T. C. Hu, Victor Klee, and David Larman. Optimization of globally convex functions. SIAM Journal on Control and Optimization, 27(5): 1026–1047, September 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hintermuller:2017:OCS

[HKW17] Michael Hintermüller, Tobias Keil, and Donat Wegner. Optimal control of a semidiscrete Cahn–Hilliard–Navier–Stokes system with nonmatched fluid densities. SIAM Journal on Control and Optimization, 55 (3):1954–1989, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hernandez-Lerma:1981:EPC

[HL81] Onésimo Hernández-Lerma. Exit probabilities for a class of perturbed degenerate systems. SIAM Journal on Control and Optimization, 19(1): 39–51, January 1981. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Han:1988:PAC

[HL88] Shih-Ping Han and Gang Lou. A parallel algorithm for a class of convex programs. SIAM Journal on Control and Optimization, 26(2): 345–355, March 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Haussmann:1990:EOC

[HL90] U. G. Haussmann and J. P. Lepeltier. On the existence of optimal controls. SIAM Journal on Control and Optimization, 28(4):851–902, July 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Haurie:1992:OOR

[HL92] Alain Haurie and Arie Leizarowitz. Overtaking optimal regulation and tracking of piecewise diffusion linear systems. SIAM Journal on Control and Optimization, 30(4):816–837, July 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hirschorn:2004:HOV

[HL04] Ronald Hirschorn and Andrew D. Lewis. High-order variations for families of vector fields. SIAM Journal on Control and Optimization, 43 (1):301–324, January 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40062.

Hintermuller:2011:OSD

[HL11] M. Hintermüller and A. Laurain. Optimal shape design subject to elliptic variational inequalities. SIAM Journal on Control and Optimization, 49(3):1015–1047, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sicon/
resource/1/sjcodc/v49/i3/
p1015_s1.

Henderson:2016:MEU

[HL16] Vicky Henderson and Gechun Liang. A multidimensional exponential utility indifference pricing model with applications to counterparty risk. SIAM Journal on Control and Optimization, 54(2):690–717, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Huang:2018:OVR

[HLCH18] Lu-Jing Huang, Yin-Ting Liao, Ting-Li Chen, and Chii-Ruey Hwang. Optimal variance reduction for Markov Chain Monte Carlo. SIAM Journal on Control and Optimization, 56(4):2977–2996, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hu:2020:ASH

[HLDM20] Junhao Hu, Wei Liu, Feiqi Deng, and Xuerong Mao. Advances in stabilization of hybrid stochastic differential equations by delay feedback control. SIAM Journal on Control and Optimization, 58 (2):735–754, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Halikias:1993:SSA

[HLG93] G. D. Halikias, D. J. N. Limebeer, and K. Glover. A statespace algorithm for the superoptimal Hankel-norm approximation problem. SIAM Journal on Control and Optimization, 31(4):960–982, July 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Huang:2021:OIT

[HLG21] Wenlin Huang, Jin Liang, and Huaying Guo. Optimal investment timing for carbon emission reduction technology with a jump-diffusion process. SIAM Journal on Control and Optimization, 59(5): 4024–4050, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hernandez-Lerma:1998:ILP

[HLGH98] Onésimo Hernández-Lerma and Juan González-Hernández. Infinite linear programming and multichain Markov control processes in uncountable spaces. SIAM Journal on Control and Optimization, 36 (1):313–335, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29223.

Hernandez-Lerma:2003:CAC

[HLGHLM03] Onésimo Hernández-Lerma, Juan González-Hernández, and Raquiel R. López-Martínez. Constrained average cost Markov control processes in Borel spaces. SIAM Journal on Control and Optimization, 42(2):442–468, March 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36162.

Hernandez-Lerma:1994:LPA

[HLL94] Onésimo Hernández-Lerma and Jean B. Lasserre. Linear programming and average optimality of Markov control processes on Borel spaces — unbounded costs. SIAM Journal on Control and Optimization, 32(2):480–500, March 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/22832.

Hernandez-Lerma:1996:AOM

[HLL96] Onésimo Hernández-Lerma and Jean B. Lasserre. Average optimality in Markov control processes via discountedcost problems and linear programming. SIAM Journal on Control and Optimization, 34(1):295–310, January 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/24530.

Hernandez-Lerma:2000:ZSS

[HLL00] Onésimo Hernández-Lerma and Jean B. Lasserre. Zerosum stochastic games in Borel spaces: Average payoff criteria. SIAM Journal on Control and Optimization, 39(5): 1520–1539, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/36196.

Hafizoglu:2017:SLQ

[HLL^{+}17] Cavit Hafizoğlu, Irena Lasiecka, Tijana Levajković, Hermann Mena, and Amjad Tuffaha. The stochastic linear quadratic control problem with singular estimates. SIAM Journal on Control and Optimization, 55 (2):595–626, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hager:2018:CRG

[HLM^{+}18] William W. Hager, Jun Liu, Subhashree Mohapatra, Anil V. Rao, and Xiang-Sheng Wang. Convergence rate for a Gauss collocation method applied to constrained optimal control. SIAM Journal on Control and Optimization, 56(2):1386–1411, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hrynkiv:2008:OCC

[HLP08] Volodymyr Hrynkiv, Suzanne Lenhart, and Vladimir Protopopescu. Optimal control of a convective boundary condition in a thermistor problem. SIAM Journal on Control and Optimization, 47(1): 20–39, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hernandez-Lerma:2001:LDC

[HLR01] Onésimo Hernández-Lerma and Rosario Romera. Limiting discounted-cost control of partially observable stochastic systems. SIAM Journal on Control and Optimization, 40(2):348–369, March 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/36519.

Hauck:2008:CDE

[HLT08] Cory D. Hauck, C. David Levermore, and André L. Tits. Convex duality and entropybased moment closures: Characterizing degenerate densities. SIAM Journal on Control and Optimization, 47(4): 1977–2015, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hu:2020:SEB

[HLT20] Ying Hu, Gechun Liang, and Shanjian Tang. Systems of ergodic BSDEs arising in regime switching forward performance processes. SIAM Journal on Control and Optimization, 58(4):2503–2534, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hernandez-Lerma:1999:SPO

[HLVAC99] Onésimo Hernández-Lerma, Oscar Vega-Amaya, and Guadalupe Carrasco. Samplepath optimality and varianceminimization of average cost Markov control processes. SIAM Journal on Control and Optimization, 38(1):79– 93, January 1999. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/34067.

Han:2024:SWE

[HLY24] Zhong-Jie Han, Zhuangyi Liu, and Kai Yu. Stabilization for wave equation with localized Kelvin–Voigt damping on cuboidal domain: A degenerate case. SIAM Journal on Control and Optimization, 62(1):441–465, February 2024. CODEN SJCODC. ISSN 1095-7138.

Huang:2015:SAB

[HLZ15] Minyi Huang, Tao Li, and Ji-Feng Zhang. Stochastic approximation based consensus dynamics over Markovian networks. SIAM Journal on Control and Optimization, 53(6): 3339–3363, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Huang:2020:ASS

[HLZ20] Shuo Huang, Gechun Liang, and Thaleia Zariphopoulou. An approximation scheme for semilinear parabolic PDEs with convex and coercive Hamiltonians. SIAM Journal on Control and Optimization, 58(1):165–191, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Han:2021:OCS

[HLZF21] Chunyan Han, Hongdan Li, Huanshui Zhang, and Minyue Fu. Optimal control and stabilization for discrete-time Markov jump linear systems with input delay. SIAM Journal on Control and Optimization, 59(5):3524–3551, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hager:1976:LDT

[HM76] William W. Hager and Sanjoy K. Mitter. Lagrange duality theory for convex control problems. SIAM Journal on Control and Optimization, 14(5):843–856, August 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hill:1983:GIR

[HM83] David J. Hill and Peter J. Moylan. General instability results for interconnected systems. SIAM Journal on Control and Optimization, 21(2): 256–279, March 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Heinricher:1988:NEL

[HM88] Arthur C. Heinricher and Victor J. Mizel. A new example of the Lavrentiev phenomenon. SIAM Journal on Control and Optimization, 26(6):1490– 1503, November 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Helton:1993:CO

[HM93] J. William Helton and Orlando Merino. Conditions for optimality over H^{∞}. SIAM Journal on Control and Optimization, 31(6):1379– 1415, November 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Holcman:2002:SAS

[HM02] David Holcman and Michael Margaliot. Stability analysis of second-order switched homogeneous systems. SIAM Journal on Control and Optimization, 41(5):1609–1625, September 2002. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38935.

Hagen:2003:SSF

[HM03] Gregory Hagen and Igor Mezić. Spillover stabilization in finite-dimensional control and observer design for dissipative evolution equations. SIAM Journal on Control and Optimization, 42(2):746–768, March 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37894.

Haurie:2008:SPP

[HM08] Alain Haurie and Francesco Moresino. Singularly perturbed piecewise deterministic games. SIAM Journal on Control and Optimization, 47 (1):73–91, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Huang:2009:CCN

[HM09] Minyi Huang and Jonathan H. Manton. Coordination and consensus of networked agents with noisy measurements: Stochastic algorithms and asymptotic behavior. SIAM Journal on Control and Optimization, 48(1):134–161, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hamadene:2014:MNS

[HM14] Said Hamadène and Hassani Mohammed. The multiplayer nonzero-sum Dynkin game in continuous time. SIAM Journal on Control and Optimization, 52(2):821–835, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hamatsuka:1981:ODL

[HMA81] Teruo Hamatsuka, Abdul-Aziz Mo’Omen, and Hajime Akashi. Observer design for linear contractive control systems on Hilbert spaces. SIAM Journal on Control and Optimization, 19(5):586– 594, September 1981. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Huijberts:1997:GCI

[HMA97] H. J. C. Huijberts, C. H.

Moog, and R. Andiarti. Generalized controlled invariance for nonlinear systems. SIAM Journal on Control and Optimization, 35(3):953–979, May 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27719.

Hintermuller:2013:OBC

[HMMS13] Michael Hintermüller, Daniel Marahrens, Peter A. Markowich, and Christof Sparber. Optimal bilinear control of Gross– Pitaevskii equations. SIAM Journal on Control and Optimization, 51(3):2509–2543, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ho-Mock-Qai:1999:SSL

[HMQD99] Bertina Ho-Mock-Qai and Wijesuriya P. Dayawansa. Simultaneous stabilization of linear and nonlinear systems by means of nonlinear state feedback. SIAM Journal on Control and Optimization, 37(6): 1701–1725, November 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31561.

Hu:2015:PAL

[HMR15] Ying Hu, Pierre-Yves Madec, and Adrien Richou. A probabilistic approach to large time behavior of mild solutions of HJB equations in infinite dimension. SIAM Journal on Control and Optimization, 53 (1):378–398, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Homberg:2010:OCT

[HMRR10] D. Hömberg, C. Meyer, J. Rehberg, and W. Ring. Optimal control for the thermistor problem. SIAM Journal on Control and Optimization, 48(5):3449–3481, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hegoburu:2018:CPC

[HMT18] Nicolas Hegoburu, Pierre Magal, and Marius Tucsnak. Controllability with positivity constraints of the Lotka– McKendrick system. SIAM Journal on Control and Optimization, 56(2):723–750, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Haraux:2005:ASI

[HMV05] A. Haraux, P. Martinez, and J. Vancostenoble. Asymptotic stability for intermittently controlled second-order evolution equations. SIAM Journal on Control and Optimization, 43(6):2089–2108, November 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43656.

Herzog:2012:CSO

[HMW12] Roland Herzog, Christian Meyer, and Gerd Wachsmuth. C-stationarity for optimal control of static plasticity with linear kinematic hardening. SIAM Journal on Control and Optimization, 50(5):3052– 3082, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Huang:1996:SRB

[HN96] L. R. Huang and K. F. Ng.

On some relations between Chaney’s generalized secondorder directional derivative and that of Ben-Tal and Zowe. SIAM Journal on Control and Optimization, 34(4):1220–1234, July 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26649.

Hermann:2000:AIR

[HN00] Ulrike Hermann and Dominikus Noll. Adaptive image reconstruction using information measures. SIAM Journal on Control and Optimization, 38(4):1223–1240, July 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/32433.

Hoppe:2022:OCQ

[HN22] Fabian Hoppe and Ira Neitzel. Optimal control of quasilinear parabolic PDEs with stateconstraints. SIAM Journal on Control and Optimization, 60(1):330–354, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1383951.

Hata:2018:OCP

[HNS18] Hiroaki Hata, Hideo Nagai, and Shuenn-Jyi Sheu. An optimal consumption problem for general factor models. SIAM Journal on Control and Optimization, 56(5):3149–3183, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Huijberts:1992:DDD

[HNv92] H. J. C. Huijberts, H. Nijmeijer, and L. L. M. van der Wegen. Dynamic disturbance decoupling for nonlinear systems. SIAM Journal on Control and Optimization, 30(2): 336–349, March 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ho:1986:SAS

[Ho86]  Lop Fat Ho.  Spectral

assignability of systems with scalar control and application to a degenerate hyperbolic system. SIAM Journal on Control and Optimization, 24(6):1212–1231, November 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hinrichsen:1990:CCO

[HO90a] D. Hinrichsen and J. O’Halloran.

A complete characterization of orbit closures of controllable singular systems under restricted system equivalence. SIAM Journal on Control and Optimization, 28(3):602–623, May 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ho:1990:ECO

[Ho90b] Lop Fat Ho. Exact controllability of the one-dimensional wave equation with locally distributed control. SIAM Journal on Control and Optimization, 28(3):733–748, May 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ho:1993:CSC

[Ho93] L. F. Ho. Controllability and stabilizability of coupled strings with control applied at the coupled points. SIAM Journal on Control and Optimization, 31(6):1416– 1437, November 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hanzon:1997:OBB

[HO97] Bernard Hanzon and Raimund J. Ober. Overlapping blockbalanced canonical forms and parametrizations: The stable SISO case. SIAM Journal on Control and Optimization, 35(1):228–242, January 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/26054.

Hanzon:2001:SSC

[HO01] Bernard Hanzon and Raimund J. Ober. A state-space calculus for rational probability density functions and applications to non-Gaussian filtering. SIAM Journal on Control and Optimization, 40(3):724–740, May 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/31610.

Hu:2008:PIL

[HO08] Yaozhong Hu and Bernt Oksendal. Partial information linear quadratic control for jump diffusions. SIAM Journal on Control and Optimization, 47(4):1744–1761, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hendrickx:2016:SCO

[HO16] Julien M. Hendrickx and Alex Olshevsky. On symmetric continuum opinion dynamics. SIAM Journal on Control and Optimization, 54(5): 2893–2918, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hartmann:2022:RSH

[HO22] Andreas Hartmann and Marcu-Antone Orsoni. Reachable space of the Hermite heat equation with boundary control. SIAM Journal on Control and Optimization, 60 (6):3409–3429, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1451634.

Hoehener:2012:VAS

[Hoe12] Daniel Hoehener. Variational approach to second-order optimality conditions for control problems with pure state constraints. SIAM Journal on Control and Optimization, 50(3):1139–1173, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hoehener:2014:VAS

[Hoe14] Daniel Hoehener. Variational approach to second-order sufficient optimality conditions in optimal control. SIAM Journal on Control and Optimization, 52(2):861–892, ???? 2014. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Hopenwasser:1984:HLO

[Hop84]  Alan Hopenwasser.  Hy-

percausal linear operators. SIAM Journal on Control and Optimization, 22(6):911– 919, November 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hopkins:1987:OSF

[Hop87] William E. Hopkins, Jr. Optimal stabilization of families of linear stochastic differential equations with jump coefficients and multiplicative noise. SIAM Journal on Control and Optimization, 25(6): 1587–1600, November 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Heath:1987:MME

[HOPS87] D. Heath, S. Orey, V. Pestien, and W. Sudderth. Minimizing or maximizing the expected time to reach zero. SIAM Journal on Control and Optimization, 25(1):195–205, January 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hager:1987:DTM

[HP87] William W. Hager and Dwayne L. Presler. Dual techniques for minimax. SIAM Journal on Control and Optimization, 25 (3):660–685, May 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hinrichsen:1994:RSL

[HP94] D. Hinrichsen and A. J.

Pritchard. Robust stability of linear evolution operators on Banach spaces. SIAM Journal on Control and Optimization, 32(6):1503–1541, November 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23040.

Hinrichsen:1996:SRS

[HP96] D. Hinrichsen and A. J.

Pritchard. Stability radii of systems with stochastic uncertainty and their optimization by output feedback. SIAM Journal on Control and Optimization, 34(6): 1972–1998, November 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27689.

Hinrichsen:1998:S

[HP98] D. Hinrichsen and A. J.

Pritchard. Stochastic H^{∞}. SIAM Journal on Control and Optimization, 36(5): 1504–1538, September 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/30133.

Hinrichsen:2009:CSU

[HP09a] Diederich Hinrichsen and Anthony J. Pritchard. Composite systems with uncertain couplings of fixed structure: Scaled Riccati equations and the problem of quadratic stability. SIAM Journal on Control and Optimization, 47(6): 3037–3075, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hu:2009:BSD

[HP09b] Yaozhong Hu and Shige Peng. Backward stochastic differential equation driven by fractional Brownian motion. SIAM Journal on Control and Optimization, 48(3):1675– 1700, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hamalainen:2010:RRD

[HP10] Timo Hämäläinen and Seppo Pohjolainen. Robust regulation of distributed parameter systems with infinitedimensional exosystems. SIAM Journal on Control and Optimization, 48(8):4846–4873, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ho:2014:SPM

[HP14] Chin Pang Ho and Panos Parpas. Singularly perturbed Markov decision processes: a multiresolution algorithm. SIAM Journal on Control and Optimization, 52(6):3854– 3886, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Haidar:2021:LKC

[HP21] Ihab Haidar and Pierdomenico Pepe. Lyapunov–Krasovskii characterization of the inputto-state stability for switching retarded systems. SIAM Journal on Control and Optimization, 59(4):2997–3016, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hermosilla:2022:OCS

[HP22] Cristopher Hermosilla and Michele Palladino. Optimal control of the sweeping process with a nonsmooth moving set. SIAM Journal on Control and Optimization, 60 (5):2811–2834, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1405472.

Hu:2003:OCM

[HPS03] Jianghai Hu, Maria Prandini, and Shankar Sastry. Optimal coordinated motions of multiple agents moving on a plane. SIAM Journal on Control and Optimization, 42(2):637–668, March 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38756.

Havarneanu:2007:EIC

[HPS07] Teodor Havârneanu, Cātālin Popa, and S. S. Sritharan. Exact internal controllability for the two-dimensional magnetohydrodynamic equations. SIAM Journal on Control and Optimization, 46(5): 1802–1830, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hu:2007:CPF

[HPT07] Jianghai Hu, Maria Prandini, and Claire Tomlin. Conjugate points in formation constrained optimal multiagent coordination: a case study. SIAM Journal on Control and Optimization, 45(6): 2119–2137, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hinrichsen:1983:GHM

[HPW83] D. Hinrichsen and D. Prätzel-Wolters. Generalized Hermite matrices and complete invariants of strict system equivalence. SIAM Journal on Control and Optimization, 21(2):

289–305, March 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Han:2010:MPB

[HPW10] Yuecai Han, Shige Peng, and Zhen Wu. Maximum principle for backward doubly stochastic control systems with applications. SIAM Journal on Control and Optimization, 48 (7):4224–4241, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Huang:2024:CBO

[HQR24] Hui Huang, Jinniao Qiu, and Konstantin Riedl. Consensusbased optimization for saddle point problems. SIAM Journal on Control and Optimization, 62(2):1093–1121, March 2024. CODEN SJCODC. ISSN 1095-7138.

Horst:2016:CCP

[HQZ16] Ulrich Horst, Jinniao Qiu, and Qi Zhang. A constrained control problem with degenerate coefficients and degenerate backward SPDEs with singular terminal condition. SIAM Journal on Control and Optimization, 54(2):946–963, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ho:1983:AIE

[HR83a] L. F. Ho and D. L. Russell. Admissible input elements for systems in Hilbert space and a Carleson measure criterion. SIAM Journal on Control and Optimization, 21 (4):614–640, July 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See erratum [HR83b].

Ho:1983:EAI

[HR83b] L. F. Ho and D. L. Russell. Erratum: “Admissible input elements for systems in Hilbert space and a Carleson measure criterion”. SIAM Journal on Control and Optimization, 21(6):985–986, November 1983. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [HR83a].

Hou:1998:PNC

[HR98] L. S. Hou and S. S. Ravindran. A penalized Neumann control approach for solving an optimal Dirichlet control problem for the Navier–Stokes equations. SIAM Journal on Control and Optimization, 36(5):1795–1814, September 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30487.

Holtmannspotter:2021:PEE

[HR21] Marita Holtmannspötter and Arnd Rösch. A priori error estimates for the finite element approximation of a nonsmooth optimal control problem governed by a coupled semilinear PDE–ODE system. SIAM Journal on Control and Optimization, 59(5):3329–3358, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Haasler:2021:MOT

[HRCK21] Isabel Haasler, Axel Ringh, Yongxin Chen, and Johan Karlsson. Multimarginal optimal transport with a tree-structured cost and the Schrödinger bridge problem. SIAM Journal on Control and Optimization, 59(4):2428– 2453, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hintermuller:2017:OSP

[HRMK17] Michael Hintermüller, Carlos N. Rautenberg, Masoumeh Mohammadi, and Martin Kanitsar. Optimal sensor placement: a robust approach. SIAM Journal on Control and Optimization, 55(6):3609– 3639, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hinz:2021:NLU

[HRPT21] Michael Hinz, Anna Rozanova-Pierrat, and Alexander Teplyaev. Non-Lipschitz uniform domain shape optimization in linear acoustics. SIAM Journal on Control and Optimization, 59 (2):1007–1032, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hannsgen:1988:ERR

[HRW88] Kenneth B. Hannsgen, Yuriko Renardy, and Robert L. Wheeler. Effectiveness and robustness with respect to time delays of boundary feedback stabilization in onedimensional viscoelasticity. SIAM Journal on Control and Optimization, 26(5):1200– 1234, September 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Helmke:2006:OFP

[HRW06] Uwe Helmke, Joachim Rosenthal, and Xiaochang Alex Wang. Output feedback pole assignment for transfer functions with symmetries. SIAM Journal on Control and Optimization, 45(5):1898–1914, ???? 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Holmaaker:1987:COP

[HS87] Kjell Holmåker and David Stewart. A class of optimization problems with noncompact constraints: General results and applications. SIAM Journal on Control and Optimization, 25(4):1032–1052, July 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Heinricher:1991:OCR

[HS91] Arthur C. Heinricher and Richard H. Stockbridge. Optimal control of the running max. SIAM Journal on Control and Optimization, 29(4): 936–953, July 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Haussmann:1995:SOSa

[HS95a] Ulrich G. Haussmann and Wulin Suo. Singular optimal stochastic controls. I. Existence. SIAM Journal on Control and Optimization, 33(3):916–936, May 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25025.

Haussmann:1995:SOSb

[HS95b] Ulrich G. Haussmann and Wulin Suo. Singular optimal stochastic controls II: Dynamic programming. SIAM Journal on Control and Optimization, 33(3):937–959, May 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25052.

Hou:1996:OPA

[HS96] L. Steven Hou and Weiwei Sun. Optimal positioning of anodes for cathodic protection. SIAM Journal on Control and Optimization, 34(3):855–873, May 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24410.

Homberg:2003:OSD

[HS03] Dietmar Hömberg and Jan Sokolowski. Optimal shape design of inductor coils for surface hardening. SIAM Journal on Control and Optimization, 42(3):1087–1117, May 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37582.

Hu:2009:SAH

[HS09] Jianghai Hu and Slobodan Simic. A study of the asymptotic holonomic efficiency problem. SIAM Journal on Control and Optimization, 48(2):1112–1133, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hante:2011:CLT

[HS11] Falk M. Hante and Mario Sigalotti. Converse Lyapunov theorems for switched systems in Banach and Hilbert spaces. SIAM Journal on Control and Optimization, 49(2): 752–770, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.

siam.org/sicon/resource/
1/sjcodc/v49/i2/p752_s1.

Hata:2012:HJBa

[HS12a] Hiroaki Hata and Shuenn-Jyi Sheu. On the Hamilton– Jacobi–Bellman equation for an optimal consumption problem: I. Existence of solution. SIAM Journal on Control and Optimization, 50(4): 2373–2400, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hata:2012:HJBb

[HS12b] Hiroaki Hata and Shuenn-Jyi Sheu. On the Hamilton– Jacobi–Bellman equation for an optimal consumption problem: II. Verification theorem. SIAM Journal on Control and Optimization, 50(4): 2401–2430, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hillar:2013:CSS

[HS13] Christopher J. Hillar and Frank Sottile. Complex static skew-symmetric output feedback control. SIAM Journal on Control and Optimization, 51(4):3011–3026, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hammouri:2015:NOB

[HS15] Hassan Hammouri and Mariem Sahnoun. Nonlinear observer based on cascade form and output injection. SIAM Journal on Control and Optimization, 53(5):3195–3227, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Haesaert:2017:VGM

[HSA17] Sofie Haesaert, Sadegh Esmaeil Zadeh Soudjani, and Alessandro Abate. Verification of general Markov decision processes by approximate similarity relations and policy refinement. SIAM Journal on Control and Optimization, 55(4):2333–2367, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hartmann:2013:BAB

[HSBTZ13] Carsten Hartmann, Boris Schäfer-Bung, and Anastasia Thöns-Zueva. Balanced averaging of bilinear systems with applications to stochastic control. SIAM Journal on Control and Optimization, 51(3): 2356–2378, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Heij:1997:SID

[HSD97] C. Heij, W. Scherrer, and M. Deistler. System identification by dynamic factor models. SIAM Journal on Control and Optimization, 35(6): 1924–1951, November 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28212.

Houssineau:2019:IMD

[HSJ19] Jeremie Houssineau, Sumeetpal S. Singh, and Ajay Jasra. Identification of multiobject dynamical systems: Consistency and Fisher information. SIAM Journal on Control and Optimization, 57(4): 2603–2627, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hernandez-Santibanez:2019:CTV

[HSM19] Nicolás Hernández-Santibáñez and Thibaut Mastrolia. Contract theory in a VUCA world. SIAM Journal on Control and Optimization, 57(4): 3072–3100, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hu:2016:GIS

[HSP16] Jianghai Hu, Jinglai Shen, and Vamsi Putta. Generalized input-to-state ℓ_{2}-gains of discrete-time switched linear control systems. SIAM Journal on Control and Optimization, 54(3):1475–1503, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hilgert:2000:NIC

[HSV00] Nadine Hilgert, Rachid Senoussi, and Jean-Pierre Vila. Nonparametric identification of controlled nonlinear time varying processes. SIAM Journal on Control and Optimization, 39(3):950–960, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/33445.

Herzog:2012:DSO

[HSW12] Roland Herzog, Georg Stadler, and Gerd Wachsmuth. Directional sparsity in optimal control of partial differential equations. SIAM Journal on Control and Optimization, 50(2): 943–963, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Herzog:2015:EDS

[HSW15] Roland Herzog, Georg Stadler, and Gerd Wachsmuth. Erratum: Directional sparsity in optimal control of partial differential equations. SIAM Journal on Control and Optimization, 53(4):2722–2723, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Han:2019:EPG

[HSW19] Yuecai Han, Qingshuo Song, and Gu Wang. Exit problems as the generalized solutions of Dirichlet problems. SIAM Journal on Control and Optimization, 57(4):2392–2414, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Helmes:2015:MAC

[HSZ15] K. L. Helmes, R. H. Stockbridge, and C. Zhu. A measure approach for continuous inventory models: Discounted cost criterion. SIAM Journal on Control and Optimization, 53(4):2100–2140, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Helmes:2024:MIC

[HSZ24] Kurt L. Helmes, Richard H.

Stockbridge, and Chao Zhu. On the modeling of impulse control with random effects for continuous Markov processes. SIAM Journal on Control and Optimization, 62(1): 699–723, February 2024. CO-DEN SJCODC. ISSN 1095-7138.

Hassin:1992:MLC

[HT92] R. Hassin and A. Tamir. Minimal length curves that are not embeddable in an open planar set: The problem of a lost swimmer with a compass. SIAM Journal on Control and Optimization, 30(3): 695–703, May 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hintermuller:2010:IPA

[HT10] Michael Hintermüller and Moulay Hicham Tber. An inverse problem in American options as a mathematical program with equilibrium constraints: C-stationarity and an active-set-newton solver. SIAM Journal on Control and Optimization, 48(7):4419– 4452, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Haberkorn:2011:CRS

[HT11] T. Haberkorn and E. Trélat. Convergence results for smooth regularizations of hybrid nonlinear optimal control problems. SIAM Journal on Control and Optimization, 49(4): 1498–1522, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicon/resource/
                                        1/sjcodc/v49/i4/p1498_s1.

Hegarty:2012:CSN

[HT12] Gareth Hegarty and Stephen Taylor. Classical solutions of nonlinear beam equations: Existence and stabilization. SIAM Journal on Control and Optimization, 50(2):703–719, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hu:2022:SLC

[HT22] Ying Hu and Shanjian Tang. Stochastic LQ control and associated Riccati equation of PDEs driven by state- and control-dependent white noise. SIAM Journal on Control and Optimization, 60(1):435–457, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1351904.

Houska:2017:SRM

[HTLV17] Boris Houska, Dries Telen, Filip Logist, and Jan Van Impe. Self-reflective model predictive control. SIAM Journal on Control and Optimization, 55(5):2959–2980, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Herings:2001:VIP

[HTY01] P. Jean-Jacques Herings, Dolf Talman, and Zaifu Yang. Variational inequality problems with a continuum of solutions: Existence and computation. SIAM Journal on Control and Optimization, 39(6): 1852–1873, ???? 2001. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hiriart-Urruty:1978:CNO

[HU78a] J. B. Hiriart-Urruty. Conditions nécessaires d’optimalité pour un programme stochastique avec recours. (French) [Necessary optimality conditions for a stochastic program with recourse]. SIAM Journal on Control and Optimization, 16(2):317–329, March 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hiriart-Urruty:1978:GGF

[HU78b] J. B. Hiriart-Urruty. Gradients généralisés de fonctions marginales. (French) [Generalized gradients of marginal functions]. SIAM Journal on Control and Optimization, 16 (2):301–316, March 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hiriart-Urruty:1982:ASO

[HU82] J. B. Hiriart-Urruty. Approximating a second-order directional derivative for nonsmooth convex functions. SIAM Journal on Control and Optimization, 20(6):783– 807, November 1982. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hiriart-Urruty:1984:CRA

[HU84] J.-B. Hiriart-Urruty. Calculus rules on the approximate second-order directional derivative of a convex function. SIAM Journal on Control and Optimization, 22(3): 381–404, May 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hu:2015:STE

[Hu15] Long Hu. Sharp time estimates for exact boundary controllability of quasilinear hyperbolic systems. SIAM Journal on Control and Optimization, 53(6):3383–3410, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hertlein:2019:IBA

[HU19] Lukas Hertlein and Michael Ulbrich. An inexact bundle algorithm for nonconvex nonsmooth minimization in Hilbert space. SIAM Journal on Control and Optimization, 57(5): 3137–3165, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Huang:1988:MML

[Hua88] Fa Lun Huang. On the mathematical model for linear elastic systems with analytic damping. SIAM Journal on Control and Optimization, 26(3): 714–724, May 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Huang:2007:UCV

[Hua07] Minyi Huang. Uniqueness of constrained viscosity solutions in hybrid control systems. SIAM Journal on Control and Optimization, 46(1): 332–355, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Huang:2010:LPL

[Hua10] Minyi Huang. Large-population LQG games involving a major player: The Nash Certainty Equivalence Principle. SIAM Journal on Control and Optimization, 48(5):3318–3353, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hughes:2016:BRU

[Hug16] Timothy H. Hughes. Behavioral realizations using companion matrices and the Smith form. SIAM Journal on Control and Optimization, 54(2): 845–865, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Huijberts:1992:NSN

[Hui92] H. J. C. Huijberts. A nonregular solution of the nonlinear dynamic disturbance decoupling problem with an application to a complete solution of the nonlinear model matching problem. SIAM Journal on Control and Optimization, 30 (2):350–366, March 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hutchinson:2009:ESM

[Hut09] Ryan Hutchinson. The existence of strongly MDS convolutional codes. SIAM Journal on Control and Optimization, 47(6):2812–2826, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Huber:2000:PHP

[HV00] Birkett Huber and Jan Verschelde. Pieri homotopies for problems in enumerative geometry applied to pole placement in linear systems control. SIAM Journal on Control and Optimization, 38(4):1265–1287, July 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/35657.

Hu:2019:BES

[HVDK19] Long Hu, Rafael Vazquez, Florent Di Meglio, and Miroslav Krstic. Boundary exponential stabilization of 1-dimensional inhomogeneous quasi-linear hyperbolic systems. SIAM Journal on Control and Optimization, 57(2):963–998, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hinderer:1999:ASM

[HW99] Karl Hinderer and Karl-Heinz Waldmann. Approximate solution of Markov renewal programs with finite time horizon. SIAM Journal on Control and Optimization, 37(2):502–520, March 1999. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/31720.

Helton:2003:GUT

[HW03] J. William Helton and Marshall A. Whittlesey. Global uniqueness tests and performance bounds for H^{∞} optima. SIAM Journal on Control and Optimization, 42 (1):363–380, January 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/38993.

Hinderer:2005:ACS

[HW05] K. Hinderer and K. H. Waldmann. Algorithms for countable state Markov decision models with an absorbing set. SIAM Journal on Control and Optimization, 43(6): 2109–2131, November 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/41102.

Hintermuller:2012:DOC

[HW12] M. Hintermüller and D. Wegner. Distributed optimal control of the Cahn–Hilliard system including the case of a double-obstacle homogeneous free energy density. SIAM Journal on Control and Optimization, 50(1):388–

418, ????  2012.  CO-

DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v50/i1/p388_s1.

Hintermuller:2014:OCS

[HW14] M. Hintermüller and D. Wegner. Optimal control of a semidiscrete Cahn–Hilliard– Navier–Stokes system. SIAM Journal on Control and Optimization, 52(1):747–772, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hu:2018:BCO

[HW18] Weiwei Hu and Jiahong Wu.

Boundary control for optimal mixing via Navier–Stokes flows. SIAM Journal on Control and Optimization, 56(4): 2768–2801, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hermosilla:2019:CMF

[HW19] Cristopher Hermosilla and Peter Wolenski. A characteristic method for fully convex Bolza problems over arcs of bounded variation. SIAM Journal on Control and Optimization, 57 (4):2873–2901, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hu:2020:MPS

[HW20] Mingshang Hu and Falei Wang. Maximum principle for stochastic recursive optimal control problem under model uncertainty. SIAM Journal on Control and Optimization, 58(3):1341–1370, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Huang:2021:OEM

[HW21] Yu-Jui Huang and Zhenhua Wang. Optimal equilibria for multidimensional timeinconsistent stopping problems. SIAM Journal on Control and Optimization, 59(2): 1705–1729, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Huang:2009:MPP

[HWX09] Jianhui Huang, Guangchen Wang, and Jie Xiong. A maximum principle for partial information backward stochastic control problems with applications. SIAM Journal on Control and Optimization, 48(4): 2106–2117, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Huang:2021:SOM

[HWY21] Jianhui Huang, Bing-Chang Wang, and Jiongmin Yong. Social optima in mean field linear-quadratic-Gaussian control with volatility uncertainty. SIAM Journal on Control and Optimization, 59(2): 825–856, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Han:2023:CRC

[HWZ23] Xia Han, Ruodu Wang, and Xun Yu Zhou. Choquet regularization for continuous-time reinforcement learning. SIAM Journal on Control and Optimization, 61(5):2777–2801, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1524734.

Hambly:2021:PGM

[HXY21] Ben Hambly, Renyuan Xu, and Huining Yang. Policy gradient methods for the noisy linear quadratic regulator over a finite horizon. SIAM Journal on Control and Optimization, 59(5):3359–3391, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hu:1995:PMP

[HY95] Bei Hu and Jiong Min Yong.

Pontryagin maximum principle for semilinear and quasilinear parabolic equations with pointwise state constraints. SIAM Journal on Control and Optimization, 33(6):1857–1880, November 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25007.

Hou:1997:DAV

[HY97a] L. S. Hou and Y. Yan. Dynamics and approximations of a velocity tracking problem for the Navier–Stokes flows with piecewise distributed controls. SIAM Journal on Control and Optimization, 35(6): 1847–1885, November 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/29286.

Hou:1997:DCN

[HY97b] L. S. Hou and Y. Yan. Dynamics for controlled Navier– Stokes systems with distributed controls. SIAM Journal on Control and Optimization, 35(2):654–677, March 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/27492.

Han:2022:PMC

[HY22] Jinhui Han and Sheung Chi Phillip Yam. A probabilistic method for a class of non-Lipschitz BSDEs with application to fund management. SIAM Journal on Control and Optimization, 60 (3):1193–1222, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M140609X.

He:2011:LDT

[HYZ11] Qi He, George Yin, and Qing Zhang. Large deviations for two-time-scale systems driven by nonhomogeneous Markov chains and associated optimal control problems. SIAM Journal on Control and Optimization, 49(4): 1737–1765, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i4/p1737_s1.

Huang:2021:IHL

[HYZ21] Jianping Huang, Jiongmin Yong, and Hua-Cheng Zhou. Infinite horizon linear quadratic overtaking optimal control problems. SIAM Journal on Control and Optimization, 59 (2):1312–1340, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Han:2024:RIP

[HYZ24] Xiaoru Han, Fahuai Yi, and Jianbo Zhang. A reversible investment problem with capacity and demand in finite horizon: Free boundary analysis. SIAM Journal on Control and Optimization, 62(2):1207– 1234, April 2024. CODEN SJ-CODC. ISSN 1095-7138.

Hansen:1995:ECS

[HZ95] Scott Hansen and Enrique Zuazua. Exact controllability and stabilization of a vibrating string with an interior point mass. SIAM Journal on Control and Optimization, 33(5):1357–1391, September 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24834.

Hu:2003:ISR

[HZ03] Ying Hu and Xun Yu Zhou. Indefinite stochastic Riccati equations. SIAM Journal on Control and Optimization, 42 (1):123–137, January 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39133.

Hu:2005:SCL

[HZ05a] Yaozhong Hu and Xun Yu Zhou. Stochastic control for linear systems driven by fractional noises. SIAM Journal on Control and Optimization, 43(6):2245–2277, November 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42604.

Hu:2005:CSL

[HZ05b] Ying Hu and Xun Yu Zhou. Constrained stochastic LQ control with random coefficients, and application to portfolio selection. SIAM Journal on Control and Optimization, 44(2):444–466, March 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/44196.

Hager:2007:ACA

[HZ07] William W. Hager and Hongchao Zhang. Asymptotic convergence analysis of a new class of proximal point methods. SIAM Journal on Control and Optimization, 46(5): 1683–1704, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hamadene:2010:CTN

[HZ10a] Said Hamadène and Jianfeng Zhang. The continuous time nonzero-sum Dynkin game problem and application in game options. SIAM Journal on Control and Optimization, 48(5):3659–3669, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Hu:2010:SSO

[HZ10b] Jianghai Hu and Wei Zhang.

Symmetry of solutions to the optimal exit time control problem. SIAM Journal on Control and Optimization, 48(8): 5488–5509, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Huang:2019:OET

[HZ19] Yu-Jui Huang and Zhou Zhou. The optimal equilibrium for time-inconsistent stopping problems — the discrete-time case. SIAM Journal on Control and Optimization, 57(1):590–609, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Heiland:2021:CST

[HZ21] Jan Heiland and Enrique Zuazua. Classical system theory revisited for turnpike in standard state space systems and impulse controllable descriptor systems. SIAM Journal on Control and Optimization, 59(5):3600–3624, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Han:2022:SDT

[HZ22] Zhong-Jie Han and Enrique Zuazua. Slow decay and turnpike for infinite-horizon hyperbolic linear quadratic problems. SIAM Journal on Control and Optimization, 60 (4):2440–2468, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1441985.

Hassan:2013:RUI

[HZB13] L. Hassan, A. Zemouche, and M. Boutayeb. Robust unknown input observers for nonlinear time-delay systems. SIAM Journal on Control and Optimization, 51(4):2735– 2752, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ichikawa:1976:LQD

[Ich76]  Akira Ichikawa.  Linear

quadratic differential games in a Hilbert space. SIAM Journal on Control and Optimization, 14(1):120–136, January 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ichikawa:1979:DPA

[Ich79] Akira Ichikawa. Dynamic programming approach to stochastic evolution equations. SIAM Journal on Control and Optimization, 17(1):152– 174, January 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ichikawa:1982:QCE

[Ich82] Akira Ichikawa. Quadratic control of evolution equations with delays in control. SIAM Journal on Control and Optimization, 20(5):645– 668, September 1982. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ichikawa:2007:NCV

[Ich07] Akira Ichikawa. Null controllability with vanishing energy for discrete-time systems in Hilbert space. SIAM Journal on Control and Optimization, 46(2):683–693, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ichihara:2011:RTO

[Ich11] Naoyuki Ichihara. Recurrence and transience of optimal feedback processes associated with Bellman equations of ergodic type. SIAM Journal on Control and Optimization, 49(5): 1938–1960, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i5/p1938_s1.

Ichihara:2016:PTC

[Ich16] Naoyuki Ichihara. Phase transitions for controlled Markov chains on infinite graphs. SIAM Journal on Control and Optimization, 54(2):450–474, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Iusem:1987:SIM

[ID87] Alfredo N. Iusem and Alvaro R. De Pierro. A simultaneous iterative method for computing projections on polyhedra. SIAM Journal on Control and Optimization, 25(1):231–243, January 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Idczak:1998:OCC

[Idc98]  Dariusz Idczak.  Optimal

control of a coercive Dirichlet problem. SIAM Journal on Control and Optimization, 36(4):1250–1267, July 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29634.

Iglesias:1996:EFT

[Igl96] Pablo A. Iglesias. An entropy formula for time-varying discrete-time control systems. SIAM Journal on Control and Optimization, 34(5): 1691–1706, September 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26433.

Ito:1997:BTD

[II97] Naoharu Ito and Hiroshi Inaba. Block triangular decoupling for linear systems over principal ideal domains. SIAM Journal on Control and Optimization, 35(3):744–765, May 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28110.

Ito:1990:ALM

[IK90] Kazufumi Ito and Karl Kunisch. The augmented Lagrangian method for parameter estimation in elliptic systems. SIAM Journal on Control and Optimization, 28(1): 113–136, January 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ito:1994:DFC

[IK94] Kazufumi Ito and Sungkwon Kang. A dissipative feedback control synthesis for systems arising in fluid dynamics. SIAM Journal on Control and Optimization, 32(3):831–854, May 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/22261.

Ito:1995:MRN

[IK95] Kazufumi Ito and Karl Kunisch. Maximizing robustness in nonlinear illposed inverse problems. SIAM Journal on Control and Optimization, 33(2):643–666, March 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23098.

Ishii:1996:NFS

[IK96a] Hitoshi Ishii and Shigeaki Koike. A new formulation of state constraint problems for first-order PDEs. SIAM Journal on Control and Optimization, 34(2):554–571, March 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25026.

Ito:1996:ALS

[IK96b] Kazufumi Ito and Karl Kunisch. Augmented Lagrangian– SQP methods for nonlinear optimal control problems of tracking type. SIAM Journal on Control and Optimization, 34(3):874–891, May 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26170.

Ito:2002:APR

[IK02a] Kazufumi Ito and Karl Kunisch. Asymptotic properties of receding horizon optimal control problems. SIAM Journal on Control and Optimization, 40(5): 1585–1610, September 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36942.

Ito:2002:OCS

[IK02b] Kazufumi Ito and Karl Kunisch. Optimal control of the solid fuel ignition model with H^{1}-cost. SIAM Journal on Control and Optimization, 40(5):1455–1472, September 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/36604.

Ito:2004:PDA

[IK04] Kazufumi Ito and Karl Kunisch. The primal-dual active set method for nonlinear optimal control problems with bilateral constraints. SIAM Journal on Control and Optimization, 43(1):357–376, January 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/41101.

Ito:2006:RHC

[IK06] Kazufumi Ito and Karl Kunisch. Receding horizon control with incomplete observations. SIAM Journal on Control and Optimization, 45(1): 207–225, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ito:2007:CPD

[IK07a] Kazufumi Ito and Karl Kunisch. Convergence of the primal-dual active set strategy for diagonally dominant systems. SIAM Journal on Control and Optimization, 46 (1):14–34, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ito:2007:OBC

[IK07b] Kazufumi Ito and Karl Kunisch. Optimal bilinear control of an abstract Schrödinger equation. SIAM Journal on Control and Optimization, 46 (1):274–287, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Isac:2008:DMA

[IK08] G. Isac and A. A. Khan.

Dubovitskii–Milyutin approach in set-valued optimization. SIAM Journal on Control and Optimization, 47(1):144–162, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ito:2009:APF

[IK09] Kazufumi Ito and Karl Kunisch. Asymptotic properties of feedback solutions for a class of quantum control problems. SIAM Journal on Control and Optimization, 48(4): 2323–2343, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ito:2010:SNM

[IK10] Kazufumi Ito and Karl Kunisch. Semismooth Newton methods for time-optimal control for a class of ODEs. SIAM Journal on Control and Optimization, 48(6):3997–4013, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ito:2011:KKT

[IK11a] Kazufumi Ito and Karl Kunisch. Karush–Kuhn–Tucker conditions for nonsmooth mathematical programming problems in function spaces. SIAM Journal on Control and Optimization, 49(5):2133– 2154, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i5/p2133_s1.

Ito:2011:MEP

[IK11b] Kazufumi Ito and Karl Kunisch. Minimal effort problems and their treatment by semismooth Newton methods. SIAM Journal on Control and Optimization, 49(5): 2083–2100, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i5/p2083_s1.

Ito:2014:OCC

[IK14] Kazufumi Ito and Karl Kunisch. Optimal control with L^{p}(Ω), p ∈ [0, 1), control cost. SIAM Journal on Control and Optimization, 52(2): 1251–1275, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ilchmann:2010:TPT

[ILR10] Achim Ilchmann, Hartmut Logemann, and Eugene P. Ryan.

Tracking with prescribed transient performance for hysteretic systems. SIAM Journal on Control and Optimization, 48(7):4731–4752, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ito:1998:ATS

[IM98] Kazufumi Ito and K. A.

Morris. An approximation theory of solutions to operator Riccati equations for H^{∞} control. SIAM Journal on Control and Optimization, 36(1):82–99, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27442.

Ilchmann:2005:BATa

[IM05a] Achim Ilchmann and Volker Mehrmann. A behavioral approach to time-varying linear systems. Part 1: General theory. SIAM Journal on Control and Optimization, 44(5):1725–1747, September 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/44223.

Ilchmann:2005:BATb

[IM05b] Achim Ilchmann and Volker Mehrmann. A behavioral approach to time-varying linear systems. Part 2: Descriptor systems. SIAM Journal on Control and Optimization, 44(5):1748–1765, September 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/60902.

Ilchmann:2008:RTG

[IM08] Achim Ilchmann and Markus Mueller. Robustness of λtracking in the gap metric. SIAM Journal on Control and Optimization, 47(5): 2724–2744, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ilchmann:2010:RFC

[IM10] Achim Ilchmann and Markus Mueller. Robustness of funnel control in the gap metric. SIAM Journal on Control and Optimization, 48(5): 3169–3190, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Immonen:2007:IMS

[Imm07] Eero Immonen. On the internal model structure for infinite-dimensional systems: Two common controller types and repetitive control. SIAM Journal on Control and Optimization, 45(6):2065–2093, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ionescu:1996:GDT

[IO96] Vlad Ionescu and Cristian Oară. Generalized discretetime Riccati theory. SIAM Journal on Control and Optimization, 34(2):601–619, March 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/26182.

Ioffe:1976:ETG

[Iof76] A. D. Ioffe. An existence theorem for a general Bolza problem. SIAM Journal on Control and Optimization, 14 (3):458–466, May 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ioffe:1977:LSIa

[Iof77a] A. D. Ioffe. On lower semicontinuity of integral functionals. I. SIAM Journal on Control and Optimization, 15(4): 521–538, July 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ioffe:1977:LSIb

[Iof77b] A. D. Ioffe. On lower semicontinuity of integral functionals. II. SIAM Journal on Control and Optimization, 15(6):991– 1000, November 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ioffe:1978:SMS

[Iof78] A. D. Ioffe. Survey of measurable selection theorems: Russian literature supplement. SIAM Journal on Control and Optimization, 16(5):728– 732, September 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ioffe:1979:NSCa

[Iof79a] A. D. Ioffe. Necessary and sufficient conditions for a local minimum — 1. a reduction theorem and first order conditions. SIAM Journal on Control and Optimization, 17 (2):245–250, March 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ioffe:1979:NSCc

[Iof79b] A. D. Ioffe. Necessary and sufficient conditions for a local minimum — 3. second order conditions and augmented duality. SIAM Journal on Control and Optimization, 17(2): 266–288, March 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ioffe:1979:NSCb

[Iof79c] A. D. Ioffe. Necessary and sufficient conditions for a local minimum. 2: Conditions of Levitin–Miljutin–Osmolovskii type. SIAM Journal on Control and Optimization, 17(2):

251–265, March 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ioffe:1983:SVR

[Iof83] A. D. Ioffe. Single-valued representation of set-valued mappings. II. Application to differential inclusions. SIAM Journal on Control and Optimization, 21(4):641–651, July 1983. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ioffe:2024:MPO

[Iof24] A. D. Ioffe. Maximum principles for optimal control problems with differential inclusions. SIAM Journal on Control and Optimization, 62(1): 271–296, January 2024. CO-DEN SJCODC. ISSN 1095-7138.

Ito:1987:CEI

[IP87] Kazufumi Ito and Robert K.

Powers. Chandrasekhar equations for infinite dimensional systems. SIAM Journal on Control and Optimization, 25 (3):596–611, May 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Immonen:2006:FFO

[IP06a] Eero Immonen and Seppo Pohjolainen. Feedback and feedforward output regulation of bounded uniformly continuous signals for infinite-dimensional systems. SIAM Journal on Control and Optimization, 45 (5):1714–1735, ???? 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Immonen:2006:WPS

[IP06b] Eero Immonen and Seppo Pohjolainen. What periodic signals can an exponentially stabilizable linear feedforward control system asymptotically track? SIAM Journal on Control and Optimization, 44(6): 2253–2268, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Isidori:2010:AEL

[IPM10] A. Isidori, L. Praly, and L. Marconi. About the existence of locally Lipschitz output feedback stabilizers for nonlinear systems. SIAM Journal on Control and Optimization, 48(5):3389–3402, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ito:2000:AKE

[IR00] Kazufumi Ito and Boris Rozovskii. Approximation of the Kushner equation for nonlinear filtering. SIAM Journal on Control and Optimization, 38(3):893–915, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34427.

Ilchmann:2002:SCF

[IRS02] A. Ilchmann, E. P. Ryan, and C. J. Sangwin. Systems of controlled functional differential equations and adaptive tracking. SIAM Journal on Control and Optimization, 40(6):1746–1764, November 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37970.

Ilchmann:2007:TPT

[IRT07] Achim Ilchmann, Eugene P.

Ryan, and Philip Townsend. Tracking with prescribed transient behavior for nonlinear systems of known relative degree. SIAM Journal on Control and Optimization, 46(1): 210–230, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Izmailov:2002:OCI

[IS02] A. F. Izmailov and M. V.

Solodov. Optimality conditions for irregular inequalityconstrained problems. SIAM Journal on Control and Optimization, 40(4):1280–1295, July 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35754.

Idan:2012:SEL

[IS12] Moshe Idan and Jason L. Speyer. State estimation for linear scalar dynamic systems with additive Cauchy noises: Characteristic function approach. SIAM Journal on Control and Optimization, 50(4):1971–1994, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Idan:2014:MCE

[IS14] Moshe Idan and Jason L. Speyer. Multivariate Cauchy estimator with scalar measurement and process noises. SIAM Journal on Control and Optimization, 52(2):1108– 1141, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Iusem:1999:CPG

[ISdCN99] Alfredo N. Iusem, B. F. Svaiter, and João Xavier da Cruz Neto. Central paths, generalized proximal point methods, and Cauchy trajectories in Riemannian manifolds. SIAM Journal on Control and Optimization, 37(2):566–588, March 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/29074.

Ishizuka:1988:OCQ

[Ish88] Yo Ishizuka. Optimality conditions for quasi-differentiable programs with application to two-level optimization. SIAM Journal on Control and Optimization, 26(6):1388– 1398, November 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Iusem:1996:MIG

[IST96] Alfredo N. Iusem, B. F.

Svaiter, and Marc Teboulle. Multiplicative interior gradient methods for minimization over the nonnegative orthant. SIAM Journal on Control and Optimization, 34 (1):389–406, January 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26294.

Ignat:2002:AON

[IST02] A. Ignat, J. Sprekels, and D. Tiba. Analysis and optimization of nonsmooth arches. SIAM Journal on Control and Optimization, 40(4):1107–1133, July 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37442.

Ilchmann:2016:BIF

[IST16] Achim Ilchmann, Tilman Selig, and Carsten Trunk.

The Byrnes–Isidori form for infinite-dimensional systems. SIAM Journal on Control and Optimization, 54(3):1504– 1534, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ito:2000:IMP

[ISW00] Naoharu Ito, Wiland Schmale, and Harald K. Wimmer. (C, A)-invariance of modules over principal ideal domains. SIAM Journal on Control and Optimization, 38(6): 1859–1873, November 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/35508.

Ito:1986:LTA

[IT86] Kazufumi Ito and Russell Teglas. Legendre–Tau approximations for functionaldifferential equations. SIAM Journal on Control and Optimization, 24(4):737–759, July 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ito:1987:LTA

[IT87] Kazufumi Ito and Russell Teglas. Legendre-tau approximation for functional differential equations. Part II: The linear quadratic optimal control problem. SIAM Journal on Control and Optimization, 25(6):1379–1408, November 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Iwata:2017:KCF

[IT17] Satoru Iwata and Mizuyo Takamatsu. On the Kronecker canonical form of singular mixed matrix pencils. SIAM Journal on Control and Optimization, 55(3):2134–2150, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ito:1990:FDC

[Ito90]  Kazufumi Ito.  Finite-

dimensional compensators for infinite-dimensional systems via Galerkin-type approximation. SIAM Journal on Control and Optimization, 28(6): 1251–1269, November 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ito:1996:AZE

[Ito96] Kazufumi Ito. Approximation of the Zakai equation for nonlinear filtering. SIAM Journal on Control and Optimization, 34(2):620–634, March 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25478.

Ito:1998:USD

[Ito98] Kazuho Ito. Uniform stabilization of the dynamic elastica by boundary feedback. SIAM Journal on Control and Optimization, 37(1):319–329, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32235.

Ito:2005:OOF

[Ito05] Kazufumi Ito. An optimal optical flow. SIAM Journal on Control and Optimization, 44(2):728–742, March 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43344.

Ilchmann:2004:ICA

[ITT04] Achim Ilchmann, Mosalagae Thuto, and Stuart Townley. Input constrained adaptive tracking with applications to exothermic chemical reaction models. SIAM Journal on Control and Optimization, 43(1):154–173, January 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39108.

Iiduka:2012:CMS

[IY12] Hideaki Iiduka and Isao Yamada. Computational method for solving a stochastic linearquadratic control problem given an unsolvable stochastic algebraic Riccati equation. SIAM Journal on Control and Optimization, 50(4):

2173–2192, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ioffe:2000:VPW

[IZ00] Alexander D. Ioffe and Alexander J. Zaslavski. Variational principles and well-posedness in optimization and calculus of variations. SIAM Journal on Control and Optimization, 38(2):566–581, March 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33563.

Izmailov:1996:AOF

[Izm96]  Rauf Izmailov.  Analysis

and optimization of feedback control algorithms for data transfers in high-speed networks. SIAM Journal on Control and Optimization, 34(5): 1767–1780, September 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27656.

Jacob:2004:WBS

[Jac04]  Birgit Jacob.  What is

the better signal space for discrete-time systems: ℓ_{2}(Z) or ℓ_{2}(N_{0})? SIAM Journal on Control and Optimization, 43(4):1521–1534, July 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/41247.

Jacobovic:2022:RCN

[Jac22]  Royi Jacobovic.  Recur-

sive construction of a Nash equilibrium in a two-player nonzero-sum stopping game with asymmetric information. SIAM Journal on Control and Optimization, 60 (6):3289–3301, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1465044.

Jafarizadeh:2017:ODS

[Jaf17] Saber Jafarizadeh. Optimizing the diffusion system based on continuous-time consensus algorithm. SIAM Journal on Control and Optimization, 55 (3):2024–2051, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jafarpour:2020:STL

[Jaf20] Saber Jafarpour. On smalltime local controllability. SIAM Journal on Control and Optimization, 58(1):425–446, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jahn:1985:CPM

[Jah85] Johannes Jahn. A characterization of properly minimal elements of a set. SIAM Journal on Control and Optimization, 23(5):649–656, September 1985. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jahn:1988:GTA

[Jah88] Johannes Jahn. A generalization of a theorem of Arrow, Barankin, and Blackwell. SIAM Journal on Control and Optimization, 26(5): 999–1005, September 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jakubczyk:1980:EUR

[Jak80] Bronisław Jakubczyk. Existence and uniqueness of realizations of nonlinear systems. SIAM Journal on Control and Optimization, 18(4):455–471, ???? 1980. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jakubczyk:1986:LRN

[Jak86] Bronisław Jakubczyk. Local realizations of nonlinear causal operators. SIAM Journal on Control and Optimization, 24 (2):230–242, March 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jenkins:2018:CPA

[JALG18] Benjamin M. Jenkins, Anuradha M. Annaswamy, Eugene Lavretsky, and Travis E. Gibson. Convergence properties of adaptive systems and the definition of exponential stability. SIAM Journal on Control and Optimization, 56(4): 2463–2484, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

James:1991:FTO

[Jam91]  Matthew R. James.  Fi-

nite time observer design by probabilistic-variational methods. SIAM Journal on Control and Optimization, 29(4): 954–967, July 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jammazi:2014:CDH

[Jam14] Chaker Jammazi. Continuous and discontinuous homogeneous feedbacks finite-time partially stabilizing controllable multichained systems. SIAM Journal on Control and Optimization, 52(1):520–544, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jarre:1993:IPM

[Jar93] Florian Jarre. An interiorpoint method for minimizing the maximum eigenvalue of a linear combination of matrices. SIAM Journal on Control and Optimization, 31(5):1360– 1377, September 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jaskiewicz:2002:ZSS

[Jas02] Anna Jaskiewicz. Zero-sum semi-Markov games. SIAM Journal on Control and Optimization, 41(3):723–739, May 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38190.

Jayawardhana:2010:SCD

[Jay10] Bayu Jayawardhana. State convergence of dissipative nonlinear systems given boundedenergy input signal. SIAM Journal on Control and Optimization, 48(6):4078–4088, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

James:1996:POD

[JB96] M. R. James and J. S. Baras.

Partially observed differential games, infinite-dimensional Hamilton–Jacobi–Isaacs equations, and nonlinear H_{∞} control. SIAM Journal on Control and Optimization, 34(4):1342–1364, July 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27333.

Jeroslow:1978:CPC

[Jer78] R. G. Jeroslow. Cuttingplanes for complementarity constraints. SIAM Journal on Control and Optimization, 16 (1):56–62, January 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jeroslow:1980:CPG

[Jer80]  Robert G. Jeroslow.  A

cutting-plane game for facial disjunctive programs. SIAM Journal on Control and Optimization, 18(3):264–281, May 1980. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jia:2017:ODS

[JFB17] Peng Jia, Noah E. Friedkin, and Francesco Bullo. Opinion dynamics and social power evolution over reducible influence networks. SIAM Journal on Control and Optimization, 55(2):1280–1301, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Janin:1999:LTS

[JG99] Robert Janin and Jacques Gauvin. Lipschitz-type stability in nonsmooth convex programs. SIAM Journal on Control and Optimization, 38 (1):124–137, January 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/29899.

Jiang:2023:AOC

[JGC^{+}23] Yi Jiang, Weinan Gao, Ci Chen, Tianyou Chai, and Frank L. Lewis. Adaptive optimal control of linear discrete-time networked control systems with two-channel stochastic dropouts. SIAM Journal on Control and Optimization, 61(5):3183–3208, October 2023. CODEN SJ-CODC. ISSN 1095-7138.

Joslin:2000:LIL

[JH00] J. A. Joslin and A. J. Heunis. Law of the iterated logarithm for a constantgain linear stochastic gradient algorithm. SIAM Journal on Control and Optimization, 39(2):533–570, March 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33100.

Jikuya:2014:KCD

[JH14] Ichiro Jikuya and Ichijo Hodaka. Kalman canonical decomposition of linear timevarying systems. SIAM Journal on Control and Optimization, 52(1):274–310, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jafarpour:2019:SKO

[JHB19] Saber Jafarpour, Elizabeth Y.

Huang, and Francesco Bullo. Synchronization of Kuramoto oscillators: Inverse Taylor expansions. SIAM Journal on Control and Optimization, 57 (5):3388–3412, ???? 2019.

CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jaimoukha:2006:GBC

[JHMG06] I. M. Jaimoukha, G. D. Halikias, U. Malik, and S. K. Gungah. On the gap between the complex structured singular value and its convex upper bound. SIAM Journal on Control and Optimization, 45(4): 1251–1278, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jiang:2006:SCD

[JK06] Shengbing Jiang and Ratnesh Kumar. Supervisory control of discrete event systems with CTL* temporal logic specifications. SIAM Journal on Control and Optimization, 44(6): 2079–2103, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jiao:2020:CPD

[JK20] Chunxi Jiao and Reiichiro Kawai. Computable primal and dual bounds for stochastic control. SIAM Journal on Control and Optimization, 58(6):3709–3733, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ji:2021:KBF

[JKSZ21] Shaolin Ji, Chuiliu Kong, Chuanfeng Sun, and Ji-Feng Zhang. Kalman–Bucy filtering and minimum mean square estimator under uncertainty. SIAM Journal on Control and Optimization, 59(4): 2669–2692, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jacob:2021:RBP

[JKZ21] Birgit Jacob, Julia T. Kaiser, and Hans Zwart. Riesz bases of port-Hamiltonian systems. SIAM Journal on Control and Optimization, 59(6): 4646–4665, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jacobs:1976:CFS

[JL76] Marc Q. Jacobs and C. E. Langenhop. Criteria for function space controllability of linear neutral systems. SIAM Journal on Control and Optimization, 14(6):1009–1048, November 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jaimoukha:1993:SSA

[JL93] I. M. Jaimoukha and D. J. N.

Limebeer. A state-space algorithm for the solution of the 2-block superoptimal distance problem. SIAM Journal on Control and Optimization, 31(5):1115–1134, September 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jeyakumar:1998:AJM

[JL98] V. Jeyakumar and D. T. Luc. Approximate Jacobian matrices for nonsmooth continuous maps and C_{1}- optimization. SIAM Journal on Control and Optimization, 36(5):1815–1832, September 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/31174.

Joly:2014:NSC

[JL14] Romain Joly and Camille Laurent. A note on the semiglobal controllability of the semilinear wave equation. SIAM Journal on Control and Optimization, 52(1):439–450, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jiang:2016:SGT

[JL16] Jifa Jiang and Xiang Lv. A small-gain theorem for nonlinear stochastic systems with inputs and outputs I: Additive white noise. SIAM Journal on Control and Optimization, 54(5):2383–2402, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jiang:2018:GSF

[JL18] Jifa Jiang and Xiang Lv. Global stability of feedback systems with multiplicative noise on the nonnegative orthant. SIAM Journal on Control and Optimization, 56(3): 2218–2247, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jackson:2023:NMC

[JL23] Joe Jackson and Gechun Liang. A new monotonicity condition for ergodic backward SDEs and ergodic control with superquadratic Hamiltonians. SIAM Journal on Control and Optimization, 61 (3):1273–1296, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1460958.

Jiang:2020:ESD

[JLL20] Bangxin Jiang, Jianquan Lu, and Yang Liu. Exponential stability of delayed systems with average-delay impulses. SIAM Journal on Control and Optimization, 58(6): 3763–3784, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jayawardhana:2009:ISS

[JLR09] Bayu Jayawardhana, Hartmut Logemann, and Eugene P. Ryan. Input-to-state stability of differential inclusions with applications to hysteretic and quantized feedback systems. SIAM Journal on Control and Optimization, 48(2):

1031–1054, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jacob:1998:CEO

[JLZ98] Birgit Jacob, Mikael Larsen, and Hans Zwart. Corrections and extensions of “Optimal Control of Linear Systems with Almost Periodic Input” by G. Da Prato and A. Ichikawa. SIAM Journal on Control and Optimization, 36(4):1473–1480, July 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/30461.  See

[DI87].

Jonsson:2005:SGT

[JM05] Ulf T. Jönsson and Alexandre Megretski. A small-gain theory for limit cycles of systems on Luré form. SIAM Journal on Control and Optimization, 44(3):909–938, May 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/43757.

Jungers:2017:FSL

[JM17] Raphaël M. Jungers and Paolo Mason. On feedback stabilization of linear switched systems via switching signal control. SIAM Journal on Control and Optimization, 55(2): 1179–1198, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jaber:2021:IOR

[JMP21] Eduardo Abi Jaber, Enzo Miller, and Huyen Pham. Integral operator Riccati equations arising in stochastic Volterra control problems. SIAM Journal on Control and Optimization, 59(2):1581– 1603, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jacob:2020:NLF

[JMPW20] Birgit Jacob, Andrii Mironchenko, Jonathan R. Partington, and Fabian Wirth. Noncoercive Lyapunov functions for inputto-state stability of infinitedimensional systems. SIAM Journal on Control and Optimization, 58(5):2952–2978, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See corrigendum [JMPW23].

Jacob:2023:CNL

[JMPW23] Birgit Jacob, Andrii Mironchenko, Jonathan R. Partington, and Fabian R. Wirth. Corrigendum: Noncoercive Lyapunov functions for inputto-state stability of infinitedimensional systems. SIAM Journal on Control and Optimization, 61(3):1910–1911, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1534559. See [JMPW20].

Jaskiewicz:2006:ZSE

[JN06] Anna Jáskiewicz and Andrzej S. Nowak. Zero-sum ergodic stochastic games with Feller transition probabilities. SIAM Journal on Control and Optimization, 45(3): 773–789, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jacka:2019:CDM

[JN19a] Saul D. Jacka and Dominykas Norgilas. On the compensator in the Doob–Meyer decomposition of the Snell envelope. SIAM Journal on Control and Optimization, 57(3): 1869–1889, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jaskiewicz:2019:CMD

[JN19b] Anna Jaśkiewicz and Andrzej S. Nowak. Constrained Markov decision processes with expected total reward criteria. SIAM Journal on Control and Optimization, 57(5): 3118–3136, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jaskiewicz:2024:MPE

[JN24] Anna Jaśkiewicz and Andrzej S. Nowak. On Markov perfect equilibria in discounted stochastic ARAT games. SIAM Journal on Control and Optimization, 62 (4):2148–2175, August 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1592365.

Joseph:2021:CNO

[JNKV21] Geethu Joseph, Buddhika Nettasinghe, Vikram Krishnamurthy, and Pramod K. Varshney. Controllability of network opinion in Erdős– Rényi graphs using sparse control inputs. SIAM Journal on Control and Optimization, 59(3):2321–2345, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jacob:2018:IDI

[JNPS18] Birgit Jacob, Robert Nabiullin, Jonathan R. Partington, and Felix L. Schwenninger. Infinite-dimensional input-to-state stability and Orlicz spaces. SIAM Journal on Control and Optimization, 56(2):868–889, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Johnson:2005:OSS

[Joh05] Stewart D. Johnson. Optimality of stasis and small switching cycles in planar systems with two-valued controls. SIAM Journal on Control and Optimization, 43(6): 1987–1999, November 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/42231.

Jonsson:2009:SCL

[Jön09] Ulf T. Jönsson. A sector condition for limit cycle robustness. SIAM Journal on Control and Optimization, 47(6): 2745–2772, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jourani:2000:EHE

[Jou00a]  Abderrahim Jourani.  Er-

ratum: “Hoffman’s Error Bound, Local Controllability, and Sensitivity Analysis”. SIAM Journal on Control and Optimization, 39 (5):1650, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/37046.  See

[Jou00b].

Jourani:2000:HEB

[Jou00b] Abderrahim Jourani. Hoffman’s error bound, local controllability, and sensitivity analysis. SIAM Journal on Control and Optimization, 38(3):947–970, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33921. See er-

ratum [Jou00a].

Jouan:2003:INS

[Jou03] Philippe Jouan. Immersion of nonlinear systems into linear systems modulo output injection. SIAM Journal on Control and Optimization, 41(6): 1756–1778, November 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39170.

Jacob:2001:BCS

[JP01] Birgit Jacob and Jonathan R.

Partington. On the boundedness and continuity of the spectral factorization mapping. SIAM Journal on Control and Optimization, 40(1):88–106, January 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/35184.

Jean:2015:CCA

[JP15] F. Jean and D. Prandi. Complexity of control-affine motion planning. SIAM Journal on Control and Optimization, 53 (2):816–844, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jacob:2014:ALC

[JPP14] Birgit Jacob, Jonathan R. Partington, and Sandra Pott. Applications of Laplace– Carleson embeddings to admissibility and controllability. SIAM Journal on Control and Optimization, 52(2): 1299–1313, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jelito:2020:LRR

[JPS20] Damian Jelito, Marcin Pitera, and Lukasz Stettner. Longrun risk-sensitive impulse control. SIAM Journal on Control and Optimization, 58(4): 2446–2468, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jaimungal:2024:MKL

[JPSB24] Sebastian Jaimungal, Silvana M. Pesenti, and Leandro Sánchez-Betancourt. Minimal Kullback–Leibler divergence for constrained Lévy–Itô processes. SIAM Journal on Control and Optimization, 62(2): 982–1005, March 2024. CO-DEN SJCODC. ISSN 1095-7138.

Jiang:1997:NNE

[JQ97] Houyuan Jiang and Liqun Qi. A new nonsmooth equations approach to nonlinear complementarity problems. SIAM Journal on Control and Optimization, 35(1):178–193, January 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27649.

Jongen:1993:NOC

[JR93] Hubertus Th. Jongen and Jan Rueckmann. Nonlinear optimization: On connected components of level sets. SIAM Journal on Control and Optimization, 31(1):86–100, January 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jakubczyk:2006:BPF

[JR06a] Bronisław Jakubczyk and Witold Respondek. Bifurcations of 1-parameter families of control-affine systems in the plane. SIAM Journal on Control and Optimization, 44(6): 2038–2062, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jobert:2006:OPM

[JR06b] A. Jobert and L. C. G.

Rogers. Option pricing with Markov-modulated dynamics. SIAM Journal on Control and Optimization, 44(6):2063– 2078, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jonckheere:1981:STL

[JS81] Edmond A. Jonckheere and Leonard M. Silverman. Spectral theory of the linearquadratic optimal control problem: Analytic factorization of rational matrix-valued functions. SIAM Journal on Control and Optimization, 19 (2):262–281, March 1981. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jurdjevic:1984:CPA

[JS84] Velimir Jurdjevic and Gauthier Sallet. Controllability properties of affine systems. SIAM Journal on Control and Optimization, 22(3):501–508, May 1984. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jahn:1986:GQM

[JS86] J. Jahn and E. Sachs. Generalized quasiconvex mappings and vector optimization. SIAM Journal on Control and Optimization, 24(2): 306–322, March 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jakubczyk:1990:CND

[JS90] Bronisław Jakubczyk and Eduardo D. Sontag. Controllability of nonlinear discretetime systems: a Lie-algebraic approach. SIAM Journal on Control and Optimization, 28 (1):1–33, January 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See erratum [JS91].

Jakubczyk:1991:ECN

[JS91] Bronisław Jakubczyk and Eduardo D. Sontag. Erratum: “Controllability of nonlinear discrete-time systems: a Liealgebraic approach”. SIAM Journal on Control and Optimization, 29(2):498, March 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [JS90].

James:1999:CRF

[JS99] François James and Mauricio Sepúlveda. Convergence results for the flux identification in a scalar conservation law. SIAM Journal on Control and Optimization, 37(3):869–891, May 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27272.

Janecek:2012:OIH

[JS12] Karel Janecek and Mihai Sîrbu. Optimal investment with high-watermark performance fee. SIAM Journal on Control and Optimization, 50 (2):790–819, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jelito:2024:LRI

[JS24] Damian Jelito and Łukasz Stettner. Long-run impulse control with generalized discounting. SIAM Journal on Control and Optimization, 62 (2):853–876, March 2024. CO-DEN SJCODC. ISSN 1095-7138.

James:2005:GMR

[JSV05] M. R. James, M. C. Smith, and G. Vinnicombe. Gap metrics, representations, and nonlinear robust stability. SIAM Journal on Control and Optimization, 43(5): 1535–1582, September 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39306.

Jourani:2011:NAL

[JT11] Abderrahim Jourani and Lionel Thibault. Noncoincidence of approximate and limiting subdifferentials of integral functionals. SIAM Journal on Control and Optimization, 49(4):1435–1453, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sicon/
resource/1/sjcodc/v49/i4/
p1435_s1.

Jung:2000:ACO

[Jun00] Michael Jung. Admissibility of control operators for solution families to Volterra integral equations. SIAM Journal on Control and Optimization, 38(5):1323–1333, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32819.

Jain:2006:SBU

[JV06] Rahul Jain and Pravin P.

Varaiya. Simulation-based uniform value function estimates of Markov decision processes. SIAM Journal on Control and Optimization, 45(5): 1633–1656, ???? 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jourani:2017:GLM

[JV17] Abderrahim Jourani and Emilio Vilches. Galerkin-like method and generalized perturbed sweeping process with nonregular sets. SIAM Journal on Control and Optimization, 55(4):2412–2436, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jakovetic:2023:DRE

[JVB^{+}23] Dusan Jakovetic, Manojlo Vukovic, Dragana Bajovic, Anit Kumar Sahu, and Soummya Kar. Distributed recursive estimation under heavy-tail communication noise. SIAM Journal on Control and Optimization, 61(3):1582–1609, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1477015.

Ji:2022:MMS

[JX22] Shaolin Ji and Rundong Xu. A modified method of successive approximations for stochastic recursive optimal control problems. SIAM Journal on Control and Optimization, 60 (5):2759–2786, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1418204.

Jin:2024:OMH

[JXZ24] Zhuo Jin, Zuo Quan Xu, and Bin Zou. Optimal moralhazard-free reinsurance under extended distortion premium principles. SIAM Journal on Control and Optimization, 62(3):1390–1416, June 2024. CODEN SJCODC. ISSN 1095-7138.

Jacob:1999:ECS

[JZ99] Birgit Jacob and Hans Zwart. Equivalent conditions for stabilizability of infinitedimensional systems with admissible control operators. SIAM Journal on Control and Optimization, 37(5): 1419–1455, September 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/33344.

Jacob:2004:CCO

[JZ04] Birgit Jacob and Hans Zwart. Counterexamples concerning observation operators for C_{0}- semigroups. SIAM Journal on Control and Optimization, 43(1):137–153, January 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42323.

Jacob:2009:HTE

[JZ09a] Birgit Jacob and Hans Zwart.

On the hautus test for exponentially stable C_{0}-groups. SIAM Journal on Control and Optimization, 48(3):1275– 1288, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jin:2009:NEP

[JZ09b] Bangti Jin and Jun Zou. Numerical estimation of piecewise constant robin coefficient. SIAM Journal on Control and Optimization, 48(3): 1977–2002, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jin:2021:NED

[JZ21] Bangti Jin and Zhi Zhou. Numerical estimation of a diffusion coefficient in subdiffusion. SIAM Journal on Control and Optimization, 59(2): 1466–1496, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jackson:2022:EUN

[JZ22] Joe Jackson and Gordan Zitković. Existence and uniqueness for non-Markovian triangular quadratic BSDEs. SIAM Journal on Control and Optimization, 60(3):1642– 1666, ???? 2022. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1435689.

Jing:2018:WRT

[JZLW18] Gangshan Jing, Guofeng Zhang, Heung Wing Joseph Lee, and Long Wang. Weak rigidity theory and its application to formation stabilization. SIAM Journal on Control and Optimization, 56(3): 2248–2273, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Klabjan:2006:EOP

[KA06] Diego Klabjan and Daniel Adelman. Existence of optimal policies for semi-Markov decision processes using duality for infinite linear programming. SIAM Journal on Control and Optimization, 44(6): 2104–2122, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kalman:1982:CRO

[Kal82] R. E. Kalman. On the computation of the reachable/ observable canonical form. SIAM Journal on Control and Optimization, 20(2):258– 260, March 1982. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kamen:1976:MSI

[Kam76] Edward W. Kamen. Module structure of infinitedimensional systems with applications to controllability. SIAM Journal on Control and Optimization, 14(3):389–408, May 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kamizono:2003:PHU

[Kam03]  Kenji Kamizono.  Partial

hedging under transaction costs. SIAM Journal on Control and Optimization, 42(5): 1545–1558, September 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40514.

Kamal:2010:CLP

[Kam10]  Sameer Kamal.  On the

convergence, lock-in probability, and sample complexity of stochastic approximation. SIAM Journal on Control and Optimization, 48(8): 5178–5192, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kaneko:1978:MPR

[Kan78] Ikuyo Kaneko. A maximization problem related to parametric linear complementarity. SIAM Journal on Control and Optimization, 16(1): 41–55, January 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kannai:1982:CSL

[Kan82] Yakar Kannai. Causality and stability of linear systems described by partial differential operators. SIAM Journal on Control and Optimization, 20(5):669–674, September 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kang:1998:BNFa

[Kan98a] Wei Kang. Bifurcation and normal form of nonlinear control systems, Part I. SIAM Journal on Control and Optimization, 36(1):193–212, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/29028.

Kang:1998:BNFb

[Kan98b] Wei Kang. Bifurcation and normal form of nonlinear control systems, Part II. SIAM Journal on Control and Optimization, 36(1):213–232, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29029.

Kang:2000:BCS

[Kan00] Wei Kang. Bifurcation control via state feedback for systems with a single uncontrollable mode. SIAM Journal on Control and Optimization, 38(5):1428–1452, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32592.

Kapitonov:1994:SEB

[Kap94]  B. V. Kapitonov.  Stabi-

lization and exact boundary controllability for Maxwell’s equations. SIAM Journal on Control and Optimization, 32(2):408–420, March 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/21848.

Karatzas:1989:OPT

[Kar89] Ioannis Karatzas. Optimization problems in the theory of continuous trading. SIAM Journal on Control and Optimization, 27(6):1221– 1259, November 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Karafyllis:2006:FTG

[Kar06] Iasson Karafyllis. Finite-time global stabilization by means of time-varying distributed delay feedback. SIAM Journal on Control and Optimization, 45(1):320–342, January 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Karafyllis:2011:SMA

[Kar11]  Iasson Karafyllis.  Stabi-

lization by means of approximate predictors for systems with delayed input. SIAM Journal on Control and Optimization, 49(3):1100– 1123, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicon/resource/
                                        1/sjcodc/v49/i3/p1100_s1.

Kaskosz:1977:NEP

[Kaś77] Barbara Kaśkosz. On a nonlinear evasion problem. SIAM Journal on Control and Optimization, 15(4):661–673, July 1977. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kassara:2002:FSC

[Kas02] K. Kassara. Feedback spreading control under speed constraints. SIAM Journal on Control and Optimization, 41(4):1281–1294, July 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36991.

Kassara:2009:USV

[Kas09] Khalid Kassara. A unified set-valued approach to control immunotherapy. SIAM Journal on Control and Optimization, 48(2):909–924, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kassara:2013:SVA

[Kas13]  Khalid Kassara.  A set-

valued approach to observability. SIAM Journal on Control and Optimization, 51(6): 4520–4543, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Katzourakis:2019:IOT

[Kat19] Nikos Katzourakis. Inverse optical tomography through PDE constrained optimization L^{∞}. SIAM Journal on Control and Optimization, 57(6): 4205–4233, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kawasaki:2001:CPT

[Kaw01] H. Kawasaki. A conjugate points theory for a nonlinear programming problem. SIAM Journal on Control and Optimization, 40(1):54–63, January 2001. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36883.

Kawan:2011:IEC

[Kaw11] Christoph Kawan. Invariance entropy of control sets. SIAM Journal on Control and Optimization, 49(2):732– 751, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i2/p732_s1.

Kort:1976:CPD

[KB76] Barry W. Kort and Dimitri P. Bertsekas. Combined primal–dual and penalty methods for convex programming. SIAM Journal on Control and Optimization, 14(2): 268–294, February 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Karatzas:1981:DMF

[KB81] Ioannis Karatzas and Václav E. Benes. A degree method for free boundaries in stochastic control. SIAM Journal on Control and Optimization, 19 (3):283–332, May 1981. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Konda:1999:ACT

[KB99] Vijaymohan R. Konda and Vivek S. Borkar. Actor– critic-type learning algorithms for Markov decision processes. SIAM Journal on Control and Optimization, 38(1):94–123, January 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31669.

Khodja:2001:SSO

[KB01] Farid Ammar Khodja and Ahmed Bader. Stabilizability of systems of one-dimensional wave equations by one internal or boundary control force. SIAM Journal on Control and Optimization, 39(6): 1833–1851, ???? 2001. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kannapan:2016:SPC

[KB16] Deepti Kannapan and Francesco Bullo. Synchronization in pulse-coupled oscillators with delayed Excitatory/Inhibitory coupling. SIAM Journal on Control and Optimization, 54 (4):1872–1894, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Khodja:2003:CTP

[KBDK03] F. Ammar Khodja, A. Benabdallah, C. Dupaix, and I. Kostin. Controllability to the trajectories of phasefield models by one control force. SIAM Journal on Control and Optimization, 42(5): 1661–1680, September 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/41782.

Karvonen:2020:SCF

[KBMS20] Toni Karvonen, Silvère Bonnabel, Eric Moulines, and Simo Särkkä. On stability of a class of filters for nonlinear stochastic systems. SIAM Journal on Control and Optimization, 58(4):2023–2049, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Koerts:2017:TGC

[KBvdSD17] Filip J. Koerts, Mathias Bürger, Arjan J. van der Schaft, and Claudio De Persis. Topological and graphcoloring conditions on the parameter-independent stability of second-order networked systems. SIAM Journal on Control and Optimization, 55 (6):3750–3778, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kim:2010:LEC

[KC10] Tujin Kim and Daomin Cao. Local exact controllability of the Navier–Stokes equations with the condition on the pressure on parts of the boundary. SIAM Journal on Control and Optimization, 48(6):

3805–3837, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kaba:2015:SCC

[KC15] M. D. Kaba and M. K. Camlibel. A spectral characterization of controllability for linear discrete-time systems with conic constraints. SIAM Journal on Control and Optimization, 53(4):2350–2372, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kitsos:2021:OFC

[KCBP21] Constantinos Kitsos, Eduardo Cerpa, Gildas Besançon, and Christophe Prieur. Output feedback control of a cascade system of linear Korteweg–de Vries equations. SIAM Journal on Control and Optimization, 59(4):2955–2976, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kalyanam:2016:PMT

[KCP16] Krishnamoorthy Kalyanam, David W. Casbeer, and Meir Pachter. Pursuit of a moving target with known constant speed on a directed acyclic graph under partial information. SIAM Journal on Control and Optimization, 54(5): 2259–2273, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Khan:2022:ETO

[KCY22] Gulam D. Khan, Zhiyong Chen, and Yamin Yan. Eventtriggered output synchronization of heterogeneous nonlinear MultiAgents. SIAM Journal on Control and Optimization, 60(1):210–232, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1371488.

Krichene:2015:OLN

[KDB15] Walid Krichene, Benjamin Drighès, and Alexandre M. Bayen. Online learning of Nash equilibria in congestion games. SIAM Journal on Control and Optimization, 53(2): 1056–1081, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Karafyllis:2021:ETG

[KEK21] Iasson Karafyllis, Nicolás Espitia, and Miroslav Krstic. Event-triggered gain scheduling of reaction-diffusion PDEs. SIAM Journal on Control and Optimization, 59(3):2047– 2067, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Keller:2024:MVT

[Kel24] Christian Keller. Mean viability theorems and secondorder Hamilton–Jacobi equations. SIAM Journal on Control and Optimization, 62 (3):1615–1642, June 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1550438.

Kennedy:1980:COP

[Ken80] D. P. Kennedy. On characterizing optimal policies in infinite-horizon stochastic control. SIAM Journal on Control and Optimization, 18(5):576– 584, September 1980. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kaneko:2003:WSF

[KF03] Osamu Kaneko and Takao Fujii. When is a storage function a state function in discrete time? SIAM Journal on Control and Optimization, 42(4): 1374–1394, January 2003. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kien:2022:OCP

[KFP22] B. T. Kien, V. E. Fedorov, and T. D. Phuong. Optimal control problems governed by fractional differential equations with control constraints. SIAM Journal on Control and Optimization, 60 (3):1732–1762, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1430728.

Kabaila:1980:EPO

[KG80] Paul V. Kabaila and Graham C. Goodwin. On the estimation of the parameter of an optimal interpolator when the class of interpolators is restricted. SIAM Journal on Control and Optimization, 18 (2):121–144, March 1980. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kumar:1995:OSC

[KG95] Ratnesh Kumar and Vijay K. Garg. Optimal supervisory control of discrete event dynamical systems. SIAM Journal on Control and Optimization, 33(2):419–439, March 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/23518.

Klintberg:2017:NSH

[KG17] Emil Klintberg and Sebastien Gros. Numerical structure of the Hessian of the Lagrange dual function for a class of convex problems. SIAM Journal on Control and Optimization, 55(1):574–593, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kipka:2019:DTG

[KG19] Robert Kipka and Rohit Gupta. The discrete-time geometric maximum principle. SIAM Journal on Control and Optimization, 57(4): 2939–2961, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kang:2022:FNN

[KG22] Wei Kang and Qi Gong.

Feedforward neural networks and compositional functions with applications to dynamical systems. SIAM Journal on Control and Optimization, 60(2):786–813, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1391596.

Krstic:2008:CTF

[KGBS08] Miroslav Krstic, Bao-Zhu Guo, Andras Balogh, and Andrey Smyshlyaev. Control of a tip-force destabilized shear beam by observer-based boundary feedback. SIAM Journal on Control and Optimization, 47(2):553–574, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kumar:1993:LSS

[KGM93] Ratnesh Kumar, Vijay Garg, and Steven I. Marcus. Language stability and stabilizability of discrete event dynamical systems. SIAM Journal on Control and Optimization, 31(5):1294–1320, September 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Krstic:2011:BCO

[KGS11] Miroslav Krstic, Bao-Zhu Guo, and Andrey Smyshlyaev. Boundary controllers and observers for the linearized Schrödinger equation. SIAM Journal on Control and Optimization, 49(4):1479–1497, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sicon/
                                        resource/1/sjcodc/v49/i4/
                                        p1479_s1.

Krasnakevich:1969:GPC

[KH69] J. R. Krasnakevich and R. A. Haddad. Generalized prediction-correction estimation. SIAM Journal on Control, 7(3):496–511, August 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic). See comments [Wei77].

Kamen:1979:ATL

[KH79a] Edward W. Kamen and Khaled M. Hafez. Algebraic theory of linear time-varying systems. SIAM Journal on Control and Optimization, 17 (4):500–510, July 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See erratum [KH79b].

Kamen:1979:EAT

[KH79b] Edward W. Kamen and Khaled M. Hafez. Erratum: Algebraic theory of linear time-varying systems. SIAM Journal on Control and Optimization, 17(6): 788, November 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [KH79a].

Kushner:1979:RCS

[KH79c] Harold J. Kushner and Hai Huang. Rates of convergence for stochastic approximation type algorithms. SIAM Journal on Control and Optimization, 17(5):607–617, September 1979. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kushner:1981:AMA

[KH81a] H. J. Kushner and Hai Huang.

Averaging methods for the asymptotic analysis of learning and adaptive systems, with small adjustment rate. SIAM Journal on Control and Optimization, 19(5):635– 650, September 1981. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kushner:1981:APS

[KH81b] Harold J. Kushner and Hai Huang. Asymptotic properties of stochastic approximations with constant coefficients. SIAM Journal on Control and Optimization, 19(1):

87–105, January 1981. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kenney:1990:SAD

[KH90] Charles Kenney and Gary Hewer. The sensitivity of the algebraic and differential Riccati equations. SIAM Journal on Control and Optimization, 28(1):50–69, January 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Khargonekar:1982:MFR

[Kha82] P. P. Khargonekar. On matrix fraction representations for linear systems over commutative rings. SIAM Journal on Control and Optimization, 20 (2):172–197, March 1982. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Khanh:1993:OCN

[Khá93] Phan Quoc Khánh. Optimality conditions via norm scalarization in vector optimization. SIAM Journal on Control and Optimization, 31(3):646–658, May 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Khapalov:1994:EOH

[Kha94] Alexander Khapalov. L^{∞}- exact observability of the heat equation with scanning pointwise sensor. SIAM Journal on Control and Optimization, 32(4):1037–1051, July 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/22737.

Khapalov:1995:EOT

[Kha95] A. Yu. Khapalov. Exact observability of the timevarying hyperbolic equation with finitely many moving internal observations. SIAM Journal on Control and Optimization, 33(4):1256–1269, July 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23621.

Khapalov:2001:MPC

[Kha01] Alexander Khapalov. Mobile point controls versus locally distributed ones for the controllability of the semilinear parabolic equation. SIAM Journal on Control and Optimization, 40(1):231–252, January 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35803.

Khapalov:2003:CSP

[Kha03] A. Y. Khapalov. Controllability of the semilinear parabolic equation governed by a multiplicative control in the reaction term: a qualitative approach. SIAM Journal on Control and Optimization, 41(6):1886–1900, November 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/39460.

Khapalov:2007:LCS

[Kha07] A. Y. Khapalov. Local controllability for a “Swimming” model. SIAM Journal on Control and Optimization, 46(2): 655–682, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kharroubi:2016:OSF

[Kha16] Idris Kharroubi. Optimal switching in finite horizon under state constraints. SIAM Journal on Control and Optimization, 54(4):2202–2233, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Korda:2014:CCM

[KHJ14] Milan Korda, Didier Henrion, and Colin N. Jones. Convex computation of the maximum controlled invariant set for polynomial control systems. SIAM Journal on Control and Optimization, 52(5): 2944–2969, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Karow:2006:ISU

[KHP06] Michael Karow, Diederich Hinrichsen, and Anthony J.

Pritchard. Interconnected systems with uncertain couplings: Explicit formulae for μ-values, spectral value sets, and stability radii. SIAM Journal on Control and Optimization, 45(3):856–884, January 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kimura:1978:DFO

[Kim78]  Hidenori Kimura.  Dead-

beat function observers for discrete-time linear systems. SIAM Journal on Control and Optimization, 16(6):880– 894, November 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kime:1990:BCM

[Kim90] Katherine A. Kime. Boundary controllability of Maxwell’s equations in a spherical region. SIAM Journal on Control and Optimization, 28(2): 294–319, March 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kim:1991:EDE

[Kim91] Jong Uhn Kim. Exponential decay of the energy of a onedimensional nonhomogeneous medium. SIAM Journal on Control and Optimization, 29 (2):368–380, March 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kim:1992:ESI

[Kim92] Jong Uhn Kim. Exact semiinternal control of an Euler– Bernoulli equation. SIAM Journal on Control and Optimization, 30(5):1001–1023, September 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kim:1993:CSO

[Kim93a]  Jong Uhn Kim.  Control

of a second-order integrodifferential equation. SIAM Journal on Control and Optimization, 31(1):101–110, January 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kim:1993:TRB

[Kim93b] Jong Uhn Kim. Trace regularity in the boundary control of a wave equation. SIAM Journal on Control and Optimization, 31(6):1479–1501, November 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kim:2021:CSH

[Kim21] Dohyun Kim. Cluster synchrony of high-dimensional Kuramoto models with higherorder couplings. SIAM Journal on Control and Optimization, 59(6):4110–4135, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kim:2022:CRA

[Kim22] Doheon Kim. Continuous rendezvous algorithm for memoryless agents with limited visibility in the Euclidean space. SIAM Journal on Control and Optimization, 60(2):732–757, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1387584.

Kindler:1984:EPT

[Kin84] Jürgen Kindler. Equilibrium point theorems for two-person games. SIAM Journal on Control and Optimization, 22 (5):671–683, September 1984. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kiwiel:1989:ETR

[Kiw89] Krzysztof C. Kiwiel. An ellipsoid trust region bundle method for nonsmooth convex minimization. SIAM Journal on Control and Optimization, 27(4):737–757, July 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kiwiel:1996:ESPa

[Kiw96a] Krzysztof C. Kiwiel. The efficiency of subgradient projection methods for convex optimization, Part I: General level methods. SIAM Journal on Control and Optimization, 34 (2):660–676, March 1996. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kiwiel:1996:ESPb

[Kiw96b] Krzysztof C. Kiwiel. The efficiency of subgradient projection methods for convex optimization, Part II: Implementations and extensions. SIAM Journal on Control and Optimization, 34(2):677–697, March 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26148.

Kiwiel:1997:PMM

[Kiw97] Krzysztof C. Kiwiel. Proximal minimization methods with generalized Bregman functions. SIAM Journal on Control and Optimization, 35(4):1142–1168, July 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28174.

Karafyllis:2007:SGT

[KJ07] Iasson Karafyllis and Zhong-Ping Jiang. A small-gain theorem for a wide class of feedback systems with control applications. SIAM Journal on Control and Optimization, 46 (4):1483–1517, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kushner:1976:SAA

[KK76] Harold J. Kushner and Milton L. Kelmanson. Stochastic approximation algorithms of the multiplier type for the sequential Monte Carlo optimization of stochastic systems. SIAM Journal on Control and Optimization, 14(5): 827–842, August 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Khalil:1979:SMP

[KK79] Hassan K. Khalil and Petar V.

Kokotovic. D-stability and multi-parameter singular perturbation. SIAM Journal on Control and Optimization, 17 (1):56–65, January 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kang:1992:EQC

[KK92] Wei Kang and Arthur J.

Krener. Extended quadratic controller normal form and dynamic state feedback linearization of nonlinear systems. SIAM Journal on Control and Optimization, 30(6): 1319–1337, November 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Korn:2002:SCA

[KK02] Ralf Korn and Holger Kraft.

A stochastic control approach to portfolio problems with stochastic interest rates. SIAM Journal on Control and Optimization, 40(4):1250–1269, July 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37779.

Krener:2003:LCN

[KK03] Arthur J. Krener and Wei Kang. Locally convergent nonlinear observers. SIAM Journal on Control and Optimization, 42(1):155–177, January 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36861.

Krishnamurthy:2006:USP

[KK06] P. Krishnamurthy and F. Khorrami. On uniform solvability of parameter-dependent Lyapunov inequalities and applications to various problems. SIAM Journal on Control and Optimization, 45(4): 1147–1164, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Karafyllis:2012:GSF

[KK12] Iasson Karafyllis and Miroslav Krstic. Global stabilization of feedforward systems under perturbations in sampling schedule. SIAM Journal on Control and Optimization, 50 (3):1389–1412, ???? 2012.

CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kovtunenko:2014:HPI

[KK14] Victor A. Kovtunenko and Karl Kunisch. High precision identification of an object: Optimality-conditions-based concept of imaging. SIAM Journal on Control and Optimization, 52(1):773–796, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Karafyllis:2016:SDS

[KK16] Iasson Karafyllis and Miroslav Krstic. Sampled-data stabilization of nonlinear delay systems with a compact absorbing set. SIAM Journal on Control and Optimization, 54 (2):790–818, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Karafyllis:2017:GDS

[KK17a] Iasson Karafyllis and Miroslav Krstic. Global dynamical solvers for nonlinear programming problems. SIAM Journal on Control and Optimization, 55(2):1302–1331, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Karafyllis:2017:IDN

[KK17b] Iasson Karafyllis and Miroslav Krstic. ISS in different norms for 1-D parabolic PDEs with boundary disturbances. SIAM Journal on Control and Optimization, 55(3):1716–1751, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kentia:2018:NEG

[KK18] Klebert Kentia and Christoph Kühn. Nash equilibria for game contingent claims with utility-based hedging. SIAM Journal on Control and Optimization, 56(6):3948–3972, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Karafyllis:2019:GSC

[KK19a] Iasson Karafyllis and Miroslav Krstic. Global stabilization of a class of nonlinear reactiondiffusion partial differential equations by boundary feedback. SIAM Journal on Control and Optimization, 57(6): 3723–3748, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Karafyllis:2019:SGB

[KK19b] Iasson Karafyllis and Miroslav Krstic. Small-gain-based boundary feedback design for global exponential stabilization of one-dimensional semilinear parabolic PDEs. SIAM Journal on Control and Optimization, 57(3):2016–2036, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kafetzis:2023:SSS

[KK23] Ioannis Kafetzis and Nicholas Karampetakis. On the stabilization of singular, strictly proper transfer function matrices. SIAM Journal on Control and Optimization, 61 (3):1297–1321, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1429199.

Karafyllis:2024:CLV

[KK24] Iasson Karafyllis and Miroslav Krstic. Control of a linearized viscous liquid-tank system with surface tension. SIAM Journal on Control and Optimization, 62(2):1034–1059, March 2024. CODEN SJ-CODC. ISSN 1095-7138.

Karafyllis:2023:DSA

[KKA23] Iasson Karafyllis, Miroslav Krstic, and Alexandros Aslanidis. On disturbance-to-state adaptive stabilization without parameter bound by nonlinear feedback of delayed state and input. SIAM Journal on Control and Optimization, 61(6): 3584–3607, December 2023. CODEN SJCODC. ISSN 1095-7138.

Kirpichnikova:2021:CAP

[KKLO21] Anna Kirpichnikova, Jussi Korpela, Matti J. Lassas, and Lauri Oksanen. Construction of artificial point sources for a linear wave equation in unknown medium. SIAM Journal on Control and Optimization, 59(5):3737–3761, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kamen:1985:TFA

[KKP85] E. W. Kamen, P. P. Khargonekar, and K. R. Poolla. A transfer-function approach to linear time-varying discretetime systems. SIAM Journal on Control and Optimization, 23(4):550–565, July 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Korobov:1993:OCP

[KKS93] V. I. Korobov, V. I. Krutin′, and G. M. Sklyar. An optimal control problem with a mixed cost function. SIAM Journal on Control and Optimization, 31(3):624–645, May 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kroner:2011:SNM

[KKV11] Axel Kröner, Karl Kunisch, and Boris Vexler. Semismooth Newton methods for optimal control of the wave equation with control constraints. SIAM Journal on Control and Optimization, 49 (2):830–858, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.

siam.org/sicon/resource/
1/sjcodc/v49/i2/p830_s1.

Komornik:1999:ITT

[KL99] Vilmos Komornik and Paola Loreti. Ingham-type theorems for vector-valued functions and observability of coupled linear systems. SIAM Journal on Control and Optimization, 37(2):461–485, March 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31750.

Krener:2002:NFB

[KL02] Arthur J. Krener and Long Li. Normal forms and bifurcations of discrete time nonlinear control systems. SIAM Journal on Control and Optimization, 40(6):1697–1723, November 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37898.

Kovtunenko:2016:STC

[KL16] Victor A. Kovtunenko and Günter Leugering. A shapetopological control problem for nonlinear crack-defect interaction: The antiplane variational model. SIAM Journal on Control and Optimization, 54(3):1329–1351, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kmit:2021:FTS

[KL21] Irina Kmit and Natalya Lyul’ko. Finite time stabilization of nonautonomous first-order hyperbolic systems. SIAM Journal on Control and Optimization, 59(5):3179– 3202, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Klein:2018:BED

[Kle18] Guillaume Klein. Best exponential decay rate of energy for the vectorial damped wave equation. SIAM Journal on Control and Optimization, 56(5):3432–3453, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kunisch:2010:OCE

[KLL10] Karl Kunisch, Kewei Liang, and Xiliang Lu. Optimal control for an elliptic system with polygonal state constraints. SIAM Journal on Control and Optimization, 48 (8):5053–5072, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Komornik:2004:NCF

[KLM04] Vilmos Komornik, Paola Loreti, and Leonardo Mazzini. Null controllability of the free surface of a liquid in a container. SIAM Journal on Control and Optimization, 43(2):685–696, March 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40528.

Kenne:2020:OCP

[KLM20a] Cyrille Kenne, Günter Leugering, and Gisèle Mophou. Optimal control of a population dynamics model with missing birth rate. SIAM Journal on Control and Optimization, 58(3):1289–1313, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kharroubi:2020:RRR

[KLM20b] Idris Kharroubi, Thomas Lim, and Thibaut Mastrolia. Regulation of renewable resource exploitation. SIAM Journal on Control and Optimization, 58(1):551–579, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kirches:2020:APT

[KLM20c] C. Kirches, F. Lenders, and P. Manns. Approximation properties and tight bounds for constrained mixed-integer optimal control. SIAM Journal on Control and Optimization, 58(3):1371–1402, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Krug:2021:TDD

[KLM^{+}21] Richard Krug, Günter Leugering, Alexander Martin, Martin Schmidt, and Dieter Weninger. Time-domain decomposition for optimal control problems governed by semilinear hyperbolic systems. SIAM Journal on Control and Optimization, 59(6):4339– 4372, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kotsalis:2021:COF

[KLN21] Georgios Kotsalis, Guanghui Lan, and Arkadi S. Nemirovski. Convex optimization for finite-horizon robust covariance control of linear stochastic systems. SIAM Journal on Control and Optimization, 59(1):296–319, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ke:2009:ATD

[KLR09] Zhenqing Ke, Hartmut Logemann, and Richard Rebarber. Approximate tracking and disturbance rejection for stable infinite-dimensional systems using sampled-data lowgain control. SIAM Journal on Control and Optimization, 48 (2):641–671, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Karatzas:1987:OPC

[KLS87] Ioannis Karatzas, John P.

Lehoczky, and Steven E. Shreve. Optimal portfolio and consumption decisions for a “Small investor” on a finite horizon. SIAM Journal on Control and Optimization, 25(6):1557–1586, November 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Karatzas:1991:MDM

[KLSX91] Ioannis Karatzas, John P.

Lehoczky, Steven E. Shreve, and Gan-Lin Xu. Martingale and duality methods for utility maximization in an incomplete market. SIAM Journal on Control and Optimization, 29(3):702–730, May 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kleptsyna:2009:LEF

[KLV09] M. L. Kleptsyna, A. Le Breton, and M. Viot. On the linear-exponential filtering problem for general Gaussian processes. SIAM Journal on Control and Optimization, 47 (6):2886–2911, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kolodziej:1986:SEC

[KM86] Wojciech J. Kolodziej and Ronald R. Mohler. State estimation and control of conditionally linear systems. SIAM Journal on Control and Optimization, 24(3):497–508, May 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kushner:1991:NMS

[KM91] Harold J. Kushner and Luiz Felipe Martins. Numerical methods for stochastic singular control problems. SIAM Journal on Control and Optimization, 29(6):1443–1475, November 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kushner:1996:HTA

[KM96] Harold J. Kushner and L. Felipe Martins. Heavy traffic analysis of a controlled multiclass queueing network via weak convergence methods. SIAM Journal on Control and Optimization, 34(5): 1781–1797, September 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27559.

Koon:1997:OCH

[KM97] Wang-Sang Koon and Jerrold E. Marsden. Optimal control for holonomic and nonholonomic mechanical systems with symmetry and Lagrangian reduction. SIAM Journal on Control and Optimization, 35(3):901–929, May 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29036.

Kurina:2004:LQO

[KM04] Galina A. Kurina and Roswitha März. On linear-quadratic optimal control problems for time-varying descriptor systems. SIAM Journal on Control and Optimization, 42(6): 2062–2077, November 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38099.

Kao:2007:NBF

[KM07a] Chung-Yao Kao and Alexandre Megretski. On the new barrier function and specialized algorithms for a class of semidefinite programs. SIAM Journal on Control and Optimization, 46(2):468–495, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kurina:2007:FSO

[KM07b] Galina A. Kurina and Roswitha März. Feedback solutions of optimal control problems with DAE constraints. SIAM Journal on Control and Optimization, 46(4):1277–1298, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kawaguchi:2012:IIP

[KM12a] Kazuhito Kawaguchi and Hiroaki Morimoto. An irreversible investment problem with maintenance expenditure. SIAM Journal on Control and Optimization, 50(3): 1244–1264, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kristalny:2012:IOS

[KM12b] Maxim Kristalny and Leonid Mirkin. Input/output stabilization in the general twosided model matching setup. SIAM Journal on Control and Optimization, 50(3):1413– 1438, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kaise:2016:IEC

[KM16] Hidehiro Kaise and William M. McEneaney. Idempotent expansions for continuous-time stochastic control. SIAM Journal on Control and Optimization, 54(1):73–98, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Krishnan:2018:DCS

[KM18] Vishaal Krishnan and Sonia Martínez. Distributed control for spatial self-organization of multi-agent swarms. SIAM Journal on Control and Optimization, 56(5):3642–3667, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kafnemer:2021:WIS

[KMC21] Meryem Kafnemer, Benmiloud Mebkhout, and Yacine Chitour. Weak input to state estimates for 2D damped wave equations with localized and nonlinear damping. SIAM Journal on Control and Optimization, 59(2):1604–1627, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kelbert:2019:HEG

[KMF19] Mark Kelbert and Harold A.

Moreno-Franco. HJB equations with gradient constraint associated with controlled jump-diffusion processes. SIAM Journal on Control and Optimization, 57(3): 2185–2213, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kapitonov:2007:BOE

[KMM07] B. Kapitonov, B. Miara, and G. Perla Menzala. Boundary observation and exact control of a quasi-electrostatic piezoelectric system in multilayered media. SIAM Journal on Control and Optimization, 46(3): 1080–1097, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Khodja:2011:EBC

[KMM11] F. Ammar Khodja, K. Mauffrey, and A. Münch. Exact boundary controllability of a system of mixed order with essential spectrum. SIAM Journal on Control and Optimization, 49(4):1857– 1879, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicon/resource/
                                        1/sjcodc/v49/i4/p1857_s1.

Kar:2013:DLP

[KMP13] Soummya Kar, José M. F. Moura, and H. Vincent Poor. Distributed linear parameter estimation: Asymptotically efficient adaptive strategies. SIAM Journal on Control and Optimization, 51(3): 2200–2229, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Krumbiegel:2010:PEA

[KMR10] K. Krumbiegel, C. Meyer, and A. Rösch. A priori error analysis for linear quadratic elliptic Neumann boundary control problems with control and state constraints. SIAM Journal on Control and Optimization, 48(8):5108–5142, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Knudsen:1998:VIR

[KMZ98] Thomas S. Knudsen, Bernhard Meister, and Mihail Zervos. Valuation of investments in real assets with implications for the stock prices. SIAM Journal on Control and Optimization, 36(6): 2082–2102, November 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31581.

Kitamura:1977:ISV

[KN77] S. Kitamura and S. Nakagiri. Identifiability of spatiallyvarying and constant parameters in distributed systems of parabolic type. SIAM Journal on Control and Optimization, 15(5):785–802, August 1977. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See comments [Cou83].

Kindermann:2003:PIR

[KN03] Stefan Kindermann and Andreas Neubauer. Parameter identification by regularization for surface representation via the moving grid approach. SIAM Journal on Control and Optimization, 42 (4):1416–1430, January 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kaya:2013:FIC

[KN13] C. Yalçin Kaya and J. Lyle Noakes. Finding interpolating curves minimizing L^{∞} acceleration in the Euclidean space via optimal control theory. SIAM Journal on Control and Optimization, 51(1): 442–464, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kien:2014:SON

[KN14a] B. T. Kien and V. H. Nhu. Second-order necessary optimality conditions for a class of semilinear elliptic optimal control problems with mixed pointwise constraints. SIAM Journal on Control and Optimization, 52(2):1166–1202, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Koch:2014:CTD

[KN14b] Ronald Koch and Ebrahim Nasrabadi. Continuous-time dynamic shortest path problems with negative transit times. SIAM Journal on Control and Optimization, 52(4): 2449–2481, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Knowles:1976:TOC

[Kno76] G. Knowles. Time optimal control of infinite-dimensional systems. SIAM Journal on Control and Optimization, 14 (5):919–933, August 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Knowles:1977:SRI

[Kno77] Greg Knowles. Some remarks on infinite-dimensional nonlinear control without convexity. SIAM Journal on Control and Optimization, 15(5): 830–840, August 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Knowles:1979:SPC

[Kno79] Greg Knowles. Some problems in the control of distributed systems, and their numerical solution. SIAM Journal on Control and Optimization, 17 (1):5–22, January 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Knowles:1982:FEA

[Kno82] Greg Knowles. Finite element approximation of parabolic time optimal control problems. SIAM Journal on Control and Optimization, 20(3): 414–427, May 1982. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Khargonekar:1991:CT

[KNP91] Pramod P. Khargonekar, Krishan M. Nagpal, and Kameshwar R. Poolla. H_{∞} control with transients. SIAM Journal on Control and Optimization, 29(6):1373–1393, November 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kurdila:1995:PEI

[KNW95] A. J. Kurdila, Francis J. Narcowich, and Joseph D. Ward. Persistency of excitation in identification using radial basis function approximants. SIAM Journal on Control and Optimization, 33(2):625–642, March 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/23255.

Kovtunenko:2018:SDL

[KO18] V. A. Kovtunenko and K. Ohtsuka. Shape differentiability of Lagrangians and application to Stokes problem. SIAM Journal on Control and Optimization, 56(5):3668–3684, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kobayashi:1976:ISD

[Kob76] Toshihiro Kobayashi. Initial state determination for distributed parameter systems. SIAM Journal on Control and Optimization, 14(5): 934–944, August 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kobayashi:1977:WPA

[Kob77] Toshihiro Kobayashi. A wellposed approximate method for initial state determination of discrete-time distributed parameter systems. SIAM Journal on Control and Optimization, 15(6):947–958, November 1977. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kobayashi:1978:SRC

[Kob78] Toshihiro Kobayashi. Some remarks on controllability for distributed parameter systems. SIAM Journal on Control and Optimization, 16 (5):733–742, September 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kobayashi:1979:DUF

[Kob79] Toshihiro Kobayashi. Determination of unknown functions for a class of distributed parameter systems. SIAM Journal on Control and Optimization, 17(4):469–476, July 1979. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kobayashi:1983:DTO

[Kob83] Toshihiro Kobayashi. Discretetime observers and parameter determination for distributed parameter systems with discrete-time input– output data. SIAM Journal on Control and Optimization, 21 (3):331–351, May 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kobayashi:1984:FSP

[Kob84] Toshihiro Kobayashi. Feedback stabilization of parabolic distributed parameter systems by discrete-time input–output data. SIAM Journal on Control and Optimization, 22(4): 509–523, July 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kobayashi:1988:DPC

[Kob88]  Toshihiro Kobayashi.  A

digital PI-controller for distributed parameter systems. SIAM Journal on Control and Optimization, 26(6):1399– 1414, November 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Koehler:1979:VCG

[Koe79]  Gary J. Koehler.  Value

convergence in a generalized Markov decision process. SIAM Journal on Control and Optimization, 17(2): 180–186, March 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Koehler:1990:RBV

[Koe90] Gary J. Koehler. Relationships between various Markovian decision problem classes.

SIAM Journal on Control and Optimization, 28(6):1452– 1460, November 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Koenig:2020:LNC

[Koe20] Armand Koenig. Lack of nullcontrollability for the fractional heat equation and related equations. SIAM Journal on Control and Optimization, 58(6):3130–3160, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kogan:1987:SEO

[Kog87] Jacob Kogan. Structure of extremals in optimal control problems. SIAM Journal on Control and Optimization, 25 (4):951–966, July 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Komornik:1991:RBS

[Kom91]  Vilmos Komornik.  Rapid

boundary stabilization of the wave equation. SIAM Journal on Control and Optimization, 29(1):197–208, January 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Komornik:1997:RBS

[Kom97]  Vilmos Komornik.  Rapid

boundary stabilization of linear distributed systems. SIAM Journal on Control and Optimization, 35(5): 1591–1613, September 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30160.

Korda:2020:CCI

[Kor20] Milan Korda. Computing controlled invariant sets from data using convex optimization. SIAM Journal on Control and Optimization, 58(5): 2871–2899, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kuusela:2004:LCD

[KOS04] Pirkko Kuusela, Daniel Ocone, and Eduardo D. Sontag. Learning complexity dimensions for a continuous-time control system. SIAM Journal on Control and Optimization, 43(3):872–898, May 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38430.

Kotta:1994:CSA

[Kot94] Ü. Kotta. Comments on a structural approach to the nonlinear model matching problem. SIAM Journal on Control and Optimization, 32(6):1555–1558, November 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23138.  See

[MPC91].

Kouhkouh:2024:VEP

[Kou24] Hicham Kouhkouh. A viscous ergodic problem with unbounded and measurable ingredients, Part 1: HJB equation. SIAM Journal on Control and Optimization, 62(1): 415–440, February 2024. CO-DEN SJCODC. ISSN 1095-7138.

Kumar:1987:STT

[KP87] P. R. Kumar and L. Praly.

Self-tuning trackers. SIAM Journal on Control and Optimization, 25(4):1053–1071, July 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kryszewski:1994:TML

[KP94a] W. Kryszewski and S. Plaskacz.

Topological methods for the local controllability of nonlinear systems. SIAM Journal on Control and Optimization, 32(1):213–223, January 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/18416.

Kunisch:1994:EIB

[KP94b] Karl Kunisch and Xiao Su Pan. Estimation of interfaces from boundary measurements. SIAM Journal on Control and Optimization, 32(6):

1643–1674, November 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/22633.

Kabanov:1997:CAS

[KP97] Yuri Kabanov and Sergei Pergamenshchikov. On convergence of attainability sets for controlled two-scale stochastic linear systems. SIAM Journal on Control and Optimization, 35(1):134–159, January 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26968.

Kim:2004:SMD

[KP04] Kyung-Soo Kim and Youngjin Park. Sliding mode design via quadratic performance optimization with pole-clustering constraint. SIAM Journal on Control and Optimization, 43(2):670–684, March 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38847.

Karamzin:2023:HOS

[KP23] Dmitry Karamzin and Fernando Lobo Pereira. On higher-order state constraints. SIAM Journal on Control and Optimization, 61(4):1913– 1933, ???? 2023. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1487497.

Karafyllis:2022:GAS

[KPCW22] Iasson Karafyllis, Pierdomenico Pepe, Antoine Chaillet, and Yuan Wang. Is global asymptotic stability necessarily uniform for time-invariant timedelay systems? SIAM Journal on Control and Optimization, 60(6):3237–3261, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1485887.

Kunisch:2014:MVD

[KPV14] Karl Kunisch, Konstantin Pieper, and Boris Vexler. Measure valued directional sparsity for parabolic optimal control problems. SIAM Journal on Control and Optimization, 52(5):3078–3108, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kroner:2018:IHS

[KPZ18] Axel Kröner, Athena Picarelli, and Hasnaa Zidani. Infinite horizon stochastic optimal control problems with running maximum cost. SIAM Journal on Control and Optimization, 56(5):3296–3319, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kien:2020:SOO

[KQWY20] Bui Trong Kien, Xiaolong Qin, Ching-Feng Wen, and Jen-Chih Yao. Second-order optimality conditions for multiobjective optimal control problems with mixed pointwise constraints and free right end point. SIAM Journal on Control and Optimization, 58(4): 2658–2677, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kaneko:1980:SDA

[KR80] Hideaki Kaneko and William H. Ruckle. Schauder decompositions, approximations and control problems. SIAM Journal on Control and Optimization, 18(4):381–390, ???? 1980. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Krener:1985:NOL

[KR85] Arthur J. Krener and Witold Respondek. Nonlinear observers with linearizable error dynamics. SIAM Journal on Control and Optimization, 23 (2):197–216, March 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kim:1987:BCT

[KR87a] Jong Uhn Kim and Yuriko Renardy. Boundary control of the Timoshenko beam. SIAM Journal on Control and Optimization, 25(6):1417– 1429, November 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kushner:1987:NOS

[KR87b] Harold J. Kushner and W. Runggaldier. Nearly optimal state feedback controls for stochastic systems with wideband noise disturbances. SIAM Journal on Control and Optimization, 25(2):298– 315, March 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kushner:1988:NOS

[KR88] H. J. Kushner and K. M. Ramachandran. Nearly optimal singular controls for wideband noise driven systems. SIAM Journal on Control and Optimization, 26(3):569–591, May 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kushner:1989:OAO

[KR89] Harold J. Kushner and K. M.

Ramachandran. Optimal and approximately optimal control policies for queues in heavy traffic. SIAM Journal on Control and Optimization, 27(6): 1293–1318, November 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kirch:2004:EHW

[KR04] Michael Kirch and Wolfgang J. Runggaldier. Efficient hedging when asset prices follow a geometric Poisson process with unknown intensities. SIAM Journal on Control and Optimization, 43(4):1174–1195, July 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/42316.

Krumbiegel:2013:SOS

[KR13] K. Krumbiegel and J. Rehberg. Second order sufficient optimality conditions for parabolic optimal control problems with pointwise state constraints. SIAM Journal on Control and Optimization, 51 (1):304–331, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kroner:2015:RIE

[KR15] Axel Kröner and Sérgio S. Rodrigues. Remarks on the internal exponential stabilization to a nonstationary solution for 1D Burgers equations. SIAM Journal on Control and Optimization, 53(2): 1020–1055, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Krastanov:2017:NSO

[KR17] M. I. Krastanov and N. K. Ribarska. Nonseparation of sets and optimality conditions. SIAM Journal on Control and Optimization, 55(3): 1598–1618, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kratz:1993:CAB

[Kra93] Werner Kratz. On a certain asymptotic behaviour of time-invariant Riccati matrix differential equations. SIAM Journal on Control and Optimization, 31(1):70–85, January 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Krastanov:2009:SCS

[Kra09] Mikhail Ivanov Krastanov.

A sufficient condition for small-time local controllability. SIAM Journal on Control and Optimization, 48(4): 2296–2322, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Krener:1977:DTD

[Kre77a] Arthur J. Krener. A decomposition theory for differentiable systems. SIAM Journal on Control and Optimization, 15(5):813–829, August 1977. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Krener:1977:HOM

[Kre77b] Arthur J. Krener. The high order maximal principle and its application to singular extremals. SIAM Journal on Control and Optimization, 15 (2):256–293, February 1977. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kreps:1978:SDP

[Kre78] David M. Kreps. Sequential decision problems with expected utility criteria. III: Upper and lower transience. SIAM Journal on Control and Optimization, 16(3):420–428, May 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Krener:1985:LIC

[Kre85] Arthur J. Krener. (Ad_{f,g}), (ad_{f,g}) and locally (ad_{f,g}) invariant and controllability distributions. SIAM Journal on Control and Optimization, 23 (4):523–549, July 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Krener:1987:ART

[Kre87] A. J. Krener. Acausal realization theory, Part I: Linear deterministic systems. SIAM Journal on Control and Optimization, 25(3):499–525, May 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Krener:2000:LSH

[Kre00] Arthur J. Krener. The local solvability of a Hamilton– Jacobi–Bellman PDE around a nonhyperbolic critical point. SIAM Journal on Control and Optimization, 39(5): 1461–1484, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36108.

Kohls:2014:PEA

[KRS14] Kristina Kohls, Arnd Rösch, and Kunibert G. Siebert. A posteriori error analysis of optimal control problems with control constraints. SIAM Journal on Control and Optimization, 52(3):1832–1861, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Krastanov:2011:PMP

[KRT11] M. I. Krastanov, N. K. Ribarska, and Ts. Y. Tsachev. A Pontryagin maximum principle for infinite-dimensional problems. SIAM Journal on Control and Optimization, 49 (5):2155–2182, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sicon/
resource/1/sjcodc/v49/i5/
p2155_s1.

Kruk:2000:OPDa

[Kru00a] Lukasz Kruk. Optimal policies for n-dimensional singular stochastic control problems Part I: The Skorokhod problem. SIAM Journal on Control and Optimization, 38(5): 1603–1622, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/34753.

Kruk:2000:OPDb

[Kru00b] Lukasz Kruk. Optimal policies for n-dimensional singular stochastic control problems. Part II: The radially symmetric case. ergodic control. SIAM Journal on Control and Optimization, 39(2):635–659, March 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/34754.

Kim:2004:PPM

[KRW04] Meeyoung Kim, Joachim Rosenthal, and Xiaochang Alex Wang. Pole placement and matrix extension problems: a common point of view. SIAM Journal on Control and Optimization, 42(6): 2078–2093, November 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/35442.

Kryazhimskii:1998:COF

[Kry98] Arkadii V. Kryazhimskii. Convex optimization via feedbacks. SIAM Journal on Control and Optimization, 37 (1):278–302, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28030.

Kumar:1981:EVR

[KS81] P. R. Kumar and T. H. Shiau.

Existence of value and randomized strategies in zerosum discrete-time stochastic dynamic games. SIAM Journal on Control and Optimization, 19(5):617–634, September 1981. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Karatzas:1984:CBO

[KS84a] Ioannis Karatzas and Steven E.

Shreve. Connections between optimal stopping and singular stochastic control. I. Monotone follower problems. SIAM Journal on Control and Optimization, 22(6): 856–877, November 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kushner:1984:IMA

[KS84b] Harold J. Kushner and Adam Shwartz. An invariant measure approach to the convergence of stochastic approximations with state dependent noise. SIAM Journal on Control and Optimization, 22(1): 13–27, January 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Karatzas:1985:CBO

[KS85a] Ioannis Karatzas and Steven E. Shreve. Connections between optimal stopping and singular stochastic control. II. Reflected follower problems. SIAM Journal on Control and Optimization, 23(3):433–451, May 1985. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kravaris:1985:IPD

[KS85b] Costas Kravaris and John H. Seinfeld. Identification of parameters in distributed parameter systems by regularization. SIAM Journal on Control and Optimization, 23(2): 217–241, March 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kushner:1985:SAH

[KS85c] H. J. Kushner and A. Shwartz. Stochastic approximation in Hilbert space: Identification and optimization of linear continuous parameter systems. SIAM Journal on Control and Optimization, 23(5):774– 793, September 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kravaris:1986:ISV

[KS86] Costas Kravaris and John H. Seinfeld. Identifiability of spatially-varying conductivity from point observation as an inverse Sturm–Liouville problem. SIAM Journal on Control and Optimization, 24(3): 522–542, May 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kappel:1987:SAR

[KS87a] F. Kappel and D. Salamon.

Spline approximation for retarded systems and the Riccati equation. SIAM Journal on Control and Optimization, 25 (4):1082–1117, July 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kelley:1987:QNM

[KS87b] C. T. Kelley and E. W. Sachs.

Quasi-Newton methods and unconstrained optimal control problems. SIAM Journal on Control and Optimization, 25(6):1503–1516, November 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kappel:1989:SPS

[KS89a] F. Kappel and D. Salamon.

On the stability properties of spline approximations for retarded systems. SIAM Journal on Control and Optimization, 27(2):407–431, March 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Krener:1989:SST

[KS89b] Arthur J. Krener and Heinz Schättler. The structure of small-time reachable sets in low dimensions. SIAM Journal on Control and Optimization, 27(1):120–147, January 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kappel:1990:ATA

[KS90a] Franz Kappel and Dietmar Salamon. An approximation theorem for the algebraic Riccati equation. SIAM Journal on Control and Optimization, 28(5):1136–1147, September 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kuijper:1990:RAE

[KS90b] M. Kuijper and J. M. Schumacher. Realization of autoregressive equations in pencil and descriptor form. SIAM Journal on Control and Optimization, 28(5):1162– 1189, September 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kelley:1992:MIG

[KS92] C. T. Kelley and E. W. Sachs. Mesh independence of the gradient projection method for optimal control problems. SIAM Journal on Control and Optimization, 30 (2):477–493, March 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Koplon:1993:LSS

[KS93] Renée Koplon and Eduardo D.

Sontag. Linear systems with sign-observations. SIAM Journal on Control and Optimization, 31(5):1245–1266, September 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kelley:1995:SOC

[KS95] C. T. Kelley and E. W.

Sachs. Solution of optimal control problems by a pointwise projected Newton method. SIAM Journal on Control and Optimization, 33(6):1731–1757, November 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24990.

Kesavan:1997:HOC

[KS97a] S. Kesavan and J. Saint Jean Paulin. Homogenization of an optimal control problem. SIAM Journal on Control and Optimization, 35(5): 1557–1573, September 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27184.

Kumar:1997:CDS

[KS97b] Ratnesh Kumar and Mark A. Shayman. Centralized and decentralized supervisory control of nondeterministic systems under partial observation. SIAM Journal on Control and Optimization, 35(2):363–383, March 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27290.

Kocan:1998:VAI

[KS98a] Maciej Kocan and Pierpaolo Soravia. A viscosity approach to infinite-dimensional Hamilton–Jacobi equations arising in optimal control with state constraints. SIAM Journal on Control and Optimization, 36(4):1348–1375, July 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30162.

Kurtz:1998:EMC

[KS98b] Thomas G. Kurtz and Richard H. Stockbridge. Existence of Markov controls and characterization of optimal Markov controls. SIAM Journal on Control and Optimization, 36(2):609–653, March 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/

dbq/article/29551. See er-

ratum [KS99b].

Kiefer:1999:PFE

[KS99a] Matthew Kiefer and Heinz Schättler. Parametrized families of extremals and singularities in solutions to the Hamilton–Jacobi–Bellman equation. SIAM Journal on Control and Optimization, 37(5):1346–1371, September 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31913.

Kurtz:1999:EEM

[KS99b] Thomas G. Kurtz and Richard H.

Stockbridge. Erratum: “Existence of Markov Controls and Characterization of Optimal Markov Controls”. SIAM Journal on Control and Optimization, 37(4):1310–1311, July 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34303.  See

[KS98b].

Kocan:2000:DGN

[KS00] Maciej Kocan and Pierpaolo Soravia. Differential games and nonlinear H_{∞} control in infinite dimensions. SIAM Journal on Control and Optimization, 39(4):1296–1322, July 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/34356.

Korn:2007:WCP

[KS07] Ralf Korn and Mogens Steffensen. On worst-case portfolio optimization. SIAM Journal on Control and Optimization, 46(6):2013–2030, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kunoth:2013:ARG

[KS13] Angela Kunoth and Christoph Schwab. Analytic regularity and GPC approximation for control problems constrained by linear parametric elliptic and parabolic PDEs. SIAM Journal on Control and Optimization, 51(3):2442–2471, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kouritzin:2015:CRD

[KS15] Michael A. Kouritzin and Samira Sadeghi. Convergence rates and decoupling in linear stochastic approximation algorithms. SIAM Journal on Control and Optimization, 53(3):1484–1508, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kanzow:2018:EBM

[KS18] Christian Kanzow and Daniel Steck. On error bounds and multiplier methods for variational problems in Banach spaces. SIAM Journal on Control and Optimization, 56(3): 1716–1738, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kerimkulov:2020:ECS

[KSS20] Bekzhan Kerimkulov, David Siska, and Lukasz Szpruch. Exponential convergence and stability of Howard’s policy improvement algorithm for controlled diffusions. SIAM Journal on Control and Optimization, 58(3):1314–1340, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Krichman:2001:IOS

[KSW01] Mikhail Krichman, Eduardo D.

Sontag, and Yuan Wang. Input-output-to-state stability. SIAM Journal on Control and Optimization, 39(6): 1874–1928, ???? 2001. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kanzow:2018:ALM

[KSW18] Christian Kanzow, Daniel Steck, and Daniel Wachsmuth. An augmented Lagrangian method for optimization problems in Banach spaces. SIAM Journal on Control and Optimization, 56(1):272–291, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kang:2023:LCP

[KSY23] Jiayi Kang, Andrew Salmon, and Stephen S.-T. Yau. Logconcave posterior densities arising in continuous filtering and a maximum A Posteriori algorithm. SIAM Journal on Control and Optimization, 61(4):2407–2424, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1508352.

Kaplan:2000:PPA

[KT00] A. Kaplan and R. Tichatschke. Proximal point approach and approximation of variational inequalities. SIAM Journal on Control and Optimization, 39(4):1136–1159, July 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33311.

Karafyllis:2003:CLT

[KT03a] I. Karafyllis and J. Tsinias. A converse Lyapunov theorem for nonuniform in time global asymptotic stability and its application to feedback stabilization. SIAM Journal on Control and Optimization, 42(3):936–965, May 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39296.

Kohlmann:2003:MRL

[KT03b] Michael Kohlmann and Shanjian Tang. Minimization of risk and linear quadratic optimal control theory. SIAM Journal on Control and Optimization, 42(3):1118–1142, May 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37246.

Kohlmann:2003:MBS

[KT03c] Michael Kohlmann and Shanjian Tang. Multidimensional backward stochastic Riccati equations and applications. SIAM Journal on Control and Optimization, 41(6): 1696–1721, November 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37876.

Konda:2003:ACA

[KT03d] Vijay R. Konda and John N.

Tsitsiklis. On actor–critic algorithms. SIAM Journal on Control and Optimization, 42(4):1143–1166, July 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38569.

Kellett:2004:WCL

[KT04] Christopher M. Kellett and Andrew R. Teel. Weak converse Lyapunov theorems and control-Lyapunov functions. SIAM Journal on Control and Optimization, 42(6): 1934–1959, November 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39818.

Kellett:2005:RSD

[KT05] Christopher M. Kellett and Andrew R. Teel. On the robustness of KL-stability for difference inclusions: Smooth discrete-time Lyapunov functions. SIAM Journal on Control and Optimization, 44(3):777–800, May 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43586.

Krejci:2016:COC

[KT16] Pavel Krejčí and Sergey A. Timoshin. Coupled ODEs control system with unbounded hysteresis region. SIAM Journal on Control and Optimization, 54(4):1934–1949, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kunisch:2016:OCU

[KTV16] Karl Kunisch, Philip Trautmann, and Boris Vexler. Optimal control of the undamped linear wave equation with measure valued controls. SIAM Journal on Control and Optimization, 54(3): 1212–1244, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kien:2012:LSS

[KTWY12] B. T. Kien, N. T. Toan, M. M.

Wong, and J. C. Yao. Lower semicontinuity of the solution set to a parametric optimal control problem. SIAM Journal on Control and Optimization, 50(5):2889–2906, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kien:2018:SOK

[KTY18] Bui Trong Kien, Nguyen Van Tuyen, and Jen-Chih Yao. Second-order KKT optimality conditions for MultiObjective optimal control problems. SIAM Journal on Control and Optimization, 56(6): 4069–4097, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kubrusly:1985:MSS

[Kub85] C. S. Kubrusly. Mean square stability for discrete bounded linear systems in Hilbert space. SIAM Journal on Control and Optimization, 23(1): 19–29, January 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kuenle:2007:MGA

[Küe07]  Heinz-Uwe Küenle.  On

Markov games with average reward criterion and weakly continuous transition probabilities. SIAM Journal on Control and Optimization, 45 (6):2156–2168, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kulldorff:1993:OCF

[Kul93] Martin Kulldorff. Optimal control of favorable games with a time limit. SIAM Journal on Control and Optimization, 31(1):52–69, January 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kumar:1982:ACC

[Kum82] P. R. Kumar. Adaptive control with a compact parameter set. SIAM Journal on Control and Optimization, 20(1):9–13, January 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kumar:1983:OAC

[Kum83]  P. R. Kumar.  Optimal

adaptive control of linearquadratic-Gaussian systems. SIAM Journal on Control and Optimization, 21(2):163– 178, March 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kumar:1985:SSR

[Kum85] P. R. Kumar. A survey of some results in stochastic adaptive control. SIAM Journal on Control and Optimization, 23 (3):329–380, May 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kumpera:1988:RPG

[Kum88] A. Kumpera. Remarks on a paper: “Global controllability of nonlinear systems” [SIAM J. Control Optimization 14 (1976), no. 4, 700– 711; MR 58 #15374] by R. M. Hirschorn. SIAM Journal on Control and Optimization, 26 (2):465–467, March 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [Hir76].

Kumar:2008:NSS

[Kum08] K. Suresh Kumar. Nonzero sum stochastic differential games with discounted payoff criterion: An approximating Markov chain approach. SIAM Journal on Control and Optimization, 47(1):374–395, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kumar:2013:SPS

[Kum13] K. Suresh Kumar. Singular perturbations in stochastic ergodic control problems. SIAM Journal on Control and Optimization, 50(6):3203–3223, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kunisch:1982:ASL

[Kun82] Karl Kunisch. Approximation schemes for the linearquadratic optimal control problem associated with delay equations. SIAM Journal on Control and Optimization, 20 (4):506–540, July 1982. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kupka:1985:FVS

[Kup85] I. A. K. Kupka. On finite Volterra series and a theorem of P. Crouch. SIAM Journal on Control and Optimization, 23(2):297–305, March 1985. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kurano:1989:EMP

[Kur89] M. Kurano. The existence of a minimum pair of state and policy for Markov decision processes under the hypothesis of Doeblin. SIAM Journal on Control and Optimization, 27 (2):296–307, March 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kushner:1978:OCA

[Kus78a] H. J. Kushner. Optimality conditions for the average cost per unit time problem with a diffusion model. SIAM Journal on Control and Optimization, 16(2):330–346, March 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kushner:1978:RCS

[Kus78b] Harold J. Kushner. Rates of convergence for sequential Monte Carlo optimization methods. SIAM Journal on Control and Optimization, 16(1):150–168, January 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kushner:1979:JDA

[Kus79] Harold J. Kushner. Jumpdiffusion approximations for ordinary differential equations with wide-band random right hand sides. SIAM Journal on Control and Optimization, 17(6):729–744, November 1979. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kushner:1989:AOC

[Kus89] Harold J. Kushner. Approximations and optimal control for the pathwise average cost per unit time and discounted problems for wideband noisedriven systems. SIAM Journal on Control and Optimization, 27(3):546–562, May 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kushner:1990:NMS

[Kus90]  Harold J. Kushner.  Nu-

merical methods for stochastic control problems in continuous time. SIAM Journal on Control and Optimization, 28(5):999–1048, September 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kushner:1995:CTL

[Kus95] Harold J. Kushner. Control of trunk line systems in heavy traffic. SIAM Journal on Control and Optimization, 33(3):765–803, May 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/22880.

Kushner:2002:NAS

[Kus02]  Harold J. Kushner.  Nu-

merical approximations for stochastic differential games. SIAM Journal on Control and Optimization, 41(2):457– 486, March 2002. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/38945.

Kushner:2004:NAS

[Kus04] Harold J. Kushner. Numerical approximations for stochastic differential games: The ergodic case. SIAM Journal on Control and Optimization, 42(6):1911–1933, November 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40034.

Kushner:2007:NAN

[Kus07]  Harold J. Kushner.  Nu-

merical approximations for nonzero-sum stochastic differential games. SIAM Journal on Control and Optimization, 46(6):1942–1971, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kushner:2013:NMS

[Kus13]  Harold J. Kushner.  Nu-

merical methods for stochastic singular control problems with state-dependent control. SIAM Journal on Control and Optimization, 51(1):685–717, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kushner:2014:MAS

[Kus14] Harold J. Kushner. Modeling and approximations for stochastic systems with statedependent singular controls and wide-band noise. SIAM Journal on Control and Optimization, 52(1):311–338, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kurzhanski:2002:RUU

[KV02] A. B. Kurzhanski and P. Varaiya. On reachability under uncertainty. SIAM Journal on Control and Optimization, 41 (1):181–216, January 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/36109.

Kurzhanski:2006:ETR

[KV06] A. B. Kurzhanski and P. Varaiya. Ellipsoidal techniques for reachability under state constraints. SIAM Journal on Control and Optimization, 45 (4):1369–1394, January 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kunisch:2007:CDB

[KV07a] K. Kunisch and B. Vexler. Constrained Dirichlet boundary control in L^{2} for a class of evolution equations. SIAM Journal on Control and Optimization, 46(5):1726–1753, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kunisch:2007:OVR

[KV07b] K. Kunisch and B. Vexler. Optimal vortex reduction for instationary flows based on translation invariant cost functionals. SIAM Journal on Control and Optimization, 46 (4):1368–1397, ???? 2007.

CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kuchler:2013:CED

[KV13] Uwe Küchler and Vyacheslav A. Vasiliev. On a certainty equivalence design of continuous-time stochastic systems. SIAM Journal on Control and Optimization, 51 (2):938–964, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Koch:2021:OIS

[KV21] Torben Koch and Tiziano Vargiolu. Optimal installation of solar panels with price impact: a solvable singular stochastic control problem. SIAM Journal on Control and Optimization, 59(4):3068–3095, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kushner:1996:SAM

[KVA96] Harold J. Kushner and Felisa J. Vázquez-Abad. Stochastic approximation methods for systems over an infinite horizon. SIAM Journal on Control and Optimization, 34(2):712–756, March 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26374.

Koerts:2021:SFC

[KvdSD21] Filip J. Koerts, Arjan van der Schaft, and Claudio De Persis. Secondary frequency control in power systems with arbitrary communication delays. SIAM Journal on Control and Optimization, 59(5): 3787–3804, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kwong:1978:EFS

[KW78] Raymond H. Kwong and Alan S. Willsky. Estimation and filter stability of stochastic delay systems. SIAM Journal on Control and Optimization, 16(4):660–681, July 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kunisch:1985:PEP

[KW85] K. Kunisch and L. White. The parameter estimation problem for parabolic equations and discontinuous observation operators. SIAM Journal on Control and Optimization, 23 (6):900–927, November 1985. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kunisch:1987:PER

[KW87a] K. Kunisch and L. W. White. Parameter estimation, regularity and the penalty method for a class of two point boundary value problems. SIAM Journal on Control and Optimization, 25(1):100–120, January 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kunisch:1987:IUA

[KW87b] Karl Kunisch and L. W.

White. Identifiability under approximation for an elliptic boundary value problem. SIAM Journal on Control and Optimization, 25(2): 279–297, March 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Karatzas:2000:UMD

[KW00] Ioannis Karatzas and Hui Wang. Utility maximization with discretionary stopping. SIAM Journal on Control and Optimization, 39 (1):306–329, January 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34632.

Kunisch:2013:TOC

[KW13] Karl Kunisch and Daniel Wachsmuth. On time optimal control of the wave equation and its numerical realization as parametric optimization problem. SIAM Journal on Control and Optimization, 51(2):1232–1262, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kwong:1980:STL

[Kwo80] R. H. Kwong. A stability theory for the linear-quadratic-Gaussian problem for systems with delays in the state, control, and observations. SIAM Journal on Control and Optimization, 18(1):49–75, January 1980. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kwong:1991:LQG

[Kwo91] R. H. Kwong. On the linear quadratic Gaussian problem with correlated noise and its relation to minimum variance control. SIAM Journal on Control and Optimization, 29(1):139–152, January 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kien:2008:NCM

[KWY08] B. T. Kien, N.-C. Wong, and J.-C. Yao. Necessary conditions for multiobjective optimal control problems with free end-time. SIAM Journal on Control and Optimization, 47(5):2251–2274, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Krener:2002:NOD

[KX02] Arthur J. Krener and MingQing Xiao. Nonlinear observer design in the Siegel domain. SIAM Journal on Control and Optimization, 41(3):932–953, May 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37533. See er-

ratum [KX04].

Krener:2004:ENO

[KX04] Arthur J. Krener and MingQing Xiao. Erratum: “Nonlinear Observer Design in the Siegel Domain”. SIAM Journal on Control and Optimization, 43(1):377–378, January 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43511.  See

[KX02].

Kawase:1987:FMS

[KY87a] Shin Kawase and Niro Yanagihara. Fine moduli spaces of infinite dimensional linear systems. SIAM Journal on Control and Optimization, 25(4): 1020–1031, July 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kushner:1987:APD

[KY87b] Harold J. Kushner and G. Yin.

Asymptotic properties of distributed and communicating stochastic approximation algorithms. SIAM Journal on Control and Optimization, 25(5):1266–1290, September 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kushner:1992:MCM

[KY92] Harold J. Kushner and Jichuan Yang. A Monte Carlo method for sensitivity analysis and parametric optimization of nonlinear stochastic systems: The ergodic case. SIAM Journal on Control and Optimization, 30(2):440–464, March 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kushner:1993:SAA

[KY93] Harold J. Kushner and Jichuan Yang. Stochastic approximation with averaging of the iterates: Optimal asymptotic rate of convergence for general processes. SIAM Journal on Control and Optimization, 31(4): 1045–1062, July 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kong:2005:GEB

[KY05] De-Xing Kong and Hui Yao. Global exact boundary controllability of a class of quasilinear hyperbolic systems of conservation laws II. SIAM Journal on Control and Optimization, 44(1):140–158, January 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/43265.

Klibanov:2007:ECT

[KY07] Michael V. Klibanov and Masahiro Yamamoto. Exact controllability for the time dependent transport equation. SIAM Journal on Control and Optimization, 46(6): 2071–2095, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kara:2019:RIP

[KY19] Ali Devran Kara and Serdar Yüksel. Robustness to incorrect priors in partially observed stochastic control. SIAM Journal on Control and Optimization, 57(3):1929– 1964, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kara:2020:RIS

[KY20] Ali D. Kara and Serdar Yüksel. Robustness to incorrect system models in stochastic control. SIAM Journal on Control and Optimization, 58(2):1144–1182, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Kim:2023:DRM

[KY23] Kihyun Kim and Insoon Yang.

Distributional robustness in minimax linear quadratic control with Wasserstein distance. SIAM Journal on Control and Optimization, 61(2):458–483, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1494105.

Kawasaki:2000:CPV

[KZ00a] Hidefumi Kawasaki and Vera Zeidan. Conjugate points for variational problems with equality and inequality state constraints. SIAM Journal on Control and Optimization, 39(2):433–456, March 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34592.

Kohlmann:2000:RBB

[KZ00b] Michael Kohlmann and Xun Yu Zhou. Relationship between backward stochastic differential equations and stochastic controls: a linear-quadratic approach. SIAM Journal on Control and Optimization, 38(5):1392–1407, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34973.

Kawan:2023:LBS

[KZ23] Christoph Kawan and Majid Zamani. A Lyapunovbased small-gain theory for infinite networks via infinitedimensional gain operators. SIAM Journal on Control and Optimization, 61(3):1778– 1804, ???? 2023. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1465500.

Lim:1996:MRL

[LA96] Wei Shi Lim and Steve Alpern.

Minimax rendezvous on the line. SIAM Journal on Control and Optimization, 34(5): 1650–1665, September 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27816.

Lacker:2017:LTC

[Lac17] Daniel Lacker. Limit theory for controlled McKean–Vlasov dynamics. SIAM Journal on Control and Optimization, 55 (3):1641–1672, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lagnese:1977:BVC

[Lag77] John Lagnese. Boundary value control of a class of hyperbolic equations in a general region. SIAM Journal on Control and Optimization, 15(6): 973–983, November 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lagnese:1978:EBV

[Lag78] John Lagnese. Exact boundary value controllability of a class of hyperbolic equations. SIAM Journal on Control and Optimization, 16(6):

1000–1017, November 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lagnese:1983:BSL

[Lag83a] John Lagnese. Boundary stabilization of linear elastodynamic systems. SIAM Journal on Control and Optimization, 21(6):968–984, November 1983. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lagnese:1983:CWP

[Lag83b] John Lagnese. Control of wave processes with distributed controls supported on a subregion. SIAM Journal on Control and Optimization, 21(1): 68–85, January 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lagnese:1988:NBS

[Lag88] John E. Lagnese. Note on boundary stabilization of wave equations. SIAM Journal on Control and Optimization, 26(5):1250–1256, September 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lagnese:1989:EBC

[Lag89]  John E. Lagnese.  Exact

boundary controllability of Maxwell’s equations in a general region. SIAM Journal on Control and Optimization, 27 (2):374–388, March 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Limebeer:1992:GTA

[LAKG92] David J. N. Limebeer, Brian D. O. Anderson, Pramod P. Khargonekar, and Michael Green. A game theoretic approach to H^{∞} control for time-varying systems. SIAM Journal on Control and Optimization, 30(2): 262–283, March 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lamm:1987:EDC

[Lam87] Patricia K. Lamm. Estimation of discontinuous coefficients in parabolic systems: Applications to reservoir simulation. SIAM Journal on Control and Optimization, 25(1): 18–37, January 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lamm:1992:PED

[Lam92] Patricia K. Lamm. Parameter estimation for distributed equations in parameter-dependent state spaces: Applications to shape identification. SIAM Journal on Control and Optimization, 30(4):894–925, July 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liang:2019:ESL

[LAM19] Weichao Liang, Nina H. Amini, and Paolo Mason. On exponential stabilization of N-level quantum angular momentum systems. SIAM Journal on Control and Optimization, 57(6):3939–3960, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liang:2021:RFS

[LAM21] Weichao Liang, Nina H. Amini, and Paolo Mason. Robust feedback stabilization of N-level quantum spin systems. SIAM Journal on Control and Optimization, 59(1): 669–692, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lizarraga:2004:RPS

[LAN04] David A. Lizárraga, Nnaedozie P. I. Aneke, and Henk Nijmeijer. Robust point stabilization of underactuated mechanical systems via the extended chained form. SIAM Journal on Control and Optimization, 42(6):2172–2199, November 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/40557.

Laraki:2001:VIS

[Lar01] Rida Laraki. Variational inequalities, system of functional equations, and incomplete information repeated games. SIAM Journal on Control and Optimization, 40(2):516–524, March 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36660.

Lasiecka:1984:RGA

[Las84] Irena Lasiecka. Ritz–Galerkin approximation of the time optimal boundary control problem for parabolic systems with Dirichlet boundary conditions. SIAM Journal on Control and Optimization, 22(3):477–500, May 1984. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lasiecka:1995:FEA

[Las95] I. Lasiecka. Finite element approximations of compensator design for analytic generators with fully unbounded controls/observations. SIAM Journal on Control and Optimization, 33(1):67–88, January 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23220.

Lasserre:1997:FLS

[Las97] Jean B. Lasserre. A Farkas lemma without a standard closure condition. SIAM Journal on Control and Optimization, 35(1):265–272, January 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/26832.

Lasiecka:1998:USF

[Las98] Irena Lasiecka. Uniform stabilizability of a full von Karman system with nonlinear boundary feedback. SIAM Journal on Control and Optimization, 36(4):1376–1422, July 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/30190.

Laumen:1999:NMI

[Lau99] Manfred Laumen. Newton’s mesh independence principle for a class of optimal shape design problems. SIAM Journal on Control and Optimization, 37(4):1070–1088, July 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/30352.

Lazimy:1988:EPF

[Laz88] Rafael Lazimy. An exact penalty function and relaxation approach for solving decomposable nonlinear programs. SIAM Journal on Control and Optimization, 26(1): 127–137, January 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Luse:1989:FSD

[LB89] D. William Luse and Joseph A.

Ball. Frequency-scale decomposition of H^{∞}-disk problems. SIAM Journal on Control and Optimization, 27(4): 814–835, July 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Luo:1998:NAS

[LB98] Xiaodong Luo and Dimitris Bertsimas. A new algorithm for state-constrained separated continuous linear programs. SIAM Journal on Control and Optimization, 37 (1):177–210, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29266.

Liberzon:2000:SAF

[LB00] Daniel Liberzon and Roger W.

Brockett. Spectral analysis of Fokker–Planck and related operators arising from linear stochastic differential equations. SIAM Journal on Control and Optimization, 38(5): 1453–1467, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33819.

Lee:2008:GOL

[LB08] C. K. Lee and P. I. Barton.

Global optimization of linear hybrid systems with varying transition times. SIAM Journal on Control and Optimization, 47(2):791–816, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2016:ECS

[LBMH^{+}16] Yuxiang Liu, Bandar Bin-Mohsin, Hichem Hajaiej, Peng-Fei Yao, and Goong Chen. Exact controllability of structural acoustic interactions with variable coefficients. SIAM Journal on Control and Optimization, 54(4): 2132–2153, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lent:1980:EHR

[LC80] Arnold Lent and Yair Censor. Extensions of Hildreth’s rowaction method for quadratic programming. SIAM Journal on Control and Optimization, 18(4):444–454, ???? 1980. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Levanony:2001:PEL

[LC01] David Levanony and Peter E. Caines. On persistent excitation for linear systems with stochastic coefficients. SIAM Journal on Control and Optimization, 40(3):882–897, May 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30045.

Leslie:2005:ILN

[LC05] David S. Leslie and E. J.

Collins. Individual Qlearning in normal form games. SIAM Journal on Control and Optimization, 44(2):495–514, March 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43797.

Li:2019:RCM

[LC19] Zhongkui Li and Jie Chen.

Robust consensus for multiagent systems communicating over stochastic uncertain networks. SIAM Journal on Control and Optimization, 57(5): 3553–3570, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Luo:2023:OCP

[LC23] Jianfeng Luo and Xiaojun Chen. An optimal control problem with terminal stochastic linear complementarity constraints. SIAM Journal on Control and Optimization, 61(6):3370–3389, November 2023. CODEN SJCODC. ISSN 1095-7138.

Laporte:2017:GSL

[LCC17] Jonathan Laporte, Antoine Chaillet, and Yacine Chitour. Global stabilization of linear systems with bounds on the feedback and its successive derivatives. SIAM Journal on Control and Optimization, 55(5):2783–2810, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lei:2018:APP

[LCF18] Jinlong Lei, Han-Fu Chen, and Hai-Tao Fang. Asymptotic properties of primal-dual algorithm for distributed stochastic optimization over random networks with imperfect communications. SIAM Journal on Control and Optimization, 56(3):2159–2188, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:1996:FGS

[LCS96] Wensheng Liu, Yacine Chitour, and Eduardo Sontag. On finite-gain stabilizability of linear systems subject to input saturation. SIAM Journal on Control and Optimization, 34(4):1190–1219, July 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26346.

Li:2018:BBC

[LCW18] Rui Li, Tianguang Chu, and Xingyuan Wang. Bisimulations of Boolean control networks. SIAM Journal on Control and Optimization, 56(1):

388–416, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:2020:CSH

[LCZ20] Shenghao Li, Min Chen, and Bingyu Zhang. Controllability and stabilizability of a higher order wave equation on a periodic domain. SIAM Journal on Control and Optimization, 58(2):1121–1143, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Langbort:2005:DCS

[LD05] Cédric Langbort and Raffaello D’Andrea. Distributed control of spatially reversible interconnected systems with boundary conditions. SIAM Journal on Control and Optimization, 44(1):1–28, January 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41580.

Lee:2006:ODA

[LD06] Ji-Woong Lee and Geir E.

Dullerud. Optimal disturbance attenuation for discretetime switched and Markovian jump linear systems. SIAM Journal on Control and Optimization, 45(4):1329– 1358, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:2019:CLF

[LD19] Haitao Li and Xueying Ding. A control Lyapunov function approach to feedback stabilization of logical control networks. SIAM Journal on Control and Optimization, 57(2): 810–831, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lv:2010:PPA

[LDZ10] Lingling Lv, Guangren Duan, and Bin Zhou. Parametric pole assignment and robust pole assignment for discretetime linear periodic systems. SIAM Journal on Control and Optimization, 48(6):3975– 3996, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

LeBalch:2019:NCT

[Le 19]  Kévin Le Balc’h.  Null-

controllability of two species reaction–diffusion system with nonlinear coupling: a new duality method. SIAM Journal on Control and Optimization, 57(4):2541–2573, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2017:CSI

[LE17] Yichen Liu and Behrouz Emamizadeh. Converse symmetry and intermediate energy values in rearrangement optimization problems. SIAM Journal on Control and Optimization, 55(3):2088–2107, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Leautaud:2012:UCS

[Léa12] Matthieu Léautaud. Uniform controllability of scalar conservation laws in the vanishing viscosity limit. SIAM Journal on Control and Optimization, 50(3):1661–1699, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lee:1986:MQC

[Lee86] Sung J. Lee. A multiresponse quadratic control problem. SIAM Journal on Control and Optimization, 24(4): 771–788, July 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lefter:2011:FSM

[Lef11] Catalin-George Lefter. Feedback stabilization of magnetohydrodynamic equations. SIAM Journal on Control and Optimization, 49(3):963– 983, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i3/p963_s1.

Leizarowitz:1985:CVS

[Lei85] Arie Leizarowitz. Convergence of viable solutions of differential inclusions with convex compact graphs. SIAM Journal on Control and Optimization, 23(4):514–522, July 1985. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Leizarowitz:1987:IHO

[Lei87]  Arie Leizarowitz.  Infinite

horizon optimization for finite state Markov chain. SIAM Journal on Control and Optimization, 25(6):1601– 1618, November 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Leibfritz:2001:LBA

[Lei01]  F. Leibfritz.  An LMI-

Based algorithm for designing suboptimal static H_{∈}/H_{∞} output feedback controllers. SIAM Journal on Control and Optimization, 39(6): 1711–1735, November 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34955.

Leitner:2006:OCF

[Lei06] Johannes Leitner. Optimal control of favorable games with expected loss constraint. SIAM Journal on Control and Optimization, 45(2):483– 495, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lepp:1990:ASP

[Lep90] Riho Lepp. Approximations to stochastic programs with complete recourse. SIAM Journal on Control and Optimization, 28(2):382–394, March 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lequeurre:2013:NCF

[Leq13] Julien Lequeurre. Null controllability of a fluid-structure system. SIAM Journal on Control and Optimization, 51 (3):1841–1872, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lestas:2012:LSH

[Les12] Ioannis Lestas. Large scale heterogeneous networks, the Davis–Wielandt shell, and graph separation. SIAM Journal on Control and Optimization, 50(4):1753–1774, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Leugering:1999:DDD

[Leu99] G. Leugering. Dynamic domain decomposition of optimal control problems for networks of strings and Timoshenko beams. SIAM Journal on Control and Optimization, 37(6):1649–1675, November 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/33198.

Levan:1985:ASA

[Lev85] N. Levan. Approximate stabilizability via the algebraic Riccati equation. SIAM Journal on Control and Optimization, 23(1):153–160, January 1985. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Levine:1987:FDF

[Lev87] J. Levine. Finite dimensional filters for a class of nonlinear systems and immersion in a linear system. SIAM Journal on Control and Optimization, 25(6):1430–1439, November 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Levy:1999:SSV

[Lev99] Adam B. Levy. Sensitivity of solutions to variational inequalities on Banach spaces. SIAM Journal on Control and Optimization, 38(1):50–60, January 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/33985.

Levy:2001:SSG

[Lev01] Adam B. Levy. Solution sensitivity from general principles. SIAM Journal on Control and Optimization, 40(1):1–38, January 2001. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35211.

Lewis:1980:DOJ

[Lew80] R. M. Lewis. Definitions of order and junction conditions in singular optimal control problems. SIAM Journal on Control and Optimization, 18(1): 21–32, January 1980. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lee:1987:DMB

[LF87] B. S. Lee and F. W. Fairman.

Direct minimal balanced realization from impulse response matrices using I/O map decomposition. SIAM Journal on Control and Optimization, 25(1):1–17, January 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2015:RSI

[LF15] Jing Liu and Mark French. Robust stability of input-output systems with initial conditions. SIAM Journal on Control and Optimization, 53(3): 1625–1653, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2021:CBG

[LFC21] Jing Liu, Mark French, and Jun-Ting Chen. Connections between Georgiou and Smith’s robust stability type theorems and the nonlinear small-gain theorems. SIAM Journal on Control and Optimization, 59 (3):1989–2010, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2015:NCS

[LFH15] Kun Liu, Emilia Fridman, and Laurentiu Hetel. Networked control systems in the presence of scheduling protocols and communication delays. SIAM Journal on Control and Optimization, 53(4): 1768–1788, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liang:2023:PCS

[LFL23] Yixing Liang, Zhenbin Fan, and Gang Li. Processcontrollability of semilinear evolution equations and applications. SIAM Journal on Control and Optimization, 61(6):3664–3694, December 2023. CODEN SJCODC. ISSN 1095-7138.

Lin:2007:SAC

[LFM07] Zhiyun Lin, Bruce Francis, and Manfredi Maggiore. State agreement for continuous-time coupled nonlinear systems. SIAM Journal on Control and Optimization, 46(1):288–307, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2007:CCN

[LG07] Xu Liu and Hang Gao. Controllability of a class of Newtonian filtration equations with control and state constraints. SIAM Journal on Control and Optimization, 46(6):2256– 2279, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Langbort:2009:MIT

[LG09] Cédric Langbort and Vijay Gupta. Minimal interconnection topology in distributed control design. SIAM Journal on Control and Optimization, 48(1):397–413, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2020:NSO

[LG20] Jiankang Liu and Bao-Zhu Guo. A new semidiscretized order reduction finite difference scheme for uniform approximation of onedimensional wave equation. SIAM Journal on Control and Optimization, 58(4):2256– 2287, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2024:ASN

[LG24] Nian Liu and Lei Guo. Adaptive stabilization of noncooperative stochastic differential games. SIAM Journal on Control and Optimization, 62(3):

1317–1342, June 2024. CO-DEN SJCODC. ISSN 1095-7138.

Lamare:2015:SRS

[LGP15] Pierre-Olivier Lamare, Antoine Girard, and Christophe Prieur. Switching rules for stabilization of linear systems of conservation laws. SIAM Journal on Control and Optimization, 53(3):1599–1624, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lovisari:2013:RBP

[LGZ13] Enrico Lovisari, Federica Garin, and Sandro Zampieri. Resistance-based performance analysis of the consensus algorithm over geometric graphs. SIAM Journal on Control and Optimization, 51(5):3918– 3945, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Limebeer:1987:APZ

[LH87] D. J. N. Limebeer and Y. S. Hung. An analysis of the polezero cancellations in H^{∞}- optimal control problems of the first kind. SIAM Journal on Control and Optimization, 25(6):1457–1493, November 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Limebeer:1988:CDB

[LH88] D. J. N. Limebeer and G. D. Halikias. A controller degree bound for H^{∞}-optimal control problems of the second kind. SIAM Journal on Control and Optimization, 26 (3):646–677, May 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2011:ICC

[LH11] Bin Liu and David J. Hill. Impulsive consensus for complex dynamical networks with nonidentical nodes and coupling time-delays. SIAM Journal on Control and Optimization, 49 (2):315–338, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i2/p315_s1.

Liu:2014:SHE

[LH14] Bin Liu and David J. Hill. Stability via hybrid-event-time Lyapunov function and impulsive stabilization for discretetime delayed switched systems. SIAM Journal on Control and Optimization, 52(2): 1338–1365, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lahbiri:2021:FAA

[LH21] Fatima Zahra Lahbiri and Said Hadd. A functional analytic approach to infinite dimensional stochastic linear systems. SIAM Journal on Control and Optimization, 59(5):

3762–3786, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2024:ESC

[LH24] Tao Liu and Jie Huang. Exponential stability for a class of discrete-time switched systems and its applications to multiagent systems. SIAM Journal on Control and Optimization, 62(2):1165–1182, April 2024. CODEN SJCODC. ISSN 1095-7138.

Leong:2016:LSS

[LHB16] Yoke Peng Leong, Matanya B. Horowitz, and Joel W. Burdick. Linearly solvable stochastic control Lyapunov functions. SIAM Journal on Control and Optimization, 54 (6):3106–3125, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lee:2023:DTS

[LHH23] Donghwan Lee, Jianghai Hu, and Niao He. A discretetime switching system analysis of Q-learning. SIAM Journal on Control and Optimization, 61(3):1861–1880, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1489976.

Lasserre:2008:NOC

[LHPT08] Jean B. Lasserre, Didier Henrion, Christophe Prieur, and Emmanuel Trélat. Nonlinear optimal control via occupation measures and LMIrelaxations. SIAM Journal on Control and Optimization, 47(4):1643–1666, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lewis:2006:DUC

[LHS06] Adrian Lewis, René Henrion, and Alberto Seeger. Distance to uncontrollability for convex processes. SIAM Journal on Control and Optimization, 45 (1):26–50, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:2012:ODW

[LHX12] Liangliang Li, Yu Huang, and Mingqing Xiao. Observer design for wave equations with van der Pol type boundary conditions. SIAM Journal on Control and Optimization, 50 (3):1200–1219, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:1994:SLC

[Li94a]  Wu Li.  Sharp Lipschitz

constants for basic optimal solutions and basic feasible solutions of linear programs. SIAM Journal on Control and Optimization, 32 (1):140–153, January 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/22723.

Logemann:1994:ASC

[LI94b] Hartmut Logemann and Achim Ilchmann. An adaptive servomechanism for a class of infinite-dimensional systems. SIAM Journal on Control and Optimization, 32(4):917–936, July 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/21748.

Li:1995:EBP

[Li95] Wu Li. Error bounds for piecewise convex quadratic programs and applications. SIAM Journal on Control and Optimization, 33(5): 1510–1529, September 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/24302.

Lee:2000:ANB

[LI00] Hyung-Chun Lee and O. Yu. Imanuvilov. Analysis of Neumann boundary optimal control problems for the stationary Boussinesq equations including solid media. SIAM Journal on Control and Optimization, 39 (2):457–477, March 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34711.

Li:2007:MDG

[Li07] Desheng Li. Morse decompositions for general dynamical systems and differential inclusions with applications to control systems. SIAM Journal on Control and Optimization, 46(1):35–60, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liang:2015:SCR

[Lia15] Gechun Liang. Stochastic control representations for penalized backward stochastic differential equations. SIAM Journal on Control and Optimization, 53(3):1440–1463, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Licea:2013:RSL

[Lic13] Gerardo Sánchez Licea. Relaxing strengthened Legendre– Clebsch condition. SIAM Journal on Control and Optimization, 51(5):3886–3902, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lidbetter:2013:SGM

[Lid13] Thomas Lidbetter. Search games with multiple hidden objects. SIAM Journal on Control and Optimization, 51 (4):3056–3074, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lidbetter:2015:CGI

[Lid15] Thomas Lidbetter. A caching game with infinitely divisible hidden material. SIAM Journal on Control and Optimization, 53(5):3040–3056, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lim:2005:MVH

[Lim05] Andrew E. B. Lim. Meanvariance hedging when there are jumps. SIAM Journal on Control and Optimization, 44(5):1893–1922, September 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/61093.

Linnemann:2007:CNC

[Lin07] Arno Linnemann. Computing the L_{1}-norm of continuoustime linear systems. SIAM Journal on Control and Optimization, 46(6):2052–2070, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lin:2014:CND

[Lin14] Feng Lin. Control of networked discrete event systems: Dealing with communication delays and losses. SIAM Journal on Control and Optimization, 52(2):1276–1298, ???? 2014. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Lipp:1997:BCF

[Lip97] Stephen C. Lipp. Brachistochrone with Coulomb friction. SIAM Journal on Control and Optimization, 35(2):562–584, March 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28795.

Lissy:2014:CFC

[Lis14] Pierre Lissy. On the cost of fast controls for some families of dispersive or parabolic equations in one space dimension. SIAM Journal on Control and Optimization, 52(4): 2651–2676, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:1988:EDP

[Liu88] Kang Sheng Liu. Energy decay problems in the design of a point stabilizer for coupled string vibrating systems. SIAM Journal on Control and Optimization, 26(6): 1348–1356, November 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:1995:SAN

[Liu95a] Jiming Liu. Sensitivity analysis in nonlinear programs and variational inequalities via continuous selections. SIAM Journal on Control and Optimization, 33(4):1040–1060, July 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24887.

Liu:1995:SSV

[Liu95b] Jiming Liu. Strong stability in variational inequalities. SIAM Journal on Control and Optimization, 33(3):725–749, May 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24052.

Liu:1997:LDC

[Liu97a] Kangsheng Liu. Locally distributed control and damping for the conservative systems. SIAM Journal on Control and Optimization, 35(5): 1574–1590, September 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28492.

Liu:1997:AAN

[Liu97b] Wensheng Liu. An approximation algorithm for nonholonomic systems. SIAM Journal on Control and Optimization, 35(4):1328–1365, July 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26050.

Liu:1997:ATH

[Liu97c] Wensheng Liu. Averaging theorems for highly oscillatory differential equations and iterated Lie brackets. SIAM Journal on Control and Optimization, 35(6):1989–2020, November 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26866.

Liu:2003:BFS

[Liu03] Weijiu Liu. Boundary feedback stabilization of an unstable heat equation. SIAM Journal on Control and Optimization, 42(3):1033–1043, May 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40241.

Liu:2008:MEO

[Liu08] Weijiu Liu. Mixing enhancement by optimal flow advection. SIAM Journal on Control and Optimization, 47(2): 624–638, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2011:FCP

[Liu11] Qing Liu. Fattening and comparison principle for level-set equations of mean curvature type. SIAM Journal on Control and Optimization, 49(6): 2518–2541, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i6/p2518_s1.

Liu:2013:GAR

[Liu13] Yungang Liu. Global asymptotic regulation via timevarying output feedback for a class of uncertain nonlinear systems. SIAM Journal on Control and Optimization, 51(6):4318–4342, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2014:CSC

[Liu14a]  Xu Liu.  Controllability

of some coupled stochastic parabolic systems with fractional order spatial differential operators by one control in the drift. SIAM Journal on Control and Optimization, 52 (2):836–860, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2014:GFT

[Liu14b] Yungang Liu. Global finitetime stabilization via timevarying feedback for uncertain nonlinear systems. SIAM Journal on Control and Optimization, 52(3):1886–1913, ???? 2014. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2018:SPM

[Liu18]  Xuan Liu.  A stochastic

Pontryagin maximum principle on the Sierpinski gasket. SIAM Journal on Control and Optimization, 56(6): 4288–4308, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2023:ETD

[Liu23] Weijiu Liu. Exponential tracking and disturbance rejection for nonlinear reaction convection diffusion equations via boundary control. SIAM Journal on Control and Optimization, 61(1):151–169, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1456546.

Lee:2008:OOR

[LK08] Ji-Woong Lee and Pramod P.

Khargonekar. Optimal output regulation for discretetime switched and Markovian jump linear systems. SIAM Journal on Control and Optimization, 47(1):40–72, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2010:CTS

[LK10] Shu-Jun Liu and Miroslav Krstic. Continuous-time stochastic averaging on the infinite interval for locally Lipschitz systems. SIAM Journal on Control and Optimization, 48(5):3589–3622, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2011:SNE

[LK11] Shu-Jun Liu and Miroslav Krstić. Stochastic Nash equilibrium seeking for games with general nonlinear payoffs. SIAM Journal on Control and Optimization, 49(4): 1659–1679, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i4/p1659_s1.

Li:2023:PTM

[LK23] Wuquan Li and Miroslav Krstic. Prescribed-time mean-nonovershooting control under finite-time vanishing noise. SIAM Journal on Control and Optimization, 61 (3):1187–1212, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1471171.

Liukkonen:1994:CIR

[LL94] John R. Liukkonen and Arnold Levine. On convergence of iterated random maps. SIAM Journal on Control and Optimization, 32(6):

1752–1762, November 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/23897.

Liu:1998:EDE

[LL98] Kangsheng Liu and Zhuangyi Liu. Exponential decay of energy of the Euler–Bernoulli beam with locally distributed Kelvin–Voigt damping. SIAM Journal on Control and Optimization, 36(3):1086–1098, May 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31070.

Lagnese:2000:DDD

[LL00a] J. E. Lagnese and G. Leugering. Dynamic domain decomposition in approximate and exact boundary control in problems of transmission for wave equations. SIAM Journal on Control and Optimization, 38(2):503–537, March 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33353.

Lewis:2000:NDS

[LL00b] A. S. Lewis and R. E. Lucchetti. Nonsmooth duality, sandwich, and squeeze theorems. SIAM Journal on Control and Optimization, 38(2):613–626, March 2000.

CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/33421.

Lawson:2007:BCF

[LL07] Jimmie Lawson and Yongdo Lim. A Birkhoff contraction formula with applications to Riccati equations. SIAM Journal on Control and Optimization, 46(3):930–951, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lekien:2009:NCC

[LL09] François Lekien and Naomi Ehrich Leonard. Nonuniform coverage and cartograms. SIAM Journal on Control and Optimization, 48(1):351–372, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:2013:SDT

[LL13] Chanying Li and James Lam. Stabilization of discrete-time nonlinear uncertain systems by feedback based on LS algorithm. SIAM Journal on Control and Optimization, 51(2): 1128–1151, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2015:ACF

[LL15] Zhenhai Liu and Xiuwen Li. Approximate controllability of fractional evolution systems with Riemann–Liouville fractional derivatives. SIAM Journal on Control and Optimization, 53(4):1920–1933, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:2018:PPC

[LL18a] Fengzhong Li and Yungang Liu. Prescribed-performance control design for purefeedback nonlinear systems with most severe uncertainties. SIAM Journal on Control and Optimization, 56(1): 517–537, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:2018:SAO

[LL18b] Xiuwen Li and Zhenhai Liu.

Sensitivity analysis of optimal control problems described by differential hemivariational inequalities. SIAM Journal on Control and Optimization, 56 (5):3569–3597, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2019:IPS

[LL19] Zhaobo Liu and Chanying Li. Is it possible to stabilize discrete-time parameterized uncertain systems growing exponentially fast? SIAM Journal on Control and Optimization, 57(3):1965–1984, ???? 2019. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Liang:2020:CAG

[LL20] Zongxia Liang and Yang Liu. A classification approach to general S-shaped utility optimization with principals’ constraints. SIAM Journal on Control and Optimization, 58 (6):3734–3762, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Laurent:2021:LDL

[LL21a] Camille Laurent and Matthieu Léautaud. Logarithmic decay for linear damped hypoelliptic wave and Schrödinger equations. SIAM Journal on Control and Optimization, 59(3): 1881–1902, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lu:2021:LGA

[LL21b] Maobin Lu and Lu Liu. A low gain approach to output consensus of networked heterogeneous linear multi-agent systems. SIAM Journal on Control and Optimization, 59(6): 4295–4313, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liao:2024:ECR

[LL24] Zhonghua Liao and Qi Lü. Exact controllability for a refined stochastic wave equation. SIAM Journal on Control and Optimization, 62(1): 563–580, February 2024. CO-DEN SJCODC. ISSN 1095-7138.

Liu:2011:CNM

[LLC11] Bo Liu, Wenlian Lu, and Tianping Chen. Consensus in networks of multiagents with switching topologies modeled as adapted stochastic processes. SIAM Journal on Control and Optimization, 49(1): 227–253, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i1/p227_s1.

Lin:2018:CCS

[LLC18] Li-Gang Lin, Yew-Wen Liang, and Li-Jen Cheng. Control for a class of second-order systems via a state-dependent Riccati equation approach. SIAM Journal on Control and Optimization, 56(1):1–18, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2018:CNM

[LLJC18] Bo Liu, Wenlian Lu, Licheng Jiao, and Tianping Chen. Consensus in networks of multiagents with stochastically switching topologies and timevarying delays. SIAM Journal on Control and Optimization, 56(3):1884–1911, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2018:DOS

[LLLT18] Chongyang Liu, Ryan Loxton, Qun Lin, and Kok Lay Teo. Dynamic optimization for switched time-delay systems with state-dependent switching conditions. SIAM Journal on Control and Optimization, 56(5):3499–3523, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2015:ACN

[LLM15] Zhenhai Liu, Xiuwen Li, and Dumitru Motreanu. Approximate controllability for nonlinear evolution hemivariational inequalities in Hilbert spaces. SIAM Journal on Control and Optimization, 53(5): 3228–3244, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lin:2024:SCT

[LLM^{+}24] Lin Lin, James Lam, Min Meng, Xiaochen Xie, Panshuo Li, Daotong Zhang, and Peng Shi. Stabilization of continuous-time probabilistic logical networks under sampling dwell time constraints. SIAM Journal on Control and Optimization, 62(2):1006– 1033, March 2024. CODEN SJCODC. ISSN 1095-7138.

Li:2002:AFE

[LLMT02] Ruo Li, Wenbin Liu, Heping Ma, and Tao Tang. Adaptive finite element approximation for distributed elliptic optimal control problems. SIAM Journal on Control and Optimization, 41(5): 1321–1349, September 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38934.

Liu:2001:ESA

[LLR01] Kangsheng Liu, Zhuangyi Liu, and Bopeng Rao. Exponential stability of an abstract nondissipative linear system. SIAM Journal on Control and Optimization, 40(1):149–165, January 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36493.

Lamnabhi-Lagarrigue:1990:SOC

[LLS90] F. Lamnabhi-Lagarrigue and G. Stefani. Singular optimal control problems: On the necessary conditions of optimality. SIAM Journal on Control and Optimization, 28(4): 823–840, July 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lagnese:1993:CPN

[LLS93] J. E. Lagnese, G. Leugering, and E. J. P. G. Schmidt. Control of planar networks of Timoshenko beams. SIAM Journal on Control and Optimization, 31(3):780–811, May 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2021:SVC

[LLS21] Kairong Liu, Meilun Li, and Zhikun She. Stability verification for a class of stochastic hybrid systems by semidefinite programming. SIAM Journal on Control and Optimization, 59(1):365–392, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2020:DET

[LLSX20] Changxin Liu, Huiping Li, Yang Shi, and Demin Xu. Distributed event-triggered model predictive control of coupled nonlinear systems. SIAM Journal on Control and Optimization, 58(2):714–734, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2018:CFS

[LLT18] Changliu Liu, Chung-Yen Lin, and Masayoshi Tomizuka. The convex feasible set algorithm for real time optimization in motion planning. SIAM Journal on Control and Optimization, 56(4):2712–2733, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2011:DCM

[LLX11] Shuai Liu, Tao Li, and Lihua Xie. Distributed consensus for multiagent systems with communication delays and limited data rate. SIAM Journal on Control and Optimization, 49(6):2239– 2262, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i6/p2239_s1.

Lions:1982:OCSa

[LM82a] Pierre-Louis Lions and José-Luis Menaldi. Optimal control of stochastic integrals and Hamilton–Jacobi– Bellman equations. I. SIAM Journal on Control and Optimization, 20(1):58–81, January 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lions:1982:OCSb

[LM82b] Pierre-Louis Lions and José-Luis Menaldi. Optimal control of stochastic integrals and Hamilton–Jacobi– Bellman equations. II. SIAM Journal on Control and Optimization, 20(1):82–95, January 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lepeltier:1984:TGD

[LM84] J. P. Lepeltier and B. Marchal. Théorie générale du contrôle impulsionnel markovien. (French) [General theory of Markovian impulsional control]. SIAM Journal on Control and Optimization, 22(4):645–665, July 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lewis:1997:CCS

[LM97a] Andrew D. Lewis and Richard M. Murray. Configuration controllability of simple mechanical control systems. SIAM Journal on Control and Optimization, 35(3):766–790, May 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28715.

Lindquist:1997:OIS

[LM97b] Anders Lindquist and György Michaletzky. Output-induced subspaces, invariant directions, and interpolation in linear discrete-time stochastic systems. SIAM Journal on Control and Optimization, 35(3):810–859, May 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26867.

Lopez:2000:SFM

[LM00] C. López and E. Martínez. Sub-Finslerian metric associated to an optimal control system. SIAM Journal on Control and Optimization, 39(3):798–811, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35756.

Logemann:2002:DTS

[LM02] H. Logemann and A. D.

Mawby. Discrete-time and sampled-data low-gain control of infinite-dimensional linear systems in the presence of input hysteresis. SIAM Journal on Control and Optimization, 41(1):113–140, January 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38577.

Losse:2008:COS

[LM08] Philip Losse and Volker Mehrmann. Controllability and observability of second order descriptor systems. SIAM Journal on Control and Optimization, 47(3):1351–1379, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Laschov:2013:MTC

[LM13a] Dmitriy Laschov and Michael Margaliot. Minimum-time control of Boolean networks. SIAM Journal on Control and Optimization, 51(4):2869– 2892, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Levinson:2013:ODF

[LM13b] Yaron Levinson and Leonid Mirkin. L^{2} optimization in discrete FIR estimation: Exploiting state-space structure. SIAM Journal on Control and Optimization, 51(1):419–441, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Laraki:2015:IGD

[LM15] Rida Laraki and Panayotis Mertikopoulos. Inertial game dynamics and applications to constrained optimization. SIAM Journal on Control and Optimization, 53(5):3141– 3170, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:2016:WSM

[LM16] Juan Li and Hui Min. Weak solutions of mean-field stochastic differential equations and application to zerosum stochastic differential games. SIAM Journal on Control and Optimization, 54(3): 1826–1858, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lefter:2022:ICP

[LM22] Catalin-George Lefter and Elena-Alexandra Melnig. Internal controllability of parabolic systems with star- and treelike couplings. SIAM Journal on Control and Optimization, 60(5):3100–3126, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1483293.

Lin:2007:MARa

[LMA07a] J. Lin, A. S. Morse, and B. D. O. Anderson. The multi-agent rendezvous problem. part 1: The synchronous case. SIAM Journal on Control and Optimization, 46(6): 2096–2119, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lin:2007:MARb

[LMA07b] J. Lin, A. S. Morse, and B. D. O. Anderson. The multi-agent rendezvous problem. part 2: The asynchronous case. SIAM Journal on Control and Optimization, 46(6): 2120–2147, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lee:2009:DAS

[LMJ09] Adrian J. Lee, Laura A.

McLay, and Sheldon H. Jacobson. Designing aviation security passenger screening systems using nonlinear control. SIAM Journal on Control and Optimization, 48(4): 2085–2105, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2017:RLQ

[LMKW17] Ren Liu, Johannes Muhle-Karbe, and Marko H. Weber. Rebalancing with linear and quadratic costs. SIAM Journal on Control and Optimization, 55(6):3533–3563, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Laugesen:2015:PEN

[LMMR15] Richard S. Laugesen, Prashant G. Mehta, Sean P. Meyn, and Maxim Raginsky. Poisson’s equation in nonlinear filtering. SIAM Journal on Control and Optimization, 53(1):501–525, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:2020:DFC

[LMMY20] Xiaoyue Li, Xuerong Mao, Denis S. Mukama, and Chenggui Yuan. Delay feedback control for switching diffusion systems based on discrete-time observations. SIAM Journal on Control and Optimization, 58(5):2900–2926, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lindquist:1995:ZSF

[LMP95] Anders Lindquist, György Michaletzky, and Giorgio Picci. Zeros of spectral factors, the geometry of splitting subspaces, and the algebraic Riccati inequality. SIAM Journal on Control and Optimization, 33(2):365–401, March 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23866.

Losse:2009:MOC

[LMPR09] Philip Losse, Volker Mehrmann, Lisa Katrin Poppe, and Timo Reis. The modified optimal H_{∞} control problem for descriptor systems. SIAM Journal on Control and Optimization, 47(6):2795–2811, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lototsky:1997:NFR

[LMR97] Sergey Lototsky, Remigijus Mikulevicius, and Boris L. Rozovskii. Nonlinear filtering revisited: a spectral approach. SIAM Journal on Control and Optimization, 35(2):435–461, March 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/24891.

Li:2023:GDT

[LMYZ23] Zhanjie Li, Yajing Ma, Dong Yue, and Jun Zhao. Guaranteed dwell-times of hysteresis switching for switched nonlinear systems via switching characteristic indices. SIAM Journal on Control and Optimization, 61(3):1610–1630, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1471699.

Li:2020:CDS

[LMZ20] Zhanjie Li, Yajing Ma, and Jun Zhao. Control design of switched nonlinear systems: an intermittent compensation switching strategy. SIAM Journal on Control and Optimization, 58(6):3684–3708, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lakshmivarahan:1982:LAT

[LN82] S. Lakshmivarahan and K. S. Narendra. Learning algorithms for two-person zerosum stochastic games with incomplete information: a unified approach. SIAM Journal on Control and Optimization, 20(4):541–552, July 1982. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lamboley:2010:POS

[LN10] Jimmy Lamboley and Arian Novruzi. Polygons as optimal shapes with convexity constraint. SIAM Journal on Control and Optimization, 48 (5):3003–3025, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Litsyn:2000:SLD

[LNP00] Elena Litsyn, Yurii V. Nepomnyashchikh, and Arcady Ponosov. Stabilization of linear differential systems via hybrid feedback controls. SIAM Journal on Control and Optimization, 38(5):1468–1480, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33255.

Liu:2002:ESO

[LNT02] W. B. Liu, P. Neittaanmäki, and D. Tiba. Existence for shape optimization problems in arbitrary dimension. SIAM Journal on Control and Optimization, 41(5): 1440–1454, September 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38814.

Li:2023:LQL

[LNW23] Min Li, Tianyang Nie, and Zhen Wu. Linear-quadratic large-population problem with partial information: Hamiltonian approach and Riccati approach. SIAM Journal on Control and Optimization, 61 (4):2114–2139, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1414152.

Leonard:2015:CLM

[LO15] Naomi Ehrich Leonard and Alex Olshevsky. Cooperative learning in multiagent systems from intermittent measurements. SIAM Journal on Control and Optimization, 53 (1):1–29, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lobry:1970:CSN

[Lob70] Claude Lobry. Contrôlabilité des systèmes non linéaires. (French) [Controllability of nonlinear systems]. SIAM Journal on Control, 8(4): 573–605, ???? 1970. CO-DEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic). See erratum [Lob76].

Lobry:1976:ECS

[Lob76] C. Lobry. Erratum: Contrôlabilité des systèmes non linéaires. (French) [Erratum: Controllability of nonlinear systems]. SIAM Journal on Control and Optimization, 14(2): 387, February 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [Lob70].

Logemann:2013:SWP

[Log13]  H. Logemann.  Stabiliza-

tion of well-posed infinitedimensional systems by dynamic sampled-data feedback.

SIAM Journal on Control and Optimization, 51(2):1203– 1231, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Loiseau:1988:MSI

[Loi88] Jean Jacques Loiseau. Sur la modification de la structure à l’infini par un retour d’état statique. (French) [On the modification of the structure at infinity by a return to static state]. SIAM Journal on Control and Optimization, 26 (2):251–273, March 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lomadze:2012:CM

[Lom12a] Vakhtang Lomadze. Controllability as minimality. SIAM Journal on Control and Optimization, 50(1):357– 367, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v50/i1/p357_s1.

Lomadze:2012:LST

[Lom12b]  Vakhtang Lomadze.  Lin-

ear systems and Taylor complexes. SIAM Journal on Control and Optimization, 50(3): 1721–1733, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lomadze:2013:RPM

[Lom13] Vakhtang Lomadze. “Reduced polynomial matrices” in several variables. SIAM Journal on Control and Optimization, 51(4):3258–3273, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lomadze:2018:KMM

[Lom18] Vakhtang Lomadze. KW models for (multivariate) linear differential systems. SIAM Journal on Control and Optimization, 56(1):456–472, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lomadze:2021:DML

[Lom21] Vakhtang Lomadze. Duality for multidimensional linear systems with homological dimension ≤ 1. SIAM Journal on Control and Optimization, 59(1):417–433, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lopez:2012:AEP

[Lóp12] Rubén López. Approximations of equilibrium problems. SIAM Journal on Control and Optimization, 50(2): 1038–1070, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See erratum [Lóp13].

Lopez:2013:EAE

[Lóp13] R. López. Erratum: Approximations of equilibrium problems. SIAM Journal on Control and Optimization, 50(6): 3374, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [Lóp12].

Loreti:1987:SPC

[Lor87] Paola Loreti. Some properties of constrained viscosity solutions of Hamilton– Jacobi–Bellman equations. SIAM Journal on Control and Optimization, 25(5):1244– 1252, September 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lorenz:2008:VTM

[Lor08] Thomas Lorenz. A viability theorem for morphological inclusions. SIAM Journal on Control and Optimization, 47(3):1591–1614, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lorenz:2010:MCP

[Lor10] Thomas Lorenz. Morphological control problems with state constraints. SIAM Journal on Control and Optimization, 48(8):5510–5546, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Loubaton:1994:RUD

[Lou94] Philippe Loubaton. Regular unitary dilation of commuting contractions and Markovian representation of Gaussian stationary processes on Z^{2}. SIAM Journal on Control and Optimization, 32 (2):332–357, March 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/18912.

Lou:2002:OCP

[Lou02] Hongwei Lou. An optimal control problem governed by quasi-linear variational inequalities. SIAM Journal on Control and Optimization, 41(4):1229–1253, July 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37503.

Lou:2003:MPO

[Lou03] Hongwei Lou. Maximum principle of optimal control for degenerate quasi-linear elliptic equations. SIAM Journal on Control and Optimization, 42(1):1–23, January 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40166.

Lou:2007:ENR

[Lou07] Hongwei Lou. Existence and nonexistence results of an optimal control problem by using relaxed control. SIAM Journal on Control and Optimization, 46(6):1923–1941, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Louisell:2015:MPS

[Lou15] James Louisell. Matrix polynomials, similar operators, and the imaginary axis eigenvalues of a matrix delay equation. SIAM Journal on Control and Optimization, 53(1): 399–413, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Levine:1979:SCK

[LP79a] P. Lévine and J. Ch. Pomerol.

Sufficient conditions for Kuhn– Tucker vectors in convex programming. SIAM Journal on Control and Optimization, 17(6):689–699, November 1979. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See erratum [LP81].

Lindquist:1979:SRP

[LP79b] Anders Lindquist and Giorgio Picci. On the stochastic realization problem. SIAM Journal on Control and Optimization, 17(3):365–389, May 1979. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Levine:1981:ESC

[LP81] P. Levine and J. Ch. Pomerol. Erratum: Sufficient conditions for Kuhn–Tucker vectors in convex programming. SIAM Journal on Control and Optimization, 19(3):431–432, May 1981. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [LP79a].

Lindquist:1985:RTM

[LP85] Anders Lindquist and Giorgio Picci. Realization theory for multivariate stationary Gaussian processes. SIAM Journal on Control and Optimization, 23(6):809–857, November 1985. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lions:1986:QVI

[LP86] P. L. Lions and B. Perthame. Quasi-variational inequalities and ergodic impulse control. SIAM Journal on Control and Optimization, 24(4):604–615, July 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lin:1987:IML

[LP87] Y. Y. Lin and J.-S. Pang. Iterative methods for large convex quadratic programs: a survey. SIAM Journal on Control and Optimization, 25(2): 383–411, March 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lewis:2010:LBR

[LP10] Adrian S. Lewis and C. H. Jeffrey Pang. Lipschitz behavior of the robust regularization. SIAM Journal on Control and Optimization, 48(5): 3080–3104, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Laakkonen:2015:FDR

[LP15] Petteri Laakkonen and Seppo Pohjolainen. Frequency domain robust regulation of signals generated by an infinitedimensional exosystem. SIAM Journal on Control and Optimization, 53(1):139–166, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Legat:2020:CUS

[LPJ20] Benoît Legat, Pablo Parrilo, and Raphaël Jungers. Certifying unstability of switched systems using sum of squares programming. SIAM Journal on Control and Optimization, 58(4):2616–2638, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Loreti:2012:BCO

[LPS12] P. Loreti, L. Pandolfi, and D. Sforza. Boundary controllability and observability of a viscoelastic string. SIAM Journal on Control and Optimization, 50(2):820–844, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lhachemi:2022:PIR

[LPT22] Hugo Lhachemi, Christophe Prieur, and Emmanuel Trélat. Proportional integral regulation control of a onedimensional semilinear wave equation. SIAM Journal on Control and Optimization, 60(1):1–21, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1346857.

Levan:1979:SSL

[LR79a] N. Levan and L. Rigby. Strong stabilizability of linear contractive control systems on Hilbert space. SIAM Journal on Control and Optimization, 17(1):23–35, January 1979. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Links:1979:FPC

[LR79b] Graham Links and David S. Rubin. Finite primal cutting plane algorithms for integer programs. SIAM Journal on Control and Optimization, 17 (1):47–55, January 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:1992:LCO

[LR92] Wen Bin Liu and J. E. Rubio. Local convergences and optimal shape design. SIAM Journal on Control and Optimization, 30(1):49–62, January 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Loewen:1994:OCU

[LR94] Philip D. Loewen and R. T.

Rockafellar. Optimal control of unbounded differential inclusions. SIAM Journal on Control and Optimization, 32(2):442–470, March 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/21749.

Loewen:1996:NNC

[LR96] P. D. Loewen and R. T. Rockafellar. New necessary conditions for the generalized problem of Bolza. SIAM Journal on Control and Optimization, 34(5):1496–1511, September 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27593.

Loewen:1997:BPG

[LR97] P. D. Loewen and R. T.

Rockafellar. Bolza problems with general time constraints. SIAM Journal on Control and Optimization, 35(6):2050–2069, November 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29880.

LeBreton:2000:AOA

[LR00a] Alain Le Breton and Marie-Christine Roubaud. Asymptotic optimality of approximate filters in stochastic systems with colored noises. SIAM Journal on Control and Optimization, 39(3):917–927, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33390.

Logemann:2000:TVA

[LR00b] H. Logemann and E. P. Ryan. Time-varying and adaptive integral control of infinitedimensional regular linear systems with input nonlinearities. SIAM Journal on Control and Optimization, 38(4):1120–1144, July 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33922.

Li:2003:EBC

[LR03] Ta-Tsien Li and Bo-Peng Rao. Exact boundary controllability for quasi-linear hyperbolic systems. SIAM Journal on Control and Optimization, 41(6):1748–1755, November 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39009.

Loreti:2006:OED

[LR06] Paola Loreti and Bopeng Rao.

Optimal energy decay rate for partially damped systems by spectral compensation. SIAM Journal on Control and Optimization, 45(5):1612–1632, ???? 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Levi:2010:NFA

[LR10] A. F. J. Levi and I. G. Rosen.

A novel formulation of the adjoint method in the optimal design of quantum electronic devices. SIAM Journal on Control and Optimization, 48(5):3191–3223, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:2016:CKT

[LR16] Tatsien Li and Bopeng Rao.

Criteria of Kalman’s type to the approximate controllability and the approximate synchronization for a coupled system of wave equations with Dirichlet boundary controls. SIAM Journal on Control and Optimization, 54 (1):49–72, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:2021:USA

[LR21] Tatsien Li and Bopeng Rao. Uniform synchronization of an abstract linear second order evolution system. SIAM Journal on Control and Optimization, 59(4):2740–2755, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:2023:ESS

[LR23a] Tatsien Li and Bopeng Rao. Exactly synchronizable state and approximate controllability for a coupled system of wave equations with locally distributed controls. SIAM Journal on Control and Optimization, 61(3):1460–1471, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1488430.

Lin:2023:ITD

[LR23b] Zhengqi Lin and Andrzej Ruszczyński. An integrated transportation distance between kernels and approximate dynamic risk evaluation in Markov systems. SIAM Journal on Control and Optimization, 61(6):3559–3583, December 2023. CODEN SJ-CODC. ISSN 1095-7138.

Liptser:1996:DAS

[LRT96] R. Sh. Liptser, W. J. Runggaldier, and M. Taksar. Deterministic approximation for stochastic control problems. SIAM Journal on Control and Optimization, 34 (1):161–178, January 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25454.

Logemann:1998:ICI

[LRT98] H. Logemann, E. P. Ryan, and S. Townley. Integral control of infinite-dimensional linear systems subject to input saturation. SIAM Journal on Control and Optimization, 36(6):1940–1961, November 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31414.

Logemann:2005:GSD

[LRT05] Hartmut Logemann, Richard Rebarber, and Stuart Townley. Generalized sampled-data stabilization of well-posed linear infinite-dimensional systems. SIAM Journal on Control and Optimization, 44(4):1345–1369, July 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43434.

Lin:2022:RHP

[LRTY22] Yiqing Lin, Zhenjie Ren, Nizar Touzi, and Junjian Yang. Random horizon principal-agent problems. SIAM Journal on Control and Optimization, 60(1):355–384, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1321620.

Logemann:1996:CRN

[LRW96] Hartmut Logemann, Richard Rebarber, and George Weiss. Conditions for robustness and nonrobustness of the stability of feedback systems with respect to small delays in the feedback loop. SIAM Journal on Control and Optimization, 34(2):572–600, March 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/25070.

Lions:1985:DGO

[LS85a] P.-L. Lions and P. E. Souganidis. Differential games, optimal control and directional derivatives of viscosity solutions of Bellman’s and Isaacs’ equations. SIAM Journal on Control and Optimization, 23 (4):566–583, July 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lojasiewicz:1985:SER

[ŁS85b] S. Łojasiewicz, Jr. and H. J. Sussmann. Some examples of reachable sets and optimal cost functions that fail to be subanalytic. SIAM Journal on Control and Optimization, 23 (4):584–598, July 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lions:1986:DGD

[LS86] P.-L. Lions and P. E. Souganidis. Differential games and directional derivatives of viscosity solutions of Isaacs’ equations. II. SIAM Journal on Control and Optimization, 24 (5):1086–1089, July 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lakner:1991:OCB

[LS91a] Peter Lakner and Eric Slud.

Optimal consumption by a bond investor: The case of random interest rate adapted to a point process. SIAM Journal on Control and Optimization, 29(3):638–655, May 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lasiecka:1991:SAO

[LS91b] Irena Lasiecka and Jan Sokołowski.

Sensitivity analysis of optimal control problems for wave equations. SIAM Journal on Control and Optimization, 29(5):1128–1149, September 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ledzewicz:1998:HOA

[LS98] Urszula Ledzewicz and Heinz Schättler. High-order approximations and generalized necessary conditions for optimality. SIAM Journal on Control and Optimization, 37(1):33–53, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/31774.

Leibfritz:1999:ISI

[LS99a] F. Leibfritz and E. W. Sachs. Inexact SQP interior point methods and large scale optimal control problems. SIAM Journal on Control and Optimization, 38(1):272–293, January 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/29879.

Liu:1999:CDR

[LS99b] Wensheng Liu and Héctor J. Sussmann. Continuous dependence with respect to the input of trajectories of controlaffine systems. SIAM Journal on Control and Optimization, 37(3):777–803, May 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/23737.

Ledzewicz:2000:HOG

[LS00] Urszula Ledzewicz and Heinz Schättler. A high-order generalized local maximum principle. SIAM Journal on Control and Optimization, 38(3):823–854, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33820.

Leugering:2002:MSF

[LS02] Guenter Leugering and J. P. Georg Schmidt. On the modelling and stabilization of flows in networks of open canals. SIAM Journal on Control and Optimization, 41 (1):164–180, January 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37566.

Laraki:2005:VZS

[LS05a] Rida Laraki and Eilon Solan.

The value of zero-sum stopping games in continuous time. SIAM Journal on Control and Optimization, 43(5): 1913–1922, September 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42902.

Lee:2005:AOC

[LS05b] Hyung-Chun Lee and Timofey Shilkin. Analysis of optimal control problems for the two-dimensional thermistor system. SIAM Journal on Control and Optimization, 44(1):268–282, January 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43476.

Ledzewicz:2007:ATC

[LS07] Urszula Ledzewicz and Heinz Schättler. Antiangiogenic therapy in cancer treatment as an optimal control problem. SIAM Journal on Control and Optimization, 46(3): 1052–1079, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Leung:2009:EHO

[LS09] Tim Leung and Ronnie Sircar. Exponential hedging with optimal stopping and application to employee stock option valuation. SIAM Journal on Control and Optimization, 48(3): 1422–1451, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:2019:SSO

[LS19a] Chen Li and Georg Stadler. Sparse solutions in optimal control of PDEs with uncertain parameters: The linear case. SIAM Journal on Control and Optimization, 57(1): 633–658, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liang:2019:DGP

[LS19b] Gechun Liang and Haodong Sun. Dynkin games with Poisson random intervention times. SIAM Journal on Control and Optimization, 57(4): 2962–2991, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Leyva:2013:GCS

[LSDS13] Horacio Leyva, Julio Solís-Daun, and Rodolfo Suárez. Global CLF stabilization of systems with control inputs constrained to a hyperbox. SIAM Journal on Control and Optimization, 51(1):745–766, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lu:2018:SBC

[LSL^{+}18] Jianquan Lu, Liangjie Sun, Yang Liu, Daniel W. C. Ho, and Jinde Cao. Stabilization of Boolean control networks under aperiodic sampled-data control. SIAM Journal on Control and Optimization, 56 (6):4385–4404, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lin:1996:SCL

[LSW96] Yuandan Lin, Eduardo D.

Sontag, and Yuan Wang. A smooth converse Lyapunov theorem for robust stability. SIAM Journal on Control and Optimization, 34 (1):124–160, January 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25998.

Lubkoll:2014:OCP

[LSW14] Lars Lubkoll, Anton Schiela, and Martin Weiser. An optimal control problem in polyconvex hyperelasticity. SIAM Journal on Control and Optimization, 52(3):1403–1422, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liang:2017:CCS

[LSW^{+}17] Quanyi Liang, Zhikun She, Lei Wang, Michael Z. Q. Chen, and Qing-Guo Wang. Characterizations and criteria for synchronization of heterogeneous networks to linear subspaces. SIAM Journal on Control and Optimization, 55(6): 4048–4071, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lasiecka:1983:DBC

[LT83a] I. Lasiecka and R. Triggiani. Dirichlet boundary control problem for parabolic equations with quadratic cost: Analyticity and Riccati’s feedback synthesis. SIAM Journal on Control and Optimization, 21(1):41–67, January 1983. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lasiecka:1983:SSA

[LT83b] I. Lasiecka and R. Triggiani.

Stabilization and structural assignment of Dirichlet boundary feedback parabolic equations. SIAM Journal on Control and Optimization, 21 (5):766–803, September 1983. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lasiecka:1986:REH

[LT86] I. Lasiecka and R. Triggiani.

Riccati equations for hyperbolic partial differential equations with L_{2}(0, T; L_{2}(Γ))– Dirichlet boundary terms. SIAM Journal on Control and Optimization, 24(5):884–925, July 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lasiecka:1989:ECE

[LT89] I. Lasiecka and R. Triggiani. Exact controllability of the Euler–Bernoulli equation with controls in the Dirichlet and Neumann boundary conditions: a nonconservative case. SIAM Journal on Control and Optimization, 27(2): 330–373, March 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Luo:1991:CMS

[LT91] Zhi-Quan Luo and Paul Tseng. On the convergence of a matrix splitting algorithm for the symmetric monotone linear complementarity problem. SIAM Journal on Control and Optimization, 29(5):1037– 1060, September 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Luo:1992:LCD

[LT92] Zhi-Quan Luo and Paul Tseng. On the linear convergence of descent methods for convex essentially smooth minimization. SIAM Journal on Control and Optimization, 30(2):408–425, March 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Logemann:1997:LGC

[LT97] Hartmut Logemann and Stuart Townley. Low-gain control of uncertain regular linear systems. SIAM Journal on Control and Optimization, 35(1):78–116, January 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/27592.

Logemann:2003:ALG

[LT03] Hartmut Logemann and Stuart Townley. Adaptive lowgain integral control of multivariable well-posed linear systems. SIAM Journal on Control and Optimization, 41(6): 1722–1732, November 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39668.

Loheac:2013:MPB

[LT13] Jérôme Lohéac and Marius Tucsnak. Maximum principle and bang–bang property of time optimal controls for Schrödinger-type systems. SIAM Journal on Control and Optimization, 51(5): 4016–4038, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2018:HSM

[LT18] Jun Liu and Andrew R. Teel.

Hybrid systems with memory: Existence and wellposedness of generalized solutions. SIAM Journal on Control and Optimization, 56(2): 1011–1037, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2023:MPO

[LT23] Guomin Liu and Shanjian Tang. Maximum principle for optimal control of stochastic evolution equations with recursive utilities. SIAM Journal on Control and Optimization, 61(6):3467–3500, November 2023. CODEN SJCODC. ISSN 1095-7138.

Lu:2021:EBS

[LtLR21] Xing Lu, Ta tsien Li, and Bopeng Rao. Exact boundary synchronization by groups for a coupled system of wave equations with coupled Robin boundary controls on a general bounded domain. SIAM Journal on Control and Optimization, 59(6):4457–4480, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:2024:DTA

[LTT24] Yunzhang Li, Xiaolu Tan, and Shanjian Tang. Discrete-time approximation of stochastic optimal control with partial observation. SIAM Journal on Control and Optimization, 62(1):326–350, January 2024. CODEN SJCODC. ISSN 1095-7138.

Lu:2013:OES

[Lü13] Qi Lü. Observability estimate for stochastic Schrödinger equations and its applications. SIAM Journal on Control and Optimization, 51(1):121–144, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lu:2014:ECS

[Lü14] Qi Lü. Exact controllability for stochastic transport equations. SIAM Journal on Control and Optimization, 52(1):

397–419, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lu:2015:SWP

[Lü15] Qi Lü. Stochastic well-posed systems and well-posedness of some stochastic partial differential equations with boundary control and observation. SIAM Journal on Control and Optimization, 53(6):3457– 3482, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lukes:1985:AFC

[Luk85] Dahlard L. Lukes. Affine feedback controllability of constant coefficient differential equations. SIAM Journal on Control and Optimization, 23 (6):952–972, November 1985. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lunardi:1991:STP

[Lun91]  Alessandra Lunardi.  Sta-

bilizability of time-periodic parabolic equations. SIAM Journal on Control and Optimization, 29(4):810–828, July 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Luque:1984:ACA

[Luq84] Fernando Javier Luque. Asymptotic convergence analysis of the proximal point algorithm. SIAM Journal on Control and Optimization, 22(2): 277–293, March 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Loreti:1997:PEC

[LV97] Paola Loreti and Vanda Valente. Partial exact controllability for spherical membranes. SIAM Journal on Control and Optimization, 35(2):641–653, March 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26962.

Leykekhman:2016:PEE

[LV16] Dmitriy Leykekhman and Boris Vexler. A priori error estimates for three dimensional parabolic optimal control problems with pointwise control. SIAM Journal on Control and Optimization, 54 (5):2403–2435, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lai:1987:AES

[LW87] T. L. Lai and C. Z. Wei. Asymptotically efficient selftuning regulators. SIAM Journal on Control and Optimization, 25(2):466–481, March 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lei:2007:OCS

[LW07] Ling Lei and Gengsheng Wang. Optimal control of semilinear parabolic equations with k-approximate periodic solutions. SIAM Journal on Control and Optimization, 46 (5):1754–1778, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Le:2010:FTH

[LW10] H. Le and C. Wang. A finite time horizon optimal stopping problem with regime switching. SIAM Journal on Control and Optimization, 48(8): 5193–5213, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lin:2011:BTO

[LW11a] Ping Lin and Gengsheng Wang. Blowup time optimal control for ordinary differential equations. SIAM Journal on Control and Optimization, 49(1):73–105, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i1/p73_s1.

Lu:2011:ETO

[LW11b] Qi Lü and Gengsheng Wang.

On the existence of time optimal controls with constraints of the rectangular type for heat equations.

SIAM Journal on Control and Optimization, 49(3):1124– 1149, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i3/p1124_s1.

Li:2012:SML

[LW12] Desheng Li and Yanling Wang. Smooth Morse– Lyapunov functions of strong attractors for differential inclusions. SIAM Journal on Control and Optimization, 50 (1):368–387, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v50/i1/p368_s1.

Li:2014:OCP

[LW14] Juan Li and Qingmeng Wei. Optimal control problems of fully coupled FBSDEs and viscosity solutions of Hamilton– Jacobi–Bellman equations. SIAM Journal on Control and Optimization, 52(3):1622– 1662, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Laurain:2015:DFC

[LW15a] Antoine Laurain and Shawn W. Walker. Droplet footprint control. SIAM Journal on Control and Optimization, 53(2): 771–799, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:2015:CAC

[LW15b] Haitao Li and Yuzhen Wang.

Controllability analysis and control design for switched Boolean networks with state and input constraints. SIAM Journal on Control and Optimization, 53(5):2955–2979, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2015:DCL

[LW15c] Zhongchang Liu and Wing Shing Wong. Distributed control of linear systems allowing choices. SIAM Journal on Control and Optimization, 53 (2):670–689, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ludovici:2015:PEE

[LW15d] Francesco Ludovici and Winnifried Wollner. A priori error estimates for a finite element discretization of parabolic optimization problems with pointwise constraints in time on mean values of the gradient of the state. SIAM Journal on Control and Optimization, 53 (2):745–770, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:2017:LBS

[LW17] Haitao Li and Yuzhen Wang. Lyapunov-based stability and construction of Lyapunov functions for Boolean networks. SIAM Journal on Control and Optimization, 55(6): 3437–3457, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lu:2022:NCF

[LW22] Qi Lü and Yu Wang. Null controllability for fourth order stochastic parabolic equations. SIAM Journal on Control and Optimization, 60 (3):1563–1590, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1472620.

Liu:2024:NSC

[LW24] Shenyu Liu and Penghui Wen. Nonconservative stability criteria for semi-Markovian impulsive switched systems. SIAM Journal on Control and Optimization, 62(3):1783– 1808, June 2024. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1564833.

Liu:2020:SCD

[LWDB20] Zhixin Liu, Lin Wang, Daoyi Dong, and John S. Baras. Stable consensus decision making for spatially distributed multiagent systems with multiple leaders. SIAM Journal on Control and Optimization, 58(3):1626–1651, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:2022:NOS

[LWfZ22] Yun Li, Fuke Wu, and Ji feng Zhang. Near optimality of stochastic control for singularly perturbed McKean– Vlasov systems. SIAM Journal on Control and Optimization, 60(5):2859–2883, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1466177.

Li:2020:PAC

[LWLZ20] Haitao Li, Shuling Wang, Xiaodong Li, and Guodong Zhao. Perturbation analysis for controllability of logical control networks. SIAM Journal on Control and Optimization, 58(6):3632–3657, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:2024:LFR

[LWR24] Xiaoyu Li, Yuhu Wu, and Lining Ru. Leader-following rendezvous control for generalized cucker-smale model on Riemannian manifolds. SIAM Journal on Control and Optimization, 62(1):724–751, February 2024. CODEN SJ-CODC. ISSN 1095-7138.

Liu:2022:CCS

[LWXY22] Hanbing Liu, Gengsheng Wang, Yashan Xu, and Huaiqiang Yu. Characterizations of complete stabilizability. SIAM Journal on Control and Optimization, 60 (4):2040–2069, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1386761.

Laurain:2021:SOS

[LWY21] Antoine Laurain, Malte Winckler, and Irwin Yousept. Shape optimization for superconductors governed by h(curl)- elliptic variational inequalities. SIAM Journal on Control and Optimization, 59(3): 2247–2272, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2023:SLS

[LWY23] Hanbing Liu, Gengsheng Wang, and Huaiqiang Yu. Stabilizability of linear systems with discrete observation mode. SIAM Journal on Control and Optimization, 61(4):2520–2545, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1518591.

Liang:2022:RMF

[LWZ22] Yong Liang, Bing-Chang Wang, and Huanshui Zhang. Robust mean field linear quadratic social control: Openloop and closed-loop strategies. SIAM Journal on Control and Optimization, 60 (4):2184–2213, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1414140.

Liu:2019:GTC

[LWZC19] Ting Liu, Jinhuan Wang, Xiao Zhang, and Daizhan Cheng. Game theoretic control of multiagent systems. SIAM Journal on Control and Optimization, 57(3):1691–1709, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:2023:CSS

[LX23] Binjie Li and Xiaoping Xie.

Convergence of a spatial semidiscretization for a backward semilinear stochastic parabolic equation. SIAM Journal on Control and Optimization, 61(1):47–71, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1453700.

Li:2014:STS

[LXY14] Zhuchun Li, Xiaoping Xue, and Daren Yu. Synchronization and transient stability in power grids based on Lojasiewicz inequalities. SIAM Journal on Control and Optimization, 52(4):2482–2511, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lenhart:2016:OCS

[LXY16] Suzanne Lenhart, Jie Xiong, and Jiongmin Yong. Optimal controls for stochastic partial differential equations with an application in population modeling. SIAM Journal on Control and Optimization, 54 (2):495–535, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:2011:OFS

[LXZ11] Wuquan Li, Xue-Jun Xie, and Siying Zhang. Outputfeedback stabilization of stochastic high-order nonlinear systems under weaker conditions. SIAM Journal on Control and Optimization, 49(3): 1262–1282, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i3/p1262_s1.

Lee:1990:QOI

[LY90] E. Bruce Lee and Yun Cheng You. Quadratic optimization for infinite-dimensional linear differential difference type systems: Syntheses via the Fredholm equation. SIAM Journal on Control and Optimization, 28(2):265–293, March 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lai:1991:PRA

[LY91a] Tze Leung Lai and Zhiliang Ying. Parallel recursive algorithms in asymptotically efficient adaptive control of linear stochastic systems. SIAM Journal on Control and Optimization, 29(5): 1091–1127, September 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lai:1991:RIA

[LY91b] Tze Leung Lai and Zhiliang Ying. Recursive identification and adaptive prediction in linear stochastic systems. SIAM Journal on Control and Optimization, 29(5): 1061–1090, September 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:1991:NCO

[LY91c] Xun Jing Li and Jiong Min Yong. Necessary conditions for optimal control of distributed parameter systems. SIAM Journal on Control and Optimization, 29(4):895–908, July 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lucet:2001:SAV

[LY01] Yves Lucet and Jane J. Ye. Sensitivity analysis of the value function for optimization problems with variational inequality constraints. SIAM Journal on Control and Optimization, 40(3):699–723, May 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36171. See er-

ratum [LY02].

Lucet:2002:ESA

[LY02] Yves Lucet and Jane J. Ye. Erratum: Sensitivity analysis of the value function for optimization problems with variational inequality constraints. SIAM Journal on Control and Optimization, 41(4):1315–1319, July 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39926.  See

[LY01].

Lou:2009:OCS

[LY09] Hongwei Lou and Jiongmin Yong. Optimality conditions for semilinear elliptic equations with leading term containing controls. SIAM Journal on Control and Optimization, 48(4):2366–2387, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:2012:VIS

[LY12] Chong Li and Jen-Chih Yao.

Variational inequalities for setvalued vector fields on Riemannian manifolds: Convexity of the solution set and the proximal point algorithm. SIAM Journal on Control and Optimization, 50(4):2486– 2514, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:2018:FBS

[LY18] Na Li and Zhiyong Yu.

Forward–backward stochastic differential equations and linear–quadratic generalized Stackelberg games. SIAM Journal on Control and Optimization, 56(6):4148–4180, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:2019:LEO

[LY19a] Tatsien Li and Lei Yu. Local exact one-sided boundary null controllability of entropy solutions to a class of hyperbolic systems of balance laws. SIAM Journal on Control and Optimization, 57(1):610–632, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2019:CES

[LY19b] Xu Liu and Yongyi Yu.

Carleman estimates of some stochastic degenerate parabolic equations and application. SIAM Journal on Control and Optimization, 57(5):3527– 3552, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liang:2020:MDP

[LY20a] Xiaoqing Liang and Virginia R. Young. Minimizing the discounted probability of exponential Parisian ruin via reinsurance. SIAM Journal on Control and Optimization, 58 (2):937–964, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lin:2020:CSV

[LY20b] Ping Lin and Jiongmin Yong. Controlled singular Volterra integral equations and Pontryagin maximum principle. SIAM Journal on Control and Optimization, 58(1):136–164, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:2023:EBC

[LY23] Tatsien Li and Lei Yu. The exact boundary controllability of nodal profile for entropy solutions to 1-D quasilinear hyperbolic systems of conservation laws with linearly degenerate negative (resp., positive) characteristic fields. SIAM Journal on Control and Optimization, 61(5):2761–2776, ???? 2023. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1395946.

Liu:2018:OIC

[LYB18] Jingzhen Liu, Ka Fai Cedric Yiu, and Alain Bensoussan. Optimal inventory control with jump diffusion and nonlinear dynamics in the demand. SIAM Journal on Control and Optimization, 56 (1):53–74, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:1999:SHE

[LYLZ99] Shengjia Li, Jingyuan Yu, Zhandong Liang, and Guangtian Zhu. Stabilization of high eigenfrequencies of a beam equation with generalized viscous damping. SIAM Journal on Control and Optimization, 37(6):1767–1779, November 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32254.

Leung:2015:OMS

[LYZ15] Tim Leung, Kazutoshi Yamazaki, and Hongzhong Zhang. Optimal multiple stopping with negative discount rate and random refraction times under Lévy models. SIAM Journal on Control and Optimization, 53(4):2373–2405, ???? 2015. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2023:SQK

[LYZ23] Jun Liu, Weiping Yan, and Can Zhang. Stabilizability for a quasilinear Klein–Gordonwave system with variable coefficients. SIAM Journal on Control and Optimization, 61 (3):1651–1678, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1508376.

Logemann:1992:RPC

[LZ92] Hartmut Logemann and Hans Zwart. On robust PI-control of infinite-dimensional systems. SIAM Journal on Control and Optimization, 30(3):573–593, May 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:1994:UES

[LZ94] Zhuangyi Liu and Song Mu Zheng. Uniform exponential stability and approximation in control of a thermoelastic system. SIAM Journal on Control and Optimization, 32(5):1226–1246, September 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/21900.

Leizarowitz:1996:IHV

[LZ96] Arie Leizarowitz and Alexander J. Zaslavski. Infinitehorizon variational problems with nonconvex integrands. SIAM Journal on Control and Optimization, 34(4):1099–1134, July 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25512.

Lim:2001:LQC

[LZ01] Andrew E. B. Lim and Xun Yu Zhou. Linearquadratic control of backward stochastic differential equations. SIAM Journal on Control and Optimization, 40(2):450–474, March 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37473.

Liu:2006:POF

[LZ06] Yun-Gang Liu and Ji-Feng Zhang. Practical outputfeedback risk-sensitive control for stochastic nonlinear systems with stable zerodynamics. SIAM Journal on Control and Optimization, 45 (3):885–926, January 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lian:2011:CDC

[LŻ11] Jianming Lian and Stanislaw H. Żak. Controller design for a class of uncertain systems with guaranteed performance. SIAM Journal on Control and Optimization, 49(3):1239– 1261, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicon/resource/
                                        1/sjcodc/v49/i3/p1239_s1.

Liu:2012:LCM

[LZ12] Xu Liu and Xu Zhang. Local controllability of multidimensional quasi-linear parabolic equations. SIAM Journal on Control and Optimization, 50 (4):2046–2064, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lokka:2013:LTO

[LZ13] Arne Løkka and Mihail Zervos. Long-term optimal investment strategies in the presence of adjustment costs. SIAM Journal on Control and Optimization, 51(2):996–1034, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Levy:2016:CAC

[LZ16a] Bernard C. Levy and Mattia Zorzi. A contraction analysis of the convergence of risksensitive filters. SIAM Journal on Control and Optimization, 54(4):2154–2173, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:2016:SDT

[LZ16b] L. Li and H. Zhang. Stabilization of discrete-time systems with multiplicative noise and multiple delays in the control variable. SIAM Journal on Control and Optimization, 54 (2):894–917, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2016:SSL

[LZ16c] Zhuangyi Liu and Qiong Zhang. Stability of a string with local Kelvin–Voigt damping and nonsmooth coefficient at interface. SIAM Journal on Control and Optimization, 54(4):1859–1871, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:2017:ECA

[LZ17] Jiexian Li and Gordan Zitković.

Existence, characterization, and approximation in the generalized monotone-follower problem. SIAM Journal on Control and Optimization, 55 (1):94–118, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:2018:CQR

[LZ18a] Yusong Li and Harry Zheng. Constrained quadratic risk minimization via forward and backward stochastic differential equations. SIAM Journal on Control and Optimization, 56(2):1130–1153, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2018:SRS

[LZ18b] Zhuangyi Liu and Qiong Zhang. Stability and regularity of solution to the Timoshenko beam equation with local Kelvin–Voigt damping. SIAM Journal on Control and Optimization, 56(6): 3919–3947, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lissy:2019:IOC

[LZ19] Pierre Lissy and Enrique Zuazua. Internal observability for coupled systems of linear partial differential equations. SIAM Journal on Control and Optimization, 57(2):832–853, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Liu:2021:DCC

[LZ21] Shuai Liu and Li Zhang. Distributed consensus and convergence rate analysis of multiagent systems with noises under G-expectation. SIAM Journal on Control and Optimization, 59(4):2624–2644, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:2021:IOS

[LZC21a] Rui Li, Qi Zhang, and Tianguang Chu. Input-output-to-state stability of systems related through simulation relations. SIAM Journal on Control and Optimization, 59(1): 614–634, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:2021:RAB

[LZC21b] Rui Li, Qi Zhang, and Tianguang Chu. Reduction and analysis of Boolean control networks by bisimulation. SIAM Journal on Control and Optimization, 59(2): 1033–1056, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:2022:BPB

[LZC22] Rui Li, Qi Zhang, and Tianguang Chu. Bisimulations of probabilistic Boolean networks. SIAM Journal on Control and Optimization, 60 (5):2631–2657, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1463471.

Lu:2016:CDB

[LZH^{+}16] Jianquan Lu, Jie Zhong, Daniel W. C. Ho, Yang Tang, and Jinde Cao. On controllability of delayed Boolean control networks. SIAM Journal on Control and Optimization, 54(2):475–494, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Li:2002:DMV

[LZL02] Xun Li, Xun Yu Zhou, and Andrew E. B. Lim. Dynamic mean-variance portfolio selection with no-shorting constraints. SIAM Journal on Control and Optimization, 40(5):1540–1555, September 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/37850.

LeGorrec:2005:DSB

[LZM05] Y. Le Gorrec, H. Zwart, and B. Maschke. Dirac structures and boundary control systems associated with skewsymmetric differential operators. SIAM Journal on Control and Optimization, 44(5): 1864–1892, September 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/61167.

Liu:2024:SEE

[LZSC24] Xinhui Liu, Kaikai Zheng, Dawei Shi, and Tongwen Chen. State estimation with event sensors: Observability analysis and multi-sensor fusion. SIAM Journal on Control and Optimization, 62(1): 167–190, January 2024. CO-DEN SJCODC. ISSN 1095-7138.

Li:2020:SPE

[LZT20] Jr-Shin Li, Wei Zhang, and Lin Tie. On separating points for ensemble controllability. SIAM Journal on Control and Optimization, 58(5): 2740–2764, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Lu:2021:SON

[LZZ21] Qi Lü, Haisen Zhang, and Xu Zhang. Second order necessary conditions for optimal control problems of stochastic evolution equations. SIAM Journal on Control and Optimization, 59(4):2924–2954, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ma:1992:PSF

[Ma92] Jin Ma. On the principle of smooth fit for a class of singular stochastic control problems for diffusions. SIAM Journal on Control and Optimization, 30(4):975–999, July 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mori:2001:FSC

[MA01] Kazuyoshi Mori and Kenichi Abe. Feedback stabilization over commutative rings: Further study of the coordinatefree approach. SIAM Journal on Control and Optimization, 39(6):1952–1973, ???? 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Moreau:2003:TBL

[MA03] Luc Moreau and Dirk Aeyels. Trajectory-based local approximations of ordinary differential equations. SIAM Journal on Control and Optimization, 41(6):1922–1945, November 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/37077.

Ma:2009:DAS

[Ma09] Hong-Bin Ma. Decentralized adaptive synchronization of a stochastic discrete-time multiagent dynamic model. SIAM Journal on Control and Optimization, 48(2):859–880, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Montenbruck:2017:SSF

[MA17] Jan Maximilian Montenbruck and Frank Allgöwer. Structured state feedback controllers and optimal submanifold stabilization. SIAM Journal on Control and Optimization, 55(4):2393–2411, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Martin:2019:ASS

[MA19] S. Martin and I. Alvarez. Anticipating shocks in the state space: Characterizing robustness and building increasingly robust evolutions. SIAM Journal on Control and Optimization, 57(1):490–509, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Machtyngier:1994:ECS

[Mac94]  Elaine Machtyngier.  Ex-

act controllability for the Schrödinger equation. SIAM Journal on Control and Optimization, 32(1):24–34, January 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/22314.

Macchelli:2014:PBC

[Mac14] Alessandro Macchelli. Passivitybased control of implicit port-Hamiltonian systems. SIAM Journal on Control and Optimization, 52(4):2422–2448, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

McConley:1998:CCL

[MADF98] Marc W. McConley, Brent D. Appleby, Munther A. Dahleh, and Eric Feron. Computational complexity of Lyapunov stability analysis problems for a class of nonlinear systems. SIAM Journal on Control and Optimization, 36(6):2176–2193, November 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32031.

Mancilla-Aguilar:2011:IPS

[MAG11] J. L. Mancilla-Aguilar and R. A. García. Invariance principles for switched systems with restrictions. SIAM Journal on Control and Optimization, 49(4):1544–1569, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sicon/
resource/1/sjcodc/v49/i4/
p1544_s1.

Mahmudov:2003:ACS

[Mah03] Nazim I. Mahmudov. Approximate controllability of semilinear deterministic and stochastic evolution equations in abstract spaces. SIAM Journal on Control and Optimization, 42(5):1604–1622, September 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39168.

Mailybaev:2003:ULA

[Mai03] Alexei A. Mailybaev. Uncontrollability for linear autonomous multi-input dynamical systems depending on parameters. SIAM Journal on Control and Optimization, 42 (4):1431–1450, January 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mainge:2008:HEV

[Mai08] Paul-Emile Maingé. A hybrid extragradient-viscosity method for monotone operators and fixed point problems. SIAM Journal on Control and Optimization, 47(3): 1499–1515, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mainge:2013:FOC

[Mai13] Paul-Emile Maingé. Firstorder continuous Newton-like systems for monotone inclusions. SIAM Journal on Control and Optimization, 51(2): 1615–1638, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Maury:2018:EPS

[MAJ18] Aymeric Maury, Grégoire Allaire, and François Jouve. Elasto-plastic shape optimization using the level set method. SIAM Journal on Control and Optimization, 56 (1):556–581, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Malanowski:1997:SOC

[Mal97] K. Malanowski. Sufficient optimality conditions for optimal control subject to state constraints. SIAM Journal on Control and Optimization, 35(1):205–227, January 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26763.

Malisoff:2002:VSB

[Mal02] Michael Malisoff. Viscosity solutions of the Bellman equation for exit time optimal control problems with vanishing Lagrangians. SIAM Journal on Control and Optimization, 40(5):1358–1383, September 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36859.

Malisani:2023:IPM

[Mal23] Paul Malisani. Interior point methods in optimal control problems of affine systems: Convergence results and solving algorithms. SIAM Journal on Control and Optimization, 61(6):3390–3414, November 2023. CODEN SJCODC. ISSN 1095-7138.

Mammadov:2014:TTI

[Mam14] Musa A. Mammadov. Turnpike theorem for an infinite horizon optimal control problem with time delay. SIAM Journal on Control and Optimization, 52(1):420–438, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Manitius:1981:NSC

[Man81] A. Manitius. Necessary and sufficient conditions of approximate controllability for general linear retarded systems. SIAM Journal on Control and Optimization, 19(4):516–532, July 1981. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mangasarian:1985:SEP

[Man85] O. L. Mangasarian. Sufficiency of exact penalty minimization. SIAM Journal on Control and Optimization, 23 (1):30–37, January 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mangasarian:1995:PGD

[Man95] O. L. Mangasarian. Parallel gradient distribution in unconstrained optimization. SIAM Journal on Control and Optimization, 33(6): 1916–1925, November 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/25022.

Mannucci:2004:NSS

[Man04] Paola Mannucci. Nonzerosum stochastic differential games with discontinuous feedback. SIAM Journal on Control and Optimization, 43(4):1222–1233, July 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/42371.

Manns:2021:RMR

[Man21] Paul Manns. Relaxed multibang regularization for the combinatorial integral approximation. SIAM Journal on Control and Optimization, 59 (4):2645–2668, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mao:1990:LFA

[Mao90] Xue Rong Mao. Lyapunov functions and almost sure exponential stability of stochastic differential equations based on semimartingales with spatial parameters. SIAM Journal on Control and Optimization, 28(6):1481–1490, November 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Martin:1977:COP

[Mar77] Jean-Claude E. Martin. Controllability and observability of parabolic systems — an addendum to two recent papers of Y. Sakawa. SIAM Journal on Control and Optimization, 15(3):363–366, May 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Marcus:1984:AGM

[Mar84] Steven I. Marcus. Algebraic and geometric methods in nonlinear filtering. SIAM Journal on Control and Optimization, 22(6):817–844, November 1984. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Martinez:1999:BSW

[Mar99] Patrick Martinez. Boundary stabilization of the wave equation in almost star-shaped domains. SIAM Journal on Control and Optimization, 37(3):673–694, May 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32372.

Marcelli:2002:VPN

[Mar02]  Cristina Marcelli.  Varia-

tional problems with nonconvex, noncoercive, highly discontinuous integrands: Characterization and existence of minimizers. SIAM Journal on Control and Optimization, 40(5):1473–1490, September 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/36141.

Marinoschi:2020:ECM

[Mar20]  Gabriela Marinoschi.  Ex-

act controllability in minimal time of the Navier– Stokes periodic flow in a 2dchannel. SIAM Journal on Control and Optimization, 58 (6):3658–3683, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mancilla-Aguilar:2023:CII

[MARRH23] José L. Mancilla-Aguilar, Jose E. Rojas-Ruiz, and Hernan Haimovich. Characterization of integral input-to-state stability for nonlinear time-varying systems of infinite dimension. SIAM Journal on Control and Optimization, 61(4):1979–2003, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1497638.

Masiero:2008:SOC

[Mas08] Federica Masiero. Stochastic optimal control problems and parabolic equations in Banach spaces. SIAM Journal on Control and Optimization, 47(1): 251–300, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Matsumoto:1976:SPM

[Mat76] T. Matsumoto. On some properties of min-max functions. SIAM Journal on Control and Optimization, 14(1): 144–155, January 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Matveev:1998:SAD

[Mat98] Alexey S. Matveev. Spectral approach to duality in nonconvex global optimization. SIAM Journal on Control and Optimization, 36(1):336–378, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27773.

Matveev:2005:IOC

[Mat05] Alexey S. Matveev. The instability of optimal control problems to time delay. SIAM Journal on Control and Optimization, 43(5): 1757–1786, September 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42321.

Maurer:1977:OCP

[Mau77] H. Maurer. On optimal control problems with bounded state variables and control appearing linearly. SIAM Journal on Control and Optimization, 15 (3):345–362, May 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mazza:1994:AFH

[Maz94] Christian Mazza. Asymptotic first hitting-time distribution of annealing processes. SIAM Journal on Control and Optimization, 32(5): 1266–1288, September 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/19717.

Mazanti:2017:RCL

[Maz17] Guilherme Mazanti. Relative controllability of linear difference equations. SIAM Journal on Control and Optimization, 55(5):3132–3153, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mazenc:2004:BBF

[MB04] Frederic Mazenc and Samuel Bowong. Backstepping with bounded feedbacks for timevarying systems. SIAM Journal on Control and Optimization, 43(3):856–871, May 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/40873.

Mirtabatabaei:2012:ODH

[MB12] Anahita Mirtabatabaei and Francesco Bullo. Opinion dynamics in heterogeneous networks: Convergence conjectures and theorems. SIAM Journal on Control and Optimization, 50(5):2763–2785, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mathkar:2016:NG

[MB16] Adwaitvedant S. Mathkar and Vivek S. Borkar. Nonlinear gossip. SIAM Journal on Control and Optimization, 54(3): 1535–1557, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mason:2006:CPL

[MBC06] Paolo Mason, Ugo Boscain, and Yacine Chitour. Common polynomial Lyapunov functions for linear switched systems. SIAM Journal on Control and Optimization, 45(1): 226–245, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Marchand:2000:SOS

[MBL00] Hervé Marchand, Olivier Boivineau, and Stéphane Lafortune. On the synthesis of optimal schedulers in discrete event control problems with multiple goals. SIAM Journal on Control and Optimization, 39(2):512–532, March 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/34196.

Matei:2013:CRL

[MBS13] Ion Matei, John S. Baras, and Christoforos Somarakis. Convergence results for the linear consensus problem under Markovian random graphs. SIAM Journal on Control and Optimization, 51(2):1574– 1591, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Meyn:1991:ABS

[MC91] S. P. Meyn and P. E. Caines. Asymptotic behavior of stochastic systems possessing Markovian realizations. SIAM Journal on Control and Optimization, 29(3):535–561, May 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mu:2012:RIW

[MC12] Bi-Qiang Mu and Han-Fu Chen. Recursive identification of Wiener–Hammerstein systems. SIAM Journal on Control and Optimization, 50(5): 2621–2658, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

McCann:1976:BOL

[McC76] Roger C. McCann. Behavior of observations of linear control systems. SIAM Journal on Control and Optimization, 14 (4):757–762, July 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

McCaffrey:2009:PIM

[McC09] D. McCaffrey. Policy iteration and the max-plus finite element method. SIAM Journal on Control and Optimization, 47(6):2912–2929, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

McEneaney:1995:UVS

[McE95] William M. McEneaney. Uniqueness for viscosity solutions of nonstationary Hamilton– Jacobi–Bellman equations under some a priori conditions (with applications). SIAM Journal on Control and Optimization, 33(5): 1560–1576, September 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26269.

McEneaney:2004:MPE

[McE04] William M. McEneaney. Maxplus eigenvector methods for nonlinear H_{∞} problems: Error analysis. SIAM Journal on Control and Optimization, 43(2):379–412, March 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41468.

McEneaney:2007:CDF

[McE07] William M. McEneaney. A curse-of-dimensionality-free numerical method for solution of certain HJB PDEs. SIAM Journal on Control and Optimization, 46(4):1239–1276, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

McEneaney:2009:CRCa

[McE09] William M. McEneaney. Convergence rate for a curseof-dimensionality-free method for Hamilton–Jacobi–Bellman PDEs represented as maxima of quadratic forms. SIAM Journal on Control and Optimization, 48(4):2651–2685, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ma:2022:DEF

[MCL^{+}22] Dan Ma, Jianqi Chen, Renquan Lu, Jie Chen, and Tianyou Chai. Delay effect on first-order consensus over directed graphs: Optimizing PID protocols for maximal robustness. SIAM Journal on Control and Optimization, 60(1):233–258, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1413171.

McNamara:1984:CDS

[McN84] J. M. McNamara. Control of a diffusion by switching between two drift-diffusion coefficient pairs. SIAM Journal on Control and Optimization, 22 (1):87–94, January 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mason:2021:WED

[MCP21] Paolo Mason, Yacine Chitour, and Dario Prandi. Worst exponential decay rate for degenerate gradient flows subject to persistent excitation. SIAM Journal on Control and Optimization, 59(4):3040–3067, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Marx:2017:GSK

[MCPA17] Swann Marx, Eduardo Cerpa, Christophe Prieur, and Vincent Andrieu. Global stabilization of a Korteweg–de Vries equation with saturating distributed control. SIAM Journal on Control and Optimization, 55(3):1452–1480, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

McShane:1989:CVB

[McS89] E. J. McShane. The calculus of variations from the beginning through optimal control theory. SIAM Journal on Control and Optimization, 27(5):916– 939, September 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mazanti:2013:STD

[MCS13] Guilherme Mazanti, Yacine Chitour, and Mario Sigalotti. Stabilization of twodimensional persistently excited linear control systems with arbitrary rate of convergence. SIAM Journal on Control and Optimization, 51(2): 801–823, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mu:2014:CIM

[MCWY14] Bi-Qiang Mu, Han-Fu Chen, Le Yi Wang, and George Yin. Characterization and identification of matrix fraction descriptions for LTI systems. SIAM Journal on Control and Optimization, 52(6): 3694–3721, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Marcotte:1989:SLP

[MD89] Patrice Marcotte and Jean-Pierre Dussault. A sequential linear programming algorithm for solving monotone variational inequalities. SIAM Journal on Control and Optimization, 27(6):1260– 1278, November 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

McEneaney:2013:PSE

[MD13] William M. McEneaney and Ameet S. Deshpande. Payoff suboptimality and errors in value induced by approximation of the Hamiltonian. SIAM Journal on Control and Optimization, 51(5):3993–4015, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

McEneaney:2015:PLA

[MD15] William M. McEneaney and Peter M. Dower. The principle of least action and fundamental solutions of mass-spring and N-body two-point boundary value problems. SIAM Journal on Control and Optimization, 53(5):2898–2933, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Maithripala:2005:IOB

[MDB05] D. H. S. Maithripala, W. P.

Dayawansa, and J. M. Berg. Intrinsic observer-based stabilization for simple mechanical systems on Lie groups. SIAM Journal on Control and Optimization, 44(5): 1691–1711, September 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43989.

McEneaney:2023:SOH

[MDW23] William M. McEneaney, Peter M. Dower, and Tao Wang. Second-order Hamilton–Jacobi PDE problems and certain related first-order problems, Part 1: Approximation. SIAM Journal on Control and Optimization, 61(6):3280–3315, November 2023. CODEN SJ-CODC. ISSN 1095-7138.

Medvedev:2012:SSC

[Med12] Georgi S. Medvedev. Stochastic stability of continuous time consensus protocols. SIAM Journal on Control and Optimization, 50(4):1859–1885, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Meeker:1989:LTO

[Mee89] L. David Meeker. Local timeoptimal feedback control of strictly normal two-input linear systems. SIAM Journal on Control and Optimization, 27 (1):53–82, January 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Menaldi:1980:OIC

[Men80a] J. L. Menaldi. On the optimal impulse control problem for degenerate diffusions. SIAM Journal on Control and Optimization, 18(6):722– 739, November 1980. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Menaldi:1980:OST

[Men80b] J. L. Menaldi. On the optimal stopping time problem for degenerate diffusions. SIAM Journal on Control and Optimization, 18(6):697– 721, November 1980. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Menaldi:1989:SEF

[Men89] José-Luis Menaldi. Some estimates for finite difference approximations. SIAM Journal on Control and Optimization, 27(3):579–607, May 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Michiels:2002:SID

[MERD02] W. Michiels, K. Engelborghs, D. Roose, and D. Dochain. Sensitivity to infinitesimal delays in neutral equations. SIAM Journal on Control and Optimization, 40(4):1134– 1158, July 2002. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/35507.

Meyer:1977:CCT

[Mey77a] Gerard G. L. Meyer. Convergence conditions for a type of algorithm model. SIAM Journal on Control and Optimization, 15(5):779–784, August 1977. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Meyer:1977:FSC

[Mey77b] Gerard G. L. Meyer. A finitely solvable class of approximating problems. SIAM Journal on Control and Optimization, 15(3):400–406, May 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Meyer:1977:CNI

[Mey77c] R. R. Meyer. A class of nonlinear integer programs solvable by a single linear program. SIAM Journal on Control and Optimization, 15(6): 935–946, November 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Meyer:1977:CFF

[Mey77d] R. R. Meyer. A comparison of the forcing function and point-to-set mapping approaches to convergence analysis. SIAM Journal on Control and Optimization, 15(4): 699–715, July 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Meyer:1983:EMF

[Mey83] Gerard G. L. Meyer. An efficient method of feasible directions. SIAM Journal on Control and Optimization, 21(1): 153–162, January 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Meyn:1989:ETD

[Mey89]  Sean P. Meyn.  Ergodic

theorems for discrete time stochastic systems using a stochastic Lyapunov function. SIAM Journal on Control and Optimization, 27(6):1409– 1439, November 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Meyn:2001:SRM

[Mey01] Sean P. Meyn. Sequencing and routing in multiclass queueing networks Part I: Feedback regulation. SIAM Journal on Control and Optimization, 40(3):741–776, May 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36272.

Meyn:2003:SRM

[Mey03] Sean P. Meyn. Sequencing and routing in multiclass queueing networks Part II: Workload relaxations. SIAM Journal on Control and Optimization, 42(1):178–217, January 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38376.

Meyn:2009:SAO

[Mey09] Sean Meyn. Stability and asymptotic optimality of generalized MaxWeight policies. SIAM Journal on Control and Optimization, 47(6):3259– 3294, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Martin:2013:CTC

[MG13] Samuel Martin and Antoine Girard. Continuous-time consensus under persistent connectivity and slow divergence of reciprocal interaction weights. SIAM Journal on Control and Optimization, 51(3):2568–2584, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Martens:2020:CAA

[MG20] Björn Martens and Matthias Gerdts. Convergence analysis for approximations of optimal control problems subject to higher index differential-algebraic equations and mixed control-state constraints. SIAM Journal on Control and Optimization, 58 (1):1–33, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Martens:2021:CAA

[MG21] Björn Martens and Matthias Gerdts. Convergence analysis for approximations of optimal control problems subject to higher index differentialalgebraic equations and pure state constraints. SIAM Journal on Control and Optimization, 59(3):1903–1926, ???? 2021. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Magron:2019:SAR

[MGHT19] Victor Magron, Pierre-Loic Garoche, Didier Henrion, and Xavier Thirioux. Semidefinite approximations of reachable sets for discrete-time polynomial systems. SIAM Journal on Control and Optimization, 57(4):2799–2820, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Martinez-Gonzalez:2019:WAA

[MGMBN^{+}19] A. Martínez-González, C.-F.

Méndez-Barrios, S.-I. Niculescu, J. Chen, and L. Félix. Weierstrass approach to asymptotic behavior characterization of critical imaginary roots for retarded differential equations. SIAM Journal on Control and Optimization, 57(1):1–22, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Martin:1978:AAG

[MH78] Clyde Martin and Robert Hermann. Applications of algebraic geometry to systems theory: The McMillan degree and Kronecker indices of transfer functions as topological and holomorphic system invariants. SIAM Journal on Control and Optimization, 16(5):743–755, September 1978. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Michel:2011:RHL

[MH11] Anthony N. Michel and Ling Hou. Relaxation of hypotheses in LaSalle–Krasovskii-type invariance results. SIAM Journal on Control and Optimization, 49(4):1383–1403, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sicon/
                                        resource/1/sjcodc/v49/i4/
                                        p1383_s1.

Mironchenko:2015:CLF

[MI15] Andrii Mironchenko and Hiroshi Ito. Construction of Lyapunov functions for interconnected parabolic systems: an iISS approach. SIAM Journal on Control and Optimization, 53(6):3364–3382, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Michel:1980:NCO

[Mic80] P. Michel. Necessary conditions for optimality of elliptic systems with positivity constraints on the state. SIAM Journal on Control and Optimization, 18(1):91–97, January 1980. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Micu:2001:CLB

[Mic01] Sorin Micu. On the controllability of the linearized Benjamin–Bona–Mahony equation. SIAM Journal on Control and Optimization, 39(6): 1677–1696, November 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36249.

Micu:2009:UBC

[Mic09] Sorin Micu. Uniform boundary controllability of a semidiscrete 1-D wave equation with vanishing viscosity. SIAM Journal on Control and Optimization, 47(6):2857–2885, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mifflin:1977:SSF

[Mif77] Robert Mifflin. Semismooth and semiconvex functions in constrained optimization. SIAM Journal on Control and Optimization, 15(6):959– 972, November 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mikkola:2007:CFD

[Mik07] Kalle M. Mikkola. Coprime factorization and dynamic stabilization of transfer functions. SIAM Journal on Control and Optimization, 45(6): 1988–2010, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mikami:2012:CKR

[Mik12a]  Toshio Mikami.  A char-

acterization of the Knothe– Rosenblatt processes by a convergence result. SIAM Journal on Control and Optimization, 50(4):1903–1920, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mikkola:2012:RSC

[Mik12b] Kalle M. Mikkola. Real solutions to control, approximation, and factorization problems. SIAM Journal on Control and Optimization, 50(3): 1071–1086, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mikami:2015:TEP

[Mik15] Toshio Mikami. Two end points marginal problem by stochastic optimal transportation. SIAM Journal on Control and Optimization, 53(4): 2449–2461, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Milman:1988:ALQ

[Mil88] Mark H. Milman. Approximating the linear quadratic optimal control law for hereditary systems with delays in the control. SIAM Journal on Control and Optimization, 26 (2):291–320, March 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Miller:1996:GSN

[Mil96] Boris M. Miller. The generalized solutions of nonlinear optimization problems with impulse control. SIAM Journal on Control and Optimization, 34(4):1420–1440, July 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26321.

Miller:2002:EFC

[Mil02] Luc Miller. Escape function conditions for the observation, control, and stabilization of the wave equation. SIAM Journal on Control and Optimization, 41(5): 1554–1566, September 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39107.

Miller:2006:CTM

[Mil06] Luc Miller. The control transmutation method and the cost of fast controls. SIAM Journal on Control and Optimization, 45(2):762–772, January 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Miller:2017:NPA

[Mil17] Christopher W. Miller. Nonlinear PDE approach to timeinconsistent optimal stopping. SIAM Journal on Control and Optimization, 55(1):557–573, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mirica:1986:SOC

[Mir86]  Ştefan Mirică. Sufficient op-

timality conditions for stratified control problems. SIAM Journal on Control and Optimization, 24(4):675–691, July 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mironchenko:2021:SGT

[Mir21] Andrii Mironchenko. Small gain theorems for general networks of heterogeneous infinite-dimensional systems. SIAM Journal on Control and Optimization, 59(2):1393– 1419, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Misztela:2019:RHJ

[Mis19] Arkadiusz Misztela. Representation of Hamilton–Jacobi equation in optimal control theory with compact control set. SIAM Journal on Control and Optimization, 57(1): 53–77, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Martins:1990:RSC

[MK90] Luiz Felipe Martins and Harold J. Kushner. Routing and singular control for queueing networks in heavy traffic. SIAM Journal on Control and Optimization, 28(5):1209– 1233, September 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

McEneaney:2010:CRC

[MK10] William M. McEneaney and L. Jonathan Kluberg. Convergence rate for a curseof-dimensionality-free method for a class of HJB PDEs. SIAM Journal on Control and Optimization, 48(5):3052– 3079, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Masubuchi:2021:LDC

[MK21] Izumi Masubuchi and Takahiro Kikuchi. Lyapunov density criteria for time-varying and periodically time-varying nonlinear systems with converse results. SIAM Journal on Control and Optimization, 59 (1):223–241, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mironchenko:2019:MMI

[MKK19] Andrii Mironchenko, Iasson Karafyllis, and Miroslav Krstic. Monotonicity methods for input-to-state stability of nonlinear parabolic PDEs with boundary disturbances. SIAM Journal on Control and Optimization, 57(1):510–532, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Michelbrink:2012:MAO

[ML12] Daniel Michelbrink and Huiling Le. A martingale approach to optimal portfolios with jump-diffusions. SIAM Journal on Control and Optimization, 50(1):583–599, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ma:2013:RPI

[ML13] Huanfei Ma and Wei Lin. Realization of parameters identification in only locally Lipschitzian dynamical systems with multiple types of time delays. SIAM Journal on Control and Optimization, 51(5): 3692–3721, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Munoz-Lecanda:2002:DCM

[MLYF02] Miguel C. Muñoz-Lecanda and F. Javier Yániz-Fernández. Dissipative control of mechanical systems: a geometric approach. SIAM Journal on Control and Optimization, 40(5):1505–1516, September 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37480.

Meng:2017:FLQ

[MLZ17] Yang Meng, Tao Li, and Ji-Feng Zhang. Finite-level quantized synchronization of discrete-time linear multiagent systems with switching topologies. SIAM Journal on Control and Optimization, 55(1): 275–299, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mangasarian:1979:NPL

[MM79] O. L. Mangasarian and R. R.

Meyer. Nonlinear perturbation of linear programs. SIAM Journal on Control and Optimization, 17(6):745– 752, November 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Miller:1980:ASS

[MM80a] R. K. Miller and A. N. Michel.

Asymptotic stability of systems: Results involving the system topology. SIAM Journal on Control and Optimization, 18(2):181–190, March 1980. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Miller:1980:EPM

[MM80b] R. K. Miller and A. N.

Michel. On existence of periodic motions in nonlinear control systems with periodic inputs. SIAM Journal on Control and Optimization, 18 (5):585–598, September 1980. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mandal:2000:BFG

[MM00] Pranab K. Mandal and V. Mandrekar. A Bayes formula for Gaussian noise processes and its applications. SIAM Journal on Control and Optimization, 39(3):852–871, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34338.

Macchelli:2004:MCT

[MM04] Alessandro Macchelli and Claudio Melchiorri. Modeling and control of the Timoshenko beam. the distributed port Hamiltonian approach. SIAM Journal on Control and Optimization, 43(2):743–767, March 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42953.

Monovich:2011:ADT

[MM11] Tal Monovich and Michael Margaliot. Analysis of discrete-time linear switched systems: a variational approach. SIAM Journal on Control and Optimization, 49 (2):808–829, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i2/p808_s1.

Mordecki:2016:OSL

[MM16] Ernesto Mordecki and Yuliya Mishura. Optimal stopping for Lévy processes with onesided solutions. SIAM Journal on Control and Optimization, 54(5):2553–2567, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Martyr:2021:NSG

[MM21] Randall Martyr and John Moriarty. Nonzero-sum games of optimal stopping and generalized Nash equilibrium problems. SIAM Journal on Control and Optimization, 59(2): 1443–1465, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mercado:2023:ECS

[MM23] Alberto Mercado and Roberto Morales. Exact controllability for a Schrödinger equation with dynamic boundary conditions. SIAM Journal on Control and Optimization, 61(6): 3501–3525, November 2023. CODEN SJCODC. ISSN 1095-7138.

Meng:2019:DHO

[MMH19] Deyuan Meng, Ziyang Meng, and Yiguang Hong. Disagreement of hierarchical opinion dynamics with changing antagonisms. SIAM Journal on Control and Optimization, 57 (1):718–742, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Miller:1986:LCF

[MMK86] R. K. Miller, A. N. Michel, and G. S. Krenz. On limit cycles of feedback systems which contain a hysteresis nonlinearity. SIAM Journal on Control and Optimization, 24(2): 276–305, March 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mehandiratta:2021:OCP

[MML21] Vaibhav Mehandiratta, Mani Mehra, and Gunter Leugering. Optimal control problems driven by time-fractional diffusion equations on metric graphs: Optimality system and finite difference approximation. SIAM Journal on Control and Optimization, 59(6):4216–4242, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Michiels:2009:CPD

[MMN09] Wim Michiels, Constantin-Irinel Morārescu, and Silviu-Iulian Niculescu. Consensus problems with distributed delays, with application to traffic flow models. SIAM Journal on Control and Optimization, 48(1):77–101, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mazenc:2017:SCD

[MMN17] Frederic Mazenc, Michael Malisoff, and Silviu-Iulian Niculescu. Stability and control design for time-varying systems with time-varying delays using a trajectory-based approach. SIAM Journal on Control and Optimization, 55 (1):533–556, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mazenc:2018:SRA

[MMÖ18] Frédéric Mazenc, Michael Malisoff, and Hitay Özbay. Stability and robustness analysis for switched systems with time-varying delays. SIAM Journal on Control and Optimization, 56(1):158–182, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Meinlschmidt:2017:OCTb

[MMR17a] H. Meinlschmidt, C. Meyer, and J. Rehberg. Optimal control of the thermistor problem in three spatial dimensions, Part 1: Existence of optimal solutions. SIAM Journal on Control and Optimization, 55(5):2876–2904, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Meinlschmidt:2017:OCTa

[MMR17b] H. Meinlschmidt, C. Meyer, and J. Rehberg. Optimal control of the thermistor problem in three spatial dimensions, Part 2: Optimality conditions. SIAM Journal on Control and Optimization, 55(4): 2368–2392, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Morgan:1977:SND

[MN77a] A. P. Morgan and K. S. Narendra. On the stability of nonautonomous differential equations ẋ = [A + B(t)]x, with skew symmetric matrix B(t). SIAM Journal on Control and Optimization, 15(1): 163–176, January 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Morgan:1977:UAS

[MN77b] A. P. Morgan and K. S. Narendra. On the uniform asymptotic stability of certain linear nonautonomous differential equations. SIAM Journal on Control and Optimization, 15(1):5–24, January 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mitter:2003:VAN

[MN03] Sanjoy K. Mitter and Nigel J. Newton. A variational approach to nonlinear estimation. SIAM Journal on Control and Optimization, 42(5): 1813–1833, September 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39389.

Michiels:2007:CDI

[MN07] Wim Michiels and Silviu-Iulian Niculescu. Characterization of delay-independent stability and delay interference phenomena. SIAM Journal on Control and Optimization, 45(6):2138–2155, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Musso:2008:RCV

[MN08] Emilio Musso and Lorenzo Nicolodi. Reduction for constrained variational problems on 3-dimensional null curves. SIAM Journal on Control and Optimization, 47(3):1399– 1414, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Micu:2022:FDE

[MN22] Sorin Micu and Constantin Nita. On a fractional diffusion equation with moving control. SIAM Journal on Control and Optimization, 60(2):871–889, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1385056.

Mazenc:2011:BNS

[MNB11] Frédéric Mazenc, Silviu-Iulian Niculescu, and Mounir Bekaik. Backstepping for nonlinear systems with delay in the input revisited. SIAM Journal on Control and Optimization, 49(6):2263–2278, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sicon/
resource/1/sjcodc/v49/i6/
p2263_s1.

Mazenc:2012:ESI

[MNB12] Frédéric Mazenc, Silviu-Iulian Niculescu, and Olivier Bernard. Exponentially stable interval observers for linear systems with delay. SIAM Journal on Control and Optimization, 50(1):286– 305, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v50/i1/p286_s1.

Mateos-Nunez:2014:TMN

[MNC14] David Mateos-Núñez and Jorge Cortés. p th Moment noise-to-state stability of stochastic differential equations with persistent noise. SIAM Journal on Control and Optimization, 52(4):2399– 2421, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mateos-Nunez:2016:NSE

[MNC16] David Mateos-Núñez and Jorge Cortés. Noise-to-state exponentially stable distributed convex optimization on weight-balanced digraphs. SIAM Journal on Control and Optimization, 54(1):266–290, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Monti:2024:FOL

[MNMS24] Andrea Monti, Benita Nortmann, Thulasi Mylvaganam, and Mario Sassano. Feedback and open-loop Nash equilibria for LQ infinite-horizon discrete-time dynamic games. SIAM Journal on Control and Optimization, 62(3):1417– 1436, June 2024. CODEN SJ-CODC. ISSN 1095-7138.

Magoules:2021:OAA

[MNORP21] Frédéric Magoulès, Thi Phuong Kieu Nguyen, Pascal Omnes, and Anna Rozanova-Pierrat. Optimal absorption of acoustic waves by a boundary. SIAM Journal on Control and Optimization, 59(1):561–583, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Matusik:2020:FTS

[MNPR20] Radosław Matusik, Andrzej Nowakowski, Sławomir Plaskacz, and Andrzej Rogowski. Finitetime stability for differential inclusions with applications to neural networks. SIAM Journal on Control and Optimization, 58(5):2854–2870, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mouallif:1991:PPD

[MNS91] Khalil Mouallif, Van Hien Nguyen, and Jean-Jacques Strodiot. A perturbed parallel decomposition method for a class of nonsmooth convex minimization problems. SIAM Journal on Control and Optimization, 29(4):829–847, July 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Morimoto:1999:SRB

[MO99] Hiroaki Morimoto and Masami Okada. Some results on the Bellman equation of ergodic control. SIAM Journal on Control and Optimization, 38(1):159–174, January 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32541.

Maurer:2002:SOS

[MO02] Helmut Maurer and Hans Joachim Oberle. Second order sufficient conditions for optimal control problems with free final time: The Riccati approach. SIAM Journal on Control and Optimization, 41(2):380–403, March 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37741.

Maurer:2004:SOS

[MO04] Helmut Maurer and Nikolai P. Osmolovskii. Second order sufficient conditions for time-optimal bang–bang control. SIAM Journal on Control and Optimization, 42(6): 2239–2263, November 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40257.

Morris:2014:MSV

[MÖ14] K. A. Morris and A. Ö. Özer.

Modeling and stabilizability of voltage-actuated piezoelectric beams with magnetic effects. SIAM Journal on Control and Optimization, 52(4): 2371–2398, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mine:1977:CIP

[MOF77] Hisashi Mine, Katsuhisa Ohno, and Masao Fukushima. A “Conjugate” interior penalty method for certain convex programs. SIAM Journal on Control and Optimization, 15(5): 747–755, August 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Molina:2019:LEB

[Mol19] Nicolás Molina. Local exact boundary controllability for the compressible Navier– Stokes equations. SIAM Journal on Control and Optimization, 57(3):2152–2184, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Montgomery:1994:AM

[Mon94] Richard Montgomery. Abnormal minimizers. SIAM Journal on Control and Optimization, 32(6):1605–1620, November 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/24494.

Moon:2021:LQS

[Moo21] Jun Moon. Linear-quadratic stochastic Stackelberg differential games for jumpdiffusion systems. SIAM Journal on Control and Optimization, 59(2):954–976, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Morgan:1979:CNS

[Mor79] A. P. Morgan. On the construction of nonautonomous stable systems with applications to adaptive identification and control. SIAM Journal on Control and Optimization, 17 (3):400–431, May 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mordukhovich:1995:DAR

[Mor95] Boris S. Mordukhovich. Discrete approximations and refined Euler–Lagrange conditions for nonconvex differential inclusions. SIAM Journal on Control and Optimization, 33(3):882–915, May 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24566.

Moreles:1998:CAU

[Mor98] Miguel Angel Moreles. A classical approach to uniform null controllability for elastic beams. SIAM Journal on Control and Optimization, 36(3):1073–1085, May 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30488.

Morimoto:2003:VIC

[Mor03] Hiroaki Morimoto. Variational inequalities for combined control and stopping. SIAM Journal on Control and Optimization, 42(2):686–708, March 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36904.

Morimoto:2010:SCP

[Mor10] Hiroaki Morimoto. A singular control problem with discretionary stopping for geometric Brownian motions. SIAM Journal on Control and Optimization, 48(6):3781–3804, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Moreau:2011:STP

[Mor11] Ludovic Moreau. Stochastic target problems with controlled loss in jump diffusion models. SIAM Journal on Control and Optimization, 49(6):2577–2607, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sicon/
                                        resource/1/sjcodc/v49/i6/
                                        p2577_s1.

Morimoto:2013:OHS

[Mor13] Hiroaki Morimoto. Optimal harvesting in the stochastic logistic growth model with finite horizon. SIAM Journal on Control and Optimization, 51(3):2335–2355, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Morris:2024:ILS

[Mor24a] Ian D. Morris. An irreducible linear switching system whose unique Barabanov norm is not strictly convex. SIAM Journal on Control and Optimization, 62(1):42–64, January 2024. CODEN SJCODC. ISSN 1095-7138.

Morris:2024:SDD

[Mor24b] Ian D. Morris. A stability dichotomy for discrete-time linear switching systems in dimension two. SIAM Journal on Control and Optimization, 62(1):400–414, January 2024. CODEN SJCODC. ISSN 1095-7138.

Mossaheb:1980:HSA

[Mos80] S. Mossaheb. Hardy spaces of analytic functions and feedback stability of scalar convolution systems. SIAM Journal on Control and Optimization, 18(5):498–510, September 1980. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mossaheb:1982:ACC

[Mos82a] S. Mossaheb. Addendum:

“The circle criterion and the L^{p} stability of feedback systems” [SIAM J. Control Optim. 20 (1982), no. 1, 144–152; MR 83a:93042]. SIAM Journal on Control and Optimization, 20(4):598–601, July 1982. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mossaheb:1982:CCS

[Mos82b] S. Mossaheb. The circle criterion and the L^{p} stability of feedback systems. SIAM Journal on Control and Optimization, 20(1):144–152, January 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Motta:1995:NOC

[Mot95]  Monica Motta.  On non-

linear optimal control problems with state constraints. SIAM Journal on Control and Optimization, 33(5): 1411–1424, September 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/24744.

Mottelet:2000:CSC

[Mot00] Stéphane Mottelet. Controllability and stabilization of a canal with wave generators. SIAM Journal on Control and Optimization, 38(3):711–735, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34713.

Maschler:1976:SSS

[MP76a] Michael Maschler and Bezalel Peleg. Stable sets and stable points of set-valued dynamic systems with applications to game theory. SIAM Journal on Control and Optimization, 14(6):985–995, November 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mereau:1976:DSC

[MP76b] P. M. Mereau and W. F. Powers. A direct sufficient condition for free final time optimal control problems. SIAM Journal on Control and Optimization, 14(4):613–622, July 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mignot:1984:OCS

[MP84] F. Mignot and J. P. Puel.

Optimal control in some variational inequalities. SIAM Journal on Control and Optimization, 22(3):466–476, May 1984. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Maurer:1994:SDN

[MP94] Helmut Maurer and Hans Josef Pesch. Solution differentiability for nonlinear parametric control problems. SIAM Journal on Control and Optimization, 32(6):1542–1554, November 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23226.

Makila:1999:SOI

[MP99] P. M. Mäkilä and J. R. Partington. Shift operator induced approximations of delay systems. SIAM Journal on Control and Optimization, 37(6):1897–1912, November 1999. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33967.

Munoz:2007:ROD

[MP07] Julio Muñoz and Pablo Pedregal. A review of an optimal design problem for a plate of variable thickness. SIAM Journal on Control and Optimization, 46(1):1–13, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Micheletti:2011:EAM

[MP11a] Stefano Micheletti and Simona Perotto. The effect of anisotropic mesh adaptation on PDE-constrained optimal control problems. SIAM Journal on Control and Optimization, 49(4):1793– 1828, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i4/p1793_s1.

Mokkadem:2011:GAP

[MP11b] Abdelkader Mokkadem and Mariane Pelletier. A generalization of the averaging procedure: The use of two-time-scale algorithms. SIAM Journal on Control and Optimization, 49(4):1523– 1543, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.

siam.org/sicon/resource/
1/sjcodc/v49/i4/p1523_s1.

Maalouf:2014:FHC

[MP14] Aline I. Maalouf and Ian R.

Petersen. Finite horizon H^{∞} control for a class of linear quantum measurement delayed systems: a dynamic game approach. SIAM Journal on Control and Optimization, 52(3):1787–1808, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mukherjee:2021:MID

[MP21] Mousumi Mukherjee and Debasattam Pal. On minimality of initial data required to uniquely characterize every trajectory in a discrete n-D system. SIAM Journal on Control and Optimization, 59(2):1520–1554, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Moog:1991:MMF

[MPC91] C. H. Moog, A. M. Perdon, and G. Conte. Model matching and factorization for nonlinear systems: a structural approach. SIAM Journal on Control and Optimization, 29(4):769–785, July 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See comments [Kot94].

Marconi:2007:OSN

[MPI07] L. Marconi, L. Praly, and A. Isidori. Output stabilization via nonlinear Luenberger observers. SIAM Journal on Control and Optimization, 45 (6):2277–2298, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Makila:1998:BPS

[MPN98] P. M. Mäkilä, J. R. Partington, and T. Norlander. Bounded power signal spaces for robust control and modeling. SIAM Journal on Control and Optimization, 37(1):92–117, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/31666.

Monache:2017:TRC

[MPR17] Maria Laura Delle Monache, Benedetto Piccoli, and Francesco Rossi. Traffic regulation via controlled speed limit. SIAM Journal on Control and Optimization, 55(5):2936–2958, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Motta:2022:UCI

[MPR22] Monica Motta, Michele Palladino, and Franco Rampazzo. Unbounded control, infimum gaps, and higher order normality. SIAM Journal on Control and Optimization, 60 (3):1436–1462, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1431692.

Morin:1999:DHT

[MPS99] Pascal Morin, Jean-Baptiste Pomet, and Claude Samson. Design of homogeneous time-varying stabilizing control laws for driftless controllable systems via oscillatory approximation of Lie brackets in closed loop. SIAM Journal on Control and Optimization, 38(1):22–49, January 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31542.

Meyer:2006:OCS

[MPT06] C. Meyer, P. Philip, and F. Tröltzsch. Optimal control of a semilinear PDE with nonlocal radiation interface conditions. SIAM Journal on Control and Optimization, 45(2): 699–721, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Monti:2018:TCL

[MPV18] Roberto Monti, Alessandro Pigati, and Davide Vittone. On tangent cones to length minimizers in Carnot– Carathéodory spaces. SIAM Journal on Control and Optimization, 56(5):3351–3369, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Marro:2003:NCA

[MPZ03] Giovanni Marro, Domenico Prattichizzo, and Elena Zattoni. A nested computational approach to the discretetime finite-horizon LQ control problem. SIAM Journal on Control and Optimization, 42(3):1002–1012, May 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/38442.

Mathis:1978:DAC

[MR78] F. H. Mathis and G. W. Reddien. Difference approximations to control problems with functional arguments. SIAM Journal on Control and Optimization, 16(3):436–449, May 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mathis:1979:RTA

[MR79] F. H. Mathis and G. W. Reddien. Ritz–Trefftz approximations in optimal control. SIAM Journal on Control and Optimization, 17(2): 307–310, March 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Minton:1986:RSA

[MR86] R. B. Minton and J. A.

Reneke. An RKH space approach to state feedback control for a class of linear stochastic systems. SIAM Journal on Control and Optimization, 24(4):700–714, July 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Menaldi:1990:ORF

[MR90] José Luis Menaldi and Maurice Robin. On the optimal reward function of the continuous time multiarmed bandit problem. SIAM Journal on Control and Optimization, 28 (1):97–112, January 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Motta:1996:DPN

[MR96] Monica Motta and Franco Rampazzo. Dynamic programming for nonlinear systems driven by ordinary and impulsive controls. SIAM Journal on Control and Optimization, 34(1):199–225, January 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25493.

Miller:1997:OOS

[MR97] Boris M. Miller and Wolfgang J. Runggaldier. Optimization of observations: a stochastic control approach. SIAM Journal on Control and Optimization, 35(3):1030–1052, May 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28787.

Meyer:2004:SPO

[MR04a] C. Meyer and A. Rösch. Superconvergence properties of optimal control problems. SIAM Journal on Control and Optimization, 43(3):970–985, May 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43160.

Mordukhovich:2004:NBC

[MR04b] Boris S. Mordukhovich and Jean-Pierre Raymond. Neumann boundary control of hyperbolic equations with pointwise state constraints. SIAM Journal on Control and Optimization, 43(4):1354–1372, July 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43117.

Meyer:2005:EAO

[MR05] C. Meyer and A. Rösch. L^{∞}-estimates for approximated optimal control problems. SIAM Journal on Control and Optimization, 44(5): 1636–1649, September 2005.

CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/61462.

Mirrahimi:2009:RTS

[MR09] Mazyar Mirrahimi and Pierre Rouchon. Real-time synchronization feedbacks for singleatom frequency standards. SIAM Journal on Control and Optimization, 48(4):2820– 2839, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Marigonda:2015:CSN

[MR15] Antonio Marigonda and Silvia Rigo. Controllability of some nonlinear systems with drift via generalized curvature properties. SIAM Journal on Control and Optimization, 53 (1):434–474, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Menaldi:2016:SOS

[MR16] J. L. Menaldi and M. Robin.

On some optimal stopping problems with constraint. SIAM Journal on Control and Optimization, 54(5):2650– 2671, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Menaldi:2017:SIC

[MR17] J. L. Menaldi and M. Robin. On some impulse control problems with constraint. SIAM Journal on Control and Optimization, 55(5):3204–3225, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Menaldi:2018:SEI

[MR18a] J. L. Menaldi and M. Robin. On some ergodic impulse control problems with constraint. SIAM Journal on Control and Optimization, 56(4):2690– 2711, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Motta:2018:ACL

[MR18b] Monica Motta and Franco Rampazzo. Asymptotic controllability and Lyapunov-like functions determined by Lie brackets. SIAM Journal on Control and Optimization, 56 (2):1508–1534, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Maggiore:2012:IKS

[MRL12] Manfredi Maggiore, Barry Rawn, and Peter Lehn. Invariance kernels of singleinput planar nonlinear systems. SIAM Journal on Control and Optimization, 50(2): 1012–1037, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mirkin:1999:DSDa

[MRP99a] Leonid Mirkin, Héctor P. Rotstein, and Zalman J. Palmor. H^{2} and H^{∞} design of sampled-data systems using lifting. Part I: General framework and solutions. SIAM Journal on Control and Optimization, 38(1):175–196, January 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32960.

Mirkin:1999:DSDb

[MRP99b] Leonid Mirkin, Héctor P. Rotstein, and Zalman J. Palmor. H^{2} and H^{∞} design of sampled-data systems using lifting. Part II: Properties of systems in the lifted domain. SIAM Journal on Control and Optimization, 38 (1):197–218, January 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32961.

Martin:2013:NCS

[MRR13] Philippe Martin, Lionel Rosier, and Pierre Rouchon. Null controllability of the structurally damped wave equation with moving control. SIAM Journal on Control and Optimization, 51(1):660–684, ???? 2013. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Martin:2016:NCO

[MRR16] Philippe Martin, Lionel Rosier, and Pierre Rouchon. Null controllability of one-dimensional parabolic equations by the flatness approach. SIAM Journal on Control and Optimization, 54(1):198–220, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Martin:2019:ECL

[MRRR19] Philippe Martin, Ivonne Rivas, Lionel Rosier, and Pierre Rouchon. Exact controllability of a linear Korteweg– de Vries equation by the flatness approach. SIAM Journal on Control and Optimization, 57(4):2467–2486, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Malisoff:2004:GAC

[MRS04] Michael Malisoff, Ludovic Rifford, and Eduardo Sontag. Global asymptotic controllability implies input-to-state stabilization. SIAM Journal on Control and Optimization, 42(6):2221–2238, November 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/42233.

Marino:1994:ENS

[MRv94] R. Marino, W. Respondek, and A. J. van der Schaft. Equivalence of nonlinear systems to input-output prime forms. SIAM Journal on Control and Optimization, 32(2):387–407, March 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/20088.

Martinez:2003:RNC

[MRV03] P. Martinez, J. P. Raymond, and J. Vancostenoble. Regional null controllability of a linearized croccotype equation. SIAM Journal on Control and Optimization, 42(2):709–728, March 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40354.

Meidner:2011:PEE

[MRV11] Dominik Meidner, Rolf Rannacher, and Boris Vexler. A priori error estimates for finite element discretizations of parabolic optimization problems with pointwise state constraints in time. SIAM Journal on Control and Optimization, 49(5):1961–1997, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sicon/

                                        resource/1/sjcodc/v49/i5/
                                        p1961_s1.

May:2013:EAF

[MRV13] S. May, R. Rannacher, and B. Vexler. Error analysis for a finite element approximation of elliptic Dirichlet boundary control problems. SIAM Journal on Control and Optimization, 51(3):2585–2611, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Martinez:2000:TNA

[MRVM00] A. Martínez, C. Rodríguez, and M. E. Vázquez-Méndez. Theoretical and numerical analysis of an optimal control problem related to wastewater treatment. SIAM Journal on Control and Optimization, 38(5):1534–1553, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/34564.

Myslinski:1985:NOP

[MS85] Andrzej Myśliński and Jan Sokołowski. Nondifferentiable optimization problems for elliptic systems. SIAM Journal on Control and Optimization, 23(4):632–648, July 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mangasarian:1987:LCS

[MS87] O. L. Mangasarian and T.-H.

Shiau. Lipschitz continuity of solutions of linear inequalities, programs and complementarity problems. SIAM Journal on Control and Optimization, 25(3):583–595, May 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Milman:1990:FSO

[MS90] Mark H. Milman and Robert E.

Scheid. Factorization and the synthesis of optimal feedback gains for distributed parameter systems. SIAM Journal on Control and Optimization, 28(5):1061–1102, September 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Makowski:1992:SAA

[MS92] Armand M. Makowski and Adam Shwartz. Stochastic approximations and adaptive control of a discrete-time single-server network with random routing. SIAM Journal on Control and Optimization, 30(6):1476–1506, November 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Moussaoui:1996:EMP

[MS96] Mohammed Moussaoui and Alberto Seeger. Epsilonmaximum principle of Pontryagin type and perturbation analysis of convex optimal control problems. SIAM Journal on Control and Optimization, 34(2):407–427, March 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24487.

Mizel:1997:ABB

[MS97a] Victor J. Mizel and Thomas I. Seidman. An abstract bang– bang principle and timeoptimal boundary control of the heat equation. SIAM Journal on Control and Optimization, 35(4):1204–1216, July 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26547.

Mordukhovich:1997:SSV

[MS97b] Boris S. Mordukhovich and Yongheng Shao. Stability of set-valued mappings in infinite dimensions: Point criteria and applications. SIAM Journal on Control and Optimization, 35(1):285–314, January 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27817.

Mendiondo:1998:AID

[MS98] Marta Susana Mendiondo and Richard H. Stockbridge. Approximation of infinitedimensional linear programming problems which arise in stochastic control. SIAM Journal on Control and Optimization, 36(4):1448–1472, July 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31336.

Morin:2002:CLA

[MS02] Pascal Morin and Claude Samson. A characterization of the Lie algebra rank condition by transverse periodic functions. SIAM Journal on Control and Optimization, 40(4):1227–1249, July 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36605.

Motta:2003:MTB

[MS03] Monica Motta and Caterina Sartori. Minimum time with bounded energy, minimum energy with bounded time. SIAM Journal on Control and Optimization, 42(3):789–809, May 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38528.

Mordukhovich:2004:AMP

[MS04a] Boris S. Mordukhovich and Ilya Shvartsman. The approximate maximum principle in constrained optimal control. SIAM Journal on Control and Optimization, 43(3):1037–1062, May 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/43301.

Morin:2004:PAS

[MS04b] Pascal Morin and Claude Samson. Practical and asymptotic stabilization of chained systems by the transverse function control approach. SIAM Journal on Control and Optimization, 43(1):32–57, January 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/42186.

Matveev:2005:MSL

[MS05] Alexey S. Matveev and Andrey V. Savkin. Multirate stabilization of linear multiple sensor systems via limited capacity communication channels. SIAM Journal on Control and Optimization, 44(2):584–617, March 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/41996.

Matveev:2007:ASI

[MS07a] Alexey S. Matveev and Andrey V. Savkin. An analogue of Shannon information theory for detection and stabilization via noisy discrete communication channels. SIAM Journal on Control and Optimization, 46(4):1323–1367, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Motta:2007:FFP

[MS07b] Monica Motta and Caterina Sartori. Finite fuel problem in nonlinear singular stochastic control. SIAM Journal on Control and Optimization, 46(4):1180–1210, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Motta:2010:WCP

[MS10] M. Motta and C. Sartori.

Weakly coercive problems in nonlinear stochastic control: Existence of optimal controls. SIAM Journal on Control and Optimization, 48(5): 3532–3561, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Meng:2011:OMI

[MS11] Hui Meng and Tak Kuen Siu. Optimal mixed impulseequity insurance control problem with reinsurance. SIAM Journal on Control and Optimization, 49(1):254–279, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sicon/
resource/1/sjcodc/v49/i1/
p254_s1.

Meyer:2017:OCN

[MS17a] Christian Meyer and Livia M. Susu. Optimal control of nonsmooth, semilinear parabolic equations. SIAM Journal on Control and Optimization, 55 (4):2206–2234, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Motee:2017:SSL

[MS17b] Nader Motee and Qiyu Sun. Sparsity and spatial localization measures for spatially distributed systems. SIAM Journal on Control and Optimization, 55(1):200–235, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mostajeran:2018:PMC

[MS18] Cyrus Mostajeran and Rodolphe Sepulchre. Positivity, monotonicity, and consensus on Lie groups. SIAM Journal on Control and Optimization, 56 (3):2436–2461, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Motee:2022:LSC

[MS22] Nader Motee and Qiyu Sun. Localized stability certificates for spatially distributed systems over sparse proximity graphs. SIAM Journal on Control and Optimization, 60(2):917–944, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/19M1298834.

Manns:2023:IOC

[MS23] Paul Manns and Annika Schiemann. On integer optimal control with total variation regularization on multidimensional domains. SIAM Journal on Control and Optimization, 61(6):3415–3441, November 2023. CODEN SJCODC. ISSN 1095-7138.

Mason:2008:LTO

[MSBC08] P. Mason, R. Salmoni, U. Boscain, and Y. Chitour. Limit time optimal synthesis for a controlaffine system on S^{2}. SIAM Journal on Control and Optimization, 47(1):111–143, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Matsumoto:1991:DFS

[MSH91] Toshihiro Matsumoto, Susumu Shindoh, and Ryuichi Hirabayashi. 1-determinacy of feasible sets. SIAM Journal on Control and Optimization, 29(2):339– 346, March 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Meng:2015:MSC

[MSJ15] Ziyang Meng, Guodong Shi, and Karl Henrik Johansson. Multiagent systems with compasses. SIAM Journal on Control and Optimization, 53(5):

3057–3080, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Michiels:2001:SPD

[MSR01] W. Michiels, R. Sepulchre, and D. Roose. Stability of perturbed delay differential equations and stabilization of nonlinear cascade systems. SIAM Journal on Control and Optimization, 40(3):661–680, May 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36504.

Martins:1996:HTC

[MSS96] L. F. Martins, S. E. Shreve, and H. M. Soner. Heavy traffic convergence of a controlled, multiclass queueing system. SIAM Journal on Control and Optimization, 34(6):2133–2171, November 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26588.

Mazziotto:1988:ICP

[MSSZ88] G. Mazziotto, Ł. Stettner, J. Szpirglas, and J. Zabczyk. On impulse control with partial observation. SIAM Journal on Control and Optimization, 26(4):964–984, July 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Meyer:2016:OCI

[MST16] C. Meyer, S. M. Schnepp, and O. Thoma. Optimal control of the inhomogeneous relativistic Maxwell–Newton–Lorentz equations. SIAM Journal on Control and Optimization, 54 (5):2490–2525, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mohammed:1986:HSS

[MSv86] S. Mohammed, M. Scheutzow, and H. von Weizsäcker. Hyperbolic state space decomposition for a linear stochastic delay equation. SIAM Journal on Control and Optimization, 24(3):543–551, May 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Minjarez-Sosa:2009:AOS

[MSVA09] J. Adolfo Minjárez-Sosa and Oscar Vega-Amaya. Asymptotically optimal strategies for adaptive zero-sum discounted Markov games. SIAM Journal on Control and Optimization, 48(3):1405–1421, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ma:2018:KBE

[MSZ18] Jin Ma, Rentao Sun, and Yonghui Zhou. Kyle–Back equilibrium models and linear conditional mean-field SDEs. SIAM Journal on Control and Optimization, 56(2):1154– 1180, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Manitius:1978:FSC

[MT78] A. Manitius and R. Triggiani. Function space controllability of linear retarded systems: a derivation from abstract operator conditions. SIAM Journal on Control and Optimization, 16(4):599–645, July 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Manitius:1994:NAH

[MT94] A. Manitius and H. T. Tran. Numerical approximations for hereditary systems with input and output delays: Convergence results and convergence rates. SIAM Journal on Control and Optimization, 32(5):1332–1363, September 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/16169.

Mania:2003:SBE

[MT03] Michael Mania and Revaz Tevzadze. A semimartingale backward equation and the variance-optimal martingale measure under general information flow. SIAM Journal on Control and Optimization, 42(5):1703–1726, September 2003. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40628.

Mikami:2008:OTP

[MT08] Toshio Mikami and Michèle Thieullen. Optimal transportation problem by stochastic optimal control. SIAM Journal on Control and Optimization, 47(3):1127–1139, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mahajan:2009:OPN

[MT09a] Aditya Mahajan and Demosthenis Teneketzis. Optimal performance of networked control systems with nonclassical information structures. SIAM Journal on Control and Optimization, 48(3):1377–1404, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mirkin:2009:IFS

[MT09b] Leonid Mirkin and Gilead Tadmor. Imposing FIR structure on H^{2} preview tracking and smoothing solutions. SIAM Journal on Control and Optimization, 48(4):2433– 2460, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Munch:2022:CEC

[MT22] Arnaud Münch and Emmanuel Trélat. Constructive exact control of semilinear 1D wave equations by a least-squares approach. SIAM Journal on Control and Optimization, 60(2):652–673, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1380661.

Maity:2023:OQS

[MT23] Dipankar Maity and Panagiotis Tsiotras. Optimal quantizer scheduling and controller synthesis for partially observable linear systems. SIAM Journal on Control and Optimization, 61(4):2682–2707, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1448707.

Mazzarella:2015:ERR

[MTS15] Luca Mazzarella, Francesco Ticozzi, and Alain Sarlette. Extending robustness and randomization from consensus to symmetrization algorithms. SIAM Journal on Control and Optimization, 53(4):2076– 2099, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Marques:2020:HPC

[MTS20] André L. Marques, Hélio V. M. Tozatti, and Josiney A. Souza. Higher prolongations of control affine systems: Absolute stability and generalized recurrence. SIAM Journal on Control and Optimization, 58(6):3019–3040, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mania:2008:MVH

[MTT08] Michael Mania, Revaz Tevzadze, and Teimuraz Toronjadze. Mean-variance hedging under partial information. SIAM Journal on Control and Optimization, 47(5):2381–2409, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Milz:2022:ASD

[MU22] Johannes Milz and Michael Ulbrich. An approximation scheme for distributionally robust PDE-constrained optimization. SIAM Journal on Control and Optimization, 60 (3):1410–1435, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M134664X.

Munoz:2000:SNC

[Muñ00] Julio Muñoz. On some necessary conditions of optimality for a nonlocal variational principle. SIAM Journal on Control and Optimization, 38(5): 1521–1533, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/34282.

Munos:2007:PBN

[Mun07] Rémi Munos. Performance bounds in L_{p}-norm for approximate value iteration. SIAM Journal on Control and Optimization, 46(2):541–561, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Munnoz:2014:ACG

[Mun14] Claudio Munñoz. On approximate controllability of generalized KdV solitons. SIAM Journal on Control and Optimization, 52(1):52–81, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Murray:1986:ETO

[Mur86] J. M. Murray. Existence theorems for optimal control and calculus of variations problems where the states can jump. SIAM Journal on Control and Optimization, 24(3):412–438, May 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Murota:1987:RSS

[Mur87] Kazuo Murota. Refined study on structural controllability of descriptor systems by means of matroids. SIAM Journal on Control and Optimization, 25 (4):967–989, July 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Murota:1989:MTA

[Mur89] Kazuo Murota. A matroidtheoretic approach to structurally fixed modes of control systems. SIAM Journal on Control and Optimization, 27(6):1381–1402, November 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Murphy:1992:EUV

[Mur92] K. A. Murphy. Estimation of unknown variable parameters in moving boundary problems. SIAM Journal on Control and Optimization, 30(3): 637–674, May 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Murota:1991:SIS

[MV91] Kazuo Murota and Jacob W.

Van Der Woude. Structure at infinity of structured descriptor systems and its applications. SIAM Journal on Control and Optimization, 29 (4):878–894, July 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Meidner:2007:AST

[MV07a] Dominik Meidner and Boris Vexler. Adaptive space–time finite element methods for parabolic optimization problems. SIAM Journal on Control and Optimization, 46(1): 116–142, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mirrahimi:2007:SFC

[MV07b] Mazyar Mirrahimi and Ramon Van Handel. Stabilizing feedback controls for quantum systems. SIAM Journal on Control and Optimization, 46(2): 445–467, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Meidner:2008:PEEa

[MV08a] Dominik Meidner and Boris Vexler. A priori error estimates for space–time finite element discretization of parabolic optimal control problems Part I: Problems without control constraints. SIAM Journal on Control and Optimization, 47(3):1150– 1177, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Meidner:2008:PEEb

[MV08b] Dominik Meidner and Boris Vexler. A priori error estimates for space–time finite element discretization of parabolic optimal control problems Part II: Problems with control constraints. SIAM Journal on Control and Optimization, 47(3):1301– 1329, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Meidner:2011:PEA

[MV11] Dominik Meidner and Boris Vexler. A priori error analysis of the Petrov–Galerkin Crank– Nicolson scheme for parabolic optimal control problems. SIAM Journal on Control and Optimization, 49(5):2183– 2211, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i5/p2183_s1.

Miranda-Villatoro:2018:SVS

[MVBC18] Félix A. Miranda-Villatoro, Bernard Brogliato, and Fernando Castaños. Set-valued sliding-mode control of uncertain linear systems: Continuous and discrete-time analysis. SIAM Journal on Control and Optimization, 56(3): 1756–1793, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Michiels:2009:SSR

[MVHN09] Wim Michiels, Tomáš Vyhlídal, Henri Huijberts, and Henk Nijmeijer. Stabilizability and stability robustness of state derivative feedback controllers. SIAM Journal on Control and Optimization, 47 (6):3100–3117, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Michiels:2009:SSN

[MVZ^{+}09] Wim Michiels, Tomáš Vyhlídal, Pavel Zítek, Henk Nijmeijer, and Didier Henrion. Strong stability of neutral equations with an arbitrary delay dependency structure. SIAM Journal on Control and Optimization, 48(2):763–786, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Meyer:1978:PCH

[MW78] R. R. Meyer and M. L. Wage. On the polyhedrality of the convex hull of the feasible set of an integer program. SIAM Journal on Control and Optimization, 16(4):682–687, July 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mordukhovich:2002:NSO

[MW02] Boris S. Mordukhovich and Bingwu Wang. Necessary suboptimality and optimality conditions via variational principles. SIAM Journal on Control and Optimization, 41(2):623–640, March 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/37481.

Mordukhovich:2004:OCN

[MW04] Boris S. Mordukhovich and Lianwen Wang. Optimal control of neutral functionaldifferential inclusions. SIAM Journal on Control and Optimization, 43(1):111–136, January 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42037.

Martin:2009:SVA

[MW09] Clyde F. Martin and Xiaochang A. Wang. Singular value assignment. SIAM Journal on Control and Optimization, 48(4):2388–2406, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Meyer:2021:OCP

[MW21] Christian Meyer and Stephan Walther. Optimal control of perfect plasticity. Part II: Displacement tracking. SIAM Journal on Control and Optimization, 59(4):2498–2523, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mehlitz:2022:SNS

[MW22] Patrick Mehlitz and Gerd Wachsmuth. Subdifferentiation of nonconvex sparsitypromoting functionals on Lebesgue spaces. SIAM Journal on Control and Optimization, 60(3):1819–1839, ????

2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1435173.

Merdon:2023:PRC

[MW23] Christian Merdon and Winnifried Wollner. Pressurerobustness in the context of optimal control. SIAM Journal on Control and Optimization, 61(1):342–360, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1482603.

Mei:2021:OEC

[MWY21] Hongwei Mei, Qingmeng Wei, and Jiongmin Yong. Optimal ergodic control of linear stochastic differential equations with quadratic cost functionals having indefinite weights. SIAM Journal on Control and Optimization, 59 (1):584–613, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ma:2019:GCF

[MXZ19] Jingtang Ma, Jie Xing, and Harry Zheng. Global closedform approximation of free boundary for optimal investment stopping problems. SIAM Journal on Control and Optimization, 57(3):2092– 2121, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mayeda:1979:SSC

[MY79] Hirokazu Mayeda and Takashi Yamada. Strong structural controllability. SIAM Journal on Control and Optimization, 17(1):123–138, January 1979. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Markus:1993:DBC

[MY93] Lawrence Markus and Yun Cheng You. Dynamical boundary control for elastic plates of general shape. SIAM Journal on Control and Optimization, 31(4):983–992, July 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Meyer:2009:SCO

[MY09] C. Meyer and I. Yousept.

State-constrained optimal control of semilinear elliptic equations with nonlocal radiation interface conditions. SIAM Journal on Control and Optimization, 48(2):734–755, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Miller:2017:OCC

[MY17] Christopher W. Miller and Insoon Yang. Optimal control of conditional value-at-risk in continuous time. SIAM Journal on Control and Optimization, 55(2):856–884, ????

2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

McDonald:2022:RIP

[MY22] Curtis McDonald and Serdar Yüksel. Robustness to incorrect priors and controlled filter stability in partially observed stochastic control. SIAM Journal on Control and Optimization, 60(2):842–870, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1417442.

Marden:2009:PBD

[MYAS09] Jason R. Marden, H. Peyton Young, Gürdal Arslan, and Jeff S. Shamma. Payoffbased dynamics for multiplayer weakly acyclic games. SIAM Journal on Control and Optimization, 48(1):373–396, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Marden:2014:APO

[MYP14] Jason R. Marden, H. Peyton Young, and Lucy Y. Pao. Achieving Pareto optimality through distributed learning. SIAM Journal on Control and Optimization, 52 (5):2753–2770, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Micu:1997:BCL

[MZ97] Sorin Micu and Enrique Zuazua. Boundary controllability of a linear hybrid system arising in the control of noise. SIAM Journal on Control and Optimization, 35(5): 1614–1637, September 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29797.

Micu:2006:CFO

[MZ06] Sorin Micu and Enrique Zuazua. On the controllability of a fractional order parabolic equation. SIAM Journal on Control and Optimization, 44 (6):1950–1972, January 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Merhi:2007:MRI

[MZ07] Amal Merhi and Mihail Zervos. A model for reversible investment capacity expansion. SIAM Journal on Control and Optimization, 46(3):839–876, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Morimoto:2009:OCG

[MZ09] Hiroaki Morimoto and Xun Yu Zhou. Optimal consumption in a growth model with the Cobb–Douglas production function. SIAM Journal on Control and Optimization, 47 (6):2991–3006, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Malisoff:2015:RAC

[MZ15] Michael Malisoff and Fumin Zhang. Robustness of adaptive control under time delays for three-dimensional curve tracking. SIAM Journal on Control and Optimization, 53(4): 2203–2236, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ma:2018:SDA

[MZ18] Hongbin Ma and Xinghong Zhang. Stabilization of decentralized adaptive control for nonlinearly parametrized coupled stochastic multiagent systems. SIAM Journal on Control and Optimization, 56(5): 3784–3815, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Mei:2020:CLE

[MZ20] Hongwei Mei and Chao Zhu. Closed-loop equilibrium for time-inconsistent McKean– Vlasov controlled problem. SIAM Journal on Control and Optimization, 58(6):3842– 3867, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

McEneaney:2022:SIS

[MZ22a] William M. McEneaney and Ruobing Zhao. Staticization and iterated staticization. SIAM Journal on Control and Optimization, 60 (3):1840–1862, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/19M1251278.

Michiels:2022:SRD

[MZ22b] Wim Michiels and Bin Zhou.

On the strong relative degree of time-delay systems with noncommensurate delays. SIAM Journal on Control and Optimization, 60 (6):3384–3408, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1442656.

Nowak:2002:ESG

[NA02] Andrzej S. Nowak and Eitan Altman. ε-equilibria for stochastic games with uncountable state space and unbounded costs. SIAM Journal on Control and Optimization, 40(6):1821–1839, November 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37837.

Nagai:1985:SCO

[Nag85]  Hideo Nagai.  Stochastic

control of one-dimensional diffusions whose generators have discontinuous coefficients. SIAM Journal on Control and Optimization, 23(3): 390–400, May 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nagase:1995:RNS

[Nag95] Noriaki Nagase. Remarks on nonlinear stochastic partial differential equations: An application of the splittingup method. SIAM Journal on Control and Optimization, 33(6):1716–1730, November 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24618.

Nagai:1996:BER

[Nag96] H. Nagai. Bellman equations of risk-sensitive control. SIAM Journal on Control and Optimization, 34(1):74–101, January 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25530.

Nagai:2003:OSR

[Nag03]  Hideo Nagai.  Optimal

strategies for risk-sensitive portfolio optimization problems for general factor models. SIAM Journal on Control and Optimization, 41(6): 1779–1800, November 2003.

CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39933.

Naito:1987:CSC

[Nai87] Koichiro Naito. Controllability of semilinear control systems dominated by the linear part. SIAM Journal on Control and Optimization, 25 (3):715–722, May 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Naito:1995:PSN

[Nai95]  Koichiro Naito.  Periodic

stability of nonlinear flexible systems with damping. SIAM Journal on Control and Optimization, 33(6): 1778–1800, November 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25352.

Nakagiri:1983:ILS

[Nak83] Shin-Ichi Nakagiri. Identifiability of linear systems in Hilbert spaces. SIAM Journal on Control and Optimization, 21(4):501–530, July 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Narukawa:1982:ACV

[Nar82] Kimiaki Narukawa. Admissible controllability of vibrating systems with constrained controls. SIAM Journal on Control and Optimization, 20(6): 770–782, November 1982. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Narukawa:1984:CCO

[Nar84] Kimiaki Narukawa. Complete controllability of onedimensional vibrating systems with bang–bang controls. SIAM Journal on Control and Optimization, 22(5):788– 804, September 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nguyen:2016:SAF

[NBF16] Le Ha Vy Nguyen, Catherine Bonnet, and André Ricardo Fioravanti. H_{∞}-stability analysis of fractional delay systems of neutral type. SIAM Journal on Control and Optimization, 54(2):740–759, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nourian:2013:NMF

[NC13] Mojtaba Nourian and Peter E. Caines. ϵ-Nash mean field game theory for nonlinear stochastic dynamical systems with major and minor agents. SIAM Journal on Control and Optimization, 51(4): 3302–3331, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Narvaez:2016:ARC

[NC16] Alfredo R. R. Narváez and Eduardo F. Costa. Average reachability of continuoustime Markov jump linear systems and the linear minimum mean square estimator. SIAM Journal on Control and Optimization, 54(4): 2063–2089, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ng:2013:ESS

[NCL13] Wai Man Ng, Dong Eui Chang, and George Labahn. Energy shaping for systems with two degrees of underactuation and more than three degrees of freedom. SIAM Journal on Control and Optimization, 51(2):881–905, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ning:2024:DGF

[NDHL24] Pengju Ning, Sergey N.

Dashkovskiy, Changchun Hua, and Kuo Li. Dual-gain function based prescribedtime output feedback control nonlinear time-delay systems. SIAM Journal on Control and Optimization, 62 (4):2254–2272, August 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1556496.

Nair:2004:SSL

[NE04] Girish N. Nair and Robin J. Evans. Stabilizability of stochastic linear systems with finite feedback data rates. SIAM Journal on Control and Optimization, 43(2):413– 436, March 2004. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/40211.

Nersesyan:2021:PAC

[Ner21] Vahagn Nersesyan. A proof of approximate controllability of the 3D Navier–Stokes system via a linear test. SIAM Journal on Control and Optimization, 59(4):2411–2427, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Newton:2000:OPQ

[New00] Nigel J. Newton. Observations preprocessing and quantization for nonlinear filters. SIAM Journal on Control and Optimization, 38(2):482–502, March 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/33114.

Newton:2006:DKB

[New06]  Nigel J. Newton.  Dual

Kalman–Bucy filters and interactive entropy production. SIAM Journal on Control and Optimization, 45(3):998– 1016, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Newton:2007:DNF

[New07] Nigel J. Newton. Dual nonlinear filters and entropy production. SIAM Journal on Control and Optimization, 46(5): 1637–1663, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ni:2013:PIA

[NF13] Yuan-Hua Ni and Hai-Tao Fang. Policy iteration algorithm for singular controlled diffusion processes. SIAM Journal on Control and Optimization, 51(5):3844–3862, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nishida:2015:SRE

[NFI15] Gou Nishida, Kenji Fujimoto, and Daiji Ichishima. Scalable reduction of elastic continuum for boundary energy control. SIAM Journal on Control and Optimization, 53(4): 2424–2448, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ngoc:2009:PSL

[Ngo09] Pham Huu Anh Ngoc. On positivity and stability of linear Volterra systems with delay. SIAM Journal on Control and Optimization, 48(3): 1939–1960, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ngoc:2011:SLV

[Ngo11] Pham Huu Anh Ngoc. Stability of linear Volterra– Stieltjes differential equations. SIAM Journal on Control and Optimization, 49(1):205– 226, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicon/resource/
                                        1/sjcodc/v49/i1/p205_s1.

Nguyen:2016:VAR

[Ngu16] Luong V. Nguyen. Variational analysis and regularity of the minimum time function for differential inclusions. SIAM Journal on Control and Optimization, 54(5): 2235–2258, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nguyen:2018:UAS

[Ngu18] Le Ha Vy Nguyen. A unified approach for the H_{∞}- Stability analysis of classical and fractional neutral systems with commensurate delays. SIAM Journal on Control and Optimization, 56(1): 538–555, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nguyen:2022:DCF

[Ngu22] Hoai-Minh Nguyen. A dependence of the cost of fast controls for the heat equation on the support of initial datum. SIAM Journal on Control and Optimization, 60(1):530–544, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1401280.

Nguyen:2012:LQG

[NH12] Son Luu Nguyen and Minyi Huang. Linear-quadratic-Gaussian mixed games with continuum-parametrized minor players. SIAM Journal on Control and Optimization, 50(5):2907–2937, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ninness:1999:GFT

[NHG99] Brett Ninness, Håkan Hjalmarsson, and Fredrik Gustafsson. Generalized Fourier and Toeplitz results for rational orthonormal bases. SIAM Journal on Control and Optimization, 37(2):429–460, March 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/30543.

Nicaise:1996:BEC

[Nic96] Serge Nicaise. Boundary exact controllability of interface problems with singularities I: Addition of the coefficients of singularities. SIAM Journal on Control and Optimization, 34(5):1512–1532, September 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/28210.

Nicaise:1997:BEC

[Nic97] Serge Nicaise. Boundary exact controllability of interface problems with singularities II: Addition of internal controls. SIAM Journal on Control and Optimization, 35(2):585–603, March 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/29203.

Nicaise:2000:EBC

[Nic00] Serge Nicaise. Exact boundary controllability of Maxwell’s equations in heterogeneous media and an application to an inverse source problem. SIAM Journal on Control and Optimization, 38(4):1145–1170, July 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34437.

Nielsen:2014:TFL

[Nie14] Christopher Nielsen. Transverse feedback linearization with partial information for single-input systems. SIAM Journal on Control and Optimization, 52(5):3002–3021, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Niederlander:2021:SOD

[Nie21] Simon K. Niederländer. Secondorder dynamics with Hessiandriven damping for linearly constrained convex minimization. SIAM Journal on Control and Optimization, 59(5): 3708–3736, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nihtila:1983:FDD

[Nih83] Markku T. Nihtilä. Finite dimensional deterministic nonlinear filters via Riccati transformation and Volterra series. SIAM Journal on Control and Optimization, 21(6): 916–930, November 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nikitin:1999:PCS

[Nik99] Sergey Nikitin. Piecewiseconstant stabilization. SIAM Journal on Control and Optimization, 37(3):911–933, May 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/32193.

Niu:2021:SCW

[Niu21] Jingrui Niu. Simultaneous control of wave systems. SIAM Journal on Control and Optimization, 59(3):2381–2409, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nurdin:2009:NSL

[NJD09] Hendra I. Nurdin, Matthew R. James, and Andrew C. Doherty. Network synthesis of linear dynamical quantum stochastic systems. SIAM Journal on Control and Optimization, 48(4):2686–2718, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Naik:1994:RIA

[NK94] Sanjeev M. Naik and P. R. Kumar. Robust indirect adaptive control of time-varying plants with unmodeled dynamics and disturbances. SIAM Journal on Control and Optimization, 32(6):1696–1725, November 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/23316.

Nair:2008:SSM

[NL08] Sujit Nair and Naomi Ehrich Leonard. Stable synchronization of mechanical system networks. SIAM Journal on Control and Optimization, 47(2): 661–683, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nesic:2009:SSS

[NLPT09] D. Nešić, A. Loría, E. Panteley, and A. R. Teel. On stability of sets for sampleddata nonlinear inclusions via their approximate discretetime models and summability criteria. SIAM Journal on Control and Optimization, 48(3):1888–1913, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nikoukhah:1991:RAW

[NLW91] Ramine Nikoukhah, Bernard C.

Levy, and Alan S. Willsky. Realization of acausal weighting patterns with boundary value descriptor systems. SIAM Journal on Control and Optimization, 29(2):420– 444, March 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ni:2019:MES

[NLZK19] Yuan-Hua Ni, Xun Li, Ji-Feng Zhang, and Miroslav Krstic. Mixed equilibrium solution of time-inconsistent stochastic linear-quadratic problem. SIAM Journal on Control and Optimization, 57(1):533–569, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nielsen:2008:LTF

[NM08] Christopher Nielsen and Manfredi Maggiore. On local transverse feedback linearization. SIAM Journal on Control and Optimization, 47(5): 2227–2250, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nielsen:2010:EPO

[NM10] Bjørn Fredrik Nielsen and Kent-Andre Mardal. Efficient preconditioners for optimality systems arising in connection with inverse problems. SIAM Journal on Control and Optimization, 48(8): 5143–5177, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nesic:1998:ODB

[NMBM98] D. Nesic, I. M. Y. Mareels, G. Bastin, and R. Mahony. Output dead beat control for a class of planar polynomial systems. SIAM Journal on Control and Optimization, 36(1):253–272, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28638.

Nishida:2015:BIM

[NMI15] Gou Nishida, Bernhard Maschke, and Ryojun Ikeura. Boundary integrability of multiple Stokes–Dirac structures. SIAM Journal on Control and Optimization, 53(2):800–815, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nagase:1990:OCS

[NN90] Noriaki Nagase and Makiko Nisio. Optimal controls for stochastic partial differential equations. SIAM Journal on Control and Optimization, 28(1):186–213, January 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nguyen:2021:SDP

[NNN21] Dang H. Nguyen, Duy Nguyen, and Son L. Nguyen. Stability in distribution of pathdependent hybrid diffusion. SIAM Journal on Control and Optimization, 59(1):434–463, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ngoc:2008:SRS

[NNSM08] Pham Huu Anh Ngoc, Toshiki Naito, Jong Son Shin, and Satoru Murakami. On stability and robust stability of positive linear Volterra equations. SIAM Journal on Control and Optimization, 47(2): 975–996, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nam:2022:EEP

[NNT22] Phan Thanh Nam, Tran Ngoc Nguyen, and Hieu Trinh. Exponential estimate of positive time-delay systems with polytopic uncertainties. SIAM Journal on Control and Optimization, 60(5):3026–3050, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1466062.

Nedelcu:2014:CCI

[NNTD14] Valentin Nedelcu, Ion Necoara, and Quoc Tran-Dinh. Computational complexity of inexact gradient augmented Lagrangian methods: Application to constrained MPC. SIAM Journal on Control and Optimization, 52(5):3109– 3134, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nakoulima:2003:PCN

[NOV03] O. Nakoulima, A. Omrane, and J. Velin. On the Pareto control and no-regret control for distributed systems with incomplete data. SIAM Journal on Control and Optimization, 42(4):1167–1184, July 2003. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38018.

Novruzi:2004:ERC

[Nov04] Arian Novruzi. C^{2}, α existence result for a class of shape optimization problems. SIAM Journal on Control and Optimization, 43(1):174–193, January 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38240.

Nowak:1985:MST

[Now85] Andrzej S. Nowak. Measurable selection theorems for minimax stochastic optimization problems. SIAM Journal on Control and Optimization, 23 (3):466–476, May 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nowak:1989:EOS

[Now89] Andrzej S. Nowak. Existence of optimal strategies in zerosum nonstationary stochastic games with lack of information on both sides. SIAM Journal on Control and Optimization, 27(2):289–295, March 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nowakowski:2008:SOC

[Now08] Andrzej Nowakowski. Sufficient optimality conditions for Dirichlet boundary control of wave equations. SIAM Journal on Control and Optimization, 47(1):92–110, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nowakowski:2017:FOS

[Now17] Andrzej Nowakowski. First order sufficient optimality conditions for Navier–Stokes flow. Dual feedback controls. SIAM Journal on Control and Optimization, 55(4):2734–2747, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nguyen:1982:IND

[NP82] Hung T. Nguyen and Tuan D. Pham. Identification of nonstationary diffusion model by the method of sieves. SIAM Journal on Control and Optimization, 20(5):603– 611, September 1982. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Neto:1996:UCA

[NP96] C. Kirjner Neto and E. Polak. On the use of consistent approximations for the optimal design of beams. SIAM Journal on Control and Optimization, 34(6):1891–1913, November 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26618.

Nicaise:2006:SIR

[NP06] Serge Nicaise and Cristina Pignotti. Stability and instability results of the wave equation with a delay term in the boundary or internal feedbacks. SIAM Journal on Control and Optimization, 45(5): 1561–1585, ???? 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nguwi:2022:CAE

[NP22] Jiang Yu Nguwi and Nicolas Privault. A constructive approach to existence of equilibria in time-inconsistent stochastic control problems. SIAM Journal on Control and Optimization, 60(2):674–698, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1370446.

Neitzel:2015:FED

[NPR15] Ira Neitzel, Johannes Pfefferer, and Arnd Rösch. Finite element discretization of stateconstrained elliptic optimal control problems with semilinear state equation. SIAM Journal on Control and Optimization, 53(2):874–904, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nandakumaran:2015:PCO

[NPS15] A. K. Nandakumaran, Ravi Prakash, and Bidhan Chandra Sardar. Periodic controls in an oscillating domain: Controls via unfolding and homogenization. SIAM Journal on Control and Optimization, 53(5): 3245–3269, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nemcova:2013:RTN

[NPvS13] Jana Nemcová, Mihály Petreczky, and Jan H. van Schuppen. Realization theory of Nash systems. SIAM Journal on Control and Optimization, 51(5):3386–3414, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Noba:2020:BDP

[NPY20] Kei Noba, José-Luis Pérez, and Xiang Yu. On the bailout dividend problem for spectrally negative Markov additive models. SIAM Journal on Control and Optimization, 58(2):1049–1076, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nagpal:1997:CEP

[NR97] Krishan M. Nagpal and R. Ravi. H_{∞} control and estimation problems with delayed measurements: Statespace solutions. SIAM Journal on Control and Optimization, 35(4):1217–1243, July 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27749.

Nguyen:2015:BSN

[NR15] Phuong Anh Nguyen and Jean-Pierre Raymond. Boundary stabilization of the Navier– Stokes equations in the case of mixed boundary conditions. SIAM Journal on Control and Optimization, 53(5): 3006–3039, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nicolau:2017:FMI

[NR17] Florentina Nicolau and Witold Respondek. Flatness of multiinput control-affine systems linearizable via one-fold prolongation. SIAM Journal on Control and Optimization, 55 (5):3171–3203, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nendel:2021:UEF

[NR21] Max Nendel and Michael Röckner. Upper envelopes of families of Feller semigroups and viscosity solutions to a class of nonlinear Cauchy problems. SIAM Journal on Control and Optimization, 59 (6):4400–4428, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nicolau:2020:FIT

[NRB20] Florentina Nicolau, Witold Respondek, and Jean-Pierre Barbot. Flat inputs: Theory and applications. SIAM Journal on Control and Optimization, 58(6):3293–3321, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nicolau:2022:LNF

[NRL22] Florentina Nicolau, Witold Respondek, and Shunjie Li. Linearization via nonregular feedback of multi-input nonlinear control systems. SIAM Journal on Control and Optimization, 60(4):2601–2630, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1450331.

Nijmeijer:1986:RLN

[NS86] H. Nijmeijer and J. M. Schumacher. The regular local noninteracting control problem for nonlinear control systems. SIAM Journal on Control and Optimization, 24(6): 1232–1245, November 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Naito:1991:IAR

[NS91] Koichiro Naito and Thomas I. Seidman. Invariance of the approximately reachable set under nonlinear perturbations.

SIAM Journal on Control and Optimization, 29(3):731–750, May 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ndiaye:2000:DSQ

[NS00] Papa Momar Ndiaye and Michel Sorine. Delay sensitivity of quadratic controllers: a singular perturbation approach. SIAM Journal on Control and Optimization, 38(6):1655–1682, November 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32985.

Ngoc:2005:SRP

[NS05] Pham Huu Anh Ngoc and Nguyen Khoa Son. Stability radii of positive linear functional differential equations under multi-perturbations. SIAM Journal on Control and Optimization, 43(6): 2278–2295, November 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43478.

Nutz:2012:SDR

[NS12] Marcel Nutz and H. Mete Soner. Superhedging and dynamic risk measures under volatility uncertainty. SIAM Journal on Control and Optimization, 50(4):2065–2089, ???? 2012. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Ntogramatzidis:2014:REA

[NS14] Lorenzo Ntogramatzidis and Robert Schmid. Robust eigenstructure assignment in geometric control theory. SIAM Journal on Control and Optimization, 52(2):960–986, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Neufeld:2018:RUM

[NS18] Ariel Neufeld and Mario Sikić. Robust utility maximization in discrete-time markets with friction. SIAM Journal on Control and Optimization, 56 (3):1912–1937, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nguyen:2020:NOI

[NS20] Thai Nguyen and Mitja Stadje. Nonconcave optimal investment with value-at-risk constraint: an application to life insurance contracts. SIAM Journal on Control and Optimization, 58(2):895–936, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nandakumaran:2023:HCD

[NS23] A. K. Nandakumaran and Abu Sufian. High contrasting diffusion in Heisenberg group: Homogenization of optimal control via unfolding. SIAM Journal on Control and Optimization, 61(3):1350–1374, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1488582.

Nandi:2019:OCF

[NSC19] Anirban Nandi, Heinz Schättler, and ShiNung Ching. Optimal control for fast, accurate threshold-hitting. SIAM Journal on Control and Optimization, 57(4):2269–2291, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nie:2017:CBM

[NSW17] Tianyang Nie, Jingtao Shi, and Zhen Wu. Connection between MP and DPP for stochastic recursive optimal control problems: Viscosity solution framework in the general case. SIAM Journal on Control and Optimization, 55 (5):3258–3294, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Neittaanmaki:1995:EDA

[NT95] P. Neittaanmäki and D. Tiba.

An embedding of domains approach in free boundary problems and optimal design. SIAM Journal on Control and Optimization, 33(5): 1587–1602, September 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/23112.

Nguyen:2016:ODO

[NT16] Minh Cuong Nguyen and Hieu Trinh. Observer design for one-sided Lipschitz discretetime systems subject to delays and unknown inputs. SIAM Journal on Control and Optimization, 54(3):1585–1601, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ngoc:2017:SAN

[NT17] Pham Huu Anh Ngoc and Hieu Trinh. Stability analysis of nonlinear neutral functional differential equations. SIAM Journal on Control and Optimization, 55(6):3947–3968, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ngoc:2018:CFD

[NT18] Pham Huu Anh Ngoc and Hieu Trinh. On contraction of functional differential equations. SIAM Journal on Control and Optimization, 56(3): 2377–2397, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ntogramatzidis:2012:SIT

[Nto12] Lorenzo Ntogramatzidis. Structural invariants of two-dimensional systems. SIAM Journal on Control and Optimization, 50 (1):334–356, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v50/i1/p334_s1.

Nam:2018:MSB

[NTP18] P. T. Nam, H. Trinh, and P. N. Pathirana. Minimization of state bounding for perturbed positive systems with delays. SIAM Journal on Control and Optimization, 56(3): 1739–1755, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nurdin:2006:SFC

[Nur06] Hendra I. Nurdin. Spectral factorization of a class of matrix-valued spectral densities. SIAM Journal on Control and Optimization, 45(5): 1801–1821, ???? 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nutz:2018:MFG

[Nut18] Marcel Nutz. A mean field game of optimal stopping. SIAM Journal on Control and Optimization, 56(2):1206– 1221, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Naghnaeian:2018:ASL

[NVD18] Mohammad Naghnaeian, Petros G. Voulgaris, and Geir E.

Dullerud. l_{p} analysis and synthesis of linear switched systems: a unified inputoutput and state-space approach. SIAM Journal on Control and Optimization, 56 (2):1181–1205, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Napp:2010:PB

[NvdPS10] Diego Napp, Marius van der Put, and Shiva Shankar. Periodic behaviors. SIAM Journal on Control and Optimization, 48(7):4652–4663, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nemcova:2009:RTR

[NvS09] Jana Nĕmcová and Jan H. van Schuppen. Realization theory for rational systems: The existence of rational realizations. SIAM Journal on Control and Optimization, 48(4): 2840–2856, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Neitzel:2017:OCP

[NWW17] I. Neitzel, T. Wick, and W. Wollner. An optimal control problem governed by a regularized phase-field fracture propagation model. SIAM Journal on Control and Optimization, 55(4):2271– 2288, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Neitzel:2019:OCP

[NWW19] Ira Neitzel, Thomas Wick, and Winnifried Wollner. An optimal control problem governed by a regularized phase-field fracture propagation model. Part II: The regularization limit. SIAM Journal on Control and Optimization, 57(3): 1672–1690, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nguyen:2016:MAS

[NY16] Dang Hai Nguyen and George Yin. Modeling and analysis of switching diffusion systems: Past-dependent switching with a countable state space. SIAM Journal on Control and Optimization, 54(5): 2450–2477, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nguyen:2018:SRS

[NY18] Dang Hai Nguyen and George Yin. Stability of regimeswitching diffusion systems with discrete states belonging to a countable set. SIAM Journal on Control and Optimization, 56(5):3893–3917, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Nyman:1994:SOM

[Nym94]  P.-O. Nyman.  A super-

optimization method for fourblock problems. SIAM Journal on Control and Optimization, 32(1):86–115, January 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/21751.

Ni:2017:DOC

[NYZZ17] Yuan-Hua Ni, Ka-Fai Cedric Yiu, Huanshui Zhang, and Ji-Feng Zhang. Delayed optimal control of stochastic LQ problem. SIAM Journal on Control and Optimization, 55(5): 3370–3407, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Napoli:1999:TDP

[NZ99] Mariateresa Napoli and Sandro Zampieri. Two-dimensional proper rational matrices and causal input/output representations of two-dimensional behavioral systems. SIAM Journal on Control and Optimization, 37(5):1538–1552, September 1999. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/33818.

Ng:2003:GWS

[NZ03] Kung Fu Ng and Xi Yin Zheng. Global weak sharp minima on Banach spaces. SIAM Journal on Control and Optimization, 41(6): 1868–1885, November 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38946.

OBrien:1979:PCS

[O’B79]  R. E. O’Brien.  Pertur-

bation of controllable systems. SIAM Journal on Control and Optimization, 17(2): 175–179, March 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ozsari:2019:PBA

[ÖB19] Türker Özsari and Ahmet Batal. Pseudo-backstepping and its application to the control of Korteweg–de Vries equation from the right endpoint on a finite domain. SIAM Journal on Control and Optimization, 57(2):1255– 1283, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ober:1991:BPC

[Obe91] Raimund Ober. Balanced parametrization of classes of linear systems. SIAM Journal on Control and Optimization, 29(6):1251–1287, November 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Oberst:2006:SSM

[Obe06] Ulrich Oberst. Stability and stabilization of multidimensional Input/Output systems. SIAM Journal on Control and Optimization, 45(4):1467– 1507, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Oberst:2007:FFT

[Obe07] Ulrich Oberst. The Fast Fourier Transform. SIAM Journal on Control and Optimization, 46(2):496–540, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Oberst:2009:RAI

[Obe09] Ulrich Oberst. Regular almost interconnection of multidimensional behaviors. SIAM Journal on Control and Optimization, 48(2):463–480, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ohsawa:2011:DHJ

[OBL11] Tomoki Ohsawa, Anthony M. Bloch, and Melvin Leok. Discrete Hamilton–Jacobi theory. SIAM Journal on Control and Optimization, 49(4): 1829–1856, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicon/resource/
                                        1/sjcodc/v49/i4/p1829_s1.

Opmeer:2004:LQG

[OC04] Mark R. Opmeer and Ruth F.

Curtain. Linear quadratic Gaussian balancing for discretetime infinite-dimensional linear systems. SIAM Journal on Control and Optimization, 43(4):1196–1221, July 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43118.

Ocanna:2012:ESS

[OC12] Eladio Ocanña and Pierre Cartigny. Explicit solutions for singular infinite horizon calculus of variations. SIAM Journal on Control and Optimization, 50(5):2573–2587, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Oner:2023:SSC

[OC23] Isil Oner and M. Kanat Camlibel. Stabilizability of strict convex processes with respect to arbitrary stability domains. SIAM Journal on Control and Optimization, 61 (3):1819–1834, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1487138.

Oostveen:2000:ATS

[OCI00] J. C. Oostveen, R. F. Curtain, and K. Ito. An approximation theory for strongly stabilizing solutions to the operator LQ Riccati equation. SIAM Journal on Control and Optimization, 38(6):1909–1937, November 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33969.

Ozer:2014:EBC

[ÖH14] A. Özkan Özer and Scott W. Hansen. Exact boundary controllability results for a multilayer Rao–Nakra sandwich beam. SIAM Journal on Control and Optimization, 52(2): 1314–1337, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ohsawa:2013:SRO

[Ohs13] Tomoki Ohsawa. Symmetry reduction of optimal control systems and principal connections. SIAM Journal on Control and Optimization, 51(1): 96–120, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ohtsuka:2008:ASN

[Oht08] Toshiyuki Ohtsuka. Algebraic structures in nonlinear systems over rings obtained by immersion. SIAM Journal on Control and Optimization, 47(4):1961–1976, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ortega:1977:DMP

[OL77] José Antonio Ortega and Richard Jeffrey Leake. Discrete maximum principle with state constrained control. SIAM Journal on Control and Optimization, 15(6):984– 990, November 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ozcaldiran:1989:GRS

[ÖL89] K. Özçaldiran and F. L. Lewis.

Generalized reachability subspaces for singular systems. SIAM Journal on Control and Optimization, 27(3):495–510, May 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Olsen:1976:MSPb

[Ols76a]  Paul Olsen.  Multistage

stochastic programming with recourse as mathematical programming in an L_{p} space. SIAM Journal on Control and Optimization, 14(3):528–537, May 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Olsen:1976:MSPa

[Ols76b]  Paul Olsen.  Multistage

stochastic programming with recourse: The equivalent deterministic problem. SIAM Journal on Control and Optimization, 14(3):495–517, May 1976. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Olsen:1976:WMS

[Ols76c] Paul Olsen. When is a multistage stochastic programming problem well-defined? SIAM Journal on Control and Optimization, 14(3):518–527, May 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Olsder:2002:OCL

[Ols02] Geert Jan Olsder. On openand closed-loop bang–bang control in nonzero-sum differential games. SIAM Journal on Control and Optimization, 40(4):1087–1106, July 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37325.

Olshevsky:2014:CTM

[Ols14]  Alex Olshevsky.  Consen-

sus with ternary Messages. SIAM Journal on Control and Optimization, 52(2):987–1009, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Olshevsky:2017:LTA

[Ols17] Alex Olshevsky. Linear time average consensus and distributed optimization on fixed graphs. SIAM Journal on Control and Optimization, 55 (6):3990–4014, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

OMalley:1976:MDS

[O’M76] R. E. O’Malley, Jr. A more direct solution of the nearly singular linear regulator problem. SIAM Journal on Control and Optimization, 14(6):1063– 1077, November 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ogura:2014:SAP

[OM14] Masaki Ogura and Clyde F.

Martin. Stability analysis of positive semi-Markovian jump linear systems with state resets. SIAM Journal on Control and Optimization, 52(3): 1809–1831, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ortigosa:2021:OCS

[OMFMC^{+}21] Rogelio Ortigosa, Jesús Martínez-Frutos, Carlos Mora-Corral, Pablo Pedregal, and Francisco Periago. Optimal control of soft materials using a Hausdorff distance functional. SIAM Journal on Control and Optimization, 59(1):393–416, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ohta:1984:ROR

[OMK84] Yoshito Ohta, Hajime Maeda, and Shinzo Kodama. Reachability, observability, and realizability of continuous-time positive systems. SIAM Journal on Control and Optimization, 22(2):171–180, March 1984. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ober:1990:BTI

[OMS90] Raimund Ober and Stephen Montgomery-Smith. Bilinear transformation of infinitedimensional state-space systems and balanced realizations of nonrational transfer functions. SIAM Journal on Control and Optimization, 28(2): 438–465, March 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

ONeill:1976:NDM

[O’N76] Richard P. O’Neill. Nested decomposition of multistage convex programs. SIAM Journal on Control and Optimization, 14(3):409–418, May 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Oza:2015:CUF

[OOS15] Harshal B. Oza, Yury V. Orlov, and Sarah K. Spurgeon. Continuous uniform finite time stabilization of planar controllable systems. SIAM Journal on Control and Optimization, 53(3):1154–1181, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ocone:1996:ASO

[OP96] Daniel Ocone and Etienne Pardoux. Asymptotic stability of the optimal filter with respect to its initial condition. SIAM Journal on Control and Optimization, 34 (1):226–243, January 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25661.

Opmeer:2007:LBC

[Opm07] Mark R. Opmeer. LQG balancing for continuous-time infinite-dimensional systems. SIAM Journal on Control and Optimization, 46(5):1831– 1848, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Orlov:2017:OFS

[OPPU17] Y. Orlov, A. Pisano, A. Pilloni, and E. Usai. Output feedback stabilization of coupled reaction-diffusion processes with constant parameters. SIAM Journal on Control and Optimization, 55(6): 4112–4155, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Orey:1987:RZR

[OPS87] Steven Orey, Victor Pestien, and William Sudderth. Reaching zero rapidly. SIAM Journal on Control and Optimization, 25(5):1253–1265, September 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Orlov:2004:FTS

[Orl04] Y. Orlov. Finite time stability and robust control synthesis of uncertain switched systems. SIAM Journal on Control and Optimization, 43(4):1253–1271, July 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42559.

Opmeer:2013:FRA

[ORW13] Mark R. Opmeer, Timo Reis, and Winnifried Wollner. Finite-rank ADI iteration for operator Lyapunov equations. SIAM Journal on Control and Optimization, 51(5): 4084–4117, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Oksendal:2002:OCP

[OS02] Bernt Oksendal and Agnès Sulem. Optimal consumption and portfolio with both fixed and proportional transaction costs. SIAM Journal on Control and Optimization, 40(6):1765–1790, November 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37601.

Oksendal:2010:MPO

[ØS10a] Bernt Øksendal and Agnès Sulem. Maximum principles for optimal control of forwardbackward stochastic differential equations with jumps. SIAM Journal on Control and Optimization, 48(5):2945– 2976, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Opmeer:2010:OIO

[OS10b] Mark R. Opmeer and Olof J.

Staffans. Optimal inputoutput stabilization of infinitedimensional discrete timeinvariant linear systems by output injection. SIAM Journal on Control and Optimization, 48(8):5084–5107, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Oksendal:2012:SSC

[ØS12a] Bernt Øksendal and Agnès Sulem. Singular stochastic control and optimal stopping with partial information of Itô–Lévy processes. SIAM Journal on Control and Optimization, 50(4):2254–2287, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Opmeer:2012:CFO

[OS12b] Mark R. Opmeer and Olof J. Staffans. Coprime factorization and optimal control on the doubly infinite discrete time axis. SIAM Journal on Control and Optimization, 50 (1):266–285, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicon/resource/
                                        1/sjcodc/v50/i1/p266_s1.

Opmeer:2014:OCD

[OS14] Mark R. Opmeer and Olof J. Staffans. Optimal control on the doubly infinite continuous time axis and coprime factorizations. SIAM Journal on Control and Optimization, 52(3):1958–2007, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Opmeer:2019:OCD

[OS19] Mark R. Opmeer and Olof J. Staffans. Optimal control on the doubly infinite time axis for well-posed linear systems. SIAM Journal on Control and Optimization, 57(3): 1985–2015, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Oshima:2006:OSP

[Osh06] Yoichi Oshima. On an optimal stopping problem of time inhomogeneous diffusion processes. SIAM Journal on Control and Optimization, 45(2): 565–579, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Osses:2001:RMA

[Oss01] Axel Osses. A rotated multiplier applied to the controllability of waves, elasticity, and tangential Stokes control. SIAM Journal on Control and Optimization, 40(3):777–800, May 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34561.

OConnor:1983:FSC

[OT83] Daniel A. O’Connor and T. J.

Tarn. On the function space controllability of linear neutral systems. SIAM Journal on Control and Optimization, 21 (2):306–329, March 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ozbay:1990:STA

[ÖT90] Hitay Özbay and Allen Tannenbaum. A skew Toeplitz approach to the H^{∞} optimal control of multivariable distributed systems. SIAM Journal on Control and Optimization, 28(3):653–670, May 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Olkin:1994:PDC

[OT94] Ingram Olkin and Y. L. Tong. Positive dependence of a class of multivariate exponential distributions. SIAM Journal on Control and Optimization, 32(4):965–974, July 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/21740.

Olshevsky:2009:CSD

[OT09] Alex Olshevsky and John N. Tsitsiklis. Convergence speed in distributed consensus and averaging. SIAM Journal on Control and Optimization, 48 (1):33–55, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

ORourke:2016:CGC

[OT16] Sean O’Rourke and Behrouz Touri. On a conjecture of Godsil concerning controllable random graphs. SIAM Journal on Control and Optimization, 54(6):3347–3378, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Outrata:2000:GMP

[Out00]  J. V. Outrata.  A gen-

eralized mathematical program with equilibrium constraints. SIAM Journal on Control and Optimization, 38(5):1623–1638, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35291.

Ouvrard:1978:LFH

[Ouv78a] Jean-Yves Ouvrard. Linear filtering in Hilbert spaces. II: An application to the smoothing theory for hereditary systems with observation delays. SIAM Journal on Control and Optimization, 16(6): 938–952, November 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ouvrard:1978:MPL

[Ouv78b] Jean Yves Ouvrard. Martingale projection and linear filtering in Hilbert spaces. I: The theory. SIAM Journal on Control and Optimization, 16(6): 912–937, November 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ouzahra:2010:EWS

[Ouz10] Mohamed Ouzahra. Exponential and weak stabilization of constrained bilinear systems. SIAM Journal on Control and Optimization, 48(6): 3962–3974, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Overkamp:2000:MSD

[OvS00] Ard Overkamp and Jan H. van Schuppen. Maximal solutions in decentralized supervisory control. SIAM Journal on Control and Optimization, 39(2):492–511, March 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/32113.

Olgac:2008:NPS

[OVS08] Nejat Olgac, Tomáš Vyhlídal, and Rifat Sipahi. A new perspective in the stability assessment of neutral systems with multiple and cross-talking delays. SIAM Journal on Control and Optimization, 47(1): 327–344, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ovseevich:2014:CMT

[Ovs14] Alexander Ovseevich. Complexity of the minimum-time damping of a physical pendulum. SIAM Journal on Control and Optimization, 52 (1):82–96, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ozveren:1992:TRD

[ÖW92] Cüneyt M. Özveren and Alan S. Willsky. Tracking and restrictability in discrete event dynamic systems. SIAM Journal on Control and Optimization, 30(6):1423–1446, November 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ober:1993:ASI

[OW93] Raimund Ober and Yuanyin Wu. Asymptotic stability of infinite-dimensional discretetime balanced realizations. SIAM Journal on Control and Optimization, 31(5):1321– 1339, September 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ober:1996:IDC

[OW96] Raimund J. Ober and Yuanyin Wu. Infinite-dimensional continuous-time linear systems: Stability and structure analysis. SIAM Journal on Control and Optimization, 34(3):757–812, May 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24531.

Ocone:2000:DVC

[OW00] Daniel Ocone and Ananda Weerasinghe. Degenerate variance control of a onedimensional diffusion. SIAM Journal on Control and Optimization, 39(1):1–24, January 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/34712.

Ortega:2000:GSS

[OZ00] Jaime H. Ortega and Enrique Zuazua. Generic simplicity of the spectrum and stabilization for a plate equation. SIAM Journal on Control and Optimization, 39(5): 1585–1614, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/35848. See ad-

dendum [OZ03].

Ortega:2003:AGS

[OZ03] Jaime H. Ortega and Enrique Zuazua. Addendum to “Generic Simplicity of the Spectrum and Stabilization for a Plate Equation”. SIAM Journal on Control and Optimization, 42(5): 1905–1910, September 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/41403.  See

[OZ00].

Ozekici:1985:ORO

[Öze85] Süleyman Özekici. Optimal replacement of one-unit systems under periodic inspection. SIAM Journal on Control and Optimization, 23(1): 122–128, January 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Potter:1993:DAL

[PA93] Lee C. Potter and K. S. Arun.

A dual approach to linear inverse problems with convex constraints. SIAM Journal on Control and Optimization, 31 (4):1080–1092, July 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Peuteman:1999:ARS

[PA99] Joan Peuteman and Dirk Aeyels. Averaging results and the study of uniform asymptotic stability of homogeneous differential equations that are not fast time-varying. SIAM Journal on Control and Optimization, 37(4):997–1010, July 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32386.

Poore:1988:ELS

[PAH88] A. B. Poore and Q. Al-Hassan.

The expanded Lagrangian system for constrained optimization problems. SIAM Journal on Control and Optimization, 26(2):417–427, March 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pandolfi:1983:CRS

[Pan83] Luciano Pandolfi. Canonical realizations of systems with delays. SIAM Journal on Control and Optimization, 21 (4):598–613, July 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pandolfi:1992:SPD

[Pan92]  Luciano Pandolfi.  Some

properties of distributed control systems with finitedimensional input space. SIAM Journal on Control and Optimization, 30(4):926–941, July 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pandolfi:1998:DLP

[Pan98]  L. Pandolfi.  Dissipativ-

ity and the Luré problem for parabolic boundary control systems. SIAM Journal on Control and Optimization, 36(6):2061–2081, November 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31543.

Pang:2013:CGD

[Pan13] C. H. Jeffrey Pang. Characterizing generalized derivatives of set-valued maps: Extending the tangential and normal approaches. SIAM Journal on Control and Optimization, 51 (1):145–171, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pandolfi:2017:CVP

[Pan17] Luciano Pandolfi. Controllability of a viscoelastic plate using one boundary control in displacement or bending. SIAM Journal on Control and Optimization, 55(6):4092– 4111, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Papavassilopoulos:1981:SSS

[Pap81] G. P. Papavassilopoulos. Solution of some stochastic quadratic Nash and leaderfollower games. SIAM Journal on Control and Optimization, 19(5):651–666, September 1981. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Papageorgiou:1986:EOA

[Pap86a] Nikolaos S. Papageorgiou. On the efficiency and optimality of allocations. II. SIAM Journal on Control and Optimization, 24(3):452–479, May 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Papavassilopoulos:1986:ITN

[Pap86b] G. P. Papavassilopoulos. Iterative techniques for the Nash solution in quadratic games with unknown parameters. SIAM Journal on Control and Optimization, 24(4):821–834, July 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Papageorgiou:1989:PRT

[Pap89] Nikolaos S. Papageorgiou. Properties of the relaxed trajectories of evolution equations and optimal control. SIAM Journal on Control and Optimization, 27(2):267– 288, March 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Peuteman:2000:BPT

[PAS00] Joan Peuteman, Dirk Aeyels, and Rodolphe Sepulchre. Boundedness properties for time-varying nonlinear systems. SIAM Journal on Control and Optimization, 39(5): 1408–1422, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/36165.

Paulsen:2008:ODP

[Pau08] Jostein Paulsen. Optimal dividend payments and reinvestments of diffusion processes with both fixed and proportional costs. SIAM Journal on Control and Optimization, 47(5):2201–2226, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Paunonen:2017:RCR

[Pau17] Lassi Paunonen. Robust controllers for regular linear systems with infinite-dimensional exosystems. SIAM Journal on Control and Optimization, 55(3):1567–1597, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Paunonen:2019:SRR

[Pau19] Lassi Paunonen. Stability and robust regulation of passive linear systems. SIAM Journal on Control and Optimization, 57(6):3827–3856, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pavon:1980:SRI

[Pav80] Michele Pavon. Stochastic realization and invariant directions of the matrix Riccati equation. SIAM Journal on Control and Optimization, 18 (2):155–180, March 1980. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pavon:1984:NRI

[Pav84a] Michele Pavon. New results on the interpolation problem for continuous-time stationary increments processes. SIAM Journal on Control and Optimization, 22(1):133–142, January 1984. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pavon:1984:OIL

[Pav84b] Michele Pavon. Optimal interpolation for linear stochastic systems. SIAM Journal on Control and Optimization, 22 (4):618–629, July 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pavel:2013:CSS

[Pav13] Lacra Pavel. Classical solutions in Sobolev spaces for a class of hyperbolic Lotka– Volterra systems. SIAM Journal on Control and Optimization, 51(3):2132–2151, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Petersen:1983:CCG

[PB83] Ian R. Petersen and B. Ross Barmish. On a certain controllability gap. SIAM Journal on Control and Optimization, 21 (1):86–94, January 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pan:1996:MSO

[PB96] Zigang Pan and Tamer Başar. Model simplification and optimal control of stochastic singularly perturbed systems under exponentiated quadratic cost. SIAM Journal on Control and Optimization, 34(5): 1734–1766, September 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25936.

Pan:1999:BCD

[PB99a] Zigang Pan and Tamer Basar.

Backstepping controller design for nonlinear stochastic systems under a risk-sensitive cost criterion. SIAM Journal on Control and Optimization, 37(3):957–995, May 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30705.

Patek:1999:SSP

[PB99b] Stephen D. Patek and Dimitri P. Bertsekas. Stochastic shortest path games. SIAM Journal on Control and Optimization, 37(3):804–824, May 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29955.

Pal:2009:DUS

[PB09] Debasattam Pal and Madhu N.

Belur. Dissipativity of uncontrollable systems, storage functions, and Lyapunov functions. SIAM Journal on Control and Optimization, 47(6): 2930–2966, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Peters:2015:SNL

[PBPB15] J. R. Peters, D. Borra, B. E.

Paden, and F. Bullo. Sensor network localization on the group of three-dimensional displacements. SIAM Journal on Control and Optimization, 53(6):3534–3561, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Prandini:2000:ALC

[PC00] M. Prandini and M. C. Campi. Adaptive LQG control of input-output systems — a cost-biased approach. SIAM Journal on Control and Optimization, 39(5):1499–1519, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36636.

Parada:2022:GWP

[PCP22] Hugo Parada, Emmanuelle Crépeau, and Christophe Prieur. Global well-posedness of the KdV equation on a star-shaped network and stabilization by saturated controllers. SIAM Journal on Control and Optimization, 60 (4):2268–2296, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1434581.

Pearlman:1979:CRB

[PD79] J. G. Pearlman and M. J. Denham. Canonical realization of bilinear input/output maps. SIAM Journal on Control and Optimization, 17(4): 451–468, July 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Phat:1992:CCL

[PD92] Vu Ngoc Phat and Trinh Cong Dieu. Constrained controllability of linear discrete nonstationary systems in Banach spaces. SIAM Journal on Control and Optimization, 30(6): 1311–1318, November 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pavon:1997:FSM

[PD97] M. Pavon and D. D’Alessandro.

Families of solutions of matrix Riccati differential equations. SIAM Journal on Control and Optimization, 35 (1):194–204, January 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27633.

Pirie:2002:RCS

[PD02] Carol Pirie and Geir E.

Dullerud. Robust controller synthesis for uncertain timevarying systems. SIAM Journal on Control and Optimization, 40(4):1312–1331, July 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36963.

Proskurnikov:2023:YLR

[PDB23] Anton V. Proskurnikov, Alexander Davydov, and Francesco Bullo. The Yakubovich Slemma revisited: Stability and contractivity in non-Euclidean norms. SIAM Journal on Control and Optimization, 61 (4):1955–1978, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1512600.

Piccoli:2019:SCH

[PDT19] Benedetto Piccoli, Nastassia Pouradier Duteil, and Emmanuel Trélat. Sparse control of Hegselmann–Krause models: Black hole and declustering. SIAM Journal on Control and Optimization, 57(4): 2628–2659, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Polyakov:2019:CDF

[PEB19] Andrey Polyakov, Denis Efimov, and Bernard Brogliato. Consistent discretization of finite-time and fixed-time stable systems. SIAM Journal on Control and Optimization, 57 (1):78–103, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pedersen:1991:PPS

[Ped91] Michael Pedersen. Pseudodifferential perturbations and stabilization of distributed parameter systems: Dirichlet feedback control problems. SIAM Journal on Control and Optimization, 29(1): 222–252, January 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pedregal:1998:ECY

[Ped98]  Pablo Pedregal.  Equilib-

rium conditions for Young measures. SIAM Journal on Control and Optimization, 36(3):797–813, May 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30080.

Peet:2020:CSO

[Pee20] Matthew M. Peet. A convex solution of the H_{∞}- Optimal controller synthesis problem for multidelay systems. SIAM Journal on Control and Optimization, 58(3): 1547–1578, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pakshin:2018:STD

[PEGR18] Pavel Pakshin, Julia Emelianova, Krzysztof Gałkowski, and Eric Rogers. Stabilization of twodimensional nonlinear systems described by Fornasini– Marchesini and Roesser models. SIAM Journal on Control and Optimization, 56(5):

3848–3866, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pelletier:2000:AAS

[Pel00] Mariane Pelletier. Asymptotic almost sure efficiency of averaged stochastic algorithms. SIAM Journal on Control and Optimization, 39(1):49–72, January 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/30816.

Penot:1984:DRD

[Pen84] Jean-Paul Penot. Differentiability of relations and differential stability of perturbed optimization problems. SIAM Journal on Control and Optimization, 22(4):529–551, July 1984. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See erratum [Pen88].

Penot:1988:EDR

[Pen88] Jean-Paul Penot. Erratum: “Differentiability of relations and differential stability of perturbed optimization problems” [SIAM J. Control Optim. 22 (1984), no. 4, 529–551; MR 85i:49041]. SIAM Journal on Control and Optimization, 26(4):997–998, July 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [Pen84].

Peng:1989:AAP

[Pen89]  Shi Ge Peng.  Analyse

asymptotique et problème homogénéisé en contrôle optimal avec vibrations rapides. (French) [Asymptotic analysis and a homogenized problem in optimal control with rapid vibrations]. SIAM Journal on Control and Optimization, 27 (4):673–696, July 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Peng:1990:GSM

[Pen90] Shi Ge Peng. A general stochastic maximum principle for optimal control problems. SIAM Journal on Control and Optimization, 28(4):966–979, July 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Peng:1992:SHJ

[Pen92]  Shi Ge Peng.  Stochas-

tic Hamilton–Jacobi–Bellman equations. SIAM Journal on Control and Optimization, 30 (2):284–304, March 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Penot:1998:SOC

[Pen98] Jean-Paul Penot. Secondorder conditions for optimization problems with constraints. SIAM Journal on Control and Optimization, 37 (1):303–318, January 1998.

CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/31109.

Penot:2013:SRS

[Pen13] Jean-Paul Penot. Semigroups of relations and subdifferentials of the minimal time function and of the geodesic distance. SIAM Journal on Control and Optimization, 51(4): 2839–2868, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Peng:2020:OOC

[Pen20] Zijia Peng. Optimal obstacle control problems involving nonsmooth cost functionals and quasilinear variational inequalities. SIAM Journal on Control and Optimization, 58(4):2236–2255, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pepe:2014:SSH

[Pep14] P. Pepe. Stabilization in the sample-and-hold sense of nonlinear retarded systems. SIAM Journal on Control and Optimization, 52(5):3053–3077, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pepe:2016:SPU

[Pep16] P. Pepe. On stability preservation under sampling and approximation of feedbacks for retarded systems. SIAM Journal on Control and Optimization, 54(4):1895–1918, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Perold:1981:EPB

[Per81] André F. Perold. Extreme points and basic feasible solutions in continuous time linear programming. SIAM Journal on Control and Optimization, 19(1):52–63, January 1981. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Perrollaz:2012:ECS

[Per12] Vincent Perrollaz. Exact controllability of scalar conservation laws with an additional control in the context of entropy solutions. SIAM Journal on Control and Optimization, 50(4):2025–2045, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Petersen:1985:SSU

[Pet85] Ian R. Petersen. Structural stabilization of uncertain systems: Necessity of the matching condition. SIAM Journal on Control and Optimization, 23(2):286–296, March 1985. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Petersen:1988:SUL

[Pet88] Ian R. Petersen. Stabilization of an uncertain linear system in which uncertain parameters enter into the input matrix. SIAM Journal on Control and Optimization, 26(6):1257– 1264, November 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Petschke:1990:TAB

[Pet90] Martin Petschke. On a theorem of Arrow, Barankin, and Blackwell. SIAM Journal on Control and Optimization, 28 (2):395–401, March 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Petersen:2002:NOU

[Pet02] Ian R. Petersen. Notions of observability for uncertain linear systems with structured uncertainty. SIAM Journal on Control and Optimization, 41(2):345–361, March 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/36807.

Petersen:2004:NSF

[Pet04] Mark A. Petersen. Nonsquare spectral factors of nonlinear control systems in terms of inner-inner factorizations. SIAM Journal on Control and Optimization, 43(3):831–855, May 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40956.

Pflug:1986:SMC

[Pfl86] Georg Ch. Pflug. Stochastic minimization with constant step-size: Asymptotic laws. SIAM Journal on Control and Optimization, 24(4):655–666, July 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pu:2018:ESF

[PG18a] Shi Pu and Alfredo Garcia. Erratum: Swarming for Faster Convergence in Stochastic Optimization. SIAM Journal on Control and Optimization, 56 (6):4488–4492, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [PG18b].

Pu:2018:SFC

[PG18b] Shi Pu and Alfredo Garcia.

Swarming for faster convergence in stochastic optimization. SIAM Journal on Control and Optimization, 56(4): 2997–3020, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See erratum [PG18a].

Pham:2014:IDP

[PGB14] Van Thang Pham, Didier Georges, and Gildas Besançon. Infinite-dimensional predictive control for hyperbolic systems. SIAM Journal on Control and Optimization, 52(6):3592–3617, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Polak:1992:RPD

[PH92] Elijah Polak and Li Min He. Rate-preserving discretization strategies for semi-infinite programming and optimal control. SIAM Journal on Control and Optimization, 30(3): 548–572, May 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pham:2000:DHD

[Pha00] Huyên Pham. Dynamic L_{p}- hedging in discrete time under cone constraints. SIAM Journal on Control and Optimization, 38(3):665–682, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34109.

Phelps:1985:MPG

[Phe85] R. R. Phelps. Metric projections and the gradient projection method in Banach spaces. SIAM Journal on Control and Optimization, 23(6): 973–977, November 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Phelps:1986:GPM

[Phe86] R. P. Phelps. The gradient projection method using Curry’s steplength. SIAM Journal on Control and Optimization, 24(4):692–699, July 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Phelps:1991:CNO

[Phe91] Andrew R. Phelps. On constructing nonlinear observers. SIAM Journal on Control and Optimization, 29(3):516–534, May 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Philpott:1994:CTS

[Phi94] A. B. Philpott. Continuoustime shortest path problems and linear programming. SIAM Journal on Control and Optimization, 32(2):538–552, March 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/19641.

Pauwels:2016:LCO

[PHL16] Edouard Pauwels, Didier Henrion, and Jean-Bernard Lasserre. Linear conic optimization for inverse optimal control. SIAM Journal on Control and Optimization, 54 (3):1798–1825, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Phung:2001:OCS

[Phu01] K. D. Phung. Observability and control of Schrödinger equations. SIAM Journal on Control and Optimization, 40 (1):211–230, January 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/36840.

Peters:1999:STO

[PI99] M. A. Peters and P. A. Iglesias. A spectral test for observability and reachability of time-varying systems. SIAM Journal on Control and Optimization, 37(5): 1330–1345, September 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32182.

Picard:1993:EQV

[Pic93]  Jean Picard.  Estimation

of the quadratic variation of nearly observed semimartingales with application to filtering. SIAM Journal on Control and Optimization, 31(2): 494–517, March 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Piccoli:1996:CGS

[Pic96] B. Piccoli. Classification of generic singularities for the planar time-optimal synthesis. SIAM Journal on Control and Optimization, 34(6): 1914–1946, November 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25614.

Pierce:1979:UIE

[Pie79] Alan Pierce. Unique identification of eigenvalues and coefficients in a parabolic problem. SIAM Journal on Control and Optimization, 17(4): 494–499, July 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pipeleers:2014:GKL

[PIH14] Goele Pipeleers, Tetsuya Iwasaki, and Shinji Hara. Generalizing the KYP lemma to multiple frequency intervals. SIAM Journal on Control and Optimization, 52(6): 3618–3638, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Piunovskiy:2015:RRS

[Piu15] Alexey Piunovskiy. Randomized and relaxed strategies in continuous-time Markov decision processes. SIAM Journal on Control and Optimization, 53(6):3503–3533, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Piunovskiy:2018:RSC

[Piu18] Alexey Piunovskiy. Realizable strategies in continuoustime Markov decision processes. SIAM Journal on Control and Optimization, 56(1): 473–495, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Polyak:1992:ASA

[PJ92] B. T. Polyak and A. B. Juditsky. Acceleration of stochastic approximation by averaging. SIAM Journal on Control and Optimization, 30(4): 838–855, July 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Jiao:2020:NCF

[pJY20] Xiao pei Jiao and Stephen S.- T. Yau. New classes of finite dimensional filters with nonmaximal rank estimation algebra on state dimension n and linear rank n−2. SIAM Journal on Control and Optimization, 58(6):3413–3427, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Poolla:1987:SSP

[PK87] Kameshwar Poolla and Pramod Khargonekar. Stabilizability and stable-proper factorizations for linear time-varying systems. SIAM Journal on Control and Optimization, 25 (3):723–736, May 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pomet:1995:FEP

[PK95] J.-B. Pomet and I. A. K.

Kupka. On feedback equivalence of a parameterized family of nonlinear systems. SIAM Journal on Control and Optimization, 33(4):1170–1207, July 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/22442.

Protasov:2023:SCT

[PK23] Vladimir Yu. Protasov and Rinat Kamalov. Stability of continuous time linear systems with bounded switching intervals. SIAM Journal on Control and Optimization, 61(5):3051–3075, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1529579.

Pan:2013:ESC

[PKL13] Yaodong Pan, Krishna Dev Kumar, and Guangjun Liu. Extremum seeking control with second-order sliding mode. SIAM Journal on Control and Optimization, 50(6): 3292–3309, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Perez:2015:TRS

[PL15] Iker Perez and Huiling Le. Time-randomized stopping problems for a family of utility functions. SIAM Journal on Control and Optimization, 53(3):1328–1345, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Platzman:1980:OIH

[Pla80] Loren K. Platzman. Optimal infinite-horizon undiscounted control of finite probabilistic systems. SIAM Journal on Control and Optimization, 18 (4):362–380, ???? 1980. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Plaksin:2021:VSH

[Pla21] Anton Plaksin. Viscosity solutions of Hamilton–Jacobi– Bellman–Isaacs equations for time-delay systems. SIAM Journal on Control and Optimization, 59(3):1951–1972, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Praly:1989:RAM

[PLK89] L. Praly, S.-F. Lin, and P. R. Kumar. A robust adaptive minimum variance controller. SIAM Journal on Control and Optimization, 27(2): 235–266, March 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Passenberg:2014:GCL

[PLSB14] Benjamin Passenberg, Marion Leibold, Olaf Stursberg, and Martin Buss. A globally convergent, locally optimal min-H algorithm for hybrid optimal control. SIAM Journal on Control and Optimization, 52 (1):718–746, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Polak:1985:AMN

[PM85] E. Polak and D. Q. Mayne. Algorithm models for nondifferentiable optimization. SIAM Journal on Control and Optimization, 23(3):477–491, May 1985. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Puiggali:2016:IPC

[PM16] M. Farré Puiggalí and T. Mestdag.

The inverse problem of the calculus of variations and the stabilization of controlled Lagrangian systems. SIAM Journal on Control and Optimization, 54(6):3297–3318, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Polak:1992:ENM

[PMH92] E. Polak, D. Q. Mayne, and J. E. Higgins. On the extension of Newton’s method to semi-infinite minimax problems. SIAM Journal on Control and Optimization, 30(2):

367–389, March 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Priscoli:2006:NAA

[PMI06] F. Delli Priscoli, L. Marconi, and A. Isidori. A new approach to adaptive nonlinear regulation. SIAM Journal on Control and Optimization, 45(3):829–855, January 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Polak:1983:ECO

[PMW83] E. Polak, D. Q. Mayne, and Y. Wardi. On the extension of constrained optimization algorithms from differentiable to nondifferentiable problems. SIAM Journal on Control and Optimization, 21 (2):179–203, March 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Padula:2020:GCD

[PNSL20] Fabrizio Padula, Lorenzo Ntogramatzidis, Robert Schmid, and Ryan Loxton. Geometric control and disturbance decoupling for fractional systems. SIAM Journal on Control and Optimization, 58(3): 1403–1428, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Portier:2000:NEA

[PO00] Bruno Portier and Abderrahim Oulidi. Nonparametric estimation and adaptive control of functional autoregressive models. SIAM Journal on Control and Optimization, 39(2):411–432, March 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31667.

Poggiolini:2017:SSB

[Pog17] Laura Poggiolini. Structural stability of bang-bang trajectories with a double switching time in the minimum time problem. SIAM Journal on Control and Optimization, 55 (6):3779–3798, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Polyak:1988:SOM

[Pol88] R. A. Polyak. Smooth optimization methods for minimax problems. SIAM Journal on Control and Optimization, 26(6):1274–1286, November 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pomet:2009:NCD

[Pom09] Jean-Baptiste Pomet. A necessary condition for dynamic equivalence. SIAM Journal on Control and Optimization, 48 (2):925–940, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pongvuthithum:2009:TVF

[Pon09] R. Pongvuthithum. A timevarying feedback approach for triangular systems with nonlinear parameterization. SIAM Journal on Control and Optimization, 48(3):1660–1674, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Poore:1996:BPS

[Poo96] Aubrey B. Poore. Bifurcation problems for some parametric nonlinear programs in Banach spaces. SIAM Journal on Control and Optimization, 34(6):1947–1971, November 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/27716.

Popa:1995:FLN

[Pop95] Cătălin Popa. Feedback laws for nonlinear distributed control problems via trotter-type product formulae. SIAM Journal on Control and Optimization, 33(4):971–999, July 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/24256.

Popa:1997:RBM

[Pop97] Cătălin Popa. The relationship between the maximum principle and dynamic programming for the control of parabolic variational inequalities. SIAM Journal on Control and Optimization, 35(5): 1711–1738, September 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/19858.

Porter:1977:CRI

[Por77] William A. Porter. Causal realization from input-output pairs. SIAM Journal on Control and Optimization, 15(1): 120–128, January 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pospisil:2017:RCN

[Pos17] M. Pospísil. Relative controllability of neutral differential equations with a delay. SIAM Journal on Control and Optimization, 55(2): 835–855, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pisano:2011:TCU

[POU11] Alessandro Pisano, Yury Orlov, and Elio Usai. Tracking control of the uncertain heat and wave equation via powerfractional and sliding-mode techniques. SIAM Journal on Control and Optimization, 49 (2):363–382, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i2/p363_s1.

Poggi:2000:NAT

[PP00] Jean-Michel Poggi and Bruno Portier. Nonlinear adaptive tracking using kernel estimators: Estimation and test for linearity. SIAM Journal on Control and Optimization, 39(3):707–727, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34961.

Pironneau:2002:CAA

[PP02] Olivier Pironneau and Elijah Polak. Consistent approximations and approximate functions and gradients in optimal control. SIAM Journal on Control and Optimization, 41(2):487–510, March 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36959.

Paunonen:2010:IMT

[PP10] L. Paunonen and S. Pohjolainen. Internal model theory for distributed parameter systems. SIAM Journal on Control and Optimization, 48(7):

4753–4775, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pal:2013:RFD

[PP13] Debasattam Pal and Harish K.

Pillai. Representation formulae for discrete 2D autonomous systems. SIAM Journal on Control and Optimization, 51(3):2406–2441, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Paunonen:2014:IMP

[PP14a] Lassi Paunonen and Seppo Pohjolainen. The internal model principle for systems with unbounded control and observation. SIAM Journal on Control and Optimization, 52(6): 3967–4000, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pennanen:2014:DCP

[PP14b] Teemu Pennanen and Ari-Pekka Perkkiö. Duality in convex problems of Bolza over functions of bounded variation. SIAM Journal on Control and Optimization, 52(3): 1481–1498, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pedersen:2018:CDO

[PP18] J. L. Pedersen and G. Peskir.

Constrained dynamic optimality and binomial terminal wealth. SIAM Journal on Control and Optimization, 56(2):1342–1357, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Peet:2009:PFS

[PPL09] Matthew M. Peet, Antonis Papachristodoulou, and Sanjay Lall. Positive forms and stability of linear time-delay systems. SIAM Journal on Control and Optimization, 47(6): 3237–3258, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Piunovskiy:2019:OIC

[PPTZ19] Alexey Piunovskiy, Alexander Plakhov, Delfim F. M. Torres, and Yi Zhang. Optimal impulse control of dynamical systems. SIAM Journal on Control and Optimization, 57(4): 2720–2752, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Puerta:2006:GSC

[PPZ06] F. Puerta, X. Puerta, and I. Zaballa. On the geometry of the solutions of the cover problem. SIAM Journal on Control and Optimization, 45(2): 389–413, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pandolfi:2013:LOI

[PPZ13] Luciano Pandolfi, Enrico Priola, and Jerzy Zabczyk. Linear operator inequality and null controllability with vanishing energy for unbounded control systems. SIAM Journal on Control and Optimization, 51 (1):629–659, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Polak:1998:GSC

[PQ98] Elijah Polak and Liqun Qi.

Globally and superlinearly convergent algorithm for minimizing a normal merit function. SIAM Journal on Control and Optimization, 36(3):1005–1019, May 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30124.

Plaskacz:2000:VFD

[PQ00] Slawomir Plaskacz and Marc Quincampoix. Value-functions for differential games and control systems with discontinuous terminal cost. SIAM Journal on Control and Optimization, 39(5): 1485–1498, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34038.

Petit:2001:FHC

[PR01] Nicolas Petit and Pierre Rouchon. Flatness of heavy chain systems. SIAM Journal on Control and Optimization, 40(2):475–495, March 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/36863.

Prajna:2007:CPT

[PR07] Stephen Prajna and Anders Rantzer. Convex programs for temporal verification of nonlinear dynamical systems. SIAM Journal on Control and Optimization, 46(3):999–1021, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Perrollaz:2014:FTS

[PR14] Vincent Perrollaz and Lionel Rosier. Finite-time stabilization of 2 × 2 hyperbolic systems on tree-shaped networks. SIAM Journal on Control and Optimization, 52(1):143–163, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Preisig:1996:CMQ

[Pre96] J. C. Preisig. Copositivity and the minimization of quadratic functions with nonnegativity and quadratic equality constraints. SIAM Journal on Control and Optimization, 34(4):1135–1150, July 1996.

CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25189.

Phelps:2016:OCU

[PRG16] Chris Phelps, Johannes O.

Royset, and Qi Gong. Optimal control of uncertain systems using sample average approximations. SIAM Journal on Control and Optimization, 54(1):1–29, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Prieto-Rumeau:2006:BOC

[PRHL06] Tomás Prieto-Rumeau and Onésimo Hernández-Lerma. Bias optimality for continuoustime controlled Markov chains. SIAM Journal on Control and Optimization, 45(1):51– 73, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Prieto-Rumeau:2008:ECC

[PRHL08] Tomás Prieto-Rumeau and Onésimo Hernández-Lerma. Ergodic control of continuoustime Markov chains with pathwise constraints. SIAM Journal on Control and Optimization, 47(4):1888–1908, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Prieur:2005:ACR

[Pri05] Christophe Prieur. Asymptotic controllability and robust asymptotic stabilizability. SIAM Journal on Control and Optimization, 43(5): 1888–1912, September 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/38551.

Propst:1990:PLA

[Pro90] Georg Propst. Piecewise linear approximation for hereditary control problems. SIAM Journal on Control and Optimization, 28(1):70–96, January 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Protasov:2022:BNG

[Pro22] Vladimir Y. Protasov. The Barabanov norm is generically unique, simple, and easily computed. SIAM Journal on Control and Optimization, 60(4):2246–2267, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1426821.

Porfiri:2008:FSA

[PRS08] M. Porfiri, D. G. Roberson, and D. J. Stilwell. Fast switching analysis of linear switched systems using exponential splitting. SIAM Journal on Control and Optimization, 47(5):2582–2597, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pathiraja:2021:MVS

[PRS21] Sahani Pathiraja, Sebastian Reich, and Wilhelm Stannat. McKean–Vlasov SDEs in nonlinear filtering. SIAM Journal on Control and Optimization, 59(6):4188–4215, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Peichl:1986:CCB

[PS86] G. Peichl and W. Schappacher. Constrained controllability in Banach spaces. SIAM Journal on Control and Optimization, 24(6):1261– 1275, November 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pritchard:1987:LQC

[PS87] A. J. Pritchard and D. Salamon. The linear quadratic control problem for infinite dimensional systems with unbounded input and output operators. SIAM Journal on Control and Optimization, 25(1):121–144, January 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pontier:1988:FOR

[PS88] Monique Pontier and Jacques Szpirglas. Filtering with observations on a Riemannian symmetric space. SIAM Journal on Control and Optimization, 26(3):609–627, May 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pillai:1999:BAC

[PS99] Harish K. Pillai and Shiva Shankar. A behavioral approach to control of distributed systems. SIAM Journal on Control and Optimization, 37(2):388–408, March 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32178.

Piccoli:2000:RSS

[PS00] Benedetto Piccoli and Héctor J. Sussmann. Regular synthesis and sufficiency conditions for optimality. SIAM Journal on Control and Optimization, 39(2):359–410, March 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32203.

Pommier:2004:TAH

[PS04] Julien Pommier and Bessem Samet. The topological asymptotic for the Helmholtz equation with Dirichlet condition on the boundary of an arbitrarily shaped hole. SIAM Journal on Control and Optimization, 43(3):899–921, May 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41616.

Plotnikov:2006:DDS

[PS06] P. I. Plotnikov and J. Sokolowski.

Domain dependence of solutions to compressible Navier– Stokes equations. SIAM Journal on Control and Optimization, 45(4):1165–1197, January 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Palczewski:2010:FHO

[PS10a] Jan Palczewski and Łukasz Stettner. Finite horizon optimal stopping of timediscontinuous functionals with applications to impulse control with delay. SIAM Journal on Control and Optimization, 48(8):4874–4909, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Plotnikov:2010:SDD

[PS10b] P. I. Plotnikov and J. Sokolowski.

Shape derivative of drag functional. SIAM Journal on Control and Optimization, 48(7): 4680–4706, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Poggiolini:2011:SLO

[PS11] Laura Poggiolini and Marco Spadini. Strong local optimality for a bang–bang trajectory in a Mayer problem. SIAM Journal on Control and Optimization, 49(1): 140–161, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i1/p140_s1.

Poggiolini:2013:SSB

[PS13] Laura Poggiolini and Gianna Stefani. Structural stability for bang-singular-bang extremals in the minimum time problem. SIAM Journal on Control and Optimization, 51 (5):3511–3531, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Palczewski:2017:ICM

[PS17a] Jan Palczewski and Lukasz Stettner. Impulse control maximizing average cost per unit time: a nonuniformly ergodic case. SIAM Journal on Control and Optimization, 55(2): 936–960, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Paunonen:2017:AIS

[PS17b] Lassi Paunonen and David Seifert. Asymptotics for infinite systems of differential equations. SIAM Journal on Control and Optimization, 55 (2):1153–1178, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ponosov:2017:SDB

[PS17c] Arcady Ponosov and Eugene Stepanov. Smooth dynamics becomes hybrid in the limit. SIAM Journal on Control and Optimization, 55(1): 29–50, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Poggiolini:2018:SLO

[PS18] Laura Poggiolini and Gianna Stefani. Strong local optimality for a bang-bang-singular extremal: The fixed-free case. SIAM Journal on Control and Optimization, 56(3):2274– 2294, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Plotnikov:2019:BCM

[PS19] Pavel I. Plotnikov and Jan Sokołowski. Boundary control of the motion of a heavy piston in viscous gas. SIAM Journal on Control and Optimization, 57(5):3166–3192, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Park:2018:RNT

[PSJ18] Gyunghoon Park, Hyungbo Shim, and Youngjun Joo. Recovering nominal tracking performance in an asymptotic sense for uncertain linear systems. SIAM Journal on Control and Optimization, 56(2): 700–722, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pietrus:2018:HOD

[PSV18] Alain Pietrus, Teresa Scarinci, and Vladimir Veliov. High order discrete approximations to Mayer’s problems for linear systems. SIAM Journal on Control and Optimization, 56 (1):102–119, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Papadimitriou:1986:IPC

[PT86] Christos H. Papadimitriou and John Tsitsiklis. Intractable problems in control theory. SIAM Journal on Control and Optimization, 24(4):639–654, July 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Panier:1987:SCF

[PT87] Eliane R. Panier and André L. Tits. A superlinearly convergent feasible method for the solution of inequality constrained optimization problems. SIAM Journal on Control and Optimization, 25(4): 934–950, July 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Puel:1995:BSK

[PT95] Jean-Pierre Puel and Marius Tucsnak. Boundary stabilization for the von Kármán equations. SIAM Journal on Control and Optimization, 33 (1):255–273, January 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/22835.

Prieur:2006:QOR

[PT06] Christophe Prieur and Emmanuel Trélat. Quasi-optimal robust stabilization of control systems. SIAM Journal on Control and Optimization, 45 (5):1875–1897, ???? 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pantz:2008:PTH

[PT08] O. Pantz and K. Trabelsi. A post-treatment of the homogenization method for shape optimization. SIAM Journal on Control and Optimization, 47(3):1380–1398, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pedregal:2009:ERO

[PT09a] Pablo Pedregal and Jorge Tiago. Existence results for optimal control problems with some special nonlinear dependence on state and control. SIAM Journal on Control and Optimization, 48(2):415–437, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pola:2009:SMN

[PT09b] Giordano Pola and Paulo Tabuada. Symbolic models for nonlinear control systems: Alternating approximate bisimulations. SIAM Journal on Control and Optimization, 48 (2):719–733, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Philip:2013:PRT

[PT13] Peter Philip and Dan Tiba. A penalization and regularization technique in shape optimization problems. SIAM Journal on Control and Optimization, 51(6):4295–4317, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Panier:1988:FGC

[PTH88] Eliane R. Panier, André L. Tits, and José N. Herskovits. A QP -free, globally convergent, locally superlinearly convergent algorithm for inequality constrained optimization. SIAM Journal on Control and Optimization, 26(4):788–811, July 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pinnau:2017:QNL

[PTT17] René Pinnau, Claudia Totzeck, and Oliver Tse. The quasineutral limit in optimal semiconductor design. SIAM Journal on Control and Optimization, 55(4):2603–2635, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Praagman:2007:PAR

[PTY07] C. Praagman, H. L. Trentelman, and R. Zavala Yoe. On the parametrization of all regularly implementing and stabilizing controllers. SIAM Journal on Control and Optimization, 45(6):2035–2053, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Privat:2017:ADP

[PTZ17] Yannick Privat, Emmanuel Trélat, and Enrique Zuazua. Actuator design for parabolic distributed parameter systems with the moment method. SIAM Journal on Control and Optimization, 55(2):1128– 1152, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pfeiffer:2021:DAS

[PTZ21] Laurent Pfeiffer, Xiaolu Tan, and Yu-Long Zhou. Duality and approximation of stochastic optimal control problems under expectation constraints.

SIAM Journal on Control and Optimization, 59(5):3231– 3260, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pfaff:2015:OBC

[PU15] Sebastian Pfaff and Stefan Ulbrich. Optimal boundary control of nonlinear hyperbolic conservation laws with switched boundary data. SIAM Journal on Control and Optimization, 53(3):1250– 1277, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Puel:2009:NAD

[Pue09] Jean-Pierre Puel. A nonstandard approach to a data assimilation problem and Tychonov regularization revisited. SIAM Journal on Control and Optimization, 48(2): 1089–1111, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pullan:1993:ACC

[Pul93] Malcolm C. Pullan. An algorithm for a class of continuous linear programs. SIAM Journal on Control and Optimization, 31(6):1558–1577, November 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pullan:1995:FOS

[Pul95] Malcolm C. Pullan. Forms of optimal solutions for separated continuous linear programs. SIAM Journal on Control and Optimization, 33(6): 1952–1977, November 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/24785.

Pullan:1996:DTS

[Pul96] Malcolm C. Pullan. A duality theory for separated continuous linear programs. SIAM Journal on Control and Optimization, 34(3):931–965, May 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25750.

Pytlak:1998:FDA

[PV98] R. Pytlak and R. B. Vinter. A feasible directions algorithm for optimal control problems with state and control constraints: Convergence analysis. SIAM Journal on Control and Optimization, 36(6):1999–2019, November 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29764.

Penot:2004:SPM

[PV04] Jean-Paul Penot and Michel Volle. Surrogate programming and multipliers in quasi-convex programming. SIAM Journal on Control and Optimization, 42(6): 1994–2003, November 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32781.

Pieper:2013:PEA

[PV13] Konstantin Pieper and Boris Vexler. A priori error analysis for discretization of sparse elliptic optimal control problems in measure space. SIAM Journal on Control and Optimization, 51(4):2788–2808, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Palladino:2014:MRM

[PV14] M. Palladino and R. B. Vinter. Minimizers that are not also relaxed minimizers. SIAM Journal on Control and Optimization, 52(4):2164–2179, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Palladino:2015:RHA

[PV15] M. Palladino and R. B. Vinter. Regularity of the Hamiltonian along optimal trajectories. SIAM Journal on Control and Optimization, 53(4): 1892–1919, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pham:2009:OSM

[PVZ09] Huyên Pham, Vathana Ly Vath, and Xun Yu Zhou. Optimal switching over multiple regimes. SIAM Journal on Control and Optimization, 48 (4):2217–2253, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pritchard:1978:UCO

[PW78] A. J. Pritchard and A. Wirth.

Unbounded control and observation systems and their duality. SIAM Journal on Control and Optimization, 16(4): 535–545, July 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Polak:1984:SMS

[PW84] E. Polak and Y. Y. Wardi.

A study of minimizing sequences. SIAM Journal on Control and Optimization, 22 (4):599–609, July 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Peng:1999:FCF

[PW99] Shige Peng and Zhen Wu.

Fully coupled forward-backward stochastic differential equations and applications to optimal control. SIAM Journal on Control and Optimization, 37(3):825–843, May 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/31354.

Pillai:2002:LDD

[PW02] Harish K. Pillai and Jan C. Willems. Lossless and dissipative distributed systems. SIAM Journal on Control and Optimization, 40(5): 1406–1430, September 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/36802.

Pham:2017:DPO

[PW17] Huyên Pham and Xiaoli Wei. Dynamic programming for optimal control of stochastic McKean–Vlasov dynamics. SIAM Journal on Control and Optimization, 55(2): 1069–1101, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ponce:2018:OCN

[PW18] Augusto C. Ponce and Nicolas Wilmet. Optimal control of nonlinear elliptic problems with sparsity. SIAM Journal on Control and Optimization, 56(4):2513–2535, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Park:2023:RRR

[PW23] Kyunghyun Park and Hoi Ying Wong. Robust retirement with return ambiguity: Optimal G-stopping time in dual space. SIAM Journal on Control and Optimization, 61 (3):1009–1037, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1459150.

Pang:1995:GNM

[PY95] Jong-Shi Pang and Jen Chih Yao. On a generalization of a normal map and equation. SIAM Journal on Control and Optimization, 33 (1):168–184, January 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/24167.

Pritchard:1996:CFO

[PY96] A. J. Pritchard and Yuncheng You. Causal feedback optimal control for Volterra integral equations. SIAM Journal on Control and Optimization, 34(6):1874–1890, November 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27594.

Pages:2016:PCL

[PY16] Gilles Pagès and Jun Yu.

Pointwise convergence of the Lloyd I algorithm in higher dimension. SIAM Journal on Control and Optimization, 54 (5):2354–2382, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Perez:2020:OPR

[PYB20] José-Luis Pérez, Kazutoshi Yamazaki, and Alain Bensoussan. Optimal periodic replenishment policies for spectrally positive Lévy demand processes. SIAM Journal on Control and Optimization, 58(6): 3428–3456, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Polak:1993:MCB

[PYM93] E. Polak, T. H. Yang, and D. Q. Mayne. A method of centers based on barrier functions for solving optimal control problems with continuum state and control constraints. SIAM Journal on Control and Optimization, 31 (1):159–179, January 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pales:1994:NOP

[PZ94] Zs. Páles and V. M. Zeidan. Nonsmooth optimum problems with constraints. SIAM Journal on Control and Optimization, 32(5): 1476–1502, September 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/22965.

Priola:2003:NCV

[PZ03] Enrico Priola and Jerzy Zabczyk. Null controllability with vanishing energy. SIAM Journal on Control and Optimization, 42(3):1013–1032, May 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40997.

Pardoux:2004:QLD

[PZ04] Étienne Pardoux and Ofer Zeitouni. Quenched large deviations for one dimensional nonlinear filtering. SIAM Journal on Control and Optimization, 43(4):1272–1297, July 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36503.

Piunovskiy:2011:DCT

[PZ11] Alexey Piunovskiy and Yi Zhang.

Discounted continuous-time Markov decision processes with unbounded rates: The convex analytic approach. SIAM Journal on Control and Optimization, 49(5):2032– 2061, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i5/p2032_s1.

Porretta:2013:LTV

[PZ13] Alessio Porretta and Enrique Zuazua. Long time versus steady state optimal control. SIAM Journal on Control and Optimization, 51(6): 4242–4273, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Pham:2014:TPZ

[PZ14] Triet Pham and Jianfeng Zhang. Two person zerosum game in weak formulation and path dependent Bellman– Isaacs equation. SIAM Journal on Control and Optimization, 52(4):2090–2121, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Piunovskiy:2020:RCG

[PZ20] Alexey Piunovskiy and Yi Zhang. On reducing a constrained gradual-impulsive control problem for a jump Markov model to a model with gradual control only. SIAM Journal on Control and Optimization, 58 (1):192–214, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Piunovskiy:2022:GIC

[PZ22] Alexey Piunovskiy and Yi Zhang. Gradual-impulsive control for continuous-time Markov decision processes with total undiscounted costs and constraints: Linear programming approach via a reduction method. SIAM Journal on Control and Optimization, 60(3):1892–1917, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1444060.

Piunovskiy:2024:EOM

[PZ24] Alexey Piunovskiy and Yi Zhang.

Extreme occupation measures in Markov decision processes with an absorbing state. SIAM Journal on Control and Optimization, 62(1):65–90, January 2024. CODEN SJCODC. ISSN 1095-7138.

Peng:2024:IPB

[PZLM24a] Xiaoxiao Peng, Shijie Zhou, Wei Lin, and Xuerong Mao. Invariance principles for G-Brownian-motion-driven stochastic differential equations and their applications to Gstochastic control. SIAM Journal on Control and Optimization, 62(3):1569–1589, June 2024. CODEN SJCODC. ISSN 1095-7138.

Peng:2024:IPG

[PZLM24b] Xiaoxiao Peng, Shijie Zhou, Wei Lin, and Xuerong Mao. Invariance principles for G-Brownian-motion-driven stochastic differential equations and their applications to Gstochastic control. SIAM Journal on Control and Optimization, 62(3):1569–1589, June 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1564936.

Qais:2023:OSL

[QBP23] Imrul Qais, Chayan Bhawal, and Debasattam Pal. Optimal singular LQR problem: a PD feedback solution. SIAM Journal on Control and Optimization, 61(4):2655–2681, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1496414.

Qi:1995:GCS

[QC95] Li Qun Qi and Xiao Jun Chen. A globally convergent successive approximation method for severely nonsmooth equations. SIAM Journal on Control and Optimization, 33(2):402–418, March 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/23619.

Qian:1994:DGI

[Qia94] Xiaojun Qian. Differential games with information lags. SIAM Journal on Control and Optimization, 32(3):808–830, May 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/20237.

Qiu:2013:SVQ

[Qiu13]  Jing-Hui Qiu.  Set-valued

quasi-metrics and a general Ekeland’s variational principle in vector optimization. SIAM Journal on Control and Optimization, 51(2):1350–1371, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Qiu:2018:VSS

[Qiu18] Jinniao Qiu. Viscosity solutions of stochastic Hamilton– Jacobi–Bellman equations. SIAM Journal on Control and Optimization, 56(5):3708– 3730, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Qiu:2022:ODS

[QJL22] Ming Qiu, Zhuo Jin, and Shuanming Li. Optimal dividend strategies with reinsurance under contagious systemic risk. SIAM Journal on Control and Optimization, 60(3):1269–1293, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1422318.

Qian:2001:SFG

[QLD01] Chunjiang Qian, Wei Lin, and W. P. Dayawansa. Smooth feedback, global stabilization, and disturbance attenuation of nonlinear systems with uncontrollable linearization. SIAM Journal on Control and Optimization, 40(1):191–210, January 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37009.

Qureshi:1982:OID

[QN82] Z. H. Qureshi and T. S. Ng. Optimal input design for dynamic system parameter estimation: The D_{s}- optimality case. SIAM Journal on Control and Optimization, 20(5):713–721, September 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Quincampoix:2011:ELV

[QR11] Marc Quincampoix and Jérôme Renault. On the existence of a limit value in some nonexpansive optimal control problems. SIAM Journal on Control and Optimization, 49(5): 2118–2132, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i5/p2118_s1.

Qiao:2013:LTI

[QS13] Wei Qiao and Rifat Sipahi. A linear time-invariant consensus dynamics with homogeneous delays: Analytical study and synthesis of rightmost eigenvalues. SIAM Journal on Control and Optimization, 51(5):3971–3992, ???? 2013.

CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Qu:2014:CRF

[Qu14] Zheng Qu. Contraction of Riccati flows applied to the convergence analysis of a maxplus curse-of-dimensionality– free method. SIAM Journal on Control and Optimization, 52(5):2677–2706, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Quadrat:1980:ESA

[Qua80] Jean Pierre Quadrat. Existence de solution et algorithme de résolution numérique, de problème de contrôle optimal de diffusion stochastique dégénérée ou non. (French) [Existence of the solution and algorithm for the numerical solution of an optimal control problem of degenerate or nondegenerate stochastic diffusion]. SIAM Journal on Control and Optimization, 18 (2):199–226, March 1980. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Quadrat:2003:FRAa

[Qua03a] A. Quadrat. The fractional representation approach to synthesis problems: An algebraic analysis viewpoint Part I: (weakly) doubly coprime factorizations. SIAM Journal on Control and Optimization, 42(1):266–299, January 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/41712.

Quadrat:2003:FRAb

[Qua03b] A. Quadrat. The fractional representation approach to synthesis problems: An algebraic analysis viewpoint Part II: Internal stabilization. SIAM Journal on Control and Optimization, 42 (1):300–320, January 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/41713.

Quadrat:2004:GSS

[Qua04] A. Quadrat. On a general structure of the stabilizing controllers based on stable range. SIAM Journal on Control and Optimization, 42(6):2264–2285, November 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/40827.

Quincampoix:1992:DIT

[Qui92] Marc Quincampoix. Differential inclusions and target problems. SIAM Journal on Control and Optimization, 30(2): 324–335, March 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Quincampoix:2004:OCU

[QV04] Marc Quincampoix and Vladimir M.

Veliov. Optimal control of uncertain systems with incomplete information for the disturbances. SIAM Journal on Control and Optimization, 43(4):1373–1399, July 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42086.

Quincampoix:2013:MRS

[QV13] M. Quincampoix and V. M.

Veliov. Metric regularity and stability of optimal control problems for linear systems. SIAM Journal on Control and Optimization, 51(5): 4118–4137, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Qin:2018:EBM

[QW18] Shulin Qin and Gengsheng Wang. Equivalence between minimal time and minimal norm control problems for the heat equation. SIAM Journal on Control and Optimization, 56(2):981–1010, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Qui:2019:FSC

[QW19] Nguyen Thanh Qui and Daniel Wachsmuth. Full stability for a class of control problems of semilinear elliptic partial differential equations. SIAM Journal on Control and Optimization, 57(4):3021–3045, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Qin:2021:SPI

[QWY21a] Shulin Qin, Gengsheng Wang, and Huaiqiang Yu. Stabilization on periodic impulse control systems. SIAM Journal on Control and Optimization, 59(2):1136–1160, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Qin:2021:SPT

[QWY21b] Shulin Qin, Gengsheng Wang, and Huaiqiang Yu. Switching properties of time optimal controls for systems of heat equations coupled by constant matrices. SIAM Journal on Control and Optimization, 59(2): 1420–1442, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Qian:2013:ESC

[QZ13] Zhongmin Qian and Xun Yu Zhou. Existence of solutions to a class of indefinite stochastic Riccati equations. SIAM Journal on Control and Optimization, 51(1):221–229, ????

2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rapisarda:2016:SSM

[RA16] P. Rapisarda and A. C. Antoulas. State-space modeling of two-dimensional vectorexponential trajectories. SIAM Journal on Control and Optimization, 54(5):2734–2753, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Raitums:1999:WCS

[Rai99] Uldis Raitums. On the weak closure of sets of feasible states for linear elliptic equations in the scalar case. SIAM Journal on Control and Optimization, 37(4):1033–1047, July 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32432.

Ramakrishna:1994:CIS

[Ram94] Viswanath Ramakrishna. Controlled invariance for singular distributions. SIAM Journal on Control and Optimization, 32(3):790–807, May 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/19772.

Rampazzo:1998:DGU

[Ram98] Franco Rampazzo. Differential games with unbounded versus bounded controls. SIAM Journal on Control and Optimization, 36(3):814–839, May 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29460.

Rampazzo:2005:FRC

[Ram05] Franco Rampazzo. Faithful representations for convex Hamilton–Jacobi equations. SIAM Journal on Control and Optimization, 44(3):867–884, May 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43685.

Rao:1995:USH

[Rao95] Bopeng Rao. Uniform stabilization of a hybrid system of elasticity. SIAM Journal on Control and Optimization, 33(2):440–454, March 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23987.

Rao:1998:SEP

[Rao98] Bopeng Rao. Stabilization of elastic plates with dynamical boundary control. SIAM Journal on Control and Optimization, 36(1):148–163, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30097.

Rapisarda:2022:SLT

[Rap22]  P. Rapisarda.  State for

linear time-varying systems, with applications to dissipative systems. SIAM Journal on Control and Optimization, 60(1):147–167, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1351126.

Ravi:2000:ITQ

[Rav00] M. S. Ravi. Interpolation theory and quantum cohomology. SIAM Journal on Control and Optimization, 39(3):981–988, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/32029.

Ravindran:2020:PDB

[Rav20] Sivaguru S. Ravindran. Penalization of Dirichlet boundary control for nonstationary magneto-hydrodynamics. SIAM Journal on Control and Optimization, 58(4):2354– 2382, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Raymond:2006:FBS

[Ray06] Jean-Pierre Raymond. Feedback boundary stabilization of the two-dimensional Navier– Stokes equations. SIAM Journal on Control and Optimization, 45(3):790–828, January 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Raymond:2010:FSF

[Ray10] Jean-Pierre Raymond. Feedback stabilization of a fluidstructure model. SIAM Journal on Control and Optimization, 48(8):5398–5443, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Reddy:2022:SLD

[RBA22] Anugu Sumith Reddy, Amarjit Budhiraja, and Amit Apte. Some large deviation asymptotics in small noise filtering problems. SIAM Journal on Control and Optimization, 60(1):385–409, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M142616X.

Ramirez:2021:SAC

[RBS21] Adrián Ramírez, Dimitri Breda, and Rifat Sipahi. A scalable approach to compute delay margin of a class of neutral-type time delay systems. SIAM Journal on Control and Optimization, 59(2): 805–824, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Richert:2016:DLP

[RC16] Dean Richert and Jorge Cortés. Distributed linear programming with eventtriggered communication. SIAM Journal on Control and Optimization, 54(3):1769–1797, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rezaei:2012:OCU

[RCA12] Farzad Rezaei, Charalambos D. Charalambous, and N. U. Ahmed. Optimal control of uncertain stochastic systems subject to total variation distance uncertainty. SIAM Journal on Control and Optimization, 50(5):2683–2725, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Reinschke:2001:RCF

[RCS01] J. Reinschke, M. W. Cantoni, and M. C. Smith. A robust control framework for linear, time-invariant, spatially distributed systems. SIAM Journal on Control and Optimization, 40(2):610–627, March 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35705.

Rebhuhn:1977:SAN

[Reb77] D. Rebhuhn. On the set of attainability of nonlinear nonautonomous control systems. SIAM Journal on Control and Optimization, 15(5): 803–812, August 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rebhuhn:1978:SES

[Reb78] Deborah Rebhuhn. On the stability of the existence of singular controls under perturbation of the control system. SIAM Journal on Control and Optimization, 16(3): 463–472, May 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rebarber:1988:CFC

[Reb88] Richard Rebarber. Canonical forms for a class of distributed parameter control systems. SIAM Journal on Control and Optimization, 26(6):1362– 1387, November 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rebarber:1989:SAD

[Reb89] Richard Rebarber. Spectral assignability for distributed parameter systems with unbounded scalar control. SIAM Journal on Control and Optimization, 27(1):148–169, January 1989. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Rebarber:1995:ESC

[Reb95] Richard Rebarber. Exponential stability of coupled beams with dissipative joints: a frequency domain approach. SIAM Journal on Control and Optimization, 33(1):1–28, January 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24032.

Reddien:1979:CGP

[Red79] G. W. Reddien. Collocation at Gauss points as a discretization in optimal control. SIAM Journal on Control and Optimization, 17(2): 298–306, March 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Redmann:2018:TIB

[Red18a] Martin Redmann. Type II balanced truncation for deterministic bilinear control systems. SIAM Journal on Control and Optimization, 56(4): 2593–2612, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Redmann:2018:TIS

[Red18b] Martin Redmann. Type II singular perturbation approximation for linear systems with Lévy noise. SIAM Journal on Control and Optimization, 56(3):2120–2158, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Redmann:2021:BSN

[Red21] Martin Redmann. Bilinear systems — a new link to H_{∈}- norms, relations to stochastic systems, and further properties. SIAM Journal on Control and Optimization, 59(4): 2477–2497, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Reinoza:1983:GBG

[Rei83] A. Reinoza. Global behavior of generalized equations: a Sard theorem. SIAM Journal on Control and Optimization, 21 (3):443–450, May 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Reid:1986:OLC

[Rei86] Russell M. Reid. Open loop control of water waves in an irregular domain. SIAM Journal on Control and Optimization, 24(4):789–796, July 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Reid:1995:CTG

[Rei95] Russell M. Reid. Control time for gravity-capillary waves on water. SIAM Journal on Control and Optimization, 33(5):

1577–1586, September 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/22635.

Restrepo:1976:CGE

[Res76] Rodrigo A. Restrepo. Controls and goals in economic equilibrium. SIAM Journal on Control and Optimization, 14(1): 104–106, January 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rossi:2015:ART

[RFF15] Wilbert Samuel Rossi, Paolo Frasca, and Fabio Fagnani. Average resistance of toroidal graphs. SIAM Journal on Control and Optimization, 53 (4):2541–2557, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rakovic:2014:PIF

[RG14] S. V. Raković and R. H. Gielen. Positively invariant families of sets for interconnected and time-delay discrete-time systems. SIAM Journal on Control and Optimization, 52 (4):2261–2283, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rakhshan:2019:PFO

[RGG19] Mohsen Rakhshan, Vijay Gupta, and Bill Goodwine.

On passivity of fractional order systems. SIAM Journal on Control and Optimization, 57(2):1378–1389, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rishel:2006:VIS

[RH06] Raymond Rishel and Kurt Helmes. A variational inequality sufficient condition for optimal stopping with application to an optimal stock selling problem. SIAM Journal on Control and Optimization, 45(2):580–598, January 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ringh:2024:GST

[RHCK24] Axel Ringh, Isabel Haasler, Yongxin Chen, and Johan Karlsson. Graph-structured tensor optimization for nonlinear density control and mean field games. SIAM Journal on Control and Optimization, 62 (4):2176–2202, August 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1571587.

Richard:1977:OIC

[Ric77] Scott F. Richard. Optimal impulse control of a diffusion process with both fixed and proportional costs of control. SIAM Journal on Control and Optimization, 15(1):

79–91, January 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rifford:2000:ELS

[Rif00] Ludovic Rifford. Existence of Lipschitz and semiconcave control-Lyapunov functions. SIAM Journal on Control and Optimization, 39(4):1043–1064, July 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/35603.

Rifford:2002:SCL

[Rif02] Ludovic Rifford. Semiconcave control-Lyapunov functions and stabilizing feedbacks. SIAM Journal on Control and Optimization, 41(3):659–681, May 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37534.

Rindler:2009:OCN

[Rin09]  Filip Rindler.  Optimal

control for nonconvex rateindependent evolution processes. SIAM Journal on Control and Optimization, 47(6): 2773–2794, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rishel:1986:EFL

[Ris86] Raymond Rishel. An exact formula for a linear quadratic adaptive stochastic optimal control law. SIAM Journal on Control and Optimization, 24 (4):667–674, July 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rishel:1994:SSP

[Ris94] Raymond Rishel. A strong separation principle for stochastic control systems driven by a hidden Markov model. SIAM Journal on Control and Optimization, 32(4):1008–1020, July 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/23223.

Rivera:1990:ECC

[Riv90] Jaime E. Muñoz Rivera. Exact controllability: Coefficient depending on the time. SIAM Journal on Control and Optimization, 28(2):498– 501, March 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rahmani:2009:CMA

[RJME09] Amirreza Rahmani, Meng Ji, Mehran Mesbahi, and Magnus Egerstedt. Controllability of multi-agent systems from a graph-theoretic perspective. SIAM Journal on Control and Optimization, 48(1):162–186, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rotea:1989:SUS

[RK89] Mario A. Rotea and Pramod P.

Khargonekar. Stabilization of uncertain systems with norm bounded uncertainty — a control Lyapunov function approach. SIAM Journal on Control and Optimization, 27(6):1462–1476, November 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ringh:2016:MRC

[RKL16] Axel Ringh, Johan Karlsson, and Anders Lindquist. Multidimensional rational covariance extension with applications to spectral estimation and image compression. SIAM Journal on Control and Optimization, 54(4): 1950–1982, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ringh:2018:MRC

[RKL18] Axel Ringh, Johan Karlsson, and Anders Lindquist. Multidimensional rational covariance extension with approximate covariance matching. SIAM Journal on Control and Optimization, 56(2): 913–944, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rodas:1978:SCF

[RL78] Hernan Rivera Rodas and C. E. Langenhop. A sufficient condition for function space controllability of a linear neutral system. SIAM Journal on Control and Optimization, 16 (3):429–435, May 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rohloff:2006:VCI

[RL06] Kurt Rohloff and Stéphane Lafortune. The verification and control of interacting similar discrete-event systems. SIAM Journal on Control and Optimization, 45(2): 634–667, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rodriguez:2020:BFS

[RL20] Charlotte Rodriguez and Günter Leugering. Boundary feedback stabilization for the intrinsic geometrically exact beam model. SIAM Journal on Control and Optimization, 58(6):3533–3558, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ren:2023:VCL

[RLXS23] Wei Ren, Jingjie Li, Junlin Xiong, and Xi-Ming Sun. Vector control Lyapunov and barrier functions for safe stabilization of interconnected systems. SIAM Journal on Control and Optimization, 61(5): 3209–3233, October 2023. CO-DEN SJCODC. ISSN 1095-7138.

Radenkovic:1992:VSS

[RM92] Miloje Radenković and Anthony N. Michel. Verification of the self-stabilization mechanism in robust stochastic adaptive control using Lyapunov function arguments. SIAM Journal on Control and Optimization, 30(6):1270– 1294, November 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rathinam:1998:CFL

[RM98] Muruhan Rathinam and Richard M.

Murray. Configuration flatness of Lagrangian systems underactuated by one control. SIAM Journal on Control and Optimization, 36(1):164–179, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30098.

Rabi:2012:ASL

[RMB12] Maben Rabi, George V. Moustakides, and John S. Baras. Adaptive sampling for linear state estimation. SIAM Journal on Control and Optimization, 50(2):672–702, ???? 2012. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Rami:2002:ISL

[RMZ02] M. Ait Rami, J. B. Moore, and Xun Yu Zhou. Indefinite stochastic linear quadratic control and generalized differential Riccati equation. SIAM Journal on Control and Optimization, 40(4):1296–1311, July 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/37108.

Reveliotis:2013:OLS

[RN13] Spyros Reveliotis and Ahmed Nazeem. Optimal linear separation of the safe and unsafe subspaces of sequential resource allocation systems as a set-covering problem: Algorithmic procedures and geometric insights. SIAM Journal on Control and Optimization, 51(2):1707–1726, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ravi:1991:CLT

[RNK91] R. Ravi, K. M. Nagpal, and P. P. Khargonekar. H^{∞} control of linear time-varying systems: a state-space approach. SIAM Journal on Control and Optimization, 29(6):1394– 1413, November 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Robin:1981:SIC

[Rob81] Maurice Robin. On some impulse control problems with long run average cost. SIAM Journal on Control and Optimization, 19(3):333–358, May 1981. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rockafellar:1976:MOP

[Roc76] R. Tyrrell Rockafellar. Monotone operators and the proximal point algorithm. SIAM Journal on Control and Optimization, 14(5):877–898, August 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rockafellar:1987:LQP

[Roc87] R. T. Rockafellar. Linearquadratic programming and optimal control. SIAM Journal on Control and Optimization, 25(3):781–814, May 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rockafellar:1989:HTD

[Roc89] R. T. Rockafellar. Hamiltonian trajectories and duality in the optimal control of linear systems with convex costs. SIAM Journal on Control and Optimization, 27(5): 1007–1025, September 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rockafellar:1996:ESV

[Roc96] R. Tyrrell Rockafellar. Equivalent subgradient versions of Hamiltonian and Euler– Lagrange equations in variational analysis. SIAM Journal on Control and Optimization, 34(4):1300–1314, July 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27346.

Rode:1982:APF

[Rod82] Gerd Rodé. Asymptotic properties of a finite state continuous time Markov decision process. SIAM Journal on Control and Optimization, 20(6): 884–892, November 1982. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rodrigues:2021:OPE

[Rod21] Sérgio S. Rodrigues. Oblique projection exponential dynamical observer for nonautonomous linear parabolic-like equations. SIAM Journal on Control and Optimization, 59 (1):464–488, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rogers:2007:PSO

[Rog07] L. C. G. Rogers. Pathwise stochastic optimal control. SIAM Journal on Control and Optimization, 46(3):

1116–1132, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rosen:1984:DES

[Ros84a] I. G. Rosen. Difference equation state approximations for nonlinear hereditary control problems. SIAM Journal on Control and Optimization, 22 (2):302–326, March 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rosen:1984:DAM

[Ros84b] I. Gary Rosen. Discrete approximation methods for parameter identification in delay systems. SIAM Journal on Control and Optimization, 22 (1):95–120, January 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rosenblueth:1992:SWR

[Ros92a] Javier F. Rosenblueth. Strongly and weakly relaxed controls for time delay systems. SIAM Journal on Control and Optimization, 30(4):856–866, July 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rosenthal:1992:NRP

[Ros92b] Joachim Rosenthal. New results in pole assignment by real output feedback. SIAM Journal on Control and Optimization, 30(1):203–211, January 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rosenthal:1994:DFC

[Ros94] Joachim Rosenthal. On dynamic feedback compensation and compactification of systems. SIAM Journal on Control and Optimization, 32 (1):279–296, January 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/22133.

Rosenberg:2000:ZSA

[Ros00a] Dinah Rosenberg. Zero sum absorbing games with incomplete information on one side: Asymptotic analysis. SIAM Journal on Control and Optimization, 39(1):208–225, January 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35443.

Rosenblueth:2000:SCP

[Ros00b] Javier F. Rosenblueth. The solution to the conjecture on properness of weakly relaxed delayed controls. SIAM Journal on Control and Optimization, 39(2):352–358, March 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33300.

Rosier:2000:EBC

[Ros00c] Lionel Rosier. Exact boundary controllability for the linear Korteweg–de Vries equation on the half-line. SIAM Journal on Control and Optimization, 39(2):331–351, March 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35322.

Rothblum:1977:NMD

[Rot77a] Uriel G. Rothblum. Normalized Markov decision chains. II: Optimality of nonstationary policies. SIAM Journal on Control and Optimization, 15(2):221–232, February 1977. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rothblum:1977:OSF

[Rot77b] Uriel G. Rothblum. Orderable set functions and continuity. II: Set functions with infinitely many null points. SIAM Journal on Control and Optimization, 15(1):144–155, January 1977. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rotstein:1996:NPA

[Rot96] Héctor Rotstein. A Nevanlinna– Pick approach to time-domain constrained H_{∞} control. SIAM Journal on Control and Optimization, 34(4):1329– 1341, July 1996. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/24988.

Roubicek:1986:GSC

[Rou86] Tomáš Roubíček. Generalized solutions of constrained optimization problems. SIAM Journal on Control and Optimization, 24(5):951–960, July 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Roubicek:2021:FMC

[Rou21]  Tomás Roubícek.  Fine

metrizable convex relaxations of parabolic optimal control problems. SIAM Journal on Control and Optimization, 59 (2):1293–1311, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Royer:1989:RSA

[Roy89] G. Royer. A remark on simulated annealing of diffusion processes. SIAM Journal on Control and Optimization, 27(6):1403–1408, November 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Reid:1985:BCS

[RR85] Russell M. Reid and David L. Russell. Boundary control and stability of linear water waves. SIAM Journal on Control and Optimization, 23(1):

111–121, January 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ran:1992:SSR

[RR92] André C. M. Ran and Leiba Rodman. Stable solutions of real algebraic matrix Riccati equations. SIAM Journal on Control and Optimization, 30 (1):63–81, January 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rachev:1994:SST

[RR94] S. T. Rachev and L. Rüschendorf.

Solution of some transportation problems with relaxed or additional constraints. SIAM Journal on Control and Optimization, 32(3):673–689, May 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/22136.

Rao:2012:AAR

[RR12] Shodhan Rao and Paolo Rapisarda. An algebraic approach to the realization of lossless negative imaginary behaviors. SIAM Journal on Control and Optimization, 50 (3):1700–1720, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Respondek:2013:EMC

[RR13] Witold Respondek and Sandra Ricardo. Equivariants of mechanical control systems. SIAM Journal on Control and Optimization, 51(4): 3027–3055, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rabi:2016:SDE

[RRJ16] Maben Rabi, Chithrupa Ramesh, and Karl H. Johansson. Separated design of encoder and controller for networked linear quadratic optimal control. SIAM Journal on Control and Optimization, 54(2):662–689, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rosenfeld:2024:OKM

[RRKJ24] Joel A. Rosenfeld, Benjamin P. Russo, Rushikesh Kamalapurkar, and Taylor T. Johnson. The occupation kernel method for nonlinear system identification. SIAM Journal on Control and Optimization, 62(3):1643–1668, June 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/19M127029X.

Ramirez:2016:OFS

[RRR16] H. Ramírez, A. Rapaport, and V. Riquelme. Optimal feedback synthesis and minimal time function for the bioremediation of water resources with two patches. SIAM Journal on Control and Optimization, 54(3):1697–1718, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Roche:2014:SRD

[RRS14] Thomas Roche, Riccarda Rossi, and Ulisse Stefanelli. Stability results for doubly nonlinear differential inclusions by variational convergence. SIAM Journal on Control and Optimization, 52(2): 1071–1107, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ravi:1996:DPA

[RRW96] M. S. Ravi, Joachim Rosenthal, and Xiaochang Wang. Dynamic pole assignment and Schubert calculus. SIAM Journal on Control and Optimization, 34(3):813–832, May 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25270.

Runggaldier:1993:NOC

[RS93] Wolfgang J. Runggaldier and Łukasz Stettner. Nearly optimal controls for stochastic ergodic problems with partial observation. SIAM Journal on Control and Optimization, 31(1):180–218, January 1993.

CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rabah:2007:AEC

[RS07a] Rabah Rabah and Grigory M. Sklyar. The analysis of exact controllability of neutral-type systems by the moment problem approach. SIAM Journal on Control and Optimization, 46(6):2148–2181, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rothblum:2007:NAC

[RS07b] Uriel G. Rothblum and Robert Shorten. Nonlinear AIMD congestion control and contraction mappings. SIAM Journal on Control and Optimization, 46(5):1882–1896, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ratschan:2010:PBA

[RS10] Stefan Ratschan and Zhikun She. Providing a basin of attraction to a target region of polynomial systems by computation of Lyapunov-like functions. SIAM Journal on Control and Optimization, 48(7): 4377–4394, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Roth:2013:SAC

[RS13] Grégory Roth and William H. Sandholm. Stochastic approximations with constant step size and differential inclusions. SIAM Journal on Control and Optimization, 51(1):525–555, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Reis:2015:FCB

[RS15a] Timo Reis and Tilman Selig.

Funnel control for the boundary controlled heat equation. SIAM Journal on Control and Optimization, 53(1):547–574, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rocca:2015:ODC

[RS15b] E. Rocca and J. Sprekels. Optimal distributed control of a nonlocal convective Cahn– Hilliard equation by the velocity in three dimensions. SIAM Journal on Control and Optimization, 53(3):1654–1680, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rivas:2020:ICN

[RS20] Ivonne Rivas and Chenmin Sun. Internal controllability of nonlocalized solution for the Kadomtsev–Petviashvili II equation. SIAM Journal on Control and Optimization, 58 (3):1715–1734, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Renaud:2021:SCP

[RS21] Jean-François Renaud and Clarence Simard. A stochastic control problem with linearly bounded control rates in a Brownian model. SIAM Journal on Control and Optimization, 59(5):3103–3117, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Reddy:2020:AOC

[RSE20] Puduru Viswanadha Reddy, Johannes M. Schumacher, and Jacob C. Engwerda. Analysis of optimal control problems for hybrid systems with one state variable. SIAM Journal on Control and Optimization, 58(6):3262–3292, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ramirez:2017:AAT

[RSMG17] Adrián Ramírez, Rifat Sipahi, Sabine Mondié, and Rubén Garrido. An analytical approach to tuning of delaybased controllers for LTI– SISO systems. SIAM Journal on Control and Optimization, 55(1):397–412, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Radenkovic:2018:ESC

[RSS18] Miloje S. Radenković, Milos S. Stanković, and Srdjan S. Stanković. Extremum seeking control with twosided stochastic perturbations. SIAM Journal on Control and Optimization, 56(5):3766– 3783, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Roig-Solvas:2022:EDB

[RSS22] Biel Roig-Solvas and Mario Sznaier. Euclidean distance bounds for linear matrix inequalities analytic centers using a novel bound on the Lambert function. SIAM Journal on Control and Optimization, 60(2):720–731, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1349928.

Rosenberg:2004:SGS

[RSV04] Dinah Rosenberg, Eilon Solan, and Nicolas Vieille. Stochastic games with a single controller and incomplete information. SIAM Journal on Control and Optimization, 43(1):86–110, January 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40710.

Reisinger:2023:LCP

[RSZ23] Christoph Reisinger, Wolfgang Stockinger, and Yufei Zhang. Linear convergence of a policy gradient method for some finite horizon continuous time control problems. SIAM Journal on Control and Optimization, 61(6):3526–3558, November 2023. CODEN SJCODC. ISSN 1095-7138.

Rebarber:1998:RRD

[RT98] Richard Rebarber and Stuart Townley. Robustness with respect to delays for exponential stability of distributed parameter systems. SIAM Journal on Control and Optimization, 37(1):230–244, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/31245.

Rosch:2003:SSO

[RT03] A. Rösch and F. Tröltzsch. Sufficient second-order optimality conditions for a parabolic optimal control problem with pointwise controlstate constraints. SIAM Journal on Control and Optimization, 42(1):138–154, January 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/40326.

Rapisarda:2004:LHB

[RT04] P. Rapisarda and H. L. Trentelman. Linear Hamiltonian behaviors and bilinear differential forms. SIAM Journal on Control and Optimization, 43(3):769–791, May 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41466.

Rosch:2006:ERL

[RT06] A. Rösch and F. Tröltzsch.

Existence of regular Lagrange multipliers for a nonlinear elliptic optimal control problem with pointwise controlstate constraints. SIAM Journal on Control and Optimization, 45(2):548–564, January 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rosch:2007:RSL

[RT07] A. Rösch and F. Tröltzsch.

On regularity of solutions and Lagrange multipliers of optimal control problems for semilinear equations with mixed pointwise control-state constraints. SIAM Journal on Control and Optimization, 46 (3):1098–1115, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Radenkovic:2016:STC

[RT16] Miloje Radenkovic and Mohsen Tadi. Self-tuning consensus on directed graphs in the case of unknown control directions. SIAM Journal on Control and Optimization, 54(5): 2339–2353, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rodrigues:2021:FSD

[RTF21] C. C. Graciani Rodrigues, M. G. Todorov, and M. D. Fragoso. Fast switching detector-based H_{2} control of Markov jump linear systems with multiplicative noises. SIAM Journal on Control and Optimization, 59(6):4243– 4267, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Richard:2020:RSS

[RTT20] Alexandre Richard, Xiaolu Tan, and Nizar Touzi. On the root solution to the Skorokhod embedding problem given full marginals. SIAM Journal on Control and Optimization, 58(4):1874–1892, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ren:2023:EOP

[RTTY23] Zhenjie Ren, Xiaolu Tan, Nizar Touzi, and Junjian Yang. Entropic optimal planning for path-dependent mean field games. SIAM Journal on Control and Optimization, 61(3):1415–1437, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1484444.

Ren:2020:CVS

[RTZ20] Zhenjie Ren, Nizar Touzi, and Jianfeng Zhang. Comparison of viscosity solutions of semilinear path-dependent PDEs. SIAM Journal on Control and Optimization, 58(1):277–302, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rauls:2019:CBG

[RU19] Anne-Therese Rauls and Stefan Ulbrich. Computation of a Bouligand generalized derivative for the solution operator of the obstacle problem. SIAM Journal on Control and Optimization, 57(5): 3223–3248, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rauls:2021:CGD

[RU21] Anne-Therese Rauls and Stefan Ulbrich. On the characterization of generalized derivatives for the solution operator of the bilateral obstacle problem. SIAM Journal on Control and Optimization, 59(5): 3683–3707, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rubio:1995:GCN

[Rub95] J. E. Rubio. The global control of nonlinear diffusion equations. SIAM Journal on Control and Optimization, 33 (1):308–322, January 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/22288.

Rudge:1978:MAL

[Rud78] J. F. Rudge. Multi-agent linear-quadratic-Gaussian systems with “Poor” observations. SIAM Journal on Control and Optimization, 16 (5):832–846, September 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Russak:1976:RAM

[Rus76] I. Bert Russak. Relations among the multipliers for problems with bounded state constraints. SIAM Journal on Control and Optimization, 14(6):1151–1155, November 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Russell:1986:DNB

[Rus86a] D. L. Russell. The Dirichlet– Neumann boundary control problem associated with Maxwell’s equations in a cylindrical region. SIAM Journal on Control and Optimization, 24(2):199–229, March 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Russell:1986:FPE

[Rus86b] D. L. Russell. Frequency/ period estimation and adaptive rejection of periodic disturbances. SIAM Journal on Control and Optimization, 24(6):1276–1308, November 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rust:1988:MLE

[Rus88] John Rust. Maximum likelihood estimation of discrete control processes. SIAM Journal on Control and Optimization, 26(5):1006–1024, September 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Russell:2006:FSS

[Rus06] David L. Russell. Forward stabilized spline interpolation. SIAM Journal on Control and Optimization, 45(4):1314– 1328, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ruszczynski:2021:SSM

[Rus21]  Andrzej Ruszczyński.  A

stochastic subgradient method for nonsmooth nonconvex multilevel composition optimization. SIAM Journal on Control and Optimization, 59(3): 2301–2320, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rivera:1987:OCB

[RV87] P. H. Rivera and C. F. Vasconcellos. Optimal control for a backward parabolic system. SIAM Journal on Control and Optimization, 25(5): 1163–1172, September 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ratschek:1990:GOU

[RV90] H. Ratschek and R. L. Voller. Global optimization over unbounded domains. SIAM Journal on Control and Optimization, 28(3): 528–539, March 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rowland:1993:DOP

[RV93] J. D. L. Rowland and R. B. Vinter. Dynamic optimization problems with free-time and active state constraints. SIAM Journal on Control and Optimization, 31(3):677–697, May 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rampazzo:2000:DOC

[RV00] Franco Rampazzo and Richard Vinter. Degenerate optimal control problems with state constraints. SIAM Journal on Control and Optimization, 39(4):989–1007, July 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34022.

Rannacher:2005:PEE

[RV05] R. Rannacher and B. Vexler.

A priori error estimates for the finite element discretization of elliptic parameter identification problems with pointwise measurements. SIAM Journal on Control and Optimization, 44(5):1844–1863, September 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/61110.

Reis:2009:PPB

[RV09] Timo Reis and Elena Virnik.

Positivity preserving balanced truncation for descriptor systems. SIAM Journal on Control and Optimization, 48(4): 2600–2619, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Reis:2019:LQO

[RV19] Timo Reis and Matthias Voigt. Linear-quadratic optimal control of differentialalgebraic systems: The infinite time horizon problem with zero terminal state. SIAM Journal on Control and Optimization, 57(3):1567–1596, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rockafellar:1976:SCP

[RW76] R. T. Rockafellar and R. J. B. Wets. Stochastic convex programming: Relatively complete recourse and induced feasibility. SIAM Journal on Control and Optimization, 14 (3):574–589, May 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rockafellar:1978:ORP

[RW78] R. T. Rockafellar and R. J.- B. Wets. The optimal recourse problem in discrete time: L^{1}- multipliers for inequality constraints. SIAM Journal on Control and Optimization, 16 (1):16–36, January 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ramadge:1987:MFL

[RW87a] P. J. Ramadge and W. M. Wonham. Modular feedback logic for discrete event systems. SIAM Journal on Control and Optimization, 25(5): 1202–1218, September 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ramadge:1987:SCC

[RW87b] P. J. Ramadge and W. M. Wonham. Supervisory control of a class of discrete event processes. SIAM Journal on Control and Optimization, 25(1): 206–230, January 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Robinson:1987:STS

[RW87c] Stephen M. Robinson and Roger J.-B. Wets. Stability in two-stage stochastic programming. SIAM Journal on Control and Optimization, 25(6): 1409–1416, November 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rockafellar:1990:GLQ

[RW90] R. T. Rockafellar and R. J.- B. Wets. Generalized linearquadratic problems of deterministic and stochastic optimal control in discrete time. SIAM Journal on Control and Optimization, 28(4):810–822, July 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rosen:1992:CDR

[RW92] I. G. Rosen and C. Wang.

On the continuous dependence with respect to sampling of the linear quadratic regulator problem for distributed parameter systems. SIAM Journal on Control and Optimization, 30(4):942–974, July 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Russell:1994:GNC

[RW94] David L. Russell and George Weiss. A general necessary condition for exact observability. SIAM Journal on Control and Optimization, 32(1):1–23, January 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/19795.

Rapisarda:1997:SML

[RW97a] Paolo Rapisarda and J. C. Willems. State maps for linear systems. SIAM Journal on Control and Optimization, 35(3):1053–1091, May 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26841.

Rocha:1997:BCD

[RW97b] Paula Rocha and Jan C. Willems. Behavioral controllability of delay-differential systems. SIAM Journal on Control and Optimization, 35 (1):254–264, January 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28305.

Rockafellar:2000:CHJa

[RW00a] R. Tyrrell Rockafellar and Peter R. Wolenski. Convexity in Hamilton–Jacobi theory I: Dynamics and duality. SIAM Journal on Control and Optimization, 39(5): 1323–1350, September 2000.

CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34536.

Rockafellar:2000:CHJb

[RW00b] R. Tyrrell Rockafellar and Peter R. Wolenski. Convexity in Hamilton–Jacobi theory II: Envelope representations. SIAM Journal on Control and Optimization, 39(5): 1351–1372, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34537.

Rocha:2001:TCI

[RW01] P. Rocha and J. Wood. Trajectory control and interconnection of 1D and nD systems. SIAM Journal on Control and Optimization, 40 (1):107–134, January 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36279.

Rosch:2008:NVO

[RW08] Arnd Rösch and Daniel Wachsmuth. Numerical verification of optimality conditions. SIAM Journal on Control and Optimization, 47(5): 2557–2581, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Robertson:2015:LTB

[RX15] Scott Robertson and Hao Xing. Large time behavior of solutions to semilinear equations with quadratic growth in the gradient. SIAM Journal on Control and Optimization, 53(1):185–212, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ren:2019:KRS

[RX19] Wei Ren and Junlin Xiong. Krasovskii and Razumikhin stability theorems for stochastic switched nonlinear timedelay systems. SIAM Journal on Control and Optimization, 57(2):1043–1067, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ren:2021:SAS

[RX21] Wei Ren and Junlin Xiong. Stability analysis of stochastic impulsive switched systems with deterministic statedependent impulses and switches. SIAM Journal on Control and Optimization, 59(3):2068– 2092, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Radenkovic:1995:UPE

[RY95] Miloje S. Radenković and B. Erik Ydstie. Using persistent excitation with fixed energy to stabilize adaptive controllers and obtain hard bounds for the parameter estimation error. SIAM Journal on Control and Optimization, 33(4):1224–1246, July 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24030.

Ryan:1994:BCS

[Rya94] E. P. Ryan. On Brockett’s condition for smooth stabilizability and its necessity in a context of nonsmooth feedback. SIAM Journal on Control and Optimization, 32(6): 1597–1604, November 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/23543.

Ryan:1998:IIP

[Rya98] E. P. Ryan. An integral invariance principle for differential inclusions with applications in adaptive control. SIAM Journal on Control and Optimization, 36(3):960–980, May 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30170.

Russell:1993:CST

[RZ93] D. L. Russell and B. Y. Zhang.

Controllability and stabilizability of the third-order linear dispersion equation on a periodic domain. SIAM Journal on Control and Optimization, 31(3):659–676, May 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Raymond:1998:PPS

[RZ98] J. P. Raymond and H. Zidani. Pontryagin’s principle for state-constrained control problems governed by parabolic equations with unbounded controls. SIAM Journal on Control and Optimization, 36(6):1853–1879, November 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/30247.

Rosier:2006:GSG

[RZ06] Lionel Rosier and Bing-Yu Zhang. Global stabilization of the generalized Korteweg– de Vries equation posed on a finite domain. SIAM Journal on Control and Optimization, 45(3):927–956, January 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rosier:2009:LEC

[RZ09] Lionel Rosier and Bing-Yu Zhang. Local exact controllability and stabilizability of the nonlinear Schrödinger equation on a bounded interval. SIAM Journal on Control and Optimization, 48(2): 972–992, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ribeiro:2014:EMN

[RZ14] Ana Margarida Ribeiro and Elvira Zappale. Existence of minimizers for NonLevel convex supremal functionals. SIAM Journal on Control and Optimization, 52(5): 3341–3370, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Reisinger:2020:EEP

[RZ20] Christoph Reisinger and Yufei Zhang. Error estimates of penalty schemes for quasivariational inequalities arising from impulse control problems. SIAM Journal on Control and Optimization, 58(1): 243–276, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Reisinger:2021:RSF

[RZ21] Christoph Reisinger and Yufei Zhang. Regularity and stability of feedback relaxed controls. SIAM Journal on Control and Optimization, 59(5): 3118–3151, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Rzymowski:2009:PGL

[Rzy09] Witold Rzymowski. A problem of guarding line segment.

SIAM Journal on Control and Optimization, 48(2):812–838, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sepulchre:1996:SDI

[SA96] Rodolphe Sepulchre and Dirk Aeyels. Stabilizability does not imply homogeneous stabilizability for controllable homogeneous systems. SIAM Journal on Control and Optimization, 34(5):1798–1813, September 1996. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26730.

Sachs:1985:CRQ

[Sac85] Ekkehard Sachs. Convergence rates of quasi-Newton algorithms for some nonsmooth optimization problems. SIAM Journal on Control and Optimization, 23(3):401–418, May 1985. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sachkov:1997:POC

[Sac97] Yuri L. Sachkov. On positive orthant controllability of bilinear systems in small codimensions. SIAM Journal on Control and Optimization, 35(1):29–35, January 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27089.

Sacchelli:2019:SGL

[Sac19]  Ludovic Sacchelli.  Short

geodesics losing optimality in contact sub-Riemannian manifolds and stability of the 5dimensional caustic. SIAM Journal on Control and Optimization, 57(4):2362–2391, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sager:2013:SDO

[Sag13] Sebastian Sager. Sampling decisions in optimum experimental design in the light of Pontryagin’s Maximum Principle. SIAM Journal on Control and Optimization, 51(4): 3181–3207, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sagara:2015:IMN

[Sag15] Nobusumi Sagara. An indirect method of nonconvex variational problems in Asplund spaces: The case for saturated measure spaces. SIAM Journal on Control and Optimization, 53(1):336–351, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sagara:2017:RPN

[Sag17] Nobusumi Sagara. Relaxation and purification for nonconvex variational problems in dual Banach spaces: The minimization principle in saturated measure spaces. SIAM Journal on Control and Optimization, 55(5):3154–3170, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sakawa:1983:FSL

[Sak83] Yoshiyuki Sakawa. Feedback stabilization of linear diffusion systems. SIAM Journal on Control and Optimization, 21(5):667–676, September 1983. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sakawa:1984:FCS

[Sak84] Yoshiyuki Sakawa. Feedback control of second order evolution equations with damping. SIAM Journal on Control and Optimization, 22(3): 343–361, May 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sakamoto:2002:AHJ

[Sak02] Noboru Sakamoto. Analysis of the Hamilton–Jacobi equation in nonlinear control theory by symplectic geometry. SIAM Journal on Control and Optimization, 40(6): 1924–1937, November 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36280.

Salamon:1985:SSF

[Sal85]  Dietmar Salamon.  Struc-

ture and stability of finite dimensional approximations for functional differential equations. SIAM Journal on Control and Optimization, 23(6): 928–951, November 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Samet:2003:TAH

[SAM03] Bessem Samet, Samuel Amstutz, and Mohamed Masmoudi. The topological asymptotic for the Helmholtz equation. SIAM Journal on Control and Optimization, 42(5):1523–1544, September 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40680.

Susca:2014:SBR

[SAMB14] Sara Susca, Pushkarini Agharkar, Sonia Martínez, and Francesco Bullo. Synchronization of beads on a ring by feedback control. SIAM Journal on Control and Optimization, 52 (2):914–938, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sand:1996:RRC

[San96]  Jan-Åke Sand.  Recipro-

cal realizations on the circle. SIAM Journal on Control and Optimization, 34 (2):507–520, March 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26590.

Saporito:2019:SCDb

[Sap19] Yuri F. Saporito. Stochastic control and differential games with path-dependent influence of controls on dynamics and running cost. SIAM Journal on Control and Optimization, 57(2):1312–1327, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sarychev:1997:FSO

[Sar97] Andrei V. Sarychev. Firstand second-order sufficient optimality conditions for bang– bang controls. SIAM Journal on Control and Optimization, 35(1):315–340, January 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24619.

Sarrazin:2022:LDV

[Sar22] Clément Sarrazin. Lagrangian discretization of variational mean field games. SIAM Journal on Control and Optimization, 60(3):1365–1392, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1377291.

Sasane:2014:RGC

[Sas14] Amol Sasane. A refinement of the generalized chordal distance. SIAM Journal on Control and Optimization, 52(6): 3538–3555, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Saveliev:2005:ALN

[Sav05] Peter Saveliev. Applications of Lefschetz numbers in control theory. SIAM Journal on Control and Optimization, 44(5):1677–1690, September 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/44224.

Shreve:1978:ATF

[SB78] Steven E. Shreve and Dimitri P. Bertsekas. Alternative theoretical frameworks for finite horizon discretetime stochastic optimal control. SIAM Journal on Control and Optimization, 16(6): 953–978, November 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Schmitendorf:1980:NCL

[SB80] W. E. Schmitendorf and B. R.

Barmish. Null controllability of linear systems with constrained controls. SIAM Journal on Control and Optimization, 18(4):327–345, ????

1980. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shapiro:1992:SAP

[SB92] A. Shapiro and J. F. Bonnans. Sensitivity analysis of parametrized programs under cone constraints. SIAM Journal on Control and Optimization, 30(6):1409–1422, November 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shklyar:1994:OIP

[SB94] Benzion Shklyar and Victor Bakhmutsky. On one identification problem for distributed controllable systems. SIAM Journal on Control and Optimization, 32(6): 1782–1791, November 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/23869.

Scheicher:2015:GTD

[SBBV15] Martin Scheicher, Ingrid Blumthaler, Mauro Bisiacco, and Maria Elena Valcher. A generalized tracking and disturbance rejection problem for multidimensional behaviors. SIAM Journal on Control and Optimization, 53(3): 1375–1405, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Saldi:2018:MNE

[SBR18] Naci Saldi, Tamer Basar, and Maxim Raginsky. Markov– Nash equilibria in mean-field games with discounted cost. SIAM Journal on Control and Optimization, 56(6):4256– 4287, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Smyrnakis:2019:GTL

[SBT19] M. Smyrnakis, D. Bauso, and H. Tembine. Game-theoretic learning and allocations in robust dynamic coalitional games. SIAM Journal on Control and Optimization, 57(4): 2902–2923, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sanjari:2023:IPO

[SBY23] Sina Sanjari, Tamer Basar, and Serdar Yuksel. Isomorphism properties of optimality and equilibrium solutions under equivalent information structure transformations: Stochastic dynamic games and teams. SIAM Journal on Control and Optimization, 61(5):3102–3130, October 2023. CODEN SJCODC. ISSN 1095-7138.

Stanford:1980:CSM

[SC80] David P. Stanford and Luther T.

Conner, Jr. Controllability and stabilizability in multipair systems. SIAM Journal on Control and Optimization, 18(5):488–497, September 1980. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Stanford:1981:ACS

[SC81] David P. Stanford and Luther T. Conner, Jr. Addendum: “Controllability and stabilizability in multipair systems” [SIAM J. Control Optim. 18 (1980), no. 5, 488–497; MR 82b:93022]. SIAM Journal on Control and Optimization, 19 (5):708–709, September 1981. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sigalotti:2006:DPS

[SC06] Mario Sigalotti and Yacine Chitour. Dubins’ problem on surfaces II: Nonpositive curvature. SIAM Journal on Control and Optimization, 45(2): 457–482, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Schuresko:2012:DTR

[SC12] Michael Schuresko and Jorge Cortés. Distributed tree rearrangements for reachability and robust connectivity. SIAM Journal on Control and Optimization, 50(5):2588–2620, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sen:2016:MFG

[SC16a] Nevroz Sen and Peter E.

Caines. Mean field game theory with a partially observed major agent. SIAM Journal on Control and Optimization, 54(6):3174–3224, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sen:2016:NFT

[SC16b] Nevroz Sen and Peter E.

Caines. Nonlinear filtering theory for McKean–Vlasov type stochastic differential equations. SIAM Journal on Control and Optimization, 54 (1):153–174, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sen:2019:MFG

[SC19] Nevroz Sen and Peter E.

Caines. Mean field games with partial observation. SIAM Journal on Control and Optimization, 57(3):2064–2091, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shang:2023:UON

[SC23] Jun Shang and Tongwen Chen. Unilateral orthogonal nonnegative matrix factorization. SIAM Journal on Control and Optimization, 61 (4):2497–2519, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1508315.

Scarpa:2019:ODC

[Sca19] Luca Scarpa. Optimal distributed control of a stochastic Cahn–Hilliard equation. SIAM Journal on Control and Optimization, 57(5):3571– 3602, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Stavrou:2018:OEN

[SCCL18] Photios A. Stavrou, Themistoklis Charalambous, Charalambos D. Charalambous, and Sergey Loyka. Optimal estimation via nonanticipative rate distortion function and applications to timevarying Gauss–Markov processes. SIAM Journal on Control and Optimization, 56(5): 3731–3765, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Szczechla:1997:PSE

[SCFV97] W. W. Szczechla, S. A. Connell, J. A. Filar, and O. J. Vrieze. On the Puiseux series expansion of the limit discount equation of stochastic games. SIAM Journal on Control and Optimization, 35(3):860–875, May 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28413.

Schittkowski:1979:APM

[Sch79] K. Schittkowski. An adaptive precision method for nonlinear optimization problems. SIAM Journal on Control and Optimization, 17(1):82–98, January 1979. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Schmidt:1980:BBP

[Sch80a] E. J. P. Georg Schmidt. The “Bang–Bang” principle for the time-optimal problem in boundary control of the heat equation. SIAM Journal on Control and Optimization, 18 (2):101–107, March 1980. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Schmidt:1980:BCH

[Sch80b] E. J. P. Georg Schmidt.

Boundary control for the heat equation with steady-state targets. SIAM Journal on Control and Optimization, 18 (2):145–154, March 1980. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Schumacher:1983:DAC

[Sch83] J. M. Schumacher. A direct approach to compensator design for distributed parameter systems. SIAM Journal on Control and Optimization, 21(6):823–836, November 1983. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Schmidt:1986:EMS

[Sch86] E. J. P. Georg Schmidt. Even more states reachable by boundary control for the heat equation. SIAM Journal on Control and Optimization, 24(6):1319–1322, November 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Schattler:1988:LSTb

[Sch88a] Heinz Schättler. The local structure of time-optimal trajectories in dimension three under generic conditions. SIAM Journal on Control and Optimization, 26(4):899–918, July 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Schattler:1988:LSTa

[Sch88b] Heinz Schättler. On the local structure of time-optimal bang–bang trajectories in R^{3}. SIAM Journal on Control and Optimization, 26(1):186– 204, January 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Schanbacher:1989:APC

[Sch89] Tilman Schanbacher. Aspects of positivity in control theory. SIAM Journal on Control and Optimization, 27(3):

457–475, May 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Scherer:1992:CSF

[Sch92a] Carsten Scherer. H_{∞}-control by state-feedback for plants with zeros on the imaginary axis. SIAM Journal on Control and Optimization, 30(1): 123–142, January 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Scherer:1992:OAF

[Sch92b]  Carsten Scherer.  H_{∞}-

optimization without assumptions on finite or infinite zeros. SIAM Journal on Control and Optimization, 30(1): 143–166, January 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Schmidt:1992:MEC

[Sch92c] E. J. P. Georg Schmidt. On the modelling and exact controllability of networks of vibrating strings. SIAM Journal on Control and Optimization, 30(1):229–245, January 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Schnaubelt:2002:FNR

[Sch02] Roland Schnaubelt. Feedbacks for nonautonomous regular linear systems. SIAM Journal on Control and Optimization, 41(4):1141–1165, July 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39169.

Schiela:2009:SCO

[Sch09] Anton Schiela. State constrained optimal control problems with states of low regularity. SIAM Journal on Control and Optimization, 48(4): 2407–2432, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Scherer:2015:GSC

[Sch15a] Carsten W. Scherer. Gainscheduling control with dynamic multipliers by convex optimization. SIAM Journal on Control and Optimization, 53(3):1224–1249, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Schmidt:2015:OSC

[Sch15b] Klaus Werner Schmidt. Optimal supervisory control of discrete event systems: Cyclicity and interleaving of tasks. SIAM Journal on Control and Optimization, 53(3):1425– 1439, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Smyshlyaev:2010:BSD

[SCK10] Andrey Smyshlyaev, Eduardo Cerpa, and Miroslav Krstic. Boundary stabilization of a 1-D wave equation with indomain antidamping. SIAM Journal on Control and Optimization, 48(6):4014–4031, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Scherrer:1998:STL

[SD98] W. Scherrer and M. Deistler.

A structure theory for linear dynamic errors-in-variables models. SIAM Journal on Control and Optimization, 36(6):2148–2175, November 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26246.

Soleimani-damaneh:2010:NOU

[Sd10] M. Soleimani-damaneh. Nonsmooth optimization using Mordukhovich’s subdifferential. SIAM Journal on Control and Optimization, 48(5): 3403–3432, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Solis–Daun:2013:GCS

[SD13] Julio Solís-Daun. Global CLF stabilization of nonlinear systems. Part I: a geometric approach — compact strictly convex CVS. SIAM Journal on Control and Optimization, 51(3):2152–2175, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Solis-Daun:2015:GCS

[SD15] Julio Solís-Daun. Global CLF stabilization of nonlinear systems. Part II: an approximation approach — closed CVS. SIAM Journal on Control and Optimization, 53(2):645–669, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sun:2019:IRP

[SD19] Chuangchuang Sun and Ran Dai. An iterative rank penalty method for nonconvex quadratically constrained quadratic programs. SIAM Journal on Control and Optimization, 57(6):3749–3766, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Solis-Daun:2000:GSN

[SDSÁR00] Julio Solís-Daun, Rodolfo Suárez, and José Álvarez-Ramírez. Global stabilization of nonlinear systems with inputs subject to magnitude and rate bounds: a parametric optimization approach. SIAM Journal on Control and Optimization, 39(3):682–706, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33023.

Salapaka:1997:MOC

[SDV97] Murti V. Salapaka, Mohammed Dahleh, and Petros Voulgaris. Mixed objective control synthesis: Optimal ℓ_{1}/H_{2} control. SIAM Journal on Control and Optimization, 35(5):1672–1689, September 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28666.

Scherer:2021:CSA

[SE21] Carsten Scherer and Christian Ebenbauer. Convex synthesis of accelerated gradient algorithms. SIAM Journal on Control and Optimization, 59 (6):4615–4645, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Seidman:1977:OPH

[Sei77] Thomas I. Seidman. Observation and prediction for the heat equation. IV: Patch observability and controllability. SIAM Journal on Control and Optimization, 15(3):412–427, May 1977. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Seidman:1978:EBC

[Sei78] Thomas I. Seidman. Exact boundary control for some evolution equations.

SIAM Journal on Control and Optimization, 16(6):979– 999, November 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Seidman:1982:ROB

[Sei82] Thomas I. Seidman. Regularity of optimal boundary controls for parabolic equations. I. Analyticity. SIAM Journal on Control and Optimization, 20 (3):428–453, May 1982. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Seidman:1987:IRS

[Sei87] Thomas I. Seidman. Invariance of the reachable set under nonlinear perturbations. SIAM Journal on Control and Optimization, 25(5):1173– 1191, September 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Seierstad:1988:SCF

[Sei88] Atle Seierstad. Sufficient conditions in free final time optimal control problems. SIAM Journal on Control and Optimization, 26(1):155–167, January 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Serea:2003:RBP

[Ser03] Oana-Silvia Serea. On reflecting boundary problem for optimal control. SIAM Journal on Control and Optimization, 42(2):559–575, March 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/39593.

Serea:2009:RDG

[Ser09] Oana-Silvia Serea. Reflected differential games. SIAM Journal on Control and Optimization, 48(4):2516–2532, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Setiono:1993:IDL

[Set93] Rudy Setiono. Interior dual least 2-norm algorithm for linear programs. SIAM Journal on Control and Optimization, 31(4):875–899, July 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shustin:2008:RSS

[SFFC08] Eugenii Shustin, Leonid Fridman, Emilia Fridman, and Fernando Castaños. Robust semiglobal stabilization of the second order system by relay feedback with an uncertain variable time delay. SIAM Journal on Control and Optimization, 47(1):196–217, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Steeves:2019:CCPa

[SGM19a] Drew Steeves, Bahman Gharesifard, and Abdol-Reza Mansouri. Controllability of coupled parabolic systems with multiple underactuations, Part 1: Algebraic solvability. SIAM Journal on Control and Optimization, 57(5): 3272–3296, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Steeves:2019:CCPb

[SGM19b] Drew Steeves, Bahman Gharesifard, and Abdol-Reza Mansouri. Controllability of coupled parabolic systems with multiple underactuations, Part 2: Null controllability. SIAM Journal on Control and Optimization, 57(5): 3297–3321, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shen:2012:SDT

[SH12] Jinglai Shen and Jianghai Hu. Stability of discretetime switched homogeneous systems on cones and conewise homogeneous inclusions. SIAM Journal on Control and Optimization, 50(4):2216–2253, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shayman:1983:GARa

[Sha83a] Mark A. Shayman. Geometry of the algebraic Riccati equation, Part I. SIAM Journal on Control and Optimization, 21 (3):375–394, May 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shayman:1983:GARb

[Sha83b] Mark A. Shayman. Geometry of the algebraic Riccati equation, Part II. SIAM Journal on Control and Optimization, 21(3):395–409, May 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shapiro:1984:OCQ

[Sha84] Alexander Shapiro. On optimality conditions in quasidifferentiable optimization. SIAM Journal on Control and Optimization, 22(4):610–617, July 1984. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shayman:1985:PPM

[Sha85] Mark A. Shayman. On the phase portrait of the matrix Riccati equation arising from the periodic control problem. SIAM Journal on Control and Optimization, 23(5):717– 751, September 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shayman:1986:PPM

[Sha86] Mark A. Shayman. Phase portrait of the matrix Riccati equation. SIAM Journal on Control and Optimization, 24 (1):1–65, January 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shapiro:1988:SAN

[Sha88a] Alexander Shapiro. Sensitivity analysis of nonlinear programs and differentiability properties of metric projections. SIAM Journal on Control and Optimization, 26(3):628–645, May 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shayman:1988:HIF

[Sha88b] Mark A. Shayman. Homogeneous indices, feedback invariants and control structure theorem for generalized linear systems. SIAM Journal on Control and Optimization, 26(2): 387–400, March 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shah:1992:PEM

[Sha92] Jayant Shah. Properties of energy-minimizing segmentations. SIAM Journal on Control and Optimization, 30(1): 99–111, January 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shapiro:1994:SAP

[Sha94] Alexander Shapiro. Sensitivity analysis of parametrized programs via generalized equations. SIAM Journal on Control and Optimization, 32(2):553–571, March 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/23029.

Shankar:2001:LSB

[Sha01] Shiva Shankar. The lattice structure of behaviors. SIAM Journal on Control and Optimization, 39(6):1817–1832, ???? 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shaikhet:2010:SNA

[Sha10] L. Shaikhet. Some new aspects of Lyapunov-type theorems for stochastic differential equations of neutral type. SIAM Journal on Control and Optimization, 48(7): 4481–4499, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shankar:2014:HTG

[Sha14] Shiva Shankar. The hautus test and genericity results for controllable and uncontrollable behaviors. SIAM Journal on Control and Optimization, 52(1):32–51, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shao:2015:SSS

[Sha15] Jinghai Shao. Strong solutions and strong Feller properties for regime-switching diffusion processes in an infinite state space. SIAM Journal on Control and Optimization, 53(4): 2462–2479, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shao:2017:SRS

[Sha17] Jinghai Shao. Stabilization of regime-switching processes by feedback control based on discrete time observations. SIAM Journal on Control and Optimization, 55(2):724–740, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shao:2018:IME

[Sha18] Jinghai Shao. Invariant measures and Euler–Maruyama’s approximations of state-dependent regime-switching diffusions. SIAM Journal on Control and Optimization, 56(5):3215– 3238, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sheu:1985:SCE

[She85] Shuenn Jyi Sheu. Stochastic control and exit probabilities of jump processes. SIAM Journal on Control and Optimization, 23(2):306– 328, March 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shi:1988:OCS

[Shi88] Shu Zhong Shi. Optimal control of strongly monotone variational inequalities. SIAM Journal on Control and Optimization, 26(2):274– 290, March 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See erratum [Shi90].

Shi:1990:EOC

[Shi90] Shu Zhong Shi. Erratum:

“Optimal control of strongly monotone variational inequalities” [SIAM J. Control Optim. 26 (1988), no. 2, 274– 290; MR 89d:49009]. SIAM Journal on Control and Optimization, 28(1):243–249, January 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [Shi88].

Shi:2002:LHJ

[Shi02] Peng Shi. Limit Hamilton– Jacobi–Isaacs equations for singularly perturbed zerosum dynamic (discrete time) games. SIAM Journal on Control and Optimization, 41(3):826–850, May 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37908.

Shreve:1981:RBM

[Shr81]  Steven E. Shreve.  Re-

flected Brownian motion in the “Bang–Bang” control of Brownian drift. SIAM Journal on Control and Optimization, 19 (4):469–478, July 1981. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sklyar:2000:MAN

[SI00] G. M. Sklyar and S. Yu. Ignatovich. Moment approach to nonlinear time optimality. SIAM Journal on Control and Optimization, 38(6): 1707–1728, November 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32976.

Sklyar:2003:ATO

[SI03] G. M. Sklyar and S. Yu. Ignatovich. Approximation of time-optimal control problems via nonlinear power moment problems. SIAM Journal on Control and Optimization, 42 (4):1325–1346, January 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Swei:2000:QCS

[SIC00] Shan-Min Swei, Tetsuya Iwasaki, and Martin Corless. Quadratic controllability, strong controllability, and a related output feedback property. SIAM Journal on Control and Optimization, 39(5): 1373–1390, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/36510.

Simionescu:1976:OCP

[Sim76] Claudia Simionescu. Optimal control problems in Sobolev spaces with weights. SIAM Journal on Control and Optimization, 14(1):137–143, January 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sirbu:2014:MPC

[Sîr14a] Mihai Sîrbu. On martingale problems with continuoustime mixing and values of zerosum games without the Isaacs condition. SIAM Journal on Control and Optimization, 52 (5):2877–2890, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sirbu:2014:SPM

[Sîr14b] Mihai Sîrbu. Stochastic Perron’s method and elementary strategies for zero-sum differential games. SIAM Journal on Control and Optimization, 52(3):1693–1711, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sirbu:2015:APM

[Sîr15]  Mihai Sîrbu.  Asymptotic

Perron’s method and simple Markov strategies in stochastic games and control. SIAM Journal on Control and Optimization, 53(4):1713–1733, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shi:2013:RCC

[SJ13] Guodong Shi and Karl Henrik Johansson. Robust consensus for continuous-time multiagent dynamics. SIAM Journal on Control and Optimization, 51(5):3673–3691, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shaked:1986:MVC

[SK86] U. Shaked and P. R. Kumar. Minimum variance control of discrete time multivariable ARMAX systems. SIAM Journal on Control and Optimization, 24(3):396–411, May 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sivashankar:1994:CIN

[SK94] N. Sivashankar and Pramod P. Khargonekar. Characterization of the L_{∈}-induced norm for linear systems with jumps with applications to sampleddata systems. SIAM Journal on Control and Optimization, 32(4):1128–1150, July 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/22312.

Shayman:1995:SCN

[SK95] Mark A. Shayman and Ratnesh Kumar. Supervisory control of nondeterministic systems with driven events via prioritized synchronization and trajectory models. SIAM Journal on Control and Optimization, 33 (2):469–497, March 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/23960.

Sridhar:2023:MFA

[SK23] Anirudh Sridhar and Soummya Kar. Mean-field approximations for stochastic population processes with heterogeneous interactions. SIAM Journal on Control and Optimization, 61(6):3442–3466, November 2023. CODEN SJ-CODC. ISSN 1095-7138.

Salapaka:1999:SMP

[SKD99] Murti V. Salapaka, M. Khammash, and M. Dahleh. Solution of MIMO H_{∈}/ℓ_{∞} problem without zero interpolation. SIAM Journal on Control and Optimization, 37(6): 1865–1873, November 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/33257.

Saberi:1990:GSP

[SKS90] A. Saberi, P. V. Kokotović, and H. J. Sussmann. Global stabilization of partially linear composite systems. SIAM Journal on Control and Optimization, 28(6): 1491–1503, November 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Swenson:2017:RPF

[SKXL17] Brian Swenson, Soummya Kar, João Xavier, and David S. Leslie. Robustness properties in fictitious-play-type algorithms. SIAM Journal on Control and Optimization, 55 (5):3295–3318, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sengupta:1998:OCT

[SL98] Raja Sengupta and Stéphane Lafortune. An optimal control theory for discrete event systems. SIAM Journal on Control and Optimization, 36(2):488–541, March 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/26095.

Sarlette:2013:SPS

[SL13] A. Sarlette and C. Lageman.

Synchronization with partial state coupling on SO(n). SIAM Journal on Control and Optimization, 50(6):3242– 3268, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sarkans:2016:ISS

[SL16] E. Sarkans and H. Logemann. Input-to-state stability of discrete-time Lur’e systems. SIAM Journal on Control and Optimization, 54(3): 1739–1768, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shang:2018:CPS

[SL18] Yunxia Shang and Shumin Li.

Control properties for secondorder hyperbolic systems in anisotropic cases with applications in inhomogeneous and anisotropic elastodynamic systems. SIAM Journal on Control and Optimization, 56(6): 4181–4202, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shu:2022:SAA

[SL22] Yadong Shu and Bo Li. Stability and attractivity analysis of uncertain switched systems under optimistic value criterion. SIAM Journal on Control and Optimization, 60 (4):2420–2439, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1450343.

Slawig:2000:EFD

[Sla00] Thomas Slawig. An explicit formula for the derivative of a class of cost functionals with respect to domain variations in Stokes flow. SIAM Journal on Control and Optimization, 39(1):141–158, January 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34787.

Slawig:2003:FDR

[Sla03] Thomas Slawig. A formula for the derivative with respect to domain variations in Navier–Stokes flow based on an embedding domain method. SIAM Journal on Control and Optimization, 42(2):495–512, March 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38570.

Slemrod:1978:SBC

[Sle78] Marshall Slemrod. Stabilization of bilinear control systems with applications to nonconservative problems in elasticity. SIAM Journal on Control and Optimization, 16(1): 131–141, January 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shreve:1984:OCG

[SLG84] S. E. Shreve, J. P. Lehoczky, and D. P. Gaver. Optimal consumption for general diffusions with absorbing and reflecting barriers. SIAM Journal on Control and Optimization, 22 (1):55–75, January 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shin:2023:NOD

[SLQ^{+}23] Sungho Shin, Yiheng Lin, Guannan Qu, Adam Wierman, and Mihai Anitescu. Near-optimal distributed linearquadratic regulator for networked systems. SIAM Journal on Control and Optimization, 61(3):1113–1135, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1489836.

Sun:2016:OLC

[SLY16] Jingrui Sun, Xun Li, and Jiongmin Yong. Open-loop and closed-loop solvabilities for stochastic linear quadratic optimal control problems. SIAM Journal on Control and Optimization, 54(5):2274– 2308, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Said:2013:FOM

[SM13] Salem Said and Jonathan H. Manton. On filtering with observation in a manifold: Reduction to a classical filtering problem. SIAM Journal on Control and Optimization, 51 (1):767–783, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shubov:1997:ECD

[SMDB97] Marianna A. Shubov, Clyde F. Martin, Jerald P. Dauer, and Boris P. Belinskiy. Exact controllability of the damped wave equation. SIAM Journal on Control and Optimization, 35(5):1773–1789, September 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/29161.

Smith:1990:WPO

[Smi90] Malcolm C. Smith. Wellposedness of H^{∞} optimal control problems. SIAM Journal on Control and Optimization, 28(2):342–358, March 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Smirnov:1996:SCC

[Smi96] Georgi V. Smirnov. Stabilization by constrained controls. SIAM Journal on Control and Optimization, 34(5): 1616–1649, September 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26210.

Swenson:2018:BRD

[SMK18] Brian Swenson, Ryan Murray, and Soummya Kar. On bestresponse dynamics in potential games. SIAM Journal on Control and Optimization, 56(4): 2734–2767, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Srikantakumar:1982:LMR

[SN82] P. R. Srikantakumar and K. S.

Narendra. A learning model for routing in telephone networks. SIAM Journal on Control and Optimization, 20(1): 34–57, January 1982. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Savla:2009:MLR

[SNB09] Ketan Savla, Giuseppe Notarstefano, and Francesco Bullo. Maintaining limitedrange connectivity among second-order agents. SIAM Journal on Control and Optimization, 48(1):187–205, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Snyders:1977:SPD

[Sny77] Jakov Snyders. Stationary probability distributions for linear time-invariant systems. SIAM Journal on Control and Optimization, 15(3):428–437, May 1977. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sipahi:2006:CSA

[SO06a] Rifat Sipahi and Nejat Olgac. Complete stability analysis of neutral-type first order two-time-delay systems with cross-talking delays. SIAM Journal on Control and Optimization, 45(3):957–971, January 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sipahi:2006:SRR

[SO06b] Rifat Sipahi and Nejat Olgac. Stability robustness of retarded LTI systems with single delay and exhaustive determination of their imaginary spectra. SIAM Journal on Control and Optimization, 45(5): 1680–1696, ???? 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Scheicher:2013:MDS

[SO13] Martin Scheicher and Ulrich Oberst. Multidimensional discrete stability by Serre categories and the construction and parametrization of observers via Gabriel localizations. SIAM Journal on Control and Optimization, 51(3): 1873–1908, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sokolowski:1987:SAC

[Sok87]  Jan Sokołowski.  Sensitiv-

ity analysis of control constrained optimal control problems for distributed parameter systems. SIAM Journal on Control and Optimization, 25(6):1542–1556, November 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sokolowski:1988:SSA

[Sok88] Jan Sokołowski. Shape sensitivity analysis of boundary optimal control problems for parabolic systems. SIAM Journal on Control and Optimization, 26(4):763–787, July 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sontag:1979:OPS

[Son79]  Eduardo D. Sontag.  On

the observability of polynomial systems, I: Finite-time problems. SIAM Journal on Control and Optimization, 17(1):139–151, January 1979. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sontag:1983:LLC

[Son83]  Eduardo D. Sontag.  A

Lyapunov-like characterization of asymptotic controllability. SIAM Journal on Control and Optimization, 21(3): 462–471, May 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Soner:1986:OCSa

[Son86a] Halil Mete Soner. Optimal control with state-space constraint. I. SIAM Journal on Control and Optimization, 24 (3):552–561, May 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Soner:1986:OCSb

[Son86b]  Halil Mete Soner.  Opti-

mal control with state-space constraint. II. SIAM Journal on Control and Optimization, 24(6):1110–1122, November 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sontag:1988:CHD

[Son88] Eduardo D. Sontag. Controllability is harder to decide than accessibility. SIAM Journal on Control and Optimization, 26(5):1106–1118, September 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Soner:1993:SPM

[Son93] H. Mete Soner. Singular perturbations in manufacturing. SIAM Journal on Control and Optimization, 31(1): 132–146, January 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Song:2023:PCL

[Son23] Yuanzhuo Song. The partial controllability of linear stochastic control systems with terminal constraints and its applications to game-based control systems with jumps. SIAM Journal on Control and Optimization, 61(6):3635– 3663, December 2023. CO-DEN SJCODC. ISSN 1095-7138.

Soravia:1993:PEP

[Sor93] Pierpaolo Soravia. Pursuit– evasion problems and viscosity solutions of Isaacs equations. SIAM Journal on Control and Optimization, 31(3): 604–623, May 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Soravia:1996:CNS

[Sor96] Pierpaolo Soravia. H_{∞} control of nonlinear systems: Differential games and viscosity solutions. SIAM Journal on Control and Optimization, 34(3):1071–1097, May 1996. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26641.

Soravia:1998:ECF

[Sor98] Pierpaolo Soravia. Estimates of convergence of fully discrete schemes for the Isaacs equation of pursuit-evasion differential games via maximum principle. SIAM Journal on Control and Optimization, 36(1):1–11, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29186.

Schwartz:1996:CAO

[SP96] A. Schwartz and E. Polak. Consistent approximations for optimal control problems based on Runge–Kutta integration. SIAM Journal on Control and Optimization, 34(4):1235–1269, July 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26735.

Senapati:2002:SFE

[SP02] K. D. Senapati and G. Panda. Solution of a functional equation arising in continuous games: a dynamic programming approach. SIAM Journal on Control and Optimization, 41(3):820–825, May 2002. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35724.

Shen:2005:LCS

[SP05] Jinglai Shen and Jong-Shi Pang. Linear complementarity systems: Zeno states. SIAM Journal on Control and Optimization, 44(3):1040–1066, May 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/61227.

Smith:2010:DVR

[SPBF10] Stephen L. Smith, Marco Pavone, Francesco Bullo, and Emilio Frazzoli. Dynamic vehicle routing with priority classes of stochastic demands. SIAM Journal on Control and Optimization, 48 (5):3224–3245, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Spingarn:1980:FVC

[Spi80] J. E. Spingarn. Fixed and variable constraints in sensitivity analysis. SIAM Journal on Control and Optimization, 18 (3):297–310, May 1980. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shi:2015:NSC

[SPJ15] Guodong Shi, Alexandre Proutiere, and Karl Henrik Johansson. Network synchronization with convexity. SIAM Journal on Control and Optimization, 53(6):3562–3583, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sayin:2022:FPZ

[SPO22] Muhammed O. Sayin, Francesca Parise, and Asuman Ozdaglar. Fictitious play in zero-sum stochastic games. SIAM Journal on Control and Optimization, 60(4):2095–2114, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1426675.

Sirbu:2004:PCB

[SPS04] Mihai Sîrbu, Igor Pikovsky, and Steven E. Shreve. Perpetual convertible bonds. SIAM Journal on Control and Optimization, 43(1):58–85, January 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41245.

Soulaimani:2009:RGQ

[SQS09] Sami As Soulaimani, Marc Quincampoix, and Sylvain Sorin. Repeated games and qualitative differential games: Approachability and comparison of strategies. SIAM Journal on Control and Optimization, 48(4):2461–2479, ???? 2009. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Sideris:1997:COC

[SR97] Athanasios Sideris and Héctor Rotstein. Constrained H_{∞} optimal control over an infinite horizon. SIAM Journal on Control and Optimization, 35(4):1244–1262, July 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25689.

Santos:2004:CPP

[SR04] Manuel S. Santos and John Rust. Convergence properties of policy iteration. SIAM Journal on Control and Optimization, 42(6): 2094–2115, November 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39982.

Stamoulas:2018:CSR

[SR18] Serap Tay Stamoulas and Muruhan Rathinam. Convergence, stability, and robustness of multidimensional opinion dynamics in continuous time. SIAM Journal on Control and Optimization, 56(3): 1938–1967, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shafieepoorfard:2016:RIC

[SRM16] Ehsan Shafieepoorfard, Maxim Raginsky, and Sean P. Meyn. Rationally inattentive control of Markov processes. SIAM Journal on Control and Optimization, 54(2):987–1016, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sklyar:2015:SAN

[SRS15] K. V. Sklyar, R. Rabah, and G. M. Sklyar. Spectral assignment for neutral-type systems and moment problems. SIAM Journal on Control and Optimization, 53(2):845–873, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Soner:1989:RVF

[SS89] H. Mete Soner and Steven E. Shreve. Regularity of the value function for a twodimensional singular stochastic control problem. SIAM Journal on Control and Optimization, 27(4):876–907, July 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shankar:1992:AGA

[SS92] Shiva Shankar and V. R. Sule. Algebraic geometric aspects of feedback stabilization. SIAM Journal on Control and Optimization, 30(1):11–30, January 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sano:1999:CDS

[SS99a] Hideki Sano and Yoshiyuki Sakawa. H_{∞} control of diffusion systems by using a finite-dimensional controller. SIAM Journal on Control and Optimization, 37(2):409–428, March 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30290.

Scholtes:1999:EPM

[SS99b] Stefan Scholtes and Michael Stöhr. Exact penalization of mathematical programs with equilibrium constraints. SIAM Journal on Control and Optimization, 37(2):617–652, March 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30612.

Solodov:1999:NPM

[SS99c] M. V. Solodov and B. F.

Svaiter. A new projection method for variational inequality problems. SIAM Journal on Control and Optimization, 37(3):765–776, May 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31747.

Stankovic:2003:IPL

[SS03a] Srdjan S. Stankovic and Dragoslav D. Siljak. Inclusion principle for linear timevarying systems. SIAM Journal on Control and Optimization, 42(1):321–341, January 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/39060.

Stein:2003:SSI

[SS03b] Oliver Stein and Georg Still. Solving semi-infinite optimization problems with interior point techniques. SIAM Journal on Control and Optimization, 42(3):769–788, May 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/39839.

Sirbu:2006:TPG

[SS06] Mihai Sîrbu and Steven E. Shreve. A two-person game for pricing convertible bonds. SIAM Journal on Control and Optimization, 45(4):1508– 1539, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sarlette:2009:COM

[SS09a] Alain Sarlette and Rodolphe Sepulchre. Consensus optimization on manifolds. SIAM Journal on Control and Optimization, 48(1):56–76, ????

2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Scheibe:2009:BAP

[SS09b] Frank Scheibe and Malcolm C.

Smith. A behavioral approach to play in mechanical networks. SIAM Journal on Control and Optimization, 47(6): 2967–2990, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Schmidt:2009:IRA

[SS09c] Stephan Schmidt and Volker Schulz. Impulse response approximations of discrete shape Hessians with application in CFD. SIAM Journal on Control and Optimization, 48(4): 2562–2580, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Suttner:2018:FSC

[SS18] Raik Suttner and Zhiyong Sun. Formation shape control based on distance measurements using Lie bracket approximations. SIAM Journal on Control and Optimization, 56(6):4405–4433, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Soltani:2019:MAL

[SS19] Mohammad Soltani and Abhyudai Singh. Moment analysis of linear time-varying dynamical systems with renewal transitions. SIAM Journal on Control and Optimization, 57(4):2660–2685, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shangerganesh:2020:OCP

[SS20] L. Shangerganesh and P. T. Sowndarrajan. An optimal control problem of nonlocal pyragas feedback controllers for convective FitzHugh– Nagumo equations with timedelay. SIAM Journal on Control and Optimization, 58(6): 3613–3631, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Santambrogio:2021:CSI

[SS21] Filippo Santambrogio and Woojoo Shim. A Cucker– Smale inspired deterministic mean field game with velocity interactions. SIAM Journal on Control and Optimization, 59(6):4155–4187, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Schmidt:2023:LVS

[SS23] Stephan Schmidt and Volker H. Schulz. A linear view on shape optimization. SIAM Journal on Control and Optimization, 61(4):2358–2378, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1488910.

Siska:2024:GFR

[SS24] David Siska and Lukasz Szpruch. Gradient flows for regularized stochastic control problems. SIAM Journal on Control and Optimization, 62 (4):2036–2070, August 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1373645.

Shoarinejad:2002:SDC

[SSK02] Kambiz Shoarinejad, Jason L.

Speyer, and Ioannis Kanellakopoulos. A stochastic decentralized control problem with noisy communication. SIAM Journal on Control and Optimization, 41(3):975–990, May 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38562.

Stefanatos:2011:MTF

[SSL11] Dionisis Stefanatos, Heinz Schaettler, and Jr-Shin Li. Minimum-time frictionless atom cooling in harmonic traps. SIAM Journal on Control and Optimization, 49(6): 2440–2462, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i6/p2440_s1.

Shen:2013:RSM

[SSO13] Yun Shen, Wilhelm Stannat, and Klaus Obermayer. Risksensitive Markov control processes. SIAM Journal on Control and Optimization, 51(5): 3652–3672, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Schulz:2015:SIM

[SSW15] Volker H. Schulz, Martin Siebenborn, and Kathrin Welker. Structured inverse modeling in parabolic diffusion problems. SIAM Journal on Control and Optimization, 53(6):3319–3338, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sanjari:2024:NEE

[SSY24] Sina Sanjari, Naci Saldi, and Serdar Yüksel. Nash equilibria for exchangeable teamagainst-team games, their mean-field limit, and the role of common randomness. SIAM Journal on Control and Optimization, 62(3):1437–1464, June 2024. CODEN SJCODC. ISSN 1095-7138.

Song:2011:OHP

[SSZ11] Qingshuo Song, Richard H. Stockbridge, and Chao Zhu. On optimal harvesting problems in random environments. SIAM Journal on Control and Optimization, 49(2):859–

889, ????  2011.  CO-

DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i2/p859_s1.

Stoorvogel:1990:QMI

[ST90] A. A. Stoorvogel and H. L.

Trentelman. The quadratic matrix inequality in singular H_{∞} control with state feedback. SIAM Journal on Control and Optimization, 28(5): 1190–1208, September 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Solodov:1996:MPT

[ST96] Michael V. Solodov and Paul Tseng. Modified projectiontype methods for monotone variational inequalities. SIAM Journal on Control and Optimization, 34(5): 1814–1830, September 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/26865.

Sprekels:1999:DAO

[ST99] J. Sprekels and D. Tiba.

A duality approach in the optimization of beams and plates. SIAM Journal on Control and Optimization, 37(2):486–501, March 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32540.

Soner:2000:SUG

[ST00] H. Mete Soner and Nizar Touzi. Superreplication under gamma constraints. SIAM Journal on Control and Optimization, 39(1):73–96, January 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/34899.

Soner:2002:STP

[ST02] H. Mete Soner and Nizar Touzi. Stochastic target problems, dynamic programming, and viscosity solutions. SIAM Journal on Control and Optimization, 41(2):404–424, March 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/37886.

Soner:2009:DPE

[ST09a] H. Mete Soner and Nizar Touzi. The dynamic programming equation for second order stochastic target problems. SIAM Journal on Control and Optimization, 48(4): 2344–2365, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sprekels:2009:CVA

[ST09b] Jürgen Sprekels and Dan Tiba.

The control variational approach for differential systems. SIAM Journal on Control and Optimization, 47(6): 3220–3236, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Soner:2013:HAS

[ST13] H. Mete Soner and Nizar Touzi. Homogenization and asymptotics for small transaction costs. SIAM Journal on Control and Optimization, 51(4):2893–2921, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Santacroce:2014:FBS

[ST14] Marina Santacroce and Barbara Trivellato. Forward backward semimartingale systems for utility maximization. SIAM Journal on Control and Optimization, 52(6): 3517–3537, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Son:2016:CRL

[ST16] Nguyen Khoa Son and Do Duc Thuan. Controllability radii of linear systems with constrained controls under structured perturbations. SIAM Journal on Control and Optimization, 54(5):2820–2843, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shao:2023:OSS

[ST23] Jinghai Shao and Taoran Tian. From the optimal singular stochastic control to the optimal stopping for regimeswitching processes. SIAM Journal on Control and Optimization, 61(3):1631–1650, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1506705.

Stalford:1978:STD

[Sta78] Harold Stalford. Sufficiency theorem for discontinuous optimal cost surfaces. SIAM Journal on Control and Optimization, 16(1):63–82, January 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Stanford:1979:SMR

[Sta79] David P. Stanford. Stability for a multi-rate sampleddata system. SIAM Journal on Control and Optimization, 17 (3):390–399, May 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Staffans:1993:FBM

[Sta93] Olof J. Staffans. The fourblock model matching problem in ℓ^{1} and infinite-dimensional linear programming. SIAM Journal on Control and Optimization, 31(3):747–779, May 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Staffans:1998:CFW

[Sta98a] Olof J. Staffans. Coprime factorizations and well-posed linear systems. SIAM Journal on Control and Optimization, 36(4):1268–1292, July 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28541.

Staffans:1998:QOC

[Sta98b] Olof J. Staffans. Quadratic optimal control of well-posed linear systems. SIAM Journal on Control and Optimization, 37(1):131–164, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31425.

Stavre:2015:DCP

[Sta15] R. Stavre. A distributed control problem for two coupled fluids in a porous medium. SIAM Journal on Control and Optimization, 53(1):313–335, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Stewart:1992:PDT

[Ste92] David E. Stewart. A problem decomposition technique with application to the optimal distribution of enzymes. SIAM Journal on Control and Optimization, 30(2):390– 407, March 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Stern:2004:CSC

[Ste04]  R. J. Stern.  Character-

ization of the state constrained minimal time function. SIAM Journal on Control and Optimization, 43(2):697–707, March 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42603.

Stefanovski:2007:SFC

[Ste07] Jovan Stefanovski. Simplified formula for a controller in optimal control problems. SIAM Journal on Control and Optimization, 45(6):2011–2034, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Stefanelli:2008:BEP

[Ste08] Ulisse Stefanelli. The Brezis– Ekeland principle for doubly nonlinear equations. SIAM Journal on Control and Optimization, 47(3):1615–1642, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Stettner:2011:PMF

[Ste11] Ł. Stettner. Penalty method for finite horizon stopping problems. SIAM Journal on Control and Optimization, 49 (3):1078–1099, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sicon/
resource/1/sjcodc/v49/i3/
p1078_s1.

Stefanovski:2013:ASB

[Ste13] Jovan Stefanovski. All solutions of a bitangential interpolation problem that includes boundary points. SIAM Journal on Control and Optimization, 51(6):4387–4413, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Stefanovski:2015:PNI

[Ste15] Jovan Stefanovski. H_{∞} problem with nonstrict inequality and all solutions: Interpolation approach. SIAM Journal on Control and Optimization, 53(4):1734–1767, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Stefanatos:2017:MTT

[Ste17] Dionisis Stefanatos. Minimumtime transitions between thermal and fixed average energy states of the quantum parametric oscillator. SIAM Journal on Control and Optimization, 55(3):1429–1451, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Stettner:2019:LRC

[Ste19] Lukasz Stettner. Long run control with degenerate observation. SIAM Journal on Control and Optimization, 57(2): 880–899, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Stettner:2022:AAC

[Ste22] Lukasz Stettner. On an approximation of average cost per unit time impulse control of Markov processes. SIAM Journal on Control and Optimization, 60(4):2115–2131, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1434143.

Stefanovski:2023:DSA

[Ste23] Jovan D. Stefanovski. Deflating subspace approach to UIO design and application to fault-tolerant control. SIAM Journal on Control and Optimization, 61(5):2876–2901, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1535619.

Stillfjord:2018:SVD

[Sti18] Tony Stillfjord. Singular value decay of operator-valued differential Lyapunov and Riccati equations. SIAM Journal on Control and Optimization, 56(5):3598–3618, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Stone:1976:ITO

[Sto76] Lawrence D. Stone. Incremental and total optimization of separable functionals with constraints. SIAM Journal on Control and Optimization, 14(5):791–802, August 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Stojanovic:1991:ODC

[Sto91a] Srdjan Stojanovic. Optimal damping control and nonlinear elliptic systems. SIAM Journal on Control and Optimization, 29(3):594–608, May 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Stoorvogel:1991:SCP

[Sto91b] A. A. Stoorvogel. The singular H_{∞} control problem with dynamic measurement feedback. SIAM Journal on Control and Optimization, 29(1): 160–184, January 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Stoorvogel:1992:DTC

[Sto92] A. A. Stoorvogel. The discrete time H_{∞} control problem with measurement feedback. SIAM Journal on Control and Optimization, 30(1): 182–202, January 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Stojanovic:1997:PFR

[Sto97] Srdjan Stojanovic. Perturbation formula for regular free boundaries in elliptic and parabolic obstacle problems. SIAM Journal on Control and Optimization, 35(6): 2086–2100, November 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28279.

Stojanovic:2004:OMH

[Sto04]  Srdjan Stojanovic.  Opti-

mal momentum hedging via hypoelliptic reduced Monge– Ampère PDE. SIAM Journal on Control and Optimization, 43(4):1151–1173, July 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/42117.

Studniarski:1986:NSC

[Stu86] Marcin Studniarski. Necessary and sufficient conditions for isolated local minima of nonsmooth functions. SIAM Journal on Control and Optimization, 24(5):1044–1049, July 1986. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Sturm:2015:MLA

[Stu15] Kevin Sturm. Minimax Lagrangian approach to the differentiability of nonlinear PDE constrained shape functions without saddle point assumption. SIAM Journal on Control and Optimization, 53 (4):2017–2039, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sturm:2016:SON

[Stu16] Kevin Sturm. Shape optimization with nonsmooth cost functions: From theory to numerics. SIAM Journal on Control and Optimization, 54 (6):3319–3346, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Song:2020:MPP

[STW20] Yuanzhuo Song, Shanjian Tang, and Zhen Wu. The maximum principle for progressive optimal stochastic control problems with random jumps. SIAM Journal on Control and Optimization, 58(4): 2171–2187, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Szpruch:2024:OSE

[STZ24] Lukasz Szpruch, Tanut Treetanthiploet, and Yufei Zhang.

Optimal scheduling of entropy regularizer for continuoustime linear-quadratic reinforcement learning. SIAM Journal on Control and Optimization, 62(1):135–166, January 2024. CODEN SJCODC. ISSN 1095-7138.

Schulze:2018:MRL

[SU18] Philipp Schulze and Benjamin Unger. Model reduction for linear systems with low-rank switching. SIAM Journal on Control and Optimization, 56 (6):4365–4384, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Schmitt:2021:OBC

[SU21] Johann M. Schmitt and Stefan Ulbrich. Optimal boundary control of hyperbolic balance laws with state constraints. SIAM Journal on Control and Optimization, 59 (2):1341–1369, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Subrahmanyam:1981:ACT

[Sub81] M. B. Subrahmanyam. On applications of control theory to integral inequalities: II. SIAM Journal on Control and Optimization, 19(4):479–489, July 1981. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sule:1994:FSC

[Sul94] V. R. Sule. Feedback stabilization over commutative rings: The matrix case. SIAM Journal on Control and Optimization, 32(6):1675–1695, November 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/22027. See cor-

rigendum [Sul98].

Sule:1998:CFS

[Sul98] V. R. Sule. Corrigendum:

Feedback stabilization over commutative rings: The matrix case. SIAM Journal on Control and Optimization, 36(6):2194–2195, November 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31966.  See

[Sul94].

Sun:1981:SDC

[Sun81] Shun Hua Sun. On spectrum distribution of completely controllable linear systems. SIAM Journal on Control and Optimization, 19(6): 730–743, November 1981. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). Translated from the Chinese by L. F. Ho.

Sun:1982:BSH

[Sun82] Shun Hua Sun. Boundary stabilization of hyperbolic systems with no dissipative conditions. SIAM Journal on Control and Optimization, 20(6): 862–883, November 1982. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sun:1996:KVE

[Sun96] S. M. Sun. The Korteweg– de Vries equation on a periodic domain with singularpoint dissipation. SIAM Journal on Control and Optimization, 34(3):892–912, May 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26949.

Sun:2021:TPZ

[Sun21]  Jingrui Sun.  Two-person

zero-sum stochastic linearquadratic differential games. SIAM Journal on Control and Optimization, 59(3):1804– 1829, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sussmann:1978:SCL

[Sus78] Héctor J. Sussmann. A sufficient condition for local controllability. SIAM Journal on Control and Optimization, 16 (5):790–802, September 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sussmann:1979:BBT

[Sus79] Héctor J. Sussmann. A bang– bang theorem with bounds on the number of switchings. SIAM Journal on Control and Optimization, 17 (5):629–651, September 1979. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sussmann:1983:LBL

[Sus83] Héctor J. Sussmann. Lie brackets and local controllability: a sufficient condition for scalar-input systems. SIAM Journal on Control and Optimization, 21 (5):686–713, September 1983. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sussmann:1987:GTL

[Sus87a] H. J. Sussmann. A general theorem on local controllability. SIAM Journal on Control and Optimization, 25(1): 158–194, January 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sussmann:1987:RST

[Sus87b] H. J. Sussmann. Regular synthesis for time-optimal control of single-input real analytic systems in the plane. SIAM Journal on Control and Optimization, 25(5):1145– 1162, September 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sussmann:1987:STOa

[Sus87c] H. J. Sussmann. The structure of time-optimal trajectories for single-input systems in the plane: The C^{∞} nonsingular case. SIAM Journal on Control and Optimization, 25 (2):433–465, March 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sussmann:1987:STOb

[Sus87d] H. J. Sussmann. The structure of time-optimal trajectories for single-input systems in the plane: The general real analytic case. SIAM Journal on Control and Optimization, 25 (4):868–904, July 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Suttner:2020:ESC

[Sut20] Raik Suttner. Extremum seeking control for a class of nonholonomic systems. SIAM Journal on Control and Optimization, 58(4):2588–2615, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Stassinopoulos:1979:CDS

[SV79] G. I. Stassinopoulos and R. B. Vinter. Continuous dependence of solutions of a differential inclusion on the right hand side with applications to stability of optimal control problems. SIAM Journal on Control and Optimization, 17 (3):432–449, May 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Silva:1997:NCO

[SV97] G. N. Silva and R. B. Vinter. Necessary conditions for optimal impulsive control problems. SIAM Journal on Control and Optimization, 35(6):1829–1846, November 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28185.

Sprekels:2017:NTI

[SV17] Jürgen Sprekels and Enrico Valdinoci. A new type of identification problems: Optimizing the fractional order in a nonlocal evolution equation. SIAM Journal on Control and Optimization, 55(1): 70–93, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Svaiter:2011:WCD

[Sva11] B. F. Svaiter. On weak convergence of the Douglas–Rachford method. SIAM Journal on Control and Optimization, 49 (1):280–287, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.

                                        siam.org/sicon/resource/
                                        1/sjcodc/v49/i1/p280_s1.

Singh:2009:FSP

[SVBZ09] Sumeetpal S. Singh, Ba-Ngu Vo, Adrian Baddeley, and Sergei Zuyev. Filters for spatial point processes. SIAM Journal on Control and Optimization, 48(4):2275–2295, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Schmidt:1978:BBS

[SW78] E. J. P. Georg Schmidt and Norbert Weck. On the boundary behavior of solutions to elliptic and parabolic equations — with applications to boundary control for parabolic equations. SIAM Journal on Control and Optimization, 16 (4):593–598, July 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Studniarski:1999:WSM

[SW99] Marcin Studniarski and Doug E. Ward. Weak sharp minima: Characterizations and sufficient conditions. SIAM Journal on Control and Optimization, 38(1):219–236, January 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/30126.

Sontag:2000:LCI

[SW00] Eduardo Sontag and Yuan Wang. Lyapunov characterizations of input to output stability. SIAM Journal on Control and Optimization, 39(1):226–249, January 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35021.

Shen:2011:EMF

[SW11] Jinglai Shen and Xiao Wang.

Estimation of monotone functions via P -splines: a constrained dynamical optimization approach. SIAM Journal on Control and Optimization, 49(2):646–671, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sicon/
resource/1/sjcodc/v49/i2/
p646_s1.

Staffans:2012:PMC

[SW12] Olof J. Staffans and George Weiss. A physically motivated class of scattering passive linear systems. SIAM Journal on Control and Optimization, 50(5):3083–3112, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Stannat:2021:PMP

[SW21] Wilhelm Stannat and Lukas Wessels. Peng’s maximum principle for stochastic partial differential equations. SIAM Journal on Control and Optimization, 59(5):3552–3573, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Singh:2023:SOS

[SW23a] Shantanu Singh and George Weiss. Second order systems on Hilbert spaces with nonlinear damping. SIAM Journal on Control and Optimization, 61(4):2630–2654, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M154199X.

Surowiec:2023:OCL

[SW23b] Thomas M. Surowiec and Shawn W. Walker. Optimal control of the Landau– de Gennes model of nematic liquid crystals. SIAM Journal on Control and Optimization, 61(4):2546–2570, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1506158.

Swiech:2020:VSH

[Świ20]  Andrzej Świech.  Viscos-

ity solutions to HJB equations for boundary-noise and boundary-control problems. SIAM Journal on Control and Optimization, 58(1):303–326, ???? 2020. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Sain:1991:EZM

[SWP91] Michael K. Sain, Bostwick F.

Wyman, and Joseph L. Peczkowski. Extended zeros and model matching. SIAM Journal on Control and Optimization, 29(3):562–593, May 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shi:2022:ERS

[SWW22] Banban Shi, Ya Wang, and Fuke Wu. Ergodicity of regime-switching functional diffusions with infinite delay and application to a numerical algorithm for stochastic optimization. SIAM Journal on Control and Optimization, 60(5):2658–2683, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1470050.

Sun:2023:ZSS

[SWW23] Jingrui Sun, Hanxiao Wang, and Jiaqiang Wen. Zero-sum Stackelberg stochastic linearquadratic differential games. SIAM Journal on Control and Optimization, 61(1):252–284, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1450458.

Sahu:2012:UHI

[SWY12] D. R. Sahu, N. C. Wong, and J. C. Yao. A unified hybrid iterative method for solving variational inequalities involving generalized pseudocontractive mappings. SIAM Journal on Control and Optimization, 50(4):2335–2354, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Schwenninger:2024:BSI

[SWZ24] Felix L. Schwenninger, Alexander A. Wierzba, and Hans Zwart. On BIBO stability of infinite-dimensional linear state-space systems. SIAM Journal on Control and Optimization, 62(1):22–41, January 2024. CODEN SJCODC. ISSN 1095-7138.

Shao:2014:SRR

[SX14a] Jinghai Shao and Fubao Xi. Stability and recurrence of regime-switching diffusion processes. SIAM Journal on Control and Optimization, 52 (6):3496–3516, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

She:2014:DML

[SX14b] Zhikun She and Bai Xue. Discovering multiple Lyapunov functions for switched hybrid systems. SIAM Journal on Control and Optimization, 52 (5):3312–3340, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shao:2019:SRS

[SX19] Jinghai Shao and Fubao Xi. Stabilization of regimeswitching processes by feedback control based on discrete time observations II: State-dependent case. SIAM Journal on Control and Optimization, 57(2):1413–1439, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sun:2023:SLQ

[SX23] Jingrui Sun and Jie Xiong.

Stochastic linear-quadratic optimal control with partial observation. SIAM Journal on Control and Optimization, 61(3):1231–1247, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M146586X.

Song:2009:RCN

[SY09] Q. S. Song and G. Yin. Rates of convergence of numerical methods for controlled regimeswitching diffusions with stopping times in the costs. SIAM Journal on Control and Optimization, 48(3):1831–1857, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sun:2014:LQS

[SY14] Jingrui Sun and Jiongmin Yong. Linear quadratic stochastic differential games: Open-loop and closed-loop saddle points. SIAM Journal on Control and Optimization, 52(6):4082–4121, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shi:2017:FDE

[SY17] Ji Shi and Stephen S.-T. Yau. Finite dimensional estimation algebras with state dimension 3 and rank 2, I: Linear structure of Wong matrix. SIAM Journal on Control and Optimization, 55(6): 4227–4246, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sanjari:2021:OPC

[SY21] Sina Sanjari and Serdar Yüksel. Optimal policies for convex symmetric stochastic dynamic teams and their mean-field limit. SIAM Journal on Control and Optimization, 59(2):777–804, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Soner:2024:VSM

[SY24a] H. Mete Soner and Qinxin Yan. Viscosity solutions for McKean–Vlasov control on a torus. SIAM Journal on Control and Optimization, 62(2): 903–923, March 2024. CO-DEN SJCODC. ISSN 1095-7138.

Sun:2024:TPM

[SY24b] Jingrui Sun and Jiongmin Yong. Turnpike properties for mean-field linear-quadratic optimal control problems. SIAM Journal on Control and Optimization, 62(1):752–775, February 2024. CODEN SJ-CODC. ISSN 1095-7138.

Song:2012:OSC

[SYZ12] Qingshuo Song, G. George Yin, and Chao Zhu. Optimal switching with constraints and utility maximization of an indivisible market. SIAM Journal on Control and Optimization, 50(2):629–651, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Song:2013:WCM

[SYZ13] Qingshuo Song, George Yin, and Qing Zhang. Weak convergence methods for approximation of the evaluation of path-dependent functionals. SIAM Journal on Control and Optimization, 51(5): 4189–4210, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Seidman:1982:EUO

[SZ82] Thomas I. Seidman and Hong Xing Zhou. Existence and uniqueness of optimal controls for a quasilinear parabolic equation.

SIAM Journal on Control and Optimization, 20(6):747– 762, November 1982. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sethi:1995:MHD

[SZ95] S. P. Sethi and Q. Zhang. Multilevel hierarchical decision making in stochastic marketing–production systems. SIAM Journal on Control and Optimization, 33(2):528–553, March 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25054.

Stefani:1996:OCC

[SZ96] Gianna Stefani and Pierluigi Zezza. Optimality conditions for a constrained control problem. SIAM Journal on Control and Optimization, 34(2):635–659, March 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26094.

Stefani:1997:CRL

[SZ97] G. Stefani and P. Zezza. Constrained regular LQ-Control problems. SIAM Journal on Control and Optimization, 35(3):876–900, May 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28684.

Sokolowski:1999:TDS

[SZ99] J. Sokolowski and A. Zochowski. On the topological derivative in shape optimization. SIAM Journal on Control and Optimization, 37(4):1251–1272, July 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32323.

Sokolowski:2003:OCS

[SZ03] J. Sokolowski and A. Zochowski. Optimality conditions for simultaneous topology and shape optimization. SIAM Journal on Control and Optimization, 42(4):1198–1221, July 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38443.

Sircar:2004:BAA

[SZ04] Ronnie Sircar and Thaleia Zariphopoulou. Bounds and asymptotic approximations for utility prices when volatility is random. SIAM Journal on Control and Optimization, 43(4):1328–1353, July 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40925.

Song:2010:FFM

[SZ10] Dong-Ping Song and Qing Zhang. A fluid flow model for empty container repositioning policy with a single port and stochastic demand. SIAM Journal on Control and Optimization, 48(5): 3623–3642, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Saporito:2019:SCDa

[SZ19] Yuri F. Saporito and Jianfeng Zhang. Stochastic control with delayed information and related nonlinear master equation. SIAM Journal on Control and Optimization, 57(1): 693–717, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Szajowski:1995:OSD

[Sza95] Krzysztof Szajowski. Optimal stopping of a discrete Markov process by two decision makers. SIAM Journal on Control and Optimization, 33(5):1392–1410, September 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/24687.

Sznaier:1993:SIN

[Szn93] Mario Sznaier. A set induced norm approach to the robust control of constrained systems. SIAM Journal on Control and Optimization, 31 (3):733–746, May 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sznaier:1995:MOA

[Szn95]  Mario Sznaier.  A mixed

l_{∞}/H_{∞} optimization approach to robust controller design. SIAM Journal on Control and Optimization, 33(4):1086–1101, July 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/23132.

Sun:2021:SDS

[SZS21] Tao Sun, Xudong Zhao, and Xi-Ming Sun. Switched dynamic systems with logic switching and its stability analysis. SIAM Journal on Control and Optimization, 59 (2):1188–1217, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Shi:2021:SSS

[SZSC21] Lei Shi, Wei Xing Zheng, Jinliang Shao, and Yuhua Cheng. Sub-super-stochastic matrix with applications to bipartite tracking control over signed networks. SIAM Journal on Control and Optimization, 59(6):4563–4589, ???? 2021. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Song:2016:CMM

[SZX16] H. T. Song, G. F. Zhang, and Z. R. Xi. Continuous-mode MultiPhoton filtering. SIAM Journal on Control and Optimization, 54(3):1602–1632, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Sun:1997:PCA

[SZZ97] Jie Sun, Jishan Zhu, and Gongyun Zhao. A predictorcorrector algorithm for a class of nonlinear saddle point problems. SIAM Journal on Control and Optimization, 35(2):532–551, March 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/27611.

Takeuchi:1981:NFF

[TA81] Yoshiki Takeuchi and Hajime Akashi. Nonlinear filtering formulas for discrete-time observations. SIAM Journal on Control and Optimization, 19 (2):244–261, March 1981. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tidball:1996:ADZ

[TA96] Mabel M. Tidball and Eitan Altman. Approximations in dynamic zero-sum games. I. SIAM Journal on Control and Optimization, 34 (1):311–328, January 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25534.

Timoshin:2021:RPD

[TA21] Sergey A. Timoshin and Toyohiko Aiki. Relaxation in population dynamics models with hysteresis. SIAM Journal on Control and Optimization, 59 (1):693–708, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tadmor:1988:TSC

[Tad88] Gilead Tadmor. Trajectory stabilizing controls in hereditary linear systems. SIAM Journal on Control and Optimization, 26(1):138–154, January 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tadmor:1989:ISC

[Tad89] Gilead Tadmor. H_{∞} interpolation in systems with commensurate input lags. SIAM Journal on Control and Optimization, 27(3):511–526, May 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tadmor:1993:SPM

[Tad93] Gilead Tadmor. The standard H_{∞} problem and the maximum principle: The general linear case. SIAM Journal on Control and Optimization, 31 (4):813–846, July 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tadmor:1997:WSM

[Tad97] Gilead Tadmor. Weighted sensitivity minimization in systems with a single input delay: a state space solution. SIAM Journal on Control and Optimization, 35(5): 1445–1469, September 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27965.

Tahraoui:1992:COE

[Tah92] R. Tahraoui. Contrôle optimal dans les équations elliptiques. (French) [Optimal control in elliptic equations]. SIAM Journal on Control and Optimization, 30(3):495–521, May 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Taksar:1997:IDL

[Tak97] Michael I. Taksar. Infinitedimensional linear programming approach to singular stochastic control. SIAM Journal on Control and Optimization, 35(2):604–625, March 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/28685.

Tallos:1991:VPN

[Tal91] Péter Tallós. Viability problems for nonautonomous differential inclusions. SIAM Journal on Control and Optimization, 29(2):253–263, March 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tall:2005:FCM

[Tal05] Issa Amadou Tall. Feedback classification of multiinput nonlinear control systems. SIAM Journal on Control and Optimization, 43(6): 2049–2070, November 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41050.

Tamir:1980:EAS

[Tam80] Arie Tamir. Efficient algorithms for a selection problem with nested constraints and its application to a productionsales planning model. SIAM Journal on Control and Optimization, 18(3):282–287, May 1980. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tanino:1988:SSA

[Tan88] Tetsuzo Tanino. Stability and sensitivity analysis in convex vector optimization. SIAM Journal on Control and Optimization, 26(3):521–536, May 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tang:1998:MPP

[Tan98] Shanjian Tang. The maximum principle for partially observed optimal control of stochastic differential equations. SIAM Journal on Control and Optimization, 36(5): 1596–1617, September 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/31310.

Tang:2000:BPC

[Tan00] Shanjian Tang. Brockett’s problem of classification of finite-dimensional estimation algebras for nonlinear filtering systems. SIAM Journal on Control and Optimization, 39(3):900–916, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33464.

Tang:2003:GLQ

[Tan03] Shanjian Tang. General linear quadratic optimal stochastic control problems with random coefficients: Linear stochastic Hamilton systems and backward stochastic Riccati equations. SIAM Journal on Control and Optimization, 42(1):53–75, January 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38755.

Tang:2015:DPG

[Tan15] Shanjian Tang. Dynamic programming for general linear quadratic optimal stochastic control with random coefficients. SIAM Journal on Control and Optimization, 53(2): 1082–1106, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tao:2018:LEC

[Tao18]  Qiang Tao.  Local exact

controllability for the planar compressible magnetohydrodynamic equations. SIAM Journal on Control and Optimization, 56(6):4461–4487, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tarres:1985:AES

[Tar85] R. Tarres. Asymptotic evolution of a stochastic control problem. SIAM Journal on Control and Optimization, 23 (4):614–631, July 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tardella:1990:NPL

[Tar90] Fabio Tardella. A new proof of the Lyapunov convexity theorem. SIAM Journal on Control and Optimization, 28(2): 478–481, March 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tanwani:2014:SOD

[TBP14] Aneel Tanwani, Bernard Brogliato, and Christophe Prieur. Stability and observer design for Lur’e systems with multivalued, nonmonotone, time-varying nonlinearities and state jumps. SIAM Journal on Control and Optimization, 52(6):3639–3672, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tanwani:2018:WPO

[TBP18] Aneel Tanwani, Bernard Brogliato, and Christophe Prieur. Well-posedness and output regulation for implicit time-varying evolution variational inequalities. SIAM Journal on Control and Optimization, 56(2):751–781, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tseng:1990:PAP

[TBT90] P. Tseng, D. P. Bertsekas, and J. N. Tsitsiklis. Partially asynchronous, parallel algorithms for network flow and other problems. SIAM Journal on Control and Optimization, 28 (3):678–710, May 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Taringoo:2013:OCI

[TC13] Farzin Taringoo and Peter E.

Caines. On the optimal control of impulsive hybrid systems on Riemannian manifolds. SIAM Journal on Control and Optimization, 51(4): 3127–3153, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Thuan:2014:CSC

[TC14] Le Quang Thuan and M. Kanat Camlibel. Controllability and stabilizability of a class of continuous piecewise affine dynamical systems. SIAM Journal on Control and Optimization, 52(3):1914–1934, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Takei:2015:OCB

[TCC^{+}15a] R. Takei, W. Chen, Z. Clawson, S. Kirov, and A. Vladimirsky. Optimal control with budget constraints and resets. SIAM Journal on Control and Optimization, 53(2):712–744, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tzortzis:2015:DPS

[TCC15b] Ioannis Tzortzis, Charalambos D. Charalambous, and Themistoklis Charalambous.

Dynamic programming subject to total variation distance ambiguity. SIAM Journal on Control and Optimization, 53 (4):2040–2075, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tzortzis:2019:IHA

[TCC19] Ioannis Tzortzis, Charalambos D. Charalambous, and Themistoklis Charalambous. Infinite horizon average cost dynamic programming subject to total variation distance ambiguity. SIAM Journal on Control and Optimization, 57 (4):2843–2872, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tian:1994:GPM

[TD94] T. Tian and J. C. Dunn. On the gradient projection method for optimal control problems with nonnegative L^{∈} inputs. SIAM Journal on Control and Optimization, 32(2):516–537, March 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/23365.

Tadic:2021:BPA

[TD21] Vladislav Z. B. Tadić and Arnaud Doucet. Bias of particle approximations to optimal filter derivative. SIAM Journal on Control and Optimization, 59(1):727–748, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tang:2022:RPD

[TDBL22] Yujie Tang, Emiliano Dall’Anese, Andrey Bernstein, and Steven Low. Running primaldual gradient method for time-varying nonconvex problems. SIAM Journal on Control and Optimization, 60 (4):1970–1990, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1371063.

Taringoo:2018:OMR

[TDNT18] Farzin Taringoo, Peter M.

Dower, Dragan Nesić, and Ying Tan. Optimization methods on Riemannian manifolds via extremum seeking algorithms. SIAM Journal on Control and Optimization, 56 (5):3867–3892, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tebou:2011:SRC

[Teb11] Louis Tebou. Some results on the controllability of coupled semilinear wave equations: The desensitizing control case. SIAM Journal on Control and Optimization, 49(3):1221– 1238, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicon/resource/
                                        1/sjcodc/v49/i3/p1221_s1.

Tebou:2022:SDE

[Teb22]  Louis Tebou.  Sharp de-

cay estimates for semigroups associated with some onedimensional fluid-structure interactions involving degeneracy. SIAM Journal on Control and Optimization, 60 (5):2787–2810, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M141854X.

Teglas:1984:CCS

[Teg84] Russell G. Teglas. On the control canonical structure of a class of scalar input systems. SIAM Journal on Control and Optimization, 22(4): 552–569, July 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Terrell:1994:ONS

[Ter94]  William J. Terrell.  The

output-nulling space, projected dynamics, and system decomposition for linear time-varying singular systems. SIAM Journal on Control and Optimization, 32(3):876–889, May 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/22561.

Tessitore:1992:SRR

[Tes92] Gianmario Tessitore. Some remarks on the Riccati equation arising in an optimal control problem with stateand control-dependent noise. SIAM Journal on Control and Optimization, 30(3):717–744, May 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Todorov:2008:OFC

[TF08] Marcos G. Todorov and Marcelo D. Fragoso. Output feedback H_{∞} control of continuous-time infinite Markovian jump linear systems via LMI methods. SIAM Journal on Control and Optimization, 47(2):950–974, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Todorov:2011:RSS

[TF11] Marcos G. Todorov and Marcelo D. Fragoso. On the robust stability, stabilization, and stability radii of continuous-time infinite Markov jump linear systems. SIAM Journal on Control and Optimization, 49(3):1171– 1196, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i3/p1171_s1.

Tse:2014:PSO

[TFL14] Shu Tong Tse, Peter A. Forsyth, and Yuying Li. Preservation of scalarization optimal points in the embedding technique for continuous time mean variance optimization. SIAM Journal on Control and Optimization, 52(3): 1527–1546, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Trentelman:2011:ORS

[TFT11] H. L. Trentelman, Shaik Fiaz, and K. Takaba. Optimal robust stabilization and dissipativity synthesis by behavioral interconnection. SIAM Journal on Control and Optimization, 49(1):288–314, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sicon/
                                        resource/1/sjcodc/v49/i1/
                                        p288_s1.

Thitsa:2012:RCI

[TG12] Makhin Thitsa and W. Steven Gray. On the radius of convergence of interconnected analytic nonlinear input-output systems. SIAM Journal on Control and Optimization, 50 (5):2786–2813, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Trentelman:2013:DBB

[TG13] H. L. Trentelman and S. V.

Gottimukkala. Distance between behaviors and rational representations. SIAM Journal on Control and Optimization, 51(6):4211–4241, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Touri:2023:UFC

[TG23] Behrouz Touri and Bahman Gharesifard. A unified framework for continuous-time unconstrained distributed optimization. SIAM Journal on Control and Optimization, 61 (4):2004–2020, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1442711.

Torrisi:2018:PGC

[TGSM18] Giampaolo Torrisi, Sergio Grammatico, Roy S. Smith, and Manfred Morari. A projected gradient and constraint linearization method for nonlinear model predictive control. SIAM Journal on Control and Optimization, 56(3): 1968–1999, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Thakur:1986:OIC

[Tha86] Lakshman S. Thakur. Optimal interpolation with convex splines of second degree. SIAM Journal on Control and Optimization, 24(1):157–168, January 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tanaka:2002:SOC

[THA02] Shunji Tanaka, Tomomichi Hagiwara, and Mituhiko Araki. A study on the optimal cost for the H_{∞} problem of discrete linear periodically time-varying systems. SIAM Journal on Control and Optimization, 41(2):362–379, March 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37651.

Tang:2007:SGS

[ThH07] Shanjian Tang and Shui hung Hou. Switching games of stochastic differential systems. SIAM Journal on Control and Optimization, 46(3):900–929, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Thibault:1991:SOV

[Thi91] Lionel Thibault. On subdifferentials of optimal value functions. SIAM Journal on Control and Optimization, 29(5): 1019–1036, September 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Thibault:1997:SCS

[Thi97] Lionel Thibault. Sequential convex subdifferential calculus and sequential Lagrange multipliers. SIAM Journal on Control and Optimization, 35(4):1434–1444, July 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28771.

Tiba:1985:OCD

[Tib85] D. Tiba. Optimality conditions for distributed control problems with nonlinear state equation. SIAM Journal on Control and Optimization, 23 (1):85–110, January 1985. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tiba:2011:FEA

[Tib11] Dan Tiba. Finite element approximation for shape optimization problems with Neumann and mixed boundary conditions. SIAM Journal on Control and Optimization, 49(3):1064–1077, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sicon/
resource/1/sjcodc/v49/i3/
p1064_s1.

Tie:2014:NCN

[Tie14] Lin Tie. On near-controllability, nearly controllable subspaces, and near-controllability index of a class of discrete-time bilinear systems: a root locus approach. SIAM Journal on Control and Optimization, 52 (2):1142–1165, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tie:2020:OCO

[Tie20] Lin Tie. Output-controllability and output-near-controllability of driftless discrete-time bilinear systems. SIAM Journal on Control and Optimization, 58(4):2114–2142, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Timoshin:2014:VSS

[Tim14] Sergey A. Timoshin. Variational stability of some optimal control problems describing hysteresis effects. SIAM Journal on Control and Optimization, 52(4):2348–2370, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Toshio:1981:ISG

[TK81] Nomura Toshio and Frtauta Katsuhisa. Invariant structures of general dynamical systems. SIAM Journal on Control and Optimization, 19(1): 154–167, January 1981. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tavakoli-Kakhki:2010:MSS

[TKH10] Mahsan Tavakoli-Kakhki and Mohammad Haeri. The minimal state space realization for a class of fractional order transfer functions. SIAM Journal on Control and Optimization, 48(7):4317–4326, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tiesler:2012:SCO

[TKXP12] Hanne Tiesler, Robert M.

Kirby, Dongbin Xiu, and Tobias Preusser. Stochastic collocation for optimal control problems with stochastic PDE constraints. SIAM Journal on Control and Optimization, 50(5):2659–2682, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tang:1994:NCO

[TL94] Shan Jian Tang and Xun Jing Li. Necessary conditions for optimal control of stochastic systems with random jumps. SIAM Journal on Control and Optimization, 32(5): 1447–1475, September 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/23385.

Tamzali-Lafond:2003:VIU

[TL03] Moulka Tamzali-Lafond. Variational inclusions under state constraints. SIAM Journal on Control and Optimization, 42(1):342–362, January 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39155.

Tatsien:2006:EBC

[TL06] Li Tatsien and Yu Lixin. Exact boundary controllability for 1-D quasilinear wave equations. SIAM Journal on Control and Optimization, 45(3): 1074–1083, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Thistle:2009:ECS

[TL09] J. G. Thistle and H. M. Lamouchi. Effective control synthesis for partially observed discrete-event systems. SIAM Journal on Control and Optimization, 48(3): 1858–1887, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tang:1999:AEP

[TLC99] Qian-Yu Tang, Pierre L’Ecuyer, and Han-Fu Chen. Asymptotic efficiency of perturbationanalysis-based stochastic approximation with averaging. SIAM Journal on Control and Optimization, 37(6): 1822–1847, November 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32124.

Terpin:2024:DPP

[TLD24] Antonio Terpin, Nicolas Lanzetti, and Florian Dörfler. Dynamic programming in probability spaces via optimal transport. SIAM Journal on Control and Optimization, 62(2): 1183–1206, April 2024. CO-DEN SJCODC. ISSN 1095-7138.

Tan:2016:EBC

[TLL16] Shaolin Tan, Jinhu Lü, and Zongli Lin. Emerging behavioral consensus of evolutionary dynamics on complex networks. SIAM Journal on Control and Optimization, 54(6): 3258–3272, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tidball:2000:COV

[TLPA00] Mabel M. Tidball, Ariel Lombardi, Odile Pourtallier, and Eitan Altman. Continuity of optimal values and solutions for control of Markov chains with constraints. SIAM Journal on Control and Optimization, 38(4):1204–1222, July 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/28029.

Li:2023:ACU

[tLR23] Ta tsien Li and Bopeng Rao. Algebraic conditions to the uniqueness theorem for a coupled system of wave equations. SIAM Journal on Control and Optimization, 61(1):135–150, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1468279.

Tudor:2015:CPD

[TO15] Sebastian F. Tudor and Cristian Oara. H_{∞} control problem for discrete-time algebraic dynamical systems. SIAM Journal on Control and Optimization, 53(5):3171–3194, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Toader:2011:TDH

[Toa11]  Anca-Maria Toader.  The

topological derivative of homogenized elastic coefficients of periodic microstructures. SIAM Journal on Control and Optimization, 49(4):1607– 1628, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicon/resource/
                                        1/sjcodc/v49/i4/p1607_s1.

Todd:1983:PLH

[Tod83] Michael J. Todd. Piecewiselinear homotopy algorithms for sparse systems of nonlinear equations. SIAM Journal on Control and Optimization, 21 (2):204–214, March 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tolstonogov:2016:RNO

[Tol16] Alexander A. Tolstonogov.

Relaxation in nonconvex optimal control problems containing the difference of two subdifferentials. SIAM Journal on Control and Optimization, 54 (1):175–197, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tolstonogov:2018:FWT

[Tol18] A. A. Tolstonogov. Filippov– Wazewski theorem for subdifferential inclusions with an unbounded perturbation. SIAM Journal on Control and Optimization, 56(4):2878–2900, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ton:2003:OSC

[Ton03]  Bui An Ton.  Optimal

shape control problem for the Navier–Stokes equations. SIAM Journal on Control and Optimization, 41(6): 1733–1747, November 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39128.

Topkis:1979:EPN

[Top79] Donald M. Topkis. Equilibrium points in nonzero-sum n-person submodular games. SIAM Journal on Control and Optimization, 17(6):773– 787, November 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Townley:1996:TAU

[Tow96] Stuart Townley. Topological aspects of universal adaptive stabilization. SIAM Journal on Control and Optimization, 34(3):1044–1070, May 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/23195.

Teel:1995:TSS

[TP95] Andrew Teel and Laurent Praly. Tools for semiglobal stabilization by partial state and output feedback. SIAM Journal on Control and Optimization, 33(5):1443–1488, September 1995. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/24143.

Tabuada:2005:QFN

[TP05] Paulo Tabuada and George J. Pappas. Quotients of fully nonlinear control systems. SIAM Journal on Control and Optimization, 43(5): 1844–1866, September 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39902.

Tidball:1997:ADZ

[TPA97] Mabel M. Tidball, Odile Pourtallier, and Eitan Altman. Approximations in dynamic zerosum games. II. SIAM Journal on Control and Optimization, 35(6):2101–2117, November 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/27246.

Trentelman:2001:PMC

[TR01] H. L. Trentelman and P. Rapisarda. Pick matrix conditions for sign-definite solutions of the algebraic Riccati equation. SIAM Journal on Control and Optimization, 40(3):969–991, May 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/36851.

Tall:2002:FCN

[TR02] Issa Amadou Tall and Witold Respondek. Feedback classification of nonlinear singleinput control systems with controllable linearization: Normal forms, canonical forms, and invariants. SIAM Journal on Control and Optimization, 41(5):1498–1531, September 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/38175.

Trentelman:1986:ADD

[Tre86] Harry L. Trentelman. Almost disturbance decoupling with bounded peaking. SIAM Journal on Control and Optimization, 24(6):1150–1176, November 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Trentelman:1989:RFE

[Tre89] Harry L. Trentelman. The regular free-endpoint linear quadratic problem with indefinite cost. SIAM Journal on Control and Optimization, 27 (1):27–42, January 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Treiman:1990:OCS

[Tre90] Jay S. Treiman. Optimal control with small generalized gradients. SIAM Journal on Control and Optimization, 28 (3):720–732, May 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Treiman:1999:LMN

[Tre99] Jay S. Treiman. Lagrange multipliers for nonconvex generalized gradients with equality, inequality, and set constraints. SIAM Journal on Control and Optimization, 37(5):1313–1329, September 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30659.

Triggiani:1976:ERC

[Tri76] Roberto Triggiani. Extensions of rank conditions for controllability and observability to Banach spaces and unbounded operators. SIAM Journal on Control and Optimization, 14 (2):313–338, February 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Triggiani:1977:NLE

[Tri77] Roberto Triggiani. A note on the lack of exact controllability for mild solutions in Banach spaces. SIAM Journal on Control and Optimization, 15 (3):407–411, May 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See addendum [Tri80].

Triggiani:1978:RBF

[Tri78] Roberto Triggiani. On the relationship between first and second order controllable systems in Banach spaces. SIAM Journal on Control and Optimization, 16(6):847– 859, November 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Triggiani:1980:ANL

[Tri80] Roberto Triggiani. Addendum: “A Note on the Lack of Exact Controllability for Mild Solutions in Banach Spaces”. SIAM Journal on Control and Optimization, 18 (1):98–99, January 1980. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [Tri77].

Trouve:1996:CDS

[Tro96]  Alain Trouvé.  Cycle de-

compositions and simulated annealing. SIAM Journal on Control and Optimization, 34(3):966–986, May 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/25858.

Troltzsch:1999:LNS

[Trö99]  Fredi Tröltzsch.  On the

Lagrange–Newton–SQP method for the optimal control of semilinear parabolic equations. SIAM Journal on Control and Optimization, 38 (1):294–312, January 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/34142.

Trentelman:1995:SDD

[TS95] H. L. Trentelman and A. A. Stoorvogel. Sampled-data and discrete-time H_{2} optimal control. SIAM Journal on Control and Optimization, 33(3):834–862, May 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24199.

Turetsky:2018:PEG

[TS18] Vladimir Turetsky and Tal Shima. Pursuit-evasion guidance in a switched system. SIAM Journal on Control and Optimization, 56(4):2613– 2633, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tang:2024:LOP

[TS24] Qing Tang and Jiahao Song.

Learning optimal policies in potential mean field games: Smoothed policy iteration algorithms. SIAM Journal on Control and Optimization, 62(1):351–375, January 2024. CODEN SJCODC. ISSN 1095-7138.

Tsai:1978:PSL

[Tsa78] Chun-Ping Tsai. Perturbed stochastic linear regulator problems. SIAM Journal on Control and Optimization, 16 (3):396–410, May 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tseng:1990:DAM

[Tse90] Paul Tseng. Dual ascent methods for problems with strictly convex costs and linear constraints: a unified approach. SIAM Journal on Control and Optimization, 28 (1):214–242, January 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tseng:1991:ASA

[Tse91] Paul Tseng. Applications of a splitting algorithm to decomposition in convex programming and variational inequalities. SIAM Journal on Control and Optimization, 29(1): 119–138, January 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tseng:1993:PFA

[Tse93] Paul Tseng. A path-following algorithm for linear programming using quadratic and logarithmic penalty functions. SIAM Journal on Control and Optimization, 31(6):1578– 1598, November 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tseng:2000:MFB

[Tse00]  Paul Tseng.  A modified

forward-backward splitting method for maximal monotone mappings. SIAM Journal on Control and Optimization, 38(2):431–446, March 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33880.

Tsinias:1991:ECL

[Tsi91] J. Tsinias. Existence of control Lyapunov functions and applications to state feedback stabilizability of nonlinear systems. SIAM Journal on Control and Optimization, 29(2): 457–473, March 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tsinias:1992:ECL

[Tsi92] John Tsinias. On the existence of control Lyapunov functions: Generalizations of Vidyasagar’s theorem on nonlinear stabilization. SIAM Journal on Control and Optimization, 30(4):879–893, July 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tsinias:1993:VSI

[Tsi93] John Tsinias. Versions of Sontag’s “Input to State Stability Condition” and the global stabilizability problem. SIAM Journal on Control and Optimization, 31(4):928–941, July 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tsinias:1995:SGS

[Tsi95] John Tsinias. Smoothly global stabilizability by dynamic feedback and generalizations of Artstein’s theorem. SIAM Journal on Control and Optimization, 33(4):1071–1085, July 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24190.

Tyukin:2008:NSG

[TSNvL08] Ivan Tyukin, Erik Steur, Henk Nijmeijer, and Cees van Leeuwen. Nonuniform smallgain theorems for systems with unstable invariant sets. SIAM Journal on Control and Optimization, 47(2):849–882, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tranninger:2022:DCS

[TSS^{+}22] Markus Tranninger, Richard Seeber, Martin Steinberger, Martin Horn, and Christian Pötzsche. Detectability conditions and state estimation for linear time-varying and nonlinear systems. SIAM Journal on Control and Optimization, 60(4):2514–2537, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1429473.

Tahar:2007:DPE

[TST07] Imen Ben Tahar, H. Mete Soner, and Nizar Touzi. The dynamic programming equation for the problem of optimal investment under capital gains taxes. SIAM Journal on Control and Optimization, 46(5): 1779–1801, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tuan:2020:QTT

[TT20] H. T. Tuan and H. Trinh. A qualitative theory of time delay nonlinear fractional-order systems. SIAM Journal on Control and Optimization, 58 (3):1491–1518, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Talbi:2023:DPE

[TTZ23a] Mehdi Talbi, Nizar Touzi, and Jianfeng Zhang. Dynamic programming equation for the mean field optimal stopping problem. SIAM Journal on Control and Optimization, 61 (4):2140–2164, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1404259.

Talbi:2023:VSO

[TTZ23b] Mehdi Talbi, Nizar Touzi, and Jianfeng Zhang. Viscosity solutions for obstacle problems on Wasserstein space. SIAM Journal on Control and Optimization, 61(3):1712–1736, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1488119.

Tucsnak:1996:REC

[Tuc96] Marius Tucsnak. Regularity and exact controllability for a beam with piezoelectric actuator. SIAM Journal on Control and Optimization, 34(3):922–930, May 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/26546.

Tudor:1990:OCI

[Tud90]  Constantin Tudor.  Opti-

mal control for an infinitedimensional periodic problem under white noise perturbations. SIAM Journal on Control and Optimization, 28(2): 253–264, March 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Touzi:2002:CTD

[TV02] Nizar Touzi and Nicolas Vieille. Continuous-time Dynkin games with mixed strategies. SIAM Journal on Control and Optimization, 41(4):1073–1088, July 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/36981.

Thistle:1994:CIB

[TW94a] J. G. Thistle and W. M. Wonham. Control of infinite behavior of finite automata. SIAM Journal on Control and Optimization, 32(4):1075–1097, July 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/21753.

Thistle:1994:SIB

[TW94b] J. G. Thistle and W. M.

Wonham. Supervision of infinite behavior of discreteevent systems. SIAM Journal on Control and Optimization, 32(4):1098–1113, July 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/21752.

Tucsnak:2000:SEC

[TW00] Marius Tucsnak and George Weiss. Simultaneous exact controllability and some applications. SIAM Journal on Control and Optimization, 38(5):1408–1427, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35271.

Tucsnak:2003:HGC

[TW03] Marius Tucsnak and George Weiss. How to get a conservative well-posed linear system out of thin air. Part II. controllability and stability. SIAM Journal on Control and Optimization, 42(3):907–935, May 2003. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/39929.

Tian:2023:HSI

[TWB23] Ye Tian, Long Wang, and Francesco Bullo. How social influence affects the wisdom of crowds in influence networks. SIAM Journal on Control and Optimization, 61 (4):2334–2357, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1492751.

Tucsnak:2016:PTO

[TWW16] Marius Tucsnak, Gensgsheng Wang, and Chi-Ting Wu. Perturbations of time optimal control problems for a class of abstract parabolic systems. SIAM Journal on Control and Optimization, 54(6): 2965–2991, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tam:1990:NSC

[TWY90] Luen-Fai Tam, Wing Shing Wong, and Stephen S.-T. Yau. On a necessary and sufficient condition for finite dimensionality of estimation algebras. SIAM Journal on Control and Optimization, 28(1): 173–185, January 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Trelat:2016:ISD

[TWZ16] Emmanuel Trélat, Lijuan Wang, and Yubiao Zhang. Impulse and sampled-data optimal control of heat equations, and error estimates. SIAM Journal on Control and Optimization, 54(5):2787–2819, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tan:2021:FSP

[TWZ21] Ken Seng Tan, Wei Wei, and Xun Yu Zhou. Failure of smooth pasting principle and nonexistence of equilibrium stopping rules under timeinconsistency. SIAM Journal on Control and Optimization, 59(6):4136–4154, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Trouve:2000:CDM

[TY00] Alain Trouvé and Laurent Younes. On a class of diffeomorphic matching problems in one dimension. SIAM Journal on Control and Optimization, 39(4):1112–1135, July 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34864.

Tian:2023:MFT

[TY23] Ran Tian and Zhiyong Yu. Mean-field type FBSDEs under domination-monotonicity conditions and application to LQ problems. SIAM Journal on Control and Optimization, 61(1):22–46, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M144219X.

Thanh:2011:SOR

[TYW11] Le Van Thanh, G. George Yin, and Le Yi Wang. State observers with random sampling times and convergence analysis of double-indexed and randomly weighted sums of mixing processes. SIAM Journal on Control and Optimization, 49(1):106–124, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sicon/
resource/1/sjcodc/v49/i1/
p106_s1.

Taksar:2007:OTW

[TZ07] Michael Taksar and Xudong Zeng. Optimal terminal wealth under partial information: Both the drift and the volatility driven by a discretetime Markov chain. SIAM Journal on Control and Optimization, 46(4):1461–1482, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tang:2009:NCF

[TZ09] Shanjian Tang and Xu Zhang.

Null controllability for forward and backward stochastic parabolic equations. SIAM Journal on Control and Optimization, 48(4):2191–2216, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Taksar:2010:MCU

[TZ10] Michael Taksar and Xudong Zeng. On maximizing CRRA utility in regime switching markets with random endowment. SIAM Journal on Control and Optimization, 48(5): 2984–3002, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tang:2011:OSO

[TZK11] Shanjian Tang, Wei Zhong, and Hyeng Keun Koo. Optimal switching of one-dimensional reflected BSDEs and associated multidimensional BS-DEs with oblique reflection. SIAM Journal on Control and Optimization, 49(6):2279– 2317, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i6/p2279_s1.

Tianqiu:2021:NRS

[TZQ21] Yu Tianqiu, Di Zhao, and Li Qiu. Networked robust stability for LTV systems with simultaneous uncertainties in plant, controller, and communication channels. SIAM Journal on Control and Optimization, 59(1):1–23, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tarn:1988:FSL

[TZZ88] Tzyh Jong Tarn, Xiaoming Zeng, and John R. Zavgren, Jr. Feedback stabilization of linear dynamic systems with multirate sampled output. SIAM Journal on Control and Optimization, 26(4): 812–833, July 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Trelat:2018:SSP

[TZZ18] Emmanuel Trélat, Can Zhang, and Enrique Zuazua. Steadystate and periodic exponential turnpike property for optimal control problems in Hilbert spaces. SIAM Journal on Control and Optimization, 56(2): 1222–1252, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Tang:2022:EHE

[TZZ22] Wenpin Tang, Yuming Paul Zhang, and Xun Yu Zhou. Exploratory HJB equations and their convergence. SIAM Journal on Control and Optimization, 60(6):3191–3216, ???? 2022. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1448185.

Uchida:1978:ENE

[Uch78] Kenko Uchida. On existence of a Nash equilibrium point in N-person nonzero sum stochastic differential games. SIAM Journal on Control and Optimization, 16(1):142– 149, January 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Uchida:1979:NEN

[Uch79] Kenko Uchida. A note on the existence of a Nash equilibrium point in stochastic differential games. SIAM Journal on Control and Optimization, 17(1):1–4, January 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Uchida:1981:SCO

[Uch81]  Kenkō Uchida.  Sufficient

conditions for observationinnovation equivalence in white Gaussian channels with feedback. SIAM Journal on Control and Optimization, 19 (5):667–675, September 1981. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Uchida:1983:SPL

[Uch83] Kenko Uchida. Structural properties of the linear– quadratic stochastic control problem. SIAM Journal on Control and Optimization, 21 (5):677–685, September 1983. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ulbrich:2002:SAC

[Ulb02] Stefan Ulbrich. A sensitivity and adjoint calculus for discontinuous solutions of hyperbolic conservation laws with source terms. SIAM Journal on Control and Optimization, 41(3):740–797, May 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37076.

Udwadia:1988:ODS

[UMD88] F. E. Udwadia, F. Mantel, and M. Dryja. On the optimal design of structures subjected to periodic base excitations. SIAM Journal on Control and Optimization, 26(6): 1306–1320, November 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Underwood:1976:FDE

[Und76] R. G. Underwood. A functional differential equation approach to solving infinite games. SIAM Journal on Control and Optimization, 14(2): 251–267, February 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ugrinovskii:1999:ASM

[UP99] V. A. Ugrinovskii and I. R.

Petersen. Absolute stabilization and minimax optimal control of uncertain systems with stochastic uncertainty. SIAM Journal on Control and Optimization, 37(4):1089–1122, July 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30996.

Ugrinovskii:2002:MLC

[UP02] Valery A. Ugrinovskii and Ian R. Petersen. Minimax LQG control of stochastic partially observed uncertain systems. SIAM Journal on Control and Optimization, 40(4):1189–1226, July 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34935.

Urquiza:2005:REF

[Urq05] Jose Manuel Urquiza. Rapid exponential feedback stabilization with unbounded control operators. SIAM Journal on Control and Optimization, 43(6):2233–2244, November 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38845.

Ursescu:1982:TSC

[Urs82] Corneliu Ursescu. Tangent sets’ calculus and necessary conditions for extremality. SIAM Journal on Control and Optimization, 20(4):563–574, July 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ulbrich:2000:SCA

[UU00] Michael Ulbrich and Stefan Ulbrich. Superlinear convergence of affine-scaling interiorpoint Newton methods for infinite-dimensional nonlinear problems with pointwise bounds. SIAM Journal on Control and Optimization, 38(6):1938–1984, November 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/32591.

Ulbrich:1999:GCT

[UUH99] Michael Ulbrich, Stefan Ulbrich, and Matthias Heinkenschloss. Global convergence of trust-region interiorpoint algorithms for infinitedimensional nonconvex minimization subject to pointwise bounds. SIAM Journal on Control and Optimization, 37(3):731–764, May 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/31954.

Underwood:1979:NCN

[UY79] Robert G. Underwood and Donald F. Young. Null controllability of nonlinear functional differential equations. SIAM Journal on Control and Optimization, 17(6):753– 772, November 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Vega-Amaya:2003:ZSA

[VA03] Oscar Vega-Amaya. Zerosum average semi-Markov games: Fixed-point solutions of the Shapley equation. SIAM Journal on Control and Optimization, 42(5): 1876–1894, September 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/40842.

Valcher:2001:NRA

[Val01] Maria Elena Valcher. Nonnegative realization of autonomous systems in the behavioral approach. SIAM Journal on Control and Optimization, 40(2):540–556, March 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/37820.

vanderSchaft:1982:OCS

[van82] A. J. van der Schaft. Observability and controllability for smooth nonlinear systems. SIAM Journal on Control and Optimization, 20(3):338–354, May 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

vanEijndhoven:1986:SLC

[van86] J. T. J. van Eijndhoven. Solving the linear complementarity problem in circuit simulation. SIAM Journal on Control and Optimization, 24(5): 1050–1062, July 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

vanderSchaft:1987:SOC

[van87] A. J. van der Schaft. Symmetries in optimal control. SIAM Journal on Control and Optimization, 25(2):245– 259, March 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Vanderbei:1992:OSA

[Van92] Robert J. Vanderbei. Optimal switching among several Brownian motions. SIAM Journal on Control and Optimization, 30(5):1150–1162, September 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

VanKeulen:1994:RLC

[Van94] Bert Van Keulen. Redheffer’s lemma and H_{∞} control for infinite-dimensional systems. SIAM Journal on Control and Optimization, 32 (1):261–278, January 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/21741.

vanKeulen:1995:ECS

[van95] Bert van Keulen. Equivalent conditions for the solvability of the nonstandard LQ-problem for Pritchard–Salamon systems. SIAM Journal on Control and Optimization, 33(5): 1326–1356, September 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/22921.

vandenHof:1998:PLO

[van98] J. M. van den Hof. Positive linear observers for linear compartmental systems. SIAM Journal on Control and Optimization, 36(2):590–608, March 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30611.

Varaiya:1976:PSD

[Var76] Pravin Varaiya. N-player stochastic differential games. SIAM Journal on Control and Optimization, 14(3):538–545, May 1976. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Vanspranghe:2023:ORI

[VB23] Nicolas Vanspranghe and Lucas Brivadis. Output regulation of infinite-dimensional nonlinear systems: a forwarding approach for contraction semigroups. SIAM Journal on Control and Optimization, 61(4):2571–2594, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1473248.

Vellekoop:2003:NFA

[VC03] M. H. Vellekoop and J. M. C. Clark. A nonlinear filtering approach to changepoint detection problems: Direct and differential-geometric methods. SIAM Journal on Control and Optimization, 42(2):469–494, March 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/37595.

Vieira:2022:BSN

[VC22] Alexandre Vieira and Pierre-Henri Cocquet. The Boussinesq system with nonsmooth boundary conditions: Existence, relaxation, and topology optimization. SIAM Journal on Control and Optimization, 60(6):3457–3484, ???? 2022. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1465159.

vonDaniels:2015:CNT

[vDHV15] Nikolaus von Daniels, Michael Hinze, and Morten Vierling. Crank–Nicolson time stepping and variational discretization of control-constrained parabolic optimal control problems. SIAM Journal on Control and Optimization, 53 (3):1182–1198, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

vanderSchaft:2013:PHS

[vdSM13] A. J. van der Schaft and B. M.

Maschke. Port-hamiltonian systems on graphs. SIAM Journal on Control and Optimization, 51(2):906–937, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

vanderSchaft:2011:SMI

[vdSR11] Arjan van der Schaft and Paolo Rapisarda. State maps from integration by parts. SIAM Journal on Control and Optimization, 49(6):2415– 2439, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v49/i6/p2415_s1.

vanderWoude:1999:GNI

[vdW99] Jacob van der Woude. The generic number of invariant zeros of a structured linear system. SIAM Journal on Control and Optimization, 38(1):1–21, January 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/31011.

Veliov:1997:TDC

[Vel97] Vladimir Veliov. On the timediscretization of control systems. SIAM Journal on Control and Optimization, 35(5): 1470–1486, September 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/28898.

Verhaegen:1997:WSM

[Ver97] Michel Verhaegen. Weighted sensitivity minimization for causal, linear, discrete Time– Varying systems. SIAM Journal on Control and Optimization, 35(3):791–809, May 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/25339.

Vest:2013:RSS

[Ves13] Ambroise Vest. Rapid stabilization in a semigroup framework. SIAM Journal on Control and Optimization, 51(5):

4169–4188, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Verriest:1995:GAM

[VG95] Erik I. Verriest and W. Steven Gray. A geometric approach to the minimum sensitivity design problem. SIAM Journal on Control and Optimization, 33(3):863–881, May 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/22007.

Vasudevan:2013:CAOa

[VGBS13a] Ramanarayan Vasudevan, Humberto Gonzalez, Ruzena Bajcsy, and S. Shankar Sastry. Consistent approximations for the optimal control of constrained switched systems — Part 1: a conceptual algorithm. SIAM Journal on Control and Optimization, 51(6): 4463–4483, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Vasudevan:2013:CAOb

[VGBS13b] Ramanarayan Vasudevan, Humberto Gonzalez, Ruzena Bajcsy, and S. Shankar Sastry. Consistent approximations for the optimal control of constrained switched systems — Part 2: an implementable algorithm. SIAM Journal on Control and Optimization, 51(6):4484–4503, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

vanHandel:2006:AGS

[vH06]  Ramon van Handel.  Al-

most global stochastic stability. SIAM Journal on Control and Optimization, 45(4): 1297–1313, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

vanHandel:2010:WDN

[vH10] Ramon van Handel. When do nonlinear filters achieve maximal accuracy? SIAM Journal on Control and Optimization, 48(5):3151–3168, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Vidyasagar:1977:ICI

[Vid77] M. Vidyasagar. L_{2}-instability criteria for interconnected systems. SIAM Journal on Control and Optimization, 15(2): 312–328, February 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Vieille:2000:SSP

[Vie00]  Nicolas Vieille.  Solvable

states in n-player stochastic games. SIAM Journal on Control and Optimization, 38(6): 1794–1804, November 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34593.

Vigodner:1997:LSP

[Vig97] Alexander Vigodner. Limits of singularly perturbed control problems with statistical dynamics of fast motions. SIAM Journal on Control and Optimization, 35(1):1–28, January 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/26420.

Villa:2004:WPN

[Vil04] Silvia Villa. Well-posedness of nonconvex integral functionals. SIAM Journal on Control and Optimization, 43(4):1298–1312, July 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43728.

Vinter:1977:FSS

[Vin77] Richard B. Vinter. Filter stability for stochastic evolution equations. SIAM Journal on Control and Optimization, 15 (3):465–485, May 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Vinter:1980:CRS

[Vin80]  Richard Vinter.  A char-

acterization of the reachable set for nonlinear control systems. SIAM Journal on Control and Optimization, 18(6): 599–610, November 1980. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Vinter:1983:NGO

[Vin83a] R. B. Vinter. New global optimality conditions in optimal control theory. SIAM Journal on Control and Optimization, 21(2):235–245, March 1983. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Vinter:1983:WCE

[Vin83b] R. B. Vinter. Weakest conditions for existence of Lipschitz continuous Krotov functions in optimal control theory. SIAM Journal on Control and Optimization, 21(2): 215–234, March 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Vinter:1993:CDN

[Vin93] Richard Vinter. Convex duality and nonlinear optimal control. SIAM Journal on Control and Optimization, 31(2): 518–538, March 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Vinter:2005:MOC

[Vin05] R. B. Vinter. Minimax optimal control. SIAM Journal on Control and Optimization, 44(3):939–968, May 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41524.

Vinter:2014:HIN

[Vin14] Richard B. Vinter. The Hamiltonian inclusion for nonconvex velocity sets. SIAM Journal on Control and Optimization, 52(2):1237–1250, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Vinter:2019:OCP

[Vin19] Richard B. Vinter. Optimal control problems with time delays: Constancy of the Hamiltonian. SIAM Journal on Control and Optimization, 57 (4):2574–2602, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Vivalda:2003:GOU

[Viv03] J. C. Vivalda. On the genericity of the observability of uncontrolled discrete nonlinear systems. SIAM Journal on Control and Optimization, 42 (4):1509–1522, January 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Vinter:1977:OCN

[VJ77] Richard B. Vinter and Timothy L. Johnson. Optimal control of nonsymmetric hyperbolic systems in n variables on the half-space. SIAM Journal on Control and Optimization, 15(1):129–143, January 1977. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Vukovic:2024:NCU

[VJBK24] Manojlo Vukovic, Dusan Jakovetic, Dragana Bajovic, and Soummya Kar. Nonlinear consensus + innovations under correlated heavy-tailed noises: Mean square convergence rate and asymptotics. SIAM Journal on Control and Optimization, 62(1):376–399, January 2024. CODEN SJ-CODC. ISSN 1095-7138.

Vinter:1981:ITQ

[VK81] R. B. Vinter and R. H. Kwong. The infinite time quadratic control problem for linear systems with state and control delays: An evolution equation approach. SIAM Journal on Control and Optimization, 19(1):139–153, January 1981. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Vinter:1978:ESW

[VL78a] Richard B. Vinter and Richard M. Lewis. The equivalence of strong and weak formulations for certain problems in optimal control. SIAM Journal on Control and Optimization, 16 (4):546–570, July 1978. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Vinter:1978:NSC

[VL78b] Richard B. Vinter and Richard M.

Lewis. A necessary and sufficient condition for optimality of dynamic programming type, making no a priori assumptions on the controls. SIAM Journal on Control and Optimization, 16(4):571–583, July 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Vicinansa:2023:EEL

[VL23] Guilherme S. Vicinansa and Daniel Liberzon. Estimation entropy, Lyapunov exponents, and quantizer design for switched linear systems. SIAM Journal on Control and Optimization, 61(1):198–224, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1411871.

Vielma-Leal:2024:ECS

[VLP24] Francisco J. Vielma-Leal and Ademir Pastor. Exact controllability and stabilization for linear dispersive PDE’s on the two-dimensional torus. SIAM Journal on Control and Optimization, 62(1):539–562, February 2024. CODEN SJ-CODC. ISSN 1095-7138.

Vancostenoble:2000:OEE

[VM00] Judith Vancostenoble and Patrick Martinez. Optimality of energy estimates for the wave equation with nonlinear boundary velocity feedbacks. SIAM Journal on Control and Optimization, 39(3):776–797, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/35421.

vanNeerven:2004:NCA

[vN04] J. M. A. M. van Neerven. Null controllability and the algebraic Riccati equation in Banach spaces. SIAM Journal on Control and Optimization, 43(4):1313–1327, July 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/43705.

Viator:2022:SEN

[VO22] Robert Viator and Braxton Osting. Steklov eigenvalues of nearly spherical domains. SIAM Journal on Control and Optimization, 60 (3):1546–1562, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1411925.

Vogel:1990:CLS

[Vog90] Curtis R. Vogel. A constrained least squares regularization method for nonlinear ill-posed problems. SIAM Journal on Control and Optimization, 28(1):34–49, January 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Volkwein:2000:MIA

[Vol00] S. Volkwein. Mesh-independence for an augmented Lagrangian– SQP method in Hilbert spaces. SIAM Journal on Control and Optimization, 38(3):767–785, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33446.

Verdu:1987:ADP

[VP87] Sergio Verdú and H. Vincent Poor. Abstract dynamic programming models under commutativity conditions. SIAM Journal on Control and Optimization, 25(4): 990–1006, July 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Vinter:1988:MPO

[VP88] R. B. Vinter and F. M. F. L.

Pereira. A maximum principle for optimal processes with discontinuous trajectories. SIAM Journal on Control and Optimization, 26(1):205–229, January 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Vath:2007:ESO

[VP07] Vathana Ly Vath and Huyên Pham. Explicit solution to an optimal switching problem in the two-regime case. SIAM Journal on Control and Optimization, 46(2):395–426, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Vladimirov:2015:SDM

[VP15] Igor G. Vladimirov and Ian R. Petersen. State distributions and minimum relative entropy noise sequences in uncertain stochastic systems: The discrete-time case. SIAM Journal on Control and Optimization, 53(3):1107–1153, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Vladimirov:2022:MDO

[VP22a] Igor G. Vladimirov and Ian R. Petersen. Moment dynamics and observer design for a class of quasilinear quantum stochastic systems. SIAM Journal on Control and Optimization, 60(3):1223–1249, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1386529.

Vogt:2022:BOC

[VP22b] Ryan H. Vogt and N. Anders Petersson. Binary optimal control of single-flux-quantum pulse sequences. SIAM Journal on Control and Optimization, 60(6):3217–3236, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M142808X.

Vladimirov:2019:EPU

[VPJ19] Igor G. Vladimirov, Ian R.

Petersen, and Matthew R. James. Effects of parametric uncertainties in cascaded open quantum harmonic oscillators and robust generation of Gaussian invariant states. SIAM Journal on Control and Optimization, 57(3): 1597–1628, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Vrdoljak:2016:COD

[Vrd16] Marko Vrdoljak. Classical optimal design in two-phase conductivity problems. SIAM Journal on Control and Optimization, 54(4):2020–2035, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

vanNieuwstadt:1998:DFA

[vRM98] M. van Nieuwstadt, M. Rathinam, and R. M. Murray. Differential flatness and absolute equivalence of nonlinear control systems. SIAM Journal on Control and Optimization, 36(4):1225–1239, July 1998. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/27402.

VanHarten:1980:WPS

[VS80] Aart Van Harten and Hans Schumacher. Well-posedness of some evolution problems in the theory of automatic feedback control for systems with distributed parameters. SIAM Journal on Control and Optimization, 18(4):391–419, ???? 1980. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Voss:2014:PHM

[VS14] Thomas Voß and Jacquelien M. A. Scherpen. Port-Hamiltonian modeling of a nonlinear Timoshenko beam with piezo actuation. SIAM Journal on Control and Optimization, 52(1):493–519, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Vieten:2018:CFE

[VS18] Martin G. Vieten and Richard H. Stockbridge. Convergence of finite element methods for singular stochastic control. SIAM Journal on Control and Optimization, 56(6):4336–4364, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Vieten:2020:SSI

[VS20] Martin G. Vieten and Richard H.

Stockbridge. On the solution structure of infinitedimensional linear problems stemming from singular stochastic control problems. SIAM Journal on Control and Optimization, 58(6):3363–3388, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Vukosavljev:2018:RIL

[VSB18] Marijan Vukosavljev, Angela P. Schoellig, and Mireille E. Broucke. The regular indefinite linear quadratic optimal control problem: Stabilizable case. SIAM Journal on Control and Optimization, 56(1): 496–516, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Vicino:1995:RCS

[VT95] Antonio Vicino and Alberto Tesi. Regularity conditions for the stability margin problem with linear dependent perturbations. SIAM Journal on Control and Optimization, 33(4):1000–1016, July 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23461.

vanderLaan:1999:EAR

[vTY99] Gerard van der Laan, Dolf Talman, and Zaifu Yang. Existence and approximation of robust solutions of variational inequality problems over polytopes. SIAM Journal on Control and Optimization, 37(2):333–352, March 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30934.

Vuong:2021:SOD

[Vuo21] Phan Tu Vuong. A second order dynamical system and its discretization for strongly pseudo-monotone variational inequalities. SIAM Journal on Control and Optimization, 59(4):2875–2897, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

vanderSchaft:1984:NPS

[vW84] A. J. van der Schaft and J. C. Willems. A new procedure for stochastic realization of spectral density matrices. SIAM Journal on Control and Optimization, 22(6): 845–855, November 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Vinter:1990:HJT

[VW90] R. B. Vinter and P. Wolenski. Hamilton–Jacobi theory for optimal control problems with data measurable in time. SIAM Journal on Control and Optimization, 28(6):

1404–1419, November 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Vogel:1995:ACD

[VW95] C. R. Vogel and J. G. Wade.

Analysis of costate discretizations in parameter estimation for linear evolution equations. SIAM Journal on Control and Optimization, 33 (1):227–254, January 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/23244.

Volkwein:2002:AIC

[VW02] S. Volkwein and M. Weiser.

Affine invariant convergence analysis for inexact augmented Lagrangian–SQP methods. SIAM Journal on Control and Optimization, 41(3):875–899, May 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38308.

Vexler:2008:AFE

[VW08] B. Vexler and W. Wollner.

Adaptive finite elements for elliptic optimization problems with control constraints. SIAM Journal on Control and Optimization, 47(1):509–534, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

vanWaarde:2023:QMI

[vWCET23] Henk J. van Waarde, M. Kanat Camlibel, Jaap Eising, and Harry L. Trentelman. Quadratic matrix inequalities with applications to data-based control. SIAM Journal on Control and Optimization, 61 (4):2251–2281, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1486807.

Vinter:1997:EEL

[VZ97a] Richard Vinter and Harry Zheng. The extended Euler– Lagrange condition for nonconvex variational problems. SIAM Journal on Control and Optimization, 35(1):56– 77, January 1997. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/28313.

Volkov:1997:GSO

[VZ97b] Y. V. Volkov and S. K. Zavriev. A general stochastic outer approximations method. SIAM Journal on Control and Optimization, 35(4):1387–1421, July 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/26320.

Vettori:2000:CSD

[VZ00] P. Vettori and S. Zampieri.

Controllability of systems described by convolutional or delay-differential equations. SIAM Journal on Control and Optimization, 39(3):728–756, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35971.

Valein:2009:SWE

[VZ09] Julie Valein and Enrique Zuazua. Stabilization of the wave equation on 1-D networks. SIAM Journal on Control and Optimization, 48(4): 2771–2797, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Venel:2016:SUV

[VZ16] Xavier Venel and Bruno Ziliotto. Strong uniform value in gambling houses and partially observable Markov decision processes. SIAM Journal on Control and Optimization, 54(4):1983–2008, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Venel:2021:HDE

[VZ21] Xavier Venel and Bruno Ziliotto. History-dependent evaluations in partially observable Markov decision process. SIAM Journal on Control and Optimization, 59(2):

1730–1755, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wolovich:1984:CDE

[WA84] W. A. Wolovich and P. J. Antsaklis. The canonical Diophantine equations with applications. SIAM Journal on Control and Optimization, 22 (5):777–787, September 1984. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wachsmuth:2007:ASM

[Wac07] Daniel Wachsmuth. Analysis of the SQP-method for optimal control problems governed by the nonstationary Navier– Stokes equations based on L^{p}- theory. SIAM Journal on Control and Optimization, 46(3): 1133–1153, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wachsmuth:2012:OCQ

[Wac12]  Gerd Wachsmuth.  Opti-

mal control of quasi-static plasticity with linear kinematic hardening, Part I: Existence and discretization in time. SIAM Journal on Control and Optimization, 50(5): 2836–2861, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wachsmuth:2016:TMS

[Wac16] Gerd Wachsmuth. Towards mstationarity for optimal control of the obstacle problem with control constraints. SIAM Journal on Control and Optimization, 54(2):964–986, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wachsmuth:2019:IHT

[Wac19] Daniel Wachsmuth. Iterative hard-thresholding applied to optimal control problems with L^{0}(Ω) control cost. SIAM Journal on Control and Optimization, 57(2):854–879, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wagner:1977:SMS

[Wag77] Daniel H. Wagner. Survey of measurable selection theorems. SIAM Journal on Control and Optimization, 15(5): 859–903, August 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wagner:1991:NNS

[Wag91] K. G. Wagner. Nonlinear noninteraction with stability by dynamic state feedback. SIAM Journal on Control and Optimization, 29(3):609–622, May 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wakaiki:2021:SSS

[Wak21] Masashi Wakaiki. Strong stability of sampled-data Rieszspectral systems. SIAM Journal on Control and Optimization, 59(5):3498–3523, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:1991:OFG

[Wan91] Xiaochang Wang. On output feedback via Grassmannians. SIAM Journal on Control and Optimization, 29(4): 926–935, July 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:1994:DPA

[Wan94]  Xiaochang Wang.  De-

centralized pole assignment and product Grassmannians. SIAM Journal on Control and Optimization, 32(3):855–875, May 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/22333.

Wang:1998:FSB

[Wan98] Hualin Wang. Feedback stabilization of bilinear control systems. SIAM Journal on Control and Optimization, 36(5): 1669–1684, September 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30549.

Wang:2000:OCP

[Wan00]  Gengsheng Wang.  Opti-

mal control of parabolic differential equations with two point boundary state constraints. SIAM Journal on Control and Optimization, 38(5):1639–1654, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/33813.

Wang:2002:OCD

[Wan02]  Gengsheng Wang.  Opti-

mal controls of 3-dimensional Navier–Stokes equations with state constraints. SIAM Journal on Control and Optimization, 41(2):583–606, March 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38576.

Wang:2008:NCH

[Wan08] Gengsheng Wang. L^{∞}-null controllability for the heat equation and its consequences for the time optimal control problem. SIAM Journal on Control and Optimization, 47 (4):1701–1720, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2021:OCS

[Wan21] Huaqiao Wang. Optimal controls for the stochastic compressible Navier–Stokes equations. SIAM Journal on Control and Optimization, 59(5): 3661–3682, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2022:BSV

[Wan22] Tianxiao Wang. Backward stochastic Volterra integrodifferential equations and applications in optimal control problems. SIAM Journal on Control and Optimization, 60 (4):2393–2419, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1371464.

Warfield:1976:SMP

[War76a]  Virginia M. Warfield.  A

stochastic maximum principle. SIAM Journal on Control and Optimization, 14(5): 803–826, August 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Warga:1976:CNC

[War76b] J. Warga. Controllability and necessary conditions in unilateral problems without differentiability assumptions. SIAM Journal on Control and Optimization, 14(3):546–573, May 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Warga:1977:SDR

[War77] J. Warga. Steepest descent with relaxed controls. SIAM Journal on Control and Optimization, 15(4):674–682, July 1977. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Warga:1978:CMR

[War78a] J. Warga. Controllability and a multiplier rule for nondifferentiable optimization problems. SIAM Journal on Control and Optimization, 16 (5):803–812, September 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Warga:1978:CNH

[War78b]  J. Warga.  Controllability

of nondifferentiable hereditary processes. SIAM Journal on Control and Optimization, 16 (5):813–831, September 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Warga:1982:IPC

[War82] J. Warga. Iterative procedures for constrained and unilateral optimization problems. SIAM Journal on Control and Optimization, 20(3):360–376, May 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Warga:1983:OCD

[War83]  J. Warga.  Optimization

and controllability without differentiability assumptions. SIAM Journal on Control and Optimization, 21(6):837– 855, November 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Warga:1984:SON

[War84] J. Warga. Second order necessary conditions in optimization. SIAM Journal on Control and Optimization, 22(4): 524–528, July 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Warga:1985:SOC

[War85] J. Warga. Second order controllability and optimization with ordinary controls. SIAM Journal on Control and Optimization, 23(1):49–60, January 1985. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Warga:1986:HOC

[War86] J. Warga. Higher order conditions with and without Lagrange multipliers. SIAM Journal on Control and Optimization, 24(4):715–730, July 1986. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Warga:1988:HOC

[War88] J. Warga. Higher-order conditions for conical controllability. SIAM Journal on Control and Optimization, 26(6):1471– 1480, November 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Warga:1989:GDC

[War89] J. Warga. Global directional controllability. SIAM Journal on Control and Optimization, 27(5):976–990, September 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Warga:1991:NPC

[War91] J. Warga. Nonsmooth problems with conflicting controls. SIAM Journal on Control and Optimization, 29(3):678–701, May 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Warma:2019:ACE

[War19]  Mahamadi Warma.  Ap-

proximate controllability from the exterior of space–time fractional diffusive equations. SIAM Journal on Control and Optimization, 57(3):2037– 2063, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Washburn:1979:BBI

[Was79] Don Washburn. A bound on the boundary input map for parabolic equations with application to time optimal control. SIAM Journal on Control and Optimization, 17(5):652– 671, September 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Watson:1979:SNC

[Wat79] Layne T. Watson. Solving the nonlinear complementarity problem by a homotopy method. SIAM Journal on Control and Optimization, 17 (1):36–46, January 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Watanabe:2024:MCA

[Wat24] Itsuki Watanabe. Markov chain approximation for Hamilton– Jacobi–Bellman equation with absorbing boundary. SIAM Journal on Control and Optimization, 62(2):1152–1164, March 2024. CODEN SJ-CODC. ISSN 1095-7138.

Ward:1987:NCF

[WB87] D. E. Ward and J. M. Borwein. Nonsmooth calculus in finite dimensions. SIAM Journal on Control and Optimization, 25(5):1312–1340, September 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Willems:1981:DDM

[WC81] Jan C. Willems and Christian Commault. Disturbance decoupling by measurement feedback with stability or pole placement. SIAM Journal on Control and Optimization, 19 (4):490–504, July 1981. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:1989:ABS

[WC89] Han Kun Wang and Goong Chen. Asymptotic behavior of solutions of the onedimensional wave equation with a nonlinear boundary stabilizer. SIAM Journal on Control and Optimization, 27 (4):758–775, July 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:1991:ALE

[WC91] Han Kun Wang and Goong Chen. Asymptotic locations of eigenfrequencies of Euler–Bernoulli beam with nonhomogeneous structural and viscous damping coefficients. SIAM Journal on Control and Optimization, 29(2): 347–367, March 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:1996:CBI

[WC96a] Chi-Jo Wang and J. Daniel Cobb. A characterization of bounded-input boundedoutput stability for linear time-invariant systems with distributional inputs. SIAM Journal on Control and Optimization, 34(3):987–1000, May 1996. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/24837.

Wei:1996:MFC

[WC96b] Yuanjun Wei and Peter E. Caines. On Markovian fragments of COCOLOG for logic control systems. SIAM Journal on Control and Optimization, 34(5): 1707–1733, September 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/25453.

Wei:2019:NSE

[WC19] Qingda Wei and Xian Chen. Nonzero-sum expected average discrete-time stochastic games: The case of uncountable spaces. SIAM Journal on Control and Optimization, 57(6):4099–4124, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2004:PAV

[WCCL04] J. N. Wang, S. H. Chou, Y. C. Chen, and W. W. Lin. Pole assignment for a vibrating system with aerodynamic effect. SIAM Journal on Control and Optimization, 42(6):2116–2129, November 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41156.

Wen:2014:WPE

[WCG14] Ruili Wen, Shugen Chai, and Bao-Zhu Guo. Well-posedness and exact controllability of fourth order Schrödinger equation with boundary control and collocated observation. SIAM Journal on Control and Optimization, 52(1):365–396, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Winkin:2004:SFS

[WCJP04] J. J. Winkin, F. M. Callier, B. Jacob, and J. R. Partington. Spectral factorization by symmetric extraction for distributed parameter systems. SIAM Journal on Control and Optimization, 43(4):1435–1466, July 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/41645.

Wu:2005:SAC

[WCLC05] Jun Wu, Sheng Chen, Gang Li, and Jian Chu. A search algorithm for a class of optimal finite-precision controller realization problems with saddle points. SIAM Journal on Control and Optimization, 44(5):1787–1810, September 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/43508.

Wu:2012:BCV

[WCS12] Zhaojing Wu, Mingyue Cui, and Peng Shi. Backstepping control in vector form for stochastic Hamiltonian systems. SIAM Journal on Control and Optimization, 50(2): 925–942, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:1976:CFL

[WD76] S. H. Wang and E. J. Davison. Canonical forms of linear multivariable systems. SIAM Journal on Control and Optimization, 14(2):236– 250, February 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wan:1979:EOC

[WD79] C. B. Wan and M. H. A. Davis. Existence of optimal controls for stochastic jump processes. SIAM Journal on Control and Optimization, 17 (4):511–524, July 1979. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Weiser:2007:ICP

[WD07] Martin Weiser and Peter Deuflhard. Inexact central path following algorithms for optimal control problems. SIAM Journal on Control and Optimization, 46(3):792–815, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2014:CEN

[WD14] Chunpeng Wang and Runmei Du. Carleman estimates and null controllability for a class of degenerate parabolic equations with convection terms. SIAM Journal on Control and Optimization, 52(3):1457–1480, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Weinert:1979:ASS

[WDS79] H. L. Weinert, U. B. Desai, and G. S. Sidhu. Arma splines, system inverses, and least-squares estimates. SIAM Journal on Control and Optimization, 17(4):525–536, July 1979. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Weck:1982:RCS

[Wec82] N. Weck. A remark on controllability for symmetric hyperbolic systems in one space dimension. SIAM Journal on Control and Optimization, 20 (1):1–8, January 1982. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Weck:1984:MSR

[Wec84]  N. Weck.  More states

reachable by boundary control of the heat equation. SIAM Journal on Control and Optimization, 22(5):699– 710, September 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Weck:2000:EBC

[Wec00] N. Weck. Exact boundary controllability of a Maxwell problem. SIAM Journal on Control and Optimization, 38(3):736–750, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/34755.

Weerasinghe:1992:FFS

[Wee92] A. P. N. Weerasinghe. A finite fuel stochastic control problem on a finite time horizon. SIAM Journal on Control and Optimization, 30(6):1395– 1408, November 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Weerasinghe:2003:MIT

[Wee03] Ananda Weerasinghe. Minimizing infinite time horizon discounted cost with mean, variance, and bounded variation controls. SIAM Journal on Control and Optimization, 42(4):1395–1415, January 2003. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Weerasinghe:2005:BVC

[Wee05]  Ananda Weerasinghe.  A

bounded variation control problem for diffusion processes. SIAM Journal on Control and Optimization, 44(2):389–417, March 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43611.

Weerasinghe:2007:ALA

[Wee07] Ananda Weerasinghe. An Abelian limit approach to a singular ergodic control problem. SIAM Journal on Control and Optimization, 46(2): 714–737, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Weerasinghe:2018:CRM

[Wee18] Ananda Weerasinghe. Controlling the running maximum of a diffusion process and an application to queueing systems. SIAM Journal on Control and Optimization, 56(2): 1412–1440, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2022:IBC

[WeFC22] Biao Wang, Jun e Feng, and Daizhan Cheng. On identification of Boolean control networks. SIAM Journal on Control and Optimization, 60 (3):1591–1612, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M1373773.

Weiner:1977:CGP

[Wei77] L. B. Weiner. Comments on: “Generalized Prediction-Correction Estimation”. SIAM Journal on Control and Optimization, 15(5):857–858, August 1977. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [KH69].

Weiss:1989:AUC

[Wei89] George Weiss. Admissibility of unbounded control operators. SIAM Journal on Control and Optimization, 27(3): 527–545, May 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Weiser:2005:IPM

[Wei05]  Martin Weiser.  Interior

point methods in function space. SIAM Journal on Control and Optimization, 44(5): 1766–1786, September 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/43727.

Wenk:2015:CSG

[Wen15] Carl J. Wenk. Control sequencing in a game of identity pursuit-evasion. SIAM Journal on Control and Optimization, 53(4):1815–1841, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wolovich:1976:ICF

[WF76] W. A. Wolovich and P. L.

Falb. Invariants and canonical forms under dynamic compensation. SIAM Journal on Control and Optimization, 14 (6):996–1008, November 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2024:SDF

[WF24] Pengfei Wang and Emilia Fridman. Sampled-data finite-dimensional observerbased control of 1D stochastic parabolic PDEs. SIAM Journal on Control and Optimization, 62(1):297–325, January 2024. CODEN SJCODC. ISSN 1095-7138.

Waeber:2013:BSN

[WFH13] Rolf Waeber, Peter I. Frazier, and Shane G. Henderson. Bisection search with noisy responses. SIAM Journal on Control and Optimization, 51 (3):2261–2279, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2020:CSS

[WFZ20] Hongxia Wang, Minyue Fu, and Huanshui Zhang. H_{∞} control for stochastic systems with disturbance preview. SIAM Journal on Control and Optimization, 58(1): 215–242, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2011:WES

[WGK11] Jun-Min Wang, Bao-Zhu Guo, and Miroslav Krstic. Wave equation stabilization by delays equal to even multiples of the wave propagation time. SIAM Journal on Control and Optimization, 49(2): 517–554, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicon/resource/
                                        1/sjcodc/v49/i2/p517_s1.

Wu:2020:OSS

[WH20] Zeyang Wu and Qie He. Optimal switching sequence for switched linear systems. SIAM Journal on Control and Optimization, 58(2):1183–1206, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Willner:1992:PCU

[WHB92] Yosef Willner, Michael Heymann, and Marc Bodson. Parameter convergence and uniqueness in nonminimal parameterizations for multivariable adaptive control. SIAM Journal on Control and Optimization, 30(4):800–815, July 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wei:2003:CMC

[WHC03] T. Wei, Y. C. Hon, and J. Cheng. Computation for MultiDimensional Cauchy problem. SIAM Journal on Control and Optimization, 42(2):381–396, March 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38939.

Wu:2022:SSM

[WHGY22] Hao Wu, Junhao Hu, Shuaibin Gao, and Chenggui Yuan. Stabilization of stochastic McKean–Vlasov equations with feedback control based on discrete-time state observation. SIAM Journal on Control and Optimization, 60 (5):2884–2901, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1454997.

White:1979:CPP

[Whi79] L. W. White. Control of a pseudo-parabolic initial-value problem to a target function. SIAM Journal on Control and Optimization, 17 (5):587–595, September 1979. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

White:1980:CPP

[Whi80a] L. W. White. Controllability properties of pseudo-parabolic boundary control problems. SIAM Journal on Control and Optimization, 18(5):534– 539, September 1980. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Whitt:1980:RAN

[Whi80b] Ward Whitt. Representation and approximation of noncooperative sequential games. SIAM Journal on Control and Optimization, 18(1):33– 48, January 1980. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

White:1982:PCS

[Whi82] L. W. White. Point control with state constraints. SIAM Journal on Control and Optimization, 20(4):490–496, July 1982. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2019:MTI

[WHS19] Tao Wang, Yu Huang, and Hai-Wei Sun. Measuretheoretic invariance entropy for control systems. SIAM Journal on Control and Optimization, 57(1):310–333, ????

2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Willemstein:1977:ORN

[Wil77a] A. P. Willemstein. Optimal regulation of nonlinear dynamical systems on a finite interval. SIAM Journal on Control and Optimization, 15(6):1050–1069, November 1977. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wilson:1977:DGN

[Wil77b]  D. J. Wilson.  Differen-

tial games with no information. SIAM Journal on Control and Optimization, 15(2): 233–246, February 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wimmer:1992:USD

[Wim92] H. K. Wimmer. Unmixed solutions of the discrete-time algebraic Riccati equation. SIAM Journal on Control and Optimization, 30(4):867–878, July 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wimmer:1994:DPS

[Wim94] Harald K. Wimmer. Decomposition and parametrization of semidefinite solutions of the continuous-time algebraic Riccati equation. SIAM Journal on Control and Optimization, 32(4):995–1007, July 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/22922.

Wimmer:2006:PSD

[Wim06]  Harald K. Wimmer.  A

parametrization of solutions of the discrete-time algebraic Riccati equation based on pairs of opposite unmixed solutions. SIAM Journal on Control and Optimization, 44 (6):1992–2005, January 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Winston:1978:SGM

[Win78] Wayne Winston. A stochastic game model of a weapons development competition. SIAM Journal on Control and Optimization, 16(3):411–419, May 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Winston:1982:HM

[Win82] Elliot Winston. Hedging and maxmin. SIAM Journal on Control and Optimization, 20 (1):14–19, January 1982. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wirth:1998:DTV

[Wir98] Fabian Wirth. Dynamics of time-varying discrete-time linear systems: Spectral theory and the projected system. SIAM Journal on Control and Optimization, 36(2):447–487, March 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/29960.

Wirth:2005:CLT

[Wir05] Fabian Wirth. A converse Lyapunov theorem for linear parameter-varying and linear switching systems. SIAM Journal on Control and Optimization, 44(1):210–239, January 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/43479.

Wu:2016:OCO

[WJE16] Xueran Wu, Birgit Jacob, and Hendrik Elbern. Optimal control and observation locations for time-varying systems on a finite-time horizon. SIAM Journal on Control and Optimization, 54(1):291–316, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2019:MVP

[WJW19] Tianxiao Wang, Zhuo Jin, and Jiaqin Wei. Meanvariance portfolio selection under a non-Markovian regimeswitching model: Timeconsistent solutions. SIAM Journal on Control and Optimization, 57(5):3249–3271, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wyss:2012:HRE

[WJZ12] C. Wyss, B. Jacob, and H. J. Zwart. Hamiltonians and Riccati equations for linear systems with unbounded control and observation operators. SIAM Journal on Control and Optimization, 50(3): 1518–1547, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wierzbicki:1977:PCP

[WK77] Andrzej P. Wierzbicki and Stanislaw Kurcyusz. Projection on a cone, penalty functionals and duality theory for problems with inequaltity constraints in Hilbert space. SIAM Journal on Control and Optimization, 15(1):25– 56, January 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Walker:2013:ASO

[WK13] Shawn W. Walker and Eric E. Keaveny. Analysis of shape optimization for magnetic microswimmers. SIAM Journal on Control and Optimization, 51(4):3093–3126, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Willems:1986:SOC

[WKS86] J. C. Willems, A. Kitapçi, and L. M. Silverman. Singular optimal control: a geometric approach. SIAM Journal on Control and Optimization, 24 (2):323–337, March 1986. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:1995:NBV

[WL95] Huai Zhong Wang and Yong Li. Neumann boundary value problems for secondorder ordinary differential equations across resonance. SIAM Journal on Control and Optimization, 33(5): 1312–1325, September 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/24532.

Wang:2017:FEO

[WL17] Shuo Wang and Jr-Shin Li.

Fixed-endpoint optimal control of bilinear ensemble systems. SIAM Journal on Control and Optimization, 55(5): 3039–3065, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wei:2019:RLI

[WL19] Yusheng Wei and Zongli Lin.

Regulation of linear input delayed systems without delay knowledge. SIAM Journal on Control and Optimization, 57 (2):999–1022, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wei:2021:MDA

[WL21] Yusheng Wei and Zongli Lin. A memoryless delay-adaptive feedback law for the regulation of discrete-time linear systems. SIAM Journal on Control and Optimization, 59(4): 2756–2773, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2024:GOF

[WL24] Yuan Wang and Yungang Liu. Global output-feedback control by exploiting high-gain dynamic-compensation mechanisms. SIAM Journal on Control and Optimization, 62 (2):1122–1151, March 2024. CODEN SJCODC. ISSN 1095-7138.

Wang:2022:DAN

[WLZ22] Ying Wang, Wuquan Li, and Ji-Feng Zhang. Distributed adaptive nonlinear control with fusion leastsquares. SIAM Journal on Control and Optimization, 60 (5):3148–3172, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1419763.

Woittennek:2010:CNS

[WM10] Frank Woittennek and Hugues Mounier. Controllability of networks of spatially onedimensional second order PDEs—an algebraic approach. SIAM Journal on Control and Optimization, 48(6):3882– 3902, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wu:2012:STB

[WM12] Zhi-Gang Wu and Mehran Mesbahi. Symplectic transformation based analytical and numerical methods for linear quadratic control with hard terminal constraints. SIAM Journal on Control and Optimization, 50(2):652–671, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wu:2022:DET

[WMWL22] Xiaoqun Wu, Bing Mao, Xiuqi Wu, and Jinhu Lu. Dynamic event-triggered leaderfollower consensus control for MultiAgent systems. SIAM Journal on Control and Optimization, 60(1):189–209, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/20M1321152.

Wakaiki:2018:LOS

[WOH18] Masashi Wakaiki, Masaki Ogura, and João P. Hespanha. LQ-optimal sampleddata control under stochastic delays: Gridding approach for stabilizability and detectability. SIAM Journal on Control and Optimization, 56(4): 2634–2661, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wolenski:1990:EFR

[Wol90] Peter R. Wolenski. The exponential formula for the reachable set of a Lipschitz differential inclusion. SIAM Journal on Control and Optimization, 28(5):1148–1161, September 1990. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wollner:2012:OCE

[Wol12] W. Wollner. Optimal control of elliptic equations with pointwise constraints on the gradient of the state in nonsmooth polygonal domains. SIAM Journal on Control and Optimization, 50(4):2117– 2129, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wong:2009:CCC

[Won09] Wing Shing Wong. Control communication complexity of distributed control systems. SIAM Journal on Control and Optimization, 48(3): 1722–1742, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wood:2000:BAP

[WORO00] J. Wood, U. Oberst, E. Rogers, and D. H. Owens. A behavioral approach to the pole structure of one-dimensional and multidimensional linear systems. SIAM Journal on Control and Optimization, 38(2):627–661, March 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33213.

Williamson:1976:RCC

[WP76] L. J. Williamson and E. Polak. Relaxed controls and the convergence of optimal control algorithms. SIAM Journal on Control and Optimization, 14 (4):737–756, July 1976. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ward:1998:CRW

[WP98] Nicholas F. Dudley Ward and Jonathan R. Partington. A construction of rational wavelets and frames in Hardy–Sobolev spaces with applications to system modeling. SIAM Journal on Control and Optimization, 36(2):654–679, March 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/29733.

Wang:2018:FTA

[WP18] Yingfei Wang and Warren B. Powell. Finite-time analysis for the knowledge-gradient policy. SIAM Journal on Control and Optimization, 56(2): 1105–1129, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2024:DGC

[WQS^{+}24] Xiaoling Wang, Juan Qian, Housheng Su, Xiujuan Lu, and James Lam. Distributed global consensus of LTI mass with heterogeneous actuator saturation and communication noises. SIAM Journal on Control and Optimization, 62(3):1690–1716, June 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/23M1584678.

Wonham:1987:SCS

[WR87] W. M. Wonham and P. J. Ramadge. On the supremal controllable sublanguage of a given language. SIAM Journal on Control and Optimization, 25(3):637–659, May 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Weiss:2000:OEI

[WR00] George Weiss and Richard Rebarber. Optimizability and estimatability for infinite-dimensional linear systems. SIAM Journal on Control and Optimization, 39(4):1204–1232, July 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33519.

Weber:2014:CSC

[WR14] Alexander Weber and Gunther Reissig. Classical and strong convexity of sublevel sets and application to attainable sets of nonlinear systems. SIAM Journal on Control and Optimization, 52(5): 2857–2876, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Worthmann:2014:RSS

[WRGA14] Karl Worthmann, Marcus Reble, Lars Grüne, and Frank Allgöwer. The role of sampling for stability and performance in unconstrained nonlinear model predictive control. SIAM Journal on Control and Optimization, 52(1): 581–605, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wright:1989:CSL

[Wri89] Stephen Wright. Convergence of SQP-like methods for constrained optimization. SIAM Journal on Control and Optimization, 27(1):13–26, January 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wright:1993:ISC

[Wri93] Stephen J. Wright. Identifiable surfaces in constrained optimization. SIAM Journal on Control and Optimization, 31 (4):1063–1079, July 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wright:1995:CPD

[Wri95] Stephen E. Wright. Consistency of primal-dual approximations for convex optimal control problems. SIAM Journal on Control and Optimization, 33(5):1489–1509, September 1995. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/24050.

Wyman:1987:MTZ

[WS87] Bostwick F. Wyman and Michael K. Sain. Module theoretic zero structures for system matrices. SIAM Journal on Control and Optimization, 25 (1):86–99, January 1987. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:1990:NSC

[WS90] Qinghong Wang and Jason L.

Speyer. Necessary and sufficient conditions for local optimality of a periodic process. SIAM Journal on Control and Optimization, 28(2): 482–497, March 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:1992:ADE

[WS92] Yuan Wang and Eduardo D.

Sontag. Algebraic differential equations and rational control systems. SIAM Journal on Control and Optimization, 30(5):1126–1149, September 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:1995:OIO

[WS95] Yuan Wang and Eduardo D.

Sontag. Orders of input/output differential equations and state-space dimensions. SIAM Journal on Control and Optimization, 33(4):1102–1126, July 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/24682.

Wang:2013:UCR

[WS13a] Xiao Wang and Jinglai Shen.

Uniform convergence and rate adaptive estimation of convex functions via constrained optimization. SIAM Journal on Control and Optimization, 51(4):2753–2787, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Weiss:2013:MES

[WS13b] George Weiss and Olof J. Staffans. Maxwell’s equations as a scattering passive linear system. SIAM Journal on Control and Optimization, 51(5):3722–3756, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wakaiki:2020:ETC

[WS20] Masashi Wakaiki and Hideki Sano. Event-triggered control of infinite-dimensional systems. SIAM Journal on Control and Optimization, 58(2): 605–635, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wood:2004:CSI

[WSR04] J. Wood, V. R. Sule, and E. Rogers. Causal and stable input/output structures on multidimensional behaviors. SIAM Journal on Control and Optimization, 43(4):1493–1520, July 2004. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42997.

Wang:2021:IOD

[WSZ^{+}21] Xiaoling Wang, Housheng Su, Fan Zhang, Ali Zemouche, and Guanrong Chen. Interval observer design and consensus of MultiAgent systems with time-varying interval uncertainties. SIAM Journal on Control and Optimization, 59(5):3392–3417, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wu:1983:OCP

[WT83] Z. S. Wu and K. L. Teo. Optimal control problems involving second boundary value problems of parabolic type. SIAM Journal on Control and Optimization, 21(5):729– 757, September 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Willems:1998:QDF

[WT98] J. C. Willems and H. L.

Trentelman. On quadratic differential forms. SIAM Journal on Control and Optimization, 36(5):1703–1749, September 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30306.

Wen:2021:ALC

[WTJY21] Liyan Wen, Gang Tao, Bin Jiang, and Hao Yang. Adaptive LQ control using reduced Hamiltonian for continuoustime systems with unmatched input disturbances. SIAM Journal on Control and Optimization, 59(5):3625–3660, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wu:2018:EVD

[Wu18] Xiaochi Wu. Existence of value for a differential game with incomplete information and revealing. SIAM Journal on Control and Optimization, 56(4):2536–2562, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Willems:1983:RSU

[WW83] Jacques L. Willems and Jan C. Willems. Robust stabilization of uncertain systems. SIAM Journal on Control and Optimization, 21(3):352–374, May 1983. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2002:SCO

[WW02] Gengsheng Wang and Lijuan Wang. State-constrained optimal control governed by Non– Well-Posed parabolic differential equations. SIAM Journal on Control and Optimization, 40(5):1517–1539, September 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37700.

Wang:2003:OTC

[WW03] Lijuan Wang and Gengsheng Wang. The optimal time control of a phase-field system. SIAM Journal on Control and Optimization, 42(4): 1483–1508, January 2003. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wachsmuth:2016:ECF

[WW16] D. Wachsmuth and J.-E.

Wurst. Exponential convergence of hp-finite element discretization of optimal boundary control problems with elliptic partial differential equations. SIAM Journal on Control and Optimization, 54(5): 2526–2552, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2023:EBC

[WW23] Ke Wang and Libin Wang.

Exact boundary controllability of nodal profile for quasilinear hyperbolic systems and its asymptotic stability. SIAM Journal on Control and Optimization, 61(1):1–21, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1464270.

Wang:2021:SIS

[WWS21] Pengfei Wang, Shuaiqi Wang, and Huan Su. Stochastic input-to-state stability of impulsive stochastic nonlinear systems in infinite dimensions. SIAM Journal on Control and Optimization, 59(4): 2774–2797, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2013:MPF

[WWX13] Guangchen Wang, Zhen Wu, and Jie Xiong. Maximum principles for forward-backward stochastic control systems with correlated state and observation noises. SIAM Journal on Control and Optimization, 51(1):491–524, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2006:SOE

[WWY06] Gengsheng Wang, Lijuan Wang, and Donghui Yang. Shape optimization of an elliptic equation in an exterior domain. SIAM Journal on Control and Optimization, 45(2): 532–547, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wei:2014:SWS

[WWZ14] Lifeng Wei, Zhen Wu, and Huaizhong Zhao. Sobolev weak solutions of the Hamilton– Jacobi–Bellman equations. SIAM Journal on Control and Optimization, 52(3):1499– 1526, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2013:ETD

[WX13] Gengsheng Wang and Yashan Xu. Equivalence of three different kinds of optimal control problems for heat equations and its applications. SIAM Journal on Control and Optimization, 51(2):848–880, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2005:ESL

[WXY05] Jun-Min Wang, Gen-Qi Xu, and Siu-Pang Yung. Exponential stabilization of laminated beams with structural damping and boundary feedback controls. SIAM Journal on Control and Optimization, 44(5):1575–1597, September 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/61000.

Wang:2015:ASB

[WXZ15] Gengsheng Wang, Yashan Xu, and Yubiao Zhang. Attainable subspaces and the bangbang property of time optimal controls for heat equations. SIAM Journal on Control and Optimization, 53(2): 592–621, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wu:2006:CEA

[WY06] Xi Wu and Stephen S.-T. Yau. Classification of estimation algebras with state dimension 2. SIAM Journal on Control and Optimization, 45(3): 1039–1073, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wu:2008:DPP

[WY08] Zhen Wu and Zhiyong Yu. Dynamic programming principle for one kind of stochastic recursive optimal control problem and Hamilton–Jacobi– Bellman equation. SIAM Journal on Control and Optimization, 47(5):2616–2641, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2016:BBP

[WY16] Lijuan Wang and Qishu Yan. Bang-bang property of time optimal null controls for some semilinear heat equation. SIAM Journal on Control and Optimization, 54(6): 2949–2964, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wei:2021:IHF

[WY21] Qingmeng Wei and Zhiyong Yu. Infinite horizon forward-backward SDEs and open-loop optimal controls for stochastic linear-quadratic problems with random coefficients. SIAM Journal on Control and Optimization, 59(4): 2594–2623, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2023:SVS

[WY23] Tianxiao Wang and Jiongmin Yong. Spike variations for stochastic Volterra integral equations. SIAM Journal on Control and Optimization, 61(6):3608–3634, December 2023. CODEN SJCODC. ISSN 1095-7138.

Wynn:2009:CDC

[Wyn09] Andrew Wynn. Counterexamples to the discrete and continuous weighted Weiss conjectures. SIAM Journal on Control and Optimization, 48(4): 2620–2635, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wei:2017:TIR

[WYY17] Qingmeng Wei, Jiongmin Yong, and Zhiyong Yu. Timeinconsistent recursive stochastic optimal control problems. SIAM Journal on Control and Optimization, 55(6):4156– 4201, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2021:CAN

[WYY21] Lijuan Wang, Qishu Yan, and Huaiqiang Yu. Constrained approximate null controllability of the coupled heat equation with impulse controls. SIAM Journal on Control and Optimization, 59(5): 3418–3446, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wong:2017:UDR

[WYZ17] K. C. Wong, S. C. P. Yam, and H. Zheng. Utility-deviation-risk portfolio selection. SIAM Journal on Control and Optimization, 55(3):1819–1861, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2023:LQO

[WYZ23] Hanxiao Wang, Jiongmin Yong, and Chao Zhou. Linearquadratic optimal controls for stochastic Volterra integral equations: Causal state feedback and path-dependent Riccati equations. SIAM Journal on Control and Optimization, 61(4):2595–2629, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1492696.

Williams:1977:RDO

[WZ77] N. S. Williams and V. Zakian. A ring of delay operators with applications to delay-differential systems. SIAM Journal on Control and Optimization, 15(2):247– 255, February 1977. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Warga:1994:EED

[WZ94] J. Warga and Q. J. Zhu.

The equivalence of extremals in different representations of unbounded control problems. SIAM Journal on Control and Optimization, 32(4):1151–1169, July 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/23401.

Wolenski:1998:PAM

[WZ98] Peter R. Wolenski and Yu Zhuang.

Proximal analysis and the minimal time function. SIAM Journal on Control and Optimization, 36(3):1048–1072, May 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/29933.

Wu:2001:SFC

[WZ01] Hanzhong Wu and Xun Yu Zhou. Stochastic frequency characteristics. SIAM Journal on Control and Optimization, 40(2):557–576, March 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37375.

Wolenski:2007:SMA

[WŽ07] Peter R. Wolenski and Stanislav Žabić. A sampling method and approximation results for impulsive systems. SIAM Journal on Control and Optimization, 46(3):983–998, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Weiss:2009:WPC

[WZ09] George Weiss and Xiaowei Zhao. Well-posedness and controllability of a class of coupled linear systems. SIAM Journal on Control and Optimization, 48(4):2719–2750, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2012:MFG

[WZ12a] Bing-Chang Wang and Ji-Feng Zhang. Mean field games for large-population multiagent systems with Markov jump parameters. SIAM Journal on Control and Optimization, 50(4):2308–2334, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2012:AOT

[WZ12b] Gengsheng Wang and Guojie Zheng. An approach to the optimal time for a time optimal control problem of an internally controlled heat equation. SIAM Journal on Control and Optimization, 50(2): 601–628, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2012:EMT

[WZ12c] Gengsheng Wang and Enrique Zuazua. On the equivalence of minimal time and minimal norm controls for internally controlled heat equations. SIAM Journal on Control and Optimization, 50(5): 2938–2958, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2017:SOM

[WZ17a] Bing-Chang Wang and Ji-Feng Zhang. Social optima in mean field linear-quadratic-Gaussian models with Markov jump parameters. SIAM Journal on Control and Optimization, 55(1):429–456, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2017:OIM

[WZ17b] Gengsheng Wang and Can Zhang. Observability inequalities from measurable sets for some abstract evolution equations. SIAM Journal on Control and Optimization, 55(3): 1862–1886, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2017:OCP

[WZ17c] Tianxiao Wang and Haisen Zhang. Optimal control problems of forward-backward stochastic Volterra integral equations with closed control regions. SIAM Journal on Control and Optimization, 55 (4):2574–2602, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2021:CLE

[WZ21] Tianxiao Wang and Harry Zheng. Closed-loop equilibrium strategies for general time-inconsistent optimal control problems. SIAM Journal on Control and Optimization, 59(5):3152–3178, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2024:SET

[WZ24] Ziyuan Wang and Zhou Zhou. Sharp equilibria for timeinconsistent mean-field stopping games. SIAM Journal on Control and Optimization, 62(4):2319–2345, August 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1625512.

Wu:2020:GST

[WZC20] Shu-Lin Wu, Tao Zhou, and Xiaojun Chen. A Gauss–Seidel type method for dynamic nonlinear complementarity problems. SIAM Journal on Control and Optimization, 58(6): 3389–3412, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wan:2023:SSN

[WZD^{+}23] Fangzhe Wan, Xueyan Zhao, Feiqi Deng, Peilin Yu, and Xiongding Liu. Stochastic stabilization of neutral stochastic delay systems based on discrete observations. SIAM Journal on Control and Optimization, 61(6):3259–3279, November 2023. CODEN SJ-CODC. ISSN 1095-7138.

Wang:2019:DSF

[WZHZ19] Yinghui Wang, Wenxiao Zhao, Yiguang Hong, and Mohsen Zamani. Distributed subgradient-free stochastic optimization algorithm for nonsmooth convex functions over time-varying networks. SIAM Journal on Control and Optimization, 57(4):2821–2842, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2006:TTS

[WZY06] J. W. Wang, Q. Zhang, and G. Yin. Two-time-scale hybrid filters: Near optimality. SIAM Journal on Control and Optimization, 45(1):298–319, January 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Wang:2009:SCC

[WZY09] Xiaoqing Wang, Shuzhong Zhang, and David D. Yao. Separated continuous conic programming: Strong duality and an approximation algorithm. SIAM Journal on Control and Optimization, 48(4): 2118–2138, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Xiang:1993:PRT

[XA93] X. Xiang and N. U. Ahmed. Properties of relaxed trajectories of evolution equations and optimal control. SIAM Journal on Control and Optimization, 31(5):1135– 1142, September 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Xiao:1999:FDC

[XB99] Mingqing Xiao and Tamer Basar. Finite-dimensional compensators for the H^{∞}- Optimal control of infinitedimensional systems via a Galerkin-type approximation. SIAM Journal on Control and Optimization, 37(5): 1614–1647, September 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33350.

Xu:2003:MRS

[XC03] Shengyuan Xu and Tongwen Chen. H_{∞} model reduction in the stochastic framework. SIAM Journal on Control and Optimization, 42(4): 1293–1309, January 2003. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Xia:2018:DCS

[XC18] Weiguo Xia and Ming Cao.

Determination of clock synchronization errors in distributed networks. SIAM Journal on Control and Optimization, 56(2):610–632, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Xiong:2020:SSI

[XCZ20] Dan Xiong, Li Chai, and Jingxin Zhang. Sparse system identification in pairs of pulse and Takenaka– Malmquist bases. SIAM Journal on Control and Optimization, 58(2):965–985, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Xia:1989:NOD

[XG89] Xiao-Hua Xia and Wei-Bin Gao. Nonlinear observer design by observer error linearization. SIAM Journal on Control and Optimization, 27(1):199–216, January 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Xu:2003:RBP

[XG03] Gen-Qi Xu and Bao-Zhu Guo. Riesz basis property of evolution equations in Hilbert spaces and application to a coupled string equation. SIAM Journal on Control and Optimization, 42(3):966–984, May 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40008.

Xie:2018:ANL

[XG18] Siyu Xie and Lei Guo. Analysis of normalized least mean squares-based consensus adaptive filters under a general information condition. SIAM Journal on Control and Optimization, 56(5):3404–3431, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Xia:2011:RAP

[Xia11] Jianming Xia. Risk aversion and portfolio selection in a continuous-time model. SIAM Journal on Control and Optimization, 49(5):1916– 1937, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.

siam.org/sicon/resource/
1/sjcodc/v49/i5/p1916_s1.

Xiang:2019:NCL

[Xia19] S. Xiang. Null controllability of a linearized Korteweg– de Vries equation by backstepping approach. SIAM Journal on Control and Optimization, 57(2):1493–1515, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Xie:2012:SCB

[Xie12] Li Xie. Stochastic comparison, boundedness, weak convergence, and ergodicity of a random Riccati equation with Markovian binary switching. SIAM Journal on Control and Optimization, 50(1):532–558, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Xing:2024:CGO

[XJ24] Yu Xing and Karl H. Johansson. Concentration in gossip opinion dynamics over random graphs. SIAM Journal on Control and Optimization, 62(3):1521–1545, June 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1545823.

Xu:2008:ASO

[XLL08] Gen Qi Xu, Dong Yi Liu, and Yan Qing Liu. Abstract second order hyperbolic system and applications to controlled network of strings. SIAM Journal on Control and Optimization, 47(4):1762–1784, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Xu:1996:SRB

[XS96] Cheng-Zhong Xu and Gauthier Sallet. On spectrum and Riesz basis assignment of infinite-dimensional linear systems by bounded linear feedbacks. SIAM Journal on Control and Optimization, 34(2):521–541, March 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25071.

Xie:2008:FHR

[XUP08] Li Xie, Valery A. Ugrinovskii, and Ian R. Petersen. Finite horizon robust state estimation for uncertain finitealphabet hidden Markov models with conditional relative entropy constraints. SIAM Journal on Control and Optimization, 47(1):476–508, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Xiao:2012:ARA

[XW12] Feng Xiao and Long Wang. Asynchronous rendezvous analysis via set-valued consensus theory. SIAM Journal on Control and Optimization, 50(1):

196–221, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v50/i1/p196_s1.

Xiang:2023:OOC

[XWPG23] Qiaomin Xiang, Ze-Hao Wu, Ju H. Park, and Bao-Zhu Guo. Observability and observers for a class of twodimensional hyperbolic PDE chaotic systems. SIAM Journal on Control and Optimization, 61(4):2282–2304, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1462829.

Xue:2021:SRD

[XWZ^{+}21] Bai Xue, Qiuye Wang, Naijun Zhan, Shijie Wang, and Zhikun She. Synthesizing robust domains of attraction for state-constrained perturbed polynomial systems. SIAM Journal on Control and Optimization, 59(2):1083–1108, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Xu:2019:SBF

[XX19] Cheng-Zhong Xu and Gen Qi Xu. Saturated boundary feedback stabilization of a linear wave equation. SIAM Journal on Control and Optimization, 57(1):290–309, ???? 2019. CODEN SJCODC.

ISSN 0363-0129 (print), 1095-7138 (electronic).

Xu:2021:ICD

[XYZ21] Fen Xu, Dacheng Yao, and Hanqin Zhang. Impulse control with discontinuous setup costs: Discounted cost criterion. SIAM Journal on Control and Optimization, 59(1): 267–295, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Xiong:2007:MVP

[XZ07] Jie Xiong and Xun Yu Zhou. Mean-variance portfolio selection under partial information. SIAM Journal on Control and Optimization, 46(1):156–175, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Xu:2015:AAF

[XZ15] Yifeng Xu and Jun Zou. Analysis of an adaptive finite element method for recovering the Robin coefficient. SIAM Journal on Control and Optimization, 53(2):622–644, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Xi:2017:FSF

[XZ17] Fubao Xi and Chao Zhu. On Feller and strong Feller properties and exponential ergodicity of regime-switching jump diffusion processes with countable regimes. SIAM Journal on Control and Optimization, 55(3):1789–1818, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Xu:2020:USD

[XZ20] Gen Qi Xu and Li Zhang.

Uniform stabilization of 1-D coupled wave equations with anti-dampings and joint delayed control. SIAM Journal on Control and Optimization, 58(6):3161–3184, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Xu:2023:SDT

[XZGT23] Jiapeng Xu, Huanshui Zhang, Guoxiang Gu, and Yang Tang. Stabilization of discrete-time systems with random delays and packet drops: a predictorlike controller. SIAM Journal on Control and Optimization, 61(3):1737–1759, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1496062.

Xiong:2020:MVP

[XZZ20] Jie Xiong, Yong Zeng, and Shuaiqi Zhang. Mean-variance portfolio selection for partially observed point processes. SIAM Journal on Control and Optimization, 58(6): 3041–3061, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yu:2009:CMP

[YADF09] Changbin Yu, Brian D. O. Anderson, Soura Dasgupta, and Baris Fidan. Control of minimally persistent formations in the plane. SIAM Journal on Control and Optimization, 48 (1):206–233, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yakowitz:1993:GCS

[Yak93] Sid Yakowitz. A globally convergent stochastic approximation. SIAM Journal on Control and Optimization, 31(1): 30–40, January 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yamamoto:1988:PRI

[Yam88] Yutaka Yamamoto. Pseudorational input/output maps and their realizations: a fractional representation approach to infinite-dimensional systems. SIAM Journal on Control and Optimization, 26(6): 1415–1430, November 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yamamoto:1989:RCI

[Yam89] Yutaka Yamamoto. Reachability of a class of infinitedimensional linear systems: An external approach with applications to general neutral systems. SIAM Journal on Control and Optimization, 27(1):217–234, January 1989. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yame:2012:SRC

[Yam12] Joseph J. Yamé. Some remarks on the controllability of sampled-data systems. SIAM Journal on Control and Optimization, 50(4):1775–1803, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yanagi:1993:OMI

[Yan93]  Kenjiro Yanagi.  Opti-

mal mutual information for coders and jammers in mismatched communication channels. SIAM Journal on Control and Optimization, 31(1): 41–51, January 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yang:1995:OSC

[Yan95]  Hong Yang.  H_{∞} opti-

mal sensitivity for a class of infinite-dimensional systems. SIAM Journal on Control and Optimization, 33(2):351–364, March 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/23945.

Yang:1996:SAC

[Yan96] Zaifu Yang. A simplicial algorithm for computing robust stationary points of a continuous function on the unit simplex. SIAM Journal on Control and Optimization, 34(2):491–506, March 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/25850.

Yang:2014:MSS

[Yan14] Yipeng Yang. A multidimensional stochastic singular control problem via Dynkin game and Dirichlet form. SIAM Journal on Control and Optimization, 52(6):3807–3832, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yao:1999:OIE

[Yao99] Peng-Fei Yao. On the observability inequalities for exact controllability of wave equations with variable coefficients. SIAM Journal on Control and Optimization, 37(5): 1568–1599, September 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/33148.

Yao:2000:OIS

[Yao00] Peng-Fei Yao. Observability inequalities for shallow shells. SIAM Journal on Control and Optimization, 38(6): 1729–1756, November 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33869.

Yakoubi:2006:LSS

[YC06] Karim Yakoubi and Yacine Chitour. Linear systems subject to input saturation and time delay: Finite-gain L^{p}- stabilization. SIAM Journal on Control and Optimization, 45(3):1084–1115, January 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yu:2024:RAS

[YC24] Hao Yu and Tongwen Chen.

RBFNN adaptive sampleddata control for nonlinear plants: a validity analysis. SIAM Journal on Control and Optimization, 62 (3):1908–1932, June 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1595035.

Yu:2013:SPC

[YCLK13] Wenwu Yu, Guanrong Chen, Jinhu Lü, and Jürgen Kurths. Synchronization via pinning control on general complex networks. SIAM Journal on Control and Optimization, 51 (2):1395–1416, ???? 2013.

CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yin:2021:IRS

[YCRZ21] Wensheng Yin, Jinde Cao, Yong Ren, and Guoqiang Zheng. Improved results on stabilization of G-SDEs by feedback control based on discrete-time observations. SIAM Journal on Control and Optimization, 59(3):1927– 1950, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yang:2013:SCR

[YCY13] Xinsong Yang, Jinde Cao, and Zhichun Yang. Synchronization of coupled reactiondiffusion neural networks with time-varying delays via pinning-impulsive controller. SIAM Journal on Control and Optimization, 51(5):3486– 3510, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yin:2003:WCH

[YD03] G. Yin and S. Dey. Weak convergence of hybrid filtering problems involving nearly completely decomposable hidden Markov chains. SIAM Journal on Control and Optimization, 41(6): 1820–1842, November 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38846.

Yang:2008:RSP

[YD08] Guang-Hong Yang and Jiuxiang Dong. Robust stability of polytopic systems via affine parameter-dependent Lyapunov functions. SIAM Journal on Control and Optimization, 47(5):2642–2662, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

You:1998:OBC

[YDZ98] Puhong You, Zhonghai Ding, and Jianxin Zhou. Optimal boundary control of the Stokes fluids with point velocity observations. SIAM Journal on Control and Optimization, 36(3):981–1004, May 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/30027.

Ye:1995:NCB

[Ye95] Jane J. Ye. Necessary conditions for bilevel dynamic optimization problems. SIAM Journal on Control and Optimization, 33(4):1208–1223, July 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/24971.

Ye:1997:OSB

[Ye97] Jane J. Ye. Optimal strategies for bilevel dynamic problems. SIAM Journal on Control and Optimization, 35(2):512–531, March 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/25615.

Ye:2000:DSH

[Ye00a]  J. J. Ye.  Discontinuous

solutions of the Hamilton– Jacobi equation for exit time problems. SIAM Journal on Control and Optimization, 38(4):1067–1085, July 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/32623.

Ye:2000:MRU

[Ye00b] Jane J. Ye. Multiplier rules under mixed assumptions of differentiability and Lipschitz continuity. SIAM Journal on Control and Optimization, 39(5):1441–1460, September 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/35847.

Yamashita:1997:EUM

[YF97] Nobuo Yamashita and Masao Fukushima. Equivalent unconstrained minimization and global error bounds for variational inequality problems. SIAM Journal on Control and Optimization, 35(1):273– 284, January 1997. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27764.

Young:1995:MUD

[YG95] Stanley Young and Vijay K.

Garg. Model uncertainty in discrete event systems. SIAM Journal on Control and Optimization, 33(1):208–226, January 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/21976.

Yan:2018:NAC

[YGC18] Fei Yan, Guoxiang Gu, and Xiang Chen. A new approach to cooperative output regulation for heterogeneous multiagent systems. SIAM Journal on Control and Optimization, 56(3):2074–2094, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yu:2021:SWE

[YH21] Kai Yu and Zhong-Jie Han.

Stabilization of wave equation on cuboidal domain via Kelvin–Voigt damping: a case without geometric control condition. SIAM Journal on Control and Optimization, 59(3):

1973–1988, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yuan:2016:CRD

[YHH16] Deming Yuan, Daniel W. C. Ho, and Yiguang Hong. On convergence rate of distributed stochastic gradient algorithm for convex optimization with inequality constraints. SIAM Journal on Control and Optimization, 54(5):2872–2892, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yu:2018:DCM

[YHJ^{+}18] Zhiyong Yu, Da Huang, Haijun Jiang, Cheng Hu, and Wenwu Yu. Distributed consensus for multiagent systems via directed spanning tree based adaptive control. SIAM Journal on Control and Optimization, 56(3):2189–2217, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Ying:1999:SSM

[Yin99]  Jiang Qian Ying.  On

the strong stabilizability of MIMO n-dimensional linear systems. SIAM Journal on Control and Optimization, 38 (1):313–335, January 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/33581.

Yin:2022:CYS

[Yin22] George Yin. Celebrating 60 years of SICON: 1962 to 2022. SIAM Journal on Control and Optimization, 60(6):i–ii, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22N97560X.

Yang:1991:MCM

[YK91] Jichuan Yang and Harold J.

Kushner. A Monte Carlo method for sensitivity analysis and parametric optimization of nonlinear stochastic systems. SIAM Journal on Control and Optimization, 29(5): 1216–1249, September 1991. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yan:2010:GPT

[YL10] XueHua Yan and YunGang Liu. Global practical tracking for high-order uncertain nonlinear systems with unknown control directions. SIAM Journal on Control and Optimization, 48(7):4453–4473, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yuksel:2012:OCO

[YL12] Serdar Yüksel and Tamás Linder. Optimization and convergence of observation channels in stochastic control. SIAM Journal on Control and Optimization, 50(2):864–887, ????

2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

You:2015:SHS

[YLL^{+}15] Surong You, Wei Liu, Jianqiu Lu, Xuerong Mao, and Qinwei Qiu. Stabilization of hybrid systems by feedback control based on discrete-time state observations. SIAM Journal on Control and Optimization, 53(2):905–925, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yu:2015:DAC

[YLYC15] Wenwu Yu, Jinhu Lü, Xinghuo Yu, and Guanrong Chen. Distributed adaptive control for synchronization in directed complex networks. SIAM Journal on Control and Optimization, 53(5):2980–3005, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yan:1992:SDO

[YM92] Di Yan and H. Mukai. Stochastic discrete optimization. SIAM Journal on Control and Optimization, 30(3): 594–612, May 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yan:1994:RAS

[YMH94] Wei Yong Yan, John B. Moore, and Uwe Helmke. Recursive algorithms for solving a class of nonlinear matrix equations with applications to certain sensitivity optimization problems. SIAM Journal on Control and Optimization, 32(6):1559–1576, November 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/22685.

Yan:2008:SAS

[YÖ08] Peng Yan and Hitay Özbay.

Stability analysis of switched time delay systems. SIAM Journal on Control and Optimization, 47(2):936–949, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yong:1988:DEG

[Yon88a] Jiong Min Yong. On differential evasion games. SIAM Journal on Control and Optimization, 26(1):1–22, January 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yong:1988:DPG

[Yon88b]  Jiong Min Yong.  On

differential pursuit games. SIAM Journal on Control and Optimization, 26(2):478– 495, March 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yong:1990:ZSD

[Yon90] Jiong Min Yong. A zero-sum differential game in a finite duration with switching strategies. SIAM Journal on Control and Optimization, 28(5):1234– 1250, September 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yong:1993:IDV

[Yon93] Jiong Min Yong. Infinitedimensional Volterra–Stieltjes evolution equations and related optimal control problems. SIAM Journal on Control and Optimization, 31(3): 539–568, May 1993. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yong:2002:LFS

[Yon02] Jiongmin Yong. A leaderfollower stochastic linear quadratic differential game. SIAM Journal on Control and Optimization, 41(4):1015– 1041, July 2002. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/39192.

Yong:2010:OVP

[Yon10] Jiongmin Yong. Optimality variational principle for controlled forward-backward stochastic differential equations with mixed initialterminal conditions. SIAM Journal on Control and Optimization, 48(6):4119–4156, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yong:2013:LQO

[Yon13]  Jiongmin Yong.  Linear–

Quadratic optimal control problems for mean-field stochastic differential equations. SIAM Journal on Control and Optimization, 51(4):2809– 2838, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

You:1987:NBP

[You87] Yun Cheng You. A nonquadratic Bolza problem and a quasi-Riccati equation for distributed parameter systems. SIAM Journal on Control and Optimization, 25(4):905–920, July 1987. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yousept:2013:OCQ

[You13] Irwin Yousept. Optimal control of quasilinear H(curl)- Elliptic partial differential equations in magnetostatic field problems. SIAM Journal on Control and Optimization, 51(5):3624–3651, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yousept:2017:OCN

[You17] Irwin Yousept. Optimal control of non-smooth hyperbolic evolution Maxwell equations in Type-II superconductivity. SIAM Journal on Control and Optimization, 55(4): 2305–2332, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yuksel:2017:CAD

[YS17] Serdar Yüksel and Naci Saldi. Convex analysis in decentralized stochastic control, strategic measures, and optimal solutions. SIAM Journal on Control and Optimization, 55 (1):1–28, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yang:2013:DGS

[YT13] Zhou Yang and Shanjian Tang. Dynkin game of stochastic differential equations with random coefficients and associated backward stochastic partial differential variational inequality. SIAM Journal on Control and Optimization, 51 (1):64–95, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yu:1995:IPD

[Yu95] Wen Huan Yu. Identification for parabolic distributed parameter systems with constraints on the parameters and the state. SIAM Journal on Control and Optimization, 33(6):1801–1815, November 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25664.

Yu:2013:LST

[Yu13] Huizhen Yu. Least squares temporal difference methods: an analysis under general conditions. SIAM Journal on Control and Optimization, 50 (6):3310–3343, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yu:2014:ATO

[Yu14] Huaiqiang Yu. Approximation of time optimal controls for heat equations with perturbations in the system potential. SIAM Journal on Control and Optimization, 52(3): 1663–1692, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yu:2015:CVI

[Yu15a] Huizhen Yu. On convergence of value iteration for a class of total cost Markov decision processes. SIAM Journal on Control and Optimization, 53 (4):1982–2016, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yu:2015:OFC

[Yu15b] Zhiyong Yu. An optimal feedback control-strategy pair for zero-sum linear-quadratic stochastic differential game: the Riccati equation approach. SIAM Journal on Control and Optimization, 53(4):2141– 2167, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yu:2020:ACO

[Yu20a] Huizhen Yu. Average cost optimality inequality for Markov decision processes with Borel spaces and universally measurable policies. SIAM Journal on Control and Optimization, 58(4):2469–2502, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yu:2020:MPA

[Yu20b] Huizhen Yu. On the minimum pair approach for average cost Markov decision processes with countable discrete action spaces and strictly unbounded costs. SIAM Journal on Control and Optimization, 58(2):660–685, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yuksel:2016:SEP

[Yük16] Serdar Yüksel. Stationary and ergodic properties of stochastic NonLinear systems controlled over communication channels. SIAM Journal on Control and Optimization, 54 (5):2844–2871, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yuksel:2017:SSC

[Yük17] Serdar Yüksel. On stochastic stability of a class of non-Markovian processes and applications in quantization. SIAM Journal on Control and Optimization, 55(2):1241– 1260, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yuksel:2019:NSO

[Yük19] Serdar Yüksel. A note on the separation of optimal quantization and control policies in networked control. SIAM Journal on Control and Optimization, 57(1):773–782, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yuksel:2020:UDP

[Yük20] Serdar Yüksel. A universal dynamic program and refined existence results for decentralized stochastic control. SIAM Journal on Control and Optimization, 58(5):2711–2739, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yuksel:2024:BYR

[Yük24] Serdar Yüksel. On Borkar and Young relaxed control topologies and continuous dependence of invariant measures on control policy. SIAM Journal on Control and Optimization, 62(4):2367–2386, August 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1571940.

Yushkevich:1997:BOB

[Yus97] Alexander A. Yushkevich. Blackwell optimality in Borelian continuous-in-action Markov decision processes. SIAM Journal on Control and Optimization, 35(6):2157–2182, November 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http://

epubs.siam.org/sam-bin/
dbq/article/29246.

Yao:2007:GSS

[YW07] Peng-Fei Yao and George Weiss. Global smooth solutions and exponential stability for a nonlinear beam. SIAM Journal on Control and Optimization, 45(6):1931–1964, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yang:2012:FTS

[YW12] Renming Yang and Yuzhen Wang. Finite-time stability and stabilization of a class of nonlinear time-delay systems. SIAM Journal on Control and Optimization, 50(5): 3113–3131, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yin:2010:SRS

[YX10] George Yin and Fubao Xi.

Stability of regime-switching jump diffusions. SIAM Journal on Control and Optimization, 48(7):4525–4549, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yang:2021:FTN

[YXC21] Junyi Yang, Feng Xiao, and Tongwen Chen. Formation tracking of nonholonomic systems on the special Euclidean group under fixed and switching topologies: an affine formation strategy. SIAM Journal on Control and Optimization, 59(4):2850–2874, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yao:2011:TSE

[YXJ11] Tao Yao, Susan H. Xu, and Bin Jiang. Trigger strategy equilibriums in stochastic differential games with information time lags: An analysis of cooperative production strategies. SIAM Journal on Control and Optimization, 49(3): 1197–1220, ???? 2011. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

                                        tronic). URL http://epubs.
                                        siam.org/sicon/resource/
                                        1/sjcodc/v49/i3/p1197_s1.

Yau:2005:SFP

[YY05] Stephen S. T. Yau and Shing-Tung Yau. Solution of filtering problem with nonlinear observations. SIAM Journal on Control and Optimization, 44(3):1019–1039, May 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/41197.

Yau:2008:RTS

[YY08] Shing-Tung Yau and Stephen S.-T. Yau. Real time solution of the nonlinear filtering problem without memory II. SIAM Journal on Control and Optimization, 47(1): 163–195, ???? 2008. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yin:2016:MBE

[YYH^{+}16] Xiuxia Yin, Dong Yue, Songlin Hu, Chen Peng, and Yusheng Xue. Model-based eventtriggered predictive control for networked systems with data dropout. SIAM Journal on Control and Optimization, 54 (2):567–586, ???? 2016. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

You:1997:CPE

[YZ97] Puhong You and Jianxin Zhou. Constrained LQR problems in elliptic distributed control systems with point observations — on convergence rates. SIAM Journal on Control and Optimization, 35(5): 1739–1754, September 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27818.

Yang:2016:OIB

[YZ16] Donghui Yang and Jie Zhong.

Observability inequality of backward stochastic heat equations for measurable sets and its applications. SIAM Journal on Control and Optimization, 54(3):1157–1175, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Yu:2022:CER

[YZ22] Yongyi Yu and Ji-Feng Zhang.

Carleman estimates of refined stochastic beam equations and applications. SIAM Journal on Control and Optimization, 60(5):2947–2970, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1463513.

Yu:2023:TMP

[YZ23] Yongyi Yu and Ji-Feng Zhang.

Two multiobjective problems for stochastic degenerate parabolic equations. SIAM Journal on Control and Optimization, 61(4):2708–2735, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1510194.

Yang:2023:ATC

[YZSS23] Dong Yang, Guangdeng Zong, Yang Shi, and Peng Shi. Adaptive tracking control of hybrid switching Markovian systems with its applications. SIAM Journal on Control and Optimization, 61(2):434–457, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1442036.

Yao:2001:SLQ

[YZZ01] David D. Yao, Shuzhong Zhang, and Xun Yu Zhou. Stochastic linear-quadratic control via semidefinite programming. SIAM Journal on Control and Optimization, 40(3):801–823, May 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35548.

Zavala:2010:RTN

[ZA10] Victor M. Zavala and Mihai Anitescu. Real-time nonlinear optimization as a generalized equation. SIAM Journal on Control and Optimization, 48(8):5444–5467, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zabczyk:1978:DG

[Zab78] Jerzy Zabczyk. On decomposition of generators. SIAM Journal on Control and Optimization, 16(4):523–534, July 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See erratum [Zab80].

Zabczyk:1980:EDG

[Zab80] Jerzy Zabczyk. Erratum: On decomposition of generators. SIAM Journal on Control and Optimization, 18(3):325, May 1980. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [Zab78].

Zaffaroni:2003:DED

[Zaf03] Alberto Zaffaroni. Degrees of efficiency and degrees of minimality. SIAM Journal on Control and Optimization, 42(3):1071–1086, May 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41153.

Zagatti:2000:MFG

[Zag00] Sandro Zagatti. Minimization of functionals of the gradient by Baire’s theorem. SIAM Journal on Control and Optimization, 38(2):384–399, March 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/33520.

Zampieri:1998:CIO

[Zam98] Sandro Zampieri. Causal input/output representation of 2D systems in the behavioral approach. SIAM Journal on Control and Optimization, 36(4):1133–1146, July 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/29257.

Zaremba:1984:EVG

[Zar84] L. S. Zaremba. Existence of value in generalized pursuitevasion games. SIAM Journal on Control and Optimization, 22(6):894–901, November 1984. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zariphopoulou:1992:ICM

[Zar92] Thaleia Zariphopoulou. Investmentconsumption models with transaction fees and Markovchain parameters. SIAM Journal on Control and Optimization, 30(3):613–636, May 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zariphopoulou:1994:CIM

[Zar94] Thaleia Zariphopoulou. Consumptioninvestment models with constraints. SIAM Journal on Control and Optimization, 32(1):59–85, January 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/21882.

Zaslavski:1995:OPIa

[Zas95a] A. J. Zaslavski. Optimal programs on infinite horizon. Part 1. SIAM Journal on Control and Optimization, 33(6): 1643–1660, November 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25726.

Zaslavski:1995:OPIb

[Zas95b] A. J. Zaslavski. Optimal programs on infinite horizon. Part 2. SIAM Journal on Control and Optimization, 33(6): 1661–1686, November 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/25727.

Zaslavski:1997:TPO

[Zas97] A. J. Zaslavski. Turnpike property of optimal solutions of infinite-horizon variational problems. SIAM Journal on Control and Optimization, 35(4):1169–1203, July 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/28935.

Zaslavski:2000:GWP

[Zas00] Alexander J. Zaslavski. Generic well-posedness of optimal control problems without convexity assumptions. SIAM Journal on Control and Optimization, 39(1):250–280, January 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/34539.

Zaslavski:2006:NLP

[Zas06] Alexander J. Zaslavski. Nonoccurrence of the Lavrentiev phenomenon for many optimal control problems. SIAM Journal on Control and Optimization, 45(3):1116–1146, January 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zelazo:2013:FTD

[ZBA13] Daniel Zelazo, Mathias Bürger, and Frank Allgöwer. A finitetime dual method for negotiation between dynamical systems. SIAM Journal on Control and Optimization, 51(1): 172–194, ???? 2013. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zheng:2007:SOD

[ZBB07] Gang Zheng, Driss Boutat, and Jean-Pierre Barbot. Single output-dependent observability normal form. SIAM Journal on Control and Optimization, 46(6):2242–2255, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhou:1991:WMD

[ZC91] Jian-Xin Zhou and Goong Chen. The wave method for determining the asymptotic damping rates of eigenmodes I: The wave equation on a rectangular or circular domain. SIAM Journal on Control and Optimization, 29 (3):656–677, May 1991. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhang:1996:ACS

[ZC96] Ji Feng Zhang and Peter E.

Caines. Adaptive control via a simple switching algorithm. SIAM Journal on Control and Optimization, 34 (1):365–388, January 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/24075.

Zhang:2006:SFC

[ZC06] Weihai Zhang and Bor-Sen Chen. State feedback H_{∞} control for a class of nonlinear stochastic systems.

SIAM Journal on Control and Optimization, 44(6):1973– 1991, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhang:2024:SMP

[ZC24] Shuaiqi Zhang and Zhen-Qing Chen. Stochastic maximum principle for subdiffusions and its applications. SIAM Journal on Control and Optimization, 62(2):953–981, March 2024. CODEN SJCODC. ISSN 1095-7138.

Zhang:2007:PRD

[ZCC07] Guofeng Zhang, Tongwen Chen, and Xiang Chen. Performance recovery in digital implementation of analogue systems. SIAM Journal on Control and Optimization, 45 (6):2207–2223, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zeiler:2010:KOO

[ZCGT10] Irmgard Zeiler, Jonathan P. Caulkins, Dieter Grass, and Gernot Tragler. Keeping options open: an optimal control model with trajectories that reach a DNSS point in positive time. SIAM Journal on Control and Optimization, 48(6): 3698–3707, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zyla:1983:NSI

[Zd83] L. V. Zyla and Rui J. P. de Figueiredo. Nonlinear system identification based on a Fock space framework. SIAM Journal on Control and Optimization, 21(6):931–939, November 1983. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhao:2014:NTS

[ZD14] Xueyan Zhao and Feiqi Deng.

New type of stability criteria for stochastic functional differential equations via Lyapunov functions. SIAM Journal on Control and Optimization, 52 (4):2319–2347, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhu:2024:GUF

[ZD24] Wenwu Zhu and Haibo Du.

Global uniform finite-time output feedback stabilization for disturbed nonlinear uncertain systems: Backsteppinglike observer and nonseparation principle design. SIAM Journal on Control and Optimization, 62(3):1717–1736, June 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1519365.

Zhang:2020:BSR

[ZDM20] Fu Zhang, Yuchao Dong, and Qingxin Meng. Backward stochastic Riccati equation with jumps associated with stochastic linear quadratic optimal control with jumps and random coefficients. SIAM Journal on Control and Optimization, 58(1):393–424, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zeidan:1984:EJS

[Zei84] Vera Zeidan. Extended Jacobi sufficiency criterion for optimal control. SIAM Journal on Control and Optimization, 22 (2):294–301, March 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zeidan:1992:SCF

[Zei92] Vera Zeidan. Sufficiency criteria via focal points and via coupled points. SIAM Journal on Control and Optimization, 30(1):82–98, January 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zeidan:1994:REO

[Zei94] Vera Zeidan. The Riccati equation for optimal control problems with mixed statecontrol constraints: Necessity and sufficiency. SIAM Journal on Control and Optimization, 32(5):1297–1321, September 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/23364.

Zeidan:2001:NSO

[Zei01] Vera Zeidan. New secondorder optimality conditions for variational problems with C^{2}- Hamiltonians. SIAM Journal on Control and Optimization, 40(2):577–609, March 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35872.

Zervos:2003:PSE

[Zer03] Mihail Zervos. A problem of sequential entry and exit decisions combined with discretionary stopping. SIAM Journal on Control and Optimization, 42(2):397–421, March 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38111.

Zerz:2005:CFP

[Zer05]  Eva Zerz.  Characteris-

tic frequencies, polynomialexponential trajectories, and linear exact modeling with multidimensional behaviors. SIAM Journal on Control and Optimization, 44(3):1148– 1163, May 2005. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/44173.

Zhang:2012:SMP

[ZES12] Xin Zhang, Robert J. Elliott, and Tak Kuen Siu. A stochastic maximum principle for a Markov regime-switching jump-diffusion model and its application to finance. SIAM Journal on Control and Optimization, 50(2):964–990, ???? 2012. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhu:2021:SEF

[ZFKZ21] Bin Zhu, Augusto Ferrante, Johan Karlsson, and Mattia Zorzi. M^{2} spectral estimation: a flexible approach ensuring rational solutions. SIAM Journal on Control and Optimization, 59(4): 2977–2996, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zabarankin:2017:FIP

[ZG17] Michael Zabarankin and Bogdan Grechuk. Forward and inverse problems in two-phase fluid dynamics. SIAM Journal on Control and Optimization, 55(6):3969–3989, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhou:2018:BFS

[ZG18] Hua-Cheng Zhou and Bao-Zhu Guo. Boundary feedback stabilization for an unstable time fractional reaction diffusion equation. SIAM Journal on Control and Optimization, 56(1):75–101, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhang:2019:CSG

[ZG19] Renren Zhang and Lei Guo.

Controllability of stochastic game-based control systems. SIAM Journal on Control and Optimization, 57(6):3799– 3826, ???? 2019. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhang:2021:SGB

[ZG21] Renren Zhang and Lei Guo.

Stabilizability of game-based control systems. SIAM Journal on Control and Optimization, 59(5):3999–4023, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhao:1999:ABP

[ZGS99] Ke-You Zhao, Michael J.

Grimble, and Jakob Stoustrup. Average H_{∈} performance and maximal parameter perturbation radius for uncertain systems. SIAM Journal on Control and Optimization, 37(6):1742–1750, November 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30172.

Zhou:2002:ECP

[ZH02] Jun Zhou and Tomomichi Hagiwara. Existence conditions and properties of the frequency response operators of continuous-time periodic systems. SIAM Journal on Control and Optimization, 40(6): 1867–1887, November 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/38235.

Zhang:2003:BSH

[ZH03] Qiong Zhang and Falun Huang. Boundary stabilization for a hybrid system of viscoelasticity. SIAM Journal on Control and Optimization, 42(5):1643–1660, September 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38991.

Zhou:2005:RNC

[ZH05] Jun Zhou and Tomomichi Hagiwara. 2-regularized Nyquist criterion in linear continuous-time periodic systems and its implementation. SIAM Journal on Control and Optimization, 44(2):618–645, March 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/41690.

Zhu:2018:SAP

[ZH18] Lijun Zhu and David J. Hill.

Stability analysis of power systems: a network synchronization perspective. SIAM Journal on Control and Optimization, 56(3):1640–1664, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zheng:2024:GMP

[ZH24] Yueyang Zheng and Yaozhong Hu. The global maximum principle for optimal control of partially observed stochastic systems driven by fractional Brownian motion. SIAM Journal on Control and Optimization, 62(1):509–538, February 2024. CODEN SJCODC. ISSN 1095-7138.

Zhang:1995:RSP

[Zha95] Qing Zhang. Risk-sensitive production planning of stochastic manufacturing systems: a singular perturbation approach. SIAM Journal on Control and Optimization, 33(2):498–527, March 1995. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/24362.

Zhang:1998:NFC

[Zha98] Q. Zhang. Nonlinear filtering and control of a switching diffusion with small observation noise. SIAM Journal on Control and Optimization, 36(5):1638–1668, September 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31544.

Zhang:1999:EBC

[Zha99]  Bing-Yu Zhang.  Exact

boundary controllability of the Korteweg–de Vries equation. SIAM Journal on Control and Optimization, 37(2):543–565, March 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/32750.

Zhang:2000:EOI

[Zha00] Xu Zhang. Explicit observability inequalities for the wave equation with lower order terms by means of Carleman inequalities. SIAM Journal on Control and Optimization, 39(3):812–834, May 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35029.

Zhang:2001:STO

[Zha01a] Q. Zhang. Stock trading: An optimal selling rule. SIAM Journal on Control and Optimization, 40(1):64–87, January 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/35632.

Zhang:2001:RNE

[Zha01b] Xu Zhang. A remark on null exact controllability of the heat equation. SIAM Journal on Control and Optimization, 40(1):39–53, January 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37169.

Zhang:2005:SRT

[Zha05] Pingjian Zhang. Some results on two-person zero-sum linear quadratic differential games. SIAM Journal on Control and Optimization, 43(6): 2157–2165, November 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/42560.

Zhang:2008:EIN

[Zha08] Zhigang Zhang. Effect of input noise on a magnetometer with quantum feedback. SIAM Journal on Control and Optimization, 47(2):639–660, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhang:2009:VIO

[Zha09] Tusheng Zhang. Variational inequalities and optimization for Markov processes associated with semi-Dirichlet forms. SIAM Journal on Control and Optimization, 48(3): 1743–1755, ???? 2009. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhang:2013:TOC

[Zha13] Can Zhang. The time optimal control with constraints of the rectangular type for linear time-varying ODEs. SIAM Journal on Control and Optimization, 51(2):1528–1542, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhang:2015:TET

[Zha15] Yubiao Zhang. Two equivalence theorems of different kinds of optimal control problems for Schrödinger equations. SIAM Journal on Control and Optimization, 53(2): 926–947, ???? 2015. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhang:2017:EGV

[Zha17a] Fu Zhang. The existence of game value for pathdependent stochastic differential game. SIAM Journal on Control and Optimization, 55(4):2519–2542, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhang:2017:CTM

[Zha17b] Yi Zhang. Continuous-time Markov decision processes with exponential utility. SIAM Journal on Control and Optimization, 55(4):2636–2660, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhang:2018:ICS

[Zha18]  Christophe Zhang.  Inter-

nal controllability of systems of semilinear coupled onedimensional wave equations with one control. SIAM Journal on Control and Optimization, 56(4):3092–3127, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhang:2021:POL

[ZHB21] Kaiqing Zhang, Bin Hu, and Tamer Basar. Policy optimization for H_{∈} linear control with H_{∞} robustness guarantee: Implicit regularization and global convergence. SIAM Journal on Control and Optimization, 59(6):4081–4109, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhang:2006:MRU

[ZHC06] Liqian Zhang, Biao Huang, and Tongwen Chen. Model reduction of uncertain systems with multiplicative noise based on balancing. SIAM Journal on Control and Optimization, 45(5):1541–1560, ???? 2006. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhang:2017:MOD

[ZHH17] Jiangbo Zhang, Yiguang Hong, and Xiaoming Hu. Multiagent opinion dynamics of bounded confidence with nonlocal aggregative interaction. SIAM Journal on Control and Optimization, 55(4): 2543–2573, ???? 2017. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhou:1983:ACC

[Zho83] Hong Xing Zhou. Approximate controllability for a class of semilinear abstract equations. SIAM Journal on Control and Optimization, 21(4): 551–565, July 1983. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhou:1984:CPL

[Zho84] Hong Xing Zhou. Controllability properties of linear and semilinear abstract control systems. SIAM Journal on Control and Optimization, 22 (3):405–422, May 1984. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhou:1992:EOR

[Zho92] Xun Yu Zhou. On the existence of optimal relaxed controls of stochastic partial differential equations. SIAM Journal on Control and Optimization, 30(2):247– 261, March 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhou:1993:NCO

[Zho93]  Xun Yu Zhou.  On the

necessary conditions of optimal controls for stochastic partial differential equations. SIAM Journal on Control and Optimization, 31(6): 1462–1478, November 1993. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhou:1998:SNO

[Zho98] Xun Yu Zhou. Stochastic near-optimal controls: Necessary and sufficient conditions for near-optimality. SIAM Journal on Control and Optimization, 36(3):929–947, May 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/30266.

Zhou:2007:HFC

[Zho07] Jun Zhou. A harmonic framework for controllability in linear continuous-time periodic systems. SIAM Journal on Control and Optimization, 46 (2):630–654, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhou:2023:PTO

[Zho23a] Xiuxiang Zhou. Perturbations of time optimal control problems for parabolic integrodifferential equations. SIAM Journal on Control and Optimization, 61(4):2069–2087, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1527064.

Zhou:2023:TIP

[Zho23b] Zhou Zhou. Time inconsistency, precommitment, and equilibrium strategies for a Stackelberg game. SIAM Journal on Control and Optimization, 61(2):361–397, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1477659.

Zhang:2014:QDC

[ZHSP14] Hongzhong Zhang, Olympia Hadjiliadis, Tobias Schäfer, and H. Vincent Poor. Quickest detection in coupled systems. SIAM Journal on Control and Optimization, 52(3): 1567–1596, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhu:1992:CDG

[Zhu92] Hang Zhu. Convex duality and generalized solutions in optimal control problem for stopped processes: The deterministic model. SIAM Journal on Control and Optimization, 30(2):465–476, March 1992. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhuang:1994:DRP

[Zhu94]  Deming Zhuang.  Density

results for proper efficiencies. SIAM Journal on Control and Optimization, 32(1):51–58, January 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/17151.

Zhu:1996:CPD

[Zhu96] Ciyou Zhu. On a certain parameter of the discretized extended linear-quadratic problem of optimal control. SIAM Journal on Control and Optimization, 34(1):62–73, January 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/25271.

Zhu:2000:HNC

[Zhu00] Qiji J. Zhu. Hamiltonian necessary conditions for a multiobjective optimal control problem with endpoint constraints. SIAM Journal on Control and Optimization, 39(1):97–112, January 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/35082.

Zhu:2021:OIC

[Zhu21] Jinxia Zhu. Optimal impulse control for growth-restricted linear diffusions with regime switching. SIAM Journal on Control and Optimization, 59 (1):185–222, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhao:2000:PMM

[ZI00] Yun-Bin Zhao and George Isac. Properties of a multivalued mapping associated with some nonmonotone complementarity problems. SIAM Journal on Control and Optimization, 39(2):571–593, March 2000. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/34519.

Zitkovic:2014:DPC

[Zit14] Gordan Zitković. Dynamic programming for controlled Markov families: Abstractly and over martingale measures. SIAM Journal on Control and Optimization, 52(3): 1597–1621, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhu:1989:NRR

[ZJ89] J. Zhu and C. D. Johnson.

New results in the reduction of linear time-varying dynamical systems. SIAM Journal on Control and Optimization, 27 (3):476–494, May 1989. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhang:2006:OCH

[ZJ06] Huan Zhang and Matthew R.

James. Optimal control of hybrid systems and a system of quasi-variational inequalities. SIAM Journal on Control and Optimization, 45(2): 722–761, January 2006. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhou:1988:SUL

[ZK88] Kemin Zhou and Pramod P.

Khargonekar. On the stabilization of uncertain linear systems via bound invariant Lyapunov functions. SIAM Journal on Control and Optimization, 26(6):1265–1273, November 1988. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhu:2020:ARP

[ZK20] Yang Zhu and Miroslav Krstic.

Adaptive and robust predictors for multi-input linear systems with distributed delays. SIAM Journal on Control and Optimization, 58(6):

3457–3485, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zad:1999:SOC

[ZKW99] S. Hashtrudi Zad, R. H. Kwong, and W. M. Wonham. Supremum operators and computation of supremal elements in system theory. SIAM Journal on Control and Optimization, 37(3):695–709, May 1999. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/31718.

Zhang:2020:GCP

[ZKZB20] Kaiqing Zhang, Alec Koppel, Hao Zhu, and Tamer Basar. Global convergence of policy gradient methods to (almost) locally optimal policies. SIAM Journal on Control and Optimization, 58(6): 3586–3612, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhao:2001:ELB

[ZL01] Yun-Bin Zhao and Duan Li. Existence and Limiting behavior of a Non–Interior-Point trajectory for nonlinear complementarity problems without strict feasibility condition. SIAM Journal on Control and Optimization, 40(3):898–924, May 2001. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37247.

Zerz:2002:CSI

[ZL02] E. Zerz and V. Lomadze. A constructive solution to interconnection and decomposition problems with multidimensional behaviors. SIAM Journal on Control and Optimization, 40(4):1072–1086, July 2002. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/37474.

Zhou:2007:LEB

[ZL07] Yi Zhou and Zhen Lei. Local exact boundary controllability for nonlinear wave equations. SIAM Journal on Control and Optimization, 46(3): 1022–1051, ???? 2007. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhang:2021:ECL

[ZL21] Wei Zhang and Jr-Shin Li. Ensemble control on Lie groups. SIAM Journal on Control and Optimization, 59(5):3805– 3827, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhu:2024:FPS

[ZL24] Shuaibing Zhu and Jinhu Lü. Further on pinning synchronization of dynamical networks with coupling delay. SIAM Journal on Control and Optimization, 62 (4):1933–1952, August 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/23M1578085.

Zhou:2010:GSG

[ZLD10] Bin Zhou, Zongli Lin, and Guang-Ren Duan. Global and semi-global stabilization of linear systems with multiple delays and saturations in the input. SIAM Journal on Control and Optimization, 48(8): 5294–5332, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhou:2012:LIC

[ZLD12] Bin Zhou, Zongli Lin, and Guang-Ren Duan. A Lyapunov inequality characterization of and a Riccati inequality approach to L_{∞} and L_{2} low gain feedback. SIAM Journal on Control and Optimization, 50(1):1–22, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (elec-

tronic). URL http://epubs.
siam.org/sicon/resource/
1/sjcodc/v50/i1/p1_s1.

Zhang:2012:CFC

[ZLHZ12] Guofeng Zhang, Heung Wing Joseph Lee, Bo Huang, and Hu Zhang. Coherent feedback control of linear quantum optical systems via squeezing and phase shift. SIAM Journal on Control and Optimization, 50(4): 2130–2150, ???? 2012. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhu:2018:OCB

[ZLLC18] Qunxi Zhu, Yang Liu, Jianquan Lu, and Jinde Cao. On the optimal control of Boolean control networks. SIAM Journal on Control and Optimization, 56(2):1321–1341, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zong:2018:CCC

[ZLZ18] Xiaofeng Zong, Tao Li, and Ji-Feng Zhang. Consensus conditions of continuous-time multiagent systems with additive and multiplicative measurement noises. SIAM Journal on Control and Optimization, 56(1):19–52, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zabarankin:2011:SOV

[ZM11] Michael Zabarankin and Anton Molyboha. 3D shape optimization in viscous incompressible fluid under Oseen approximation. SIAM Journal on Control and Optimization, 49(3):1358–1382, ???? 2011. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sicon/
                                        resource/1/sjcodc/v49/i3/
                                        p1358_s1.

Zhu:2013:DCG

[ZM13] Minghui Zhu and Sonia Martínez. Distributed coverage games for energy-aware mobile sensor networks. SIAM Journal on Control and Optimization, 51(1):1–27, ???? 2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhu:2014:ARD

[ZM14] Minghui Zhu and Sonia Martínez. On attack-resilient distributed formation control in operator-vehicle networks. SIAM Journal on Control and Optimization, 52(5): 3176–3202, ???? 2014. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zeng:2021:NQH

[ZMK21] Shengda Zeng, Stanisław Migórski, and Akhtar A. Khan. Nonlinear quasihemivariational inequalities: Existence and optimal control. SIAM Journal on Control and Optimization, 59(2): 1246–1274, ???? 2021. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zolezzi:1978:CSV

[Zol78] T. Zolezzi. Characterizations of some variational perturbations of the abstract linearquadratic problem. SIAM Journal on Control and Optimization, 16(1):106–121, January 1978. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zolezzi:1981:CWP

[Zol81] T. Zolezzi. A characterization of well-posed optimal control systems. SIAM Journal on Control and Optimization, 19(5):604–616, September 1981. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zolezzi:1992:WLP

[Zol92] Tullio Zolezzi. Wellposedness and the Lavrentiev phenomenon. SIAM Journal on Control and Optimization, 30 (4):787–799, July 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zorzi:2017:CAF

[Zor17] Mattia Zorzi. Convergence analysis of a family of robust Kalman filters based on the contraction principle. SIAM Journal on Control and Optimization, 55(5):3116–3131, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zevin:2003:NAL

[ZP03a] A. A. Zevin and M. A. Pinsky. A new approach to the Lur’e problem in the theory of absolute stability. SIAM Journal on Control and Optimization, 42(5):1895–1904, September 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/40985.

Zevin:2003:SCL

[ZP03b] Alexandr Zevin and Mark Pinsky. Stability criteria for linear systems with partly uncertain periodic coefficients. SIAM Journal on Control and Optimization, 42(4):1185–1197, July 2003. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

//epubs.siam.org/sam-bin/
dbq/article/38552.

Zevin:2005:ASC

[ZP05] A. A. Zevin and M. A. Pinsky. Absolute stability criteria for a generalized Lur’e problem with delay in the feedback. SIAM Journal on Control and Optimization, 43(6): 2000–2008, November 2005. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/43759.

Zhang:2020:LPI

[ZPQ20] Liguo Zhang, Christophe Prieur, and Junfei Qiao. Local proportional-integral boundary feedback stabilization for quasilinear hyperbolic systems of balance laws. SIAM Journal on Control and Optimization, 58(4):2143–2170, ???? 2020. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zheng:2018:EMS

[ZQ18] Jianying Zheng and Li Qiu.

Existence of a mean-square stabilizing solution to a modified algebraic Riccati equation. SIAM Journal on Control and Optimization, 56(1): 367–387, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zazo:2023:BCS

[ZR23] Miriam Steinherr Zazo and Jens D. M. Rademacher. Bifurcation control for a ship maneuvering model with nonsmooth nonlinearities. SIAM Journal on Control and Optimization, 61(1):225–251, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1417259.

Zhu:1997:CAO

[ZRS97] G. Zhu, M. A. Rotea, and R. Skelton. A convergent algorithm for the output covariance constraint control problem. SIAM Journal on Control and Optimization, 35 (1):341–361, January 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

                                        /epubs.siam.org/sam-bin/
                                        dbq/article/26397.

Zavala-Rio:2022:CFT

[ZRSZG22] Arturo Zavala-Río, Tonametl Sanchez, and Griselda I. Zamora-Gómez. On the continuous finite-time stabilization of the double integrator. SIAM Journal on Control and Optimization, 60(2):699–719, ???? 2022. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/20M136459X.

Zhou:2019:FLQ

[ZS19] Bonan Zhou and Jason L. Speyer. Fractional linear quadratic regulators using Wiener–Hopf spectral factorization. SIAM Journal on Control and Optimization, 57 (6):4011–4032, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zheng:2023:GMP

[ZS23] Yueyang Zheng and Jingtao Shi. The global maximum principle for progressive optimal control of partially observed forward-backward stochastic systems with random jumps. SIAM Journal on Control and Optimization, 61(3):1063–1094, ???? 2023.

CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/21M1452585.

Zhang:2010:PSE

[ZSM10] Xin Zhang, Tak Kuen Siu, and Qingbin Meng. Portfolio selection in the enlarged Markovian regime-switching market. SIAM Journal on Control and Optimization, 48(5): 3368–3388, ???? 2010. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhou:2020:SLE

[ZSS^{+}20] Zhengyuan Zhou, Jonathan R.

Shewchuk, Dusan Stipanović, Haomiao Huang, and Claire J. Tomlin. Smarter lions: Efficient cooperative pursuit in general bounded arenas. SIAM Journal on Control and Optimization, 58(2):1229– 1256, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhang:2018:GSM

[ZSX18] Xin Zhang, Zhongyang Sun, and Jie Xiong. A general stochastic maximum principle for a Markov regime switching jump-diffusion model of meanfield type. SIAM Journal on Control and Optimization, 56(4):2563–2592, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhu:2016:MTC

[ZTC16] Jiamin Zhu, Emmanuel Trélat, and Max Cerf. Minimum time control of the rocket attitude reorientation associated with orbit dynamics. SIAM Journal on Control and Optimization, 54(1):391–422, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zuazua:1990:USW

[Zua90] Enrike Zuazua. Uniform stabilization of the wave equation by nonlinear boundary feedback. SIAM Journal on Control and Optimization, 28(2): 466–477, March 1990. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zuazua:1996:URL

[Zua96] Enrique Zuazua. A uniqueness result for the linear system of elasticity and its control theoretical consequences. SIAM Journal on Control and Optimization, 34(5):1473–1495, September 1996. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:

                                        //epubs.siam.org/sam-bin/
                                        dbq/article/26007. See er-

ratum [Zua98].

Zuazua:1998:EUR

[Zua98] Enrique Zuazua. Erratum: a uniqueness result for the linear system of elasticity and its control theoretical consequences. SIAM Journal on Control and Optimization, 37 (1):330–331, January 1998. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/33412.  See

[Zua96].

Zuyev:2016:ESN

[Zuy16] Alexander Zuyev. Exponential stabilization of nonholonomic systems by means of oscillating controls. SIAM Journal on Control and Optimization, 54(3):1678–1696, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zarrouati-Vissiere:2014:ISB

[ZVBR14] Nadège Zarrouati-Vissière, Karine Beauchard, and Pierre Rouchon. Inertial-sensor bias estimation from brightness/depth images and based on SO(3)-invariant integro/ partial differential equations on the unit sphere. SIAM Journal on Control and Optimization, 52(6):3463–3495, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhou:2018:OFE

[ZW18] Hua-Cheng Zhou and George Weiss. Output feedback exponential stabilization for onedimensional unstable wave equations with boundary control matched disturbance. SIAM Journal on Control and Optimization, 56(6):4098– 4129, ???? 2018. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zheng:2021:HMV

[ZW21] Xiangcheng Zheng and Hong Wang. A hidden-memory variable-order time-fractional optimal control model: Analysis and approximation. SIAM Journal on Control and Optimization, 59(3):1851–1880, ???? 2021. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zwart:1988:EBO

[Zwa88a] Hans Zwart. Equivalence between open-loop and closedloop invariance for infinitedimensional systems: a frequency domain approach. SIAM Journal on Control and Optimization, 26(5):1175– 1199, September 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zwart:1988:CAC

[Zwa88b] Hansc Zwart. Characterization of all controlled invariant subspaces for spectral systems. SIAM Journal on Control and Optimization, 26(2): 369–386, March 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zong:2014:AST

[ZWYJ14] Xiaofeng Zong, Fuke Wu, George Yin, and Zhuo Jin. Almost sure and p th-moment stability and stabilization of regime-switching jump diffusion systems. SIAM Journal on Control and Optimization, 52(4):2595–2622, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhou:2003:MMV

[ZY03] Xun Yu Zhou and G. Yin.

Markowitz’s mean-variance portfolio selection with regime switching: a continuous-time model. SIAM Journal on Control and Optimization, 42(4): 1466–1482, January 2003. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhu:2007:APH

[ZY07] C. Zhu and G. Yin. Asymptotic properties of hybrid diffusion systems. SIAM Journal on Control and Optimization, 46(4):1155–1179, ???? 2007. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhang:2008:GSS

[ZY08] Zhi-Fei Zhang and Peng-Fei Yao. Global smooth solutions of the quasi-linear wave equation with internal velocity feedback. SIAM Journal on Control and Optimization, 47(4):2044–2077, ???? 2008. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhu:2009:SFR

[ZY09] C. Zhu and G. Yin. On strong Feller, recurrence, and weak stabilization of regimeswitching diffusions. SIAM Journal on Control and Optimization, 48(3):2003–2031, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhao:2024:TAA

[ZY24] Cheng Zhao and Shuo Yuan. On tracking and antidisturbance ability of PID controllers. SIAM Journal on Control and Optimization, 62 (3):1857–1883, June 2024. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/22M1522498.

Zhao:2023:FTS

[ZYGpJ23] Zhi-Liang Zhao, Ruonan Yuan, Bao-Zhu Guo, and Zhong ping Jiang. Finite-time stabilization of nonlinear uncertain MIMO systems. SIAM Journal on Control and Optimization, 61(4):2305–2333, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

//epubs.siam.org/doi/10.
1137/22M1496086.

Zeng:2018:DCT

[ZYHX18] Xianlin Zeng, Peng Yi, Yiguang Hong, and Lihua Xie. Distributed continuous-time algorithms for nonsmooth extended monotropic optimization problems. SIAM Journal on Control and Optimization, 56(6):3973–3993, ???? 2018. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhou:1997:SVT

[ZYL97] Xun Yu Zhou, Jiongmin Yong, and Xunjing Li. Stochastic verification theorems within the framework of viscosity solutions. SIAM Journal on Control and Optimization, 35 (1):243–253, January 1997. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL http:/

/epubs.siam.org/sam-bin/
dbq/article/27997.

Zeidan:1988:NCO

[ZZ88] V. Zeidan and P. Zezza. Necessary conditions for optimal control problems: Conjugate points. SIAM Journal on Control and Optimization, 26 (3):592–608, May 1988. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zeitouni:1992:OTP

[ZZ92] Ofer Zeitouni and Moshe Zakai. On the optimal tracking problem. SIAM Journal on Control and Optimization, 30 (2):426–439, March 1992. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See erratum [ZZ94].

Zeitouni:1994:EOT

[ZZ94] Ofer Zeitouni and Moshe Zakai. Erratum: “On the optimal tracking problem” [SIAM J. Control Optim. 30 (1992), no. 2, 426–439; MR 92m:93054]. SIAM Journal on Control and Optimization, 32(4):1194, July 1994. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [ZZ92].

Zhang:2009:VSL

[ZZ09] Qing Zhang and Xun Yu Zhou. Valuation of stock loans with regime switching. SIAM Journal on Control and Optimization, 48(3):1229–1250, ???? 2009. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhang:2013:QDB

[ZZ13] Qiang Zhang and Ji-Feng Zhang. Quantized data–based distributed consensus under directed time-varying communication topology. SIAM Journal on Control and Optimization, 51(1):332–352, ????

2013. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhang:2015:PSO

[ZZ15] Haisen Zhang and Xu Zhang.

Pointwise second-order necessary conditions for stochastic optimal controls, Part I: The case of convex control constraint. SIAM Journal on Control and Optimization, 53 (4):2267–2296, ???? 2015. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhang:2017:PSO

[ZZ17] Haisen Zhang and Xu Zhang.

Pointwise second-order necessary conditions for stochastic optimal controls, Part II: The general case. SIAM Journal on Control and Optimization, 55(5):2841–2875, ???? 2017. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zheng:2020:ISS

[ZZ20] Jun Zheng and Guchuan Zhu.

Input-to-state stability for a class of one-dimensional nonlinear parabolic PDEs with nonlinear boundary conditions. SIAM Journal on Control and Optimization, 58(4): 2567–2587, ???? 2020. CO-DEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhu:2023:WPM

[ZZ23] Bin Zhu and Mattia Zorzi. A well-posed multidimensional rational covariance and generalized cepstral extension problem. SIAM Journal on Control and Optimization, 61 (3):1532–1556, ???? 2023. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). URL https:

                                        //epubs.siam.org/doi/10.
                                        1137/21M1452937.

Zhu:2019:ERA

[ZZCaL19] Shuaibing Zhu, Jin Zhou, Guanrong Chen, and Jun an Lu. Estimating the region of attraction on a complex dynamical network. SIAM Journal on Control and Optimization, 57(2):1189–1208, ???? 2019. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zi-zong:2010:SER

[ZzJh10] Yan Zi-zong and Guo Jinhai. Some equivalent results with Yakubovich’s Slemma. SIAM Journal on Control and Optimization, 48 (7):4474–4480, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhang:2014:HOL

[ZZS14] Kuize Zhang, Lijun Zhang, and Fuchun Sun. High-order S-lemma with application to stability of a class of switched nonlinear systems. SIAM Journal on Control and Optimization, 52(1):120–142, ???? 2014. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhang:2016:WPG

[ZZS16] Kuize Zhang, Lijun Zhang, and Rong Su. A weighted pair graph representation for reconstructibility of Boolean control networks. SIAM Journal on Control and Optimization, 54(6):3040–3060, ???? 2016. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).

Zhao:2010:IHS

[ZZWY10] Yanlong Zhao, Ji-Feng Zhang, Le Yi Wang, and G. George Yin. Identification of Hammerstein systems with quantized observations. SIAM Journal on Control and Optimization, 48(7):4352–4376, ???? 2010. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic).